A Complete Bibliography of Publications in the Journal of Applied Probability

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
     WWW URL: https://www.math.utah.edu/~beebe/

25 March 2025 Version 1.24

Title word cross-reference

(1 + 1) [CL23]. (k_{1}, k_{2}) [Vel04]. (n − k + 1) [BBH14]. (n − r + 1) [BHB18]. (R, Q) [BPS23]. (T, P_{j}) [Jaw84]. 0 [C.12]. 0.234 [She13]. 10 [HM97]. 2 [Fad74, Had05, MS10]. 2 × 2 [Nei01, Van86]. 3 [Ors87b]. ^{k} [CP92, Fal94]. ^{X} [Abr01b, BT73, Con94, LLPC94, LKT90, LS96, Miy90]. ^{Y} [BT73, Miy90]. _{k} [Fin69, HF70, Hok77]. _{λ}, τ^{M} [AHN90]. _{λ}^{M} [LA98]. _{m} [De 85c, SP92]. _{r} [ASS85]. α [DN04, Shi06]. α(P^{k}) > 0 [Mad75b]. AR(1) [BMR16]. B [MS05a, MR11]. β [MS21, AD13]. C^{2} [Ken92]. · [Web79]. χ^{2} [Cam73, Lin89]. cμ [CCPW92]. D [Che81b, Dsh01a, Fer95, RS15, Aff89, CMW16, DH89, Mar86, Van89b, Van06, Yun00]. D[0, ∞) [Lin73]. d ≥ 2 [FNV11]. E_{k}/M^{b}/1/m [Van22]. EARMA(1, 1) [Ano81e, Che80b]. ℓ^{p} [KR18]. ℓ_{1} [CM98]. ℓ_{p} [Oes18]. ϵ [MT12]. F [AARJ10, PH88, PR95, SM78]. F^{α} [HN00]. G [Gel93, HP95, LF08, KOH23]. γ [HJ14]. Garch(1, 1) [BG05]. GARCH(P, Q) [Kan23]. h < 1/2 [AG12]. H = λG [Miy95]. i [Tav81]. ∞ [ALRW99, C.12, CP92, Con94, Dvu84, Fal94, FG98, FY83, GW11, GMS96, GP98b, Huf87, Hun86, IJB18, KS94a, LKT90, LS96, Nar68, OP86, Pre97a, Pre97b, Pre02, Sta85, BNW13]. ∞ + 1 [Zha96]. K [BT73, CK00c, GG91b, Sau23, AI94, BBH18, BEN98c, CM92, CHO22, CP90, Cse79, Dav86, DJ95, Hai96, Hal84, Hun76b, KBBW07, Kol15, LZ02, LD10, Mad75b, Mi99, Mus00, NE07, Nue10, PH88, PR95, Pri19, Sie89, Sta81, Tho85, VB02, Wan92, WJ95, WX08, Wei69b, ZLB08]. km [Sud83]. L^{1} [RT01]. L^{2} [GM97, RT01, Sim19, Win90, HL16b]. L^{p} [DLMR17, Hua95]. L_{2} [BR95]. Λ [CS20a, GIM11]. L exp(μ\sqrt{2 log(1 + L)}) [KOH23]. L log L [LRS09]. LU [GdlI18]. M [ABGH00, Bil76, Çin68, GNRW94, Hun76b, Jan15, Kwe75, LL82, MW08, Pla85, SPS80, Sud83, Whi72c]. M^{[X]} [LT04b]. M^{b}/E_{k}/1/m [Van22]. M^{X}/G/1 [KNR05]. N [JMZ11]. R^{d} [KP23]. Z [DP12b, HK22, LMM10, Ros18]. Z^{d} [BR18, DSS11, LSWX20, dL23]. L [Kla02, Sen95]. μ [GPZ05]. N [CNO^{+}84, Coh88, FH91, FNP19, Kap71, KL17, LLPC94, OS91, Pin71, SB06, Adl76, AG80, Ano86a, BBH14, BBH18, BdHK93, BHB18, BEN98c, CM92, CH91, CP90, Dvu84, Gra85, Hai96, Hal82a, IKS70, JM86, LZ02, LD10, Mi99, Mus00, NSK08, NE07, PH88, PR95, Pla85, QW84, Rub77, SPS80, Sie89, Sim97, Tam70, WJ95, ZLB08]. O(|E| log^{∗} |V |) [HWX18]. p [AAO90, Bie98, BM20, San06, Saw77]. p_{ij}(∞) [JK71]. P_{λ}, τ^{M} [AH00, BKL03, LBL05]. PF_{r} [Ano87g]. Q [GPZ05, Pol90, WZ07]. R [Ano93f, FH91, Hen82, Kij92b, Mul72, BJM93, CK00b, CMW16, Mul73, TL69]. R^{2n} [PM03]. R^{d} [Bél92, Lop14, Tan93]. R^{N} [KKS15, PM03, Ray70]. rho [HL13a]. s [NJS96b, Van06]. S_{1}, S_{2}, ⋯ S_{k}, 2 ≤ k < m [Kwe75]. S_{n}/n [CF22a]. T [BS98, Ano91f, Dur86, Gue91, Hai92]. T^{2} [Dav82b]. τ [De 94]. θ [Byt81]. U [BHTLV20, BS79c, NS22, Owa23, PT23a, RR10]. V [FHL13, GW17, Mat04]. x log x [Olo98]. Y_{t+1} = A_{t}Y_{t} + B_{t} [Hor01]. Z [Lam14]. Z^{2} [BS08]. Z^{d} [Bie98]. Z_{+} [Sim00]. Z_{+}^{3} [Kot94].

-actions [BR18]. -affine [CL23]. -Ary [MR11, CMW16, MW08]. -Ball [Hal82a, Rub77]. -Capacity [GG91b]. -Class [Kla02, Sen95]. -coalescent [AD13]. -Coalescents [GIM11]. -Convergence [GM97]. -Correlated [Bie98]. -Corrélées [Bie98]. -Cube [DH89]. -CUSUM [Had05]. -demimartingales [KL17]. -dependent [Jan15]. -Dimensional [Adl76, Aff89, Çin68, Dav86, Hal84, JM86, Mar86, QW84, Whi72c, Cse79, Coh88]. -Events [Vel04]. -Exchangeable [AG80]. -Expectation [KOH23]. -Fold [IKS70]. -geometrically [FHL13]. -Harmonic [Luo00]. -hop [Pri19]. -integrable [KOH23]. -Invariant [GPZ05]. -Invariants [BS98]. -Isotropic [CM98]. -max [NS22]. -mixing [HL13a, MS21]. -Molecular [TL69]. -Networks [Gel93, HP95, LF08]. -Node [BJM93]. -norm [GW17]. -norms [Oes18]. -of- [PR95]. -Optimal [San06]. -Optimality [MT12]. -Order [NJS96b]. -Out-Of- [LD10, BBH14, BBH18, BHB18, CM92, CP90, LZ02, Mi99, Mus00, NE07, PH88, ZLB08, FH91]. -Periodic [OS91]. -Player [SB06]. -Point [Hai96, Sie89]. -Policy [Dsh01a, LLPC94, LA98]. -Polytope [Van89b]. -Processes [GPZ05, Lin89, Pol90, WZ07]. -reflected [HJ14]. -Runs [WX08]. -SAT [MS10]. -Scan [CK00b]. -Server [Tam70, Tho85].

-Smallest [Wei69b]. -Spacings [GNRW94]. -spectral [HL16b]. -Sphere [Ors87b]. -spheres [KR18]. -Stability [MS05a]. -Stable [DN04, Shi06]. -Stage [AI94, Fad74]. -star [FNP19]. -Stationary [Byt81]. -Statistics [Dur86, BS79c, RR10, BHTLV20, Owa23, PT23a]. -Step [Ano91f, Gue91, Hai92]. -th [BM20, Kol15]. -th-nearest [CHO22]. -th-Order [AAO90, Wan92]. -Theorem [Win90]. -tons [Sau23]. -Topology [De 94]. -Transient [Ano93f, Kij92b]. -Uniformly [Mat04]. -Valued [Fer95]. -Word [KBBW07].

/1 [BT73, Bhu06, Fin69, SP92, VB02]. /1/k [Miy90]. /G [BT73]. /G/ [Con94, Fal94, LKT90, LS96]. /G/1 [HF70, Hok77, LLPC94, LT04b]. /GI/1/k [Miy90]. /GI/1/n [Abr01b]. /m/ [CP92]. /M/1 [Web79]. /PH [VB02]. /s [De 85c].

1 [Ano72g, AF83, AT96, AK97, AIH88, BM89a, BT73, BS02, Bäu97, BGT98, BS76, BKNZ91, Bhu06, BM94, BY97, BPSZ10, BG00, CP01, Cha94, CY90, CR97, Coh67a, Coh71, CL93, CT76, DN80a, Dal65, DS94a, DT80, Dos85, DGM04, DN99, DS04, Enn69, FF91, FT80, Fin69, FR10, Gil76, GU92, Gre89, HL76, Hea65, HF70, Hey64, Hey74, HKN89, Hok77, Hoo83, Hoo70, HL04, HPR86, HAPV12, Ino19, JS96, Jen66, Jen72, Kat12, KS86b, KR88, KZB05, Ken74, KKK07, KM93, Kyp72a, LLPC94, Lem74, LZ94, LO95, LSC97, LH02, LT71, LT04b, Lou78, LN94, MH76a, MT00, Min79, NG95, Neu68a, Neu81, Neu94a, Ott79, Ott84a, Ott84b, Pak70, Pak73c, Ram84, RL88, RHK90, Ros87a, Ros73, Ros74, RS99, Rou96, Sch88a, SGC16, SP92, SZ93]. 1 [SDH94, Tak92b, Tam71, Tan79a, Tan94, VB02, VD69, Wak91, Wan99a, Web79, Whi72a, Wil88c, YC01, YT83, ZZ04]. 1-FIFO [CN86]. 1-Type [He03, LT00, MNS01, LH06]. 1/e [Pfe89]. 105D [Yan90].

2 [Han87, RS90, Wie89].

= [Van91b].

A. [vDB87]. Abelian [Bin69]. above [Adl76, Bon13, Fro05, Has76]. Abscissa [HTV92]. Absolute [Bra02, Cha17, DN19, GS79, KNT10, MKN11, ZY08, Luo14]. Absolutely [CV09, KLP12]. Absorbing [Bru78, DS65, DS67, HS69a, Kas66, LRS94, Nak98, Nak06, Rus78, Sch92c]. Absorption [DP00, Dir75, Gal66a, Gal93, Gre76a, HS69a, Kem98, Neu78, Pak78b, PS80, PS83, Pol99, Wal98, AM16, Zhu23]. Abundance [SS83b]. acceleration [DR15]. Acceptable [Pre06]. Acceptance [AF87, Haw68, Miy92, Rob99, She13]. Accepted [Gil76, The79]. Access [BFS09, Sto08]. Accesses [LLS84]. Accessibility [Agr91, HLF19]. Accident [Ano86i, Bar86]. According [KX09]. Account [DDM98]. Accumulated [FK89, Ris86]. Accumulation [BSW07, FKK91]. Accuracy [GL83].

Accurate [GW85, XS03]. Accurately [DFSD01]. Acid [SS85b, ZS92]. Acknowledgement [Ano83a, Ano86a, Ano90a, Ano92a, Ano00a, Bau84a, HRF85, Sta87a]. Acquaintance [BS13a, BS15b]. Acquired [MD96]. across [Ano66d, Hor65]. Action [Har09]. Actions [Las94b, Wan76a, BR18]. Active [LD10, Mi99, MDG09]. Activity [FS07, FSW12, Hey99, Whi88]. Actor [AR05a]. Actual [HL76, Min79]. actuarial [YL16]. Actuated [Cow78]. Acute [Hal82a]. acyclic [DO16]. AdapFail [LR14a]. Adapted [CJ08b]. Adaptive [Agr91, BCP00, Dos81, HS04, Her94, Her99, Pap05, RR07, Sol04, Sta74, Tam85, VOZ95, Vis86, YT83, AW15, Dai14]. Added [Rus83]. Addendum [Pia00a]. Addiction [BD88a]. Additional [FH72, Ott84a]. additions [Sim21]. Additive [Ano07e, Bre08, Bre10, Che98, ÇP72, DIKM10, Das96, Eba84, Jac12, KF70, KS02a, Kha77, Liu11, Mac01, PP77, Spo07, Ste06, Zaj95, FVZ19, Iva14, dL23, vdHKPS23]. additive-increase [vdHKPS23]. adjacent [SP22]. adjacent-swap [SP22]. Adjusted [BC73, Hoo87, Tro83]. Adjustment [AN95]. Admission [CL11, Emm72, Les06, Lew01, LAF02, LR04, Ye07]. Adsorption [Orr69]. Advances [PWW12]. Advancing [Dan77]. Advantage [MRP04]. Advantages [JP69]. Adventures [Gan88a]. Affects [SC75]. Affine [BR95, CL23, GGZ19, KS17]. after [Pie07, PP07]. Aftereffect [Mej65]. Aftereffects [GM84a]. against [KS05, vLR13]. Age [ADLN82, Ano78f, APS74, Ber76, BJ79, BBS85, BK73a, Coh89b, CM69, Don76a, Don76b, Ebr97, Fea80, GD74, Gol78b, Gre76c, Kap74, Kap75a, Kap75b, Kat02, KNS73, KMR75, KBM06, KS02b, LY95, Lev77, Mi02a, MA69, MJP87, Mur80, Pak78a, Qui78, Rad73a, Sau76a, Sav88a, Sgi01, SR70b, SC81, Tav80a, Wau74a, Wei66b, Wei66a, Wei69a, Wei70, Wei72, Wei74, Wei75a, Wei76, Wol88, AFP22, AA17, Hon13, ISK18, JL22, KS18, LM20, MC13, OT16]. age-biased [AFP22]. Age-Dependent [Ano78f, APS74, BBS85, Coh89b, CM69, Don76a, Don76b, Fea80, GD74, Gol78b, Gre76c, Kap74, Kap75a, Kap75b, LY95, MJP87, Qui78, Rad73a, Sau76a, Sgi01, Wei70, Wei72, Wei75a, Wei76, Hon13, OT16]. Age-Replacement [Mi02a]. age-structured [JL22]. Ageing [BI05, DKS86, FGL00, Fin07, HL05, Kle83b, KM92, KL89b, LS97a, LZ02, Li04, MP04, PS02a, SS99, SD94, AMV21, BSC18, BBH14, GM18, Szy19b]. agent [KADK19]. agent-based [KADK19]. Ages [Eth92b]. Aggregate [Tan07, YWK14]. Aggregated [FR86, Lar98, Ryd96]. Aggregates [Ano78g, Dav76]. aggregating [LP23]. Aggregation [Cao95, HI01, KR08, LRS94, BDT22, FL19, PL17]. Aging [CM98]. Ahead [AH77, Kab83, Yas88]. Ahsanullah [Hua74b]. AIC [Oga80]. Akaike [Han82, Ton75]. Alabama [JL12]. alarm [LPU20]. Aléatoires [HK94, Ray70]. Alfréd [Ken70]. Algae [ML77]. Algebra [Gor70, Hou12, Kin66]. Algebraic [Daw81, FHTZ10, HMPS03, LT73, SCZ77, Sch86, LLL20]. Algebras [Hol68, Hol92]. Algorithm [BB80, DJ09, DS04, Grü98, Irl06, KMW11, LR93, Mah03, MT12, Osb82, OS86, OS07, Sad01, Sol04, BS14a, BHZ22, CM16b, DLMR17, HMBL14, MV21, RRT22, SK23]. Algorithmic [ALH01, Ste03]. Algorithms [AR05b, BG06, Bél92, CD98, DG99, FT78, GG87, Kur87, Mar15, MSS09, Ram81, Rob99, RR07, RS00, Suz10, AW15, KSS16, LR14a, LY20, MRSA21, PLRB23, RR16]. Aligned [Coo89]. Alignment [Spo04]. All-Optical [Mor06]. Allele [CG76, Kle88, Saw77, MK15, MK17]. Alleles [Ano77g, Avn79, Eth92a, Eth92b, Fea03, Gla77, Gri92, Joy94, Sch91b, Wat76, Zho15]. Allelic [Möh10]. Allelopathy [Lan05]. Allen [KLL15]. Allocating [Rig96]. Allocation [Cow92, DG93a, ENPS86, Gan04a, Git72, Git74, Gla90, GADL04, HW96, LMS72, LD10, Loi05, Men96, Mi99, MDG09, MWCY01, Rev93, SS97, SG93, Tsi07, Whi88, GR24, KA22, KS16, KSC13, SZY19a, YZZ21]. Allocations [CY08, Gri98, MMD11, SM94a, AGP13]. Allowing [Cer88, Kap74, Pak74, Pak76, Rad73a, Rey70, Tes84, Yan72]. Allozyme [GWMO79]. Almost [Bag86a, Bèr04, BCF09, Ber13a, CD92, Dav73, GD74, Gre77b, IM10, Iso82, KS78, Kuc82, LK77, Lin94, LHT17, Mei10, Pur87, RHK90, RT73, dHR73, HMBL14]. Almost-Lack-of-Memory [CD92, Lin94]. Almost-Sure [GD74, IM10, Mei10, HMBL14]. Aloft [Jac99]. ALOHA [Ana92, Sha88]. alpha [AES19]. alpha-mixture [AES19]. ALPS [RRT22]. Alternate [Die82]. Alternating [ACSS11, CY65, May05, Pou92, Yad70, DM18, RDM19]. Alternative [Ano64c, Gla79b, Gla82b, Hos09, Hos86, Kei64a, RV08, Sly07, Wei80a, KL16]. Alternatives [Had05, Yao93]. Always [CSZ04]. AM [Ano04f, GS08, Sen17]. American [Dol10, Eks03, Eks06, Iva07, Pau01]. Amid [Han67]. among [Bha91, Cer88, CM69, ES23, Gra00, Sam71, TK15]. Amount [Ano04g, Ste04]. amounts [Zha21]. Amplitudes [McC97]. Analogue [Di 99, Hey80, Rit80, Psa22]. Analogues [BD88b, Hey71]. Analyses [Sha80]. Analysing [Lot84, MH76b, Str75a]. Analysis [Agn72, Agn75, AR05b, AS04, AMAKAH97, AC02, Ano05e, ALH01, BM89a, BMY02, Ber10, BKNZ91, BN94, Bin78, Bla84, BM94, BDO07, BC92, Bre69, BCMW83, Byt81, Car67, CG96, CH97, Cow81, CX01, De 85a, DJ09, DO97, Dsh01b, FS80b, FH95, GM05, GXS12, Ger23, Git73, GP98b, GMDB03, GP04b, Hal69, HST83, Hey88, HCK02, ISK95, JC06a, Jen66, KKLO86, KK96a, Kat02, Kim00, Kro97, KG02, LR09, Lan87, LLPC94, Lee97, LSC97, MSS92, MW08, MF01, MP92b, McN70, Mia13, Miu77, Møl88, Mor79, NSK08, NNY20, Oga01, Ors89, Oza86, OI01, Par86, PS99, PT97, PW96, Pro96, Pur75a, RS89a, Res04, Rip76, Rou96, Roy72, RS00, Sch75a, Ses96, SS85b, Sol04, Sta01, Suz10, TTH93, TH11, VK00, VK10, WK65, Wes82, ZS92, Att19, BPS23]. analysis [BJ24, DO16, FS16, KSBK23, KSS16, LN13, LC16, LZ14, LY20, MRSA21, Ros14b, SGC16, SDH16, Spi15]. Analytic [DP10, Mas84, Sri74, SB06, ZY70, But19]. Analytical [BG00, GBG02, Van22]. Analyzed [Aas85]. Analyzing [Fei93]. Ancestor [Fea02]. Ancestors [BM01b, DY97]. Ancestral [Man13a, PS09, Sag99, BCKS13, Man13b]. Andersen [ABI14, Iva16, PP07].

Andrei [Par88a]. Angle [BF91, Mor66b]. angular [BK21]. Animal [MM94]. Anisotropic [Hor98, PS16, HKT24]. Anisotropically [Moo74]. Anisotropy [RS89a]. Annealing [ABD01, AF87, Bél92, Bor92, CD98, FF03, Kol95, Loc96, Nie95, Win90]. Annihilating [OS01, Sud02, Sud04, LTZ18]. Annual [SLZ01]. Anomalous [AK23]. Anomaly [Bur79]. ANOVA [Spe86]. Anti [Hud73]. Anti-Commuting [Hud73]. Antistarshapedness [Lyn87]. Antiviral [Ber94]. any [CR17]. Apollonian [CMW16]. Appendix [Hör05]. Applicant [Tam09]. Applicants [Rog87]. Application [Agn72, Agn75, Ano86i, AB86b, Bar86, Bes72, BK64, BM01a, Bro64, BM66, Bro04b, Cam73, Cam97, CK00b, CLS09, Che76b, CJ98, DS94a, DJ05, DM01, Dek96, Dia73, FI87, GWMO79, Gao04, GW94a, GJ80, GM73, GMSB10, Gri98, HTW73, Hey73, HNF85, Hop86b, Imh78, Kat77b, Kim94b, Ksi89, KMSP12, Lai77, LJ78, Lin89, MZ01, Mak94, MN79, MK73, Mi94, Miy90, MT02, Mod66, Mod67b, Moo73, NH82, Naw83, Nie85, OK86, Pak87, Phi81, PS76b, RC73, Rao66, Rob05, RMV97, Sah78, San71, Sch80a, Sch82, Sen87, Sim00, SD76, SM94b, SV71, Ton85, Toy97, Via95, Wan76c, Wat64, Wes82, Woo83a, AH13, BM17, CH13, FS16, Fin13, Kan23, Kol15, Mar16a, PLRB23, PS16, SS17]. Applications [Alp07, AKP91, And70, AG80, Att89, Bal99, BR87a, Bal01, BS04a, BCF09, BCPR11, BCY97, Bha91, Bil81, BS69a, BENP94, BSC04, Bre69, BS92b, BS93, BGP80, CT04, CJ06, Che03b, CV09, Coh81b, Coh83, CC96, Con71, Coo74, CH70, Dar66, Del78, DLU09, Di 99, DYY01, EMR71, Emb83, FG99, FS97, Gal73, Gal74, GY95a, GR05, Hal69, He03, Imh84, JL98, Kap72, Kar88, KL03a, KM97, LY95, LX00, Lin87, Lin04, Liu11, Lo86, Loy88, Lyn00, MP92a, Mar00, Mat84, Men82, Mil79, Miy79, Miy82, MY88, Miy94, Möh05, Mor73a, NB03, Orr69, Ōsa85, Pic80, Pic84, PP73b, Pur71, Pur78, Res71b, Rob90, Rui08, Ser92, Sev67, Sev69, SSSL10, SS89, SY91, SS83b, Sri98, SW88, VK85, WCW05, Wei66b, WC04, YZ82, YSS91]. Applications [YZ06, Zha10, AHO20, AA17, BBN16, BS14b, BS16b, BKM19, CC15, DM21, DDS20, FGH^{+}21, FvLB13, Ino23, JKL^{+}20, KK16, KL17, LL13, LWM20, MK15, MPS15, PL17, PT23a, SST18, Wir18, YBL23, YL16]. Applied [Ano13, Ano14, Ano15, Bon02, DVJ08, Gan88a, Han65, Hit88, Ken88, Leh73, Neu88, Ano22a, Gan10, Ano89e]. Applying [CL98, FCvS15]. Appointment [And75]. Appreciation [GG94b, Nas08, Reu82]. Approach [ARM98, Ano68d, Ano85g, ALH01, AT87, BD12, BJM93, BN02, Bro00b, BS09, CG99, CLS09, CD09, Deh84, Dev79, DGNR95, Don86, DP10, DR81, GNR89, Grü98, GZ06, Hei01, Hey94, Hol79b, HNG93, Kal92, Ken91, Kou96, LT76, LP85, LZ94, LS97a, Lil68, LT73, MH76a, Mas84, McN82, McQ67, MW90a, Mil94, Miy79, Mor68, MS10, Nic86, OI01, PS89, Pol82, Sch93, Sch86, Spe86, Ste03, Sti94, Sto91, Suz10, SB06, Tam85, Web80, Yao96, YT83, Yos84, Dai14, De 22, EOV20, FL19, HK15, Ker20, KK24, LLZ17, LC16, LP18, LWW18, PSZ13, SZY19a, Yam22, YW21, vdHKPS23]. Approaches [AK86, BS01a, Wit90]. Approaching [BL00, BT23]. Approximate [BB96, Bar72b, BS97, BP99b, Del85, Dur86, Gir82, Her74, KADK19, Ruk00, Whi73].

Approximately [FW91]. Approximates [Gri72]. Approximating [AD94, BK12a, Ben98b, BZ09, Cla12, Cra88, DS94a, Hey91, Kur71, MH76a, Wan92, MS23]. Approximation [BJ07, Bal95, BB90, BP10, BC72, BP96b, BN05, BGT04, Bre81, Che96, Che98, CV01, CK07, CT90, Dal10, Dan82a, DJ00, EV82, Erh00, FF96, FWL03, GK18, GL83, Har73b, HJ14, Hey74, Hor98, KR92, KPW94, LF08, Lud73, MCC03, Miy86, MH76b, Nak08, NO85, Nor77, O’C93, Pek96, Pek06, PRCS09, PW67, Pós09, Rad76a, RR10, Roo03, Ruz04, Ruz10, RS10, Sel91, Sez18, SS96, Sim00, Ski12, SW87, TSCD08, TG99, Tuf99, Twe75b, Vel04, Wan77a, WX08, WM77, Wit90, WPN80, Xia97, AM16, BN17, BY14, BS22, CL14, Dal16b, DJ18, Gau22, GPW22, HMBL14, JY21, KLL15, LMS18, LY22, LX20, Pen18]. Approximations [Ano74e, AR01b, BR96, CK00b, Cov09, Del78, Dol10, DR81, GY95a, Gav68, GL76, GGS^{+}95, GNRW94, GR96, Gri73, GG77b, GMS96, HGR90, Hog86, Hyn94, Igl65, Igl69, JC06b, KSF80, LG81, LL90, Maz86, NFK99, NR78, O’N96, QM86, San80, Sie79, SB09, SS80, Spo01, Tho75a, Twe98, VC96, Wan90, Wan93, Wes82, vHS86, ADPR18, Bor21, Bro15, But19, LOR19, Leh21, LB23, PT23a, SWW21, YML17]. arbitrage [DN13]. Arbitrary [Bar72a, Can74, LS02b, Miy92, Rog87, Ros07, Sau76a, Ros14a]. Arc [BD88b, GS94, Man77]. Arc-Sine [BD88b, GS94]. Arch [KL03a, Ano06f, GDP04]. Archaeological [Nak91]. Archimedean [Bal94, FPW19, JP16]. Arcs [Hen86b, Hol80a, HS87b, Huf88b, Jan88, Sie78a, SH82]. Arcsine [FGH^{+}21]. Area [Bra82, Dow72a, Dow72b, Lef79, RJ65, DPW21]. Argmax [RR94]. Argument [Wei80a]. Arguments [NTW89]. Arising [Ath71, Bri71, BPR08, Che76b, Cow01, DR81, FI98, GG74, HN01, HS93, Jag69, Jen66, Kar76c, Law79, Mor71a, Mor06, Pap94, PR67, Rou90, VAW04, Wic79, Wil96, YI99, ZR01, CDP19, MN16]. arithmetic [Kev17]. Arm [FW91]. ARMA [Bro04a, HZZG86, Kab83, Lin99, LS92b]. Armed [O’F89, CLZ24]. Army [Bré93, Kon96]. Array [And85, Sri98]. Arrayed [LV72]. Arrays [AM10, CK00b, Dzi86, Kad74, SKF79, KLP22]. Arrival [AK93, Bor04, Cha97, DN99, DS04, Eco03, Fak82, Fal94, Gil76, Had74, HW84b, KS08, Kul86, KA03, LR80, Lee80, Li96, NA04, New68c, New68a, New68b, RCLW10, RCC12, RF88, Ros76b, SP92, SDH94, Wau74b, FW21, HL13a, Ros14a]. Arrivals [AMR01, AGC99, BM97, BS94, BD94b, CR97, Çin67a, Coh76, De 85a, EF90, GGS91, GR74, HW84a, Hey82b, HCK02, KMS83, Kyp71, Lel05, McN68, MW90a, Nar68, NC72, NR78, OM72, RN80, Ros78, Sha66, SS85a, Sha75, SM97, SG88, Tho85, Tin85, VS90, WT05, YC01, Ros14a, ZL15]. Arriving [PRS08, SWW21]. Artificial [Rob90]. Ary [MR11, CMW16, MW08]. Asadi [AA15]. Ashrafi [AA15]. Asian [BRY04, Ger11, HW02, Pat13, RS95b]. Asmussen [GR12]. Aspects [DKS86, DF81a, Ewe82, Ger90, Kin02, Luo00, McD78, SMS94, Sri67]. Assemble [Cha97, Eco00, Eco03]. Assemble-Transfer [Cha97, Eco00, Eco03]. Assembly [Cra74a, Har73a]. Assembly-like [Har73a]. Assessing [Ebr01, Ebr08, Ebr11, EY11, MCC03, NMG00]. Assessment [Bro04b, GXY18]. Asset [BJV12, Hey99, Tan11, Des15, KSC13, SZY19a]. Assets [EJT05]. Assignment [DLR80, Rig90, Web78, BJ14, BKJ16, LM24]. Associated [BM05, BLBL10, Ber77b, BS04b, GG91a, GJ78, GN06, Gir82, Gla92, Gue97, HG81, Ito89, JKP11, Kel12, KK69, Lin87, Lit92b, MS94, MV95, Ros80a, SS83a, Wat88, Xu09, Mar16a, Men76, Per23, Wir18]. Association [Ano10e, HNF85, Mar07, Vrb85, Wil82c, LS19, Tu19]. assumption [Sei16]. Assumptions [PW81, Twe88]. ASTA [MW90b]. Asymmetric [FF96, Lef06, LW19, LR14b]. asympotic [KMG19]. Asymptotic [Abr94, AC94a, Ail68, Ali99, Ano73g, Ano74g, Ano78e, Ano78f, Ano91h, Ano92f, Ant78a, AS09, APZ11, BM15a, Bag82, Bag86b, BR89, BdHK93, BC23, Bal01, BMZ07, Bee75, Bel23, BS06, Ber85, BP16, BAAB00, BLL80a, Bla84, Blo86, Bor96, BK06, BLT11, Bro64, BH75a, But19, CM92, CCC03, CLCZ14, CCG90, Cra69, CCFR00, Cur22, DG74, Dav74, De 85a, DT80, Dip98, DG99, DF81a, Eng80, GA99, Gal66a, GL71, GW85, GXS12, Gir86, GMD05, Gre74, GS95, HS74a, Hai78, Han73a, HT12, HW84b, HI20, HH96, HM99, Hok77, HST75, HJL19, Hüs82, IKP01, IPV03, JY21, KMW11, Kal90, KS72, KBBW07, KK96a, KS13, Kel77, Kha68, KL77, KL78, Kle91, Kle94, Lad05, LN13, LK77, Len81, Lew69, Li05, LS91, LNTZ99, Lo86, LR07, MW15, Mar86]. Asymptotic [Mas82, MRR02, MP92b, MQ00, MPHL10, MI83, MR03, MD18, Mor79, Nag01, Nag82b, Nar68, NT04, Nic84, Ols67, OS07, PKO16, Pag71, Pic76, Pic79, PC16, Rad73a, Rob05, Roe85, SC75, ST86a, Sch80b, Sel83, Sgi01, SS89, SK98, Sri98, Sta87b, Sta01, Sve95, Szc97, Tan86, TV88, TH72, UB74, Vas82a, Vas84, VGT90, Ver14, Voi96, Vrb85, WF99, Wan14, War76, Wau72, Wau74a, Wei76, Wei69b, Wei82, Wes73, Wil92, XY09, XS03, Yan73, Yas84, ZJ04, BHTLV20, DN13, FLM15, GP13, KP22, LPN13, NR20, PT22, VKR14]. Asymptotically [DN04, Don80b, Fen93, Iso82, Kle94, Rub77, Sen73a, WY22]. Asymptotics [BF08, BCPR16, CW96, Dai04, DESM10, DM03, DSZ13, DS13, DYY01, Ger11, GW94b, GIM08, Gol95, GHW22, HM09, JL98, KKK07, Lel07, LM99, MS99, MKN11, PZ07, PPBT99, Pia08, Pia09, Pro96, RRS05, THP12, Teh09, VP10, WW06b, YZ06, AB16, DPW21, KY17b, KLY16, KM15b, RNJ16, SUY21, SWZ18, SRN21, YWK14, Zhe09]. Asynchronous [Sag99]. At-the-Money [Dur08b]. atlas [FF20]. ATM [DYY01, HW96]. Attachment [Bha68a, Mor71a, Sri67, AFP22, ARS21, BKK19, Bri20, CF22b, Jor18, LM20, SRT^{+}16]. Attainable [Teh10, Tsa94, Tsa96]. Attained [MY91, YM91, Yao13]. Attaining [Abd87]. Attempts [Han87]. Attracted [Res71a]. Attracting [BM90, Mut87]. Attraction [Ano77e, Gre76b, LS00, RVJ68, Gra18, ZZ13]. attractive [Gén23, MZ20]. Attributes [Stö74]. Auctions [Hin04]. Augmentation [ZL96]. Augmented [GS87]. Author [Ano89e]. Autocorrelation [DM01, God77, Han82, HT78, Rey68]. Autocovariance [Ren88]. Autocovariances [LZK06b]. Automata [Gra99, Nue08]. Automated [GKS14]. Automatic [BCIM01, ZRU68]. Autopsy [ADH93, Gås98, Mei81]. Autoregression [GS79, Kan75, McK82]. Autoregressive [AAO90, And84b, Ano75f, AR89, AP86, BH73, CFZ96, Chr12, Cip88, Col75, CS82, DG93b, Dur86, Fel91, HK77, HB75, HW83, Hey73, HB99, HCK02, Hyn94, JP93, Kul87, Law80, Law92, MP92b, McK81, Nic86, Oza81, Pap94, Peg80, Sim90, Sim93, Tho86, Wol75, ZG04, BLM20]. Autoregressive-Moving-Average [McK81]. Autostrada [Con84]. Availability [AB00, Bax81, Bax83, CX01, FH90, FH91, KFM06, KKY02, Mi95, SM82, GN17]. Available [Tam09]. Avalanches [GMV06, Ste86]. Average [AHN90, AAL77, BKL03, CG97a, CCFG96, CCdO05, Cos93, CD09, FT78, God77, Gro80, GHL03, GZ06, IK70, KPSC08, Ksi89, Lac05, LS87, Las94b, LL77, Lew01, MW08, MW71, McK81, McK88, Mil81, Nar68, Pea67, PRS08, PT86, Ros70a, Ros78, Tho86, YM91, Yeh85, Zha14, ADPR18, CCdOS15, CCCSdO24, GH21, MRSA21]. average-based [MRSA21]. Average-Case [MW08]. Average-Cost [Las94b, Mil81, Yeh85]. Averages [Bru71, Kul06, MW90a, NP09, PR08, SY72, BS15d]. Averaging [CS77, Dah85, Gén23, Sol04, Spi22]. averse [Sha13]. avoid [LW19]. Avoidance [OS07]. Avoiding [BJM93, Di 98, KLO99, Law83]. Axes [Bar75c]. Axiomatic [Ano99a, KL16].

Back [Ano64a, Ano64b, Ano65a, Ano65b, Ano66a, Ano66b, Ano67a, Ano67b, Ano67c, Ano68a, Ano68b, Ano68c, Ano69a, Ano69b, Ano69c, Ano70a, Ano70b, Ano70c, Ano71a, Ano71b, Ano71c, Ano71d, Ano72a, Ano72b, Ano72c, Ano72d, Ano73a, Ano73b, Ano73c, Ano73d, Ano74a, Ano74b, Ano74c, Ano74d, Ano75a, Ano75b, Ano75c, Ano75d, Ano76a, Ano76b, Ano76c, Ano76d, Ano77a, Ano77b, Ano77c, Ano77d, Ano78a, Ano78b, Ano78c, Ano78d, Ano79a, Ano79b, Ano79c, Ano79d, Ano80a, Ano80b, Ano80c, Ano80d, Ano81a, Ano81b, Ano81c, Ano81d, Ano82e, Ano82a, Ano82b, Ano82c, Ano82d, Ano83b, Ano83c, Ano83d, Ano83e, Ano84a, Ano84b, Ano84c, Ano84d, Ano85a, Ano85b, Ano85c, Ano85d, Ano86f, Ano86b, Ano86c, Ano86d, Ano86e, Ano87a, Ano87b, Ano87c, Ano87d]. Back [Ano88e, Ano88a, Ano88b, Ano88c, Ano88d, Ano89a, Ano89b, Ano89c, Ano89d, Ano90b, Ano90c, Ano90d, Ano90e, Ano91a, Ano91b, Ano91c, Ano91d, Ano92b, Ano92c, Ano92d, Ano92e, Ano93a, Ano93b, Ano93c, Ano93d, Ano94e, Ano94a, Ano94b, Ano94c, Ano94d, Ano95a, Ano95b, Ano95c, Ano95d, Ano96a, Ano96b, Ano96c, Ano96d, Ano97a, Ano97b, Ano97c, Ano97d, Ano98a, Ano98b, Ano98c, Ano98d, Ano99b, Ano99c, Ano99d, Ano99e, Ano00b, Ano00c, Ano00d, Ano00e, Ano01e, Ano01a, Ano01b, Ano01c, Ano01d, Ano02a, Ano02b, Ano02c, Ano02d, Ano03a, Ano03b, Ano03c, Ano03d, Ano04e, Ano04a, Ano04b, Ano04c, Ano04d, Ano05a, Ano05b, Ano05c, Ano05d, Ano06a, Ano06b, Ano06c, Ano06d, Ano07a, Ano07b, Ano07c, Ano07d, Ano08a, Ano08b, Ano08c, Ano08d]. Back [Ano09a, Ano09b, Ano09c, Ano09d, Ano10a, Ano10b, Ano10c, Ano10d, Ano17a, Ano18b, Ano19a, Ano19c, Ano19e, Ano19g, Ano20a, Ano20c, Ano20e, Ano21a, Ano21c, Ano21e, Ano21g, Ano22b, Ano22d, Ano22f, Ano22h, FI87]. Back-off [FI87]. Backbone [KMSP12]. Background [AA06, MS85, Sch88a]. Backlog [ALRW99, KR88, Sez18]. Backward [Bou95, CB83, DP12a, ESC15, KS77, Lam91, LS04, Möh94, Whi82b, Win89, Yan74, Jas15, KLL15, WZ13]. Backwards [CS85b, Käm89, Möh10]. Bacteria [Bha68a, GKRV94, Man74, Sri67]. Bacteriophage [Gan65]. Bak [GMV06]. Balance [AG07, Fak90, Kin78, Whi85c]. Balanced [BDM09, BJT20]. Balances [Cha97]. balancing [HLVM22, LY20, MBvLW16]. Balking [ZAFP06]. Ball [Aff89, BM84, Hal82a, Rub77, GHW19]. Ballot [Eng64, Men82, Tak64b, Tak64c, Tam01]. Ballots [Tak89]. Balls [DNW07, Mas86a, CHO22, CR17]. Band [GHK04, Has82b, Tho86, HL16b]. Band-Limited [Tho86]. Banded [BP99b]. Bandit [BS92a, Gla79b, Gla82b, GG95, Pek03, CLZ24, NR18, WZ13]. Bandits [CG00b, O’F89, WW90, Whi88]. Bands [HT71]. Bandwidth [APZ11, GRD^{+}00, HW96, GM15]. Bandwidth-Sharing [APZ11]. ’Bang [All10, All10]. Bank [KKL01, BCKS13, CPSJ22]. bar [Möh21]. Bargaining [FG08]. Barrier [Cab84, CL03, CL09, Cre74, DW11, Fro05, Ham05, HS69a, KMSP12, ML98, PS76a, PS83, Sch03, Sch92c, Shi65, Zha92, DL18, DL17]. Barrier-Absorption [PS83]. Barrier-Crossing [PS76a, DL17]. Barriers [BH12, Dom95, Dom96, DKS76, Iva10, KS04, LS04, PB83, SH70b, TW84, BCPR16, Iva14, PG13, Pih15]. Bartlett [Bes72, Gan02]. Base [FG87]. Based [ALS04, Arn75, BBC01, Bas71, BAAB00, BEN98c, BS97, Bou95, Bro04b, CG87, DO12, DJ09, DGNR95, DL02, Dix93, EP95, FPKMP98, GR87, GR89, GM05, GNR89, GM11, Hou12, HL05, HS10, Kle83b, LS07, Les06, LZ02, Li04, MJP87, Nat90, NSB11, OO07, Oga01, Pla04, SW97b, Sha72b, Sim00, SW87, Sta74, SAS04, Vrb85, War76, Yan74, ZC95, AA23, AZA18, BN13, CC15, GM18, HR15, HMBL14, HL13b, KBM21, KADK19, MRSA21, NFS20, RNJ19, Roy14, SK23, WZ18, vD15]. Baseline [Li05]. Basic [MM82, Miy86, MY88, SA72, Sch75a, ZG88]. Baskets [NR11]. Basu [FG12]. Batch [AMR01, BLL80a, BS94, CPS96, Cha97, CP92, CHPT97, Con94, Eco00, Eco03, Fal94, HT91, Kul86, Lan81, Men82, Pea65, Sha66, VS90, vD89, ZX17]. Batch-Arrival [Eco03]. Batch-Ballot [Men82]. Batch-Movement [CHPT97]. Batch-Service [Eco03]. Batches [Moh72, Sha75]. Bathtub [BLSW08, BLSW10, Sav03]. Bathtub-Shaped [BLSW08, BLSW10, Sav03]. Battles [Gol77]. Bayes [AES17, JFS12, Nag01]. Bayesian [AMAKAH97, CCH06, EG02, GG77a, Pap96, Win90]. Be [Rus83, VX04, EL19, Sco82, Was69]. Beat [KPSC08]. Been [Tan79b]. Before [Bau09, Hol76, PP07]. Behavior [Ail68, AT06, Ano65d, Ano66c, Ano91h, Ano92f, Bha68b, BS93, CG97a, Eva71, HW84b, KBBW07, Kar68a, Kar68b, Kei64b, Kei65, Kle91, LPLG08, LNTZ99, Lo86, Mas86b, Mi95, NT04, OS07, PRS08, Pit87, RR75b, Sah71a, SS89, Sum87, Tan93, Vas82a, Wu93, ZS04, BJ14, Cur22, FLM15, FVZ19, JY21, KP22, LS20, MD18, PT22, Rek23, Tav20]. Behaviors [BZ08, KHY03].

Behaviour [Ali99, Bal83a, Bal84, BB06, BS06, BD94a, BLL80a, BLS03, Boo89, BL09b, BCMW83, DT80, Deh85, DG93b, DF81a, DF88, Eme72, FF03, GN06, GG00, GM96, GMD05, GR94, Gre74, Grü85, Hai78, Ham95, Har81, Hok77, HST75, Hun95, Kal90, Kel77, Kle94, KLM04, KK06, LPSV00, LZ02, Llo71, Mat84, Meh77b, NS84, Red03, SA72, Sau76a, Sch80b, SK98, Sta87b, SZ93, Vas82b, Vas84, WF99, ZJ04, Bel23, BP16, BL15, CLCZ14, GT18, HJL19, Jas15, KMG19, MP14, Neu23, PKO16, VKR14, Ver14]. Behind [YYZ09]. beliefs [EGL17]. Bell [Fin88]. Bellman [Ang11, Gan84b, GMSB10, Hon13, Kuc82, MY85, Mur80, PK74]. Bellman-Harris [Hon13]. Belonging [Nis74b]. below [KW03, WW06a]. Belyaev [Tam70]. Benchmark [Pla04]. Benefit [Dub81]. Benefits [SM82]. Benes [CN86, GM94]. Benford [All97, JdlR04]. Bergomi [GJP^{+}21, JPS18]. Berkelmans [BBvdMP23]. Berman [Hay00]. Bernoulli [AM10, Bra83, CL85, Che01, Csá94, Gan82, Har90b, HL14, Hol07, Hol08a, Hol09, HS12, Kar83, KS86b, KPS81, KM88, KLY16, Kub24, Lee97, LU96, Lit92b, Nea03, PT23a, Ruz04, SD88, Taï96, Tak95, TTH91, WCC03, Xia97, XB11]. Bernstein [OT15]. Berry [Gol05, Jia12, Jin88, KLP19]. Bessel [Hal83b, Imh84, Lin04]. Best [ES87, EF90, Gil87, Gli00, Gne94, Gne96, GM07, HW79, KS72, KA09, MT06, Pet81, Pet83, Tam86, Tam09, TR89, Yas84, ZL96, BS17]. Best-Choice [EF90, Gli00, Gne94, Gne96, GM07, Pet81, Pet83, Tam86, Yas84, BS17]. Bet [CH06, Pon07, Rut99]. Bet-Dependent [CH06, Pon07]. Beta [GM73, GR13]. Better [Cha00, KPSC08, Lam92, VG92]. Betting [FM97, Got85, Sub79]. Bettor [ET83]. Between [CV09, JKM12, Ala76, AGSC00, BBvdMP23, Bha90, Bra72, Bra82, Bru71, CSC13, CDL88, CHPT97, CT76, DDS22b, DDM98, Dom95, Fak90, FH17, Gro80, Gup84, HN71, HS69a, Har65, Hol68, HS69b, HJL14, KBBW07, KP89, Kin78, KS82, KS80b, Kub24, Kün86, Lam68, Lau80, Law71a, Law71b, LWB08, LT81, LM04b, Lit74, Mac01, Man13b, Mar15, Mas95, Miz06, Mor80, Mor66b, Nag01, Neu67, RH81b, RR78, RP03, Sav88a, Sch92c, SR70a, Ste03, Tam70, The79, Web99, WG93, Yao13]. Beurling [Kal91b]. Beyond [Pol82, DJ18, FHI17, HWX18, KJS15]. Bhattacharya [Sha72b]. Bias [GR96, HP98b, LP00, LAF02, ZPR08, FS16, Man13a]. Biased [ALM06, Che76b, Don91, Gne98, Hop86b, LN07, OC03, AFP22, Bow18, Hu16, Lei23, LSWX20]. Biasing [Gol05]. Bidding [Pak83]. Bidimensional [Lef97]. Bienaymé [Coh82, Hey81, Le14]. Bienaymé-Galton [Le14]. Bifurcating [BZ04, HB99]. bifurcation [Yam17]. Biggins [IK16]. Biharmonic [Hel67]. Bilinear [Ano82h, LB88, PT81, TR89, WW04, ZT01]. Bin [CW86, Hal89, ST08, CR17]. Bin-Packing [CW86]. Binary [BPS12, DJ95, DGK98, Ery10, FD84, GW76, GW80, GW82, KSF80, Mak10, Mon82, Pia08, Pic77, SKF79, Str75a, WM77, XS03, AD13, GN17, HS22]. Binary-Valued [Mon82]. Bingham [De 82, Zho08]. Binomial [AD94, Ano70d, Bal83b, BNY69, BBS90, Bro64, CP82, Coo98, DK11, Eco04, EZ80, LV12b, MP92b, PRCS09, Sie86, Ski12, VP01, VU09, Wal76, Wan81, WX08, Web99, BEER15, BM15b, ESC15, FHIM14, KY17c, MV16]. binomials [FS16]. Bioequivalence [YI99]. Biological [BL09a]. Biologie [Jof82]. Biology [Jag69, Jof82]. Biometric [Hol92]. Biomolecular [CK00b]. bipartite [Lop14]. bipolar [CM16a]. Bipower [BNCPW09]. Birds [HS78]. Birnbaum [Raa80]. Birth [AHP75, Adk69, Ale01, Ano65d, Ano66c, Ano74f, Ano74e, Ano83a, Ano86g, BS01a, Bil81, BGJ12, BG79b, Bra86, Bro86, CP04, CL03, CP03, Cof95, CT81, CSV01, Cor87, CI78, Dan77, Dan82a, Dav65, DW77, Di 98, Don84b, DKM93, FY82, FDVF93, GY65, GR94, Gri72, Gri73, Gri75b, GP99, Hai78, Haj82, HN01, Iva81, KS86a, KT82, Kei64b, Kei65, Kei71, KS91a, Kij92a, Kle81, LJ78, LR93, LPSV00, Lin79, Liu11, LS74, Man74, MZ04, MV95, Mit04a, Mor74, MM03b, Nat75, Nat77, Pfe82b, Phe96a, Phe96b, PZC07, Pur68, Pur75b, RD97, RJ94, RJP97, Sau76b, SV97, SZP07, Sum84, Sum87, Tor79, VSW09, Wau72, Wau74a, Zei91, Zei95, Zha01, dSS04, vDZ05, BBN16, BC23, Cur22, HZ22a, HCY21, JF23, KS18, ZZ13, vD15, Cla12]. Birth-and-Death [Ano74f, Ano74e, Ano83a, Cof95, Di 98, Gri73, Haj82, Iva81, Mor74, MM03b, Nat75, Nat77, Sau76b, KS18, ZZ13]. Birth-and-Mutation [Man74]. Birth-Death [Ano65d, Ano66c, BS01a, BG79b, CL03, CP03, CSV01, Don84b, GY65, GR94, Gri72, Kei64b, Kei65, Kei71, KS91a, Kij92a, LPSV00, Liu11, MV95, Mit04a, SV97, SZP07, Sum84, Sum87, vDZ05, BC23, HZ22a, vD15]. Birth-Death-Migration [RD97]. Birth-Type [Ale01]. Birthday [GWMO79]. Bisection [Bac01]. Bisexual [Bag86b, Bru84, DHT86, GM96, GM97, Hul82, Hul84, Hul01, Käm89, Käm91, Möh94, Möh97, MMR02, MMR07, XW05]. Bispectra [Ros80b]. Bistate [Fu01]. Bit [MZ16]. Bivariate [AB00, Ano90f, BCL09, BKS02, Chu75, Ebr87, GL94, Gol76, Gri72, GG10, HT12, KZ88, KD78, LS87, LL75, MO67, MO93, Roy89, Sch86, Wan90, WY93, Zha94, GK22]. BK [Gol23]. Black [Che76a, CH06, CCHY08, CH10, Kea67, Pon07, Rut99, Sec97, Dol10, FPY04, Whi01]. Blackwell [DM19, JFHL09, LS08b, Lin08]. Blackwell-Type [Lin08]. Block [Ban07, DKK08, GH90, JKP11, MN82, Sav88a, SZ92, She92, She96, Tan78, ZLA99, LLL20]. Block-Associated [JKP11]. Block-Structured [GH90, ZLA99, LLL20]. Blockages [Kal88]. Blocking [AIH93, AI94, BLL84, Fal96, LC84, Lan86b, LR07, Van91b, Xu98]. Blomqvist [Spe73b]. Blow [LMW98]. Blow-Up [LMW98]. Blue [ML77]. Blue-Green [ML77]. BMAP [DN99, DS04, SGC16]. BMAP/G/1 [SGC16]. BMAP/SM/1 [DN99, DS04]. Bodies [Col69, EE93, Sch19]. Body [EE81, EE88, Kin69b, Sch85]. Bold [CSZ04, CSYZ05, CCHY08, Sch05, Yao07b, YC08]. Bolthausen [FSJ14]. Boltzmann [CG90]. bomber [KSC13]. Bond [AW93, Che01, CLS09, DWW11, KZ84, YW21]. Bond-Decorated [AW93]. Bonds [FNV11, GWS80]. Bonferroni [CS96, CS00, GL94, KS89, MS85, Mór96, Spe94, Wal81]. Bonferroni-Type [CS96, CS00, GL94]. Bookie [BH96]. Books [Nel77, Sim14]. Boolean [AB16, DS17, Sch91a]. Bootstrapping [Sri98]. Borel [Bru80, GZ06, Tas97]. Boris [Kal96]. Borne [Bal90, BC95a, Bec73, DG93a, DD68, GM91, Hen79, Jer74, KS76, Lef78, Lef79, LM88, Rad73b, Rou82, Sau75b, SLK79, Wic75]. Borovkov [Tan97]. Bose [KP23]. Both [DD68, LBL05, O’F89, SG89, Tam79, ZZ04]. Bottle [BG02]. Bottleneck [SD12]. Bound [Ano86a, Ano95e, Cha04, Che03b, Eco00, Eva78, Gal66b, Gol76, Gra85, GW87, Gui94, Hun76a, Jan81, Lou78, Pap94, Pla85, Rhe98, Tan79a, Tho11, Wol77a, AI16, Bis23, Bon13, Bro15, MA14, RY20, YW21, ZZ16]. Boundaries [BWY11, Böh00, Bor94, GS01, Haj82, KZ88, Mor71b, Mor71c, Par85, WP97, Whi73, SWZ18]. Boundary [Ano88h, Bal95, BK64, BN05, BGNR90, C.12, CY65, Dur71a, Dur85, Dur88, DW92, FW10, FW16, Gal93, GJ88, Hyn94, IKP01, JS04b, NFK99, OC03, PP73a, Pol78, PW01, RS87, Sat77, Sch92a, SV19, Smi72, ST86b, Tsi07, WP97, Wil92, Yao93, BDL19, CD13, Gén23, HZ22a, LN19, LB23]. Boundary-Crossing [Dur71a, Sch92a, Smi72, Yao93]. Bounded [Ano64c, Ano82f, BS15d, Bil81, Coh68, Gol75, Har94, Hir16, JK71, Kei64a, LT72, PS99, Pro82, Tra81, Tuf99, CCFM17, KKS15, MB16]. Bounded-hop [Hir16]. Boundedly [CCH06]. Bounding [BK86b, BK87a, HS78, Mil08, GPZ17]. Bounds [Ave85, BMR19, BTH72, BS76, BJV12, BDM21, BCS93, BC01b, BS97, BN05, BK00, Bov99, Bra75, Bro66, BL98, BD94b, Cha75, CDD91, CS96, CV09, Coo74, CT95, DN80b, DIJ04, DJ95, DR09, Dur92, Eks06, Epl83, FS80a, FF03, Fre07, GY95a, Gat84a, Gau22, Gla91, Gol05, GWS80, Has82a, He21, HR88b, Hör03, How00a, HL05, HW79, JNR09, Jia12, Kan85, Kim94a, Kol00, KS89, Kwe75, LHN86, LX00, LL90, Mas86b, Mi02b, Mit04a, Mit04b, Mun12, Mus00, Niu80, Pau01, Pek99, PC85, Pin11, Pol69, Qui76, Ros87a, Ros74, SC02, Sen94, Sen95, Sha81, SZP07, Sud01, Sud05, Tan11, Van89a, Web99, Wei00, WL94, WL97, Wol87, XL98, XS03, Yu74, ZLW06, Bro15, DIÖ23, GN17, HL16b, HLVM22, KW13, PR13, RR16, Ros16, Wan14, WBS16]. Box [Kea67, Nic86, TV04]. Boxes [Col81]. Braess [BKT97a, CSZ97, KLO99]. branch [Yam17]. Branchement [HK94]. Branches [O’B80]. Branching [ADLN82, Agr75, Ale01, AL12, Ano67e, Ano75g, Ano78f, Ano91h, Ano92f, AGSC00, Ath71, APS74, AK76, Ath12, Bag86b, Bag86a, Bar75a, BS79a, BJ79, BZ08, BZ09, BFM08, Ber11, Big77a, Big77b, BG79a, BG87a, Bin88, BW19, BV96, BM01b, BK73a, BH75b, Bru78, CG97b, CG02, CG05, Che97, Che02, CPZL04, CPLZ07, CY84, CS72, Coh77b, CP78, CS80, Coh89b, CS12, CM69, Cru70, CG76, DN80b, DHT86, Dal79b, Dal81b, Dav78, Dek87, Dio74, DY97, Don71, Don72, Don73, Don76a, Don76b, DF04, Dur71b, Dur74, Dur75, Dwa69, DLVZ20, Est76, Fan12, FG99, Fea80, GR89, Gad89, GG91a, GS76, GD74, GL67, Gol78b, GM96, GM97, GMD02, GMD03, GMD05, GMSB10, Gre77a, Gre81, Gre74, Gre76c, Gre77b, Gre78, GZ94, Ham92, Ham95, Hau12, HS66, Hei09, Hey81]. Branching [Hop75, Hop76, HS77, Hul82, Hul84, Hul01, IM10, Iva80, IS85, Jag69, Jag74a, JS00, JZ02, JK04, JL06, Jon99b, Jon04, Kab12, Kap72, Kap74, KSN75, Kap75a, Kap75b, Kap82, KN75, KK69, KS82, Kle83a, Kle84, Kle85, Kle91, KP83, Kyp98, KP08, LML06, LS08a, LR09, Lai77, Lam70, LBC81, LY95, Les73, Lew69, Li00b, Li06a, Li06b, LPLG08, Lin72, Lip77, MM09, MW10, MP00, MM03a, May05, Mei10, MY85, MJP87, MMR02, MMR07, MS79, MS04b, MS10, MS05b, NSS04, Nan79, Ner84, Neu78, O’C93, Olo96, Olo98, Ols77, Pak71, Pak76, PH82, Pak87, Pel07, Pia08, Pia09, Pol69, Pol71, PW07a, Pru85, Qui78, Qui76, QD77, QS99, Rad72a, Rad73a, Rad76a, Raf73, RY99, Rév04, RRS05, Rey70, Rey72c, RVJ68, Sam71, San71, Sau76a, Sch82]. Branching [Sen71, Sen73b, ST83, Sgi01, ST02, Shi06, Slu77, Spa91, SO11, Sri98, SC73, Sud81, Tai95, TW99, Wan76b, Wan80, WF99, Wan99b, Wei66b, Wei66a, Wei69a, Wei70, Wei72, Wei74, Wei75a, Wei76, Wei84, Wei71, WF80, Wil69, Wil88a, Wu93, XW05, Yan72, Yan73, ZJ04, AH13, BBN16, Ber13b, BY22, BP16, BIMR22, Bra22, BH20, CS20b, CLCZ14, CNZ14, CCGW23, Chu18, DHKP18, Eri16, FFK19, GZ16, GvdH18, GMdP^{+}21, GH16, HM14, HL16a, Hon13, Hon16, HK94, HMY17, IK16, IRS23, JL22, Ker20, LPU20, LP21, MZ24, MK15, MK17, OS14, PC16, RW21, Rah17, Rob22, Roy24, Sag23, Sim14, VL15, VD17, WY22, Win15, Zha22]. Branching-Process [Sam71]. branching-stable [BY22]. Branford [vDB87]. Breakage [CCG90, Ols77, SS83b, Tet02]. Breakage-Mechanism [Ols77]. Breakdown [CL11, Fak83, Wu68]. Breakdowns [LSC97]. Breaking [Bit69, HM23]. Breakpoint [ST01]. Breeding [Hol68]. Breiman [DZ07, FM12]. Bridge [EW09, LH07]. bridges [Dai14, GLS19, SWZ18]. Bridging [Har80]. broadcast [SK23]. broadcast-based [SK23]. Broken [Hol80b]. Brokered [Rig99a]. Brownian [ADW73, AS12, AG12, AC94b, BDM21, Bis23, BM18, BW19, BM20, Bre12a, Bre12b, Bre13, CD13, CMY98, CM24, Cor80, DHM20, DG89, Don93, DY98, DF04, DYY01, Dur71a, Dur88, DW92, EW09, ESE84, FI98, Fen93, FKZ13, Föl72, FW10, FW16, FPY04, GLS19, GL99, GP97, GP98a, GN96, GS95, GL05, Had05, HL81b, Hel67, HL99, Hon02, HMBL14, Hör03, HL13b, HK09, Imh78, Imh84, Imh99, Iva10, JP09, KS06, KW78, KS04, Ken76, Kim16, KK24, KMSP12, Lac93, LLZ15, Le94, Lee11a, Lee11b, LH07, Lóp16, Lou84, Luo14, MIAPAM04, Mar11, MS94, MS99, Mat90, Met13, Nád71, NP22, NFK99, Par75, Par85, PS01a, PSZ04, Phe96a, Phe96b, PW01, RRT22, RS06, Sch92a, Sch03, Sly07, Smi72, Wan80, WP97, Was69, Wen86, Wil92, Xu17, YCW15]. Brownian [Yor92, Yor95, Zha04, Zha05, Zhu23]. Brownien [Yor92, Lac93]. Bruijn [KA16]. Bruss [Rog23]. BSDEs [Coh14]. Bubble [BM66]. Bucket [GAS03, GMRY07]. Bucy [AB15, DJ05]. Buffer [AS04, CW96, GAS03, KG02, MS99, O’C97, OS91, PRX03, SM05, SG93, vD89, GS15]. Buffered [O’C98]. Buffers [ABGH00, LM99, Tet98]. Buffon [Dia76]. Built [GSS11]. Bulk [BT73, Dsh98, Dsh01a, GHL71, Nar68, Sel97, SP92, SM97]. bulking [ZX17]. Bunching [Haw70]. Bundle [Sen73a, SBG70]. Burgers [BNL05, Ros87b, SW94]. Burn [BMS93, Cha00, Cha01, Cha03, CS97, Ebr04]. Burn-In [BMS93, Cha00, Cha01, Cha03, CS97, Ebr04]. Burst [BMY02].

Busy [Ano74h, ALH01, BKNZ91, BC92, BPSZ10, Coh67a, Coh71, Con74, CT76, Dal69b, DJ69, DS94a, DS98, DT80, DGM04, Dvu84, DKO84, Enn69, FH10, GT74, Gha86, Gha90, GW05, Had74, Hey74, Ken72b, Lan87, LS96, Lou78, NSS04, Nat75, Neu68a, Pak73c, PRX03, PS01a, Pur75a, RS90, Ros75, SS85a, Sha71, Sta85, SZ93, Wie89, IJB18]. Busy-Period [BC92, DKO84, Lan87, SS85a]. Buy [AM08, YYZ09]. ‘Buy-and-Hold [YYZ09]. Buy/Sell [AM08]. Buying [BF97, SKK06]. Buying-Selling [SKK06].

C [IN86, BB12, Sim98, Gha86, Ses96, ALH01, DO16, FG84, Gha89, Gha90, GW05, Gra83, KT15, LC83, Sim98, ZL15, ZAFP06, vHS86]. c-Like [BB12]. Cône [Let94]. Caches [JRS06]. Cahn [KLL15]. Calculated [Dur86, Gue97]. Calculating [BN02, Bro00b, Con75, JK71, Mas87, Tor79, Wil69]. Calculation [DS04, HK85, LDL95, Twe73]. Calculations [BV11]. Calculus [Bré93, Ken91, Kon96, Mod66, Mod67b]. Calibration [Kru04]. Call [ELTW10, Kol15]. Calls [Fal86, Lee80]. Cambanis [KH96]. Cameron [Yas93]. can [EL19]. Cancer [Sud12, MK15]. Candidate [MC90, RR75a]. Candidates [AHBT82]. Cannibal [Pit87]. Cannings [Möh10]. Cannot [VX04]. Canonical [GJ80, Lar98]. Cantelli [Bru80, Tas97]. Capacities [Ahl73, KS10]. Capacity [Blo91, BG00, Con71, DGM04, GH10, Gil76, GNN85, GG91b, Had69, New86, NR78, PS99, RO99, STC93, Tom72, Weg76, Woo83b, GT18]. Capital [ES11, NPYY18]. Caps [SS86]. Capture [Sta74, War76]. Car [Gre80b]. Caractérisation [Dio78, Dio76]. Carboxyhemoglobin [MC78]. Card [And12]. Carlo [AW15, BG06, BSS81, Bro00b, BNS90, DPR19, DR15, FCvS15, Gor70, GH14, IK11, Kam17, NR22, PS19, PW07b, RR07, RR16, RB91]. Carousel [LA01, PPF03, VAW04]. Carpet [Shi02]. Carrier [Bal90, BC95a, Bec73, DG93a, DD68, GM91, Hen79, Jer74, KS76, Lef78, Lef79, LM88, Rad73b, Rou82, Sau75b, SLK79, Wic75, Boo89]. Carrier-Borne [Bal90, BC95a, Bec73, DG93a, DD68, GM91, Hen79, Jer74, KS76, Lef78, Lef79, LM88, Rad73b, Rou82, Sau75b, SLK79, Wic75]. Carriers [Die66, DD68, Dow67, Lef79, PW67]. Carry [Koç95]. Cars [AR82]. Cascade [GG00]. cascades [Hua16]. Cascading [Swi08, LC16, Lee20]. Case [Ano71e, Boo89, Bri86, CG97a, Dri75, GR74, Grü89, GMW90, Hel71, Hey68a, Kal93, KS93a, Kel97, MW08, MY01, Mod67a, Mur87, Pur71, Smy11, VX04, Voi00, Web99, Whi05, Yeh85, Zuc83, ABI17, HH23, Kev17]. Cases [AHN90, San06]. cash [Jia15]. Casino [GCS07, Yao07a]. Casinos [CSYZ05]. catalytic [BW19]. Catastrophe [BBMB08, CP04, Nie85, PZC07]. Catastrophes [Bro86, Eco04, KA89, Kyr99]. catastrophic [CFRS22]. Catchability [SC75]. caterpillars [BM17]. Causality [PT86]. Cause [Pré94]. Cause-Effect [Pré94]. Caused [Rei67, Mia13]. Cayley [Con22]. Cdfs [VT11]. Cell [Ale01, Bro88, CS72, GWMO79, Gre81, HB99, LM99, MT00, Rat92, SC73, Sud81, Tai95, Wat97, BBBN14]. Cell-Size-Dependent [CS72, SC73, Sud81]. Cells [AAKR05, Cow04, CM69, Ken93, DGLW17]. Cellular [Gra99]. Censored [MZ93, WG93, ZL96]. Center [BM90]. Centers [Bau83]. Central [Bal99, BK12b, Ben98a, Bèr04, BCF09, BCPR11, Bra80, BCMW83, Che03a, Che06, Che76b, CJ98, Cri16, CH71, Eag75, Fat15, Gol05, Han90, Hey71, Hey77, HMPS06, HL99, Hud73, Joh06, Lan84, LPP12, McL76, MP12, Mor73a, Mor79, Pla83, RNJ19, Roo76b, Ros76a, Via94, Web80, WW20, WW00, Xu09, Yam17, ZWV10, Zhu13, Bow18, IK16, Lam14, RY20, Rub23]. Centre [Mut87]. Centred [CJ98]. Century [Bin05]. Certain [Ano73e, AG81, BSS72, BCMW83, Cla99, Dal69a, Dal79a, DP00, Dav74, Don77, Eme72, FP94, Gal72, Gav68, Ger90, GC71b, HN01, Hog86, Hol07, Hol08a, Hol09, Hör03, Hul84, Ito89, Kan85, Kat77a, KN90, Khm11, KM92, Mad75a, Mil92, NH82, Ori79, PS84, Pol69, Pol99, RR88, Roe70c, Sie79, Tal80, Tan80b, Tay99, Vit74, Vit76, Wal98, Wol84, DNY16, Teh13, Vid20]. Certaines [CTKR83, Yor92]. Certainty [Whi86]. Chaînes [CTKR83, Die87, Let78]. Chain [Ano77f, AT87, Bal83b, Ban96, BCY97, BD74, Bro72, Car99, CL85, Cra88, Dav75, FN70, GMM05, Gor75, Grü98, Hel70, Hen72, Hen73a, Jen66, Kat77b, KR70, Kem92, Ken78, KG68, Kha70, Kot94, Kou96, Lam68, Leh71, Leh72, Let74, Lil68, Liu94, MV95, Mei99, Mon82, Neu94a, Nor97, Nue08, O’B74, Ols77, Ols78, Pak78a, PS04, Pur71, QL99, Raj74, RR07, Rob80, Sen82a, Smi92, Suz10, Tan97, Teu72, Ton75, Tor79, TSCD08, TO92, Vas81, Vrb85, Wol77b, YY11, Yun00, ZL96, AM16, AW15, Coh14, Ger23, GH14, Jan21, Kam17, LT23, LB23, PT22, RR16, SP22, dL23]. Chain-Binomial [Bal83b]. Chain-Dependent [Kat77b, O’B74, TO92]. Chain-Transformations [BCY97]. chaining [Leh14]. Chains [AML82, AML84, Agr91, Ano91i, Ano93f, AP86, BCP00, BY93, Ban88, Bas70, BB80, BS97, BRPS82, BHX98, Buc94, BNS90, Cal74, Cao95, CD07, CCFG96, CCdO05, Che91, CDL88, Coh83, DIKM10, Dal69a, DS92, DP00, DS65, DS67, DM67, DM68, DGM08, DP12a, DS80, DKK08, EG02, ELNT11, Eva71, FP05b, FP05a, Fra06, Fra00, GJQ06, Ger83, GKPS06, GMM06, GH90, Gri75a, GHL03, GHL05, HK95, Han92, Har91, Har94, HN01, He03, HJ11, Hol00, HR88b, Hou12, Hsi97, IG01, IT78, IA78, IC82, Jac09, JT01, Kal89, Kal90, KZ88, Kij87, Kij88, Kij92b, Kij93, Kij98, Kim70a, Kol00, Ksi89, KSHH98, Las94c, LT00, Led97, Led05, Lin87, LH06, Liu11, LS79, LPP12, LL90, LZK06a, Mao04, MV95, Mat04, MSCZ78, Mit03, Mit04b]. Chains [Mit05, Nag01, Nak74, Nel77, Nie95, Nue08, OA12, Ōsa85, Paw89, Ped71, Peg80, Pen96a, PS92, Pit77, Pol94, Pol99, Pol01, PS02b, PRHL12, PS01b, RM76, Rey72b, Ris95, RRS98, RT01, Ros81a, Ros81b, RS89b, Rui08, Ryd96, SCZ77, Sch68, SAI82, SVJ66, ST85, Sen01, Sen87, Sév95, Sie86, SV03a, Suo79, Szp85, Tav80a, Teu68a, Tru00, Twe74, Twe75a, Twe83, Twe88, Twe98, VS06, Vit76, Wan92, Wüb12, Wyn73, Yap09, Zei89, ABS15, AFV15, AK23, BS15d, CCdOS15, CCCSdO24, CTKR83, Die87, FHL13, GS14, GT14, Gue13, HL16b, HM17, JS14, KS19, KW14, LR13, Let78, LLL20, MPS21, MZZ16, Mar21, MB16, MV21, MN16, Neu23, PU16, Rub23, dCGG16]. Chamber [BM66]. Chances [AKM10]. Change [AI00, Bra70, BN88, DK01, Dob78, Kle94, MW91, Sie86, Yos84, dSS04, CF13]. Change-Point [DK01, Yos84, dSS04]. Changed [Miy85]. Changeover [DT01]. Changes [Ano65e, Bha65a, De 77, Eva71, GER95, Gor75, FH13, GER95]. Changing [DW77, GG91b, Whi88]. Channel [Ahl73, BMY93, HD71, HGR90, Lou73, TK84, Wol77a, Wol75]. Channels [BMY02, KS72, KA16]. Chaos [Fen97, Kim17]. Chaoticity [Gra00, Gra08]. Character [MP04, Bra14]. Characterisation [Hay00, MO92, Whi73, Whi80c, KL16, McK14]. Characteristic [BH91, Dav64, Gla77, Ken75b, KMRV00, Lau80, Mej65, Rie80, Yao96, Fin13]. Characteristics [ALO05, BS01a, GMS96, Han87, KS80a, KS80b, KS85, MT95, SC81, Wes99]. Characterization [Aal98, Acz72, AR72, Ahs76, Ahs78b, All97, Ano73f, Ano85g, Ano90f, Aut75, BLBL10, Bas70, Bha80, Bin78, Bre75, CW97, CCCSdO24, CHM^{+}22, Dal74, DH85, Deh84, DK90, DS23, Eat66, EZ80, Fei79a, FPS18, Gal72, GH94, GESK95, GS72, Gup84, GG86, Kad74, KPS81, KT83, KH75, LR90, LR92, Lib85, Luk69, Mak94, McK82, Meh67, Nie06, PH87, Phi81, Pur82, Roy89, SCZ77, Sam74, Sen01, Sha74a, Sha77, SA73, SS96, Sre83, SS70, Sri71, Tal80, Tol88, VS91, Voo87, Wan76c, Yao96, ZY70, Zuc81, DD15, Dio76, Dio78, Pyc21]. Characterizations [Ano82g, Arn80, BN06, BL73, Ber77a, Che07, CM98, GKJ80, Hin80b, Hua74a, Hua89, HS93, HLS93, Jun94, KR86, KM03, NB03, Pfe82a, Rao79, Sha72b, Sri86, THB23, WD99, Wit93, Xu98, Zij83]. Characterize [ÇJ73, MC70, MC71, Nag93]. Characterized [BS90, New70]. Characterizing [CP82, Dal87, LO95, PRX03, Teh10, Wes81]. Charge [Gre80b]. Charles [Ano04f, Ano04m, GS08]. Chase [CHJW24]. Chase-escape [CHJW24]. Chebyshev [Tur72]. Chebyshev-Like [Tur72]. Checking [Gou98]. Chemical [Ano68d, DR81, Git74, Hal83a, Hal83b, HG80, McQ67, Orr69]. Chemistry [BPR08]. Chemotherapy [SW88]. Chen [Kim00, Ros16, Vel04]. Chernoff [Big77a, KW14]. chi [Gau22]. chi-square [Gau22]. China [MZ01]. Chinese [SLZ01]. Choice [EF90, Gli00, Gne94, Gne96, GM07, Pet81, Pet83, She92, Tam86, Tam09, Yas84, BS17]. Choices [ASC00]. Choosing [Rog87, AI16]. Choquet [FS97]. Choquet-Deny [FS97]. Chord [Ano93e, Gat82a, Gat87, MS92, Nag93]. Chords [EE88, EE93, Gat82b, Gat84a]. Chris [Nas08]. Christopher [Ano04f, Ano04m, GS08]. Chromosomal [Ano04g, Ste04]. Chromosome [DFSD01, Hol92]. Chronic [Ano67d, Ber65]. Cigarette [MC78]. CIR [DGK00]. Circle [Bin78, CH91, Cre77, Gat82b, Gat84a, Hen86b, Hol80a, HH84, Huf86, HS87b, Huf88b, Hüs82, Hüs86b, Ito89, IMT00, Jan88, Kap78, KPP95, Nea08b, Roy72, Sie78a, SH82, Wak88, Wat82a, YZ82, CM24]. Circles [Aff89, Ors87a, Ros93, Sil78]. Circuit [BLL84, Kal92, Kal93, Sha89a, Zie87]. Circuit-Switched [Zie87]. Circuit-Switching [BLL84, Sha89a]. Circuits [Che03b, Kal89, Kal90, Kal91a]. Circular [Lit79, PR95]. Circularly [Mak10]. clade [FSJ14]. Claim [HPSV04, IK04, ACVZ20, Zha21]. Claims [Bré00, GT06, HT12, IK06, Kar74, OO07, Tan07, Teh10, Ven79, YGN05, DZ13, Rob14, YWK14]. Clarifies [ML85]. Class [And70, Ano65c, Ano66e, BS92a, Bél92, Bie98, Bro79, CW81, CK00a, Can77, Cao95, CCFG96, Che97, Coh94, Dav73, DJ00, DZ07, DP12b, Don72, Dow66, Eva71, FW07, FH07, GT06, GT85, GESK95, Gla78, GMD02, GMM05, GMM06, GBR75, Gra12, Hav91, HMM93, HM95, Höp85a, HLZ05, HL05, HM09, JR82a, Joe96, Kal91a, Kel12, Kla02, KM03, KKRT12, Kra64, Lan81, LK01, Lie77, Lie78, LT90, Meh67, Mur89, Nak74, NRS85a, Peg73, PP71, Pis06, Pla04, Res66, Sav85, Sen95, SSSL10, Sum87, Sve95, Ton97, Tor02, TW83, Wan97, WCW05, Wau74a, Wil99, WZ07, XW05, Zuc81, ACK20, BW23, BHTLV20, BMR16, Fat15, FX16, Fuj13, GM18, HKT24, JY21, LRS09, MB16, NvdBOdI24, NV20]. Classe [Bie98]. Classes [Ano82i, Ano91i, Ano92h, CW91, CF11a, Cha93, Che94, Daw74, FP94, FRR01, Gra08, GP04b, Höp85b, Kei78, Kij87, KM03, MR02, SAS04, Sto04, Wil82a, XZW09, BLM20, CF22b, ISK18, KBM21, Mar16a]. Classical [Ano78h, Che91, Gei99, Kar77, Par01, RF88, Wal81, KS19]. Classification [He03, Kim70a, Pol90]. Clearing [AMR01]. Clinical [WZ65]. Clique [LP94b]. Cliques [JLN10, RS10]. clock [LT13]. Clone [Esh73, SC81]. Clone-Selection [Esh73]. Closed [Ano90g, Bam92, Bau84a, Bau84b, BS98, Cao89, DS02, EV82, GXS12, Joe96, LR80, McK89, Ser90, SY86, Sig89, SD12, Tet98, Wan77a]. Closed-Form [Ser90]. closing [EL13]. Closure [CT04, FRR01, HMM93, HX02, Les89, MO92, NJS96a, SY91, WCW05]. Clouds [Fox65, Ray70]. CLTs [Rob99]. Clump [MC11]. Clumping [Sud77, Wil77]. Clumps [Kel83]. Cluster [AT79, Bau81, Bau83, Che01, HO74, JZG11, KS83b, Las78, MM10, Tha81, Wes73, Wes77b, Yun00, BMZ23, Ber23, MC18]. Clustered [Amm83, SF83]. Clustering [ARS21, BM09, Boo92, BDLL08, CCGP91, Gla81, KK69, Sap75, SF77, Set03, Str77, VE98, WC95, Che18, DS21]. Clusters [CS08, Les69, Oga01, Ste99, BMZ23, Ber13a]. Co [CJ08b, Jor18]. Co-Adapted [CJ08b]. co-existing [Jor18]. coadapted [Con13]. Coagulation [GK12]. Coalescence [Ath12, CPS96, CZ98, DP12a, Hon13, Hon16, Le14, AH13, GH16]. Coalescent [JS04a, KKL01, KK03, Möh98, Möh99, Not00, Not01, Pol10, Sag99, Sag03, Sam06, Tri92, AD13, FSJ14, KJS15, KSN16, LM15, Roy14, SP22]. coalescent-based [Roy14]. Coalescents [GIM08, GIM11]. Coalescing [Ana88, Ngu90, LTZ18]. cocktail [Möh21]. Codes [Ahl73, Par01, Sut89]. Coding [AS75, CT00, Dek96, IS86, KS72, Rat70]. Coefficient [Doh87, FZ93, LPS06, LS04, Nic86, Tan82, Vrb85, DS21, He21, Sön23].

Coefficients [AN95, Bro79, Dur86, GY95b, GG77b, Hor01, Kul06, PS92, SH03, Tan83a, KY17c]. Coexistence [BB14, HK22]. Coffman [VE98]. Cohen [BS01b, Lan79a]. Coherent [Bax83, BK12b, BK00, CDD91, ENPS78, GSS11, GAB12, JNR09, Men96, NSB10, Now90, SN97, Zha10, ASZB18, ASKK20, AZA18, BB18, BN13, Ery15, ET21, FZL13, Gup13, HF19, KA22, LL24, LSW19, Nat14, NFS20, PB14, PB20, TSN17, YZZ21, YBL23]. Coiflets [JS04b]. Coin [DW72, PP88, FHI17, HK15, Hu16]. Coin-Tossing [DW72]. Coincidences [IKS70]. Coincident [Kar88]. Coinciding [KMS83]. Collapse [Kel09, LS11]. Collapsing [Dal79b, GvdH18]. Collecting [Pós09]. Collection [AR01a, GT08, MC11, Smy11, BR16]. Collective [Igl69, MZLSRS02]. Collector [AOR03, Nea08a, SC75, XT11, ABS15, ABSGS16, BT23, Sau23, SH21, Yu17, KS08]. Colliding [Kar76b]. Collingwood [Gan71]. Collision [CPZL04, Ito79, Nis74a, Nis74b, Nis75, PVS88, Wei89, CLCZ14]. Collisions [GIM08, Har65, Nak08]. color [BJT20, KS17]. Colored [DNW07, Let78]. Colorées [Let78]. Coloring [YY11]. colour [LMS18, Thö16]. Coloured [Str77]. Columns [IM74]. Combination [Llo71]. Combinations [Huf88a, XH11]. Combinatorial [BJM93, Böh00, Gol05, MNS01, Res66, Tak64a, EDSM15]. Combinatorics [Kar88]. Come [Fak86, Fak87, FR10, VB02]. coming [LP21, SF17, Sag23]. Commande [Lef89]. Comment [Boy87, Gre86, Hin95, Hog86, Pra82]. Comments [Bal78, Fei79b, Hor80, IN86, Kem94, Tan80a, Ton73, TP79, vDB87]. Common [Ale96, CS75, CS88, Fea02, GR98, Li00a, LS86, WR97, HK19]. Communicating [BT12b]. Communication [AIH93, CL98, GL76, Sen87, Hir16]. Communications [AI94, Top86]. Communications-Type [AI94]. Communities [Gan94]. Community [Bec73, HWX18]. Commuting [Hud73]. Comonotonicity [Che07]. Compact [Ber68, RS89a, Rui08, Wil92, AM18]. Comparability [AHO20, DR84]. Comparative [Hic86, KSS16]. Compare [Ebr91, MA17]. Comparing [GR98, HF19, LU96]. Comparison [BPSS14, Bon89, Cam86, CSC13, CL97, Cla94, Day83, Den85, DM72, DG92, GK97, HD80, Hai90, Ino19, Lam84, LS98, Les86, LS97b, Mac01, MI83, MDG09, Miy90, Niu81, Pel07, PP70, Rig11, Rob91, RU18, Rüs82, Rüs04, Rüs08, Sco82, Sha87, Sim98, Tak71b, UB74, VX04, WR97, BMRRS16, GR21, KR21]. Comparisons [Ano02e, BOR01, BAAB00, Bla53, BLS90, FL97, GG10, Hu96, KK00, KX09, Lan88, LM88, PS02a, PP73b, SW97a, Sri91, Tru00, AMV21, ASZB18, ASKK20, BSC18, BBH18, DFP13, Gup13, LWM20, SP17]. Compartment [Boo92]. Compartmental [DW89, LP85, Sap87]. Compensator [KS91b]. Competing [AM05, AIGR06, CK05, DG06a, HP98a, Hal83a, HPSV04, LV12b, CC15]. Competing-Species [CK05]. Competition [Dig76, Guo02, RR78, LP21, MZ24, SV19]. Competitive [BL98, EF90, Pak83]. Complementary [Miy90]. Complements [Whi72a].

Complete [AS96, Ano89e, DD04, Gat84a, Hea67, KS10, Pla04, PWW12, SW05, Con13]. Completely [LL78, LW19, Loe09, LL90]. Completion [Git81]. Complex [Bel66, Bre81, FS80b, Hin80a, Hin82b, SK91]. Complexity [RR16, RRT22]. Complicated [Nak86c]. Component [Ano83f, Ano87e, Ery15, GIM11, JK87, MDG09, Nat82, Pin80, Rev93, SS97, BB18, De 22, GN17, HE16, LL24, PB20, Ros14b]. Components [Ano90h, Bax83, BP01, Bor95, Bur09, CDD91, CS97, CS88, Ebr08, ENPS86, Gås98, GR98, GAB12, Kup90, LM89, LD10, Lin89, LS86, Men96, NSB10, NSB11, Nor86, Pie07, Qui86, ZLB08, ASB21, AA17, AZA18, BBH14, BBH18, BHB18, DFST15, Ery15, FCP18, FHC18, HF19, JKL^{+}20, KA22, NFS20, PB20, Sei16]. Composable [Sch70]. Composite [Kim94b, KP87, Les67, Ste99]. Composition [FG87, Mac01]. Compositions [Kap72]. Compound [AH00, BKL03, BMZ07, CG99, CV01, Coh89a, CT90, Dal10, Dal16b, Del78, DMZ15, Eco04, Erh00, FS98, FC02, Gal93, GGS^{+}95, HJ02, Hsi97, IA07, Kal75, KS01b, LA98, Lin96, LS02a, McN68, Min78, Möh05, Møl96, Nea05, Pek06, PS01a, PSZ04, Sah83, Ser84, Tan05, VC96, VU09, WCW05, WL94, WL97, Wil02, WC04, ZW02, ZY08, BM14, Ges79]. Compounded [BT92]. Computable [ADPR18, HL16b, SZP07]. Computation [FT80, GM11]. Computational [CL93, KJS15, RN80, SGC16]. Computations [LBC81]. Compute [MK73]. Computer [Neu88]. Computers [Lew64]. Computing [Asm09, BS01a, BS13b, BP99b, Dur71a, ET21, Hen86a, RWN02, RS94, SV03b, VT11, Wan76a, GH14, LB23]. Concave [Reh77, Yu11, Leh21, YCW15]. Concavity [Ano10e, Kei71, Kel92, KS94b, Mar07, Sze86, AA23, GK22]. Concentrated [Fra00, Mil92]. Concentration [Rob90, Rub64]. Conception [LJ78]. Concepts [BS04a, Hic86, GK22]. concerning [Amb65, Dom96, Eng65, Fei79b, Hen73a, Hop76, Huf88b, Kal93]. Concomitant [KSC12]. Concomitants [DG74]. Condensation [Kni71, KP23]. Condition [Agr91, Che81b, CD05, Don77, Don80b, Fer93, Gla82a, GZ06, Har67, Har80, Hos09, Hul82, LM04a, NMG00, Olo98, Pur82, Tal80, TW75, Whi79, CC15, LR14a, SUY21]. condition-based [CC15]. Conditional [Aas77, Ano82i, AF83, BLSGS13, Ber77a, BG05, CL05, Cam97, CG02, CK00b, Che07, DS02, DS94a, Don84b, FPW19, FMS99, Fin13, FKZ13, FWL12, GSV11, GST00, Gre77a, GM15, Jam66, Jan01, Kij88, KH75, Lin99, Miz06, OC03, Pak73a, Pak10, Pek99, PW07a, Ser72a, Ser72b, SS88a, SS93, SK90, Sko87, SAS04, Sum84, Sum87, Wan93, Wei75a, Whi80b, Whi82a, Whi85b, Wil82a, XL98, YA88, Zha10, ZLB08, FZL13, KR18, KLP19, KS20, NR22, PB14, WGF13, YML17]. Conditionally [BG05, CJ98, CS97, Dzi86, Gås98, IPR11, Liu04, dSS04]. Conditioned [CC06, Day83, Don83, Est76, FG98, HJ11, Hoo83, Igl74, JR95, Jac09, Kao78, Ken74, Ken75a, Rob91, RJ94, RJP97, SV97, SK90, ABD17, LW19]. Conditioning [DH08]. Conditions [Aba80, Ano71e, BE73, CCFG96, CW86, CD98, DMS08, DNY16, FHT78, FT78, GT06, GR98, Ger83, Ger76, HK95, Hey68a, Hul84, Hyn94, JI88, Kle88, KS01c, KS91b, Leh72, Leh73, Lew69, LT95, MSCZ78, Mor73c, NMG00, Reh79, Rob80, SKF79, Sen87, Szp85, Was69, Whi80c, Wie80, BK21, BCK16, BMRRS16, DJ18, GW17, Jas15, MPS21, MW21, RY20, Zha14]. Conductivity [Ham88]. Confidence [GRD^{+}00, BVW21, BS15c]. Configuration [Ito89]. Configurations [Gir82, Kel85, Les10]. confirmation [CLZ24]. Confusum [Ano81f, Ren80]. Congested [Pre02, Wal11]. Congestion [CK90, Gav68, McN70, McN73, OS07]. Congruential [Law92]. Conjecture [ANW82, BS70, BS82a, ELNT11, Hor79, LU95, PT80, Tan79b, ABSKT19, CLZ24, NR18, Rob14]. Conjectures [Ano86h, Hai83, Hey82b]. Conjunction [AARJ10]. Connected [Ban88, BGJ12, Elf67, FS80a, HP17, HPP14, Wat23]. Connection [AGSC00, FW07, KS82, LT81, DT15, FHIM14, Pri19]. Connections [AES17, DSS11]. Connective [Law83]. Connectivity [Dor82, Sha23]. Consecutive [CM92, CP90, FKM10, Hol08b, Hol09, Mus00, NE07, PH88, PR95, Tav81, WJ95, WCC03]. Consecutive- [CP90, Mus00, NE07, PH88, PR95]. Consensus [CS77, HP17, SK23]. Consequences [JKM12]. Conservation [Aye07, BG87b, Bré91, FL91, MTV99, MKR91, Miy85, Miy94, YSS91]. Conservative [RS69, Sau75a, Sau76c, Rub23]. Conserving [Miy01, Wol70]. Considered [CG76]. Consistency [DS80, RS95a, Sha13, Shi86, Sol84, Hen18, JP16]. Consistent [Cra11, Kol82, Loc82, BPS15, FPY20, HS22]. Consistently [NTaYY04, WW07]. Consisting [Bec73]. Constant [Ano73e, BSS72, BH12, CZ99, CK81, De 85b, ES11, Fro05, Hal79a, Ham95, Har90a, KM01, KNT10, Law83, Lin04, Lu10, Sch96, Tan05, XS03, Ahl20, BHL^{+}18, HJ13b, WY22, CW69]. Constant-Elasticity-Of-Variance [BH12]. Constants [Bar72a, Tet98, DHM22]. Constrained [Föl72, HW88, Ken82b, Sah78, KY17b]. Constraint [CNO^{+}84, Cow93, LR09, RT91, Sta04, Dal16b, HBHH23, HM23]. Constraints [CPT99, KR80, RO99, Owa23]. Constructed [Sha79, Sut89]. Construction [Ano86a, Gra85, Ish77, JL22, JGR91, Ma04, MM71, Moy10, NH82, Par01, Whi90, AD13, HS14]. Constructive [DGNR95, GNR89]. Consulting [Rus01]. Consumers [Rig99a]. Consumption [BMP87, CCFM17, Mos18]. Contact [And03, Ano93e, Ber83, DS17, Lan05, Lan11, MSZ94, MS92, Ren79, Sud01, Sud05, Tzi11, ZG88, Ber23, Fox15, Fox16, FL20, Tzi16]. Contacts [DDM98]. Contagion [MO93]. Contagious [Cha83, RW21]. Containment [SO11, LR14a, SBM22]. Contaminated [CH97]. Contamination [GWMO79]. Content [GAS03, LM04b, Mas95, Tak67]. Context [San71]. Context-Free [San71]. Contingent [OO07]. continuation [But19]. Continued [GP98b, GP99, Han92]. Continuity [BG79a, CS80, CK81, GMSB10, Jon99b, Ken72a, Sar97, Sha88, Zha96, dCGG16, Cha17, DHM22, LS17, Zha14]. Continuous [AI89, And12, Ano88h, AB86b, BDM12, Bal95, BY93, Bat66, BKMS04, BFM08, BMP87, BLSW08, BLSW10, BP96a, Bro04a, Bro00b, BS09, Car99, Che91, Che97, CV09, CT81, CD06, CD09, DP00, DS67, DM67, DM68, Dav85, De 85c, Dur85, Fel78, FP05b, FWL12, GT85, Gao04, GG87, Ger90, Ger83, GC71b, Gre76a, Gre74, GZ02, GHL03, GHL05, Hal70, HC07, Her74, Hic84, HJ11, HLZ05, Hyn94, IK04, Irl03, IA78, JZG11, Kap10, KZ88, Ken92, Kij88, Kij93, Kij98, KT01a, KLM04, KPT99, Kro97, KP08, KLP12, LS04, LMS76, Li06a, LBL05, Lin82, Loc96, Ma03a, Mac01, Mao04, McC76, McC78, MS04a, Meh77a, Meh79, MT70, Mit03, Mit04b, Mor67, Mor69, MM03b, Pak73a, Pak78b, Pen96a, Pir80, Pol94, Pol99, PS02b, PRHL12, PS01b, Pru83, Ram81]. Continuous [Rit80, RT00, Ros93, Ryd96, Sah83, SAI82, Sha75, Sob73, Sta91, Ste03, Tan07, Tsa96, Tsa99, VS06, XY09, Yap09, YT83, Zei89, ZPR08, Zuc83, Zuc86, ADPR18, BVW21, BS16a, BEER15, BPS15, BP16, DS21, FFK19, GvdH18, GMdP^{+}21, GPZ17, GH21, HM14, HL16a, LL13, LP21, MZ24, MPS21, MK17, MV21, Pim14, WGF13, YML17, Zha14]. Continuous-Review [Ram81, Sah83]. Continuous-State [BFM08, KP08, Li06a, BP16, FFK19, HL16a, LP21, MZ24]. Continuous-State-Dependent [XY09]. Continuous-Time [BDM12, BY93, BKMS04, Bro04a, BS09, Car99, CD06, DP00, DS67, DM67, DM68, Dav85, FP05b, GT85, Gao04, GZ02, GHL03, GHL05, HC07, HJ11, HLZ05, Hyn94, Irl03, IA78, Kap10, Kij98, KT01a, KLM04, KPT99, Lin82, Ma03a, Mac01, Mao04, McC76, McC78, MS04a, Mit03, Mit04b, MM03b, Pol94, PS02b, PRHL12, PS01b, Rit80, Ryd96, SAI82, Sta91, Ste03, Tan07, VS06, Yap09, YT83, Zei89, ZPR08, BVW21, BS16a, BEER15, GvdH18, GPZ17, GH21, HM14, MV21, Zha14]. Continuous/Discrete [Sha75]. continuously [MC13]. Continuum [Bax84, BK86b, BK87a, FM06, JZG11, KB87, Mee95, Ort08, Sar97, Tan93, YCG11, DP18]. Continuum-Time [Ort08]. Contracting [JT01]. Contractive [Ben98a, ABI17, DN19]. Contracts [DJ05, RZ08]. Contrast [Lán79b]. Contributions [Jam66, LMS72, MMR07]. Control [Aas87, AHN90, AH00, Abr07, Agr91, AT74, AB82, Att87, Att89, BC67, Bat68, CL11, Cos93, CD09, Cow78, Dav75, Dav82a, De 85a, DJ09, Dos81, Dsh89, ES11, Fad74, FS71, Ger76, GM73, GMD02, GMD03, GMD05, HW84a, HS04, Jac99, KLP12, KA89, Kyr99, Leh72, Les06, Lew01, LAF02, LR04, MW96, MPHL10, MP04, Mor71b, Mor71c, PRHL12, Sel97, Sha80, VGT90, VAK92, Wak91, Wan76a, WR74, Wic75, Ye07, YT83, Yeh85, YH87, ZY70, ZRU68, Zuc81, Des15, Gag17, Kan23, Lef89, OT16, PS19, SL21, YCW15, ZC14]. Control-Limit [Ger76]. Controllable [RS90]. Controlled [Bag86a, BB80, CCFG96, CCdO05, CI78, DW77, DS80, GT08, GMD02, GMD03, GMD05, GMdP^{+}21, HP98b, JFHL09, KO05, Kol82, KPW94, KSHH98, Lal86, LP00, Lew01, PRHL12, Via94]. Controller [Tit74, CCCSdO24]. Controlling [McN85]. Controls [LT95]. Convergence [AB18, ADY07, And75, AS96, ABD01, Ank06, Ano83a, AT96, BCP00, Bag86a, BD90, BHZ22, Bar88a, BB84, Bél92, BCF09, BL01, Big77b, BP01, BGSS03, BNT02, BRPS82, BS79c, BR95, CCFG96, CS77, CG90, Che96, CC01, CK00c, Coh77b, Coh81b, CMS06, CSV01, CD98, CS82, DMS08, Dav78, Dav86, DM01, DS86a, Del98, DR00, Dur08b, FRT99, Gab16a, GD74, Gne98, GM97, Gor75, Gou97, Gre77b, Gre01, Gut75, GKL85, Hal79b, HTV92, HN01, Hea67, Hey70, Hey77, Hol00, HKQ88, HLZ05, Hua95, IK70, IM10, Iva81, JR95, Jag74a, JS04a, Jan01, JOC13, JW90, JT92, Kal92, KF70, KMR75, Ken75b, Khm11, Kol00, KSN16, KR97, Lan81, Lie77, Lie78, Lin95, Lin97, Lin72, Lin73, Loc96, Lóp16, Loy88, LS97b, Lun97, Mac01, MSS09, Mit03, Mit04a]. Convergence [Mit05, Möh99, Möh05, Mur80, Nei01, Ort08, PS86a, PRHL12, PS01b, Pur87, RT90, Rad76a, RV00, RJ94, RRS98, RT00, RT01, Rod95a, RR98, RS15, Sad01, Sag03, San80, Sar97, Sau76b, SV97, Sen81, SS96, Sim97, SP99, SD12, Tan97, Tom72, Tor02, Tro83, Unk79, VT89, VT75, VT76, Voi00, Wan80, WCC03, Whi71a, Whi71b, Win79, Zei91, ZNF10, AFV15, AW15, BJT20, BJ24, Che18, DI20, FPR20, GPW22, GS19, HZ22a, HK19, IK18, Jan21, Jan20, Joh24, Kim17, MN16, NP22, RY20, ST16]. Convergent [Dip98, Pap05, Men23]. Converging [Fin10]. Converse [Hey68b]. conversion [Man13a, Mar15]. Convex [AFGP91, BTH72, Bla92, BdH87, CH91, CV09, Col69, DM10a, DP10, Dwy88, Edd80, EE81, EE78, EE88, EE93, EV82, Gol75, Gro80, Ham05, Hen83c, Jan78, JR82b, Kin65, Kin69b, KX09, Mak94, Mas95, Mat90, MN79, MY91, Mol86, Nag93, Rog78, Sca98, Sch85, SS87, Van89b, Wei89, AHO20, ESC15, FV18, MO23, PL17, Ray70, Sch19]. Convex-Ordered [CV09]. Convexe [Ray70]. Convexifications [DZ03]. Convexity [AL00, Gra83, HZ87, Har90a, Kei71, LC83, MP92a, SY91, SG88, WW02]. Conveyor [Naw83]. Convolution [CT04, HX02, Joe96, Les89, MSS92, Pak04, Pak07, Wal76, WCW05, XB11, BMR19, BLS15, RNJ16, RNJ17, SRN21]. Convolution-Closed [Joe96]. Convolutions [FKZ09, WC04]. Cooling [Kol95]. Cooperation [EW88, CP17]. Cooperative [KLO99, LG81]. Coordinates [Wie80]. Coordination [CCH06]. Copula [KR08, AA23, DFST15, FPW19]. Copulas [Bal94, HJL14]. Core [Wes82, DDS20]. corner [FHIM14]. Corollaries [Pak07]. Corpuscle [Zäh90]. correct [GP13]. Corrected [Sie79, YML17, Hog86]. Correcting [Ano87i, Par01, VJO84]. Correction [Ano64c, Ano65c, Ano65d, Ano65e, Ano66d, Ano66c, Ano66e, Ano67e, Ano67d, Ano70d, Ano70e, Ano72e, Ano72f, Ano72g, Ano73g, Ano73f, Ano73e, Ano74e, Ano75e, Ano77e, Ano77g, Ano77f, Ano78e, Ano78f, Ano78h, Ano78g, Ano80e, Ano81e, Ano81f, Ano82f, Ano82g, Ano82h, Ano82i, Ano83f, Ano84e, Ano85g, Ano85e, Ano85f, Ano86g, Ano86i, Ano86h, Ano86j, Ano87g, Ano87e, Ano87f, Ano88h, Ano88g, Ano88f, Ano89e, Ano89f, Ano90f, Ano90g, Ano90h, Ano91e, Ano91f, Ano91g, Ano92f, Ano92g, Ano92h, Ano93e, Ano93f, Ano93g, Ano93h, Ano94f, Ano95e, Ano98e, Ano98f, Ano02e, Ano04g, Ano05e, Ano06e, Ano06f, Ano06g, Ano07e, Ano10e, BS15b, BS16b, BL15, Dow72b, Tsu18, Zho08]. Corrections [Ano68d, Ano71e, Ano74f, Ano74h, Ano74g, Ano75f, Ano75g, Ano79e, Ano79f, Ano87h, Ano87i, Ano91h, Ano91i, Dio78, Gre78, Hör05, MC71, Mod67b, Pak07, Spo01]. Correlated [All04a, All04b, AM08, Ano87h, BGM15, BH98, Bie98, Böh00, BC92, CR92, CR94b, CH69, Gan88b, GS06, HRF83, HRF84, Ios86, Iso82, KK01, LB89, Lan86a, Lan87, Lat81a, LX00, ML67, NS80, RH81a, Roe85, RS06, Rou90, SS83a, Sha75, SM97, SDH94, Zha92]. Correlation [Cra69, DS08a, DKM93, ESM08, Fin88, GM11, Lee80, Ma02, Ma03a, Mar69, McK81, PL72, RH81b, RS06, SSV72, SP92, WK65, WPN80, FS14, MC18]. Correlations [CM69, Gne00, Ma03b, Smi82, HM15, LTZ18, PKO16, Ste19]. Corrélées [Bie98]. Correspondence [CHPT97]. Corresponding [Bel71, Wal81]. Corrigendum [GJP^{+}21]. Cosine [Dav01]. Cost [AHN90, AH00, All04b, Att87, BBC01, BKL03, BP04, CL97, FS71, FT78, GJ72, GBW92, HST75, Jer70, Jer74, JCM98, KS76, Ksi89, LR09, Las94b, LM04b, Mil81, Mor71c, Pra82, Ros70a, Sav88a, Sch98, She96, WP80, Yeh85, CCCSdO24, GH21, Lin14, YCW15]. Costa [AA23]. Costly [Bal75, Mon82, BEG22]. Costs [CCFG96, Gla80, IKP01, KPT99, KR00, Rob80]. Count [AT79, KS01c, Nue10]. Countable [CD07, Cox77, JR95, Mil81, Kun22]. Countable-State [Mil81]. Counter [DO85, KS93a, Tak76]. Counterexample [ANW82, Ano91f, Gue91, Wei82]. Counterexamples [Ano86h, Hai83, Hai92]. Counterpart [Bru80, BKK19]. Counterparts [Kel09]. Counters [Ber81b, Pha74, Pha75]. Counting [Ada75, And12, Bad11, BM05, Che94, Ebr93, Kel85, KS86c, McD79, OT15, Pel94, RJ65, Tak64b, VSW09, WW06a, WY93, Wau74b, BS15a, BGM15, DM19, KSS16, PLRB23]. Counts [AR82, BL09a, Hol07, HS12, RR10, GK18, Kur22, Pri19]. Coupled [Fen93, Tzi16]. Coupling [CR17, CJ08b, Gri75a, GKL85, HKQ88, Ken83, LS98, Lin79, Lin82, Lin83, LR96, Mao02, Mei99, Mou93, Nea07, PR08, RR07, Sim14, Bis23, Con13, CM24, DJ18]. Couplings [GY04, GR96, LS02b, MO21, OC22]. Coupon [AR01a, AOR03, DWW11, Nea08a, Pós09, SC75, Smy11, XT11, ABS15, ABSGS16, BT23, BR16, Sau23, SH21, Yu17]. Coupon-Collector [AOR03, SC75]. Courant [vD15]. Courtois [VE98]. Cousins [MW08]. Covariance [BD75, Ber81b, Bha90, Dai04, DL96, Kei82, Rey72b, FPKM22, Psa22, UY23]. Covariances [Cab91, Dal69a, TV88, VG85, VT89, AH15]. Covariate [Ebr01, Mye81, PS02a]. Covariates [Ebr04]. Covariation [Wil82b]. Cover [Ano17b, Ano17a, Ano18a, Ano18c, Ano18b, Ano19b, Ano19a, Ano19d, Ano19c, Ano19f, Ano19e, Ano19h, Ano19g, Ano20b, Ano20a, Ano20d, Ano20c, Ano20f, Ano20e, Ano21b, Ano21a, Ano21d, Ano21c, Ano21f, Ano21e, Ano21h, Ano21g, Ano22c, Ano22b, Ano22e, Ano22d, Ano22g, Ano22f, Ano22i, Ano22h, Kap78, Leh14, Rob22]. Coverage [Ail66, Bri98, BS93, Coo74, Gre80a, HH84, HKQ88, Huf86, Hüs82, Hüs88, JR82b, Sch80a, Sie78b, YZ82, BFP17, LPS17]. Covered [Sta81]. Covering [Hol80a, HS87b, Lit79, SH82, Att19]. Coverings [Jan88, Kap78]. Cowan [Ano88f, Bru87, Hai03, Liu94]. Cox [Ano86i, Bar86, Bhu06, DJ05, Koo04, Sch96, TV04, Yan88b, Zhu14]. Cramér [BSX19, DM21, FS23, God77, TTI22]. Credit [CJ06, CLS09, DWW11, ZY08, DM21, Fin13]. Criteria [ABD01, Bru84, FI87, Fra06, Gne94, HL04, IT78, Rob99, Twe81, Kun22, She13]. Criterion [CCdO05, GZ02, HD80, Han82, Ksi89, MS02, PP70, Ros80a, Sch82, Ton75, Wan97, CCdOS15, LRS09]. Critical [Ano78f, Ano91h, Ano92f, Ano92g, Ath12, Bar74, BZ08, BZ09, Ber11, BS97, BV96, Bri98, BCD93, DL06, Dur71b, Dur92, Est76, Fea76, Fea80, FMS04, Gol78b, GMD05, GHW95, GLS00, GWS80, Ham82, Hey70, Hey71, Hoo13, Höp82, IS85, JK04, JRS06, JW11, KS86a, Kle89, Kle91, Lin72, MW10, ML98, Mee95, MN79, MY84, MSZ04, MMR07, MSZ94, Mur80, NW07, Ner84, Pak87, Qui78, Rad72a, Sag23, Sau76a, Spa91, Sud01, The79, Tok08, Wei70, Wei72, Wei74, Wei75a, Wei76, Wei84, Wu93, WDLC09, ZG88, Ber13b, De 22, DP18, Hon16, VD17]. Criticality [Nan79, PH82]. Criticism [Tan79b]. Critique [PT80]. Crofton [Mat76, RR73]. Crop [TG87]. Cross [Rob86, Vrb85, BRRN16, DS23]. Cross-Association [Vrb85]. Cross-Sections [Rob86]. cross-variograms [DS23]. Crossing [Ano73e, BSS72, Bor95, BN05, Bri79, BGNR90, Dur71a, Ebr99, FP05b, FP05a, FW10, GJ88, Kat02, Nád71, NFK99, Nyr94, PS76a, PF06, PW01, Ryc87, Sch92a, Smi72, WP97, Yao93, CD13, DL17, FW16, LB23]. Crossings [BD12, DR94a, Roe70c, RC80, Sel91, Sha80, Sha79]. Crudely [Ser92]. Crump [OS14, Sag20]. CS [Gol23]. Cube [DH89, Har67]. Cubes [Coo89, Lan11]. cubic [YW21]. Culture [GWMO79, Man74]. Cumulant [Ano05e, MF01, LR23]. Cumulative [And88, DM05a, Ebr99, FY83, Sah78, Sta87b, Ste78, Tro83, Bon13, DT15, TSN17]. Cumulus [Fox65]. Cups [Kla94]. Curious [Fei79b, Pic77]. Curvature [Bal95]. Curve [Lie77]. Curved [Dur88, DW92, AK23]. Curves [DZ03, EV82, GHK04, Mor66b]. Customer [BPSZ10, Bru71, CPS96, CZ98, Dal65, FG87, Gel93, KS80b, KMS83, MW71, Min79, Miy77, PR08, Ros73, VB02]. Customer-Stationary [KS80b, KMS83]. Customers [AN84, Bal72, BE91, CP92, Coh67a, Con94, DT85, Enn69, Gel91, Gil76, GW87, Gre89, GP04b, HP93, Hen93, Kei78, Kel75, Lat81b, LR80, LH02, Nat74, NR84, PRS08, PS99, Van81, Web78, ZZ04]. CUSUM [Had05]. Cut [DF04, KB91, Lac05, Lit69, Lit79]. Cut-off [DF04, KB91, Lac05]. Cutoff [KS01a, MZZ16]. cutoffs [CSC13]. Cuts [Lit74]. Cycle [Ale01, Cow81, McN68, MY01, PS01a, Sis70, Wol02]. Cycles [Hol78, Tak88]. Cyclic [APS04, BG87b, Chi68, DMS08, Jac78, Mah03, SD12, Vas84, VW86]. Cyclic-Service [BG87b]. Cyclically [Pod72]. Cyclostationary [KP93]. Cylinders [Mol86]. Czechoslovak [ZRU68].

D [BG00, Dvu84, FY83, GBG02, Pak70, Reu82, CP01, DC77, Dsh98, KS71, KT15, Pak70, Pou90]. D-Policy [CP01, Dsh98]. D. [Lin16, Wei79]. D/G/1 [Pak70]. D/G/K [Pou90]. Dagum [FPKM22]. Daley [Pit90]. Daley-Kendall [Pit90]. Dam [AHN90, AH00, Abr07, AB82, Att87, Att89, BKL03, Bat68, Blo73, BC75, Her72, KP89, LA98, LBL05, Lun94, Pha69, Pha79, Roe70a, Roe70b, Sri74, Sta93, TT79, Tur72, Yeh85, YH87, Yeo74, Yeo75, Zuc77, Zuc81]. Damage [AH84, Arn75, BSW07, Ebr99]. Damaged [AHS78a]. Damped [DM10b, Els96, KO05]. Dams [AK96, Bal78, Bri79, BPS80, ÇP72, Col74, Col75, Fad74, KS01b, Ken78, KG68, Kha70, Kha77, LS97b, Mor67, Mor69, Pak73b, PP77, PM73, Pha93, RC80, Tak64a, Tak67, TP74, TP76, HPP14]. dans [PL03, Ray70]. Darboux [GdlI18]. Darwin [Ald95]. Data [Ana88, Asm09, Ban07, Bin78, Bre69, Bro00b, Car99, DS92, Hen86a, Hou91, JKM12, Jon86, Lee80, LG81, LSC04, OK86, RS95a, RMV97, ST01, Sen04, Smi82, Str75a, SW94, VJ82, CH13, JKL^{+}20]. David [Ano01k, DVJ08, Phi01]. day [Nea16]. Dead [DO85]. Deadlines [WB03, Ino19]. Death [Adk69, Ano65d, Ano66c, Ano74f, Ano74e, Ano83a, BS01a, Bil81, BG79b, Bra86, Bro86, CP04, CL03, CP03, Cof95, CT81, CSV01, Cor87, Dav65, DSR80, Dei10, DW77, Di 98, Don84b, FE87, FDVF93, GY65, GR94, Gri72, Gri73, GP99, Hai78, Haj82, HN01, Iva81, KS86a, KT82, Kei64b, Kei65, Kei71, KS91a, Kij92a, LR93, LM90, LPSV00, Lin79, Liu11, LS74, MV95, Mit04a, Mor74, MM03b, Nat75, Nat77, Phe96a, Phe96b, PZC07, Pur68, Pur75b, RD97, RJ94, RJP97, Sau76b, SV97, SZP07, Sum84, Sum87, Tor79, VSW09, Wau72, Zei91, Zei95, vDZ05, BBN16, BC23, Cur22, HZ22a, HCY21, JF23, KS18, SF17, Sag23, ZZ13, vD15, Cla12]. Debit [ZY08]. Debt [Guo02]. Debugging [Sie85]. Decay [Cal74, Kij92a, SZP07, Zei98, CNZ14, CCGW23, vD15]. decaying [DN13]. Deceits [MZLSRS02]. Decidability [Mad75b]. Decision [AR05a, Ass83, BR87b, CRS97, CST08, Dos79, Fas90, FT78, Gla78, GBW92, GZ02, GZ06, HST75, JC06a, Kur87, Las94b, Las94a, MT70, Rob80, Ros82, Ros70a, Sob82, TH11, YT83, ZPR08, ADPR18, CCdOS15, CCCSdO24, FVZ19, GPZ17, GH21, WG15, Zha14]. Decisions [BC67, BB88, DZ00, FS71]. Decks [Got87]. Decomposable [Hau12, LL78, LY95, LL90, VL15]. Decomposition [BS82b, Dos85, Fen93, Hey95, Ish04, Ito78, KM88, KMSP12, LT90, SS89, Tor69, Tor69, BM23, Che18, Gol23, PLRB23, Sim19]. Decompositions [CD06]. Decorated [AW93]. Decorrelation [BHTLV20]. decrease [vdHKPS23]. Decreasing [Ath88, GE82, MY84, AM16]. Deep [Pha79, FPKM22]. Default [SE12, PKO16, Tar15]. Defaultable [CLS09]. Defense [AMAKAH97]. Defensive [Cow92]. deficiency [AN22]. deficiency-zero [AN22]. Defined [Bax98, Ery08, FN70, Hel70, MS91, Miy94, Reh77, Wol77b, BK21]. Defining [HY97]. Definition [Dal71, Mal74]. Deformation [GP04a]. Degeneracy [Ath92]. degenerate [Jan15]. Degradation [KKY02, Ols77, SM72]. Degree [FHTZ10, HKO94, How00b, KKP10, SZ91, Ste19, WDLC09, BS13a, BS15b, BvdHK17, BJMK18, Bri20, DS21, SRT^{+}16]. degree-wise [BJMK18]. Degrees [CM16a, HJ13a]. Delaunay [BK21, Rat92]. Delay [CY90, CST08, CP11, Li96, NO74, OV03, Ros74, Ros78, Ses96, Van99, VB02]. Delayed [BM97, Bré00, CS08, GT06, HF91, KS85, KM88, LL75, YGN05, DZ13]. Delayed-Claims [YGN05]. Delays [BBK08, BDMZ12, Gav69, LJ78, LP85, Niu80, Rei67, Wan99a, YW64]. Delbrück [Gol95, Kem94, MSS92, Möh05, Pak93, Pro96, Zhe09]. Deleted [RR98]. Deletion [RE01]. Delta [Mar00, SS07]. Demand [BDO07, DW77, GHL71, Lan87, Sah83]. Demands [De 85a, Fal94, LR07]. demimartingales [KL17]. Demographic [CW69]. Demography [Lan79a]. Denominations [Van06]. Dense [GH10, Pis06, BKK19, BCN24, HM23]. Densities [Ano87g, Ano91i, ASS85, Bor94, Car02, DGNR95, Di 98, GNR89, GRRT97, IK11, JGR91, Kei71, Kij87, Mee95, NRS85a, NRS85b, Pau87, RS83, RSS84, RS87, SS80, Sut89]. Density [Ano78e, Ano88h, Ant78a, Cab84, CW93, Cra73b, Dah85, DG93b, DS75, Dur85, Dur88, DW92, FI98, GS72, GL99, Hsi76, Imh84, JRR95, JK81, Kle88, Loe09, Meh67, Mil92, MO70, Mod67a, New66, Pap71, PP07, RRS05, RS88, Sch91a, She79, Sri91, Tan80b, Tho11, WW06b, Zip11, Zho23]. Density-Dependent [Kle88]. Denumerable [Ano93f, FHT78, FT78, GZ02, HST75, HY94, JK71, Kal89, Kij92b, SVJ66, Twe73, Twe75b]. Denumerable-State [GZ02, HY94]. Deny [FS97]. Departure [BM97, Fak82, Fol86, HPR86, ST73, Tan94, VD69]. Departures [McN74, O’C97, Sha75]. Dependence [AT06, ALS04, BS04a, BPT89, Bra83, BN92, Cao84, CD07, CJ06, Çin67b, CMS06, DS06, Den10, ER88, ER89, Fea80, Fea81, Fer93, Fin82, FH17, Gne00, HY97, Hey99, Hey02, KS01c, Kün86, LM90, Leh72, LS07, LX00, LS09a, Ma02, Ma03b, MW97, Miz06, MR01, RCC12, Rüs04, Sca98, SS99, Seg06, Sub79, BBvdMP23, BS14b, BS16b, BCS23, KS21, MV16, MR23, Tu19]. Dependency [Law71a, BBvdMP23]. Dependent [Ano74h, Ano78f, Ano92g, Ano05e, APS74, ALRW99, BCL09, BB05, BG87a, BBS85, Bra80, Bur09, Cai02, CSZ97, CL85, CZ98, CH06, CCN10, Che11, CS72, Coh89b, Con71, Con74, Con75, CM69, Dai04, Dal82, DW72, DO12, Don76a, Don76b, Dsh01b, Ebr87, Eks03, Fal94, Fea76, Fea80, FWL03, FC03, FPY04, GD74, Gao04, Gne94, Gne07, Gol78b, GAS03, Gre76c, GHL71, GMW90, Had74, HH03, Her72, Höp85a, Höp85b, Höp86a, HY02, HB02, Ish80, Jac78, Jac75, JS00, JRS06, Kap74, Kap75a, Kap75b, Kar76a, KP89, Kat77b, KR70, Kel00, KZB05, KN90, Kim94a, Kle83a, Kle84, Kle85, Kle88, Kle89, KT08, KM92, KN00, KP83, Kuu82, Lai78, LLS84, LY95, Leh73, LF08, LPSV00, Li00a, Li06a, Lin89, Lip77, LNTZ99, Llo71, Lon09, Mar05, MF01, Mer80, MY84, MY85]. Dependent [Miy94, MJP87, MMR02, Nag93, New68c, New68a, New68b, O’B74, O’B75, OA12, OM72, Pes68, PPD97, Pon07, Pra64, Pra80, PH86, Pru85, Qui78, Rad73a, Rad76a, Rah98, RF88, RS95a, Ruz10, Sau76a, SS07, Sel97, Sgi01, SS85a, Sha89b, SD76, STC93, SZ06, SR70b, SC73, Sud81, Tak92b, THP12, Tin85, Ton74, TO92, WF99, Wan99b, Wau74a, Wei66b, Wei66a, Wei69a, Wei70, Wei72, Wei74, Wei75a, Wei76, XY09, XW05, YH04, YC08, ZJ04, ABI14, ASB21, AA23, AF14, BJ14, BKJ16, BS13b, BMR19, BB18, Bre13, BN15, CS20a, CC15, DHKP18, FCP20, FW21, GOP15, GS19, GKS14, HF19, Hon13, Jan15, JF23, JF16, KA22, KA16, Kli14, NFS20, OT16, PSZ13, PR13, Vid19, Wan14, WW20, YWK14, ZC14]. Depending [Fak87, SG89]. Depletion [BCJ04, DW89, KR88]. Depletion-Time [KR88]. Deposition [Sud02]. Depths [MR11]. Derandomization [Mah03]. Derivation [Oga80, PR95]. Derivations [BM05, Miy82, Wol88]. Derivative [Hei01, VAK92, SZY19a]. Derivatives [Ano65c, BD75, CJ06, CY90, Kra64, Lau80, San06, HK94]. Derived [Cal74, Car02, PT86, Sav88b, Teu68a, WC95, HJL21]. Dérivees [HK94]. deriving [HK15]. descendant [AFP22]. Descendants [Käm91]. Descent [Bro00b]. Describing [Rad76a, SM82, Wan80]. Descriptions [Eth92b]. Descriptive [Fin82, Rob90]. Design [Ano82f, Che06, De 85b, KS78, PPBT99, Pro82, SE12, Hu16]. Designs [AG07, AGP13]. Destroyed [Hüs95]. Destruction [Lan11]. Detailed [DO97, SGC16]. Detecting [Had05, Las94b]. Detection [Bal75, Bal76, IG01, BEG22, Ber23, CF13]. detections [LP17]. Deteriorating [Gla93, JZ02, MR03, OKM86, Qiu91, Yeh91, Yeh03]. deterioration [Nat14]. Determinacy [LS02a]. determinantal [BO18, Gol23, LGD20, MO21]. Determination [Bau83, DKO84, GL84b, SC81, Ton75]. Determine [Kot94]. determined [Ros14b]. Determining [Tai95]. Déterministe [HK94]. Deterministic [Cos90, Cos93, CD09, DDM98, Gan94, GY95a, Nue08, SW11, Wu68, CdSJ20, Fuj13, Gén23, HK94, RDM19, WBS16]. Deviation [Bon02, Duf96, DS08b, FWL03, GS79, Hon02, Mas95, Suz10, Zha05, ADJ13, Chu18, FX16, GR21, KLP22, KLY16, LSWX20, Nyq17, PSK17, SUY21, Sta20, Zha22, Zhu23]. Deviations [AL12, AMW88, BL09a, Bro74, BNS90, Cai03, De 94, DM05b, DMT11, DT11, DNW07, FLP99, GY08, Gol76, HS10, Jon04, KM97, Li12, Mac01, MS06, Mór96, NTYY03, NTaYY04, Nyr99, O’C97, O’C98, Pap96, SW87, WW07, Web07, YCG11, Zha04, BSX19, Bon13, BN15, FS23, GJP^{+}21, JPS18, JS20, Leh17, LR24, LN23, LX20, SX18, SB17, UY23]. Device [Whi90, CHJW24]. device-to-device [CHJW24]. Devices [AHS78a, AH84]. DFR [BS15a, Yu11]. Diaconis [PT22]. Diaconis-Freedman [PT22]. Diagnostics [Liu04]. Diagraph [Ano65e]. Dialysis [Kas66]. diameter [MS20]. Dice [HR00, Rot98, CM16b]. Dichotomised [KN90]. dichotomy [DL17]. Dickman [Cov09]. Dickman-Type [Cov09]. Dietz [Dow67]. Difference [Dar89, DZ07, GY95b, GG77b, Gue81, Sev67, Sev69, SH03, Tol88, ABI17, ESC15, GK18]. Differences [Huf88a, Kub24]. Different [BC95b, Bar75c, Bec73, Dur88, Jag69, Kel75, Mar11, ASZB18, FLM15, Jor18]. Differentiability [How01, JZG11, DLMR17]. Differentiable [Dav73, KR97]. Differential [AT74, AS09, Ara94, Bro00a, CJ08a, Dar66, GK76, GBR75, Hey82a, HMPS03, IS83, Kur70, Kur71, Møl95, Mor71a, Nag82b, Sev67, Sev69, SH03, Tai95, Tso71, Zen10, BY14, CF17, Eri16, FX16, FFK19, FFK20, FCKS16, KP22, LZ14, NNY20, Rek23, Sön23, Wan15]. Differential-Difference [Sev67, Sev69]. Differentiated [Ye07]. Differentiation [Loi05]. Diffusion [Aas76, Aas77, AI00, Ano65d, Ano66c, Ano77g, Ara04, BMN09, BMT10, Bal88, Bar72a, Bat68, BH75a, BH75b, BGNR90, Cai11, CFS11, CK07, Dav86, Dav65, DGNR95, Doh87, Eks06, Est76, Eth79, Eth92a, FW07, FZ93, Gap07, Gav68, GNR89, GL83, GS06, GG77b, GRRT97, Har65, HRF84, Hey94, Hey74, Hog86, Hor98, IK11, Igl65, Igl69, Iva80, JFHL09, Jen74, JGR91, JC06b, Kao78, KS06, Kei64b, Kei65, Kei71, KN01, Kim64, KPSS96, Lán79b, LK77, LG81, LF08, LY22, Lin83, McK84, McN73, Mil92, Mon67, Mor71b, Mor71c, NRS85a, Nor77, OC03, Par88b, Pli75, Rad72a, Rub64, SWW21, Sch91b, ST12, Sie79, SW67, SH70b, Taï96, Wat64, Wat76, WPN80, ZW02, Zuc80a, ABSKT19, AS17, BS22, CM15a, CF13, CdOR16, Dai14, Eri16, GR14, Gor14]. diffusion [Jia15, JSW19, Kun22, Lef24, LTZ13, LB23, Man13b, MBvLW16, RR16, Rug14, SW20, Zho15, Zho23]. Diffusion-Reaction [Bal88]. Diffusion-Type [Hey94, CF13, GR14]. Diffusions [BR00, BK87b, BGT04, BHPZ06, Day83, KPS04, Ken74, Lip77, Pau87, Sat83, SL21, Spi15, Yam22]. Diffusive [Sud01]. digital [FH17]. Digits [Ano85f, BT82, Blo82, Blo84, CHM^{+}22]. Digraph [Bha65a, BGJ12, Pen16]. Digraphs [Ber73, Jaw99]. Dilation [Bor96, MPSG90]. Dimension [Bal88, FNV11, ÓT21, RWZ13]. Dimensional [Adl76, Aff89, Ano78g, BCR11, BF91, Bèr04, Bil81, BD88b, Bla84, BHPZ06, Çin68, Coh88, Coh94, Cox77, Cre74, Dal71, Dal79a, Dar89, Dav86, Dav76, DGNR95, DF81a, FW07, FG98, GW76, GW80, GSS12, GNR89, GL99, Gra99, Hal84, HRF84, IK11, Imh84, Ito80, JM86, Kar76b, Kel00, Kol06, Lai77, Lal86, Lie78, MP00, Mar86, McC74, Mil94, Mod67a, Mul68b, OW94, Ors86, Pag71, Pau87, PP71, Pic71, QW84, Sad01, Sha89a, Sta82, Tzi11, Ver69, Ver98, WM77, Whi72c, Zha92, Zie87, ZG88, AG16, Cha17, CM15a, Cse79, DHM20, FW16, Gor14, HJ13b, HS14, JP22, KR18, LTZ18, PT22, VR22, Yam22, Zho23]. Dimensioning [BP97]. Dimensions [BS82a, BS97, Coo89, DS86b, ST86b, Zha96, BO18]. d’Invariance [Bie98]. Diploid [CH84a, CD85, CH90, Hin82b, HB83a, HB84]. Direct [Jen72, SS85b, ZS92]. Directed [Bri20, Kal91a, LW93, CP14, Gab16a, KS23b]. Direction [Gor75]. Directional [Bin78]. Directionally [DM10a]. Directions [KO05]. Dirichlet [Car02, Gri88, HL14, Hop86b, LP14, McK14, Rug14, Xu09]. Disaggregation [Ben70]. Disasters [AK76, DN99, DS04, JS96, KSN75, CNZ14]. Disc [Lyl88, FV18]. disc-polygons [FV18]. Discharge [HF71]. Discipline [AA06, Fak86, Fak87, KR88, Li96, RS92, Sen89, Tan94, Whi71b, Win77]. Disciplines [RSY90, SS85a, SS87]. Discontinuous [Moy65, Sön23]. Discount [All04a, CD09, KR00, WG15]. Discounted [AHN90, Che76a, Coo98, FF09, Gap07, GA22, Gla91, GZ02, GHL05, LY22, MM94, Mor71c, Ped00, PRHL12, Rig11, Sob82, Tan07, Yao07b, YC08, BS22, Fro15, GPZ17, YWK14]. Discounting [FB88, KSHH98, Mor71b]. Discouraged [Nat74, Van81]. Discovering [AI89]. Discret [PL03, PLC03]. Discrete [AC02, And70, AW90, Ano94f, AG87, Bal78, Bal72, BLL80a, BL11, Bli89, BM94, BP99b, BK06, BBS90, C.12, Cao95, CD07, CG96, Cow01, DS65, DM67, Dav01, Daw73, DR09, Die82, DP10, Ell93, Ery10, Fad74, FF11, FH94, FZ93, FWL12, GT74, Gly89, GL83, GR01, GH09, Her72, Hic83, Hos06, HM09, HCK02, Imh74, Ish04, Kal95, KS72, Kan75, KP96, Kha77, KL77, KL78, KT01b, KPT99, KZ95, LOR19, Li06b, Lit92a, Liu11, MSS92, MS93, MP12, Meh77b, MPHL10, Min78, Mór96, NP91, Oht91, Ols67, PK80, PLC03, QM86, Ren87, Rey70, Ros93, SV97, SVJ66, Set03, Sha75, SV71, Ste03, Tak71a, Tal80, Tru00, Wal83b, Wat88, WGF13, BM14, Gol23, Gra18, GH13, Gue13, HBHH23, HL16b, KSC13, LL13, Lei23]. discrete [Mia13, Ott20, PL03, YML17, Tav80a]. Discrete- [Ste03]. Discrete-State [Cao95]. Discrete-Time [AC02, BL11, CG96, DS65, DM67, DR09, DP10, Ell93, GT74, Gly89, HCK02, Ish04, KT01b, KZ95, Liu11, MP12, Meh77b, MPHL10, Min78, Oht91, SVJ66, Tru00, Wal83b, Gue13, HBHH23, PL03, Tav80a]. Discrete-type [LOR19]. Discretely [JKM12, KPS04]. Discretionary [BHPZ06]. Discretization [KPSS96]. Discretizations [CJ08a, Cos93]. Discriminating [Nag01, Ros14a]. Discrimination [AEHS05, Kün86, MK73, Nis74a]. discs [FV18]. Discussion [Bar72b]. Disease [Ano67d, Ber65, Dub81, FG99, Gri72, Kat77a, Sap87]. Disjoint [Rog78, Van96]. Disjoint-Occurrence [Van96]. disk [Gaá20]. Disks [Cow93, Ari14]. Disorder [BEG22, JW11, Wic79, BM15b]. Disordered [CMY98]. Dispatched [Cra74b]. Dispatching [Ros69, XMKW90]. Dispersal [Rou90]. Dispersion [Ano98f, CG76, Hic86, JS98, Mar86, SM94b]. dispersions [MR15]. Dispersive [Ano85e, BR02, LT81, MPSG90, Sha82]. Dispersiveness [Lyn87]. Dispersivity [DW85, LT81]. Displaced [Col76]. Displacements [Man83, NW70]. Dissipation [Fox65]. Dissociated [BE87]. Distance [Ala76, Bau81, Ber77b, CV09, Daw81, FS71, Gro80, IS86, Lit74, Mal74, RP03, SF83, Sha81, Sud77, Tho75b, Web99, Wil77, YCG11, ANT17]. Distances [ALM06, Dob96, Gil77, Hen82, KS83b, MR11, Ryc87, Sil78, SB78, Ste03, WG93, AB18, BJ24, DDS22b, HH23, Sal13]. Distinguishable [Ito77]. Distinguishes [CI72]. Distorted [Sob67]. Distortion [Ori79]. Distributed [Ano86a, Ano90h, BN06, CCGP91, CT95, DN04, Gra85, KBBW07, KS78, Law92, LM89, Rev93, Sto08, VE98, Wei75c, ZJ04, McK14, MRSA21]. Distribution [AH84, Abr94, ALM06, AR72, Ahs76, Ahs78b, Ail68, Ala76, Ala77, Ale01, AARJ10, AA90, AT91a, Ano65c, Ano74f, Ano77g, Ano78h, Ano81f, Ano86g, Ano86j, Arn80, Avn79, BdHK93, Bal86, Bal90, BMY93, BO99, Ban96, BLBL10, BMZ07, BP04, BRY04, BL73, BS76, Ber85, BJ79, Ber92, Bin78, BV96, BK06, Bra83, BBS90, Bro64, Bro04b, BG05, CW81, CP82, CM83, Can74, CGR08, CP01, Cha75, Che76b, CN80, CK00c, Chu75, Coh67a, CS80, CT76, Cow01, CD81, Dal74, Dal81a, Dal87, Dal16a, Dav86, Dav82b, DFSD01, DK11, DH89, DM10b, DK90, Dip98, Dob96, Dom96, DG99, DF81b, DS04, Eat66, Edd80, Emb83, EZ80, Enn69, Ewe64, FN70, FG99, Fea03, FT80, FY82, Fel66, Fer95, FTV90, Fu01, FWL03, FPY04, Gab16b, Gal71a].

Distribution [Gal73, GH94, Gat82a, Ger11, GNRW94, Gol95, GH90, Gri81, Gri88, Gri92, GKJ80, GM94, GL84b, GRD^{+}00, Had69, HTW73, Hai88, Hal81, Hal84, Ham92, HST83, HA96, Hen72, Hen82, Hey64, Hey91, HH96, HM99, HK85, HR88b, Hos06, Hua74a, Huf88a, HB02, JP99, KS01a, KH83, Kap74, Kar77, Kei82, KK69, Kem94, Kem98, Kem79, Ken91, Ken76, KKK07, KT83, Kin77, KS02b, KS83a, Kra64, Kup90, Kyp71, LR90, LR92, LRS94, LP94a, Les10, Lev77, Lew67, LO95, Lit94, Lit69, Lit74, LH02, LT71, LT72, MSS92, Man77, MC78, Mar73b, MO67, MC11, MV95, MQ00, McK82, Mei81, Min78, Miy86, Miy01, Möh05, Møl96, Mor74, MZLSRS02, Nar68, Nay87, Neu68a, Nic84, Nis74a, Nis74b, NRS85a, OL65, OV03, Ott84b, Ova01, Pak78a, Pak93]. Distribution [PP73a, PP69, Phi81, Pie07, Pod72, PT11, Pro96, Pyk94, Raa80, Rat92, RZ07, Ren80, Res71a, RS90, RD99, Ros74, Ros12, Rud75, Rud76b, Ryc87, Sad01, SC75, Sam71, SG89, ST86a, Sch88a, Sel83, Ser90, Sha81, Sha74a, Sha77, SK23, SS96, Sim00, Sre83, SS70, Sri86, ST86b, Ste04, Stö74, SAS04, SC81, SBG70, Sze86, Tak67, Tak71b, Tak88, Tam70, Tan79a, Tav81, Ten75, Tol88, VS91, VB02, VP01, Ven79, VT75, VT76, Vrb85, Wan81, WCW05, Wat80a, Wat76, Web99, Web07, Wei66b, Wei69b, Wes81, Wie89, Win79, Xu98, Xu09, Yan75, YI99, Yor95, Yos84, Zac04, Zij83, ZNF10, BS13a, BS15b, BPPR16, BB18, Bon13, BAR14, CL14, CPN18, DS17, GR13, GKS14, GM20, GHW19, HZ22a, Joh24, KMG19, Lei23, Let19]. distribution [LPN13, MSV14, Mar16a, Möh21, Per23, Sal13, Wan14, ZZ13, Zha22, ZW17, Ano04g]. Distribution-Free [GRD^{+}00, Pyk94]. Distribution-Invariant [Web07]. Distributional [EW00, JS08, KP07, McK88, Pap94, ZT01, Zij83]. Distributions [Acz72, AD94, AC94a, AT79, And70, AA93, Ang11, Ano70e, Ano75g, Ano79f, Ano82g, Ano85g, Ano85e, Ano90f, Ano92h, Ano93e, Ano93f, Ano98e, Arn75, AL85, AFG11, AM05, ALRW99, BFL05, BCL09, BMZ07, BP10, BB84, BK86a, Bee75, BT92, BL00, Ber83, BM90, Ber77a, Ber77b, BD94a, Bha80, BS90, Bha91, BAAB00, BS69a, BS69b, BSW03, BP99b, BVJ00, BK73a, Bui75, C.12, CW97, CG99, CT04, CL05, CMW95, CW91, CS86, Cha83, Che81a, Che94, Che03a, Che70, CDL88, CY65, Chr12, CV09, CD92, CP03, Cla12, Coh73, CC96, Cos90, Csá94, CT90, DS65, DS67, DFSD01, DT80, Deh84, Del78, DL02, Die82, DT11, DR81, Dur86, Dur74, Eat66, EO84, Eme72, Eng64, Epl83, FHTZ10, FP94, FGL00, Fei94, FPKMP98, FMS99]. Distributions [FH72, FH94, FKZ09, FS92, FC02, FC03, GA99, Gas64, GW85, Gat87, Gay04, Gel77, Gol77, Gol78a, Got80, Gra12, Grü85, GG78, HN71, Hal78, Hal83b, HL76, Har90c, HP95, HL16a, Hic83, Hic86, HL05, HS93, HS12, Hüs82, Ish74, JdlR04, Jen66, JW82, Jun94, Kat02, KL03a, KMR75, KN75, Kel12, Kem92, KS91a, Kij92b, Kij93, KD78, KPS81, Kla02, KM03, KKRT12, Klü88, KMS83, Koo04, KTW69, Kru04, KP87, Kyp72a, Kyp72b, LS87, Lem74, LT81, LK01, LS09a, Lin94, Lin95, Lin96, Lin97, LS02a, LS09c, LH07, LS86, LV72, Liu04, LS00, LS79, Mac80, Mac81, Mae80, Mah03, MO92, MC70, MC71, Man74, MM71, MO93, MS94, Mas86a, MS91, Mil67, Miy92, Moh76, MR02, Möh05, Mor93, Mor80, Mór96, MS92, Mul68a, Mur80, Nag82a, Nag93]. Distributions [NJS96b, Nat75, Neu81, New66, Orr69, Pak75, PP70, PK80, Pfe82a, Pha93, PS86a, Pol99, PG79, QM86, Rao79, Reh77, Ren79, RWN02, RP03, Ros08, RM88, Roy89, RC80, Ryd96, SM78, Sav88b, SVJ66, ST85, SD94, Sgi99, Sha82, SS93, SS85a, SMS94, SS83b, Sim97, Sim98, Ski67, Ski68, Sko87, SV71, Sta82, Ste06, Sto04, SV03b, Tai95, Tak95, Tal69, Tal80, Tav80a, TV04, The79, Twe74, Voi96, Voo87, Wan76c, Wan90, Wat82a, Wat88, Whi85b, Whi68, WL94, WL97, Wil02, Wit93, Wol88, Woo82, Woo83a, XZW09, Yu09, Yun00, ZL10, Zie87, AB19, BY14, BLS15, BDG14, Bri20, BCdlP17, Cha17, Cur22, DDS22b, ET21, FKZ13, Gor14, Gra18, GK22, He21, HS22, HK15, HM15, ISK18, KBM21, KS20, LL13, MB16, MC18, MO21, NvdBOdI24, Neu23]. distributions [PD20, ST16, Sei16, SGC16, TK15, VKR14, Vid23, Yam22, YML17, Yin21, Zhe09]. districts [AM18]. Divergence [Dob70, VK10, Gau22]. Divergent [Hai78, Hol67]. Diversify [Ibr07]. Diversity [Mar00]. Divided [Huf88a, PL73, Rén64]. Dividend [CL09, KP07, RZ07, CY16, Jia15, NPYY18]. Dividends [Fro05, Loe09, Fro15]. Divisibility [AP94, Ano79f, Ano06e, BR77a, BGSS03, Iks04, Pak04, Pak07, PG79]. Divisible [Joe96, Mil67, VT11, CM15c, RNJ16, RNJ17, SRN21]. Division [Ale01]. Divisions [Jan78]. Dixie [Kla94]. DMRL [BKS02, FP97]. DNA [CCG90, Cow91, CC94, CCH95, Cow01, Gri98, Lit69, Lit74, Lit79, Pia00a]. Do [Kap78, MC70, MC71, PRS08, Yu11]. Dobrushin [HZ22b]. Dodecahedron [LT80]. Does [Ano10e, CNS99, DJ09, Mar07, PS18]. Domain [LS00, ZZ13, BM20, MO23]. Domains [Gra18, Kel86, RVJ68]. Domb [CR92]. Dominance [DC10, Luo14, LHT17, MWH22]. Dominant [CH84a]. Dominated [Gol78a]. Dominated-Variation [Gol78a]. dominating [PW16, Thö16]. Domination [DH08, Eco03, Hos09, ST86b]. Doney [De 82, Zho08]. Doob [Bin05, Sne05]. Double [DW11, Eli09, Kar09, KPV21, Lin99, DL18, SW20]. Double-Barrier [DW11, DL18]. Doubling [KS05]. Doubly [Gra71, Gui02, Roo76b]. down [LP21, LVZ19, LH07, SF17, Sag23]. Downcrossings [Coh77a]. Downtime [HA96, Pie07, SV03b]. Downtimes [GA99]. Downton [Boo89]. Downward [MT02, BS23]. Dr. [Gan84a]. Draft [Peg75b]. Draft-Patterns [Peg75b]. Dragged [Fra07]. Drainage [CFD09]. drains [Dio76, Dio78]. draw [LVZ19]. draw-down [LVZ19]. Drawdown [MIAPAM04, ALL21, HBHH23, LLZ17, WZ18]. drawdown-based [WZ18]. drawdowns [BS22, GR14, LLZ15]. drawing [LMS18]. drawings [BT23, KS17, SH21]. drawup [LLZ17]. Drift [AH00, Ano78e, Ant78a, AM78, AC94b, BG06, Bro74, Can73, Can74, Dom95, Dom96, DG89, DY98, Fer83, FMS99, Had05, Hal69, Hey67, HL13b, Igl74, Imh99, Jia12, Kao78, KPSS96, Kot94, Lee11a, Lin04, MS94, Pol68, RB05, Rot95, She79, BDM21, KW13, LY23, RWZ13, SWW21, SGS17, Sön23]. Driftless [Cho94]. Drifts [AFM95, BHZ22]. Driven [AHL02, AS09, BC01a, BDY07, Cai03, DJ05, DESM10, ESM08, GM11, GBR75, HT12, KW03, KLM04, LR12, Ped02, SC82, WW06a, Wan12, CCCSdO24, DM18, FCKS16, GS19, LS20, MLB16, Rek23, Sim23, TTI22, Wan15, ZTML21]. driving [JP18]. Drug [BD88a]. Dry [Hut90]. Dryden [Ken91]. Dryness [Bal78]. DSc [Ano04f, GS08, Sen17]. Dual [Lew01, LPLG08, YSL17]. Duality [BMN09, Bar75b, CG00b, Gre69, HM09, JR82a, KP89, Man13b, ZLA99, Kol15, PSZ13, Yao13]. Dubins [Yao07b, YC08, LS17]. Dubins-Savage [Yao07b, YC08]. Due [Ano83f, Arn68, Gla82a, MC78, Nat82, Nat90, Pur87, Tra81]. Duels [KL77, KL78]. d’une [Let78]. d’Unicité [Mat74]. Duplication [Arn72, CJV10]. duplications [BM15a]. Duration [AARJ10, Ano78h, Bil77a, Enn69, GP98b, Kar77, KA09, Rus78, Sve95, Yeo97, ZW02, GS15]. durations [BBN16]. Durbin [Zip11]. during [Coh67a, PRX03, Ros73, Ser90]. Dust [GIM11]. Dvoretzky [KW14]. Dynamic [AHL02, AEHS04, AEHS05, AA14, ASB94, BOP07, Bel66, BN94, BLT11, BN13, Cam97, DDS22a, Dei86, DC77, FG08, FS16, GW94a, Gre84, Har80, HKN89, HW96, Mor68, Rie76, SS91, SS93, SS89, SE12, Sto08, VS00, CHJW24, ESC15, Sta20, THB23, YBL23, ZC14]. Dynamical [Bax98, Ben98a, CF22b, Fra07, GT05, Hor04, LR95, Oza86, SC82]. Dynamically [Sie81]. Dynamics [Gra99, How00b, Jag92, KL00, Mor93, Zac00, BS19, BVW21, CFRS22, GT18, PG13, Rug14, ZTML21]. Dynkin [EGL17, KY17a, Mar16b, Oht91].

E. [Ano86p, Wei79]. Each [Min79, Pol78]. Earlier [Gre86]. Early [BL09b, IRS23, BL15]. Earthquake [Lai77, OK86, Oga01, SLZ01]. Easy [Ros08]. EATA [KK95]. Economic [Agn72, SZY19a]. Economies [Duf96]. Edge [KS10, KL03b, RR77, Ros12, ARS21, DIÖ23, Ger23]. edge-step [ARS21]. Editor [AH19, Ano73l, Che82, CH84b, Fal89, HS87a, Koo04, KKC82, Lai87, Mad76, Sha72a, TH71, Tod85, Wil88b, Wil77]. Education [Phi01]. Edward [Gan71, Ano94l]. Effect [AB98, Ano84e, BGW83, GG10, Hey82a, HLT00, Ish74, KJ81, KNR05, MA69, Pré94, SM05, Sco82, SY86, TG87, Whi80a, BJMK18, EL19, Mia13, SV19]. Effective [GRD^{+}00, RNR18]. Effectiveness [Con75]. Effects [BSW07, CW69, GT18, KM04, KPSS96, Spe86, Sud77, Wil77, Nea16, WZ13]. Efficacy [Ber10]. Efficiency [APZ11, FKM10, Jen72]. Efficient [ADJ13, AR05b, BL10, BL11, BGT04, DG06b, IK11, NR22, Sen73a, Lei23, LX18]. Efron [Hu16]. Eggenberger [GR13]. Ehrenfest [BMT10, Che06, Det94, Kap10, KM65, KS93b, LGC16, Sev70, Voi96]. eigendecomposition [LGD20]. Eigenstructure [Eth92a]. Eigentime [CM15a, Mao04]. Eigenvalues [DS94b, Fra00, Mar86, Mat92a, PS04, Men23, UY23]. Eigenvectors [Mar86]. Einstein [KP23]. Either [Zho07]. Elaborations [Bra81]. Elastic [Dom95, Dom96]. Elasticity [BH12]. Election [KMW11, Tak70]. Electricity [Hin04]. Electronic [Win05]. Electrons [Ste86]. Elementary [Ano66d, Bré91, Gei99, Hor65, Mak94, Pfe82b, Reh77, Sha74a, De 22, KK24]. Elements [YZ85]. elephant [GS21]. Elimination [FE87]. Ellipses [DS93, EV82]. Ellipsoidal [Møl88]. Ellipsoids [DS93, JM86]. Elliptical [Liu04, Sei16, Yin21]. EMA1 [LL77]. Embedded [Bri79, Cao95, EE88, HL04, Whi72b]. Embedding [BR87a, Hal69, Hol08a, LR00, Nue08, RR10, Gue13]. emergence [BCK16]. emergency [BPS23]. Emergent [AR13]. Emerging [Ors89, SO11]. Emigration [Tes84, Bra22]. Emigrations [Tra81]. Emperor [KR08]. Emphasis [Cov09]. Empirical [Ail68, BH91, CT90, De 94, KR92, Ken75b, PP70, PS01b, PS04, Tak71b, Win79, CS20b, Zha22]. Employees [Ber10]. Employer [MC90]. Employer-Candidate [MC90]. Employment [CB83, Pak83, Smi75]. Emptiness [BC75, Leh71, Leh73, Pha69, Ros76b]. Empty [CR97, Ken93, ZL15]. Emptying [FH10]. Encounters [Gar74]. Endemic [Kes08, Rad74b]. Endpoint [HJ11]. Endpoint-Conditioned [HJ11]. Endpoints [Jan88]. Energies [Lai77]. Energy [BC01a, FKM10, Has82a, HK85, JZG11, Tsa99, CGR08]. engineering [Tav20, YL16]. Enrique [Gan98]. ensemble [Ari22, HM23]. Entering [AR82]. entirely [Tzi16]. Entrance [C.12, Coh88, GG78, VK00, HZ22a]. Entrance-Exit [GG78]. Entries [Van86]. Entropies [Eba84]. Entropy [AEHS04, DVJ04, DL02, GJQ06, GL03, JFS12, Kru04, MTV99, Mar21, Rat70, Rob90, Ruk00, Sen82a, Web07, BRRN16, DT15, GT14, LR13, TSN17]. Entropy-Based [DL02]. Entry [CST08, DZ00, Naw83, PS86b]. Enumerative [Dev79]. Envelope [Adl78, Has79, Lin89, EG15, Ray70]. Enveloppe [Ray70]. Environment [Agr75, Bèr04, BSB99, CM07, CP90, CS88, Dek87, Fal96, Fel78, FG76, GR98, Gla90, GZ94, Ham92, HW84a, KK01, LM90, LS86, MM09, MP00, Nak79b, Nak79a, Nor75, OP86, PT97, PZ90, Pra80, Råd76c, Raf73, Rit80, Tor79, TV93, WF80, WK94, Bar15, DLVZ20, GPSY16, Ker20, Lam14, LP21, MZ24, PT23b, VL15, VD17]. Environmental [Raf73, CF14]. Environments [Ali99, CT81, Cor87, DP12b, Jag74b, JZ02, JK04, Jon99b, Kap72, KS82, LN07, Li00a, Nan79, PH82, Pel07, WF99, Wan99b, Wei71, Wil69, WR97, ZJ04, ASZB18, HF19, WY22]. EOQ [Bar15]. Épaisses [Mat76]. Epidemic [Aba73, Abr94, AD97, AB98, AD98, AS96, Ano74e, Ano79e, Bal83a, Bal83b, Bal84, Bal86, Bal90, BB90, BC95b, BC95a, BO99, BB05, BOP07, Bil73, Bil74, Bil75, Bil76, Bil77a, BLL79, BLL80a, BLL80b, Boo89, BL09b, Cla94, Cla99, CP03, DDM98, DL06, Dow72a, Dun80, FS80a, Fad83, GJ72, GM91, Gan94, GY95a, Ger77, GJ80, GMSB10, Gri73, Hen79, Jaw99, Jer70, Jer74, Kes08, Kry72, Kry75, KS76, KL89a, Kuu82, Lef78, Lef79, LP05, LPS17, Lud73, ML86, ML98, Nea07, Nea08b, O’N96, Pic80, Pic84, Rad76a, Rad76b, Raj81, RR67, Rou82, Sel83, Sev67, Sta01, Sve95, Via95, Wan77a, Wan80, Wat72, Wat80a, Wat80b, Wie79, YC82, Yao93, BS13a, BS15b, Ber23, BL15, KM15a, LP18, MLB16, Nea16, NNY20, PSK17, WBS16, Win15]. Epidemic [Dow72b]. Epidemics [Bec73, Bri98, BDLL08, DG93a, Die66, DD68, Dow67, LL82, Lan81, LM88, Mil08, Nea03, Nea05, Pen96b, PW67, Rad73b, Rad74a, Sau75b, SLK79, SO11, Ton97, Wic75, BBN16, LP17, Sim23]. Epidemiology [RC73]. Episodes [Pre02]. Epoch [BR02]. Epochs [Cho94, DF81b]. Equal [Har09, Raf73]. Equal-Sized [Har09]. Equalities [Imh99]. Equality [Mee95, Sca98, YM91]. equalized [Rob14]. Equalizer [Had05]. Equals [Fei94]. Equation [Aas76, BJ07, AM07, Ath71, Bar82b, Boy69, Boy70, CG87, CLS09, Del85, Don73, FT78, GK76, GC74, GL84b, Hor01, IS83, Ito89, JS08, JRR95, Jen74, Keb97, Ken83, KO05, Kyp98, Mil94, Ors86, PT71, RSS84, Ros87b, Rot94, SW94, SW88, Tol88, VAW04, Wan76c, CGR08, Eri16, KLL15, LZ14, NNY20, Ver14, Wan15]. Equations [AS09, Ara94, Bar72b, Bou95, BS98, BE73, CJ08a, Coh67b, DM75, Dar89, DZ07, FDVF93, FL98, GST00, GY95b, GBR75, GKJ80, GG77b, Gue81, GM94, Hal70, HMPS03, KN80, Kol98, Kur70, Les10, Miy86, MY88, Møl95, Møl96, Mor71a, Nak98, Ors87b, Pur73, Rüs06, Sev67, Sev69, SH03, Tai95, TH72, WD99, YZ06, ABI17, BY14, ESC15, FX16, FFK19, FFK20, FCKS16, JL22, KP22, MST13, Rek23, Sön23]. Equilibria [Kab12, BPG20, CF17]. Equilibrium [AIGR06, BLBL10, BNT02, CM97, CG90, CC08, DFSD01, Dix93, GH90, HK85, Kar68a, Kar68b, KF70, Lew69, LV72, NJS96b, New66, New68a, OL65, Orr69, Rub64, Sah71a, Tor02, Unk79, Voi00, Wes73, Whi68, Wil88c, SDH16]. Equipment [FP66]. Equiprobable [Die82, PK80]. Equivalence [AK86, CT04, Can73, Daw81, Eth92b, HS00, HW02, Hin04, Mat92b, Pak04, Pak07, Res71b, Vit74, Vit76, Wal82, Whi86, FVZ19, HM23]. Equivalent [Bla53, Oes18, RT01, Yah75, Cri18, RNJ16, RNJ17, SRN21]. Erased [Law83]. Erdos [AN22, BT76, DS86a, Kap77, Ste78]. Ergodic [CD06, GW80, GM84b, Her99, Lan79a, Lel05, Mao04, Mat04, Mit05, Nor75, SL21, Sév95, Tan97, Tho85, Win90, YY11, Zho15, Ari22, CL23, CF17, FHL13, GW17]. Ergodicity [Ano91g, BS94, Che02, CK05, FI87, FMRR03, Ger91, Her03, HL04, IT78, IA78, JI88, Kam17, KL06, LS92b, LH06, Mao02, MZ04, Mor75, Nak86c, Nor77, OT16, PS92, RR07, Ros81a, SAI82, SI83, Sen87, Szp85, Tan82, Tan83a, Teu68a, Teu68b, Teu72, TT79, Twe81, Wan12, Zei89, Zha01, BJT20, BW23, DNY16, Kun22, MW21, MS21, VR22]. Erlang [BKT98, CI72, FRT99, MH76a, Mej68, ST73, VM21, Voi00, WW76, Yu74]. Erlangian [Ano74h, Con74, Con75, KTW69]. Erosion [Pur87, TG87, Mat78]. Érosions [Mat78]. Erratum [KZN15]. Error [CFS11, GY95a, Kab83, KS72, Par01, Sch76, ZNF10, BS15d, HLVM22, VKR14]. Errors [Dei86, Kul87, Ris86, Sol86]. Errors-in-Variables [Dei86]. Escape [AC94b, Ser06, CHJW24]. ESS [Hai89, HH85b]. Esseen [Gol05, Jia12, Jin88, KLP19]. ESSs [CH90]. Estimate [Bré00, CW81, Dah85, DM05a, Eng80, Ham05, Lo86, MK73, Sgi01]. Estimates [Dob70, GT06, GS79, Hal82b, Jen72, Loc82, Mei99, Oga80, Pia00b, Zei91, ZY08, FV18, Sha13, Pia00a]. Estimating [BCD93, CO85, DR94b, Doh87, DM05b, FZ93, Hey94, Leh21, LSC04, McD79, Tsu17, Tsu18, Var80, ZR01]. Estimation [AB00, ADLN82, ALS04, AR15, Bar82b, BZ04, BS69a, BD74, Bri73, BDY07, Bur76, Cam97, CG00a, Cor81, DK01, Dio74, DY97, Eli09, EW00, FLP99, GHK04, GLS00, Gra71, Gra72, Han73b, HW83, HS72, Hey73, IK11, JG85, JS04b, Kab81, KS01a, KPS04, KN01, Kol82, Lán79b, LM04a, MW10, McK84, McV08, Mei81, Mul72, Mul73, Nan79, Nan82, NP09, NP00, PH82, Pap96, Pré94, PW07b, QD77, Ris86, Ros86, Sch91a, Shi86, Sie85, Ska71, ST88, SK91, Tan79c, Tan80b, TT86, Tho86, VAK92, VJ82, Zei95, BNW13, Fuj13, HL13b, JF16, ZZ16, dRdC18, HS74b]. Estimator [Agu03, DL96, DS80, Fer83, RS95a, Sol84]. Estimators [DE11, GR12, Gil77, Jia12, KPSS96, Lie77, Lie78, LS91, Mas82, SW87, AB18, FPY20, Nyq17, RNJ19]. ETAS [MZ01]. Etiologies [Kat77a]. Euclidean [BS82a, How00a, How01, SY72, Sta82, SSE87]. Eugene [Roh88]. Euler [BGT04, FCKS16, Hol13, Kem92, KLL15, Wan15]. Eulerian [WS96]. Evacuation [BRS75]. Evaluating [CG00b, Rog00, Sur08]. Evaluation [AM07, Gla82b, Gla90, Kij92a, RS83, Sat77, SLZ01]. Evaluations [PS76a, PS80]. Evasion [Sch75b, Sch76]. Event [LR09, NA04, BRRN16, GH14]. Events [ARW94, CJ06, Col76, Dar66, Eag79, Gla81, Heu87, Hsu79, Hüs93, HS96b, Kwe75, Lam68, Les67, Nak06, NW70, Nie95, Pla85, Pod72, SPS80, Sch88b, Seg06, Spo01, Ste99, Tas97, Tso92, Vel04, Win05, Yun00, ZR01, ADJ13, CFRS22, DL17, Hon17, Owa23]. eventual [PL03]. Éventuelle [PL03]. Every [Ibr07]. evidence [CF22b]. Evolution [CC01, CW69, DK99, GK76, GK12, HS78, JPR95, Kol98, Nor75, Tsa94, Tsa96, FNP19, MK15, Ban07]. Evolutionarily [Aba80, CH84a, CD85, Hai88, Hai90, Hin81, Hin82a, HB83a, HB83b, HB84, Hin95, Tay79]. Evolutionary [Bac01, CC08, Hol66, Lad05, CDP19]. Evolving [WDLC09]. Ewens [Han90, Joy95, Saw83, Tsu17, Tsu18]. Exact [Ano86j, BMY93, Bar72b, BM18, BDL19, BJ24, Buc94, CM83, CS86, Dai14, DW72, DQL19, DM03, DS86a, DYY01, Eng64, Fen93, FWL03, GR89, GXS12, JW82, LL82, LGD20, LRS94, LLT^{+}22, Les10, NO85, PR95, Pin11, QDZ21, QL96, RD99, SX18, Ste06, TP76, CM16b]. Exact-Order [GXS12]. Exactly [Ino23]. Example [Cox77, GL76, Lai78, Teu72, Tit74]. Examples [HS00, KS99]. Excavation [Nak91]. Exceed [Rus83]. Exceedance [HTW09, Rob05, AF14]. Exceedances [Ano06f, Coh89a, Fer93, GDP04, Hüs93, Ser80]. Exceedences [Zaj95]. Excess [Ano02e, AG81, BOR01, Kat02, Lin96, Whi85a, Wol88, BMRRS16]. excessive [Sez18]. Exchange [HN76, KZ95, NR11, Win05, Gag17]. Exchangeability [Bal94, Gra08, Sau76c, CD19]. Exchangeable [AG80, BC73, Eag79, Ery10, GIM08, HS22, Hsu79, LU96, Mor73a, RR10, Tro83, Web80, BAR14, Ery15, HJ13a, NFS20, PB20]. Exchanges [BS79b, Hai81]. Excitable [Gra99]. Excited [Bra85]. Exciting [Ada75, HO74, Oak75, CH13, FLM15, HLMOL21]. Exclusion [FF96, FNV11, LV12a, Sud02]. Excursion [GL99, GP98b, Ken76, Lou84, RNJ17, DPW21, Wiś20]. Excursions [Adl76, AK97, Csá94, GM94, GP99, Har90b, Imh78, LH07, Sch03, SD88]. exemplar [GT18]. Exhaustive [KR88, LZ94, Tak92b, Tan94]. Existence [Ale01, Ano93f, Boy70, CD05, DV20, GPZ05, KL03a, Kij92b, LU95, LS04, Loc82, MSV14, OW94, QS99, Sam67, Sgi99, Shi02, STW94, Twe83, FPP19]. existing [Jor18]. Exit [Ano86j, Bal95, Bau09, Bre12a, CS86, CST08, DS94a, DZ00, GG78, LVZ19, MC78, Mas87, PT11, Tuc76, Zho07, BPG20, BM20, GLS19, KKS15, LYZ17, vdHKPS23]. Exit-Time [MC78]. Exodus [CR97, ZL15]. Exotic [GS01]. exp [FHI17]. Expansion [ARM98, Bro64, IKP01, IS83, Mar73b, Sgi01]. Expansions [BMZ07, BR96, Bor96, CI78, DN80b, Sko87, Sri98, Tan86, But19, LPN13]. ’Expectation [Fei94, Aas77, BTH72, Cam97, KOH23, Lam92, LU95, Miz06, Sam67, Sgi01]. Expectations [Ano98e, BV96, BN02, CL05, CMW95, Dal87, Hen73b, Kad74, KH75, Ruz10, AB18, BS13b]. Expected [BG87a, BC73, BF93, BS09, CHL98, CY90, Cow91, Fro05, GL83, HST75, How01, Jam66, Kul86, KA09, Lef78, Mor93, Mor66b, Niu80, Pol78, Roe70c, Ros87a, SLK79, Sud81, TK84, Wal98, WVW86, Wic79, BBN16, BDM21, Bis23, Bre13, KS13, Pap17]. Expenditure [KS71]. Experimentation [Neu88]. Experiments [Bla53, BK73a, Fin88]. Expert [BS06, SWW21]. Explained [CN86]. Explicit [BVJ00, BN05, CS97, FH07, Guo01, Han87, HL04, IMS98, KS20, Kol00, LB89, Lou78, Man74, MT06, Neu94a, RN80, SC02, Van90, Yao07a, GG14, HZ22b, JSW19, MB16]. Exploiting [DS92]. Exploration [BPS12, BB88]. Exploratory [Oga01]. Explosion [Ano86g, CGR08, FY82, LMW00, MZ24, Sch82, Eri16]. Explosion-Time [Ano86g, FY82]. Explosions [HD80]. Explosive [DF89]. Explosivity [RJP97]. Exponent [Ber92, KS72]. Exponential [AI00, AR72, Ahs76, Ahs78b, AT06, Ano82g, Ano90f, Ano90h, Ano98f, Arn75, Bac86, BCP00, BW23, Bau09, BM09, BS76, BJ79, Bha68b, Bha80, Bon89, BD94b, Cal74, Cha93, CDL88, CMY98, CS97, Dal81a, De 85a, DK90, DGM09, FI87, Fei94, FL97, GH94, Gla79a, GS19, Gol78b, GKJ80, GS95, GP04b, GMW90, HC07, HZ22a, Hea67, HS94, Her94, Her99, Hua74a, HS93, IM10, Iva07, Jac78, JS98, Kab12, Kan23, Kei78, KS01b, KM90, LM12, Lan87, LS87, LR90, LR92, LL77, Law80, LM89, Leh86, Leh88, Let94, Lit94, Lit01, Mac01, MZ04, MO67, McK81, Mit03, MT02, NSS04, NC72, NRS85a, NRS85b, OS86, Pfe82a, Phi81, Pin80, Pin11, Pra82, Ros12, RM88, Roy89, SST18, Sen82b, ST12, Sim90, SG88, SS03]. Exponential [Tav80b, Teu68a, Teu68b, TH72, TT79, Twe81, VS90, Van91a, Van91b, VS91, VT75, VT76, Wan12, WP80, Wes81, Wil96, Wit93, Xu98, Yeo74, Yu09, Bro15, DN13, KY17b, Kim16, KS20, Kun22, LR13, LY22, LY23, MPS15, NR22, Per23, Ros14a, SW20, Yor92]. Exponentiality [GR05, GW11]. exponentially [DN13]. Exponentials [BLSW10, WH04]. Exponentielles [Yor92, Let94]. Exponents [How00a, Kub24]. Exposure [BB05, Ish74]. Exposure-Dependent [BB05]. Expression [Lag05]. Expressions [Dur88, ABI14]. Extended [BPS12, Che97, CCN10, Cse79, Kry72, MN82, PP70, Tan78, Van91a, VX04, AMV21]. Extending [McK81]. Extension [Ano71e, Bar64, Bha68a, Bre75, Cha03, CI72, DT15, Fre07, GR05, Hen73a, Hey68a, KA16, Mou99, RS66, Sha77, WW76, Yas93, MN16, NFS20]. Extensions [Ben98b, Dio78, FM12, Gne98, HB99, Let78, Miy95, Sha74b, DM19, Sta20]. Extent [LP94a]. External [BNL05, Mer80, KP23]. Extinction [Agr75, Ald95, BG79b, Bro86, Bro66, Bru84, CP04, DN80b, Est76, FG99, Fea76, Fea80, Fox21, GMSB10, GL83, Hau12, Hei09, Hit86, Höp85a, Hul82, Hul84, Hul01, Jir76, Käm91, KL89a, LP21, Len81, Li06a, Pol71, PZC07, Qui76, Raf73, RR88, Shi06, Slu77, Sud05, Tok08, Tor79, Wan76b, Wan99b, Wau72, Wil69, XW05, ZJ04, vDZ05, BP16, BH20, CLCZ14, Win15]. Extraneous [PL81]. Extrema [JP22, LM24]. Extremal [AC94a, Alp89, AA90, BR87a, Fer06, GJ78, HM97, Kin04, Pfe89, Pic71, Res73, RR94, RSY90, RS92, Seg06, SS87, Sho75, Smi92, STC93, TO92, Tur06, Wei75c, Wei82, WC95, UY23]. Extreme [AFGP91, And70, AG16, Ano86g, Ass83, BD90, BT92, BRPS82, BR77b, CF11a, DN04, DV87, EMR71, FY82, GMW90, Hal78, Hüs86a, KS99, Kar76c, KK06, KP93, KSC12, LZ14, MR15, Nag82a, Res71a, San06, Ski68, Voo87, Yun00, Joh24]. Extreme-Point [Ass83]. Extreme-Value [Ano86g, BD90, DV87, FY82]. Extremes [AC94a, Dan82b, DHSK15, Hal79b, HM95, Jun94, KR97, MP92b, McC97, Voo89, Wei82, ZNF10, BMZ23, HE16, KM15b, LX18, Tan17, VKR14]. Extremum [Tit74]. extrinsic [LLT^{+}22]. Extropy [THB23].

F [Mus00, Mak10]. F-S [Mak10]. FAA [Ano04f, GS08, Sen17]. Face [BB88]. Facteurs [Tor69]. Factoid [Pak08]. Factor [Wei00, KMG19, dCGG16]. Factorial [Ber83, Bil73, Eng93]. Factorisation [Lew67]. Factorization [BH98, BKNZ91, FG12, FvLB13, God77, Ken94, Kuz10, LM08, Vid20]. Factorizations [Imh84, GdlI18]. Factors [PS86b, Tor69, Xia99, WG15]. Failed [BP01]. Failing [FP66]. Failure [And84a, Ano83f, Asm09, BLS03, BSW03, BLSW08, BLSW10, Bur09, CM07, CG12, CN80, Ebr01, GE82, Got80, GL84a, Hol07, Ish74, LT76, Lew64, LM04a, MR03, Nat82, Nat90, PS02a, Pie07, RSZ05, She92, AA17, BB18, BLS15, Cha13, CC15, Dal16b, DN13, GK22, ISK18, KBM21, LC16, Tav20]. Failure-Dependent [Bur09, BB18]. Failure-Prone [LM04a]. Failures [Nád71, Swi08, Yeh97, Lee20]. Fair [DWW11, Fox65, Gal71a, Gre80b, GMDB03]. fake [Mah20]. Families [BF78, BS90, Gla79b, Gla82b, Gri92, LPSV00, MR02, Tes84, Let94]. familles [Let94]. Family [BB84, BS69a, BS69b, Car02, Kuz10, Les89, Lin95, Lin97, Mas86b, MM03b, New70, RY99, Smi82, Yu09, FPKM22, Gau22]. Far [Teh09]. FASSA [Ano04f, GS08, Sen17]. Fast [Ana92, LT23, Rid96, Gén23, MSV14]. Fat [Gay04, IFN19]. Fatal [Sav88b]. Fatigue [BS69a]. Faults [Ano87e, Bel66, JK87]. Favorable [Rut99]. FBM [HJ14]. FCFS [HAPV12, KG02, SS85a, Yao13]. FCFS/LCFS [SS85a]. FCFS/LCFS-P [SS85a]. Features [MCC03, Yor92]. Featuring [BOP07]. Fed [BS02, LM99, LNTZ99]. Feed [KK96b]. Feed-Forward [KK96b]. Feedback [AT74, AN84, BY97, Cor87, KS85, KM88, MWCY01, PJ02, RHK90, TTH91, Wol75]. Feedforward [Kel97]. Feldman [CLZ24, NR18]. Félix [Coh81a]. Feller [AI00, Ben68, CL23, MN13, Phe96a, Sim23, Tu19]. Ferretti [DFSD01]. Ferromagnetic [Pic79]. Ferry [Gre80b]. Fever [Rad74a]. Feynman [Gil68, Jas15]. Fibonacci [Hey80, Hey81]. Fibre [JK81]. Fibrous [KP87]. Field [Adl78, Bar88b, BS73, CM97, Fen97, Gra71, Pic79, Ska71, XY09, YZ85, BDT22, BS14a, Men23, RNJ16, VM21]. Fields [Adl76, ALS04, CJ98, Cox77, CL92, GK97, Has76, KKP10, Ors87a, PF06, Ros86, RMV97, Wu93, Yao93, ZZ99, BSX19, BR18, DHSK15, Erh23, FZ17, FS14, LRM23, LX18, NR22, RNJ17, SX18, SK19, Tan17, Wir18]. Fifo [Li96, Moy10, CN86, RS92, SS87, VG92]. fighter [KSC13]. Figures [Ber77b, Kel83, Kel85]. Filaments [Sen73a, SBG70]. Files [Ban07, LLS84]. Filling [BCJ04, Mul68b, Sie78b]. Filter [Ell93, Pap05, AB15]. Filtered [And79, And81, Mar73a]. Filtering [AR82, Bré78, CG97b, DJ05, Del98, GST00, GT05, Heu87, Hop86b, LWB08]. Filters [Led05, OW94, GW17]. Final [BC95b, BC95a, BO99, BLS03, Dun80, LP94a, ML98, ST86a]. Finance [HS00, KM97, Lin04]. Financial [BDM12, Che11, Gao04, PR96, Sen04, BL13, CH13, HBHH23, MPS15, SWW21, YL16]. FIND [Pau97]. Finding [BG06, RR77]. Finetti [KP07, WZ18]. Fingerprinting [Gri98]. Fini [PLC03]. Finite [AHN90, AML82, AML84, AS04, ACVZ20, Ary72, AFG11, AB82, Att87, Att89, BKL03, BCL98, Bas70, BC67, BL89, Ben68, BvdHvL17, Blo73, Blo91, Bra86, BG00, Buc94, CG12, Che11, Coh71, CS80, Coh82, Coh89b, Cse91, DS65, DS67, DM67, DM68, DJ95, DGM04, Eks04, Emm72, Fad74, GW05, Gil76, GNN85, Gor14, GPZ05, HTW73, HD71, Her85, Hil67, Hok75, Hok77, Hol81a, HS12, IK04, Jac99, JP69, Kap71, KK97, Kha70, Kla94, Kol06, Kol07, Led97, Led05, LWB08, LBL05, MM03a, MW96, McK88, Mil92, Mou93, MS05b, Nak74, NSS04, Neu68b, New86, NR78, Nue08, OW94, Osb82, PRX03, Pen96a, Pen96b, dNPG96, Pha69, PLC03, Roe70a, Roe70b, RR12, Ros86, Ros73, Ros75, Ros76b, RS89c, SM05, Sch68, Sen01, Ser90, STC93, Spo01]. Finite [Sri74, Tak67, Tam86, Tan05, Tet98, Teu72, Tom72, Wak91, Woo83b, Yeh85, YH87, Yeo74, Yeo75, Zuc77, Zuc81, vD89, ASB21, BGM15, Ber13b, Cin23, CM16b, GS15, HM17, KS23b, LS19, Leh17, Mar16b, Neu23, PD20]. Finite-Buffer [AS04]. Finite-Capacity [DGM04, GNN85, New86, NR78, STC93, Woo83b]. Finite-Dimensional [OW94, Gor14]. Finite-Horizon [RR12]. Finite-Parameter [Ros86]. Finite-pool [BvdHvL17]. Finite-Size [Spo01]. Finite-State [Bra86]. finite-support [PD20]. Finite-Time [Che11, IK04, MW96, Tan05, ACVZ20, BGM15]. finite-velocity [Cin23]. finitely [FVZ19]. Finiteness [Gha86, Gha89, AW21]. Fires [Lan11, CMV23]. Firing [CG69, LS74]. Firms [Guo02]. First [Ano82h, Ano87g, Ano88h, Ano91i, AS12, Arn72, ASS85, Bel71, BWY11, BH12, BH91, Bre08, BGNR90, CR72, CW93, CN80, CK81, DIKM10, Di 98, Dom95, Dom96, Dur85, Dur88, DW92, Fak86, Fak87, Fel91, FR10, FL98, Gal93, GP91, Gut75, GRRT97, HP98a, Ham05, Her85, HG80, How00a, How01, HK09, IK11, JRR95, Jan81, JP93, JGR91, Kao74, Kar83, Ken78, Kij87, KKS15, Lal86, Lam68, Law70a, Law80, Law92, Lef97, Lef06, Leh71, Leh73, LN19, LS97a, ML98, MS91, Mou99, MKN11, NRS85a, NRS85b, NRS85c, Nue10, Pau87, PSZ04, PT81, Pha69, PT11, RS69, Reh79, RS83, RSS84, RS88, Rog00, Sat77, Sch11, SS88b, SS76, Sim90, Sim93, SWZ18, Ste85, Ste81, Sum87, SR88, TR67, Tho75a, Tho11, Tuc76, TW84, VB02]. First [Vid19, Was69, Whi71a, Wic79, Wie80, Woo82, WG15, Yao85, ZT01, Zip11, Bel23, CdLH^{+}23, DS13, GA22, GLS19, KK24, LT13, Lef24, SW20, Vee15, Lef95]. First-Come-First-Served [VB02]. First-Crossing-Time [BGNR90]. First-Emptiness [Leh73]. First-Exit [Tuc76]. First-Hit [Fel91]. First-Order [Ano82h, GP91, HG80, JP93, Law70a, Law80, Law92, Leh73, PT81, RS69, Sim90, Sim93, ZT01]. First-Passage [Ano87g, Ano88h, ASS85, CW93, CK81, Dom95, Dur85, Dur88, DW92, Gal93, Her85, How00a, How01, Jan81, Lal86, LS97a, NRS85c, Pau87, Reh79, Rog00, Sat77, SS88b, SS76, Ste85, Ste81, Sum87, Tho75a, TW84, Wic79, Wie80, Woo82, SWZ18, CdLH^{+}23, GA22, Lef95]. first-passage-place [Lef24]. First-Passage-Time [Ano91i, Di 98, GRRT97, JRR95, JGR91, Kij87, MS91, NRS85a, NRS85b, RS83, RSS84, RS88, Yao85, KK24]. First-Served [FR10]. Fischer [vD15]. Fisher [AES17, CR92, GC74, Kle88, Man13b, Web86a]. Fishing [Kar09]. Fit [Hal89, LM12, Vil07]. fitness [CF22b, FL20]. fitnesses [Jor18]. Fitting [OS86, Pap94, ST01, Sen04]. Fix [EW00]. Fix-Points [EW00]. Fixation [Hol76, Kle88, Möh97, PW07a]. Fixed [Adl76, AM07, Cow81, Fea81, HW02, Käm91, KT01a, Lit79, McN68, Möh94, NP09, San80, Sis70, Tsa94, Tsa96, Tsa99, Yah75, HM15]. Fixed-Cycle [Cow81, McN68]. Fixed-Point [AM07, Fea81]. Fixed-Strike [HW02]. Flag [IK70]. flash [FPP19]. Flat [SB78]. Flats [Sch85]. Fleming [CS20a, AFG11]. Flexible [Cai11, PV12, SW05]. flight [KKS15]. flipping [DIÖ23, MZ16]. Floating [HW02]. Flooding [MS20]. Flow [AR05b, Ana88, AI94, Bri71, Cow80, Daw73, Daw74, Gav69, Hut90, KS10, MO70, PJ02, Phe96a, Phe96b, Sta99]. Flow-Stopping [Gav69]. Flowing [Gar74]. Flows [BM82, Dar89, GMRY07, Wal82, HM14]. Flowshop [FA85]. Flowtime [Web82, WVW86]. Fluctuation [Coh73, Coh83, DW72, Zho04, Sah16]. Fluctuations [Hal83a, Rub64, Tak64b, Tak70, VM21]. Fluid [AR05b, AS04, Ano95e, APZ11, BS02, Che96, FH10, Gui94, GMDB03, KK96b, KW99, Kel00, KM01, KS02a, KL00, LT04a, Lel05, MT02, PS76b, RS04, Rab06, Rid96, RS94, BCK16, Bar15, DSZ13, GS15, KSBK23]. Focus [BLS12]. Fold [IKS70]. Folding [Lad05]. Föllmer [Des15]. follow [Jia15]. Following [Ano04g, Ste04, RDM19]. Fonctionnelles [Yor92]. Foraging [MM94]. Force [CM97, HH07, Kar76a, KNT10, Tan05]. Forced [Yas83]. Forcing [BL09b, BL15]. Forecasts [DVJ04]. Foreground [AA06, Sch88a]. Foreground-Background [AA06, Sch88a]. Foreign [Win05]. Foreign-Exchange [Win05]. Forest [Lan11, RR77, CMV23]. Fork [FL97, KW89]. Fork-Join [FL97, KW89]. Form [BKT97b, Che80a, CHPT97, Gel91, GJ80, Kel00, Lam68, Mas86a, McK89, MT02, RR73, Ser90, Tan82, MB16, MBM21]. Formal [Miy79]. Formation [BR79, Bor04, CC94, Fox65, Gal71b, Hel73, HMM93, NA04]. Formed [Bli89, Cow04, LA98, SY72, Tan83b, Tan83c]. Forming [Ken78, KG68, Kha70, Leh71, TP74]. Forms [Bou95, Eat66, Rod95b, Pit18, Tu19]. Formula [Aas77, AL00, ADLN82, Ano78e, Ant78a, Ber85, BL12, Bré93, BGP80, CL93, CN86, Cow91, DY98, FS98, Han90, HA96, IK04, JS96, JZG11, Joy95, Kel12, Ken92, KZ95, Kon96, Koo04, Las94c, Las94a, Lou84, MT06, Miy95, Miy01, Nag01, Saw83, ST73, WW76, DL18, Let78, Mat76, Mat78, Tsu17, Tsu18, Vee15]. Formulae [BVJ00, JM86, Mil79, Whi01, BC23, Jas15, KY17a, Mar15]. Formulas [Bac86, CI72, Han87, LL82, Mej68, PH88, PR95, RN80]. Formule [Let78, Mat76, Mat78]. Fortes [Dio78, Dio76]. Fortune [Eth04]. Forty [Gan10]. Forty-Six [Gan10]. ’Forward [PR08, CFZ96, Irl06, KK96b, Lam91, Möh94, Rud75, Whi82b, Win89, MV21]. Forwards [Käm89, Möh10]. Foster [ABD01, Fra06, Gre78, Kun22]. Foster-Type [Fra06]. Foundations [Res04]. Founder [Hey82a, Hol66]. Four [BS82a, Kol06, MO23]. Four-Dimensional [Kol06]. Fourier [BD12, DHKR19, Ors89]. Fourth [Dur08a, Gat82b]. Fractal [Dek96, Hey99, Rui08]. Fractals [TW99, XY92]. Fraction [GP98b, Mas82]. Fractional [AG12, BMN09, BM14, BS14b, DYY01, GZ94, GN96, HW79, JL06, Lau80, Lee11a, Lee11b, LHN86, MS99, MST13, Ros76a, Sly07, BGM15, BDM21, Bis23, CF13, Cur22, DM18, Fin13, FKZ13, GOP15, HJL21, HMBL14, HJL19, HL13b, LST19, MV16, MWH22, Sön23, Ver14, WY22, Xu17, BS16b]. Fractions [GP99, Han92]. Fracture [Kim94b]. Fragment [CC94]. Fragmentation [FKM10, BDT22, Rub23, Wat23]. fragmentation-interaction-aggregation [BDT22]. fragmentations [HJL21]. Fragments [BB06, Lit79]. Framework [DGK00, DG06a, Rui08, AN22, ALL21]. Franken [Ano90m]. Fréchet [BJV12, SC02, Tan11, AB18]. Free [BJV12, Bor95, FP05b, GRD^{+}00, JZG11, Kat05, Kij93, Kin65, Peg73, Pyk94, San71, Tan11, The69, UW74, Yuk06, DS21, FH13, FM06, HH23, MZZ16, Ste19]. Freedman [PT22]. Freely [Gar74]. Freeway [McN70]. Freidlin [Gag17]. French [Bie98, CTKR83, Die87, Dio76, Dio78, Hua95, HK94, Jof82, Lac93, Lef89, Lef95, Let78, Let94, Mat74, Mat76, Mat78, PL03, PLC03, Ray70, Taï96, Tor69, Yor92]. Frequencies [Cow91, Rad73a]. Frequency [Gar74, Han73b, Saw77, VJ82, CS20a, CH13, LLZ15, PC16]. frequency-dependent [CS20a]. Freshness [LM04b]. Friedman [Sah16]. Fringe [FM10]. Frobenius [Sag20, War99]. Frog [FMS04, GR18, Ros18]. frogs [HK22]. Front [Ano64d, Ano64e, Ano65f, Ano65g, Ano66f, Ano66g, Ano67f, Ano67g, Ano67h, Ano68e, Ano68f, Ano68g, Ano69d, Ano69e, Ano69f, Ano70f, Ano70g, Ano70h, Ano71f, Ano71g, Ano71h, Ano71i, Ano72h, Ano72i, Ano72j, Ano72k, Ano73h, Ano73i, Ano73j, Ano73k, Ano74i, Ano74j, Ano74k, Ano74l, Ano75h, Ano75i, Ano75j, Ano75k, Ano76e, Ano76f, Ano76g, Ano76h, Ano77h, Ano77i, Ano77j, Ano77k, Ano78i, Ano78j, Ano78k, Ano78l, Ano79g, Ano79h, Ano79i, Ano79j, Ano80f, Ano80g, Ano80h, Ano80i, Ano81g, Ano81h, Ano81i, Ano81j, Ano82n, Ano82j, Ano82k, Ano82l, Ano82m, Ano83g, Ano83h, Ano83i, Ano83j, Ano84f, Ano84g, Ano84h, Ano84i, Ano85h, Ano85i, Ano85j, Ano85k, Ano86o, Ano86k, Ano86l, Ano86m, Ano86n, Ano87j, Ano87k, Ano87l, Ano87m]. Front [Ano88m, Ano88i, Ano88j, Ano88k, Ano88l, Ano89g, Ano89h, Ano89i, Ano89j, Ano90i, Ano90j, Ano90k, Ano90l, Ano91j, Ano91k, Ano91l, Ano91m, Ano92i, Ano92j, Ano92k, Ano92l, Ano93i, Ano93j, Ano93k, Ano93l, Ano94k, Ano94g, Ano94h, Ano94i, Ano94j, Ano95f, Ano95g, Ano95h, Ano95i, Ano96e, Ano96f, Ano96g, Ano96h, Ano97e, Ano97f, Ano97g, Ano97h, Ano98g, Ano98h, Ano98i, Ano98j, Ano99f, Ano99g, Ano99h, Ano99i, Ano00f, Ano00g, Ano00h, Ano00i, Ano01j, Ano01f, Ano01g, Ano01h, Ano01i, Ano02f, Ano02g, Ano02h, Ano02i, Ano03e, Ano03f, Ano03g, Ano03h, Ano04l, Ano04h, Ano04i, Ano04j, Ano04k, Ano05f, Ano05g, Ano05h, Ano05i, Ano06h, Ano06i, Ano06j, Ano06k, Ano07f, Ano07g, Ano07h, Ano07i, Ano08e, Ano08f, Ano08g, Ano08h]. Front [Ano09e, Ano09f, Ano09g, Ano09h, Ano10f, Ano10g, Ano10h, Ano10i, Ano17b, Ano18a, Ano18c, Ano19b, Ano19d, Ano19f, Ano19h, Ano20b, Ano20d, Ano20f, Ano21b, Ano21d, Ano21f, Ano21h, Ano22c, Ano22e, Ano22g, Ano22i, BP04, Pha94, PPD97, PP98, Rod95a, Rod95b, Wie79]. Frost [ST86a, Bal83b, LP05]. Frozen [Lan05]. FRS [Reu82]. Fuel [BC67, Jac99]. Full [BF93, GSV11, Gha90, GW05, Gne96, GM07, KA03, MT06, Pak10, STC93, Tam86, Tam09]. Full-Information [GM07, KA03, MT06, Tam86, Tam09]. Function [And79, Ano65c, Att87, BTH72, BKT98, BM90, BS76, Bhu06, BH91, BC75, BHPZ06, BSS81, CS88, Dia73, DM05a, Ebr11, EW88, FMS04, Gan82, GS72, GM73, GM11, GMD03, GZ94, GL84b, HF91, HT78, Hsi76, HW79, HAPV12, JS04b, KS94b, Ken75b, KMRV00, Kra64, Kup90, Li04, Li05, Lie78, Lin73, Mar73b, McC85, Pon07, PR67, Rie80, Sar97, Sol84, Tak71b, TK84, ZLW06, ZL10, BBvdMP23, Bon13, BFP17, GM18, GPZ17, HL23, Kan23, KBM21, MP14]. Functional [Ath71, BS79a, BIM17, CG87, Dav64, DL96, Don73, Fin82, GY08, GL84b, Han90, Hea82, HW83, Hey69, IJB18, JKM12, KN80, KPW94, Kyp98, Li96, LPP12, McL76, RR75b, SS03, Tai95, Tan82, Wan76c, Yao96, IR20, SB17, Yor92]. Functionals [Bor96, CMY98, CCFR00, Dav74, Dur08a, FPY04, Grü89, GS95, Hoo87, Hör03, Hör05, Jag74a, JT92, KO12, Liu11, Loi05, Pur71, RT90, Yun00, FCvS15, HK15, Kim16, LP18, LRM23, Wiś20]. Functions [Ail68, AT70, Ano05e, BK86b, BK87a, BL89, BF08, BK03, BLS03, Bra75, Bro79, Bro66, BG05, Cam73, CH06, Cla99, Cra88, DN80b, DHT86, Dar66, DM67, DM68, Det96, DM72, DGM04, DM05b, EP95, FPKMP98, FC02, GR98, GMD02, GMD05, Hel67, HIR03, HS78, HK90, Hul82, IA07, JT01, Kap72, KPS04, Ken92, KB87, Kuz10, Lau80, Law79, LHN86, LSC04, Lin08, Lin73, MF01, Meh67, Miy90, Møl96, Mye81, OA12, Pap96, PP71, Ren82, Res71a, Rey68, Ruz10, Sav03, Sgi96, SS91, Sol96, SSV72, Sri86, Ste06, Sur08, Swi76, Tan80b, VSW09, WP80, Wei82, Win79, WR97, WW00, WS04, YA88, Yu11, AES19, Bel23, But19, DFP13, FPKM22, FLM15, Fin13, HJL21, KS21, Leh21, LHT17, Mar15, MC18, Pra82]. Fundamental [BS01a, CY08, LN94, Mod66, Mod67b, Tay99]. Further [Ano75f, BS01a, Bar75b, BV11, Boy70, Cha03, Cow80, GL94, GW82, Hai92, HB75, HKN89, Llo89, MWH22, MW90b, NW07, Nak79b, Tan83b, Bar72b]. fusion [DPR19]. Future [Est76, Ken88]. Fuzziness [HKO94]. Fuzzy [Gue99].

G [Ano84e, AF83, AT96, AK97, ALRW99, BT73, BS76, BKNZ91, BNW13, BY97, BPSZ10, CP01, Cha94, CY90, Coh67a, Coh71, Con94, CT76, CN86, Dal65, DT80, Dos85, Enn69, Fak82, Fak83, FF91, Fal94, FT80, FG84, FG98, FR10, Gha86, Gha89, Gha90, GW05, Gil76, GU92, Gre89, GW11, HL76, HF70, Hey64, Hey74, HKN89, Hok77, Hoo83, Hoo70, HL04, Huf87, Hun86, HAPV12, IJB18, Ino19, IN86, JS96, Jen66, Jin88, JW90, Kat12, KS86b, KR88, KS94a, KZB05, Ken72b, Ken74, KKK07, KM93, LLPC94, Lem74, LZ94, LO95, LSC97, LKT90, LS96, LH06, LH02, LT71, LT04b, Lou78, LN94, MH76a, MT00, MNS01, Min79, NG95, Nar68, Neu68a, Neu94a, Ott79, Ott84a, Ott84b, Pak70, Pak73c, Pou90, RHK90, Reu82, Ros87a, Ros73, Ros70b]. G [Ros74, RS99, Rou96, Sch88a, Ses96, SS85a, SGC16, Sta85, SZ93, Sze86, Tak92b, TTH93, Tan79a, Tan94, VD69, Wak91, Wan99a, Whi72a, Whi72b, Whi80a, Wie89, Wil88c, Yao13, vHS86, Cha94, Coh67a, HKN89, Hok75, IJB18, KM93, LZ94, Yao13, ZAFP06]. G. [Bin92]. G/G/ [IJB18]. G/G/1 [Cha94, HKN89, KM93, LZ94]. G/G/s [Yao13]. G/GI/c [ZAFP06]. G/M/1 [Coh67a]. G/M/m [Hok75]. Gain [DVJ04, MT06]. Galambos [Wal81, MS85]. Galambos-Type [MS85]. Galton [Coh82, Hey81, ABD17, Ano92g, Ath71, Ath12, AS16, Att19, Bag82, Bow18, Bru84, Bui75, CB91, DHT86, Dal16a, Dar70, Est75, FS23, Fea76, Fea81, FG76, GR87, GR89, GK96, Gei99, GM96, GM97, Gre78, Hey70, Hey71, Höp82, Höp85a, Höp85b, Höp86a, Hul82, Hul84, Jag74b, Jir76, JW82, JW85, KS86a, Ken75a, Kle89, KS02b, Le14, Li06b, Lin72, Loc82, LMW98, Lyn00, MM03a, MY84, MMR02, MS04b, NW07, Nan82, O’B80, Pak74, Qui70, Red03, RVJ68, Sen67, ST02, Tok08, Via94]. Gambler [Har09, Sch05, Bar64]. Gambling [GKPS06, Yao07b, YC08]. Game [AM07, Ano78h, BV11, Che76a, CH10, Cow92, CYF96, EW88, HR00, HG81, Kar77, MU00, MC90, Nak86b, Ort08, Ros71, Rot98, Sub81, ZC95, Zen10, CDP19, CM16b, Luo14, MN13, Tav22, Gut10]. Games [Bro00a, CCH06, CH06, CC08, Dix93, Got85, GHL03, GHL05, Hai96, Kam79, KK69, LS92a, Mam87, Ork72, Pon07, Res66, Sch75b, Sch76, Sec97, Sie89, SS01, Sob73, CF17, EGL17, FPS18, GR21, Mar16b]. Gamma [AD94, Ano80e, Ano06g, Bha80, Çin80, EZ80, FS06, FS07, FSW12, Hen83b, Lan86b, LS79, McK82, Pha88, SM78, Sen04, Sim90, Tol88, VU09, Yan88b, AMV21, BDIM18, MC13]. Gamma-like [LS79, BDIM18]. Gamma-Type [Lan86b]. Gammas [BLSW10]. Gani [Sen17, Bha68a, Boy87, Bra81, Dav64, Pur87]. Gap [Haw68, Mei99, Mit04b, The79, FHI17, HL16b]. Gap-Acceptance [Haw68]. Gapped [Spo04]. Gaps [GH09, Ito80]. GARCH [BPSS14, DN19, HC07, KLM04]. GARCH-type [DN19]. Gas [Ste86]. Gases [BMT10]. Gated [BY97]. Gaussian [Adl78, Ano88h, AA23, ALO05, Bal95, BD75, BNCPW09, BF78, BZ04, BC01a, Ber70, Ber85, BL11, Bor21, Bre81, CW81, Cab91, Cai03, CD81, Dav73, De 77, DM03, DHSK15, DHJ15, Deh85, DJ00, DDS22b, Deo73, Dri75, Dur85, Edd80, Emb83, Epl83, FCP20, FI98, GA93, Gao04, GSV11, GL76, Has76, HH07, HZ92, HM99, Hoo13, JP22, JY21, KA16, KP93, KY17c, LRM23, LX18, LR24, Mal67, Mas77, MQ00, Mor73b, Mor86, NR22, Ors89, Oza86, OI01, Ryc87, SWW21, Sel91, SS76, THP12, Tan17, Tan79c, Wan77a, Wol75]. Gaussianity [Pap71]. Gem [Zei98]. gender [FS16]. Gene [Hey82a, JP69, Pol66, Man13a]. Genealogical [Don86, Mal74]. Genealogy [Kin82, LML06, GH16]. General [Ale01, All10, And87, AF87, Ano75g, Ano88h, Ano95e, AB86a, AJ00, ALL21, Bad11, Bal79, BO99, Bel71, Bil73, Bil75, Bil76, Bil77a, BM94, Boo89, BC75, Bro86, CP04, Che81a, CR94b, CT72, ÇP72, Cla99, Dal65, Dal81b, Dan82a, DG93b, Don73, Don76b, Dow72a, Dow72b, Dun80, Dur85, Dur71b, Dur74, Ell93, Fad83, FG87, Fin07, Fra06, Gal93, GJ72, GR98, GESK95, GKL01, Gly89, GL84a, Gre77a, Gui94, GM94, Haj82, HZ87, Has64, HA96, HN76, HB83a, Jag74a, JS00, Kap71, Kij89, KN01, Kry75, Kuu82, KZ84, KP07, KZ09, Kyr99, LDL95, Leh73, LS07, LT95, LB88, Lud73, MT95, ML86, Min78, MW90b, Miy01, MJP87, Mor69, MR01, Nat75, Nat77, Olo96, Olo98, PJ02, Pic82a, PZC07, PH86, Raj81, RR73, RC80, Sag99]. General [Sch72, Sch92b, SS83a, SY89, She96, Sko00, Sri74, SZ91, Sta01, Ste03, SC73, TL69, Twe74, WW06a, WP97, WZ18, War99, WF80, ZNF10, BIMR22, CS20a, DM19, FS17, HK15, IR20, IRS23, LS19, LC16, LP18, LVZ19, MS20, MBM21, She13, VKR14, WBS16, Yam17, ZZ16]. Generalisation [Mej68, Taï96]. Generalised [Ano74h, Che02, Con74, CT95, FW91, GM73, GG95, HM20, Nea08a, NMG00, Rob92, RM88, Rub64, DQL19]. Generalization [Ano80e, Çin80, Dal71, Det94, Fei94, FS98, Kap77, KS86c, Lan79a, McK89, MO71, OM72, RF88, Wil96, DDS20, EDSM15, Lin16, Pyc21, Taï96]. Generalizations [Ber77c, BLL80b, CF11a, CMV23, Miy95, Reh77]. Generalized [AL00, Ano73f, AG07, BFL05, Bar82a, Bha91, Bil81, BK00, BDMZ12, CW97, Cha01, Cha03, Che96, Con75, CG00b, DIKM10, DZ03, DFSD01, DM10a, DIMZ13, Don76a, Dos79, DG06a, Emb83, Fan12, FKK91, FL20, GWMO79, GS06, GP04b, Gut10, HD71, Han92, Her94, Her99, HMPS06, HK90, Hu16, Ito89, JK76, Kap75b, Kel12, KSC13, Lel05, LZ94, Li09, LT71, LT72, Mar00, MO67, MM82, NS22, OS70, OS07, PW81, Rat70, Ren82, Sav03, SA73, Smy11, Spe94, Tal80, VOZ95, VK10, Woo83b, XT11, XS03, Zac04, ASB21, ABS15, ABSGS16, BSC18, CK13, FHZ15, FZ17, GXY18, HLVM22, KM23, KS20, LGC16, MN13, PL17, RWZ13, Sei16, TTI22]. Generalizing [KM03]. Generally [CT95]. Generated [BL03, Cra11, Deh85, EE81, EE88, Ery10, Gra72, IS83, Ito80, IMT00, Nue10, Smi73, Ste03, Wei75c, FGH^{+}21]. Generating [BF08, Bra75, BCIM01, Bro66, DN80b, DM67, DM68, Det96, Dia73, FC02, Gan82, GZ94, HS78, HL71, HW79, IA07, Kap72, LHN86, Li04, Miy90, PP71, ZL10, But19, GM18, Lei23]. Generation [Ano04g, BG87a, Can73, Fea76, Fea80, Fea81, QW84, SG89, Ste04]. Generation-Dependent [BG87a, Fea76]. Generations [Sam71, BIMR22, IRS23]. Generator [Phe96a]. Generators [BP99b, JK71, Law92]. Generic [CM02, BBvdMP23, Leh14]. Genesis [BS81]. Genetic [Ano78e, Ano04g, Ant78a, AM78, Can73, Can74, CW69, Daw81, Dia73, Fei93, Fel66, Hol68, Kar68a, Kar68b, Mar69, MSS09, Pol68, Ste04, Suz10, Wat64, WK65]. Genetics [Aas76, CS85b, Don86, Eth79, Ewe64, Ewe82, FW07, Hol66, Hol79a, Hop86b, Jen74, Kel76, Kim64, Les10, Lil68, PW07a, MN16]. genomics [FGH^{+}21, Mar16a]. Genotypes [CW69]. Geo [KMRV00]. Geo-Stable [KMRV00]. geodesic [AB18]. Geographical [Bon02, Rad73b, SC81]. geographically [AM18]. Geom [CG96]. Geometric [Aas87, Arn80, BL01, Ber77b, BO90, Cal74, CG90, CCFR00, Dal74, Dal87, Dal10, Dav83, Del78, Fin10, FMRR03, Fra00, FB88, GL67, GMD03, GMM05, GP98a, GL05, HL13a, Hör03, HS93, HLS93, JM86, KN75, Kle84, KS02b, Kol00, KM92, LS87, Led97, Lee11a, LS02a, LL90, Lun97, MSS09, Mil79, Mor73a, Neu80, Pek96, Ris95, RT01, SS80, Sre83, Sri86, Teu72, Wil02, WC04, Wit93, Woo82, Yuk06, Zij83, CdLH^{+}23, Dal16b, ET21, KK18, LLL20, Met13, Owa23]. Geometric-Type [Fin10]. Geometric/Probabilistic [Dav83]. Geometrical [Ano66d, ALO05, Hor65, Mor66a, Sch80b, Sie78b, Swi76]. geometrically [FHL13]. Geometricity [Pha93]. Geometry [Rui08]. Geophysics [RMV97]. George [DVJ08]. Gerber [BPPR16, BFP17]. gerrymandering [AM18]. Geske [FS98]. Geske-Formula [FS98]. Geyser [Boo79, Ste81]. GI [Abr01b, Bäu97, DS94a, DGM04, Miy86, Miy90, Rig99b, Sim97, Sim00, SDH94, ZZ04, ZAFP06, Abr01a, Ano84e, ALRW99, BS76, BKNZ91, CG96, CY90, CK00c, DN80a, Dal65, De 85c, Dos85, Dvu84, Fin69, Gha89, Gha90, GW05, GU92, HL76, He03, Hea65, Hey64, Hoo70, HL04, Jin88, JW90, KS86b, Ken72b, Kyp72a, LT00, Lem74, LO95, LKT90, LS96, LT71, LT73, Lou78, MH76a, Min79, Miy90, NG95, Nar68, Neu81, Pak70, Pak73c, Ram84, RL88, Ros87a, Ros70b, Ros74, Sim97, Sim98, SP92, SZ93, Sze86, Tan79a, VD69, Whi72a, Whi72b, Whi80a, YC01, vHS86]. GI/D/ [Dvu84]. GI/D/1 [Pak70]. GI/E [Fin69, SP92]. GI/G/ [ALRW99, Nar68]. GI/G/1 [BS76, BKNZ91, CY90, Dal65, Dos85, GU92, HL76, Hey64, Hoo70, KS86b, Lem74, LO95, LT71, Lou78, MH76a, Min79, NG95, Pak73c, Ros87a, Ros74, SZ93, Tan79a, VD69, Whi72a]. GI/G/c [Gha89, Gha90, GW05, vHS86]. GI/G/I [Sze86]. GI/G/K [Jin88, JW90, Ros70b]. GI/G/s [Ano84e, Ken72b, Whi72b, Whi80a]. GI/Geom/1/N [CG96]. GI/H [De 85c]. GI/M [Miy90]. GI/M/1 [DN80a, He03, Hea65, Kyp72a, LT00, Ram84, SP92, YC01]. GI/M/1-Type [He03, LT00]. GI/M/1/K [CK00c]. GI/M/1/n [Abr01a, Sim97]. GI/M/c [Sim98]. GI/M/c/n [Sim98]. GI/M/n [HL04]. GI/M/S [LT73]. GI/PH/1 [Neu81, RL88]. Giant [AFP22, BGJ12, BT12a, Ber13a, JKL^{+}20, Pen16]. Gibbs [BN17, BHTLV20, DV20, HJL21, HZ22b, KKP10, WD99]. Gibbsian [GK12, Mas86a]. Gifts [Gal71a]. Gilbert [VE98]. Gillis [CR92, GP05]. Gillis-Domb-Fisher [CR92]. Gini [AES17, BM17]. Ginibre [DFV15]. Giorno [IN86]. Girsanov [Phe96b]. Gittins [FH07]. Given [Ano85f, BT82, Blo82, Bor94, CG97b, CG05, GPZ05, Miz06, Pol94, Rus83, SK86, BM20]. Glaser [BLS12]. Global [AIGR06, Bes72, Loc96]. Globally [Dip98]. Gnedenko [Kal96]. Goettge [Gre78]. Golosov [Fin13]. goods [Mos18]. Goofspiel [Ros71]. Goovaerts [ABSKT19, Rob14]. Gordin [Eag75, Roo76a]. Gorillas [ML77]. Governed [Hul82, KS86c, Ors86, Ors87b, Sri74]. Grade [Ano91f, Dav82a, Gue91, Gue97, VG85]. Gradual [KM93]. Grain [And84a, KS83a, PS16]. Grammars [San71]. Graph [ABE99, BR89, Che01, DD04, DFM^{+}06, FMS04, Ger77, HH23, JLN10, KS10, LP94b, Nea03, PS09, Ros81c, Tak88, WDLC09, YCG11, BM15a, BKK19, BS13a, BS15b, BJMK18, Con13, HP17, KS23b, Man13a, Ste19, VI14, Win15]. Graph-Related [BR89]. graphical [HE16, NvdBOdI24]. graphon [BKK19]. Graphs [AW93, BDLL08, CJV10, Dor82, DGM09, GP05, Ito79, IMT00, KKP10, Kuu82, Mat92b, RS10, Tak89, AFP22, ANT17, ARS21, BvdHK17, BCN24, CF22b, CdLH^{+}23, De 22, DIÖ23, Erh23, Gab16a, GvdH18, HM23, JKL^{+}20, Jor18, KY17b, Kur22, LN19, Lop14, MS20, ÓT21, OC22, Pen18, Sha23, Vid23]. Great [Nea08b]. greedy [Con22, Gab16b]. Green [ML77]. Gromov [DDS22b]. Group [Bin69, Coh76, Fak82, Han65, Mor93, BT23, SH21]. Group-Arrival [Fak82]. Group-Departure [Fak82]. grouped [Mar16a]. groupes [Let94]. Groups [Ano66e, Bec73, Dow66, Let94]. Grow [Sud83, IFN19]. Growing [Gan94, PS19]. grown [LM20]. Growth [Ano92g, BJ79, BK64, Bro72, BCD93, CS82, Dei10, Dur92, Fos75, FG76, GMD03, GMM05, GMM06, Gre81, Gre76c, Gri75b, HP98a, HS72, HS74b, Hit88, JS00, KR70, Kel09, Ker86, Kle83a, Kle84, Kle89, LH81, Len81, LPP12, LS11, Mac80, MSZ04, Mon67, Nie85, Pra80, QR90, Qui91, QL99, Ren74, Ren77, Tok08, Tra81, Tzi11, Wau72, Wau74a, Wau74b, ZWV10, BBBN14, Hös19, Wat23, YW21]. Growth-Catastrophe [Nie85]. Growth-Collapse [Kel09, LS11]. growth-fragmentation [Wat23]. Growth-Rate [Bro72]. GSMP [MY88]. Guessers [Pit80]. Guesses [HZ82]. Guessing [HZ82]. Guided [HZ82, KS23a]. Gumbel [Ano90f, Joh24, Roy89].

H [Bin92, Cri18, De 85c]. H-SII [Cri18]. H. [Hör05, Tan79b, Tan80a, TP79]. Hájek [Ano71e, Hey68a]. Half [Bin05, Bre75, Mul68a]. Half-Century [Bin05]. Half-Space [Mul68a]. Halfin [LY20]. Hall [BN17]. Hamilton [Che03b]. Hamilton-Type [Che03b]. Hamiltonicity [ELNT11].

Hamiltonicity-Trace [ELNT11]. Hammersley [OW79]. Hamming [IS86]. Handling [CW96]. Hankel [Sch80a]. Hannan [Ano86p, Ano94l]. Hanski [MP14]. Haploid [CD85]. Haplotypes [Fea03]. Hard [Gir82, Lyl88, Wes82, Ari14, DDS20, Gaá20, Roy24]. Hard-Core [Wes82, DDS20]. Hardware [Ban07]. harmonic [MS23, Luo00]. Harper [War99]. Harris [Ang11, AP86, GMSB10, Hon13, Kuc82, MY85, Mur80, PK74, TW75]. Hartman [BRY04, Ger11]. Harvesting [Gre84]. Hashing [BP97]. Haskey [Kry72]. Hastings [BJ24, DJ09, Rob99]. Having [BLSW10, CD92, Duf96, LM08, Lin94, PH82, Raa80, ZW17]. Hawkes [BM01b, BNT02, FLM15, KSBM18, LR23, LP23, Zhu13, Zhu14]. Hazard [Bad11, BL01, BM01a, BENP94, FH95, Kij98, KK18, KL89b, Li05, Mac81, NS06, SS88a, BMR19, Bro15, Leh21, LWM20]. hazards [AES17]. Head [Sch75a]. Head-of-the-Line [Sch75a]. Heaps [Don91]. Hearing [Sud85]. Heat [Ors87b]. Heavy [AT96, AK97, Blo73, CT04, Che03a, DMS08, DT11, GT06, GN06, GMS96, HGR90, HT12, Har73b, Har73c, HS10, JTW^{+}13, Jin88, JW90, KM97, Köl74, Köl79, Kro97, Kul06, Kyp72b, LS08a, LPS06, LF08, LS09a, Lou73, McC97, O’B80, OV03, Ott79, PS99, PW96, Res04, SW11, SZ93, Tan07, TV04, Tur00, WW06b, Whi72a, ZG04, AB19, ACVZ20, BvdHvL17, DPW21, GH14, HLVM22, KS21, Leh17, Zha22]. Heavy-Tail [Res04]. Heavy-Tailed [CT04, Che03a, DT11, GN06, HT12, KM97, Kul06, LPS06, LS09a, TV04, WW06b, ACVZ20, BvdHvL17, DPW21, GH14, KS21, Leh17]. Heavy-Traffic [SW11]. Hedging [Cai11, KPT99, SS07, DHKR19, Tar15]. Height [BD94a, FS92, Grü85, Her85, PS09]. Heights [Don80a, Tak92a, Ber13a]. Heine [Kem92]. Her [Van06]. Heritability [Wil82d]. Hermite [Bai20, IS83, Mar13]. Hessenberg [DKK08]. Heston [AKK20, JS20, dRdC18]. Heterogeneity [AB98, Hou91, Ton97, Cha13]. Heterogeneous [CM82, CF11b, CP92, Con94, Fal94, Fal96, Fin04, GST00, GT05, GT08, KAR11, LN07, LM88, LD10, NE81, NE82, NSB11, PS86b, Rig96, Rig99a, Ros12, Sob90, XMKW90, XB11, Yu74, DFP13, EGL17, HR15, KA22, Ros14a]. Heterogeneous-Server [Yu74]. Heteroskedastic [Lin99, Liu04]. Heterozygotes [Hol76]. Heuristic [LA01, Pol82]. Heuristics [GBW92, PP98, Top86]. Heyde [Ano04f, GS08, Ano04m, MU11, Nas08, Tom71, Won11]. Hidden [Her99, KN01, MLB05, Rui08, ST01, HWX18, HK19, HJL14, SZY19a]. Hide [Sub81]. Hide-Search [Sub81]. Hierarchical [Arn75, BN94, Car99, Pol67]. Hierarchically [SF83]. Hierarchies [Hel73]. High [And84a, Coh89a, DE11, Fer93, Has76, Pap71, Ren88, RV00, Ser80, Sta82, WC95, BO18, CH13, FW16, HS14, KR18]. High-Dimensional [Sta82, FW16, HS14, KR18]. High-Intensity [And84a]. High-Level [Ser80]. High-Order [Ren88]. High-Quantile [DE11]. High-Speed [RV00]. Higher [ALS04, AM05, BD88b, Pap96, ST86b, Wil88a, Wil02, GT14].

Higher-Dimensional [BD88b]. Higher-Order [ALS04, AM05, Pap96, Wil02, GT14]. Highly [ML67, Tuf99]. Highway [Gav69, New66, Rén64]. Hilbert [BKNZ91]. Hilbertian [Gui02]. Hill [Lo86, RS95a]. Histograms [Ken93]. History [Saw77]. Hit [Fel91, BRSC22, RU18]. hit-and-run [BRSC22, RU18]. Hitchcock [GH92, Gan04b]. Hits [Rud76b]. Hitting [Ath88, Ber99, BB80, DGM08, DS93, ER88, ER89, ESE84, KS01b, Lac05, Lad05, Lef97, LS11, Mei99, MT02, RS99, Met13]. HIV [Ber94]. HNBUE [BB84, BK86a]. Hoare [Grü98]. Hog [Sri91]. Hold [YYZ09]. holds [ABSKT19]. Home [DR94b]. Homing [Kuh85]. Homogeneous [Ano91g, Ber68, Car99, Coh88, Dar66, FPY04, Ger90, Ger91, GMM05, GMM06, Gri75a, GRRT97, Har94, Has64, HLT00, How00b, JGR91, Jir76, KZ88, KS94a, KT01b, KT01a, LM88, MSZ94, Paw89, SAI82, Ser90, Ska71, TV88, Tsa94, Tsa96, Tsa99, TV93, Tuc76, Vas81, Vas82b, Vas84, VT89, VGT90, VP92, Wes77b, XS03, Zei89, ALL21, BW19, CHR15, DHSK15, GP13, LLZ17, LTZ13, MZ20, RW21]. Homogeneously [GJ78]. Homozygosity [Kar68b]. Honest [Har90b, SD88]. hop [Hir16, Pri19]. Hope [LM08]. Hopf [Ano64c, FCvS15, FG12, Kei64a, Ken94, KMG19, Kuz10]. Hopfield [Bov99]. Hoppe [HI20, Hol08b, LN13]. Hori [Wei80b]. Horizon [BL11, Eks04, Emm72, IK06, PLC03, RR12, Zhe91, BDL19, Mar16b]. Horse [Hen81]. Horton [Yam17]. Hospital [SV71]. Host [Rad73b, Rad76b, FL20]. Host-Vector [Rad73b, Rad76b]. hosts [BB14]. Hot [Lyl88]. Hotelling [Dav82b]. Hour [New68b]. Hourglass [CT00]. House [Che76a, CCHY08, MS02]. House-Selling [MS02]. Household [Nea05, Nea16]. Hull [AFGP91, BdH87, CH91, Dwy88, Edd80, Mas95, Van89b]. Hulls [Bla92, JR82b, Rog78]. Human [Ben82]. Hunting [Aba79]. Hurst [BGW83, OA12]. Hybrid [DKK08, FMRR03, DSZ13]. Hydrodynamic [FNV11]. Hydrology [Kar76c]. hyper [Con13, SBM22]. hyper-complete [Con13]. hyper-recursive [SBM22]. Hyperbolic [SY72]. Hypercube [CJ08b]. hyperfluctuating [KL22]. Hypergeometric [Che76b, VSW09, Cri16, KPV21]. hypergraphs [BCKK17, Spi22]. Hyperplane [CO85, Sch19]. Hyperplanes [Lan11, Mat74]. Hyperplans [Mat74]. Hypotheses [GGS^{+}95, Hos86]. Hypothesis [Pri00].

I.I.D. [CV09, Deh85, MP92a, Sch82]. Ideal [Bar82b]. Identical [Eat66, EE76, KK96a, Pra82, WP80]. Identically [Ano86a, DN04, Gra85, Wei75c, ZJ04]. Identifiability [ADH93, Gås98, Mor75, Now90, PM99, Ros70b, Roy14, Ryd96, Sol86, Tal69]. Identifiable [Can77]. Identification [GKRV94, HKY96, HZZG86, Iso82, KMS83, Pol94, PH86]. Identities [Dar66, Peg73, Rao79, Whi82b, Zho04, Con22, EDSM15, FvLB13, Hol13, Psa22]. Identity [Bro00b, Coh88, Man69, Mao04, Sch86, CM15a, Iva16, KA16]. Idle [DS98, FY83, LO95, FS17, HR15]. Idle-Period [LO95]. If [CNS99]. IFR [BKS02, BLSW10, FRR01, Sav85]. IFRA [CDD91]. Ignatov [ETV88]. ignition [CMV23]. II [Ban07, Bar72b, BK87a, BF08, Ber81b, BP96a, Cra74b, DJ69, KN80, Kar68b, Kei65, Köl79, Mul73, New68a, Pic84, Roe70b, Sch76, Spe73a, VT76]. III [Bar74, New68b, Pic79]. Image [Dek96, Win90]. Imbedded [Bac91, Jen66, KS80a]. Imbedding [BPS80]. Immediate [BBMB08]. Immigration [Abr94, Ano75g, APS74, BRS75, CY84, Coh77b, DD68, Dur74, Eco04, Gre76c, HS72, HS74b, HL99, Hon02, Höp82, Iva80, IS85, Kap74, Kap75a, KP83, Kyr99, Lai77, Li00b, Li06b, MW10, MY84, MY85, Olo96, Pak71, PK74, Pak74, Pak76, Pru85, Qui70, Rad72a, Rad73a, Rah98, Rey70, ST83, Wei72, Wei74, Yan72, Zha04, Zha05, DI20, DLVZ20, HL16a, HMY17, RW21, Rah17]. Immigration-Branching [Lai77]. Immunity [MD96]. Immunization [Tzi11]. Impact [And79, DS08a, FG87, JTW^{+}13]. Impatience [BPSZ10, Dal65, Moy10]. Impatient [BE91, DT85, GADL04]. Imperfect [BP83, CCH06, CC08, DG06a, Hu96, LS03, MWCY01, SG92, MC13]. Imperfectly [Wak91]. Implementation [CST08]. Implications [AI00, CL93, God77, Lew64, NSK08, Wen86]. Implicit [Kev17, JOC13]. Implied [Dur08b, Teh09, EM24]. Implies [BJ79]. Imply [Ano10e, Mar07, Yu11]. Importance [Asm09, BCP00, BL10, KB87, Nor86, Ery15, GKS14, Kim16, Nyq17, PB20, Spi15]. Improve [DJ09]. Improved [BE87, Cer88, CL14, Cow78, FF03, GN17, MCC03, Mus00, Ros16, Tan11, BJV12]. Improvement [Aut75]. Improvements [BP96b, GM20]. Improving [BR85, GR12, Sch05]. Impulse [Cos93, Gag17]. IMRL [AA06]. in-degree [SRT^{+}16]. in-degrees [HJ13a]. Inaccessible [Mor71b]. Inaccuracy [Ano73f, AS75, SA73]. Inactivity [GAB12, Zha10, Gup13, PB14]. Inanimate [Ano74f, Mor74]. INAR [AAO90, IPV03]. Inbreeding [ST88]. incidence [MP14]. Incidents [Gav69]. Including [Khm11, Hen18]. Inclusion [Sud02]. Inclusion-Exclusion [Sud02]. Income [Kei82, ZC14]. Incomplete [BE83, GM73, Moy65, Mos18, Tar15]. Incompressible [PS76b]. inconsistent [LHT17]. Incorporating [Cla05, Nea16]. Incorrect [Oga80]. increase [vdHKPS23]. Increasing [Bal95, BSW03, CV01, DM10a, DY97, GE82, Git81, Gne99, KÖ89, Mak94, Mor78, PL17, Ros76b, RSZ05, SY86, Son96, Sys94, Wau74b, Yan75, CS23, GK22]. Increasing-Failure-Rate [RSZ05]. Increment [Ano06g, CA98, FS06, IPR11, Ros72]. Increment-Vector [CA98]. Increments [Ano82g, BNCPW09, Ber85, Bin69, CF09, DLU09, FK89, FKK91, GSV11, Gla92, JKP11, KS94b, KS86c, Pfe82a, RR75b, Ser72b, WW06b, AF14, DPW21]. Independence [Blo86, Gal72, LT00, MRR02, Sri86, BJ24, FPW19, Sei16]. Independent [AARJ10, And85, Ano75e, Ano82g, Ano90h, Ano91e, Bad11, BdHK93, Bas71, Bin69, BPT92, Bor04, BR77b, BP00, BL03, CF09, CP88, DN04, DLU09, Dri75, FI98, GS99, IPR11, Ish75, KK97, Lam91, LM89, LM99, Meh77b, NP91, Oak79, Pfe82a, Pie07, Pol75, RR75b, RH81b, Ros72, Ros12, Ser72b, SSSL10, Sha88, Slu77, SKK06, SR70a, TR65, VG92, VS90, Vas82b, Wei75c, Wei95, Win05, Wit96, XH11, ZLB08, ZJ04, AFP22, AI16, BS15a, Con22, FCP18, GN17]. Independently [Cra74b, Leh21]. Independently-Dispatched [Cra74b]. Indeterminate [LS09c, Sto04]. Index [Ano89e, Ano13, Ano14, Ano15, Dal01a, FW91, FH11, Fer06, GG95, HM97, OA12, Pau01, Seg06, Smi92, TO92, Tur06, VPT92, WW90, WC95, BM17]. Indexed [TW99, GZ16]. Indication [Kha68]. Indicator [CH97, Eng93]. Indices [FH07, FH13]. indifference [Kli14]. Indirect [BPT89, Bol01, Cla05]. Individual [CI72, MP12, KS18]. Individually [KW85]. Individuals [BFM08, Cer88, DDM98, GS76, Jag69, Mer80, Sag20]. Induced [Imo01, Li00a, MR01, Mye81, Nag82b, BKK19]. Inducing [Par01]. Inductive [Bas10]. inegro [KP22]. inegro-differential [KP22]. Inequalities [Abr01a, Agn72, Agn75, Ano67e, ASC00, Bro79, CS00, CZ99, DF81b, FP97, Fra00, GL94, Gol74, Gol75, GP98a, Gri98, HS66, HR87, HL05, Huf87, JR16, KS80b, MSCZ78, MS85, Mór96, O’B75, Pek99, PC85, SPS80, Sch96, Sch85, Spe94, Sun92, Tur72, VK85, Van96, Wal81, Wit96, XL98, XH11, KW14, KL17, LR24]. Inequality [Ano71e, Dek96, Dur08a, FF09, Fin88, Hey68a, Kal91b, KK97, Lee11a, Nis75, Pek06, Tom71, Ton85, Wil96, YI99, dSS04, AA23, Gol23, KLP19, Rog23, Zho15]. Infection [AB98, Ber94, Boo89, CM82, Cla99, Cla05, DD04, FS80a, GM91, Ren81, Sap87, Sud01, ZG88]. Infections [Kes08]. Infectious [FG99, Sta01]. Infective [BC95b]. Infectives [Abr94, Dow72a, Dow72b]. Infectivities [Nea07]. Infer [DS92]. Inference [Ano82i, BH75b, BSB99, CH13, EG02, Fin64, HT71, IPV03, Pic76, Pic79, Pic82a, Pru85, Toy97, Wan93, Wil82a, BKM19, KJS15, Rah17]. Infimum [Zho07]. Infinite [AS04, AP94, And87, Ano79f, Ano06e, Avn79, BR77a, BS79a, BM97, BM89b, Ber11, BL11, BGSS03, BR69, BS93, Che01, Coh71, CP78, CS80, Con71, Dal82, Dar70, Dar89, Dav78, DO12, DP12a, Dsh89, Ell66, FW07, GS87, Gre77b, Gri81, GS79, Had69, Ham82, HK77, HL81a, Her72, Hey91, Hey95, HS77, HB02, IK06, Iks04, JZG11, JW11, Joy94, KG68, Kha77, KO05, KSBM18, Leh71, Llo71, MS04a, MY01, Mol86, NC72, OL65, Pak04, Pak07, PL81, PG79, RN80, SCZ77, Sen73b, Sgi96, Sha66, SS96, TP74, Zhe91, AB15, AR15, BN17, BJT20, BvdHK17, BS15d, BDL19, BCdlP17, CGR08, CHR15, CR17, DV20, EOV20, Gab16a, MK15, MK17, SDH16, Thö16, Zho23]. Infinite- [AS04]. infinite-allele [MK15, MK17]. Infinite-Alleles [Avn79]. infinite-bin [CR17]. Infinite-Dimensional [Dar89, FW07, Zho23]. Infinite-Horizon [IK06, Zhe91]. infinite-order [BN17]. Infinite-Server [BM89b, BS93, HL81a, PL81, RN80, KSBM18]. Infinite-Sites [Gri81]. Infinitely [Ano77g, Ano87h, DKO84, Eth92a, Joe96, Lan86a, Mil67, Nel77, Pha79, Sch91b, SM97, VT11, Wat76, RNJ16, RNJ17, SRN21, Zho15]. Infinitely-Many [Ano77g, Wat76]. infinitely-many-alleles [Zho15]. Infinitely-Many-Neutral-Alleles [Eth92a]. Infinitesimal [BP99b, Twe73]. Infinity [CK00c, SVJ66, Sim97, HZ22a, LP21, MSV14, SF17, Sag23]. Inflation [CSZ04, CSYZ05]. Inflow [Ben70]. Inflow-Stock-Outflow [Ben70]. Inflows [KM01, Llo71]. Influence [Phi01, Pur75b]. Information [Alp07, Ano64f, Ano65h, Ano66h, Ano67i, Ano68h, Ano69g, Ano70i, Ano71j, Ano72l, Ano73f, Ano73m, Ano74m, Ano75l, Ano76i, Ano77l, Ano78m, Ano79k, Ano80j, Ano81k, Ano82p, Ano83k, Ano84j, Ano85l, Ano86q, Ano87n, Ano88o, Ano89k, Ano90n, Ano91n, Ano92m, Ano93m, Ano94n, Ano94o, Ano95j, Ano96i, Ano97i, Ano98k, Ano99j, Ano00k, Ano01l, Ano01m, Ano02j, Ano03i, Ano04n, Ano05j, Ano06l, Ano06m, Ano07j, Ano07k, Ano08i, Ano08j, AEHS05, Aut75, Bal75, BN94, BEN98c, BS79b, BF93, DVJ04, Daw73, Daw74, DS75, Fin64, Gne96, GM07, Hai81, Hol79a, Joh24, Kab81, KN80, KD78, Kot66, KA03, MT06, MK73, MWCY01, Mon82, Nat90, Nis74a, Sch75b, Sch76, SA73, SAG73, Tam86, Tam09, Ton75, BNW13, EG15, EV16, VI14]. Information-Based [Nat90]. Information-theoretic [Joh24]. information-theoretically [VI14]. informed [Tar15]. Ingersoll [Zhu14]. Inheritance [BCS23, CH84a, HB83a]. inhomogeneity [Sha23]. Inhomogeneous [AB19, BGJ12, GJQ06, Hou12, Meh77b, MO70, Min78, Pen18, NNY20]. Initial [BN06, BLS03, CGR08, Dio74, Rad73b, Sco82, BCK16, GW17, Mar16a, SK23]. injection [NPYY18]. Injections [ES11]. Innovation [OI01, CdOR16]. Innovations [ZG04]. Input [AH00, Att87, BBK08, Bha68b, Blo91, BC75, BPS80, Cha94, DN99, DS04, DYY01, FH10, GAS03, Has82a, Has64, Has65, Hun86, Ish04, KP89, KK96b, KW92, KS02a, KM93, Lan86b, LA98, Leh72, Leh73, Lem81, Lem89, Lun94, Lun96, MS99, Meh77b, Min78, Mor69, Nat77, Neu66, OS91, Pea65, Pha79, Pur75a, Sah71a, Sch75a, Sha88, Sri74, Vas82b, Wak91, YH87, vD89, DSZ13, HJ14, KB13, KSBM18]. Input-Output [Cha94]. Inputs [BKL03, BW89, ÇP72, Col75, DM03, Her72, Her74, Kel00, Ken78, KG68, Kha70, Kha77, Leh71, LBL05, Miy77, OL65, Ott84a, Pak73b, PP77, Pha69, PM73, TP74, TP76, Wil97, Wu68]. Insensitive [HR88b, Van89a]. Insensitivity [Bré00, BLL84, Fak90, FS92, Hav91, Hen83a, HT89, Whi85c, Zac07]. Insider [HØ07]. Insight [Hai03]. Inspected [CX01, KFM06, KKY02, BM23]. Inspection [BBC01, Die99, OKM86, Par94, Qiu91, Yeh95, Yeh03, Zuc80b, Zuc89]. Inspection-Repair-Replacement [Yeh95, Yeh03]. Inspired [Pós09]. Instabilities [Jag92]. Instantaneous [CL03, GPZ05, Qui91]. Instants [LR80]. Insurance [ARZ08, Bal01, BBMB08, Che11, KM97, KK06, KP07, KZ09, Ven79, Wil96, ZLW06, BPPR16, LPS17]. insurer [LY23]. insurers [ZC14]. Integer [AAO90, Gue97, IK06, Ruz10, Ver69]. Integer-Valued [AAO90, IK06, Ruz10]. Integers [AT70, Det96, Haj82, Neu94b, Rut99, PP14, Sim21]. Integrability [HK95]. integrable [AB15, KOH23]. Integral [BBP03, CLS09, Coh67b, CCFR00, Dal69a, Del85, Gil68, JRR95, JP09, JM86, Mil79, Møl96, PT71, Pur73, Pur71, Rap08, RSS84, Rie80, SW88, Tok08, TH72, Wiś20, BFK16, Lac93, LP18]. Integral-Equation [CLS09]. intégrale [Lac93]. Integrals [Ben98b, BRPS82, HZ92, ML67, Ors87a, Pak10, PS02b, RW64, Ros76a, Tan80b, AR15, Jan21, NR22, Sim19]. Integrated [AC94b, Dav01, DO97, EY11, HH07, Klü88, McN82, Sly07, Zac04, Lef24, Met13]. Integrating [BM09]. Integration [Mat04, Pri00, PW07b, HS14, PS19]. Intelligence [Rob90]. Intensities [DM75, Gra72, SI83, YA88]. Intensity [And84a, BCIM01, CF11b, Dal82, DJ05, Fei94, Gra71, Gra72, Gra83, JI88, Kab12, LS91, Miy85, Miy01, OK86, Pla04, Pra70, Ros87a, Sch96, dSS04, PLRB23]. Intensity-Based [Pla04]. Inter [Ano72e, Dig76, Gil76, Sho72a, Sho73, SH70a, Yan75]. Inter-Arrival [Gil76]. Inter-Plant [Dig76]. Inter-Record [Ano72e, Sho72a, Sho73, SH70a, Yan75]. Interacting [Bec70, Fen97, GLS00, IMS98, KS23b, Tor02, BCN24, CLCZ14, DLM14, JLX22, Men23, MZ20, Sah16]. Interaction [Bas71, Cer88, CG00a, CC01, CG69, Hol68, Law70a, Law70b, Law71a, Law71b, Law79, Råd72c, Sch19, SR70a, TR65, BDT22, BN17, DV20]. Interactions [BFS09, CG90, Ver98, BW23, HZ22b, KP23]. Interannual [ZR01]. Interarrival [HB02, KS93a, Lat81a, O’B75, YC82, HL13a]. Intercept [Col81, MC70, MC71]. Interchange [NTW89]. Interchangeability [KM90, Web79]. Interchangeable [OS91]. Interconnected [Pur68, Ren73]. Interest [Cai02, ES11, FS98, GCS07, HH07, KNT10, Muc02, Sch98, Tan05, ZY08, HJ13b]. Interesting [Hen72, Hen73a, Let74, Neu94b]. Interfaces [CW93, CH22]. interfailure [SP17]. Interference [DFM^{+}06, KJ81, VI14]. Interior [CH90, Mas95, Van89b]. Intermediate [KR92, Neu68b, BIMR22, CCFM17, IK18]. Intermittent [Ish74, Nád71]. Internal [Guo02, MM03a, dL23]. Interparticle [LM90]. Interpenetrating [Mor73e]. Interpoint [KS83b]. interpolating [HS14]. Interpolation [Pou92, FHZ15]. Interpolator [Bha90, Kab81]. Interpretation [Hal83a, HKO94, Jas15, KY17c]. Interrenewal [Yu11]. Interruptions [HT89, Ish04, Sah71b]. Intersecting [Mas86a, Bai20]. Intersection [Gat84a, JK81, Kwe75, Leh72, Rei67, RS10, Str75b, Weg76, YW64, MO23]. Intersections [Cow78, Gat82b, Mor66b, RS22, WG93]. Intershock [And87]. Intersite [Ste03]. intertimes [OT15]. Interval [Ada75, Ave09, BV11, Dob70, GHK04, IM03, Llo89, Man83, MW96, Mil92, MP93, Qui91, Rhe98, Wil92, BGM15, KS13, LW19]. Intervals [BBC01, Dal82, GL67, GRD^{+}00, Lai78, Lam68, Law71a, Law71b, PS80, Ros81b, Tam70, BS15c, JP22]. intervention [SL21]. Intrinsic [IM10, Mar00]. Intuitive [Hai03, NMG00]. Invariance [AS89, AT79, Bie98, Bré91, Bro69a, CVW84, Cse79, Gue81, HIR03, Mas95, Miy82, LSWX20, PU16]. Invariant [All97, And03, BK12b, Ben68, BPS80, CC96, CD05, ES87, FMS99, GPZ05, Kal91a, KL03a, Led97, Lop14, Neu80, Pol94, RS94, Sen71, Suo79, Toy97, Twe88, Twe98, Web07, Let94, Neu23]. Invariant-Sum [All97]. invariantes [Let94]. Invariants [BS98, IMS98]. Invasion [DF81a, Kel77, Sud76]. Inventory [Agn75, AB86b, Bat66, Boy69, Boy70, De 85a, Dos81, Fel78, GHL71, HNG93, Ram81, Sah83, Zhe91, BPS23, OT16, YCW15]. Inverse [Amb65, ARM98, BKT98, Bha90, CR72, DC10, DGM08, Emb83, FSW12, KR86, Lag05, Lef97, Pit77, RM76, Ski68, Kli14, KK24]. Inverses [GG74, Yan88a, KM23]. Inversion [Har90c]. Inversions [CCC03, Ros08]. Inverting [Wak88]. Inverts [Lit92b]. Investigations [CGR08, Cse91]. Investing [DR14]. Investment [DGK00, DC77, DZ00, ET83, RR12, Zuc80a, CCFM17]. Investments [HT12, NS05, KP22]. investors [Tar15]. Involving [BK86a, Don93, Gri72, LK01, Moh72, Mon67, Pet81, RE01, Yas83]. Ion [BMY93, BMY02]. Irreducibility [Twe88]. Irreducible [AML84, Bas70, Rab06]. Irregular [CS80]. Ising [How00b, Pic76, Pic79, Pic82a]. Islands [Spo04]. Isolated [Pic82b, Pen18]. Isolation [Aba73, BOP07, Ols78, SF83, Sud77, Wil77, Ber23]. Isomorphism [BF09]. Isotropic [CM98, Rub79, FPKM22]. issue [Ano17b, Ano17a, Ano18a, Ano18c, Ano18b, Ano19b, Ano19a, Ano19d, Ano19c, Ano19f, Ano19e, Ano19h, Ano19g, Ano20b, Ano20a, Ano20d, Ano20c, Ano20f, Ano20e, Ano21b, Ano21a, Ano21d, Ano21c, Ano21f, Ano21e, Ano21h, Ano21g, Ano22c, Ano22b, Ano22e, Ano22d, Ano22g, Ano22f, Ano22i, Ano22h]. Issuing [BS73, EE76]. Itô [BHPZ06, IS83, Ken92, Rap08, TR89]. Item [Bha69, LA01]. Items [MN82, Tan78]. Iterated [Deo73, GK96, HIR03, Hey68b, Hey73, Hey77, JT01, Jia12, KM04, Par75, Pia09, Sen81, SH70a, WW00, WS04, BIMR22, HL23, IK16, IRS23, Kan23, MC18]. Iteration [CCFG96, CCdO05, Las94a]. Iterative [LL90, TG99]. Iversen [GNN86].

J [Ano86p, Tan79b, Tan80a, TP79, vDB87]. J. [OW79]. jack [Zho23]. Jackson [BFL05, BS94, Cao89, Che96, GM84b, Lel05, Lem87, Mas84, Mas86b, Mas87, SM05, SDH16, Wei05]. Jacob [BS01b]. Jagers [OS14, Sag20]. James [Ano94l]. Janken [MU00]. Jenkins [Nic86]. Jesse [Ano00j]. Jiashi [MZ01]. Jigsaw [BCKK17]. Jirina [DS12, Li09]. Job [Cha93, DS02, MC90, WB03, Zuc83]. Job-Observer [DS02]. Jobs [Git81, Gla93, Leh88, PC85, Web82, WVW86, XMKW90]. Johnson [Erh00]. Johnson-Mehl [Erh00]. Join [FL97, KW89, Tur00, FS17]. join-idle-queue [FS17]. Joint [AT91a, Ano90h, Ano06f, BSDG20, CCC03, CPN18, Csá94, Fea03, Fer95, Fer93, GDP04, HH96, Hsi76, Huf88a, HS12, Jen66, KT83, Lam91, LM89, NSB10, Neu68a, PP07, Ryc87, Sal13, Sau76c, She79, Wei69b, Win89, BAR14, HJ13b]. Jones [Ano01k, Hop76, Phi01]. Jordan [DIKM10, GJ80]. José [Gan98, Tur92]. Joseph [Sen17, Sne05]. Journal [Ano15, Ano13, Ano14]. JPR [Ano17b, Ano17a, Ano18a, Ano18c, Ano18b, Ano19b, Ano19a, Ano19d, Ano19c, Ano19f, Ano19e, Ano19h, Ano19g, Ano20b, Ano20a, Ano20d, Ano20c, Ano20f, Ano20e, Ano21b, Ano21a, Ano21d, Ano21c, Ano21f, Ano21e, Ano21h, Ano21g, Ano22c, Ano22b, Ano22e, Ano22d, Ano22g, Ano22f, Ano22i, Ano22h]. jth [Hen86c]. Juggler [LV12a]. Jump [AH84, Ara04, BWY11, Cai11, CFS11, Eks06, Gap07, JC06b, Kur70, Kur71, Li00a, Loe09, SS88b, ST12, SH03, Yam89, ZPR08, DM18, GKS14, Jia15, Kun22, Per23, Wan15]. Jump-Diffusion [Gap07, ST12, Jia15]. jump-telegraph [DM18]. Jumps [AR01b, BD12, Bre08, Bre10, CJ08a, Cov09, DWW11, Fra07, GHK04, Hor04, JKM12, KS01b, LM08, LR12, MT02, Par85, Pla04, Zho04, DIMZ13, HL16a, KMG19, OT15, SW20, ZW17, Zhu14]. Jumpwise [LBL05]. jumpy [CM24].

Kac [Ara94, CGR08, Jas15, Lou84]. Kac-Type [Ara94]. Kalikow [PLRB23]. Kalman [AB15, DJ05]. Karlin [Khm11]. Kato [RD80]. Katona [DD15]. Katz [Olk77]. Keeping [Jac99]. Keilson [GR05]. Kelly [Eth04]. Kemeny [BHL^{+}18]. Kendall [Reu82, Ano75f, BH73, Csö80, DVJ08, HB75, Pit90]. Kent [Csö80]. Kermack [WBS16]. Kernel [Eli09, KS23a]. Kernels [Sch11, Sag20]. Kesten [BB80, HWX18]. Kesten-Spitzer [BB80]. Key [Ser92]. Khinchin [DM75]. Khinchine [Kel12]. Khintchine [JS96]. Kiefer [KW14]. Kiefer-Wolfowitz-type [KW14]. killed [KK24, Mar21, Vid19]. killed-resurrected [Mar21]. Killing [KT82, Wen86, Wen90, vDZ05, HCY21]. Kimura [EH95]. Kin [JW82, JW85]. kind [JSW19]. Kinds [Gas64, Jag69, LL82]. Kinetics [Ano68d, Bro88, DR81, Hal83b, McQ67, MV86, Ols78]. Kingman [KJS15, Yao96]. Kiu [AW21]. Klaus [Sto99]. Klimov [Cha93]. Knight [BPS12]. Known [BF08, Bru05, Saw77]. Kohonen [Sad01]. Kolmogorov [Dur71a, Gel77, Par88a, Pol82, Pyk94]. Kolmogorov-Type [Pyk94]. Kroese [GR12]. Kroese-Type [GR12]. Kronecker [DO12]. Kronecker-Based [DO12]. Kruskal [DD15]. kth [Dzi84, Dzi86, Hen86c]. Kumar [NMG00].

L [Hör05, HL05]. L-Like [HL05]. labeled [HLF19]. Labelling [BK73a]. Labels [NA04]. Laboratory [WR97]. Lack [CD92, GM84a, Lin94]. Ladder [BD94a, Cho94, Don80a, DF81b, FS92, FL98, Grü85, MM03a, Pis06, Sch11, Sta77]. Ladder-Height [Grü85]. Lag [Ben70, Kar74, KS78, Sch75b, Sch76]. Lagged [Hey81]. Lagrangian [BM90, SMS94]. Lajos [Ano94m, Bin94, GG94b]. Lampard [Pha74, Pha75]. Lamperti [AW21, CC06, NP00]. Lanchester [Gol77, Kin02]. landscape [Dal16a]. Lane [BR79, Cow80, McN70]. Language [Bra70, Dob70]. Laplace [AKP91, CT90, GG14, HTV92, Har90c, HL05, Keb97, Kim16, Kle83b, KLY16, Met13, Ros08, ST16, Sen95, SW97b, Taï96, Wil99, ZW17]. Large [Abr07, AMW88, BC95b, Bon02, Bre69, Bro79, BNS90, BN15, Cai03, CCN10, De 94, Duf96, DM05b, DS08b, DMT11, DT11, DNW07, FG08, FLP99, FF03, FS17, FWL03, GY08, Gol76, Hai89, Heu87, Hey75, HS10, JLN10, Jon04, Kin82, KLP22, KM97, Leh17, LN23, LM99, Mac01, MS06, MS99, NTYY03, NTaYY04, Nyr99, O’C97, O’C98, Pap96, PSK17, PS80, Pen96b, PW67, Red03, RRS05, SUY21, SW87, SB17, Suz10, UY23, Van99, Ver98, WW07, Web07, YCG11, Zac00, Zaj95, Zha04, Zha05, ADJ13, ACVZ20, AR13, Ber13a, BR18, CHO22, Chu18, GJP^{+}21, JPS18, KLY16, LSWX20, Sal13, SX18, Tsu17, Tsu18, Zha22, Zhu23]. large-claim [ACVZ20]. Large-deviation [KLP22, SUY21, Zha22]. Large-scale [FS17]. Large-Time [FF03, RRS05].

Largest [DH89, MZ93, ST86b, Ber13b, CMW16, KK18, LWM20, Zha21]. largest-order [LWM20]. Last [Ano86j, Bau09, BP00, CS86, DS94a, Fak86, Fak87, FM97, FR10, HY02, Iva16, LYZ17]. Last-Come [FR10]. Last-Come-First-Served [Fak86, Fak87]. Last-Exit [DS94a]. Latent [RS69, Fox16]. Later [FC03, HH03]. Lattice [BMZ82, GR94, KSF80, Ma03b, Mar86, McD80, Mor73b, Mor73c, Mor79, Pic76, Pic77, Pic79, RH81b, RB98, Sha89a, Str75a, Sud04, VSW09, Ver69, WG93, Wie84, Yad70, CdOR16, Gaá20, JX13, KSS16, YW21]. Lattices [Ham88, Hor98, SKF79, Shi02, Str77, CP14]. Law [Aye07, BK12b, BD88b, BT76, Bre10, CG02, Che80a, CCN10, DS86a, Deo73, Duf96, FG08, FR10, GST00, Gne00, Gol77, Gol78b, Hey68b, Hey75, Hos06, JLN10, Jia12, KM04, Ma03b, MW97, McC85, McD79, McK89, Miy85, Par75, Pfe89, RS94, Set03, Ski12, Ste78, SH70a, Tor69, Ver98, Via94, WH75, YSS91, Yos84, ZZ09, BVW21, BSDG20, BCdlP17, DDS22a, DS17, FF20, IK16, Pyc21, All97, JdlR04, Khm11]. Law-Invariant [BK12b]. Laws [Ano90h, BDM09, BG87b, BdH87, Dzi84, Dzi86, EMR71, GS94, Imh99, LM89, MS94, Miy94, Res71a, TW83, TO92, BIM17, CHO22, FGH^{+}21, HL14, Mar16a, RS15, Tan17, Xu17]. Layered [AMAKAH97]. lazy [CSC13, HK22]. LCFS [HAPV12, Miy92, vD89]. LCFS-P [Miy92, SS85a]. Lead [Ano82f, Eng64, Pro82, Sah83]. Leadbetter [Che81b]. Leader [KMW11, FHI17]. Leading [GE82]. Leadtimes [GHL71]. Leaky [GAS03, GMRY07]. Leaky-Bucket [GAS03]. Learning [Agr91, CCH06, Dob77, HB83b, Kur87, MT12, Pru83, BKM19]. Least [GS79, Hea82, HW83, Pla85, SPS80, Sol84, dRdC18]. least-squares [dRdC18]. Leavable [Fas90]. Leaves [SSE87]. Leaving [AR82, MW96]. Lebesgue [KY17a]. Leffler [HJL21, JP93, Möh21]. Left [Gre76a, Kij93, LMS76, Pak73a, Pak78b]. Left-Continuous [Gre76a, LMS76, Pak73a, Pak78b]. Lemma [Bes72, Bru80, Mou99, RS66, SD76, Tas97, Wei80a]. Length [Ano93e, CP01, CT76, DS92, DS94a, Enn69, GR87, Gat82a, Gat87, Haa92, HT78, Jin88, JW90, KS85, LMW98, Man77, MN79, Miy86, Miy92, MP04, MS92, Nag93, Nat74, PS09, Van81, Wak91, Wil88c, BR18, FH17, HK19]. Lengths [CK00b, Col81, GR89, Hen93, Hol80b, KST99, KNR05, Li00a, LS86, Lit79, MW71, McK89, O’C98, Ros12, WCC03, BBH14]. Leo [Olk77]. Leonard [Sne05]. LEPT [CCPW92, CR94a]. Leslie [Men76]. Less [Ebr93]. Letter [AH19, Che82, CH84b, Fal89, GNN86, HS87a, Hor80, Koo04, Lai87, Mad76, Sha72a, Tod85, Wil88b, Wil77]. Letters [KKC82, RD99, TH71, Tan80a, Wei79]. Level [Adl76, BD12, Bri79, Cli83, Coh89a, DO12, Eks03, FP05b, FP05a, Hon04, Kat02, LT00, Nyr94, Roe70c, RC80, Rus83, Ser80, Sha80, Sha79, SS97, BS13b, DR15, GS19, JF23]. Level-Crossing [FP05a, Kat02, Nyr94]. Level-Dependent [DO12, Eks03, BS13b, GS19, JF23]. Level-Phase [LT00]. Levels [Fer93, Has76, IC82, OL65, IK18, Kli14, LL24]. Levy [Kal93, KK06, Lou84, ARZ08, AHL02, AS09, AB15, AR01b, Bau09, BV11, BPPR16, BS23, BBK08, BM09, BM23, Bra02, BDY07, BIM17, BL21, CC06, Cha17, CS20b, CPN18, CP11, DESM10, DSZ13, DS13, Doo72, DGM04, ESM08, FCvS15, FCKS16, Fin13, Fra07, Fro15, GS94, GM11, GS19, HT12, HK09, Iva07, JP18, KK96b, Kel92, KW92, Kel00, KBM06, KB20, KM23, KKS15, KLM04, KMG19, Kuz10, KP07, KZ09, KLP12, KO12, KPV21, LM12, LP21, LM08, LYZ17, LVZ19, LS20, Loe09, MZ24, MPS15, NPYY18, Pat13, Ped02, Pis06, Rek23, RZ07, Ren14, Rog00, RNJ16, RNJ17, RNJ19, SRN21, Sur08, Tho75b, Vid19, Wan12, WZ18, YSL17, Zho04, Zho07, ZW17, KB13]. Lévy-based [RNJ19]. Lévy-driven [GS19, LS20, Rek23]. Lévy-type [KB13]. Lewis [DMR02, Tor69]. Lexical [Dob78]. Liberation [Dav84]. Libraries [Don91]. Library [Nel77]. Library-Type [Nel77]. Lie [Hou12]. Liées [Hua95]. Life [AH84, Ano92h, AG81, BB84, BS90, Bha91, BAAB00, BS69a, BS69b, CW97, CW91, Che94, FP94, GESK95, GE82, GL84b, Jun94, KNS73, Kla02, Li00a, LK01, LZ02, Lin97, LS86, MP70, MO93, MR02, SD94, SS91, SAS04, ZLW06, ZL10, BBH14, BLS15, KBM21, LSW19, PB14, Ros14b]. Life-Testing [GESK95]. Lifemonitored [Gås98]. Lifespan [Ben82]. Lifetime [Ano83f, CG87, CT95, DL02, FGL00, KN05, KM03, KP87, Mei81, Mer80, Nat82, Nat90, Raa80, BB18, DT15, FZL13]. Lifetimes [Ano90h, Ano02e, BOR01, Bur09, EE76, GST00, JNR09, LM89, Mor78, NE07, NBS08, Pin80, Wei69a, WC04, BV14, BHB18, BMR19, BB18, Gup13, HF19, KZN15, MR15, RNR18]. Lifo [Li96, CN86, SS87]. Lift [Kap81]. Light [Cow81, Ken72b, McN68, MO71, NG95, OM72, Sis70]. Like [BB12, HL05, Tur72, BDIM18, GK18, Har73a, LS79, OC22]. Likelihood [ADLN82, ALS04, ARW94, BZ04, Bau81, BH75a, DS80, Hey94, Jen72, Joy94, LK77, MK73, Mul72, Mul73, Oga80, Oga01, Whi82a, Yao93]. Likelihood-Based [Oga01]. Limit [ALM06, Agu09, Amm83, And87, AD97, AD98, Ano72e, Ano81e, Ano91e, ANT78b, AK76, Avn79, BKP^{+}22, Bal79, Bal99, Bar75a, BE87, Bas10, BKMS04, BT92, BK12b, Ben98a, Bèr04, BCF09, BE83, BCPR11, BG79a, Bin88, BP96b, Bra80, CM97, Che76a, Che81a, Che03a, Che06, CCHY08, CHO22, Che80b, CY84, CP88, Cli83, Coh89a, CS80, CFD09, Col76, CJ98, Cov09, CCH95, CH71, DL96, DH89, DY97, Don72, Dur71b, Dzi84, Dzi86, Eag75, Eag79, Ell66, Eth79, Eva78, FTV90, FS07, FSW12, FPY20, FNV11, Gal74, GH10, GZ16, Gei99, Gel77, Ger76, Gol78b, Gol05, GM96, Gre77a, GS06, Gut10, Hal78, Han90, Har73c, HL77, HL81b, HM14, Hen82, Hey67, Hey71, Hey77, HMPS06, Hol78, Hol79b, HM95, HL99, Hoo83, Hoo70, Höp82, Hsi97, Hud73]. Limit [HZ82, Jan83, Jan01, JCM98, Joh06, JW11, Kap78, KS06, Kim70b, KS82, Kle85, KPT99, Kuc82, Kul06, KP87, Kur71, Lan84, Lel05, LST19, Les73, Li06b, LPLG08, Li09, LPP12, LR95, ML85, ML86, MS94, Mat84, MN13, McL76, MP12, MS04a, Moh76, Mor73a, Mor79, MR11, Nak74, Nyr94, O’B74, O’B80, Ort08, Owa23, Pak73a, Pak76, Pla83, Pós09, Pra70, Pur78, Qui78, QS99, Råd72b, RHK90, Roo76b, Ros76a, RD80, SF17, Sat83, Saw79, Sel91, Sév95, Sha74b, Sho72a, SB09, Sta81, Sta82, Szá74, Tak88, Tak95, Tan17, Tom72, TW83, TO92, Twe73, Twe75b, VL15, Via94, Wan81, WJ95, Web80, Whi72a, Wol77b, WW00, WS04, Xu09, Yam89, ZWV10, Zhu14, dHR73, ABSKT19, AG12, BLSGS13, Bow18, Cri16, Fat15, GHW19, IK16, IJB18, IR20, KM15a]. limit [KR18, Kur22, Lam14, NS22, RY20, RNJ19, Rub23, SH21, SGS17, Sei16, Sim14, WW20, Xu17, Yam17, Zhu13, dL23]. Limited [AIH87, AIH88, SCZ77, Tak74, Tho86, Yao07a, Yao07b]. Limiting [AB00, AT91a, Ang11, Ano75g, BJ14, Boo89, Bra83, Bri20, BdH87, CM83, CJ06, CP78, CdLH^{+}23, Deh85, Dur74, Ebr08, Eme72, Eng64, FN70, FMS99, Ham92, Ham95, Igl65, Kao78, Kap74, Ken74, Kij93, Kru04, Lip77, LMX04, MM71, Meh77b, Mi95, Nis74a, PF06, Pha93, RWN02, SA72, Sau75b, SKF79, SZ93, Sum87, Tam70, Vas82b, Woo83a, Wyn73, ZLW06, BLS15, Cur22, GT18, MP14, MC18, Vid23]. Limits [AK93, BMZ82, CG00a, DMS08, De 74, Est76, Har91, HN01, Kal75, Kat11, Kat12, Kur70, McN73, Nea05, RJP97, SF77, SW11, SB78, SD12, Whi72c, ABD17, BDT22, BCK16, BY22, McK14, RR16]. Lindley [Ano90f, Lun97, Roy89, SS89, VAW04, VP10]. Lindley-Type [VAW04]. Line [Alp07, Bre75, Cre77, CH70, ESE84, Fad83, Fal86, FS78, GM73, Kol98, KS01c, LM04a, Mei10, PC76, Pre06, Sch75a, Win77, BPG20, Gab16b, HKT24, KS16]. Lineage [HB99]. Linéaires [Die87]. Linear [AL00, Ano72f, Ano73g, Ano74f, Ano92g, AGC97, BNL05, Bar71, Bar72b, Bar74, Bha90, BSW03, CJ08a, Cha74, CS72, Cof95, DJ05, De 85b, Dei86, DG93b, DM72, Eat66, Ebr08, FDVF93, Fen95, FW10, GC71b, Gir86, GG00, Gra71, Gre80a, GZ94, GG77b, Hai78, Han71, Han73a, Hea82, Heu87, Hey88, HH85b, Huf88a, HW79, IKP01, Jia12, JC06b, JL06, KT82, KW99, KS01b, KS02a, Kle83a, Kle89, Kle94, Lam84, LHN86, LM90, Lit79, LS92b, LMW00, MJ85, Mal67, MC70, Mas77, MT70, MSZ04, Mor74, NP00, Oza80, Oza81, Oza86, PR95, Pha91, PH86, Pur73, Pur75b, QR90, QL99, RS04, Rob72, Ros80b, SC82, TR89, Ton74, VOZ95, Wei75b, WR74, XH11, Die87, FHZ15, FZ17, Hoo13, KS18, NR20, SX18, WY22, WW20, MC71]. linear-fractional [WY22]. Linear-Intercept [MC70, MC71]. Linearization [BW98]. Linearly [Con71, LV72, Mak10, Mor78, OK86, Rey72a]. Lines [Hal81, KMM81, PV12, Pin71, Sag99, SY72, Tan83b, Tan83c, MO23]. Link [DH89, LRS07, ST86b]. linkage [ACK20]. Linked [BH01, Fea03, Hol81a]. Linnik [AA93, Mor73d]. Liouville [CCGW23, HJL21, HM20]. Lipschitz [HL23]. Liquidation [ELTW10, EV16]. List [Ahl73, ANW82, KR80, VOZ95]. Literary [Bra72]. Little [FR10, McK89]. Littlewood [Agu03]. Litvin [Jof77]. live [BBH14, BHB18]. Lives [BS73, Ros14b]. Load [HST83, LDL95, Sch72, HLVM22, LY20, MBvLW16]. load-balancing [LY20]. Load-Sharing [HST83]. Local [ABD17, AN95, BD94a, BP99a, CJ08a, CC01, Fer93, GP04a, GL99, HB84, Höp82, JC06b, Kon96, Lou84, Mor79, NT04, Oza86, Pih15, Qui78, RB98, SS07, Sév95, SB09, SW94, TW01, Tet98, WCW05, Wil92, dL23, Bai20, Bre13, HL13b, KADK19, Kur22, Pit18, SDH16]. Locally [JT01, Kuu82, OC22]. Located [Str75b]. Location [Bel66, Dav80, Fer95, SSSL10, She79, Pim14]. Location-Independent [SSSL10]. Locations [AW90, Ish75]. Loci [Fea03, Hol81a]. Locomotion [Wat97]. Locus [Eth79, Fea02, KF70, Kin88, Man13b]. Log [Ald95, Ano10e, GK22, Kei71, Mar07, MK73, FHI17]. Log-Concavity [Ano10e, Kei71, Mar07, GK22]. Log-Convexity [Kei71]. log-exp [FHI17]. Log-Likelihood [MK73]. Logarithm [Deo73, Hey68b, Hey73, Hey77, Jia12, KM04, Par75, SH70a, IK16]. Logarithmic [GW94b, Gri75b, HLVM22, KM15b, LR93, LS09c, Ros08, Ber13a, HH23, LPN13]. Logistic [Aba79, AD98, AR89, CP03, DM10b, KL89a, Ova01, Pra80, Sim93, Voo87, Fox21]. Loglogistic [Voo87]. Lognormal [BS81]. Loi [Tor69]. Lomax [Ano90f, Nay87, Roy89]. Long [AHN90, AHL02, ALS04, BS08, CD07, CV01, Dai04, DSS11, Dia76, FKZ09, Gaá20, Gao04, GKL01, GG00, Gne00, GAS03, HY97, Hey02, Jac09, Ksi89, KS01c, KS21, Kün86, Lam70, LS07, LS20, LNTZ99, Ma02, Ma03a, Ma03b, MW97, OA12, Ols77, Par86, BS15d, BS14b, BS16b, BCKS13, LRM23, MV16, Neu23, WW20]. Long-Chain [Ols77]. Long-Memory [AHL02, Ma03a, Par86]. Long-Range [ALS04, BS08, CD07, DSS11, Gne00, GAS03, HY97, Hey02, KS01c, Kün86, Lam70, LS07, LNTZ99, Ma02, Ma03b, MW97, Gaá20, BS14b, BS16b, BCKS13, MV16]. Long-Range-Dependent [Dai04, Gao04, OA12]. Long-Run [Ksi89, BS15d]. Long-Tailed [FKZ09]. long-term [Neu23]. Long-time [LS20]. Longest [CS75, CMW16, FWL03, LH07, STC93, Ton85, HK19, KLY16, PW16]. longest-dominating-queue [PW16]. longest-queue [PW16]. Longtailed [EO84]. Look [AH77, Ano79e, BLL79, Coh82, FR10, Kap81, Sen94, Sta93, Yan90, Yas88, FPKM22, HPP14, LGC16]. Look-Ahead [AH77]. Looking [CS85b, Käm89, Möh10]. Loop [ABGH00, Law83]. Loop-Erased [Law83]. Lorenz [DZ03, Lin95]. ’Loss [BKT97b, Abr01a, Agr91, BKT97a, BBMB08, CD95, Fak82, Fak83, FG98, Gla77, Gne07, Hun95, JFS12, LM99, LR04, LR07, MT00, Miy90, NE81, NE82, NTYY03, PS84, PS86b, Pou90, ST73, Sob90, SZ06, SG88, Wil97, WW76, WW02, ZZ99, Zac00, Fro15, HR15, LHT17, OS14, Ros14a, VM21]. Losses [Abr01b, Erl67, PRX03, Rig99b, Tam70, Wol02, GS15]. Lost [KK69]. Lost-Games [KK69]. Lotka [Bil77b, Hul01, Ito89]. Low [BC72, MO70, New66, Ros87a, BHZ22]. Lower [Ano06e, Ave85, Bis23, BC01b, BGSS03, Cha75, CCC03, CS96, Che03b, Eco00, GWS80, HST83, How00a, Iks04, Kan85, Pin11, SC02, Sud01, Sud05, Tho11, WW04, AI16, Bon13, GPZ17, MA14, ZZ16]. LRU [JRS06, VG92]. Lu [NMG00]. Lu-Kumar [NMG00]. Lukacs [Roh88]. Lumpability [AML82, AML84, BY93, Buc94, Led97, LWB08, RS89b, SR88, GT14, KADK19]. Lumped [Nor97, Tay71]. lumping [KSS16]. Lumpings [GT14]. Lundberg [Bré00, DM21, Nie06, Sch96, TTI22, WL94]. Luria [Gol95, Kem94, MSS92, Möh05, Pak93, Pro96, Zhe09]. Lyapunov [ABD01, Kub24, Kun22, MW21]. Lyapunov-type [MW21].

M [Abr01a, Ano72g, ALH01, AIH88, BM89a, BS02, BGT98, BM94, BB12, CR97, CK00c, Coh67a, CL93, DN80a, DS98, Gra83, GMS96, GP98b, Han87, He03, Hea65, Hok75, HL04, HPR86, Jen72, KT15, Kyp72a, LT00, LC83, LT73, OP86, OW79, Pre97a, Pre97b, Pre02, Ram84, RS90, Sim97, Sim98, SP92, Tam71, Tso71, Web79, Wei79, Wei80b, YC01, YT83, ZL15, CP92, Abr01b, Ano72g, ALH01, AF83, AT96, AK97, AIH88, BM89a, BT73, BS02, Bhu06, BNW13, BM94, BY97, BPSZ10, BG00, CP01, CR97, CP92, Coh67a, Coh71, Con94, CL93, CT76, CN86, DS94a, DS98, DT80, DGM04, Enn69, Fak82, Fak83, FF91, Fal94, FT80, FG84, FG98, FY83, FR10, GBG02, Gha86, Gil76, Gra83, Gre89, GW11, GMS96, GP98b, Han87, Hey74]. M [Hoo83, HL04, HPR86, Huf87, Hun86, HAPV12, Ino19, JS96, Jen66, Jen72, Kat12, KR88, KS94a, KZB05, Ken74, KKK07, LC83, LLPC94, LSC97, LH06, LH02, MT00, MNS01, Miy86, Miy90, Neu68a, Neu94a, OP86, Ott79, Ott84a, Ott84b, Pre97a, Pre97b, Pre02, RHK90, RS90, Rig99b, Ros73, RS99, Rou96, Sch88a, Ses96, SS85a, Sim97, Sim00, Sta85, Tak92b, TTH93, Tam71, Tan94, Wak91, Wan99a, Wie89, Wil88c, YT83, ZL15, ZZ04, Hok75]. M. [Wei79]. M/Cox [Bhu06]. M/D/ [FY83]. M/D/1 [BG00]. M/D/1/N [GBG02]. M/G/ [BNW13, FG98, GW11, Huf87, KS94a, Sta85]. M/G/1 [AF83, AT96, AK97, BY97, BPSZ10, CP01, Coh67a, Coh71, CT76, CN86, DT80, Enn69, FF91, FT80, FR10, Gil76, Gre89, Hey74, Hoo83, HL04, HAPV12, Ino19, JS96, Jen66, Kat12, KR88, KZB05, Ken74, KKK07, LSC97, LH06, LH02, MT00, MNS01, Neu68a, Neu94a, Ott79, Ott84a, Ott84b, RHK90, Ros73, RS99, Rou96, Sch88a, Tak92b, Tan94, Wak91, Wan99a, Wil88c]. M/G/1-FIFO [CN86]. M/G/1-Type [MNS01, LH06]. M/G/1/ [Hun86]. M/G/1/K [SS85a]. M/G/1/K/N [TTH93]. M/G/1/n [CT76]. M/G/2 [Wie89]. M/G/C [Gha86, Ses96, FG84]. M/G/k [Fak82, Fak83]. M/GI/1 [DS94a, DGM04, ZZ04]. M/GI/1/n [Rig99b, Sim97, Sim00]. M/GI/s [Miy86]. M/M/ [GMS96, GP98b, OP86, Pre97a, Pre97b, Pre02]. M/M/1 [Ano72g, AIH88, BM89a, BS02, BM94, CR97, CL93, HPR86, Jen72, Tam71, YT83]. M/M/2 [RS90]. M/M/2/2 [Han87]. M/M/c [ALH01, Gra83, LC83, ZL15]. M/M/k [DS98]. MA [BH98]. Mabinogion [Lin16]. Machine [AC02, CR94a, Fro93, Gla93, KJ81, Mei81, Rig97, Son96, Zuc89, BKM19]. Machines [CCPW92, Käm87, Leh88, PC85, Web82, WVW86]. MacLaurin [GU92]. macro [Sag20]. macro-individuals [Sag20]. Macroparasite [HI01]. Made [MH76b]. Magnitude [FS71]. Maintainability [Ano86h, Dav75, Hai83, VP92]. Maintainable [Ano91f, Dav81, Gue91, Hai92]. Maintained [Bax83, Kin88, MT95, WK94]. Maintaining [Abd87]. Maintenance [BLS90, DG06a, Hei01, Lew64, LS03, MR03, Nak86a, OA70, SZ91, Yeh05, CC15]. Majorisation [Hic83, Hic84]. Majority [BS06, BCS08, BPT89, Bol01]. Majorization [DC06, DC10, MN79, Ton97, GR24]. Make [CNS99]. Makers [CRS97]. Makespan [Web82]. making [PS18]. Malaria [Rad74b]. Malthusian [Sau76a]. Mammal [Hol67]. Management [Guo02, Des15, HBHH23]. Managing [KAR11]. Mandelbrot [Hua16]. Manifold [CM97]. manifolds [AK23, Hen18]. Manpower [Ano86h, Ano91f, Dav81, Dav82a, Dav83, Dav85, Gue91, Gue99, Hai83, Meh77a, Meh77b, Meh79, Vas81, Vas82b, VG85, Woo83a]. Manufacturing [AIH93]. Many [Ano77g, Ano85f, Ano87h, BT82, CW96, DR84, DM03, DKO84, Eth92a, Igl65, Lan86a, Naw83, Nel77, PS89, Sch91b, SP99, SM97, Wat76, Zho15]. Many-Server [DR84, Naw83, PS89, SP99]. Map [Cha94, CL09, Ski68, BFK16, AW21, DO16, LN94]. MAP/G/1 [LN94]. MAP/PH/c [DO16]. Mapping [Fei93, Jaw84]. Mappings [Ber81a, BM90, BGP80, BR95, Eng93, HJ02, Jaw99, Mut87, HJ13a]. Maps [BCIM01, TW01, HL23]. Marches [Bie98]. Mardia [Ken91]. Marginal [SS17, SS93, SM82]. Marginalisability [BY93]. Marginals [Kal95, HM15, Teh13]. Margins [Ano98f, Joe96, JS98]. Mark [Sen17]. Marked [AK93, BNT02, Dsh01b, FS80b, Hol08a, KK95, LRS07, Miy94, Rüs77, Tor02, Van96, BMZ23, GP13, Ros14b]. Marker [CK00b, Lit79]. Market [Ano99a, FPY04, Rig99a, PS18, SWW21, Tar15]. Markets [FG08, Iva07, PR96, Win05, BL13, HBHH23, Mos18]. Markov [AML82, AML84, AC94a, Agr91, AR05a, AS04, AM16, ABS15, And81, AFV15, Ano70e, Ano73e, Ano77f, Ano91g, Ano91i, Ano93f, Ano07e, ADPR18, ANT78b, AK23, AT87, Ass83, AW15, Att87, Att89, APZ13, ALL21, BCP00, Bal79, BY93, Bal99, BMY02, BM05, Ban88, Ban96, BW23, Bar75a, Bar82a, Bas70, BR98, BS15d, BSS72, BM82, BR87b, BB80, BR96, BD74, BM01a, BS97, BP01, Bon81, BP99b, Bra83, Bre08, Bre10, Bre12a, Bre12b, Bre13, Bro72, BRPS82, BHX98, Buc94, BT12b, BNS90, Bur76, Cal74, Cam73, Cao95, Car99, CD07, CCFG96, CCdO05, CCdOS15, CCCSdO24, CL85, Che91, Che02, CPLZ07, CNZ14, CCGW23, Che70, CT72, CDL88, CK07, CdSJ20, CTKR83, Coh14, Coh83, CJ98, Cos90, CD05, CD06, CD09, CDD10, Cox77]. Markov [CT95, Cra11, CM15c, Cra88, Cra73b, CL92, Cse91, DIKM10, DS06, Dal69a, Dal82, DS92, Dan69, DP00, DS65, DS67, DM67, DM68, DW72, Dav75, Dav81, Dav82a, Dav83, Dav85, Daw73, Daw74, DGM08, DK11, Die87, Dos79, DS80, Du95, DKK08, DP10, EG02, ELNT11, Eva71, FN70, Fas90, FT78, Fei93, Fel76, Fer95, FHL13, FP05b, FP05a, FH72, FKK91, Fra06, Fra00, FR86, FWL03, FC03, FL98, Fuj13, Gad89, GG91a, Gan82, GT85, GY95a, GY08, GW94a, GS14, GJQ06, GT14, Gén23, Ger23, Ger90, Ger91, Ger83, Ger76, GL03, GKPS06, Gla78, GBW92, GMM05, GMM06, GKS14, GH90, Gre77b, Gri75a, GS85, Grü98, GH14, Gue13, GZ02, GHL03, GHL05, GZ06, GPZ17, GH21, GG78, HK95, HH03, Han92, Har90c, Har91, Har94, Haw70, HN01]. Markov [He03, Hei09, Hel70, Hen72, Hen73a, Her99, Her03, HL16b, HNF85, HJ11, Hol00, HST75, HR88b, HLZ05, HK19, Hou12, Hsi97, HY02, HY94, HM17, ISK95, IG01, IT78, IA78, IC82, Ish77, Iva10, Iva14, Jac09, JP16, JC06a, Jan21, JR82a, JS14, JT01, JL98, Jen66, JK71, Jia15, Kal89, Kal90, Kal93, Kam17, KS19, Kao74, KÖ89, KZ88, KS02a, KS04, KM23, Kem92, Ken90, Ken78, KG68, Kha70, KM66, Kha68, Kij87, Kij88, Kij92b, Kij93, Kij98, Kim70a, Kin69a, KT01b, KT01a, Kle94, Kol00, Kol82, KW14, KR21, Kot94, Kou96, Ksi89, Kur87, KSHH98, Kur70, Kur71, Lai78, LS92a, Lam68, LLZ17, Lar98, Las94b, Las94c, Las94a, LT00, LT04a, LRS94, Led97, Led05, LR13, Leh71, Leh72, LP85, Les73, Let74, Lev77]. Markov [LT23, LS97a, LPLG08, LB23, Lil68, Lin87, Lit92b, Liu94, LH06, Liu11, LLL20, LS79, LPP12, LL90, LZK06a, Mac01, MPS21, Mao02, Mao04, MW21, Mar73a, MV95, MSV14, Mar21, MS91, Mas93, Mat04, Mat92b, McC80, MB16, McN85, McN74, Mei99, MPHL10, MR03, MSCZ78, MV21, Mit03, Mit04b, Mit05, MLB05, MT02, MN16, Mon82, MP04, Mor86, Mou99, Moy65, Nag01, Nak74, Nel77, Neu23, Neu66, Neu78, Neu94a, NP22, Nie85, Nie95, Nor97, Nue08, Nue10, Oga80, OA12, OW94, Ols78, Ōsa85, Pak78a, Pap96, Paw89, Ped71, Peg80, PU16, Pen96a, Per23, Pes68, PS92, Pha88, PPD97, Pin71, Pit77, PS86a, PT11, Pol94, Pol99, Pol01, PS02b, PT94, PRHL12, PS01b, PS04, Pru85, Pur75a, Pur71, QL99, RS04, Rab06, Raj74, RG88]. Markov [RM76, Rey72b, Rey72c, Rid96, RB05, Ris95, Rob91, RRS98, RT00, RT01, RR07, RR16, Rob80, Rod95a, Rod95b, Ros81a, Ros82, RD80, Ros81b, Ros70a, RS89c, RS89b, Rud76a, Rui08, Rüs08, Rus78, Ryd96, Sah78, SCZ77, ST01, Sar73, Sch11, Sch82, Sch70, Sch68, SAI82, SI83, SVJ66, Sen82a, ST85, Sen01, Sen87, Ser90, Sév95, Sgi96, SH03, Sie86, SP22, Smi73, Smi92, Sob82, SB09, SV03a, SK98, Spo07, SV71, Ste06, SR88, Suo79, Suz10, Szá74, Szp85, TH11, Tan97, Tav80a, Teu68a, Teu68b, Teu72, Tin85, Ton75, TSCD08, Tru00, TV88, Tsa94, Tsa96, Tsa99, TV93, Tuc76, TW84, TG99, Twe73, Twe74, Twe75a, Twe75b, Twe81, Twe83, Twe88, Twe98, VK00, VS06, Vas81, Vas82a, Vas84, VT89, VGT90, VP92]. Markov [Vid20, Vit76, Vrb85, Wan76b, Wan81, Wan92, WCC03, WZ65, Wol77b, Wu93, WG15, Wüb12, Wyn73, XS03, Yan73, YH04, YY11, Yao85, Yap09, Yas88, YT83, Yun00, ZZ99, Zaj95, Zei89, Zha14, ZL96, ZPR08, Zuc78, dCGG16, dL23]. Markov-Additive [Zaj95]. Markov-Chain [Vrb85]. Markov-Dependent [Dal82, DW72, HY02, Lai78, Pes68, Tin85, YH04]. Markov-Modulated [BR98, BR96, Bre12a, Bre12b, Du95, GY08, Iva10, JL98, KS02a, KS04, MT02, OW94, PT94, RS04, Rab06, RB05, Ryd96, Bre13, NP22]. Markov-regime [Jia15]. Markov-Renewal [LT04a]. Markov-Type [MP04]. Markovian [Aal98, Alp89, AA90, AK93, AM05, Bar15, Bas71, BEER15, BB14, BPS80, BD94b, CSC13, Çin67a, Çin67b, Cra73a, DN99, DS04, GGS^{+}95, Gri75b, GM20, HL13a, KP89, KFM06, KADK19, KKY02, KS86c, Kol98, Kol95, Lat81a, LOR19, Len81, Lit94, Llo71, LS02b, LS11, MP00, Man69, Man71, Meh77a, Meh77b, Meh79, MT70, Mor73b, Mor73c, NC72, Neu79, NP91, NA04, Oak75, OL65, Pak73b, PM73, Pha79, PT23b, RCLW10, RCC12, Sgi99, Sha89b, Sie81, Tav80b, Tay71, TP76, TW89, Tuf99, TT79, Vas82b, WJ95, WF99, Wau72, WF80, WT05, ZTML21]. Marks [Ano74f, Mor74]. Marriage [VJ01]. martensitic [CH22]. Martin [BK64, Yas93]. Martingale [Ara04, Big77b, Hey77, Lyn00, MW90a, Nie06, Pak87, PH87, Phe96a, Pic80, Pic84, Rou96, San06, Sch93, Sim19, Sto91, Tan79a, Web80, Wes99, WH04, Cri18, DM19, Des15, IK16, KR21, KS17, LR13, SZY19a]. Martingales [Bac86, BCF09, BD94b, Dek96, HS00, Hey73, Hua95, IM10, KW92, Rie76, Sen82a, KB13, Pih15, Teh13]. Masking [Sie85]. Mass [CR97, RO99, ZL15]. Matchbox [BG02]. Matched [Oak79]. Matches [KBBW07, MRP04]. Matching [Ano74g, CZ99, HS74a, RS00, TW01, MM16, MBM21]. matchings [AFP22]. Matchmaking [GW11]. Material [Ano04g, Ste04]. Materials [Kim94b]. Matérn [DDS20]. Mathematical [Hol66, Pur69, Rei67, Rén64, WK65]. Mating [Chi68, CD85, DHT86, Ell66, Hin82b, Hul82, Kar68a, Kar68b, MM09, MMR02, Pol10]. Matrices [BF78, Ber73, DS94b, FHT78, GG74, GS87, Gre84, Hey95, IM74, JI88, KS93b, NS80, Nei01, Peg75a, Pen96a, SS96, Ste95, Twe71, Van86, ZLA99, HL16b, McK14, Men23, SUY21, UY23, WY22]. Matrix [Bha90, DKK08, FDVF93, Har94, HN01, He03, Hey91, Kam79, KS78, LS97a, Mar86, Mil94, MT02, Neu80, Ork72, Pha91, PS04, Pur73, Res66, Sha72b, ST12, SB06, Tay99, ET21, Men76, Per23]. Matrix-Analytic [SB06]. Matrix-Exponential [ST12, Per23]. Matrix-Geometric [Neu80, ET21]. Matter [Ano64a, Ano64b, Ano64d, Ano64e, Ano65a, Ano65b, Ano65f, Ano65g, Ano66a, Ano66b, Ano66f, Ano66g, Ano67a, Ano67b, Ano67c, Ano67f, Ano67g, Ano67h, Ano68a, Ano68b, Ano68c, Ano68e, Ano68f, Ano68g, Ano69a, Ano69b, Ano69c, Ano69d, Ano69e, Ano69f, Ano70a, Ano70b, Ano70c, Ano70f, Ano70g, Ano70h, Ano71a, Ano71b, Ano71c, Ano71d, Ano71f, Ano71g, Ano71h, Ano71i, Ano72a, Ano72b, Ano72c, Ano72d, Ano72h, Ano72i, Ano72j, Ano72k, Ano73a, Ano73b, Ano73c, Ano73d, Ano73h, Ano73i, Ano73j, Ano73k, Ano74a, Ano74b, Ano74c, Ano74d, Ano74i, Ano74j, Ano74k, Ano74l, Ano75a, Ano75b, Ano75c, Ano75d, Ano75h, Ano75i, Ano75j, Ano75k, Ano76a, Ano76b, Ano76c, Ano76d, Ano76e, Ano76f, Ano76g, Ano76h, Ano77a, Ano77b, Ano77c, Ano77d]. Matter [Ano77h, Ano77i, Ano77j, Ano77k, Ano78a, Ano78b, Ano78c, Ano78d, Ano78i, Ano78j, Ano78k, Ano78l, Ano79a, Ano79b, Ano79c, Ano79d, Ano79g, Ano79h, Ano79i, Ano79j, Ano80a, Ano80b, Ano80c, Ano80d, Ano80f, Ano80g, Ano80h, Ano80i, Ano81a, Ano81b, Ano81c, Ano81d, Ano81g, Ano81h, Ano81i, Ano81j, Ano82e, Ano82a, Ano82b, Ano82c, Ano82d, Ano82n, Ano82j, Ano82k, Ano82l, Ano82m, Ano83b, Ano83c, Ano83d, Ano83e, Ano83g, Ano83h, Ano83i, Ano83j, Ano84a, Ano84b, Ano84c, Ano84d, Ano84f, Ano84g, Ano84h, Ano84i, Ano85a, Ano85b, Ano85c, Ano85d, Ano85h, Ano85i, Ano85j, Ano85k, Ano86f, Ano86b, Ano86c, Ano86d, Ano86e, Ano86o, Ano86k, Ano86l, Ano86m, Ano86n, Ano87a, Ano87b, Ano87c, Ano87d, Ano87j, Ano87k, Ano87l, Ano87m]. Matter [Ano88e, Ano88a, Ano88b, Ano88c, Ano88d, Ano88m, Ano88i, Ano88j, Ano88k, Ano88l, Ano89a, Ano89b, Ano89c, Ano89d, Ano89g, Ano89h, Ano89i, Ano89j, Ano90b, Ano90c, Ano90d, Ano90e, Ano90i, Ano90j, Ano90k, Ano90l, Ano91a, Ano91b, Ano91c, Ano91d, Ano91j, Ano91k, Ano91l, Ano91m, Ano92b, Ano92c, Ano92d, Ano92e, Ano92i, Ano92j, Ano92k, Ano92l, Ano93a, Ano93b, Ano93c, Ano93d, Ano93i, Ano93j, Ano93k, Ano93l, Ano94e, Ano94a, Ano94b, Ano94c, Ano94d, Ano94k, Ano94g, Ano94h, Ano94i, Ano94j, Ano95a, Ano95b, Ano95c, Ano95d, Ano95f, Ano95g, Ano95h, Ano95i, Ano96a, Ano96b, Ano96c, Ano96d, Ano96e, Ano96f, Ano96g, Ano96h, Ano97a, Ano97b, Ano97c, Ano97d, Ano97e, Ano97f, Ano97g, Ano97h, Ano98a, Ano98b, Ano98c, Ano98d]. Matter [Ano98g, Ano98h, Ano98i, Ano98j, Ano99b, Ano99c, Ano99d, Ano99e, Ano99f, Ano99g, Ano99h, Ano99i, Ano00b, Ano00c, Ano00d, Ano00e, Ano00f, Ano00g, Ano00h, Ano00i, Ano01e, Ano01a, Ano01b, Ano01c, Ano01d, Ano01j, Ano01f, Ano01g, Ano01h, Ano01i, Ano02a, Ano02b, Ano02c, Ano02d, Ano02f, Ano02g, Ano02h, Ano02i, Ano03a, Ano03b, Ano03c, Ano03d, Ano03e, Ano03f, Ano03g, Ano03h, Ano04e, Ano04a, Ano04b, Ano04c, Ano04d, Ano04l, Ano04h, Ano04i, Ano04j, Ano04k, Ano05a, Ano05b, Ano05c, Ano05d, Ano05f, Ano05g, Ano05h, Ano05i, Ano06a, Ano06b, Ano06c, Ano06d, Ano06h, Ano06i, Ano06j, Ano06k, Ano07a, Ano07b, Ano07c, Ano07d, Ano07f, Ano07g, Ano07h, Ano07i, Ano08a, Ano08b, Ano08c, Ano08d, Ano08e, Ano08f, Ano08g, Ano08h]. Matter [Ano09a, Ano09b, Ano09c, Ano09d, Ano09e, Ano09f, Ano09g, Ano09h, Ano10a, Ano10b, Ano10c, Ano10d, Ano10f, Ano10g, Ano10h, Ano10i, Ano17b, Ano17a, Ano18a, Ano18c, Ano18b, Ano19b, Ano19a, Ano19d, Ano19c, Ano19f, Ano19e, Ano19h, Ano19g, Ano20b, Ano20a, Ano20d, Ano20c, Ano20f, Ano20e, Ano21b, Ano21a, Ano21d, Ano21c, Ano21f, Ano21e, Ano21h, Ano21g, Ano22c, Ano22b, Ano22e, Ano22d, Ano22g, Ano22f, Ano22i, Ano22h]. Matthes [Sto99]. Maturity [Teh09]. Maurice [Gan02]. Mavis [GH92, Gan04b]. Max [AP94, BR77a, Bal94, CT04, DR94a, Rüs06, DDS20, FHZ15, FZ17, Hoo13, NS22, Oes18, SS17]. Max-Infinite [AP94, BR77a]. max-linear [FHZ15, FZ17]. max-plus [Hoo13]. Max-Stability [Bal94]. Max-Stable [DR94a, DDS20, FZ17, Oes18, SS17]. Max-Sum [CT04]. Maxima [AT91a, AT91b, Ano77e, Ano94f, BR87a, BD75, BV96, CM83, Dek78, Deo73, Eme72, Fan12, Gap07, Gre76b, Has76, Hen82, MS93, O’B80, Pak87, Pis06, RT73, RS06, Sel91, SW97a, THP12, TW83, TO92, ZS04, FPR20, GR14, Pap17, Pit18, SK19, VKR14]. Maximal [Ber11, CT76, Deh85, GMV06, GP98a, Hüs88, Lam70, SGS15, GM20, KL17]. Maximin [AM07]. Maximization [Mi94, Rev93, Sob90]. maximize [LY23]. Maximizes [Eth04]. Maximizing [BM20, BT12a, JK81, KA09, Rig90, Yos84]. Maximum [Abr94, ADLN82, All04b, All10, AG80, Ano81e, Ano86a, AEHS04, Ave85, BdHK93, BG06, BF91, Ber85, BL11, Bra02, BO90, CG97a, Che80b, Coh67a, Don93, DS80, Enn69, FN70, Fer95, FH94, Gal73, Gal66b, Gra85, Gre76a, GL05, Haa92, HTW09, Hea65, Hey69, HH96, HM99, Jen72, KS86a, KP92, Ken76, KS82, KNR05, LS11, MIAPAM04, Mat84, MQ00, Mor68, Mul72, Mul73, New68a, Oga80, Ors87a, Ors89, Pak78b, Ped00, RY99, Ros72, She79, SD76, Spa91, Tam79, Tav81, VT75, VT76, Wei84, Yan74, AI16, AM14, BDL19, GHW19, JF23, KW13, LPN13, PP14, Pim14, Vid20]. May [Sco82]. McArthur [Gan84a]. McKean [BV11, Gir86, Kni71]. McKendrick [WBS16]. MCMC [NP09]. McNeil [Sis70]. McShane [AT74]. Mean [And87, Ano86a, Ave85, BD75, BS79a, Bas10, Bil77a, Blo82, Bri73, BHX98, ÇJ73, CP78, CS80, Coh82, Coh89b, CT76, Dai04, Dar70, Dav78, DC06, DC10, Di 99, Dio74, Eva78, Fen97, GD74, GE82, Gra83, Gra85, Gre77b, Haa92, Hal84, Has76, HS77, Hun76b, HZ82, Jen72, Jun94, KS94b, KM87, Kin65, KS93b, KR97, KPSC08, Lou78, MP70, MS04a, MY01, Mor78, NG95, NE07, Sen67, Sen73b, SS91, Ska71, Ste95, Tho75a, VX04, XY09, BDT22, BS14a, Bon13, BCdlP17, CS20b, DLMR17, GP13, He21, ISK18, KBM21, LGC16, LSW19, Men23, Psa22, RNJ19, VM21, WY22]. Mean-Field [XY09, BS14a, Men23, VM21]. Mean-Guided [HZ82]. Mean-Square [GD74]. Mean-Value [KM87]. Meander [Imh84, Tak95]. Means [Bar74, Eng93, FS80b, Hol68, RS89a, Sud12, GLS19, Leh17]. Measurable [Dob77]. Measure [AI00, Ano73f, CHL98, CD05, DL02, GER95, GPZ05, HZ87, Mol86, MS04b, MW91, Nor86, Pol94, SA73, ST02, VAK92, Yan73, BBvdMP23, CdSJ20, Des15, DGLW17, Jan21, LR13, RNJ16, RNJ17, SRN21, Wiś20]. measure-valued [CdSJ20]. Measurement [Che98, GRD^{+}00, Han67, Hic82, JPR95, Les06, RJ65]. Measurement-Based [Les06]. Measurements [BM66, Lee80]. Measures [Ail68, And03, Ara04, Ath92, Bax81, Bax83, BK12b, Ben68, Con75, Daw81, De 94, FM12, Gre01, Hop75, KN80, KL89b, LC83, Lin73, Mi95, PS76b, San06, Sen71, SE12, SAG73, Sho75, Suo79, Swi76, Szy19b, Twe88, Twe98, VK10, Web07, AR15, Che18, Cri18, ESC15, Iva14, KL22, Nyq17, Sta20]. Measuring [Wil82c]. Mechanical [CH71, Hud73, Wat97]. Mechanism [Boo89, HI01, Ols77]. Mecke [Liu94, Miy95]. Median [Eth04, GL84b, NP09]. Medium [Web99]. Medium-Sized [Web99]. Meeting [Sch85]. Meetings [UW74]. Mehl [Erh00]. Melamed [BB96]. Memories [Ano89f, Han89]. Memory [AHL02, CD92, DP12a, DG76, GM84a, GKL01, KR80, Lin94, Ma03a, Mas77, Nel82, Par86, Tam86, Whi90, BN17, BR18, GT18, KK16, LRM23, NR20]. Memoryless [KS72, Pek03]. Mendelian [KM66, Kha68]. Mendelssohn [Gre86]. Merge [Cow79]. Mergers [Sag99, Sag03]. Mesoadditive [Ham88]. Messages [LG81]. Metal [Lyl88]. metapopulation [MP14]. metapopulations [LM15]. metastability [MPS21]. Meteorology [Kat77b]. Method [AB86b, Bar88a, Cra88, Dal10, Fei93, FDVF93, GR96, He01, Jen66, Kim00, Mar00, Nak06, Not00, Pek96, Pur71, Res66, Sen89, SV03b, Tso71, Vel04, Ari22, BT23, DR15, GR13, Kim17, KR21, Lef95, Lei23, MS23, Wir18]. Méthode [Lef95]. Methodology [CA98]. Methods [Aas85, AF87, Ano64c, BS04b, BCIM01, Kei64a, Ken83, Mao02, RN80, Sud02, Tak64a, MPS15]. Métivier [Jof89]. Metric [Bas10, Dob77, Nis75, ÓT21]. Metropolis [BJ24, DJ09, DLMR17, KS23a, Rob99, RU18, She13]. Michel [Jof89]. Micropulses [Mar11]. Microsatellite [DK99]. MIFR [SS88a]. MIFRA [SS88a]. Migration [Adk69, Bra85, Dur75, Hol68, Mar69, Not00, Not01, RC73, RD97, Sam06, Saw79, Wat97, Whi68, JLX22]. Migratory [Cru70]. MIL [Yan90]. MIL-STD-105D [Yan90]. Mild [New68b]. Miloslav [DS12]. Minification [Kal95, LM91, Lit92a]. Minima [Dal01a, SW97a]. Minimal [Ara04, AJ00, BL89, BB80, BBS85, BEN98c, CF11b, Epl83, Fin04, FSJ14, Har67, HST75, MV95, Nat90, Ros12, She96, SSE87, Des15, ET21, JP18, LR13, Yam22]. Minimally [BM97]. Minimax [BH96, Cor81, Gou98, Pyk94]. minimisation [WG15]. Minimising [Kab83]. Minimization [BP99a, GS99]. Minimize [Pra82, Web82, WVW86, WP80]. Minimizing [Agr91, BF93, BS09, KST99]. Minimum [AEHS05, Ave85, Dan69, Ito80, Kru04, Lán79b, MK73, Rut99, Ten75, Yu17, JF23]. Minority [Ort08]. Miss [FS71]. Misspecified [DK01, McK84]. Mittag [HJL21, JP93, Möh21]. Mittag-Leffler [HJL21, JP93, Möh21]. Mixability [PWW12]. Mixed [Ano86i, Bar86, BMS93, CG12, DH85, Hay00, HS94, JNR09, KS99, Law80, Mah03, PH87, Wie84, BM14, FVZ19, LSW19, MPS15, MR15, Sön23, TTI22, TSN17]. mixed-exponential [MPS15]. Mixing [And84b, And85, AP86, BS97, Cha74, Che81b, DIÖ23, Dzi84, Dzi86, KL06, Las78, Mat90, Tas97, Zaj95, BCS23, HL13a, HM17, LT23, MS21]. Mixture [Ano90h, BLS12, CY08, JP95, LM89, NBS08, Sha81, Zha10, ASB21, AES19, FZL13, KZN15]. Mixtures [Ano79f, AL85, BS04a, Bha80, BLS03, BSW03, BLSW08, BLSW10, Die82, Li05, LS09a, Lyn87, NJS96a, PK80, PG79, Roo03, RM88, Tal69, Yu09, KS20]. Miyazawa [Bré91]. MMAP [VB02]. MMPP [BGT98]. MMPP/M/1 [BGT98]. Mod [Kan75]. Modal [Whi90]. Mode [OS14, Sag20]. Model [Aal98, Aba79, AAKR05, Agu03, Ald95, Ale01, And84a, AS96, Ano67d, Ano77g, Ano82h, Ano98f, Ano99a, ADH93, AG87, AJ00, Avn79, BDM12, BCR11, Bal79, BMT10, BC95b, BC95a, BB05, Ban07, BN06, Bat66, Bat68, BH01, Bec70, Ber65, BJV12, Bha68a, BD88a, Bit69, Boo92, Bov99, Bra81, Bro72, Bro69b, BPR08, BK73b, Cai11, Cer88, Cha01, Cha03, CM07, CL11, CPTW85, Cha93, CW05, CK05, CY08, CM98, Cla05, CS85b, CFD09, CT00, CDD10, Cow78, CC94, Cow01, CX01, DG93a, DD04, Dav81, Dav82a, Dav83, Dav85, DFSD01, De 85a, DT01, DW77, Det94, DDM98, Die66, Dig76, Dob78, Dol10, DWW11, DR00, DZ00, Dun80, DK99, Ebr99, Ebr01, EY11, Eks03, Eks06, Erh00, Ery08, Eth92a, FLP99, FMS04, Fox65, FM06, FRT99, Gås98, Gal71a]. Model [Gal71b, GT05, GMV06, GNN85, Gri81, GP91, Had05, HP98a, HJ99, HT12, HN76, HRF84, Her99, Hey99, Hey74, HK85, HST75, HR87, HS96a, HB99, Hüs95, JS08, Jof82, JK76, Jon99a, JS98, Joy94, Kab83, Kal88, Käm89, KF70, Kar74, KJ81, Kes08, Kin78, Kle81, Kol82, Kol95, KNT10, KL06, KS93b, KN00, Kry72, Kry75, KZ09, Lad05, LR09, Lai77, LM12, Lan05, LN07, LJ78, Law80, LL93, Lee97, LP05, Lew64, LS09b, Lin99, LB88, LS92b, LS74, MZ01, MO93, MW97, Mar69, MM10, MRR02, McC76, McC80, MP12, Möh10, ML77, Mor06, Muc02, NB03, Nat74, Nat75, Nat77, Nea08b, NR84, New86, Nie85, Ols78, Ova01, Pak83, Pap94, Pap10, Peg75a, Peg80, PS02a, PT86, PW96, PW84, PT81, Pit87, Pit90]. Model [PP07, Pla83, Pli75, Pra80, PW81, QR90, Qui91, Råd72c, Rad76a, RG88, Rei67, RZ07, RR67, Rob86, RS94, RR12, Ros73, Rou82, Sah83, SF77, Sav88b, Saw79, Sch96, Sch91a, Sch91b, Sen04, Sev70, Sha73, Shi86, Sie85, Sko00, Sol84, SM72, Sri67, Sta91, Sta99, Tan93, Tan94, Tan05, Tan07, Tan11, Tri92, Tur00, Unk79, VG92, Van81, Vas81, Vas82b, Via95, Voi00, Vrb85, Wan80, WW04, War99, Wat80b, Wat76, Web86a, WK65, WZ65, Wei95, Wit93, Woo83b, XY09, Yeh91, Yeh03, Yeh05, YGN05, ZT01, ZY08, Zuc78, Zuc80a, Zuc81, Zuc86, vD89, AKK20, AB16, ALL21, Bar15, Ber23, BBBN14, BFP17, BPS23, Bra14, CH22, CC15, CS23, CHY14, CR17, CdOR16, CP17, DZ13, DHM20, DM21, DS17, FNP19, Fox16, GJP^{+}21, GXY18, GKP20]. model [HR15, Hös19, HJ13b, JP18, JPS18, JS20, KSBK23, KK16, LGC16, Lee20, LTZ13, LWW18, LWM20, Mah20, MM16, Man13b, MP14, MBM21, Nea16, NNY20, OS14, PL03, PS16, Pri19, Ros18, Ros14a, Roy14, Rug14, SRT^{+}16, SRN21, SB17, Taï96, Thö16, TTI22, Wat23, WBS16, YSL17, Zho15]. Model-Free [BJV12, Tan11]. Modèle [Jof82, PL03, Taï96]. modeled [ACK20]. Modeling [CS82, Fei93, GK76, TG87, AA14, AZA18, CF14, CH13, HE16, LC16]. Modelled [Rou90]. Modelling [Ana88, BPT89, Car99, Fin88, Gao04, GMSB10, Gre81, Hou91, Imo01, Kap10, McN82, Nic86, Oza86, Pia00a, PR96, QL99, Sap87, MC13]. Models [AHP75, ADY07, Abr94, AR05b, Amm83, And87, And88, AB98, AHL02, Ano73g, Ano98f, AEHS04, AEHS05, AB86b, BCL09, Bal83a, Bal84, Bal86, Bal90, BB90, BMY93, BO99, BOP07, BSW07, Bar71, Bar72b, Bar74, Bar88b, Bas71, BZ04, Bat92, BF08, BM09, BK03, Bil77b, BS78, Blo91, Bro88, BG05, BCD93, Can77, Can73, CG87, CF11a, CH10, CL09, Cla94, Cla99, CG69, CNS99, CL92, CPT99, Dav75, Dav84, Dei86, Don86, DW89, Dow98, DGM09, Dur92, Els96, Eth79, FP94, FF11, FG87, Fel76, Fen95, GA93, GY65, Gan65, GY95a, Gan04a, Gao04, GY08, Gav68, Gav76, GESK95, GE82, GG77a, Gne00, God77, Got80, GL84a, GWS80, Gri80, GS85, GHL71, Grü89, GG10, HKY96, Han73a, Has82b, Hav91, Hen81, Hey86, HMPS03, HMPS06, Hin81, HH85b, Hit86]. Models [HF71, HN00, Hol79b, Hou12, HM09, Hut90, ISK95, IPV03, Ito79, Jac99, Jac12, JP95, Joe96, JS98, Kab83, KKL01, Käm91, KS93a, KM65, Kar68a, Kar68b, Kat77a, KS86b, KR88, Kel76, KK69, Kem92, Ker86, Kim64, Kle88, KX11, KS83a, Kro97, KS76, LS87, Law92, LP85, Les10, Li00a, LS03, Lil68, Lin89, Lit94, Liu04, Lot84, Lun96, LS97b, Ma02, Ma03a, Ma03b, Ma04, Mac80, MT95, MW97, McC78, McK84, McN73, Meh77a, Meh77b, Meh79, MD96, MLB05, Möh94, Möh97, Möh99, Mor71a, Mor93, Nea07, O’N96, Oga80, OK86, Oga01, Ōsa85, Oza80, Oza81, Pel94, Pic80, Pic84, Pic82a, Pla04, Pol67, PH86, PRHL12, PW07a, Pru83, PL81, Ren74, Ren77, Ren81, Rén64, RV00, Rid96, RT00, ST01, Sat83, SS07, Sen82a, ST83, SS83a, Sim90]. Models [Sol86, Spe86, SO11, SR70b, Ste03, Sud77, Sud12, Sve95, Swi08, Tak92b, TL69, Tam09, Tha81, TT79, Tur72, Wan77a, WW07, Wat97, Wat80a, Wat64, Wes82, Wie80, Wil77, Won97, Yam89, YC82, dSS04, ABI14, ASB21, BKK19, BEER15, BCKS13, Bor21, Bri20, CK13, CF22b, CdLH^{+}23, Cri18, DHKR19, DNY16, FHZ15, FZ17, FPW19, FF20, GR18, GR14, GA22, GHW22, GdlI18, HJL19, HK19, Ino23, JR16, JF23, Kan23, KADK19, KS17, LC16, LPU20, MPS15, NvdBOdI24, NPYY18, PSK17, Pat13, RNJ19, SST18, Ste19, VM21, YWK14]. Moderate [BL09a, Dav82b, FX16, GY08, Hon02, Li12, LR24, Nyq17, BSX19, FS23, JS20, LX20, SX18]. Modes [Hey02, SD87, Wau74b]. Modification [Arn68]. Modified [Ano87e, Che06, DS80, DKS76, DO85, JK87, MH76a, Pak73c, Rou96, SS89, She96]. Modulated [BR98, BR96, Bre12a, Bre12b, BS92b, Du95, GY08, Iva10, JL98, KS02a, KS04, MT02, OW94, PT94, RS04, Rab06, RB05, Ryd96, Bre13, NP22]. Modulation [AK96]. Module [Nat90]. Modules [GSS11]. Molchan [Fin13]. Moléculaire [Jof82]. Molecular [Bit69, Jof82, MCC03, Orr69, TL69]. Molecule [CCH95, Gir86]. Molecules [Gor75]. Moment [Ano71e, BF08, Bro66, Dal01a, Dur08a, FP97, Gut03, Hör05, Jan01, Kim94a, LBC81, Li04, LS02a, LS09c, Mac80, New70, PW81, Pur82, Reh79, Sen94, Ski67, Ski68, Sto04, Wie80, Wol84, ZL10, AB15, BM20, But19, GM18, RY20]. Moment-Generating-Function [Li04]. Moment-Indeterminate [LS09c, Sto04]. Moments [Ail66, Ano87i, BR00, BCF09, BS90, Bil73, BBS90, Bro79, Cho94, DS94a, Dom95, Don80a, Dun80, Eng93, FY83, FL98, Gat82b, Gha86, Gha89, GW05, GNRW94, GI11, Gra12, Gre80a, Hen86b, Hua16, JG85, Jun94, Kao74, KNS73, KHY03, Kup90, Lag05, Lau80, LV08, Mac81, MC78, McK89, MKN11, NW07, NB03, NSS04, NJS96b, Neu78, Nue10, Pau97, Pin11, Pri19, Rab06, RL88, RS88, ST85, Sgi99, Sie78b, SS83b, Sim00, Ski68, SC82, Ste85, SBG70, Tay71, Tsi07, Twe83, VJO84, Wei75a, Wei84, WF80, Wil88a, Wyn73, Yao85, BS23, BM23, Fro15, Leh17, MST13, Vee15]. Momentum [EV16, EL13]. Money [Dur08b]. monitored [MC13]. Monitoring [Bal76, CCH06, CC08, CFS11, LM04a]. Monoecious [Pol10]. Monotone [Ank06, Ave09, BM01a, BS82b, Coh81b, Coh83, GA99, Lel07, Lin87, Loe09, Mam87, NS84, PS92, PS86a, RT00, Ros07, SS96, Whi82a, Yeh05, DJ18, GN17, KS19, KM23, LL13, Tam15]. Monotone-Separable [Lel07]. Monotonic [FMS04, Kün86, Mac81]. Monotonicities [Kij98]. Monotonicity [Ano91f, Bäu97, BR98, BE91, Bro74, CD95, Du95, EM24, GMSB10, Gue91, Hai92, HF91, How01, LZK06a, Oht91, SY89, TW89, ZAFP06, Kol15]. Monro [MD18]. Monte [AW15, BG06, BSS81, Bro00b, BNS90, DPR19, DR15, FCvS15, Gor70, GH14, IK11, Kam17, NR22, PS19, PW07b, RR07, RR16, RB91]. Moran [Ano82o, Ano89f, Arn68, CS85b, DR00, Han89, HM09, Käm89, Sta93, Wat64]. Mortal [ASC00]. Mortality [Rob92]. Mosaic [Cow04]. Mosaics [Moo74]. Most [Kwe75]. Motion [ADW73, AS12, AC94b, Bre12a, Bre12b, CM97, CMY98, DG89, Don93, DY98, DF04, Dur71a, Dur88, DW92, ESE84, Föl72, FD84, FW10, GP97, GP98a, GN96, GS95, GL05, Had05, HL81b, Hel67, HL99, Hon02, Hör03, HK09, Imh78, Imh84, Imh99, Iva10, JP09, Kar76a, KW78, KS04, KO05, Kol06, Kol07, KMSP12, Lee11a, Lee11b, MIAPAM04, Mar11, MS94, Mat90, NFK99, Ors86, Par85, PS01a, PSZ04, PS76b, PW01, Sar73, Sch92a, Sly07, Smi72, SK86, Wan80, WP97, Was69, Wen86, Wil92, Zha04, AG12, BDM21, Bis23, BM18, BW19, BM20, Bre13, CD13, CM24, FKZ13, FW16, HMBL14, HL13b, KK24, Lac93, LLZ15, Met13, NP22, RRT22, Xu17, Yor92, Zhu23]. Motions [Car67, FI98, Le94, Ors87b, RS06, Cin23, JP18]. Motivated [BRY04]. Motoo [EH95]. Mouvement [Lac93, Yor92]. Move [BP04, Pha94, PPD97, PP98, Rod95a, Rod95b]. Move-to-Front [BP04, Pha94, PPD97, PP98, Rod95a, Rod95b]. Movement [CHPT97, Fox65, Gel93, HT91, OC03, SCZ77, Wat97]. Moves [Ori79]. Moving [ASB94, Bar75c, Cor81, DS86a, GJ88, God77, Hel71, Kul06, LS87, LL77, McD79, McK81, McK88, Pea67, PT86, Tho86, TW84, Web86b, YZ85, ZS04, SWZ18]. Moving-Average [LS87, LL77, McK88, PT86, Tho86]. Moyal [Gan98]. MR [Bäu97, SDH94]. MR/GI/1 [Bäu97, SDH94]. much [PS18]. Multi [AR05a, Ano78g, BJV12, Bha68b, Bui75, Cra74a, Dal71, Dav76, DJ09, DHKP18, Don76b, GL99, GM73, Hin81, Jon99b, Lou73, LTZ18, Mod67a, Möh10, NR78, O’C98, PP98, Qui70, Sad01, Tan11, WW07, Wei70, Wei72, Whi90, Wol77a, XMKW90, Zuc89, DR15, JP22, LMS18, PT22]. Multi-Actor [AR05a]. Multi-Allelic [Möh10]. Multi-Asset [BJV12, Tan11]. Multi-Buffered [O’C98]. Multi-Channel [Lou73, Wol77a]. multi-colour [LMS18]. Multi-Dimensional [Ano78g, Dal71, Dav76, GL99, Mod67a, Sad01, JP22, PT22]. multi-level [DR15]. Multi-Line [GM73]. Multi-Modal [Whi90]. Multi-point [LTZ18]. Multi-Priority [XMKW90]. Multi-Proposal [DJ09]. Multi-Request [PP98]. Multi-Risk [WW07]. Multi-Server [Bha68b, Cra74a, NR78]. Multi-Species [Hin81]. Multi-Type [Bui75, Don76b, Jon99b, Qui70, Wei70, Wei72, DHKP18]. Multi-Unit [Zuc89]. Multichannel [Dsh89, Wol87]. Multiclass [Gra08, HKN89, Kul86, KG02, LAF02, Ye07, DO16, PW16]. Multicolor [BCPR11, Ery08, Gou97, Bor21]. Multicommodity [KS10]. Multicomponent [FGL00, Hei01]. Multiconductance [BMY02]. Multidimensional [BR00, Bau81, Bau83, Bau84a, Bau84b, BL10, Bor95, Cab84, Dol10, GMM06, Ish77, Kap74, Ros80a, Tan97, WZ07, BM18, BN15, Hua16, KM15b, MS23]. multifactor [AKK20]. Multifocal [EV82]. Multifractal [BDM12, MS04b, ST02, Att19]. Multifractional [Sly07, HLMOL21, Mar13]. Multigene [Gri92]. Multigrade [McC78, McC80]. Multigraphs [BZ08]. Multihit [HKY96]. Multilag [Peg80]. Multilinear [Fre07]. Multimodality [DK11]. Multinomial [Bof65, CL85, Gri98, KM66, KT83, Peg75a, LM24]. Multiobjective [DP10]. Multiparameter [CW81]. Multiparticle [Bur76]. Multiperiod [BG05]. Multiphase [BK73a]. Multiple [AKM10, Ano74g, ASC00, BP97, BFS09, BY97, BF97, CPTW85, CG90, CK00b, Dow98, FH10, GM91, Gli00, GIM08, Got87, HS74a, HGR90, Hol80a, Ish04, Kal89, Kim94b, KX11, LMS18, LR80, LLPC94, MD96, RS69, Rob72, Sag03, Sch88b, She92, SD12, Wol84, KS17, Mos18, WGF13]. Multiple-Access [BFS09]. Multiple-Outlier [KX11]. Multiple-Server [Wol84]. Multiple-Threshold [CPTW85]. Multiplexer [LNTZ99, XS03]. Multiplexers [BD94b]. Multiplication [Lit92b]. Multiplicative [HS72, HS74b, Law92, Moy64, Tam10, MA14, vdHKPS23]. multiplicative-decrease [vdHKPS23]. multiply [LL13]. Multiserver [BE91, Emm72, Git81, Hal79a, HZ87, Har90a, HW84a, Hem90, Mar05, NE81, NE82, SD87, Sig88]. Multistage [Sud12, Vis86]. Multistate [Ant01, BS82b, ENPS78, FC02, FC03, Gri80, HNF85, IA07, NS84, Nue10, Pie07, Yeh05, BAR14, GN17, Nat14]. Multitrap [And84a]. Multitype [Ano91h, Ano92f, Ano92g, Bri98, CG05, Cla94, Coh89b, DS02, Fen97, GR87, GR89, Gad89, Gol78b, GMM05, Hei09, Hop75, JW85, JL06, Jon04, Kle89, Kle91, Lan05, LY95, Lot84, MK17, MV86, Mur80, Nan82, QD77, RW21, Rad73a, Rah98, ST86a, Spa91, SO11, Wei74, Wei71, BC23, Hon16, MK15]. Multiurn [Sev70]. Multivariate [AP94, Ano74e, Arn75, ASC98, BK12a, Bal99, BMY02, BR02, Ber77c, BM82, Bla94, BC01b, BL12, BN88, CL05, CS96, CS00, Che76b, DM75, Dsh01b, ER89, Ebr97, FM12, GM84a, GR96, Gri73, Hos06, Hu96, HM15, JR82a, JK76, LS09a, LS86, MS91, ML67, Nay87, Nic84, NO85, PD20, PT23a, Pru83, RR10, Res04, Roo03, Rüs04, Sav85, Sav88b, Sca98, SS88a, SS91, SG92, Sob67, SM94b, Wei75c, Whi82a, Won97, ZS04, AA23, BSC18, BHB18, FCP18, FPR20, Hen18, JP18, PKO16, Yin21]. Munro [McL76]. Must [Rus83]. Mutant [Wat82b]. Mutants [FTV90, Man74]. Mutation [Esh73, GY65, Kin78, Man74, CS23, Tsu17, Tsu18]. Mutations [CC01, Hin82a, Hin95]. Mutual [KD78]. Mutually [Ada75, Ano75e, JLX22, Mas86a, Pol75]. Mutually-Exciting [Ada75]. My [Mod67b]. Myopic [BS92a, DGK98, FB88].

N [RS90, Ski67, CG96, GBG02, TTH93, Abr01a, Abr01b, CT76, HL04, Rig99b, Sim97, Sim98, Sim00]. N-Like [BB12]. Nanocomponent [EY11]. Nanocomponents [Ebr11]. Nanosystem [Ebr08]. Narrow [HT71, Has82b]. Narrow-Band [Has82b]. Nash [CF17]. Natural [Mod66, Mod67b, Mor71c]. Nature [CG87, Hun86, Kha68, VP01]. Natvig [Nor86]. Navier [He01]. Navigation [FM06]. NBU [FRR01, HX02, LK01]. NBUC [Ano92h, CW91, HMM93, HX02]. NBUE [BS78, KÖ89, KM03]. Near [Ano91h, Ano92f, Ano92g, JK04, Kle89, Kle91, MMR07, BP16, Est76, Gaá20, JKL^{+}20, RR75b, VKR14]. Near-Critical [Ano91h, Ano92f, Ano92g, JK04, Kle89, Kle91, MMR07]. near-lattice [Gaá20]. near-neighbor [JKL^{+}20]. Nearest [Bar71, Bar72b, Bar74, Bau81, Bèr04, Bes72, Coh94, DH89, GR94, Hen82, Hen86c, Ken90, LA01, Pic82b, Ren79, ST86b, Str75a, BS14a, CHO22, GHW19]. Nearest-Neighbor [Bau81, ST86b, Str75a, BS14a]. Nearest-Neighbour [Bar71, Bar72b, Bar74, Bèr04, Bes72, Coh94, DH89, GR94, Ken90, GHW19]. Nearly [And85, IPV03, LL78, LL90, Mal67, ML98, PRS08, SW11, WW20]. Nearly-Completely [LL78]. Necessarily [JGR91]. Necessary [CCFG96, CW86, Har67, Hul82, LT95, MSCZ78, Mor73c]. Necessity [BL21].

Necklace [MS08, Nak09]. need [PS18]. Needle [Dia76]. Negative [AD94, Ano70d, Ano10e, AGC99, BG06, BNY69, Bau09, BV11, BS04a, CP82, Coo98, CP11, Det96, Ebr87, EZ80, FMS99, Gel91, GGS91, Gre84, HP93, Hen93, Hut90, Igl74, KS86c, KLP12, KO12, Loe09, MP92a, Mar07, MP92b, NC72, Neu94b, Par85, Phi81, Sur08, Twe71, VS00, VU09, WX08, Whi79, Whi80c, YC01, ZW02, Zho07, ZZ04, BS23, BM15b, Fro15, KMG19, KW13, LS19, LYZ17, LVZ19, MV16, MR23, NPYY18, Vid19, WZ18]. Negatively [BS04b, CCN10, GN06, AK23, Wir18]. Negri [IN86]. Neighbor [Bau81, ST86b, Str75a, BS14a, CHO22, JKL^{+}20, PP14]. Neighbors [Pic82b, BK21]. Neighbour [Bar71, Bar72b, Bar74, Bèr04, Bes72, Coh94, DH89, GR94, Hen82, Hen86c, Ken90, Ren79, GHW19]. Nelson [Let78]. Nested [Hos86, RB91, Sun92]. Net [Bro72]. Network [Ary72, BKT97a, BKT97b, BN94, BS94, BLT11, BLL84, CSZ97, Cao89, CK90, CFD09, DS08a, Dar89, GH10, GM84b, Gra00, Jac78, Les06, Mad75a, Mas84, Mas86b, MRR02, Nak74, NMG00, PW96, Rab06, RV00, SM05, SY86, Ton85, Wal83b, Ye07, FNP19, GPSY16, PS16]. Networks [Ano90g, Ano05e, AT98, APZ11, APS04, Bac86, BFL05, Bam92, Bar82a, BM89b, CPS96, Cha97, CZ98, Che96, CHPT97, DS02, Dei10, DG06b, DGM09, DYY01, Eco00, Eco03, FG98, FM06, Gel91, Gel93, GXS12, GMRY07, HRZ07, HL81a, Har81, HP95, Hem90, HT91, Hen93, HLT00, Hun95, Kan85, KK96b, KS94a, KW96, Kel97, KW99, Kel00, KS02a, Kel75, KW89, KLO99, LR80, LF08, Lel05, Lel07, Lem87, LR07, Mas87, MF01, McK89, Mil08, Mor06, RWN02, RS92, SD87, Sha89a, SY89, Sig89, SD12, Sto08, Tet98, Top86, TK84, TW89, Van91a, Van91b, Wal82, Wal83a, Wal11, WB03, Wei05, Yuk06, ZZ99, Zac00, Zac07, Zie87, ACK20, AN22, AR13, AA14, AA17, BDG14, CHJW24, CM16a, CMW16, GM15, PW16, Sal13, SDH16]. Neural [AT98, BDG14]. Neuron [CG69, HF71, LS74]. Neuronal [CT00]. Neutral [Ano77g, CG76, Eth92a, Gri81, Möh99, Möh10, Saw79, Sch91b, Wat76]. Neutron [Mul68a]. Never [Jac09, Sud85]. Neveu [KZ95]. news [Mah20]. Next [Hal89, PP14]. Next-Fit [Hal89]. Nielsen [GNN86]. Nikodym [BD75]. Nikolaevich [Par88a]. Nilpotency [Ano66e, Dow66]. No [CM98, FA85, ABGH00, Ano89e, CCFM17, CNS99, Gri81, Mor71b, Twe88]. No-Aging [CM98]. No-Wait [FA85]. Nobile [IN86]. Node [BJM93, SM05]. Nodes [Hem90, SD12, ABD17, MW15]. Noise [AAL77, AHL02, AS09, BC01a, BD12, Cai03, Che98, Dal71, DJ05, Dri75, DT11, Han67, Heu87, HM95, Huf87, JFS12, Kel09, Kul06, Lan84, LPS06, LS91, Lun96, LBP99, LMX04, Mar73a, McC97, Pap71, Pré94, Sly07, Smi73, Sob67, SW94, Wol75, BN15, Hoo13, IR20, KA16, PT23b, Spi15]. Noise-Distorted [Sob67]. Noisy [JS08, KSC12, RWN02]. Non [AML84, And84b, And85, Ano72f, Ano73e, Ano82g, Ano91e, Ano91g, AHO20, ANT78b, Ary72, Ath92, AC94b, BM97, Bar71, Bar72b, Bar74, BZ04, BSS72, Blo91, Blo86, BGH71, Bre81, Can77, Can73, CA98, Che98, Che76b, CP88, Dav82b, Det96, DG93b, DM72, DF81a, EE76, FI87, Fei94, FI98, Fen95, Fol86, GA93, GJ78, Ger90, Ger91, Gir86, Gla78, GM84b, Got87, Gre84, Gri75a, Gri75b, GRRT97, GKRV94, Han71, Has64, Hen83a, Heu87, Hey82b, HH85b, Höp82, HR88b, Hor01, Hun86, Hüs86a, Hüs88, Hüs93, Hut90, IK06, IKP01, JP99, Jir76, Joh06, KS23a, Kar68a, Kar68b, KP89, KS94a, Kel00, KS86c, Kim94a, KL77, KS80a, KLO99, Las94b, LU95, Les89, LS92b, Loc82, LMW00, Mac81, MJ85, MP92a, Mar69, Mas86a, Mas77, MKR91]. Non [Moo73, Neu94b, Nie85, Oht91, Ols77, Oza80, Oza81, Oza86, OI01, Paw89, Pea65, PM99, Pfe82a, Pra82, PH86, Pur73, RG88, Ren79, RJP97, Rob72, Ros78, Rot95, Rud75, Sav88b, SAI82, Sen87, Sév95, Sha88, Sha89b, SC82, Sob67, STC93, Sta81, SW94, SG88, Szp85, Tho85, Ton73, Ton74, TV88, TV93, TW89, Twe71, VS90, Van91a, Van91b, VS00, Vas81, Vas82b, Vas84, VT89, VGT90, VP92, Ver98, Weg76, WP80, Wei75c, Wes77b, Wil96, Wit93, Zei89, Ger23, KA16, LOR19, MW21, NFS20, RW21, Yam22, Pic79]. Non-Additive [Che98]. Non-Asymptotic [DF81a]. Non-Central [Che76b]. Non-Closure [Les89]. Non-Comparability [AHO20]. Non-Constant [Ano73e, BSS72]. Non-Cooperative [KLO99]. Non-Critical [Höp82]. Non-Degeneracy [Ath92]. Non-Discrete [KL77]. Non-Ergodic [GM84b]. Non-Ergodicity [FI87, Sen87, Szp85]. non-exchangeable [NFS20]. Non-Existence [LU95, Loc82]. Non-Explosion [LMW00]. Non-Explosivity [RJP97]. Non-Exponential [VS90, Van91a, Van91b, Wil96]. Non-Fatal [Sav88b]. Non-Full [STC93]. Non-Gaussian [BZ04, FI98, OI01, KA16]. Non-Homogeneous [Ano91g, Ger90, Ger91, Gri75a, Has64, Jir76, KS94a, Paw89, SAI82, TV88, TV93, Vas81, Vas82b, Vas84, VT89, VGT90, VP92, Wes77b, Zei89, RW21]. Non-Homogeneously [GJ78]. Non-Identical [EE76, Pra82, WP80]. Non-Identically [Wei75c]. Non-Identifiable [Can77]. Non-Imbedded [KS80a]. Non-Independent [Sha88]. Non-Intersecting [Mas86a]. Non-Irreducible [AML84]. Non-Linear [Ano72f, Bar71, Bar72b, Bar74, DG93b, DM72, Fen95, Gir86, Han71, Heu87, HH85b, IKP01, LS92b, LMW00, MJ85, Mas77, Oza80, Oza81, Oza86, PH86, Pur73, Rob72, SC82]. Non-Local [SW94]. Non-Markov [Ano73e, BSS72, Gla78, RG88]. Non-Markovian [Gri75b, KP89, Sha89b, TW89, LOR19]. non-minimal [Yam22]. Non-Monotonic [Mac81]. Non-Nearest-Neighbour [Ren79]. Non-Negative [Det96, Gre84, Hut90, KS86c, MP92a, Neu94b, Twe71, VS00]. Non-Normal [JP99]. Non-Normality [Dav82b]. Non-Optimal [Las94b]. Non-Overlapping [Ano91e, Can73, CP88, Joh06]. Non-Poisson [FI98]. Non-Product [Kel00]. Non-Random [Got87, Kar68a, Kar68b, Sta81]. Non-Recurrent [Pea65]. Non-Renewal [Blo91, Hun86]. Non-Reversible [HR88b, KS23a, Ger23]. Non-Serial [Ary72]. Non-Singular [ANT78b, Nie85]. Non-Singularity [Blo86]. Non-Standard [Hen83a]. Non-Stationary [Ano82g, BM97, BGH71, Bre81, CA98, Fei94, Fol86, GA93, Hey82b, Hor01, Hüs86a, Hüs93, Kim94a, MKR91, Moo73, PM99, Pfe82a, Ros78, Rud75, Sob67, SG88, Tho85, Ton73, Ton74, Weg76, Wit93]. Non-Strong [And84b, And85, MW21]. Non-Symmetric [Ver98]. Non-Symmetrical [Mar69]. Non-Uniform [Hüs88, Ols77].

Non-Uniformly [Sév95]. Non-Uniqueness [GKRV94]. Non-Zero [AC94b, Bar74, Rot95, Pic79]. Non-Zero-Sum [Oht91]. Nonasymptotic [Spi15]. noncoexistence [BB14]. Nonconservative [Sam06]. Nonconstant [Pon07]. Nondecreasing [Bon13]. Nonergodic [SDH16]. Nonexponential [YZ06]. Nonhomogeneous [BR02, LMM10, HMY17, LST19, Ros18, Tan17]. Nonidentical [BP01, ZLB08]. Nonidentically [BN06]. Nonidentifiability [RCLW10]. Noninteger [FS07, FSW12]. Nonlinear [BJ07, Del98, FW10, LPP12, Mon67, Neu23, Sol96, ZWV10, BDL19, Eri16, GW17, KM15b, Sim19, SB17, Zhu13]. Nonnegative [BDY07, KT08, PT11, VT11, ZT01, PP14]. Nonneutral [Fea02, HM09]. Nonoptimality [AA06, CCHY08]. Nonparametric [CG00a, Fuj13, ISK18]. nonrepairable [Nat14]. nonreversible [HM17]. nonsingular [BR18]. Nonstandard [SRT^{+}16]. Nonstationarity [LP05, PM03]. Nonstationary [CCdO05, Deo73, GP04a, LAF02, CH13]. Nontrivial [GS01]. Nonuniversality [BvdHK17]. Nonzero [GHL05, CF17]. Nonzero-Sum [GHL05, CF17]. Norm [Hol00, GW17]. Norma [Gan84a]. Normal [BL73, BG05, CFS11, Dan82b, DR81, GR96, Hal79b, Hen81, HH96, JP99, KR97, KSF80, Lan81, LS09c, ML67, NO85, Pol78, QM86, RR10, Rub77, Tuf99, YI99, BPS23, CL14, LPN13]. Normalised [Mac81]. Normality [BR89, CCC03, Dav82b, Eng80, MPHL10, MI83, NR20]. Normalization [Tet98]. Normalizations [BS79a]. Normalized [AD94, Chu18, FS23]. Norms [Tan83a, Oes18]. Northeast [KL06]. Norton [Wal83a]. Note [Aas76, And81, And88, Ano73l, Ano02e, Ath71, BT73, Bal86, Bal90, Bar75b, Bau83, BOR01, Ben70, BJV12, Bin69, Bla87, Boe69, Boo92, BG79b, BH73, Bru84, Cam73, Cha04, CG12, CP03, CH90, Dav78, Dav83, Dav85, Dav64, DSR80, Don76a, Don77, Don83, Don84a, Dos85, Dow67, DW85, Dsh01a, DH08, Dun80, Erl67, FHT78, Fin69, Fin10, FS98, Gad89, GT86, GG87, Gat84b, GS72, Gra71, Gre01, GS85, GPZ17, HS69a, HL81a, Hem90, HF70, HRF83, HIR03, Hic82, Hic86, HS69b, Hol08a, Hon04, Höp85a, Hsu79, HS77, Hüs93, Hüs95, HS96b, JRR95, Jer70, Kao74, Kap75a, Kap75b, Kap82, Kat02, Kat12, Kea67, Ken72b, Kla02, Kol07, KS76, KA03, LC83, Leh71, LMS76, LP00, Lin94, Lin79, Llo71, LS79, LT04b, Ma97, Mac81, Mar86, Mas77, Meh77b]. Note [MU11, Mil67, Mit04a, Miy95, MP04, Mor66a, MS85, MS05b, MN82, Nar68, Naw83, NRS85c, OL65, Pak70, Pek99, Pha69, PSW71, Pha74, Pha79, Pit77, Pol90, PZC07, Pou90, Qiu91, Rad74a, RCC12, RT91, RS83, Rig99b, Rob99, Roe85, Roo76b, Ros81a, Rud76a, Sap75, SLK79, Sch03, ST83, Sev70, STW94, Sis70, Ska71, Smi72, SD76, Spe73a, SZ93, SM82, Swe88, Tal80, Tam70, Tho75b, TW75, Tur06, Van96, Van91a, VS00, Voi00, Wal83a, Wes77b, Wes77a, Wic75, Wil97, Wil88c, YC01, YA88, Zac07, ZS92, ZJ04, BFK16, DFV15, DN13, FZL13, Gol23, KT15, LGC16, MA14, Tav22, VR22]. Notes [Ano90g, EE93, HS74b, Hol79a, Nor86, PS84, Sig89]. Notion [CG00b, EP95]. Notions [BKS02, Li04]. Nowhere [Dav73]. nth [Ski68]. Nuages [Ray70]. Nucleation [QL99]. Nucleic [SS85b, ZS92]. Nucleotides [Wat82b]. Null [ABI17, AG07, GP91, IC82, CTKR83, Eri16, dL23]. Nulle [CTKR83].

Number [AHBT82, Abr94, AFGP91, Ano91e, AM10, Ave85, AIH87, AIH88, Blo82, BP01, CP88, Coh67a, DY97, Eng64, Eng80, Enn69, FG84, Fel66, FTV90, Gal73, Gal66b, Gat82b, Gri81, Gri92, Hal82b, Has76, Hol76, Hol08b, JW85, KP92, Kel83, Ken72b, KO05, Kup90, Law92, LM99, LH02, LP94b, Man74, McN74, MI83, Moh76, Mor66b, MZLSRS02, NH82, Nak84, PP69, Pet83, Pod72, Pol78, Red03, Rig90, Roe70c, Rus83, RS10, SLK79, SK91, SP92, SSE87, Tak88, TSCD08, UW74, UB74, Var80, WJ95, Yas84, ABD17, Att19, CR17, IJB18, IK18, JP18, JX13, KS13, PS19, Sau23]. Numberings [Har67]. Numbers [CCC03, FG08, GT08, Hos06, IA07, JW82, LJ78, Ver98, WS96, SUY21, CCN10, Hey75]. Numéraire [GER95, BP99a]. Numerical [ADW73, BY14, BRY04, Del98, Del78, Dob74, GC74, GKRV94, Mat04, ZG88]. NWU [CK00a]. NWUC [Ano92h, CW91]. NWUE [KÖ89].

Obituary [Ano89f, Ano90m, Ano94l, Ano00j, Bin92, BS01b, Coh81a, Doo72, DMR02, EH95, Gan71, Gan84a, Gan84b, Gan98, Gan02, Gan04b, GS08, Han89, Jof89, Kal96, KH96, Ken70, Mor73d, Olk77, Par88a, Roh88, Sen17, Sne05, Sto99, Tur92, DS12]. Object [Hel71, KA09, MT06, Nak86b, Web86b]. Objective [HK90]. Objects [Yas84]. Observable [Led05, Mas93, Mon82]. Observables [Hud73]. Observation [All04b, Gil87, HH85a, IKP01, KR00, MZ93, Maz86, Sch75b, Ser90, Sch76]. Observations [FZ93, Git73, HTW73, Her99, HS69b, KW03, KA09, Neu67, Pet81, Pet83, SKK06, Swe94, Wei95, BS16a, BEG22, BEER15, GW17, VKR14]. Observed [BMY93, JKM12, KPS04, LT95, Wak91, Yos84, Rah17]. Observer [DS02]. Observing [CG02, CG05]. Obstacles [KN01]. Obtained [Ano85f, BT82, Dah85]. Occupancy [Amm83, ALRW99, Cha83, DNW07, Fan82, Hol81b, Jan83, JK76, Khm11, Nak08, OC03, Swe88, GKP20]. Occupation [Ber70, Bon81, Bre12b, DF04, GT08, Hsi76, Ped71, Jan21, Xu17]. Occupied [Sud76]. Occurrence [Lai77, Pla85, SPS80, Ste99, Van96, Yeh97]. Occurrences [Ano91e, CP88, Gol05, IA07, Mak10, RD99, Sch88b, Ste03, WJ95]. Occurring [BJ07, Bha65b, Jen74, Pha91, Pod72]. Occurs [Blo82]. Odds [AKM10, Tam10, MA14, MA17, vLR13]. Off [LS07, Ano95e, DF04, DS08b, FI87, Gui94, KB91, Lac05, LM04b]. Offensive [Cow92]. Offered [The79]. Offers [BF97]. offline [KSC13]. Offspring [Ber11, Can74, Fel66, KN75, KS02b, MM09, SG89, Ber13b]. Oil [BB88]. Okazaki [BB06]. Old [Man74]. Oliveira [Tur92]. Omitted [Gre86]. On-Line [LM04a, Pre06]. on-off [Ano95e, DS08b, Gui94]. On/Off [LS07]. Once [All12, YYZ09]. Onchocerciasis [Sap87]. One [Abd87, AH77, AT70, Ano86j, Bal88, Bèr04, Boo92, BHPZ06, Cab84, CS86, Cox77, Dav82b, FG98, Gel77, GNR89, GG91b, HE16, IK11, Imh78, Ito80, IM03, Kab83, Kar76b, Kas66, KS06, MP00, McC74, Mul68b, Nak86b, Ori79, OT77, Pau87, PP71, Pra70, Pyk94, Rog00, SCZ77, Sen67, Sha89a, Sud76, Tak71a, Tam10, Tzi11, Ver69, Ver98, Yas88, Zha92, Zie87, ZG88, Cha17, CM15a, DvdHS24, Iva14, JSW19, MA17, MZ20, Pat13, RWZ13, VR22, Yam22]. One- [Gel77]. One-Compartment [Boo92]. One-component [HE16]. One-Dimensional [Bèr04, BHPZ06, Cox77, FG98, GNR89, IK11, Ito80, Kar76b, MP00, McC74, Pau87, PP71, Sha89a, Tzi11, Ver98, Zha92, Zie87, ZG88, Cha17, CM15a, VR22, Yam22]. One-Sample [Dav82b]. One-Sided [Ano86j, CS86, Imh78, IM03, KS06, Pyk94, Rog00, Iva14, Pat13]. One-Step [Yas88]. One-Step-Ahead [Kab83]. One-to-One [AT70]. Online [ACSS11]. Open [ABGH00, BMY93, Bar82a, CDD10, O’N96, Rig97, SV71, Whi68, WZ13]. Open-Loop [ABGH00]. Operating [RS90, WR97, WK94, HF19]. Operation [Lit92b, Pli75]. operational [FHC18]. Operations [Sam70]. Operator [Mas84, Whi85a]. Operator-Analytic [Mas84]. Operators [AL00, Att89, HJL21, KL16]. opinion [BVW21]. opinions [SWW21, SK23]. Opponent [KS05]. opportunity [Lin14]. Opposite [Lit74]. Optical [Mor06]. Optimal [Aba73, Aba79, AHS78a, AHN90, AH00, Abr07, AKLL18, AT74, All04a, AM08, AI89, AM14, ANW82, Ano88f, ASC98, AGSC00, AB86a, AG87, Ave09, BS92a, BL13, Ber94, BMP87, BCS93, Boy69, Boy70, BS22, BS73, Bru87, BL03, Cha04, CL11, CCFM17, CNO^{+}84, CSZ04, CY16, CY08, Chr12, CF13, CdSJ20, CJ08b, Con13, CM24, CS97, Day70, De 85b, DGK00, DLU09, DLR80, DW77, Dip98, Dob77, DGK98, Dur75, DLMR17, ES11, EW09, ELTW10, EL13, ENPS86, Emm72, EE76, Esh73, Fad74, Fas90, FW91, Fel76, FP66, Föl72, Fos75, Fro93, Gap07, GR14, Gil87, GY04, GADL04, Gne07, GI11, GL84a, Got85, Gou98, GP97, GP98a, GCS07, Guo01, GS01, HR00, Hel71, HW84a, HH85a, HK90, HØ07, IKP01, Irl06, JFS12, Jia15, Kab81, KR80]. Optimal [KS77, Kar09, KS05, KA22, Ken82b, KS10, KL03b, KN01, Kim16, KOH23, KR00, KW85, KLP12, KA89, Kyr99, LR00, LMS72, Las94b, Lew01, LAF02, LT95, LD10, LY23, LR04, LM04b, Loe09, Loi05, MT06, Maz86, MW96, McN85, Men96, MM94, Mi99, Mi02a, Mi02b, MDG09, MWCY01, Mon82, Mor71b, Mor71c, Mos18, Muc02, MR01, Nak84, Nak91, NE81, NE82, NR11, Nue08, OKM86, PH82, PV12, Ped00, Pli75, PZ90, Qiu91, Rie76, Rig96, Ros69, RZ08, Ros14a, Rot95, Rot98, San06, Sar73, SW05, Sch98, Sel97, SE12, She13, She92, ST12, Sie85, SKK06, SG93, SZ06, SZ91, Sta74, Sun92, Tam82, Tam88, Tam01, Tam09, TT86, Ven79, Vil07, Wak91, Wan76a, Wan77b, War76, Web78, Web86b, Whi73, WR74, Wic75, YZZ21, YWR82, Yao07a, YCW15]. Optimal [Yas88, Yeh85, YH87, Yeh91, Yeh95, ZY70, ZC95, ZX17, Zha94, ZC14, Zuc78, Zuc80a, Zuc81, Zuc86, Zuc89, AS17, BS15c, Buo15, CCdOS15, CCCSdO24, DHKR19, EG15, EFS14, FHC18, FPS18, GA22, Kli14, KSC13, LOR19, Mar16b, MA14, NPYY18, Tam15, Lef89, SC02]. Optimale [Lef89]. Optimality [AH77, All12, APZ11, BS06, Ber76, Ber80, BR16, CCPW92, CR94a, CSYZ05, FT78, FB88, Ger76, GZ06, Had05, HP98b, JFHL09, JCM98, Käm87, LP00, LT94, MT12, Rob80, VPT92, Win77, Yao07b, YC08, Zhe91, ZPR08, WG15, Zha14]. Optimisation [Whi82b]. Optimization [ABSGS16, BBC01, Cip88, Dip98, DG99, Hei01, LBL05, Loc96, Men76, RB05, Wal11, WZ18]. Optimized [Sie81]. Optimizing [Swe94]. Optimum [AH86, Ano88g, AIGR06, KKLO86, NO74, RG88]. Option [BN02, DIJ04, Eks06, GER95, LR12, Lu10, MKN11, RS95b]. Options [BRY04, BBMB08, Cai11, DW11, Dol10, Eks03, Eks04, EJT05, FS98, FPY04, Ger11, GS01, HW02, Iva07, SS07, Sch03, Tan11, EM24, Ges79, Kol15, Pat13, BJV12]. Order [Ahs76, AAO90, And03, ALS04, Ano75e, Ano82h, Ano85g, AM05, Bal72, BLR91, BT92, Bla94, BCS93, BENP94, BP96b, Bra83, Byt81, Dal87, DG74, DC06, DC10, DH85, Deh84, DJ95, DG93b, DH08, DMT11, Dvu84, Dzi84, Dzi86, Fei79a, Gal72, Gal74, GXS12, Gil87, GP91, Gup84, Hen81, Her03, HG80, HH85a, Hu96, Hua74a, HCK02, JP93, Kad74, KR80, KK16, KT01b, KT01a, KLM04, KX09, KX11, Lam68, Law70a, Law80, Law92, Leh73, Li04, Lib85, Ma97, Mar86, Meh67, Mil76, Mul72, Mul73, NSK08, NJS96b, NS06, NO85, Pap96, PM03, PT81, Pol75, Pur82, RS04, RS69, Ren88, Rip76, Ris86, Ryd96, Sen95, Set03, SK90, SAG73, Sim90, Sim93, Sri86, Tak64c, Ton75, Wan92, Whi80b, Whi82a, Wil02, Yun00, ZT01, ZL10]. order [AHO20, AZA18, BN17, BMRRS16, BN13, Gaá20, GT14, GM18, HJL14, JR16, KK18, Kol15, LWM20, LSW19, Ott20, PL17, Sim14, TK15, Wir18, Xu17, Yam17]. Ordered [CV09, DS06, FC03, GNRW94, Mak10, Naw83, PS86b, Tak77, Tru00, WVW86, BV14, YBL23]. Ordering [AKP91, ANW82, Ano85e, Ano93g, Bac91, BLR91, BR02, BM01a, Blo84, BPT92, BENP94, BSC04, CY08, CS95, DS06, Ebr90, EP95, Epl83, FP05b, FP05a, GMRY07, IG01, Irl03, KS94b, KS91b, Leh86, Li96, Lin95, Ma97, Mak94, MY91, MPSG90, NS06, Sha82, SS87, Whi85b, YL16, Yu09, ZLB08, HR15, HCY21, KK18, RNR18, YCW15]. Orderings [Bha91, CW97, FPKMP98, Huf86, NJS96a, Sca98, Ton97, BBH14, BHB18, Yin21]. Orderliness [AT79, DM75]. Orders [BSC02, Bor92, CC79, DC06, DM10a, GS99, SW97b, SSSL10, Won97, BPS23, Leh17, Zha21]. Ordinary [Bon89, Buc94, Kur70, Kur71, Sum87, KP22, LZ14]. Organizations [Stö74]. Organizing [CL97, Lam84, LLS84]. Orientation [Nag93]. Orientation-Dependent [Nag93]. orientational [Gaá20]. Oriented [BS97, GWS80, Ito79, IMT00, AG16]. Origin [Hit88, Jac09, RR75b, CMV23]. originating [JR16]. Orlicz [SST18]. Ornstein [Bee75, BSDG20, BWY11, BDY07, CM97, Dir75, DM72, FLP99, JP99, Jia12, Lac05, Li00b, LS20, LS00, Mia13, NRS85b, Ped02, QDZ21, Ren87, RS87, RS88, Tho75a, Tho11, Vee15, Wan12, Wen90, XZW09, ZZ16, ZZ09]. Orthant [BS04a, Lee11a, Lee11b, Mor86]. Orthogonal [GP99, RMV97, Sen01, Wat88]. Oscillating [KW78, KS86b]. Oscillations [HG80, Ott20, Pap10]. Oscillator [Luo00, MS23]. oscillators [BCN24]. Oscillatory [Has79]. Other [Ma03b, MW97, GK22, Tu19, vLR13]. out-degree [SRT^{+}16]. Out-Guesses [HZ82]. Outbreaks [FG99, Hol67]. Outcome [BC95b, BC95a, Fad83, LP17]. Outcomes [Cla99]. outdegrees [AH15]. Outflow [Ben70]. Outlier [KX11, Mat84]. Outlier-Resistance [Mat84]. Output [Aal98, BGT98, Boe69, Box75, Bré78, Cha94, Fad74, GR74, KP89, KG02, Leh73, Nat77, PL73, Zuc77, Zuc81]. Outputs [Col74, Wu68]. Overflow [Ano95e, CW96, DM03, Gui94, HR87, Ken78, Van89a, GHW22]. Overlap [Kel83, LP94b]. Overlapping [Ano91e, Can73, CP88, Joh06]. Overlay [DGM09]. Overload [Kan85]. Overloaded [Les06]. Overshoots [Ber99]. Overtaking [The69, ZPR08]. Owen [Fre07]. Own [Hen86c]. Owning [KA09, MT06].

P [Ano82o, AAO90, AHN90, LA98, Miy92, SS85a]. P. [Ano82o, Git73, Let78]. Paced [BN92, GJ78, HN00]. Packet [GMRY07, NP91, GM15]. Packetized [OS91]. Packing [BS70, BS82a, Coo89, CW86, Cow93, DS75, Hal89, Ito80, Kim94b, Rhe98, RB91, ST08, Sut89, TP79, Wei78]. Paid [Fro05]. Paintbox [Cra11]. Pair [AIGR06, Bor04, MC18, NA04, Pre06, KP23]. Paired [Lat81b, MI83, UB74]. Pairing [FE87]. Pairs [Dek96, JPR95, Lin14, Oak79, RR10]. Pairwise [CG00a]. Palásti [BS70, BS82a, Hor79, PT80, Tan79b]. Palm [Bré93, DM75, KS83a, Kon96, MO21, Sti94]. Palm-Type [KS83a]. Panel [Car99]. Paper [Bal78, Boy87, Dow67, Gre78, Gre86, IN86, Jof77, MU11, Mod67b, Sis70, Tan79b, vDB87, FHI17]. Parabolic [BJ07, Cab91, GBR75, ML98]. paraboloid [Let94]. paraboloïde [Let94]. Paradox [BKT97a, CSZ97, CK90, JL12, KLO99, ML85, Ye07, Ber23]. Paradoxes [SS99]. Parallel [ABGH00, Bar75c, BMPR07, Bla87, CCPW92, CR94a, ENPS86, Gla79a, HD71, HA96, HMM93, Käm87, KS99, KM01, KK00, KX09, KMM81, KST99, KW85, Leh88, LX00, MP70, MMD11, Nak79b, Nak79a, Nak84, Pin80, PC85, SZ06, Web78, Web82, WVW86, HMBL14, TK15]. Parallel-Processing [KS99]. Parallel-Series [ENPS86]. parallelism [JTW^{+}13]. Parallelotope [Rub79]. Parameter [BC01a, Bon89, Cab84, Cai03, Car02, FS07, Kab81, Kab83, KS01a, Kij92a, Kli14, KS93b, LM04a, Nan79, PH82, PP73a, PRCS09, Pru83, Ros86, Shi86, SH03, SZP07, VAK92, Xu09, CCGW23, HL13b, Möh21, Rug14, Tsu17, Tsu18, ZZ16, Zho15, Zho23, vD15]. Parameters [BZ09, BLS12, Che91, FSW12, Hei01, Hey94, Jia12, KPSS96, TH11, Tan79c, Tso71, Wil92, YC08]. Parametric [Pru85]. Parametrized [OK86]. Parasite [KN90, MD96]. Parent [Sha81]. Pareto [Ano93h, KS01a, KR08, Pil91, SW87, YAR88]. Parisian [BPPR16, CP11, DW11, DL18, DHJ15, LWW18, LY22, Sch03, YSL17]. Parities [Sau76b]. Parity [GS76]. Parking [AR82, Tam82, Tam88, YH04, Bel23, FHIM14]. Parrondo [EL19]. Part [CS80, Ros73, BP96a, Kar68b, KP96, Kei65, LOR19, VT76]. Partial [AG80, Ano75e, Ano77e, Ara94, BC73, BPT92, Cha97, EP95, Eme72, Epl83, Fak90, FPKMP98, Gha86, GS99, Gre76b, KR80, Maz86, NJS96a, Nel82, PS83, PP69, Pol75, Pol78, SD76, Tak77, Whi85c, Woo82, BY14, BNW13, EG15, EV16, Eri16, FX16, NNY20]. Partially [Cre74, DS06, Dav82a, DKS76, Kwe75, Led05, LT95, Mas93, Mon82, Tar15, Rah17]. Particle [Ano78g, CM97, CG90, Dar89, Dav76, Fen97, FF96, GM05, Gil67, GLS00, IMS98, JG85, Kar76a, Mac80, Mou93, Ori79, Pap05, SS83b, SK86, Zha05, BHTLV20, FL19, GPW22, Hös19, MZ20]. Particles [CG05, GR87, GR89, Gan88b, Har65, Kar76b, McD79, Møl88, KP23]. Particular [Ano81e, BD75, Che80b, Cov09, Neu94a, Rod95b, SS76, Tan82, Woo83a]. Partisan [AM18]. Partition [BSC02, Cra11, Hop86b, Joy98]. Partitioning [KKLO86, RS00]. Partitions [JT92, CM15c, HJL21]. Parts [Mei81, PW07b]. Passage [Ano87g, Ano88h, Ano91i, AS12, ASS85, BR00, Bel71, BWY11, BH12, Bre08, CR72, CW93, CK81, DIKM10, Dar66, DM67, DM68, Di 98, Dom95, Dom96, Don84b, Dur85, Dur88, DW92, FL98, Gal93, Gut75, GRRT97, HP98a, Ham05, Har90c, Her85, How00a, How01, HK09, IK11, JRR95, Jan81, JGR91, Kei71, Kij87, Kij88, KS93b, Lal86, Lef95, Lef06, LS97a, ML98, MS91, Mil67, NRS85a, NRS85b, NRS85c, Par88b, Pau87, Reh79, RS83, RSS84, RS88, Rog00, Sat77, Sch11, SS88b, SS76, Ste85, Ste95, Ste81, Sum84, Sum87, SR88, TR67, Tho75a, Tho11, TW84, Was69, Whi71a, Wic79, Wie80, Win05, Woo82, Yao85, Zip11, BS23, CdLH^{+}23, DS13, GA22, HJ14, Iva16, KK24, Lef24, LN19, SWZ18, SW20, Vee15, Vid19, WG15]. Passage-Time [DM67, DM68, Don84b, Har90c]. Passages [BD75, KS13]. Passing [Unk79]. Past [Ken88, PR08, DT15, DL02, Saw77]. Pasta [Sch93]. Pat [Ano89f, Han89]. Patch [Saw79]. Path [BM89a, DMT11, DT11, FPY04, Gra00, PV12, PS02b, Ros12, SS07, Ses96, SK86, Spo04, Sti94, Ten75, VSW09, VD17, WG93, Wol88, CCdOS15, FW21, FCvS15, FH17, Hoo13, KSS16]. Path-Dependent [FPY04, SS07, FW21]. path-length [FH17]. Pathological [Hun95]. Paths [Ano66d, Col69, Di 98, EE78, Hor65, Kin65, MN79, PS76a, AG16, BS17, CMW16, JX13]. Pathwise [Bor92, GJP^{+}21, JPS18]. Pattern [Ano91e, Bel66, CP88, Gol05, Kar83, Moo73, Nue08, Nue10, Ste03]. Patterned [BF78, GG74]. Patterns [Ant01, AM05, Blo84, BWZ85, BL03, Cow91, CH84a, FC02, GGS^{+}95, GKPS06, HH03, HB83a, IA07, KS71, Lot84, Peg75b, Rou90, GS14, RDM19]. Paul [Doo72]. Payment [EW88, GM15]. Payments [GCS07, RZ07]. Payoff [GHL03]. Payoffs [GHL05, OO07]. PDE [LMW00]. PDEs [GPW22, HS00]. Pdfs [VT11]. peak [Dal16a]. Peaks [WS96]. Peculiarities [Lit01]. Pedersen [Hör05]. Pedigrees [Kar68b, Ner84]. Penalties [VPT92]. Penalty [AHN90, AH00, Att87, BFP17]. Pensionable [Sah78]. percentages [SGS15]. Percolation [AW93, App94, Ari14, BS17, BS97, BS08, Che01, CS08, CK81, DSS11, Dek87, DFM^{+}06, GWS80, HP98a, Her85, How00a, How01, Jan81, JZG11, KZ84, Lam70, McD80, Mee95, Ngu90, OW79, Reh79, Rüs82, Sar97, Sch11, Sch80b, Shi02, Smy76, Tan93, VK85, VI14, Wie80, Wie84, Wu93, YCG11, Zha96, AG16, Ber13a, BCKK17, CdLH^{+}23, DS21, DP18, HH23, Hir16, HLF19, LN19, YW21]. Perfect [Bov99, DP12a, HS04, Mon82, Sch75b]. Perfection [Ken82a]. Performance [BK86b, BK87a, CG96, GW87, HZ87, KL00, LC83, PPF03, PPBT99, Rod95b, JTW^{+}13, Spi15]. Period [Ano74h, ALH01, BKNZ91, BC92, BPSZ10, Coh67a, Con74, CT76, Dal69b, DJ69, DT80, Dvu84, DKO84, Enn69, GT74, Had74, Hey74, Lan87, LO95, LS96, Lou78, LN94, Nat75, Neu68a, Pak73c, Pur75a, RS90, Ros75, Ser90, SS85a, Son96, Sta85, Wie89]. Periodic [Ano91g, AT87, BW89, Ber10, Ger91, HL77, KS78, Lem81, Lem89, Nak86a, OV03, OS91, Ped02, PL73, SS76, NPYY18]. Periodically [BH98, CX01, Iso82, KFM06, KKY02]. Periodicity [Rad74b, TV88]. Periodogram [Dav74, Ols67, Pag71, Roo76a]. Periodograms [Dah85]. Periods [Coh71, DS94a, DS98, DGM04, FH10, Gha86, Gha90, GW05, NSS04, PRX03, Sha71, Sta01, SZ93]. peripatric [LM15]. permanence [DNY16]. Permanent [Nád71]. Permutation [BPR08, Gla93, Hen83b]. Permutations [Die87, Don91, Gne98, Mah03, FGH^{+}21]. Perpetual [Eks03, Eks06, Lu10, Pau01, EM24]. Perpetuities [Hit10, AW21, BDIM18, DQL19, HL14]. Perpetuity [Pre97b]. Perron [Sag20]. Persistence [AM78, CDP19, GHW95, Kab12, Wan97, BC23]. Persistency [Elf67, IC82]. Person [CH06, CH10, LS92a, Sec97, ZC95]. Perturbation [FH95, Has65, Mit04b, Not00, Sch68, Sen89, Tit74, FHL13]. Perturbation-Type [Tit74]. Perturbations [Hin82a, Hin95, Las94c, Mer80, Twe75a, Via95, Gag17]. Perturbed [CL09, ELNT11, Els95, Hol78, KO12, LR95, RRS98, SV03a, Win79, ZW02, BIMR22, CTKR83, IRS23, PZ10]. Perturbées [CTKR83]. Pest [KA89]. Peter [Ano90m, BN17]. ’Petersburg [ML85, Gut10]. PF [ASS85]. PH [DO16, Neu81, RL88, VB02, DO16, KT15, BB12]. PH/M/c [KT15, BB12]. Ph/M/c-Like [BB12]. Ph/M/c/N [BB12]. Ph/M/c/N-Like [BB12]. Phage [GY65, Pur69, SR70b]. Phage-Reproduction [Pur69]. Phages [Bha68a, Sri67]. Phase [ANT17, Ary72, AL85, BCL09, Bre08, Bre10, BK73a, CL05, Chr12, CC96, CM02, Cox77, FHTZ10, Fen95, FH13, Gil68, HJ99, KS99, KS93a, LT00, MO92, MS79, PL81, RN80, Ros75, Ryd96, Shi02, Tan93, Ver98, AB19, Bre13, CH22, He21, PD20]. Phase-Age [BK73a]. phase-dependent [Bre13]. Phase-Type [AL85, BCL09, Bre10, CL05, Chr12, CC96, CM02, FHTZ10, KS99, KS93a, MO92, RN80, Ryd96, AB19, He21, PD20]. Phases [BMT10, Haj82]. Phatarfod [Bal78]. Phenomena [DS75, Ham82, Kin04, Pra88, FL19]. Phenomenon [Imh74, KB91, WDLC09]. Philology [NH82]. Photographic [And84a]. Phylogenetic [BS15c, BF09, LS09b]. Physical [Bar71, Bar72b, Bar74, BEN98c, CG87]. Pickands [DHM22]. Pickard [Wei79]. picked [KS18]. Piece [Hal89]. Pieces [Hol80b]. Piecewise [BL12, Cos90, Cos93, CD09, Kol95, Lu10, RDM19, Sch96, CdSJ20, Fuj13, Gén23]. Piecewise-Deterministic [Cos90, Fuj13]. pinned [Men23]. Piret [VE98]. Place [AR82, Hen81, Lef24]. Placed [Kel83]. Planar [Kel85, KM87, KO05, Kol07, KS83b, Ors86, RS89a, SK90, BM20, MO23]. Plane [App94, AFM95, BS70, DF88, EV82, Kel83, MC70, MC71, Mor66b, Neu82, Rog78, Tan83b, Tan83c, Wei78, Yad70, YZ85, AH15, vLR13, TP79]. Planes [SY72, Wu93]. Planetary [Gal71b]. Planning [Gue99, SV71, Woo83a]. Plant [Dig76, Rou90]. Plate [FS78]. Platoon [BR79]. Play [CSYZ05, CCHY08, Sch05, Yao07b, YC08]. Played [DN04]. Player [Bar64, CD85, SB06, Tay79]. Playing [Yao07a, Yao07b]. plus [Hoo13]. Point [AM07, Amb65, AK93, Ass83, AB86b, Bac91, Bas71, Bau84a, Bau84b, BV11, Bax98, Ber77c, BMZ82, Bor96, BM01b, Bri71, BN88, BHX98, Byt81, Cao95, CL85, CG00a, Çin68, CYF96, DK01, DM75, Dal82, DVJ04, DM01, DH85, Den85, Elf67, Fea81, Fei79a, FS80b, GK97, Gha90, GBR75, HD80, Hai96, Hen86c, How00a, HS10, Hüs86b, HS96b, IPR11, Imo01, Ish75, Ish77, Ish80, Ito77, Ito78, Kal75, Ken90, KK95, KS80a, KS80b, KS01c, KS91b, Lai78, Las78, Law70a, Law70b, LL77, Law79, LC16, Lie77, Lot84, MZ01, Mas95, Miy94, Mor75, Neu79, New70, OK86, PW07b, PL73, Pru83, Pur78, Pur82, Råd72b, Rao66, Rip76, Rud75, Rüs77, San80, Ser84, Sol04, SR70a, Tor02, Van96, VJ82, Vit74, Whi72c, Wil82b, Wil82c, Yah75, Yos84, dSS04]. point [ADJ13, BO18, BK21, BN17, BW19, CH13, DFV15, DV20, Gab16b, GP13, GOP15, Hös19, HZ22b, LS19, LGD20, LTZ18, MC18, MO21, Ros14b, Sie89]. point-catalytic [BW19]. Point-Process [OK86, VJ82]. Point-to-Point [How00a]. Points [Ala76, AFGP91, CH91, Dix93, Dwy88, EW00, Gri88, Ito89, IMT00, KD78, Les69, Man83, Ray70, Ste99, VW86, dHR73, BKP^{+}22]. Pointwise [CCFG96]. Poisson [BS16b, AH00, Acz72, AMR01, AT79, AAL77, Ano79f, Ano93e, BK21, BN17, BKL03, BB90, BE87, Bar88a, BM97, Bau81, Bau83, BM14, Bel71, BR02, Ber77b, BE83, BM82, BS14b, Bla92, Bon13, BP96b, Bor04, BM23, Bre75, Bro69a, BR69, BS79c, BN92, CG00a, CK00b, CV01, ÇJ73, Coh89a, CM15b, DK01, DJ18, DH85, DMZ15, DM18, Die99, Dix93, DG06b, Du95, DT11, Dur71a, Dwa67, Eag79, Eco04, EF90, Erh00, FP94, FI98, FCKS16, FMS99, Fol86, Gab16a, Gal66a, Gal93, GK18, GZ16, GP13, GGS^{+}95, GL67, GNRW94, Gra71, Gri88, GR74, GG86, Hal83b, Has64, Has65, Hay00, HKT24, Hey82b, HT78, HN00, HKQ88, Hol08a, Hop86b, Hsi97, Hua89, HLS93, HMY17, Jac75, KR92, KS01b, KBM06, Kel86, Kem79, Ken93, KP23, KPS81]. Poisson [KT83, KW03, KMS83, KS83b, KA03, KA09, Kyp71, Lan84, Law70a, Law71a, Law71b, LL75, LA98, Lem81, Lem89, LST19, Les69, Lew69, Lib85, Lie78, LS91, LN23, LX20, LMX04, MV16, Mar05, MM10, McC76, McC80, McN68, Meh77b, Mil76, Min78, Möh05, MS92, Nak08, Nea05, Neu66, NR78, Oak79, PS89, Paw89, PJ02, Pek06, Pen18, PS01a, PSZ04, PH87, Pha69, PT23a, Pic71, Pol67, Pós09, Pur75a, PG79, Pyc21, QW84, Råd72c, RW21, Rat92, RR77, Roo03, Roo76b, Ros69, Ros78, Rug14, Ruz04, RS10, Ryd96, Sam74, SF77, SH21, Sch19, Ser72a, Sha74a, Sha77, SB78, Ski12, SC82, Spo01, SS70, Sri71, SP99, Tan05, Tho85, TTI22, TSCD08, UW74, Van96, Vas82b, VX04, VC96, Vel04, WCW05, Web99, Wes99]. Poisson [Wes73, Wes77b, Wic79, Wit90, Won97, Xia97, Xu09, YZ85, YC01, Yao96, ZW02, ZY08]. Poisson-Binomial [Ski12]. Poisson-Driven [SC82]. Poisson-like [GK18]. Poisson-Paced [BN92]. Poisson-saddlepoint [BN17]. Poisson-Type [PSZ04]. Poisson-weighted [Gab16a]. Poissonian [FF96, Mat74]. Poissoniens [Mat74]. polar [Sei16]. Policies [AHN90, AH00, Ber76, BR87b, BK03, BLS90, BLS93, BS73, CG00b, EE76, Fro93, Gla93, JFS12, JCM98, Käm87, Lan88, Les86, Lew01, LAF02, Mi02b, Miy92, NO74, Nak86a, PV12, Pek03, Rie76, Rig96, Sav88a, SZ92, She92, STC93, VPT92, Wan76a, Zhe91, Zuc80b, Zuc81, CCdOS15, HR15]. Policy [Aba73, Aba79, ADY07, APZ11, BKL03, Bha69, CL11, CP01, Dsh98, Dsh01a, Emm72, Ger76, Las94a, LLPC94, LA98, LBL05, LT94, McN85, Mi02a, MN82, Nel82, OKM86, Qiu91, RS90, She96, Tan78, Wan77b, WW90, Yas88, Yeh95, ZC95, Zha94, Zuc89, AA17, BFP17, Jia15, SP17, YCW15]. Polish [ZPR08]. Pollaczek [Coh81a, JS96, Kel12]. Polling [CNS99, Dow98, Kro97, OV03, TTH91, Van99, AKLL18, KSBK23]. Pólya [AG07, BSC18, BK64, Bor21, CK13, FS14, GR13, Gou97, HMPS06, Hol08b, Jan20, MO93, Nak09, Nie06, RW79, SGS15, Ski12, Thö16, Tri92, Wal76]. Polygamy [Can74]. Polygons [Geo87, MM82, Nag93, Neu82, SY72, Tan83b, FV18]. Polyhedron [Mat90]. Polymer [MV86, SM72]. Polymers [JW11, LV72, Ols77]. Polymorphism [HH85b]. Polymorphisms [Hai90, Kin88]. Polynomial [BLBL10, Bra75, DM72, Ort08, Ros82, Sen01, CDP19, Kim17, VR22]. Polynomial-Rate [Ort08]. Polynomially [Eba84]. Polynomials [GP99, Wat88, Mar13, MA17]. Polytope [Dwy88, Van89b]. Polytopes [BM84, QZ19]. Pontriagin [Mor68]. pool [BvdHvL17]. Pooled [PL73]. Pooling [PPBT99]. Popov [Tso71]. Populate [BSB99]. Population [Aas76, Ano92g, BC95b, BG87a, BK64, BLT11, BM66, Bro70, CM82, Chi68, CS85b, CDD10, CH90, CS82, Don86, Eth79, Ewe82, Fel66, FW07, GT86, GST00, GT05, Hey82a, Hey86, Hin80b, Hin81, HB84, Hol67, Hol79a, Höp85a, Höp85b, Höp86a, Hou12, Jag92, JS00, JPR95, JP69, Jen74, JW85, KK03, Käm91, Kar68a, Kar68b, Kel76, Kim64, Kin69a, Kin77, Kla94, Kle83a, Kle84, Kle85, Kle89, Kle94, KS02b, LH81, Les10, Li06a, Lil68, Lip77, Man74, McC76, McC78, McC80, MJP87, Möh94, Möh97, Möh99, Möh10, MMR02, Mon67, Moy64, O’N96, PW67, Pol66, Pol68, Pol10, Pol67, Pra80, PW07a, RR75a, Ren73, Ren74, Ren77, Sam71, SG89, SF83, ST88, SB09, Sta99, Sud81, Sud83, Sud85, Tes84, Tra81, WF99, Wan99b, Wat72, Wat64]. Population [Wil88a, Woo83a, XW05, Yan75, ZTML21, ZJ04, BS19, BEER15, BBBN14, CFRS22, Cha13, MN16, Roy14]. Population-Size [Lip77]. Population-Size-Dependent [Ano92g, Höp85a, Höp85b, Höp86a, JS00, Kle83a, Kle84, Kle85, Kle89, Li06a, MMR02, WF99, Wan99b, XW05, ZJ04]. Populations [BOP07, BR85, Bec70, Bil76, BMS93, CF11b, CG12, CW69, CH84a, Cru70, Ell66, FF11, Fin04, Gal72, Gla77, Hin80a, Hin82b, HB83a, Hol81a, Jag69, JS04a, Kab12, KN90, Kin82, LM88, Len81, Ner84, Rad74a, Tai95]. Portfolio [Bro00a, CY08, CPT99, HBHH23, HØ07, RB05, ZLW06]. portfolios [BL13]. POS [CL98]. Position [Rad76a, CS20b]. Position-Dependent [Rad76a]. Positions [AIH87, AIH88, Eng64, Bel23]. Positive [AH00, AL00, AT70, BS04a, Bre75, Bre08, CMS06, Den10, DKS86, Ebr87, Gel91, GY04, Igl74, IM74, Kao78, Lee11a, Lee11b, LM08, McV08, Miz06, Ros80a, Rüs04, TO92, Twe75a, ZZ04, FPS18, Gra18, KMG19, Pro18, Tu19, Vid20]. Positively [SDH94]. Positiveness [Kal93]. Positivity [AP94, CS80, Day83, QS99]. Possible [ALS04, Gre76c, Mas86a]. Posteriors [Pia08]. Potential [BNL05, Iva14, KS06, Nat74, Van81, Whi90, KP23, Pit18]. Potentially [Had69]. potentials [Kub24]. Pour [CTKR83, Mat74, Mat76, Mat78, Bie98]. Power [BNGS04, Duf96, Dur71a, Gne00, JLN10, Kal91b, Ma03b, MW97, NW07, Wei95, BVW21, DS17, Gau22]. Power-Law [Duf96, Gne00, JLN10, Ma03b, MW97, BVW21, DS17]. powered [Kol15]. Powers [MSS92, BS17]. Precedence [BSC04, Gla80]. Precise [NTYY03, NTaYY04, WW07]. Precision [NP09]. Predation [Che76b]. Predator [Bil77b, CB91, Cof95, GK76, Hit86, RR88, DNY16, SB17]. Predator-Prey [CB91, Cof95, Hit86, RR88, DNY16, SB17]. Predecessors [HJ13a]. Predictable [Ebr93]. Predicting [All10, All12, Dub81]. Prediction [CFZ96, DM72, GN96, Kab83, LBP99, MM10, Rév04, Ris86, Sob67, SM94b, Tor10]. Predictions [Gue97, SLZ01, Cha13]. Predictors [TR89]. Predicts [DFSD01]. Preemptive [CN86, Fak86, Fak87, FR10, Sah71b, Sch72, Whi71b]. Preemptive-Resume [CN86, Fak86, Fak87, Whi71b]. Preface [Ano88n, Dal01b, GG94a, GH82, Gan86]. Preferential [CJV10, Dei10, Jor18, ARS21, BKK19, Bri20, CF22b, LM20, SRT^{+}16]. Preliminary [Ger77]. Premier [Lef95]. Premiers [Tor69]. Premium [DWW11, Jac12, Ven79]. Prescribed [Bro04b, Teh13]. Presence [AT06, AK97, CSZ04, CSYZ05, Cre74, DV87, DKS76, FGL00, GKL01, Hos86, Qui86, CY16, Roy24]. Present [Coh67a, Kei82, Ken88, RZ07, Sch88a, ZLW06]. Presentation [AL85]. Preservation [AL00, BS04a, BMPR07, BAAB00, FP94, LSW19, Sav03, Won97, GT14]. Preserve [DG99]. Preserving [SAG73, Cri18]. prespecified [AGP13]. Pressure [Tra81]. Prevalence [AN22]. Preventive [LS03, MR03, Nak86a, OA70]. Previous [Hör05]. Prey [Bil77b, CB91, Cof95, GK76, Hit86, RR88, DNY16, SB17]. Prey-Predator [GK76]. Price [JKM12, BPS15, JP18]. Prices [BMP87, CLS09, Eks06, Lu10, MKN11, OO07]. Pricing [BRY04, BJV12, BBMB08, BN02, Cai11, DJ05, DW77, DWW11, FPY04, Gay04, GER95, Ger11, GS01, Iva07, LR12, San06, Tan11, KL16]. Pries [BBvdMP23]. Prime [Tor69]. Primitive [CSYZ05]. Principe [Bie98]. Principle [All10, Bar72a, Bie98, Bré91, Cse79, DS08b, DNW07, Gue81, Hol66, MTV99, Mas95, Mor68, Pir80, Suz10, Vil07, Whi86, Zha05, Cin23, FX16, LSWX20, PSK17, PU16]. Principles [AS89, Ano99a, HIR03, Mac01, PR96, KLY16, Nyq17]. prion [OS14]. Prior [Kab81, NvdBOdI24]. Priorities [Sha89a, Wol70, Yad70]. Priority [Bau84a, DN80a, DT01, Har73c, Käm87, Sch72, Sch75a, Whi71b, XMKW90]. prizes [Ano22a]. Probabiliste [Jof82]. Probabilistic [AAKR05, BJ07, Ber10, BDO07, CGR08, Dav83, Di 99, DO97, EY11, GKRV94, HKO94, He01, Hey80, Hol13, Kal92, KKLO86, Lin99, LMW00, Luo00, Neu80, Xia99, CH22, KY17c, Psa22, Jof82]. Probabilité [PL03]. Probabilities [AF87, Ano78f, AN95, Asm09, BK87b, Bil73, Bof65, BJM93, BN05, Bro66, Bro00b, BL98, Bur76, BLL84, CG99, Cai02, CL93, Coo74, Dal69a, Dal79a, DS94a, Di 98, Dio74, DS93, Dur71a, Fel91, FW10, Gal66a, GW94b, HT12, HS69a, Hau12, Hen81, Hen83b, Hit86, HPSV04, Ios86, Kal93, Kar88, KM66, KNT10, Koo04, Kry72, Kry75, LL82, LR80, LV12b, LR07, MM71, Miy90, MT02, Møl95, Mor86, Nat74, NT04, Nis74b, NFK99, Nyr94, PS76a, PS80, PS83, PF06, Pha94, Pit80, Pol71, Raf73, Rob05, Rog00, Rog87, RR12, Rüs82, Smi72, SK86, Sti94, Sud05, Tor79, Twe73, Twe75a, Twe75b, Van81, Web99, Wei89, Wei76, Wil69, WW02, Yao93, Zei95, ACVZ20, BH20, CD13, FW16, GS14, GS19, GH14, HJ13b, KP22, LB23, MO23, PKO16]. probabilities [PT23b, Sez18, Zha22]. Probability [Abd87, Ail68, Ano66d, Ano78h, Ano95e, Ano22a, AC94b, Bal95, BS06, Bin94, BR96, BCS93, BS97, Bra75, Bro79, Cao89, CR72, CH06, Che11, CD05, Cra88, CP11, DN80b, DVJ04, DVJ08, Dav84, Dev79, DJ95, DGNR95, DR09, Dir75, DKK08, Fea76, Fea80, Fin88, FMS04, FC02, GT06, Gan82, Gan88a, Gan10, Gat84a, GER95, GS72, Gil68, GNR89, Gol75, Gol76, GWS80, GZ94, Gre01, Gui94, Han65, HH07, Hen86c, HS78, Hit88, Höp85a, Hor65, Hsi76, HS87b, Hul01, Hun76a, IK04, IK06, JRR95, JL12, Käm91, Kan85, Kap72, Kar77, Ken88, Kwe75, LHN86, Leh73, Len81, Les10, Lin73, MZ93, Mas95, MT00, Mi94, Mil92, Mil08, Mor66a, Mul68a, MZLSRS02, NS80, Neu88, New66, PVS88, Par88b, Peg75a, Pen96a, Pla85, Pon07, PW01, Qui76, RS83]. Probability [RSS84, RS87, Rog78, Ros08, SPS80, Sat77, SMS94, Sie78b, SS80, SO11, Sto04, Sub79, Tan05, WP97, Whi85b, Yos84, Zha96, ZY08, ZJ04, vDZ05, AA17, CDP19, CLCZ14, Che18, DHM20, DN13, Eri16, ES23, GHW19, LY22, Ros16, TTI22, Ano15, PL03, Ano14, Ano89e, Ano13]. Probes [Col81]. Problem [AHBT82, AR01a, AOR03, AW90, ARM98, Ano78h, Ano88f, Bal75, Bal76, BG02, BNL05, Bar64, BKNZ91, Bon02, Boo79, Bru87, BL98, Bru05, BS09, BGNR90, CR72, CW93, CRS97, CB83, Con84, CS97, CST08, Cow81, CCG90, DK01, Dek87, Del98, Dev79, Dia76, Dia73, DS86b, Dri75, DP10, DKO84, Elf67, EF90, Fan82, FDVF93, GWMO79, GT05, GC71a, GJ80, Gli00, Gne94, Gne96, GM07, GI11, Gre84, Guo01, Gut03, HTW73, Hal85, Has79, Hen86a, Hör05, Hsi76, Hüs88, Igl65, Iso82, Jag69, JW85, Kap81, Kar77, Kar09, Kea67, KS08, Ken82b, KSC12, Kuh85, KA03, KP07, Lal86, Lef97, LRS07, Loe09, MS02, MT06, Maz86, MC90, McN68, Men82, MM94, Mul68b, Nak79b, Nak79a, Nak91, Nak08, Nea08a, Oht91, Osb82, OS91, PVS88]. Problem [Pet83, Pha69, Pha91, Phe96a, Pit77, PL72, QL96, RC73, RR75a, RM76, RF88, Rév04, Rig90, Rog87, Rot95, Rub85a, Rub85b, Sap75, Shi65, Sis70, Smi75, SKK06, Ste81, Sto91, SB06, Swe88, Tam82, Tam86, Tam88, Tam01, TR67, Tuc76, VAW04, WR74, Wic79, XT11, YZ82, Yas83, Yas84, Yeo97, Yos84, Zäh90, ABS15, ABSGS16, BJ14, BKJ16, BS14a, BS17, BT23, BM15b, CLZ24, CF22a, EG15, EDSM15, GR24, GKP20, Gue13, KK24, Lef24, Lin16, NR18, RWZ13, Sau23, SH21, Yu17]. Problèmes [Lef95, PLC03]. Problems [Amb65, Ano73e, AT96, AK96, BS92a, BRY04, BSS72, Ber81a, BL10, BDO07, Bra82, Bre12a, Cam97, Che91, CL98, CY65, CP78, Coo74, DGM08, DNW07, Eng64, Eng65, Eva71, FW91, FT78, Fro93, Gav68, GL76, Gil87, Git72, Gla91, GG95, GS01, Guo02, HH03, HH85a, HK90, Hol81b, HKQ88, Huf86, IKP01, Irl06, Jan83, JR82b, KR86, Khm11, Kin04, KA97, Lef06, Leh73, Mil81, MT70, MW90b, MY91, Nak74, NRS85c, Now90, Ped00, Pet81, Pfe89, RO99, RS00, SC75, Sam70, Sch80a, ST12, Sie79, Ste00, Sun92, Tam85, Vil07, Whi05, YH04, Yao07b, YC08, Yas88, Zha92, Zho07, AS17, Buo15, EOV20, FVZ19, FHIM14, GR14, GA22, HZ20, Kli14, KS16, KSC13, LOR19, Lef95, LVZ19, Men76, MV21, Tam15, vdHKPS23, PLC03, Hog86]. Procedure [Cha00, Fad74, SKK06, Zuc77, HK94, HK94, RNR18]. Procedures [Ano82i, Cha01, FP66, Gla90, Her94, ML86, Sta87b, Twe71, Wil82a]. Procesess [QW84]. Process [AHP75, Aal98, AH84, AH00, AB00, ADLN82, Adk69, ARZ08, Ale01, AAO90, And03, Ano74f, Ano75g, Ano78f, Ano79e, Ano83a, Ano86g, Ano87i, Ano88h, Ano06f, AK76, Bac91, Bad11, Bag86b, Bal95, BD75, Bar75a, BS79a, Bar88a, BM97, Bau83, BV11, BGT98, Bee75, Bel71, Ber85, BJ79, BFM08, Ber11, BMZ82, Bha65b, BG87a, Bil77a, BLL79, BLL80a, BLL80b, Bin88, Bla92, BBP03, BN05, Bra70, Bra85, Bra02, BM66, Bro74, Bro86, BRS75, BH75b, Bru78, BDMZ12, Cab84, Cab91, CP04, Car02, CG97b, CG02, CF09, Cha75, Che94, CPZL04, CFS11, CY84, ÇJ73, CS72, Cli83, CB91, Cof95, CCGP91, CP78, CS80, CI78, CS12, Cra11, Cra73b, CD81, CM69, CG76, CS82, CP11, DIKM10, DK01, DN80b, Dal81b, Dan69, Dan77, Dan82a, Dar70]. Process [DJ05, Dav78, Dav64, DM01, De 77, DSR80, DW77, DM10b, DR09, Dio74, Dir75, Dix93, Dom95, Dom96, DM72, Don71, Don76a, Don91, DF81a, Du95, DM05a, DS08b, Dun82, Dur85, DW92, Dur71b, Dur74, Dur75, Dwa69, Eco04, Elf67, Est75, FS80a, FE87, Fal86, FY82, Fer83, FMS99, FS07, FSW12, FK89, Fos75, FG76, FNV11, Fra07, GR87, GR89, Gad89, GG91a, GW76, GW80, Gal66a, Gal93, GM05, GS76, GT86, GW94a, GK96, Ger76, Ger77, GL67, GNRS86, Gla81, Gly89, GDP04, GM96, GM97, GHW95, Gra72, GR94, GZ94, GS85, Gri72, Gri88, Gri75b, GG86, Had69, Hai78, Ham92, Ham95, Has82a, Has79, HC07, Haw70, HO74, HH07, HS94, Hey70, HS72, HT78, HNF85, HF71, Hoo87, HLZ05, Hua89, HLS93, HPR86, Hul01, Hun86, HK09]. Process [HCK02, Imh84, Imo01, Ish75, Ish77, IS83, Iva81, IS85, Jac75, Jag69, Jam66, JP09, JP93, JGR91, Jia12, JW90, JW82, KK03, Kao74, KSN75, Kap75a, Kap75b, Kap10, KP89, KS06, KR70, KR88, KBM06, Kel86, Kel77, Ken75a, KM66, Kha68, KT83, KS82, Kle83a, KLM04, KP93, Kuc82, KP83, KA09, KP07, Kyr99, Lag05, Lam68, Lan05, Lan11, Lan84, Lan86b, Law70a, Law70b, Law71a, Law71b, LL75, LL77, LY95, LA98, Lee11a, Lee11b, LM90, LP94a, LPS06, LV12a, Li06a, Lib85, Lie77, Lie78, LNTZ99, LS00, LR12, LV08, LT73, Lyn00, MZ01, MJ85, MS08, Man74, Mar73a, ML98, Mar86, MS91, MW96, McD78, ML67, McL76, McN85, Mer80, Mil92, Miu77, Miy94, MSZ04, MMR02, MS79, Mor73b, Mor74, MSZ94, MS10, Muc02]. Process [NW07, Nak98, Nan79, Nan82, Neu78, Neu79, NA04, NTYY03, Nis75, Not01, O’C93, OS01, Oak75, OK86, Ols67, OS91, PK74, Pak76, PH82, Pak87, PZ07, PP73a, PS80, PS83, PB83, Ped00, Pel94, PS01a, Pic71, Pic77, Pis06, Pli75, PT11, PZC07, PW07a, PL81, Pur75b, Pur87, Qui78, Qui70, Qui76, Råd72c, RT90, Rad72a, Rad73a, Rad76a, RC73, RCLW10, RCC12, Red03, Ren87, RR75b, RR94, RRS05, RS87, RS88, RB05, Ros69, Ros72, Roy72, RVJ68, Rus78, Ruz04, SA72, Sam71, Sam74, SG89, Sar73, Sat77, Sau76b, ST86a, Sch72, Sch92b, Sch98, Sch92c, Sen81, Sen67, Sen73b, ST83, Ser90, Sgi01, ST73, SS89, Sha79, SS76, She79, Sie81, SZP07, Smi73, Sol04, Spa91, SO11, Sri91, Sri71, Sud76, Sud81, Sud01, Sud04]. Process [Sud05, Sys94, Tak76, Tak92b, Tan94, Tav80b, Tav81, Tho75a, Tho11, TP74, TW84, VE98, VJ82, VJO84, VD69, WW06a, WCW05, Was69, WY93, Wau72, Wei70, Wei72, Wei74, Wei75a, Wei76, Wei84, Wei82, WM77, Wen90, Wes99, Wes77b, Whi85a, WF80, Whi68, Wil88a, Wil82b, Wil82c, Win89, WDLC09, XY09, Xia97, Yan73, YT83, YH87, Yeh05, Yos84, ZRU68, Zei98, ZW02, ZG88, Zuc80a, vDZ05, AMV21, AM14, AS16, BSC18, Ber13b, BS14b, BS16b, BSDG20, BCKS13, BIMR22, Bor21, BM15b, CS20b, CH13, CL23, CPN18, DM19, Dal16a, DFV15, DIMZ13, DMZ15, DM18, FS23, FLM15, Fox15, Fox16, FL20, Fox21, Fro15, FHIM14, GZ16, GP13, GG14, GH16, HL16a, HL13a, Hös19, IRS23, Jia15, KM15a, KOH23, KSN16, KS18, Le14, LC16, LST19, LR24, LM24]. process [Man13b, MK15, MC13, Mia13, MD18, Möh21, OS14, Per23, PT23a, Pim14, RW21, Ren14, Ros14b, Sag23, Spi22, VR22, Vee15, Vid23, Wat23, Win15, YW21, Zhu14, dRdC18, dL23, vD15, HS74b]. Processes [Aas77, Aas87, Ada75, Agr75, AR05a, AARJ10, Amb65, AL12, AT79, And79, And81, And70, AA93, ADW73, And84b, Ang11, AHL02, Ano64c, Ano65c, Ano65d, Ano66c, Ano67e, Ano70d, Ano70e, Ano73e, Ano74e, Ano80e, Ano86i, Ano91h, Ano92f, Ano92g, Ano93h, Ano94f, Ano02e, Ano06g, Ano07e, Ara04, ANT78b, AG81, AR89, AK93, AR01b, Ass83, AGSC00, AFG11, Ath71, APS74, AP86, Ath88, Ath12, Att89, BPS12, Bac91, BDM12, BK12a, Bag82, Bag86a, BdHK93, Bal99, BS01a, BM05, BR87a, Bar75a, BNY69, BNCPW09, Bar86, Bas71, Bau81, Bau84a, Bau84b, Bau09, Bax98, BSS72, BOR01, BR02, Ben68, Ben98b, BH98, Ber83, Ber70, Ber77b, Ber77c, BMZ82, BM82, BR87b, BD12, Bil81, Bin69, Bla94, BL11, Bli89, BM01a, BGH71, BWY11, BH12]. Processes [BPT92, Bon81, BP99b, Bor96, BV96, Bor99, BVJ00, Bor04, BL12, BG79b, Bra86, Bre75, BM01b, BNT02, Bre81, Bre08, Bre10, Bri71, Bro70, BK73a, Bro04a, BDY07, Bro69a, BH75a, BR77b, BN88, BHX98, BS92b, BS93, Bru78, Bru84, BT12b, BCMW83, Bui75, BGNR90, Byt81, CC06, CW81, Cam73, Cam97, Cao95, CG05, Cha74, CL85, CG00a, CG90, Che97, CK00b, CC01, Che02, CL03, CPLZ07, Che70, CT72, Chr12, Çin68, Çin80, Cip88, CP03, Cla12, CK07, CT81, Coh77b, Coh89b, CG69, Col76, CSV01, Coo98, Cor87, Cos90, Cos93, CD05, CD06, CD09, CT95, Cru70, CCFR00, Cse91, DS06, DS08a, DM75, Dal82, DHT86, DVJ04, Dal79b, Das96, Dav73, Dav86, Dav01, Dav65, DR94a, DM01, DZ03, DO12, DH85, Dek87, DJ00, DLU09, Den85]. Processes [Deo73, DGNR95, Di 98, DG93b, DY97, Don72, Don73, Don76b, Don84b, DF04, DKM93, Dos79, Dsh94, Dsh98, Dsh01b, DGM04, DP10, Dur71a, Dur08a, Dzi84, Eag75, Ebr87, ER88, ER89, Ebr93, Epl83, Est76, FG99, Fas90, FS78, FI87, Fea76, Fea80, Fea81, Fei79a, Fei93, Fel91, FI98, FW07, Fer95, FL91, FP05a, FH07, Fin07, Fin10, FS06, FH72, Fol86, FKK91, FKM10, FS80b, FR86, FW10, GR87, GR89, GG91a, GA93, GW82, GT06, GT85, GY95a, GD74, Gao04, Gap07, GSV11, GJ78, Gei99, GP04a, GS94, Gha90, GM84a, GNR89, GL03, Gir86, Gla92, Gla78, Gla79b, Gla82b, GBW92, Gol78b, GMD02, GMD03, GMD05, GMSB10, GBR75, Gor75, GK12, GL83, GPZ05, Gre77a, Gre81, Gre74, Gre76c, Gre77b, Gre78, Gri73, Gui02]. Processes [GM94, GP99, GZ02, GZ06, GRRT97, GG78, HD80, Haj82, Hal69, Ham88, HK77, HL81b, Har90c, Has82b, Hau12, Hay00, HS66, HW83, Hei09, HT89, Hen73b, Her03, Her85, Hey69, Hey71, Hey73, Hey81, Hey88, Hey94, HM99, Hol67, Hol08a, HM95, HZZG86, Hoo83, Höp82, Höp85a, Höp85b, Höp86a, Hop75, Hop76, HY94, HS77, Huf87, Hul82, Hul84, HS10, Hüs86b, HS96b, HPSV04, Hyn94, IPR11, IK70, Irl03, Ish80, Iso82, Ito77, Ito78, JR95, Jac75, Jag74a, Jag74b, JS00, JZ02, JK04, JKP11, JC06a, JK81, JR82a, JFHL09, Jaw99, JK71, JP99, JC06b, Jir76, JL06, Jon99b, Jon04, JMZ11, Kal95, Kal75, Kal92, Kal93, KS86a, Kan75, KS93a, Kap72, Kap74, KÖ89, KT82, Kar76c, KNS73, Kat77b, KR92, KL03a, KN75, Kei64a, Kei64b]. Processes [Kei65, Kei71, Kel92, KW92, KS94b, KS08, Kel09, KN90, Ken90, KS86c, KS91a, Kij92a, KN01, Kin69a, Kle84, Kle85, Kle89, Kle91, Kle94, KPSS96, KK06, Kni71, KK95, KHY03, KS80a, KS80b, KPW94, Kra64, KR97, KS83b, KS01c, Kul87, Kur87, Kur70, Kur71, KZ84, Kuz10, KS91b, KP08, KLP12, KO12, Lac05, LML06, LS08a, Lai78, Lam91, Lam92, Lam70, LJ78, LBC81, Lán79b, Lar98, Las78, Las94b, Las94a, LL78, LR93, Law79, LRS94, LK77, LW93, Lef97, Lef06, LS07, LPSV00, Les73, Lev77, Lew69, LM91, LM08, LS97a, Li00b, Li06b, LPLG08, Li09, LS91, Lin89, Lin72, Lin79, Lin82, Lin83, LR96, Lin04, Lip77, Lit92a, Lit92b, LH06, Liu11, Loc82, Loe09, Loi05, LS11, Lot84, Luk69, LMX04, MM09, MW10, Mac01, MS06]. Processes [Mal67, Man69, Man71, Mao02, MZ04, ML86, Mas86a, Mas77, Mas93, May05, MKR91, MP92b, MS93, McC97, MQ00, McC84, McK84, McK81, McK88, McV08, MPHL10, Mil76, MY84, MY85, MY01, Mit04a, Miy77, Miy79, MW90b, MJP87, Möh94, MMR07, Mon67, Moo73, Mor73c, Mor86, MH76b, Mor75, Mor71b, Mor71c, Mou99, Moy64, Moy65, MM03b, MS86, Mur87, Mur80, Mye81, Nak09, Nak06, NSS04, Nat77, New70, Nic84, Nie85, Nie06, Nis74a, Nis74b, NRS85a, NP00, O’B74, Oak79, Olo96, Olo98, Ols77, OC03, Pak71, Pak74, PC76, Ped02, Pel07, PSZ04, dNPG96, Pfe82b, PH87, Pfe89, Pia08, Pia09, Pic71, Pil91, Pin71, Pis06, PP73b, Pol69, Pol71, PS86a, Pol90, Pou92, PW07b, PL73, Pru83, Pru85, Pur68, Pur71, Pur78, Pur82, QD77]. Processes [QS99, Råd72b, RS69, Raf73, Rah98, RY99, Rao66, Ren73, RD97, Res73, RR94, Rey70, Rey72a, Rey72c, RH81b, RB98, RR88, Rip76, Rob91, RJ94, RJP97, Roe70c, Rog00, Roo76b, Ros80a, Ros82, Ros76a, Ros80b, RD80, Ros70a, RE01, RS89c, Rud75, Rüs77, Rüs08, Ryc87, Ryd96, Sah78, San71, Sap87, Sau75a, Sau76c, Sau76a, Sch86, SV97, Sch82, Sch80b, Sch70, SI83, Sel91, Sen71, Ser72a, Ser72b, Ser80, Ser84, Ser92, Sev69, SS88b, Sha80, ST12, Shi06, Sho75, Sim90, Sim93, Slu77, Smy76, Sob82, SB09, SM94b, Sol96, SK98, Spo01, Spo07, SR70a, Sri98, SZ04, SV71, Ste06, Ste78, SC73, Sud02, Sum84, Sum87, SR88, Sur08, SH70b, Szá74, Szc97, Tai95, TH11, THP12, Tan79c, TR65, Tes84, Teu68b, TW75]. Processes [Tho86, Tok08, Ton73, Ton74, Tor02, Tor10, Tsi07, Tuc76, TG99, Tur06, Twe73, Twe75b, Twe81, Tzi11, Van96, VSW09, Via94, Vit74, Wan76b, WF99, Wan99b, Wan12, Wau74a, Wau74b, Wei66b, Wei66a, Wei69a, Wei75b, Wei75c, Wei71, Wes73, Whi71a, Whi72b, Whi72c, Whi90, Wil69, Wil99, WZ07, XW05, XZW09, Yan72, Yan88b, Yan88a, Yao85, Yao96, YAR88, Zac04, Zaj95, ZG04, Zei91, Zei95, Zha01, ZS04, ZZ09, ZWV10, Zho04, Zho07, ZPR08, ZJ04, dSS04, ADPR18, AB15, AS17, AF14, APZ13, BO18, BS19, BK21, BDT22, BN17, BS15a, Bai20, BBN16, BC23, BW23, BMZ23, BS13b, BPPR16, BM14, BGM15, BS23, BPSS14, BPS15, BHTLV20, BY22, BP16, BMR16, BLM20, BM23, Bra22, BIM17, BL21, BN15, CS20a, Cha13, Cha17, CLCZ14, CNZ14, CCGW23, CM15a]. processes [CF13, Chu18, CdSJ20, CM15b, Cur22, DHJ15, DS13, DV20, DDS20, DHKP18, DI20, DN19, DLVZ20, FFK19, FW21, FCvS15, Fin13, FKZ13, FvLB13, Fuj13, Gaá20, GN17, GvdH18, GOP15, Gén23, GS19, Gol23, GMdP^{+}21, GKS14, GPZ17, GH21, HLMOL21, HJ14, HZ22a, Hon13, Hon16, HZ22b, HM20, HCY21, HMY17, IR20, Iva14, JP16, JF23, JL22, JP22, JY21, KB13, KM23, Ker20, KMG19, KW14, KR21, Kun22, KPV21, LLZ15, LLZ17, LS19, LGD20, LLT^{+}22, LR23, Lef89, Lef24, LP18, LP21, LW19, LYZ17, LVZ19, LP23, LB23, LS20, LN23, LR14b, MZ24, MV16, MW21, MSV14, Mar13, MK17, MC18, MO21, NR20, Oes18, OT15, Ott20, PLRB23, Pro18, QDZ21, Rah17, RDM19, RY20, SS17, SF17, Sag20, Sch19, SV19, Sim23, SW20, TTI22]. processes [Tu19, Tzi16, VL15, VD17, Vid19, Vid20, WZ18, WY22, WW20, WZ13, WG15, ZZ16, ZZ13, Zha14, ZTML21, ZW17, Zhu13, vdHKPS23]. Processing [Cha93, FL97, KS99, Leh88, Web82, WVW86]. Processor [Ana88, AF83, Coh84, FK89, GP04b, HRZ07, KMM81, Moy08, Ott84b, Ram84, RHK90, Sch88a, Sen89, WT05]. Processor-Sharing [AF83, FK89, GP04b, KMM81, Moy08, Sch88a, WT05]. Processors [Gla79a, Pra82, WP80, XMKW90]. Processus [Lef89]. Product [BKT97b, BCS93, BC01b, Bou95, CHPT97, DW77, Gel91, Kel00, Mae80, McK89, McK82, MBM21]. Product-Form [BKT97b, CHPT97, Gel91, McK89]. Product-Type [BC01b]. Production [De 85a, Dos81, GJQ06, KP89, TG87]. Production-Inventory [De 85a, Dos81]. Products [Nei01, NP09, Pak08, Res71a, Van86, YL16]. Professor [Reu82, Wei80b]. profile [BKM19, SBM22]. Profiles [Mar00, Sri91]. profit [CCFM17]. profits [FPP19]. Progeny [Don71, Dwa69, GR87, GR89, Ken75a, KP83, Wei74, Let19]. Program [Ano82f, Pro82]. Programming [Bel66, Eva71, Gre84, Har80, MT70, Mor68, Rie76, Whi79, Whi80c]. Programs [PZ90, Sha13]. Project [DC77, Zuc80a]. Projected [Ano78g, Dav76]. Projection [MJP87]. Projections [Pou92]. Projects [Aas85, KS71, O’F89]. Prolific [BFM08]. Prolonging [DO85]. Promotion [Kal88]. Prone [LM04a]. Proof [Ano06e, Ber76, Bré91, BN88, Die82, DG92, ELNT11, ETV88, FG12, GKL85, Har80, Hor79, Iks04, Ken92, Kim70b, KZ95, Kon96, LMW00, Sha74a, Teh10, Van89a, Wan97, Xia99, Zhe91]. Proofreading [CS85a, FH89, YWR82]. Proofs [Gei99, Hör05, HX02, LN94, Raj81, Hol13]. Propagation [Fen97, Ren77]. Proper [AL85, Whi01]. Properties [AH84, Ada75, Agu03, Ail68, Ano86g, AT98, AP86, Bad11, Bag82, Bag86b, BMY02, BM05, BKS02, Bau81, BMPR07, BE91, BI05, BLT11, Cha74, CD06, CI78, Cre77, CCFR00, Dav74, Ebr99, FY82, Fin07, GG91a, GW80, GW82, Gat87, GMSB10, Gue97, HF91, Hav91, KMW11, Kal91a, KÖ89, KMR75, KW96, KZB05, Khm11, KL77, KL78, KK95, KL06, KM92, LL75, Lew69, LZ02, Li04, LS91, Lin99, Loc96, LN94, LMX04, Mas82, Mat92b, MR02, NJS96b, Nay87, New70, Ols67, Pag71, PC76, RS92, RRS98, Roe85, Ros07, Sch92b, SS88a, SS87, SY91, STC93, SAS04, Tho85, Twe73, UB74, Wil02, Woo83a, XY09, Yan73, ZG04, ZT01, ZLW06, Zij83, AMV21, ASB21, BM15a, BSC18, BBH14, BHTLV20, CNZ14, CL23, Erh23, HM20, Hu16]. properties [LPN13, MW15, MWH22, TSN17, dCGG16]. Property [Abr97, Ano91f, Bas70, Bra86, Bré00, Bro69a, CK00a, Che06, CM98, CD92, CM02, DS02, DP12b, Els07, Emb83, EO84, Fea81, Fei79a, FS92, Gue91, GMRY07, Hai92, Hal79a, Hay00, HX02, Ish04, Ken90, Kle83b, LM12, Lin94, Mej65, Pur82, Saw83, Sch93, Str75a, Tam68, WH04, BS15a, HL13a]. Prophet [ASC00, KK97]. Proportion [Mas95, Sud85, Tan83c, HKT24]. Proportional [BCR11, ET83, FS71, SI83, AES17, BS22, LWM20, LY23]. Proportions [Gou97, Jag69]. Proposal [DJ09, Kam17]. Prospective [NTYY03]. Prospective-Loss [NTYY03]. protected [ABD17, MW15]. Protein [Lad05]. Protocol [BFS09, FI87]. Proved [Tan79b]. Pruned [GH13]. pruning [AD13]. Pseudo [BG87b, Cha83, Ewe64, DS23]. Pseudo-Conservation [BG87b]. Pseudo-Contagious [Cha83]. Pseudo-Transient [Ewe64]. Publications [Ano82o, Ano86p, Ano94m, Ano01k, Ano04m]. Pure [AH84, Ano86g, FY82, Kle81, Kur70, Li00a, Pfe82b, Ros71, SS88b, Van89a, dSS04, SF17, Sag23, Wan15]. Pursuit [MT12, Sch75b, Sch76]. Pursuit-Evasion [Sch75b, Sch76]. Put [Eks03, Lu10, Pau01, EM24, Kol15].

QBD [BS13b, DO12, dNPG96]. Quadratic [BNL05, Wen86, Wen90, CCGW23, Pit18]. Quadrilaterals [Gat82a, Tan83c]. Quality [MP04, Wan76a]. Quality-Control [MP04]. quantifying [MO21]. Quantile [Cai11, DE11, JF16, Par86]. Quantiles [BCY97, BT76, Boo79, Das96, DS86a, SM78, Yor95]. Quantitative [AFV15].

Quantities [KM87]. quantization [Che18]. Quantogram [Csö80]. Quantum [CH71, Fin88, Hud73, Par88b, Par01, MS23]. Quantum-Mechanical [CH71, Hud73]. Quarter [AFM95, vLR13]. Quasi [ALS04, Ano70e, Ano79f, Ano93f, BP10, BZ04, Bon81, Bui75, CCH06, Che70, CP03, DS65, DS67, DKK08, FDVF93, FH72, HN01, Hey94, Hun86, KH83, KS91a, Kij92b, Kij93, Kyp71, Kyp72a, Kyp72b, LR93, LRS94, LSC04, LS00, MS94, MV95, Pia00b, PS86a, PG79, SVJ66, ST85, TW83, Twe74, VSW09, Zei89, Zie87, Cur22, HZ22a, HL14, JF23, Kim17, Yam22, Pia00a]. Quasi-Bayesian [CCH06]. quasi-Bernoulli [HL14]. Quasi-Birth-and-Death [HN01, VSW09]. Quasi-Birth-Death [FDVF93, LR93, JF23]. Quasi-Ergodicity [Zei89]. Quasi-Infinite [Ano79f, PG79]. Quasi-Input [Hun86]. Quasi-Likelihood [BZ04, Hey94]. Quasi-Likelihood-Based [ALS04]. Quasi-Limiting [Kij93, Cur22]. Quasi-Markov [Bon81]. quasi-regression [Kim17]. Quasi-Renewal [Pia00b, Pia00a]. Quasi-Score [LSC04]. Quasi-Stationary [Ano70e, Ano93f, BP10, Bui75, Che70, CP03, DS65, DS67, FH72, KH83, KS91a, Kij92b, Kyp71, Kyp72a, Kyp72b, LRS94, LS00, MS94, MV95, PS86a, SVJ66, ST85, TW83, Twe74, Zie87, HZ22a, Yam22]. Quasi-Toeplitz [DKK08]. quasispecies [Dal16a]. Quasistationarity [DP00]. Quasistationary [AFG11, C.12, Cla12, Ova01, PT11, Pol99, MSV14, ZZ13]. quenched [Bow18, DS21, Lam14]. Query [VOZ95]. Question [Bra80, BSB99, LS17]. Questions [JFS12]. Queue [Ano72g, Ano74h, Ano84e, Ano87h, AGC97, AGC99, ALH01, ALRW99, Bac86, BM89a, Bal72, BS02, BM97, BGT98, BKNZ91, Bhu06, BJM93, BBP03, BC92, Box75, BY97, BPSZ10, BS93, CP01, Cha94, CG96, CY90, CR97, CP92, Coh68, Coh71, Coh76, Con94, CH69, Con74, CT76, Cra69, DN80a, Dal65, Dal69b, DJ69, DS92, DS94a, De 85c, Dos85, Du95, DYY01, Dvu84, Emm72, Enn69, ESM08, Fak86, Fak87, Fal86, FF91, Fal94, FT80, FG84, FY83, Gil76, GW94b, GM11, GU92, Gra00, Gra83, GW11, HN71, Hal85, Han87, HL76, Has64, Has65, Hea65, Hen93, Hey64, Hey74, Hil67, HT78, Hok75, Hok77, Hoo83, Hoo70, HPR86, Huf87, HB02, Hun86, HCK02, Igl65, Ish04, Jen66, Jen72, Jin88, JW90, Kap71, Ken72a, Ken72b, Ken74, KKK07, KM93]. Queue [KS85, KST99, KNR05, Kul86, KG02, Kyp72a, Lan86a, Lan87, Lat81a, LLPC94, Lee97, Lem74, LZ94, LR04, LH02, Lou78, LN94, Mar05, MW71, MS99, McK89, McN74, Min78, Min79, Miy86, Miy92, MT02, Mor80, Moy08, Moy10, Nai71, NG95, Nat74, Neu66, Neu68a, NC72, Neu81, New68a, OP86, O’C98, Ott79, Ott84a, Ott84b, Pak73c, PS89, PJ02, Pes68, Pra70, Pur75a, PL81, RS04, Ram84, RL88, RHK90, RS90, Ros87a, Ros75, Ros76b, RS99, Rou96, Rub85b, Sah71b, Sch72, Sch75a, Sch88a, SW05, Sel97, STC93, SP99, Tak74, TTH93, Tam68, Tam71, Tho85, Tom72, Tur00, Van81, VB02, VD69, Wak91, Whi72a, Whi72b, Whi80a, Wil88c, Wol02, YC01, ZAFP06, vHS86, AKLL18, BNW13, DO16, FS17, IJB18, KSBK23, KT15, PW16]. queue [SGC16, ZL15]. Queue-Length [KS85, Miy86]. Queue-Size [Mor80]. Queueing [AS89, Ano90g, Ano05e, Ary72, APS04, BT73, Bam92, Bar82a, BKT97b, Bha65b, BN94, BK03, Bla84, Bré78, Bro64, CSZ97, Cha97, CC79, Coh67a, Coh67b, Coh73, CHPT97, Con84, DS02, DS08a, DT85, Del78, DT01, DK90, Dsh89, Dsh98, Dsh01a, DN99, DS04, Eco00, Eco03, Enn69, FT80, FG87, Fin69, Gel91, GXS12, GNN85, Gra00, Gre89, GMDB03, Had69, Har81, Hav91, HP98b, He03, HF70, HT91, Hen93, HS93, Jac78, JL98, KR86, Kan85, KK96a, KS80a, KS80b, KM88, KB91, KTW69, Kul83, LC84, Lan86b, LR80, LC83, Lel05, Les06, LP00, Lew01, LAF02, LSC97, Lin04, LT71, LT72, LT90, LT04b, MF01, McK89, Mej68, MO71, Miy79, Nak86c, Nat74, Nat75, Nat77, Naw83, New86, OM72, Pak70, Pea65, PS99, PS86b]. Queueing [PT94, Pur73, Rao66, RSY90, RS92, Ros80a, Sch86, SY86, SY89, Sha75, Sha89b, Sig89, Sim00, Sob90, SG93, SSV72, Sta85, Szc97, Tan94, Tet98, Tin85, TT79, VS90, Van90, Van91b, Van81, Vis86, Wal83b, Wal11, WB03, Wie89, Wil96, Wil99, Xu98, Yad70, Ye07, ZY70, BvdHvL17, GPSY16, HR15, JTW^{+}13, PW16, Ros14a, SGC16]. queueing-time [SGC16]. Queues [Abr01b, AS04, Ali70, Ano95e, AT87, AT96, AIH87, Aye07, BF95, BW89, Bäu97, BR98, BBK08, Bha68b, BE91, Bla87, BC72, BM94, BB12, Bri79, BR69, BS92b, Bru71, BG00, Cao84, CPS96, CZ98, CK00c, Çin67a, Çin67b, Con75, CL93, Cra73a, Cra74a, Cra74b, DMS08, DR84, DS94a, DS98, DESM10, DM03, DMO05, DGM04, Duf96, DKO84, FW21, FG98, FH10, Fol86, FK89, FKK91, FR10, FL97, GBG02, GGS91, GT74, Gha86, Gha89, Gha90, GW05, GG91b, Gre69, GW87, GR74, Gui94, GMS96, Had74, HD71, HGR90, HZ87, Har90a, Har73a, Har73b, Har73c, HL77, HL81a, HP93, HW84a, Hey82b, HW84b, HKN89, HL04, HAPV12, JS96, KM01, KZB05, Kel75, KM90, KAR11, Kin66, Köl74, Köl79, KMS83, KS85, KST99, KNR05, Kyp71, Kyp72b]. Queues [Lat81b, LT04a, Leh72, Leh86, Lem81, Lem87, Lem89, LO95, LX00, Lou73, Mad75a, MH76a, MT00, MNS01, MO71, Miy77, Miy85, MY88, Miy90, MY91, Miy92, Miy01, Moh72, NSS04, Neu68b, New68c, New68a, New68b, Niu80, Niu81, NR78, O’B75, OM72, Pea67, PR08, Pha93, Rab06, RN80, RHK90, Rey68, Rig99b, Ros70b, Ros74, Ros78, Sen89, ST08, Sha66, Sha80, SS85a, SS87, Sha73, Sha71, Sig88, SW11, SZ06, Sta87b, SP92, SZ93, SD12, SM97, Sze86, SDH94, Tak71a, Tak89, Tan79a, Tso92, Wal82, Wan99a, Web79, Whi71b, Wol77a, Wol87, Wol84, WPN80, WT05, YSS91, YT83, Yu74, ZZ04, CHR15, DSZ13, GS15, Hon17, Ino19, KB20, KSBM18, Lóp16, PSZ13, Sez18, Van22, Yao13, ZX17]. Queuing [Boe69, Bur79, CK90, Hem90, Råd72c, SD87, Wal83a]. Queuing-Theoretic [Bur79]. Quick [RB98]. Quickest [Bal75, Bal76]. Quiescence [MZ01]. Quintuple [Bre10]. Quite [Kap78].

r [Bhu06, DC77, KS71]. R&D [Aas85, PZ90, Zuc80a]. Races [Hen81]. Radial [QW84]. Radially [Gor75]. Radiative [Bro70]. radii [DS17]. Radioactive [Lit69]. Radon [BD75]. Rahman [Hua74b]. Rainfall [ST01]. Random [Adl76, Adl78, Aff89, Agr75, ALM06, Agu09, Ail68, Ala76, Ala77, AFGP91, Ale96, Ali99, All04a, All04b, AM08, All10, ALS04, AG80, Ano66d, Ano66e, Ano75e, Ano78e, Ano85f, Ano86a, Ano86j, Ano98e, Ant78a, AM78, ACSS11, AFM95, Ave85, Bac91, Bac01, Bad11, BO99, Bal72, Bal01, BMZ07, BE87, BG06, BN06, BCL98, Bar75b, Bar75c, BP04, Bau84a, Bau84b, BKMS04, BF91, BL00, BTH72, Bèr04, BM89b, Ber81a, Ber83, BM90, BZ08, BZ09, Ber99, Big77a, Big77b, BG79a, BS04b, BS70, BS82a, BL10, BT82, Blo82, Blo84, BGJ12, BWY11, BC73, BP96a, Böh00, BPT92, BSC02, BF09, BFS09, Bor94, Box75, BS79b, Bra80, BDLL08, BN88, BHX98, Bro04b, BPR08, Bru78, BO90, BL03, BDMZ12, BM84, BGP80, BR95, CK00a, CG97a, CMW95, CH91]. Random [CFRS22, CM07, CS86, CHL98, CCC03, CNO^{+}84, CR92, CR94b, CCN10, Chi68, Cho94, CZ99, CV09, CP90, CS75, CT81, Coh84, Coh88, Coh94, CS95, CP78, Coh81b, Col69, Col81, Col74, CJ98, CS08, CJ08b, Coo89, Coo98, Cor87, Cow93, Cox77, Cre74, CL92, CX01, Dal79a, DG93a, DS94a, DD04, DN04, Deh85, Dei10, Dek87, DZ07, DP12b, Det96, DM10b, DS94b, DY97, DS75, Don77, Don80a, Don80b, Don83, DGK98, Dor82, DG99, DF88, Dow66, DGM09, DKS76, DS93, Dwa69, Dwy88, Dzi86, EE81, Ell66, Eme72, Eng65, Eng93, EE78, EE93, EW88, FN70, FP97, Fal96, Fan12, FG84, Fel78, Fel91, FM10, FH11, FS14, FD84, FL98, Gal73, Gal66b, Gan82, Gan88b, Gan04a, GPSY16, Gat82b, Gat84a, Gat84b, Gav76, GJ78, Geo87, GK97, Ger77, GC71b, Gir82]. Random [Gne99, GI11, Gol75, GMD02, GMD03, GMD05, Got87, GGZ19, GS99, Gra85, Gre76a, GZ94, GR74, Grü85, GH09, GS06, GP05, GG10, Gut03, Haa92, HJ99, Hai81, Hai90, Hal89, Hal81, Hal83a, Ham92, Ham95, HJ02, HTW09, HS69a, Has76, Hel70, HN76, HW84a, HRF83, HRF84, Hen83b, Hen83c, Hen86b, Hen86c, HIR03, Hey67, Hey69, HG81, HH96, HN00, Hog86, HW88, Hol80b, Hol80a, HH84, Hor04, Hor65, Hor98, HS96a, HS14, HS87b, Huf88b, Hüs82, Hüs86a, HL71, IFN19, Igl74, IM10, Ios86, Ito79, Ito80, IS86, Ito89, IMT00, IKS70, Jac12, JZ02, JK04, Jan78, Jan88, JLN10, Jaw84, Jaw99, JL98, JP69, JR82b, Jon99a, JT92, KS72, KS99, Kao78, Kap72, Kap78, Kap82, Kar68a, Kar68b, Kar76a, KP92, KP96, KN75, Kei82, KS86b]. Random [KM01, Kel85, Ken94, KK97, KS10, KL03b, KN05, Kim94b, Kim70b, Kin69b, KW03, KS82, KKRT12, KM97, KT08, Kol98, KO05, Kol06, Kol07, KK01, KKP10, KP93, Kul06, Kup90, Kuu82, Kyp98, LB89, Lal86, LJ78, Law83, Law92, LMM10, LL93, LM90, LU96, LPS06, LMS76, LT80, Li12, LV12b, LH07, Lin73, LR96, Lit01, Lit74, Llo89, LR12, LP94b, Lun97, MM09, MP00, Mae80, MP92a, MO01, Man83, MM03a, Mar73b, Mas82, Mas95, Mat84, McC74, McL76, MS04a, Mei10, MSS09, Mil94, Moo73, Moo74, Mor73e, Mor75, MW91, Mor66b, MP93, MS05b, Mul68b, MR11, Mun12, Mut87, NB03, NS80, NH82, Nak79b, Nak79a, NG95, Nan79, Nea03, Nei01, NW70, Neu94b, NT04, NTaYY04, Ngu90, Nic86, Nis74a, Nis74b, Nis75, Nor75]. Random [Nue10, OP86, Ors86, Ors87a, Ors89, Oza80, Oza81, Oza86, Pak73a, Pak78b, PH82, Pak08, Pap10, Par88b, Peg73, Pel07, PP71, PF06, PT97, Pet83, PSW71, Pin11, Pit90, PP73b, Pod72, Pol75, Pol68, Pol10, PS76b, PZ90, Pra80, PL72, Pru85, QR90, Qui91, QL99, Råd76c, Raf73, RR10, RH81a, Rév04, RT91, Rhe98, RB98, Rit80, Rob92, RD99, RB91, Roe85, RP03, Ros86, Ros87b, Ros93, Ros81c, RSZ05, Ros12, Rot95, RE01, Rou90, Roy72, Rub77, Rub79, RMV97, Rus83, Ruz10, RS10, Sah83, SY72, SKF79, Sch85, Sch82, Sel91, Sgi99, SW97a, Sha87, She96, Sho75, Sie78a, Sie78b, SH82, Sie79, Sim00, Sim21, Ska71, SD76, SS86, SM94b, Spe73a, Spe86, Spo04, SZ04, Sta77, Sta81, Sta04, Sto08, Sud04, Swi76]. Random [SS03, Tak77, Tak88, Tak89, Tak92a, TW01, Tan83b, Tan83c, Tay99, TW99, TW75, Tol88, Tor79, TP79, Tro83, TK84, Tsi07, Tso71, Van86, VT11, VW86, VC96, VU09, VT75, VT76, Ver69, Ver98, Via95, VP10, Voi96, Voo87, Voo89, Wal76, WW06a, WF99, Wan99b, WW06b, WW07, WS96, Wat80b, Web80, Wei00, Wei78, Wei75c, Wei71, WG93, Wil69, Wit96, Wol77b, WK94, WW00, WS04, WT05, WDLC09, XY92, XB11, XH11, Yad70, YZ82, Yan74, YY11, Yao93, YA88, Yas84, Yos84, ZZ99, Zha92, ZJ04, vLR13, AH15, ADJ13, ANT17, ABI17, ARS21, ASZB18, AH13, BKP^{+}22, BM15a, BS15a, BKJ16, BM17, BS13a, BS15b, BLSGS13, BCK16, BvdHK17, BS16a, BMZ23, BSX19, BMR19, BJMK18, BT18, BBvdMP23, BCPR16, BY22, BCN24, BR18, BDL19, BRSC22]. random [BIMR22, BCKK17, Bow18, BH20, CP14, CDP19, CM16a, CL14, Cin23, CdLH^{+}23, CMW16, Con13, CHM^{+}22, De 22, DHSK15, DS13, DPW21, DIMZ13, DFP13, DL17, DI20, DLMR17, DLVZ20, Erh23, FZ17, FCP18, FGH^{+}21, FPY20, FV18, FH17, Gaá20, Gab16a, GZ16, GA22, GH13, GH14, GS21, HJ13a, HF19, HM23, HLF19, HK94, IK16, IK18, IRS23, Ino23, JP22, JX13, KS13, KY17b, KP23, KK18, KL22, KLP22, Kub24, KW13, KS21, Kur22, Lam14, Leh17, Leh21, Lei23, LRM23, LN19, LP14, LX18, LZ14, LSWX20, Lop14, LTZ18, MW15, MWW19, McK14, Men23, MS20, NV20, NR22, ÓT21, OC22, OT15, PZ10, PL17, Pen16, Pen18, PS16, Pit18, Pri19, Ray70, RS22, Rob22, RNJ17, Roy24, RU18, SL21, Sal13, SX18, SGS17, Sha23]. random [She13, SK23, Sim14, SUY21, SWZ18, Ste19, Tan17, VKR14, VL15, VD17, Wan14, Win15, Wir18, Zha22]. Random-Allocation [Gan04a]. Random-Coefficient [LPS06]. random-connection [Pri19]. Random-Service [WT05]. Randomisation [MPS15]. Randomised [HZ20, HJL19, JS20]. Randomization [LP05, Mah03]. Randomized [Fel91, HKY96, JKL^{+}20, LW93, Swe88, LT23, VM21]. Randomly [AGP13, BCPR11, Bri86, Col76, CD85, Els95, Hel71, Hin82b, Kel83, Lee80, LU95, Lit69, Lit79, McD79, Miy85, RR98, Sam67, Str75b, Sut89, Web86b, BS16a, BCPR16, CP14, Cri16, KS18, SWW21]. Randomly-Scanned [Lee80]. Randomness [EE81, Hic82, Hic83, Hic84, Ruk00]. Range [ALS04, BS06, BC73, BS08, BRPS82, CD07, Dai04, De 74, DR94b, DSS11, Gao04, Gne00, GAS03, HY97, Hey02, HW88, Hoo87, KM87, KN00, KS01c, Kün86, Lam70, LS07, LNTZ99, Ma02, Ma03b, MW97, NMG00, OA12, Pen96b, Ski67, Tro83, BS14b, BS16b, BCKS13, DV20, Gaá20, KP23, KS21, MV16, WW20]. Range-Dependent [KN00]. ranges [Pap17]. Rank [Ano10e, BF93, BL98, BS09, Gne07, Mar07, Osb82, RW79, Rog87, Yan74]. Rank-Dependent [Gne07]. Rankable [KA09]. Ranked [Pre97b]. Ranking [BCS08]. Rao [Sha74a, Sha77, Tal80]. Rapidly [BMZ07, BS97]. Rare [ARW94, FTV90, Heu87, Hon17, Hüs93, HS96b, LR09, Seg06, Spo01, Tso92, BRRN16, GH14, Owa23]. rare-event [BRRN16, GH14]. Rarefactions [Ser84]. Rarity [GR05]. Rate [Aal98, AK96, Att87, Bad11, BD90, BM01a, BLS03, BENP94, BNT02, Bro72, BPS80, CG90, Che96, CW69, Cla99, De 85b, DS86a, DM05a, DM05b, ES11, FT80, FL91, GPW22, GE82, Gla77, GBW92, GM07, GL84a, Hal79b, Hey70, IM10, Jan20, Jin88, KR70, KS01b, Kij98, Kle83a, Kle84, Kle94, KS80a, KR97, Law79, LT76, Li05, Lun97, Mac81, MS02, MKR91, MR03, Mit04a, Miy85, Miy94, NS06, Neu80, NP22, Ort08, Pru85, RS04, RT90, Rob99, Ros76b, RSZ05, Ruz04, Sim97, Sud01, Tan97, Tom72, Tor02, VT89, VT75, VT76, Voi00, Win79, Xia97, Yeh97, Yeo75, Zei91, ZG88, ZNF10, BHZ22, BMR19, BLS15, Bro15, Che18, Dal16b, GT14, GS19, GK22, HK19, HJ13b, KK18, LWM20, RY20, Tav20]. Rates [AT96, Bar75c, BL01, BSW03, BLSW08, BLSW10, BN05, BG79b, BS79c, Cai02, CG12, CK00c, Dav86, Esh73, FG76, FS98, FRT99, FH95, Git81, GMM06, Gre76c, GZ02, GHL03, Had74, HS72, HS74b, HS78, Hey82a, HW84b, HKQ88, HLZ05, Jac12, LS97b, Mac01, Mar69, Mat90, Muc02, NE81, NE82, Nei01, New68c, New68a, New68b, PS01b, Rob92, RT00, Sau75a, Sau76c, Sch98, SS88a, SY86, SS96, STC93, Ven79, Wau74b, AB18, AFV15, AW15, BKJ16, Bre13, CS23, DHKP18, Gag17, GH21, JOC13, PSZ13]. Ratio [ASC00, BHX98, DFM^{+}06, Heu87, Peg73, Tak64b, Whi82a, Yao93, MA17, VI14, Yam17]. Rational [BT12b, CCH06, CCHY08, LM08, KMG19, ZW17]. Rational-Valued [CCHY08]. Rationales [YYZ09]. Ratios [ARW94, Joy94, Mac80, MK73, Rob92, SW97b, BCdlP17]. Raw [Miu77]. Ray [BPS12]. Rayleigh [GNRS86]. Reach [BO18]. Reaches [KP92]. Reaching [HP17, Tsi07, vLR13]. Reactant [Ols78]. Reaction [Bal88, DF89, DR81, ACK20, AN22]. Reactions [HG80, MV86, Mul72, Mul73, Orr69, TL69]. Real [Bre75, KS01c, Tak77, KS16, Yor92]. Realization [Bau83, Cao89, YY11]. Really [DJ09]. Rearrangement [Ton85]. Recall [Pet81, Swe94]. Received [GL05]. Reciprocal [Erh23, Fei94]. Reciprocity [And84a]. Recognition [Bel66, Gat82a]. Recognizing [Tam79, Yan74]. Recombination [Gri81, PS09, Man13b]. Reconciliation [Dur88]. Reconfigurability [BN94]. Reconsideration [Hul01]. Reconstruction [De 77, Moo73, Win90, BR23]. Record [Ano72e, Ano82g, Ano85g, BN06, Ber92, Dal81a, Deh84, Gav76, Gup84, HN00, HS69b, Nag82a, NB03, Neu67, Pfe82a, Res73, Sho72a, Sho72b, Sho73, SH70a, Wes77a, Wit93, Yan75, dHR73]. Records [Ano06e, BR85, Bor99, BGSS03, BN92, CCC03, FCP20, GJ78, GS99, GR01, HTW73, Hol09, Iks04, Wei95, FCP18]. Recouvrement [Hua95]. recover [MZ16]. Recovering [HWX18, Lu10]. Recovery [Bil75, Hai70, Hua95, Lit69]. Recruitment [Dav75, McC76, McC80, Pol67]. Rectangles [CHL98]. Rectangulaire [Tor69]. Rectangular [Cha75, Lew67, SKF79, Tor69]. Rectifiable [Man77]. Recurrence [ADLN82, Ano82f, AFM95, Bor92, Bre76, Dwa67, FHT78, GP91, HRF83, Hun76b, Ker86, Kot94, Lam91, Les69, Les10, Let74, McC84, MSCZ78, Mil94, Peg75b, Pin71, Pro82, Rad74a, Reh77, Ros80a, Rud75, TW75, Twe75a, War99, Win89, XY92, ZG04, CTKR83, ABI17, AK23, CTKR83, MST13]. Recurrency [FT78]. Recurrent [Amb65, Bha68b, Bil74, BB12, BNS90, Dar66, Gra99, Kal90, Les67, Pea65, Ros76b, SR70a, TR65, Zaj95, IFN19, dL23]. Recurring [Koç95]. Recurring-with-Carry [Koç95]. Recursion [Sev69, VP10, MPS15, Vee15]. Recursions [KP87]. Recursive [Ban96, BD74, DL18, Dob96, Ell93, EW00, FM10, FH11, JS08, Led05, MR11, Mun12, Nak98, Ork72, Rüs06, Tet02, Via95, AH15, CMW16, IK18, MW15, SBM22]. Recycling [DJ09]. Red [Che76a, CH06, CCHY08, CH10, Pon07, Rut99, Sec97, Ren14]. Red-and-Black [Che76a, CH06, CCHY08, CH10, Pon07, Sec97]. Reduced [LS08a, JX13, MO21]. Reduces [Son96]. Reducible [ADY07, Pit77, RM76]. Reducing [GP04a]. Reduction [Ano83f, LR93, Nat82, Nat90, Tru00, MC13]. Reductions [KM01]. Redundancies [LD10, MDG09]. Redundancy [Mi99, MMD11, SM94a, KA22, LL24]. Redundant [KH83, MA69, OA70]. Reed [Bal83b, LP05, ST86a]. Reed-Frost [ST86a]. Réel [Yor92]. Reeling [Miu77]. Reestimation [BBMB08]. Reference [Sap87, VG92]. Refinement [CS00]. Reflected [BWY11, BGT04, Bre12a, Kel92, KS94b, KBM06, Lee11a, Lee11b, PS01a, Wil92, XZW09, Zho07, BM18, BMR16, BLM20, FvLB13, HJ14, HL13b, KM23, ZZ16]. Reflecting [BH12, Cre74, DKS76, HS69a, Iva10, Kas66, KS04, Kin65, RS87, Sch92c, Iva14, PG13, Pih15, SW20]. Reflection [CL03, Cin23, McV08]. refracted [Ren14, ZW17]. Refresh [LM04b]. Refusal [Yas83]. Refusals [Abr97]. Regeneration [Chu75, Sig88, STW94]. Regenerative [ARW94, Ath88, Gzy77, Imh74, Kin04, LS07, Mil81, MY01, MS86, PZ07, Ser80, Ser92, Sha80, Tho85, AR15, Bai20]. Regime [Guo01, AB16, GKP20, Jia15, Kun22, LS20, LY20]. regime-switching [GKP20, Kun22]. regimes [Kun22]. Region [Cha75, EE78, Hüs95, Kin65, RW64, Sud76, HZ22b, KKS15, Tzi16]. Regions [Bor94, Dav81, Hen83c, SK90]. register [LT23]. Regression [Ano72f, BM90, Bri86, CG87, Dur86, Han71, Hea82, Jon86, KS78, Kul87, MJ85, Osb82, Rob72, Kim17, LR24]. Regressions [Smi82]. Regressive [Rey72a]. Regular [Ano65d, Ano66c, Ben68, Ber73, Coh73, Col76, FHL13, Gra12, Kei64b, Kei65, Pol90, BHZ22, CP14, DIÖ23, HE16, HJL14, Rob22, Sal13, SRT^{+}16]. Regularity [Ger83, Len81, RD97, Whi79, DN19, Kim17]. Regularly [Ber92, BL10, NW70, DL17]. Regulated [GMRY07]. Reinforced [BCPR11, AGP13, BCPR16, Cri16]. Reinforcement [DW89, DLM14]. Reinforcement-Depletion [DW89]. Reinsurance [BCR11, DJ05, Day70, ES11, GY08, MS06, RR12, Zen10, ACVZ20, BFP17, BS22, CY16, LY23]. reinsurers [CY16]. Related [AM07, BR89, BLBL10, BRY04, BKS02, Ber81a, Bor99, Bra70, BT12b, CP78, CG69, DK90, Dwa69, FW10, Gan94, Gne00, Hea67, Hua95, Huf86, Huf87, Kuz10, MW90b, MY91, Nak74, NRS85c, PL81, Sta87b, Sun92, Voi96, Voo87, Wes99, Yao93, Zac04, ZL10, AB16, FKZ13, Gol23, JF23, KW14, Van22, Win15, Zhe09]. Relating [KS93a, PR08, Whi82b]. Relation [Bra72, Bra82, Bré91, Bru71, Fak90, Gha90, HN71, Lau80, Miz06, Råd72c, The79, Yao13]. Relational [CD19]. Relations [Ban96, Gla80, HJL14, Mil94, Miy82, Mor80, Reh77, Möh21, TK15]. Relationship [Bha90, Bra81, CDL88, CT76, Dav83, Lit94, Pré94, Sav88a, Toy97]. Relationships [Gup84, JKM12, KM88, LWB08, Lyn00]. Relative [Hol00, Jen72, Kru04, MZ01, SD94, Slu77, Yan74]. Relatively [KA09, MT06, RB98]. Relatives [Pak08]. Relativistic [Gil67]. Relaxation [DJ18]. Release [BH01, BVJ00, BC75, ÇP72, De 85b, Gly89, KS01b, Mor67, Mor69, TR67, Yeo74, Yeo75, ZC95, LLT^{+}22]. Relevation [GKJ80, LR90, LR92, LL24, SP17]. Relevation-Type [GKJ80]. relevations [YZZ21]. Reliability [Agu03, ADH93, AZA18, Bha91, BM01a, BENP94, BK00, CM92, CDD91, CP90, CS88, Di 99, Die99, Ebr01, Ebr08, Ebr11, EY11, FS80b, Gås98, GR98, GXY18, Gou98, Gri80, HNF85, HL05, KB87, Kou96, LDL95, LSC97, Lin87, MT95, Mi02b, Mus00, NSK08, NSB11, Nay87, PH88, PR95, Pek99, PT97, Phi81, PP73b, Råd76c, Rev93, Sen94, Sen95, SS93, Sko00, SK91, Son96, Ton85, Tor10, VK85, VK10, WR97, XL98, AA14, KK16, Mar15]. Reliable [Tuf99]. Remainder [Eng80, Sgi01]. Remaining [Ano83f, BAAB00, Gha90, KNS73, Nat82, Nat90]. Remains [Nak91, Ros73]. Remark [Bra80, Bré91, CPLZ07, Jof77, Ken92, Buo15, Pro18]. Remarks [Bra02, DG89, GNRS86, Gol75, Grü89, Kin77, Kul87, Lau80, LS08b, Lot84, Pak83, Pak87, Pak93, Smy76, Spe73b, Tak76, Wei80b, Zhe09]. Removable [LT04b]. Removal [Cla99]. Rendezvous [ABE99, Alp07, AW90, PSZ04]. Reneging [ZAFP06]. Renewable [SK91]. Renewal [AM16, And79, And81, Ano67d, Ano02e, ANT78b, AG81, BS19, BM05, BOR01, BL01, Ber65, Bli89, Blo91, BGH71, BIMR22, Bor04, BWZ85, BSS81, CL85, Che94, CK00b, CK07, CG69, CH70, Cse79, Cse91, Das96, De 85a, Del85, Dsh01b, Grü89, HF91, HTW09, HF71, HLS93, Hun86, ISK95, IRS23, Jam66, KS93a, KNS73, Keb97, KS08, KS86c, Kol82, KNT10, KPW94, Lag05, Lam91, LJ78, LT04a, Law70a, Law70b, Law71a, Law71b, LPS06, Lib85, Lin08, Lin82, LR96, Mar73a, McN85, Mei10, Mod67a, Mur87, Neu78, Nic84, NO85, Nie85, Pia00b, Råd72b, Råd72c, Reh77, Rob05, Sah83, Sch75a, Sen81, Ser92, Sgi96, SZ92, SS83a, Sie81, Smy76, Spo01, Ste78, Tan07, Teu68b, Tor10, Whi72b, Whi85a, Wil99, Win89, Yan90, Yan88b, Yan88a, YZ06]. Renewal [Yu11, BMZ23, BCS23, DM19, EOV20, GR18, JOC13, Kev17, RY20, Pia00a]. Renewal-Process [Whi85a]. Renewal-Process-Type [Lag05]. Renewal-Theoretic [Cse91]. Renewals [Eng80, Moh76]. Renormalization [Ros87b]. Rényi [AN22, Ano71e, BT76, Bro69b, DS86a, Hey68a, Kap77, Ken70, Pyc21, Ste78, Yao96]. Reordering [Top86]. Reorganization [VOZ95]. Repair [Ano87e, AJ00, BBS85, BLS93, BEN98c, BP83, FH90, Fin04, Fin07, Hu96, JCM98, JK87, Kij89, LS03, Nat90, OA70, SG92, She96, SK91, SZ91, Yeh95, Yeh03]. Repair-Cost-Limit [JCM98]. Repairable [AC02, Ano93g, Ebr90, FS80b, Kij89, LS98, Mil76, SZ91, SV03b, Yeh91, Zha94].

Repairing [GL84a, RG88]. Repairman [DLR80, Igl65]. Repairmen [Fro93]. Repairs [CF11b, FH91, Rig96, MC13]. Repeat [Van91b]. Repeated [AGC97, AGC99, CS77, CC79, DDM98, EW88, Fal86, Got85, Han87, Kam79]. Repeating [DKK08, GH90]. Repeats [Ano07e, SS85b, Spo07, TSCD08, ZS92]. Repetition [Kar83]. Repetitive [Ano82f, Pro82, TG99]. Replacement [AHS78a, AH86, ADY07, Ano88g, AB86a, AG87, Ber76, Ber80, Bha69, BLS93, Dav64, Ebr97, FH90, Fel76, GW94a, GL84a, Kla94, Lan88, Les86, Man69, Man71, Mi02a, MA69, MN82, NO74, Nak79b, Nak79a, Nel82, OKM86, RG88, Sav88a, SZ92, SS89, She92, She96, Sie81, Tan78, Wan77b, Yeh91, Yeh95, Yeh03, Zha94, Zuc80b, AA17, Glo23, Ino23, ISK18, SP17]. Replacements [BAAB00]. Replica [BDT22]. Replica-mean-field [BDT22]. Replicates [CC94]. replicating [Bra14]. Replication [CCH95, Cow01, CPT99, Pia00a]. Replications [MI83, UB74]. Reply [GNN86, Wei79, Wei80b]. Reporting [Kar74]. Representation [BBS90, CC06, Cam97, CP78, Die82, Gar74, HO74, Kal89, Kat77a, KMRV00, Lar98, Lin04, PK80, Rie80, TR89, Wil99, DHKR19, FZL13]. Representations [Bai20, Bro04a, CC96, CM02, Dal69a, Hal78, Han65, Hey86, NBS08, NSB11, Pen96a, PB20, RMV97, Ton73, Zha10, vD15, ADJ13, BM14, BFK16, KZN15, NFS20, Oes18]. Represented [Hey82a, Ors87a]. Representing [Kal90]. Reproducing [Kin77, MM94]. Reproduction [GY65, HS78, Pur69, Rah98, SR70b]. Reproduction-Dependent [Rah98]. Reproductive [Hey82a]. repulsion [BO18]. repulsive [HZ22b, KP23]. repulsiveness [MO21]. Request [AGC99, PP98]. Requests [AGC97, JRS06, PPD97, Rod95b]. Required [Ano85f, BT82, FG84, NR84]. Requirements [HRZ07, Web82]. resampling [QZ19]. Research [Git74, Mor66a, Rao66, Sam70]. Resequencing [HB02]. Reserve [Loi05, Sch98]. Reserves [Fos75, Jia15]. Reservoir [Imo01, Leh71, Llo71, OL65, Pli75]. Reservoir-Induced [Imo01]. Reservoirs [GG00, Peg75b]. Residual [GST00, GE82, GAB12, GL84b, Kul87, LZ02, MJ85, Mor78, NE07, NBS08, Neu68a, SS91, WC04, BBH14, BHB18, FZL13, Gup13, KZN15, LSW19, PB14, TSN17]. Residuals [Dur86]. Resistance [Mat84]. Résolubles [Lef95]. Resolvent [Att89, Bre13]. Resonance [RW64]. Resource [CVW84, Git72, Git74, Gla90, HRZ07, LMS72, O’C98]. Resource-Sharing [CVW84]. Resources [Cow92]. Respect [BM01a, Gra83, Tan83a, VAK92, AHO20, AR15, GW17, Jan21]. Respondent [MLB16]. Respondent-driven [MLB16]. Responding [Kal88]. Response [AF83, AIH88, HP95, KST99, MM71, SB17]. Responses [Cam86, Law71a, Law71b]. Restart [MSS09, APZ13]. Restarting [Miy92]. restaurant [Möh21]. Restless [WW90, Whi88]. Restoration [Tet02]. Restoring [Hüs95]. Restricted [Cha83, Eng65, Fan82, Jan88, LSC04, Sch05, Unk79, Van90]. Restriction [TW01]. Restrictions [CYF96, PSZ13]. Result [Arn68, BR98, Blo73, CD95, CZ98, Du95, GL71, Git73, Hai03, HS00, Hey70, Hey73, Hop76, Kap77, LS04, LT90, Sch92a, Sch91b, Spe73b, SG88, Yas93, ZAFP06, Rüs08]. Resulting [Gav69, Hos06, Kel83]. Results [Ali70, Ano74g, Ano75f, Ano86h, Bal01, BNGS04, BK86a, Bäu97, Ber77c, BAAB00, BP96a, BR69, Bro69b, CL85, CK00c, CC79, Cla94, Coh73, Cow80, CCG90, DR84, DW72, Day83, Don84a, Don93, GW85, Ger77, Gir86, Gre69, GS95, Gup84, HS74a, Hai83, HB75, Har90a, Has65, HT91, Her74, HKN89, HM95, Höp86a, Huf88b, JR82a, Kat11, KP96, KK00, Kij89, KS91a, Kot66, KSC12, Kuc82, LB89, Law70a, LK01, Llo89, Man71, McC85, MW90b, Möh98, Möh05, Mur89, Nag82b, Nak79b, Nyr94, O’F89, Ott79, OW79, Pel07, PM73, Pra64, Pra68, Pur69, QL96, RD97, Sau75b, SKF79, SD94, SZ92, ST73, Str75b, SC73, Tan83b, TP76, War76, Wes73, Wol84, Yas84, ZL10, ZLA99, ZG88, ASKK20, ABS15, ABSGS16, BPSS14, BR16, Fox15, HI20, Ino19]. results [JSW19, KLP22, KS20, Sta20]. Resume [CN86, Fak86, Fak87, Sch72, Whi71b]. resurrected [Mar21]. Resurrection [Che02]. Retail [Agn75]. Retirement [GH92, EFS14]. Retraction [NE82]. Retrial [FF91, KKK07, Kul86, LH02]. Retrials [Kul83, Pou90, DO16]. Retrieval [Bov99]. Return [BCMW83, Dal79a, ET83, Ios86, Joh06, LL78, Lee11b, New68a, MSV14]. Returning [GW87, Gre89]. Returns [BDM12, Tak70]. Reuter [Bin92]. Revealed [Alp07, Ano87e, JK87]. Revenue [Hin04]. Revenue-Equivalence [Hin04]. Reversal [GW94a, Spo04, Tav80a]. Reverse [Est75, Tzi11]. Reversed [BMPR07, BM01a, CFZ96, Kij98, Sta77, BMR19, LWM20, Tav20]. Reversibility [CDL88, GJQ06, Lit92a, Lit94, Nat77, Ōsa85, RD97, Wei75b]. Reversible [CM15c, HR88b, Lit92b, LPP12, LZK06a, Mit04b, Pha74, Pha75, RT01, Suo79, Wüb12, ZWV10, ACK20, Ger23, Kam17, KS23a, Lam14, PU16]. Review [Ano65d, Ano66c, AB86b, BNGS04, Fel78, Kei64b, Kei65, Ram81, Sah83]. Revisited [AOR03, Ger11, KS89, Lyn00, Neu82, OS86, PP77, PP88, Pre97a, BLSGS13]. revolution [Let94, Let94]. Reward [BHPZ06, Coo98, Das96, GZ02, KL03b, MS91, Mas93, MS02, MM94, Ros82, Sol96, SK98, Ste06, TH11, VK00, VS06, CCdOS15, KOH23]. Rewards [FF09, Rig11]. Rheology [BJ07]. Rice [BL12]. Richard [DMR02, Gan84b]. ricocheted [Wat23]. Riddle [Sen82b]. Riemann [HJL21]. Riemannian [Rap08]. Right [FPKMP98, FP05b, KS76, Rei67, SCZ77]. Right-Shift [KS76]. Right-Spread [FPKMP98]. Right-Turning [Rei67]. rigid [KL22]. Rigorous [Sud01, Sud05]. Ring [Bam92, Gra99]. rings [LT13]. Risk [Agn72, AT06, ARZ08, BCL09, BCR11, Bal01, BK12b, BP99a, CCdO05, CL09, CP11, DR09, Dol10, DWW11, FM12, GT06, GY08, Gay04, Guo02, GH21, HT12, Igl69, Ish74, Jac12, KK06, KNT10, Kuh85, KP07, KZ09, Loi05, MS06, Muc02, Nag01, PLC03, RZ07, Rob05, RR12, Sch92b, Sch98, SE12, WW07, Web07, Whi86, Wil96, Wil99, YGN05, ZY08, ABI14, AES17, BPPR16, BFP17, CCCSdO24, CCFM17, DZ13, DHJ15, DHM20, DM21, Des15, ESC15, Fin13, Fro15, GS19, HJ13b, LPS17, LWW18, Nyq17, PL03, Ren14, Sha13, SZY19a, WZ18, YWK14, YSL17, ZC14]. risk-seeking [CCCSdO24]. Risk-Sensitive [CCdO05, Kuh85, Whi86, GH21, Des15]. Risks [Che11, CY08, DG06a, KR08, Rüs04, CC15, PR13]. Risky [DC77, HT12, Hey99, KS71]. Risque [PL03, PLC03]. River [Hut90]. Road [HTW73, Haw68, Sri91, The69, YY11, YW64]. Road-Hog [Sri91]. Roads [BR79]. Robbins [Bru05, BS09, GI11, McL76, MD18]. Robbins-Munro [McL76]. Robertson [Rog23]. Robin [LT94]. Robotics [MZLSRS02]. Robust [BKM19, HW83]. Robustness [HL23, Joy95, Möh98, SS07, Twe75a, Sha23]. rock [FHI17]. rock-paper-scissors [FHI17]. Role [DN04, GT08, HI01, PT86, Des15]. Roles [Hal83b]. Room [Coh71, Dsh89, Hok75, Hok77, Ros73]. Root [Hai70, Sat77, Gor14]. Rooted [Tak92a, HS22, HLF19, Jon19]. Roots [RS69, Rud76a, WCW05, SGC16]. Ross [Hey82b, Zhu14]. Rouault [Khm11]. Rough [Nyr94, GJP^{+}21, JPS18]. Roulette [HG81]. Round [FM97, LT94]. Round-Robin [LT94]. Routeing [BDO07, DS08a, DG16, DG06b, OV03, SZ06, AKLL18]. Routine [JK71]. Routing [ABGH00, BM89b, CSZ97, KW85, LT94, Top86, VM21]. Row [KZ88]. Row-Continuous [KZ88]. Rows [DKK08, GH90]. Rubin [Hop76, Sha74a, Sha77, Tal80]. Ruin [Ano78h, AN95, AT96, AK96, Bar64, BL10, BR96, Cai02, Che11, CP11, DR09, Fro05, GT06, HT12, HH07, HPSV04, IK04, Jac12, KP22, Kar77, KNT10, Møl95, Nyr99, PLC03, PP07, PT23b, RR12, SB06, Tan05, YGN05, ZY08, ABI14, ACVZ20, BPPR16, BS14b, BS16b, DHJ15, DHM20, GS19, HJ13b, IK06, LY22, PL03, Rob14, TTI22, YSL17]. Ruine [PL03, PLC03]. Ruinous [KK06]. Rule [BF95, BP04, ÇP72, GM05, Gly89, Gou98, HD71, Mor69, She13]. Rules [All04a, AM08, ASC98, BS06, BCS08, CCPW92, CL97, CL98, CS85a, Cow04, ES87, FH89, Had05, KPSC08, Lam84, YWR82, Yan90, HZ20]. Rumor [JMZ11]. Rumour [Dun82, LP94a, OT77, Sud85, ES23]. Rumours [Pit90]. Run [AHN90, Ery08, FWL03, Ksi89, MP04, SK86, Ton85, BS15d, BRSC22, GXY18, RU18]. Running [Fer95, KO12, LS11, CPN18, GR14, JF23]. Runoff [Jon19]. Runs [Ano77f, AM10, Ban88, HK15, Imh74, KPP95, MO01, Raj74, WX08, KLY16]. Rush [New68b]. Russian [Eks04]. Russo [JZG11]. Ruth [Gan84a].

s [Fel78, HNG93, Ram81, Sah83, Zhe91, KL89a, LT73, Mak10, Ano84e, De 85c, Ken72b, Miy86, Whi72b, Whi80a, Yao13]. S-I-S [KL89a]. S. [Jof77]. Sabr [RV08]. Saddle [CYF96, KD78, ADJ13]. Saddle-Point [CYF96, ADJ13]. Saddlepoint [Dan82a, Sko87, Wan90, Wan93, BN17, But19]. Safety [Ave09]. SAI [PL17]. salesman [BS14a]. Same [DDM98, Nis74b, Zho04]. Sample [AFGP91, ACSS11, BM89a, BT76, Boo79, Bre69, BdH87, CCdOS15, Dai04, Das96, Dav82b, DM01, De 74, DH85, Di 98, DMT11, DT11, Edd80, Gel77, Gne99, Jun94, Kal90, Lac05, Les10, LSC04, O’B80, PV12, RH81b, Ses96, Sti94, Van89b, Voo87, Voo89, Wei69b, Wei82, Wol88, BR18, GP13, Sha13, UY23]. Sample-Path [Sti94, Wol88, CCdOS15]. Sampled [Bri86]. Sampler [WD99, RU18]. Samplers [FMRR03, BJ24]. Samples [BO90, GH09, Hop86b, ML67, Rog78, ST88, GH13]. Sampling [Asm09, BCP00, Bel71, Ber81a, Ber83, BL10, BSC04, Bri73, Che76b, DF89, Geo87, Han90, Joy95, Kem98, Kla94, ML86, SC75, Saw83, Yan90, BDL19, BCS23, DL17, GKS14, HS22, Kim16, LGD20, MLB16, Nyq17, Spi15, Tsu17, Tsu18, Glo23]. sampling-consistent [HS22]. Sampling-Tagging [SC75]. Sankoff [DFSD01]. Sankoff-Ferretti [DFSD01]. SAT [MS10]. Satellite [Jac99]. Satisfied [Don77, Eri16]. Sattolo [Mah03]. Saturation [Ana92, BF95, New68c]. Saunders [Raa80]. Savage [LS17, Yao07b, YC08]. scalar [YL16]. Scale [DS21, DGM04, Duf96, FM06, Kat05, Ken82a, LS09a, Ros87b, Sur08, Wat80b, Yuk06, FH13, FS17, HH23, MBvLW16, Ste19]. Scale-Free [Kat05, Yuk06, DS21, FM06, FH13, HH23, Ste19]. Scales [BKMS04, Ros08]. Scaling [BY22, CFD09, DE11, DL06, DLMR17, PS16, She13]. Scalings [Duf96, KM15b]. Scan [BK06, CK00b, Cre77, Ery10, Fu01, FWL12, GW85, KL17, WGF13, YML17]. Scanned [Lee80]. Scattering [FF11]. Scenery [Li12]. sceptics [ES23]. Scheduled [Dek78, AG12]. Schedules [KS86b, Kol95, Lee97]. Scheduling [CR94a, FA85, GG77a, Git81, Gla79a, Gla80, Gla91, Gla93, HKN89, Käm87, Leh88, NTW89, Pin80, PC85, Rig96, Rig97, SW05, Sto08, VPT92, Web82, WVW86, WP80, PW16, Pra82]. Scheme [BP97, BK64, BGT04, Las94a, LS07, Pha94, PPD97, Rod95a, Rod95b, TG99, FCKS16, Glo23, Wan15]. Schemes [Agu09, AIH88, Bar82a, DG99, FKK91, MP04, GM15, Ino23, MBvLW16, QZ19, Spi15]. Scherk [Che01]. Scholes [Dol10, FPY04, Whi01]. School [Bro04b]. School-Based [Bro04b]. Schuh [Don84a]. Schweitzer [Las94a]. Schweizer [Des15]. science [JKL^{+}20, YL16]. scissors [FHI17]. Score [CZ99, DVJ04, LSC04]. Scores [AMW88, Tak64b, GM20]. Scoring [DVJ04]. Scrabble [AMW88]. Screaming [Tav22]. Screening [Ano82f, Dub81, GWMO79, Pro82, CF14]. SDEs [BHZ22, LS04]. Sea [ALO05, Car67]. Search [ABE99, Alp07, ASB94, BP04, Bat92, BK73b, CG97a, Cor80, Cor81, DGK98, DG99, GG87, Hel71, Lam84, MC90, Nak86b, NR84, Pha91, PR67, Sar73, Sta91, Sub81, Web86b, Zuc83, Zuc86]. Searchers [MM06, Nak86b]. Searching [MM06, McC74]. Seasonal [God77, LS79, Lun94]. Secants [EE81, Kin69b]. Second [Bla94, BCS93, Bra83, Byt81, Hor80, HZ82, KLM04, Mar86, Meh67, Neu78, New70, PM03, Pit80, RS04, Rip76, Tam79, Tas97, Wyn73, Xu17, AB15, BJMK18, RY20, Wir18]. Second-Order [Bla94, BCS93, Bra83, Byt81, KLM04, Mar86, Meh67, PM03, RS04, Rip76, Xu17, Wir18]. Secondary [Sah71a]. Seconds [HM97]. Secretary [AHBT82, CRS97, CB83, KSC12, Pfe89, QL96, RF88, Rog87, Smi75, Yeo97, Yos84]. Sections [Ano78g, CO85, Dav76, Mat76, Rob86]. secure [VI14]. Security [Pla83, SZY19a]. Sedimentation [Gan88b]. Sedimented [Lit69]. See [MW90a, PRS08]. Seed [KKL01, BCKS13, CPSJ22]. seeking [CCCSdO24]. Segment [Ano04g, FS78, PC76, Ste04]. segmental [GM20]. Segments [Bit69]. Segregating [Gri81]. SEIS [Fox16]. Seismic [BCIM01]. Seismicity [Imo01]. Seismology [VJ01]. Selected [Bla92]. Selecting [BP00, HY02, FHI17]. Selection [AKM10, Ano88f, ACSS11, Bru87, BL98, CNO^{+}84, CC08, Esh73, Gne99, Gra00, Grü98, HD71, Hen73b, HH85a, HØ07, Kin78, Kin88, KPSC08, KA03, Mod66, Nag82b, Pet81, Pre06, RT91, Shi86, Sta04, Swe94, Web86a, CS20a, CP17, Ros14a, Mod67b]. Selective [Che76b, JP69, Law70a, Law70b, Law71a, Law71b, SR70a, TR65]. Selectively [Saw79]. Selectively-Neutral [Saw79]. Selectors [Gli00]. Self [Ano87i, Bla94, Bra85, CL97, Chu18, FS23, FS07, FSW12, GL84a, HLMOL21, HO74, KP08, Lam84, LLS84, Law83, NP00, Oak75, RG88, VJO84, Ver98, CH13, CM16a, DHJ15, KPV21, Vid20]. Self-Avoiding [Law83]. Self-Correcting [Ano87i, VJO84]. Self-Excited [Bra85]. Self-Exciting [HO74, Oak75, HLMOL21, CH13]. Self-Interactions [Ver98]. Self-normalized [Chu18, FS23]. Self-Organizing [CL97, Lam84, LLS84]. Self-Repairing [GL84a, RG88]. Self-Similar [Bla94, FS07, FSW12, NP00, CM16a, DHJ15, Vid20]. Self-Similarity [KP08, KPV21]. Sell [AM08, YYZ09]. Sell-at-Once’ [YYZ09]. Selling [BF97, MS02, SKK06]. Semi [Ano66e, Ano70e, Ano93h, Bal79, Bal99, BMY02, Bar82a, BR87b, Bur76, Che70, CT72, Çin67a, Çin67b, CT95, CH84a, Cse91, DHKR19, Dos79, Dow66, FT78, Fel76, FP05a, FH72, FKK91, Ger83, Ger76, GL03, GBW92, Ham82, Haw70, JR82a, Kao74, Kas66, LS92a, LP85, Llo71, MS91, Mas93, McC80, McN74, Meh79, MSZ04, Neu66, NC72, OL65, Pil91, Pur75a, Ros82, Ros70a, Sah78, ST01, Ser80, SH03, Smi73, SK98, Ste06, SR88, Szá74, Tor10, Twe75b, VP92, WZ65, WF80, WT05, Zuc78, DN19, Kan23]. semi-contractive [DN19]. Semi-Dominant [CH84a]. Semi-Groups [Ano66e, Dow66]. Semi-Infinite [Ham82, Llo71, OL65]. Semi-Linear [MSZ04]. Semi-Markov [Ano70e, Bal79, Bal99, BMY02, Bar82a, BR87b, Bur76, Che70, CT72, CT95, Cse91, Dos79, FT78, Fel76, FP05a, FH72, FKK91, Ger83, Ger76, GL03, GBW92, Haw70, JR82a, Kao74, LS92a, LP85, MS91, Mas93, McC80, McN74, Neu66, Pur75a, Ros82, Ros70a, Sah78, ST01, SH03, Smi73, SK98, Ste06, SR88, Szá74, Twe75b, VP92, WZ65, Zuc78]. Semi-Markovian [Çin67a, Çin67b, Meh79, NC72, WF80, WT05]. Semi-Pareto [Ano93h, Pil91]. Semi-Reflecting [Kas66]. Semi-Renewal [Tor10]. Semi-static [DHKR19]. Semi-Stationary [Ser80]. semi-strong [Kan23]. semicoherent [MR15]. semicontinuous [SS17]. Semiexponential [DT11, KLP22]. Semigroup [Phe96a]. Semigroups [BCL98, CSC13]. Semilinear [LMW98]. Semimartingale [Aas85]. Semimartingales [Ken92, TT86, PG13, Wan15]. Semiparametric [BRRN16]. Semiregenerative [Pra88]. Sengupta [Bré91, Toy97]. Sense [CNS99]. Sensitive [CCdO05, Kuh85, Whi86, Des15, GH21]. Sensitivities [Eli09].

Sensitivity [Cao89, CT95, Mit05, MLB05, TH11, VK00, VM21, JF16]. Sensor [KS16]. Separable [Lel07, WW04]. Separate [GSS11]. Separately [Bax83]. Separation [MZZ16, Lef95, Lef95]. Sequence [Alp89, Ano81e, Ano91e, Ano10e, Ant01, Blo82, BP00, Cam73, CCC03, Che80b, CP88, Deh85, DJ95, FN70, Fu01, FC02, FWL03, FC03, GS99, Gra72, GR01, Haa92, HTW73, HTW09, Hey80, HL71, IA07, Kar83, KW03, KA97, LL77, Mar07, MCC03, Nue10, Pak87, Pha88, Pol78, RD99, Sel91, Spa91, Tam79, WS96, Wol77b, Woo82, WDLC09, Yan74, AI16, FHC18, GM20]. Sequences [AC94a, AA90, AT91a, AT91b, Ano85f, Ano07e, AMW88, AM05, Bal99, BR87a, BL09a, BL00, BT82, Blo84, Bor99, CFZ96, CK00b, CZ99, CV01, CS75, Coh81b, Dek78, HN01, Hol07, Hol08a, Hol09, HS12, HS10, Hüs86a, Hüs93, Kal92, KBBW07, Kim94a, Koç95, Kur71, Mak10, RF88, SS83a, SS85b, Spo04, Spo07, Tas97, TW83, TO92, Via95, Wal76, WCC03, WC95, AR15, FCP20, JF16, Pap17, ZS92]. Sequential [BKJ16, BC67, BB88, BS70, BS82a, Bre69, CNO^{+}84, Coo74, Coo89, Cow04, FS71, Git73, Gne99, Hal69, Har91, Hen73b, Hol81b, IS86, Jan83, KPSC08, Man83, McN85, MWCY01, Nak86a, ÓT21, PR67, Rig90, RW79, SS83b, SS01, SKK06, Sta04, Wei78, AZA18, BJ14, BHB18, BN13, CF13, TP79]. Serial [Ary72, Dal69a, PV12, SP92]. Series [ARM98, Ano72f, Ano73g, Ano75f, Ano82h, Ano98f, Bar71, Bar72b, Bar74, BMPR07, Ben70, Ber83, Bli89, BI05, Bri73, Bri86, BH73, Car67, CA98, CI78, Dai04, DN80b, ENPS86, FH90, FC03, GG74, GU92, Han71, Han73a, HB75, Har73b, Hen86a, Hey86, Hey88, Hut90, JP95, Joe96, Jon86, JS98, Kim00, Law92, Lit94, LB88, Liu04, Ma04, MP70, Mas87, McK82, McN82, MDG09, MMD11, Mor67, Neu68b, Oga80, Ors89, Oza80, OI01, Par86, PT81, Qiu91, Rob72, Rob86, Sha73, Sol86, Tan86, TR89, Tso92, Web79, Wil82b, Wil82c, ZT01, Zip11, HPP14]. Series-Parallel [ENPS86]. Served [Bal72, Enn69, Fak86, Fak87, FR10, LH02, VB02]. Server [AGC97, AGC99, AIH87, Aye07, Bal72, BM97, BM89b, Bha68b, BM94, BBP03, Box75, BR69, BS93, CL11, CD95, Coh68, Coh76, Con71, Cra74a, Dal69b, DJ69, DR84, DK90, Du95, Fak86, Fak87, GH94, GW94b, Har90a, Har73b, HL81a, HP93, Has64, Has65, Hil67, HT78, HS93, HB02, Igl65, Kap71, Kei78, Ken72a, KM90, KTW69, LSC97, LR04, LT04b, MRP04, Min78, MY91, Miy01, Nai71, Naw83, Neu66, NC72, NR84, NR78, OV03, PS89, PS86b, Pra70, PT94, Pre97b, PL81, RN80, Rub85a, Rub85b, Sah71b, Sha66, SS87, Sta87b, SP99, Sto91, SG88, Tak71a, Tak74, Tam70, Tan94, Tho85, Wil97, Wol02, Wol84, Xu98, YSS91, Yu74, KSBM18, Ros14a]. Server-Systems [Kei78]. Servers [ABGH00, Ano87h, Ary72, AIH93, AI94, AIGR06, CI72, DLR80, Dow98, DKO84, Fak83, FG84, Ken72b, KAR11, Köl74, Köl79, Lan86a, Leh86, NE81, NE82, PS86b, Rig96, SW05, Sob90, SM97, Web78, CHR15, HR15, IJB18, Ros14a]. Service [AA06, AMR01, AN84, Ary72, AIH87, AIH88, APS04, BFL05, Bha68b, Box75, BG87b, Cao84, Cha97, Coh84, Con71, DT80, Eco00, Eco03, Fak87, FT80, Gas64, GL76, Gla79a, GADL04, GR74, Had69, Had74, Hal79a, Har90a, Has64, HB02, Ish04, Jac78, KK96a, KR88, KS80a, KNR05, KG02, Lan87, LG81, Lel05, LZ94, Li96, Min78, Miy85, NE81, NE82, Neu66, NR84, O’B75, PS84, Pea67, Pes68, Pra82, Pur75a, RS04, RSY90, Ros73, Ros75, Sah71b, Sah78, Sch88a, Sel97, SS85a, SY86, SS87, STC93, SM97, Tak92b, Tan94, Tom72, WP80, WT05, Ye07, Yu74, FS17, PSZ13, ZX17]. Service-Time [BFL05]. Services [CPS96, Çin67b, GP04b, NC72, ZZ04, BvdHvL17]. Servicing [Pea65, VS90, Van91b]. ses [HK94]. Set [AB86a, CNS99, Dos85, FHT78, Gre80a, Har91, Jan78, Kab83, KN05, KSHH98, Mol86, Moo73, Con22, CM15c, Mar16a]. Set-Chains [Har91, KSHH98]. Set-up [AB86a, Dos85]. SETAR [GP91]. Sets [Ail68, Ath88, Bau84a, Bau84b, BL89, Gro80, Gue99, Gzy77, MC70, Mas82, MS05b, RR98, Str75b, Wal76, Wei89, AFP22, Bai20, HS14, RS22, RNJ17, MC71]. Setup [HN00, LZ94]. Setups [CNS99]. Sevastyanov [HMY17]. Several [BC95b, BWZ85, EJT05, Gan94, Gas64, Gra00, HL04, Kei78, Köl74, Köl79, LH06, MS85, PL73, Sch88b]. Severities [BB05, IK04]. Severity [Pre02, Rad76b]. Sex [GT86, MM09]. Shadowing [YZ82, YZ85]. Shalt [Ibr07]. Shannon [AB16, AES17]. Shape [Ken91, Le94, Møl88, CPSJ22, CdLH^{+}23]. Shaped [Bal95, BLSW08, BLSW10, Dan69, GS72, Sav03]. Shapes [Ano66d, FM10, Hor65, LZK06b, PC16]. Shapley [GR21, GR24]. Shared [NSB10, O’C97, AZA18]. Sharing [AF83, AIH87, AIH88, APZ11, Coh84, CVW84, CS88, FK89, GR98, GP04b, HRZ07, HST83, KMM81, LDL95, LS86, Moy08, Ott84b, Ram84, RHK90, Sch88a, Sen89, WT05, GM15]. Sharp [Bov99, Bro15, BL98, CK00c, HL05, Pla85, PR13, AI16, Dal16a]. sheep [Lin16]. Sheets [Par75]. Shelf [Nel77]. Shepp [Glo23, Ksi89]. Shift [KS76, LT23]. shift-register [LT23]. Shifting [Cli83]. Shifts [Rud76a]. Ship [BC67, FS71]. Shiu [BPPR16, BFP17]. Shock [AHP75, And87, And88, AG87, BS78, CF09, CF11a, FP94, Fel76, GE82, GM84a, Got80, GL84a, ISK95, Kle81, Li00a, Pel94, Sav88b, SS83a, Sko00, Tha81, Won97, Yam89, Zuc78, GXY18]. Shocks [AH86, Ano88g, KFM06, LL93, MO01, RG88]. Shop [Rig97]. Short [ETV88, Rek23, SB78, KP23, NR20]. short-memory [NR20]. short-range [KP23]. Short-time [Rek23]. Shorter [Tur00]. Shortest [Ale96, Gra00, Hal85, Lit74, STC93, Win77]. Shortest/Longest [STC93]. Shortfall [Dol10]. Shot [BD12, Dal71, DJ05, DT11, HM95, Huf87, Kel09, Lan84, LS91, Lun96, LBP99, LMX04, McC97, Pap71, Smi73, SW94, BN15, IR20, PT23b]. Shot-Noise [Kel09, Lan84, LS91, BN15, PT23b]. Showing [MS79]. Shrinkage [JS04b]. Shuffling [Got87]. Sibuya [Let19]. Side [Pol78, vLR13]. Sided [And79, Ano86j, BH12, CS86, CY65, Had05, Imh78, IM03, KS06, Pyk94, Rog00, Shi65, Iva14, Pat13]. Sidney [Ano00j]. Sierpiński [Shi02]. Sign [De 77, VP10, Zho04]. Sign-Changes [De 77]. Signal [DO97, DFM^{+}06, Dri75, Han67, Heu87, Miz06, Sob67, VI14]. Signal-Noise [Dri75]. Signal-to-Noise [Heu87]. Signaling [CZ98]. Signalised [Cow78]. Signalling [Wol75]. Signals [Den10, GL05]. Signature [NSB10, NSB11, ET21, NFS20, YBL23]. Signature-Based [NSB11, NFS20]. Signatures [Bol01, GSS11, GSS12, ASKK20, BV14, BN13, DD15, Mar15]. Significance [MCC03, Neu80]. significands [BT18]. SII [Cri18]. Silk [Miu77]. Similar [Bla94, Car67, FS07, FSW12, LPSV00, NP00, Pol01, Yap09, CM16a, DHJ15, Vid20]. Similarity [KP08, KPV21]. Simple [ADLN82, Ano66e, Ano79e, Ano83a, Ano86a, Ano86j, Ano06e, ASC00, BM05, Bar75c, BLL79, BLL80a, BLL80b, BP96a, BPT89, BN88, BGP80, CS86, CD98, Dav78, DG92, DF88, Dow66, FF96, Gan94, Gra85, Har80, Has82b, HS77, Iks04, Iva81, Jen72, Jer70, KP92, KP96, Ken94, Kim70b, Kin78, Kry72, KP83, LL82, Lad05, Lan81, LH07, Mac80, MPS21, Miy82, Pak87, Pap10, PS89, PR95, Qui76, Sen71, Sen73b, Sys94, Van89a, Van22, Van89b, Vrb85, Whi79, Wie79, XY92, YC82, Zhe91, ABI14, CP14, Kub24]. Simpler [BE73, LN94, Raj81]. Simplex [Rub77, BKP^{+}22, PT22]. Simplexes [Gat84b, GGZ19, NV20]. Simplices [Mae80]. simplification [FPS18]. Simplified [Fox65, WL97]. Simpson [SS99]. Simpson-Type [SS99]. Simulate [RB98]. Simulated [ABD01, AF87, Bél92, Bor92, CD98, FF03, Kol95, Loc96, Nie95, Win90]. Simulation [Ana92, ARW94, AR01b, BL11, Bor94, BSS81, BNS90, DP12a, DS86b, GR12, GM11, Gne00, HS04, IK11, Lam91, Rid96, Ros14b, SS86, Win89, ADJ13, BM18, BRRN16, Dai14, DQL19, DFV15, FCvS15, LLT^{+}22, LX18, PLRB23, QDZ21]. Simulation-Based [GM11]. Simulations [Coo89, Tuf99, NR22]. Simultaneous [DHM20, FHT78, HRZ07, Hos86, KM01, KPS81, Sag03]. Simultaneously [Coh67a, DH85]. Sine [BD88b, GS94]. Single [AGC97, AGC99, Aye07, BMY93, BMY02, Bal72, BBP03, Box75, CD95, Coh68, Coh76, CHPT97, Con71, Cow80, CT95, Dal69b, DJ69, DR84, Don76b, Du95, Fak86, Fak87, Fal86, Gla93, GW94b, Har90a, Har73b, HP93, Has64, Has65, HF71, Ken72a, KM90, Kin88, KTW69, KG02, LR04, MZ04, Min78, MY91, Miy01, Mut87, Nai71, Nel77, Neu66, Pra70, PT94, RSY90, Sah71b, SS87, Sta87b, SG88, Swe94, Tak74, Wil97, Wol02, YSS91, Zha01, ACK20, Ger23, HM23]. Single- [DR84, Don76b]. Single-Birth [MZ04, Zha01]. Single-Buffer [KG02]. Single-Lane [Cow80]. Single-Line [Fal86]. Single-Movement [CHPT97]. Single-Server [AGC97, Aye07, Box75, CD95, Coh76, Dal69b, DJ69, Du95, Fak86, Fak87, GW94b, Har90a, HP93, Has64, Has65, KM90, KTW69, LR04, Min78, MY91, Miy01, PT94, SS87, Sta87b, SG88, Tak74, Wol02, YSS91]. Single-Shelf [Nel77]. Single-Stage [RSY90]. Singpurwalla [Ano90f, Roy89]. Singular [Aas76, Ano98e, ANT78b, C.12, CMW95, Hal79a, Jen74, Las94c, Nie85, DFST15]. Singularity [Blo86, Fei79b, Pro96]. Singularly [ELNT11]. Sinusoidal [Qui86]. SIR [BL09b, BL15, Nea03, Nea07, Nea16, Sim23, Win15, Gan71]. SIS [BBN16, CP03, Cla05, DL06, Kes08, MP12, Nea08b]. Sister [CM69]. Sites [CH97, Gri81, KK01, Lan05, LV72]. Six [Gan10]. Size [AD97, Ano92g, Bal86, Bal90, BSW07, BG87a, BT12a, Cha04, Che81a, Chi68, CW69, CS72, DFSD01, Don86, Don91, Dun80, Fak87, FT80, Gne98, Gne99, GR96, Gol05, Gra83, Höp85a, Höp85b, Höp86a, Hop86b, JS00, JS04a, KK03, Käm91, Kes08, KKK07, KT01b, KT01a, Kle83a, Kle84, Kle85, Kle89, KS83a, Le94, Lef78, Li06a, Lip77, LV08, Mac80, ML98, Mas86a, Mi02a, Mil08, MMR02, Mor93, Mor80, Pol10, RY99, Rap08, RR75a, SG89, ST86a, Sel83, Sha75, SS83b, Spo01, SC73, Sud76, Sud81, Sud83, Sve95, Tai95, Tsa94, Tsa96, Tsa99, Ver69, Voo87, Voo89, WF99, Wan99b, Wat80a, Wau74a, Wil88a, XW05, ZJ04, FSJ14, FH17, Mia13]. Size-and-Age-Dependent [Wau74a]. Size-and-Shape [Le94]. Size-Biased [Don91, Gne98, Hop86b]. Sized [Har09, Web99]. Sizes [Dav82a, Gue97, Hal89, Möh94, Möh97, Saw79, SH82, TV88, VG85, VP92, ACVZ20, De 22]. Sizing [JRS06]. Skeletal [FFK19, FFK20]. Skeleton [O’C93]. Skellam [GK18]. Skew [AS12, LS09c, RRT22, CL14, LPN13]. Skew-Normal [LS09c, LPN13]. Skewed [BDM12]. Skewness [KX11]. skill [HR15]. Skip [Bor95, FP05b, Kij93, Peg73, MZZ16]. Skip-Free [Bor95, Kij93, MZZ16]. Skip-Free-to-the-Right [FP05b]. Skorokhod [BFK16, Pri00, RWZ13]. Slepian [GA93, Lin89]. slice [RU18]. slim [IFN19]. Slotted [Ana92, MT00, Sha88]. Slow [FNV11, Kyp98, Rub85a, Rub85b, Sto91, Gén23]. slow-fast [Gén23]. SM [DN99, DS04]. Small [AKK20, AR01b, Cai03, Cov09, DG06b, FM06, Gat84b, Hos06, JS20, MKN11, RWN02, Ros81b, SGS17, Yuk06, Spi15]. Small-Time [AKK20, MKN11, JS20]. Small-World [DG06b, FM06]. Smallest [Wei69b, Gab16b]. Smile [BF08, AKK20]. Smirnov [Dur71a, Gel77, Pol82]. Smoking [MC78]. Smoluchowski [McD78]. Smooth [BL12, HØ07, LM12, Pau87, Vil07, WG93, FV18]. Smooth-Fit [LM12, Vil07]. Smoothed [FH95]. Smoothing [AK86, Bon89, Her94, Her99, HLT00]. Smoothness [Swi76, KP22]. Sneppen [GMV06]. Snowball [Ber81a, Ber83]. Socks [KR08]. soft [KK24]. soft-killed [KK24]. Software [Agu03, BSC02, ZC95]. Sojourn [BNW13, Cao84, Cse91, DMS08, DS08b, GC71a, HP93, Hem90, HB02, Kat11, Kat12, Lem87, LRM23, LT72, Ott84b, Ram84, RS89c, Sah78, SD87, SS87, TTH91, Wau72, GM15, Ino19, Pro18]. Sojourn-Time [Ott84b]. Solar [HK85]. Solicitation [CB83, KS77, Yan74]. Solidarity [CT72, Sgi96]. Solution [ADW73, Ara94, AGC97, BC95a, BS02, Böh00, BB12, BG00, CS97, CYF96, Del85, Dia73, Dri75, GWMO79, GBG02, GC74, Guo01, Hen79, Hou12, Law80, Mar05, McN68, Mor71a, Neu94a, PS89, dNPG96, Pra80, RC73, Sap87, SW05, Sel97, Ser90, SH03, Sri74, Van90, WR74, Eri16]. Solutions [Ass83, Bar72b, Bes72, Boy69, CGR08, CY65, Gue81, HY94, Kur70, Kyp98, Ros87b, Sev67, SH70b, SW88, YZ06, MST13, Rek23, Van22]. Solvable [GNN85, New86, Ino23, Lef95]. Solve [FDVF93, CM16b]. Solving [Del98, Irl06, Jen74, Mar16b, Mil94, Sen89, Sev69, vdHKPS23]. Some [Ada75, AI00, AA90, ASKK20, Amm83, And70, Ano74g, Ano82i, Ano86j, Ano89f, Bal79, BR89, BB90, Bal01, Ban88, BNGS04, Bas71, BKS02, BK86a, BK87b, Ber77c, Bha91, Bil81, BPT92, Bra81, BG79b, Bre81, Bro79, BR69, Bro69b, CW97, Can73, CF11a, Cha75, CL85, CS86, CS77, CDD91, CY65, CY84, Cla94, Coh73, CP78, Coh81b, Coh83, CG69, CC96, Con71, Cre77, CH70, Csá94, Cse91, DR84, DM67, Daw74, Don84a, DG89, Don93, DF81a, DF81b, EMR71, Emb83, Eng65, Fei79b, Fen95, Fin07, GG74, GY65, Gap07, Gat87, Ger77, GNRS86, Git72, Gol74, Gol75, GL83, Gre69, Grü89, GS95, GS01, Guo02, Gup84, HS74a, Han89, HF91, Has65, Hea67, HT91, Her74, Hey71, Hey88, Hic83, HF71, HKQ88, Hoo70, Höp85b, Höp86a]. Some [Hör05, Hua89, HS93, HLS93, Huf88b, Imh99, Jac12, Jan83, JR82a, JSW19, Kat11, KZB05, KM04, KK00, Kij89, KS91a, Kij92a, Kni71, KK95, KB91, KSC12, Kul87, LB89, Les73, LK01, LZ02, Li04, Lil68, Lin08, Lit01, Llo89, Loi05, Lot84, Loy88, LN94, Lyn00, Mac80, Mad75b, Man71, MT95, McD78, Mi95, Mi02b, Mil79, Mod66, Mod67b, MR02, MMR07, Mor71a, Mul68a, Mur89, Nag82b, Nak74, NJS96a, Niu80, NRS85c, O’F89, O’N96, Ott79, OW79, Pag71, Pak76, Pak08, PC76, PS84, Peg73, Pek03, PM73, PP98, Pic80, Pic84, Pol71, Pra68, Pur69, RT90, RN80, Rao79, Rey72b, Rod95b, Rud75, Sam70, San06, Sat83, Sav88b, Sch86, Sen89, SD94, Set03, Sev69, Sgi99, SZ92, Sha72b, ST73, Sha74b, Sha80, Spe73b, Sri67, Ste85, Ste00]. Some [SC73, SH70b, Sze86, Szp85, Tak76, Tam85, Tan80a, Tan83b, Tho75a, TP76, Ton73, Ton97, TSN17, TP79, Wan77a, War76, Wat80a, Wat88, Wen86, Wil82a, Wit90, Wit93, Wol84, XY92, XH11, Yao93, ZRU68, Zei91, Zha92, ZT01, ZLW06, ZL10, Zho04, AF14, AW15, BCdlP17, BL21, CTKR83, EDSM15, HM17, LS19, MZ20, Ott20, Yor92]. Somewhere [Dav73]. Sooner [HH03, KA97]. Sort [Ibr07]. Source [Nue10, Wak91, vD89, Lei23]. Sources [Ano95e, CW96, Duf96, Gui94, Hal82b, HLT00, LM99, Var80, XS03]. Space [Ary72, BMN09, BCJ04, BC67, Ber68, Bèr04, Bla84, CD07, FS71, Gil68, Gra00, Gre74, Hal84, JR95, Lin73, LS02b, Mor73e, Mul68a, Mul68b, Sau76a, Sie78b, Ski68, Tru00, Tur06, Twe74, DS21, GPW22, Neu23, RS15, Sim19, SRN21]. Space-Filling [Mul68b]. Space-Ship [BC67, FS71]. Space-Time [Bèr04, Tur06, SRN21]. Spaced [Jon86, NW70, BS16a]. Spaces [Ail66, AFG11, BS82a, DS06, Dob77, Fra06, GZ06, HZ92, Le94, Sta82, Zäh90, ZPR08, LS19, SST18]. Spacing [Deh85]. Spacings [Ahs78b, GNRW94, Huf88a, Hüs88, McC85, Wei69b, BMRRS16, Hen18]. Spanning [SSE87]. Spares [FH91, MDG09, Pin80, SS97, SM82]. Sparre [ABI14, Iva16, PP07]. Sparse [KL03b, Kur22, LN19]. Sparseness [SKF79]. Spatial [AD97, Ano81f, BFS09, Bri86, Dan77, Dig76, Fad83, HP98a, Ken90, Kin77, KN00, Kuu82, KZ84, Lan05, Lot84, Mil92, Pen96b, PM99, Ren80, Ren81, Rou90, CS23, CP17]. Spatially [FS80a, BW19, NNY20]. spatio [Hös19]. spatio-temporal [Hös19]. spatiotemporal [LRM23]. SPDEs [Fat15]. Special [Ano72g, BK87b, Hel71, Tam71]. Specialized [Kij92a]. Speciation [Ald95]. Species [AI89, Bra82, CK05, CD85, DFSD01, Hin81, RR78, Rug14, SS83b]. Species-Area [Bra82]. Specific [Ano83f, Ham88, Nat82, Nat90]. Specified [Kwe75, BS13a, BS15b]. Spectra [C.12, Kei78, ST02]. Spectral [ALS04, Ano91i, BH98, Cra73b, Dah85, DL96, HT71, HR88a, Kij87, Lin04, Mei99, Mit04b, Par86, Pen96a, Rey68, Roy72, Tan80b, Tho86, Ton73, Wüb12, HL16b]. Spectrally [Bau09, BV11, CP11, KLP12, KO12, Loe09, Rog00, Sur08, Zho07, BS23, Fro15, LYZ17, LVZ19, NPYY18, Vid19, WZ18]. Spectrum [CL97, GL67, Law71b, MS04b]. Speculations [Con75]. Speculative [Git74]. Speed [JW90, KS04, Kol06, Kol07, RV00]. speedup [Ger23]. Spends [Nor97]. Spent [GR01, KW03, Sch88a, Ser90, WW06a, Ren14]. Sphere [Ors87b, RB91, SS86, Wat82a, Gaá20]. Spheres [ESE84, Gir82, Hal84, Mor73e, Sil78, Sta82, KR18]. Spherical [Fra07, Zäh90]. Spherically [Sch80a]. Spidemics [KZ84]. Spin [How00b, Pir80]. Spins [KKP10]. Spitzer [BB80, Don80b, Sch86, SD76]. Spline [AK86]. Split [AL12, CG97b]. Splitting [Cow04, LR09, LT13, Llo89, FHIM14, LP23]. Splittings [MP93]. Spot [Dur08b]. Spread [AB98, CM82, CDD10, Die66, ELTW10, FS80a, FPKMP98, Gri72, Kat02, Rad73b, Rad76a, Wan80, Wol88]. spreading [Mah20]. Spreads [CLS09]. Square [GD74, Gol77, KR97, McD80, Mor73b, Sat77, Sud04, WG93, Wie84, Gau22, Gor14]. Square-Law [Gol77]. square-root [Gor14]. squared [He21]. Squares [Hea82, HW83, Sol84, dRdC18]. Srivastava [Pyc21]. St. [Gut10]. Stabilité [CTKR83]. Stabilities [Jag92]. Stability [AMR01, Ano90g, AS09, BF95, Bal94, BM89b, Boy69, Che02, CW86, CH90, Dow98, Els07, GW17, HRZ07, Hil67, Hin80a, Hin80b, Hin82b, HB84, JT01, KR86, Kar76b, KK96b, KW96, KW89, Lun96, MM16, MS05a, Mit03, Moy08, NMG00, PW16, RT73, Saw83, Sha88, Sig89, TH72, Vas81, Vas84, Voo87, Voo89, WB03, Whi80c, KS19, CTKR83]. Stabilization [Les06, SDH16]. Stable [Aba80, Ale01, BK12a, BJ79, BC75, CC06, CH84a, CD85, Dal69b, DJ69, DS98, DR94a, DN04, DS04, EMR71, Ebr93, Eva71, Gra12, Hai88, Hai90, Hey69, Hin81, Hin82a, HB83a, HB83b, HB84, Hin95, Kim70b, KMRV00, KKRT12, LH81, Lem74, Li12, Luk69, Ner84, Ott79, Shi06, SM94b, Tay79, Wan76c, ZZ09, Bai20, BY22, BR18, CCdOS15, DV20, DDS20, FZ17, Gra18, HJL21, LW19, Oes18, QDZ21, RS15, SS17, Wat23, ZTML21]. stacked [FL20]. Stage [AI94, BL09b, Fad74, LC84, RSY90, SW05, SG93, Zuc77, BL15, CdOR16, Fox15, Fox16]. Stages [AN84, GM91, Sap87]. Staircase [BHPZ06]. Stalling [Rub85b]. Stamatis [KH96]. Stand [Par94]. Stand-by [Par94]. Standard [Hen83a, Möh05, Wat80b, Bon13, GR21, MWW19]. standardization [SS17]. Standby [KH83, MA69, MMD11, OA70, Raa80, Yeh95, FHC18]. star [FNP19]. Starr [Hal82b, Var80]. Started [PT11]. Starting [Sun92]. Starting-Stopping [Sun92]. State [Ano74h, Ano77f, Ano05e, AGC97, BO99, Bax81, BBK08, BFM08, Bla84, Bra86, BCMW83, CSZ97, Cao95, CD07, CZ98, Con71, Con74, Con75, CL93, Dek87, FT78, FWL03, GOP15, GC71a, GKPS06, GW94b, Gre74, GHL71, GZ02, Had74, Hel70, HST75, Hsi76, HY94, JR95, KP89, Kas66, KN90, Kem92, KT01b, KT01a, KMS83, KP83, KP08, LL82, LR80, LF08, Li06a, Lon09, LS02b, MT95, MF01, Mer80, Mil81, MY84, MY85, Miy79, Nat74, NS84, NRS85a, Nor97, Ort08, Ped71, PH86, Raj74, RCLW10, RCC12, Rod95a, Sau76a, Sch88a, SS85a, STC93, SZ06, Tru00, TV88, Twe74, Van81, VP92, Wan92, XY09, Yeh03, AA14, AA17, BVW21, BP16, BN15, DO16, EOV20, FFK19, GKS14, HL16a, Jan21, LS19, LP21, LY20, MZ24, Neu23, PSZ13]. state [Vid19, ZC14]. State-Dependent [Ano74h, Ano05e, CSZ97, CZ98, Con74, Con75, GHL71, KP89, KN90, KP83, LF08, Lon09, MF01, Mer80, MY84, MY85, PH86, SS85a, STC93, SZ06, GOP15, BN15, GKS14, PSZ13, ZC14]. state-dependent-killed [Vid19]. State-Space [Gre74]. States [FH10, Kal89, Kao74, Kim70a, Pic79, SVJ66, Ser90, Tay71, Wan77b, Wyn73, Yap09, Mar16a, WZ13]. Static [Käm87, DHKR19, OC22]. Stationarity [AT98, CSV01, DR00, GP04a, Kel97, KLM04, LS92b, PM03, SP99, Sta99, Szc97]. Stationary [AB00, AA90, AT91a, AT91b, Ano70e, Ano77g, Ano82g, Ano93f, Ano94f, Ano98f, Ano06g, AK97, Avn79, BS02, BP10, BNCPW09, BM97, Ber80, Ber70, Ber77c, Ber85, BR87b, Bha90, BGH71, BP99b, BVJ00, BL12, BPS23, Bre81, Bri73, Bui75, CW81, Cam97, CA98, Che70, CDL88, CP03, Cos90, DS08a, Dal82, Dan69, DS65, DS67, Dav73, Dav86, De 77, DP12b, DM10b, DG93b, DS04, Dur08a, Eag75, Fas90, FT80, Fei94, FL91, FS06, FH72, Fol86, FK89, FKK91, GA93, GG74, GSV11, Gha90, GY95b, GG77b, HN71, Has82a, Has82b, Hen72, Hey64, Hey88, Hey82b, Hey91, HH96, Hop75, HR88b, Hor01, HS10, Hüs86a, Hüs93, Ito78, Jen66, JS98, KH83, KS94b, Kel09, KS91a, Kij92b, Kim94a, KM93, KS80a, KS80b, KMS83, KS83a, KS85]. Stationary [KW89, KR97, KS01c, Kyp71, Kyp72a, Kyp72b, LR80, Law70a, Law70b, Law79, LRS94, Lel05, Lem74, Lie77, LM99, Lit92b, Lit94, LS00, Llo71, LZK06b, Ma04, Mar86, MS94, MV95, Mas82, Mat92a, MKR91, MS93, Mej65, Miy77, MW90b, Miy94, Miy01, Moo73, Mor75, Mor71b, Mor71c, Moy10, Neu81, Ols67, Ort08, PM99, Pfe82a, PS86a, Pou92, RL88, Rao66, RR75b, RWN02, Rip76, Rod95a, Ros76a, Ros86, Ros72, Ros78, RC80, Rud75, Ryc87, Sch91a, SVJ66, ST85, Ser72b, Ser80, Sgi99, Sha88, SS76, SS96, Sim98, Sim00, SC82, Sob67, SR70a, SV71, SP92, SG88, Tan79c, Tav80b, Tho85, TP74, Ton73, Ton74, TSCD08, TW83, Twe74, Twe83, VAK92, Wat76, Weg76, WC95, Whi85a, Wit93, XZW09, YZ85, Yan73, Zaj95, Zie87]. stationary [Bai20, BY14, BS13b, BR18, BDG14, CCdOS15, CHM^{+}22, FCP20, HZ22a, HL13a, KT15, Mar16a, Mar21, MB16, SK19, Wir18, Yam22, Byt81]. Stationary-Excess [Whi85a]. Stationary-Increment [Ano06g, FS06]. Stations [Sig88]. Statistic [BK06, Cre77, DH85, Fei79a, Fu01, Ksi89, Lib85, Pur82, Wic79, KL17, YML17]. Statistical [Agu03, BCF09, BEN98c, BSB99, Fin64, Gil67, MCC03, Phi01, Rah17, RW64, Rus01, SS85b, Wes82, XS03, ZRU68, ZS92, FHZ15, Let94]. Statistics [Ahs76, Ano67d, Ano75e, Ano85g, Ano86i, BR89, Ban88, BLR91, Bar86, BH01, BT92, Ber65, BE83, BP96a, BENP94, BS79c, Cam73, Dal87, DG74, DC06, DC10, Deh84, DH08, DMT11, Dur86, Dzi84, Dzi86, Ery08, Ery10, FWL12, Gal72, Gal74, GW85, Gel77, Gup84, Hen81, HJ11, Hu96, Hua74a, Joy98, Kad74, KP96, KX11, Loy88, Ma97, MC11, Mei81, Mil76, NS06, Pol75, RR10, Rob90, Set03, Sri86, Tak64c, Tan86, VJ01, Wak88, AZA18, BHTLV20, BN13, FPY20, Gau22, Hoo13, JR16, KK16, KK18, LWM20, NS22, Ott20, Owa23, PT23a, WGF13]. statistiques [Let94]. Stay [Igl74, Kao78, Rus78]. Staying [LS02b]. STD [Yan90]. Steady [AGC97, DO16, GW94b, Kem92, LY20, Miy79, NS84, NRS85a, Sch88a, Tai95, BVW21]. Steady-Size [Tai95]. Steady-State [GW94b, Kem92, NS84, NRS85a, Sch88a, DO16, LY20]. Steck [Bre76]. Steele [Rog23]. stein [Psa22, Bar88a, BT23, Dal10, GR96, GR13, Kim00, MS23, Pek96, Ros16, Vel04, Wei00, Wir18, Xia99]. stein-type [Psa22]. Steiner [Mat78, RS89a]. Stemma [NH82]. Step [Abd87, AH77, Ano91f, BM82, GC71b, Gue91, Hai92, Kab83, LR09, Rap08, SCZ77, Yas88, ARS21, BJMK18]. Stepping [Mar69, Ren74, Ren77, Sat83, WK65]. Stepping-Stone [Ren77, Sat83]. Stereological [JG85, JM86, Mas82, Møl88]. Stereology [Dav80]. Stevenson [Gan02]. Stick [Hol80b]. Sticky [BPS15, HL81b, JSW19, SW20]. Stieltjes [KY17a, Sto04]. Stigum [HWX18]. Still [Jac75]. Stimuli [Cam86]. Stochastic [Aas87, Aba79, AT74, AR05b, AMV21, AM07, ASZB18, ASB21, And70, AD98, And03, Ano68d, Ano02e, AK86, AS09, AMW88, Bac91, BBH14, BdHK93, BO99, BOP07, BLR91, BNGS04, Bar88b, BV11, BK86b, BK87a, Bec70, BOR01, Ben98a, Ben98b, Ben70, Ber73, BW98, Bha65b, BE91, Bil73, Bil75, Bil76, Bin69, Bit69, BMP87, BK64, Blo84, BPT92, BSC02, BSC04, Boo79, Bra70, BL09b, BM66, Bro70, Bro88, BS73, BR77b, Bro00a, BCD93, CW97, Can77, CJ08a, Cer88, CM07, CL11, CF11b, CF14, Cha74, CD13, Che98, CK05, CH97, Cla05, CS95, CC94, CCFR00, CPT99, Dar89, Dav82a, DM01, DZ03, DF89, DC06, DC10, DFP13, Dig76, Dip98, Dob78, DL06, Dos85, Dow72a, Dow72b, Dsh89, DH08, DR81, Dur92, DK99, Ebr99, Ebr01]. Stochastic [Eco00, Eco03, Els96, Els07, EE76, Epl83, FE87, FI87, FHT78, Fen95, FS98, FA85, FL97, Gag17, Gal71b, Gan65, GJ72, GK76, GP04a, GK97, GS87, Git72, GG77a, GY95b, Gla81, Gla80, Gla90, Gla91, Gla93, GMSB10, GBR75, Gra71, Gra72, GL83, Gre81, GHL71, Grü89, GdlI18, GG77b, Gue81, Gui02, GMRY07, GG10, Gup13, Hel73, HG80, Hey91, Hey95, HR87, Hor01, Hos09, HCY21, Hu96, Irl03, IM74, IS83, Jac99, Kal91a, Käm87, Kar68b, Kar74, Kat77a, KW92, KS94b, KW96, Kel97, KW99, Kel76, Ken91, KK00, Kin02, Kol15, Kol95, KS80b, KL00, Kry72, Kry75, KL89a, KS91b, Lam68, Lan79a, LH81, LS98, LL78, LWB08, LW93, LT95, LX00, Li00a, LWM20, LS09b, Lud73, Luo14, Ma97, MP92a, Man77, McC78, McQ67, Mil79]. Stochastic [Miu77, Miy01, MRSA21, Mod66, Mod67b, MJP87, Møl95, MH76b, Mul72, Mul73, Mye81, NTW89, Nak98, Nei01, Ors87a, Ors87b, Ova01, Oza86, PT71, Pak10, PB83, PS02a, PC85, PP73b, Pri00, Pur71, RT90, Raj81, RR67, Rie80, Rig97, Rig11, Roe70c, RV08, Roo76b, RE01, RS00, Rüs06, Sah71a, SA72, San80, Sau75a, Sau76c, Sca98, SW05, Sch75b, Sch76, Sec97, Sel83, Sev69, SW97a, SW97b, Sha87, SY89, SK90, SY91, SS96, Sob73, SP17, SR70b, SM72, Sri91, Sri67, Sta99, Sta01, Ste81, Sud12, Swi08, SW88, TR67, Ton73, Ton74, TK15, Tru00, TV93, TH72, Van86, Van91a, VPT92, WD99, Wan77a, Web82, Wei75b, Whi80b, Whi82a, Won97, WH04, XMKW90, Yah75, YC82, Yin21, Yu74, Yu09, Zac07, Zen10, dSS04, ASKK20]. stochastic [BVW21, BSC18, BJ14, BKJ16, BBH18, BY14, BHB18, BBBN14, BDG14, BL15, Cha13, CC15, CF17, CdOR16, DHKR19, DNY16, ESC15, FX16, FFK19, FFK20, FCKS16, FPS18, Jan21, JL22, KB13, KLL15, LMS18, Lee20, LHT17, MM16, MWH22, McK14, MST13, MBM21, NNY20, PLRB23, Pim14, Rek23, RNR18, Sha13, Sim19, Sön23, SB17, Wan15, WBS16]. Stochastically [ACK20, Coh81b, Coh83, DW77, FP66, Hol78, KK03, LR95, PS92, PS86a, Qiu91, Rig90, RT00, WVW86, KM23]. Stock [Ben70, GM73, Pau01]. Stokes [He01]. Stone [Mar69, Ren74, Ren77, Sat83, WK65]. ’Stop [All12, Gou98, Tam10, Van91b, MA17]. Stoppable [Gla79b]. Stopped [Bad11, Fer83, IS85, LU95, Pak08, Sam67, ST83, BS15a, BAR14, DLVZ20, LT13, Leh21]. Stopping [AH77, All04a, All04b, ASC98, AGSC00, BHPZ06, BL03, CS85a, Chr12, Cla99, CS97, DLU09, DP10, EW09, ES87, FH89, Föl72, FB88, Gap07, GS14, Gav69, Gil87, Gne07, GI11, GP97, GP98a, Guo01, GS01, GL05, Hen73b, HH85a, HK90, IKP01, Irl06, KS77, Kar09, Ken82b, KR00, LR00, LRS07, Mam87, Maz86, Mil81, Mon82, Muc02, MR01, Nak91, Ped00, Rot95, Rot98, Sch98, ST12, Sie85, Sta87b, Sta74, Sun92, Tam85, Tam88, Tam01, Var80, Vil07, War76, Whi73, YWR82, Yas83, Yas88, Zac04, Zuc78, Zuc86, AM14, AS17, BAR14, Buo15, CdSJ20, Coh14, DL18, DR14, GR14, GA22, HZ20, KOH23, Kli14, LOR19, MA14, MV21, Tam15]. Storage [Aal98, Bal79, Ban07, BG02, Bro72, BS92b, BK73b, De 85b, Gly89, GM94, HL81b, Has82a, Her74, HK85, KP89, KW92, KW96, Kel97, Ken83, KHY03, Lun96, Ōsa85, Pra64, Pra68, PW81, TR67, Woo83b, Wu68, HPP14, KB13, PSZ13]. Storage/Production [KP89]. Stores [DMO05]. Strahler [Yam17]. Straight [ESE84]. Straight-Line [ESE84]. Strains [MD96]. Strands [Lit74]. Strategic [AIGR06]. Strategies [Aba80, AI89, BS92a, BBK08, Ber80, CL09, CH84a, CD85, DGK00, DC77, Dob77, Fas90, FW91, Gla82b, Gla90, Hai88, Hai90, Hin81, Hin82a, HB83a, KLP12, SZ91, Sub79, Tay79, Vil07, Zuc80a, AS17, FPS18, FVZ19, FPP19, NPYY18, Hin95]. Strategy [Ano87e, CSZ04, Got85, GCS07, HR00, Hin80a, Hin80b, Hin82b, HB83b, HB84, JK87, KS05, Ros71, RZ08, YYZ09, Yao07a, EG15]. Stratified [Wat72]. Strauss [SF77]. Stream [Abr97, AAL77, Kei82, NS05, SDH94]. Streams [Ana88, AK93, Ish04, Mej65, NP91]. Strength [Ano87f, DS08a, Ebr85, HST83, KP87, Sen73a, SBG70]. Stress [Ano87f, BH01, BVJ00, Ebr85, Tsa99, CF14, LLT^{+}22, PKO16]. stress-release [LLT^{+}22]. Stress-Strength [Ano87f, Ebr85]. Strict [LS92b]. Strictly [KMRV00, PL73]. Strictly-Periodic [PL73]. Strike [HW02]. String [BC01a, Cai03, Bra14]. Stringent [Kwe75]. Strings [Ale96, BL03, Hol07, HS12, Mak10, Ste03]. strip [BH20]. Strong [ADY07, And84b, And85, Ano91g, BJT20, BT76, BDM09, Bre81, Cha74, CL23, Che81b, DS86a, DS80, DW85, Dzi84, FPR20, Ger91, Gou97, HGR90, Hey99, IT78, IA78, JI88, JW11, Lan79a, LT81, Mao02, Mat92a, McC85, Not00, O’N96, OC22, Pir80, SAI82, SI83, SY91, Ste78, SZ93, Twe81, Via94, YC08, Zaj95, Zha01, BCS23, Dio76, Dio78, EL19, Kan23, MW21, NP22, Pen16, CCN10, Hey75]. Strong-Mixing [Dzi84, Zaj95]. Strongly [BGJ12, Kol82, OS91, THP12, Tas97, BL21, KL22, Nat14]. Structural [DWW11, KW96]. Structure [Abd87, AAO90, Ano91i, BLBL10, BK86b, BK87a, BL89, BK03, BN92, Cri18, Don84b, DKK08, ER88, ER89, ESM08, God77, He03, Kap10, Kei82, Kel09, Kij87, KB87, Law71a, Lee80, LX00, MM03a, McK81, McK88, McV08, PM99, Pfe82b, QS99, RCC12, Ren88, RZ08, Sum84, Whi01, FS14]. Structure-preserving [Cri18]. Structured [BOP07, GH90, Not00, Not01, Sam06, ZLA99, JL22, KSN16, LLL20]. Structures [Ano86h, Ano91f, Bax84, BK00, CJ06, Dav75, Ebr87, Gue91, Hai83, Hai92, Hen86a, Hop86b, Joy98, Kou96, KZ09, MM82, MR01, SN97, Tsa94, Tsa96, VK00, AR13, CM15c]. Student [Ano06g, Dur86, FS06, ST16]. Studied [Hol68]. Studies [GL94, HB99, YI99]. Study [Ano64c, Ano72g, Ary72, BRY04, CG99, Can73, CM07, DS86b, HI01, Hol66, Kei64a, Kei78, Kou96, KP07, RB91, SS86, Tam71, JF23]. Studying [Pur71]. Sturm [CCGW23]. Sub [Bar74, Bil76, LY20]. Sub-Critical [Bar74]. sub-Halfin [LY20]. Sub-Populations [Bil76]. Subadditive [AMR01]. Subarcs [Man77]. Subcritical [Ath12, Bag82, CY84, Dek87, Don76a, FG99, Hop76, HMY17, LBC81, PK74, RVJ68, Sau76a, Voi00, Wil88a, Hon13, Hon16, Win15, dRdC18]. Subdivided [Gla77, Pol66, Pol68]. Subexponential [APS04, BFL05, BMZ07, FKZ09, Gol78a, JL98, Klü88, KT08, Les89, LR07, Mur87, Mur89, NT04, WCW05]. Subfair [Che76a, CCHY08]. Subgeometric [HLZ05, MS21, AFV15]. Subgraph [RR10, Kur22]. Subject [AH84, AH86, Ano88g, Ano89e, Du95, Eco04, Fak83, KKY02, LL93, LM04a, Llo71, Mer80, OL65, PL81, RG88, Wu68, YZ85, YY11]. Subjected [LM90, Tra81]. Sublinear [Hör03]. Submartingales [Hal69]. Submultiplicative [Sgi99, Sgi01]. Suboptimal [Gla82b, KS05]. Subordinated [HK09, BL21, WW20]. subordination [BL21]. subordinator [DMZ15]. Subordinators [Cov09, Gzy77, Lag05, Win05, HJL21]. Subpopulation [CG02]. Subsequence [Gne99]. Subsequences [ACSS11, CS75, HK19]. Subset [AR01a, Ser90]. Subsets [Bla92, LS02b, McD80]. Substability [Nak86c]. Substantially [JS04a]. Substitution [Wat82b]. subtraces [BR23]. subtrees [CMW16]. Success [Ano77f, HY02, Pit80, Raj74, Web99, BKJ16, KLY16]. Successes [BP00, Hol08b, Hun76b, Rig90, WJ95, WCC03]. Succession [KN00]. Successive [Arn75, BR87a, Lam68, Ryc87, San80, KSS16]. successors [HJ13a]. Such [Bar75a]. Sufficiency [Bof65]. Sufficient [CCFG96, CD05, CW86, Ger76, Gla82a, Hav91, IMS98, Joy98, Kot94, KS01c, MSCZ78, Szp85, Was69, Wic79, BMRRS16]. Suitable [BL73]. Suits [Van06]. Sum [All97, CT04, DN04, DJ95, Gan82, GHL03, GHL05, HH96, HM99, Jac75, LS92a, Mar73b, MQ00, Oht91, Pon07, Rap08, RT91, Rüs06, Sta87b, Sta04, Tam10, CL14, CF17, Wan14]. Summability [BS04b]. Summands [DN04]. Summary [HJ11]. Summation [KKRT12]. Sums [AT91a, AT91b, AG80, Ano75e, Ano94f, Ano06e, BMZ07, BE87, BT92, BC73, BPT92, Bon02, CK00a, CY65, CV09, Coo98, Eme72, FP97, GN06, GJ88, GI11, Gol75, Gut03, Hel70, Hey69, Iks04, KM97, KT08, KM92, LU95, LU96, LS02a, MP92a, MO01, MWW19, MS93, NT04, NTaYY04, O’B74, PP69, Pha88, Pol75, Pol78, RP03, Sam67, Sch82, SD76, Tak77, Tro83, VC96, VU09, WW07, Wei69a, Wol77b, Woo82, ADJ13, BAR14, Jan15, KS17, NR20, PT23a, PR13]. Super [CPT99, DF04, GHW95, Hey70, Hey71, HL99, Hon02, KMSP12, Rad72a, RR94, ZG04, Zha04, Zhu23]. Super-2 [GHW95]. Super-Brownian [DF04, HL99, Hon02, KMSP12, Zha04, Zhu23]. Super-Critical [Hey70, Hey71, Rad72a]. Super-Extremal [RR94]. Super-Heavy-Tailed [ZG04]. Super-Replication [CPT99]. Superadditive [DHT86, GM96, GM97, Hul82]. Supercritical [APS74, Bag86b, BFM08, Bin88, Coh77b, Coh82, Coh89b, Dal79b, Don72, GR87, GD74, Gre78, Ham92, Ham95, Hau12, IM10, Jon04, Kap75b, Kuc82, KMSP12, LML06, Pak74, Pak76, PH82, SLK79, Tan93, YCG11, ZJ04, Bow18, Bra22, Chu18, FS23, GZ16, IK16, PC16, Tzi16, Zhu23]. superdiffusions [LRS09]. Superdiffusive [FH07]. Superexponential [Zei98]. Superimposed [BGH71, Dan69]. Superinfection [MD96]. Superitical [Nan79]. Superlinear [Smy11]. supermodular [DFP13]. supermodularity [KM23]. Superposition [Amb65, Bli89, CL85, Çin68, HLT00, Ito77, Ito78, Lam91, NP91, PS01a, Whi72c, Wil97, CM15b]. Superprocesses [AI00, Gla82a, Hon04, Wan97, FFK20, JLX22]. Superprophet [Wit96]. Superreplication [EJT05, OO07]. Supersession [OT77]. Superstrings [Ale96]. Supplement [Hey75]. Supplemented [MY88]. Supplements [Hey77]. Suppliers [Rig99a]. Supply [Pla83, Wei05, SDH16]. supply-local [SDH16]. Support [BS98, GSV11, Hai89, Pak10, PD20]. supports [Pit18]. Suprathreshold [Tho11]. Supremum [All12, CW81, Cha75, CFS11, CD81, Hey69, Hol00, JP09, KO12, PZ07, PP73a, RR75b, WW06b, Zho07, BDM21, Bis23, BSDG20, CPN18, RNJ16]. Sure [Bag86a, Bèr04, BCF09, GD74, Gre77b, IM10, Kuc82, LK77, Mei10, Pur87, RHK90, RT73, dHR73, FPP19, HMBL14]. Surface [Car67, EE81]. Surplus [Fro05, PP07, ZW02]. Surprising [Con22]. Survey [Gre86, LSC04, ZRU68]. Survivability [AAKR05]. Survival [BS78, CS12, EP95, GE82, GS85, GG10, Hey82a, Hou91, Kle81, Mi94, Mye81, Pol66, Pol71, VK10, YA88, AES19, ES23, JP16, VD17]. Surviving [Ano04g, Ste04]. Survivors [SLK79]. Survivorship [Bre69]. Susceptibilities [Nea07]. Susceptibles [DD68, LL82]. Sustained [CM83]. swap [SP22]. Swarm [MZ01]. Swing [RZ08]. Swiss [Bré93, Kon96]. Switch [Leh72, Van99, HLVM22]. Switch-Over [Van99]. Switched [Zie87]. Switches [HW96]. Switching [BLL84, DT01, Gla80, Guo01, LPS06, Sha89a, VPT92, XY09, Yan90, EOV20, GKP20, Jia15, Kun22, LS20, Mar16b, ZTML21]. Sylvester [DS86b]. Symmetric [AIH87, BE83, CJ08b, Di 98, Don80b, Gor75, JT92, Kar83, KZB05, ML86, Ros07, Sch80a, Shi06, Ver98, WR74, BR18, DFP13, MA17, PP14, Pro18]. Symmetrical [Mar69]. Symmetries [BF09, KADK19]. Symmetrized [BGT04]. Symmetry [DP12b, DGNR95, GNR89, Spe86, Kol15]. Symmetry-Based [DGNR95, GNR89]. Symptotic [LL82]. Synchronization [DLM14, Sah16]. Synchronized [KB20, LX00]. Synchronizing [YY11]. System [AH86, Abr01a, Agu03, AMR01, AN84, And75, Ano83f, Ano87f, Ano88g, Ano90h, AF83, Ave09, AB86b, BT73, Bax81, Bla84, Boe69, BP01, Bur76, CL11, CD95, CN80, CP90, Con71, CW86, Cra73a, Cra74b, CS88, DF89, DT85, De 85b, Die99, DK90, Dos81, DN99, DS04, Ebr85, Ebr01, Enn69, Eth04, Fak82, Fak83, Fal96, FH90, FH91, FT80, Fel78, Fin69, Fra07, Gal71b, Gas64, GL76, GNN85, GAS03, GP98b, GMDB03, Had69, HA96, Hei01, HF70, HS93, IMS98, Jin88, JW90, Kab83, KH83, Kas66, KM93, KT01a, KM88, LC84, Lan86b, LL93, LM89, Lew01, LAF02, LM04a, LS86, LKT90, LMW00, LT71, LT72, MR03, Mi95, Mi99, MA69, MDG09, Miy92, MJP87, Nak79b, Nak79a, Nak84, NE81, NE82, Nat82, Nat90, New86, Nor86]. System [OKM86, OA70, PH88, PPF03, Pea65, PS99, Pha74, Pha75, Pie07, Pin80, Pli75, Pou90, Pre97b, Raa80, Ram81, Rev93, Sah71a, Sen87, Ses96, ST73, Sha75, Sha89b, Sim00, SS97, Sko00, Sob90, SK91, SG93, Sta85, SP92, Tam70, Tin85, Tsa94, Tsa96, Tsa99, TV93, VOZ95, Van89a, VS90, Vis86, WW06a, Wie89, Wol87, WR97, Xu98, Yan90, Yeh03, Yeh05, ZY70, Zha94, Zha05, BV14, DD15, FZL13, FHC18, FL19, FS17, GXY18, LL24, Mar15, MR15, Ros14b, RNR18, YCW15, YBL23]. System-Point [AB86b]. Systematic [CO85]. Systems [AT74, AMAKAH97, Ano87e, Ano91f, Ano91g, Ano93g, Bax83, Bax98, Bel66, BMPR07, Ben98a, Bes72, Bha65b, Bla87, BR96, BS82b, BLS03, BPT89, BEN98c, Bol01, BI05, BG87b, Bur09, CM92, Car99, CDD91, CC79, Coh67a, CMY98, Con71, CVW84, Cra73a, Cra74b, CX01, Dar89, DJ05, Daw73, Daw74, Dob77, Dsh89, Dsh98, Dsh01a, Ebr90, Ebr91, ENPS78, ENPS86, FI98, Fen93, Fen97, FS80b, Gås98, GA99, GR98, Ger90, Ger91, GSS11, GAB12, GLS00, Gra08, Gre89, Gue91, GP04b, Hal79a, Hal83a, Ham82, HST83, HP98b, HMM93, Her74, Hil67, Hor04, HNG93, JNR09, JK87, KK96a, Kar76b, Kat11, Kat12, Kei78, KS86b, KK00, KFM06, KKY02, Kij89, KT01b, KX09, KS80a, KS80b, KB91, KL00, KST99, KTW69, KSF80, Kul83, KW85]. Systems [Kwe75, LS98, LMM10, LWB08, LC83, LDL95, LP00, LSC97, LX00, LZ02, LD10, LA01, LMW98, LT90, LT04b, LR95, MT95, MP70, Men96, Mil76, MDG09, MMD11, Mou93, Mus00, Nak86c, NS84, NS06, NE07, NBS08, NSB10, NSB11, Naw83, Now90, OV03, Orr69, Oza86, Pak70, PS84, PR95, PRX03, PT97, Pir80, PS86b, PT94, PH86, Qiu91, Råd76c, RG88, RSY90, Sen01, SH03, SM94a, SC82, STC93, SZ91, SV03b, SG88, TTH91, Tay71, TR67, Tom72, TV88, Tso71, Tuf99, Van99, Van90, VS00, Vas81, Vas82a, Vas84, VG85, VT89, VGT90, VP92, Wil97, WK94, Yeh91, Yeh95, Zha10, ZLB08, Zhe91, AKLL18, ASZB18, ASKK20, AZA18, BBH14, BBH18, BHB18, BPS15, BB18, BN13, Ery15, ET21, GN17, Gup13, HF19, HL23]. systems [Hoo13, HLVM22, JTW^{+}13, KZN15, Kan23, KA22, LL24, LSW19, MZ20, Nat14, NFS20, OT16, PB14, PSZ13, PB20, Tav20, TSN17, TK15, YZZ21, YBL23]. Sznitman [FSJ14].

Tableaux [DMO05]. Taboo [Wau72]. Tagged [FF96, Gir86]. Tagged-Molecule [Gir86]. Tagging [SC75]. Tail [Ber92, Bro64, CG99, CL05, CJ06, DPW21, DG93b, GR24, GT06, GN06, GW94b, Grü85, Ham95, HST83, KKK07, KHY03, Koo04, Lel07, LS09a, LPN13, Lin96, Mun12, NT04, Nie95, PZ07, Res71b, Res04, RNJ16, SW87, SRN21, VP10, FPW19, HJL14, JS14, JTW^{+}13, KS17]. Tailed [CT04, Che03a, DT11, FKZ09, Gay04, GN06, HT12, KM97, Kul06, LPS06, LS09a, Lit94, McC97, TV04, WW06b, ZG04, ACVZ20, BvdHvL17, DPW21, GH14, KS21, Leh17]. Tails [AT96, AK97, BFL05, BPG20, BV96, CW81, Gol78a, Hit10, HS10, Klü88, KT08, LS08a, NTaYY04, Pak75, Pak08, Tan07, WW07, WL94, WL97, AB19, AW21, Bri20, BDIM18, DL17, JTW^{+}13, Leh21, Zha22]. Takác [Sha74b]. Takács [Bin94, DY98, GG94b]. Takácz [Ano94m]. Taken [Sud83]. takes [DvdHS24]. Taking [Bin69, DDM98]. Tandem [AIH93, AI94, BJM93, Cao84, HP95, Hil67, HLT00, KM90, Leh86, Nai71, Niu80, Niu81, SW05, Sig88, SZ06, SM97, Wol84, CHR15, Lóp16, Sez18]. Tanemura [Wei79]. Tangential [Mol86]. Tankov [BJV12]. Tapered [KS01a]. Target [ASB94, Cor80, Cor81, Sar73, BJ14]. target-dependent [BJ14]. Targets [YZ85]. Tasks [Gla79a, GADL04, Pra82, WP80]. Tax [ARZ08, KZ09, Whi05, ALL21, LTZ13]. Taxonomy [Dob74]. Taylor [AL00, BCdlP17, DDS22a]. TCP [LV08, OS07]. Teaching [Rus01]. Team [WR74]. Teamball [CYF96]. Teams [GKPS06]. Technique [Bri79, Rou96, VT11, War99, FCvS15]. Techniques [Dur71a, MNS01, Tru00]. Telecom [LN23]. Telegraph [BDMZ12, DM10b, DO97, LR12, Ors86, Ren88, SZ04, Zac04, DIMZ13, DM18, LR14b, Ver14]. Telephone [Erl67, Gas64, PT71]. Teletraffic [Lee80]. Temperature [How00b]. Temperatures [Bar74]. Tempered [BK12a, Gra12, MSS09, ZZ09, Gra18, QDZ21]. Temporal [BC95a, Nor77, Seg06, Hös19, LWW18, Vid20]. Temporally [Tuc76]. Temps [PL03, PLC03]. tendency [Szy19b]. Tends [CK00c, Sim97]. Tennis [MRP04]. Tensor [Tsa99]. Term [Ano81e, BL09b, Che80b, Eng80, FN70, Kap10, BL15, Neu23]. Term-Time [BL09b, BL15]. Terminal [FS71, Gol77, Loe09, NH82, San71]. Terminating [CF09, Iva14]. Termination [Dsh94, Bre13]. Terms [Kal91b, MSCZ78, BLSGS13, ISK18]. Territorial [Cow92]. Tertiles [Ahl20]. Tessellation [Ano93e, BK21, MS92]. Tessellations [KM87, Ori79, FPY20, HKT24]. Test [Ave09, Dur71a, Hos86, NSK08, Sen73a, BR18, Hen18]. Testing [Ber68, Ber10, BSC02, BSC04, GESK95, GKL01, GW11, Han82, Pri00, Qui86, Ruk00]. Tests [Pyk94, Yao93]. Text [Bra72]. Textile [Rao66]. th [AAO90, BM20, CHO22, Kol15, Mul73, Ski67, Wan92, Yun00, Hen82, Mul72]. th-Nearest-Neighbour [Hen82]. th-Order [Mul73, Yun00, Mul72]. Their [Att89, BSW07, BCY97, Bor94, Byt81, CT04, Gha90, Gne00, HKY96, JKM12, KR86, Kel09, Kul87, KO12, LM91, LX00, ML86, Miy79, Miy82, Miy90, Miy94, NSK08, New70, Nis74a, PS76b, Ros08, Sap87, Wes82, Yu09, Zha10, BKK19, BB14, Bri20, FHZ15, GLS19, HI20, HM20, Ino23, MO21]. Theorem [AD98, AS96, AKM10, Ano72e, Ano81e, Ano91e, AS75, BPS12, Bal83b, BE87, BB96, BF78, BKMS04, Ben98a, Bèr04, BCF09, BE83, BCPR11, Big77a, BP96b, Bre75, Bré78, BN88, BR95, Cam97, Che03a, Che80b, CP88, CJ98, CCH95, CH71, De 82, Del98, DZ07, Di 99, Don72, Eag75, Eag79, ETV88, Eth79, FS97, Gal74, Gol78b, GR05, Han90, Har73c, Hay00, Hen73a, Hey67, Hey77, Hey95, Hin04, HL99, Hua74b, Hud73, Joh06, JT92, Kal93, Kap72, Kap78, KL03a, Ken75b, Kim70b, Kle85, Lan84, Lan79a, LS08b, Li06b, Li09, LPP12, ML85, MP12, Mei10, Mod67a, Mor73a, Nor75, O’B80, Paw89, Pla83, Pós09, Pur78, Qui78, Rat70, Reh77, Rey70, Rey72c, Rob90, Rob91, Roo76a, Roo76b, Ros76a, RD80, RR73, Sau76b, Saw79, Ser92, Sév95, Sev69]. Theorem [Sgi96, Sho72a, Szá74, Tak64c, VX04, Via94, Wal83a, Wei80a, Win90, WW00, Xu09, Yah75, ZWV10, Ari22, Bow18, DD15, DM19, Fat15, FHI17, IK16, IR20, JOC13, KM15a, KR18, Lam14, Mat74, MA14, MN16, Psa22, RY20, RNJ19, Rub23, Tor69, Yam17, Zhu13, Zhu23, dL23, vD15, FM12, FG12]. Théorème [Mat74, Tor69]. Theorems [Agu09, Amm83, And87, AD97, ANT78b, AK76, Bal79, Bal99, Bar75a, Bas10, Bél92, BK12b, Bon02, Boy69, Boy70, CS77, Che81a, CT72, CY84, Cli83, Col76, DY97, Dur71b, Ell66, Gei99, Gol05, Gre77a, Gut10, Hal78, HL77, HMPS06, Hol79b, Hoo83, Hoo70, Höp82, Hsi97, HZ82, Jan83, Jan01, KS77, KS06, Kim70a, KPW94, KH75, Kul06, KP87, Kur71, Les73, Lin08, LR95, ML86, McL76, MS04a, Mod66, Mod67b, Mor79, MR11, Nak74, O’B74, Pak73a, Pak76, Pol82, Pra70, Råd72b, Raj81, Rey72b, RD80, Sel91, Sev67, Sha74b, Sta82, Tom72, WJ95, WCC03, Web80, Whi71b, Whi72a, Wol77b, WS04, Yam89, BKP^{+}22, BLSGS13, Cri16, DS23, GZ16, HM14, IJB18, LST19, MN13, NS22, SF17, Sah16, SH21, SGS17, Sei16, VL15, WW20, Zhu14]. Theoretic [Ano67d, Ber65, Bur79, Cse91, Joh24]. Theoretical [McN70, Tak71b, Hu16]. theoretically [VI14]. Théorie [PLC03]. Theory [AT06, AS89, Ali70, And70, Ano73g, Ano73f, ADH93, BD90, BW98, Bin94, Bla94, BWZ85, Bré78, Bro64, Bro70, BH75a, BNS90, CC79, Coh67b, Coh73, Con71, CH70, Cse79, Cse91, DG74, Dav65, DV87, DL96, Del78, Del85, Di 99, DY97, Ewe82, Fin64, GG74, GL71, Gre69, Han73a, HS72, Hey73, HS74b, HNF85, Igl69, KKL01, KS86c, Kin02, Köl74, Köl79, Kot66, LMS72, LK77, Lin87, LS97b, Mej68, MO71, MMR07, Mor69, MW91, Naw83, Nay87, NO85, OM72, Orr69, PT71, Phi81, Pic80, Pic84, PLC03, Pol10, PR67, Pra64, Pra68, Pra88, Pur73, Pur69, Råd72c, Rao66, RW64, Rob05, Sch86, Sch68, Sev67, SZ92, SS93, SA73, Smy76, SSV72, Ste78, Tak64a, Wat64, Wes73, Wüb12, AH13, BS14b, BIMR22, CDP19, CCGW23, EOV20, FPY20]. theory [GR18, IRS23, KS19, Kev17, Owa23, Sag20, BS16b]. Therapy [Ber94]. There [Sco82]. Thereof [Jag74a]. Thermal [Has82a, Kal91b]. Theta [Kuz10]. Thick [Ano78g, Dav76, Mat76]. Thinned [Kal75, Yan88a, DS17]. Thinning [CL85, Ish80, NG95, Råd72b]. Third [Gat82b]. Those [Hul82]. Thou [Ibr07]. Thoughts [Hey88]. Three [Ano91f, Bar64, BS82a, Car02, Che91, Coo89, Dal79a, DO97, Dix93, DS86b, GC71a, Gue91, Hin80b, How00b, Hsi76, Imh84, LS97b, PRCS09, AB16, AA14, AA17, Let78, Möh21]. Three-Dimensional [Dal79a, Imh84]. Three-Grade [Ano91f, Gue91]. Three-Parameter [Car02, PRCS09, Möh21]. Three-Player [Bar64]. Three-State [GC71a, Hsi76, AA14, AA17]. Three-Valued [DO97]. Threshold [AD98, Bal83a, Bal83b, Bal84, Bel71, BK03, CPTW85, DT01, Ebr99, HTW09, Hei01, Hyn94, KJ81, KW03, Lin99, ML86, Oza81, PW84, Raj81, Ros07, Vil07, WW06a, AS17, AS16, BFP17, DP18, EG15, HWX18, MZ20, YW21, ZX17]. threshold-one [MZ20]. Thresholds [Ano73e, BSS72, MS10, Rob05, AB16, GA22, Kli14, Tam15]. Throughput [Cao89, Moy08, Sob90, TW89]. Tiago [Tur92]. tick [Mia13]. Tides [Car67]. Tied [LH07]. Tight [CT95, Rhe98]. Tightness [Fan12]. Time [AKK20, AC02, AI89, Ana92, And12, Ano65e, Ano72f, Ano73g, Ano82f, Ano82h, Ano86g, Ano86j, Ano91i, Ano98f, AS12, Arn72, AM05, AG87, AIH88, ALRW99, BFL05, BDM12, Bad11, BMN09, BR00, BY93, Bal72, Bar71, Bar72b, Bar74, BB06, Bat66, Bau09, BKMS04, BC01a, Ben70, Bèr04, BS76, Bha65a, BB80, Bil77a, BMP87, BL11, Bli89, BM94, BP96a, Bri73, BL09b, Bro86, Bro04a, BN88, Bru71, BS09, BGNR90, Cai03, CFZ96, CR72, Car99, CD07, Car67, CS86, CG96, Che91, Che97, CC01, Che11, CDL88, CN80, Çin67b, CS85b, CT81, Coh68, CD06, CCH95, CK81, Dal69b, DJ69, DP00, DS65, DS67, DM67, DM68, Dav85, DGM08, DT80, De 85c, Di 98, DR09, Dom95, Dom96, Don84b, DF04, Dsh94, Dsh01b, DP10]. Time [DO85, Eks04, Ell93, Fal86, FF91, FG99, FY82, FP05b, FS07, FSW12, FY83, FF03, Fro05, FC02, FC03, GG74, GT85, Gao04, GP04a, GT74, Ger90, Ger83, GC71b, GC71a, GL99, GG91b, Gly89, GMSB10, GK12, GL83, Gre76a, Gre74, GR01, GZ02, GHL03, GHL05, GRRT97, Had69, HN71, Hal79a, Hal70, Ham05, HH03, Han71, Han73a, HL76, Har90c, HP95, Has64, HC07, Hen86a, Hey64, Hey86, Hey88, Hey99, HW84b, HNF85, HJ11, Hoo83, HLZ05, Hou12, How01, Hsi76, HB02, HK09, Hut90, HCK02, Hyn94, IK11, IK04, Imh74, Irl03, IA78, Ish74, Ish04, Jac12, JRR95, JP95, Jen72, JGR91, Joe96, Jon86, JS98, Kap10, Kar74, Kat11, Kat12, KP96, Kei71, KR88, KZB05, Ken78, Kij87, Kij88, Kij93, Kij98, Kim00, KT01b, KT01a]. Time [KW03, KLM04, KPT99, KS80b, KMS83, KZ95, KNR05, KA97, Kyp71, Lad05, LC84, Lan86b, Law92, LW93, Lee11b, Lef97, Leh71, Lem87, Lem89, LPSV00, LZ94, Lie77, Lin96, Lin82, Lit94, LA01, LB88, Liu04, Liu11, Llo71, LT71, LT72, LT73, LT04b, Lou84, Ma03a, Ma04, Mac01, MU00, Mao04, MC78, Mar05, MW71, ML98, MS91, MW96, McC76, McC78, MT00, McK82, McN82, MP12, MS04a, Meh77a, Meh77b, Meh79, MW90a, MPHL10, Men23, Min78, Min79, Mit03, Mit04b, Miy77, Miy94, Mor80, MW91, MKN11, MM03b, NSK08, NG95, Nar68, Nat75, Neu81, New68c, New68a, New68b, NR11, NRS85a, NRS85b, NRS85c, Nor97, NS05, Nyr99, Oga80, Oht91, Ort08, Ott84b, Oza80, OI01, Pak75, Pak78b, PS84, Par86, PRS08, PR08]. Time [Pen96a, PSZ04, PT81, PLC03, PT11, Pol94, Pol99, PS02b, Pra64, Pra80, PRHL12, PS01b, Pro82, Pru83, Pru85, Ram84, RL88, RRS05, Rey70, RS83, RSS84, RS88, Rit80, RT00, Rob72, Rob86, Ros87a, Ros81b, RS99, Rud75, Ryd96, Sat77, SV97, SAI82, Sel97, SVJ66, Ser90, Ser06, Sha75, Sha89b, SS76, Sie86, Sol86, Sta87b, Sta91, SV71, Ste00, Ste03, Sud83, Sum84, Suo79, Sze86, Tak74, Tak92b, Tam68, Tan79a, Tan86, Tan05, Tan07, Tav80a, TR89, Tho75a, Tho11, Ton74, Toy97, Tru00, Tsa96, Tsa99, Tuc76, TW84, Tur06, VS06, Var80, Wal98, Wal83b, WW06a, Wan76b, Was69, Wat80b, Wei75b, Whi71a, Wic79, Wie80, Wil82b, Wil82c, Xu98, YM91, Yao85, Yao07a, YC08, Yap09, YT83, Zei89, ZT01, ZPR08, Ahl20, ACVZ20, AM16, ADPR18]. time [ALL21, BVW21, BS15a, BKJ16, Bai20, BC23, BS16a, BEER15, BGM15, BNW13, BM20, BL15, Cur22, DL18, DIÖ23, DHKP18, EDSM15, FCP20, Fox21, GA22, GvdH18, GMdP^{+}21, Gue13, GPZ17, GH21, HWX18, HR15, HM14, HBHH23, HL13b, Ino19, Iva16, ISK18, JS20, KT15, KBM21, KKS15, KK24, LLZ17, LTZ13, LS20, Mar16b, MK17, MV21, MZ16, Nea16, PB14, PL03, Rek23, Ren14, Rob22, SGC16, SWZ18, SRN21, Vee15, Ver14, Win15, WZ13, Yao13, Zha14]. Time-Average [PRS08]. time-backward [WZ13]. Time-Changing [GG91b]. Time-Continuous [Che97]. Time-convergent [Men23]. Time-Dependent [ALRW99, KZB05, LPSV00, Llo71, Mar05, Miy94, New68c, New68a, New68b, Pra64, Pra80, Pru85, Sel97, Sha89b, Tak92b, Ton74, YC08, BKJ16, DHKP18, FCP20]. time-fractional [Cur22, Ver14]. Time-Homogeneous [JGR91, ALL21, LLZ17, LTZ13]. Time-Inhomogeneous [Hou12, Min78]. Time-Non-Homogeneous [GRRT97]. Time-Reversibility [CDL88, Wei75b]. Time-Reversible [Suo79]. Time-Scale [Wat80b]. Time-Series [Ano73g, Bar71, Bar72b, Bar74, Car67, Han73a, McK82, Rob72]. Time-Sharing [AIH88]. Time-Slotted [MT00]. Time-Stationary [KS80b, KMS83]. Time-Varying [CC01, HW84b]. Times [AA06, AH77, Agr75, AMR01, AL12, And87, Ano72e, Ano82f, Ano87g, Ant01, AF83, ASS85, Ath88, AB86a, AIH88, APS04, BMY93, Ban96, Bel71, BR02, Ber70, Ber99, Bha68b, Bla87, Blo84, BWY11, BH12, Bon81, Bor95, Bre08, Bre12b, BC75, BSB99, CP04, Cao84, CG97b, Cha93, Cla99, Coh88, Cse91, DMS08, DN80a, DR84, Dar66, DP12a, DT01, Die99, Dob70, Dos85, DG76, Dwa67, ER88, ER89, FB88, FL97, FL98, Gal93, Gil76, GKPS06, Gla79a, GAB12, Gut75, Har90a, HP93, Hei09, Hem90, HS69b, HB02, Hun76b, IA07, Jac78, Joh06, Kao74, KS93a, KP92, KS01b, Kij88, KST99, KS93b, Kul86, Lac05, Lai77, Lam91, LL78, Lat81a, Leh88, Lel05, Les69, LS97a, LS11, LT72, MW71, Mas87, Mat92a, McK89, MS04a, Mei99, Mil67]. Times [MY91, Moh76, Nád71, Neu66, Neu78, Niu81, O’B75, OC03, Par94, Pea67, Ped71, Peg75b, Pes68, Pol71, PZC07, Pra82, Pro82, Pur75a, Res73, RR88, RS89c, Sah78, Sah83, SD87, Sch88a, Sch88b, SS88b, SS87, Sho72a, Sho72b, Sho73, STW94, Slu77, Sta81, Sta74, Ste85, Ste95, Ste81, SH70a, Sum87, SR88, SM97, TTH91, Van99, VK00, War76, Wat82b, Wau72, WVW86, WP80, Wes77a, Wil92, Win89, Woo82, Yan75, YC82, Yao07b, Yu74, Yu11, Zac04, Zha10, Zip11, AF14, BPG20, BS23, BAR14, Bre13, Coh14, DS13, DI20, FS17, GLS19, GM15, Gup13, HJ14, HL13a, HM17, LGC16, Le14, Leh14, LYZ17, LVZ19, Met13, Pro18, Ros14a, SL21, SW20, SP17, Xu17]. Timing [Ber94, ZC95]. Tin [Bal78, Boy87]. Todorovic [Pur87]. Toeplitz [DKK08]. Toes [Tav22]. Token [Bra81, Bra82]. Tokens [Bra72]. Tollbooth [CHR15]. tons [Sau23]. Too [Jac09]. Tool [Kei78]. Tools [Ham05]. top [BRSC22]. top-to-random [BRSC22]. Topics [Hea67]. topological [FH13, Owa23]. Topologically [TK84]. Topology [De 94]. Tornadoes [Lan11]. torus [ANT17]. Toruses [Sil78]. Tory [Wei79]. Tossing [DW72, PP88, FHI17, HK15]. Total [AHN90, AP94, AD97, Bal86, Bal90, BP10, Dal69b, DJ69, Dwa69, GR01, Hol76, HST75, Ken75a, KP83, Kyr99, NSK08, PS84, Ser90, SV03b, Tak92a, TR67, VS06, Web99, Wei74, YM91, ZW02]. Totally [GY04, Tru00]. tournaments [MR23]. Toy [Ald95]. Trace [ELNT11, BR23]. tracing [Ber23]. Tracking [APZ11]. Trade [LM04b, EL13]. Trade-off [LM04b]. trading [Lin14]. traditional [Jan20]. Traffic [Blo73, BC72, Bou95, BS98, Bri71, Bro69b, CW96, Cow79, Cow80, Cow81, DMS08, Erl67, GL71, Gas64, GAS03, Gra83, GMS96, HTW73, HGR90, Har73b, Har73c, Haw68, Jin88, JW90, Ken72b, Köl74, Köl79, Kro97, Kyp72b, Leh72, LF08, Lou73, Mar73a, MRR02, McN68, MO70, MO71, NG95, Nar68, New66, NMG00, OM72, OW94, OV03, Ott79, PT71, PJ02, PS99, PW96, Pra70, Rén64, RV00, Ros87a, Sha88, SW11, Sis70, SZ93, The69, Tur00, UW74, Unk79, Weg76, Whi72a, YW64, AG12, HLVM22, PS16]. Traffic-Light [Cow81]. Traffic-Time-Average [Nar68]. Train [GW94a]. trait [BS15c]. Trajectories [Tam01, Kan23]. Trajectory [Dow72a, Dow72b, Lef79]. Transaction [KPT99]. Transactional [DS92]. Transfer [Bro70, Cha97, Eco00, Eco03, LP85, Sol84]. Transform [CT90, Dav01, HTV92, Har90c, HL05, KPS04, Kle83b, KX09, LR90, LR92, Wak88, Wil99, AHO20, GG14, Kim16, KLY16, Met13, Zha21]. Transformation [Dan82b, NP00, Phe96b, Sch80a, Wat80b, LT13]. Transformations [BL73, BCY97, LM91, Mas77, TV04, Ton74, VW86, GdlI18]. Transformed [Xu09]. Transforming [Bro04b, CA98, GG95]. Transforms [Keb97, LM08, NSK08, Ros08, SW97b, Tok08, LR23, ZW17]. Transience [AFM95, Die87, GP91, Ker86, Let74, MSCZ78, Pre97a, ABI17, LLL20, Die87]. Transient [AR05b, Ano65d, Ano66c, Ano93f, Ano05e, Bal01, BKNZ91, Bha68b, BM94, BS93, BCMW83, Cam86, CL93, DN80a, DESM10, DF88, Ewe64, FR10, GBG02, Gol77, GR94, Gri81, GMS96, Har81, HY94, IFN19, JC06a, Kao74, Kei64b, Kei65, Kij92b, Mar05, Mas86b, MF01, MS86, Mur87, Nat74, PS89, Red03, Rey72b, Van81, Wan99a, Wyn73, DM21, McK14]. Transition [Ano78e, Ant78a, BK87b, Bof65, BJM93, Bur76, Cab84, Cox77, Dal69a, DKK08, GZ02, GHL03, HJ99, Har94, Kal93, KM66, KS93b, Kry72, Kry75, Mil92, MS79, NS80, New68c, Nis74b, Pap96, Peg75a, Pen96a, Pha94, PS04, Sch11, Shi02, Ste95, Twe75a, Ver98, ZZ09, ZLA99, Zho23, ANT17, CP14, CH22, GH21, HL16b, LGC16]. Transitions [Fen95]. transitory [Hon17]. Translations [Mor75]. Translocation [Wat97]. Transmission [Asm09, Cla05, Hal82b, HI01, Var80]. Transport [Gan88b, GH10, Gor75, Mul68a, Vid23]. Transportation [Cra73a, Cra74b, PW96, RO99]. Transposition [GM05, Nel82]. Trasient [LV08]. Travel [LA01]. traveling [BS14a]. Treated [Yad70]. treaties [ACVZ20]. Treating [Sud02]. Treatments [Sco82, Whi82b]. Tree [AM07, BR23, FM10, GMW90, Hai70, He03, How00b, JW11, Kat05, KL03b, MSZ94, MS04b, ST02, SSE87, TW99, Att19, BIMR22, CPSJ22, ESC15, Gab16a, IRS23, OC22, Roy14]. Tree-Dependent [GMW90]. tree-like [OC22]. Trees [BPS12, Bas10, BF09, Dek87, Dob96, Dob74, FH11, FD84, IM03, LS09b, LMW98, LM20, MM03a, MM03b, MR11, Mun12, NH82, O’B80, Sud04, Tak92a, Tet02, Ver69, AD13, ABD17, AH15, AFP22, BS19, BM17, Ber13a, Bow18, CMW16, Con22, DG16, FH13, FH17, GR18, GvdH18, HI20, HS22, HLF19, IFN19, IK18, JOC13, Jon19, LT13, LN13, MW15, MZ20, PC16, Rob22, SP22, SBM22]. Trend [Dan69, DV87, GKL01]. Trends [BGW83, Bor99, Kün86, NRS85a, NRS85b]. Triadic [RS90]. Trials [Ano91e, Ant01, Bar88b, Bra83, BP00, CP88, Ery10, Fu01, FC02, FWL03, FC03, HH03, HY02, Hun76b, IA07, KA97, WJ95, WZ65, BAR14, KLY16]. Triangle [Ala77, HJ99, Jon99a, SW67]. Triangles [Gat82a, Hal82a, Hen83c, Ken91, SB78, HKT24]. Triangular [Agu09, And85, BDM09, Cow04, Kad74, WW04, KLP22]. Triangulations [App94]. Tribolium [Ano81f, Ren80]. Tridiagonal [KS93b, Ste95]. Tries [ALM06, MW08]. Trigger [Oga01]. Triggered [Gel93]. triggerings [LPU20]. Trigonometric [Bri86, Has82b]. Trimmed [Kem79, BIM17]. Trimming [KM04]. Trinary [KA97]. Trivariate [Enn69]. Trivial [FS98]. Trois [Let78]. Trotter [RD80]. True [Kab83]. Truncated [BS08, HS94, Kyr99, Pin11, Vit74, ZTML21]. Truncating [GG91b]. Truncation [SM05, Twe71, Twe75b, Twe98, BS15d]. Truncations [GS87]. Trust [Hit88, Ano22a, Gan10]. Tunable [BDLL08]. Tuple [Bor04]. Turbulence [BNL05]. Turn [Tam88]. Turning [Rei67]. turns [JX13]. Tweedie [DMR02]. Twenty [JFS12]. Twist [BLS12]. ’Two [Ibr07, Ali70, AN84, AARJ10, Amb65, And79, Ano66e, Ano77f, Ano87e, AK86, Arn80, Bac86, BCR11, BSW07, BR79, Bas71, Bax81, BKMS04, Bec70, BBK08, BC01a, BF91, Ber81b, BCY97, BS90, Bil81, Bla84, BH12, BS97, BGNR90, Cai03, Car99, Cha93, CRS97, CH06, CH10, CDL88, CW69, CY65, CS75, ÇJ73, Coh67b, Coh94, Cre74, CD85, Daw81, Dek87, DGNR95, Dom95, Dom96, DF81a, Dow66, Dur88, Ebr91, Eth79, Eth92b, FWL03, GW76, GW80, GT86, Gel77, GSS12, GKPS06, Gra99, Gro80, GP04b, Had05, Hav91, Hel70, HRF84, HT78, Hol08b, Hon04, Iva10, Jac99, Jac75, Jag69, JK87, KH83, KBBW07, Kao74, Kap81, KF70, KS77, Kar76c, Kel00, KS04, Kim70a, KMM81, KS93b, LR09, Lai77, LS92a, Lal86, LC84, Law79, Lee97, LS08b]. Two [Lem74, LS04, Les86, Lie78, Lit94, MM09, MT95, McN70, Mil94, MR02, Nag01, Nak86b, Nea07, Neu68b, NA04, NR11, O’F89, Ors86, OA70, Pag71, PP73a, Ped71, PS99, Pic71, Pin80, Raj74, Raj81, RCLW10, RCC12, Rén64, RH81b, RR78, Rig97, Rog78, Rüs77, SM05, SW05, SH21, Sco82, Sec97, SD94, Sev67, She96, Shi65, SG93, SR70a, Sta20, SH70b, Tay79, TR65, Ver69, Wan76a, Wan92, WY22, Wat64, WM77, XZW09, Xu09, YYZ09, ZC95, Zuc77, AKLL18, AG16, BLM20, CHY14, CY16, CLZ24, CdOR16, DHM20, EG15, EOV20, Fox15, GK18, HJ13b, KSBK23, KS17, LTZ18, Man13b, PG13, Pih15, RDM19, Rug14, Sez18, Zho15, CD85]. Two-Armed [O’F89, CLZ24]. Two-Boundary [BGNR90]. two-color [KS17]. Two-Component [Ano87e, JK87, Pin80]. Two-Dimensional [BCR11, BF91, Bil81, Bla84, Coh94, Cre74, DGNR95, DF81a, GW76, GW80, GSS12, Gra99, HRF84, Kel00, Lai77, Lal86, Lie78, Mil94, Ors86, Pag71, Pic71, WM77, AG16, DHM20, HJ13b, LTZ18]. two-envelope [EG15]. Two-Lane [BR79, McN70]. Two-Level [Hon04]. Two-Lift [Kap81]. Two-Locus [Eth79, Man13b]. Two-Machine [Rig97]. Two-Node [SM05]. Two-Parameter [BC01a, Cai03, KS93b, PP73a, Xu09, Rug14, Zho15]. Two-Person [CH06, CH10, LS92a, Sec97, ZC95]. Two-Queue [Bac86, Lee97, AKLL18, KSBK23]. Two-Sample [Gel77]. Two-Server [HT78]. Two-Sex [GT86, MM09]. Two-Sided [And79, BH12, CY65, Had05]. Two-Species [CD85]. Two-Stage [LC84, SW05, SG93, Zuc77, CdOR16, Fox15]. Two-State [Ano77f, Bax81, BBK08, Dek87, FWL03, GKPS06, Hel70, MT95, Raj74, RCLW10, RCC12, Wan92, EOV20]. Two-Step [LR09]. Two-Type [Jag69, WY22]. Two-Unit [KH83, OA70]. two-urn [CHY14]. Two-Wave [Car99]. Type [Ale01, Ara94, Ary72, AL85, AI94, BCL09, Ber81b, BC01b, Bra81, Bra82, Bre08, Bre10, Bui75, CL05, CP14, CG05, CS96, CS00, Che03b, Chr12, CC96, CM02, Cov09, CD98, Don76b, FHTZ10, Fin10, Fra06, FNV11, GL94, Gel77, GR12, Gir86, GP97, GKJ80, Hey94, Jag69, JP99, Jon99b, KS99, KS93a, Kni71, KS83a, Lag05, Lan86b, Li00b, Lin08, Mad75a, MO92, MNS01, MP04, MS85, Mur80, Nel77, Neu94a, Nis74b, PT71, PSZ04, Pha69, Pur73, Pyk94, Qui70, RN80, Ros75, Rüs06, Ryd96, SS99, Sch75a, Ste78, Sud76, SW88, Tit74, VB02, VAW04, Wei70, Wei72, AB19, BSX19, CF13, DHKP18, DN19, GR14, He21, KW14, LOR19, LH06, MW21, PD20, Psa22, WY22, He03, LT00, KB13]. Type-Token [Bra81, Bra82]. Types [BC95b, BY97, Bra72, CD85, HL04, Kel75, LR80, LH06, Mar11, PS99, Rad73a, Red03, San71, Tay79, Zha92, Jor18]. Typical [Ban07, Kin88, Rat92].

U [Bal95, Dan69, GS72, Tam88]. U-Shaped [Bal95, Dan69, GS72]. U-Turn [Tam88]. Uhlenbeck [Bee75, BSDG20, BWY11, BDY07, CM97, Dir75, DM72, FLP99, JP99, Jia12, Lac05, Li00b, LS20, LS00, Mia13, NRS85b, Ped02, QDZ21, Ren87, RS87, RS88, Tho75a, Tho11, Vee15, Wan12, Wen90, XZW09, ZZ16, ZZ09]. Ultimate [Lef78, YGN05]. Ultra [Yuk06, CH13]. ultra-high [CH13]. Ultra-Small [Yuk06]. Unbiased [Fer83, FPY20]. Unbounded [CCFG96, FH95, GZ02, GHL03, GHL05, KKP10, OO07, Rob80, Ruz10, SK91, CCFM17, GH21]. Uncertain [CB83, Smi75, Swe94, TH11]. Uncertainty [BB88, DL02, EP95, Pet81, PR67, Glo23]. Unconditional [Eng64, YML17]. Uncontrolled [Cow79, Rei67, Zac00]. uncountable [WZ13]. Underlying [AHP75, EJT05, Pel94]. Understanding [Ken94]. Undirected [Mat92b, Erh23]. Undiscounted [Coh14, Whi05]. Unequal [Ale01]. Unequally [Jon86]. Unexpected [AGSC00]. Uni [GPZ05]. Uni-Instantaneous [GPZ05]. Unification [Wit90]. Unified [AS04, Ano85g, CG99, Deh84, Hol79b, LT76, vdHKPS23, KS19, LLZ17]. Uniform [ANT78b, Bin78, Che18, Del98, Gol78b, Gri75a, GR01, Her03, HS78, Hüs88, JdlR04, KW14, NR20, Ols77, Rus83, SK90, Sta82, Van89b, Whi80b, Whi82a, Whi85b, YWK14, BJT20, Bel23, BT18, HMBL14, Lei23, MWW19]. Uniformity [Ber68, Nor77, Hen18]. Uniformization [BR87b, Van91a, VS00]. Uniformly [BNS90, Coh68, ES87, KBBW07, Law92, LX18, LT72, Mat04, Mit05, Pap05, Sév95, Zaj95]. Unifying [Aye07, KY17a, Ker20, Yam22]. Unilateral [GW82]. Unimodal [App94, QM86]. Unimodality [DW85, LT81, OS70, SN97, BB18]. unimodular [BS19]. Union [CHL98, Hun76a, Kwe75, Wal76, Ros16]. Unique [Cam97]. Uniqueness [BO90, C.12, CPLZ07, CNZ14, Eri16, FH94, GKRV94, Has79, JZG11, Kyp98, HZ22b, MSV14, Mat74, Teh13]. Unit [Aff89, BDO07, DH89, GC71b, KH83, OL65, OA70, Zuc89]. Unit-Demand [BDO07]. Units [EE76, Nak84, Par94, Raa80, SK91]. Univariate [BSC18, BHB18, Ber77c, ER88, Joe96]. Universality [MBvLW16, CF22b]. Unknown [AHBT82, RR75a, Ven79]. Unlimited [Wei05]. Unobservable [RB05, Wan77b, Yeh03, LY23]. Unreliable [CL11, DK90, GH94, HS93, Xu98]. Unrestricted [GC71b]. Unrevealed [Ano87e, JK87]. Unrooted [Dob74]. Unstable [IPV03, BPG20]. until [Ano85f, Arn72, BT82, Blo82, VK00, Wal98]. Untraceable [Nak06]. unusual [MLB16]. Up- [Coh77a]. Upcrossings [Fer06, Lin89]. Upgrading [Ros87a]. upon [Bra81, LZ02]. Upper [AL12, Ano86a, Ano95e, Ave85, Ber92, BS97, Bov99, BD94b, Cha04, CCC03, DKK08, Eva78, Gal93, Gra85, GW87, Gui94, Hun76a, Jan81, KS01a, KW13, Lou78, Pla85, Wol77a, Wol87, Bro15, SS17, YW21]. Upper-Hessenberg [DKK08]. upward [MZZ16]. upwards [Vid19]. Urban [Rad74a, HPP14]. Urn [ADY07, Agu09, AG07, BMT10, BK64, CW05, Det94, DW89, Ery08, Gan04a, Gou97, HMPS03, Hol79b, HKQ88, Hol08b, KM65, KS93b, MO93, Pit87, RW79, Tri92, Voi96, AGP13, Bor21, Che06, CK13, CHY14, Cri16, Glo23, GR13, GHW22, GdlI18, HMPS06, Ino23, JR16, KS17, LGC16, Mor06, SGS15, Thö16]. Urns [BCPR11, BDM09, Nak09, BJT20, BCPR16, Jan20, KS23b, LMS18, Sah16, Sim21]. Use [CT00, Hey86, Kab81, Kim00, KTW69, Lew64, Ma04, Sap87, Tak64c, Var80]. Used [Bha69, CS82, GAB12, Lam92, MN82, NBS08, Rad76a, Tan78, Wan80, KZN15]. Useful [Jof82, KW92, KB13, KS86c, Kle83b, Wat80b]. Usefulness [Nay87]. User [SZ06]. User-Optimal [SZ06]. Uses [Ewe64]. Using [AHN90, AH00, Agu03, ABD01, Car99, Che03b, Cow91, Dal87, Gol05, Gre81, Gue99, HS04, Hey74, HKQ88, Hul01, JI88, KPS04, KH75, Las94a, LRS94, Les06, MSS92, Mat04, NB03, Nis74a, PR08, PH86, Rou96, ST83, Sha80, Tru00, ADJ13, ASKK20, ABS15, AR15, Bai20, Bis23, BR23, But19, KADK19, Leh21, Lei23, SGC16]. Usually [BJ79]. Utile [Jof82]. Utility [Ank06, Lon09, OO07, Wal11, LY23, LHT17]. Utility-Based [OO07]. Utilization [FDVF93, PS86b].

V [Mor73d]. V. [IN86, Tso71]. Vacancy [Hal84, Hen86b]. Vacation [Blo91, Kat11, Kat12, KS86b, KR88, LT90, Tak92b, TTH91, Tan94]. Vacations [BM94, BY97, Dos85, LLPC94, LSC97]. Vaccination [BOP07, Bri98, BS13a, BS15b]. Validity [Rot94]. valuation [Ges79]. Value [And70, Ano86g, BD90, BK03, Bhu06, CCFG96, CCdO05, DV87, Di 99, FY82, Hal78, HAPV12, Jam66, KS72, Kar76c, Kei82, KM87, Loe09, Nag82a, NS05, RZ07, Res71a, Res73, RS95b, RZ08, Voo87, Wit93, Yas88, ZLW06, BS15c, GR24, Psa22]. Valued [AAO90, CCHY08, DO97, Fer95, Gue97, IK06, Mon82, Ruz10, CdSJ20]. Values [Ano85g, Ber92, BZ08, Bin69, BRPS82, BR77b, BCD93, ÇJ73, Dal81a, Deh84, Dur92, FH94, Gla77, GLS00, Gup84, GMW90, Hüs86a, Kab83, KS99, KP93, MR01, Nag82a, Sho72b, Sho73, Tav81, WC95, dHR73, EOV20, GR21, Joh24]. Vanishing [CD09]. Vardi [GCS07, Hal82b, Yao07a]. Variability [Ano84e, Bha91, CI78, DKM93, Huf86, KNR05, Miz06, SSSL10, VGT90, Whi80a, Whi85b, ZR01]. Variable [Bal76, BTH72, Bro04b, CP78, Don86, FT80, Ham95, KS82, Möh97, Nea07, Rap08, Sha75, Sta01, Yeo75, FL20]. Variable-Population-Size [Don86]. Variables [AG80, Ano75e, Ano86a, Ano98e, Ave85, BO99, BMZ07, BE87, BG79a, BS04b, BC73, BPT92, Bra80, BHX98, BL03, CMW95, CNO^{+}84, CCN10, CY65, CV09, Coh81b, Coo98, Deh85, Dei86, Dzi86, Eng93, FN70, FP97, FL98, Gal73, Gal66b, Gan82, Gol75, GS99, Gra85, HTW09, Hel70, Hen83b, Hen73b, Hey69, HH96, KS99, KK97, KT08, KSC12, Lef95, LU96, LV12b, Lit01, MP92a, MM03a, Mat84, NT04, NTaYY04, Pin11, Pol75, RT91, Rob92, RSZ05, Ros12, Rus83, Ruz10, SKF79, Sch82, She96, SD76, Sol86, Tak77, Tro83, VT11, VC96, VU09, WW07, Web80, Wei00, Wei75c, Wit96, Wol77b, XB11, XH11, AA23, BT18, BBvdMP23, CHM^{+}22, Jan15, KK18, KLP22, Leh17, MWW19, PL17, Wan14, Wir18]. Variance [AG80, Ano06g, Ber11, BH12, BHX98, FS06, FV18, GS79, Hal84, HK77, HS94, HG80, Kel86, Kem79, Law79, Lie77, Lie78, LPP12, Nan82, RH81b, Rus78, Sen04, Sgi96, Sob82, SAS04, Tam68, VS06, Wil92, ZWV10, BvdHK17, Ber13b, Buo15, DHKR19, GR21, GR24, GP13, GKS14, WG15, Yu17]. Variance-Gamma [Ano06g, FS06, Sen04]. variance-optimal [DHKR19]. Variance-Time [Lie77]. Variances [FF09, Gir82, Gol74, Hen73b, JNR09, KK97, MW10, Sud12, TV88, VG85, VT89, JR16]. Variant [RD80]. Variants [FS97]. Variates [DJ09, HS04, Mor73a, PS19]. Variation [BP10, BNGS04, BNCPW09, Bro79, Coh73, Gol78a, Gra12, Kyp98, MP12, Rou82, Web99, He21, HE16, HJL14, SRT^{+}16]. Variational [Pir80, Sun92]. Variations [GS21, IM03, HI20]. Variety [DFSD01, FM10]. Variogram [Ma04, RNJ19]. variograms [DS23]. Various [Pra82, WP80, Zha92]. Vary [O’F89]. Varying [Agr75, BMZ07, Ber92, BL10, Bon89, CC01, FG76, HW84b, Jag74b, JS04a, Jon99b, KK03, KN75, KS82, KS80a, Møl88, Moy08, NTaYY04, Pol10, WW07, Web86a, DL17, Ker20, WY22, WG15]. Vector [BL89, BCF09, BF97, CA98, DGK98, GY95b, Gri72, KT01b, KT01a, Rad73b, Rad76b, Tan83a, VT89]. Vectorial [Gar74]. Vectors [Bor94, CS95, DN04, Gue97, KMRV00, Kot94, Neu80, Sha87, SM94b, Sta04, ADJ13, CL14, DFP13, FCP18]. Vehicle [BDO07, Cow78, Sha71]. Vehicle-Actuated [Cow78]. Vehicles [AAL77, Gar74, Rei67]. Vehicular [Bri71]. Velocities [Ren77, Ren79, Ren81, SZ04]. Velocity [FS80a, Jan81, Kar76a, Wie79, Cin23]. Velocity-Dependent [Kar76a]. Vere [Hop76, Ano01k, Phi01]. Vere-Jones [Hop76, Ano01k, Phi01]. verifiable [Jas15]. verifying [RNR18]. Versatile [Neu79]. version [BRSC22, YBL23]. Versions [BB96, FR10, JM86, THB23]. Versus [Lan11, AIGR06, DM72, HS00, IFN19, SS97]. Vertex [Dei10]. Vertices [NH82, IK18, Pen18]. Vervaat [DQL19, LT13]. Very [LS08a]. Veto [JS08]. Veto-Voter [JS08]. Via [BMT10, Bha80, Bré78, FG12, IK11, Li96, Mah03, Mei99, Nak09, NMG00, RR10, AD13, Ano05e, ARW94, AN95, Ber77a, BT23, BR18, BCdlP17, BD94b, CG00b, DLM14, FPP19, GKS14, Hay00, KM23, MF01, Miy01, NG95, NP00, Oes18, Pek99, Pro96, RR16, RD80, San06, Spe86, Vel04, Wes99, XL98, Xu98]. Viability [KF70]. Vibration [Els07, HS96a]. Vibrations [Els95, Els96, Ors89, Oza80, Oza81, Oza86]. Vidanges [Dio78, Dio76]. View [AR01b, DM19, MO21]. Viewpoint [Fin82]. Viot [AFG11, CS20a]. Viral [Rad76b]. Virology [Mor71a]. Virtual [Coh68, FF91, Gha89, HL76, Kyp71, Neu68a, Tak74, KT15]. Virus [CS12, BBBN14]. Viruses [CDD10, BB14]. Visibility [YZ85]. visited [Gab16b]. Visits [Pol78]. Vitality [KL89b]. Vladimirovich [Kal96]. Voice [LG81, OS91]. Volatilities [Dur08b]. Volatility [BNGS04, BM09, CPT99, DM01, Dur08b, Eks03, JKM12, Lu10, RV08, SS07, Teh09, AKK20, BL13, DHKR19, EM24, GM15, HJL19]. Volterra [Bil77b, Ito89, JRR95, PT71, Pur73, SW88]. Volume [Ano64f, Ano65h, Ano66h, Ano67i, Ano68h, Ano69g, Ano70i, Ano71j, Ano72l, Ano73m, Ano74m, Ano75l, Ano76i, Ano77l, Ano78m, Ano79k, Ano80j, Ano81k, Ano82p, Ano83k, Ano84j, Ano85l, Ano86q, Ano87n, Ano88o, Ano89k, Ano90n, Ano91n, Ano92m, Ano93m, Ano94n, Ano94o, Ano95j, Ano96i, Ano97i, Ano98k, Ano99j, Ano00k, Ano01l, Ano01m, Ano02j, Ano03i, Ano04n, Ano05j, Ano06l, Ano06m, Ano07j, Ano07k, Ano08i, Ano08j, Ano13, Ano14, Ano15, Gat84b, GS15, JG85, JW11, Mas82, Mor73e, Rat92, Rub77, Rub79, Sta81, Ano17b, Ano17a, Ano18a, Ano18c, Ano18b, Ano19b, Ano19a, Ano19d, Ano19c, Ano19f, Ano19e, Ano19h, Ano19g, Ano20b, Ano20a, Ano20d, Ano20c, Ano20f, Ano20e, Ano21b, Ano21a, Ano21d, Ano21c, Ano21f, Ano21e, Ano21h, Ano21g]. volume [Ano22c, Ano22b, Ano22e, Ano22d, Ano22g, Ano22f, Ano22i, Ano22h]. Volumes [CO85]. Voronoi [Ano93e, BK21, DGLW17, MM82, MS92, Ori79]. Vortices [He01]. Voter [JS08, LN07]. Vylder [ABSKT19, Rob14].

Wagering [Har09]. Wait [FA85, Jac09]. Waiting [AN84, Ant01, Arn72, AM05, Ban96, BS76, Bla87, Blo84, BSB99, Coh68, Coh71, CNS99, DN80a, Dal69b, DJ69, DR84, Dsh89, DG76, Fal86, FF91, FC02, FC03, GKPS06, Had69, HN71, HH03, HL76, Hey64, Hok75, Hok77, HS69b, Hoo83, HCK02, IA07, Jen72, KA97, Kul86, Kyp71, LC84, Lan86b, Lem89, LT71, LT72, LT73, LT04b, MU00, MW71, McK89, MS04a, Min79, MY91, Moh76, Mor80, NG95, Nat75, Neu81, Niu81, Pak75, RL88, Ros87a, Ros73, Sch88b, Ser06, Ste00, Sze86, Tak74, Tam68, Tan79a, Toy97, YM91, EDSM15, Ino19, KT15, Yao13]. Waiting-Time [BS76, Fal86, Hoo83, Lan86b, Mor80, Nat75, Neu81, Pak75, Sze86, Tan79a]. Waitingroom [Neu68b]. Waitingtime [Neu68a]. Waitingtimes [Neu67]. Wald [Bac86, BE73, FL98, GP97, Hal70, Mou99, RS66, Rot94]. Wald-Type [GP97]. Walk [All04b, All10, Ano86j, BG06, Bar75c, BF91, Big77a, Big77b, BG79a, BP96a, Böh00, CS86, CR92, CR94b, Coh88, CS08, CJ08b, Cre74, CX01, DS94a, Det96, Don80b, Don83, DF88, DKS76, Dwa69, Eng65, Gan88b, Gre76a, HS69a, HRF83, HRF84, HG81, Hog86, HW88, JL98, Kap82, KP92, KP96, KS86b, Ken94, KK01, Kyp98, Lal86, Law83, LH07, Lun97, Neu94b, Pak73a, Pak78b, Par88b, PL72, RH81a, Rév04, Rit80, Rou90, Sie79, Sim00, Sta77, Tay99, TW75, VT75, VT76, Ver98, WW06b, Yad70, YY11, Zha92, BLSGS13, BJMK18, BDL19, BH20, Con13, DPW21, DLMR17, Gab16b, GH14, GS21, IK16, KS13, Kub24, KW13, LSWX20, PP14, Roy24, RU18, She13, SWZ18]. Walker [McC74]. Walks [Ali99, All04a, AM08, Ano66e, AFM95, Bac01, Bal01, BCL98, Bar75b, BKMS04, Bèr04, BZ08, BZ09, Ber99, Bie98, BL10, Bor95, Cho94, Coh94, Dal79a, DD04, DP12b, Don77, Don80a, Dow66, Eme72, Fan12, Fel91, FL98, GC71b, Grü85, GS06, GP05, Hey67, Hor98, Igl74, IM10, Ios86, Kao78, LB89, LMM10, LMS76, LT80, Li12, LR96, MP00, MM03a, MS04a, Mei10, Mil94, MS05b, NS80, Ngu90, Peg73, PP71, PSW71, Roe85, Ros93, Rot95, Sgi99, Spe73a, SS03, TW99, Voi96, XY92, AH13, BS16a, BY22, BIMR22, Bow18, CP14, CSC13, DS13, DL17, FGH^{+}21, GPSY16, IFN19, IRS23, Lam14, Leh21, LP14, LTZ18, McK14, NV20, PZ10, Rob22, SGS17, Sim14, Zha22, vLR13]. wall [Roy24]. Wandering [Han67, How00a]. Warfare [Kin02]. Warm [KH83, FHC18]. Warm-Standby [KH83]. Wasserstein [BKM19, BJ24, DDS22b, GPW22]. Waste [DJ09]. Watanabe [Bre75]. water [HPP14]. Watson [ABD17, Ano92g, Ath71, Ath12, AS16, Att19, Bag82, BRY04, Bow18, Bru84, Bui75, CB91, Coh82, DHT86, Dal16a, Dar70, Est75, FS23, Fea76, Fea81, FG76, GR87, GR89, GK96, Gei99, Ger11, GM96, GM97, Gre78, Hey70, Hey71, Hey81, Höp82, Höp85a, Höp85b, Höp86a, Hul82, Hul84, Jag74b, Jir76, JW82, JW85, KS86a, Ken75a, Kle89, KS02b, Le14, Li06b, Lin72, Loc82, LMW98, Lyn00, MM03a, MY84, MMR02, MS04b, NW07, Nan82, O’B80, Pak74, Qui70, Red03, RVJ68, Sen67, ST02, Tok08, Via94]. Wave [Car99, Dan77, KO05, Ors87b, Ren88]. Waveforms [Ren79, Ren81]. Wavelet [JS04b, HMBL14]. wavelet-based [HMBL14]. Waves [ALO05]. Way [RB98]. Weak [AML82, Agr91, And75, Ano83a, Avn79, BB84, BRPS82, CJ08a, Che80a, CMS06, CS82, DMS08, De 74, DJ00, DI20, Gut75, GKL85, Hei01, HM95, IK70, IK18, Iva81, JR95, Jan21, Kal92, Ken75b, Led97, Li09, Lie77, Lie78, Lin95, Lin97, Lin73, Loy88, McC85, Möh99, Not01, Reh79, RV00, RJ94, RR98, RS89b, ST16, SV97, Sen81, SD12, Tro83, Whi71a, Whi71b, Wie80, BCS23, BL21, JY21, Kur22, Zha14]. Weakening [Sei16]. Weakest [LRS07]. Weakly [BL00, BCN24, Eag79, Lat81a, Lie77, Wüb12, ACK20]. wealth [BMRRS16]. Wear [CF09, KFM06, MC13]. Weather [Fox65, ZR01]. Web [CL11]. Weibull [BBH18, DS08b, TK15, Wan76c]. ’Weighted [Hen82, BCS08, Bon02, CPLZ07, GMDB03, Kal89, Kal90, Lin08, MR11, Pha88, dRdC18, Ari22, BvdHK17, Gab16a, MS20, NR20]. Weights [BP04, KL03b, LN19]. Weiner [Hor79, Hor80, PT80, Tan79b, Tan80a, TP79]. Weiss [Die66, KM15a, Rou82]. Wentzell [Gag17]. Wheels [GW94a]. Where [DP00, KP92, Nat74, SCZ77, Van81]. Which [Ano65c, Bro72, CI72, ÇJ73, Dav73, FG84, Gri72, HL05, JGR91, Kad74, Kap78, Kra64, Lit92b, LV72, ML85, NR84, New70, Ros73, Rus83]. while [Fro05]. White [BC01a, Cai03, Sly07]. Whitt [LY20]. Whittle [Cow91, Git73, Gla82a, Har80]. Whole [CH70]. Whose [Fra07, Jac75, Lam68, Sar73, Mar16a]. Wicksell [Zäh90]. Wide [DFSD01]. wider [NvdBOdI24]. Width [HR88a, Kat05, TK84]. Widths [Ros07]. Wiener [Ano64c, BN05, Bro74, Cab84, Cha75, Dom95, Dom96, Fer83, FCvS15, Fos75, FG12, Hal69, HZ92, Hör05, IS83, Kei64a, Ken94, KMG19, Kuz10, Lef89, Lef06, LM08, ML98, ML67, PP73a, PS76a, PS80, PS83, PB83, RRS05, Sat77, Sch92c, She79, TR89, YH87]. Wiener-Hope [LM08]. Will [Rog78]. Willem [BS01b]. Williams [Lin16]. Win [CH06, Hai96, Pon07, Sie89]. Win-by- [Hai96, Sie89]. Winding [BF91]. Window [HR88a, LV08, WGF13]. winner [DvdHS24]. Winning [BL98, GM07, Sub79]. Wireless [GH10, Hir16]. wise [BJMK18]. Wishart [BL13, GG14]. Withdrawals [OL65]. within [ABSKT19, BB84, DS94a, EE88, Gra08, Hal84, Hun76b, Lin95, Lin97, MZ01]. without [Ano71e, BM01b, DG76, Glo23, Hey68a, Höp82, LGD20, Mej65, NW07, Pur82]. Wold [God77, Lai78]. Wolfowitz [KW14]. Wong [MU11, Won11]. Word [BL09a, GGS^{+}95, KBBW07, RD99]. Words [DSS11, IG01, MC11]. Work [Bin94, BS92b, Cha04, CNS99, FK89, Gha89, Gre86, KM93, Mi02a, Miy01, PT80, Sch72, Son96, Wei05, Wol70, Wol87, ZRU68]. Work-Conserving [Miy01, Wol70]. Work-Modulated [BS92b]. Workers [PV12]. Workload [BR98, BBP03, GM11, KSBK23, Lem89, Miy01, Moy10, Van90, YM91, Yao13]. World [DG06b, FM06, Whi88]. Wright [Kle88, Man13b, Web86a].

Xinjiang [MZ01].

Yaglom [Eva78]. Yakowitz [Ano00j]. Year [JS00]. years [Gan10]. Yeast [Gre81, OS14]. Yeh [Cha75, PP73a]. Yellow [Rad74a]. Yielding [BS92a]. yields [Psa22]. young [LM20]. young-age [LM20]. Yu. [Mor73d]. Yule [BRS75, O’C93].

z [Bon13]. Zabczyk [Ano88f, Bru87]. Zag [VR22]. Zagreb [FH11]. Zeifman [Mit04a]. Zero [AFM95, AC94b, Bar74, BJM93, BH91, DWW11, Gol05, GHL03, How00b, IS85, LS92a, Leh72, Mas77, Oht91, PP69, Pic79, Rot95, Ryc87, ST83, AN22, DLVZ20, GKS14]. Zero-Avoiding [BJM93]. Zero-Coupon [DWW11]. Zero-Memory [Mas77]. Zero-Sum [GHL03, LS92a]. Zero-Switch [Leh72]. Zero-Temperature [How00b]. zero-variance [GKS14]. Zig [VR22]. Zig-Zag [VR22]. Zipf [Che80a, FF20, WH75].

References

Alpuim:1990:SDS

[AA90] M. T. Alpuim and E. Athayde. On the stationary distribution of some extremal Markovian sequences. Journal of Applied Probability, 27(2):291–302, June 1990. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

             jstor.org/stable/3214648.

Anderson:1993:LDP

[AA93] Dale N. Anderson and Barry C. Arnold. Linnik distributions and processes. Journal of Applied Probability, 30(2):330–340, June 1993. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

             (electronic). URL http://www.jstor.org/stable/3214842.

Aalto:2006:NFB

[AA06] S. Aalto and U. Ayesta. On the nonoptimality of the foregroundbackground discipline for IMRL service times. Journal of Applied Probability, 43(2):523–534, June 2006. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/27595748.

Ashrafi:2014:DRM

[AA14] S. Ashrafi and M. Asadi. Dynamic reliability modeling of three-state networks. Journal of Applied Probability, 51(4):999– 1020, December 2014. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://projecteuclid.

org/euclid.jap/1421763324.

Ashrafi:2015:AA

[AA15] S. Ashrafi and M. Asadi. On Ashrafi and Asadi (2014). Journal of Applied Probability, 52(1):305–??, March 2015. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://projecteuclid.org/euclid.jap/1429282625.

Ashrafi:2017:FPC

[AA17] S. Ashrafi and M. Asadi. The failure probability of components in three-state networks with applications to age replacement policy. Journal of Applied Probability, 54(4): 1051–1070, December 2017. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL https:/

/www.cambridge.org/core/journals/journal-of-applied-
probability/article/failure-probability-of-components-
in-threestate-networks-with-applications-to-age-replacement-
policy/B486FDCB2A9FD697DC62B8F0284B6D81.

Asgari:2023:CCI

[AA23] Fatemeh Asgari and Mohammad Hossein Alamatsaz. Costa’s concavity inequality for dependent variables based on the multivariate Gaussian copula. Journal of Applied Probability, 60(4):1136–1156, December 2023. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL https:/

/www.cambridge.org/core/journals/journal-of-applied-
probability/article/costas-concavity-inequality-for-
dependent-variables-based-on-the-multivariate-gaussian-
copula/0A308F87FC5B9BC1A118DF65E7D91236.

Adler:2005:PMS

[AAKR05] Ilan Adler, Hyun-Soo Ahn, Richard M. Karp, and Sheldon M. Ross. A probabilistic model for the survivability of cells. Journal of Applied Probability, 42(4):919–931, December 2005. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/30040818.

Andersen:1977:ANP

[AAL77] Per K. Andersen, Søren Andersen, and Steffen L. Lauritzen. The average noise from a Poisson stream of vehicles. Journal of Applied Probability, 14(4):817–828, December 1977. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3213354.

Aalto:1998:COR

[Aal98] Samuli Aalto. Characterization of the output rate process for a Markovian storage model. Journal of Applied Probability, 35 (1):184–199, March 1998. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3215556.

Alzaid:1990:IVO

[AAO90] A. A. Alzaid and M. Al-Osh. An integer-valued p-th-order autoregressive structure (INAR(p)) process. Journal of Applied Probability, 27(2):314–324, June 1990. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3214650.

Alodat:2010:DDC

[AARJ10] M. T. Alodat, M. Al-Rawwash, and M. A. Jebrini. Duration distribution of the conjunction of two independent F processes. Journal of Applied Probability, 47(1):179–190, March 2010. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/25662508.

Aase:1976:NSD

[Aas76] Knut Aase. A note on a singular diffusion equation in population genetics. Journal of Applied Probability, 13(1):1–8, March 1976. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3212659.

Aase:1977:CEF

[Aas77] Knut K. Aase. A conditional expectation formula for diffusion processes. Journal of Applied Probability, 14(3):626–629, September 1977. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3213468.

Aase:1985:RPA

[Aas85] Knut K. Aase. R&D projects analyzed by semimartingale methods. Journal of Applied Probability, 22(2):288–299, June 1985. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/3213773.

Aase:1987:SCG

[Aas87] Knut K. Aase. Stochastic control of geometric processes. Journal of Applied Probability, 24(1):97–104, March 1987. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3214062.

Attia:1982:CFD

[AB82] F. A. Attia and P. J. Brockwell. The control of a finite dam. Journal of Applied Probability, 19(4):815–825, December 1982. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/3213834.

Aven:1986:ORT

[AB86a] Terje Aven and Bo Bergman. Optimal replacement times: A general set-up. Journal of Applied Probability, 23(2):432– 442, June 1986. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3214185.

Azoury:1986:ASP

[AB86b] K. Azoury and P. H. Brill. An application of the system-point method to inventory models under continuous review. Journal of Applied Probability, 23(3):778–789, September 1986. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3214015.

Andersson:1998:HEM

[AB98] Håkan Andersson and Tom Britton. Heterogeneity in epidemic models and its effect on the spread of infection. Journal of Applied Probability, 35(3):651–661, September 1998. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3215640.

Abraham:2000:ELA

[AB00] B. Abraham and N. Balakrishna. Estimation of limiting availability for a stationary bivariate process. Journal of Applied Probability, 37(3):696–704, September 2000. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/3215606.

Applebaum:2015:KBF

[AB15] David Applebaum and Stefan Blackwood. The Kalman–Bucy filter for integrable Lévy processes with infinite second moment. Journal of Applied Probability, 52(3):636–648, September 2015. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-

6072 (electronic). URL http://projecteuclid.org/euclid.
jap/1445543837.

Anantharam:2016:BMS

[AB16] Venkat Anantharam and François Baccelli. The Boolean model in the Shannon regime: three thresholds and related asymptotics. Journal of Applied Probability, 53(4):1001–1018, December 2016. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-

6072 (electronic). URL http://projecteuclid.org/euclid.
jap/1481132832.

Aaron:2018:CRE

[AB18] Catherine Aaron and Olivier Bodart. Convergence rates for estimators of geodesic distances and Frèchet expectations. Journal of Applied Probability, 55(4):1001–1013, December 2018. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL https://www.cambridge.org/core/journals/journal-
of-applied-probability/article/convergence-rates-for-
estimators-of-geodesic-distances-and-frechet-expectations/
B52D9E022580EDF9041C7235C41315A4.

Albrecher:2019:IPT

[AB19] Hansjörg Albrecher and Mogens Bladt. Inhomogeneous phasetype distributions and heavy tails. Journal of Applied Probability, 56(4):1044–1064, December 2019. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL https:/

/www.cambridge.org/core/journals/journal-of-applied-

probability/article/inhomogeneous-phasetype-distributions-
and-heavy-tails/5403F9884F9D6402CDA7EC51794DDAC4.

Abakuks:1973:OIP

[Aba73] Andris Abakuks. An optimal isolation policy for an epidemic. Journal of Applied Probability, 10(2):247–262, June 1973. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/3212343.

Abakuks:1979:OHP

[Aba79] Andris Abakuks. An optimal hunting policy for a stochastic logistic model. Journal of Applied Probability, 16(2):319–331, June 1979. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3212900.

Abakuks:1980:CES

[Aba80] Andris Abakuks. Conditions for evolutionarily stable strategies. Journal of Applied Probability, 17(2):559–562, June 1980. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/3213046.

Abdallaoui:1987:PMA

[Abd87] G. Abdallaoui. Probability of maintaining or attaining a structure in one step. Journal of Applied Probability, 24(4):1006–1011, December 1987. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3214223.

Andrieu:2001:CSA

[ABD01] Christophe Andrieu, Laird A. Breyer, and Arnaud Doucet. Convergence of simulated annealing using Foster–Lyapunov criteria. Journal of Applied Probability, 38(4):975–994, December 2001. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/3215778.

Abraham:2017:LLG

[ABD17] Romain Abraham, Aymen Bouaziz, and Jean-François Delmas. Local limits of Galton–Watson trees conditioned on the number of protected nodes. Journal of Applied Probability, 54(1):55–65, March 2017. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL https:/

/www.cambridge.org/core/journals/journal-of-applied-

probability/article/local-limits-of-galtonwatson-trees-
conditioned-on-the-number-of-protected-nodes/F2C9E7E126708F98E769AC9CDA66A91B.

Alpern:1999:RSG

[ABE99] Steve Alpern, V. J. Baston, and Skander Essegaier. Rendezvous search on a graph. Journal of Applied Probability, 36(1):223– 231, March 1999. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3215416.

Altman:2000:OLR

[ABGH00] Eitan Altman, Sandjai Bhulai, Bruno Gaujal, and Arie Hordijk. Open-loop routing to M parallel servers with no buffers. Journal of Applied Probability, 37(3):668–684, September 2000. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3215604.

Albrecher:2014:SRE

[ABI14] Hansjörg Albrecher, Onno J. Boxma, and Jevgenijs Ivanovs. On simple ruin expressions in dependent Sparre Andersen risk models. Journal of Applied Probability, 51(1):293–296, March 2014. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://projecteuclid.org/euclid.jap/
1395771431.

Alsmeyer:2017:NRT

[ABI17] Gerold Alsmeyer, Dariusz Buraczewski, and Alexander Iksanov. Null recurrence and transience of random difference equations in the contractive case. Journal of Applied Probability, 54(4):1089–1110, December 2017. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL https:/

/www.cambridge.org/core/journals/journal-of-applied-
probability/article/null-recurrence-and-transience-of-
random-difference-equations-in-the-contractive-case/
B35A86428326E26F434254BFC4A6D9A5.

Abramov:1994:ADM

[Abr94] V. M. Abramov. On the asymptotic distribution of the maximum number of infectives in epidemic models with immigration. Journal of Applied Probability, 31(3):606–613, September 1994. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/3215141.

Abramov:1997:PRS

[Abr97] Vyacheslav M. Abramov. On a property of a refusals stream. Journal of Applied Probability, 34(3):800–805, September 1997. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/3215106.

Abramov:2001:IGM

[Abr01a] Vyacheslav M. Abramov. Inequalities for the GI/M/1/n loss system. Journal of Applied Probability, 38(1):232–234, March 2001. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3215754.

Abramov:2001:LMG

[Abr01b] Vyacheslav M. Abramov. On losses in M^{X} /GI/1/n queues. Journal of Applied Probability, 38(4):1079–1080, December 2001. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/3215787.

Abramov:2007:OCL

[Abr07] Vyacheslav M. Abramov. Optimal control of a large dam. Journal of Applied Probability, 44(1):249–258, March 2007. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/27595835.

Anceaume:2015:NRG

[ABS15] Emmanuelle Anceaume, Yann Busnel, and Bruno Sericola. New results on a generalized coupon collector problem using Markov chains. Journal of Applied Probability, 52(2):405–418, June 2015. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://projecteuclid.org/euclid.jap/
1437658606.

Anceaume:2016:ORG

[ABSGS16] Emmanuelle Anceaume, Yann Busnel, Ernst Schulte-Geers, and Bruno Sericola. Optimization results for a generalized coupon collector problem. Journal of Applied Probability, 53(2):622–629, June 2016. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-

6072 (electronic). URL http://projecteuclid.org/euclid.
jap/1466172881.

Ankirchner:2019:VGC

[ABSKT19] Stefan Ankirchner, Christophette Blanchet-Scalliet, and Nabil Kazi-Tani. The De Vylder–Goovaerts conjecture holds within the diffusion limit. Journal of Applied Probability, 56(2):546–557, June 2019. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL https:/

/www.cambridge.org/core/journals/journal-of-applied-
probability/article/de-vyldergoovaerts-conjecture-holds-
within-the-diffusion-limit/7219A3637FBA53E50C53528B333F1C7D.

Adke:1994:ADE

[AC94a] S. R. Adke and C. Chandran. Asymptotic distributions of extremes of extremal Markov sequences. Journal of Applied Probability, 31(1):256–261, March 1994. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3215252.

Atkinson:1994:EPI

[AC94b] R. A. Atkinson and Peter Clifford. The escape probability for integrated Brownian motion with non-zero drift. Journal of Applied Probability, 31(4):921–929, December 1994. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3215317.

Alfa:2002:DTA

[AC02] Attahiru Sule Alfa and I. T. Castro. Discrete-time analysis of a repairable machine. Journal of Applied Probability, 39(3): 503–516, September 2002. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3216074.

Anderson:2020:SMW

[ACK20] David F. Anderson, Daniele Cappelletti, and Jinsu Kim. Stochastically modeled weakly reversible reaction networks with a single linkage class. Journal of Applied Probability, 57(3):792–810, September 2020. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL https:/

/www.cambridge.org/core/journals/journal-of-applied-
probability/article/stochastically-modeled-weakly-reversible-
reaction-networks-with-a-single-linkage-class/5E9F7DAE999525F8D4A3392F80B3F0C9.

Arlotto:2011:OSA

[ACSS11] Alessandro Arlotto, Robert W. Chen, Lawrence A. Shepp, and J. Michael Steele. Online selection of alternating subsequences from a random sample. Journal of Applied Probability, 48(4):

1114–1132, December 2011. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.

org/stable/23066446.

Albrecher:2020:FTR

[ACVZ20] Hansjörg Albrecher, Bohan Chen, Eleni Vatamidou, and Bert Zwart. Finite-time ruin probabilities under large-claim reinsurance treaties for heavy-tailed claim sizes. Journal of Applied Probability, 57(2):513–530, June 2020. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL https:/

/www.cambridge.org/core/journals/journal-of-applied-
probability/article/finitetime-ruin-probabilities-under-
largeclaim-reinsurance-treaties-for-heavytailed-claim-
sizes/B0C53BD00FC978770074B79DD3264037.

Aczel:1972:CPD

[Acz72] J. Aczél. On a characterization of Poisson distributions. Journal of Applied Probability, 9(4):852–856, December 1972. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3212622.

Adell:1994:AGD

[AD94] José A. Adell and Jesús De la Cal. Approximating gamma distributions by normalized negative binomial distributions. Journal of Applied Probability, 31(2):391–400, June 1994. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3215032.

Andersson:1997:LTT

[AD97] Håkan Andersson and Boualem Djehiche. Limit theorems for the total size of a spatial epidemic. Journal of Applied Probability, 34(3):698–710, September 1997. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.

org/stable/3215095.

Andersson:1998:TLT

[AD98] Håkan Andersson and Boualem Djehiche. A threshold limit theorem for the stochastic logistic epidemic. Journal of Applied Probability, 35(3):662–670, September 1998. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/3215641.

Abraham:2013:CSC

[AD13] Romain Abraham and Jean-François Delmas. A construction of a β-coalescent via the pruning of binary trees. Journal of Applied Probability, 50(3):772–790, September 2013. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

projecteuclid.org/euclid.jap/1378401235.

Adamopoulos:1975:SCI

[Ada75] L. Adamopoulos. Some counting and interval properties of the mutually-exciting processes. Journal of Applied Probability, 12 (1):78–86, March 1975. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3212409.

Antoine:1993:IAM

[ADH93] Robin Antoine, Hani Doss, and Myles Hollander. On identifiability in the autopsy model of reliability theory. Journal of Applied Probability, 30(4):913–930, December 1993. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/3214522.

Agarwal:2013:ESL

[ADJ13] Ankush Agarwal, Santanu Dey, and Sandeep Juneja. Efficient simulation of large deviation events for sums of random vectors using saddle-point representations. Journal of Applied Probability, 50(3):703–720, September 2013. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http:/

/projecteuclid.org/euclid.jap/1378401231.

Adke:1969:BDM

[Adk69] S. R. Adke. A birth, death and migration process. Journal of Applied Probability, 6(3):687–691, December 1969. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3212113.

Adler:1976:EAF

[Adl76] Robert J. Adler. Excursions above a fixed level by n-dimensional random fields. Journal of Applied Probability, 13(2):276–289, June 1976. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3212831.

Adler:1978:EGR

[Adl78] R. J. Adler. On the envelope of a Gaussian random field. Journal of Applied Probability, 15(3):502–513, September 1978. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3213113.

Ades:1982:RFM

[ADLN82] M. Adès, J.-P. Dion, G. Labelle, and K. Nanthi. Recurrence formula and the maximum likelihood estimation of the age in a simple branching process. Journal of Applied Probability, 19(4): 776–784, December 1982. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3213830.

Anselmi:2018:CAA

[ADPR18] Jonatha Anselmi, François Dufour, and Tomás Prieto-Rumeau. Computable approximations for average Markov decision processes in continuous time. Journal of Applied Probability, 55(2):571–592, June 2018. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL https:/

/www.cambridge.org/core/journals/journal-of-applied-
probability/article/computable-approximations-for-average-
markov-decision-processes-in-continuous-time/C1221FCBAAF97FDE4AE051B6576BCB1B.

Anderssen:1973:NSB

[ADW73] R. S. Anderssen, F. R. De Hood, and R. Weiss. On the numerical solution of Brownian motion processes. Journal of Applied Probability, 10(2):409–418, June 1973. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/3212357.

Abraham:2007:SCU

[ADY07] R. Abraham, J. S. Dhersin, and B. Ycart. Strong convergence for urn models with reducible replacement policy. Journal of Applied Probability, 44(3):652–660, September 2007. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/27595872.

Asadi:2004:MDE

[AEHS04] Majid Asadi, Nader Ebrahimi, G. G. Hamedani, and Ehsan S. Soofi. Maximum dynamic entropy models. Journal of Applied Probability, 41(2):379–390, June 2004. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3216023.

Asadi:2005:MDD

[AEHS05] Majid Asadi, Nader Ebrahimi, G. G. Hamedani, and Ehsan S. Soofi. Minimum dynamic discrimination information models. Journal of Applied Probability, 42(3):643–660, September 2005. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/30040847.

Asadi:2017:CGF

[AES17] Majid Asadi, Nader Ebrahimi, and Ehsan S. Soofi. Connections of Gini, Fisher, and Shannon by Bayes risk under proportional hazards. Journal of Applied Probability, 54 (4):1027–1050, December 2017. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL https:/

/www.cambridge.org/core/journals/journal-of-applied-
probability/article/connections-of-gini-fisher-and-shannon-
by-bayes-risk-under-proportional-hazards/CCCA43E06A752650C3D936372D807DB8.

Asadi:2019:AMS

[AES19] Majid Asadi, Nader Ebrahimi, and Ehsan S. Soofi. The alphamixture of survival functions. Journal of Applied Probability, 56(4):1151–1167, December 2019. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL https:/

/www.cambridge.org/core/journals/journal-of-applied-
probability/article/alphamixture-of-survival-functions/
E8EC6FCC74001B664BC9E02C66E8584C.

Asare:1983:CRT

[AF83] B. K. Asare and F. G. Foster. Conditional response times in the M/G/1 processor-sharing system. Journal of Applied Probability, 20(4):910–915, December 1983. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3213604.

Anily:1987:SAM

[AF87] S. Anily and A. Federgruen. Simulated annealing methods with general acceptance probabilities. Journal of Applied Probability, 24(3):657–667, September 1987. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.

org/stable/3214097.

Asmussen:2014:ETS

[AF14] Søren Asmussen and Sergey Foss. On exceedance times for some processes with dependent increments. Journal of Applied Probability, 51(1):136–151, March 2014. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http:/

/projecteuclid.org/euclid.jap/1395771419.

Affentranger:1989:RCD

[Aff89] Fernando Affentranger. Random circles in the d-dimensional unit ball. Journal of Applied Probability, 26(2):408–412, June 1989. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3214047.

Asselah:2011:QDF

[AFG11] Amine Asselah, Pablo A. Ferrari, and Pablo Groisman. Quasistationary distributions and Fleming–Viot processes in finite spaces. Journal of Applied Probability, 48(2):322–332, June 2011. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/23065888.

Aldous:1991:NEP

[AFGP91] David J. Aldous, Bert Fristedt, Philip S. Griffin, and William E. Pruitt. The number of extreme points in the convex hull of a random sample. Journal of Applied Probability, 28(2):287– 304, June 1991. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3214867.

Asymont:1995:RWQ

[AFM95] I. M. Asymont, G. Fayolle, and M. V. Menshikov. Random walks in a quarter plane with zero drifts: Transience and recurrence. Journal of Applied Probability, 32(4):941–955, December 1995. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3215207.

Acan:2022:GDT

[AFP22] Hüseyin Acan, Alan Frieze, and Boris Pittel. Giant descendant trees, matchings, and independent sets in agebiased attachment graphs. Journal of Applied Probability, 59(2):299–324, June 2022. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL https:/

/www.cambridge.org/core/journals/journal-of-applied-

probability/article/giant-descendant-trees-matchings-
and-independent-sets-in-agebiased-attachment-graphs/
E2A5B93BA6B7790471940BCEB7AD2028.

Andrieu:2015:QCR

[AFV15] Christophe Andrieu, Gersende Fort, and Matti Vihola. Quantitative convergence rates for subgeometric Markov chains. Journal of Applied Probability, 52(2):391–404, June 2015. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://projecteuclid.org/euclid.jap/1437658605.

Anis:1980:VMP

[AG80] A. A. Anis and M. Gharib. On the variance of the maximum of partial sums of n-exchangeable random variables with applications. Journal of Applied Probability, 17(2):432–439, June 1980. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/3213032.

Arnold:1981:ELC

[AG81] Barry C. Arnold and Richard A. Groeneveld. On excess life in certain renewal processes. Journal of Applied Probability, 18 (2):378–389, June 1981. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3213284.

Aven:1987:ORS

[AG87] Terje Aven and Simen Gaarder. Optimal replacement in a shock model: Discrete time. Journal of Applied Probability, 24(1):281– 287, March 1987. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3214082.

Antognini:2007:GPU

[AG07] Alessandro Baldi Antognini and Simone Giannerini. Generalized Pólya urn designs with null balance. Journal of Applied Probability, 44(3):661–669, September 2007. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/27595873.

Araman:2012:FBM

[AG12] Victor F. Araman and Peter W. Glynn. Fractional Brownian motion with h < 1/2 as a limit of scheduled traffic. Journal of Applied Probability, 49(3):710–718, September 2012. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/41713800.

Andjel:2016:EPO

[AG16] E. D. Andjel and L. F. Gray. Extreme paths in oriented twodimensional percolation. Journal of Applied Probability, 53 (2):369–380, June 2016. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://projecteuclid.

org/euclid.jap/1466172860.

Artalejo:1997:SSS

[AGC97] J. R. Artalejo and A. Gomez-Corral. Steady state solution of a single-server queue with linear repeated requests. Journal of Applied Probability, 34(1):223–233, March 1997. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3215189.

Artalejo:1999:SSQ

[AGC99] J. R. Artalejo and A. Gomez-Corral. On a single server queue with negative arrivals and request repeated. Journal of Applied Probability, 36(3):907–918, September 1999. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/3215450.

Agnew:1972:IAE

[Agn72] Robert A. Agnew. Inequalities with application in economic risk analysis. Journal of Applied Probability, 9(2):441–444, June 1972. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3212814.

Agnew:1975:IAR

[Agn75] Robert A. Agnew. Inequalities with application in retail inventory analysis. Journal of Applied Probability, 12(4):852–858, December 1975. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3212739.

Aletti:2013:RRU

[AGP13] Giacomo Aletti, Andrea Ghiglietti, and Anna Maria Paganoni. Randomly reinforced urn designs with prespecified allocations. Journal of Applied Probability, 50(2):486–498, June 2013. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://projecteuclid.org/euclid.jap/1371648956.

Agresti:1975:ETV

[Agr75] Alan Agresti. On the extinction times of varying and random environment branching processes. Journal of Applied Probability, 12(1):39–46, March 1975. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3212405.

Agrawal:1991:MLL

[Agr91] Rajeev Agrawal. Minimizing the learning loss in adaptive control of Markov chains under the weak accessibility condition. Journal of Applied Probability, 28(4):779–790, December 1991. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3214681.

Assaf:2000:UCB

[AGSC00] David Assaf, Larry Goldstein, and Ester Samuel-Cahn. An unexpected connection between branching processes and optimal stopping. Journal of Applied Probability, 37(3):613–626, September 2000. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3215599.

Agustin:2003:SPS

[Agu03] Marcus A. Agustin. Statistical properties of a system reliability estimator using the Littlewood software reliability model. Journal of Applied Probability, 40(3):766–778, September 2003. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/3215949.

Aguech:2009:LTR

[Agu09] Rafik Aguech. Limit theorems for random triangular urn schemes. Journal of Applied Probability, 46(3):827–843, September 2009. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/25662463.

Abdel-Hameed:1977:OOS

[AH77] M. Abdel-Hameed. Optimality of the one step look-ahead stopping times. Journal of Applied Probability, 14(1):162–169, March 1977. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3213269.

Abdel-Hameed:1984:LDP

[AH84] Mohamed Abdel-Hameed. Life distribution properties of devices subject to a pure jump damage process. Journal of Applied Probability, 21(4):816–825, December 1984. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/3213698.

Abdel-Hameed:1986:ORS

[AH86] Mohamed Abdel-Hameed. Optimum replacement of a system subject to shocks. Journal of Applied Probability, 23(1):107– 114, March 1986. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/

3214120. See correction [Ano88g].

Abdel-Hameed:2000:OCD

[AH00] Mohamed Abdel-Hameed. Optimal control of a dam using P_{λ}, τ^{M} policies and penalty cost when the input process is a compound Poisson process with positive drift. Journal of Applied Probability, 37(2):408–416, June 2000. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/3215717.

Athreya:2013:ACT

[AH13] K. B. Athreya and Jyy-I Hong. An application of the coalescence theory to branching random walks. Journal of Applied Probability, 50(3):893–899, September 2013. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

projecteuclid.org/euclid.jap/1378401245.

Acan:2015:COR

[AH15] Hüseyin Acan and Pawel Hitczenko. On the covariances of outdegrees in random plane recursive trees. Journal of Applied Probability, 52(3):904–907, September 2015. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

projecteuclid.org/euclid.jap/1445543856.

Angel:2019:LE

[AH19] Omer Angel and Mark Holmes. Letter to the Editor. Journal of Applied Probability, 56(4):1269–1270, December 2019. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL https://www.cambridge.org/core/journals/journal-
of-applied-probability/article/letter-to-the-editor/
4E525B66C2DC8129CE24C1B4F8E6A6FA.

Abdel-Hamid:1982:SPU

[AHBT82] A. R. Abdel-Hamid, J. A. Bather, and G. B. Trustrum. The secretary problem with an unknown number of candidates. Journal of Applied Probability, 19(3):619–630, September 1982. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3213519.

Ahlswede:1973:CCL

[Ahl73] Rudolf Ahlswede. Channel capacities for list codes. Journal of Applied Probability, 10(4):824–836, December 1973. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3212384.

Anh:2002:DML

[AHL02] V. V. Anh, C. C. Heyde, and N. N. Leonenko. Dynamic models of long-memory processes driven by Lévy noise. Journal of Applied Probability, 39(4):730–747, December 2002. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3215999.

Ahlberg:2020:TTC

[Ahl20] Daniel Ahlberg. Tertiles and the time constant. Journal of Applied Probability, 57(2):407–408, June 2020. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL https://www.cambridge.org/core/journals/
journal-of-applied-probability/article/tertiles-and-
the-time-constant/78916D4AD14FCDE446B3D979D1605D55.

Abdel-Hameed:1990:OCF

[AHN90] M. Abdel-Hameed and Y. Nakhi. Optimal control of a finite dam using P_{λ}, τ^{M} policies and penalty cost: Total discounted and long run average cases. Journal of Applied Probability, 27(4): 888–898, December 1990. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3214831.

Arab:2020:NCR

[AHO20] Idir Arab, Milto Hadjikyriakou, and Paulo Eduardo Oliveira. Non-comparability with respect to the convex transform order with applications. Journal of Applied Probability, 57 (4):1339–1348, December 2020. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL https:/

/www.cambridge.org/core/journals/journal-of-applied-
probability/article/noncomparability-with-respect-to-
the-convex-transform-order-with-applications/225BCD4D829A2889711671B7C3E1819B.

A-Hameed:1975:SMU

[AHP75] M. S. A-Hameed and F. Proschan. Shock models with underlying birth process. Journal of Applied Probability, 12(1):18–28, March 1975. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3212403.

Ahsanullah:1976:CED

[Ahs76] M. Ahsanullah. On a characterization of the exponential distribution by order statistics. Journal of Applied Probability, 13(4): 818–822, December 1976. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3212540.

Abdel-Hameed:1978:ORD

[AHS78a] M. Abdel-Hameed and I. N. Shimi. Optimal replacement of damaged devices. Journal of Applied Probability, 15(1):153– 161, March 1978. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3213245.

Ahsanullah:1978:CED

[Ahs78b] M. Ahsanullah. A characterization of the exponential distribution by spacings. Journal of Applied Probability, 15(3):650–653, September 1978. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3213130.

Alsmeyer:1989:OSD

[AI89] Gerold Alsmeyer and Albrecht Irle. Optimal strategies for discovering new species in continuous time. Journal of Applied Probability, 26(4):695–706, December 1989. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/3214375.

Avi-Itzhak:1994:STS

[AI94] Benjamin Avi-Itzhak. Servers in tandem with k-stage blocking and communications-type flow. Journal of Applied Probability, 31(4):1061–1069, December 1994. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3215329.

Adler:2000:ECM

[AI00] Robert J. Adler and Srikanth K. Iyer. An exponential change of measure for the Feller diffusion and some implications for superprocesses. Journal of Applied Probability, 37(1):236–245, March 2000. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3215671.

Allaart:2016:SLB

[AI16] Pieter C. Allaart and José A. Islas. A sharp lower bound for choosing the maximum of an independent sequence. Journal of Applied Probability, 53(4):1041–1051, December 2016. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://projecteuclid.org/euclid.jap/1481132835.

Avi-Itzhak:2006:SEV

[AIGR06] Benjamin Avi-Itzhak, Boaz Golany, and Uriel G. Rothblum. Strategic equilibrium versus global optimum for a pair of competing servers. Journal of Applied Probability, 43(4):1165–1172, December 2006. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
27595807.

Avi-Itzhak:1987:SSL

[AIH87] Benjamin Avi-Itzhak and Shlomo Halfin. Server sharing with a limited number of service positions and symmetric queues. Journal of Applied Probability, 24(4):990–1000, December 1987. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/3214221.

Avi-Itzhak:1988:RTM

[AIH88] Benjamin Avi-Itzhak and Shlomo Halfin. Response times in M/M/1 time-sharing schemes with limited number of service positions. Journal of Applied Probability, 25(3):579–595, September 1988. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3213986.

Avi-Itzhak:1993:STC

[AIH93] Benjamin Avi-Itzhak and Shlomo Halfin. Servers in tandem with communication and manufacturing blocking. Journal of Applied Probability, 30(2):429–437, June 1993. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3214851.

Ailam:1966:MCC

[Ail66] Gedalia Ailam. Moments of coverage and coverage spaces. Journal of Applied Probability, 3(2):550–555, December 1966. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/3212138.

Ailam:1968:PPM

[Ail68] Gedalia Ailam. On probability properties of measures of random sets and the asymptotic behavior of empirical distribution functions. Journal of Applied Probability, 5(1):196–202, April 1968. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3212086.

Aven:2000:GMR

[AJ00] Terje Aven and Uwe Jensen. A general minimal repair model. Journal of Applied Probability, 37(1):187–197, March 2000. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/3215667.

Athreya:1976:LTB

[AK76] K. B. Athreya and N. Kaplan. Limit theorems for a branching process with disasters. Journal of Applied Probability, 13(3): 466–475, September 1976. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3212466.

Ansley:1986:ETS

[AK86] Craig F. Ansley and Robert Kohn. On the equivalence of two stochastic approaches to spline smoothing. Journal of Applied Probability, 23A:391–405, 1986. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.

org/stable/3214367.

Asmussen:1993:MPP

[AK93] Søren Asmussen and Ger Koole. Marked point processes as limits of Markovian arrival streams. Journal of Applied Probability, 30 (2):365–372, June 1993. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3214845.

Asmussen:1996:RMD

[AK96] Søren Asmussen and Offer Kella. Rate modulation in dams and ruin problems. Journal of Applied Probability, 33(2):523– 535, June 1996. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3215076.

Asmussen:1997:SMG

[AK97] Søren Asmussen and Claudia Klüppelberg. Stationary M/G/1 excursions in the presence of heavy tails. Journal of Applied Probability, 34(1):208–212, March 1997. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http:/

/www.jstor.org/stable/3215187.

Armstrong:2023:ARM

[AK23] John Armstrong and Tim King. Anomalous recurrence of Markov chains on negatively curved manifolds. Journal of Applied Probability, 60(1):204–222, March 2023. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL https://www.cambridge.org/core/journals/journal-
of-applied-probability/article/anomalous-recurrence-
of-markov-chains-on-negatively-curved-manifolds/4EB47ED1DBC5AF88D9B9E8B6BEA50B1B.

Ahn:2020:STS

[AKK20] Dohyun Ahn, Kyoung-Kuk Kim, and Younghoon Kim. Smalltime smile for the multifactor volatility Heston model. Journal of Applied Probability, 57(4):1070–1087, December 2020. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL https://www.cambridge.org/core/journals/journal-
of-applied-probability/article/smalltime-smile-for-the-
multifactor-volatility-heston-model/41693B714DE9DF40BFE48F8F8E09C0D9.

Adan:2018:ORT

[AKLL18] I. J. B. F. Adan, V. G. Kulkarni, N. Lee, and E. Lefeber. Optimal routeing in two-queue polling systems. Journal of Applied Probability, 55(3):944–967, September 2018. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL https:/

/www.cambridge.org/core/journals/journal-of-applied-
probability/article/optimal-routeing-in-twoqueue-polling-
systems/6853C7EEADD088F77DE53E4B58AA2F2E.

Ano:2010:OTM

[AKM10] Katsunori Ano, Hideo Kakinuma, and Naoto Miyoshi. Odds theorem with multiple selection chances. Journal of Applied Probability, 47(4):1093–1104, December 2010. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/29777429.

Alzaid:1991:LOA

[AKP91] Abdulhamid Alzaid, Jee Soo Kim, and Frank Proschan. Laplace ordering and its applications. Journal of Applied Probability, 28 (1):116–130, March 1991. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3214745.

Assaf:1985:PPT

[AL85] David Assaf and Naftali A. Langberg. Presentation of phasetype distributions as proper mixtures. Journal of Applied Probability, 22(1):247–250, March 1985. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3213766.

Adell:2000:TFP

[AL00] José A. Adell and Alberto Lekuona. Taylor’s formula and preservation of generalized convexity for positive linear operators. Journal of Applied Probability, 37(3):765–777, September 2000. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/3215612.

Amini:2012:UDS

[AL12] Hamed Amini and Marc Lelarge. Upper deviations for split times of branching processes. Journal of Applied Probability, 49 (4):1134–1143, December 2012. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.

org/stable/41713833.

Alagar:1976:DDB

[Ala76] Vangalur S. Alagar. The distribution of the distance between random points. Journal of Applied Probability, 13(3):558–566, September 1976. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3212475.

Alagar:1977:DRT

[Ala77] V. S. Alagar. On the distribution of a random triangle. Journal of Applied Probability, 14(2):284–297, June 1977. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3212999.

Aldous:1995:DLT

[Ald95] David Aldous. Darwin’s log: A toy model of speciation and extinction. Journal of Applied Probability, 32(2):279–295, June 1995. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3215288.

Alexander:1996:SCS

[Ale96] Kenneth S. Alexander. Shortest common superstrings of random strings. Journal of Applied Probability, 33(4):1112–1126, December 1996. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3214990.

Alexandersson:2001:ESB

[Ale01] Marina Alexandersson. On the existence of the stable birthtype distribution in a general branching process cell cycle model with unequal cell division. Journal of Applied Probability, 38(3): 685–695, September 2001. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3216121.

Artalejo:2001:ABP

[ALH01] J. R. Artalejo and M. J. Lopez-Herrero. Analysis of the busy period for the M/M/c queue: An algorithmic approach. Journal of Applied Probability, 38(1):209–222, March 2001. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3215752.

Ali:1970:TRT

[Ali70] H. Ali. Two results in the theory of queues. Journal of Applied Probability, 7(1):219–226, April 1970. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3212161.

Alili:1999:ABR

[Ali99] S. Alili. Asymptotic behaviour for random walks in random environments. Journal of Applied Probability, 36(2):334–349, June 1999. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3215460.

Allaart:1997:ISC

[All97] Pieter C. Allaart. An invariant-sum characterization of Benford’s Law. Journal of Applied Probability, 34(1):288–291, March 1997. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3215195.

Allaart:2004:OSR

[All04a] Pieter Allaart. Optimal stopping rules for correlated random walks with a discount. Journal of Applied Probability, 41(2):483– 496, June 2004. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3216030.

Allaart:2004:SMC

[All04b] Pieter Allaart. Stopping the maximum of a correlated random walk, with cost for observation. Journal of Applied Probability, 41(4):998–1007, December 2004. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.

org/stable/4141373.

Allaart:2010:GBB

[All10] Pieter Allaart. A general ’Bang–Bang’ principle for predicting the maximum of a random walk. Journal of Applied Probability, 47(4):1072–1083, December 2010. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/29777427.

Allaart:2012:PSO

[All12] Pieter C. Allaart. Predicting the supremum: Optimality of ’stop at once or not at all’. Journal of Applied Probability, 49(3): 806–820, September 2012. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/41713807.

Avram:2021:GDG

[ALL21] Florin Avram, Bin Li, and Shu Li. General drawdown of general tax model in a time-homogeneous Markov framework. Journal of Applied Probability, 58(4):1131–1151, December 2021. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL https://www.cambridge.org/core/journals/journal-
of-applied-probability/article/general-drawdown-of-general-
tax-model-in-a-timehomogeneous-markov-framework/EE04B5EC61DEA30D92F8CFEF597A4241.

Aguech:2006:LDD

[ALM06] Rafik Aguech, Nabil Lasmar, and Hosam Mahmoud. Limit distribution of distances in biased random tries. Journal of Applied Probability, 43(2):377–390, June 2006. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/27595738.

Azais:2005:GCG

[ALO05] Jean-Marc Azaïs, José R. León, and Joaquïn Ortega. Geometrical characteristics of Gaussian sea waves. Journal of Applied Probability, 42(2):407–425, June 2005. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/30040801.

Alpuim:1989:EMS

[Alp89] M. Teresa Alpuim. An extremal Markovian sequence. Journal of Applied Probability, 26(2):219–232, June 1989. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3214030.

Alpern:2007:RSR

[Alp07] Steve Alpern. Rendezvous search with revealed information: Applications to the line. Journal of Applied Probability, 44(1): 1–15, March 2007. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
27595819.

Ayhan:1999:TDO

[ALRW99] H. Ayhan, J. Limon-Robles, and M. A. Wortman. On the timedependent occupancy and backlog distributions for the GI/G/∞ queue. Journal of Applied Probability, 36(2):558–569, June 1999. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/3215474.

Anh:2004:QLB

[ALS04] V. V. Anh, N. N. Leonenko, and L. M. Sakhno. Quasi-likelihood-based higher-order spectral estimation of random fields with possible long-range dependence. Journal of Applied Probability, 41A:35–53, 2004. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3215967.

Antonelli:1978:PRG

[AM78] Peter L. Antonelli and Kenneth Morgan. On the persistence of random genetic drift. Journal of Applied Probability, 15(1):179– 183, March 1978. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3213248.

Aston:2005:WTD

[AM05] John A. D. Aston and Donald E. K. Martin. Waiting time distributions of competing patterns in higher-order Markovian sequences. Journal of Applied Probability, 42(4):977–988, December 2005. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/30040822.

Alsmeyer:2007:SMF

[AM07] Gerold Alsmeyer and Matthias Meiners. A stochastic maximin fixed-point equation related to game tree evaluation. Journal of Applied Probability, 44(3):586–606, September 2007. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/27595868.

Allaart:2008:OBS

[AM08] Pieter Allaart and Michael Monticino. Optimal buy/sell rules for correlated random walks. Journal of Applied Probability, 45 (1):33–44, March 2008. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/27595921.

Aoudia:2010:NRB

[AM10] Djilali Ait Aoudia and Éric Marchand. On the number of runs for Bernoulli arrays. Journal of Applied Probability, 47(2):367– 377, June 2010. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
20778485.

Alvarez:2014:OSM

[AM14] Luis H. R. Alvarez and Pekka Matomäki. Optimal stopping of the maximum process. Journal of Applied Probability, 51(3): 818–836, September 2014. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://projecteuclid.

org/euclid.jap/1409932676.

Alsmeyer:2016:RAA

[AM16] Gerold Alsmeyer and Alexander Marynych. Renewal approximation for the absorption time of a decreasing Markov chain. Journal of Applied Probability, 53(3):765–782, September 2016. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://projecteuclid.org/euclid.jap/1476370775.

Alexeev:2018:PGG

[AM18] Boris Alexeev and Dustin G. Mixon. Partisan gerrymandering with geographically compact districts. Journal of Applied Probability, 55(4):1046–1059, December 2018. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL https:/

/www.cambridge.org/core/journals/journal-of-applied-
probability/article/partisan-gerrymandering-with-geographically-
compact-districts/CAA387454B3BCCCCA57FF47CC0E209A1.

Al-Mutairi:1997:BAL

[AMAKAH97] Dhaifalla K. Al-Mutairi, Talal M. Al-Khamis, and Mohamed S. Abdel-Hameed. A Bayesian analysis of layered defense systems. Journal of Applied Probability, 34(2):449–457, June 1997. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/3215384.

Ambartzumian:1965:TIP

[Amb65] R. V. Ambartzumian. Two inverse problems concerning the superposition of recurrent point processes. Journal of Applied Probability, 2(2):449–454, December 1965. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/3212205.

Abdel-Moneim:1982:WLF

[AML82] Atef M. Abdel-Moneim and Frederick W. Leysieffer. Weak lumpability in finite Markov chains. Journal of Applied Probability, 19(3):685–691, September 1982. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3213528.

Abdel-Moneim:1984:LNI

[AML84] Atef M. Abdel-Moneim and Frederick W. Leysieffer. Lumpability for non-irreducible finite Markov chains. Journal of Applied Probability, 21(3):567–574, September 1984. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/3213618.

Ammann:1983:SLT

[Amm83] Larry P. Ammann. Some limit theorems for clustered occupancy models. Journal of Applied Probability, 20(4):788–802, December 1983. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3213590.

Aldous:2001:SBC

[AMR01] David Aldous, Masakiyo Miyazawa, and Tomasz Rolski. On the stability of a batch clearing system with Poisson arrivals and subadditive service times. Journal of Applied Probability, 38(3): 621–634, September 2001. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3216116.

AlMasry:2021:SCA

[AMV21] Zeina Al Masry, Sophie Mercier, and Ghislain Verdier. Stochastic comparisons and ageing properties of an extended gamma process. Journal of Applied Probability, 58(1):140–163, March 2021. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL https:/

/www.cambridge.org/core/journals/journal-of-applied-
probability/article/stochastic-comparisons-and-ageing-
properties-of-an-extended-gamma-process/BCE8E62A01BA950AF98D119AF07AE65F.

Arratia:1988:SSL

[AMW88] Richard Arratia, Pricilla Morris, and Michael S. Waterman. Stochastic Scrabble: Large deviations for sequences with scores. Journal of Applied Probability, 25(1):106–119, March 1988. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/3214238.

Ali:1984:SST

[AN84] Osman M. E. Ali and Marcel F. Neuts. A service system with two stages of waiting and feedback of customers. Journal of Applied Probability, 21(2):404–413, June 1984. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3213649.

Asmussen:1995:RPL

[AN95] Søren Asmussen and Hanne Mandrup Nielsen. Ruin probabilities via local adjustment coefficients. Journal of Applied Probability, 32(3):736–755, September 1995. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.

org/stable/3215126.

Anderson:2022:PDZ

[AN22] David F. Anderson and Tung D. Nguyen. Prevalence of deficiency-zero reaction networks in an Erdős–Rényi framework. Journal of Applied Probability, 59(2):384–398, June 2022. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL https://www.cambridge.org/core/journals/journal-
of-applied-probability/article/prevalence-of-deficiencyzero-
reaction-networks-in-an-erdosrenyi-framework/822134FD0755E1B42C50BAB4FDF5FF90.

Anantharam:1988:MFC

[Ana88] V. Anantharam. Modelling the flow of coalescing data streams through a processor. Journal of Applied Probability, 25(1):184– 193, March 1988. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3214244.

Anantharam:1992:FST

[Ana92] V. Anantharam. On fast simulation of the time to saturation of slotted ALOHA. Journal of Applied Probability, 29(3):682–690, September 1992. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3214904.

Anderson:1970:EVT

[And70] C. W. Anderson. Extreme value theory for a class of discrete distributions with applications to some stochastic processes. Journal of Applied Probability, 7(1):99–113, April 1970. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3212152.

Anderson:1975:WCA

[And75] C. W. Anderson. Weak convergence in an appointment system. Journal of Applied Probability, 12(1):188–194, March 1975. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/3212428.

Andersen:1979:FRP

[And79] Per Kragh Andersen. Filtered renewal processes with a two-sided impact function. Journal of Applied Probability, 16(4):813–821, December 1979. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3213147.

Andersen:1981:NFM

[And81] Per Kragh Andersen. A note on filtered Markov renewal processes. Journal of Applied Probability, 18(3):752–756, September 1981. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3213332.

Anderson:1984:MMH

[And84a] William J. Anderson. A multitrap model for high-intensity reciprocity failure in the photographic grain. Journal of Applied Probability, 21(3):464–478, September 1984. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/3213610.

Andrews:1984:NSM

[And84b] Donald W. K. Andrews. Non-strong mixing autoregressive processes. Journal of Applied Probability, 21(4):930–934, December 1984. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3213710.

Andrews:1985:NIN

[And85] Donald W. K. Andrews. A nearly independent, but non-strong mixing, triangular array. Journal of Applied Probability, 22(3): 729–731, September 1985. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3213877.

Anderson:1987:LTG

[And87] Kevin K. Anderson. Limit theorems for general shock models with infinite mean intershock times. Journal of Applied Probability, 24(2):449–456, June 1987. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3214268.

Anderson:1988:NCS

[And88] Kevin K. Anderson. A note on cumulative shock models. Journal of Applied Probability, 25(1):220–223, March 1988. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3214249.

Andjel:2003:SOI

[And03] E. D. Andjel. Stochastic order for invariant measures of the contact process. Journal of Applied Probability, 40(4):1155–1157, December 2003. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3216067.

Andersson:2012:CCC

[And12] Patrik Andersson. Card counting in continuous time. Journal of Applied Probability, 49(1):184–198, March 2012. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/41713759.

Angerer:2011:NLD

[Ang11] Wolfgang P. Angerer. New limiting distributions for Bellman– Harris processes. Journal of Applied Probability, 48(2):313– 321, June 2011. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
23065887.

Ankirchner:2006:MUC

[Ank06] Stefan Ankirchner. Monotone utility convergence. Journal of Applied Probability, 43(3):622–633, September 2006. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/27595760.

Anonymous:1964:BMa

[Ano64a] Anonymous. Back matter. Journal of Applied Probability, 1 (1):??, June 1964. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3212070.

Anonymous:1964:BMb

[Ano64b] Anonymous. Back matter. Journal of Applied Probability, 1(2): ??, December 1964. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3211871.

Anonymous:1964:CAW

[Ano64c] Anonymous. Correction: An alternative to Wiener–Hopf methods for the study of bounded processes. Journal of Applied Probability, 1(2):397, December 1964. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3211870. See [Kei64a].

Anonymous:1964:FMa

[Ano64d] Anonymous. Front matter. Journal of Applied Probability, 1 (1):??, June 1964. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3212058.

Anonymous:1964:FMb

[Ano64e] Anonymous. Front matter. Journal of Applied Probability, 1(2): ??, December 1964. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3211855.

Anonymous:1964:VI

[Ano64f] Anonymous. Volume information. Journal of Applied Probability, 1(2):399, December 1964. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.

org/stable/3211854.

Anonymous:1965:BMa

[Ano65a] Anonymous. Back matter. Journal of Applied Probability, 2 (1):??, June 1965. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3211889.

Anonymous:1965:BMb

[Ano65b] Anonymous. Back matter. Journal of Applied Probability, 2 (2):477, December 1965. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3212211.

Anonymous:1965:CCD

[Ano65c] Anonymous. Correction: A class of distribution function processes which have derivatives. Journal of Applied Probability, 2 (1):223, June 1965. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/

3211888. See [Kra64].

Anonymous:1965:CRT

[Ano65d] Anonymous. Correction: A review of transient behavior in regular diffusion and birth-death processes. Journal of Applied Probability, 2(1):223, June 1965. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3211887. See [Kei64b].

Anonymous:1965:CTC

[Ano65e] Anonymous. Correction: On time changes in a diagraph. Journal of Applied Probability, 2(2):475, December 1965. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3212210. See [Bha65a].

Anonymous:1965:FMa

[Ano65f] Anonymous. Front matter. Journal of Applied Probability, 2 (1):??, June 1965. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3211872.

Anonymous:1965:FMb

[Ano65g] Anonymous. Front matter. Journal of Applied Probability, 2(2): ??, December 1965. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3212192.

Anonymous:1965:VI

[Ano65h] Anonymous. Volume information. Journal of Applied Probability, 2(2):476, December 1965. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.

org/stable/3212191.

Anonymous:1966:BMa

[Ano66a] Anonymous. Back matter. Journal of Applied Probability, 3 (1):??, June 1966. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3212057.

Anonymous:1966:BMb

[Ano66b] Anonymous. Back matter. Journal of Applied Probability, 3(2): ??, December 1966. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3212142.

Anonymous:1966:CRT

[Ano66c] Anonymous. Correction: A review of transient behavior in regular diffusion and birth-death processes. Journal of Applied Probability, 3(1):??, June 1966. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3212056. See [Kei64b].

Anonymous:1966:CPR

[Ano66d] Anonymous. Correction: Probability of random paths across elementary geometrical shapes. Journal of Applied Probability, 3(1):??, June 1966. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/

3212055. See [Hor65].

Anonymous:1966:CSR

[Ano66e] Anonymous. Correction: Simple random walks on semi-groups of nilpotency class two. Journal of Applied Probability, 3(2):562, December 1966. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/

3212141. See [Dow66].

Anonymous:1966:FMa

[Ano66f] Anonymous. Front matter. Journal of Applied Probability, 3 (1):??, June 1966. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3212038.

Anonymous:1966:FMb

[Ano66g] Anonymous. Front matter. Journal of Applied Probability, 3(2): ??, December 1966. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3212121.

Anonymous:1966:VI

[Ano66h] Anonymous. Volume information. Journal of Applied Probability, 3(2):??, December 1966. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.

org/stable/3212120.

Anonymous:1967:BMa

[Ano67a] Anonymous. Back matter. Journal of Applied Probability, 4(1): ??, April 1967. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3212318.

Anonymous:1967:BMb

[Ano67b] Anonymous. Back matter. Journal of Applied Probability, 4(2): ??, August 1967. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3212037.

Anonymous:1967:BMc

[Ano67c] Anonymous. Back matter. Journal of Applied Probability, 4 (3):614, December 1967. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3212226.

Anonymous:1967:CRT

[Ano67d] Anonymous. Correction: A renewal theoretic model for chronic disease statistics. Journal of Applied Probability, 4(1):214, April 1967. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3212316.

See [Ber65].

Anonymous:1967:CIB

[Ano67e] Anonymous. Correction: Inequalities for branching processes. Journal of Applied Probability, 4(1):215, April 1967. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3212317. See [HS66].

Anonymous:1967:FMa

[Ano67f] Anonymous. Front matter. Journal of Applied Probability, 4 (1):??, April 1967. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3212294.

Anonymous:1967:FMb

[Ano67g] Anonymous. Front matter. Journal of Applied Probability, 4(2): ??, August 1967. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3212018.

Anonymous:1967:FMc

[Ano67h] Anonymous. Front matter. Journal of Applied Probability, 4(3): ??, December 1967. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3212213.

Anonymous:1967:VI

[Ano67i] Anonymous. Volume information. Journal of Applied Probability, 4(3):612–613, December 1967. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3212212.

Anonymous:1968:BMa

[Ano68a] Anonymous. Back matter. Journal of Applied Probability, 5(1): ??, April 1968. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3212092.

Anonymous:1968:BMb

[Ano68b] Anonymous. Back matter. Journal of Applied Probability, 5(2): ??, August 1968. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3212271.

Anonymous:1968:BMc

[Ano68c] Anonymous. Back matter. Journal of Applied Probability, 5(3): ??, December 1968. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3211933.

Anonymous:1968:CSA

[Ano68d] Anonymous. Corrections: Stochastic approach to chemical kinetics. Journal of Applied Probability, 5(2):484–485, August 1968. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3212270.

See [McQ67].

Anonymous:1968:FMa

[Ano68e] Anonymous. Front matter. Journal of Applied Probability, 5 (1):??, April 1968. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3212071.

Anonymous:1968:FMb

[Ano68f] Anonymous. Front matter. Journal of Applied Probability, 5(2): ??, August 1968. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3212253.

Anonymous:1968:FMc

[Ano68g] Anonymous. Front matter. Journal of Applied Probability, 5(3): ??, December 1968. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3211919.

Anonymous:1968:VI

[Ano68h] Anonymous. Volume information. Journal of Applied Probability, 5(3):??, December 1968. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.

org/stable/3211918.

Anonymous:1969:BMa

[Ano69a] Anonymous. Back matter. Journal of Applied Probability, 6(1): ??, April 1969. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3212293.

Anonymous:1969:BMb

[Ano69b] Anonymous. Back matter. Journal of Applied Probability, 6(2): ??, August 1969. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3212017.

Anonymous:1969:BMc

[Ano69c] Anonymous. Back matter. Journal of Applied Probability, 6(3): ??, December 1969. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3212119.

Anonymous:1969:FMa

[Ano69d] Anonymous. Front matter. Journal of Applied Probability, 6 (1):??, April 1969. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3212272.

Anonymous:1969:FMb

[Ano69e] Anonymous. Front matter. Journal of Applied Probability, 6(2): ??, August 1969. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3211994.

Anonymous:1969:FMc

[Ano69f] Anonymous. Front matter. Journal of Applied Probability, 6(3): ??, December 1969. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3212094.

Anonymous:1969:VI

[Ano69g] Anonymous. Volume information. Journal of Applied Probability, 6(3):??, December 1969. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.

org/stable/3212093.

Anonymous:1970:BMa

[Ano70a] Anonymous. Back matter. Journal of Applied Probability, 7(1): ??, April 1970. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3212167.

Anonymous:1970:BMb

[Ano70b] Anonymous. Back matter. Journal of Applied Probability, 7(2): ??, August 1970. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3211993.

Anonymous:1970:BMc

[Ano70c] Anonymous. Back matter. Journal of Applied Probability, 7(3): ??, December 1970. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3211960.

Anonymous:1970:CNB

[Ano70d] Anonymous. Correction: Negative binomial processes. Journal of Applied Probability, 7(1):??, April 1970. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/3212166. See [BNY69].

Anonymous:1970:CQS

[Ano70e] Anonymous. Correction: Quasi-stationary distributions in semi-Markov processes. Journal of Applied Probability, 7(3):788, December 1970. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/

3211959. See [Che70].

Anonymous:1970:FMa

[Ano70f] Anonymous. Front matter. Journal of Applied Probability, 7 (1):??, April 1970. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3212143.

Anonymous:1970:FMb

[Ano70g] Anonymous. Front matter. Journal of Applied Probability, 7(2): ??, August 1970. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3211961.

Anonymous:1970:FMc

[Ano70h] Anonymous. Front matter. Journal of Applied Probability, 7(3): ??, December 1970. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3211935.

Anonymous:1970:VI

[Ano70i] Anonymous. Volume information. Journal of Applied Probability, 7(3):??, December 1970. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.

org/stable/3211934.

Anonymous:1971:BMa

[Ano71a] Anonymous. Back matter. Journal of Applied Probability, 8(1): ??, March 1971. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3211853.

Anonymous:1971:BMb

[Ano71b] Anonymous. Back matter. Journal of Applied Probability, 8 (2):??, June 1971. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3211917.

Anonymous:1971:BMc

[Ano71c] Anonymous. Back matter. Journal of Applied Probability, 8 (3):??, September 1971. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3212190.

Anonymous:1971:BMd

[Ano71d] Anonymous. Back matter. Journal of Applied Probability, 8(4): ??, December 1971. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3212252.

Anonymous:1971:CEH

[Ano71e] Anonymous. Corrections: An extension of the Hájek–Rényi inequality for the case without moment conditions. Journal of Applied Probability, 8(2):??, June 1971. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3211916. See [Hey68a].

Anonymous:1971:FMa

[Ano71f] Anonymous. Front matter. Journal of Applied Probability, 8(1): ??, March 1971. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3211832.

Anonymous:1971:FMb

[Ano71g] Anonymous. Front matter. Journal of Applied Probability, 8 (2):??, June 1971. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3211890.

Anonymous:1971:FMc

[Ano71h] Anonymous. Front matter. Journal of Applied Probability, 8 (3):??, September 1971. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3212168.

Anonymous:1971:FMd

[Ano71i] Anonymous. Front matter. Journal of Applied Probability, 8(4): ??, December 1971. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3212228.

Anonymous:1971:VI

[Ano71j] Anonymous. Volume information. Journal of Applied Probability, 8(4):??, December 1971. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.

org/stable/3212227.

Anonymous:1972:BMa

[Ano72a] Anonymous. Back matter. Journal of Applied Probability, 9(1): ??, March 1972. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3212657.

Anonymous:1972:BMb

[Ano72b] Anonymous. Back matter. Journal of Applied Probability, 9 (2):??, June 1972. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3212821.

Anonymous:1972:BMc

[Ano72c] Anonymous. Back matter. Journal of Applied Probability, 9 (3):??, September 1972. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3212341.

Anonymous:1972:BMd

[Ano72d] Anonymous. Back matter. Journal of Applied Probability, 9(4): ??, December 1972. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3212629.

Anonymous:1972:CLT

[Ano72e] Anonymous. Correction: A limit theorem for inter-record times. Journal of Applied Probability, 9(4):877, December 1972. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/3212628. See [Sho72a].

Anonymous:1972:CNL

[Ano72f] Anonymous. Correction: Non-linear time series regression. Journal of Applied Probability, 9(3):686, September 1972. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3212340. See [Han71].

Anonymous:1972:CSS

[Ano72g] Anonymous. Correction: Study of a special M/M/1 queue. Journal of Applied Probability, 9(1):234, March 1972. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3212656. See [Tam71].

Anonymous:1972:FMa

[Ano72h] Anonymous. Front matter. Journal of Applied Probability, 9(1): ??, March 1972. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3212630.

Anonymous:1972:FMb

[Ano72i] Anonymous. Front matter. Journal of Applied Probability, 9 (2):??, June 1972. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3212795.

Anonymous:1972:FMc

[Ano72j] Anonymous. Front matter. Journal of Applied Probability, 9 (3):??, September 1972. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3212319.

Anonymous:1972:FMd

[Ano72k] Anonymous. Front matter. Journal of Applied Probability, 9(4): ??, December 1972. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3212607.

Anonymous:1972:VI

[Ano72l] Anonymous. Volume information. Journal of Applied Probability, 9(4):??, December 1972. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.

org/stable/3212606.

Anonymous:1973:BMa

[Ano73a] Anonymous. Back matter. Journal of Applied Probability, 10(1): ??, March 1973. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3212516.

Anonymous:1973:BMb

[Ano73b] Anonymous. Back matter. Journal of Applied Probability, 10 (2):??, June 1973. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3212371.

Anonymous:1973:BMc

[Ano73c] Anonymous. Back matter. Journal of Applied Probability, 10 (3):??, September 1973. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3212794.

Anonymous:1973:BMd

[Ano73d] Anonymous. Back matter. Journal of Applied Probability, 10(4): ??, December 1973. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3212400.

Anonymous:1973:CCP

[Ano73e] Anonymous. Correction: Crossing problems for non-constant thresholds and certain non-Markov processes. Journal of Applied Probability, 10(3):702, September 1973. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http:/

/www.jstor.org/stable/3212793. See [BSS72].

Anonymous:1973:CCG

[Ano73f] Anonymous. Correction: On characterization of a generalized inaccuracy measure in information theory. Journal of Applied Probability, 10(4):913, December 1973. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http:/

/www.jstor.org/stable/3212398. See [SA73].

Anonymous:1973:CAT

[Ano73g] Anonymous. Correction: The asymptotic theory of linear timeseries models. Journal of Applied Probability, 10(4):913, December 1973. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/

3212397. See [Han73a].

Anonymous:1973:FMa

[Ano73h] Anonymous. Front matter. Journal of Applied Probability, 10 (1):??, March 1973. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3212490.

Anonymous:1973:FMb

[Ano73i] Anonymous. Front matter. Journal of Applied Probability, 10 (2):??, June 1973. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3212342.

Anonymous:1973:FMc

[Ano73j] Anonymous. Front matter. Journal of Applied Probability, 10 (3):??, September 1973. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3212769.

Anonymous:1973:FMd

[Ano73k] Anonymous. Front matter. Journal of Applied Probability, 10 (4):??, December 1973. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3212373.

Anonymous:1973:NE

[Ano73l] Anonymous. Note to the Editor. Journal of Applied Probability, 10(4):914, December 1973. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3212399.

Anonymous:1973:VI

[Ano73m] Anonymous. Volume information. Journal of Applied Probability, 10(4):??, December 1973. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.

org/stable/3212372.

Anonymous:1974:BMa

[Ano74a] Anonymous. Back matter. Journal of Applied Probability, 11(1): ??, March 1974. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3212605.

Anonymous:1974:BMb

[Ano74b] Anonymous. Back matter. Journal of Applied Probability, 11 (2):??, June 1974. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3212768.

Anonymous:1974:BMc

[Ano74c] Anonymous. Back matter. Journal of Applied Probability, 11 (3):??, September 1974. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3212714.

Anonymous:1974:BMd

[Ano74d] Anonymous. Back matter. Journal of Applied Probability, 11(4): ??, December 1974. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3212576.

Anonymous:1974:CMB

[Ano74e] Anonymous. Correction: Multivariate birth-and-death processes as approximations to epidemic processes. Journal of Applied Probability, 11(1):223, March 1974. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3212604. See [Gri73].

Anonymous:1974:CDI

[Ano74f] Anonymous. Corrections: On the distribution of inanimate marks over a linear birth-and-death process. Journal of Applied Probability, 11(4):862, December 1974. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/3212575. See [Mor74].

Anonymous:1974:CSA

[Ano74g] Anonymous. Corrections: Some asymptotic results on multiple matching. Journal of Applied Probability, 11(4):862, December 1974. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/

3212574. See [HS74a].

Anonymous:1974:CGS

[Ano74h] Anonymous. Corrections: The generalised state-dependent Erlangian queue: The busy period. Journal of Applied Probability, 11(4):862, December 1974. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3212573. See [Con74].

Anonymous:1974:FMa

[Ano74i] Anonymous. Front matter. Journal of Applied Probability, 11 (1):??, March 1974. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3212577.

Anonymous:1974:FMb

[Ano74j] Anonymous. Front matter. Journal of Applied Probability, 11 (2):??, June 1974. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3212743.

Anonymous:1974:FMc

[Ano74k] Anonymous. Front matter. Journal of Applied Probability, 11 (3):??, September 1974. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3212686.

Anonymous:1974:FMd

[Ano74l] Anonymous. Front matter. Journal of Applied Probability, 11 (4):??, December 1974. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3212546.

Anonymous:1974:VI

[Ano74m] Anonymous. Volume information. Journal of Applied Probability, 11(4):863–865, December 1974. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3212545.

Anonymous:1975:BMa

[Ano75a] Anonymous. Back matter. Journal of Applied Probability, 12(1): ??, March 1975. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3212433.

Anonymous:1975:BMb

[Ano75b] Anonymous. Back matter. Journal of Applied Probability, 12 (2):??, June 1975. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3212461.

Anonymous:1975:BMc

[Ano75c] Anonymous. Back matter. Journal of Applied Probability, 12 (3):??, September 1975. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3212890.

Anonymous:1975:BMd

[Ano75d] Anonymous. Back matter. Journal of Applied Probability, 12(4): ??, December 1975. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3212742.

Anonymous:1975:COS

[Ano75e] Anonymous. Correction: Order statistics of partial sums of mutually independent random variables. Journal of Applied Probability, 12(4):864, December 1975. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3212741. See [Pol75].

Anonymous:1975:CFR

[Ano75f] Anonymous. Corrections: Further results on Kendall’s autoregressive series. Journal of Applied Probability, 12(3):658–659, September 1975. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/

3212889. See [HB75].

Anonymous:1975:CLD

[Ano75g] Anonymous. Corrections: Limiting distributions for the general branching process with immigration. Journal of Applied Probability, 12(3):657, September 1975. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3212888. See [Dur74].

Anonymous:1975:FMa

[Ano75h] Anonymous. Front matter. Journal of Applied Probability, 12 (1):??, March 1975. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3212401.

Anonymous:1975:FMb

[Ano75i] Anonymous. Front matter. Journal of Applied Probability, 12 (2):??, June 1975. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3212434.

Anonymous:1975:FMc

[Ano75j] Anonymous. Front matter. Journal of Applied Probability, 12 (3):??, September 1975. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3212855.

Anonymous:1975:FMd

[Ano75k] Anonymous. Front matter. Journal of Applied Probability, 12 (4):??, December 1975. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3212716.

Anonymous:1975:VI

[Ano75l] Anonymous. Volume information. Journal of Applied Probability, 12(4):865–868, December 1975. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3212715.

Anonymous:1976:BMa

[Ano76a] Anonymous. Back matter. Journal of Applied Probability, 13(1): ??, March 1976. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3212685.

Anonymous:1976:BMb

[Ano76b] Anonymous. Back matter. Journal of Applied Probability, 13 (2):??, June 1976. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3212854.

Anonymous:1976:BMc

[Ano76c] Anonymous. Back matter. Journal of Applied Probability, 13 (3):??, September 1976. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3212489.

Anonymous:1976:BMd

[Ano76d] Anonymous. Back matter. Journal of Applied Probability, 13(4): ??, December 1976. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3212544.

Anonymous:1976:FMa

[Ano76e] Anonymous. Front matter. Journal of Applied Probability, 13 (1):??, March 1976. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3212658.

Anonymous:1976:FMb

[Ano76f] Anonymous. Front matter. Journal of Applied Probability, 13 (2):??, June 1976. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3212822.

Anonymous:1976:FMc

[Ano76g] Anonymous. Front matter. Journal of Applied Probability, 13 (3):??, September 1976. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3212462.

Anonymous:1976:FMd

[Ano76h] Anonymous. Front matter. Journal of Applied Probability, 13 (4):??, December 1976. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3212518.

Anonymous:1976:VI

[Ano76i] Anonymous. Volume information. Journal of Applied Probability, 13(4):837–840, December 1976. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3212517.

Anonymous:1977:BMa

[Ano77a] Anonymous. Back matter. Journal of Applied Probability, 14(1): ??, March 1977. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3213278.

Anonymous:1977:BMb

[Ano77b] Anonymous. Back matter. Journal of Applied Probability, 14 (2):??, June 1977. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3213018.

Anonymous:1977:BMc

[Ano77c] Anonymous. Back matter. Journal of Applied Probability, 14 (3):??, September 1977. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3213476.

Anonymous:1977:BMd

[Ano77d] Anonymous. Back matter. Journal of Applied Probability, 14(4): ??, December 1977. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3213367.

Anonymous:1977:CPA

[Ano77e] Anonymous. Correction: Partial attraction of maxima. Journal of Applied Probability, 14(4):897, December 1977. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3213365. See [Gre76b].

Anonymous:1977:CSR

[Ano77f] Anonymous. Correction: Success runs in a two-state Markov chain. Journal of Applied Probability, 14(3):661, September 1977. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3213474.

See [Raj74].

Anonymous:1977:CSD

[Ano77g] Anonymous. Correction: The stationary distribution of the infinitely-many neutral alleles diffusion model. Journal of Applied Probability, 14(4):897, December 1977. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/3213366. See [Wat76].

Anonymous:1977:FMa

[Ano77h] Anonymous. Front matter. Journal of Applied Probability, 14 (1):??, March 1977. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3213255.

Anonymous:1977:FMb

[Ano77i] Anonymous. Front matter. Journal of Applied Probability, 14 (2):??, June 1977. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3212992.

Anonymous:1977:FMc

[Ano77j] Anonymous. Front matter. Journal of Applied Probability, 14 (3):??, September 1977. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3213445.

Anonymous:1977:FMd

[Ano77k] Anonymous. Front matter. Journal of Applied Probability, 14 (4):??, December 1977. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3213340.

Anonymous:1977:VI

[Ano77l] Anonymous. Volume information. Journal of Applied Probability, 14(4):898–901, December 1977. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3213339.

Anonymous:1978:BMa

[Ano78a] Anonymous. Back matter. Journal of Applied Probability, 15(1): ??, March 1978. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3213254.

Anonymous:1978:BMb

[Ano78b] Anonymous. Back matter. Journal of Applied Probability, 15 (2):??, June 1978. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3213421.

Anonymous:1978:BMc

[Ano78c] Anonymous. Back matter. Journal of Applied Probability, 15 (3):??, September 1978. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3213135.

Anonymous:1978:BMd

[Ano78d] Anonymous. Back matter. Journal of Applied Probability, 15(4): ??, December 1978. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3213444.

Anonymous:1978:CAF

[Ano78e] Anonymous. Correction: An asymptotic formula for the transition density of random genetic drift. Journal of Applied Probability, 15(3):661, September 1978. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3213133. See [Ant78a].

Anonymous:1978:CAP

[Ano78f] Anonymous. Correction: Asymptotic probabilities in a critical age-dependent branching process. Journal of Applied Probability, 15(3):661, September 1978. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3213134. See [Wei76].

Anonymous:1978:CPT

[Ano78g] Anonymous. Correction: Projected thick sections through multidimensional particle aggregates. Journal of Applied Probability, 15(2):456, June 1978. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3213420. See [Dav76].

Anonymous:1978:CPD

[Ano78h] Anonymous. Correction: The probability distribution of the duration of the game in the classical ruin problem. Journal of Applied Probability, 15(4):865, December 1978. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/3213443. See [Kar77].

Anonymous:1978:FMa

[Ano78i] Anonymous. Front matter. Journal of Applied Probability, 15 (1):??, March 1978. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3213230.

Anonymous:1978:FMb

[Ano78j] Anonymous. Front matter. Journal of Applied Probability, 15 (2):??, June 1978. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3213394.

Anonymous:1978:FMc

[Ano78k] Anonymous. Front matter. Journal of Applied Probability, 15 (3):??, September 1978. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3213108.

Anonymous:1978:FMd

[Ano78l] Anonymous. Front matter. Journal of Applied Probability, 15 (4):??, December 1978. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3213423.

Anonymous:1978:VI

[Ano78m] Anonymous. Volume information. Journal of Applied Probability, 15(4):866–869, December 1978. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3213422.

Anonymous:1979:BMa

[Ano79a] Anonymous. Back matter. Journal of Applied Probability, 16(1): ??, March 1979. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3213393.

Anonymous:1979:BMb

[Ano79b] Anonymous. Back matter. Journal of Applied Probability, 16 (2):??, June 1979. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3212917.

Anonymous:1979:BMc

[Ano79c] Anonymous. Back matter. Journal of Applied Probability, 16 (3):??, September 1979. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3213107.

Anonymous:1979:BMd

[Ano79d] Anonymous. Back matter. Journal of Applied Probability, 16(4): ??, December 1979. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3213160.

Anonymous:1979:CNL

[Ano79e] Anonymous. Corrections: A new look at the simple epidemic process. Journal of Applied Probability, 16(3):696, September 1979. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3213102.

See [BLL79].

Anonymous:1979:CPM

[Ano79f] Anonymous. Corrections: Poisson mixtures and quasi-infinite divisibility of distributions. Journal of Applied Probability, 16 (3):696, September 1979. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3213101. See [PG79].

Anonymous:1979:FMa

[Ano79g] Anonymous. Front matter. Journal of Applied Probability, 16 (1):??, March 1979. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3213368.

Anonymous:1979:FMb

[Ano79h] Anonymous. Front matter. Journal of Applied Probability, 16 (2):??, June 1979. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3212891.

Anonymous:1979:FMc

[Ano79i] Anonymous. Front matter. Journal of Applied Probability, 16 (3):??, September 1979. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3213075.

Anonymous:1979:FMd

[Ano79j] Anonymous. Front matter. Journal of Applied Probability, 16 (4):??, December 1979. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3213137.

Anonymous:1979:VI

[Ano79k] Anonymous. Volume information. Journal of Applied Probability, 16(4):927–930, December 1979. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3213136.

Anonymous:1980:BMa

[Ano80a] Anonymous. Back matter. Journal of Applied Probability, 17(1): ??, March 1980. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3212951.

Anonymous:1980:BMb

[Ano80b] Anonymous. Back matter. Journal of Applied Probability, 17 (2):??, June 1980. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3213051.

Anonymous:1980:BMc

[Ano80c] Anonymous. Back matter. Journal of Applied Probability, 17 (3):??, September 1980. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3212991.

Anonymous:1980:BMd

[Ano80d] Anonymous. Back matter. Journal of Applied Probability, 17(4): ??, December 1980. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3213229.

Anonymous:1980:CGG

[Ano80e] Anonymous. Correction: On a generalization of gamma processes. Journal of Applied Probability, 17(3):893, September 1980. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3212990.

See [Çin80].

Anonymous:1980:FMa

[Ano80f] Anonymous. Front matter. Journal of Applied Probability, 17 (1):??, March 1980. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3212918.

Anonymous:1980:FMb

[Ano80g] Anonymous. Front matter. Journal of Applied Probability, 17 (2):??, June 1980. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3213019.

Anonymous:1980:FMc

[Ano80h] Anonymous. Front matter. Journal of Applied Probability, 17 (3):??, September 1980. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3212952.

Anonymous:1980:FMd

[Ano80i] Anonymous. Front matter. Journal of Applied Probability, 17 (4):??, December 1980. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3213199.

Anonymous:1980:VI

[Ano80j] Anonymous. Volume information. Journal of Applied Probability, 17(4):1145–1149, December 1980. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3213198.

Anonymous:1981:BMa

[Ano81a] Anonymous. Back matter. Journal of Applied Probability, 18(1): ??, March 1981. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3213197.

Anonymous:1981:BMb

[Ano81b] Anonymous. Back matter. Journal of Applied Probability, 18 (2):??, June 1981. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3213310.

Anonymous:1981:BMc

[Ano81c] Anonymous. Back matter. Journal of Applied Probability, 18 (3):??, September 1981. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3213338.

Anonymous:1981:BMd

[Ano81d] Anonymous. Back matter. Journal of Applied Probability, 18(4): ??, December 1981. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3213074.

Anonymous:1981:CLT

[Ano81e] Anonymous. Correction: A limit theorem for the maximum term in a particular EARMA(1, 1) sequence. Journal of Applied Probability, 18(2):569, June 1981. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3213308. See [Che80b].

Anonymous:1981:CSD

[Ano81f] Anonymous. Correction: The spatial distribution of Tribolium Confusum. Journal of Applied Probability, 18(2):569, June 1981. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/3213309. See

[Ren80].

Anonymous:1981:FMa

[Ano81g] Anonymous. Front matter. Journal of Applied Probability, 18 (1):??, March 1981. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3213161.

Anonymous:1981:FMb

[Ano81h] Anonymous. Front matter. Journal of Applied Probability, 18 (2):??, June 1981. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3213279.

Anonymous:1981:FMc

[Ano81i] Anonymous. Front matter. Journal of Applied Probability, 18 (3):??, September 1981. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3213311.

Anonymous:1981:FMd

[Ano81j] Anonymous. Front matter. Journal of Applied Probability, 18 (4):??, December 1981. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3213053.

Anonymous:1981:VI

[Ano81k] Anonymous. Volume information. Journal of Applied Probability, 18(4):964–967, December 1981. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3213052.

Anonymous:1982:BMa

[Ano82a] Anonymous. Back matter. Journal of Applied Probability, 19(1): ??, March 1982. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3213939.

Anonymous:1982:BMb

[Ano82b] Anonymous. Back matter. Journal of Applied Probability, 19 (2):??, June 1982. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3213505.

Anonymous:1982:BMc

[Ano82c] Anonymous. Back matter. Journal of Applied Probability, 19 (3):??, September 1982. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3213542.

Anonymous:1982:BMd

[Ano82d] Anonymous. Back matter. Journal of Applied Probability, 19(4): ??, December 1982. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3213853.

Anonymous:1982:BM

[Ano82e] Anonymous. Back matter. Journal of Applied Probability, 19A: 305–434, 1982. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3213581.

Anonymous:1982:CBR

[Ano82f] Anonymous. Correction: Bounded recurrence times and lead time in the design of a repetitive screening program. Journal of Applied Probability, 19(4):906, December 1982. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3213851. See [Pro82].

Anonymous:1982:CCE

[Ano82g] Anonymous. Correction: Characterizations of exponential distributions by independent non-stationary record increments. Journal of Applied Probability, 19(4):906, December 1982. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/3213850. See [Pfe82a].

Anonymous:1982:CFO

[Ano82h] Anonymous. Correction: On the first-order bilinear time series model. Journal of Applied Probability, 19(3):744, September 1982. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3213541.

See [PT81].

Anonymous:1982:CSC

[Ano82i] Anonymous. Correction: Some classes of conditional inference procedures. Journal of Applied Probability, 19(4):907, December 1982. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3213852.

See [Wil82a].

Anonymous:1982:FMa

[Ano82j] Anonymous. Front matter. Journal of Applied Probability, 19 (1):??, March 1982. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3213909.

Anonymous:1982:FMb

[Ano82k] Anonymous. Front matter. Journal of Applied Probability, 19 (2):??, June 1982. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3213477.

Anonymous:1982:FMc

[Ano82l] Anonymous. Front matter. Journal of Applied Probability, 19 (3):??, September 1982. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3213506.

Anonymous:1982:FMd

[Ano82m] Anonymous. Front matter. Journal of Applied Probability, 19 (4):??, December 1982. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3213826.

Anonymous:1982:FM

[Ano82n] Anonymous. Front matter. Journal of Applied Probability, 19A:7–185, 1982. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3213543.

Anonymous:1982:PPP

[Ano82o] Anonymous. Publications of P. A. P. Moran. Journal of Applied Probability, 19A:1–6, 1982. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3213545.

Anonymous:1982:VI

[Ano82p] Anonymous. Volume information. Journal of Applied Probability, 19(4):908–911, December 1982. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3213825.

Anonymous:1983:AWC

[Ano83a] Anonymous. Acknowledgement: Weak convergence of the simple birth-and-death process. Journal of Applied Probability, 20(2): 448, June 1983. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3213823.

Anonymous:1983:BMa

[Ano83b] Anonymous. Back matter. Journal of Applied Probability, 20(1): ??, March 1983. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3213741.

Anonymous:1983:BMb

[Ano83c] Anonymous. Back matter. Journal of Applied Probability, 20 (2):??, June 1983. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3213824.

Anonymous:1983:BMc

[Ano83d] Anonymous. Back matter. Journal of Applied Probability, 20 (3):??, September 1983. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3213908.

Anonymous:1983:BMd

[Ano83e] Anonymous. Back matter. Journal of Applied Probability, 20(4): ??, December 1983. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3213607.

Anonymous:1983:CRR

[Ano83f] Anonymous. Correction: On the reduction in remaining system lifetime due to the failure of a specific component. Journal of Applied Probability, 20(3):713, September 1983. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3213907. See [Ano83f].

Anonymous:1983:FMa

[Ano83g] Anonymous. Front matter. Journal of Applied Probability, 20 (1):??, March 1983. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3213714.

Anonymous:1983:FMb

[Ano83h] Anonymous. Front matter. Journal of Applied Probability, 20 (2):??, June 1983. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3213796.

Anonymous:1983:FMc

[Ano83i] Anonymous. Front matter. Journal of Applied Probability, 20 (3):??, September 1983. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3213882.

Anonymous:1983:FMd

[Ano83j] Anonymous. Front matter. Journal of Applied Probability, 20 (4):??, December 1983. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3213583.

Anonymous:1983:VI

[Ano83k] Anonymous. Volume information. Journal of Applied Probability, 20(4):924–927, December 1983. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3213582.

Anonymous:1984:BMa

[Ano84a] Anonymous. Back matter. Journal of Applied Probability, 21(1): ??, March 1984. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3213683.

Anonymous:1984:BMb

[Ano84b] Anonymous. Back matter. Journal of Applied Probability, 21 (2):??, June 1984. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3213657.

Anonymous:1984:BMc

[Ano84c] Anonymous. Back matter. Journal of Applied Probability, 21 (3):??, September 1984. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3213632.

Anonymous:1984:BMd

[Ano84d] Anonymous. Back matter. Journal of Applied Probability, 21(4): ??, December 1984. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3213713.

Anonymous:1984:CEV

[Ano84e] Anonymous. Correction: The effect of variability in the GI/G/s queue. Journal of Applied Probability, 21(2):445–446, June 1984.

CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/3213656. See

[Whi80a].

Anonymous:1984:FMa

[Ano84f] Anonymous. Front matter. Journal of Applied Probability, 21 (1):??, March 1984. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3213658.

Anonymous:1984:FMb

[Ano84g] Anonymous. Front matter. Journal of Applied Probability, 21 (2):??, June 1984. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3213633.

Anonymous:1984:FMc

[Ano84h] Anonymous. Front matter. Journal of Applied Probability, 21 (3):??, September 1984. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3213608.

Anonymous:1984:FMd

[Ano84i] Anonymous. Front matter. Journal of Applied Probability, 21 (4):??, December 1984. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3213685.

Anonymous:1984:VI

[Ano84j] Anonymous. Volume information. Journal of Applied Probability, 21(4):938–941, December 1984. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3213684.

Anonymous:1985:BMa

[Ano85a] Anonymous. Back matter. Journal of Applied Probability, 22(1): ??, March 1985. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3213768.

Anonymous:1985:BMb

[Ano85b] Anonymous. Back matter. Journal of Applied Probability, 22 (2):??, June 1985. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3213795.

Anonymous:1985:BMc

[Ano85c] Anonymous. Back matter. Journal of Applied Probability, 22 (3):??, September 1985. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3213881.

Anonymous:1985:BMd

[Ano85d] Anonymous. Back matter. Journal of Applied Probability, 22(4): ??, December 1985. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3213973.

Anonymous:1985:CDO

[Ano85e] Anonymous. Correction: Dispersive ordering of distributions. Journal of Applied Probability, 22(3):745, September 1985. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/3213880. See [Sha82].

Anonymous:1985:CHM

[Ano85f] Anonymous. Correction: How many random digits are required until given sequences are obtained? Journal of Applied Probability, 22(2):485, June 1985. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3213794. See [BT82].

Anonymous:1985:CCD

[Ano85g] Anonymous. Correction: The characterization of distributions by order statistics and record values: A unified approach. Journal of Applied Probability, 22(4):997, December 1985. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3213972. See [Deh84].

Anonymous:1985:FMa

[Ano85h] Anonymous. Front matter. Journal of Applied Probability, 22 (1):??, March 1985. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3213742.

Anonymous:1985:FMb

[Ano85i] Anonymous. Front matter. Journal of Applied Probability, 22 (2):??, June 1985. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3213769.

Anonymous:1985:FMc

[Ano85j] Anonymous. Front matter. Journal of Applied Probability, 22 (3):??, September 1985. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3213854.

Anonymous:1985:FMd

[Ano85k] Anonymous. Front matter. Journal of Applied Probability, 22 (4):??, December 1985. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3213941.

Anonymous:1985:VI

[Ano85l] Anonymous. Volume information. Journal of Applied Probability, 22(4):998–1002, December 1985. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3213940.

Anonymous:1986:ASC

[Ano86a] Anonymous. Acknowledgement: A simple construction of an upper bound for the mean of the maximum of n identically distributed random variables. Journal of Applied Probability, 23 (3):850, September 1986. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3214027.

Anonymous:1986:BMa

[Ano86b] Anonymous. Back matter. Journal of Applied Probability, 23(1): ??, March 1986. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3214140.

Anonymous:1986:BMb

[Ano86c] Anonymous. Back matter. Journal of Applied Probability, 23 (2):??, June 1986. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3214202.

Anonymous:1986:BMc

[Ano86d] Anonymous. Back matter. Journal of Applied Probability, 23 (3):??, September 1986. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3214028.

Anonymous:1986:BMd

[Ano86e] Anonymous. Back matter. Journal of Applied Probability, 23(4): ??, December 1986. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3214483.

Anonymous:1986:BM

[Ano86f] Anonymous. Back matter. Journal of Applied Probability, 23A: 229–438, 1986. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3214371.

Anonymous:1986:CEV

[Ano86g] Anonymous. Correction: Extreme-value properties of the explosion-time distribution in a pure birth process. Journal of Applied Probability, 23(2):562, June 1986. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/3214201. See [FY82].

Anonymous:1986:CMM

[Ano86h] Anonymous. Correction: Maintainability of manpower structures: Counterexamples, results and conjectures. Journal of Applied Probability, 23(3):849, September 1986. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3214025. See [Hai83].

Anonymous:1986:CMC

[Ano86i] Anonymous. Correction: Mixed Cox processes, with an application to accident statistics. Journal of Applied Probability, 23(3): 848–849, September 1986. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3214024. See [Bar86].

Anonymous:1986:CSE

[Ano86j] Anonymous. Correction: Some exact distribution of a last onesided exit time in the simple random walk. Journal of Applied Probability, 23(3):849, September 1986. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http:/

/www.jstor.org/stable/3214026. See [CS86].

Anonymous:1986:FMa

[Ano86k] Anonymous. Front matter. Journal of Applied Probability, 23 (1):??, March 1986. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3214111.

Anonymous:1986:FMb

[Ano86l] Anonymous. Front matter. Journal of Applied Probability, 23 (2):??, June 1986. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3214171.

Anonymous:1986:FMc

[Ano86m] Anonymous. Front matter. Journal of Applied Probability, 23 (3):??, September 1986. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3213997.

Anonymous:1986:FMd

[Ano86n] Anonymous. Front matter. Journal of Applied Probability, 23 (4):??, December 1986. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3214458.

Anonymous:1986:FM

[Ano86o] Anonymous. Front matter. Journal of Applied Probability, 23A:i–157, 1986. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3214336.

Anonymous:1986:PJH

[Ano86p] Anonymous. Publications of E. J. Hannan. Journal of Applied Probability, 23A:1–5, 1986. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3214338.

Anonymous:1986:VI

[Ano86q] Anonymous. Volume information. Journal of Applied Probability, 23(4):1051–1055, December 1986. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3214457.

Anonymous:1987:BMa

[Ano87a] Anonymous. Back matter. Journal of Applied Probability, 24(1): ??, March 1987. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3214087.

Anonymous:1987:BMb

[Ano87b] Anonymous. Back matter. Journal of Applied Probability, 24 (2):??, June 1987. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3214283.

Anonymous:1987:BMc

[Ano87c] Anonymous. Back matter. Journal of Applied Probability, 24 (3):??, September 1987. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3214110.

Anonymous:1987:BMd

[Ano87d] Anonymous. Back matter. Journal of Applied Probability, 24(4): ??, December 1987. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3214226.

Anonymous:1987:CMR

[Ano87e] Anonymous. Correction: A modified repair strategy for twocomponent systems with revealed and unrevealed faults. Journal of Applied Probability, 24(4):1014, December 1987. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3214225. See [JK87].

Anonymous:1987:CSS

[Ano87f] Anonymous. Correction: A stress-strength system. Journal of Applied Probability, 24(3):786, September 1987. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3214109. See [Ebr85].

Anonymous:1987:CFP

[Ano87g] Anonymous. Correction: First-passage times with PF_{r} densities. Journal of Applied Probability, 24(2):555, June 1987. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3214282. See [ASS85].

Anonymous:1987:CCQ

[Ano87h] Anonymous. Corrections: A correlated queue with infinitely many servers. Journal of Applied Probability, 24(1):291, March 1987. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3214086.

See [Lan86a].

Anonymous:1987:CMS

[Ano87i] Anonymous. Corrections: On the moments of a self-correcting process. Journal of Applied Probability, 24(1):291, March 1987. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/3214085. See

[VJO84].

Anonymous:1987:FMa

[Ano87j] Anonymous. Front matter. Journal of Applied Probability, 24 (1):??, March 1987. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3214053.

Anonymous:1987:FMb

[Ano87k] Anonymous. Front matter. Journal of Applied Probability, 24 (2):??, June 1987. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3214253.

Anonymous:1987:FMc

[Ano87l] Anonymous. Front matter. Journal of Applied Probability, 24 (3):??, September 1987. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3214088.

Anonymous:1987:FMd

[Ano87m] Anonymous. Front matter. Journal of Applied Probability, 24 (4):??, December 1987. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3214204.

Anonymous:1987:VI

[Ano87n] Anonymous. Volume information. Journal of Applied Probability, 24(4):1015–1018, December 1987. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3214203.

Anonymous:1988:BMa

[Ano88a] Anonymous. Back matter. Journal of Applied Probability, 25(1): ??, March 1988. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3214252.

Anonymous:1988:BMb

[Ano88b] Anonymous. Back matter. Journal of Applied Probability, 25 (2):??, June 1988. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3214456.

Anonymous:1988:BMc

[Ano88c] Anonymous. Back matter. Journal of Applied Probability, 25 (3):??, September 1988. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3213996.

Anonymous:1988:BMd

[Ano88d] Anonymous. Back matter. Journal of Applied Probability, 25(4): ??, December 1988. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3214309.

Anonymous:1988:BM

[Ano88e] Anonymous. Back matter. Journal of Applied Probability, 25A: 377–387, 1988. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3214170.

Anonymous:1988:COS

[Ano88f] Anonymous. Correction: On an optimal selection problem of Cowan and Zabczyk. Journal of Applied Probability, 25(3):647, September 1988. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/

3213995. See [Bru87].

Anonymous:1988:COR

[Ano88g] Anonymous. Correction: Optimum replacement of a system subject to shocks. Journal of Applied Probability, 25(3):647, September 1988. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/

3213994. See [AH86].

Anonymous:1988:CFP

[Ano88h] Anonymous. Correction: The first-passage density of a continuous Gaussian process to a general boundary. Journal of Applied Probability, 25(4):840, December 1988. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/3214308. See [Dur85].

Anonymous:1988:FMa

[Ano88i] Anonymous. Front matter. Journal of Applied Probability, 25 (1):??, March 1988. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3214227.

Anonymous:1988:FMb

[Ano88j] Anonymous. Front matter. Journal of Applied Probability, 25 (2):??, June 1988. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3214431.

Anonymous:1988:FMc

[Ano88k] Anonymous. Front matter. Journal of Applied Probability, 25 (3):??, September 1988. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3213974.

Anonymous:1988:FMd

[Ano88l] Anonymous. Front matter. Journal of Applied Probability, 25 (4):??, December 1988. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3214285.

Anonymous:1988:FM

[Ano88m] Anonymous. Front matter. Journal of Applied Probability, 25A:i–320, 1988. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3214141.

Anonymous:1988:P

[Ano88n] Anonymous. Preface. Journal of Applied Probability, 25A:v–viii, 1988. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3214142.

Anonymous:1988:VI

[Ano88o] Anonymous. Volume information. Journal of Applied Probability, 25(4):841–844, December 1988. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3214284.

Anonymous:1989:BMa

[Ano89a] Anonymous. Back matter. Journal of Applied Probability, 26(1): ??, March 1989. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3214335.

Anonymous:1989:BMb

[Ano89b] Anonymous. Back matter. Journal of Applied Probability, 26 (2):??, June 1989. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3214052.

Anonymous:1989:BMc

[Ano89c] Anonymous. Back matter. Journal of Applied Probability, 26 (3):??, September 1989. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3214430.

Anonymous:1989:BMd

[Ano89d] Anonymous. Back matter. Journal of Applied Probability, 26(4): ??, December 1989. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3214400.

Anonymous:1989:CBP

[Ano89e] Anonymous. Correction: Applied Probability complete author and subject index no. 7. Journal of Applied Probability, 26 (4):906, December 1989. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3214399.

Anonymous:1989:COS

[Ano89f] Anonymous. Correction: Obituary: Some memories of Pat Moran. Journal of Applied Probability, 26(4):906, December 1989. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3214398.

See [Han89].

Anonymous:1989:FMa

[Ano89g] Anonymous. Front matter. Journal of Applied Probability, 26 (1):??, March 1989. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3214310.

Anonymous:1989:FMb

[Ano89h] Anonymous. Front matter. Journal of Applied Probability, 26 (2):??, June 1989. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3214029.

Anonymous:1989:FMc

[Ano89i] Anonymous. Front matter. Journal of Applied Probability, 26 (3):??, September 1989. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3214401.

Anonymous:1989:FMd

[Ano89j] Anonymous. Front matter. Journal of Applied Probability, 26 (4):??, December 1989. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3214373.

Anonymous:1989:VI

[Ano89k] Anonymous. Volume information. Journal of Applied Probability, 26(4):907–910, December 1989. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3214372.

Anonymous:1990:A

[Ano90a] Anonymous. Acknowledgement. Journal of Applied Probability, 27(1):237, March 1990. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3214613.

Anonymous:1990:BMa

[Ano90b] Anonymous. Back matter. Journal of Applied Probability, 27(1): ??, March 1990. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3214615.

Anonymous:1990:BMb

[Ano90c] Anonymous. Back matter. Journal of Applied Probability, 27 (2):??, June 1990. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3214672.

Anonymous:1990:BMc

[Ano90d] Anonymous. Back matter. Journal of Applied Probability, 27 (3):??, September 1990. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3214561.

Anonymous:1990:BMd

[Ano90e] Anonymous. Back matter. Journal of Applied Probability, 27(4): ??, December 1990. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3214836.

Anonymous:1990:CCG

[Ano90f] Anonymous. Correction: A characterization of Gumbel’s bivariate exponential and Lindley and Singpurwalla’s bivariate Lomax distributions. Journal of Applied Probability, 27(3):736, September 1990. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/

3214560. See [Roy89].

Anonymous:1990:CNS

[Ano90g] Anonymous. Correction: Notes on the stability of closed queueing networks. Journal of Applied Probability, 27(3):735, September 1990. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3214559.

See [Sig89].

Anonymous:1990:CSC

[Ano90h] Anonymous. Correction: On a system of components with joint lifetimes distributed as a mixture of independent exponential laws. Journal of Applied Probability, 27(1):238, March 1990. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/3214614. See

[LM89].

Anonymous:1990:FMa

[Ano90i] Anonymous. Front matter. Journal of Applied Probability, 27 (1):??, March 1990. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3214590.

Anonymous:1990:FMb

[Ano90j] Anonymous. Front matter. Journal of Applied Probability, 27 (2):??, June 1990. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3214642.

Anonymous:1990:FMc

[Ano90k] Anonymous. Front matter. Journal of Applied Probability, 27 (3):??, September 1990. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3214533.

Anonymous:1990:FMd

[Ano90l] Anonymous. Front matter. Journal of Applied Probability, 27 (4):??, December 1990. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3214818.

Anonymous:1990:OPF

[Ano90m] Anonymous. Obituary: Peter Franken 1937–1989. Journal of Applied Probability, 27(4):919–923, December 1990. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3214835.

Anonymous:1990:VI

[Ano90n] Anonymous. Volume information. Journal of Applied Probability, 27(4):924–927, December 1990. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3214817.

Anonymous:1991:BMa

[Ano91a] Anonymous. Back matter. Journal of Applied Probability, 28(1): ??, March 1991. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3214758.

Anonymous:1991:BMb

[Ano91b] Anonymous. Back matter. Journal of Applied Probability, 28 (2):??, June 1991. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3214886.

Anonymous:1991:BMc

[Ano91c] Anonymous. Back matter. Journal of Applied Probability, 28 (3):??, September 1991. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3214507.

Anonymous:1991:BMd

[Ano91d] Anonymous. Back matter. Journal of Applied Probability, 28(4): ??, December 1991. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3214707.

Anonymous:1991:CLT

[Ano91e] Anonymous. Correction: A limit theorem for the number of non-overlapping occurrences of a pattern in a sequence of independent trials. Journal of Applied Probability, 28(1):251, March 1991. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3214757.

See [CP88].

Anonymous:1991:CMP

[Ano91f] Anonymous. Correction: Monotonicity property of t-step maintainable structures in three-grade manpower systems: A counterexample. Journal of Applied Probability, 28(2):487, June 1991. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3214885.

See [Gue91].

Anonymous:1991:CPS

[Ano91g] Anonymous. Correction: Periodic strong ergodicity in nonhomogeneous Markov systems. Journal of Applied Probability, 28(2):487, June 1991. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3214884. See [Ger91].

Anonymous:1991:CAB

[Ano91h] Anonymous. Corrections: Asymptotic behavior of near-critical multitype branching processes. Journal of Applied Probability, 28(4):962, December 1991. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/ stable/3214705. See [Kle91, Ano92f].

Anonymous:1991:CSS

[Ano91i] Anonymous. Corrections: Spectral structure of the first-passage-time densities for classes of Markov chains. Journal of Applied Probability, 28(4):962, December 1991. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/3214706. See [Kij87].

Anonymous:1991:FMa

[Ano91j] Anonymous. Front matter. Journal of Applied Probability, 28 (1):??, March 1991. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3214734.

Anonymous:1991:FMb

[Ano91k] Anonymous. Front matter. Journal of Applied Probability, 28 (2):??, June 1991. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3214863.

Anonymous:1991:FMc

[Ano91l] Anonymous. Front matter. Journal of Applied Probability, 28 (3):??, September 1991. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3214484.

Anonymous:1991:FMd

[Ano91m] Anonymous. Front matter. Journal of Applied Probability, 28 (4):??, December 1991. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3214674.

Anonymous:1991:VI

[Ano91n] Anonymous. Volume information. Journal of Applied Probability, 28(4):963–966, December 1991. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3214673.

Anonymous:1992:A

[Ano92a] Anonymous. Acknowledgement. Journal of Applied Probability, 29(1):245, March 1992. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3214813.

Anonymous:1992:BMa

[Ano92b] Anonymous. Back matter. Journal of Applied Probability, 29(1): ??, March 1992. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3214816.

Anonymous:1992:BMb

[Ano92c] Anonymous. Back matter. Journal of Applied Probability, 29 (2):??, June 1992. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3214589.

Anonymous:1992:BMc

[Ano92d] Anonymous. Back matter. Journal of Applied Probability, 29 (3):??, September 1992. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3214914.

Anonymous:1992:BMd

[Ano92e] Anonymous. Back matter. Journal of Applied Probability, 29(4): ??, December 1992. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3214733.

Anonymous:1992:CAB

[Ano92f] Anonymous. Correction: Asymptotic behavior of near-critical multitype branching processes. Journal of Applied Probability, 29(1):246, March 1992. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/ stable/3214815. See [Kle91, Ano91h].

Anonymous:1992:CLG

[Ano92g] Anonymous. Correction: Linear growth in near-critical population-size-dependent multitype Galton–Watson processes. Journal of Applied Probability, 29(1):246, March 1992. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3214814. See [Kle89].

Anonymous:1992:CNN

[Ano92h] Anonymous. Correction: The NBUC and NWUC classes of life distributions. Journal of Applied Probability, 29(3):753, September 1992. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/

3214912. See [CW91].

Anonymous:1992:FMa

[Ano92i] Anonymous. Front matter. Journal of Applied Probability, 29 (1):??, March 1992. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3214785.

Anonymous:1992:FMb

[Ano92j] Anonymous. Front matter. Journal of Applied Probability, 29 (2):??, June 1992. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3214562.

Anonymous:1992:FMc

[Ano92k] Anonymous. Front matter. Journal of Applied Probability, 29 (3):??, September 1992. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3214887.

Anonymous:1992:FMd

[Ano92l] Anonymous. Front matter. Journal of Applied Probability, 29 (4):??, December 1992. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3214709.

Anonymous:1992:VI

[Ano92m] Anonymous. Volume information. Journal of Applied Probability, 29(4):1007–1010, December 1992. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3214708.

Anonymous:1993:BMa

[Ano93a] Anonymous. Back matter. Journal of Applied Probability, 30(1): ??, March 1993. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3214641.

Anonymous:1993:BMb

[Ano93b] Anonymous. Back matter. Journal of Applied Probability, 30 (2):??, June 1993. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3214862.

Anonymous:1993:BMc

[Ano93c] Anonymous. Back matter. Journal of Applied Probability, 30 (3):??, September 1993. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3214784.

Anonymous:1993:BMd

[Ano93d] Anonymous. Back matter. Journal of Applied Probability, 30(4): ??, December 1993. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3214532.

Anonymous:1993:CCC

[Ano93e] Anonymous. Correction: Contact and chord length distributions of the Poisson Voronoi tessellation. Journal of Applied Probability, 30(3):749, September 1993. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3214782. See [MS92].

Anonymous:1993:CEQ

[Ano93f] Anonymous. Correction: On the existence of quasi-stationary distributions in denumerable R-transient Markov chains. Journal of Applied Probability, 30(2):496, June 1993. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3214861. See [Kij92b].

Anonymous:1993:COR

[Ano93g] Anonymous. Correction: Ordering of repairable systems. Journal of Applied Probability, 30(3):749, September 1993. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3214783. See [Ebr90].

Anonymous:1993:CSP

[Ano93h] Anonymous. Correction: Semi-Pareto processes. Journal of Applied Probability, 30(4):995, December 1993. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/3214531. See [Pil91].

Anonymous:1993:FMa

[Ano93i] Anonymous. Front matter. Journal of Applied Probability, 30 (1):??, March 1993. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3214616.

Anonymous:1993:FMb

[Ano93j] Anonymous. Front matter. Journal of Applied Probability, 30 (2):??, June 1993. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3214837.

Anonymous:1993:FMc

[Ano93k] Anonymous. Front matter. Journal of Applied Probability, 30 (3):??, September 1993. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3214759.

Anonymous:1993:FMd

[Ano93l] Anonymous. Front matter. Journal of Applied Probability, 30 (4):??, December 1993. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3214509.

Anonymous:1993:VI

[Ano93m] Anonymous. Volume information. Journal of Applied Probability, 30(4):996–999, December 1993. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3214508.

Anonymous:1994:BMa

[Ano94a] Anonymous. Back matter. Journal of Applied Probability, 31(1): ??, March 1994. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3215257.

Anonymous:1994:BMb

[Ano94b] Anonymous. Back matter. Journal of Applied Probability, 31 (2):??, June 1994. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3215053.

Anonymous:1994:BMc

[Ano94c] Anonymous. Back matter. Journal of Applied Probability, 31 (3):??, September 1994. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3215167.

Anonymous:1994:BMd

[Ano94d] Anonymous. Back matter. Journal of Applied Probability, 31(4): ??, December 1994. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3215340.

Anonymous:1994:BM

[Ano94e] Anonymous. Back matter. Journal of Applied Probability, 31A: 199–375, 1994. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3214970.

Anonymous:1994:CSM

[Ano94f] Anonymous. Correction: Sums and maxima of discrete stationary processes. Journal of Applied Probability, 31(3):867, September 1994. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/

3215166. See [MS93].

Anonymous:1994:FMa

[Ano94g] Anonymous. Front matter. Journal of Applied Probability, 31 (1):??, March 1994. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3215229.

Anonymous:1994:FMb

[Ano94h] Anonymous. Front matter. Journal of Applied Probability, 31 (2):??, June 1994. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3215023.

Anonymous:1994:FMc

[Ano94i] Anonymous. Front matter. Journal of Applied Probability, 31 (3):??, September 1994. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3215139.

Anonymous:1994:FMd

[Ano94j] Anonymous. Front matter. Journal of Applied Probability, 31 (4):??, December 1994. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3215312.

Anonymous:1994:FM

[Ano94k] Anonymous. Front matter. Journal of Applied Probability, 31A:i–97, 1994. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3214940.

Anonymous:1994:OEJ

[Ano94l] Anonymous. Obituary: Edward James Hannan. Journal of Applied Probability, 31(3):857–866, September 1994. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3215165.

Anonymous:1994:PLT

[Ano94m] Anonymous. Publications of Lajos Takácz. Journal of Applied Probability, 31A:xv–xxiv, 1994. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.

org/stable/3214943.

Anonymous:1994:VIa

[Ano94n] Anonymous. Volume information. Journal of Applied Probability, 31(4):1143–1147, December 1994. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3215311.

Anonymous:1994:VIb

[Ano94o] Anonymous. Volume information. Journal of Applied Probability, 31A:377–381, 1994. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3214939.

Anonymous:1995:BMa

[Ano95a] Anonymous. Back matter. Journal of Applied Probability, 32(1): 1–3, March 1995. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3214938.

Anonymous:1995:BMb

[Ano95b] Anonymous. Back matter. Journal of Applied Probability, 32 (2):??, June 1995. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3215310.

Anonymous:1995:BMc

[Ano95c] Anonymous. Back matter. Journal of Applied Probability, 32 (3):??, September 1995. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3215138.

Anonymous:1995:BMd

[Ano95d] Anonymous. Back matter. Journal of Applied Probability, 32(4): ??, December 1995. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3215228.

Anonymous:1995:COP

[Ano95e] Anonymous. Correction: Overflow probability upper bound in fluid queues with general on-off sources. Journal of Applied Probability, 32(1):278, March 1995. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.

org/stable/3214937. See [Gui94].

Anonymous:1995:FMa

[Ano95f] Anonymous. Front matter. Journal of Applied Probability, 32 (1):??, March 1995. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3214915.

Anonymous:1995:FMb

[Ano95g] Anonymous. Front matter. Journal of Applied Probability, 32 (2):??, June 1995. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3215287.

Anonymous:1995:FMc

[Ano95h] Anonymous. Front matter. Journal of Applied Probability, 32 (3):??, September 1995. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3215112.

Anonymous:1995:FMd

[Ano95i] Anonymous. Front matter. Journal of Applied Probability, 32 (4):??, December 1995. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3215198.

Anonymous:1995:VI

[Ano95j] Anonymous. Volume information. Journal of Applied Probability, 32(4):1145–1149, December 1995. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3215197.

Anonymous:1996:BMa

[Ano96a] Anonymous. Back matter. Journal of Applied Probability, 33(1): ??, March 1996. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3215286.

Anonymous:1996:BMb

[Ano96b] Anonymous. Back matter. Journal of Applied Probability, 33 (2):??, June 1996. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3215082.

Anonymous:1996:BMc

[Ano96c] Anonymous. Back matter. Journal of Applied Probability, 33 (3):??, September 1996. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3215369.

Anonymous:1996:BMd

[Ano96d] Anonymous. Back matter. Journal of Applied Probability, 33(4): ??, December 1996. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3214997.

Anonymous:1996:FMa

[Ano96e] Anonymous. Front matter. Journal of Applied Probability, 33 (1):??, March 1996. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3215258.

Anonymous:1996:FMb

[Ano96f] Anonymous. Front matter. Journal of Applied Probability, 33 (2):??, June 1996. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3215054.

Anonymous:1996:FMc

[Ano96g] Anonymous. Front matter. Journal of Applied Probability, 33 (3):??, September 1996. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3215341.

Anonymous:1996:FMd

[Ano96h] Anonymous. Front matter. Journal of Applied Probability, 33 (4):??, December 1996. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3214972.

Anonymous:1996:VI

[Ano96i] Anonymous. Volume information. Journal of Applied Probability, 33(4):1201–1204, December 1996. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3214971.

Anonymous:1997:BMa

[Ano97a] Anonymous. Back matter. Journal of Applied Probability, 34(1): ??, March 1997. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3215196.

Anonymous:1997:BMb

[Ano97b] Anonymous. Back matter. Journal of Applied Probability, 34 (2):??, June 1997. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3215396.

Anonymous:1997:BMc

[Ano97c] Anonymous. Back matter. Journal of Applied Probability, 34 (3):??, September 1997. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3215111.

Anonymous:1997:BMd

[Ano97d] Anonymous. Back matter. Journal of Applied Probability, 34(4): ??, December 1997. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3215022.

Anonymous:1997:FMa

[Ano97e] Anonymous. Front matter. Journal of Applied Probability, 34 (1):??, March 1997. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3215168.

Anonymous:1997:FMb

[Ano97f] Anonymous. Front matter. Journal of Applied Probability, 34 (2):??, June 1997. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3215370.

Anonymous:1997:FMc

[Ano97g] Anonymous. Front matter. Journal of Applied Probability, 34 (3):??, September 1997. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3215083.

Anonymous:1997:FMd

[Ano97h] Anonymous. Front matter. Journal of Applied Probability, 34 (4):??, December 1997. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3214999.

Anonymous:1997:VI

[Ano97i] Anonymous. Volume information. Journal of Applied Probability, 34(4):1083–1087, December 1997. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3214998.

Anonymous:1998:BMa

[Ano98a] Anonymous. Back matter. Journal of Applied Probability, 35(1): ??, March 1998. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3215566.

Anonymous:1998:BMb

[Ano98b] Anonymous. Back matter. Journal of Applied Probability, 35 (2):??, June 1998. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3215707.

Anonymous:1998:BMc

[Ano98c] Anonymous. Back matter. Journal of Applied Probability, 35 (3):??, September 1998. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3215653.

Anonymous:1998:BMd

[Ano98d] Anonymous. Back matter. Journal of Applied Probability, 35(4): ??, December 1998. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3215539.

Anonymous:1998:CDE

[Ano98e] Anonymous. Correction: Distributions and expectations of singular random variables. Journal of Applied Probability, 35(1): 253, March 1998. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/

3215565. See [CMW95].

Anonymous:1998:CST

[Ano98f] Anonymous. Correction: Stationary time series models with exponential dispersion model margins. Journal of Applied Probability, 35(2):516, June 1998. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3215706. See [JS98].

Anonymous:1998:FMa

[Ano98g] Anonymous. Front matter. Journal of Applied Probability, 35 (1):??, March 1998. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3215540.

Anonymous:1998:FMb

[Ano98h] Anonymous. Front matter. Journal of Applied Probability, 35 (2):??, June 1998. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3215682.

Anonymous:1998:FMc

[Ano98i] Anonymous. Front matter. Journal of Applied Probability, 35 (3):??, September 1998. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3215627.

Anonymous:1998:FMd

[Ano98j] Anonymous. Front matter. Journal of Applied Probability, 35 (4):??, December 1998. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3215514.

Anonymous:1998:VI

[Ano98k] Anonymous. Volume information. Journal of Applied Probability, 35(4):1018–1022, December 1998. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3215513.

Anonymous:1999:APM

[Ano99a] Anonymous. Axiomatic principles for a market model. Journal of Applied Probability, 36(1):295–300, March 1999. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3215425.

Anonymous:1999:BMa

[Ano99b] Anonymous. Back matter. Journal of Applied Probability, 36(1): ??, March 1999. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3215426.

Anonymous:1999:BMb

[Ano99c] Anonymous. Back matter. Journal of Applied Probability, 36 (2):??, June 1999. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3215481.

Anonymous:1999:BMc

[Ano99d] Anonymous. Back matter. Journal of Applied Probability, 36 (3):??, September 1999. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3215455.

Anonymous:1999:BMd

[Ano99e] Anonymous. Back matter. Journal of Applied Probability, 36(4): ??, December 1999. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3215597.

Anonymous:1999:FMa

[Ano99f] Anonymous. Front matter. Journal of Applied Probability, 36 (1):??, March 1999. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3215397.

Anonymous:1999:FMb

[Ano99g] Anonymous. Front matter. Journal of Applied Probability, 36 (2):??, June 1999. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3215456.

Anonymous:1999:FMc

[Ano99h] Anonymous. Front matter. Journal of Applied Probability, 36 (3):??, September 1999. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3215427.

Anonymous:1999:FMd

[Ano99i] Anonymous. Front matter. Journal of Applied Probability, 36 (4):??, December 1999. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3215568.

Anonymous:1999:VI

[Ano99j] Anonymous. Volume information. Journal of Applied Probability, 36(4):1258–1262, December 1999. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3215567.

Anonymous:2000:A

[Ano00a] Anonymous. Acknowledgement. Journal of Applied Probability, 37(3):924, September 2000. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3215625.

Anonymous:2000:BMa

[Ano00b] Anonymous. Back matter. Journal of Applied Probability, 37(1): ??, March 2000. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3215681.

Anonymous:2000:BMb

[Ano00c] Anonymous. Back matter. Journal of Applied Probability, 37 (2):??, June 2000. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3215735.

Anonymous:2000:BMc

[Ano00d] Anonymous. Back matter. Journal of Applied Probability, 37 (3):??, September 2000. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3215626.

Anonymous:2000:BMd

[Ano00e] Anonymous. Back matter. Journal of Applied Probability, 37(4): ??, December 2000. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3215512.

Anonymous:2000:FMa

[Ano00f] Anonymous. Front matter. Journal of Applied Probability, 37 (1):??, March 2000. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3215654.

Anonymous:2000:FMb

[Ano00g] Anonymous. Front matter. Journal of Applied Probability, 37 (2):??, June 2000. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3215708.

Anonymous:2000:FMc

[Ano00h] Anonymous. Front matter. Journal of Applied Probability, 37 (3):??, September 2000. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3215598.

Anonymous:2000:FMd

[Ano00i] Anonymous. Front matter. Journal of Applied Probability, 37 (4):??, December 2000. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3215483.

Anonymous:2000:OSJ

[Ano00j] Anonymous. Obituary: Sidney Jesse Yakowitz. Journal of Applied Probability, 37(2):606–611, June 2000. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/3215734.

Anonymous:2000:VI

[Ano00k] Anonymous. Volume information. Journal of Applied Probability, 37(4):1188–1192, December 2000. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3215482.

Anonymous:2001:BMa

[Ano01a] Anonymous. Back matter. Journal of Applied Probability, 38(1): ??, March 2001. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3215763.

Anonymous:2001:BMb

[Ano01b] Anonymous. Back matter. Journal of Applied Probability, 38 (2):??, June 2001. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3215911.

Anonymous:2001:BMc

[Ano01c] Anonymous. Back matter. Journal of Applied Probability, 38 (3):??, September 2001. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3216136.

Anonymous:2001:BMd

[Ano01d] Anonymous. Back matter. Journal of Applied Probability, 38(4): ??, December 2001. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3215789.

Anonymous:2001:BM

[Ano01e] Anonymous. Back matter. Journal of Applied Probability, 38A: ??, 2001. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3215887.

Anonymous:2001:FMa

[Ano01f] Anonymous. Front matter. Journal of Applied Probability, 38 (1):??, March 2001. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3215736.

Anonymous:2001:FMb

[Ano01g] Anonymous. Front matter. Journal of Applied Probability, 38 (2):??, June 2001. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3215888.

Anonymous:2001:FMc

[Ano01h] Anonymous. Front matter. Journal of Applied Probability, 38 (3):??, September 2001. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3216114.

Anonymous:2001:FMd

[Ano01i] Anonymous. Front matter. Journal of Applied Probability, 38 (4):??, December 2001. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3215765.

Anonymous:2001:FM

[Ano01j] Anonymous. Front matter. Journal of Applied Probability, 38A:i–viii, 2001. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3215858.

Anonymous:2001:PDV

[Ano01k] Anonymous. Publications of David Vere-Jones. Journal of Applied Probability, 38A:ix–xvi, 2001. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3215860.

Anonymous:2001:VIa

[Ano01l] Anonymous. Volume information. Journal of Applied Probability, 38(4):1086–1089, December 2001. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3215764.

Anonymous:2001:VIb

[Ano01m] Anonymous. Volume information. Journal of Applied Probability, 38A:289–292, 2001. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3215857.

Anonymous:2002:BMa

[Ano02a] Anonymous. Back matter. Journal of Applied Probability, 39(1): ??, March 2002. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3215932.

Anonymous:2002:BMb

[Ano02b] Anonymous. Back matter. Journal of Applied Probability, 39 (2):??, June 2002. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3216113.

Anonymous:2002:BMc

[Ano02c] Anonymous. Back matter. Journal of Applied Probability, 39 (3):??, September 2002. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3216090.

Anonymous:2002:BMd

[Ano02d] Anonymous. Back matter. Journal of Applied Probability, 39(4): ??, December 2002. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3216015.

Anonymous:2002:CNS

[Ano02e] Anonymous. Correction: A note on stochastic comparisons of excess lifetimes of renewal processes. Journal of Applied Probability, 39(1):231, March 2002. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3215931. See [BOR01].

Anonymous:2002:FMa

[Ano02f] Anonymous. Front matter. Journal of Applied Probability, 39 (1):??, March 2002. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3215912.

Anonymous:2002:FMb

[Ano02g] Anonymous. Front matter. Journal of Applied Probability, 39 (2):??, June 2002. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3216091.

Anonymous:2002:FMc

[Ano02h] Anonymous. Front matter. Journal of Applied Probability, 39 (3):??, September 2002. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3216069.

Anonymous:2002:FMd

[Ano02i] Anonymous. Front matter. Journal of Applied Probability, 39 (4):??, December 2002. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3215995.

Anonymous:2002:VI

[Ano02j] Anonymous. Volume information. Journal of Applied Probability, 39(4):910–913, December 2002. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3215994.

Anonymous:2003:BMa

[Ano03a] Anonymous. Back matter. Journal of Applied Probability, 40(1): ??, March 2003. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3215856.

Anonymous:2003:BMb

[Ano03b] Anonymous. Back matter. Journal of Applied Probability, 40 (2):??, June 2003. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3215809.

Anonymous:2003:BMc

[Ano03c] Anonymous. Back matter. Journal of Applied Probability, 40 (3):??, September 2003. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3215960.

Anonymous:2003:BMd

[Ano03d] Anonymous. Back matter. Journal of Applied Probability, 40(4): ??, December 2003. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3216068.

Anonymous:2003:FMa

[Ano03e] Anonymous. Front matter. Journal of Applied Probability, 40 (1):??, March 2003. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3215837.

Anonymous:2003:FMb

[Ano03f] Anonymous. Front matter. Journal of Applied Probability, 40 (2):??, June 2003. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3215790.

Anonymous:2003:FMc

[Ano03g] Anonymous. Front matter. Journal of Applied Probability, 40 (3):??, September 2003. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3215933.

Anonymous:2003:FMd

[Ano03h] Anonymous. Front matter. Journal of Applied Probability, 40 (4):??, December 2003. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3216044.

Anonymous:2003:VI

[Ano03i] Anonymous. Volume information. Journal of Applied Probability, 40(4):1158–1161, December 2003. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3216043.

Anonymous:2004:BMa

[Ano04a] Anonymous. Back matter. Journal of Applied Probability, 41(1): ??, March 2004. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3215836.

Anonymous:2004:BMb

[Ano04b] Anonymous. Back matter. Journal of Applied Probability, 41 (2):??, June 2004. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3216042.

Anonymous:2004:BMc

[Ano04c] Anonymous. Back matter. Journal of Applied Probability, 41 (3):??, September 2004. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/4141365.

Anonymous:2004:BMd

[Ano04d] Anonymous. Back matter. Journal of Applied Probability, 41(4): ??, December 2004. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
4141401.

Anonymous:2004:BM

[Ano04e] Anonymous. Back matter. Journal of Applied Probability, 41A: 229–414, 2004. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3215993.

Anonymous:2004:CCH

[Ano04f] Anonymous. Christopher Charles Heyde, AM, DSc, FAA, FASSA. Journal of Applied Probability, 41A:vii–x, 2004. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/3215962.

Anonymous:2004:CDA

[Ano04g] Anonymous. Correction: [Distribution of the amount of genetic material from a chromosomal segment surviving to the following generation]. Journal of Applied Probability, 41(4):1248–1249, December 2004. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/

4141400. See [Ste04].

Anonymous:2004:FMa

[Ano04h] Anonymous. Front matter. Journal of Applied Probability, 41 (1):??, March 2004. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3215810.

Anonymous:2004:FMb

[Ano04i] Anonymous. Front matter. Journal of Applied Probability, 41 (2):??, June 2004. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3216016.

Anonymous:2004:FMc

[Ano04j] Anonymous. Front matter. Journal of Applied Probability, 41 (3):??, September 2004. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/4141340.

Anonymous:2004:FMd

[Ano04k] Anonymous. Front matter. Journal of Applied Probability, 41 (4):??, December 2004. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/4141367.

Anonymous:2004:FM

[Ano04l] Anonymous. Front matter. Journal of Applied Probability, 41A:i–189, 2004. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3215961.

Anonymous:2004:PCC

[Ano04m] Anonymous. Publications of Christopher Charles Heyde. Journal of Applied Probability, 41A:xi–xviii, 2004. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3215963.

Anonymous:2004:VI

[Ano04n] Anonymous. Volume information. Journal of Applied Probability, 41(4):1250–1254, December 2004. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/4141366.

Anonymous:2005:BMa

[Ano05a] Anonymous. Back matter. Journal of Applied Probability, 42(1): ??, March 2005. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
30040791.

Anonymous:2005:BMb

[Ano05b] Anonymous. Back matter. Journal of Applied Probability, 42 (2):??, June 2005. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
30040813.

Anonymous:2005:BMc

[Ano05c] Anonymous. Back matter. Journal of Applied Probability, 42 (3):??, September 2005. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/30040869.

Anonymous:2005:BMd

[Ano05d] Anonymous. Back matter. Journal of Applied Probability, 42(4): ??, December 2005. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
30040840.

Anonymous:2005:CTA

[Ano05e] Anonymous. Correction: Transient analysis of state-dependent queueing networks via cumulant functions. Journal of Applied Probability, 42(1):302, March 2005. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/30040790. See [MF01].

Anonymous:2005:FMa

[Ano05f] Anonymous. Front matter. Journal of Applied Probability, 42 (1):??, March 2005. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
30040764.

Anonymous:2005:FMb

[Ano05g] Anonymous. Front matter. Journal of Applied Probability, 42 (2):??, June 2005. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
30040793.

Anonymous:2005:FMc

[Ano05h] Anonymous. Front matter. Journal of Applied Probability, 42 (3):??, September 2005. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/30040842.

Anonymous:2005:FMd

[Ano05i] Anonymous. Front matter. Journal of Applied Probability, 42 (4):??, December 2005. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/30040816.

Anonymous:2005:VI

[Ano05j] Anonymous. Volume information. Journal of Applied Probability, 42(4):??, December 2005. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.

org/stable/30040815.

Anonymous:2006:BMa

[Ano06a] Anonymous. Back matter. Journal of Applied Probability, 43(1): ??, March 2006. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
27595730.

Anonymous:2006:BMb

[Ano06b] Anonymous. Back matter. Journal of Applied Probability, 43 (2):??, June 2006. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
27595756.

Anonymous:2006:BMc

[Ano06c] Anonymous. Back matter. Journal of Applied Probability, 43 (3):??, September 2006. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/27595784.

Anonymous:2006:BMd

[Ano06d] Anonymous. Back matter. Journal of Applied Probability, 43(4): ??, December 2006. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
27595815.

Anonymous:2006:CID

[Ano06e] Anonymous. Correction: Infinite divisibility of infinite sums of lower records: A simple proof. Journal of Applied Probability, 43 (2):602, June 2006. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/

27595755. See [Iks04].

Anonymous:2006:CJE

[Ano06f] Anonymous. Correction: Joint exceedances of the ARCH process. Journal of Applied Probability, 43(4):1206, December 2006. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/27595813. See

[GDP04].

Anonymous:2006:CSI

[Ano06g] Anonymous. Correction: Stationary-increment Student and variance-gamma processes. Journal of Applied Probability, 43 (4):1207, December 2006. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/27595814. See [FS06].

Anonymous:2006:FMa

[Ano06h] Anonymous. Front matter. Journal of Applied Probability, 43 (1):??, March 2006. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
27595705.

Anonymous:2006:FMb

[Ano06i] Anonymous. Front matter. Journal of Applied Probability, 43 (2):??, June 2006. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
27595732.

Anonymous:2006:FMc

[Ano06j] Anonymous. Front matter. Journal of Applied Probability, 43 (3):??, September 2006. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/27595758.

Anonymous:2006:FMd

[Ano06k] Anonymous. Front matter. Journal of Applied Probability, 43 (4):??, December 2006. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/27595787.

Anonymous:2006:VIa

[Ano06l] Anonymous. Volume information. Journal of Applied Probability, 43(1):??, March 2006. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/27595704.

Anonymous:2006:VIb

[Ano06m] Anonymous. Volume information. Journal of Applied Probability, 43(4):??, December 2006. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.

org/stable/27595786.

Anonymous:2007:BMa

[Ano07a] Anonymous. Back matter. Journal of Applied Probability, 44(1): ??, March 2007. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
27595838.

Anonymous:2007:BMb

[Ano07b] Anonymous. Back matter. Journal of Applied Probability, 44 (2):??, June 2007. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
27595864.

Anonymous:2007:BMc

[Ano07c] Anonymous. Back matter. Journal of Applied Probability, 44 (3):??, September 2007. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/27595887.

Anonymous:2007:BMd

[Ano07d] Anonymous. Back matter. Journal of Applied Probability, 44(4): ??, December 2007. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
27595915.

Anonymous:2007:CMA

[Ano07e] Anonymous. Correction: Markov additive processes and repeats in sequences. Journal of Applied Probability, 44(4):1122, December 2007. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/27595914.

See [Spo07].

Anonymous:2007:FMa

[Ano07f] Anonymous. Front matter. Journal of Applied Probability, 44 (1):??, March 2007. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
27595818.

Anonymous:2007:FMb

[Ano07g] Anonymous. Front matter. Journal of Applied Probability, 44 (2):??, June 2007. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
27595840.

Anonymous:2007:FMc

[Ano07h] Anonymous. Front matter. Journal of Applied Probability, 44 (3):??, September 2007. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/27595866.

Anonymous:2007:FMd

[Ano07i] Anonymous. Front matter. Journal of Applied Probability, 44 (4):??, December 2007. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/27595890.

Anonymous:2007:VIa

[Ano07j] Anonymous. Volume information. Journal of Applied Probability, 44(1):??, March 2007. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/27595817.

Anonymous:2007:VIb

[Ano07k] Anonymous. Volume information. Journal of Applied Probability, 44(4):??, December 2007. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.

org/stable/27595889.

Anonymous:2008:BMa

[Ano08a] Anonymous. Back matter. Journal of Applied Probability, 45(1): ??, March 2008. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
27595942.

Anonymous:2008:BMb

[Ano08b] Anonymous. Back matter. Journal of Applied Probability, 45 (2):??, June 2008. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
27595968.

Anonymous:2008:BMc

[Ano08c] Anonymous. Back matter. Journal of Applied Probability, 45 (3):??, September 2008. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/27595994.

Anonymous:2008:BMd

[Ano08d] Anonymous. Back matter. Journal of Applied Probability, 45(4): ??, December 2008. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
27596020.

Anonymous:2008:FMa

[Ano08e] Anonymous. Front matter. Journal of Applied Probability, 45 (1):??, March 2008. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
27595918.

Anonymous:2008:FMb

[Ano08f] Anonymous. Front matter. Journal of Applied Probability, 45 (2):??, June 2008. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
27595944.

Anonymous:2008:FMc

[Ano08g] Anonymous. Front matter. Journal of Applied Probability, 45 (3):??, September 2008. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/27595970.

Anonymous:2008:FMd

[Ano08h] Anonymous. Front matter. Journal of Applied Probability, 45 (4):??, December 2008. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/27595997.

Anonymous:2008:VIa

[Ano08i] Anonymous. Volume information. Journal of Applied Probability, 45(1):??, March 2008. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/27595917.

Anonymous:2008:VIb

[Ano08j] Anonymous. Volume information. Journal of Applied Probability, 45(4):??, December 2008. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.

org/stable/27595996.

Anonymous:2009:BMa

[Ano09a] Anonymous. Back matter. Journal of Applied Probability, 46(1): ??, March 2009. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
25662423.

Anonymous:2009:BMb

[Ano09b] Anonymous. Back matter. Journal of Applied Probability, 46 (2):??, June 2009. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
25662448.

Anonymous:2009:BMc

[Ano09c] Anonymous. Back matter. Journal of Applied Probability, 46 (3):??, September 2009. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/25662469.

Anonymous:2009:BMd

[Ano09d] Anonymous. Back matter. Journal of Applied Probability, 46(4): ??, December 2009. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
25662494.

Anonymous:2009:FMa

[Ano09e] Anonymous. Front matter. Journal of Applied Probability, 46 (1):??, March 2009. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
25662403.

Anonymous:2009:FMb

[Ano09f] Anonymous. Front matter. Journal of Applied Probability, 46 (2):??, June 2009. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
25662425.

Anonymous:2009:FMc

[Ano09g] Anonymous. Front matter. Journal of Applied Probability, 46 (3):??, September 2009. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/25662450.

Anonymous:2009:FMd

[Ano09h] Anonymous. Front matter. Journal of Applied Probability, 46 (4):??, December 2009. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/25662471.

Anonymous:2010:BMa

[Ano10a] Anonymous. Back matter. Journal of Applied Probability, 47(1): ??, March 2010. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
25662517.

Anonymous:2010:BMb

[Ano10b] Anonymous. Back matter. Journal of Applied Probability, 47 (2):??, June 2010. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
20778503.

Anonymous:2010:BMc

[Ano10c] Anonymous. Back matter. Journal of Applied Probability, 47 (3):??, September 2010. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/20778525.

Anonymous:2010:BMd

[Ano10d] Anonymous. Back matter. Journal of Applied Probability, 47(4): ??, December 2010. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
29777439.

Anonymous:2010:CNA

[Ano10e] Anonymous. Correction: Negative association does not imply log-concavity of the rank sequence. Journal of Applied Probability, 47(2):608–609, June 2010. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/20778502. See [Mar07].

Anonymous:2010:FMa

[Ano10f] Anonymous. Front matter. Journal of Applied Probability, 47 (1):??, March 2010. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
25662496.

Anonymous:2010:FMb

[Ano10g] Anonymous. Front matter. Journal of Applied Probability, 47 (2):??, June 2010. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
20778480.

Anonymous:2010:FMc

[Ano10h] Anonymous. Front matter. Journal of Applied Probability, 47 (3):??, September 2010. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/20778505.

Anonymous:2010:FMd

[Ano10i] Anonymous. Front matter. Journal of Applied Probability, 47 (4):??, December 2010. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/29777413.

Anonymous:2013:BAP

[Ano13] Anonymous. Journal of Applied Probability volume 50 (2013): Index. Journal of Applied Probability, 50(4):1213–1216, December 2013. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-

6072 (electronic). URL http://projecteuclid.org/euclid.
jap/1389370109.

Anonymous:2014:BAP

[Ano14] Anonymous. Journal of Applied Probability Volume 51 (2014): Index. Journal of Applied Probability, 51(4):1195–1198, December 2014. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-

6072 (electronic). URL http://projecteuclid.org/euclid.
jap/1421763335.

Anonymous:2015:JAP

[Ano15] Anonymous. Journal of Applied Probability volume 52 (2015): Index. Journal of Applied Probability, 52(4):1204–1208, December 2015. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-

6072 (electronic). URL http://projecteuclid.org/euclid.
jap/1450802765.

Anonymous:2017:JVIb

[Ano17a] Anonymous. JPR volume 54 issue 4 cover and back matter. Journal of Applied Probability, 54(4):b1–b6, December 2017. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-

6072 (electronic). URL https://www.cambridge.org/core/
journals/journal-of-applied-probability/article/jpr-
volume-54-issue-4-cover-and-back-matter/06B9B91128166057B16DCC3B8D31E6FA.

Anonymous:2017:JVIa

[Ano17b] Anonymous. JPR volume 54 issue 4 cover and front matter. Journal of Applied Probability, 54(4):f1–f2, December 2017. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-

6072 (electronic). URL https://www.cambridge.org/core/
journals/journal-of-applied-probability/article/jpr-
volume-54-issue-4-cover-and-front-matter/1DB41255027579D7A671866D66517CDE.

Anonymous:2018:JVIa

[Ano18a] Anonymous. JPR volume 55 issue 1 cover and front matter. Journal of Applied Probability, 55(1):f1–f4, March 2018. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-

6072 (electronic). URL https://www.cambridge.org/core/
journals/journal-of-applied-probability/article/jpr-
volume-55-issue-1-cover-and-front-matter/AA94C4742B1EFCF2009A9CBE547346A3.

Anonymous:2018:JVIc

[Ano18b] Anonymous. JPR volume 55 issue 3 cover and back matter. Journal of Applied Probability, 55(3):b1–b2, September 2018. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-

6072 (electronic). URL https://www.cambridge.org/core/
journals/journal-of-applied-probability/article/jpr-
volume-55-issue-3-cover-and-back-matter/8D111F6B73D9D051054F5B9F74997AA1.

Anonymous:2018:JVIb

[Ano18c] Anonymous. JPR volume 55 issue 3 cover and front matter. Journal of Applied Probability, 55(3):f1–f2, September 2018. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-

6072 (electronic). URL https://www.cambridge.org/core/
journals/journal-of-applied-probability/article/jpr-
volume-55-issue-3-cover-and-front-matter/75720950428FE14F2057EB704F29715A.

Anonymous:2019:JVIb

[Ano19a] Anonymous. JPR volume 56 issue 1 cover and back matter. Journal of Applied Probability, 56(1):b1–b4, March 2019. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-

6072 (electronic). URL https://www.cambridge.org/core/
journals/journal-of-applied-probability/article/jpr-
volume-56-issue-1-cover-and-back-matter/ADCF2FFF8A2D54285797FDF454939C91.

Anonymous:2019:JVIa

[Ano19b] Anonymous. JPR volume 56 issue 1 cover and front matter. Journal of Applied Probability, 56(1):f1–f4, March 2019. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-

6072 (electronic). URL https://www.cambridge.org/core/
journals/journal-of-applied-probability/article/jpr-
volume-56-issue-1-cover-and-front-matter/9952A19EA10F8ED53A6FBB52BF06B09C.

Anonymous:2019:JVId

[Ano19c] Anonymous. JPR volume 56 issue 2 cover and back matter. Journal of Applied Probability, 56(2):b1–b2, June 2019. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-

6072 (electronic). URL https://www.cambridge.org/core/
journals/journal-of-applied-probability/article/jpr-
volume-56-issue-2-cover-and-back-matter/1F45105249F9FB068CE7A53258B86638.

Anonymous:2019:JVIc

[Ano19d] Anonymous. JPR volume 56 issue 2 cover and front matter. Journal of Applied Probability, 56(2):f1–f2, June 2019. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-

6072 (electronic). URL https://www.cambridge.org/core/
journals/journal-of-applied-probability/article/jpr-
volume-56-issue-2-cover-and-front-matter/DEA9714C8CBBFCD8809D0CE0B359E97D.

Anonymous:2019:JVIf

[Ano19e] Anonymous. JPR volume 56 issue 3 cover and back matter. Journal of Applied Probability, 56(3):b1–b4, September 2019. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-

6072 (electronic). URL https://www.cambridge.org/core/
journals/journal-of-applied-probability/article/jpr-
volume-56-issue-3-cover-and-back-matter/E776E035D1A1E22553651704C2569C05.

Anonymous:2019:JVIe

[Ano19f] Anonymous. JPR volume 56 issue 3 cover and front matter. Journal of Applied Probability, 56(3):f1–f2, September 2019. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-

6072 (electronic). URL https://www.cambridge.org/core/
journals/journal-of-applied-probability/article/jpr-
volume-56-issue-3-cover-and-front-matter/2A22125447614BC56619A9098EC65006.

Anonymous:2019:JVIh

[Ano19g] Anonymous. JPR volume 56 issue 4 cover and back matter. Journal of Applied Probability, 56(4):b1–b2, December 2019. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-

6072 (electronic). URL https://www.cambridge.org/core/
journals/journal-of-applied-probability/article/jpr-
volume-56-issue-4-cover-and-back-matter/523844D6842C96F693AAA83642A5C5BB.

Anonymous:2019:JVIg

[Ano19h] Anonymous. JPR volume 56 issue 4 cover and front matter. Journal of Applied Probability, 56(4):f1–f2, December 2019. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-

6072 (electronic). URL https://www.cambridge.org/core/
journals/journal-of-applied-probability/article/jpr-
volume-56-issue-4-cover-and-front-matter/23C0B3C347C52CD5EE2A4E5E92B946AD.

Anonymous:2020:JVIb

[Ano20a] Anonymous. JPR volume 57 issue 2 cover and back matter. Journal of Applied Probability, 57(2):b1–b2, June 2020. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-

6072 (electronic). URL https://www.cambridge.org/core/
journals/journal-of-applied-probability/article/jpr-
volume-57-issue-2-cover-and-back-matter/EA5084C82CB86CFEA4FC583C534AA7ED.

Anonymous:2020:JVIa

[Ano20b] Anonymous. JPR volume 57 issue 2 cover and front matter. Journal of Applied Probability, 57(2):f1–f2, June 2020. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-

6072 (electronic). URL https://www.cambridge.org/core/
journals/journal-of-applied-probability/article/jpr-
volume-57-issue-2-cover-and-front-matter/67CA0A2791A1CDB03BB4441B1BBC1E0D.

Anonymous:2020:JVId

[Ano20c] Anonymous. JPR volume 57 issue 3 cover and back matter. Journal of Applied Probability, 57(3):b1–b2, September 2020. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-

6072 (electronic). URL https://www.cambridge.org/core/
journals/journal-of-applied-probability/article/jpr-
volume-57-issue-3-cover-and-back-matter/620980C9ACAE9D2210F3C905D4803728.

Anonymous:2020:JVIc

[Ano20d] Anonymous. JPR volume 57 issue 3 cover and front matter. Journal of Applied Probability, 57(3):f1–f2, September 2020. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-

6072 (electronic). URL https://www.cambridge.org/core/
journals/journal-of-applied-probability/article/jpr-
volume-57-issue-3-cover-and-front-matter/A3ED680C7444B1C49471AC1FD52768AE.

Anonymous:2020:JVIf

[Ano20e] Anonymous. JPR volume 57 issue 4 cover and back matter. Journal of Applied Probability, 57(4):b1–b2, December 2020. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-

6072 (electronic). URL https://www.cambridge.org/core/
journals/journal-of-applied-probability/article/jpr-
volume-57-issue-4-cover-and-back-matter/D324E06B12344742E46307F4EC8584C5.

Anonymous:2020:JVIe

[Ano20f] Anonymous. JPR volume 57 issue 4 cover and front matter. Journal of Applied Probability, 57(4):f1–f2, December 2020. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-

6072 (electronic). URL https://www.cambridge.org/core/
journals/journal-of-applied-probability/article/jpr-
volume-57-issue-4-cover-and-front-matter/211C32F328ACD270E9FFF989DB906DB5.

Anonymous:2021:JVIb

[Ano21a] Anonymous. JPR volume 58 issue 1 cover and back matter. Journal of Applied Probability, 58(1):b1–b2, March 2021. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-

6072 (electronic). URL https://www.cambridge.org/core/
journals/journal-of-applied-probability/article/jpr-
volume-58-issue-1-cover-and-back-matter/B0A643DFA750B7D70DF0776DB50D4EF5.

Anonymous:2021:JVIa

[Ano21b] Anonymous. JPR volume 58 issue 1 cover and front matter. Journal of Applied Probability, 58(1):f1–f2, March 2021. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-

6072 (electronic). URL https://www.cambridge.org/core/
journals/journal-of-applied-probability/article/jpr-
volume-58-issue-1-cover-and-front-matter/9EF7F6DEEE98143E9ED2E86936C73CD4.

Anonymous:2021:JVId

[Ano21c] Anonymous. JPR volume 58 issue 2 cover and back matter. Journal of Applied Probability, 58(2):b1–b2, June 2021. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-

6072 (electronic). URL https://www.cambridge.org/core/
journals/journal-of-applied-probability/article/jpr-
volume-58-issue-2-cover-and-back-matter/0DDE5A95CE669E038842EB5094B08741.

Anonymous:2021:JVIc

[Ano21d] Anonymous. JPR volume 58 issue 2 cover and front matter. Journal of Applied Probability, 58(2):f1–f2, June 2021. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-

6072 (electronic). URL https://www.cambridge.org/core/
journals/journal-of-applied-probability/article/jpr-
volume-58-issue-2-cover-and-front-matter/537F403F938DD5605AFE7803C7DAA7B3.

Anonymous:2021:JVIf

[Ano21e] Anonymous. JPR volume 58 issue 3 cover and back matter. Journal of Applied Probability, 58(3):b1–b2, September 2021. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-

6072 (electronic). URL https://www.cambridge.org/core/
journals/journal-of-applied-probability/article/jpr-
volume-58-issue-3-cover-and-back-matter/6D78386A77A8C44428323B62FC8BABF8.

Anonymous:2021:JVIe

[Ano21f] Anonymous. JPR volume 58 issue 3 cover and front matter. Journal of Applied Probability, 58(3):f1–f2, September 2021. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-

6072 (electronic). URL https://www.cambridge.org/core/
journals/journal-of-applied-probability/article/jpr-
volume-58-issue-3-cover-and-front-matter/6777C1C9133B4111E94653E5872C2AFB.

Anonymous:2021:JVIh

[Ano21g] Anonymous. JPR volume 58 issue 4 cover and back matter. Journal of Applied Probability, 58(4):b1–b2, December 2021. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-

6072 (electronic). URL https://www.cambridge.org/core/
journals/journal-of-applied-probability/article/jpr-
volume-58-issue-4-cover-and-back-matter/F5EBEDF7F6F1D9F5D216947B9B16669D.

Anonymous:2021:JVIg

[Ano21h] Anonymous. JPR volume 58 issue 4 cover and front matter. Journal of Applied Probability, 58(4):f1–f2, December 2021. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-

6072 (electronic). URL https://www.cambridge.org/core/
journals/journal-of-applied-probability/article/jpr-
volume-58-issue-4-cover-and-front-matter/1EE0215807F08E0F3B8B90613869AF5F.

Anonymous:2022:APT

[Ano22a] Anonymous. Applied Probability Trust prizes 2021. Journal of Applied Probability, 59(1):297, March 2022. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL https:/

/www.cambridge.org/core/journals/journal-of-applied-
probability/article/applied-probability-trust-prizes-
2021/43EDDAD3F041519F84764A3264D65572.

Anonymous:2022:JVIb

[Ano22b] Anonymous. JPR volume 59 issue 1 cover and back matter. Journal of Applied Probability, 59(1):b1–b2, March 2022. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-

6072 (electronic). URL https://www.cambridge.org/core/
journals/journal-of-applied-probability/article/jpr-
volume-59-issue-1-cover-and-back-matter/D28F225786FBF99B216BFAD07365D0C0.

Anonymous:2022:JVIa

[Ano22c] Anonymous. JPR volume 59 issue 1 cover and front matter. Journal of Applied Probability, 59(1):f1–f2, March 2022. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-

6072 (electronic). URL https://www.cambridge.org/core/
journals/journal-of-applied-probability/article/jpr-
volume-59-issue-1-cover-and-front-matter/853D1CDD5EBFC6FA6C7296E43C225C13.

Anonymous:2022:JVId

[Ano22d] Anonymous. JPR volume 59 issue 2 cover and back matter. Journal of Applied Probability, 59(2):b1–b2, June 2022. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-

6072 (electronic). URL https://www.cambridge.org/core/
journals/journal-of-applied-probability/article/jpr-
volume-59-issue-2-cover-and-back-matter/4289D31F8B5C14AE6788F7528C1E0F24.

Anonymous:2022:JVIc

[Ano22e] Anonymous. JPR volume 59 issue 2 cover and front matter. Journal of Applied Probability, 59(2):f1–f2, June 2022. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-

6072 (electronic). URL https://www.cambridge.org/core/
journals/journal-of-applied-probability/article/jpr-
volume-59-issue-2-cover-and-front-matter/87348F2E2243440D7372EE80CCC40983.

Anonymous:2022:JVIf

[Ano22f] Anonymous. JPR volume 59 issue 3 cover and back matter. Journal of Applied Probability, 59(3):b1–b2, September 2022. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-

6072 (electronic). URL https://www.cambridge.org/core/
journals/journal-of-applied-probability/article/jpr-
volume-59-issue-3-cover-and-back-matter/DD4890B1943995F22BAE0D232EF97F60.

Anonymous:2022:JVIe

[Ano22g] Anonymous. JPR volume 59 issue 3 cover and front matter. Journal of Applied Probability, 59(3):f1–f2, September 2022. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-

6072 (electronic). URL https://www.cambridge.org/core/
journals/journal-of-applied-probability/article/jpr-
volume-59-issue-3-cover-and-front-matter/2F27EC38DF36243A6FC4D2D858C86502.

Anonymous:2022:JVIh

[Ano22h] Anonymous. JPR volume 59 issue 4 cover and back matter. Journal of Applied Probability, 59(4):b1–b2, December 2022. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-

6072 (electronic). URL https://www.cambridge.org/core/
journals/journal-of-applied-probability/article/jpr-
volume-59-issue-4-cover-and-back-matter/12DA5EDBB06B4310C21D0FA30E12B218.

Anonymous:2022:JVIg

[Ano22i] Anonymous. JPR volume 59 issue 4 cover and front matter. Journal of Applied Probability, 59(4):f1–f2, December 2022. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-

6072 (electronic). URL https://www.cambridge.org/core/
journals/journal-of-applied-probability/article/jpr-
volume-59-issue-4-cover-and-front-matter/1B66D4722CD74265F31AFB1DA345A0B0.

Antonelli:1978:AFT

[Ant78a] Peter L. Antonelli. An asymptotic formula for the transition density of random genetic drift. Journal of Applied Probability, 15(1):184–186, March 1978. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/ stable/3213249. See correction [Ano78e].

Arjas:1978:ULT

[ANT78b] Elja Arjas, Esa Nummelin, and Richard L. Tweedie. Uniform limit theorems for non-singular renewal and Markov renewal processes. Journal of Applied Probability, 15(1):112–125, March 1978. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3213241.

Antzoulakos:2001:WTP

[Ant01] Demetrios L. Antzoulakos. Waiting times for patterns in a sequence of multistate trials. Journal of Applied Probability, 38 (2):508–518, June 2001. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3215903.

Ajazi:2017:PTR

[ANT17] Fioralba Ajazi, George M. Napolitano, and Tatyana Turova. Phase transition in random distance graphs on the torus. Journal of Applied Probability, 54(4):1278–1294, December 2017. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL https://www.cambridge.org/core/journals/journal-
of-applied-probability/article/phase-transition-in-random-
distance-graphs-on-the-torus/3FCA30272C362FE9E2511F1EDE807DBF.

Anderson:1982:CCO

[ANW82] E. J. Anderson, P. Nash, and R. R. Weber. A counterexample to a conjecture on optimal list ordering. Journal of Applied Probability, 19(3):730–732, September 1982. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/3213536.

Adler:2003:CCP

[AOR03] Ilan Adler, Shmuel Oren, and Sheldon M. Ross. The couponcollector’s problem revisited. Journal of Applied Probability, 40 (2):513–518, June 2003. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3215807.

Athreya:1986:MPH

[AP86] Krishna B. Athreya and Sastry G. Pantula. Mixing properties of Harris chains and autoregressive processes. Journal of Applied Probability, 23(4):880–892, December 1986. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/3214462.

Alzaid:1994:MID

[AP94] Abdulhamid A. Alzaid and Frank Proschan. Max-infinite divisibility and multivariate total positivity. Journal of Applied Probability, 31(3):721–730, September 1994. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/3215150.

Appel:1994:PPT

[App94] Martin J. B. Appel. AB percolation on plane triangulations is unimodal. Journal of Applied Probability, 31(1):193–204, March 1994. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3215246.

Athreya:1974:SAD

[APS74] K. B. Athreya, P. R. Parthasarathy, and G. Sankaranarayanan. Supercritical age-dependent branching processes with immigration. Journal of Applied Probability, 11(4):695–702, December 1974. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3212553.

Ayhan:2004:CQN

[APS04] H. Ayhan, Z. Palmowski, and S. Schlegel. Cyclic queueing networks with subexponential service times. Journal of Applied Probability, 41(3):791–801, September 2004. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/4141354.

Avrachenkov:2011:AFO

[APZ11] Konstantin Avrachenkov, Alexey Piunovskiy, and Yi Zhang. Asymptotic fluid optimality and efficiency of the tracking policy for bandwidth-sharing networks. Journal of Applied Probability, 48(1):90–113, March 2011. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/29777448.

Avrachenkov:2013:MPR

[APZ13] Konstantin Avrachenkov, Alexey Piunovskiy, and Yi Zhang. Markov processes with restart. Journal of Applied Probability, 50(4):960–968, December 2013. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http:/

/projecteuclid.org/euclid.jap/1389370093.

Ahsanullah:1972:CED

[AR72] M. Ahsanullah and M. Rahman. A characterization of the exponential distribution. Journal of Applied Probability, 9(2):457– 461, June 1972. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3212817.

Andersen:1982:FCC

[AR82] Søren Andersen and Mats Rudemo. Filtering counts of cars entering and leaving a parking place. Journal of Applied Probability, 19(1):64–71, March 1982. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3213916.

Arnold:1989:ALP

[AR89] Barry C. Arnold and C. A. Robertson. Autoregressive logistic processes. Journal of Applied Probability, 26(3):524–531, September 1989. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3214410.

Adler:2001:CSC

[AR01a] Ilan Adler and Sheldon M. Ross. The coupon subset collection problem. Journal of Applied Probability, 38(3):737–746, September 2001. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3216125.

Asmussen:2001:ASJ

[AR01b] Søren Asmussen and Jan Rosiński. Approximations of small jumps of Lévy processes with a view towards simulation. Journal of Applied Probability, 38(2):482–493, June 2001. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3215901.

Ahn:2005:MAM

[AR05a] Hyun-Soo Ahn and Rhonda Righter. Multi-actor Markov decision processes. Journal of Applied Probability, 42(1):15–26, March 2005. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
30040766.

Ahn:2005:EAT

[AR05b] Soohan Ahn and V. Ramaswami. Efficient algorithms for transient analysis of stochastic fluid flow models. Journal of Applied Probability, 42(2):531–549, June 2005. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/30040808.

Aristoff:2013:ESL

[AR13] David Aristoff and Charles Radin. Emergent structures in large networks. Journal of Applied Probability, 50(3):883– 888, September 2013. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://projecteuclid.

org/euclid.jap/1378401243.

Athreya:2015:EIR

[AR15] Krishna B. Athreya and Vivekananda Roy. Estimation of integrals with respect to infinite measures using regenerative sequences. Journal of Applied Probability, 52(4):1133–1145, December 2015. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-

6072 (electronic). URL http://projecteuclid.org/euclid.
jap/1450802757.

Arato:1994:SKT

[Ara94] Mátyás Arató. On the solution of Kac-type partial differential equations. Journal of Applied Probability, 31A:311–324, 1994. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/3214964.

Arai:2004:MMM

[Ara04] Takuji Arai. Minimal martingale measures for jump diffusion processes. Journal of Applied Probability, 41(1):263–270, March 2004. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3215830.

Aristoff:2014:PHD

[Ari14] D. Aristoff. Percolation of hard disks. Journal of Applied Probability, 51(1):235–246, March 2014. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http:/

/projecteuclid.org/euclid.jap/1395771426.

Aristoff:2022:ETW

[Ari22] David Aristoff. An ergodic theorem for the weighted ensemble method. Journal of Applied Probability, 59(1):152–166, March 2022. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-

6072 (electronic). URL https://www.cambridge.org/core/
journals/journal-of-applied-probability/article/an-ergodic-
theorem-for-the-weighted-ensemble-method/C032C0720A9295B562CCEC138AB147CB.

Angulo:1998:SEA

[ARM98] J. M. Angulo and M. D. Ruiz-Medina. A series expansion approach to the inverse problem. Journal of Applied Probability, 35(2):371–382, June 1998. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3215691.

Arnold:1968:MRD

[Arn68] Barry C. Arnold. A modification of a result due to Moran. Journal of Applied Probability, 5(1):220–223, April 1968. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/3212090.

Arnold:1972:WTU

[Arn72] Barry C. Arnold. The waiting time until first duplication. Journal of Applied Probability, 9(4):841–846, December 1972. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/3212620.

Arnold:1975:MED

[Arn75] Barry C. Arnold. Multivariate exponential distributions based on hierarchical successive damage. Journal of Applied Probability, 12(1):142–147, March 1975. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3212416.

Arnold:1980:TCG

[Arn80] Barry C. Arnold. Two characterizations of the geometric distribution. Journal of Applied Probability, 17(2):570–573, June 1980. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3213048.

Alves:2021:CPA

[ARS21] Caio Alves, Rodrigo Ribeiro, and Rémy Sanchis. Clustering in preferential attachment random graphs with edge-step. Journal of Applied Probability, 58(4):890–908, December 2021. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL https://www.cambridge.org/core/journals/journal-
of-applied-probability/article/clustering-in-preferential-
attachment-random-graphs-with-edgestep/1BA037BE65D328F0923DBD5062E210D0.

Asmussen:1994:RRE

[ARW94] Søren Asmussen, Reuven Y. Rubinstein, and Chia-Li Wang. Regenerative rare events simulation via likelihood ratios. Journal of Applied Probability, 31(3):797–815, September 1994. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3215157.

Arya:1972:SNS

[Ary72] Krishan Lall Arya. Study of a network of serial and non-serial servers with phase type service and finite queueing space. Journal of Applied Probability, 9(1):198–201, March 1972. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3212649.

Albrecher:2008:LIR

[ARZ08] Hansjörg Albrecher, Jean-François Renaud, and Xiaowen Zhou. A Lévy insurance risk process with tax. Journal of Applied Probability, 45(2):363–375, June 2008. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/27595950.

Autar:1975:ICT

[AS75] Ram Autar and Raminder Singh Soni. Inaccuracy and a coding theorem. Journal of Applied Probability, 12(4):845–851, December 1975. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3212738.

Alex:1989:IPQ

[AS89] Michael Alex and Josef Steinebach. Invariance principles in queueing theory. Journal of Applied Probability, 26(4):845–857, December 1989. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3214389.

Andjel:1996:CCT

[AS96] Enrique Andjel and Rinaldo Schinazi. A complete convergence theorem for an epidemic model. Journal of Applied Probability, 33(3):741–748, September 1996. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.

org/stable/3215355.

Akar:2004:IFB

[AS04] Nail Akar and Khosrow Sohraby. Infinite- and finite-buffer Markov fluid queues: A unified analysis. Journal of Applied Probability, 41(2):557–569, June 2004. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3216036.

Applebaum:2009:ASS

[AS09] David Applebaum and Michailina Siakalli. Asymptotic stability of stochastic differential equations driven by Lévy noise. Journal of Applied Probability, 46(4):1116–1129, December 2009. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/25662484.

Appuhamillage:2012:FPT

[AS12] Thilanka Appuhamillage and Daniel Sheldon. First passage time of skew Brownian motion. Journal of Applied Probability, 49(3): 685–696, September 2012. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/41713798.

Athreya:2016:GWP

[AS16] K. B. Athreya and H.-J. Schuh. A Galton–Watson process with a threshold. Journal of Applied Probability, 53(2):614–621, June 2016. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://projecteuclid.org/euclid.jap/
1466172880.

Arkin:2017:TSO

[AS17] Vadim Arkin and Alexander Slastnikov. On the threshold strategies in optimal stopping problems for diffusion processes. Journal of Applied Probability, 54(3):963–969, September 2017. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL https://www.cambridge.org/core/journals/journal-
of-applied-probability/article/on-the-threshold-strategies-
in-optimal-stopping-problems-for-diffusion-processes/
144FF966851C5020C987498B5916E697.

Assaf:1994:DSM

[ASB94] David Assaf and Ariela Sharlin-Bilitzky. Dynamic search for a moving target. Journal of Applied Probability, 31(2):438– 457, June 1994. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3215037.

Amini-Seresht:2021:SPG

[ASB21] Ebrahim Amini-Seresht and Narayanaswamy Balakrishnan. Stochastic properties of generalized finite mixture models with dependent components. Journal of Applied Probability, 58 (3):794–804, September 2021. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL https:/

/www.cambridge.org/core/journals/journal-of-applied-
probability/article/stochastic-properties-of-generalized-
finite-mixture-models-with-dependent-components/40B484106EC5FBAC23ED360DA94434FD.

Assaf:1998:OMS

[ASC98] David Assaf and Ester Samuel-Cahn. Optimal multivariate stopping rules. Journal of Applied Probability, 35(3):693–706, September 1998. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3215644.

Assaf:2000:SRP

[ASC00] David Assaf and Ester Samuel-Cahn. Simple ratio prophet inequalities for a mortal with multiple choices. Journal of Applied Probability, 37(4):1084–1091, December 2000. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3215496.

Amini-Seresht:2020:SNR

[ASKK20] Ebrahim Amini-Seresht, Baha-Eldin Khaledi, and Subhash Kochar. Some new results on stochastic comparisons of coherent systems using signatures. Journal of Applied Probability, 57(1):156–173, March 2020. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL https:/

/www.cambridge.org/core/journals/journal-of-applied-
probability/article/some-new-results-on-stochastic-comparisons-
of-coherent-systems-using-signatures/261FBB21E553D38297CABBD27CE56E27.

Asmussen:2009:ISF

[Asm09] Søren Asmussen. Importance sampling for failure probabilities in computing and data transmission. Journal of Applied Probability, 46(3):768–790, September 2009. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/25662460.

Assaf:1983:EPS

[Ass83] David Assaf. Extreme-point solutions in Markov decision processes. Journal of Applied Probability, 20(4):835–842, December 1983. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3213594.

Assaf:1985:FPT

[ASS85] David Assaf, Moshe Shaked, and J. George Shanthikumar. Firstpassage times with PF_{r} densities. Journal of Applied Probability, 22(1):185–196, March 1985. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/ stable/3213758. See correction [Ano87g].

Amini-Seresht:2018:SCC

[ASZB18] Ebrahim Amini-Seresht, Yiying Zhang, and Narayanaswamy Balakrishnan. Stochastic comparisons of coherent systems under different random environments. Journal of Applied Probability, 55(2):459–472, June 2018. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL https:/

/www.cambridge.org/core/journals/journal-of-applied-
probability/article/stochastic-comparisons-of-coherent-
systems-under-different-random-environments/CB9D98B1F7622DE74D596D90F7D8AD5F.

Ailam:1970:OOF

[AT70] Gedalia Ailam and Mahabanoo N. Tata. One-to-one functions on the positive integers. Journal of Applied Probability, 7(2): 505–507, August 1970. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3211991.

Ahmed:1974:OFC

[AT74] N. U. Ahmed and K. L. Teo. Optimal feedback control of stochastic McShane differential systems. Journal of Applied Probability, 11(2):302–309, June 1974. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3212751.

Ammann:1979:CDO

[AT79] Larry P. Ammann and Peter F. Thall. Count distributions, orderliness and invariance of Poisson cluster processes. Journal of Applied Probability, 16(2):261–273, June 1979. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3212895.

Asmussen:1987:MCA

[AT87] Søren Asmussen and Hermann Thorisson. A Markov chain approach to periodic queues. Journal of Applied Probability, 24 (1):215–225, March 1987. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3214072.

Anderson:1991:JLD

[AT91a] C. W. Anderson and K. F. Turkman. The joint limiting distribution of sums and maxima of stationary sequences. Journal of Applied Probability, 28(1):33–44, March 1991. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/3214738.

Anderson:1991:SMS

[AT91b] C. W. Anderson and K. F. Turkman. Sums and maxima in stationary sequences. Journal of Applied Probability, 28(3):715–716, September 1991. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3214506.

Asmussen:1996:CRM

[AT96] Søren Asmussen and Jozef L. Teugels. Convergence rates for M/G/1 queues and ruin problems with heavy tails. Journal of Applied Probability, 33(4):1181–1190, December 1996. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3214995.

Asmussen:1998:SPN

[AT98] Søren Asmussen and Tatyana S. Turova. Stationarity properties of neural networks. Journal of Applied Probability, 35(4): 783–794, December 1998. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3215515.

Albrecher:2006:EBP

[AT06] Hansjörg Albrecher and Jef L. Teugels. Exponential behavior in the presence of dependence in risk theory. Journal of Applied Probability, 43(1):257–273, March 2006. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http:/

/www.jstor.org/stable/27595725.

Athreya:1971:NFE

[Ath71] Krishna B. Athreya. A note on a functional equation arising in Galton–Watson branching processes. Journal of Applied Probability, 8(3):589–598, September 1971. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3212181.

Athreya:1988:HTD

[Ath88] K. B. Athreya. Hitting times of decreasing sets for regenerative processes. Journal of Applied Probability, 25(1):89–96, March 1988. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3214236.

Athreya:1992:MND

[Ath92] K. B. Athreya. On measures of non-degeneracy. Journal of Applied Probability, 29(3):733–739, September 1992. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3214908.

Athreya:2012:CCS

[Ath12] K. B. Athreya. Coalescence in critical and subcritical Galton– Watson branching processes. Journal of Applied Probability, 49 (3):627–638, September 2012. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.

org/stable/41713794.

Attia:1987:CFD

[Att87] F. A. Attia. The control of a finite dam with penalty cost function: Markov input rate. Journal of Applied Probability, 24 (2):457–465, June 1987. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3214269.

Attia:1989:ROM

[Att89] F. A. Attia. Resolvent operators of Markov processes and their applications in the control of a finite dam. Journal of Applied Probability, 26(2):314–324, June 1989. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3214038.

Attia:2019:MAC

[Att19] Najmeddine Attia. On the multifractal analysis of the covering number on the Galton–Watson tree. Journal of Applied Probability, 56(1):265–281, March 2019. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL https:/

/www.cambridge.org/core/journals/journal-of-applied-
probability/article/on-the-multifractal-analysis-of-
the-covering-number-on-the-galtonwatson-tree/523EB4F3942F195CEE3235909619013D.

Autar:1975:CII

[Aut75] Ram Autar. On a characterization of information improvement. Journal of Applied Probability, 12(2):407–411, June 1975. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/3212459.

Aven:1985:ULB

[Ave85] Terje Aven. Upper (lower) bounds on the mean of the maximum (minimum) of a number of random variables. Journal of Applied Probability, 22(3):723–728, September 1985. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/3213876.

Aven:2009:OTI

[Ave09] Terje Aven. Optimal test interval for a monotone safety system. Journal of Applied Probability, 46(2):330–341, June 2009. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/25662427.

Avni:1979:WLS

[Avn79] Haim Avni. Weak limit for the stationary distribution of the infinite-alleles model. Journal of Applied Probability, 16(3):459– 472, September 1979. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3213076.

Anderson:1990:RPD

[AW90] E. J. Anderson and R. R. Weber. The rendezvous problem on discrete locations. Journal of Applied Probability, 27(4):839–851, December 1990. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3214827.

Appel:1993:PBD

[AW93] Martin J. B. Appel and John C. Wierman. AB percolation on bond-decorated graphs. Journal of Applied Probability, 30 (1):153–166, March 1993. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3214628.

Atchade:2015:CRS

[AW15] Yves Atchadé and Yizao Wang. On the convergence rates of some adaptive Markov chain Monte Carlo algorithms. Journal of Applied Probability, 52(3):811–825, September 2015. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://projecteuclid.org/euclid.jap/1445543848.

Alili:2021:FTP

[AW21] Larbi Alili and David Woodford. On the finiteness and tails of perpetuities under a Lamperti–Kiu MAP. Journal of Applied Probability, 58(4):1086–1113, December 2021. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL https://www.cambridge.org/core/journals/journal-
of-applied-probability/article/on-the-finiteness-and-
tails-of-perpetuities-under-a-lampertikiu-map/8211100C523DF0E609F0EDB4948093C5.

Ayesta:2007:UCL

[Aye07] Urtzi Ayesta. A unifying conservation law for single-server queues. Journal of Applied Probability, 44(4):1078–1087, December 2007. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/27595909.

Ashrafi:2018:RMC

[AZA18] S. Ashrafi, S. Zarezadeh, and M. Asadi. Reliability modeling of coherent systems with shared components based on sequential order statistics. Journal of Applied Probability, 55(3):845–861, September 2018. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL https:/

/www.cambridge.org/core/journals/journal-of-applied-
probability/article/reliability-modeling-of-coherent-
systems-with-shared-components-based-on-sequential-order-
statistics/34C7E8A0D6C626FD55C93F932FE04E9B.

Bhattacharjee:2000:PRL

[BAAB00] M. C. Bhattacharjee, A. M. Abouammoh, A. N. Ahmed, and A. M. Barry. Preservation results for life distributions based on comparisons with asymptotic remaining life under replacements. Journal of Applied Probability, 37(4):999–1009, December 2000. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/3215490.

Baccelli:1986:EMW

[Bac86] François Baccelli. Exponential martingales and Wald’s formulas for two-queue networks. Journal of Applied Probability, 23(3): 812–819, September 1986. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3214018.

Baccelli:1991:SOR

[Bac91] François Baccelli. Stochastic ordering of random processes with an imbedded point process. Journal of Applied Probability, 28 (3):553–567, September 1991. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.

org/stable/3214491.

Bach:2001:RBE

[Bac01] Eric Bach. Random bisection and evolutionary walks. Journal of Applied Probability, 38(2):582–596, June 2001. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3215908.

Badia:2011:HRP

[Bad11] F. G. Badia. Hazard rate properties of a general counting process stopped at an independent random time. Journal of Applied Probability, 48(1):56–67, March 2011. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/29777446.

Bagley:1982:APS

[Bag82] J. H. Bagley. Asymptotic properties of subcritical Galton– Watson processes. Journal of Applied Probability, 19(3):510–517, September 1982. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3213510.

Bagley:1986:ASC

[Bag86a] J. H. Bagley. On the almost sure convergence of controlled branching processes. Journal of Applied Probability, 23(3):827– 831, September 1986. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3214020.

Bagley:1986:APS

[Bag86b] J. H. Bagley. On the asymptotic properties of a supercritical bisexual branching process. Journal of Applied Probability, 23 (3):820–826, September 1986. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.

org/stable/3214019.

Bai:2020:RHP

[Bai20] Shuyang Bai. Representations of Hermite processes using local time of intersecting stationary stable regenerative sets. Journal of Applied Probability, 57(4):1234–1251, December 2020. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL https://www.cambridge.org/core/journals/journal-
of-applied-probability/article/representations-of-hermite-
processes-using-local-time-of-intersecting-stationary-
stable-regenerative-sets/3135D07C01DB4E9DE93DCA4976F1A448.

Balmer:1972:SSQ

[Bal72] D. W. Balmer. A single server queue in discrete time with customers served in random order. Journal of Applied Probability, 9(4):862–867, December 1972. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.

org/stable/3212624.

Balmer:1975:QDP

[Bal75] D. W. Balmer. On a quickest detection problem with costly information. Journal of Applied Probability, 12(1):87–97, March 1975. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3212410.

Balmer:1976:QDP

[Bal76] D. W. Balmer. On a quickest detection problem with variable monitoring. Journal of Applied Probability, 13(4):760–767, December 1976. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3212531.

Balagopal:1978:DDD

[Bal78] K. Balagopal. Dryness of discrete dams: Comments on a paper by Tin and Phatarfod. Journal of Applied Probability, 15(4): 858–864, December 1978. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3213442. See [TP76].

Balagopal:1979:SLT

[Bal79] K. Balagopal. Some limit theorems for the general semi-Markov storage model. Journal of Applied Probability, 16(3):607–617, September 1979. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3213088.

Ball:1983:TBE

[Bal83a] Frank Ball. The threshold behaviour of epidemic models. Journal of Applied Probability, 20(2):227–241, June 1983. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3213797.

Ball:1983:TTR

[Bal83b] Frank Ball. A threshold theorem for the Reed–Frost chainbinomial epidemic. Journal of Applied Probability, 20(1):153– 157, March 1983. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3213729.

Ball:1984:TBE

[Bal84] Frank Ball. The threshold behaviour of epidemic models. Journal of Applied Probability, 21(2):443–444, June 1984. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3213655.

Ball:1986:NTS

[Bal86] Frank Ball. A note on the total size distribution of epidemic models. Journal of Applied Probability, 23(3):832–836, September 1986. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3214021.

Balding:1988:DRO

[Bal88] David Balding. Diffusion-reaction in one dimension. Journal of Applied Probability, 25(4):733–743, December 1988. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3214294.

Ball:1990:NTS

[Bal90] Frank Ball. A note on the total size distribution of carrier-borne epidemic models. Journal of Applied Probability, 27(4):908–912, December 1990. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3214833.

Ballerini:1994:ACE

[Bal94] Rocco Ballerini. Archimedean copulas, exchangeability, and max-stability. Journal of Applied Probability, 31(2):383–390, June 1994. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3215031.

Balabushkin:1995:AEP

[Bal95] A. N. Balabushkin. Approximation to the exit probability of a continuous Gaussian process over a U-shaped boundary of increasing curvature. Journal of Applied Probability, 32(2):429– 442, June 1995. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3215298.

Ball:1999:CLT

[Bal99] Frank Ball. Central limit theorems for multivariate semi-Markov sequences and processes, with applications. Journal of Applied Probability, 36(2):415–432, June 1999. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3215465.

Baltrunas:2001:SAR

[Bal01] Aleksandras Baltrunas. Some asymptotic results for transient random walks with applications to insurance risk. Journal of Applied Probability, 38(1):108–121, March 2001. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3215745.

Bambos:1992:CRQ

[Bam92] Nicholas Bambos. On closed ring queueing networks. Journal of Applied Probability, 29(4):979–995, December 1992. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3214729.

Banjevic:1988:SSC

[Ban88] Dragan Banjević. On some statistics connected with runs in Markov chains. Journal of Applied Probability, 25(4):815–821, December 1988. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3214303.

Banjevic:1996:RRD

[Ban96] Dragan Banjevic. Recursive relations for the distribution of waiting times in a Markov chain. Journal of Applied Probability, 33 (1):48–56, March 1996. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3215263.

Bansaye:2007:MSF

[Ban07] Vincent Bansaye. On a model for the storage of files on a hardware. II. Evolution of a typical data block. Journal of Applied Probability, 44(4):901–927, December 2007. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/27595896.

Barnett:1964:TPE

[Bar64] V. D. Barnett. A three-player extension of the Gambler’s ruin problem. Journal of Applied Probability, 1(2):321–334, December 1964. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3211863.

Bartlett:1971:PNN

[Bar71] M. S. Bartlett. Physical nearest-neighbour models and nonlinear time-series. Journal of Applied Probability, 8(2):222– 232, June 1971. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3211892.

Barbour:1972:PDA

[Bar72a] Andrew D. Barbour. The principle of the diffusion of arbitrary constants. Journal of Applied Probability, 9(3):519–541, September 1972. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3212323.

Bartlett:1972:PNN

[Bar72b] M. S. Bartlett. Physical nearest-neighbour models and nonlinear time-series. II. Further discussion of approximate solutions and exact equations. Journal of Applied Probability, 9(1): 76–86, March 1972. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3212638.

Bartlett:1974:PNN

[Bar74] M. S. Bartlett. Physical nearest-neighbour models and nonlinear time-series: III non-zero means and sub-critical temperatures. Journal of Applied Probability, 11(4):715–725, December 1974. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3212555.

Barbour:1975:LTP

[Bar75a] Andrew D. Barbour. Limit theorems for processes such as the Markov branching process. Journal of Applied Probability, 12 (2):289–297, June 1975. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3212442.

Barnett:1975:FND

[Bar75b] N. S. Barnett. A further note on duality of random walks. Journal of Applied Probability, 12(3):635–638, September 1975. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/3212884.

Barnett:1975:SRW

[Bar75c] V. Barnett. A simple random walk on parallel axes moving at different rates. Journal of Applied Probability, 12(3):466–476, September 1975. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3212861.

Barbour:1982:GSM

[Bar82a] A. D. Barbour. Generalized semi-Markov schemes and open queueing networks. Journal of Applied Probability, 19(2):469– 474, June 1982. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3213503.

Bartlett:1982:IEE

[Bar82b] M. S. Bartlett. The ‘ideal’ estimation equation. Journal of Applied Probability, 19A:187–200, 1982. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3213559.

Bartlett:1986:MCP

[Bar86] M. S. Bartlett. Mixed Cox processes, with an application to accident statistics. Journal of Applied Probability, 23A:321–333, 1986. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3214361.

See correction [Ano86i].

Barbour:1988:SMP

[Bar88a] A. D. Barbour. Stein’s method and Poisson process convergence. Journal of Applied Probability, 25A:175–184, 1988. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/3214155.

Bartlett:1988:SMF

[Bar88b] M. S. Bartlett. Stochastic models and field trials. Journal of Applied Probability, 25A:79–89, 1988. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3214148.

Boutsikas:2014:JDS

[BAR14] M. V. Boutsikas, D. L. Antzoulakos, and A. C. Rakitzis. On the joint distribution of stopping times and stopped sums in multistate exchangeable trials. Journal of Applied Probability, 51(2):483–491, June 2014. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://projecteuclid.

org/euclid.jap/1402578638.

Barron:2015:FEM

[Bar15] Yonit Barron. A fluid EOQ model with Markovian environment. Journal of Applied Probability, 52(2):473–489, June 2015. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://projecteuclid.org/euclid.jap/1437658610.

Basawa:1970:CPF

[Bas70] I. V. Basawa. On a characterization property of finite irreducible Markov chains. Journal of Applied Probability, 7(3):771–775, December 1970. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3211955.

Basawa:1971:SMB

[Bas71] I. V. Basawa. Some models based on the interaction of two independent Markovian point processes. Journal of Applied Probability, 8(1):193–197, March 1971. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3211848.

Basrak:2010:LTI

[Bas10] Bojan Basrak. Limit theorems for the inductive mean on metric trees. Journal of Applied Probability, 47(4):1136–1149, December 2010. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/29777432.

Bather:1966:CTI

[Bat66] J. A. Bather. A continuous time inventory model. Journal of Applied Probability, 3(2):538–549, December 1966. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3212137.

Bather:1968:DMC

[Bat68] J. A. Bather. A diffusion model for the control of a dam. Journal of Applied Probability, 5(1):55–71, April 1968. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3212077.

Bather:1992:SM

[Bat92] J. A. Bather. Search models. Journal of Applied Probability, 29 (3):605–615, September 1992. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.

org/stable/3214897.

Baudin:1981:LNN

[Bau81] Michel Baudin. Likelihood and nearest-neighbor distance properties of multidimensional Poisson cluster processes. Journal of Applied Probability, 18(4):879–888, December 1981. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3213062.

Baudin:1983:NDC

[Bau83] Michel Baudin. Note on the determination of cluster centers from a realization of a multidimensional Poisson cluster process. Journal of Applied Probability, 20(1):136–143, March 1983. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/3213727.

Baudin:1984:APM

[Bau84a] Michel Baudin. Acknowledgement of priority: Multidimensional point processes and random closed sets. Journal of Applied Probability, 21(3):675, September 1984. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3213631.

Baudin:1984:MPP

[Bau84b] Michel Baudin. Multidimensional point processes and random closed sets. Journal of Applied Probability, 21(1):173–178, March 1984. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3213675.

Bauerle:1997:MRM

[Bäu97] Nicole Bäuerle. Monotonicity results for MR/GI/1 queues. Journal of Applied Probability, 34(2):514–524, June 1997. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3215390.

Baurdoux:2009:LEB

[Bau09] E. J. Baurdoux. Last exit before an exponential time for spectrally negative Lévy processes. Journal of Applied Probability, 46(2):542–558, June 2009. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/25662442.

Baxter:1981:AMT

[Bax81] Laurence A. Baxter. Availability measures for a two-state system. Journal of Applied Probability, 18(1):227–235, March 1981. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/3213182.

Baxter:1983:AMC

[Bax83] Laurence A. Baxter. Availability measures for coherent systems of separately maintained components. Journal of Applied Probability, 20(3):627–636, September 1983. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3213898.

Baxter:1984:CS

[Bax84] Laurence A. Baxter. Continuum structures I. Journal of Applied Probability, 21(4):802–815, December 1984. CODEN JPRBAM.

ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/3213697.

Baxter:1998:DSD

[Bax98] Laurence A. Baxter. Dynamical systems defined on point processes. Journal of Applied Probability, 35(3):581–588, September 1998. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3215633.

Bhattacharjee:1980:KSA

[BB80] Manish C. Bhattacharjee and Sujit K. Basu. On the kesten-Spitzer algorithm for minimal hitting time in controlled Markov chains. Journal of Applied Probability, 17(3):716–725, September 1980. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3212965.

Basu:1984:WCW

[BB84] Sujit K. Basu and Manish C. Bhattacharjee. On weak convergence within the HNBUE family of life distributions. Journal of Applied Probability, 21(3):654–660, September 1984. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3213627.

Benkherouf:1988:OES

[BB88] L. Benkherouf and J. A. Bather. Oil exploration: Sequential decisions in the face of uncertainty. Journal of Applied Probability, 25(3):529–543, September 1988. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3213982.

Ball:1990:PAS

[BB90] Frank Ball and A. D. Barbour. Poisson approximation for some epidemic models. Journal of Applied Probability, 27(3):479–490, September 1990. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3214534.

Barbour:1996:AVM

[BB96] A. D. Barbour and Timothy C. Brown. Approximate versions of Melamed’s theorem. Journal of Applied Probability, 33(2):472– 489, June 1996. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3215072.

Ball:2005:EME

[BB05] Frank Ball and Tom Britton. An epidemic model with exposuredependent severities. Journal of Applied Probability, 42(4):932– 949, December 2005. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/30040819.

Bartoszek:2006:TBO

[BB06] Krzysztof Bartoszek and Wojciech Bartoszek. On the time behaviour of okazaki fragments. Journal of Applied Probability, 43 (2):500–509, June 2006. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/27595746.

Brandwajn:2012:RSP

[BB12] Alexandre Brandwajn and Thomas Begin. A recurrent solution of Ph/M/c/N-like and Ph/M/c-like queues. Journal of Applied Probability, 49(1):84–99, March 2012. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/41713753.

Bjornberg:2014:CNM

[BB14] Jakob E. Björnberg and Erik I. Broman. Coexistence and noncoexistence of Markovian viruses and their hosts. Journal of Applied Probability, 51(1):191–208, March 2014. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://projecteuclid.org/euclid.jap/1395771423.

Bieniek:2018:ULD

[BB18] M. Bieniek and M. Burkschat. On unimodality of the lifetime distribution of coherent systems with failure-dependent component lifetimes. Journal of Applied Probability, 55(2):473–487, June 2018. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-

6072 (electronic). URL https://www.cambridge.org/core/
journals/journal-of-applied-probability/article/on-unimodality-
of-the-lifetime-distribution-of-coherent-systems-with-
failuredependent-component-lifetimes/B1EB8B6DD4909E7367A05C6AC630BE33.

Bjornberg:2014:SMV

[BBBN14] J. E. Björnberg, T. Britton, E. I. Broman, and E. Natan. A stochastic model for virus growth in a cell population. Journal of Applied Probability, 51(3):599–612, September 2014. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://projecteuclid.org/euclid.jap/1409932661.

Badia:2001:OII

[BBC01] F. G. Badía, M. D. Berrade, and Clemente A. Campos. Optimization of inspection intervals based on cost. Journal of Applied Probability, 38(4):872–881, December 2001. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/3215770.

Balakrishnan:2014:SOA

[BBH14] Narayanaswamy Balakrishnan, Ghobad Barmalzan, and Abedin Haidari. Stochastic orderings and ageing properties of residual life lengths of live components in (n − k + 1)-out-of-n systems. Journal of Applied Probability, 51(1):58–68, March 2014. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://projecteuclid.org/euclid.jap/1395771413.

Balakrishnan:2018:SCK

[BBH18] Narayanaswamy Balakrishnan, Ghobad Barmalzan, and Abedin Haidari. On stochastic comparisons of k-out-of-n systems with Weibull components. Journal of Applied Probability, 55(1):216–232, March 2018. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL https:/

/www.cambridge.org/core/journals/journal-of-applied-
probability/article/on-stochastic-comparisons-of-koutofn-
systems-with-weibull-components/B3E0B5416144E048D2EFBED422DE9C50.

Bekker:2008:QDT

[BBK08] R. Bekker, O. J. Boxma, and O. Kella. Queues with delays in two-state strategies and Lévy input. Journal of Applied Probability, 45(2):314–332, June 2008. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/27595947.

Biagini:2008:PCI

[BBMB08] Francesca Biagini, Yuliya Bregman, and Thilo Meyer-Brandis. Pricing of catastrophe insurance options under immediate loss reestimation. Journal of Applied Probability, 45(3):831–845, September 2008. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
27595989.

Ball:2016:EDB

[BBN16] Frank Ball, Tom Britton, and Peter Neal. On expected durations of birth–death processes, with applications to branching processes and SIS epidemics. Journal of Applied Probability, 53 (1):203–215, March 2016. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://projecteuclid.

org/euclid.jap/1457470569.

Borovkov:2003:IWP

[BBP03] A. A. Borovkov, O. J. Boxma, and Z. Palmowski. On the integral of the workload process of the single server queue. Journal of Applied Probability, 40(1):200–225, March 2003. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3215849.

Block:1985:ADM

[BBS85] Henry W. Block, Wagner S. Borges, and Thomas H. Savits. Agedependent minimal repair. Journal of Applied Probability, 22 (2):370–385, June 1985. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3213780.

Brandt:1990:RDD

[BBS90] Andreas Brandt, Manfred Brandt, and Hannelore Sulanke. A representation of a discrete distribution by its binomial moments. Journal of Applied Probability, 27(1):208–214, March 1990. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3214608.

Berkelmans:2023:BPD

[BBvdMP23] Guus Berkelmans, Sandjai Bhulai, Rob van der Mei, and Joris Pries. The Berkelmans–Pries dependency function: a generic measure of dependence between random variables. Journal of Applied Probability, 60(4):1115–1135, December 2023. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL https://www.cambridge.org/core/journals/journal-
of-applied-probability/article/berkelmanspries-dependency-
function-a-generic-measure-of-dependence-between-random-
variables/35623DB91C22CA64F3878EB7F134EC27.

Bather:1967:SDC

[BC67] John Bather and Herman Chernoff. Sequential decisions in the control of a space-ship (finite fuel). Journal of Applied Probability, 4(3):584–604, December 1967. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3212223.

Bloomfield:1972:LTA

[BC72] P. Bloomfield and D. R. Cox. A low traffic approximation for queues. Journal of Applied Probability, 9(4):832–840, December 1972. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3212619.

Boes:1973:ERE

[BC73] D. C. Boes and J. D. Salas-La Cruz. On the expected range and expected adjusted range of partial sums of exchangeable random variables. Journal of Applied Probability, 10(3):671–677, September 1973. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3212787.

Brockwell:1975:ETD

[BC75] P. J. Brockwell and K. L. Chung. Emptiness times of a dam with stable input and general release function. Journal of Applied Probability, 12(1):212–217, March 1975. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http:/

/www.jstor.org/stable/3212432.

Borst:1992:BPA

[BC92] S. C. Borst and M. B. Combé. Busy-period analysis of a correlated queue. Journal of Applied Probability, 29(2):482–483, June 1992. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3214587.

Ball:1995:FOT

[BC95a] Frank Ball and Damian Clancy. The final outcome and temporal solution of a carrier-borne epidemic model. Journal of Applied Probability, 32(2):304–315, June 1995. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3215290.

Ball:1995:FOE

[BC95b] Frank Ball and Damian Clancy. The final outcome of an epidemic model with several different types of infective in a large population. Journal of Applied Probability, 32(3):579–590, September 1995. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3215114.

Belinskiy:2001:ESD

[BC01a] Boris P. Belinskiy and Peter Caithamer. Energy of a string driven by a two-parameter Gaussian noise white in time. Journal of Applied Probability, 38(4):960–974, December 2001. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3215777.

Block:2001:MPT

[BC01b] Henry W. Block and Tuhao Chen. Multivariate producttype lower bounds. Journal of Applied Probability, 38(2):407– 420, June 2001. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3215896.

Ball:2023:APT

[BC23] Frank G Ball and Damian Clancy. Asymptotic persistence time formulae for multitype birth-death processes. Journal of Applied Probability, 60(3):895–920, September 2023. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL https:/

/www.cambridge.org/core/journals/journal-of-applied-
probability/article/asymptotic-persistence-time-formulae-
for-multitype-birthdeath-processes/4F78DE9122149576DF56D44CA65F276B.

Buttell:1993:ECV

[BCD93] L. Buttell, J. T. Cox, and R. Durrett. Estimating the critical values of stochastic growth models. Journal of Applied Probability, 30(2):455–461, June 1993. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3214854.

Brown:2017:TLR

[BCdlP17] Mark Brown, Joel E. Cohen, and Victor H. de la Peña. Taylor’s law, via ratios, for some distributions with infinite mean. Journal of Applied Probability, 54(3):657–669, September 2017. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL https://www.cambridge.org/core/journals/journal-
of-applied-probability/article/taylors-law-via-ratios-
for-some-distributions-with-infinite-mean/10A3BCFDE83F61CBB2BAE25C964C21C4.

Bercu:2009:ASC

[BCF09] Bernard Bercu, Peggy Cénac, and Guy Fayolle. On the almost sure central limit theorem for vector martingales: Convergence of moments and statistical applications. Journal of Applied Probability, 46(1):151–169, March 2009. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http:/

/www.jstor.org/stable/25662413.

Brillinger:2001:AMG

[BCIM01] David R. Brillinger, Chang Chiann, Rafael A. Irizarry, and Pedro A. Morettin. Automatic methods for generating seismic intensity maps. Journal of Applied Probability, 38A:188–201, 2001. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3215878.

Baryshnikov:2004:SFD

[BCJ04] Yuliy Baryshnikov, E. G. Coffman, Jr., and Predrag Jelenković. Space filling and depletion. Journal of Applied Probability, 41(3): 691–702, September 2004. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/4141347.

Barbour:2016:ERI

[BCK16] A. D. Barbour, P. Chigansky, and F. C. Klebaner. On the emergence of random initial conditions in fluid limits. Journal of Applied Probability, 53(4):1193–1205, December 2016. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://projecteuclid.org/euclid.jap/1481132846.

Bollobas:2017:JPR

[BCKK17] Béla Bollobás, Oliver Cooley, Mihyun Kang, and Christoph Koch. Jigsaw percolation on random hypergraphs. Journal of Applied Probability, 54(4):1261–1277, December 2017. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL https://www.cambridge.org/core/journals/journal-
of-applied-probability/article/jigsaw-percolation-on-
random-hypergraphs/AB26271025355F7025D18C83C94787E0.

Blath:2013:APL

[BCKS13] Jochen Blath, Adrián González Casanova, Noemi Kurt, and Dario Spanò. The ancestral process of long-range seed bank models. Journal of Applied Probability, 50(3):741–759, September 2013. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-

6072 (electronic). URL http://projecteuclid.org/euclid.
jap/1378401233.

Barnes:1998:RWF

[BCL98] George R. Barnes, Patricia B. Cerrito, and Inessa Levi. Random walks on finite semigroups. Journal of Applied Probability, 35(4): 824–832, December 1998. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3215519.

Badescu:2009:DRM

[BCL09] Andrei L. Badescu, Eric C. K. Cheung, and David Landriault. Dependent risk models with bivariate phase-type distributions. Journal of Applied Probability, 46(1):113–131, March 2009. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/25662411.

Buckholtz:1983:ATB

[BCMW83] Peter G. Buckholtz, L. Lorne Campbell, Ross D. Milbourne, and M. T. Wasan. Analysis of transient behaviour of certain processes with return to a central state. Journal of Applied Probability, 20(1):61–70, March 1983. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3213720.

Bet:2024:WIO

[BCN24] Gianmarco Bet, Fabio Coppini, and Francesca Romana Nardi. Weakly interacting oscillators on dense random graphs. Journal of Applied Probability, 61(1):255–278, March 2024. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL https://www.cambridge.org/core/journals/journal-
of-applied-probability/article/weakly-interacting-oscillators-
on-dense-random-graphs/CC20B2407EA73401277EBE7DC134A3C8.

Baggerly:2000:ECA

[BCP00] Keith Baggerly, Dennis Cox, and Rick Picard. Exponential convergence of adaptive importance sampling for Markov chains. Journal of Applied Probability, 37(2):342–358, June 2000. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/3215712.

Berti:2011:CLT

[BCPR11] Patrizia Berti, Irene Crimaldi, Luca Pratelli, and Pietro Rigo. A central limit theorem and its applications to multicolor randomly reinforced urns. Journal of Applied Probability, 48(2):527– 546, June 2011. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
23065900.

Berti:2016:ARR

[BCPR16] Patrizia Berti, Irene Crimaldi, Luca Pratelli, and Pietro Rigo. Asymptotics for randomly reinforced urns with random barriers. Journal of Applied Probability, 53(4):1206–1220, December 2016. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://projecteuclid.org/euclid.jap/
1481132847.

Badescu:2011:TDR

[BCR11] Andrei L. Badescu, Eric C. K. Cheung, and Landy Rabehasaina. A two-dimensional risk model with proportional reinsurance. Journal of Applied Probability, 48(3):749–765, September 2011. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/23065934.

Block:1993:OSO

[BCS93] Henry W. Block, Timothy M. Costigan, and Allan R. Sampson. Optimal second-order product probability bounds. Journal of Applied Probability, 30(3):675–691, September 1993. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3214774.

Berend:2008:RWM

[BCS08] Daniel Berend, Yuri Chernyavsky, and Luba Sapir. Ranking of weighted majority rules. Journal of Applied Probability, 45(4): 994–1006, December 2008. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/27596004.

Brandes:2023:ISM

[BCS23] Dirk-Philip Brandes, Imma Valentina Curato, and Robert Stelzer. Inheritance of strong mixing and weak dependence under renewal sampling. Journal of Applied Probability, 60(2):435–451, June 2023. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL https:/

/www.cambridge.org/core/journals/journal-of-applied-
probability/article/inheritance-of-strong-mixing-and-
weak-dependence-under-renewal-sampling/FD7406930688CEB524204D9F98AB758A.

Bertoin:1997:TCT

[BCY97] J. Bertoin, L. Chaumont, and M. Yor. Two chain-transformations and their applications to quantiles. Journal of Applied Probability, 34(4):882–897, December 1997. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3215004.

Blight:1974:REM

[BD74] B. J. N. Blight and J. L. Devore. The recursive estimation of a Markov chain. Journal of Applied Probability, 11(2):394– 400, June 1974. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3212763.

Bar-David:1975:RND

[BD75] Israel Bar-David. Radon–Nikodym derivatives, passages and maxima for a Gaussian process with particular covariance and mean. Journal of Applied Probability, 12(4):724–733, December 1975. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3212723.

Billard:1988:DAM

[BD88a] L. Billard and P. W. A. Dayananda. A drug addiction model. Journal of Applied Probability, 25(4):649–662, December 1988. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/3214286.

Bingham:1988:HDA

[BD88b] N. H. Bingham and R. A. Doney. On higher-dimensional analogues of the arc-sine law. Journal of Applied Probability, 25 (1):120–131, March 1988. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/ stable/3214239. See correction [Zho08].

Balkema:1990:CRE

[BD90] A. A. Balkema and L. De Haan. A convergence rate in extremevalue theory. Journal of Applied Probability, 27(3):577–585, September 1990. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3214542.

Bertoin:1994:LBL

[BD94a] J. Bertoin and R. A. Doney. On the local behaviour of ladder height distributions. Journal of Applied Probability, 31(3): 816–821, September 1994. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3215158.

Buffet:1994:EUB

[BD94b] E. Buffet and N. G. Duffield. Exponential upper bounds via martingales for multiplexers with Markovian arrivals. Journal of Applied Probability, 31(4):1049–1060, December 1994. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3215328.

Bierme:2012:FAL

[BD12] Hermine Biermé and Agnès Desolneux. A Fourier approach for the level crossings of shot noise processes with jumps. Journal of Applied Probability, 49(1):100–113, March 2012. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/41713754.

Borovkov:2014:SDS

[BDG14] K. Borovkov, G. Decrouez, and M. Gilson. On stationary distributions of stochastic neural networks. Journal of Applied Probability, 51(3):837–857, September 2014. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

projecteuclid.org/euclid.jap/1409932677.

Brozius:1987:LLC

[BdH87] Henk Brozius and Laurens de Haan. On limiting laws for the convex hull of a sample. Journal of Applied Probability, 24(4): 852–862, December 1987. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3214210.

Balkema:1993:ADM

[BdHK93] A. A. Balkema, L. de Haan, and R. L. Karandikar. Asymptotic distribution of the maximum of n independent stochastic processes. Journal of Applied Probability, 30(1):66–81, March 1993. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3214622.

Buraczewski:2018:PGL

[BDIM18] Dariusz Buraczewski, Piotr Dyszewski, Alexander Iksanov, and Alexander Marynych. On perpetuities with gamma-like tails. Journal of Applied Probability, 55(2):368–389, June 2018. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL https://www.cambridge.org/core/journals/journal-
of-applied-probability/article/on-perpetuities-with-
gammalike-tails/E42C37BC0EB0AE1C58C7CA77414ADA0D.

Blanchet:2019:ESI

[BDL19] Jose Blanchet, Jing Dong, and Zhipeng Liu. Exact sampling of the infinite horizon maximum of a random walk over a nonlinear boundary. Journal of Applied Probability, 56(1):116–138, March 2019. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL https:/

/www.cambridge.org/core/journals/journal-of-applied-
probability/article/exact-sampling-of-the-infinite-horizon-
maximum-of-a-random-walk-over-a-nonlinear-boundary/AFF6F369133273B7FC5A1FEF65C4F221.

Britton:2008:ERG

[BDLL08] Tom Britton, Maria Deijfen, Andreas N. Lagerås, and Mathias Lindholm. Epidemics on random graphs with tunable clustering. Journal of Applied Probability, 45(3):743–756, September 2008. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/27595984.

Bose:2009:SLB

[BDM09] Arup Bose, Amites Dasgupta, and Krishanu Maulik. Strong laws for balanced triangular urns. Journal of Applied Probability, 46 (2):571–584, June 2009. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/25662444.

Bacry:2012:CTS

[BDM12] Emmanuel Bacry, Laurent Duvernet, and Jean-François Muzy. Continuous-time skewed multifractal processes as a model for financial returns. Journal of Applied Probability, 49(2):482– 502, June 2012. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
41713781.

Bisewski:2021:BES

[BDM21] Krzysztof Bisewski, Krzysztof Debicki, and Michel Mandjes. Bounds for expected supremum of fractional Brownian motion with drift. Journal of Applied Probability, 58(2):411–427, June 2021. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL https:/

/www.cambridge.org/core/journals/journal-of-applied-
probability/article/bounds-for-expected-supremum-of-
fractional-brownian-motion-with-drift/0E6F93F947F6E389FC7C3CF13FB734AB.

Bshouty:2012:GTP

[BDMZ12] Daoud Bshouty, Antonio Di Crescenzo, Barbara Martinucci, and Shelemyahu Zacks. Generalized telegraph process with random delays. Journal of Applied Probability, 49(3):850–865, September 2012. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/41713810.

Bompadre:2007:PAU

[BDO07] Agustín Bompadre, Moshe Dror, and James B. Orlin. Probabilistic analysis of unit-demand vehicle routeing problems. Journal of Applied Probability, 44(1):259–278, March 2007. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/27595836.

Baccelli:2022:RMF

[BDT22] François Baccelli, Michel Davydov, and Thibaud Taillefumier. Replica-mean-field limits of fragmentation-interaction-aggregation processes. Journal of Applied Probability, 59(1):38– 59, March 2022. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL https://www.cambridge.org/
core/journals/journal-of-applied-probability/article/
replicameanfield-limits-of-fragmentationinteractionaggregation-
processes/1B25A957ED6CF6EF363281F32E364DA0.

Brockwell:2007:ENL

[BDY07] Peter J. Brockwell, Richard A. Davis, and Yu Yang. Estimation for nonnegative Lévy–driven Ornstein–Uhlenbeck processes. Journal of Applied Probability, 44(4):977–989, December 2007. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/27595901.

Brown:1973:SCW

[BE73] B. M. Brown and G. K. Eagleson. Simpler conditions for Wald equations. Journal of Applied Probability, 10(2):451–455, June 1973. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3212363.

Berman:1983:PLT

[BE83] M. Berman and G. K. Eagleson. A Poisson limit theorem for incomplete symmetric statistics. Journal of Applied Probability, 20(1):47–60, March 1983. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3213719.

Barbour:1987:IPL

[BE87] A. D. Barbour and G. K. Eagleson. An improved Poisson limit theorem for sums of dissociated random variables. Journal of Applied Probability, 24(3):586–599, September 1987. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3214091.

Bhattacharya:1991:SMP

[BE91] Partha P. Bhattacharya and Anthony Ephremides. Stochastic monotonicity properties of multiserver queues with impatient customers. Journal of Applied Probability, 28(3):673–682, September 1991. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3214501.

Becker:1970:SMT

[Bec70] Niels G. Becker. A stochastic model for two interacting populations. Journal of Applied Probability, 7(3):544–564, December 1970. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3211937.

Becker:1973:CBE

[Bec73] Niels Becker. Carrier-borne epidemics in a community consisting of different groups. Journal of Applied Probability, 10(3): 491–501, September 1973. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3212770.

Beekman:1975:ADO

[Bee75] John A. Beekman. Asymptotic distributions for the Ornstein– Uhlenbeck process. Journal of Applied Probability, 12(1):107– 114, March 1975. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3212412.

Bean:2015:BOC

[BEER15] N. G. Bean, R. Elliott, A. Eshragh, and J. V. Ross. On binomial observations of continuous-time Markovian population models. Journal of Applied Probability, 52(2):457–472, June 2015. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://projecteuclid.org/euclid.jap/1437658609.

Bayraktar:2022:DDC

[BEG22] Erhan Bayraktar, Erik Ekström, and Jia Guo. Disorder detection with costly observations. Journal of Applied Probability, 59(2):338–349, June 2022. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL https:/

/www.cambridge.org/core/journals/journal-of-applied-
probability/article/disorder-detection-with-costly-observations/
982D0805ECB8C6827BE2FC1700E592D7.

Bellman:1966:DPP

[Bel66] Richard Bellman. Dynamic programming, pattern recognition and location of faults in complex systems. Journal of Applied Probability, 3(1):268–271, June 1966. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3212051.

Belkin:1971:FPG

[Bel71] Barry Belkin. First passage to a general threshold for a process corresponding to sampling at Poisson times. Journal of Applied Probability, 8(3):573–588, September 1971. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/3212180.

Belisle:1992:CTC

[Bél92] Claude J. P. Bélisle. Convergence theorems for a class of simulated annealing algorithms on R^{d}. Journal of Applied Probability, 29(4):885–895, December 1992. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3214721.

Bellin:2023:ABF

[Bel23] Etienne Bellin. Asymptotic behaviour of the first positions of uniform parking functions. Journal of Applied Probability, 60(4):1201–1218, December 2023. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL https:/

/www.cambridge.org/core/journals/journal-of-applied-
probability/article/asymptotic-behaviour-of-the-first-
positions-of-uniform-parking-functions/746E246F39BF78159299E9138A73CB17.

Benes:1968:FRI

[Ben68] V. E. Benes. Finite regular invariant measures for Feller processes. Journal of Applied Probability, 5(1):203–209, April 1968. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/3212087.

Benishay:1970:ISO

[Ben70] Haskel Benishay. Inflow-stock-outflow time series, stochastic lag and disaggregation: A note. Journal of Applied Probability, 7 (1):234–239, April 1970. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3212163.

Benjamin:1982:HL

[Ben82] Bernard Benjamin. The human lifespan. Journal of Applied Probability, 19A:159–172, 1982. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.

org/stable/3213557.

Benda:1998:CLT

[Ben98a] Martin Benda. A central limit theorem for contractive stochastic dynamical systems. Journal of Applied Probability, 35(1):200– 205, March 1998. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3215557.

Benhenni:1998:AIS

[Ben98b] Karim Benhenni. Approximating integrals of stochastic processes: Extensions. Journal of Applied Probability, 35(4):843– 855, December 1998. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3215521.

Boland:1998:SIB

[BEN98c] Philip J. Boland and Emad El-Neweihi. Statistical and information based (physical) minimal repair for k out of n systems. Journal of Applied Probability, 35(3):731–740, September 1998. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/3215647.

Boland:1994:AHR

[BENP94] Philip J. Boland, Emad El-Neweihi, and Frank Proschan. Applications of the hazard rate ordering in reliability and order statistics. Journal of Applied Probability, 31(1):180–192, March 1994. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3215245.

Berman:1965:RTM

[Ber65] Simeon M. Berman. A renewal theoretic model for chronic disease statistics. Journal of Applied Probability, 2(1):119– 128, June 1965. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/

3211878. See correction [Ano67d].

Beran:1968:TUC

[Ber68] R. J. Beran. Testing for uniformity on a compact homogeneous space. Journal of Applied Probability, 5(1):177–195, April 1968. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/3212085.

Berman:1970:OTS

[Ber70] Simeon M. Berman. Occupation times of stationary Gaussian processes. Journal of Applied Probability, 7(3):721–733, December 1970. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3211949.

Berman:1973:RSM

[Ber73] Elizabeth Berman. Regular stochastic matrices and digraphs. Journal of Applied Probability, 10(1):241–243, March 1973. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/3212514.

Berg:1976:POA

[Ber76] Menachem Berg. A proof of optimality for age replacement policies. Journal of Applied Probability, 13(4):751–759, December 1976. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3212530.

Berk:1977:CCD

[Ber77a] Robert H. Berk. Characterizations via conditional distributions. Journal of Applied Probability, 14(4):806–816, December 1977. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/3213353.

Berman:1977:DDA

[Ber77b] Mark Berman. Distance distributions associated with Poisson processes of geometric figures. Journal of Applied Probability, 14 (1):195–199, March 1977. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3213273.

Berman:1977:SMG

[Ber77c] Mark Berman. Some multivariate generalizations of results in univariate stationary point processes. Journal of Applied Probability, 14(4):748–757, December 1977. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3213348.

Bergman:1980:OSR

[Ber80] Bo Bergman. On the optimality of stationary replacement strategies. Journal of Applied Probability, 17(1):178–186, March 1980. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3212935.

Berg:1981:SSR

[Ber81a] Sven Berg. On snowball sampling, random mappings and related problems. Journal of Applied Probability, 18(1):283–290, March 1981. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3213190.

Berman:1981:CTT

[Ber81b] Mark Berman. The covariance of two type II counters. Journal of Applied Probability, 18(3):782–787, September 1981. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3213337.

Berg:1983:RCP

[Ber83] Sven Berg. Random contact processes, snowball sampling and factorial series distributions. Journal of Applied Probability, 20 (1):31–46, March 1983. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3213718.

Berman:1985:AFD

[Ber85] Simeon M. Berman. An asymptotic formula for the distribution of the maximum of a Gaussian process with stationary increments. Journal of Applied Probability, 22(2):454–460, June 1985. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3213789.

Berred:1992:RVE

[Ber92] Mohamed Berred. On record values and the exponent of a distribution with regularly varying upper tail. Journal of Applied Probability, 29(3):575–586, September 1992. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/3214894.

Berman:1994:OTA

[Ber94] Simeon M. Berman. Optimal timing of antiviral therapy in HIV infection. Journal of Applied Probability, 31A:3–15, 1994. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/3214944.

Bertoin:1999:OHT

[Ber99] Jean Bertoin. On overshoots and hitting times for random walks. Journal of Applied Probability, 36(2):593–600, June 1999. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/3215477.

Berard:2004:ASC

[Bèr04] Jean Bèrard. The almost sure central limit theorem for onedimensional nearest-neighbour random walks in a space-time random environment. Journal of Applied Probability, 41(1):83– 92, March 2004. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3215816.

Berman:2010:PAE

[Ber10] Simeon M. Berman. Probabilistic analysis of the efficacy of periodic testing of employees. Journal of Applied Probability, 47(4): 1174–1190, December 2010. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/29777435.

Bertoin:2011:MOC

[Ber11] Jean Bertoin. On the maximal offspring in a critical branching process with infinite variance. Journal of Applied Probability, 48 (2):576–582, June 2011. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/23065904.

Bertoin:2013:AGC

[Ber13a] Jean Bertoin. Almost giant clusters for percolation on large trees with logarithmic heights. Journal of Applied Probability, 50(3): 603–611, September 2013. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://projecteuclid.

org/euclid.jap/1378401225.

Bertoin:2013:LOC

[Ber13b] Jean Bertoin. On largest offspring in a critical branching process with finite variance. Journal of Applied Probability, 50(3): 791–800, September 2013. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://projecteuclid.

org/euclid.jap/1378401236.

Bertoin:2023:MEC

[Ber23] Jean Bertoin. A model for an epidemic with contact tracing and cluster isolation, and a detection paradox. Journal of Applied Probability, 60(3):1079–1095, September 2023. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL https://www.cambridge.org/core/journals/
journal-of-applied-probability/article/model-for-an-
epidemic-with-contact-tracing-and-cluster-isolation-
and-a-detection-paradox/BBF7F5CF6DF61822407D6B515449DD59.

Besag:1972:NNS

[Bes72] J. E. Besag. Nearest-neighbour systems: A lemma with application to Bartlett’s global solutions. Journal of Applied Probability, 9(2):418–421, June 1972. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3212810.

Barrett:1978:GFT

[BF78] Wayne W. Barrett and Philip J. Feinsilver. Gaussian families and a theorem on patterned matrices. Journal of Applied Probability, 15(3):514–522, September 1978. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3213114.

Belisle:1991:WAM

[BF91] Claude Bélisle and Julian Faraway. Winding angle and maximum winding angle of the two-dimensional random walk. Journal of Applied Probability, 28(4):717–726, December 1991. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/3214675.

Bruss:1993:MER

[BF93] F. Thomas Bruss and Thomas S. Ferguson. Minimizing the expected rank with full information. Journal of Applied Probability, 30(3):616–626, September 1993. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3214770.

Baccelli:1995:SRS

[BF95] François Baccelli and Serguei Foss. On the saturation rule for the stability of queues. Journal of Applied Probability, 32(2):494– 507, June 1995. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3215303.

Bruss:1997:MBS

[BF97] F. Thomas Bruss and Thomas S. Ferguson. Multiple buying or selling with vector offers. Journal of Applied Probability, 34(4): 959–973, December 1997. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3215010.

Benaim:2008:SAI

[BF08] Shalom Benaim and Peter Friz. Smile asymptotics II: Models with known moment generating functions. Journal of Applied Probability, 45(1):16–32, March 2008. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/27595920.

Bona:2009:ISR

[BF09] Miklós Bóna and Philippe Flajolet. Isomorphism and symmetries in random phylogenetic trees. Journal of Applied Probability, 46(4):1005–1019, December 2009. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/25662477.

Buckingham:2016:NIR

[BFK16] Patrick Buckingham, Brian Fralix, and Offer Kella. A note on integral representations of the Skorokhod map. Journal of Applied Probability, 53(1):293–298, March 2016. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http:/

/projecteuclid.org/euclid.jap/1457470575.

Baccelli:2005:TGJ

[BFL05] François Baccelli, Serguei Foss, and Marc Lelarge. Tails in generalized Jackson networks with subexponential service-time distributions. Journal of Applied Probability, 42(2):513–530, June 2005. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/30040807.

Bertoin:2008:PIS

[BFM08] Jean Bertoin, Joaquin Fontbona, and Servet Martínez. On prolific individuals in a supercritical continuous-state branching process. Journal of Applied Probability, 45(3):714–726, September 2008. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/27595982.

Boxma:2017:RRM

[BFP17] Onno J. Boxma, Esther Frostig, and David Perry. A reinsurance risk model with a threshold coverage policy: the Gerber–Shiu penalty function. Journal of Applied Probability, 54(1):267–285, March 2017. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL https:/

/www.cambridge.org/core/journals/journal-of-applied-
probability/article/reinsurance-risk-model-with-a-threshold-
coverage-policy-the-gerbershiu-penalty-function/5E8CCE1C852954471B18E29F9FAB7CA0.

Bordenave:2009:RMA

[BFS09] Charles Bordenave, Serguei Foss, and Vsevolod Shneer. A random multiple-access protocol with spatial interactions. Journal of Applied Probability, 46(3):844–865, September 2009. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/25662464.

Biggins:1979:CLR

[BG79a] J. D. Biggins and D. R. Grey. Continuity of limit random variables in the branching random walk. Journal of Applied Probability, 16(4):740–749, December 1979. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3213141.

Bramson:1979:NER

[BG79b] Maury Bramson and David Griffeath. A note on the extinction rates of some birth-death processes. Journal of Applied Probability, 16(4):897–902, December 1979. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3213155.

Biggins:1987:EPS

[BG87a] J. D. Biggins and Thomas Götz. Expected population size in the generation-dependent branching process. Journal of Applied Probability, 24(2):304–314, June 1987. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3214255.

Boxma:1987:PCL

[BG87b] O. J. Boxma and W. P. Groenendijk. Pseudo-conservation laws in cyclic-service systems. Journal of Applied Probability, 24(4): 949–964, December 1987. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3214218.

Brun:2000:ASF

[BG00] Olivier Brun and Jean-Marie Garcia. Analytical solution of finite capacity M/D/1 queues. Journal of Applied Probability, 37(4): 1092–1098, December 2000. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3215497.

Baringhaus:2002:MBS

[BG02] Ludwig Baringhaus and Rudolf Grübel. From matchbox to bottle: A storage problem. Journal of Applied Probability, 39(3): 650–656, September 2002. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3216087.

Brummelhuis:2005:MCD

[BG05] Raymond Brummelhuis and Dominique Guégan. Multiperiod conditional distribution functions for conditionally normal Garch(1, 1) models. Journal of Applied Probability, 42(2):426– 445, June 2005. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
30040802.

Baringhaus:2006:MCA

[BG06] Ludwig Baringhaus and Rudolf Grübel. Monte Carlo algorithms for finding the maximum of a random walk with negative drift. Journal of Applied Probability, 43(1):74–86, March 2006. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/27595711.

Blumenthal:1971:SNS

[BGH71] S. Blumenthal, J. A. Greenwood, and L. Herbach. Superimposed non-stationary renewal processes. Journal of Applied Probability, 8(1):184–192, March 1971. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3211847.

Bloznelis:2012:BSC

[BGJ12] Mindaugas Bloznelis, Friedrich Götze, and Jerzy Jaworski. Birth of a strongly connected giant in an inhomogeneous random digraph. Journal of Applied Probability, 49(3):601–611, September 2012. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/41713792.

Beghin:2015:CFC

[BGM15] Luisa Beghin, Roberto Garra, and Claudio Macci. Correlated fractional counting processes on a finite-time interval. Journal of Applied Probability, 52(4):1045–1061, December 2015. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://projecteuclid.org/euclid.jap/1450802752.

Buonocore:1990:TBF

[BGNR90] A. Buonocore, V. Giorno, A. G. Nobile, and L. M. Ricciardi. On the two-boundary first-crossing-time problem for diffusion processes. Journal of Applied Probability, 27(1):102–114, March 1990. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3214598.

Burtin:1980:SFR

[BGP80] Y. D. Burtin, I. Gertsbakh, and B. Pittel. On a simple formula for random mappings and its applications. Journal of Applied Probability, 17(2):403–414, June 1980. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3213029.

Bose:2003:CLR

[BGSS03] Arup Bose, Sreela Gangopadhyay, Anish Sarkar, and Arindam Sengupta. Convergence of lower records and infinite divisibility. Journal of Applied Probability, 40(4):865–880, December 2003. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/3216047.

Bean:1998:OPM

[BGT98] Nigel Bean, David Green, and Peter Taylor. The output process of an MMPP/M/1 queue. Journal of Applied Probability, 35(4): 998–1002, December 1998. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3215535.

Bossy:2004:SES

[BGT04] Mireille Bossy, Emmanuel Gobet, and Denis Talay. A symmetrized Euler scheme for an efficient approximation of reflected diffusions. Journal of Applied Probability, 41(3):877–889, September 2004. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
4141359.

Bhattacharya:1983:HEU

[BGW83] R. N. Bhattacharya, Vijay K. Gupta, and Ed Waymire. The Hurst effect under trends. Journal of Applied Probability, 20(3): 649–662, September 1983. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3213900.

Brown:1973:NKA

[BH73] T. J. Brown and J. C. Hardin. A note on Kendall’s autoregressive series. Journal of Applied Probability, 10(2):475–478, June 1973. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3212368.

Brown:1975:ALT

[BH75a] B. M. Brown and J. I. Hewitt. Asymptotic likelihood theory for diffusion processes. Journal of Applied Probability, 12(2):228– 238, June 1975. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3212436.

Brown:1975:IDB

[BH75b] B. M. Brown and J. I. Hewitt. Inference for the diffusion branching process. Journal of Applied Probability, 12(3):588– 594, September 1975. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3212875.

Braker:1991:FZE

[BH91] H. U. Bräker and J. Hüsler. On the first zero of an empirical characteristic function. Journal of Applied Probability, 28(3): 593–601, September 1991. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3214494.

Barry:1996:MB

[BH96] Daniel Barry and John A. Hartigan. The minimax bookie. Journal of Applied Probability, 33(4):1093–1107, December 1996. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/3214988.

Bentarzi:1998:SFP

[BH98] Mohamed Bentarzi and Marc Hallin. Spectral factorization of periodically correlated MA(1) processes. Journal of Applied Probability, 35(1):46–54, March 1998. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3215545.

Bebbington:2001:SLS

[BH01] Mark Bebbington and David S. Harte. On the statistics of the linked stress release model. Journal of Applied Probability, 38A: 176–187, 2001. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3215877.

Bo:2012:FPT

[BH12] Lijun Bo and Chen Hao. First passage times of constantelasticity-of-variance processes with two-sided reflecting barriers. Journal of Applied Probability, 49(4):1119–1133, December 2012. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/41713832.

Braunsteins:2020:PEB

[BH20] Peter Braunsteins and Sophie Hautphenne. The probabilities of extinction in a branching random walk on a strip. Journal of Applied Probability, 57(3):811–831, September 2020. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL https://www.cambridge.org/core/journals/journal-
of-applied-probability/article/probabilities-of-extinction-
in-a-branching-random-walk-on-a-strip/C69BB72B6B4A0CDEA7324E21BA97BC4E.

Bhargava:1965:TCD

[Bha65a] T. N. Bhargava. On time changes in a digraph. Journal of Applied Probability, 2(1):101–118, June 1965. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/3211877. See correction [Ano65e].

Bhat:1965:SPO

[Bha65b] U. N. Bhat. On a stochastic process occurring in queueing systems. Journal of Applied Probability, 2(2):467–469, December 1965. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3212208.

Bhat:1968:EGM

[Bha68a] B. R. Bhat. On an extension of Gani’s model for attachment of phages to bacteria. Journal of Applied Probability, 5(3):572–578, December 1968. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3211922.

Bhat:1968:TBM

[Bha68b] U. Narayan Bhat. Transient behavior of multi-server queues with recurrent input and exponential service times. Journal of Applied Probability, 5(1):158–168, April 1968. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/3212083.

Bhat:1969:UIR

[Bha69] B. R. Bhat. Used item replacement policy. Journal of Applied Probability, 6(2):309–318, August 1969. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http:/

/www.jstor.org/stable/3212002.

Bhattacharjee:1980:CGD

[Bha80] Manish C. Bhattacharjee. On a characterization of gamma distributions via exponential mixtures. Journal of Applied Probability, 17(2):574–576, June 1980. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3213049.

Bhansali:1990:RBI

[Bha90] R. J. Bhansali. On a relationship between the inverse of a stationary covariance matrix and the linear interpolator. Journal of Applied Probability, 27(1):156–170, March 1990. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3214603.

Bhattacharjee:1991:SGV

[Bha91] M. C. Bhattacharjee. Some generalized variability orderings among life distributions with reliability applications. Journal of Applied Probability, 28(2):374–383, June 1991. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3214873.

Barmalzan:2018:UMS

[BHB18] Ghobad Barmalzan, Abedin Haidari, and Narayanaswamy Balakrishnan. Univariate and multivariate stochastic orderings of residual lifetimes of live components in sequential (n − r + 1)-out-of-n systems. Journal of Applied Probability, 55(3):834–844, September 2018. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL https:/

/www.cambridge.org/core/journals/journal-of-applied-
probability/article/univariate-and-multivariate-stochastic-
orderings-of-residual-lifetimes-of-live-components-in-
sequential-1outof-systems/E5913FA58DA2CE5A9B2EE8516EB192EC.

Bini:2018:WKC

[BHL^{+}18] Dario Bini, Jeffrey J. Hunter, Guy Latouche, Beatrice Meini, and Peter Taylor. Why is Kemeny’s constant a constant?

Journal of Applied Probability, 55(4):1025–1036, December 2018. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-

6072 (electronic). URL https://www.cambridge.org/core/
journals/journal-of-applied-probability/article/why-
is-kemenys-constant-a-constant/90B8A0AA175FCF4B0FA594A09590DC8A.

Bronstein:2006:DSO

[BHPZ06] Anne Laure Bronstein, Lane P. Hughston, Martijn R. Pistorius, and Mihail Zervos. Discretionary stopping of one-dimensional Itô diffusions with a staircase reward function. Journal of Applied Probability, 43(4):984–996, December 2006. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/27595794.

Benes:2020:DCG

[BHTLV20] Viktor Benes, Christoph Hofer-Temmel, Günter Last, and Jakub Vecera. Decorrelation of a class of Gibbs particle processes and asymptotic properties of U-statistics. Journal of Applied Probability, 57(3):928–955, September 2020. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL https:/

/www.cambridge.org/core/journals/journal-of-applied-
probability/article/decorrelation-of-a-class-of-gibbs-
particle-processes-and-asymptotic-properties-of-ustatistics/
9637E597627350238F4E33E31AEFD4E3.

Bhulai:2006:VFM

[Bhu06] Sandjai Bhulai. On the value function of the M/Cox(r)/1 queue. Journal of Applied Probability, 43(2):363–376, June 2006. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/27595737.

Brown:1998:VMR

[BHX98] Timothy C. Brown, Kais Hamza, and Aihua Xia. On the variance to mean ratio for random variables from Markov chains and point processes. Journal of Applied Probability, 35(2):303– 312, June 1998. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3215687.

Bao:2022:CRE

[BHZ22] Jianhai Bao, Xing Huang, and Shao-Qin Zhang. Convergence rate of the EM algorithm for SDEs with low regular drifts. Journal of Applied Probability, 59(2):447–470, June 2022. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL https://www.cambridge.org/core/journals/journal-
of-applied-probability/article/convergence-rate-of-the-
em-algorithm-for-sdes-with-low-regular-drifts/740D6EDD78AAA1BE8EC05C86A734E95B.

Bon:2005:APS

[BI05] Jean-Louis Bon and Abbas Illayk. Ageing properties and series systems. Journal of Applied Probability, 42(1):279–286, March 2005. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/30040787.

Bienvenue:1998:PIP

[Bie98] Alexis Bienvenüe. Un principe d’invariance pour une classe de marches p-corrélées sur Z^{d}. (French) [an invariance principle for a class of p-correlated walks on Z^{d}]. Journal of Applied Probability, 35(3):557–565, September 1998. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3215631.

Biggins:1977:CTB

[Big77a] J. D. Biggins. Chernoff’s theorem in the branching random walk. Journal of Applied Probability, 14(3):630–636, September 1977. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/3213469.

Biggins:1977:MCB

[Big77b] J. D. Biggins. Martingale convergence in the branching random walk. Journal of Applied Probability, 14(1):25–37, March 1977. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3213258.

Billard:1973:FMP

[Bil73] L. Billard. Factorial moments and probabilities for the general stochastic epidemic. Journal of Applied Probability, 10(2):277– 288, June 1973. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3212345.

Billard:1974:RE

[Bil74] L. Billard. The recurrent epidemic. Journal of Applied Probability, 11(3):568–571, September 1974. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3212701.

Billard:1975:GSE

[Bil75] L. Billard. General stochastic epidemic with recovery. Journal of Applied Probability, 12(1):29–38, March 1975. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3212404.

Billard:1976:SGE

[Bil76] L. Billard. A stochastic general epidemic in m sub-populations. Journal of Applied Probability, 13(3):567–572, September 1976. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/3212476.

Billard:1977:MDT

[Bil77a] L. Billard. Mean duration time for a general epidemic process. Journal of Applied Probability, 14(2):232–240, June 1977. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/3212994.

Billard:1977:LVP

[Bil77b] L. Billard. On Lotka–Volterra predator prey models. Journal of Applied Probability, 14(2):375–381, June 1977. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3213008.

Billard:1981:GTD

[Bil81] L. Billard. Generalized two-dimensional bounded birth and death processes and some applications. Journal of Applied Probability, 18(2):335–347, June 1981. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3213281.

Buchmann:2017:FLT

[BIM17] Boris Buchmann, Yuguang F. Ipsen, and Ross Maller. Functional laws for trimmed Lévy processes. Journal of Applied Probability, 54(3):873–889, September 2017. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL https:/

/www.cambridge.org/core/journals/journal-of-applied-
probability/article/functional-laws-for-trimmed-levy-
processes/BB595C5302D23CFEB57B88048AB47C98.

Bohun:2022:RTI

[BIMR22] Vladyslav Bohun, Alexander Iksanov, Alexander Marynych, and Bohdan Rashytov. Renewal theory for iterated perturbed random walks on a general branching process tree: intermediate generations. Journal of Applied Probability, 59(2):421–446, June 2022. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL https:/

/www.cambridge.org/core/journals/journal-of-applied-
probability/article/renewal-theory-for-iterated-perturbed-
random-walks-on-a-general-branching-process-tree-intermediate-
generations/EB7A6C82BA0DFECA4E3430E48DF7141B.

Bingham:1969:NSP

[Bin69] M. S. Bingham. A note on stochastic processes with independent increments taking values in an Abelian group. Journal of Applied Probability, 6(2):449–452, August 1969. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3212014.

Bingham:1978:CUD

[Bin78] M. S. Bingham. A characterization of the uniform distribution on the circle in the analysis of directional data. Journal of Applied Probability, 15(4):852–857, December 1978. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3213441.

Bingham:1988:LSB

[Bin88] N. H. Bingham. On the limit of a supercritical branching process. Journal of Applied Probability, 25A:215–228, 1988. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/3214158.

Bingham:1992:OGH

[Bin92] N. H. Bingham. Obituary: G. E. H. Reuter. Journal of Applied Probability, 29(3):754–757, September 1992. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/3214913.

Bingham:1994:WLT

[Bin94] N. H. Bingham. The work of Lajos Takács on probability theory. Journal of Applied Probability, 31A:29–39, 1994. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3214946.

Bingham:2005:DHC

[Bin05] N. H. Bingham. Doob: A half-century on. Journal of Applied Probability, 42(1):257–266, March 2005. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http:/

/www.jstor.org/stable/30040784.

Bisewski:2023:LBE

[Bis23] Krzysztof Bisewski. Lower bound for the expected supremum of fractional Brownian motion using coupling. Journal of Applied Probability, 60(4):1232–1248, December 2023. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL https://www.cambridge.org/core/journals/journal-
of-applied-probability/article/lower-bound-for-the-expected-
supremum-of-fractional-brownian-motion-using-coupling/
E7C7E6D07D8842EFE9C077AA2F28BB8E.

Bithell:1969:SMB

[Bit69] J. F. Bithell. A stochastic model for the breaking of molecular segments. Journal of Applied Probability, 6(1):59–73, April 1969. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3212276.

Berndtsson:1979:EGB

[BJ79] Bo Berndtsson and Peter Jagers. Exponential growth of a branching process usually implies stable age distribution. Journal of Applied Probability, 16(3):651–656, September 1979. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/3213093.

Alaya:2007:PAN

[BJ07] Mohamed Ben Alaya and Benjamin Jourdain. Probabilistic approximation of a nonlinear parabolic equation occurring in rheology. Journal of Applied Probability, 44(2):528–546, June 2007. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/27595859.

Baharian:2014:LBT

[BJ14] Golshid Baharian and Sheldon H. Jacobson. Limiting behavior of the target-dependent stochastic sequential assignment problem. Journal of Applied Probability, 51(4):943–953, December 2014. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://projecteuclid.org/euclid.jap/
1421763320.

Brown:2024:ECA

[BJ24] Austin Brown and Galin L. Jones. Exact convergence analysis for Metropolis–Hastings independence samplers in Wasserstein distances. Journal of Applied Probability, 61(1):33–54, March 2024. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-

6072 (electronic). URL https://www.cambridge.org/core/
journals/journal-of-applied-probability/article/exact-
convergence-analysis-for-metropolishastings-independence-
samplers-in-wasserstein-distances/89EC23E4455F6A7B5C112B96D0E98192.

Bohm:1993:ZAT

[BJM93] Walter Böhm, J. L. Jain, and S. G. Mohanty. On zero-avoiding transition probabilities of an r-node tandem queue: A combinatorial approach. Journal of Applied Probability, 30(3):737–741, September 1993. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3214780.

Berenhaut:2018:DWE

[BJMK18] Kenneth S. Berenhaut, Hongyi Jiang, Katelyn M. McNab, and Elizabeth J. Krizay. The degree-wise effect of a second step for a random walk on a graph. Journal of Applied Probability, 55(4):1203–1210, December 2018. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL https://www.cambridge.org/core/journals/journal-
of-applied-probability/article/degreewise-effect-of-
a-second-step-for-a-random-walk-on-a-graph/C6C7A881478934AA156AED8080765A55.

Bandyopadhyay:2020:SCI

[BJT20] Antar Bandyopadhyay, Svante Janson, and Debleena Thacker. Strong convergence of infinite color balanced urns under uniform ergodicity. Journal of Applied Probability, 57(3): 853–865, September 2020. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL https:/

/www.cambridge.org/core/journals/journal-of-applied-
probability/article/strong-convergence-of-infinite-color-
balanced-urns-under-uniform-ergodicity/836BF39100DA2C1882B5604246FB5503.

Bernard:2012:NBF

[BJV12] Carole Bernard, Xiao Jiang, and Steven Vanduffel. A note on ‘Improved Fréchet Bounds and Model-Free Pricing of Multi-Asset Options’ by Tankov (2011). Journal of Applied Probability, 49(3):866–875, September 2012. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/41713811. See [Tan11].

Blackwell:1964:MBP

[BK64] David Blackwell and David Kendall. The Martin boundary for Pólya’s urn scheme, and an application to stochastic population growth. Journal of Applied Probability, 1(2):284–296, December 1964. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3211860.

Brockwell:1973:LEP

[BK73a] P. J. Brockwell and W. H. Kuo. Labelling experiments and phase-age distributions for multiphase branching processes. Journal of Applied Probability, 10(4):739–747, December 1973. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/3212377.

Burville:1973:MSS

[BK73b] P. J. Burville and J. F. C. Kingman. On a model for storage and search. Journal of Applied Probability, 10(3):697–701, September 1973. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3212792.

Basu:1986:SRI

[BK86a] A. P. Basu and S. N. U. A. Kirmani. Some results involving HNBUE distributions. Journal of Applied Probability, 23(4): 1038–1044, December 1986. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.

org/stable/3214478.

Baxter:1986:BSP

[BK86b] Laurence A. Baxter and Chul Kim. Bounding the stochastic performance of continuum structure functions. I. Journal of Applied Probability, 23(3):660–669, September 1986. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3214004.

Baxter:1987:BSP

[BK87a] Laurence A. Baxter and Chul Kim. Bounding the stochastic performance of continuum structure functions. II. Journal of Applied Probability, 24(3):609–618, September 1987. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3214093.

Benes:1987:TPS

[BK87b] Václav E. Benes and Ioannis Karatzas. Transition probabilities for some ‘special’ diffusions. Journal of Applied Probability, 24 (4):888–898, December 1987. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.

org/stable/3214213.

Boutsikas:2000:GRB

[BK00] Michael V. Boutsikas and Markos V. Koutras. Generalized reliability bounds for coherent structures. Journal of Applied Probability, 37(3):778–794, September 2000. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3215613.

Bhulai:2003:SVF

[BK03] Sandjai Bhulai and Ger Koole. On the structure of value functions for threshold policies in queueing models. Journal of Applied Probability, 40(3):613–622, September 2003. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3215939.

Boutsikas:2006:ADD

[BK06] Michael V. Boutsikas and Markos V. Koutras. On the asymptotic distribution of the discrete scan statistic. Journal of Applied Probability, 43(4):1137–1154, December 2006. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/27595805.

Baeumer:2012:AMT

[BK12a] Boris Baeumer and Mihály Kovács. Approximating multivariate tempered stable processes. Journal of Applied Probability, 49 (1):167–183, March 2012. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/41713758.

Belomestny:2012:CLT

[BK12b] Denis Belomestny and Volker Krätschmer. Central limit theorems for law-invariant coherent risk measures. Journal of Applied Probability, 49(1):1–21, March 2012. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/41713749.

Baccelli:2021:PPD

[BK21] François Baccelli and Sanket S. Kalamkar. On point processes defined by angular conditions on Delaunay neighbors in the Poisson–Voronoi tessellation. Journal of Applied Probability, 58(4):952–965, December 2021. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL https:/

/www.cambridge.org/core/journals/journal-of-applied-
probability/article/on-point-processes-defined-by-angular-
conditions-on-delaunay-neighbors-in-the-poissonvoronoi-
tessellation/75010305F987F5C913679DB332155448.

Baharian:2016:SSA

[BKJ16] Golshid Baharian, Arash Khatibi, and Sheldon H. Jacobson. Sequential stochastic assignment problem with time-dependent random success rates. Journal of Applied Probability, 53(4): 1052–1063, December 2016. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://projecteuclid.

org/euclid.jap/1481132836.

Backhausz:2019:DPA

[BKK19] Ágnes Backhausz and Dávid Kunszenti-Kovács. On the dense preferential attachment graph models and their graphon induced counterpart. Journal of Applied Probability, 56(2):590–601, June 2019. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-

6072 (electronic). URL https://www.cambridge.org/core/
journals/journal-of-applied-probability/article/on-the-
dense-preferential-attachment-graph-models-and-their-
graphon-induced-counterpart/D9B3F5F9EB619CD1C91AA5DF4B94F2C3.

Bae:2003:ACU

[BKL03] Jongho Bae, Sunggon Kim, and Eui Yong Lee. Average cost under the P_{λ}, τ^{M} policy in a finite dam with compound Poisson inputs. Journal of Applied Probability, 40(2):519–526, June 2003. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3215808.

Blanchet:2019:RWP

[BKM19] Jose Blanchet, Yang Kang, and Karthyek Murthy. Robust Wasserstein profile inference and applications to machine learning. Journal of Applied Probability, 56(3):830–857, September 2019. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-

6072 (electronic). URL https://www.cambridge.org/core/

journals/journal-of-applied-probability/article/robust-
wasserstein-profile-inference-and-applications-to-machine-
learning/4024D05DE4681E67334E45D039295527.

Becker-Kern:2004:LTC

[BKMS04] Peter Becker-Kern, Mark M. Meerschaert, and Hans-Peter Scheffler. Limit theorem for continuous-time random walks with two time scales. Journal of Applied Probability, 41(2):455– 466, June 2004. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3216028.

Bertsimas:1991:TBP

[BKNZ91] Dimitris J. Bertsimas, Julian Keilson, Daisuke Nakazato, and Hongtao Zhang. Transient and busy period analysis of the GI/G/1 queue as a Hilbert factorization problem. Journal of Applied Probability, 28(4):873–885, December 1991. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3214690.

Baci:2022:LTR

[BKP^{+}22] Anastas Baci, Zakhar Kabluchko, Joscha Prochno, Mathias Sonnleitner, and Christoph Thäle. Limit theorems for random points in a simplex. Journal of Applied Probability, 59(3):685–701, September 2022. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL https:/

/www.cambridge.org/core/journals/journal-of-applied-
probability/article/limit-theorems-for-random-points-
in-a-simplex/EDA62840DD403AFD1669C159B5AA009C.

Bassan:2002:SBN

[BKS02] Bruno Bassan, Subhash Kochar, and Fabio Spizzichino. Some bivariate notions of IFR and DMRL and related properties. Journal of Applied Probability, 39(3):533–544, September 2002. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/3216076.

Bean:1997:BPL

[BKT97a] N. G. Bean, F. P. Kelly, and P. G. Taylor. Braess’s paradox in a loss network. Journal of Applied Probability, 34(1):155– 159, March 1997. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3215183.

Berezner:1997:PFL

[BKT97b] S. A. Berezner, A. E. Krzesinski, and P. G. Taylor. A productform ’loss network’ with a form of queueing. Journal of Applied Probability, 34(4):1075–1078, December 1997. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3215020.

Berezner:1998:IEF

[BKT98] S. A. Berezner, A. E. Krzesinski, and P. G. Taylor. On the inverse of Erlang’s function. Journal of Applied Probability, 35 (1):246–252, March 1998. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3215564.

Beer:1973:CND

[BL73] S. Beer and E. Lukacs. Characterizations of the normal distribution by suitable transformations. Journal of Applied Probability, 10(1):100–108, March 1973. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3212498.

Baxter:1989:SFF

[BL89] Laurence A. Baxter and Seung Min Lee. Structure functions with finite minimal vector sets. Journal of Applied Probability, 26(1):196–201, March 1989. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3214331.

Bruss:1998:SBW

[BL98] F. Thomas Bruss and Guy Louchard. Sharp bounds for winning probabilities in the competitive rank selection problem. Journal of Applied Probability, 35(4):1007–1011, December 1998. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/3215537.

Belyaev:2000:WAS

[BL00] Yuri Belyaev and Sara Sjöstedt-De Luna. Weakly approaching sequences of random distributions. Journal of Applied Probability, 37(3):807–822, September 2000. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3215615.

Berenhaut:2001:GRC

[BL01] Kenneth S. Berenhaut and Robert Lund. Geometric renewal convergence rates from hazard rates. Journal of Applied Probability, 38(1):180–194, March 2001. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3215750.

Bruss:2003:OSP

[BL03] F. Thomas Bruss and Guy Louchard. Optimal stopping on patterns in strings generated by independent random variables. Journal of Applied Probability, 40(1):49–72, March 2003. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/3215841.

Behrens:2009:MDW

[BL09a] Sarah Behrens and Matthias Löwe. Moderate deviations for word counts in biological sequences. Journal of Applied Probability, 46(4):1020–1037, December 2009. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/25662478.

Britton:2009:ESB

[BL09b] Tom Britton and Mathias Lindholm. The early stage behaviour of a stochastic SIR epidemic with term-time forcing. Journal of Applied Probability, 46(4):975–992, December 2009. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/25662475. See correc-

tion [BL15].

Blanchet:2010:EIS

[BL10] Jose Blanchet and Jingchen Liu. Efficient importance sampling in ruin problems for multidimensional regularly varying random walks. Journal of Applied Probability, 47(2):301–322, June 2010. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/20778481.

Blanchet:2011:ESM

[BL11] Jose Blanchet and Chenxin Li. Efficient simulation for the maximum of infinite horizon discrete-time Gaussian processes. Journal of Applied Probability, 48(2):467–489, June 2011. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/23065897.

Borovkov:2012:RFS

[BL12] K. Borovkov and G. Last. On Rice’s formula for stationary multivariate piecewise smooth processes. Journal of Applied Probability, 49(2):351–363, June 2012. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/41713772.

Bauerle:2013:OPF

[BL13] Nicole Bäuerle and Zejing Li. Optimal portfolios for financial markets with Wishart volatility. Journal of Applied Probability, 50(4):1025–1043, December 2013. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

projecteuclid.org/euclid.jap/1389370097.

Britton:2015:CES

[BL15] T. Britton and M. Lindholm. Correction: The early stage behaviour of a stochastic SIR epidemic with term-time forcing. Journal of Applied Probability, 52(4):1202–1203, December 2015. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-

6072 (electronic). URL http://projecteuclid.org/euclid.
jap/1450802764. See [BL09b].

Buchmann:2021:NWS

[BL21] Boris Buchmann and Kevin W. Lu. Necessity of weak subordination for some strongly subordinated Lévy processes. Journal of Applied Probability, 58(4):868–879, December 2021. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL https://www.cambridge.org/core/journals/
journal-of-applied-probability/article/necessity-of-
weak-subordination-for-some-strongly-subordinated-levy-
processes/8921C32CB91AD5594F213A378D11348E.

Blackwell:1953:ECE

[Bla53] David Blackwell. Equivalent comparisons of experiments. Annals of Mathematical Statistics, 24(2):265–272, June 1953. CO-DEN AASTAD. ISSN 0003-4851. URL http://projecteuclid.

org/euclid.aoms/1177729032. See remarks [LS08b].

Blanc:1984:AAQ

[Bla84] J. P. C. Blanc. Asymptotic analysis of a queueing system with a two-dimensional state space. Journal of Applied Probability, 21(4):870–886, December 1984. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.

org/stable/3213703.

Blanc:1987:NWT

[Bla87] J. P. C. Blanc. A note on waiting times in systems with queues in parallel. Journal of Applied Probability, 24(2):540–546, June 1987. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3214278.

Blackwell:1992:CHS

[Bla92] Paul Blackwell. Convex hulls of selected subsets of a Poisson process. Journal of Applied Probability, 29(4):814–824, December 1992. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3214714.

Bladt:1994:MSS

[Bla94] M. Bladt. Multivariate self-similar processes: Second-order theory. Journal of Applied Probability, 31(1):139–147, March 1994. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/3215241.

Bar-Lev:2010:CRE

[BLBL10] Shaul K. Bar-Lev, Onno Boxma, and Gérard Letac. A characterization related to the equilibrium distribution associated with a polynomial structure. Journal of Applied Probability, 47 (1):293–299, March 2010. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/25662516.

Blight:1989:TSF

[Bli89] P. A. Blight. Time series formed from the superposition of discrete renewal processes. Journal of Applied Probability, 26 (1):189–195, March 1989. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3214330.

Billard:1979:NLS

[BLL79] L. Billard, H. Lacayo, and N. A. Langberg. A new look at the simple epidemic process. Journal of Applied Probability, 16 (1):198–202, March 1979. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/ stable/3213387. See corrections [Ano79e].

Billard:1980:DAB

[BLL80a] L. Billard, H. Lacayo, and N. A. Langberg. The discrete asymptotic behaviour of a simple batch epidemic process. Journal of Applied Probability, 17(1):25–32, March 1980. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3212921.

Billard:1980:GSE

[BLL80b] L. Billard, H. Lacayo, and N. A. Langberg. Generalizations of the simple epidemic process. Journal of Applied Probability, 17 (4):1072–1078, December 1980. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.

org/stable/3213216.

Burman:1984:IBP

[BLL84] D. Y. Burman, J. P. Lehoczky, and Y. Lim. Insensitivity of blocking probabilities in a circuit-switching network. Journal of Applied Probability, 21(4):850–859, December 1984. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3213701.

Boxma:2020:TCR

[BLM20] Onno Boxma, Andreas Löpker, and Michel Mandjes. On two classes of reflected autoregressive processes. Journal of Applied Probability, 57(2):657–678, June 2020. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL https:/

/www.cambridge.org/core/journals/journal-of-applied-
probability/article/on-two-classes-of-reflected-autoregressive-
processes/A33FEF774254A290C1998746676D277D.

Blomqvist:1973:HTR

[Blo73] Nils Blomqvist. A heavy traffic result for the finite dam. Journal of Applied Probability, 10(1):223–228, March 1973. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/3212511. See remarks

[Spe73b].

Blom:1982:MNR

[Blo82] Gunnar Blom. On the mean number of random digits until a given sequence occurs. Journal of Applied Probability, 19(1):136– 143, March 1982. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3213923.

Blom:1984:SOW

[Blo84] Gunnar Blom. On the stochastic ordering of waiting times for patterns in sequences of random digits. Journal of Applied Probability, 21(4):730–737, December 1984. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/3213691.

Bloomfield:1986:NSA

[Blo86] Peter Bloomfield. Non-singularity and asymptotic independence. Journal of Applied Probability, 23A:9–21, 1986. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3214339.

Blondia:1991:FCV

[Blo91] C. Blondia. Finite capacity vacation models with non-renewal input. Journal of Applied Probability, 28(1):174–197, March 1991. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/3214749.

Barbour:1991:SOO

[BLR91] A. D. Barbour, T. Lindvall, and L. C. G. Rogers. Stochastic ordering of order statistics. Journal of Applied Probability, 28 (2):278–286, June 1991. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3214866.

Block:1990:MCB

[BLS90] Henry W. Block, Naftali A. Langberg, and Thomas H. Savits. Maintenance comparisons: Block policies. Journal of Applied Probability, 27(3):649–657, September 1990. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/3214548.

Block:1993:RRP

[BLS93] Henry W. Block, Naftali A. Langberg, and Thomas H. Savits. Repair replacement policies. Journal of Applied Probability, 30 (1):194–206, March 1993. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3214632.

Block:2003:IFB

[BLS03] Henry W. Block, Yulin Li, and Thomas H. Savits. Initial and final behaviour of failure rate functions for mixtures and systems. Journal of Applied Probability, 40(3):721–740, September 2003. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3215946.

Block:2012:GTF

[BLS12] Henry W. Block, Naftali A. Langberg, and Thomas H. Savits. A Glaser twist: Focus on the mixture parameters. Journal of Applied Probability, 49(4):1144–1155, December 2012. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/41713834.

Block:2015:LFR

[BLS15] Henry W. Block, Naftali A. Langberg, and Thomas H. Savits. The limiting failure rate for a convolution of life distributions. Journal of Applied Probability, 52(3):894–898, September 2015. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://projecteuclid.org/euclid.jap/
1445543854.

Bar-Lev:2013:CLT

[BLSGS13] Shaul K. Bar-Lev, Ernst Schulte-Geers, and Wolfgang Stadje. Conditional limit theorems for the terms of a random walk revisited. Journal of Applied Probability, 50(3):871–882, September 2013. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-

6072 (electronic). URL http://projecteuclid.org/euclid.
jap/1378401242.

Block:2008:CMB

[BLSW08] Henry W. Block, Yulin Li, Thomas H. Savits, and Jie Wang. Continuous mixtures with bathtub-shaped failure rates. Journal of Applied Probability, 45(1):260–270, March 2008. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/27595938.

Block:2010:CME

[BLSW10] Henry W. Block, Naftali A. Langberg, Thomas H. Savits, and Jie Wang. Continuous mixtures of exponentials and IFR gammas having bathtub-shaped failure rates. Journal of Applied Probability, 47(4):899–907, December 2010. CODEN JPRBAM.

ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/29777414.

Britton:2011:DND

[BLT11] Tom Britton, Mathias Lindholm, and Tatyana Turova. A dynamic network in a dynamic population: Asymptotic properties. Journal of Applied Probability, 48(4):1163–1178, December 2011. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/23066449.

Brockwell:1966:SPP

[BM66] P. J. Brockwell and J. E. Moyal. A stochastic population process and its application to bubble chamber measurements. Journal of Applied Probability, 3(1):280–284, June 1966. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3212054.

Beutler:1982:MPF

[BM82] Frederick J. Beutler and Benjamin Melamed. Multivariate Poisson flows on Markov step processes. Journal of Applied Probability, 19(2):289–300, June 1982. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3213481.

Buchta:1984:RPB

[BM84] C. Buchta and J. Müller. Random polytopes in a ball. Journal of Applied Probability, 21(4):753–762, December 1984. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3213693.

Baccelli:1989:SPA

[BM89a] François Baccelli and William A. Massey. A sample path analysis of the M/M/1 queue. Journal of Applied Probability, 26(2):418– 422, June 1989. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3214049.

Berezner:1989:SIS

[BM89b] S. A. Berezner and V. A. Malyshev. The stability of infiniteserver networks with random routing. Journal of Applied Probability, 26(2):363–371, June 1989. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3214041.

Berg:1990:RMA

[BM90] Sven Berg and Ljuben Mutafchiev. Random mappings with an attracting center: Lagrangian distributions and a regression function. Journal of Applied Probability, 27(3):622–636, September 1990. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3214546.

Bohm:1994:TAM

[BM94] W. Böhm and S. G. Mohanty. Transient analysis of M/M/1 queues in discrete time with general server vacations. Journal of Applied Probability, 31A:115–129, 1994. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/3214952.

Barnes:1997:SPD

[BM97] John A. Barnes and Richard Meili. A stationary Poisson departure process from a minimally delayed infinite server queue with non-stationary Poisson arrivals. Journal of Applied Probability, 34(3):767–772, September 1997. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3215101.

Bloch-Mercier:2001:MMP

[BM01a] Sophie Bloch-Mercier. Monotone Markov processes with respect to the reversed hazard rate ordering: An application to reliability. Journal of Applied Probability, 38(1):195–208, March 2001. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/3215751.

Bremaud:2001:HBP

[BM01b] Pierre Brémaud and Laurent Massoulié. Hawkes branching point processes without ancestors. Journal of Applied Probability, 38 (1):122–135, March 2001. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3215746.

Ball:2005:SDP

[BM05] Frank Ball and Robin K. Milne. Simple derivations of properties of counting processes associated with Markov renewal processes. Journal of Applied Probability, 42(4):1031–1043, December 2005. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/30040826.

Bender:2009:IVC

[BM09] Christian Bender and Tina Marquardt. Integrating volatility clustering into exponential Lévy models. Journal of Applied Probability, 46(3):609–628, September 2009. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/25662451.

Beghin:2014:FDP

[BM14] Luisa Beghin and Claudio Macci. Fractional discrete processes: compound and mixed Poisson representations. Journal of Applied Probability, 51(1):19–36, March 2014. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

projecteuclid.org/euclid.jap/1395771411.

Backhausz:2015:APR

[BM15a] Ágnes Backhausz and Tamás F. Móri. Asymptotic properties of a random graph with duplications. Journal of Applied Probability, 52(2):375–390, June 2015. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

projecteuclid.org/euclid.jap/1437658604.

Buonaguidi:2015:DPN

[BM15b] Bruno Buonaguidi and Pietro Muliere. On the disorder problem for a negative binomial process. Journal of Applied Probability, 52(1):167–179, March 2015. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://projecteuclid.

org/euclid.jap/1429282613.

Balaji:2017:GIR

[BM17] Hrishikesh Balaji and Hosam Mahmoud. The Gini index of random trees with an application to caterpillars. Journal of Applied Probability, 54(3):701–709, September 2017. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL https:/

/www.cambridge.org/core/journals/journal-of-applied-
probability/article/gini-index-of-random-trees-with-
an-application-to-caterpillars/4EE9272B5F5383D1C3512DE9DAB35427.

Blanchet:2018:ESM

[BM18] Jose Blanchet and Karthyek Murthy. Exact simulation of multidimensional reflected Brownian motion. Journal of Applied Probability, 55(1):137–156, March 2018. CODEN JPRBAM.

ISSN 0021-9002 (print), 1475-6072 (electronic). URL https:/

/www.cambridge.org/core/journals/journal-of-applied-
probability/article/exact-simulation-of-multidimensional-
reflected-brownian-motion/EFD1DCA10291E41D40F80CD6449AA9A7.

Boudabra:2020:MME

[BM20] Maher Boudabra and Greg Markowsky. Maximizing the p-th moment of the exit time of planar Brownian motion from a given domain. Journal of Applied Probability, 57 (4):1135–1149, December 2020. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL https:/

/www.cambridge.org/core/journals/journal-of-applied-
probability/article/maximizing-the-pth-moment-of-the-
exit-time-of-planar-brownian-motion-from-a-given-domain/
692F3E1354805A548E7BE91FC87745E7.

Boxma:2023:DLP

[BM23] Onno Boxma and Michel Mandjes. A decomposition for Lévy processes inspected at Poisson moments. Journal of Applied Probability, 60(2):557–569, June 2023. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL https:/

/www.cambridge.org/core/journals/journal-of-applied-
probability/article/decomposition-for-levy-processes-
inspected-at-poisson-moments/6793F2CD7EAE7A7145BA73D78A9BD9C6.

Baeumer:2009:STD

[BMN09] Boris Baeumer, Mark M. Meerschaert, and Erkan Nane. Space– time duality for fractional diffusion. Journal of Applied Probability, 46(4):1100–1115, December 2009. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/25662483.

Bjork:1987:OCS

[BMP87] T. Björk, J. Myhrman, and M. Persson. Optimal consumption with stochastic prices in continuous time. Journal of Applied Probability, 24(1):35–47, March 1987. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3214057.

Belzunce:2007:RPP

[BMPR07] Félix Belzunce, Helena Martínez-Puertas, and José M. Ruiz. Reversed preservation properties for series and parallel systems. Journal of Applied Probability, 44(4):928–937, December 2007.

CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/27595897.

Boxma:2016:CRA

[BMR16] Onno Boxma, Michel Mandjes, and Josh Reed. On a class of reflected AR(1) processes. Journal of Applied Probability, 53(3): 818–832, September 2016. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://projecteuclid.

org/euclid.jap/1476370778.

Belzunce:2019:BHR

[BMR19] Félix Belzunce and Carolina Martínez-Riquelme. Bounds for the hazard rate and the reversed hazard rate of the convolution of dependent random lifetimes. Journal of Applied Probability, 56(4):1033–1043, December 2019. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL https:/

/www.cambridge.org/core/journals/journal-of-applied-
probability/article/bounds-for-the-hazard-rate-and-the-
reversed-hazard-rate-of-the-convolution-of-dependent-
random-lifetimes/8B873D450B8B031F8740D639C75E0CEA.

Belzunce:2016:SCC

[BMRRS16] Félix Belzunce, Carolina Martínez-Riquelme, José M. Ruiz, and Miguel A. Sordo. On sufficient conditions for the comparison in the excess wealth order and spacings. Journal of Applied Probability, 53(1):33–46, March 2016. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

projecteuclid.org/euclid.jap/1457470556.

Block:1993:BMP

[BMS93] Henry W. Block, Jie Mi, and Thomas H. Savits. Burn-in and mixed populations. Journal of Applied Probability, 30(3):692– 702, September 1993. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3214775.

Balaji:2010:PDG

[BMT10] Srinivasan Balaji, Hosam Mahmoud, and Zhang Tong. Phases in the diffusion of gases via the Ehrenfest urn model. Journal of Applied Probability, 47(3):841–855, September 2010. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/20778520.

Ball:1993:EDO

[BMY93] Frank Ball, Robin K. Milne, and Geoffrey F. Yeo. On the exact distribution of observed open times in single ion channel models. Journal of Applied Probability, 30(3):529–537, September 1993. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3214763.

Ball:2002:MSM

[BMY02] F. Ball, R. K. Milke, and G. F. Yeo. Multivariate semi-Markov analysis of burst properties of multiconductance single ion channels. Journal of Applied Probability, 39(1):179–196, March 2002. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/3215926.

Besag:1982:PPL

[BMZ82] Julian Besag, Robin Milne, and Stan Zachary. Point process limits of lattice processes. Journal of Applied Probability, 19 (1):210–216, March 1982. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3213930.

Barbe:2007:AED

[BMZ07] Ph. Barbe, W. P. McCormick, and C. Zhang. Asymptotic expansions for distributions of compound sums of random variables with rapidly varying subexponential distribution. Journal of Applied Probability, 44(3):670–684, September 2007. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/27595874.

Basrak:2023:ERC

[BMZ23] Bojan Basrak, Nikolina Milincević, and Petra Zugec. On extremes of random clusters and marked renewal cluster processes. Journal of Applied Probability, 60(2):367–381, June 2023. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL https://www.cambridge.org/core/journals/journal-
of-applied-probability/article/on-extremes-of-random-
clusters-and-marked-renewal-cluster-processes/BAE3340A3FCE9EC1D2CA422B2CDBDDEF.

Brown:1988:SPM

[BN88] Timothy C. Brown and M. Gopalan Nair. A simple proof of the multivariate random time change theorem for point processes.

Journal of Applied Probability, 25(1):210–214, March 1988. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/3214247.

Bunge:1992:DSP

[BN92] J. A. Bunge and H. N. Nagaraja. Dependence structure of Poisson-paced records. Journal of Applied Probability, 29(3): 587–596, September 1992. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3214895.

Bhat:1994:QNA

[BN94] U. Narayan Bhat and Richard E. Nance. A queueing network analysis of dynamic reconfigurability in a hierarchical information network. Journal of Applied Probability, 31A:99–114, 1994. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/3214951.

Borovkov:2002:NAC

[BN02] K. Borovkov and A. Novikov. On a new approach to calculating expectations for option pricing. Journal of Applied Probability, 39(4):889–895, December 2002. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.

org/stable/3216011.

Borovkov:2005:EBA

[BN05] K. Borovkov and A. Novikov. Explicit bounds for approximation rates of boundary crossing probabilities for the Wiener process. Journal of Applied Probability, 42(1):82–92, March 2005. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/30040771.

Barlevy:2006:CRR

[BN06] Gadi Barlevy and H. N. Nagaraja. Characterizations in a random record model with a nonidentically distributed initial record. Journal of Applied Probability, 43(4):1119–1136, December 2006. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/27595804.

Burkschat:2013:DSC

[BN13] M. Burkschat and J. Navarro. Dynamic signatures of coherent systems based on sequential order statistics. Journal of Applied Probability, 50(1):??, March 2013. CODEN JPRBAM.

ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

projecteuclid.org/euclid.jap/1363784438.

Budhiraja:2015:LDM

[BN15] Amarjit Budhiraja and Pierre Nyquist. Large deviations for multidimensional state-dependent shot-noise processes. Journal of Applied Probability, 52(4):1097–1114, December 2015. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://projecteuclid.org/euclid.jap/1450802755.

Baddeley:2017:PSA

[BN17] Adrian Baddeley and Gopalan Nair. Poisson-saddlepoint approximation for Gibbs point processes with infinite-order interaction: in memory of Peter Hall. Journal of Applied Probability, 54(4):1008–1026, December 2017. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL https:/

/www.cambridge.org/core/journals/journal-of-applied-
probability/article/poissonsaddlepoint-approximation-
for-gibbs-point-processes-with-infiniteorder-interaction-
in-memory-of-peter-hall/36141CB9DD4480D7B8567560D2641F49.

Barndorff-Nielsen:2009:BVG

[BNCPW09] Ole E. Barndorff-Nielsen, José Manuel Corcuera, Mark Podolskij, and Jeannette H. C. Woerner. Bipower variation for Gaussian processes with stationary increments. Journal of Applied Probability, 46(1):132–150, March 2009. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http:/

/www.jstor.org/stable/25662412.

Barndorff-Nielsen:2004:PVS

[BNGS04] Ole E. Barndorff-Nielsen, Svend Erik Graversen, and Neil Shephard. Power variation and stochastic volatility: A review and some new results. Journal of Applied Probability, 41A:133–143, 2004. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3215973.

Barndorff-Nielsen:2005:BTP

[BNL05] O. E. Barndorff-Nielsen and N. N. Leonenko. Burgers’ turbulence problem with linear or quadratic external potential. Journal of Applied Probability, 42(2):550–565, June 2005. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/30040809.

Bucklew:1990:MCS

[BNS90] James A. Bucklew, Peter Ney, and John S. Sadowsky. Monte Carlo simulation and large deviations theory for uniformly recurrent Markov chains. Journal of Applied Probability, 27(1): 44–59, March 1990. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3214594.

Bremaud:2002:RCE

[BNT02] P. Brémaud, G. Nappo, and G. L. Torrisi. Rate of convergence to equilibrium of marked Hawkes processes. Journal of Applied Probability, 39(1):123–136, March 2002. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http:/

/www.jstor.org/stable/3215923.

Blanghaps:2013:STE

[BNW13] Nafna Blanghaps, Yuval Nov, and Gideon Weiss. Sojourn time estimation in an M/G/∞ queue with partial information. Journal of Applied Probability, 50(4):1044–1056, December 2013. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-

6072 (electronic). URL http://projecteuclid.org/euclid.
jap/1389370098.

Barndorff-Nielsen:1969:NBP

[BNY69] Ole Barndorff-Nielsen and G. F. Yeo. Negative binomial processes. Journal of Applied Probability, 6(3):633–647, December 1969. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3212108.

See correction [Ano70d].

Bruss:1990:MUG

[BO90] F. Thomas Bruss and Colm Art O’Cinneide. On the maximum and its uniqueness for geometric random samples. Journal of Applied Probability, 27(3):598–610, September 1990. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3214544.

Ball:1999:DGF

[BO99] Frank Ball and Philip O’Neill. The distribution of general final state random variables for stochastic epidemic models. Journal of Applied Probability, 36(2):473–491, June 1999. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3215469.

Baccelli:2018:RRD

[BO18] François Baccelli and Eliza O’Reilly. Reach of repulsion for determinantal point processes in high dimensions. Journal of Applied Probability, 55(3):760–788, September 2018. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL https://www.cambridge.org/core/journals/journal-
of-applied-probability/article/reach-of-repulsion-for-
determinantal-point-processes-in-high-dimensions/E4212D64EEC3E9A39D265DAE6C2EEC0B.

Boes:1969:NOQ

[Boe69] Duane C. Boes. Note on the output of a queuing system. Journal of Applied Probability, 6(2):459–461, August 1969. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3212016.

Bofinger:1965:SMT

[Bof65] Eve Bofinger. Sufficiency for multinomial and transition probabilities. Journal of Applied Probability, 2(2):470–474, December 1965. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3212209.

Bohm:2000:CRW

[Böh00] W. Böhm. The correlated random walk with boundaries: A combinatorial solution. Journal of Applied Probability, 37(2):470– 479, June 2000. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3215721.

Boland:2001:SIM

[Bol01] Philip J. Boland. Signatures of indirect majority systems. Journal of Applied Probability, 38(2):597–603, June 2001. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3215909.

Bondesson:1981:OTQ

[Bon81] Lennart Bondesson. On occupation times for quasi-Markov processes. Journal of Applied Probability, 18(1):297–301, March 1981. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3213192.

Bonsdorff:1989:COV

[Bon89] Heikki Bonsdorff. A comparison of the ordinary and a varying parameter exponential smoothing. Journal of Applied Probability, 26(4):784–792, December 1989. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3214383.

Bonin:2002:LDT

[Bon02] Oliver Bonin. Large deviation theorems for weighted sums applied to a geographical problem. Journal of Applied Probability, 39(2):251–260, June 2002. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3216093.

Bondareva:2013:NLB

[Bon13] M. Bondareva. Nondecreasing lower bound on the Poisson cumulative distribution function for z standard deviations above the mean. Journal of Applied Probability, 50(4):909–917, December 2013. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-

6072 (electronic). URL http://projecteuclid.org/euclid.
jap/1389370089.

Book:1979:SGP

[Boo79] Stephen A. Book. The stochastic geyser problem for sample quantiles. Journal of Applied Probability, 16(2):445–448, June 1979. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3212914.

Booth:1989:LBD

[Boo89] James G. Booth. On the limiting behaviour of Downton’s Carrier epidemic in the case of a general infection mechanism. Journal of Applied Probability, 26(3):625–630, September 1989. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3214419.

Booth:1992:NOC

[Boo92] James G. Booth. A note on a one-compartment model with clustering. Journal of Applied Probability, 29(3):535–542, September 1992. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3214891.

Ball:2007:SEM

[BOP07] Frank Ball, Philip D. O’Neill, and James Pike. Stochastic epidemic models in structured populations featuring dynamic vaccination and isolation. Journal of Applied Probability, 44(3): 571–585, September 2007. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/27595867.

Borkar:1992:PRO

[Bor92] Vivek S. Borkar. Pathwise recurrence orders and simulated annealing. Journal of Applied Probability, 29(2):472–476, June 1992. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3214585.

Borovkov:1994:SRV

[Bor94] K. A. Borovkov. On simulation of random vectors with given densities in regions and on their boundaries. Journal of Applied Probability, 31(1):205–220, March 1994. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http:/

/www.jstor.org/stable/3215247.

Borovkov:1995:CTM

[Bor95] K. A. Borovkov. On crossing times for multidimensional walks with skip-free components. Journal of Applied Probability, 32(4): 991–1006, December 1995. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3215213.

Borovkov:1996:AEF

[Bor96] A. A. Borovkov. Asymptotic expansions for functionals of dilation of point processes. Journal of Applied Probability, 33 (2):573–591, June 1996. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3215080.

Borovkov:1999:RRP

[Bor99] K. Borovkov. On records and related processes for sequences with trends. Journal of Applied Probability, 36(3):668–681, September 1999. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3215432.

Belzunce:2001:NSC

[BOR01] Félix Belzunce, Eva M. Ortega, and José M. Ruiz. A note on stochastic comparisons of excess lifetimes of renewal processes. Journal of Applied Probability, 38(3):747–753, September 2001. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/3216126. See

correction [Ano02e].

Borovkov:2004:PTF

[Bor04] K. Borovkov. On pair and tuple formation under independent Poisson or renewal arrival processes. Journal of Applied Probability, 41(4):1138–1144, December 2004. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/4141385.

Borovkov:2021:GPA

[Bor21] Konstantin Borovkov. Gaussian process approximations for multicolor Pólya urn models. Journal of Applied Probability, 58(1):274–286, March 2021. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL https:/

/www.cambridge.org/core/journals/journal-of-applied-
probability/article/gaussian-process-approximations-
for-multicolor-polya-urn-models/7DADBB7C6C46CFD07A70C0F6A6FF50CC.

Boucherie:1995:PFB

[Bou95] Richard J. Boucherie. Product forms based on backward traffic equations. Journal of Applied Probability, 32(2):508–518, June 1995. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3215304.

Bovier:1999:SUB

[Bov99] Anton Bovier. Sharp upper bounds on perfect retrieval in the Hopfield model. Journal of Applied Probability, 36(3):941–950, September 1999. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3215454.

Bowditch:2018:QCL

[Bow18] Adam Bowditch. A quenched central limit theorem for biased random walks on supercritical Galton–Watson trees. Journal of Applied Probability, 55(2):610–626, June 2018. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL https://www.cambridge.org/core/journals/journal-
of-applied-probability/article/quenched-central-limit-
theorem-for-biased-random-walks-on-supercritical-galtonwatson-
trees/9FB68B57082048E39D3577CBADA18EF9.

Boxma:1975:SSQ

[Box75] O. J. Boxma. The single-server queue with random service output. Journal of Applied Probability, 12(4):763–778, December 1975. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3212727.

Boylan:1969:STS

[Boy69] Edward S. Boylan. Stability theorems for solutions to the optimal inventory equation. Journal of Applied Probability, 6 (1):211–217, April 1969. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3212289.

Boylan:1970:FET

[Boy70] Edward S. Boylan. Further existence theorems for the optimal inventory equation. Journal of Applied Probability, 7(2):469– 472, August 1970. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3211984.

Boyko:1987:CPG

[Boy87] Roman Boyko. Comment on paper by Gani and Tin. Journal of Applied Probability, 24(2):554, June 1987. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/3214281. See [GT85].

Brown:1983:IR

[BP83] Mark Brown and Frank Proschan. Imperfect repair. Journal of Applied Probability, 20(4):851–859, December 1983. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3213596.

Bohm:1996:SRW

[BP96a] W. Böhm and W. Panny. Simple random walk statistics. Part II: Continuous time results. Journal of Applied Probability, 33 (2):331–339, June 1996. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3215057.

Borovkov:1996:IOA

[BP96b] K. Borovkov and D. Pfeifer. On improvements of the order of approximation in the Poisson limit theorem. Journal of Applied Probability, 33(1):146–155, March 1996. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http:/

/www.jstor.org/stable/3215272.

Barbour:1997:DMH

[BP97] A. D. Barbour and R. M. Phatarfod. Dimensioning a multiple hashing scheme. Journal of Applied Probability, 34(2):477– 486, June 1997. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3215386.

Biagini:1999:LRM

[BP99a] F. Biagini and M. Pratelli. Local risk minimization and Numéraire. Journal of Applied Probability, 36(4):1126–1139, December 1999. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3215583.

Borges:1999:CAS

[BP99b] Carlos F. Borges and Craig S. Peters. Computing approximate stationary distributions for discrete Markov processes with banded infinitesimal generators. Journal of Applied Probability, 36(4):1086–1100, December 1999. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3215580.

Bruss:2000:SSL

[BP00] F. Thomas Bruss and Davy Paindaveine. Selecting a sequence of last successes in independent trials. Journal of Applied Probability, 37(2):389–399, June 2000. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3215715.

Bon:2001:CNF

[BP01] Jean-Louis Bon and Eugen Păltănea. Convergence of the number of failed components in a Markov system with nonidentical components. Journal of Applied Probability, 38(4):882–897, December 2001. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3215771.

Barrera:2004:DSC

[BP04] Javiera Barrera and Christian Paroissin. On the distribution of the search cost for the move-to-front rule with random weights. Journal of Applied Probability, 41(1):250–262, March 2004. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/3215829.

Barbour:2010:TVA

[BP10] A. D. Barbour and P. K. Pollett. Total variation approximation for quasi-stationary distributions. Journal of Applied Probability, 47(4):934–946, December 2010. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/29777417.

Berzunza:2016:ABN

[BP16] Gabriel Berzunza and Juan Carlos Pardo. Asymptotic behaviour near extinction of continuous-state branching processes. Journal of Applied Probability, 53(2):381–391, June 2016. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://projecteuclid.org/euclid.jap/1466172861.

Bakhtin:2020:TET

[BPG20] Yuri Bakhtin and Zsolt Pajor-Gyulai. Tails of exit times from unstable equilibria on the line. Journal of Applied Probability, 57(2):477–496, June 2020. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL https:/

/www.cambridge.org/core/journals/journal-of-applied-
probability/article/tails-of-exit-times-from-unstable-
equilibria-on-the-line/D123EC7CC1FBD51CB04D464062F0648F.

Baurdoux:2016:GSD

[BPPR16] Erik J. Baurdoux, Juan Carlos Pardo, José Luis Pérez, and Jean-François Renaud. Gerber–Shiu distribution at Parisian ruin for Lévy insurance risk processes. Journal of Applied Probability, 53(2):572–584, June 2016. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://projecteuclid.

org/euclid.jap/1466172875.

Brown:2008:RPM

[BPR08] Mark Brown, Erol A. Peköz, and Sheldon M. Ross. A random permutation model arising in chemistry. Journal of Applied Probability, 45(4):1060–1070, December 2008. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/27596008.

Brockwell:1980:IID

[BPS80] P. J. Brockwell and N. Pacheco-Santiago. Invariant imbedding and dams with Markovian input rate. Journal of Applied Probability, 17(3):778–789, September 1980. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3212971.

Ba:2012:BTE

[BPS12] Mamadou Ba, Étienne Pardoux, and Ahmadou Bamba Sow. Binary trees, exploration processes, and an extended Ray–Knight theorem. Journal of Applied Probability, 49(1):210–225, March 2012. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/41713761.

Bender:2015:SCP

[BPS15] Christian Bender, Mikko S. Pakkanen, and Hasanjan Sayit. Sticky continuous processes have consistent price systems. Journal of Applied Probability, 52(2):586–594, June 2015. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://projecteuclid.org/euclid.jap/1437658617.

Boxma:2023:SAR

[BPS23] Onno Boxma, David Perry, and Wolfgang Stadje. Stationary analysis of an (R, Q) inventory model with normal and emergency orders. Journal of Applied Probability, 60(1):106–126, March 2023. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL https:/

/www.cambridge.org/core/journals/journal-of-applied-
probability/article/stationary-analysis-of-an-r-q-inventory-
model-with-normal-and-emergency-orders/3524335CC52C17BC22D51AB87225716B.

Bellini:2014:CRG

[BPSS14] Fabio Bellini, Franco Pellerey, Carlo Sgarra, and Salimeh Yasaei Sekeh. Comparison results for GARCH processes. Journal of Applied Probability, 51(3):685–698, September 2014. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://projecteuclid.org/euclid.jap/1409932667.

Boxma:2010:BPM

[BPSZ10] Onno Boxma, David Perry, Wolfgang Stadje, and Shelley Zacks. The busy period of an M/G/1 queue with customer impatience. Journal of Applied Probability, 47(1):130–145, March 2010. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/25662505.

Boland:1989:MDS

[BPT89] Philip J. Boland, Frank Proschan, and Y. L. Tong. Modelling dependence in simple and indirect majority systems. Journal of Applied Probability, 26(1):81–88, March 1989. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3214318.

Boland:1992:SOP

[BPT92] Philip J. Boland, Frank Proschan, and Y. L. Tong. A stochastic ordering of partial sums of independent random variables and of some random processes. Journal of Applied Probability, 29(3): 645–654, September 1992. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3214901.

Brown:1969:SRI

[BR69] Mark Brown and Sheldon M. Ross. Some results for infinite server Poisson queues. Journal of Applied Probability, 6(3): 604–611, December 1969. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3212106.

Balkema:1977:MID

[BR77a] A. A. Balkema and S. I. Resnick. Max-infinite divisibility. Journal of Applied Probability, 14(2):309–319, June 1977. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3213001.

Brown:1977:EVI

[BR77b] Bruce M. Brown and Sidney I. Resnick. Extreme values of independent stochastic processes. Journal of Applied Probability, 14(4):732–739, December 1977. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.

org/stable/3213346.

Barzily:1979:PFT

[BR79] Zeev Barzily and Michael Rubinovitch. On platoon formation on two-Lane roads. Journal of Applied Probability, 16(2):347– 361, June 1979. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3212902.

Ballerini:1985:RIP

[BR85] Rocco Ballerini and Sidney Resnick. Records from improving populations. Journal of Applied Probability, 22(3):487–502, September 1985. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3213855.

Ballerini:1987:ESS

[BR87a] Rocco Ballerini and Sidney I. Resnick. Embedding sequences of successive maxima in extremal processes, with applications. Journal of Applied Probability, 24(4):827–837, December 1987. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/3214208.

Beutler:1987:USM

[BR87b] Frederick J. Beutler and Keith W. Ross. Uniformization for semi-Markov decision processes under stationary policies. Journal of Applied Probability, 24(3):644–656, September 1987. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/3214096.

Baldi:1989:ANS

[BR89] Pierre Baldi and Yosef Rinott. Asymptotic normality of some graph-related statistics. Journal of Applied Probability, 26(1): 171–175, March 1989. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3214327.

Burton:1995:CTR

[BR95] Robert M. Burton and Uwe Rösler. An L_{2} convergence theorem for random affine mappings. Journal of Applied Probability, 32 (1):183–192, March 1995. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3214928.

Blaszczyszyn:1996:EMM

[BR96] Bartlomiej Blaszczyszyn and Tomasz Rolski. Expansions for Markov-modulated systems and approximations of ruin probability. Journal of Applied Probability, 33(1):57–70, March 1996. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/3215264.

Bauerle:1998:MRW

[BR98] Nicole Bäuerle and Tomasz Rolski. A monotonicity result for the workload in Markov-modulated queues. Journal of Applied Probability, 35(3):741–747, September 1998. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/3215648.

Balaji:2000:PTM

[BR00] S. Balaji and S. Ramasubramanian. Passage time moments for multidimensional diffusions. Journal of Applied Probability, 37 (1):246–251, March 2000. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3215672.

Belzunce:2002:MDO

[BR02] Felíx Belzunce and José-María Ruiz. Multivariate dispersive ordering of epoch times of nonhomogeneous Poisson processes. Journal of Applied Probability, 39(3):637–643, September 2002. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/3216085.

Brown:2016:ORC

[BR16] Mark Brown and Sheldon M. Ross. Optimality results for coupon collection. Journal of Applied Probability, 53(3):930– 937, September 2016. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://projecteuclid.

org/euclid.jap/1476370787.

Bhattacharya:2018:LST

[BR18] Ayan Bhattacharya and Parthanil Roy. A large sample test for the length of memory of stationary symmetric stable random fields via nonsingular Z^{d}-actions. Journal of Applied Probability, 55(1):179–195, March 2018. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL https://www.cambridge.org/core/journals/
journal-of-applied-probability/article/large-sample-
test-for-the-length-of-memory-of-stationary-symmetric-
stable-random-fields-via-nonsingular-dactions/B1D3F9E79A428DA6995F73A67559D7A5.

Brailovskaya:2023:TTR

[BR23] Tatiana Brailovskaya and Miklós Z. Rácz. Tree trace reconstruction using subtraces. Journal of Applied Probability, 60(2):629–641, June 2023. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL https:/

/www.cambridge.org/core/journals/journal-of-applied-
probability/article/tree-trace-reconstruction-using-
subtraces/15198285F1EBAAA28398BE0817929633.

Brainerd:1970:SPR

[Bra70] Barron Brainerd. A stochastic process related to language change. Journal of Applied Probability, 7(1):69–78, April 1970. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/3212149.

Brainerd:1972:RBT

[Bra72] Barron Brainerd. On the relation between types and tokens in literary text. Journal of Applied Probability, 9(3):507–518, September 1972. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3212322.

Braun:1975:PBP

[Bra75] Henry Braun. Polynomial bounds for probability generating functions. Journal of Applied Probability, 12(3):507–514, September 1975. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3212865.

Bradley:1980:RCL

[Bra80] Richard C. Bradley, Jr. A remark on the central limit question for dependent random variables. Journal of Applied Probability, 17(1):94–101, March 1980. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3212927.

Brainerd:1981:SEU

[Bra81] Barron Brainerd. Some elaborations upon Gani’s model for the type-token relationship. Journal of Applied Probability, 18 (2):452–460, June 1981. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3213291.

Brainerd:1982:RBT

[Bra82] Barron Brainerd. On the relation between the type-token and species-area problems. Journal of Applied Probability, 19(4): 785–793, December 1982. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3213831.

Brainerd:1983:LDB

[Bra83] Barron Brainerd. A limiting distribution for Bernoulli trials with second-order Markov dependence. Journal of Applied Probability, 20(2):419–422, June 1983. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.

org/stable/3213817.

Branford:1985:SEM

[Bra85] A. J. Branford. A self-excited migration process. Journal of Applied Probability, 22(1):58–67, March 1985. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/3213748.

Branford:1986:PFS

[Bra86] A. J. Branford. On a property of finite-state birth and death processes. Journal of Applied Probability, 23(4):859–866, December 1986. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/

3214460. See comments [vDB87].

Bratiychuk:2002:RAM

[Bra02] Mykola Bratiychuk. Remarks on the absolute maximum of a Lévy process. Journal of Applied Probability, 39(2):282– 295, June 2002. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3216096.

Bradley:2014:RCS

[Bra14] Richard C. Bradley. On a ‘replicating character string’ model. Journal of Applied Probability, 51(2):512–527, June 2014. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://projecteuclid.org/euclid.jap/1402578640.

Braun:2022:SBP

[Bra22] Georg Braun. On supercritical branching processes with emigration. Journal of Applied Probability, 59(3):734–754, September 2022. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-

6072 (electronic). URL https://www.cambridge.org/core/
journals/journal-of-applied-probability/article/on-supercritical-
branching-processes-with-emigration/2D9449FAAD2EF5497EAE6E5938D67E31.

Breslow:1969:LSS

[Bre69] Norman Breslow. On large sample sequential analysis with applications to survivorship data. Journal of Applied Probability, 6 (2):261–274, August 1969. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3211997.

Bremaud:1975:EWT

[Bre75] P. Bremaud. An extension of Watanabe’s theorem of characterization of Poisson processes over the positive real half line. Journal of Applied Probability, 12(2):396–399, June 1975. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/3212457.

Breth:1976:RS

[Bre76] M. Breth. On a recurrence of Steck. Journal of Applied Probability, 13(4):823–825, December 1976. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3212541.

Bremaud:1978:OTQ

[Bré78] P. Brémaud. On the output theorem of queueing theory, via filtering. Journal of Applied Probability, 15(2):397–405, June 1978. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3213410.

Breuer:1981:SAS

[Bre81] Peter Breuer. A strong approximation for some non-stationary complex Gaussian processes. Journal of Applied Probability, 18 (2):390–402, June 1981. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3213285.

Bremaud:1991:EPS

[Bré91] P. Brémaud. An elementary proof of Sengupta’s invariance relation and a remark on Miyazawa’s conservation principle. Journal of Applied Probability, 28(4):950–954, December 1991. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3214703. See [Sen89, Miy85].

Bremaud:1993:SAF

[Bré93] P. Brémaud. A Swiss Army formula of Palm calculus. Journal of Applied Probability, 30(1):40–51, March 1993. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3214620.

Bremaud:2000:IPL

[Bré00] Pierre Brémaud. An insensitivity property of Lundberg’s estimate for delayed claims. Journal of Applied Probability, 37(3): 914–917, September 2000. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3215623.

Breuer:2008:FPT

[Bre08] Lothar Breuer. First passage times for Markov additive processes with positive jumps of phase type. Journal of Applied Probability, 45(3):779–799, September 2008. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/27595986.

Breuer:2010:QLM

[Bre10] Lothar Breuer. A quintuple law for Markov additive processes with phase-type jumps. Journal of Applied Probability, 47(2): 441–458, June 2010. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/20778490.

Breuer:2012:EPR

[Bre12a] Lothar Breuer. Exit problems for reflected Markov-modulated Brownian motion. Journal of Applied Probability, 49(3):697–709, September 2012. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
41713799.

Breuer:2012:OTM

[Bre12b] Lothar Breuer. Occupation times for Markov-modulated Brownian motion. Journal of Applied Probability, 49(2):549–565, June 2012. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/41713785.

Breuer:2013:REL

[Bre13] Lothar Breuer. The resolvent and expected local times for Markov-modulated Brownian motion with phase-dependent termination rates. Journal of Applied Probability, 50(2):430–438, June 2013. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-

6072 (electronic). URL http://projecteuclid.org/euclid.
jap/1371648951.

Brill:1971:PPA

[Bri71] Edward A. Brill. Point processes arising in vehicular traffic flow. Journal of Applied Probability, 8(4):809–814, December 1971. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/3212244.

Brillinger:1973:EMS

[Bri73] David R. Brillinger. Estimation of the mean of a stationary time series by sampling. Journal of Applied Probability, 10(2):419– 431, June 1973. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3212358.

Brill:1979:ELC

[Bri79] P. H. Brill. An embedded level crossing technique for dams and queues. Journal of Applied Probability, 16(1):174–186, March 1979. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3213385.

Brillinger:1986:RRS

[Bri86] David R. Brillinger. Regression for randomly sampled spatial series: The trigonometric case. Journal of Applied Probability, 23A:275–289, 1986. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3214358.

Britton:1998:CVC

[Bri98] Tom Britton. On critical vaccination coverage in multitype epidemics. Journal of Applied Probability, 35(4):1003–1006, December 1998. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3215536.

Britton:2020:DPA

[Bri20] Tom Britton. Directed preferential attachment models: Limiting degree distributions and their tails. Journal of Applied Probability, 57(1):122–136, March 2020. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL https:/

/www.cambridge.org/core/journals/journal-of-applied-
probability/article/directed-preferential-attachment-
models-limiting-degree-distributions-and-their-tails/
FA12F1802D36778DC1B57E6E0C7A48B4.

Brockwell:1964:AET

[Bro64] P. J. Brockwell. An asymptotic expansion for the tail of a binomial distribution and its application in queueing theory. Journal of Applied Probability, 1(1):161–167, June 1964. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3212067.

Brook:1966:BMG

[Bro66] D. Brook. Bounds for moment generating functions and for extinction probabilities. Journal of Applied Probability, 3(1):171– 178, June 1966. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3212045.

Brown:1969:IPP

[Bro69a] Mark Brown. An invariance property of Poisson processes. Journal of Applied Probability, 6(2):453–458, August 1969. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3212015.

Brown:1969:SRT

[Bro69b] Mark Brown. Some results on a traffic model of Rényi. Journal of Applied Probability, 6(2):293–300, August 1969. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3212000.

Brockwell:1970:SPP

[Bro70] P. J. Brockwell. Stochastic population processes in the theory of radiative transfer. Journal of Applied Probability, 7(2):272– 290, August 1970. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3211964.

Brockwell:1972:SMW

[Bro72] P. J. Brockwell. A storage model in which the net growth-rate is a Markov chain. Journal of Applied Probability, 9(1):129– 139, March 1972. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3212642.

Brockwell:1974:DMW

[Bro74] P. J. Brockwell. Deviations from monotonicity of a Wiener process with drift. Journal of Applied Probability, 11(1):206– 210, March 1974. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3212600.

Bronk:1979:SIM

[Bro79] Burt V. Bronk. Some inequalities for moments and coefficients of variation for a large class of probability functions. Journal of Applied Probability, 16(3):665–670, September 1979. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3213096.

Brockwell:1986:ETG

[Bro86] P. J. Brockwell. The extinction time of a general birth and death process with catastrophes. Journal of Applied Probability, 23(4): 851–858, December 1986. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3214459.

Brockwell:1988:SMC

[Bro88] Peter J. Brockwell. Stochastic models in cell kinetics. Journal of Applied Probability, 25A:91–111, 1988. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/3214149.

Browne:2000:SDP

[Bro00a] Sid Browne. Stochastic differential portfolio games. Journal of Applied Probability, 37(1):126–147, March 2000. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3215664.

Browning:2000:MCA

[Bro00b] Sharon Browning. A Monte Carlo approach to calculating probabilities for continuous identity by descent data. Journal of Applied Probability, 37(3):850–864, September 2000. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3215618.

Brockwell:2004:RCT

[Bro04a] Peter J. Brockwell. Representations of continuous-time ARMA processes. Journal of Applied Probability, 41A:375–382, 2004. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/3215990.

Brown:2004:TRV

[Bro04b] Timothy C. Brown. Transforming a random variable to a prescribed distribution: An application to school-based assessment. Journal of Applied Probability, 41A:239–252, 2004. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3215980.

Brown:2015:SBE

[Bro15] Mark Brown. Sharp bounds for exponential approximations under a hazard rate upper bound. Journal of Applied Probability, 52(3):841–850, September 2015. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http:/

/projecteuclid.org/euclid.jap/1445543850.

Brockwell:1982:EVR

[BRPS82] P. J. Brockwell, S. I. Resnick, and N. Pacheco-Santiago. Extreme values, range and weak convergence of integrals of Markov chains. Journal of Applied Probability, 19(2):272–288, June 1982. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/3213480.

Botev:2016:SCE

[BRRN16] Z. I. Botev, A. Ridder, and L. Rojas-Nandayapa. Semiparametric cross entropy for rare-event simulation. Journal of Applied Probability, 53(3):633–649, September 2016. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

projecteuclid.org/euclid.jap/1476370767.

Brown:1975:EYP

[BRS75] Mark Brown, Sheldon Ross, and Richard Shorrock. Evacuation of a Yule process with immigration. Journal of Applied Probability, 12(4):807–811, December 1975. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3212731.

Boardman:2022:HRV

[BRSC22] Samuel Boardman, Daniel Rudolf, and Laurent Saloff-Coste. The hit-and-run version of top-to-random. Journal of Applied Probability, 59(3):860–879, September 2022. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL https://www.cambridge.org/core/journals/journal-
of-applied-probability/article/hitandrun-version-of-
toptorandom/D0DF93D7FA1876ECFE78B1A7CB9A01C8.

Brumelle:1971:RBC

[Bru71] Shelby L. Brumelle. On the relation between customer and time averages in queues. Journal of Applied Probability, 8(3):508–520, September 1971. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3212174.

Bruss:1978:BPR

[Bru78] F. Thomas Bruss. Branching processes with random absorbing process. Journal of Applied Probability, 15(1):54–64, March 1978. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3213236.

Bruss:1980:CBC

[Bru80] F. Thomas Bruss. A counterpart of the Borel–Cantelli lemma. Journal of Applied Probability, 17(4):1094–1101, December 1980. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/3213220.

Bruss:1984:NEC

[Bru84] F. Thomas Bruss. A note on extinction criteria for bisexual Galton–Watson processes. Journal of Applied Probability, 21(4): 915–919, December 1984. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3213707.

Bruss:1987:OSP

[Bru87] F. Thomas Bruss. On an optimal selection problem of Cowan and Zabczyk. Journal of Applied Probability, 24(4):918–928, December 1987. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/

3214216. See correction [Ano88f].

Bruss:2005:WKA

[Bru05] F. Thomas Bruss. What is known about Robbins’ problem? Journal of Applied Probability, 42(1):108–120, March 2005. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/30040773.

Barrieu:2004:SHW

[BRY04] P. Barrieu, A. Rouault, and M. Yor. A study of the Hartman– Watson distribution motivated by numerical problems related to the pricing of Asian options. Journal of Applied Probability, 41 (4):1049–1058, December 2004. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.

org/stable/4141377.

Birnbaum:1969:EFL

[BS69a] Z. W. Birnbaum and S. C. Saunders. Estimation for a family of life distributions with applications to fatigue. Journal of Applied Probability, 6(2):328–347, August 1969. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http:/

/www.jstor.org/stable/3212004.

Birnbaum:1969:NFL

[BS69b] Z. W. Birnbaum and S. C. Saunders. A new family of life distributions. Journal of Applied Probability, 6(2):319–327, August 1969. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3212003.

Blaisdell:1970:RSP

[BS70] B. Edwin Blaisdell and Herbert Solomon. On random sequential packing in the plane and a conjecture of Palásti. Journal of Applied Probability, 7(3):667–698, December 1970. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3211946.

Brown:1973:OIP

[BS73] Mark Brown and Herbert Solomon. Optimal issuing policies under stochastic field lives. Journal of Applied Probability, 10 (4):761–768, December 1973. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.

org/stable/3212379.

Bergmann:1976:EBW

[BS76] R. Bergmann and D. Stoyan. On exponential bounds for the waiting-time distribution function in GI/G/1. Journal of Applied Probability, 13(2):411–417, June 1976. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/3212850.

Block:1978:SMN

[BS78] Henry W. Block and Thomas H. Savits. Shock models with NBUE survival. Journal of Applied Probability, 15(3):621–628, September 1978. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3213125.

Barbour:1979:FNB

[BS79a] Andrew D. Barbour and H.-J. Schuh. Functional normalizations for the branching process with infinite mean. Journal of Applied Probability, 16(3):513–525, September 1979. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/3213081.

Boyd:1979:REI

[BS79b] David W. Boyd and J. Michael Steele. Random exchanges of information. Journal of Applied Probability, 16(3):657–661, September 1979. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3213094.

Brown:1979:RPC

[BS79c] T. C. Brown and B. W. Silverman. Rates of Poisson convergence for U-statistics. Journal of Applied Probability, 16(2):428– 432, June 1979. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3212911.

Brown:1981:LG

[BS81] Gavin Brown and J. W. Sanders. Lognormal Genesis. Journal of Applied Probability, 18(2):542–547, June 1981. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3213303.

Blaisdell:1982:RSP

[BS82a] B. Edwin Blaisdell and Herbert Solomon. Random sequential packing in Euclidean spaces of dimensions three and four and a conjecture of Palásti. Journal of Applied Probability, 19(2):382– 390, June 1982. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3213489.

Block:1982:DMM

[BS82b] Henry W. Block and Thomas H. Savits. A decomposition for multistate monotone systems. Journal of Applied Probability, 19(2):391–402, June 1982. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3213490.

Bhattacharjee:1990:FLD

[BS90] Manish C. Bhattacharjee and Jayaram Sethuraman. Families of life distributions characterized by two moments. Journal of Applied Probability, 27(3):720–725, September 1990. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3214556.

Banks:1992:CBP

[BS92a] Jeffrey S. Banks and Rangarajan K. Sundaram. A class of bandit problems yielding myopic optimal strategies. Journal of Applied Probability, 29(3):625–632, September 1992. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/3214899.

Browne:1992:WMQ

[BS92b] Sid Browne and Karl Sigman. Work-modulated queues with applications to storage processes. Journal of Applied Probability, 29(3):699–712, September 1992. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.

org/stable/3214906.

Browne:1993:TBC

[BS93] Sid Browne and J. Michael Steele. Transient behavior of coverage processes with applications to the infinite-server queue. Journal of Applied Probability, 30(3):589–601, September 1993. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3214768.

Borovkov:1994:EJN

[BS94] A. A. Borovkov and R. Schassberger. Ergodicity of a Jackson network with batch arrivals. Journal of Applied Probability, 31 (3):847–853, September 1994. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.

org/stable/3215163.

Bollabas:1997:AUB

[BS97] Béla Bollabás and Alan Stacey. Approximate upper bounds for the critical probability of oriented percolation in two dimensions based on rapidly mixing Markov chains. Journal of Applied Probability, 34(4):859–867, December 1997. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/3215002.

Boucherie:1998:CSI

[BS98] Richard J. Boucherie and Matteo Sereno. On closed support T -invariants and the traffic equations. Journal of Applied Probability, 35(2):473–481, June 1998. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3215700.

Ball:2001:FAC

[BS01a] Frank Ball and Valeri T. Stefanov. Further approaches to computing fundamental characteristics of birth-death processes. Journal of Applied Probability, 38(4):995–1005, December 2001. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/3215779.

Boxma:2001:OJW

[BS01b] Onno J. Boxma and Ryszard Syski. Obituary: Jacob Willem Cohen. Journal of Applied Probability, 38(2):604–608, June 2001. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/3215910.

Barbot:2002:SSF

[BS02] N. Barbot and B. Sericola. Stationary solution to the fluid queue fed by an M/M/1 queue. Journal of Applied Probability, 39 (2):359–369, June 2002. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3216101.

Belzunce:2004:PPN

[BS04a] Félix Belzunce and Patrizia Semeraro. Preservation of positive and negative orthant dependence concepts under mixtures and applications. Journal of Applied Probability, 41(4):961–974, December 2004. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
4141370.

Bingham:2004:SMN

[BS04b] N. H. Bingham and H. R. Nili Sani. Summability methods and negatively associated random variables. Journal of Applied Probability, 41A:231–238, 2004. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3215979.

Berend:2006:RAB

[BS06] Daniel Berend and Luba Sapir. Range of asymptotic behaviour of the optimality probability of the expert and majority rules. Journal of Applied Probability, 43(1):16–31, March 2006. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/27595707.

BorgesDeLima:2008:TLR

[BS08] Bernardo Nuñes Borges De Lima and Artëm Sapozhnikov. On the truncated long-range percolation on Z^{2}. Journal of Applied Probability, 45(1):287–291, March 2008. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http:/

/www.jstor.org/stable/27595941.

Bruss:2009:CTA

[BS09] F. Thomas Bruss and Yvik C. Swan. A continuous-time approach to Robbins’ problem of minimizing the expected rank. Journal of Applied Probability, 46(1):1–18, March 2009. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/25662404.

Ball:2013:AVE

[BS13a] Frank Ball and David Sirl. Acquaintance vaccination in an epidemic on a random graph with specified degree distribution. Journal of Applied Probability, 50(4):1147–1168, December 2013. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-

6072 (electronic). URL http://projecteuclid.org/euclid.

jap/1389370105. See correction [BS15b].

Baumann:2013:CSE

[BS13b] Hendrik Baumann and Werner Sandmann. Computing stationary expectations in level-dependent QBD processes. Journal of Applied Probability, 50(1):151–165, March 2013. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://projecteuclid.org/euclid.jap/1363784430.

Bandyopadhyay:2014:NNA

[BS14a] Antar Bandyopadhyay and Farkhondeh Sajadi. On the nearestneighbor algorithm for the mean-field traveling salesman problem. Journal of Applied Probability, 51(1):106–117, March 2014. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://projecteuclid.org/euclid.jap/
1395771417.

Biard:2014:FPP

[BS14b] Romain Biard and Bruno Saussereau. Fractional Poisson process: long-range dependence and applications in ruin theory. Journal of Applied Probability, 51(3):727–740, September 2014. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://projecteuclid.org/euclid.jap/

1409932670. See correction [BS16b].

Badia:2015:DPC

[BS15a] F. G. Badía and C. Sangüesa. The DFR property for counting processes stopped at an independent random time. Journal of Applied Probability, 52(2):574–585, June 2015. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://projecteuclid.org/euclid.jap/1437658616.

Ball:2015:CAV

[BS15b] F. Ball and D. Sirl. Correction: Acquaintance vaccination in an epidemic on a random graph with specified degree distribution. Journal of Applied Probability, 52(3):908, September 2015. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://projecteuclid.org/euclid.jap/

1445543857. See [BS13a].

Bartoszek:2015:PCI

[BS15c] Krzysztof Bartoszek and Serik Sagitov. Phylogenetic confidence intervals for the optimal trait value. Journal of Applied Probability, 52(4):1115–1132, December 2015. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

projecteuclid.org/euclid.jap/1450802756.

Baumann:2015:BTE

[BS15d] Hendrik Baumann and Werner Sandmann. Bounded truncation error for long-run averages in infinite Markov chains. Journal of Applied Probability, 52(3):609–621, September 2015. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://projecteuclid.org/euclid.jap/1445543835.

Basrak:2016:RSO

[BS16a] Bojan Basrak and Drago Spoljarić. On randomly spaced observations and continuous-time random walks. Journal of Applied Probability, 53(3):888–898, September 2016. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

projecteuclid.org/euclid.jap/1476370783.

Biard:2016:CFP

[BS16b] R. Biard and B. Saussereau. Correction: Fractional Poisson process: long-range dependence and applications in ruin theory. Journal of Applied Probability, 53(4):1271–1272, December 2016. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://projecteuclid.org/euclid.jap/

1481132852. See [BS14b].

Benevides:2017:PBC

[BS17] Fabricio Siqueira Benevides and Malgorzata Sulkowska. Percolation and best-choice problem for powers of paths. Journal of Applied Probability, 54(2):343–362, June 2017. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL https:/

/www.cambridge.org/core/journals/journal-of-applied-
probability/article/percolation-and-bestchoice-problem-
for-powers-of-paths/3193A3E0A65EA0352B9B795A9B5BD8D2.

Baccelli:2019:RPP

[BS19] François Baccelli and Antonio Sodre. Renewal processes, population dynamics, and unimodular trees. Journal of Applied Probability, 56(2):339–357, June 2019. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL https:/

/www.cambridge.org/core/journals/journal-of-applied-
probability/article/renewal-processes-population-dynamics-
and-unimodular-trees/1F76AEB836FDF7FD0DF8DD2D85C390C1.

Brinker:2022:ODD

[BS22] Leonie Violetta Brinker and Hanspeter Schmidli. Optimal discounted drawdowns in a diffusion approximation under proportional reinsurance. Journal of Applied Probability, 59(2):527–540, June 2022. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL https:/

/www.cambridge.org/core/journals/journal-of-applied-
probability/article/optimal-discounted-drawdowns-in-
a-diffusion-approximation-under-proportional-reinsurance/
46882FEA07871CC9A18726A420B9632F.

Behme:2023:MDP

[BS23] Anita Behme and Philipp Lukas Strietzel. On moments of downward passage times for spectrally negative Lévy processes.

Journal of Applied Probability, 60(2):452–464, June 2023. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL https://www.cambridge.org/core/journals/journal-
of-applied-probability/article/on-moments-of-downward-
passage-times-for-spectrally-negative-levy-processes/
060B88071BBEBC19608086699204268B.

Bruss:1999:WTP

[BSB99] F. Thomas Bruss and M. Slavtchova-Bojkova. On waiting times to populate an environment and a question of statistical inference. Journal of Applied Probability, 36(1):261–267, March 1999. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/3215419.

Boland:2002:SOP

[BSC02] Philip J. Boland, Harshinder Singh, and Bojan Cukic. Stochastic orders in partition and random testing of software. Journal of Applied Probability, 39(3):555–565, September 2002. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3216078.

Boland:2004:SPO

[BSC04] Philip J. Boland, Harshinder Singh, and Bojan Cukic. The stochastic precedence ordering with applications in sampling and testing. Journal of Applied Probability, 41(1):73–82, March 2004. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/3215815.

Badia:2018:UMS

[BSC18] F. G. Badía, C. Sangüesa, and Ji Hwan Cha. Univariate and multivariate stochastic comparisons and ageing properties of the generalized Pólya process. Journal of Applied Probability, 55(1):233–253, March 2018. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL https:/

/www.cambridge.org/core/journals/journal-of-applied-
probability/article/univariate-and-multivariate-stochastic-
comparisons-and-ageing-properties-of-the-generalized-
polya-process/4EF231E5728862960236EC78ADE613F7.

Blanchet-Scalliet:2020:JLO

[BSDG20] Christophette Blanchet-Scalliet, Diana Dorobantu, and Laura Gay. Joint law of an Ornstein–Uhlenbeck process and its supremum. Journal of Applied Probability, 57(2):541–558, June 2020.

CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL https://www.cambridge.org/core/journals/
journal-of-applied-probability/article/joint-law-of-
an-ornsteinuhlenbeck-process-and-its-supremum/85CBDAA1D3ABF14DF90747FF7A41ABCA.

Belkin:1972:CPN

[BSS72] Barry Belkin, Lawrence D. Stone, and Martin Avery Snyder. Crossing problems for non-constant thresholds and certain non-Markov processes. Journal of Applied Probability, 9(4):769–788, December 1972. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/

3212614. See correction [Ano73e].

Brown:1981:MCS

[BSS81] Mark Brown, Herbert Solomon, and Michael A. Stephens. Monte Carlo simulation of the renewal function. Journal of Applied Probability, 18(2):426–434, June 1981. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3213288.

Block:2003:MDI

[BSW03] Henry W. Block, Thomas H. Savits, and Eshetu T. Wondmagegnehu. Mixtures of distributions with increasing linear failure rates. Journal of Applied Probability, 40(2):485–504, June 2003. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/3215805.

Ballani:2007:TDA

[BSW07] F. Ballani, D. Stoyan, and S. Wolf. On two damage accumulation models and their size effects. Journal of Applied Probability, 44 (1):99–114, March 2007. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/27595825.

Beknazaryan:2019:CTM

[BSX19] Aleksandr Beknazaryan, Hailin Sang, and Yimin Xiao. Cramér type moderate deviations for random fields. Journal of Applied Probability, 56(1):223–245, March 2019. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL https:/

/www.cambridge.org/core/journals/journal-of-applied-
probability/article/cramer-type-moderate-deviations-
for-random-fields/8223F019C07E27ACDF6855DFC923F563.

Bagchi:1973:NMG

[BT73] Tapan P. Bagchi and J. G. C. Templeton. A note on the M^{X} /G^{Y} /1, K bulk queueing system. Journal of Applied Probability, 10(4):901–906, December 1973. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.

org/stable/3212395.

Book:1976:ERS

[BT76] Stephen A. Book and Donald R. Truax. An Erdős–Rényi strong law for sample quantiles. Journal of Applied Probability, 13(3): 578–583, September 1976. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3212478.

Blom:1982:HMR

[BT82] Gunnar Blom and Daniel Thorburn. How many random digits are required until given sequences are obtained? Journal of Applied Probability, 19(3):518–531, September 1982. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/3213511. See correction

[Ano85f].

Beirlant:1992:LDC

[BT92] Jan Beirlant and Jozef L. Teugels. Limit distributions for compounded sums of extreme order statistics. Journal of Applied Probability, 29(3):557–574, September 1992. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/3214893.

Britton:2012:MSG

[BT12a] Tom Britton and Pieter Trapman. Maximizing the size of the giant. Journal of Applied Probability, 49(4):1156–1165, December 2012. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/41713835.

Buchholz:2012:RPR

[BT12b] Peter Buchholz and Miklós Telek. Rational processes related to communicating Markov processes. Journal of Applied Probability, 49(1):40–59, March 2012. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/41713751.

Berger:2018:SUR

[BT18] Arno Berger and Isaac Twelves. On the significands of uniform random variables. Journal of Applied Probability, 55(2):353–367, June 2018. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL https:/

/www.cambridge.org/core/journals/journal-of-applied-
probability/article/on-the-significands-of-uniform-random-
variables/0D08D0EEC38F52E431CF41290651603A.

Betken:2023:ACC

[BT23] Carina Betken and Christoph Thäle. Approaching the coupon collector’s problem with group drawings via Stein’s method. Journal of Applied Probability, 60(4):1352–1366, December 2023. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL https://www.cambridge.org/core/journals/
journal-of-applied-probability/article/approaching-the-
coupon-collectors-problem-with-group-drawings-via-steins-
method/8DE6ADB0D1E527392B1A05539DE6CF3A.

Ben-Tal:1972:MBE

[BTH72] A. Ben-Tal and E. Hochman. More bounds on the expectation of a convex function of a random variable. Journal of Applied Probability, 9(4):803–812, December 1972. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/3212616.

Buchholz:1994:EOL

[Buc94] Peter Buchholz. Exact and ordinary lumpability in finite Markov chains. Journal of Applied Probability, 31(1):59–75, March 1994. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/3215235.

Buiculescu:1975:QSD

[Bui75] Mioara Buiculescu. On quasi-stationary distributions for multitype Galton–Watson processes. Journal of Applied Probability, 12(1):60–68, March 1975. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3212407.

Buonaguidi:2015:ROV

[Buo15] Bruno Buonaguidi. A remark on optimal variance stopping problems. Journal of Applied Probability, 52(4):1187–1194, December 2015. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-

6072 (electronic). URL http://projecteuclid.org/euclid.
jap/1450802762.

Burman:1976:ETP

[Bur76] David Burman. Estimation of transition probabilities in a multiparticle semi-Markov system. Journal of Applied Probability, 13(4):696–706, December 1976. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.

org/stable/3212524.

Burke:1979:QTA

[Bur79] P. J. Burke. A queuing-theoretic anomaly. Journal of Applied Probability, 16(2):373–383, June 1979. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3212904.

Burkschat:2009:SFD

[Bur09] M. Burkschat. Systems with failure-dependent lifetimes of components. Journal of Applied Probability, 46(4):1052–1072, December 2009. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
25662480.

Butler:2019:AES

[But19] Ronald W. Butler. Asymptotic expansions and saddlepoint approximations using the analytic continuation of moment generating functions. Journal of Applied Probability, 56(1):307–338, March 2019. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL https:/

/www.cambridge.org/core/journals/journal-of-applied-
probability/article/asymptotic-expansions-and-saddlepoint-
approximations-using-the-analytic-continuation-of-moment-
generating-functions/63F9AB40061ABF664BAA64AB97F16361.

Borovkov:1996:DTE

[BV96] K. A. Borovkov and V. A. Vatutin. On distribution tails and expectations of maxima in critical branching processes. Journal of Applied Probability, 33(3):614–622, September 1996. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3215343.

Baurdoux:2011:FCM

[BV11] E. J. Baurdoux and K. Van Schaik. Further calculations for the McKean stochastic game for a spectrally negative Lévy process: from a point to an interval. Journal of Applied Probability, 48 (1):200–216, March 2011. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/29777455.

Balakrishnan:2014:SOS

[BV14] N. Balakrishnan and William Volterman. On the signatures of ordered system lifetimes. Journal of Applied Probability, 51 (1):82–91, March 2014. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://projecteuclid.

org/euclid.jap/1395771415.

Baroni:2017:NWR

[BvdHK17] Enrico Baroni, Remco van der Hofstad, and Júlia Komjáthy. Nonuniversality of weighted random graphs with infinite variance degree. Journal of Applied Probability, 54(1):146–164, March 2017. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-

6072 (electronic). URL https://www.cambridge.org/core/
journals/journal-of-applied-probability/article/nonuniversality-
of-weighted-random-graphs-with-infinite-variance-degree/
51D4CBD04E812E1C1918B659AEBFADF5.

Bet:2017:FPQ

[BvdHvL17] Gianmarco Bet, Remco van der Hofstad, and Johan S. H. van Leeuwaarden. Finite-pool queueing with heavy-tailed services. Journal of Applied Probability, 54(3):921–942, September 2017. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL https://www.cambridge.org/core/journals/
journal-of-applied-probability/article/finitepool-queueing-
with-heavytailed-services/410D89569EB3E55D67E457500C29F4F5.

Borovkov:2000:EFS

[BVJ00] K. Borovkov and D. Vere-Jones. Explicit formulae for stationary distributions of stress release processes. Journal of Applied Probability, 37(2):315–321, June 2000. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3215709.

Baccelli:2021:SSC

[BVW21] François Baccelli, Sriram Vishwanath, and Jae Oh Woo. On the steady state of continuous-time stochastic opinion dynamics with power-law confidence. Journal of Applied Probability, 58(3):746–772, September 2021. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL https:/

/www.cambridge.org/core/journals/journal-of-applied-
probability/article/on-the-steady-state-of-continuoustime-
stochastic-opinion-dynamics-with-powerlaw-confidence/
2D0D3C699DAE187E86DDD5050391D8B7.

Bambos:1989:QPI

[BW89] Nicholas Bambos and Jean Walrand. On queues with periodic inputs. Journal of Applied Probability, 26(2):381–389, June 1989. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/3214043.

Bernard:1998:SLT

[BW98] Pierre Bernard and Liming Wu. Stochastic linearization: The theory. Journal of Applied Probability, 35(3):718–730, September 1998. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3215646.

Bocharov:2019:BBM

[BW19] Sergey Bocharov and Li Wang. Branching Brownian motion with spatially homogeneous and point-catalytic branching. Journal of Applied Probability, 56(3):891–917, September 2019. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL https://www.cambridge.org/core/journals/journal-
of-applied-probability/article/branching-brownian-motion-
with-spatially-homogeneous-and-pointcatalytic-branching/
4A5F3B1CAC94A6247CBCF70236C33BBF.

Bao:2023:EEC

[BW23] Jianhai Bao and Jian Wang. Exponential ergodicity for a class of Markov processes with interactions. Journal of Applied Probability, 60(2):465–478, June 2023. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL https:/

/www.cambridge.org/core/journals/journal-of-applied-
probability/article/exponential-ergodicity-for-a-class-
of-markov-processes-with-interactions/9EFFA5427933C5D535BE1085F01A224B.

Bo:2011:FPT

[BWY11] Lijun Bo, Yongjin Wang, and Xuewei Yang. First passage times of (reflected) Ornstein–Uhlenbeck processes over random jump boundaries. Journal of Applied Probability, 48(3):723–732, September 2011. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
23065932.

Breen:1985:RTS

[BWZ85] Stephen Breen, Michael S. Waterman, and Ning Zhang. Renewal theory for several patterns. Journal of Applied Probability, 22 (1):228–234, March 1985. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3213763.

Ball:1993:LMC

[BY93] Frank Ball and Geoffrey F. Yeo. Lumpability and marginalisability for continuous-time Markov chains. Journal of Applied Probability, 30(3):518–528, September 1993. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/3214762.

Boxma:1997:MGQ

[BY97] Onno J. Boxma and Uri Yechiali. An M/G/1 queue with multiple types of feedback and gated vacations. Journal of Applied Probability, 34(3):773–784, September 1997. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/3215102.

Bao:2014:NAS

[BY14] Jianhai Bao and Chenggui Yuan. Numerical approximation of stationary distributions for stochastic partial differential equations. Journal of Applied Probability, 51(3):858–873, September 2014. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://projecteuclid.org/euclid.jap/
1409932678.

Bertoin:2022:SLB

[BY22] Jean Bertoin and Hairuo Yang. Scaling limits of branching random walks and branching-stable processes. Journal of Applied Probability, 59(4):1009–1025, December 2022. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL https://www.cambridge.org/core/journals/journal-
of-applied-probability/article/scaling-limits-of-branching-
random-walks-and-branchingstable-processes/750483FF4B00031B5718B5E505B1249E.

Byth:1981:SPP

[Byt81] Karen Byth. θ-Stationary point processes and their second-order analysis. Journal of Applied Probability, 18(4):864–878, December 1981. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3213061.

Basawa:2004:NGB

[BZ04] I. V. Basawa and J. Zhou. Non-Gaussian bifurcating models and quasi-likelihood estimation. Journal of Applied Probability, 41A:55–64, 2004. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3215968.

Bertacchi:2008:CBC

[BZ08] Daniela Bertacchi and Fabio Zucca. Critical behaviors and critical values of branching random walks on multigraphs. Journal of Applied Probability, 45(2):481–497, June 2008. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/27595959.

Bertacchi:2009:ACP

[BZ09] Daniela Bertacchi and Fabio Zucca. Approximating critical parameters of branching random walks. Journal of Applied Probability, 46(2):463–478, June 2009. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/25662437.

C:2012:UQD

[C.12] Jorge Littin C. Uniqueness of quasistationary distributions and discrete spectra when ∞ is an entrance boundary and 0 is singular. Journal of Applied Probability, 49(3):719–730, September 2012. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/41713801.

Chen:1998:IVM

[CA98] Zhao-Guo Chen and Oliver D. Anderson. Increment-vector methodology: Transforming non-stationary series to stationary series. Journal of Applied Probability, 35(1):64–77, March 1998.

CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/3215547.

Cabana:1984:TDM

[Cab84] E. M. Cabaña. On the transition density of a multidimensional parameter Wiener process with one barrier. Journal of Applied Probability, 21(1):197–200, March 1984. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http:/

/www.jstor.org/stable/3213679.

Cabana:1991:GPP

[Cab91] Enrique M. Cabaña. A Gaussian process with parabolic covariances. Journal of Applied Probability, 28(4):898–902, December 1991. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3214693.

Cai:2002:RPD

[Cai02] Jun Cai. Ruin probabilities with dependent rates of interest. Journal of Applied Probability, 39(2):312–323, June 2002. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/3216098.

Caithamer:2003:LSD

[Cai03] Peter Caithamer. Large and small deviations of a string driven by a two-parameter Gaussian noise white in time. Journal of Applied Probability, 40(4):946–960, December 2003. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3216052.

Cai:2011:PHQ

[Cai11] Ning Cai. Pricing and hedging of quantile options in a flexible jump diffusion model. Journal of Applied Probability, 48(3): 637–656, September 2011. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/23065926.

Callaert:1974:GED

[Cal74] Herman Callaert. Geometric and exponential decay in derived Markov chains. Journal of Applied Probability, 11(2):388– 393, June 1974. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3212762.

Cameron:1973:NFM

[Cam73] Murray A. Cameron. A note on functions of Markov processes with an application to a sequence of χ^{2} statistics. Journal of Applied Probability, 10(4):895–900, December 1973. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3212394.

Cameron:1986:CTR

[Cam86] Murray A. Cameron. The comparison of transient responses to stimuli. Journal of Applied Probability, 23A:159–172, 1986. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/3214350.

Campi:1997:URT

[Cam97] Marco Campi. A unique representation theorem for the conditional expectation of stationary processes and application to dynamic estimation problems. Journal of Applied Probability, 34(2):372–380, June 1997. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3215377.

Cannings:1973:ESO

[Can73] C. Cannings. The equivalence of some overlapping and nonoverlapping generation models for the study of genetic drift. Journal of Applied Probability, 10(2):432–436, June 1973. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/3212359.

Cannings:1974:GDP

[Can74] C. Cannings. Genetic drift with polygamy and arbitrary offspring distribution. Journal of Applied Probability, 11(4):633– 641, December 1974. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3212547.

Cane:1977:CNI

[Can77] Violet R. Cane. A class of non-identifiable stochastic models. Journal of Applied Probability, 14(3):475–482, September 1977. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/3213450.

Cao:1984:DST

[Cao84] Xi-Ren Cao. The dependence of sojourn times on service times in tandem queues. Journal of Applied Probability, 21(3):661–667, September 1984. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3213628.

Cao:1989:RPT

[Cao89] Xi-Ren Cao. Realization probability and throughput sensitivity in a closed Jackson network. Journal of Applied Probability, 26 (3):615–624, September 1989. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.

org/stable/3214418.

Cao:1995:SAD

[Cao95] Xi-Ren Cao. State aggregation and discrete-state Markov chains embedded in a class of point processes. Journal of Applied Probability, 32(1):39–51, March 1995. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3214919.

Cartwright:1967:TSA

[Car67] D. E. Cartwright. Time-series analysis of tides and similar motions of the sea surface. Journal of Applied Probability, 4(1):103– 112, April 1967. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3212302.

Carette:1999:MHS

[Car99] Philippe Carette. Modelling hierarchical systems by a continuous-time homogeneous Markov chain using two-wave panel data. Journal of Applied Probability, 36(3):644–653, September 1999. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3215430.

Carlton:2002:FDD

[Car02] Matthew A. Carlton. A family of densities derived from the three-parameter Dirichlet process. Journal of Applied Probability, 39(4):764–774, December 2002. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3216001.

Choe:1983:SPB

[CB83] K. I. Choe and D. S. Bai. A secretary problem with backward solicitation and uncertain employment. Journal of Applied Probability, 20(4):891–896, December 1983. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3213601.

Coffey:1991:GWP

[CB91] John Coffey and Wolfgang J. Bühler. The Galton–Watson predator-prey process. Journal of Applied Probability, 28(1): 9–16, March 1991. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3214736.

Choo:1979:NRT

[CC79] Qui Hoon Choo and Brian Conolly. New results in the theory of repeated orders queueing systems. Journal of Applied Probability, 16(3):631–640, September 1979. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3213090.

Cowan:1994:SMF

[CC94] Richard Cowan and S. N. Chiu. A stochastic model of fragment formation when DNA replicates. Journal of Applied Probability, 31(2):301–308, June 1994. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3215025.

Commault:1996:IRP

[CC96] C. Commault and J. P. Chemla. An invariant of representations of phase-type distributions and some applications. Journal of Applied Probability, 33(2):368–381, June 1996. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3215060.

Chen:2001:CEP

[CC01] Hsiao-Chi Chen and Yunshyoung Chow. On the convergence of evolution processes with time-varying mutations and local interaction. Journal of Applied Probability, 38(2):301–323, June 2001. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3215890.

Caballero:2006:CSL

[CC06] M. E. Caballero and L. Chaumont. Conditioned stable Lévy processes and the Lamperti representation. Journal of Applied Probability, 43(4):967–983, December 2006. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/27595793.

Chen:2008:ESE

[CC08] Hsiao-Chi Chen and Yunshyong Chow. Equilibrium selection in evolutionary games with imperfect monitoring. Journal of Applied Probability, 45(2):388–402, June 2008. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/27595952.

Cha:2015:SFM

[CC15] Ji Hwan Cha and Inma T. Castro. A stochastic failure model with dependent competing risks and its applications to condition-based maintenance. Journal of Applied Probability, 52 (2):558–573, June 2015. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://projecteuclid.

org/euclid.jap/1437658615.

Chao:2003:AJN

[CCC03] Chern-Ching Chao, Yi-Liang Chen, and Wei-Hou Cheng. The asymptotic joint normality of the numbers of upper records, lower records and inversions in a random sequence. Journal of Applied Probability, 40(1):257–263, March 2003. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3215853.

Cavazos-Cadena:2024:COA

[CCCSdO24] Rolando Cavazos-Cadena, Hugo Cruz-Suárez, and Raúl Montes de Oca. Characterization of the optimal average cost in Markov decision chains driven by a risk-seeking controller. Journal of Applied Probability, 61(1):340–367, March 2024. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL https://www.cambridge.org/core/journals/journal-
of-applied-probability/article/characterization-of-the-
optimal-average-cost-in-markov-decision-chains-driven-
by-a-riskseeking-controller/510C72A5527CCE8392545FE99A36512A.

Cavazos-Cadena:2005:NVI

[CCdO05] Rolando Cavazos-Cadena and Raúl Montes de Oca. Nonstationary value iteration in controlled Markov chains with risksensitive average criterion. Journal of Applied Probability, 42(4): 905–918, December 2005. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/30040817.

Cavazos-Cadena:2015:SPO

[CCdOS15] Rolando Cavazos-Cadena, Raúl Montes de Oca, and Karel Sladký. Sample-path optimal stationary policies in stable Markov decision chains with the average reward criterion. Journal of Applied Probability, 52(2):419–440, June 2015. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://projecteuclid.org/euclid.jap/1437658607.

Cavazos-Cadena:1996:VIC

[CCFG96] Rolando Cavazos-Cadena and Emmanuel Fernández-Gaucherand. Value iteration in a class of average controlled Markov chains with unbounded costs: Necessary and sufficient conditions for pointwise convergence. Journal of Applied Probability, 33(4): 986–1002, December 1996. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3214980.

Chau:2017:OII

[CCFM17] Huy N. Chau, Andrea Cosso, Claudio Fontana, and Oleksii Mostovyi. Optimal investment with intermediate consumption under no unbounded profit with bounded risk. Journal of Applied Probability, 54(3):710–719, September 2017. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL https://www.cambridge.org/core/journals/journal-
of-applied-probability/article/optimal-investment-with-
intermediate-consumption-under-no-unbounded-profit-with-
bounded-risk/5C2162DD333945A3333C9A8279B91A80.

Csaki:2000:API

[CCFR00] Endre Csáki, Miklós Csörgő, Antónia Földes, and Pál Révész. Asymptotic properties of integral functionals of geometric stochastic processes. Journal of Applied Probability, 37(2):480– 493, June 2000. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3215722.

Cowan:1990:ARP

[CCG90] Richard Cowan, David Culpin, and David Gates. Asymptotic results for a problem of DNA breakage. Journal of Applied Probability, 27(2):433–439, June 1990. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3214663.

Coffman:1991:DCP

[CCGP91] E. G. Coffman, Jr., P.-J. Courtois, E. N. Gilbert, and Ph. Piret. A distributed clustering process. Journal of Applied Probability, 28(4):737–750, December 1991. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.

org/stable/3214677.

Chen:2023:SLT

[CCGW23] Anyue Chen, Yong Chen, Wu-Jun Gao, and Xiaohan Wu. Sturm–Liouville theory and decay parameter for quadratic Markov branching processes. Journal of Applied Probability, 60(3):737–764, September 2023. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL https:/

/www.cambridge.org/core/journals/journal-of-applied-
probability/article/sturmliouville-theory-and-decay-
parameter-for-quadratic-markov-branching-processes/D74BD2F9A11FD056955EF7854C1646FB.

Cowan:1995:LTR

[CCH95] Richard Cowan, S. N. Chiu, and Lars Holst. A limit theorem for the replication time of a DNA molecule. Journal of Applied Probability, 32(2):296–303, June 1995. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3215289.

Chen:2006:BRQ

[CCH06] Hsiao-Chi Chen, Yunshyong Chow, and June Hsieh. Boundedly rational quasi-Bayesian learning in coordination games with imperfect monitoring. Journal of Applied Probability, 43(2):335– 350, June 2006. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
27595735.

Chen:2008:NBP

[CCHY08] May-Ru Chen, Pei-Shou Chung, Shoou-Ren Hsiau, and Yi-Ching Yao. On nonoptimality of bold play for subfair redand-black with a rational-valued house limit. Journal of Applied Probability, 45(4):1024–1038, December 2008. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/27596006.

Chen:2010:SLL

[CCN10] Yiqing Chen, Anyue Chen, and Kai W. Ng. The Strong Law of Large Numbers for extended negatively dependent random variables. Journal of Applied Probability, 47(4):908–922, December 2010. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/29777415.

Chang:1992:OLR

[CCPW92] Cheng-Shang Chang, Xiuli Chao, Michael Pinedo, and Richard Weber. On the optimality of LEPT and cμ rules for machines in parallel. Journal of Applied Probability, 29(3):667–681, September 1992. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3214903.

Cressie:1981:SDA

[CD81] Noel Cressie and Robert W. Davis. The supremum distribution of another Gaussian process. Journal of Applied Probability, 18 (1):131–138, March 1981. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3213173.

Cressman:1985:ESS

[CD85] R. Cressman and A. T. Dash. Evolutionarily stable strategies with two types of player. I. Two-species haploid or randomly mating diploid. Journal of Applied Probability, 22(1): 1–14, March 1985. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3213743.

Chukova:1992:DHA

[CD92] S. Chukova and B. Dimitrov. On distributions having the almost-lack-of-memory property. Journal of Applied Probability, 29(3):691–698, September 1992. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3214905.

Chao:1995:MRS

[CD95] Xiuli Chao and Liyi Dai. A monotonicity result for a singleserver loss system. Journal of Applied Probability, 32(4):1112– 1117, December 1995. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3215224.

Cruz:1998:SCC

[CD98] J. R. Cruz and C. C. Y. Dorea. Simple conditions for the convergence of simulated annealing type algorithms. Journal of Applied Probability, 35(4):885–892, December 1998. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/3215524.

Costa:2005:SCE

[CD05] O. L. V. Costa and F. Dufour. A sufficient condition for the existence of an invariant probability measure for Markov processes. Journal of Applied Probability, 42(3):873–878, September 2005. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/30040864.

Costa:2006:EPE

[CD06] O. L. V. Costa and F. Dufour. Ergodic properties and ergodic decompositions of continuous-time Markov processes. Journal of Applied Probability, 43(3):767–781, September 2006. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/27595771.

Carpio:2007:LRD

[CD07] K. J. E. Carpio and D. J. Daley. Long-range dependence of Markov chains in discrete time on countable state space. Journal of Applied Probability, 44(4):1047–1055, December 2007. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/27595906.

Costa:2009:VDA

[CD09] O. L. V. Costa and F. Dufour. The vanishing discount approach for the average continuous control of piecewise deterministic Markov processes. Journal of Applied Probability, 46 (4):1157–1183, December 2009. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.

org/stable/25662487.

Che:2013:SBC

[CD13] Xiaonan Che and Angelos Dassios. Stochastic boundary crossing probabilities for the Brownian motion. Journal of Applied Probability, 50(2):419–429, June 2013. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

projecteuclid.org/euclid.jap/1371648950.

Crane:2019:RE

[CD19] Harry Crane and Walter Dempsey. Relational exchangeability. Journal of Applied Probability, 56(1):192–208, March 2019. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL https://www.cambridge.org/core/journals/journal-
of-applied-probability/article/relational-exchangeability/
35142F1C41D764CA572061A8C3C4CC0C.

Chaudhuri:1991:SBR

[CDD91] Gopal Chaudhuri, Jayant V. Deshpande, and Avinash D. Dharmadhikari. Some bounds on reliability of coherent systems of IFRA components. Journal of Applied Probability, 28(3):709– 714, September 1991. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3214505.

Coutin:2010:MMS

[CDD10] L. Coutin, L. Decreusefond, and J. S. Dhersin. A Markov model for the spread of viruses in an open population. Journal of Applied Probability, 47(4):976–996, December 2010. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/29777420.

Chernick:1988:TRR

[CDL88] M. R. Chernick, D. J. Daley, and R. P. Littlejohn. A timereversibility relationship between two Markov chains with exponential stationary distributions. Journal of Applied Probability, 25(2):418–422, June 1988. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3214450.

Coletti:2023:LSF

[CdLH^{+}23] Cristian F. Coletti, Lucas R. de Lima, Alexander Hinsen, Benedikt Jahnel, and Daniel Valesin. Limiting shape for firstpassage percolation models on random geometric graphs. Journal of Applied Probability, 60(4):1367–1385, December 2023. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL https://www.cambridge.org/core/journals/
journal-of-applied-probability/article/limiting-shape-

for-firstpassage-percolation-models-on-random-geometric-
graphs/3FA004F83BC3A3CE1DCF43159081F223.

Coletti:2016:STS

[CdOR16] Cristian F. Coletti, Karina B. E. de Oliveira, and Pablo M. Rodriguez. A stochastic two-stage innovation diffusion model on a lattice. Journal of Applied Probability, 53(4):1019–1030, December 2016. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-

6072 (electronic). URL http://projecteuclid.org/euclid.
jap/1481132833.

Can:2019:PPR

[CDP19] Van Hao Can, Manh Hong Duong, and Viet Viet Hung Pham. Persistence probability of a random polynomial arising from evolutionary game theory. Journal of Applied Probability, 56(3):870–890, September 2019. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL https:/

/www.cambridge.org/core/journals/journal-of-applied-
probability/article/persistence-probability-of-a-random-
polynomial-arising-from-evolutionary-game-theory/83E0D1B3EC7EADD56B4933AF7CF7FE26.

Cloez:2020:OSM

[CdSJ20] Bertrand Cloez, Benoîte de Saporta, and Maud Joubaud. Optimal stopping for measure-valued piecewise deterministic Markov processes. Journal of Applied Probability, 57(2):497–512, June 2020. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL https:/

/www.cambridge.org/core/journals/journal-of-applied-
probability/article/optimal-stopping-for-measurevalued-
piecewise-deterministic-markov-processes/37EBBE16049A145A81A7DFBA8853B26F.

Cerrito:1988:ISM

[Cer88] P. B. Cerrito. An improved stochastic model allowing interaction among individuals. Journal of Applied Probability, 25(3): 492–500, September 1988. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3213978.

Cha:2009:TSP

[CF09] Ji Hwan Cha and Maxim Finkelstein. On a terminating shock process with independent wear increments. Journal of Applied Probability, 46(2):353–362, June 2009. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/25662429.

Cha:2011:NCE

[CF11a] Ji Hwan Cha and Maxim Finkelstein. On new classes of extreme shock models and some generalizations. Journal of Applied Probability, 48(1):258–270, March 2011. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/29777458.

Cha:2011:SIM

[CF11b] Ji Hwan Cha and Maxim Finkelstein. Stochastic intensity for minimal repairs in heterogeneous populations. Journal of Applied Probability, 48(3):868–876, September 2011. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/23065943.

Chronopoulou:2013:OSC

[CF13] Alexandra Chronopoulou and Georgios Fellouris. Optimal sequential change detection for fractional diffusion-type processes. Journal of Applied Probability, 50(1):29–41, March 2013. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://projecteuclid.org/euclid.jap/1363784422.

Cha:2014:SME

[CF14] Ji Hwan Cha and Maxim Finkelstein. Stochastic modeling for environmental stress screening. Journal of Applied Probability, 51(2):387–399, June 2014. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://projecteuclid.

org/euclid.jap/1402578632.

Cohen:2017:NEN

[CF17] Samuel N. Cohen and Victor Fedyashov. Nash equilibria for nonzero-sum ergodic stochastic differential games. Journal of Applied Probability, 54(4):977–994, December 2017. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL https://www.cambridge.org/core/journals/journal-
of-applied-probability/article/nash-equilibria-for-nonzerosum-
ergodic-stochastic-differential-games/D5445AF7AB2FFB9417480BDA83B49515.

Christensen:2022:P

[CF22a] Sören Christensen and Simon Fischer. On the S_{n}/n problem. Journal of Applied Probability, 59(2):571–583, June 2022. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL https://www.cambridge.org/core/journals/journal-
of-applied-probability/article/on-the-snn-problem/8CE3A5834AC35E7ADD137749F61E27A2.

Cipriani:2022:DFM

[CF22b] Alessandra Cipriani and Andrea Fontanari. Dynamical fitness models: evidence of universality classes for preferential attachment graphs. Journal of Applied Probability, 59 (3):609–630, September 2022. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL https:/

/www.cambridge.org/core/journals/journal-of-applied-
probability/article/dynamical-fitness-models-evidence-
of-universality-classes-for-preferential-attachment-
graphs/061292E02BDE3478CF09391D059FD6A9.

Coletti:2009:SLD

[CFD09] C. F. Coletti, L. R. G. Fontes, and E. S. Dias. Scaling limit for a drainage network model. Journal of Applied Probability, 46(4): 1184–1197, December 2009. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/25662488.

Cattiaux:2022:RPD

[CFRS22] Patrick Cattiaux, Jens Fischer, Sylvie Roelly, and Samuel Sindayigaya. Random population dynamics under catastrophic events. Journal of Applied Probability, 59(4):962–982, December 2022. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-

6072 (electronic). URL https://www.cambridge.org/core/
journals/journal-of-applied-probability/article/random-
population-dynamics-under-catastrophic-events/5121309AFD018FF445041D432ED387AB.

Chen:2011:MES

[CFS11] Ao Chen, Liming Feng, and Renming Song. On the monitoring error of the supremum of a normal jump diffusion process. Journal of Applied Probability, 48(4):1021–1034, December 2011. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/23066440.

Cambanis:1996:FRT

[CFZ96] Stamatis Cambanis and Issa Fakhre-Zakeri. Forward and reversed time prediction of autoregressive sequences. Journal of Applied Probability, 33(4):1053–1060, December 1996. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3214985.

Coleman:1969:SMI

[CG69] R. Coleman and J. L. Gastwirth. Some models for interaction of renewal processes related to neuron firing. Journal of Applied Probability, 6(1):38–58, April 1969. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3212275.

Crump:1976:DNA

[CG76] Kenny S. Crump and John H. Gillespie. The dispersion of a neutral allele considered as a branching process. Journal of Applied Probability, 13(2):208–218, June 1976. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/3212824.

Castillo:1987:LRM

[CG87] Enrique Castillo and Janos Galambos. Lifetime regression models based on a functional equation of physical nature. Journal of Applied Probability, 24(1):160–169, March 1987. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3214067.

Chauvin:1990:GRC

[CG90] Brigitte Chauvin and Gaston Giroux. A geometric rate of convergence to the equilibrium for Boltzmann processes with multiple particle interactions. Journal of Applied Probability, 27(3): 510–520, September 1990. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3214537.

Chaudhry:1996:PAD

[CG96] M. L. Chaudhry and U. C. Gupta. Performance analysis of the discrete-time GI/Geom/1/N queue. Journal of Applied Probability, 33(1):239–255, March 1996. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3215281.

Calvin:1997:ACB

[CG97a] James M. Calvin and Peter W. Glynn. Average case behavior of random search for the maximum. Journal of Applied Probability, 34(3):632–642, September 1997. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.

org/stable/3215090.

Ceci:1997:FBP

[CG97b] Claudia Ceci and Anna Gerardi. Filtering of a branching process given its split times. Journal of Applied Probability, 34(3): 565–574, September 1997. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3215084.

Cai:1999:UAS

[CG99] Jun Cai and José Garrido. A unified approach to the study of tail probabilities of compound distributions. Journal of Applied Probability, 36(4):1058–1073, December 1999. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3215578.

Chang:2000:PLN

[CG00a] Wei-Bin Chang and John A. Gubner. Poisson limits and nonparametric estimation for pairwise interaction point processes. Journal of Applied Probability, 37(1):252–260, March 2000. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/3215673.

Crosbie:2000:EPG

[CG00b] J. H. Crosbie and K. D. Glazebrook. Evaluating policies for generalized bandits via a notion of duality. Journal of Applied Probability, 37(2):540–546, June 2000. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3215728.

Ceci:2002:CLB

[CG02] Claudia Ceci and Anna Gerardi. Conditional law of a branching process observing a subpopulation. Journal of Applied Probability, 39(1):112–122, March 2002. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3215922.

Ceci:2005:MBP

[CG05] Claudia Ceci and Anna Gerardi. Multitype branching processes observing particles of a given type. Journal of Applied Probability, 42(2):446–462, June 2005. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/30040803.

Cha:2012:NFR

[CG12] Ji Hwan Cha and Massimiliano Giorgio. A note on the failure rates in finite mixed populations. Journal of Applied Probability, 49(2):405–415, June 2012. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/41713775.

Carlen:2008:PIE

[CGR08] Eric Carlen, Ester Gabetta, and Eugenio Regazzini. Probabilistic investigations on the explosion of solutions of the Kac equation with infinite energy initial distribution. Journal of Applied Probability, 45(1):95–106, March 2008. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/27595926.

Conolly:1969:CQ

[CH69] B. W. Conolly and N. Hadidi. A correlated queue. Journal of Applied Probability, 6(1):122–136, April 1969. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/3212280.

Crump:1970:SAR

[CH70] Kenny S. Crump and David G. Hoel. Some applications of renewal theory on the whole line. Journal of Applied Probability, 7(3):734–746, December 1970. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.

org/stable/3211950.

Cushen:1971:QMC

[CH71] C. D. Cushen and R. L. Hudson. A quantum-mechanical central limit theorem. Journal of Applied Probability, 8(3):454–469, September 1971. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3212170.

Cressman:1984:ESS

[CH84a] R. Cressman and W. G. S. Hines. Evolutionarily stable strategies of diploid populations with semi-dominant inheritance patterns. Journal of Applied Probability, 21(1):1–9, March 1984. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3213659.

Cressman:1984:LE

[CH84b] R. Cressman and W. G. S. Hines. Letter to the Editor. Journal of Applied Probability, 21(1):213–214, March 1984. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3213682.

Cressman:1990:NSI

[CH90] R. Cressman and W. G. S. Hines. A note on the stability of interior ESSs in a diploid population. Journal of Applied Probability, 27(2):476–477, June 1990. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3214671.

Carnal:1991:CHR

[CH91] H. Carnal and J. Hüsler. On the convex hull of n random points on a circle. Journal of Applied Probability, 28(1):231–237, March 1991. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3214754.

Christakos:1997:SIA

[CH97] George Christakos and Dionissios T. Hristopulos. Stochastic indicator analysis of contaminated sites. Journal of Applied Probability, 34(4):988–1008, December 1997. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http:/

/www.jstor.org/stable/3215012.

Chen:2006:TPR

[CH06] May-Ru Chen and Shoou-Ren Hsiau. Two-person red-and-black games with bet-dependent win probability functions. Journal of Applied Probability, 43(4):905–915, December 2006. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/27595788.

Chen:2010:TNM

[CH10] May-Ru Chen and Shoou-Ren Hsiau. Two new models for the two-person red-and-black game. Journal of Applied Probability, 47(1):97–108, March 2010. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/25662503.

Chen:2013:INS

[CH13] Feng Chen and Peter Hall. Inference for a nonstationary selfexciting point process with an application in ultra-high frequency financial data modeling. Journal of Applied Probability, 50(4):1006–1024, December 2013. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http:/

/projecteuclid.org/euclid.jap/1389370096.

Cesana:2022:PMI

[CH22] Pierluigi Cesana and Ben M. Hambly. A probabilistic model for interfaces in a martensitic phase transition. Journal of Applied Probability, 59(4):1081–1105, December 2022. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL https://www.cambridge.org/core/journals/journal-
of-applied-probability/article/probabilistic-model-for-
interfaces-in-a-martensitic-phase-transition/4DA2126B1EE438A0B5C106340458DB9A.

Chanda:1974:SMP

[Cha74] K. C. Chanda. Strong mixing properties of linear stochastic processes. Journal of Applied Probability, 11(2):401–408, June 1974. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3212764.

Chan:1975:SLB

[Cha75] Arthur H. C. Chan. Some lower bounds for the distribution of the supremum of the Yeh–Wiener process over a rectangular region. Journal of Applied Probability, 12(4):824–830, December 1975. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3212734.

Charalambides:1983:RPC

[Cha83] Ch. A. Charalambides. On restricted and pseudo-contagious occupancy distributions. Journal of Applied Probability, 20(4): 872–876, December 1983. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3213598.

Chao:1993:KMT

[Cha93] Xiuli Chao. On Klimov’s model with two job classes and exponential processing times. Journal of Applied Probability, 30(3): 716–724, September 1993. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3214777.

Chang:1994:IOM

[Cha94] Cheng-Shang Chang. On the input-output map of a G/G/1 queue. Journal of Applied Probability, 31(4):1128–1133, December 1994. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3215337.

Chao:1997:PBB

[Cha97] Xiuli Chao. Partial balances in batch arrival batch service and assemble-transfer queueing networks. Journal of Applied Probability, 34(3):745–752, September 1997. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3215099.

Cha:2000:BBP

[Cha00] Ji Hwan Cha. On a better burn-in procedure. Journal of Applied Probability, 37(4):1099–1103, December 2000. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3215498.

Cha:2001:BPG

[Cha01] Ji Hwan Cha. Burn-in procedures for a generalized model. Journal of Applied Probability, 38(2):542–553, June 2001. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3215905.

Cha:2003:FEG

[Cha03] Ji Hwan Cha. A further extension of the generalized burn-in model. Journal of Applied Probability, 40(1):264–270, March 2003. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3215854.

Cha:2004:NUB

[Cha04] Ji Hwan Cha. A note on the upper bound for the optimal work size. Journal of Applied Probability, 41(1):277–280, March 2004. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/3215832.

Cha:2013:SPF

[Cha13] Ji Hwan Cha. On stochastic predictions of failure processes under population heterogeneity. Journal of Applied Probability, 50(3):810–826, September 2013. CODEN JPRBAM.

ISSN 0021-9002 (print), 1475-6072 (electronic). URL http:/

/projecteuclid.org/euclid.jap/1378401238.

Chang:2017:ACD

[Cha17] Tongkeun Chang. Absolute continuity of distributions of one-dimensional Lévy processes. Journal of Applied Probability, 54(3):852–872, September 2017. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL https:/

/www.cambridge.org/core/journals/journal-of-applied-
probability/article/absolute-continuity-of-distributions-
of-onedimensional-levy-processes/832D8CD7DDF0E5A750C98707790427C1.

Cheong:1970:QSD

[Che70] C. K. Cheong. Quasi-stationary distributions in semi-Markov processes. Journal of Applied Probability, 7(2):388–399, August 1970. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3211972.

See correction [BNY69].

Chen:1976:SDR

[Che76a] Robert Chen. Subfair discounted red-and-black game with a house limit. Journal of Applied Probability, 13(3):608–613, September 1976. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3212483.

Chesson:1976:NCM

[Che76b] Jean Chesson. A non-central multivariate hypergeometric distribution arising from biased sampling with application to selective predation. Journal of Applied Probability, 13(4):795–797, December 1976. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3212535.

Chen:1980:WFZ

[Che80a] Wen-Chen Chen. On the weak form of Zipf’s law. Journal of Applied Probability, 17(3):611–622, September 1980. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3212955.

Chernick:1980:LTM

[Che80b] Michael R. Chernick. A limit theorem for the maximum term in a particular EARMA(1, 1) sequence. Journal of Applied Probability, 17(3):869–873, September 1980. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3212983. See correction [Ano81e].

Chen:1981:LTG

[Che81a] Wen-Chen Chen. Limit theorems for general size distributions. Journal of Applied Probability, 18(1):139–147, March 1981. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/3213174.

Chernick:1981:SML

[Che81b] Michael R. Chernick. On strong mixing and Leadbetter’s D condition. Journal of Applied Probability, 18(3):764–769, September 1981. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3213334.

Chernick:1982:LE

[Che82] Michael R. Chernick. Letter to the Editor. Journal of Applied Probability, 19(1):250, March 1982. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3213937.

Chen:1991:TCP

[Che91] Mu-Fa Chen. On three classical problems for Markov chains with continuous time parameters. Journal of Applied Probability, 28 (2):305–320, June 1991. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3214868.

Chen:1994:CLD

[Che94] Yi-Hau Chen. Classes of life distributions and renewal counting process. Journal of Applied Probability, 31(4):1110–1115, December 1994. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3215334.

Chen:1996:RCF

[Che96] Hong Chen. Rate of convergence of the fluid approximation for generalized Jackson networks. Journal of Applied Probability, 33 (3):804–814, September 1996. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.

org/stable/3215360.

Chen:1997:ECT

[Che97] Rong-Rong Chen. An extended class of time-continuous branching processes. Journal of Applied Probability, 34(1):14–23, March 1997. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3215170.

Chen:1998:SAN

[Che98] Han-Fu Chen. Stochastic approximation with non-additive measurement noise. Journal of Applied Probability, 35(2):407– 417, June 1998. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3215694.

Chen:2001:ICB

[Che01] Dayue Chen. On the infinite cluster of Bernoulli bond percolation in Scherk’s graph. Journal of Applied Probability, 38(4): 828–840, December 2001. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3215767.

Chen:2002:ESG

[Che02] Anyue Chen. Ergodicity and stability of generalised Markov branching processes with resurrection. Journal of Applied Probability, 39(4):786–803, December 2002. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3216003.

Chen:2003:CLT

[Che03a] Jinwen Chen. From the central limit theorem to heavy-tailed distributions. Journal of Applied Probability, 40(3):803–806, September 2003. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3215954.

Chen:2003:LBU

[Che03b] John T. Chen. A lower bound using Hamilton-type circuits and its applications. Journal of Applied Probability, 40(4):1121–1132, December 2003. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3216064.

Chen:2006:CLP

[Che06] Yung-Pin Chen. A central limit property under a modified Ehrenfest urn design. Journal of Applied Probability, 43(2):409– 420, June 2006. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
27595740.

Cheung:2007:CCC

[Che07] Ka Chun Cheung. Characterizations of conditional comonotonicity. Journal of Applied Probability, 44(3):607–617, September 2007. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/27595869.

Chen:2011:FTR

[Che11] Yiqing Chen. The finite-time ruin probability with dependent insurance and financial risks. Journal of Applied Probability, 48 (4):1035–1048, December 2011. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.

org/stable/23066441.

Chevallier:2018:UDP

[Che18] Julien Chevallier. Uniform decomposition of probability measures: quantization, clustering and rate of convergence. Journal of Applied Probability, 55(4):1037–1045, December 2018. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL https://www.cambridge.org/core/journals/journal-
of-applied-probability/article/uniform-decomposition-
of-probability-measures-quantization-clustering-and-
rate-of-convergence/DBBC1764874C2F56EA934D6DDB6EF194.

Chia:1968:RMP

[Chi68] A. B. Chia. Random mating in a population of cyclic size. Journal of Applied Probability, 5(1):21–30, April 1968. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3212074.

Cali:2024:CED

[CHJW24] Elie Cali, Alexander Hinsen, Benedikt Jahnel, and Jean-Philippe Wary. Chase-escape in dynamic device-to-device networks. Journal of Applied Probability, 61(1):311–339, March 2024. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL https://www.cambridge.org/core/journals/journal-
of-applied-probability/article/chaseescape-in-dynamic-
devicetodevice-networks/2EE1DA862BC4C36CAD24524E74C9689F.

Chao:1998:EMU

[CHL98] Chern-Ching Chao, Hsien-Kuei Hwang, and Wen-Qi Liang. Expected measure of the union of random rectangles. Journal of Applied Probability, 35(2):495–500, June 1998. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3215703.

Cornean:2022:CRV

[CHM^{+}22] Horia D. Cornean, Ira W. Herbst, Jesper Møller, Kasper S. Sørensen, and Benjamin B. Støttrup. Characterization of random variables with stationary digits. Journal of Applied Probability, 59(4):931–947, December 2022. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL https:/

/www.cambridge.org/core/journals/journal-of-applied-
probability/article/characterization-of-random-variables-
with-stationary-digits/12099A8D7108EA17AC08E8F1060BAF91.

Chow:1994:MLE

[Cho94] Yuan S. Chow. On the moments of ladder epochs for driftless random walks. Journal of Applied Probability, 31A:201–205, 1994. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3214957.

Chenavier:2022:LLL

[CHO22] Nicolas Chenavier, Norbert Henze, and Moritz Otto. Limit laws for large k-th-nearest neighbor balls. Journal of Applied Probability, 59(3):880–894, September 2022. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL https://www.cambridge.org/core/journals/journal-
of-applied-probability/article/limit-laws-for-large-
k-thnearest-neighbor-balls/AF39D4EE12343BC06E932F24562CC554.

Coleman:1997:CBP

[CHPT97] J. L. Coleman, W. Henderson, C. E. M. Pearce, and P. G. Taylor. A correspondence between product-form batch-movement queueing networks and single-movement networks. Journal of Applied Probability, 34(1):160–175, March 1997. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3215184.

Christensen:2012:PTD

[Chr12] Sören Christensen. Phase-type distributions and optimal stopping for autoregressive processes. Journal of Applied Probability, 49(1):22–39, March 2012. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/41713750.

Chao:2015:TTQ

[CHR15] Xiuli Chao, Qi-Ming He, and Sheldon Ross. Tollbooth tandem queues with infinite homogeneous servers. Journal of Applied Probability, 52(4):941–961, December 2015. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

projecteuclid.org/euclid.jap/1450802745.

Chung:1975:BDR

[Chu75] Kai Lai Chung. A bivariate distribution in regeneration. Journal of Applied Probability, 12(4):837–839, December 1975. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3212736.

Chu:2018:SNL

[Chu18] Weijuan Chu. Self-normalized large deviation for supercritical branching processes. Journal of Applied Probability, 55(2):450–458, June 2018. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL https:/

/www.cambridge.org/core/journals/journal-of-applied-
probability/article/selfnormalized-large-deviation-for-
supercritical-branching-processes/D401E4D1AECF4CE28958D0C07EE51FAA.

Chen:2014:NTU

[CHY14] May-Ru Chen, Shoou-Ren Hsiau, and Ting-Hsin Yang. A new two-urn model. Journal of Applied Probability, 51(2):590–597, June 2014. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-

6072 (electronic). URL http://projecteuclid.org/euclid.
jap/1402578645.

Chaiken:1972:EEF

[CI72] Jan M. Chaiken and Edward Ignall. An extension of Erlang’s formulas which distinguishes individual servers. Journal of Applied Probability, 9(1):192–197, March 1972. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/3212648.

Cox:1978:SEP

[CI78] D. R. Cox and V. Isham. Series expansions for the properties of a birth process of controlled variability. Journal of Applied Probability, 15(3):610–616, September 1978. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/3213123.

Cinlar:1967:QSM

[Çin67a] Erhan Çinlar. Queues with semi-Markovian arrivals. Journal of Applied Probability, 4(2):365–379, August 1967. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3212030.

Cinlar:1967:TDQ

[Çin67b] Erhan Çinlar. Time dependence of queues with semi-Markovian services. Journal of Applied Probability, 4(2):356–364, August 1967. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3212029.

Cinlar:1968:SDP

[Çin68] Erhan Çinlar. On the superposition of m-dimensional point processes. Journal of Applied Probability, 5(1):169–176, April 1968. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/3212084.

Cinlar:1980:GGP

[Çin80] Erhan Çinlar. On a generalization of gamma processes. Journal of Applied Probability, 17(2):467–480, June 1980. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/3213036. See correction

[Ano80e].

Cinque:2023:RPF

[Cin23] Fabrizio Cinque. Reflection principle for finite-velocity random motions. Journal of Applied Probability, 60(2):479–492, June 2023. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-

6072 (electronic). URL https://www.cambridge.org/core/
journals/journal-of-applied-probability/article/reflection-
principle-for-finitevelocity-random-motions/6E2D69267193209958209FDDBCA4275B.

Cipra:1988:APO

[Cip88] Tomás Cipra. Autoregressive processes in optimization. Journal of Applied Probability, 25(2):302–312, June 1988. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3214438.

Cinlar:1973:TMV

[ÇJ73] Erhan Çinlar and Peter Jagers. Two mean values which characterize the Poisson process. Journal of Applied Probability, 10(3): 678–681, September 1973. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3212788.

Comets:1998:CLT

[CJ98] Francis Comets and Martin Janzura. A central limit theorem for conditionally centred random fields with an application to Markov fields. Journal of Applied Probability, 35(3):608–621, September 1998. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3215636.

Charpentier:2006:LDS

[CJ06] Arthur Charpentier and Alessandro Juri. Limiting dependence structures for tail events, with applications to credit derivatives. Journal of Applied Probability, 43(2):563–586, June 2006. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/27595751.

Carbonell:2008:WLL

[CJ08a] F. Carbonell and J. C. Jimenez. Weak local linear discretizations for stochastic differential equations with jumps. Journal of Applied Probability, 45(1):201–210, March 2008. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/27595934.

Connor:2008:OCA

[CJ08b] Stephen Connor and Saul Jacka. Optimal co-adapted coupling for the symmetric random walk on the hypercube. Journal of Applied Probability, 45(3):703–713, September 2008. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/27595981.

Cohen:2010:PDG

[CJV10] Netta Cohen, Jonathan Jordan, and Margaritis Voliotis. Preferential duplication graphs. Journal of Applied Probability, 47 (2):572–585, June 2010. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/20778498.

Cox:1981:CTC

[CK81] J. Theodore Cox and Harry Kesten. On the continuity of the time constant of first-passage percolation. Journal of Applied Probability, 18(4):809–819, December 1981. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/3213056.

Cohen:1990:PCQ

[CK90] Joel E. Cohen and Frank P. Kelly. A paradox of congestion in a queuing network. Journal of Applied Probability, 27(3):730–734, September 1990. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3214558.

Cai:2000:NPC

[CK00a] Jun Cai and Vladimir Kalashnikov. NWU property of a class of random sums. Journal of Applied Probability, 37(1):283– 289, March 2000. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3215677.

Chen:2000:PAC

[CK00b] Chingfer Chen and Samuel Karlin. Poisson approximations for conditional r-scan lengths of multiple renewal processes and application to marker arrays in biomolecular sequences. Journal of Applied Probability, 37(3):865–880, September 2000. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3215619.

Choi:2000:SRC

[CK00c] Bong Dae Choi and Bara Kim. Sharp results on convergence rates for the distribution of GI/M/1/K queues as K tends to infinity. Journal of Applied Probability, 37(4):1010–1019, December 2000. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3215491.

Chen:2005:SCS

[CK05] Zengjing Chen and Reg Kulperger. A stochastic competingspecies model and ergodicity. Journal of Applied Probability, 42 (3):738–753, September 2005. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.

org/stable/30040853.

Clark:2007:DAM

[CK07] Steven P. Clark and Peter C. Kiessler. A diffusion approximation for Markov renewal processes. Journal of Applied Probability, 44(2):366–378, June 2007. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/27595847.

Chen:2013:GPU

[CK13] May-Ru Chen and Markus Kuba. On generalized Pólya urn models. Journal of Applied Probability, 50(4):1169–1186, December 2013. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-

6072 (electronic). URL http://projecteuclid.org/euclid.
jap/1389370106.

Chandramohan:1985:BMM

[CL85] J. Chandramohan and Lung-Kuang Liang. Bernoulli, multinomial and Markov chain thinning of some point processes and some results about the superposition of dependent renewal processes. Journal of Applied Probability, 22(4):828–835, December 1985. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3213950.

Cressie:1992:NMM

[CL92] Noel Cressie and Subhash Lele. New models for Markov random fields. Journal of Applied Probability, 29(4):877–884, December 1992. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3214720.

Conolly:1993:NFT

[CL93] B. W. Conolly and Christos Langaris. On a new formula for the transient state probabilities for M/M/1 queues and computational implications. Journal of Applied Probability, 30(1):237– 246, March 1993. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3214635.

Chong:1997:CCS

[CL97] K. S. Chong and K. Lam. Cost comparison of a spectrum of selforganizing rules. Journal of Applied Probability, 34(3):583–592, September 1997. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3215086.

Chong:1998:APR

[CL98] K. S. Chong and K. Lam. Applying POS(i) rules to communication problems. Journal of Applied Probability, 35(3):762–769, September 1998. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3215650.

Chen:2003:BDP

[CL03] Anyue Chen and Kai Liu. Birth-death processes with an instantaneous reflection barrier. Journal of Applied Probability, 40 (1):163–179, March 2003. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3215847.

Cai:2005:CTE

[CL05] Jun Cai and Haijun Li. Conditional tail expectations for multivariate phase-type distributions. Journal of Applied Probability, 42(3):810–825, September 2005. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.

org/stable/30040858.

Cheung:2009:PMR

[CL09] Eric C. K. Cheung and David Landriault. Perturbed map risk models with dividend barrier strategies. Journal of Applied Probability, 46(2):521–541, June 2009. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/25662441.

Cha:2011:SBM

[CL11] Ji Hwan Cha and Eui Yong Lee. A stochastic breakdown model for an unreliable Web server system and an optimal admission control policy. Journal of Applied Probability, 48(2):453– 466, June 2011. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
23065896.

Christiansen:2014:IAS

[CL14] Marcus C. Christiansen and Nicola Loperfido. Improved approximation of the sum of random vectors by the skew normal distribution. Journal of Applied Probability, 51(2):466–482, June 2014. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-

6072 (electronic). URL http://projecteuclid.org/euclid.
jap/1402578637.

Chen:2023:SFE

[CL23] Shukai Chen and Zenghu Li. Strong Feller and ergodic properties of the (1 + 1)-affine process. Journal of Applied Probability, 60(3):812–834, September 2023. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL https:/

/www.cambridge.org/core/journals/journal-of-applied-
probability/article/strong-feller-and-ergodic-properties-
of-the-11affine-process/9AEDE0A703122A4D4A6525597C923535.

Clancy:1994:SCR

[Cla94] Damian Clancy. Some comparison results for multitype epidemic models. Journal of Applied Probability, 31(1):9–21, March 1994. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/3215231.

Clancy:1999:OEM

[Cla99] Damian Clancy. Outcomes of epidemic models with general infection and removal rate functions at certain stopping times. Journal of Applied Probability, 36(3):799–813, September 1999. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/3215442.

Clancy:2005:SSI

[Cla05] Damian Clancy. A stochastic SIS infection model incorporating indirect transmission. Journal of Applied Probability, 42(3): 726–737, September 2005. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/30040852.

Clancy:2012:AQD

[Cla12] Damian Clancy. Approximating quasistationary distributions of Birth–Death processes. Journal of Applied Probability, 49(4): 1036–1051, December 2012. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/41713827.

Chen:2014:ABE

[CLCZ14] Anyue Chen, Junping Li, Yiqing Chen, and Dingxuan Zhou. Asymptotic behaviour of extinction probability of interacting branching collision processes. Journal of Applied Probability, 51 (1):219–234, March 2014. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://projecteuclid.

org/euclid.jap/1395771425.

Cline:1983:LTS

[Cli83] Daren B. H. Cline. Limit theorems for the shifting level process. Journal of Applied Probability, 20(2):322–337, June 1983. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/3213805.

Chen:2009:IEA

[CLS09] Yu-Ting Chen, Cheng-Few Lee, and Yuan-Chung Sheu. An integral-equation approach for defaultable bond prices with application to credit spreads. Journal of Applied Probability, 46 (1):71–84, March 2009. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/25662408.

Chen:2024:CCF

[CLZ24] Zengjing Chen, Yiwei Lin, and Jichen Zhang. A confirmation of a conjecture on Feldman’s two-armed bandit problem. Journal of Applied Probability, 61(1):121–136, March 2024. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL https://www.cambridge.org/core/journals/
journal-of-applied-probability/article/confirmation-
of-a-conjecture-on-feldmans-twoarmed-bandit-problem/
5EDA51A210FDC01CC7917550AA8CA901.

Crump:1969:ADB

[CM69] Kenny S. Crump and Charles J. Mode. An age-dependent branching process with correlations among sister cells. Journal of Applied Probability, 6(1):205–210, April 1969. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3212288.

Cane:1982:SIH

[CM82] Violet Cane and Rosanne McNamee. The spread of infection in a heterogeneous population. Journal of Applied Probability, 19A:173–184, 1982. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3213558.

Canfield:1983:ELD

[CM83] E. R. Canfield and W. P. McCormick. Exact and limiting distribution of sustained maxima. Journal of Applied Probability, 20(4):803–813, December 1983. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.

org/stable/3213591.

Canfield:1992:ARC

[CM92] E. R. Canfield and W. P. McCormick. Asymptotic reliability of consecutive k-out-of-n systems. Journal of Applied Probability, 29(1):142–155, March 1992. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3214798.

Calzolari:1997:LMO

[CM97] Antonella Calzolari and Federico Marchetti. Limit motion of an Ornstein–Uhlenbeck particle on the equilibrium manifold of a force field. Journal of Applied Probability, 34(4):924–938, December 1997. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3215007.

Chick:1998:NCN

[CM98] Stephen E. Chick and Max B. Mendel. New characterizations of the no-aging property and the ℓ_{1}-isotropic model. Journal of Applied Probability, 35(4):903–910, December 1998. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3215526.

Commault:2002:GPP

[CM02] Christian Commault and Stéphane Mocanu. A generic property of phase-type representations. Journal of Applied Probability, 39 (4):775–785, December 2002. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.

org/stable/3216002.

Cha:2007:SSF

[CM07] Ji Hwan Cha and Jie Mi. Study of a stochastic failure model in a random environment. Journal of Applied Probability, 44 (1):151–163, March 2007. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/27595828.

Cheng:2015:EIO

[CM15a] Li-Juan Cheng and Yong-Hua Mao. Eigentime identity for onedimensional diffusion processes. Journal of Applied Probability, 52(1):224–237, March 2015. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://projecteuclid.

org/euclid.jap/1429282617.

Crane:2015:PSP

[CM15b] Harry Crane and Peter McCullagh. Poisson superposition processes. Journal of Applied Probability, 52(4):1013–1027, December 2015. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-

6072 (electronic). URL http://projecteuclid.org/euclid.
jap/1450802750.

Crane:2015:RMS

[CM15c] Harry Crane and Peter McCullagh. Reversible Markov structures on divisible set partitions. Journal of Applied Probability, 52(3):622–635, September 2015. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http:/

/projecteuclid.org/euclid.jap/1445543836.

Chen:2016:DRS

[CM16a] Chen Chen and Hosam Mahmoud. Degrees in random selfsimilar bipolar networks. Journal of Applied Probability, 53 (2):434–447, June 2016. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://projecteuclid.

org/euclid.jap/1466172865.

Crocce:2016:FEA

[CM16b] Fabián Crocce and Ernesto Mordecki. A finite exact algorithm to solve a dice game. Journal of Applied Probability, 53(1):91–105, March 2016. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-

6072 (electronic). URL http://projecteuclid.org/euclid.
jap/1457470561.

Connor:2024:OCJ

[CM24] Stephen B. Connor and Roberta Merli. Optimal coupling of jumpy Brownian motion on the circle. Journal of Applied Probability, 61(1):18–32, March 2024. CODEN JPRBAM.

ISSN 0021-9002 (print), 1475-6072 (electronic). URL https:/

/www.cambridge.org/core/journals/journal-of-applied-
probability/article/optimal-coupling-of-jumpy-brownian-
motion-on-the-circle/14C65CFBA7588133DC3E87DB860BBCDF.

Colangelo:2006:PDW

[CMS06] A. Colangelo, A. Müller, and M. Scarsini. Positive dependence and weak convergence. Journal of Applied Probability, 43(1):48– 59, March 2006. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
27595709.

Comets:2023:GFF

[CMV23] Francis Comets, Mikhail Menshikov, and Stanislav Volkov. Generalizations of forest fires with ignition at the origin. Journal of Applied Probability, 60(2):418–434, June 2023. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL https://www.cambridge.org/core/journals/journal-
of-applied-probability/article/generalizations-of-forest-
fires-with-ignition-at-the-origin/C510265CB2DEEF68620A8B97F0DA31AE.

Campbell:1995:DES

[CMW95] L. L. Campbell, A. L. McKellips, and P. H. Wittke. Distributions and expectations of singular random variables. Journal of Applied Probability, 32(4):1032–1040, December 1995. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/3215217. See correction

[Ano98e].

Collevecchio:2016:LPR

[CMW16] Andrea Collevecchio, Abbas Mehrabian, and Nick Wormald. Longest paths in random Apollonian networks and largest rary subtrees of random d-ary recursive trees. Journal of Applied Probability, 53(3):846–856, September 2016. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

projecteuclid.org/euclid.jap/1476370780.

Comtet:1998:EFB

[CMY98] Alain Comtet, Cécile Monthus, and Marc Yor. Exponential functionals of Brownian motion and disordered systems. Journal of Applied Probability, 35(2):255–271, June 1998. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3215683.

Chiang:1980:DTF

[CN80] Dalen T. Chiang and Shun-Chen Niu. On the distribution of time to first system failure. Journal of Applied Probability, 17 (2):481–489, June 1980. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3213037.

Cooper:1986:BFM

[CN86] Robert B. Cooper and Shun-Chen Niu. Beneš’s formula for M/G/1-FIFO ‘explained’ by preemptive-resume LIFO. Journal of Applied Probability, 23(2):550–554, June 1986. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3214199.

Chen:1984:OSS

[CNO^{+}84] R. W. Chen, V. N. Nair, A. M. Odlyzko, L. A. Shepp, and Y. Vardi. Optimal sequential selection of N random variables under a constraint. Journal of Applied Probability, 21(3):537–547, September 1984. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3213615.

Cooper:1999:SPM

[CNS99] Robert B. Cooper, Shun-Chen Niu, and Mandyam M. Srinivasan. Setups in polling models: Does it make sense to set up if no work is waiting? Journal of Applied Probability, 36(2):585– 592, June 1999. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3215476.

Chen:2014:UDP

[CNZ14] Anyue Chen, Kai Wang Ng, and Hanjun Zhang. Uniqueness and decay properties of Markov branching processes with disasters. Journal of Applied Probability, 51(3):613–624, September 2014. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://projecteuclid.org/euclid.jap/
1409932662.

Cruz-Orive:1985:EVS

[CO85] Luis-M. Cruz-Orive. Estimating volumes from systematic hyperplane sections. Journal of Applied Probability, 22(3):518–530, September 1985. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3213857.

Coffey:1995:LBD

[Cof95] John Coffey. A linear birth-and-death predator-prey process. Journal of Applied Probability, 32(1):274–277, March 1995. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/3214936.

Cohen:1967:DMN

[Coh67a] J. W. Cohen. The distribution of the maximum number of customers present simultaneously during a busy period for the queueing systems M/G/1 and G/M/1. Journal of Applied Probability, 4(1):162–179, April 1967. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3212309.

Cohen:1967:TIE

[Coh67b] J. W. Cohen. On two integral equations of queueing theory. Journal of Applied Probability, 4(2):343–355, August 1967. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/3212028.

Cohen:1968:SSQ

[Coh68] J. W. Cohen. Single server queue with uniformly bounded virtual waiting time. Journal of Applied Probability, 5(1):93–122, April 1968. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3212080.

Cohen:1971:BPM

[Coh71] J. W. Cohen. On the busy periods for the M/G/1 queue with finite and with infinite waiting room. Journal of Applied Probability, 8(4):821–827, December 1971. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3212246.

Cohen:1973:SRR

[Coh73] J. W. Cohen. Some results on regular variation for distributions in queueing and fluctuation theory. Journal of Applied Probability, 10(2):343–353, June 1973. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3212351.

Cohen:1976:SSQ

[Coh76] J. W. Cohen. On a single-server queue with group arrivals. Journal of Applied Probability, 13(3):619–622, September 1976. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/3212485.

Cohen:1977:D

[Coh77a] J. W. Cohen. On up- and downcrossings. Journal of Applied Probability, 14(2):405–410, June 1977. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3213014.

Cohn:1977:CSB

[Coh77b] Harry Cohn. On the convergence of the supercritical branching processes with immigration. Journal of Applied Probability, 14 (2):387–390, June 1977. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3213010.

Cohen:1981:OFP

[Coh81a] J. W. Cohen. Obituary: Félix Pollaczek. Journal of Applied Probability, 18(4):959–963, December 1981. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/3213073.

Cohn:1981:CSM

[Coh81b] Harry Cohn. On the convergence of stochastically monotone sequences of random variables and some applications. Journal of Applied Probability, 18(3):592–605, September 1981. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3213314.

Cohn:1982:ALF

[Coh82] Harry Cohn. Another look at the finite mean supercritical Bienaymé–Galton–Watson. Journal of Applied Probability, 19A: 307–312, 1982. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3213569.

Cohn:1983:FSM

[Coh83] Harry Cohn. On the fluctuation of stochastically monotone Markov chains and some applications. Journal of Applied Probability, 20(1):178–184, March 1983. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3213733.

Cohen:1984:PSR

[Coh84] J. W. Cohen. On processor sharing and random service. Journal of Applied Probability, 21(4):937, December 1984. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3213712.

Cohen:1988:ETH

[Coh88] J. W. Cohen. On entrance times of a homogeneous Ndimensional random walk: An identity. Journal of Applied Probability, 25A:321–333, 1988. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3214166.

Cohen:1989:CPL

[Coh89a] Jonathan Cohen. On the compound Poisson limit for high level exceedances. Journal of Applied Probability, 26(3):458–465, September 1989. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3214404.

Cohn:1989:MFM

[Coh89b] Harry Cohn. Multitype finite mean supercritical age-dependent branching processes. Journal of Applied Probability, 26(2):398– 403, June 1989. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3214045.

Cohen:1994:CTD

[Coh94] J. W. Cohen. On a class of two-dimensional nearest-neighbour random walks. Journal of Applied Probability, 31A:207–237, 1994. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3214958.

Cohen:2014:UMC

[Coh14] Samuel N. Cohen. Undiscounted Markov chain BSDEs to stopping times. Journal of Applied Probability, 51(1):262–281, March 2014. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://projecteuclid.org/euclid.jap/
1395771428.

Coleman:1969:RPT

[Col69] Rodney Coleman. Random paths through convex bodies. Journal of Applied Probability, 6(2):430–441, August 1969. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3212012.

Collings:1974:DRO

[Col74] P. S. Collings. Dams with random outputs. Journal of Applied Probability, 11(4):858–861, December 1974. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/3212572.

Collings:1975:DAI

[Col75] P. S. Collings. Dams with autoregressive inputs. Journal of Applied Probability, 12(3):533–541, September 1975. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3212868.

Collings:1976:LTP

[Col76] P. S. Collings. Limit theorems for processes of randomly displaced regular events. Journal of Applied Probability, 13(3): 530–537, September 1976. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3212472.

Coleman:1981:ILR

[Col81] Rodney Coleman. Intercept lengths of random probes through boxes. Journal of Applied Probability, 18(1):276–282, March 1981. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3213189.

Conolly:1971:SAT

[Con71] B. W. Conolly. Some applications of the theory of infinite capacity service systems to a single server system with linearly state dependent service. Journal of Applied Probability, 8(1): 202–207, March 1971. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3211850.

Conolly:1974:GSD

[Con74] B. W. Conolly. The generalised state-dependent Erlangian queue: The busy period. Journal of Applied Probability, 11(3):

618–623, September 1974. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/ stable/3212711. See corrections [Ano74h].

Conolly:1975:GSD

[Con75] B. W. Conolly. Generalized state-dependent Erlangian queues: Speculations about calculating measures of effectiveness. Journal of Applied Probability, 12(2):358–363, June 1975. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3212451.

Conolly:1984:AQP

[Con84] B. W. Conolly. The autostrada queueing problem. Journal of Applied Probability, 21(2):394–403, June 1984. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3213648.

Cong:1994:MGQ

[Con94] Tang Dac Cong. On the M^{X} /G/∞ queue with heterogeneous customers in a batch. Journal of Applied Probability, 31(1):280– 286, March 1994. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3215256.

Connor:2013:OCC

[Con13] Stephen Connor. Optimal coadapted coupling for a random walk on the hyper-complete graph. Journal of Applied Probability, 50(4):1117–1130, December 2013. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

projecteuclid.org/euclid.jap/1389370103.

Contat:2022:SIG

[Con22] Alice Contat. Surprising identities for the greedy independent set on Cayley trees. Journal of Applied Probability, 59 (4):1042–1058, December 2022. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL https:/

/www.cambridge.org/core/journals/journal-of-applied-
probability/article/surprising-identities-for-the-greedy-
independent-set-on-cayley-trees/31BF7C2FBD046EAD182F02CE43C56170.

Cooke:1974:BCP

[Coo74] Peter J. Cooke. Bounds for coverage probabilities with applications to sequential coverage problems. Journal of Applied Probability, 11(2):281–293, June 1974. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3212749.

Cooper:1989:RSP

[Coo89] Douglas W. Cooper. Random sequential packing simulations in three dimensions for aligned cubes. Journal of Applied Probability, 26(3):664–670, September 1989. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3214426.

Cooper:1998:NBS

[Coo98] William L. Cooper. Negative binomial sums of random variables and discounted reward processes. Journal of Applied Probability, 35(3):589–599, September 1998. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.

org/stable/3215634.

Corwin:1980:SBT

[Cor80] Thomas L. Corwin. On search for a Brownian target. Journal of Applied Probability, 17(1):243–247, March 1980. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3212941.

Corwin:1981:SMT

[Cor81] Thomas L. Corwin. Search for a moving target under minimax estimation. Journal of Applied Probability, 18(1):167– 180, March 1981. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3213177.

Cornez:1987:BDP

[Cor87] Richard Cornez. Birth and death processes in random environments with feedback. Journal of Applied Probability, 24(1): 25–34, March 1987. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3214056.

Costa:1990:SDP

[Cos90] O. L. V. Costa. Stationary distributions for piecewisedeterministic Markov processes. Journal of Applied Probability, 27(1):60–73, March 1990. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3214595.

Costa:1993:DAI

[Cos93] O. L. V. Costa. Discretizations for the average impulse control of piecewise deterministic processes. Journal of Applied Probability, 30(2):405–420, June 1993. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3214849.

Covo:2009:ASJ

[Cov09] Shai Covo. On approximations of small jumps of subordinators with particular emphasis on a Dickman-type limit. Journal of Applied Probability, 46(3):732–755, September 2009. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/25662458.

Cowan:1978:IMS

[Cow78] Richard Cowan. An improved model for signalised intersections with vehicle-actuated control. Journal of Applied Probability, 15 (2):384–396, June 1978. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3213409.

Cowan:1979:UTM

[Cow79] Richard Cowan. The uncontrolled traffic merge. Journal of Applied Probability, 16(2):384–392, June 1979. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/3212905.

Cowan:1980:FRS

[Cow80] Richard Cowan. Further results on single-Lane traffic flow. Journal of Applied Probability, 17(2):523–531, June 1980. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3213041.

Cowan:1981:AFC

[Cow81] Richard Cowan. An analysis of the fixed-cycle traffic-light problem. Journal of Applied Probability, 18(3):672–683, September 1981. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3213321.

Cowan:1991:EFD

[Cow91] Richard Cowan. Expected frequencies of DNA patterns using Whittle’s formula. Journal of Applied Probability, 28(4):

886–892, December 1991. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3214691.

Cowan:1992:AOD

[Cow92] Richard Cowan. The allocation of offensive and defensive resources in a territorial game. Journal of Applied Probability, 29 (1):190–195, March 1992. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3214803.

Cowan:1993:CRP

[Cow93] Richard Cowan. A constraint on the random packing of disks. Journal of Applied Probability, 30(1):263–268, March 1993. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/3214639.

Cowan:2001:NDD

[Cow01] Richard Cowan. A new discrete distribution arising in a model of DNA replication. Journal of Applied Probability, 38(3):754–760, September 2001. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3216127.

Cowan:2004:MTC

[Cow04] Richard Cowan. A mosaic of triangular cells formed with sequential splitting rules. Journal of Applied Probability, 41A:3– 15, 2004. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3215964.

Cox:1977:EPT

[Cox77] Ted Cox. An example of phase transition in countable onedimensional Markov random fields. Journal of Applied Probability, 14(1):205–211, March 1977. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3213275.

Cinlar:1972:DAI

[ÇP72] E. Çinlar and M. Pinsky. On dams with additive inputs and a general release rule. Journal of Applied Probability, 9(2):422– 429, June 1972. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3212811.

Cohn:1978:RLR

[CP78] Harry Cohn and Anthony G. Pakes. A representation for the limiting random variable of a branching process with infinite mean and some related problems. Journal of Applied Probability, 15(2):225–234, June 1978. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3213396.

Cacoullos:1982:CNB

[CP82] Theophilos Cacoullos and H. Papageorgiou. Characterizing the negative binomial distribution. Journal of Applied Probability, 19(3):742–743, September 1982. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.

org/stable/3213540.

Chryssaphinou:1988:LTN

[CP88] O. Chryssaphinou and S. Papastavridis. A limit theorem for the number of non-overlapping occurrences of a pattern in a sequence of independent trials. Journal of Applied Probability, 25(2):428–431, June 1988. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3214452. See [Ano91e].

Chryssaphinou:1990:RCS

[CP90] Ourania Chryssaphinou and Stavros Papastavridis. Reliability of a consecutive-k-out-of-n system in a random environment. Journal of Applied Probability, 27(2):452–458, June 1990. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/3214666.

Choi:1992:MMQ

[CP92] Bong Dae Choi and Kwang Kyu Park. The M^{k} /m/∞ queue with heterogeneous customers in a batch. Journal of Applied Probability, 29(2):477–481, June 1992. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3214586.

Chae:2001:QLD

[CP01] Kyung C. Chae and Yonil Park. The queue length distribution for the M/G/1 queue under the D-policy. Journal of Applied Probability, 38(1):278–279, March 2001. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http:/

/www.jstor.org/stable/3215761.

Clancy:2003:NQS

[CP03] Damian Clancy and Philip K. Pollett. A note on quasi-stationary distributions of birth-death processes and the SIS logistic epidemic. Journal of Applied Probability, 40(3):821–825, September 2003. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3215957.

Cairns:2004:ETG

[CP04] Ben Cairns and P. K. Pollett. Extinction times for a general birth, death and catastrophe process. Journal of Applied Probability, 41(4):1211–1218, December 2004. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/4141394.

Czarna:2011:RPP

[CP11] Irmina Czarna and Zbigniew Palmowski. Ruin probability with Parisian delay for a spectrally negative Lévy risk process. Journal of Applied Probability, 48(4):984–1002, December 2011. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/23066438.

Campanino:2014:TTS

[CP14] Massimo Campanino and Dimitri Petritis. Type transition of simple random walks on randomly directed regular lattices. Journal of Applied Probability, 51(4):1065–1080, December 2014. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-

6072 (electronic). URL http://projecteuclid.org/euclid.
jap/1421763328.

Czuppon:2017:SMS

[CP17] Peter Czuppon and Peter Pfaffelhuber. A spatial model for selection and cooperation. Journal of Applied Probability, 54(2):522–539, June 2017. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL https:/

/www.cambridge.org/core/journals/journal-of-applied-
probability/article/spatial-model-for-selection-and-
cooperation/6AE45F14AA471867A90FDC1997CFC2DE.

Chen:2007:RUW

[CPLZ07] Anyue Chen, Phil Pollett, Junping Li, and Hanjun Zhang. A remark on the uniqueness of weighted Markov branching processes.

Journal of Applied Probability, 44(1):279–283, March 2007. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/27595837.

Coutin:2018:JDL

[CPN18] Laure Coutin, Monique Pontier, and Waly Ngom. Joint distribution of a Lévy process and its running supremum. Journal of Applied Probability, 55(2):488–512, June 2018. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL https:/

/www.cambridge.org/core/journals/journal-of-applied-
probability/article/joint-distribution-of-a-levy-process-
and-its-running-supremum/3B4E2CCA384068C02F19ABD486FCBD8D.

Chao:1996:NQB

[CPS96] Xiuli Chao, Michael Pinedo, and Dequan Shaw. Networks of queues with batch services and customer coalescence. Journal of Applied Probability, 33(3):858–869, September 1996. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3215364.

Casanova:2022:SSB

[CPSJ22] Adrián González Casanova, Lizbeth Peñaloza, and Arno Siri-Jégousse. The shape of a seed bank tree. Journal of Applied Probability, 59(3):631–651, September 2022. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL https://www.cambridge.org/core/journals/journal-
of-applied-probability/article/shape-of-a-seed-bank-
tree/20087F4BE69EC0548100E0336E5FAFF8.

Cvitanic:1999:SRS

[CPT99] Jaksa Cvitanić, Huyên Pham, and Nizar Touzi. Superreplication in stochastic volatility models under portfolio constraints. Journal of Applied Probability, 36(2):523–545, June 1999. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3215472.

Chan:1985:MTA

[CPTW85] K. S. Chan, Joseph D. Petruccelli, H. Tong, and Samuel W. Woolford. A multiple-threshold AR(1) model. Journal of Applied Probability, 22(2):267–279, June 1985. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/3213771.

Chen:2004:CBP

[CPZL04] Anyue Chen, Phil Pollett, Hanjun Zhang, and Junping Li. The collision branching process. Journal of Applied Probability, 41 (4):1033–1048, December 2004. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.

org/stable/4141376.

Capocelli:1972:IFP

[CR72] R. M. Capocelli and L. M. Ricciardi. On the inverse of the first passage time probability problem. Journal of Applied Probability, 9(2):270–287, June 1972. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.

org/stable/3212798.

Chen:1992:GDF

[CR92] Anyue Chen and Eric Renshaw. The gillis-domb-Fisher correlated random walk. Journal of Applied Probability, 29(4): 792–813, December 1992. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3214713.

Chang:1994:OLP

[CR94a] Cheng-Shang Chang and Rhonda Righter. The optimality of LEPT in parallel machine scheduling. Journal of Applied Probability, 31(3):788–796, September 1994. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3215156.

Chen:1994:GCR

[CR94b] Anyue Chen and Eric Renshaw. The general correlated random walk. Journal of Applied Probability, 31(4):869–884, December 1994. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3215313.

Chen:1997:MMQ

[CR97] Anyue Chen and Eric Renshaw. The M/M/1 queue with mass exodus and mass arrivals when empty. Journal of Applied Probability, 34(1):192–207, March 1997. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3215186.

Chernysh:2017:CNB

[CR17] Ksenia Chernysh and Sanjay Ramassamy. Coupling any number of balls in the infinite-bin model. Journal of Applied Probability, 54(2):540–549, June 2017. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL https:/

/www.cambridge.org/core/journals/journal-of-applied-
probability/article/coupling-any-number-of-balls-in-
the-infinitebin-model/D524BE9DD246BD8B222F2D850975AD98.

Craven:1969:ACQ

[Cra69] B. D. Craven. Asymptotic correlation in a queue. Journal of Applied Probability, 6(3):573–583, December 1969. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3212103.

Crane:1973:QTS

[Cra73a] Michael A. Crane. Queues in transportation systems, I: A Markovian system. Journal of Applied Probability, 10(3):630– 643, September 1973. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3212782.

Craven:1973:SDM

[Cra73b] B. D. Craven. The spectral density of a Markov process. Journal of Applied Probability, 10(3):520–527, September 1973. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3212773.

Crane:1974:MSA

[Cra74a] Michael A. Crane. Multi-server assembly queues. Journal of Applied Probability, 11(3):629–632, September 1974. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3212713.

Crane:1974:QTS

[Cra74b] Michael A. Crane. Queues in transportation systems, II: An independently-dispatched system. Journal of Applied Probability, 11(1):145–158, March 1974. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.

org/stable/3212591.

Crank:1988:MAP

[Cra88] Keith N. Crank. A method for approximating the probability functions of a Markov chain. Journal of Applied Probability, 25 (4):808–814, December 1988. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.

org/stable/3214302.

Crane:2011:CMP

[Cra11] Harry Crane. A consistent Markov partition process generated from the paintbox process. Journal of Applied Probability, 48(3): 778–791, September 2011. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/23065936.

Cressie:1974:TDR

[Cre74] Noel Cressie. A two-dimensional random walk in the presence of a partially reflecting barrier. Journal of Applied Probability, 11 (1):199–205, March 1974. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3212599.

Cressie:1977:SPS

[Cre77] Noel Cressie. On some properties of the scan statistic on the circle and the line. Journal of Applied Probability, 14(2):272– 283, June 1977. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3212998.

Crimaldi:2016:CLT

[Cri16] Irene Crimaldi. Central limit theorems for a hypergeometric randomly reinforced urn. Journal of Applied Probability, 53(3): 899–913, September 2016. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://projecteuclid.

org/euclid.jap/1476370784.

Criens:2018:SPE

[Cri18] David Criens. Structure-preserving equivalent martingale measures for H-SII models. Journal of Applied Probability, 55(1):1–14, March 2018. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL https:/

/www.cambridge.org/core/journals/journal-of-applied-
probability/article/structurepreserving-equivalent-martingale-
measures-for-sii-models/1B28060A162BE278BF59C9BAABA148CC.

Chen:1997:SPT

[CRS97] Robert W. Chen, Burton Rosenberg, and Larry A. Shepp. A secretary problem with two decision makers. Journal of Applied Probability, 34(4):1068–1074, December 1997. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3215019.

Crump:1970:MPB

[Cru70] Kenny S. Crump. Migratory populations in branching processes. Journal of Applied Probability, 7(3):565–572, December 1970. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/3211938.

Clifford:1972:LCS

[CS72] Peter Clifford and Aidan Sudbury. The linear cell-size-dependent branching process. Journal of Applied Probability, 9(4):687–696, December 1972. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3212608.

Chvatal:1975:LCS

[CS75] Vacláv Chvátal and David Sankoff. Longest common subsequences of two random sequences. Journal of Applied Probability, 12(2):306–315, June 1975. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3212444.

Chatterjee:1977:TCS

[CS77] S. Chatterjee and E. Seneta. Towards consensus: Some convergence theorems on repeated averaging. Journal of Applied Probability, 14(1):89–97, March 1977. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3213262.

Cohn:1980:CPF

[CS80] H. Cohn and H.-J. Schuh. On the continuity and the positivity of the finite part of the limit distribution of an irregular branching process with infinite mean. Journal of Applied Probability, 17(3): 696–703, September 1980. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3212963.

Cumberland:1982:WCA

[CS82] W. G. Cumberland and Z. M. Sykes. Weak convergence of an autoregressive process used in modeling population growth. Journal of Applied Probability, 19(2):450–455, June 1982. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3213500.

Chow:1985:SRP

[CS85a] Chung-Wen Chow and Zvi Schechner. On stopping rules in proofreading. Journal of Applied Probability, 22(4):971–977, December 1985. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3213967.

Clifford:1985:LBT

[CS85b] Peter Clifford and Aidan Sudbury. Looking backwards in time in the Moran model in population genetics. Journal of Applied Probability, 22(2):437–442, June 1985. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3213786.

Chao:1986:SED

[CS86] Chern-Ching Chao and John Slivka. Some exact distributions of a last one-sided exit time in the simple random walk. Journal of Applied Probability, 23(2):332–340, June 1986. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/3214177. See correction

[Ano86j].

Currit:1988:RFS

[CS88] Allen Currit and Nozer D. Singpurwalla. On the reliability function of a system of components sharing a common environment. Journal of Applied Probability, 25(4):763–771, December 1988. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/3214297.

Cohen:1995:SOR

[CS95] Arthur Cohen and Harold B. Sackrowitz. On stochastic ordering of random vectors. Journal of Applied Probability, 32(4): 960–965, December 1995. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3215209.

Chen:1996:MBT

[CS96] Tuhao Chen and E. Seneta. Multivariate Bonferroni-type lower bounds. Journal of Applied Probability, 33(3):729–740, September 1996. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3215354.

Costantini:1997:ESO

[CS97] C. Costantini and F. Spizzichino. Explicit solution of an optimal stopping problem: The burn-in of conditionally exponential components. Journal of Applied Probability, 34(1):267– 282, March 1997. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3215193.

Chen:2000:RMB

[CS00] Tuhao Chen and E. Seneta. A refinement of multivariate Bonferroni-type inequalities. Journal of Applied Probability, 37 (1):276–282, March 2000. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3215676.

Comets:2008:RWD

[CS08] Francis Comets and François Simenhaus. Random walk delayed on percolation clusters. Journal of Applied Probability, 45(3): 689–702, September 2008. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/27595980.

Cox:2012:BPV

[CS12] J. Theodore Cox and Rinaldo B. Schinazi. A branching process for virus survival. Journal of Applied Probability, 49(3):888–894, September 2012. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
41713814.

Casanova:2020:FVP

[CS20a] Adrian Gonzalez Casanova and Charline Smadi. On Λ–Fleming– Viot processes with general frequency-dependent selection. Journal of Applied Probability, 57(4):1162–1197, December 2020. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL https://www.cambridge.org/core/journals/
journal-of-applied-probability/article/on-flemingviot-

processes-with-general-frequencydependent-selection/
27BC3C7F6A4409ED450083D2EC317C67.

Cheek:2020:EMP

[CS20b] David Cheek and Seva Shneer. The empirical mean position of a branching Lévy process. Journal of Applied Probability, 57(4):1252–1259, December 2020. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL https:/

/www.cambridge.org/core/journals/journal-of-applied-
probability/article/empirical-mean-position-of-a-branching-
levy-process/40845A366B0DBE94AA256DD4BDAC4B29.

Chao:2023:SMM

[CS23] Brian Chao and Jason Schweinsberg. A spatial mutation model with increasing mutation rates. Journal of Applied Probability, 60(4):1157–1180, December 2023. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL https:/

/www.cambridge.org/core/journals/journal-of-applied-
probability/article/spatial-mutation-model-with-increasing-
mutation-rates/821CB2186A990C80694CE1965E6176E4.

Csaki:1994:SJD

[Csá94] Endre Csáki. Some joint distributions in Bernoulli excursions. Journal of Applied Probability, 31A:239–250, 1994. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3214959.

Chen:2013:CCB

[CSC13] Guan-Yu Chen and Laurent Saloff-Coste. Comparison of cutoffs between lazy walks and Markovian semigroups. Journal of Applied Probability, 50(4):943–959, December 2013. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://projecteuclid.org/euclid.jap/1389370092.

Csenki:1979:IPD

[Cse79] Attila Csenki. An invariance principle in k-dimensional extended renewal theory. Journal of Applied Probability, 16(3):567–574, September 1979. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3213085.

Csenki:1991:SRT

[Cse91] Attila Csenki. Some renewal-theoretic investigations in the theory of sojourn times in finite semi-Markov processes. Journal of Applied Probability, 28(4):822–832, December 1991. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3214685.

Csorgo:1980:QKK

[Csö80] Sándor Csörgő. On the quantogram of Kendall and Kent. Journal of Applied Probability, 17(2):440–447, June 1980. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3213033.

Costeniuc:2008:EED

[CST08] Marius Costeniuc, Michaela Schnetzer, and Luca Taschini. Entry and exit decision problem with implementation delay. Journal of Applied Probability, 45(4):1039–1059, December 2008. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/27596007.

Coolen-Schrijner:2001:CSB

[CSV01] Pauline Coolen-Schrijner and Erik A. Van Doorn. On the convergence to stationarity of birth-death processes. Journal of Applied Probability, 38(3):696–706, September 2001. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/3216122.

Chen:2005:OBP

[CSYZ05] Robert W. Chen, Larry A. Shepp, Yi-Ching Yao, and Cun-Hui Zhang. On optimality of bold play for primitive casinos in the presence of inflation. Journal of Applied Probability, 42(1):121– 137, March 2005. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
30040774.

Calvert:1997:BPQ

[CSZ97] Bruce Calvert, Wiremu Solomon, and Ilze Ziedins. Braess’s paradox in a queueing network with state-dependent routing. Journal of Applied Probability, 34(1):134–154, March 1997. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3215182.

Chen:2004:BSA

[CSZ04] Robert W. Chen, Larry A. Shepp, and Alan Zame. A bold strategy is not always optimal in the presence of inflation. Journal of Applied Probability, 41(2):587–592, June 2004. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3216039.

Cheong:1972:GST

[CT72] Choong K. Cheong and Jozef L. Teugels. General solidarity theorems for semi-Markov processes. Journal of Applied Probability, 9(4):789–802, December 1972. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3212615.

Cooper:1976:RBD

[CT76] Robert B. Cooper and Borge Tilt. On the relationship between the distribution of maximal queue length in the M/G/1 queue and the mean busy period in the M/G/1/n queue. Journal of Applied Probability, 13(1):195–199, March 1976. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3212684.

Cogburn:1981:BDP

[CT81] Robert Cogburn and William C. Torrez. Birth and death processes with random environments in continuous time. Journal of Applied Probability, 18(1):19–30, March 1981. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3213163.

Csorgo:1990:ELT

[CT90] Sándor Csörgő and Jef L. Teugels. Empirical Laplace transform and approximation of compound distributions. Journal of Applied Probability, 27(1):88–101, March 1990. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/3214597.

Coyle:1995:TBS

[CT95] A. J. Coyle and P. G. Taylor. Tight bounds on the sensitivity of generalised semi-Markov processes with a single generally distributed lifetime. Journal of Applied Probability, 32(1):63– 73, March 1995. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3214921.

Cottrell:2000:UHM

[CT00] M. Cottrell and T. S. Turova. Use of an Hourglass model in neuronal coding. Journal of Applied Probability, 37(1):168– 186, March 2000. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3215666.

Cai:2004:MSE

[CT04] Jun Cai and Qihe Tang. On max-sum equivalence and convolution closure of heavy-tailed distributions and their applications. Journal of Applied Probability, 41(1):117–130, March 2004. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/3215819.

Cocozza-Thivent:1983:SRN

[CTKR83] C. Cocozza-Thivent, C. Kipnis, and M. Roussignol. Stabilité de la récurrence nulle pour certaines chaînes de Markov perturbées. (French) [Stability of the null recurrence for some perturbed Markov chains]. Journal of Applied Probability, 20(3):482–504, September 1983. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3213886.

Curinao:2022:ABQ

[Cur22] Jorge Littin Curinao. Asymptotic behavior and quasi-limiting distributions on time-fractional birth and death processes. Journal of Applied Probability, 59(4):1199–1227, December 2022. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL https://www.cambridge.org/core/journals/
journal-of-applied-probability/article/asymptotic-behavior-
and-quasilimiting-distributions-on-timefractional-birth-
and-death-processes/80D2F20B1370B7DADD24ECD263A6A3DA.

Chryssaphinou:2001:CPA

[CV01] Ourania Chryssaphinou and Eutichia Vaggelatou. Compound Poisson approximation for long increasing sequences. Journal of Applied Probability, 38(2):449–463, June 2001. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3215899.

Christofides:2009:BDB

[CV09] Tasos C. Christofides and Eutichia Vaggelatou. Bounds for the distance between the distributions of sums of absolutely continuous I.I.D. convex-ordered random variables with applications. Journal of Applied Probability, 46(1):255–271, March 2009. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/25662419.

Courcoubetis:1984:IRS

[CVW84] C. Courcoubetis, P. Varaiya, and J. Walrand. Invariance in resource-sharing systems. Journal of Applied Probability, 21(4): 777–785, December 1984. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3213695.

Chia:1969:DER

[CW69] A. B. Chia and G. A. Watterson. Demographic effects on the rate of genetic evolution: I. Constant size populations with two genotypes. Journal of Applied Probability, 6(2):231–248, August 1969. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3211995.

Cabana:1981:ETD

[CW81] E. M. Cabaña and M. Wschebor. An estimate for the tails of the distribution of the supremum for a class of stationary multiparameter Gaussian processes. Journal of Applied Probability, 18(2):536–541, June 1981. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3213302.

Courcoubetis:1986:NSC

[CW86] C. Courcoubetis and R. R. Weber. Necessary and sufficient conditions for stability of a bin-packing system. Journal of Applied Probability, 23(4):989–999, December 1986. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/3214471.

Cao:1991:NNC

[CW91] Jinhua Cao and Yuedong Wang. The NBUC and NWUC classes of life distributions. Journal of Applied Probability, 28(2):473– 479, June 1991. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/

3214882. See correction [Ano92h].

Chayes:1993:DIN

[CW93] L. Chayés and C. Winfield. The density of interfaces: A new first-passage problem. Journal of Applied Probability, 30(4): 851–862, December 1993. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3214517.

Courcoubetis:1996:BOA

[CW96] Costas Courcoubetis and Richard Weber. Buffer overflow asymptotics for a buffer handling many traffic sources. Journal of Applied Probability, 33(3):886–903, September 1996. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3215366.

Cai:1997:CLD

[CW97] Jun Cai and Yanhong Wu. Characterization of life distributions under some generalized stochastic orderings. Journal of Applied Probability, 34(3):711–719, September 1997. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/3215096.

Chen:2005:NUM

[CW05] May-Ru Chen and Ching-Zong Wei. A new urn model. Journal of Applied Probability, 42(4):964–976, December 2005. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/30040821.

Cui:2001:AAP

[CX01] Lirong Cui and M. Xie. Availability analysis of periodically inspected systems with random walk model. Journal of Applied Probability, 38(4):860–871, December 2001. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/3215769.

Chover:1965:SST

[CY65] J. Chover and G. Yeo. Solutions of some two-sided boundary problems for sums of variables with alternating distributions. Journal of Applied Probability, 2(2):377–395, December 1965. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/3212200.

Chow:1984:SLT

[CY84] Y. S. Chow and K. F. Yu. Some limit theorems for a subcritical branching process with immigration. Journal of Applied Probability, 21(1):50–57, March 1984. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3213663.

Chen:1990:DED

[CY90] Hong Chen and David D. Yao. Derivatives of the expected delay in the GI/G/1 queue. Journal of Applied Probability, 27(4): 899–907, December 1990. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3214832.

Cheung:2008:OOP

[CY08] Ka Chun Cheung and Hailiang Yang. Ordering of optimal portfolio allocations in a model with a mixture of fundamental risks. Journal of Applied Probability, 45(1):55–66, March 2008. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/27595923.

Chen:2016:ODR

[CY16] Mi Chen and Kam Chuen Yuen. Optimal dividend and reinsurance in the presence of two reinsurers. Journal of Applied Probability, 53(2):554–571, June 2016. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

projecteuclid.org/euclid.jap/1466172874.

Cowan:1996:RSP

[CYF96] Richard Cowan, Philip L. H. Yu, and Thomas S. Ferguson. Restrictions on the saddle-point solution in the game of teamball. Journal of Applied Probability, 33(2):443–447, June 1996. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/3215069.

Chao:1998:RNQ

[CZ98] Xiuli Chao and Shaohui Zheng. A result on networks of queues with customer coalescence and state-dependent signaling. Journal of Applied Probability, 35(1):151–164, March 1998. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3215554.

Chow:1999:ISC

[CZ99] Yunshyong Chow and Yu Zhang. Inequalities for the score constant in matching random sequences. Journal of Applied Probability, 36(2):601–606, June 1999. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3215478.

Dahlhaus:1985:SDE

[Dah85] Rainer Dahlhaus. On a spectral density estimate obtained by averaging periodograms. Journal of Applied Probability, 22(3): 598–610, September 1985. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3213863.

Dai:2004:ASM

[Dai04] Wen Dai. Asymptotics of the sample mean and sample covariance of long-range-dependent series. Journal of Applied Probability, 41A:383–392, 2004. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3215991.

Dai:2014:ESD

[Dai14] Hongsheng Dai. Exact simulation for diffusion bridges: an adaptive approach. Journal of Applied Probability, 51(2):346–358, June 2014. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-

6072 (electronic). URL http://projecteuclid.org/euclid.
jap/1402578629.

Daley:1965:GCI

[Dal65] D. J. Daley. General customer impatience in the queue GI/G/1. Journal of Applied Probability, 2(1):186–205, June 1965. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/3211884.

Daley:1969:IRT

[Dal69a] D. J. Daley. Integral representations of transition probabilities and serial covariances of certain Markov chains. Journal of Applied Probability, 6(3):648–659, December 1969. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3212109.

Daley:1969:TWTa

[Dal69b] D. J. Daley. The total waiting time in a busy period of a stable single-server queue, I. Journal of Applied Probability, 6(3): 550–564, December 1969. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3212101.

Daley:1971:DMD

[Dal71] D. J. Daley. The definition of a multi-dimensional generalization of shot noise. Journal of Applied Probability, 8(1):128– 135, March 1971. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3211843.

Dallas:1974:CGD

[Dal74] A. C. Dallas. A characterization of the geometric distribution. Journal of Applied Probability, 11(3):609–611, September 1974. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/3212709.

Daley:1979:RPC

[Dal79a] D. J. Daley. Return probabilities for certain three-dimensional random walks. Journal of Applied Probability, 16(1):45–53, March 1979. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3213373.

Daly:1979:CSB

[Dal79b] Fergus Daly. Collapsing supercritical branching processes. Journal of Applied Probability, 16(4):732–739, December 1979. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/3213140.

Dallas:1981:RVE

[Dal81a] A. C. Dallas. Record values and the exponential distribution. Journal of Applied Probability, 18(4):949–951, December 1981. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/3213070.

Daly:1981:GBP

[Dal81b] Fergus Daly. On a general branching process. Journal of Applied Probability, 18(1):253–255, March 1981. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http:/

/www.jstor.org/stable/3213185.

Daley:1982:SPP

[Dal82] D. J. Daley. Stationary point processes with Markov-dependent intervals and infinite intensity. Journal of Applied Probability, 19A:313–320, 1982. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3213570.

Dallas:1987:CGD

[Dal87] A. C. Dallas. Characterizing the geometric distribution using expectations of order statistics. Journal of Applied Probability, 24(2):534–539, June 1987. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3214277.

Daley:2001:MIM

[Dal01a] D. J. Daley. The moment index of minima. Journal of Applied Probability, 38A:33–36, 2001. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.

org/stable/3215865.

Daley:2001:P

[Dal01b] Daryl Daley. Preface. Journal of Applied Probability, 38A:v–vi, 2001. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3215859.

Daly:2010:SMC

[Dal10] Fraser Daly. Stein’s method for compound geometric approximation. Journal of Applied Probability, 47(1):146–156, March 2010. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/25662506.

Dalmau:2016:DQG

[Dal16a] Joseba Dalmau. Distribution of the quasispecies for a Galton– Watson process on the sharp peak landscape. Journal of Applied Probability, 53(2):606–613, June 2016. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

projecteuclid.org/euclid.jap/1466172879.

Daly:2016:CGA

[Dal16b] Fraser Daly. Compound geometric approximation under a failure rate constraint. Journal of Applied Probability, 53(3):700– 714, September 2016. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://projecteuclid.

org/euclid.jap/1476370771.

Daniels:1969:MSM

[Dan69] H. E. Daniels. The minimum of a stationary Markov process superimposed on a U-shaped trend. Journal of Applied Probability, 6(2):399–408, August 1969. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3212009.

Daniels:1977:AWS

[Dan77] H. E. Daniels. The advancing wave in a spatial birth process. Journal of Applied Probability, 14(4):689–701, December 1977. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/3213343.

Daniels:1982:SAG

[Dan82a] H. E. Daniels. The saddlepoint approximation for a general birth process. Journal of Applied Probability, 19(1):20–28, March 1982. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3213912.

Daniels:1982:TNE

[Dan82b] H. E. Daniels. A transformation for normal extremes. Journal of Applied Probability, 19A:201–206, 1982. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/3213560.

Darroch:1966:IPT

[Dar66] John N. Darroch. Identities for passage times with applications to recurrent events and homogeneous differential functions. Journal of Applied Probability, 3(2):435–444, December 1966. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/3212129.

Darling:1970:GWP

[Dar70] D. A. Darling. The Galton–Watson process with infinite mean. Journal of Applied Probability, 7(2):455–456, August 1970. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/3211981.

Darling:1989:IDS

[Dar89] R. W. R. Darling. Infinite-dimensional stochastic difference equations for particle systems and network flows. Journal of Applied Probability, 26(2):325–344, June 1989. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/3214039.

Dassios:1996:SQA

[Das96] Angelos Dassios. Sample quantiles of additive renewal reward processes. Journal of Applied Probability, 33(4):1018–1032, December 1996. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3214983.

Davis:1964:NCF

[Dav64] A. W. Davis. A note on the characteristic functional for a replacement process of Gani. Journal of Applied Probability, 1 (1):157–160, June 1964. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3212066.

Davis:1965:TBD

[Dav65] A. W. Davis. On the theory of birth, death and diffusion processes. Journal of Applied Probability, 2(2):293–322, December 1965. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3212196.

Davies:1973:CAN

[Dav73] P. L. Davies. A class of almost nowhere differentiable stationary Gaussian processes which are somewhere differentiable. Journal of Applied Probability, 10(3):682–684, September 1973. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3212789.

Davis:1974:APC

[Dav74] Herbert T. Davis. Asymptotic properties of certain functionals of the periodogram. Journal of Applied Probability, 11(3):578–581, September 1974. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3212703.

Davies:1975:MSM

[Dav75] G. S. Davies. Maintainability of structures in Markov chain models under recruitment control. Journal of Applied Probability, 12 (2):376–382, June 1975. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3212454.

Davy:1976:PTS

[Dav76] Pamela Davy. Projected thick sections through multidimensional particle aggregates. Journal of Applied Probability, 13(4):714–722, December 1976. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3212526. See correction [Ano78g].

Davies:1978:SBP

[Dav78] P. L. Davies. The simple branching process: A note on convergence when the mean is infinite. Journal of Applied Probability, 15(3):466–480, September 1978. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.

org/stable/3213110.

Davy:1980:SL

[Dav80] Pamela J. Davy. The stereology of location. Journal of Applied Probability, 17(3):860–864, September 1980. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/3212981.

Davies:1981:MRM

[Dav81] G. S. Davies. Maintainable regions in a Markov manpower model. Journal of Applied Probability, 18(3):738–742, September 1981. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3213329.

Davies:1982:CGS

[Dav82a] G. S. Davies. Control of grade sizes in a partially stochastic Markov manpower model. Journal of Applied Probability, 19 (2):439–443, June 1982. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3213498.

Davis:1982:DHO

[Dav82b] A. W. Davis. On the distribution of Hotelling’s one-sample T^{2} under moderate non-normality. Journal of Applied Probability, 19A:207–216, 1982. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3213561.

Davies:1983:NGP

[Dav83] G. S. Davies. A note on the geometric/probabilistic relationship in a Markov manpower model. Journal of Applied Probability, 20(2):423–428, June 1983. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3213818.

Davy:1984:PML

[Dav84] Pamela J. Davy. Probability models for liberation. Journal of Applied Probability, 21(2):260–269, June 1984. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3213638.

Davies:1985:NCT

[Dav85] G. S. Davies. A note on a continuous-time Markov manpower model. Journal of Applied Probability, 22(4):932–938, December 1985. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3213961.

Davies:1986:RCS

[Dav86] P. L. Davies. Rates of convergence to the stationary distribution for k-dimensional diffusion processes. Journal of Applied Probability, 23(2):370–384, June 1986. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3214180.

Davies:2001:IPD

[Dav01] Robert B. Davies. Integrated processes and the discrete cosine transform. Journal of Applied Probability, 38A:105–121, 2001. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/3215872.

Dawson:1973:IFD

[Daw73] D. A. Dawson. Information flow in discrete Markov systems. Journal of Applied Probability, 10(1):63–83, March 1973. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/3212496.

Dawson:1974:IFS

[Daw74] D. A. Dawson. Information flow in some classes of Markov systems. Journal of Applied Probability, 11(3):594–600, September 1974. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3212706.

Dawson:1981:AET

[Daw81] Jeremy E. Dawson. The algebraic equivalence of two measures of genetic distance. Journal of Applied Probability, 18(2):520– 522, June 1981. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3213299.

Dayananda:1970:OR

[Day70] P. W. A. Dayananda. Optimal reinsurance. Journal of Applied Probability, 7(1):134–156, April 1970. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3212155.

Day:1983:CRD

[Day83] Martin V. Day. Comparison results for diffusions conditioned on positivity. Journal of Applied Probability, 20(4):766–777, December 1983. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3213588.

Deshmukh:1977:DIS

[DC77] S. D. Deshmukh and S. D. Chikte. Dynamic investment strategies for a risky R and D project. Journal of Applied Probability, 14(1):144–152, March 1977. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3213267.

DeLaCal:2006:SOM

[DC06] Jesús De La Cal and Javier Cárcamo. Stochastic orders and majorization of mean order statistics. Journal of Applied Probability, 43(3):704–712, September 2006. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/27595766.

DeLaCal:2010:ISD

[DC10] Jesús De La Cal and Javier Cárcamo. Inverse stochastic dominance, majorization, and mean order statistics. Journal of Applied Probability, 47(1):277–292, March 2010. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/25662515.

deCarvalho:2016:CPF

[dCGG16] Walter A. F. de Carvalho, Sandro Gallo, and Nancy L. Garcia. Continuity properties of a factor of Markov chains. Journal of Applied Probability, 53(1):216–230, March 2016. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://projecteuclid.org/euclid.jap/1457470570.

Dietz:1968:CBE

[DD68] K. Dietz and F. Downton. Carrier-borne epidemics with immigration. I: Immigration of both susceptibles and carriers. Journal of Applied Probability, 5(1):31–42, April 1968. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3212075.

Datta:2004:RWC

[DD04] Nilanjana Datta and Tony C. Dorlas. Random walks on a complete graph: A model for infection. Journal of Applied Probability, 41(4):1008–1021, December 2004. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/4141374.

DAndrea:2015:KKT

[DD15] Alessandro D’Andrea and Luca De Sanctis. The Kruskal– Katona theorem and a characterization of system signatures. Journal of Applied Probability, 52(2):508–518, June 2015. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://projecteuclid.org/euclid.jap/1437658612.

Diekmann:1998:DEM

[DDM98] O. Diekmann, M. C. M. De Jong, and J. A. J. Metz. A deterministic epidemic model taking account of repeated contacts between the same individuals. Journal of Applied Probability, 35(2):448–462, June 1998. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3215698.

Dirrler:2020:GMH

[DDS20] Martin Dirrler, Christopher Dörr, and Martin Schlather. A generalization of Matérn hard-core processes with applications to max-stable processes. Journal of Applied Probability, 57(4):1298–1312, December 2020. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL https:/

/www.cambridge.org/core/journals/journal-of-applied-
probability/article/generalization-of-matern-hardcore-
processes-with-applications-to-maxstable-processes/CD259FD043405FB9EE4F1B487978F85B.

DeLaPena:2022:DTL

[DDS22a] Victor De La Pena, Paul Doukhan, and Yahia Salhi. A dynamic Taylor’s law. Journal of Applied Probability, 59(2):584–607, June 2022. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-

6072 (electronic). URL https://www.cambridge.org/core/
journals/journal-of-applied-probability/article/dynamic-
taylors-law/DCBA159835EF36828CA13F83A4102FD2.

Delon:2022:GWD

[DDS22b] Julie Delon, Agnes Desolneux, and Antoine Salmona. Gromov– Wasserstein distances between Gaussian distributions. Journal of Applied Probability, 59(4):1178–1198, December 2022. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL https://www.cambridge.org/core/journals/journal-
of-applied-probability/article/gromovwasserstein-distances-
between-gaussian-distributions/DE7E314ED7ACE13FBB4A0E7983A7F84A.

DeHaan:1974:WLS

[De 74] Laurens De Haan. Weak limits of sample range. Journal of Applied Probability, 11(4):836–841, December 1974. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3212568.

DeMare:1977:RSG

[De 77] Jacques De Maré. Reconstruction of a stationary Gaussian process from its sign-changes. Journal of Applied Probability, 14(1): 38–57, March 1977. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3213259.

DeMeyer:1982:TBD

[De 82] A. De Meyer. On a theorem of Bingham and Doney. Journal of Applied Probability, 19(1):217–220, March 1982. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3213931.

DeKok:1985:AAP

[De 85a] A. G. De Kok. Asymptotic analysis for a production-inventory control model with renewal arrivals and exponential demands. Journal of Applied Probability, 22(3):653–667, September 1985. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/3213868.

DeMorais:1985:ODS

[De 85b] Paulo Renato De Morais. Optimal design of a storage system with constant or linear release rate. Journal of Applied Probability, 22(3):732–738, September 1985. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3213878.

DeSmit:1985:QGH

[De 85c] Jos H. A. De Smit. The queue GI/H_{m} /s in continuous time. Journal of Applied Probability, 22(1):214–222, March 1985. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/3213761.

DeAcosta:1994:LDE

[De 94] A. De Acosta. On large deviations of empirical measures in the τ-topology. Journal of Applied Probability, 31A:41–47, 1994. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/3214947.

DeAmbroggio:2022:EAC

[De 22] Umberto De Ambroggio. An elementary approach to component sizes in critical random graphs. Journal of Applied Probability, 59(4):1228–1242, December 2022. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL https:/

/www.cambridge.org/core/journals/journal-of-applied-
probability/article/an-elementary-approach-to-component-
sizes-in-critical-random-graphs/EF3E6E9A081945622F763481ED23CE62.

Degen:2011:SHQ

[DE11] Matthias Degen and Paul Embrechts. Scaling of high-quantile estimators. Journal of Applied Probability, 48(4):968–983, December 2011. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/23066437.

Deheuvels:1984:CDO

[Deh84] Paul Deheuvels. The characterization of distributions by order statistics and record values: A unified approach. Journal of Applied Probability, 21(2):326–334, June 1984. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/3213643. See correction [Ano85g].

Deheuvels:1985:LBM

[Deh85] Paul Deheuvels. The limiting behaviour of the maximal spacing generated by an I.I.D. sequence of Gaussian random variables. Journal of Applied Probability, 22(4):816–827, December 1985. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/3213949.

Deistler:1986:LDE

[Dei86] M. Deistler. Linear dynamic errors-in-variables models. Journal of Applied Probability, 23A:23–39, 1986. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/3214340.

Deijfen:2010:RNP

[Dei10] Maria Deijfen. Random networks with preferential growth and vertex death. Journal of Applied Probability, 47(4):1150–1163, December 2010. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
29777433.

Deken:1978:SMS

[Dek78] Joseph G. Deken. Scheduled maxima sequences. Journal of Applied Probability, 15(3):543–551, September 1978. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3213117.

Dekking:1987:SBP

[Dek87] F. M. Dekking. Subcritical branching processes in a two-state random environment, and a percolation problem on trees. Journal of Applied Probability, 24(4):798–808, December 1987. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/3214206.

Dekking:1996:IPM

[Dek96] F. M. Dekking. An inequality for pairs of martingales and its application to fractal image coding. Journal of Applied Probability, 33(4):968–973, December 1996. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3214978.

Delbrouck:1978:NAC

[Del78] L. E. N. Delbrouck. Numerical approximations for compound geometric distributions with applications in queueing theory. Journal of Applied Probability, 15(1):202–208, March 1978. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3213253.

Deligonul:1985:ASI

[Del85] Z. Seyda Deligönül. An approximate solution of the integral equation of renewal theory. Journal of Applied Probability, 22 (4):926–931, December 1985. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.

org/stable/3213960.

DelMoral:1998:UCT

[Del98] P. Del Moral. A uniform convergence theorem for the numerical solving of the nonlinear filtering problem. Journal of Applied Probability, 35(4):873–884, December 1998. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/3215523.

Deng:1985:CPP

[Den85] Y. L. Deng. On the comparison of point processes. Journal of Applied Probability, 22(2):300–313, June 1985. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3213774.

Denuit:2010:PDS

[Den10] Michel Denuit. Positive dependence of signals. Journal of Applied Probability, 47(3):893–897, September 2010. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/20778524.

Deo:1973:ILL

[Deo73] Chandrakant M. Deo. An iterated logarithm law for maxima of nonstationary Gaussian processes. Journal of Applied Probability, 10(2):402–408, June 1973. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3212356.

Deshpande:2015:RFS

[Des15] Amogh Deshpande. On the role of Föllmer–Schweizer minimal martingale measure in risk-sensitive control asset management. Journal of Applied Probability, 52(3):703–717, September 2015. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://projecteuclid.org/euclid.jap/
1445543841.

DeBicki:2010:TAL

[DESM10] Krzysztof Debicki, Abdelghafour Es-Saghouani, and Michel Mandjes. Transient asymptotics of Lévy–driven queues. Journal of Applied Probability, 47(1):109–129, March 2010. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/25662504.

Dette:1994:GEU

[Det94] Holger Dette. On a generalization of the Ehrenfest urn model. Journal of Applied Probability, 31(4):930–939, December 1994. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/3215318.

Dette:1996:GFR

[Det96] Holger Dette. On the generating functions of a random walk on the non-negative integers. Journal of Applied Probability, 33(4): 1033–1052, December 1996. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3214984.

Devitt:1979:EAP

[Dev79] John S. Devitt. An enumerative approach to a problem in probability. Journal of Applied Probability, 16(1):129–137, March 1979. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3213381.

Downham:1981:SAN

[DF81a] D. Y. Downham and S. B. Fotopoulos. Some aspects of the non-asymptotic behaviour of a two-dimensional invasion process. Journal of Applied Probability, 18(3):732–737, September 1981. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/3213328.

Downham:1981:SID

[DF81b] D. Y. Downham and S. B. Fotopoulos. Some inequalities on the distribution of ladder epochs. Journal of Applied Probability, 18 (3):770–775, September 1981. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.

org/stable/3213335.

Downham:1988:TBS

[DF88] D. Y. Downham and S. B. Fotopoulos. The transient behaviour of the simple random walk in the plane. Journal of Applied Probability, 25(1):58–69, March 1988. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3214233.

Dawson:1989:SER

[DF89] Donald A. Dawson and Klaus Fleischmann. A stochastic explosive reaction system with sampling. Journal of Applied Probability, 26(1):9–22, March 1989. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.

org/stable/3214312.

Dong:2004:OTP

[DF04] Zhao Dong and Shui Feng. Occupation time processes of super-Brownian motion with cut-off branching. Journal of Applied Probability, 41(4):984–997, December 2004. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/4141372.

Dousse:2006:PSI

[DFM^{+}06] Olivier Dousse, Massimo Franceschetti, Nicolas Macris, Ronald Meester, and Patrick Thiran. Percolation in the signal to interference ratio graph. Journal of Applied Probability, 43(2):552– 562, June 2006. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
27595750.

DiCrescenzo:2013:SCS

[DFP13] Antonio Di Crescenzo, Esther Frostig, and Franco Pellerey. Stochastic comparisons of symmetric supermodular functions of heterogeneous random vectors. Journal of Applied Probability, 50(2):464–474, June 2013. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://projecteuclid.

org/euclid.jap/1371648954.

De:2001:EDG

[DFSD01] Arkendra De, Michael Ferguson, Suzanne Sindi, and Richard Durrett. The equilibrium distribution for a generalized sankoff-Ferretti model accurately predicts chromosome size distributions in a wide variety of species. Journal of Applied Probability, 38 (2):324–334, June 2001. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3215891.

Durante:2015:SCC

[DFST15] Fabrizio Durante, Juan Fernández-Sánchez, and Wolfgang Trutschnig. On the singular components of a copula. Journal of Applied Probability, 52(4):1175–1182, December 2015. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://projecteuclid.org/euclid.jap/1450802760.

Decreusefond:2015:NSG

[DFV15] Laurent Decreusefond, Ian Flint, and Anais Vergne. A note on the simulation of the Ginibre point process. Journal of Applied Probability, 52(4):1003–1012, December 2015. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://projecteuclid.org/euclid.jap/1450802749.

David:1974:ATC

[DG74] H. A. David and J. Galambos. The asymptotic theory of concomitants of order statistics. Journal of Applied Probability, 11 (4):762–770, December 1974. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.

org/stable/3212559.

Durrett:1976:WTM

[DG76] Richard T. Durrett and S. G. Ghurye. Waiting times without memory. Journal of Applied Probability, 13(1):65–75, March 1976. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3212666.

Doney:1989:SRB

[DG89] R. A. Doney and D. R. Grey. Some remarks on Brownian motion with drift. Journal of Applied Probability, 26(3):659–663, September 1989. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3214425.

Doney:1992:SCP

[DG92] R. A. Doney and Charles M. Goldie. A simple comparison proof. Journal of Applied Probability, 29(1):244, March 1992. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3214812.

Daley:1993:RAM

[DG93a] D. J. Daley and J. Gani. A random allocation model for carrierborne epidemics. Journal of Applied Probability, 30(4):751–765, December 1993. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3214510.

Diebolt:1993:TBS

[DG93b] Jean Diebolt and Dominique Guégan. Tail behaviour of the stationary density of general non-linear autoregressive processes of order 1. Journal of Applied Probability, 30(2):315–329, June 1993. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3214841.

Dorea:1999:SSR

[DG99] Chang C. Y. Dorea and Cátia R. Gonçalves. Search schemes for random optimization algorithms that preserve the asymptotic distribution. Journal of Applied Probability, 36(3):825–836, September 1999. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3215444.

Doyen:2006:IMG

[DG06a] Laurent Doyen and Olivier Gaudoin. Imperfect maintenance in a generalized competing risks framework. Journal of Applied Probability, 43(3):825–839, September 2006. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/27595775.

Draief:2006:ERP

[DG06b] M. Draief and A. Ganesh. Efficient routeing in Poisson smallworld networks. Journal of Applied Probability, 43(3):678–686, September 2006. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
27595764.

Deijfen:2016:RT

[DG16] Maria Deijfen and Nina Gantert. Routeing on trees. Journal of Applied Probability, 53(2):475–488, June 2016. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://projecteuclid.org/euclid.jap/1466172868.

Dor:1998:OMS

[DGK98] Avner Dor, Eitan Greenshtein, and Ephraim Korach. Optimal and myopic search in a binary random vector. Journal of Applied Probability, 35(2):463–472, June 1998. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3215699.

Deelstra:2000:OIS

[DGK00] Griselda Deelstra, Martino Grasselli, and Pierre-François Koehl. Optimal investment strategies in a CIR framework. Journal of Applied Probability, 37(4):936–946, December 2000. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3215485.

Devroye:2017:MVC

[DGLW17] Luc Devroye, László Györfi, Gábor Lugosi, and Harro Walk. On the measure of Voronoi cells. Journal of Applied Probability, 54(2):394–408, June 2017. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL https:/

/www.cambridge.org/core/journals/journal-of-applied-
probability/article/on-the-measure-of-voronoi-cells/
6C211D33E80A85FA8D67D8E413AC534F.

Dube:2004:SFL

[DGM04] Parijat Dube, Fabrice Guillemin, and Ravi R. Mazumdar. Scale functions of Lévy processes and busy periods of finite-capacity M/GI/1 queues. Journal of Applied Probability, 41(4):1145– 1156, December 2004. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/4141386.

DeLaPena:2008:HTI

[DGM08] Victor De La Peña, Henryk Gzyl, and Patrick McDonald. Hitting time and inverse problems for Markov chains. Journal of Applied Probability, 45(3):640–649, September 2008. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/27595977.

Draief:2009:ERG

[DGM09] M. Draief, A. Ganesh, and L. Massoulié. Exponential random graphs as models of overlay networks. Journal of Applied Probability, 46(1):199–220, March 2009. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/25662416.

DiCrescenzo:1995:SBC

[DGNR95] A. G. Di Crescenzo, V. Giorno, A. G. Nobile, and L. M. Ricciardi. On a symmetry-based constructive approach to probability densities for two-dimensional diffusion processes. Journal of Applied Probability, 32(2):316–336, June 1995. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/3215291.

Deffner:1985:COS

[DH85] A. Deffner and E. Haeusler. A characterization of order statistic point processes that are mixed Poisson processes and mixed sample processes simultaneously. Journal of Applied Probability, 22(2):314–323, June 1985. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3213775.

Dette:1989:LDL

[DH89] Holger Dette and Norbert Henze. The limit distribution of the largest nearest-neighbour link in the unit d-cube. Journal of Applied Probability, 26(1):67–80, March 1989. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/3214317.

Dubhashi:2008:NCS

[DH08] Devdatt Dubhashi and Olle Häggström. A note on conditioning and stochastic domination for order statistics. Journal of Applied Probability, 45(2):575–579, June 2008. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/27595966.

Debicki:2015:PRS

[DHJ15] Krzysztof Debicki, Enkelejd Hashorva, and Lanpeng Ji. Parisian ruin of self-similar Gaussian risk processes. Journal of Applied Probability, 52(3):688–702, September 2015. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

projecteuclid.org/euclid.jap/1445543840.

Dolgopyat:2018:MTB

[DHKP18] D. Dolgopyat, P. Hebbar, L. Koralov, and M. Perlman. Multitype branching processes with time-dependent branching rates.

Journal of Applied Probability, 55(3):701–727, September 2018. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL https://www.cambridge.org/core/journals/
journal-of-applied-probability/article/multitype-branching-
processes-with-timedependent-branching-rates/90A94780E218D1C045AE972BB6606AD7.

DiTella:2019:SSV

[DHKR19] Paolo Di Tella, Martin Haubold, and Martin Keller-Ressel. Semi-static variance-optimal hedging in stochastic volatility models with Fourier representation. Journal of Applied Probability, 56(3):787–809, September 2019. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL https:/

/www.cambridge.org/core/journals/journal-of-applied-
probability/article/semistatic-varianceoptimal-hedging-
in-stochastic-volatility-models-with-fourier-representation/
9552BEEF3DADB7E5288C1A7924158E17.

Debicki:2020:SRP

[DHM20] Krzysztof Debicki, Enkelejd Hashorva, and Zbigniew Michna. Simultaneous ruin probability for two-dimensional Brownian risk model. Journal of Applied Probability, 57(2):597–612, June 2020. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL https://www.cambridge.org/core/journals/
journal-of-applied-probability/article/simultaneous-
ruin-probability-for-twodimensional-brownian-risk-model/
F185030CABD20B265695CD56FFCFEFEA.

Debicki:2022:CPC

[DHM22] Krzysztof Debicki, Enkelejd Hashorva, and Zbigniew Michna. On the continuity of Pickands constants. Journal of Applied Probability, 59(1):187–201, March 2022. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL https://www.cambridge.org/core/journals/journal-
of-applied-probability/article/on-the-continuity-of-
pickands-constants/283D7F9DF31C9BA4B743910DA603DFA1.

deHaan:1973:ASL

[dHR73] Laurens de Haan and Sidney I. Resnick. Almost sure limit points of record values. Journal of Applied Probability, 10(3):528–542, September 1973. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3212774.

Debicki:2015:EHG

[DHSK15] Krzysztof Debicki, Enkelejd Hashorva, and Natalia Soja-Kukiela. Extremes of homogeneous Gaussian random fields. Journal of Applied Probability, 52(1):55–67, March 2015. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://projecteuclid.org/euclid.jap/1429282606.

Daley:1986:BGW

[DHT86] D. J. Daley, David M. Hull, and James M. Taylor. Bisexual Galton–Watson branching processes with superadditive mating functions. Journal of Applied Probability, 23(3):585–600, September 1986. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3213999.

DiCrescenzo:1998:FPT

[Di 98] Antonio Di Crescenzo. First-passage-time densities and avoiding probabilities for birth-and-death processes with symmetric sample paths. Journal of Applied Probability, 35(2):383–394, June 1998. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3215692.

DiCrescenzo:1999:PAM

[Di 99] Antonio Di Crescenzo. A probabilistic analogue of the mean value theorem and its applications to reliability theory. Journal of Applied Probability, 36(3):706–719, September 1999. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3215435.

Dong:2020:WCR

[DI20] Congzao Dong and Alexander Iksanov. Weak convergence of random processes with immigration at random times. Journal of Applied Probability, 57(1):250–265, March 2020. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL https://www.cambridge.org/core/journals/journal-
of-applied-probability/article/weak-convergence-of-random-
processes-with-immigration-at-random-times/738A2AC85A949D1D0A0304D89C93799F.

Diamantis:1973:AGF

[Dia73] Basil Diamantis. An application of the generating function to the solution of a genetic problem. Journal of Applied Probability, 10(2):437–440, June 1973. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3212360.

Diaconis:1976:BPL

[Dia76] Persi Diaconis. Buffon’s problem with a long needle. Journal of Applied Probability, 13(3):614–618, September 1976. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3212484.

Dietz:1966:MWS

[Die66] Klaus Dietz. On the model of Weiss for the spread of epidemics by carriers. Journal of Applied Probability, 3(2):375–382, December 1966. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/

3212126. See note [Dow67].

Dieter:1982:APR

[Die82] U. Dieter. An alternate proof for the representation of discrete distributions by equiprobable mixtures. Journal of Applied Probability, 19(4):869–872, December 1982. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/3213842.

Dies:1987:TCM

[Die87] Jacques-Édouard Dies. Transience des chaînes de Markov linéaires sur les permutations. (French) [Transience of Markov chains on linear permutations]. Journal of Applied Probability, 24(4):899–907, December 1987. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.

org/stable/3214214.

Dieulle:1999:RSP

[Die99] Laurence Dieulle. Reliability of a system with Poisson inspection times. Journal of Applied Probability, 36(4):1140–1154, December 1999. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3215584.

Diggle:1976:SSM

[Dig76] Peter J. Diggle. A spatial stochastic model of inter-plant competition. Journal of Applied Probability, 13(4):662–671, December 1976. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3212521.

DeLaPena:2004:OB

[DIJ04] Victor H. De La Peña, Rustam Ibragimov, and Steve Jordan. Option bounds. Journal of Applied Probability, 41A:145–156, 2004. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3215974.

DAuria:2010:FPM

[DIKM10] Bernardo D’Auria, Jevgenijs Ivanovs, Offer Kella, and Michel Mandjes. First passage of a Markov additive process and generalized Jordan chains. Journal of Applied Probability, 47(4): 1048–1057, December 2010. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.

org/stable/29777425.

DiCrescenzo:2013:GTP

[DIMZ13] Antonio Di Crescenzo, Antonella Iuliano, Barbara Martinucci, and Shelemyahu Zacks. Generalized telegraph process with random jumps. Journal of Applied Probability, 50(2):450–463, June 2013. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://projecteuclid.org/euclid.jap/
1371648953.

Dion:1974:EMI

[Dio74] Jean-Pierre Dion. Estimation of the mean and the initial probabilities of a branching process. Journal of Applied Probability, 11(4):687–694, December 1974. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.

org/stable/3212552.

Dion:1976:CVF

[Dio76] Jean-Guy Dion. Sur une caractérisation des vidanges fortes. (French) [On a characterization of strong drains]. Advances in Applied Probability, 8(1):105–126, March 1976. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic).

URL http://www.jstor.org/stable/1426024. See corrections

[Dio78].

Dion:1978:CVF

[Dio78] Jean-Guy Dion. Sur une caractérisation des vidanges fortes: Corrections and extensions. (French) [On a characterization of strong drains: corrections and extensions]. Journal of Applied Probability, 15(2):268–279, June 1978. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3213400. See [Dio76].

Demirci:2023:MTB

[DIÖ23] Yunus Emre Demirci, Ümit Islak, and Alperen Özdemir. Mixing time bounds for edge flipping on regular graphs. Journal of Applied Probability, 60(4):1317–1332, December 2023. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL https://www.cambridge.org/core/journals/journal-
of-applied-probability/article/mixing-time-bounds-for-
edge-flipping-on-regular-graphs/8F06FFBA5440514574E326C9A6994F0B.

Dippon:1998:GCS

[Dip98] Jürgen Dippon. Globally convergent stochastic optimization with optimal asymptotic distribution. Journal of Applied Probability, 35(2):395–406, June 1998. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3215693.

Dirkse:1975:APO

[Dir75] John P. Dirkse. An absorption probability for the Ornstein– Uhlenbeck process. Journal of Applied Probability, 12(3):595– 599, September 1975. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3212876.

Dixon:1993:EPT

[Dix93] M. T. Dixon. Equilibrium points for three games based on the Poisson process. Journal of Applied Probability, 30(3):627–638, September 1993. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3214771.

Daley:1969:TWTb

[DJ69] D. J. Daley and D. R. Jacobs, Jr. The total waiting time in a busy period of a stable single-server queue, II. Journal of Applied Probability, 6(3):565–572, December 1969. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/3212102.

Dhar:1995:PBF

[DJ95] Sunil K. Dhar and Xulun Jiang. Probability bounds on the finite sum of the binary sequence of order k. Journal of Applied Probability, 32(4):1014–1027, December 1995. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/3215215.

Delgado:2000:WAC

[DJ00] Rosario Delgado and Maria Jolis. Weak approximation for a class of Gaussian processes. Journal of Applied Probability, 37 (2):400–407, June 2000. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3215716.

Dassios:2005:KBF

[DJ05] Angelos Dassios and Ji-Wook Jang. Kalman–Bucy filtering for linear systems driven by the Cox process with shot noise intensity and its application to the pricing of reinsurance contracts. Journal of Applied Probability, 42(1):93–107, March 2005. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/30040772.

Delmas:2009:DWR

[DJ09] Jean-François Delmas and Benjamin Jourdain. Does waste recycling really improve the multi-proposal Metropolis–Hastings algorithm? An analysis based on control variates. Journal of Applied Probability, 46(4):938–959, December 2009. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/25662473.

Daly:2018:RMC

[DJ18] Fraser Daly and Oliver Johnson. Relaxation of monotone coupling conditions: Poisson approximation and beyond. Journal of Applied Probability, 55(3):742–759, September 2018. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL https://www.cambridge.org/core/journals/journal-
of-applied-probability/article/relaxation-of-monotone-
coupling-conditions-poisson-approximation-and-beyond/
3E235F5222DE6D31D532570876977ABA.

Dimitrov:1990:NCE

[DK90] B. Dimitrov and Z. Khalil. On a new characterization of the exponential distribution related to a queueing system with an unreliable server. Journal of Applied Probability, 27(1):221– 226, March 1990. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3214610.

Durrett:1999:NSM

[DK99] Richard Durrett and Semyon Kruglyak. A new stochastic model of microsatellite evolution. Journal of Applied Probability, 36(3): 621–631, September 1999. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3215428.

Dabye:2001:MCP

[DK01] Ali S. Dabye and Yury A. Kutoyants. Misspecified change-point estimation problem for a Poisson process. Journal of Applied Probability, 38A:122–130, 2001. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.

org/stable/3215873.

Dekking:2011:MMB

[DK11] Michel Dekking and Derong Kong. Multimodality of the Markov binomial distribution. Journal of Applied Probability, 48(4): 938–953, December 2011. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/23066435.

Dudin:2008:MCH

[DKK08] Alexander Dudin, Chesoong Kim, and Valentina Klimenok. Markov chains with hybrid repeating rows: Upper-Hessenberg, quasi-Toeplitz structure of the block transition probability matrix. Journal of Applied Probability, 45(1):211–225, March 2008. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/27595935.

Donnelly:1993:CVB

[DKM93] Peter Donnelly, Thomas Kurtz, and Paul Marjoram. Correlation and variability in birth processes. Journal of Applied Probability, 30(2):275–284, June 1993. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3214838.

Dvurecenskij:1984:PBP

[DKO84] A. Dvurecenskij, L. A. Kuljukina, and G. A. Ososkov. On a problem of the busy-period determination in queues with infinitely many servers. Journal of Applied Probability, 21(1):201– 206, March 1984. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3213680.

Dua:1976:MRW

[DKS76] Saroj Dua, Shobha Khadilkar, and Kanwar Sen. A modified random walk in the presence of partially reflecting barriers. Journal of Applied Probability, 13(1):169–175, March 1976. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3212680.

Deshpande:1986:APA

[DKS86] Jayant V. Deshpande, Subhash C. Kochar, and Harshinder Singh. Aspects of positive ageing. Journal of Applied Probability, 23(3):748–758, September 1986. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3214012.

Dehay:1996:FLT

[DL96] Dominique Dehay and Jacek Leśkow. Functional limit theory for the spectral covariance estimator. Journal of Applied Probability, 33(4):1077–1092, December 1996. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3214987.

DiCrescenzo:2002:EBM

[DL02] Antonio Di Crescenzo and Maria Longobardi. Entropy-based measure of uncertainty in past lifetime distributions. Journal of Applied Probability, 39(2):434–440, June 2002. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3216111.

Dolgoarshinnykh:2006:CSS

[DL06] R. G. Dolgoarshinnykh and Steven P. Lalley. Critical scaling for the SIS stochastic epidemic. Journal of Applied Probability, 43 (3):892–898, September 2006. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.

org/stable/27595782.

Dieker:2017:DSB

[DL17] A. B. Dieker and Guido R. Lagos. A dichotomy for sampling barrier-crossing events of random walks with regularly varying tails. Journal of Applied Probability, 54(4): 1213–1232, December 2017. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL https:/

/www.cambridge.org/core/journals/journal-of-applied-

probability/article/dichotomy-for-sampling-barriercrossing-
events-of-random-walks-with-regularly-varying-tails/
48A7188C2207E44D16E2FBC2E8506A73.

Dassios:2018:RFD

[DL18] Angelos Dassios and Jia Wei Lim. Recursive formula for the double-barrier Parisian stopping time. Journal of Applied Probability, 55(1):282–301, March 2018. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL https:/

/www.cambridge.org/core/journals/journal-of-applied-
probability/article/recursive-formula-for-the-doublebarrier-
parisian-stopping-time/087967DA1806B848585B115C2CF654EA.

deLoynes:2023:LLT

[dL23] Basile de Loynes. Local limit theorem for a Markov additive process on Z^{d} with a null recurrent internal Markov chain. Journal of Applied Probability, 60(2):570–588, June 2023. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL https://www.cambridge.org/core/journals/journal-
of-applied-probability/article/local-limit-theorem-for-
a-markov-additive-process-on-mathbbzd-with-a-null-recurrent-
internal-markov-chain/610FE3130A5E9FC863013D50A36DB185.

DaiPra:2014:SIR

[DLM14] Paolo Dai Pra, Pierre-Yves Louis, and Ida G. Minelli. Synchronization via interacting reinforcement. Journal of Applied Probability, 51(2):556–568, June 2014. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http:/

/projecteuclid.org/euclid.jap/1402578643.

Durmus:2017:OSR

[DLMR17] Alain Durmus, Sylvain Le Corff, Eric Moulines, and Gareth O. Roberts. Optimal scaling of the random walk Metropolis algorithm under L^{p} mean differentiability. Journal of Applied Probability, 54(4):1233–1260, December 2017. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL https:/

/www.cambridge.org/core/journals/journal-of-applied-
probability/article/optimal-scaling-of-the-random-walk-
metropolis-algorithm-under-l-p-mean-differentiability/
53894E10E1CE92715CFA6C29FD503747.

Derman:1980:OAS

[DLR80] C. Derman, G. J. Lieberman, and S. M. Ross. On the optimal assignment of servers and a repairman. Journal of Applied Probability, 17(2):577–581, June 1980. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3213050.

Deligiannidis:2009:OSP

[DLU09] G. Deligiannidis, H. Le, and S. Utev. Optimal stopping for processes with independent increments, and applications. Journal of Applied Probability, 46(4):1130–1145, December 2009. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/25662485.

Dyakonova:2020:BPR

[DLVZ20] Elena Dyakonova, Doudou Li, Vladimir Vatutin, and Mei Zhang. Branching processes in a random environment with immigration stopped at zero. Journal of Applied Probability, 57(1):237–249, March 2020. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL https:/

/www.cambridge.org/core/journals/journal-of-applied-
probability/article/branching-processes-in-a-random-
environment-with-immigration-stopped-at-zero/1A37441E30969BCB97A89E661238182E.

Darroch:1967:SPT

[DM67] J. N. Darroch and K. W. Morris. Some passage-time generating functions for discrete-time and continuous-time finite Markov chains. Journal of Applied Probability, 4(3):496–507, December 1967. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3212217.

Darroch:1968:PTG

[DM68] J. N. Darroch and K. W. Morris. Passage-time generating functions for continuous-time finite Markov chains. Journal of Applied Probability, 5(2):414–426, August 1968. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3212262.

Donelson:1972:CLV

[DM72] John Donelson III and Fred Maltz. A comparison of linear versus non-linear prediction for polynomial functions of the Ornstein– Uhlenbeck process. Journal of Applied Probability, 9(4):725–744, December 1972. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3212611.

Daley:1975:OIP

[DM75] D. J. Daley and R. K. Milne. Orderliness, intensities and Palm–Khinchin equations for multivariate point processes. Journal of Applied Probability, 12(2):383–389, June 1975. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3212455.

Davis:2001:PPC

[DM01] Richard A. Davis and Thomas Mikosch. Point process convergence of stochastic volatility processes with application to sample autocorrelation. Journal of Applied Probability, 38A:93–104, 2001. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3215871.

Debicki:2003:EOA

[DM03] Krzysztof Debicki and Michel Mandjes. Exact overflow asymptotics for queues with many Gaussian inputs. Journal of Applied Probability, 40(3):704–720, September 2003. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/3215945.

Duffy:2005:HER

[DM05a] Ken Duffy and Anthony P. Metcalfe. How to estimate the rate function of a cumulative process. Journal of Applied Probability, 42(4):1044–1052, December 2005. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/30040827.

Duffy:2005:LDE

[DM05b] Ken Duffy and Anthony P. Metcalfe. The large deviations of estimating rate functions. Journal of Applied Probability, 42 (1):267–274, March 2005. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/30040785.

Denuit:2010:GIC

[DM10a] Michel M. Denuit and Mhamed Mesfioui. Generalized increasing convex and directionally convex orders. Journal of Applied Probability, 47(1):264–276, March 2010. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/25662514.

DiCrescenzo:2010:DTR

[DM10b] Antonio Di Crescenzo and Barbara Martinucci. A damped telegraph random process with logistic stationary distribution. Journal of Applied Probability, 47(1):84–96, March 2010. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/25662502.

DiCrescenzo:2018:JTP

[DM18] Antonio Di Crescenzo and Alessandra Meoli. On a jumptelegraph process driven by an alternating fractional Poisson process. Journal of Applied Probability, 55(1):94–111, March 2018. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-

6072 (electronic). URL https://www.cambridge.org/core/
journals/journal-of-applied-probability/article/on-a-
jumptelegraph-process-driven-by-an-alternating-fractional-
poisson-process/D7166DDFD95CAE4502F5BF04816EB661.

Daley:2019:MVB

[DM19] Daryl J. Daley and Masakiyo Miyazawa. A martingale view of Blackwell’s renewal theorem and its extensions to a general counting process. Journal of Applied Probability, 56(2):602–623, June 2019. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL https:/

/www.cambridge.org/core/journals/journal-of-applied-
probability/article/martingale-view-of-blackwells-renewal-
theorem-and-its-extensions-to-a-general-counting-process/
D29B76979631C93881048926B4056E89.

Delsing:2021:TCL

[DM21] Guusje Delsing and Michel Mandjes. A transient Cramér– Lundberg model with applications to credit risk. Journal of Applied Probability, 58(3):721–745, September 2021. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL https://www.cambridge.org/core/journals/journal-
of-applied-probability/article/transient-cramerlundberg-
model-with-applications-to-credit-risk/284130A946A935AA053326E8933350E1.

Draief:2005:QST

[DMO05] Moez Draief, Jean Mairesse, and Neil O’Connell. Queues, stores, and tableaux. Journal of Applied Probability, 42(4):1145–1167, December 2005. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
30040835.

Dunsmuir:2002:ORL

[DMR02] William T. M. Dunsmuir, Sean P. Meyn, and Gareth Roberts. Obituary: Richard Lewis Tweedie. Journal of Applied Probability, 39(2):441–454, June 2002. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3216112.

Daduna:2008:WCL

[DMS08] Hans Daduna, Christian Malchin, and Ryszard Szekli. Weak convergence limits for sojourn times in cyclic queues under heavy traffic conditions. Journal of Applied Probability, 45(2):333– 346, June 2008. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
27595948.

Duffy:2011:SPLa

[DMT11] Ken R. Duffy, Claudio Macci, and Giovanni Luca Torrisi. Sample path large deviations for order statistics. Journal of Applied Probability, 48(1):238–257, March 2011. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http:/

/www.jstor.org/stable/29777457.

DiCrescenzo:2015:CPP

[DMZ15] Antonio Di Crescenzo, Barbara Martinucci, and Shelemyahu Zacks. Compound Poisson process with a Poisson subordinator. Journal of Applied Probability, 52(2):360–374, June 2015. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://projecteuclid.org/euclid.jap/1437658603.

Dalen:1980:TWT

[DN80a] G. Dalen and B. Natvig. On the transient waiting times for a GI/M/1 priority queue. Journal of Applied Probability, 17 (1):227–234, March 1980. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3212939.

Daley:1980:SEP

[DN80b] D. J. Daley and Prakash Narayan. Series expansions of probability generating functions and bounds for the extinction probability of a branching process. Journal of Applied Probability, 17(4): 939–947, December 1980. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3213203.

Dudin:1999:BSQ

[DN99] Alexander Dudin and Shoichi Nishimura. A BMAP/SM/1 queueing system with Markovian arrival input of disasters. Journal of Applied Probability, 36(3):868–881, September 1999. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/3215447.

Davydov:2004:RPE

[DN04] Yu. Davydov and A. V. Nagaev. On the role played by extreme summands when a sum of independent and identically distributed random vectors is asymptotically α-stable. Journal of Applied Probability, 41(2):437–454, June 2004. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3216027.

Du:2013:NAE

[DN13] Kai Du and Ariel David Neufeld. A note on asymptotic exponential arbitrage with exponentially decaying failure probability. Journal of Applied Probability, 50(3):801–809, September 2013. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://projecteuclid.org/euclid.jap/
1378401237.

Doukhan:2019:ARS

[DN19] Paul Doukhan and Michael H. Neumann. Absolute regularity of semi-contractive GARCH-type processes. Journal of Applied Probability, 56(1):91–115, March 2019. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL https:/

/www.cambridge.org/core/journals/journal-of-applied-
probability/article/absolute-regularity-of-semicontractive-
garchtype-processes/633871BEE02EEEA0CFAFD1EB637B1AFA.

Persone:1996:SFQ

[dNPG96] Vittoria de Nitto Personè and Vincenzo Grassi. Solution of finite QBD processes. Journal of Applied Probability, 33(4):1003–1010, December 1996. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3214981.

Dupuis:2007:LDP

[DNW07] Paul Dupuis, Carl Nuzman, and Phil Whiting. Large deviations principle for occupancy problems with colored balls. Journal of Applied Probability, 44(1):115–141, March 2007. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/27595826.

Du:2016:CPE

[DNY16] Nguyen Huu Du, Dang Hai Nguyen, and G. George Yin. Conditions for permanence and ergodicity of certain stochastic predator-prey models. Journal of Applied Probability, 53(1): 187–202, March 2016. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://projecteuclid.

org/euclid.jap/1457470568.

Dvurecenskij:1985:MCP

[DO85] A. Dvurecenskij and G. A. Ososkov. On a modified counter with prolonging dead time. Journal of Applied Probability, 22 (3):678–687, September 1985. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.

org/stable/3213870.

DiMatteo:1997:DPA

[DO97] Ilaria Di Matteo and Enzo Orsingher. Detailed probabilistic analysis of the integrated three-valued telegraph signal. Journal of Applied Probability, 34(3):671–684, September 1997. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3215093.

Dayar:2012:KBI

[DO12] Tugrul Dayar and M. Can Orhan. Kronecker-based infinite leveldependent QBD processes. Journal of Applied Probability, 49(4): 1166–1187, December 2012. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/41713836.

Dayar:2016:SSA

[DO16] Tuğrul Dayar and M. Can Orhan. Steady-state analysis of a multiclass MAP/PH/c queue with acyclic PH retrials. Journal of Applied Probability, 53(4):1098–1110, December 2016. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://projecteuclid.org/euclid.jap/1481132839.

Dobson:1970:IEL

[Dob70] A. J. Dobson. Interval estimates of language divergence times. Journal of Applied Probability, 7(3):781–784, December 1970.

CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/3211957.

Dobson:1974:UTN

[Dob74] Annette J. Dobson. Unrooted trees for numerical taxonomy. Journal of Applied Probability, 11(1):32–42, March 1974. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/3212580.

Doberkat:1977:OSM

[Dob77] Ernst-Erich Doberkat. Optimal strategies in measurable learning systems on metric spaces. Journal of Applied Probability, 14(4): 795–805, December 1977. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3213352.

Dobson:1978:SML

[Dob78] Annette J. Dobson. A stochastic model of lexical change. Journal of Applied Probability, 15(1):38–45, March 1978. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3213234.

Dobrow:1996:DDR

[Dob96] Robert P. Dobrow. On the distribution of distances in recursive trees. Journal of Applied Probability, 33(3):749–757, September 1996. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3215356.

Dohnal:1987:EDC

[Doh87] Gejza Dohnal. On estimating the diffusion coefficient. Journal of Applied Probability, 24(1):105–114, March 1987. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3214063.

Dolinsky:2010:SRA

[Dol10] Yan Dolinsky. Shortfall risk approximations for American options in the multidimensional Black–Scholes model. Journal of Applied Probability, 47(4):997–1012, December 2010. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/29777421.

Domine:1995:MFP

[Dom95] Marco Dominé. Moments of the first-passage time of a Wiener process with drift between two elastic barriers. Journal of Applied Probability, 32(4):1007–1013, December 1995. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3215214.

Domine:1996:FPT

[Dom96] Marco Dominé. First passage time distribution of a Wiener process with drift concerning two elastic barriers. Journal of Applied Probability, 33(1):164–175, March 1996. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http:/

/www.jstor.org/stable/3215274.

Doney:1971:PBP

[Don71] R. A. Doney. The progeny of a branching process. Journal of Applied Probability, 8(2):407–412, June 1971. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/3211911.

Doney:1972:LTC

[Don72] R. A. Doney. A limit theorem for a class of supercritical branching processes. Journal of Applied Probability, 9(4):707–724, December 1972. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3212610.

Doney:1973:FEG

[Don73] R. A. Doney. On a functional equation for general branching processes. Journal of Applied Probability, 10(1):198–205, March 1973. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3212507.

Doney:1976:NSG

[Don76a] R. A. Doney. A note on the subcritical generalized agedependent branching process. Journal of Applied Probability, 13(4):798–803, December 1976. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.

org/stable/3212536.

Doney:1976:SMT

[Don76b] R. A. Doney. On single- and multi-type general age-dependent branching processes. Journal of Applied Probability, 13(2):239– 246, June 1976. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3212827.

Doney:1977:NCS

[Don77] R. A. Doney. A note on a condition satisfied by certain random walks. Journal of Applied Probability, 14(4):843–849, December 1977. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3213357.

Doney:1980:MLH

[Don80a] R. A. Doney. Moments of ladder heights in random walks. Journal of Applied Probability, 17(1):248–252, March 1980. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3212942.

Doney:1980:SCA

[Don80b] R. A. Doney. Spitzer’s condition for asymptotically symmetric random walk. Journal of Applied Probability, 17(3):856–859, September 1980. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3212980.

Doney:1983:NCR

[Don83] R. A. Doney. A note on conditioned random walk. Journal of Applied Probability, 20(2):409–412, June 1983. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/3213815.

Doney:1984:NSR

[Don84a] R. A. Doney. A note on some results of Schuh. Journal of Applied Probability, 21(1):192–196, March 1984. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http:/

/www.jstor.org/stable/3213678. See [Sch82].

Doney:1984:CPT

[Don84b] R. A. Doney. On conditional passage-time structure of birthdeath processes. Journal of Applied Probability, 21(3):673–674, September 1984. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3213630.

Donnelly:1986:GAV

[Don86] Peter Donnelly. A genealogical approach to variable-population-size models in population genetics. Journal of Applied Probability, 23(2):283–296, June 1986. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3214173.

Donnelly:1991:HPL

[Don91] Peter Donnelly. The heaps process, libraries, and size-biased permutations. Journal of Applied Probability, 28(2):321–335, June 1991. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3214869.

Doney:1993:SRI

[Don93] R. A. Doney. Some results involving the maximum of Brownian motion. Journal of Applied Probability, 30(4):805–818, December 1993. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3214514.

Doob:1972:OPL

[Doo72] J. L. Doob. Obituary: Paul Lévy. Journal of Applied Probability, 9(4):871–872, December 1972. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.

org/stable/3212626.

Dorea:1982:CRG

[Dor82] Chang C. Y. Dorea. Connectivity of random graphs. Journal of Applied Probability, 19(4):880–884, December 1982. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3213844.

Doshi:1979:GSM

[Dos79] Bharat T. Doshi. Generalized semi-Markov decision processes. Journal of Applied Probability, 16(3):618–630, September 1979. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/3213089.

Doshi:1981:ACP

[Dos81] Bharat T. Doshi. Adaptive control of a production-inventory system. Journal of Applied Probability, 18(1):204–215, March 1981. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3213180.

Doshi:1985:NSD

[Dos85] Bharat T. Doshi. A note on stochastic decomposition in a GI/G/1 queue with vacations or set-up times. Journal of Applied Probability, 22(2):419–428, June 1985. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/3213784.

Dowson:1966:SRW

[Dow66] D. C. Dowson. Simple random walks on semi-groups of nilpotency class two. Journal of Applied Probability, 3(1):156– 170, June 1966. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/

3212044. See correction [Kei64b].

Downton:1967:ECN

[Dow67] F. Downton. Epidemics with carriers: A note on a paper of Dietz. Journal of Applied Probability, 4(2):264–270, August 1967. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/3212021. See

[Die66].

Downton:1972:AUI

[Dow72a] F. Downton. The area under the infectives trajectory of the general stochastic epidemic. Journal of Applied Probability, 9 (2):414–417, June 1972. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/ stable/3212809. See correction [Dow72b].

Downton:1972:CBA

[Dow72b] F. Downton. A correction to “The Area under the Infectives Trajectory of the General Stochastic Epidemic”. Journal of Applied Probability, 9(4):873–876, December 1972. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/3212627. See [Dow72a].

Down:1998:SPM

[Dow98] D. Down. On the stability of polling models with multiple servers. Journal of Applied Probability, 35(4):925–935, December 1998. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3215529.

Darlington:2000:QCT

[DP00] S. J. Darlington and P. K. Pollett. Quasistationarity in continuous-time Markov chains where absorption is not certain. Journal of Applied Probability, 37(2):598–600, June 2000. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/3215732.

Dufour:2010:MSP

[DP10] F. Dufour and A. B. Piunovskiy. Multiobjective stopping problem for discrete-time Markov processes: Convex analytic approach. Journal of Applied Probability, 47(4):947–966, December 2010. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/29777418.

DeSantis:2012:BCT

[DP12a] Emilio De Santis and Mauro Piccioni. Backward coalescence times for perfect simulation of chains with infinite memory. Journal of Applied Probability, 49(2):319–337, June 2012. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/41713770.

Derrien:2012:SPC

[DP12b] Jean-Marc Derrien and Frédérique Plantevin. A symmetry property for a class of random walks in stationary random environments on Z. Journal of Applied Probability, 49(2):338–350, June 2012. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/41713771.

Dereudre:2018:CTC

[DP18] David Dereudre and Mathew Penrose. On the critical threshold for continuum AB percolation. Journal of Applied Probability, 55(4):1228–1237, December 2018. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL https:/

/www.cambridge.org/core/journals/journal-of-applied-
probability/article/on-the-critical-threshold-for-continuum-
ab-percolation/71228F7C59B77A291B5CCEA0FE1F8AC1.

Dai:2019:MCF

[DPR19] Hongsheng Dai, Murray Pollock, and Gareth Roberts. Monte Carlo fusion. Journal of Applied Probability, 56(1):174–191, March 2019. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-

6072 (electronic). URL https://www.cambridge.org/core/

journals/journal-of-applied-probability/article/monte-
carlo-fusion/2C7099791C6073E37F98A7FFD91C331E.

Denisov:2021:TAA

[DPW21] Denis Denisov, Elena Perfilev, and Vitali Wachtel. Tail asymptotics for the area under the excursion of a random walk with heavy-tailed increments. Journal of Applied Probability, 58(1):217–237, March 2021. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL https://www.cambridge.org/core/journals/journal-
of-applied-probability/article/tail-asymptotics-for-
the-area-under-the-excursion-of-a-random-walk-with-heavytailed-
increments/9E66D019DF7C21D72E15D394B40D053A.

Dassios:2019:ESG

[DQL19] Angelos Dassios, Yan Qu, and Jia Wei Lim. Exact simulation of generalised Vervaat perpetuities. Journal of Applied Probability, 56(1):57–75, March 2019. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL https:/

/www.cambridge.org/core/journals/journal-of-applied-
probability/article/exact-simulation-of-generalised-
vervaat-perpetuities/9D6ABDEC436968F3129D00BE9F0BE110.

Dunstan:1981:NAD

[DR81] F. D. J. Dunstan and J. F. Reynolds. Normal approximations for distributions arising in the stochastic approach to chemical reaction kinetics. Journal of Applied Probability, 18(1):263– 267, March 1981. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3213187.

Daley:1984:SCR

[DR84] D. J. Daley and T. Rolski. Some comparability results for waiting times in single- and many-server queues. Journal of Applied Probability, 21(4):887–900, December 1984. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/3213704.

Davis:1994:CMS

[DR94a] Richard A. Davis and Sidney I. Resnick. Crossings of max-stable processes. Journal of Applied Probability, 31(1):130–138, March 1994. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3215240.

DeHaan:1994:EHR

[DR94b] L. De Haan and Sidney Resnick. Estimating the home range. Journal of Applied Probability, 31(3):700–720, September 1994. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/3215149.

Donnelly:2000:CSM

[DR00] Peter Donnelly and Eliane R. Rodrigues. Convergence to stationarity in the Moran model. Journal of Applied Probability, 37 (3):705–717, September 2000. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.

org/stable/3215607.

Diasparra:2009:BRP

[DR09] Maikol A. Diasparra and Rosario Romera. Bounds for the ruin probability of a discrete-time risk process. Journal of Applied Probability, 46(1):99–112, March 2009. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/25662410.

Duembgen:2014:IS

[DR14] Moritz Duembgen and L. C. G. Rogers. Investing and stopping. Journal of Applied Probability, 51(4):898–909, December 2014. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://projecteuclid.org/euclid.jap/
1421763316.

Debrabant:2015:AML

[DR15] Kristian Debrabant and Andreas Rössler. On the acceleration of the multi-level Monte Carlo method. Journal of Applied Probability, 52(2):307–322, June 2015. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

projecteuclid.org/euclid.jap/1437658600.

deChaumaray:2018:WLS

[dRdC18] M. du Roy de Chaumaray. Weighted least-squares estimation for the subcritical Heston process. Journal of Applied Probability, 55(2):543–558, June 2018. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL https:/

/www.cambridge.org/core/journals/journal-of-applied-
probability/article/weighted-leastsquares-estimation-
for-the-subcritical-heston-process/96999FC9BEFA1088C827DD88E015A473.

Driscoll:1975:SNP

[Dri75] Michael F. Driscoll. The signal-noise problem: A solution for the case that signal and noise are Gaussian and independent. Journal of Applied Probability, 12(1):183–187, March 1975. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/3212427.

Darroch:1965:QSD

[DS65] J. N. Darroch and E. Seneta. On quasi-stationary distributions in absorbing discrete-time finite Markov chains. Journal of Applied Probability, 2(1):88–100, June 1965. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/3211876.

Darroch:1967:QSD

[DS67] J. N. Darroch and E. Seneta. On quasi-stationary distributions in absorbing continuous-time finite Markov chains. Journal of Applied Probability, 4(1):192–196, April 1967. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3212311.

Dolby:1975:IDP

[DS75] James L. Dolby and Herbert Solomon. Information density phenomena and random packing. Journal of Applied Probability, 12 (2):364–370, June 1975. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3212452.

Doshi:1980:SCM

[DS80] Bharat Doshi and Steven E. Shreve. Strong consistency of a modified maximum likelihood estimator for controlled Markov chains. Journal of Applied Probability, 17(3):726–734, September 1980. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3212966.

Deheuvels:1986:ECR

[DS86a] Paul Deheuvels and Josef Steinebach. Exact convergence rate of an Erdős–Rényi strong law for moving quantiles. Journal of Applied Probability, 23(2):355–369, June 1986. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/3214179.

Do:1986:SSS

[DS86b] Kim-Anh Do and Herbert Solomon. A simulation study of Sylvester’s problem in three dimensions. Journal of Applied Probability, 23(2):509–513, June 1986. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3214192.

Daley:1992:EMC

[DS92] D. J. Daley and L. D. Servi. Exploiting Markov chains to infer queue length from transactional data. Journal of Applied Probability, 29(3):713–732, September 1992. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/3214907.

Duma:1993:HPR

[DS93] Andrei Duma and Marius Stoka. Hitting probabilities for random ellipses and ellipsoids. Journal of Applied Probability, 30(4): 971–974, December 1993. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3214526.

Daley:1994:ALE

[DS94a] D. J. Daley and L. D. Servi. Approximating last-exit probabilities of a random walk, with application to conditional queue length moments within busy periods of M/GI/1 queues. Journal of Applied Probability, 31A:251–267, 1994. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/3214960.

Diaconis:1994:ERM

[DS94b] Persi Diaconis and Mehrdad Shahshahani. On the eigenvalues of random matrices. Journal of Applied Probability, 31A:49–62, 1994. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3214948.

Daley:1998:IBP

[DS98] D. J. Daley and L. D. Servi. Idle and busy periods in stable M/M/k queues. Journal of Applied Probability, 35(4):950–962, December 1998. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3215531.

Daduna:2002:CJO

[DS02] Hans Daduna and Ryszard Szekli. Conditional job-observer property for multitype closed queueing networks. Journal of Applied Probability, 39(4):865–881, December 2002. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3216009.

Dudin:2004:SAS

[DS04] Alexander Dudin and Olga Semenova. A stable algorithm for stationary distribution calculation for a BMAP/SM/1 queueing system with Markovian arrival input of disasters. Journal of Applied Probability, 41(2):547–556, June 2004. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/3216035.

Daduna:2006:DOM

[DS06] Hans Daduna and Ryszard Szekli. Dependence ordering for Markov processes on partially ordered spaces. Journal of Applied Probability, 43(3):793–814, September 2006. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/27595773.

Daduna:2008:IRC

[DS08a] Hans Daduna and Ryszard Szekli. Impact of routeing on correlation strength in stationary queueing network processes. Journal of Applied Probability, 45(3):846–878, September 2008. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/27595990.

Duffy:2008:LDP

[DS08b] Ken R. Duffy and Artem Sapozhnikov. The large deviation principle for the on-off Weibull sojourn process. Journal of Applied Probability, 45(1):107–117, March 2008. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http:/

/www.jstor.org/stable/27595927.

Darroch:2012:OMJ

[DS12] John Darroch and Eugene Seneta. OBITUARY: Miloslav Jiřina. Journal of Applied Probability, 49(2):595–599, June 2012. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/41713789.

Denisov:2013:AFP

[DS13] Denis Denisov and Vsevolod Shneer. Asymptotics for the first passage times of Lévy processes and random walks. Journal of Applied Probability, 50(1):64–84, March 2013. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://projecteuclid.org/euclid.jap/1363784425.

Dong:2017:CDT

[DS17] Yinghua Dong and Gennady Samorodnitsky. Contact distribution in a thinned Boolean model with power-law radii. Journal of Applied Probability, 54(3):670–684, September 2017. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL https://www.cambridge.org/core/journals/journal-
of-applied-probability/article/contact-distribution-
in-a-thinned-boolean-model-with-powerlaw-radii/64E3C7584E832625BD9B3966135C657E.

Dalmau:2021:SFP

[DS21] Joseba Dalmau and Michele Salvi. Scale-free percolation in continuous space: quenched degree and clustering coefficient. Journal of Applied Probability, 58(1):106–127, March 2021. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL https://www.cambridge.org/core/journals/journal-
of-applied-probability/article/scalefree-percolation-
in-continuous-space-quenched-degree-and-clustering-coefficient/
FC8F980E5C0115AA4F8B5A75FAFB82D2.

Dorr:2023:CTP

[DS23] Christopher Dörr and Martin Schlather. Characterization theorems for pseudo cross-variograms. Journal of Applied Probability, 60(4):1219–1231, December 2023. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL https:/

/www.cambridge.org/core/journals/journal-of-applied-
probability/article/characterization-theorems-for-pseudo-
crossvariograms/94776E48865034E67DDEA9D8273C6951.

Dshalalow:1989:MQS

[Dsh89] Jewgeni Dshalalow. Multichannel queueing systems with infinite waiting room and stochastic control. Journal of Applied Probability, 26(2):345–362, June 1989. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3214040.

Dshalalow:1994:TTP

[Dsh94] Jewgeni H. Dshalalow. On termination time processes. Journal of Applied Probability, 31A:325–336, 1994. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/3214965.

Dshalalow:1998:QPB

[Dsh98] Jewgeni H. Dshalalow. Queueing processes in bulk systems under the D-policy. Journal of Applied Probability, 35(4):976–989, December 1998. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3215533.

Dshalalow:2001:NPB

[Dsh01a] Jewgeni H. Dshalalow. A note on D-policy bulk queueing systems. Journal of Applied Probability, 38(1):280–283, March 2001. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3215762.

Dshalalow:2001:TDA

[Dsh01b] Jewgeni H. Dshalalow. Time dependent analysis of multivariate marked renewal processes. Journal of Applied Probability, 38(3): 707–721, September 2001. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3216123.

DeVries:1980:NDP

[DSR80] S. A. De Vries, A. J. Stam, and J. Reddingius. Note on a death process. Journal of Applied Probability, 17(3):853–855, September 1980. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3212979.

deSantis:2004:CPM

[dSS04] Emilio de Santis and Fabio Spizzichino. Change-point models and conditionally pure birth processes: An inequality on the stochastic intensity. Journal of Applied Probability, 41(4): 939–952, December 2004. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/4141368.

DeLima:2011:PWL

[DSS11] B. N. B. De Lima, R. Sanchis, and R. W. C. Silva. Percolation of words on Z^{d} with long-range connections. Journal of Applied Probability, 48(4):1152–1162, December 2011. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/23066448.

Debicki:2013:AHF

[DSZ13] Krzysztof Deabicki, Iwona Sierpińska, and Bert Zwart. Asymptotics of hybrid fluid queues with Lévy input. Journal of Applied Probability, 50(1):103–113, March 2013. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http:/

/projecteuclid.org/euclid.jap/1363784427.

DeMeyer:1980:ABD

[DT80] A. De Meyer and J. L. Teugels. On the asymptotic behaviour of the distributions of the busy period and service time in M/G/1. Journal of Applied Probability, 17(3):802–813, September 1980. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/3212973.

DeKok:1985:QSI

[DT85] A. G. De Kok and H. G. Tijms. A queueing system with impatient customers. Journal of Applied Probability, 22(3):688–696, September 1985. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3213871.

Deng:2001:PQM

[DT01] Yonglu Deng and Jiqing Tan. Priority queueing model with changeover times and switching threshold. Journal of Applied Probability, 38A:263–273, 2001. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.

org/stable/3215885.

Duffy:2011:SPLb

[DT11] Ken R. Duffy and Giovanni Luca Torrisi. Sample path large deviations of Poisson shot noise with heavy-tailed semiexponential distributions. Journal of Applied Probability, 48(3):688–698, September 2011. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
23065929.

Crescenzo:2015:EPL

[DT15] Antonio Di Crescenzo and Abdolsaeed Toomaj. Extension of the past lifetime and its connection to the cumulative entropy. Journal of Applied Probability, 52(4):1156–1174, December 2015. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-

6072 (electronic). URL http://projecteuclid.org/euclid.
jap/1450802759.

Du:1995:MRS

[Du95] Qing Du. A monotonicity result for a single-server queue subject to a Markov-modulated Poisson process. Journal of Applied Probability, 32(4):1103–1111, December 1995. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3215223.

Dubin:1981:PBS

[Dub81] Neil Dubin. Predicting the benefit of screening for disease. Journal of Applied Probability, 18(2):348–360, June 1981. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3213282.

Duffield:1996:ESQ

[Duf96] N. G. Duffield. Economies of scale in queues with sources having power-law large deviation scalings. Journal of Applied Probability, 33(3):840–857, September 1996. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3215363.

Dunstan:1980:NMF

[Dun80] Ross Dunstan. A note on the moments of the final size of the general epidemic model. Journal of Applied Probability, 17(2):532– 538, June 1980. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3213042.

Dunstan:1982:RP

[Dun82] Ross Dunstan. The rumour process. Journal of Applied Probability, 19(4):759–766, December 1982. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3213828.

Durbin:1971:BCP

[Dur71a] J. Durbin. Boundary-crossing probabilities for the Brownian motion and Poisson processes and techniques for computing the power of the Kolmogorov–Smirnov test. Journal of Applied Probability, 8(3):431–453, September 1971. CODEN JPRBAM.

ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/3212169.

Durham:1971:LTG

[Dur71b] Stephen D. Durham. Limit theorems for a general critical branching process. Journal of Applied Probability, 8(1):1–16, March 1971. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3211833.

Durham:1974:LDG

[Dur74] S. D. Durham. Limiting distributions for the general branching process with immigration. Journal of Applied Probability, 11(4): 809–813, December 1974. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/ stable/3212563. See corrections [Ano75g].

Durham:1975:OBM

[Dur75] S. D. Durham. An optimal branching migration process. Journal of Applied Probability, 12(3):569–573, September 1975. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3212872.

Durbin:1985:FPD

[Dur85] J. Durbin. The first-passage density of a continuous Gaussian process to a general boundary. Journal of Applied Probability, 22(1):99–122, March 1985. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/ stable/3213751. See correction [Ano88h].

Durbin:1986:ADS

[Dur86] J. Durbin. Approximate distributions of Student’s t-statistics for autoregressive coefficients calculated from regression residuals. Journal of Applied Probability, 23A:173–185, 1986. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3214351.

Durbin:1988:RTD

[Dur88] J. Durbin. A reconciliation of two different expressions for the first-passage density of Brownian motion to a curved boundary. Journal of Applied Probability, 25(4):829–832, December 1988. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/3214305.

Durrett:1992:SGM

[Dur92] R. Durrett. Stochastic growth models: Bounds on critical values. Journal of Applied Probability, 29(1):11–20, March 1992. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/3214787.

Durieu:2008:FMI

[Dur08a] Oliver Durieu. A fourth moment inequality for functionals of stationary processes. Journal of Applied Probability, 45(4):1086– 1096, December 2008. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/27596010.

Durrleman:2008:CMI

[Dur08b] Valdo Durrleman. Convergence of at-the-money implied volatilities to the spot volatility. Journal of Applied Probability, 45 (2):542–550, June 2008. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/27595963.

DeHaan:1987:EVT

[DV87] L. De Haan and E. Verkade. On extreme-value theory in the presence of a trend. Journal of Applied Probability, 24(1):62– 76, March 1987. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3214059.

Dereudre:2020:EGP

[DV20] David Dereudre and Thibaut Vasseur. Existence of Gibbs point processes with stable infinite range interaction. Journal of Applied Probability, 57(3):775–791, September 2020. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL https://www.cambridge.org/core/journals/
journal-of-applied-probability/article/existence-of-
gibbs-point-processes-with-stable-infinite-range-interaction/
D7D4C9A9F1A048ED123758114056CF7D.

Deijfen:2024:WTI

[DvdHS24] Maria Deijfen, Remco van der Hofstad, and Matteo Sfragara. The winner takes it all but one. Journal of Applied Probability, 61(1):137–152, March 2024. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL https:/

/www.cambridge.org/core/journals/journal-of-applied-
probability/article/winner-takes-it-all-but-one/0B985E5130F237A33C5E4BCD724B18A4.

Daley:2004:SPF

[DVJ04] Daryl J. Daley and David Vere-Jones. Scoring probability forecasts for point processes: The entropy score and information gain. Journal of Applied Probability, 41A:297–312, 2004. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/3215984.

Daley:2008:DGK

[DVJ08] D. J. Daley and D. Vere-Jones. David George Kendall and applied probability. Journal of Applied Probability, 45(2):293– 296, June 2008. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
27595945.

Dvurecenskij:1984:BPO

[Dvu84] A. Dvurecenskij. The busy period of order n in the GI/D/∞ queue. Journal of Applied Probability, 21(1):207–212, March 1984. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3213681.

Darroch:1972:EFR

[DW72] J. N. Darroch and Heather J. Whitford. Exact fluctuation results for Markov-dependent coin-tossing. Journal of Applied Probability, 9(1):158–168, March 1972. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3212644.

Deshmukh:1977:CBD

[DW77] S. D. Deshmukh and Wayne Winston. A controlled birth and death process model of optimal product pricing under stochastically changing demand. Journal of Applied Probability, 14 (2):328–339, June 1977. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3213003.

Droste:1985:NSU

[DW85] W. Droste and W. Wefelmeyer. A note on strong unimodality and dispersivity. Journal of Applied Probability, 22(1):235– 239, March 1985. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3213764.

Donnelly:1989:RDC

[DW89] Peter Donnelly and Ward Whitt. On reinforcement-depletion compartmental urn models. Journal of Applied Probability, 26 (3):477–489, September 1989. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.

org/stable/3214406.

Durbin:1992:FPD

[DW92] J. Durbin and D. Williams. The first-passage density of the Brownian motion process to a curved boundary. Journal of Applied Probability, 29(2):291–304, June 1992. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/3214567.

Dassios:2011:DBP

[DW11] Angelos Dassios and Shanle Wu. Double-barrier Parisian options. Journal of Applied Probability, 48(1):1–20, March 2011. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/29777442.

Dwass:1967:PRT

[Dwa67] Meyer Dwass. Poisson recurrence times. Journal of Applied Probability, 4(3):605–608, December 1967. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/3212224.

Dwass:1969:TPB

[Dwa69] Meyer Dwass. The total progeny in a branching process and a related random walk. Journal of Applied Probability, 6(3): 682–686, December 1969. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3212112.

Dong:2011:PZC

[DWW11] Yinghui Dong, Guojing Wang, and Rong Wu. Pricing the zerocoupon bond and its fair premium under a structural credit risk model with jumps. Journal of Applied Probability, 48(2):404– 419, June 2011. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
23065893.

Dwyer:1988:CHR

[Dwy88] Rex A. Dwyer. On the convex hull of random points in a polytope. Journal of Applied Probability, 25(4):688–699, December 1988. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3214289.

Dion:1997:LTE

[DY97] J. P. Dion and N. M. Yanev. Limit theorems and estimation theory for branching processes with an increasing random number of ancestors. Journal of Applied Probability, 34(2):309–327, June 1997. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3215372.

Doney:1998:FTB

[DY98] R. A. Doney and M. Yor. On a formula of Takács for Brownian motion with drift. Journal of Applied Probability, 35(2):272– 280, June 1998. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3215684.

Duncan:2001:EAQ

[DYY01] Tyrone E. Duncan, Yi Yan, and Peng Yan. Exact asymptotics for a queue with fractional Brownian input and applications in ATM networks. Journal of Applied Probability, 38(4):932–945, December 2001. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3215775.

Duckworth:2000:IME

[DZ00] J. Kate Duckworth and Mihail Zervos. An investment model with entry and exit decisions. Journal of Applied Probability, 37 (2):547–559, June 2000. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3215729.

Davydov:2003:GLC

[DZ03] Youri Davydov and Ricardas Zitikis. Generalized Lorenz curves and convexifications of stochastic processes. Journal of Applied Probability, 40(4):906–925, December 2003. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/3216050.

Denisov:2007:TBC

[DZ07] Denis Denisov and Bert Zwart. On a theorem of breiman and a class of random difference equations. Journal of Applied Probability, 44(4):1031–1046, December 2007. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/27595905.

Dassios:2013:RMD

[DZ13] Angelos Dassios and Hongbiao Zhao. A risk model with delayed claims. Journal of Applied Probability, 50(3):686–702, September 2013. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-

6072 (electronic). URL http://projecteuclid.org/euclid.
jap/1378401230.

Dziubdziela:1984:LLK

[Dzi84] W. Dziubdziela. Limit laws for kth order statistics from strongmixing processes. Journal of Applied Probability, 21(4):720–729, December 1984. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3213690.

Dziubdziela:1986:LLK

[Dzi86] W. Dziubdziela. Limit laws for kth order statistics from conditionally mixing arrays of random variables. Journal of Applied Probability, 23(3):679–687, September 1986. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/3214006.

Eagleson:1975:GCL

[Eag75] G. K. Eagleson. On Gordin’s central limit theorem for stationary processes. Journal of Applied Probability, 12(1):176–179, March 1975. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3212425.

Eagleson:1979:PLT

[Eag79] G. K. Eagleson. A Poisson limit theorem for weakly exchangeable events. Journal of Applied Probability, 16(4):794–802, December 1979. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3213145.

Eaton:1966:CDI

[Eat66] Morris L. Eaton. Characterization of distributions by the identical distribution of linear forms. Journal of Applied Probability, 3(2):481–494, December 1966. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.

org/stable/3212133.

Ebanks:1984:PAE

[Eba84] Bruce R. Ebanks. Polynomially additive entropies. Journal of Applied Probability, 21(1):179–185, March 1984. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3213676.

Ebrahimi:1985:SSS

[Ebr85] Nader Ebrahimi. A stress-strength system. Journal of Applied Probability, 22(2):467–472, June 1985. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3213791. See correction [Ano87f].

Ebrahimi:1987:BPP

[Ebr87] Nader Ebrahimi. Bivariate processes with positive or negative dependent structures. Journal of Applied Probability, 24(1):115– 122, March 1987. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3214064.

Ebrahimi:1990:ORS

[Ebr90] Nader Ebrahimi. Ordering of repairable systems. Journal of Applied Probability, 27(1):193–201, March 1990. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/3214606. See correction

[MS92].

Ebrahimi:1991:HCT

[Ebr91] Nader Ebrahimi. How to compare two systems. Journal of Applied Probability, 28(4):940–946, December 1991. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3214701.

Ebrahimi:1993:MPL

[Ebr93] Nader Ebrahimi. More predictable, less predictable and stable counting processes. Journal of Applied Probability, 30(2):341– 352, June 1993. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3214843.

Ebrahimi:1997:MAR

[Ebr97] Nader Ebrahimi. Multivariate age replacement. Journal of Applied Probability, 34(4):1032–1040, December 1997. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3215015.

Ebrahimi:1999:SPC

[Ebr99] Nader Ebrahimi. Stochastic properties of a cumulative damage threshold crossing model. Journal of Applied Probability, 36(3): 720–732, September 1999. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3215436.

Ebrahimi:2001:SCF

[Ebr01] Nader Ebrahimi. A stochastic covariate failure model for assessing system reliability. Journal of Applied Probability, 38(3): 761–767, September 2001. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3216128.

Ebrahimi:2004:BC

[Ebr04] Nader Ebrahimi. Burn-in and covariates. Journal of Applied Probability, 41(3):735–745, September 2004. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/4141350.

Ebrahimi:2008:ALN

[Ebr08] Nader Ebrahimi. Assessing a linear Nanosystem’s limiting reliability from its components. Journal of Applied Probability, 45(3): 879–887, September 2008. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/27595991.

Ebrahimi:2011:ARF

[Ebr11] Nader Ebrahimi. Assessing the reliability function of nanocomponents. Journal of Applied Probability, 48(1):31–42, March 2011. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/29777444.

Economou:2000:SLB

[Eco00] Antonis Economou. A stochastic lower bound for assembletransfer batch service queueing networks. Journal of Applied Probability, 37(3):881–889, September 2000. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/3215620.

Economou:2003:SDB

[Eco03] Antonis Economou. On the stochastic domination for batcharrival, batch-service and assemble-transfer queueing networks. Journal of Applied Probability, 40(4):1103–1120, December 2003. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/3216063.

Economou:2004:CPI

[Eco04] Antonis Economou. The compound Poisson immigration process subject to binomial catastrophes. Journal of Applied Probability, 41(2):508–523, June 2004. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3216032.

Eddy:1980:DCH

[Edd80] William F. Eddy. The distribution of the convex hull of a Gaussian sample. Journal of Applied Probability, 17(3):686–695, September 1980. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3212962.

El-Desouky:2015:GWT

[EDSM15] B. S. El-Desouky, F. A. Shiha, and A. M. Magar. On a generalization of a waiting time problem and some combinatorial identities. Journal of Applied Probability, 52(4):981–989, December 2015. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://projecteuclid.org/euclid.jap/
1450802747.

Enns:1976:OIP

[EE76] E. G. Enns and P. F. Ehlers. Optimal issuing policies for nonidentical units with stochastic lifetimes. Journal of Applied Probability, 13(1):183–189, March 1976. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3212682.

Enns:1978:RPT

[EE78] E. G. Enns and P. F. Ehlers. Random paths through a convex region. Journal of Applied Probability, 15(1):144–152, March 1978. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3213244.

Ehlers:1981:RSC

[EE81] P. Frank Ehlers and Ernest G. Enns. Random secants of a convex body generated by surface randomness. Journal of Applied Probability, 18(1):157–166, March 1981. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3213176.

Enns:1988:CTC

[EE88] E. G. Enns and P. F. Ehlers. Chords through a convex body generated from within an embedded body. Journal of Applied Probability, 25(4):700–707, December 1988. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/3214290.

Enns:1993:NRC

[EE93] E. G. Enns and P. F. Ehlers. Notes on random chords in convex bodies. Journal of Applied Probability, 30(4):889–897, December 1993. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3214520.

Enns:1990:CBC

[EF90] E. G. Enns and E. Z. Ferenstein. A competitive best-choice problem with Poisson arrivals. Journal of Applied Probability, 27(2):333–342, June 1990. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3214652.

Ernst:2014:OR

[EFS14] Philip A. Ernst, Dean P. Foster, and Larry A. Shepp. On optimal retirement. Journal of Applied Probability, 51(2):333–345, June 2014. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-

6072 (electronic). URL http://projecteuclid.org/euclid.
jap/1402578628.

Eichelsbacher:2002:BIM

[EG02] Peter Eichelsbacher and Ayalvadi Ganesh. Bayesian inference for Markov chains. Journal of Applied Probability, 39(1):91– 99, March 2002. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3215920.

Egozcue:2015:OTS

[EG15] Martin Egozcue and Luis Fuentes García. An optimal threshold strategy in the two-envelope problem with partial information. Journal of Applied Probability, 52(1):298–304, March 2015. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://projecteuclid.org/euclid.jap/1429282624.

Ekstrom:2017:DGH

[EGL17] Erik Ekström, Kristoffer Glover, and Marta Leniec. Dynkin games with heterogeneous beliefs. Journal of Applied Probability, 54(1):236–251, March 2017. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL https:/

/www.cambridge.org/core/journals/journal-of-applied-
probability/article/dynkin-games-with-heterogeneous-
beliefs/BFFFA20FF978F1F4A2EDADE1BD8751BF.

Ewens:1995:OMK

[EH95] W. J. Ewens and C. C. Heyde. Obituary: Motoo Kimura. Journal of Applied Probability, 32(4):1142–1144, December 1995. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/3215227.

Ekstrom:2005:SOS

[EJT05] Erik Ekström, Svante Janson, and Johan Tysk. Superreplication of options on several underlying assets. Journal of Applied Probability, 42(1):27–38, March 2005. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/30040767.

Ekstrom:2003:PAP

[Eks03] Erik Ekström. Perpetual American put options in a leveldependent volatility model. Journal of Applied Probability, 40 (3):783–789, September 2003. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.

org/stable/3215951.

Ekstrom:2004:ROF

[Eks04] Erik Ekström. Russian options with a finite time horizon. Journal of Applied Probability, 41(2):313–326, June 2004. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3216018.

Ekstrom:2006:BPA

[Eks06] Erik Ekström. Bounds for perpetual American option prices in a jump diffusion model. Journal of Applied Probability, 43(3): 867–873, September 2006. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/27595778.

Ekstrom:2013:OCM

[EL13] Erik Ekström and Carl Lindberg. Optimal closing of a momentum trade. Journal of Applied Probability, 50(2):374–387, June 2013. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-

6072 (electronic). URL http://projecteuclid.org/euclid.
jap/1371648947.

Ethier:2019:HSC

[EL19] S. N. Ethier and Jiyeon Lee. How strong can the Parrondo effect be? Journal of Applied Probability, 56(4):1198–1216, December 2019. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-

6072 (electronic). URL https://www.cambridge.org/core/
journals/journal-of-applied-probability/article/how-
strong-can-the-parrondo-effect-be/B5B6E6D52502426C519E7C27A23C9600.

Elfving:1967:PPC

[Elf67] G. Elfving. A persistency problem connected with a point process. Journal of Applied Probability, 4(1):77–89, April 1967. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/3212300.

Elie:2009:DKE

[Eli09] Romuald Elie. Double kernel estimation of sensitivities. Journal of Applied Probability, 46(3):791–811, September 2009. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/25662461.

Ellison:1966:LTR

[Ell66] B. E. Ellison. Limit theorems for random mating in infinite populations. Journal of Applied Probability, 3(1):94–114, June 1966. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3212040.

Elliott:1993:GRD

[Ell93] Robert J. Elliott. A general recursive discrete-time filter. Journal of Applied Probability, 30(3):575–588, September 1993. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/3214767.

Ejov:2011:PHT

[ELNT11] Vladimir Ejov, Nelly Litvak, Giang T. Nguyen, and Peter G. Taylor. Proof of the Hamiltonicity-trace conjecture for singularly perturbed Markov chains. Journal of Applied Probability, 48(4): 901–910, December 2011. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/23066432.

Elshamy:1995:RPV

[Els95] M. Elshamy. Randomly perturbed vibrations. Journal of Applied Probability, 32(2):417–428, June 1995. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3215297.

Elshamy:1996:SMD

[Els96] Maged Elshamy. Stochastic models of damped vibrations. Journal of Applied Probability, 33(4):1159–1168, December 1996. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/3214993.

Elshamy:2007:SPS

[Els07] M. Elshamy. A stability property of stochastic vibration. Journal of Applied Probability, 44(2):444–457, June 2007. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/27595853.

Ekstrom:2010:OLC

[ELTW10] Erik Ekström, Carl Lindberg, Johan Tysk, and Henrik Wanntorp. Optimal liquidation of a call spread. Journal of Applied Probability, 47(2):586–593, June 2010. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/20778499.

Ekstrom:2024:MIV

[EM24] Erik Ekström and Ebba Mellquist. Monotonicity of implied volatility for perpetual put options. Journal of Applied Probability, 61(1):301–310, March 2024. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL https:/

/www.cambridge.org/core/journals/journal-of-applied-
probability/article/monotonicity-of-implied-volatility-
for-perpetual-put-options/9E1321BACC3A4B45CCC1B19390A8041F.

Embrechts:1983:PGI

[Emb83] Paul Embrechts. A property of the generalized inverse Gaussian distribution with some applications. Journal of Applied Probability, 20(3):537–544, September 1983. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3213890.

Emery:1972:LBD

[Eme72] D. J. Emery. Limiting behaviour of the distributions of the maxima of partial sums of certain random walks. Journal of Applied Probability, 9(3):572–579, September 1972. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/3212326.

Emmons:1972:OAP

[Emm72] H. Emmons. The optimal admission policy to a multiserver queue with finite horizon. Journal of Applied Probability, 9 (1):103–116, March 1972. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3212640.

Eaton:1971:ESL

[EMR71] Morris L. Eaton, Carl Morris, and Herman Rubin. On extreme stable laws and some applications. Journal of Applied Probability, 8(4):794–801, December 1971. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3212242.

Engelberg:1964:ELD

[Eng64] Ora Engelberg. Exact and limiting distributions of the number of lead positions in “unconditional” ballot problems. Journal of Applied Probability, 1(1):168–172, June 1964. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3212068.

Engelberg:1965:SPC

[Eng65] Ora Engelberg. On some problems concerning a restricted random walk. Journal of Applied Probability, 2(2):396–404, December 1965. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3212201.

Englund:1980:RTE

[Eng80] Gunnar Englund. Remainder term estimate for the asymptotic normality of the number of renewals. Journal of Applied Probability, 17(4):1108–1113, December 1980. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/3213222.

English:1993:FMR

[Eng93] Brian J. English. Factorial moments for random mappings by means of indicator variables. Journal of Applied Probability, 30 (1):167–174, March 1993. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3214629.

Enns:1969:TDM

[Enn69] E. G. Enns. The trivariate distribution of the maximum queue length, the number of customers served and the duration of the busy period for the M/G/1 queueing system. Journal of Applied Probability, 6(1):154–161, April 1969. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3212282.

El-Neweihi:1978:MCS

[ENPS78] E. El-Neweihi, F. Proschan, and J. Sethuraman. Multistate coherent systems. Journal of Applied Probability, 15(4):675–688, December 1978. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3213425.

El-Neweihi:1986:OAC

[ENPS86] Emad El-Neweihi, Frank Proschan, and Jayaram Sethuraman. Optimal allocation of components in parallel-series and seriesparallel systems. Journal of Applied Probability, 23(3):770–777, September 1986. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3214014.

Embrechts:1984:PLD

[EO84] Paul Embrechts and Edward Omey. A property of longtailed distributions. Journal of Applied Probability, 21(1):80–87, March 1984. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3213666.

Ekstrom:2020:RTA

[EOV20] Erik Ekström, Marcus Olofsson, and Martin Vannestål. A renewal theory approach to two-state switching problems with infinite values. Journal of Applied Probability, 57(1):1–18, March 2020. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL https:/

/www.cambridge.org/core/journals/journal-of-applied-
probability/article/renewal-theory-approach-to-twostate-
switching-problems-with-infinite-values/6B3755DEEE6780F4E2B2AA3D53E9DD23.

Ebrahimi:1995:NPO

[EP95] Nader Ebrahimi and Franco Pellerey. New partial ordering of survival functions based on the notion of uncertainty. Journal of Applied Probability, 32(1):202–211, March 1995. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3214930.

Eplett:1983:MDS

[Epl83] W. J. R. Eplett. Minimal distributions in a stochastic partial ordering and bounds for Gaussian processes. Journal of Applied Probability, 20(3):529–536, September 1983. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/3213889.

Ebrahimi:1988:DSH

[ER88] Nader Ebrahimi and T. Ramalingam. On the dependence structure of hitting times of univariate processes. Journal of Applied Probability, 25(2):355–362, June 1988. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/3214443.

Ebrahimi:1989:DSH

[ER89] Nader Ebrahimi and T. Ramallingam. On the dependence structure of hitting times of multivariate processes. Journal of Applied Probability, 26(2):287–295, June 1989. CODEN JPRBAM.

ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/3214035.

Erhardsson:2000:CPA

[Erh00] Torkel Erhardsson. Compound Poisson approximation for the Johnson-mehl model. Journal of Applied Probability, 37(1):101– 117, March 2000. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3215662.

Erhardsson:2023:RPR

[Erh23] Torkel Erhardsson. Reciprocal properties of random fields on undirected graphs. Journal of Applied Probability, 60 (3):781–796, September 2023. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL https:/

/www.cambridge.org/core/journals/journal-of-applied-
probability/article/reciprocal-properties-of-random-
fields-on-undirected-graphs/A70BF6C2D47E71418912AC5B20EB48EA.

Erickson:2016:UNS

[Eri16] K. Bruce Erickson. Uniqueness of the null solution to a nonlinear partial differential equation satisfied by the explosion probability of a branching diffusion. Journal of Applied Probability, 53(3): 938–945, September 2016. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://projecteuclid.

org/euclid.jap/1476370788.

Erlander:1967:NTT

[Erl67] Sven Erlander. A note on telephone traffic with losses. Journal of Applied Probability, 4(2):406–408, August 1967. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3212035.

Eryilmaz:2008:RSD

[Ery08] Serkan Eryilmaz. Run statistics defined on the multicolor urn model. Journal of Applied Probability, 45(4):1007–1023, December 2008. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/27596005.

Eryilmaz:2010:DSS

[Ery10] Serkan Eryilmaz. Discrete scan statistics generated by exchangeable binary trials. Journal of Applied Probability, 47(4):1084– 1092, December 2010. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/29777428.

Eryilmaz:2015:CIC

[Ery15] Serkan Eryilmaz. Component importance in coherent systems with exchangeable components. Journal of Applied Probability, 52(3):851–863, September 2015. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http:/

/projecteuclid.org/euclid.jap/1445543851.

Engen:1987:UBI

[ES87] Steinar Engen and Eva Seim. Uniformly best invariant stopping rules. Journal of Applied Probability, 24(1):77–87, March 1987. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3214060.

Eisenberg:2011:OCC

[ES11] Julia Eisenberg and Hanspeter Schmidli. Optimal control of capital injections by reinsurance with a constant rate of interest. Journal of Applied Probability, 48(3):733–748, September 2011. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/23065933.

Esmaeeli:2023:PRS

[ES23] Neda Esmaeeli and Farkhondeh Alsadat Sajadi. On the probability of rumour survival among sceptics. Journal of Applied Probability, 60(3):1096–1111, September 2023. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL https://www.cambridge.org/core/journals/journal-
of-applied-probability/article/on-the-probability-of-
rumour-survival-among-sceptics/67F277C30BE6EEB7F946C02B097E5A0A.

Elliott:2015:BSD

[ESC15] Robert J. Elliott, Tak Kuen Siu, and Samuel N. Cohen. Backward stochastic difference equations for dynamic convex risk measures on a binomial tree. Journal of Applied Probability, 52(3):771–785, September 2015. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http:/

/projecteuclid.org/euclid.jap/1445543845.

Enns:1984:HSS

[ESE84] E. G. Enns, B. R. Smith, and P. F. Ehlers. Hitting spheres with straight-line motion or Brownian motion. Journal of Applied Probability, 21(1):70–79, March 1984. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3213665.

Eshel:1973:CSO

[Esh73] Ilan Eshel. Clone-selection and optimal rates of mutation. Journal of Applied Probability, 10(4):728–738, December 1973. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/3212376.

Es-Saghouani:2008:CSL

[ESM08] Abdelghafour Es-Saghouani and Michel Mandjes. On the correlation structure of a Lévy–driven queue. Journal of Applied Probability, 45(4):940–952, December 2008. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/27596000.

Esty:1975:RGW

[Est75] Warren W. Esty. The reverse Galton–Watson process. Journal of Applied Probability, 12(3):574–580, September 1975. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3212873.

Esty:1976:DLC

[Est76] Warren W. Esty. Diffusion limits of critical branching processes conditioned on extinction in the near future. Journal of Applied Probability, 13(2):247–254, June 1976. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3212828.

Ethier:1983:PBR

[ET83] S. N. Ethier and S. Tavaré. The proportional Bettor’s return on investment. Journal of Applied Probability, 20(3):563–573, September 1983. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3213893.

Eryilmaz:2021:CMS

[ET21] Serkan Eryilmaz and Fatih Tank. Computing minimal signature of coherent systems through matrix-geometric distributions. Journal of Applied Probability, 58(3):621–636, September 2021.

CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL https://www.cambridge.org/core/journals/
journal-of-applied-probability/article/computing-minimal-
signature-of-coherent-systems-through-matrixgeometric-
distributions/758B5B92CDDB0143046BC48D266B9929.

Ethier:1979:LTT

[Eth79] S. N. Ethier. A limit theorem for two-locus diffusion models in population genetics. Journal of Applied Probability, 16(2):402– 408, June 1979. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3212907.

Ethier:1992:EIM

[Eth92a] S. N. Ethier. Eigenstructure of the infinitely-many-neutral-alleles diffusion model. Journal of Applied Probability, 29(3): 487–498, September 1992. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3214888.

Ethier:1992:ETD

[Eth92b] S. N. Ethier. Equivalence of two descriptions of the ages of alleles. Journal of Applied Probability, 29(1):185–189, March 1992. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3214802.

Ethier:2004:KSM

[Eth04] S. N. Ethier. The Kelly system maximizes median fortune. Journal of Applied Probability, 41(4):1230–1236, December 2004. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/4141397.

Engelen:1988:ITN

[ETV88] Ron Engelen, Paul Tommassen, and Wim Vervaat. Ignatov’s theorem: A new and short proof. Journal of Applied Probability, 25A:229–236, 1988. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3214159.

Erdos:1982:ACC

[EV82] P. Erdős and I. Vincze. On the approximation of convex, closed plane curves by multifocal ellipses. Journal of Applied Probability, 19A:89–96, 1982. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3213552.

Ekstrom:2016:MLU

[EV16] Erik Ekström and Martin Vannestål. Momentum liquidation under partial information. Journal of Applied Probability, 53 (2):341–359, June 2016. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://projecteuclid.

org/euclid.jap/1466172858.

Evans:1971:PPC

[Eva71] Richard V. Evans. Programming problems and changes in the stable behavior of a class of Markov chains. Journal of Applied Probability, 8(3):543–550, September 1971. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/3212177.

Evans:1978:UBM

[Eva78] Lawrence S. Evans. An upper bound for the mean of Yaglom’s limit. Journal of Applied Probability, 15(1):199–201, March 1978. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3213252.

Eshel:1988:CRG

[EW88] Ilan Eshel and Daphna Weinshall. Cooperation in a repeated game with random payment function. Journal of Applied Probability, 25(3):478–491, September 1988. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3213977.

Embrechts:2000:RED

[EW00] Paul Embrechts and Harro Walk. Recursive estimation of distributional fix-points. Journal of Applied Probability, 37(1):73– 87, March 2000. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3215660.

Ekstrom:2009:OSB

[EW09] Erik Ekström and Henrik Wanntorp. Optimal stopping of a Brownian bridge. Journal of Applied Probability, 46(1):170– 180, March 2009. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
25662414.

Ewens:1964:PTD

[Ewe64] W. J. Ewens. The pseudo-transient distribution and its uses in genetics. Journal of Applied Probability, 1(1):141–156, June 1964. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3212065.

Ewens:1982:APG

[Ewe82] W. J. Ewens. Aspects of population genetics theory. Journal of Applied Probability, 19A:9–17, 1982. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3213546.

Ebrahimi:2011:IPM

[EY11] Nader Ebrahimi and Yarong Yang. An integrated probabilistic model for assessing a nanocomponent’s reliability. Journal of Applied Probability, 48(3):832–842, September 2011. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/23065940.

Engel:1980:CGD

[EZ80] Jan Engel and Mynt Zijlstra. A characterization of the gamma distribution by the negative binomial distribution. Journal of Applied Probability, 17(4):1138–1144, December 1980. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3213228.

Frostig:1985:SFN

[FA85] E. Frostig and I. Adiri. Stochastic flowshop no-wait scheduling. Journal of Applied Probability, 22(1):240–246, March 1985. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/3213765.

Faddy:1974:OCF

[Fad74] M. J. Faddy. Optimal control of finite dams: Discrete (2-stage) output procedure. Journal of Applied Probability, 11(1):111– 121, March 1974. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3212588.

Faddy:1983:OGS

[Fad83] M. J. Faddy. The outcome of a general spatial epidemic on the line. Journal of Applied Probability, 20(4):715–727, December 1983. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3213584.

Fakinos:1982:MGK

[Fak82] D. Fakinos. The M/G/k group-arrival group-departure loss system. Journal of Applied Probability, 19(4):826–834, December 1982. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3213835.

Fakinos:1983:MGK

[Fak83] D. Fakinos. The M/G/k loss system with servers subject to breakdown. Journal of Applied Probability, 20(3):706–712, September 1983. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3213906.

Fakinos:1986:SSQ

[Fak86] D. Fakinos. On the single-server queue with the preemptiveresume last-come-first-served queue discipline. Journal of Applied Probability, 23(1):243–248, March 1986. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3214136.

Fakinos:1987:SSQ

[Fak87] D. Fakinos. The single-server queue with service depending on queue size and with the preemptive-resume last-come-first-served queue discipline. Journal of Applied Probability, 24(3): 758–767, September 1987. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3214105.

Fakinos:1990:RBP

[Fak90] D. Fakinos. On the relation between partial balance and insensitivity. Journal of Applied Probability, 27(2):445–451, June 1990. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3214665.

Falin:1986:WTP

[Fal86] G. I. Falin. On the waiting-time process in a single-line queue with repeated calls. Journal of Applied Probability, 23(1):185– 192, March 1986. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3214127.

Falin:1989:LE

[Fal89] Gennadij Falin. Letter to the Editor. Journal of Applied Probability, 26(3):683–684, September 1989. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3214429.

Falin:1994:MGB

[Fal94] Gennadi Falin. The M^{k} /G/∞ batch arrival queue with heterogeneous dependent demands. Journal of Applied Probability, 31 (3):841–846, September 1994. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.

org/stable/3215162.

Falin:1996:HBS

[Fal96] G. Falin. A heterogeneous blocking system in a random environment. Journal of Applied Probability, 33(1):211–216, March 1996. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3215278.

Fang:1982:ROP

[Fan82] Kai-Tai Fang. A restricted occupancy problem. Journal of Applied Probability, 19(3):707–711, September 1982. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3213532.

Fang:2012:TMG

[Fan12] Ming Fang. Tightness for maxima of generalized branching random walks. Journal of Applied Probability, 49(3):652–670, September 2012. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
41713796.

Fassbender:1990:OSS

[Fas90] Matthias Fassbender. Optimal stationary strategies in leavable Markov decision processes. Journal of Applied Probability, 27 (1):134–145, March 1990. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3214601.

Fatheddin:2015:CLT

[Fat15] Parisa Fatheddin. Central limit theorem for a class of SPDEs. Journal of Applied Probability, 52(3):786–796, September 2015. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-

6072 (electronic). URL http://projecteuclid.org/euclid.
jap/1445543846.

Fristedt:1988:OMS

[FB88] Bert Fristedt and Donald A. Berry. Optimality of myopic stopping times for geometric discounting. Journal of Applied Probability, 25(2):437–443, June 1988. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3214454.

Fu:2002:PGF

[FC02] James C. Fu and Y. M. Chang. On probability generating functions for waiting time distributions of compound patterns in a sequence of multistate trials. Journal of Applied Probability, 39 (1):70–80, March 2002. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3215918.

Fu:2003:OSL

[FC03] James C. Fu and Yung-Ming Chang. On ordered series and later waiting time distributions in a sequence of Markov dependent multistate trials. Journal of Applied Probability, 40(3):623–642, September 2003. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3215940.

Ferreiro-Castilla:2016:EPS

[FCKS16] A. Ferreiro-Castilla, A. E. Kyprianou, and R. Scheichl. An Euler–Poisson scheme for Lévy driven stochastic differential equations. Journal of Applied Probability, 53(1):262–278, March 2016. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://projecteuclid.org/euclid.jap/
1457470573.

Falk:2018:MRR

[FCP18] M. Falk, A. Khorrami Chokami, and S. A. Padoan. On multivariate records from random vectors with independent components. Journal of Applied Probability, 55(1):43–53, March 2018. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL https://www.cambridge.org/core/journals/journal-
of-applied-probability/article/on-multivariate-records-
from-random-vectors-with-independent-components/564EAFB2CE36527C2A015DCF307D2105.

Falk:2020:RTD

[FCP20] Michael Falk, Amir Khorrami Chokami, and Simone A. Padoan. Records for time-dependent stationary Gaussian sequences. Journal of Applied Probability, 57(1):78–96, March 2020. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL https://www.cambridge.org/core/journals/journal-
of-applied-probability/article/records-for-timedependent-
stationary-gaussian-sequences/ADE2A427AEF890F2E60619189EE7DB55.

Ferreiro-Castilla:2015:AWH

[FCvS15] Albert Ferreiro-Castilla and Kees van Schaik. Applying the Wiener–Hopf Monte Carlo simulation technique for Lévy processes to path functionals. Journal of Applied Probability, 52 (1):129–148, March 2015. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://projecteuclid.

org/euclid.jap/1429282611.

Frank:1984:RMB

[FD84] Douglas H. Frank and Stephen Durham. Random motion on binary trees. Journal of Applied Probability, 21(1):58–69, March 1984. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3213664.

Feldman:1993:UML

[FDVF93] Richard M. Feldman, Bryan L. Deuermeyer, and Ciriaco Valdez-Flores. Utilization of the method of linear matrix equations to solve a quasi-birth-death problem. Journal of Applied Probability, 30(3):639–649, September 1993. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3214772.

Faddy:1987:SDP

[FE87] M. J. Faddy and Warren W. Esty. A stochastic death process with elimination by pairing. Journal of Applied Probability, 24 (1):241–245, March 1987. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3214075.

Fearn:1976:PEC

[Fea76] Dean H. Fearn. Probability of extinction of critical generationdependent Galton–Watson processes. Journal of Applied Probability, 13(3):573–577, September 1976. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3212477.

Fearn:1980:EPC

[Fea80] Dean H. Fearn. Extinction probability for critical age-dependent branching processes with generation dependence. Journal of Applied Probability, 17(1):16–24, March 1980. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/3212920.

Fearn:1981:FPP

[Fea81] Dean H. Fearn. A fixed-point property for Galton–Watson processes with generation dependence. Journal of Applied Probability, 18(2):514–519, June 1981. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3213298.

Fearnhead:2002:CAN

[Fea02] Paul Fearnhead. The common ancestor at a nonneutral locus. Journal of Applied Probability, 39(1):38–54, March 2002. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/3215916.

Fearnhead:2003:HJD

[Fea03] Paul Fearnhead. Haplotypes: The joint distribution of alleles at linked loci. Journal of Applied Probability, 40(2):505–512, June 2003. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3215806.

Feigin:1979:CPP

[Fei79a] Paul D. Feigin. On the characterization of point processes with the order statistic property. Journal of Applied Probability, 16 (2):297–304, June 1979. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3212898.

Feigin:1979:SCC

[Fei79b] Paul D. Feigin. Some comments concerning a curious singularity. Journal of Applied Probability, 16(2):440–444, June 1979. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/3212913.

Feingold:1993:MPM

[Fei93] Eleanor Feingold. Markov processes for modeling and analyzing a new genetic mapping method. Journal of Applied Probability, 30(4):766–779, December 1993. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.

org/stable/3214511.

Feinberg:1994:GEE

[Fei94] Eugene A. Feinberg. A generalization of ’expectation equals reciprocal of intensity’ to non-stationary exponential distributions. Journal of Applied Probability, 31(1):262–267, March 1994. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/3215253.

Feldman:1966:OND

[Fel66] M. W. Feldman. On the offspring number distribution in a genetic population. Journal of Applied Probability, 3(1):129– 141, June 1966. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3212042.

Feldman:1976:ORS

[Fel76] Richard M. Feldman. Optimal replacement with semi-Markov shock models. Journal of Applied Probability, 13(1):108–117, March 1976. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3212670.

Feldman:1978:CRI

[Fel78] Richard M. Feldman. A continuous review (s, s) inventory system in a random environment. Journal of Applied Probability, 15(3):654–659, September 1978. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.

org/stable/3213131.

Feldman:1991:APF

[Fel91] Raisa Epstein Feldman. Autoregressive processes and first-hit probabilities for randomized random walks. Journal of Applied Probability, 28(2):347–359, June 1991. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3214871.

Fendick:1993:AED

[Fen93] Kerry W. Fendick. An asymptotically exact decomposition of coupled Brownian systems. Journal of Applied Probability, 30 (4):819–834, December 1993. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.

org/stable/3214515.

Feng:1995:PTS

[Fen95] Shui Feng. Phase transitions of some non-linear stochastic models. Journal of Applied Probability, 32(1):193–201, March 1995. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/3214929.

Feng:1997:PCM

[Fen97] Shui Feng. The propagation of chaos of multitype mean field interacting particle systems. Journal of Applied Probability, 34 (2):346–362, June 1997. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3215375.

Ferebee:1983:UED

[Fer83] Brooks Ferebee. An unbiased estimator for the drift of a stopped Wiener process. Journal of Applied Probability, 20(1):94–102, March 1983. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3213723.

Ferreira:1993:JEH

[Fer93] Helena Ferreira. Joint exceedances of high levels under a local dependence condition. Journal of Applied Probability, 30(1):112– 120, March 1993. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3214625.

Ferger:1995:JDR

[Fer95] Dietmar Ferger. The joint distribution of the running maximum and its location of D-valued Markov processes. Journal of Applied Probability, 32(3):842–845, September 1995. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3215136.

Ferreira:2006:UIE

[Fer06] H. Ferreira. The upcrossings index and the extremal index. Journal of Applied Probability, 43(4):927–937, December 2006. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/27595790.

Falin:1991:VWT

[FF91] G. Falin and C. Fricker. On the virtual waiting time in an M/G/1 retrial queue. Journal of Applied Probability, 28(2):446– 460, June 1991. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3214879.

Ferrari:1996:PAT

[FF96] P. A. Ferrari and L. R. G. Fontes. Poissonian approximation for the tagged particle in asymmetric simple exclusion. Journal of Applied Probability, 33(2):411–419, June 1996. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3215064.

Fontenas:2003:IBL

[FF03] Eric Fontenas and Olivier François. Improved bounds for the large-time behaviour of simulated annealing. Journal of Applied Probability, 40(4):961–969, December 2003. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/3216053.

Feinberg:2009:IVD

[FF09] Eugene A. Feinberg and Jun Fei. An inequality for variances of the discounted rewards. Journal of Applied Probability, 46(4): 1209–1212, December 2009. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/25662491.

Fayard:2011:DMS

[FF11] Patrick Fayard and Timothy R. Field. Discrete models for scattering populations. Journal of Applied Probability, 48(1):285– 292, March 2011. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
29777460.

Fernholz:2020:ZLA

[FF20] Ricardo T. Fernholz and Robert Fernholz. Zipf’s law for atlas models. Journal of Applied Probability, 57(4): 1276–1297, December 2020. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL https:/

/www.cambridge.org/core/journals/journal-of-applied-
probability/article/zipfs-law-for-atlas-models/5D6B730DDEE4C05CF494213FDA57B064.

Fekete:2019:SSD

[FFK19] D. Fekete, J. Fontbona, and A. E. Kyprianou. Skeletal stochastic differential equations for continuous-state branching processes. Journal of Applied Probability, 56(4):1122–1150, December 2019. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL https://www.cambridge.org/core/journals/
journal-of-applied-probability/article/skeletal-stochastic-
differential-equations-for-continuousstate-branching-
processes/0F6409F07D9EA14484CC4EBDC7F17C35.

Fekete:2020:SSD

[FFK20] Dorottya Fekete, Joaquin Fontbona, and Andreas E. Kyprianou. Skeletal stochastic differential equations for superprocesses. Journal of Applied Probability, 57(4):1111–1134, December 2020. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL https://www.cambridge.org/core/journals/
journal-of-applied-probability/article/skeletal-stochastic-
differential-equations-for-superprocesses/401EFD8362722DE4DE7941090CA1236C.

Foster:1976:RGG

[FG76] James H. Foster and Robert T. Goettge. The rates of growth of the Galton–Watson process in varying environment. Journal of Applied Probability, 13(1):144–147, March 1976. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/3212675. See correction

[Gre78].

Federgruen:1984:MGC

[FG84] Awi Federgruen and Linda Green. An M/G/c queue in which the number of servers required is random. Journal of Applied Probability, 21(3):583–601, September 1984. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/3213620.

Federgruen:1987:ICC

[FG87] A. Federgruen and H. Groenevelt. The impact of the composition of the customer base in general queueing models. Journal of Applied Probability, 24(3):709–724, September 1987. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3214101.

Ferrari:1998:ODL

[FG98] Pablo A. Ferrari and Nancy Lopes Garcia. One-dimensional loss networks and conditioned M/G/∞ queues. Journal of Applied Probability, 35(4):963–975, December 1998. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/3215532.

Farrington:1999:DTE

[FG99] C. P. Farrington and A. D. Grant. The distribution of time to extinction in subcritical branching processes: Applications to outbreaks of infectious disease. Journal of Applied Probability, 36(3):771–779, September 1999. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.

org/stable/3215440.

Ferland:2008:LLN

[FG08] René Ferland and Gaston Giroux. Law of large numbers for dynamic bargaining markets. Journal of Applied Probability, 45 (1):45–54, March 2008. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/27595922.

Fralix:2012:NPW

[FG12] Brian Fralix and Colin Gallagher. A new proof of the Wiener– Hopf factorization via Basu’s Theorem. Journal of Applied Probability, 49(3):876–882, September 2012. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http:/

/www.jstor.org/stable/41713812.

Fang:2021:ALR

[FGH^{+}21] Xiao Fang, Han L. Gan, Susan Holmes, Haiyan Huang, Erol Peköz, Adrian Röllin, and Wenpin Tang. Arcsine laws for random walks generated from random permutations with applications to genomics. Journal of Applied Probability, 58 (4):851–867, December 2021. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL https:/

/www.cambridge.org/core/journals/journal-of-applied-
probability/article/arcsine-laws-for-random-walks-generated-
from-random-permutations-with-applications-to-genomics/
8138272A2116DDD209544DA303FF6E3E.

Fan:2000:MLD

[FGL00] Juanjuan Fan, S. G. Ghurye, and Richard A. Levine. Multicomponent lifetime distributions in the presence of ageing. Journal of Applied Probability, 37(2):521–533, June 2000. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3215726.

Flaspohler:1972:AQS

[FH72] David C. Flaspohler and Paul T. Holmes. Additional quasistationary distributions for semi-Markov processes. Journal of Applied Probability, 9(3):671–676, September 1972. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3212336.

Ferguson:1989:SRP

[FH89] T. S. Ferguson and J. P. Hardwick. Stopping rules for proofreading. Journal of Applied Probability, 26(2):304–313, June 1989. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/3214037.

Fawzi:1990:ASS

[FH90] B. B. Fawzi and A. G. Hawkes. Availability of a series system with replacement and repair. Journal of Applied Probability, 27 (4):873–887, December 1990. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.

org/stable/3214830.

Fawzi:1991:ASS

[FH91] B. B. Fawzi and Alan G. Hawkes. Availability of an R-out-of-N system with spares and repairs. Journal of Applied Probability, 28(2):397–408, June 1991. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3214875.

Floreen:1994:UMV

[FH94] Patrik Floréen and Taneli Huuskonen. Uniqueness of maximum values in discrete distributions. Journal of Applied Probability, 31(3):757–764, September 1994. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.

org/stable/3215153.

Fu:1995:UHR

[FH95] Michael C. Fu and Jian-Qiang Hu. On unbounded hazard rates for smoothed perturbation analysis. Journal of Applied Probability, 32(3):659–667, September 1995. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3215120.

Filliger:2007:EGI

[FH07] Roger Filliger and Max-Olivier Hongler. Explicit Gittins indices for a class of superdiffusive processes. Journal of Applied Probability, 44(2):554–559, June 2007. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/27595861.

Field:2010:BPF

[FH10] A. J. Field and P. G. Harrison. Busy periods in fluid queues with multiple emptying input states. Journal of Applied Probability, 47(2):474–497, June 2010. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/20778492.

Feng:2011:ZIR

[FH11] Qunqiang Feng and Zhishui Hu. On the Zagreb index of random recursive trees. Journal of Applied Probability, 48(4):1189–1196, December 2011. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
23066451.

Feng:2013:PCT

[FH13] Qunqiang Feng and Zhishui Hu. Phase changes in the topological indices of scale-free trees. Journal of Applied Probability, 50 (2):516–532, June 2013. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://projecteuclid.

org/euclid.jap/1371648958.

Fuchs:2017:DBP

[FH17] Michael Fuchs and Hsien-Kuei Hwang. Dependence between path-length and size in random digital trees. Journal of Applied Probability, 54(4):1125–1143, December 2017. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL https://www.cambridge.org/core/journals/journal-
of-applied-probability/article/dependence-between-pathlength-
and-size-in-random-digital-trees/CA7EA5182AEC91E4E1A19DB1CFF6F874.

Finkelstein:2018:OOS

[FHC18] Maxim Finkelstein, Nil Kamal Hazra, and Ji Hwan Cha. On optimal operational sequence of components in a warm standby system. Journal of Applied Probability, 55(4): 1014–1024, December 2018. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL https:/

/www.cambridge.org/core/journals/journal-of-applied-
probability/article/on-optimal-operational-sequence-
of-components-in-a-warm-standby-system/71EE057C4EE36F1BBD85363ACA522B24.

Fuchs:2017:CTR

[FHI17] Michael Fuchs, Hsien-Kuei Hwang, and Yoshiaki Itoh. From coin tossing to rock-paper-scissors and beyond: a logexp gap theorem for selecting a leader. Journal of Applied Probability, 54(1):213–235, March 2017. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL https://www.cambridge.org/core/journals/journal-
of-applied-probability/article/from-coin-tossing-to-
rockpaperscissors-and-beyond-a-logexp-gap-theorem-for-
selecting-a-leader/5B40CFB42B2540AC499EF2C582283830.

Fuchs:2014:BSP

[FHIM14] Michael Fuchs, Hsien-Kuei Hwang, Yoshiaki Itoh, and Hosam H. Mahmoud. A binomial splitting process in connection with corner parking problems. Journal of Applied Probability, 51(4): 971–989, December 2014. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://projecteuclid.

org/euclid.jap/1421763322.

Ferre:2013:RPG

[FHL13] Déborah Ferré, Loïc Hervé, and James Ledoux. Regular perturbation of V -geometrically ergodic Markov chains. Journal of Applied Probability, 50(1):184–194, March 2013. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://projecteuclid.org/euclid.jap/1363784432.

Federgruen:1978:NSR

[FHT78] A. Federgruen, A. Hordijk, and H. C. Tijms. A note on simultaneous recurrence conditions on a set of denumerable stochastic matrices. Journal of Applied Probability, 15(4):842–847, December 1978. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3213439.

Fackrell:2010:ADP

[FHTZ10] Mark Fackrell, Qi-Ming He, Peter Taylor, and Hanqin Zhang. The algebraic degree of phase-type distributions. Journal of Applied Probability, 47(3):611–629, September 2010. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/20778506.

Falk:2015:GML

[FHZ15] Michael Falk, Martin Hofmann, and Maximilian Zott. On generalized max-linear models and their statistical interpolation. Journal of Applied Probability, 52(3):736–751, September 2015. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://projecteuclid.org/euclid.jap/
1445543843.

Fayolle:1987:CNE

[FI87] Guy Fayolle and Rudolph Iasnogorodski. Criteria for the nonergodicity of stochastic processes: Application to the exponential back-off protocol. Journal of Applied Probability, 24(2):347– 354, June 1987. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3214259.

Feldman:1998:NGD

[FI98] Raisa E. Feldman and Srikanth K. Iyer. Non-Gaussian density processes arising from non-Poisson systems of independent Brownian motions. Journal of Applied Probability, 35(1):213– 220, March 1998. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3215559.

Finch:1964:TIS

[Fin64] P. D. Finch. The theory of information and statistical inference. I. Journal of Applied Probability, 1(1):121–140, June 1964. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/3212064.

Finch:1969:NQS

[Fin69] P. D. Finch. A note on the queueing system GI/E_{k} /1. Journal of Applied Probability, 6(3):708–710, December 1969. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3212117.

Finch:1982:FDD

[Fin82] P. D. Finch. Functional dependence from a descriptive viewpoint. Journal of Applied Probability, 19A:217–224, 1982. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/3213562.

Finch:1988:BIP

[Fin88] P. D. Finch. Bell’s inequality, probability modelling and quantum correlation experiments. Journal of Applied Probability, 25A:139–150, 1988. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3214152.

Finkelstein:2004:MRH

[Fin04] M. S. Finkelstein. Minimal repair in heterogeneous populations. Journal of Applied Probability, 41(1):281–286, March 2004. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/3215833.

Finkelstein:2007:SAP

[Fin07] Maxim Finkelstein. On some ageing properties of general repair processes. Journal of Applied Probability, 44(2):506–513, June 2007. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/27595857.

Finkelstein:2010:NCG

[Fin10] Maxim Finkelstein. A note on converging geometric-type processes. Journal of Applied Probability, 47(2):601–607, June 2010. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/20778501.

Fink:2013:CCF

[Fin13] Holger Fink. Conditional characteristic functions of Molchan– Golosov fractional Lévy processes with application to credit risk. Journal of Applied Probability, 50(4):983–1005, December 2013. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-

6072 (electronic). URL http://projecteuclid.org/euclid.
jap/1389370095.

Foley:1989:SIA

[FK89] R. D. Foley and Georgia-Ann Klutke. Stationary increments in the accumulated work process in processor-sharing queues. Journal of Applied Probability, 26(3):671–677, September 1989. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/3214427.

Foley:1991:SIA

[FKK91] R. D. Foley, Georgia-Ann Klutke, and Dieter König. Stationary increments of accumulation processes in queues and generalized semi-Markov schemes. Journal of Applied Probability, 28 (4):864–872, December 1991. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.

org/stable/3214689.

Fontbona:2010:EEC

[FKM10] Joaquín Fontbona, Nathalie Krell, and Servet Martínez. Energy efficiency of consecutive fragmentation processes. Journal of Applied Probability, 47(2):543–561, June 2010. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/20778496.

Foss:2009:CLT

[FKZ09] Sergey Foss, Dmitry Korshunov, and Stan Zachary. Convolutions of long-tailed and subexponential distributions. Journal of Applied Probability, 46(3):756–767, September 2009. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/25662459.

Fink:2013:CDP

[FKZ13] Holger Fink, Claudia Klüppelberg, and Martina Zähle. Conditional distributions of processes related to fractional Brownian motion. Journal of Applied Probability, 50(1):166–183, March 2013. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://projecteuclid.org/euclid.jap/
1363784431.

Ferrandiz:1991:RCS

[FL91] Josep M. Ferrandiz and Aurel A. Lazar. Rate conservation for stationary processes. Journal of Applied Probability, 28(1):146– 158, March 1991. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3214747.

Frostig:1997:SCF

[FL97] Esther Frostig and Tapani Lehtonen. Stochastic comparisons for fork-join queues with exponential processing times. Journal of Applied Probability, 34(2):487–497, June 1997. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3215387.

Fuh:1998:WEF

[FL98] Cheng Der Fuh and Tze Leung Lai. Wald’s equations, first passage times and moments of ladder variables in Markov random walks. Journal of Applied Probability, 35(3):566–580, September 1998. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3215632.

Flandoli:2019:PSA

[FL19] Franco Flandoli and Marta Leocata. A particle system approach to aggregation phenomena. Journal of Applied Probability, 56(1):282–306, March 2019. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL https:/

/www.cambridge.org/core/journals/journal-of-applied-
probability/article/particle-system-approach-to-aggregation-
phenomena/43C53396E961BF330EEDEDF12E85B74C.

Foxall:2020:GSC

[FL20] Eric Foxall and Nicolas Lanchier. Generalized stacked contact process with variable host fitness. Journal of Applied Probability, 57(1):97–121, March 2020. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL https:/

/www.cambridge.org/core/journals/journal-of-applied-
probability/article/generalized-stacked-contact-process-
with-variable-host-fitness/AE9F48CB4115C81E27F0C52206E41833.

Fierro:2015:HPD

[FLM15] Raúl Fierro, Víctor Leiva, and Jesper Møller. The Hawkes process with different exciting functions and its asymptotic behavior. Journal of Applied Probability, 52(1):37–54, March 2015. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://projecteuclid.org/euclid.jap/
1429282605.

Florens-Landais:1999:LDE

[FLP99] Danielle Florens-Landais and Huyên Pham. Large deviations in estimation of an Ornstein–Uhlenbeck model. Journal of Applied Probability, 36(1):60–77, March 1999. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3215402.

Ferguson:1997:LRB

[FM97] Thomas S. Ferguson and C. Melolidakis. Last round betting. Journal of Applied Probability, 34(4):974–987, December 1997. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/3215011.

Franceschetti:2006:NSW

[FM06] Massimo Franceschetti and Ronald Meester. Navigation in small-world networks: A scale-free continuum model. Journal of Applied Probability, 43(4):1173–1180, December 2006. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/27595808.

Feng:2010:VSF

[FM10] Qunqiang Feng and Hosam M. Mahmoud. On the variety of shapes on the fringe of a random recursive tree. Journal of Applied Probability, 47(1):191–200, March 2010. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/25662509.

Fougeres:2012:RMM

[FM12] Anne-Laure Fougeres and Cecile Mercadier. Risk measures and multivariate extensions of Breiman’s Theorem. Journal of Applied Probability, 49(2):364–384, June 2012. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/41713773.

Fort:2003:GEH

[FMRR03] G. Fort, E. Moulines, G. O. Roberts, and J. S. Rosenthal. On the geometric ergodicity of hybrid samplers. Journal of Applied Probability, 40(1):123–146, March 2003. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http:/

/www.jstor.org/stable/3215845.

Fierro:1999:LCC

[FMS99] Raúl Fierro, Servet Martínez, and Jaime San Martín. Limiting conditional and conditional invariant distributions for the Poisson process with negative drift. Journal of Applied Probability, 36(4):1194–1209, December 1999. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/3215588.

Fontes:2004:CPF

[FMS04] L. R. Fontes, F. P. Machado, and A. Sarkar. The critical probability for the frog model is not a monotonic function of the graph. Journal of Applied Probability, 41(1):292–298, March 2004. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/3215835.

Fabens:1970:LDM

[FN70] Augustus J. Fabens and Marcel F. Neuts. The limiting distribution of the maximum term in a sequence of random variables defined on a Markov chain. Journal of Applied Probability, 7(3): 754–760, December 1970. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3211952.

Fazekas:2019:SNE

[FNP19] István Fazekas, Csaba Noszály, and Attila Perecsényi. The N-star network evolution model. Journal of Applied Probability, 56(2):416–440, June 2019. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL https:/

/www.cambridge.org/core/journals/journal-of-applied-
probability/article/nstar-network-evolution-model/BBEF64BC1C7D8D0D16CEB1CA28A3A501.

Franco:2011:HLT

[FNV11] Tertuliano Franco, Adriana Neumann, and Glauco Valle. Hydrodynamic limit for a type of exclusion process with slow bonds in dimension d ≥ 2. Journal of Applied Probability, 48(2):333– 351, June 2011. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
23065889.

Follmer:1972:OSC

[Föl72] Hans Föllmer. Optimal stopping of constrained Brownian motion. Journal of Applied Probability, 9(3):557–571, September 1972. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3212325.

Foley:1986:SPD

[Fol86] Robert D. Foley. Stationary Poisson departure processes from non-stationary queues. Journal of Applied Probability, 23(1): 256–260, March 1986. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3214138.

Foster:1975:ORG

[Fos75] W. F. Foster. Optimal reserves growth under a Wiener process. Journal of Applied Probability, 12(3):457–465, September 1975. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3212860.

Fox:1965:SMF

[Fox65] Martin Fox. A simplified model for the formation, movement, and dissipation of fair weather Cumulus clouds. Journal of Applied Probability, 2(1):178–185, June 1965. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/3211883.

Foxall:2015:NRT

[Fox15] Eric Foxall. New results for the two-stage contact process. Journal of Applied Probability, 52(1):258–268, March 2015. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://projecteuclid.org/euclid.jap/1429282620.

Foxall:2016:SMC

[Fox16] Eric Foxall. The SEIS model, or, the contact process with a latent stage. Journal of Applied Probability, 53(3):783– 801, September 2016. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://projecteuclid.

org/euclid.jap/1476370776.

Foxall:2021:ETL

[Fox21] Eric Foxall. Extinction time of the logistic process. Journal of Applied Probability, 58(3):637–676, September 2021. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL https://www.cambridge.org/core/journals/journal-
of-applied-probability/article/extinction-time-of-the-
logistic-process/C823AF1805CB006A16B28DC095851075.

Folkman:1966:OPS

[FP66] Jon Folkman and Sidney Port. Optimal procedures for stochastically failing equipment. Journal of Applied Probability, 3(2): 521–537, December 1966. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3212136.

Fagiuoli:1994:PCC

[FP94] Enrico Fagiuoli and Franco Pellerey. Preservation of certain classes of life distributions under Poisson shock models. Journal of Applied Probability, 31(2):458–465, June 1994. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3215038.

Fagiuoli:1997:MIS

[FP97] Enrico Fagiuoli and Franco Pellerey. Moment inequalities for sums of DMRL random variables. Journal of Applied Probability, 34(2):525–535, June 1997. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3215391.

Ferreira:2005:LCOb

[FP05a] Fátima Ferreira and António Pacheco. Level-crossing ordering of semi-Markov processes and Markov chains. Journal of Applied Probability, 42(4):989–1002, December 2005. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/30040823.

Ferreira:2005:LCOa

[FP05b] Fátima Ferreira and António Pacheco. Level crossing ordering of skip-free-to-the-right continuous-time Markov chains. Journal of Applied Probability, 42(1):52–60, March 2005. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/30040769.

Faouzi:2022:DLD

[FPKM22] Tarik Faouzi, Emilio Porcu, Igor Kondrashuk, and Anatoliy Malyarenko. A deep look into the Dagum family of isotropic covariance functions. Journal of Applied Probability, 59(4):1026–1041, December 2022. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL https:/

/www.cambridge.org/core/journals/journal-of-applied-

probability/article/deep-look-into-the-dagum-family-
of-isotropic-covariance-functions/155F2A83A963782A7792F108B18D9502.

Fernandez-Ponce:1998:POD

[FPKMP98] J. M. Fernandez-Ponce, S. C. Kochar, and J. Muñoz-Perez. Partial orderings of distributions based on right-spread functions. Journal of Applied Probability, 35(1):221–228, March 1998. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/3215560.

Fontana:2019:ESP

[FPP19] Claudio Fontana, Markus Pelger, and Eckhard Platen. On the existence of sure profits via flash strategies. Journal of Applied Probability, 56(2):384–397, June 2019. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL https:/

/www.cambridge.org/core/journals/journal-of-applied-
probability/article/on-the-existence-of-sure-profits-
via-flash-strategies/B3FFE70881C9297726E552148003328D.

Falk:2020:SCM

[FPR20] Michael Falk, Simone A. Padoan, and Stefano Rizzelli. Strong convergence of multivariate maxima. Journal of Applied Probability, 57(1):314–331, March 2020. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL https:/

/www.cambridge.org/core/journals/journal-of-applied-
probability/article/strong-convergence-of-multivariate-
maxima/A99B1BC58FB247E8982E63E8534C77BB.

Flesch:2018:CSO

[FPS18] János Flesch, Arkadi Predtetchinski, and William Sudderth. Characterization and simplification of optimal strategies in positive stochastic games. Journal of Applied Probability, 55(3):728–741, September 2018. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL https:/

/www.cambridge.org/core/journals/journal-of-applied-
probability/article/characterization-and-simplification-
of-optimal-strategies-in-positive-stochastic-games/0D7328E54E9AE404F18FA8FD74D3D49B.

Falk:2019:CTI

[FPW19] Michael Falk, Simone A. Padoan, and Florian Wisheckel. Conditional tail independence in Archimedean copula models. Journal of Applied Probability, 56(3):858–869, September 2019. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL https://www.cambridge.org/core/journals/journal-
of-applied-probability/article/conditional-tail-independence-
in-archimedean-copula-models/CA1D26B414CA2596EE48BA68AD11455E.

Fujita:2004:PPD

[FPY04] T. Fujita, F. Petit, and M. Yor. Pricing path-dependent options in a Black–Scholes market from the distribution of homogeneous Brownian functionals. Journal of Applied Probability, 41(1):1– 18, March 2004. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3215811.

Flimmel:2020:LTU

[FPY20] Daniela Flimmel, Zbynek Pawlas, and J. E. Yukich. Limit theory for unbiased and consistent estimators of statistics of random tessellations. Journal of Applied Probability, 57(2):679–702, June 2020. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL https:/

/www.cambridge.org/core/journals/journal-of-applied-
probability/article/limit-theory-for-unbiased-and-consistent-
estimators-of-statistics-of-random-tessellations/62A19963DBBBF68462F56D94204F8C76.

Fredkin:1986:AMP

[FR86] Donald R. Fredkin and John A. Rice. On aggregated Markov processes. Journal of Applied Probability, 23(1):208–214, March 1986. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3214130.

Fralix:2010:NLT

[FR10] Brian H. Fralix and Germán Riaño. A new look at transient versions of Little’s law, and M/G/1 preemptive last-come– first-served queues. Journal of Applied Probability, 47(2):459– 473, June 2010. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
20778491.

Francois:2000:GIE

[Fra00] Olivier François. Geometric inequalities for the eigenvalues of concentrated Markov chains. Journal of Applied Probability, 37 (1):15–28, March 2000. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3215656.

Fralix:2006:FTC

[Fra06] Brian H. Fralix. Foster-type criteria for Markov chains on general spaces. Journal of Applied Probability, 43(4):1194–1200, December 2006. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
27595811.

Franke:2007:LPW

[Fra07] Brice Franke. A Lévy process whose jumps are dragged by a spherical dynamical system. Journal of Applied Probability, 44 (3):732–741, September 2007. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.

org/stable/27595878.

Freixas:2007:BOM

[Fre07] Josep Freixas. Bounds for Owen’s multilinear extension. Journal of Applied Probability, 44(4):852–864, December 2007. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/27595892.

Frostig:1993:OPM

[Fro93] Esther Frostig. Optimal policies for machine repairmen problems. Journal of Applied Probability, 30(3):703–715, September 1993. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3214776.

Frostig:2005:ETR

[Fro05] Esther Frostig. On the expected time to ruin and the expected dividends when dividends are paid while the surplus is above a constant barrier. Journal of Applied Probability, 42(3):595–607, September 2005. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
30040843.

Frostig:2015:MDL

[Fro15] Esther Frostig. The moments of the discounted loss and the discounted dividends for a spectrally negative Lévy risk process. Journal of Applied Probability, 52(3):665–687, September 2015. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://projecteuclid.org/euclid.jap/
1445543839.

Franco:2001:CIN

[FRR01] Manuel Franco, José M. Ruiz, and M. Carmen Ruiz. On closure of the IFR(2) and NBU(2) classes. Journal of Applied Probability, 38(1):235–241, March 2001. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3215755.

Fricker:1999:RCE

[FRT99] Christine Fricker, Philippe Robert, and Danielle Tibi. On the rates of convergence of Erlang’s model. Journal of Applied Probability, 36(4):1167–1184, December 1999. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/3215586.

Feder:1971:SDC

[FS71] Paul I. Feder and Thomas Stroud. Sequential decisions in the control of a space-ship (terminal cost proportional to magnitude of miss distance). Journal of Applied Probability, 8(2):285– 297, June 1971. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3211899.

Fava:1978:PLS

[FS78] N. A. Fava and L. A. Santaló. Plate and line segment processes. Journal of Applied Probability, 15(3):494–501, September 1978. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/3213112.

Faddy:1980:BVS

[FS80a] M. J. Faddy and I. H. Slorach. Bounds on the velocity of spread of infection for a spatially connected epidemic process. Journal of Applied Probability, 17(3):839–845, September 1980. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3212977.

Franken:1980:RAC

[FS80b] Peter Franken and Arnfried Streller. Reliability analysis of complex repairable systems by means of marked point processes. Journal of Applied Probability, 17(1):154–167, March 1980. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/3212933.

Frenz:1992:IPL

[FS92] Michael Frenz and Volker Schmidt. An insensitivity property of ladder height distributions. Journal of Applied Probability, 29 (3):616–624, September 1992. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.

org/stable/3214898.

Fosam:1997:VCD

[FS97] E. B. Fosam and D. N. Shanbhag. Variants of the Choquetdeny theorem with applications. Journal of Applied Probability, 34(1):101–106, March 1997. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3215179.

Frey:1998:GGF

[FS98] Rüdiger Frey and Daniel Sommer. The generalization of the Geske-formula for compound options to stochastic interest rates is not trivial — a note. Journal of Applied Probability, 35(2):501– 509, June 1998. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/

3215704. See [Ges79].

Finlay:2006:SIS

[FS06] Richard Finlay and Eugene Seneta. Stationary-increment Student and variance-gamma processes. Journal of Applied Probability, 43(2):441–453, June 2006. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/27595742. See correction [Ano06g].

Finlay:2007:GAT

[FS07] Richard Finlay and Eugene Seneta. A gamma activity time process with noninteger parameter and self-similar limit. Journal of Applied Probability, 44(4):950–959, December 2007. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/27595899.

Finlay:2014:RFP

[FS14] Richard Finlay and Eugene Seneta. Random fields with Pólya correlation structure. Journal of Applied Probability, 51(4):1037– 1050, December 2014. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://projecteuclid.

org/euclid.jap/1421763326.

Ferreira:2016:DBA

[FS16] Eva Ferreira and Winfried Stute. Dynamic binomials with an application to gender bias analysis. Journal of Applied Probability, 53(1):82–90, March 2016. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://projecteuclid.

org/euclid.jap/1457470560.

Foss:2017:LSJ

[FS17] S. Foss and A. L. Stolyar. Large-scale join-idle-queue system with general service times. Journal of Applied Probability, 54(4):995–1007, December 2017. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL https:/

/www.cambridge.org/core/journals/journal-of-applied-
probability/article/largescale-joinidlequeue-system-
with-general-service-times/67206ECCD5DE08D67B2F82FFBC420303.

Fan:2023:SNC

[FS23] Xiequan Fan and Qi-Man Shao. Self-normalized Cramér moderate deviations for a supercritical Galton–Watson process. Journal of Applied Probability, 60(4):1281–1292, December 2023. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL https://www.cambridge.org/core/journals/
journal-of-applied-probability/article/selfnormalized-
cramer-moderate-deviations-for-a-supercritical-galtonwatson-
process/0DE14C76DC19CAEEC08001097CAD8AD4.

Freund:2014:MCS

[FSJ14] Fabian Freund and Arno Siri-Jégousse. Minimal clade size in the Bolthausen–Sznitman coalescent. Journal of Applied Probability, 51(3):657–668, September 2014. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http:/

/projecteuclid.org/euclid.jap/1409932665.

Finlay:2012:IGA

[FSW12] Richard Finlay, Eugene Seneta, and Dingcheng Wang. An inverse gamma activity time process with noninteger parameters and a self-similar limit. Journal of Applied Probability, 49 (2):441–450, June 2012. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/41713778.

Federgruen:1978:OEA

[FT78] A. Federgruen and H. C. Tijms. The optimality equation in average cost denumerable state semi-Markov decision problems, recurrency conditions and algorithms. Journal of Applied Probability, 15(2):356–373, June 1978. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3213407.

Federgruen:1980:CSD

[FT80] A. Federgruen and H. C. Tijms. Computation of the stationary distribution of the queue size in an M/G/1 queueing system with variable service rate. Journal of Applied Probability, 17(2):515– 522, June 1980. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3213040.

Finkelstein:1990:LDN

[FTV90] Mark Finkelstein, Howard G. Tucker, and Jerry Alan Veeh. The limit distribution of the number of rare mutants. Journal of Applied Probability, 27(2):239–250, June 1990. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/3214643.

Fu:2001:DSS

[Fu01] James C. Fu. Distribution of the scan statistic for a sequence of bistate trials. Journal of Applied Probability, 38(4):908–916, December 2001. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3215773.

Fujii:2013:NEC

[Fuj13] Takayuki Fujii. Nonparametric estimation for a class of piecewise-deterministic Markov processes. Journal of Applied Probability, 50(4):931–942, December 2013. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

projecteuclid.org/euclid.jap/1389370091.

Fodor:2018:VER

[FV18] Ferenc Fodor and Viktor Vígh. Variance estimates for random disc-polygons in smooth convex discs. Journal of Applied Probability, 55(4):1143–1157, December 2018. CODEN JPRBAM.

ISSN 0021-9002 (print), 1475-6072 (electronic). URL https:/

/www.cambridge.org/core/journals/journal-of-applied-
probability/article/variance-estimates-for-random-discpolygons-
in-smooth-convex-discs/1C78CAB92958ADAB7205CE56170175DB.

Fralix:2013:FIR

[FvLB13] Brian H. Fralix, Johan S. H. van Leeuwaarden, and Onno J. Boxma. Factorization identities for reflected processes, with applications. Journal of Applied Probability, 50(3):632–653, September 2013. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://projecteuclid.org/
euclid.jap/1378401227.

Flesch:2019:EMB

[FVZ19] János Flesch, Dries Vermeulen, and Anna Zseleva. On the equivalence of mixed and behavior strategies in finitely additive decision problems. Journal of Applied Probability, 56 (3):810–829, September 2019. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL https:/

/www.cambridge.org/core/journals/journal-of-applied-
probability/article/on-the-equivalence-of-mixed-and-
behavior-strategies-in-finitely-additive-decision-problems/
58F01DEEF37D2E9F5F7DBD4B005EE324.

Fay:1991:AOI

[FW91] N. A. Fay and J. C. Walrand. On approximately optimal index strategies for generalised arm problems. Journal of Applied Probability, 28(3):602–612, September 1991. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/3214495.

Feng:2007:CID

[FW07] Shui Feng and Feng-Yu Wang. A class of infinite-dimensional diffusion processes with connection to population genetics. Journal of Applied Probability, 44(4):938–949, December 2007. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/27595898.

Fu:2010:LNB

[FW10] James C. Fu and Tung-Lung Wu. Linear and nonlinear boundary crossing probabilities for Brownian motion and related processes. Journal of Applied Probability, 47(4):1058–1071, December 2010. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/29777426.

Fu:2016:BCP

[FW16] James C. Fu and Tung-Lung Wu. Boundary crossing probabilities for high-dimensional Brownian motion. Journal of Applied Probability, 53(2):543–553, June 2016. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

projecteuclid.org/euclid.jap/1466172873.

Fendick:2021:QPD

[FW21] Kerry Fendick and Ward Whitt. Queues with path-dependent arrival processes. Journal of Applied Probability, 58(2):484–504, June 2021. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-

6072 (electronic). URL https://www.cambridge.org/core/
journals/journal-of-applied-probability/article/queues-
with-pathdependent-arrival-processes/283AC69B2FE9EA968E9A54D2CDF11D23.

Fu:2003:ELD

[FWL03] James C. Fu, Liqun Wang, and W. Y. Wendy Lou. On exact and large deviation approximation for the distribution of the longest run in a sequence of two-state Markov dependent trials. Journal of Applied Probability, 40(2):346–360, June 2003. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3215795.

Fu:2012:CDC

[FWL12] James C. Fu, Tung-Lung Wu, and W. Y. Wendy Lou. Continuous, discrete, and conditional scan statistics. Journal of Applied Probability, 49(1):199–209, March 2012. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http:/

/www.jstor.org/stable/41713760.

Fatheddin:2016:MDP

[FX16] Parisa Fatheddin and Jie Xiong. Moderate deviation principle for a class of stochastic partial differential equations. Journal of Applied Probability, 53(1):279–292, March 2016. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://projecteuclid.org/euclid.jap/1457470574.

Feigin:1982:EVP

[FY82] Paul D. Feigin and Emmanuel Yashchin. Extreme-value properties of the explosion-time distribution in a pure birth process. Journal of Applied Probability, 19(3):500–509, September 1982.

CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/3213509. See

correction [Ano86g].

Foley:1983:MCI

[FY83] R. D. Foley and M. Yadin. On the moments of the cumulative idle time in an M/D/∞ queue. Journal of Applied Probability, 20(2):445–447, June 1983. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3213822.

Florens-Zmirou:1993:EDC

[FZ93] Danielle Florens-Zmirou. On estimating the diffusion coefficient from discrete observations. Journal of Applied Probability, 30(4): 790–804, December 1993. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3214513.

Falk:2017:GML

[FZ17] Michael Falk and Maximilian Zott. On generalized max-linear models in max-stable random fields. Journal of Applied Probability, 54(3):797–810, September 2017. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL https:/

/www.cambridge.org/core/journals/journal-of-applied-
probability/article/on-generalized-maxlinear-models-
in-maxstable-random-fields/EB09CB558EBBAB8EBB343280651DE20B.

Feng:2013:NMR

[FZL13] Xiuying Feng, Shuhong Zhang, and Xiaohu Li. A note on the mixture representation of the conditional residual lifetime of a coherent system. Journal of Applied Probability, 50(2):475–485, June 2013. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-

6072 (electronic). URL http://projecteuclid.org/euclid.
jap/1371648955.

Gadrich:1993:SMN

[GA93] Tamar Gadrich and Robert J. Adler. Slepian models for nonstationary Gaussian processes. Journal of Applied Probability, 30(1):98–111, March 1993. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3214624.

Gaasemyr:1999:ADD

[GA99] J. Gåsemyr and Terje Aven. Asymptotic distributions for downtimes of monotone systems. Journal of Applied Probability, 36 (3):814–824, September 1999. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.

org/stable/3215443.

Gapeev:2022:DOS

[GA22] Pavel V. Gapeev and Hessah Al Motairi. Discounted optimal stopping problems in first-passage time models with random thresholds. Journal of Applied Probability, 59(3): 714–733, September 2022. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL https:/

/www.cambridge.org/core/journals/journal-of-applied-
probability/article/discounted-optimal-stopping-problems-
in-firstpassage-time-models-with-random-thresholds/3B7B5A7F2F2DE6289456C12198B396D4.

Gaal:2020:LRO

[Gaá20] Alexisz Tamás Gaál. Long-range orientational order of random near-lattice hard sphere and hard disk processes. Journal of Applied Probability, 57(2):559–577, June 2020. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL https:/

/www.cambridge.org/core/journals/journal-of-applied-
probability/article/longrange-orientational-order-of-
random-nearlattice-hard-sphere-and-hard-disk-processes/
7CCBCE1E3FEBA8FC822EBD4056DDB763.

Goliforushani:2012:RIT

[GAB12] S. Goliforushani, M. Asadi, and N. Balakrishnan. On the residual and inactivity times of the components of used coherent systems. Journal of Applied Probability, 49(2):385–404, June 2012. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/41713774.

Gabrysch:2016:CDR

[Gab16a] Katja Gabrysch. Convergence of directed random graphs to the Poisson-weighted infinite tree. Journal of Applied Probability, 53(2):463–474, June 2016. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://projecteuclid.

org/euclid.jap/1466172867.

Gabrysch:2016:DSV

[Gab16b] Katja Gabrysch. Distribution of the smallest visited point in a greedy walk on the line. Journal of Applied Probability, 53(3): 880–887, September 2016. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://projecteuclid.

org/euclid.jap/1476370782.

Gadag:1989:NMM

[Gad89] V. G. Gadag. A note on the multitype Markov branching process. Journal of Applied Probability, 26(3):631–636, September 1989. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3214420.

Glazebrook:2004:OAS

[GADL04] K. D. Glazebrook, P. S. Ansell, R. T. Dunn, and R. R. Lumley. On the optimal allocation of service to impatient tasks. Journal of Applied Probability, 41(1):51–72, March 2004. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/3215814.

Gagnon:2017:SIC

[Gag17] Gregory Gagnon. Stochastic impulse control of exchange rates with Freidlin–Wentzell perturbations. Journal of Applied Probability, 54(1):23–41, March 2017. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL https:/

/www.cambridge.org/core/journals/journal-of-applied-
probability/article/stochastic-impulse-control-of-exchange-
rates-with-freidlinwentzell-perturbations/2CBEAB4C3FB430B944D7666B3358FDA7.

Gallot:1966:AAP

[Gal66a] S. Gallot. Asymptotic absorption probabilities for a Poisson process. Journal of Applied Probability, 3(2):445–452, December 1966. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3212130.

Gallot:1966:BMN

[Gal66b] S. Gallot. A bound for the maximum of a number of random variables. Journal of Applied Probability, 3(2):556–558, December 1966. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3212139.

Galambos:1971:MFD

[Gal71a] János Galambos. On a model for a fair distribution of gifts. Journal of Applied Probability, 8(4):681–690, December 1971. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/3212232.

Galley:1971:SMF

[Gal71b] S. J. Galley. A stochastic model for the formation of a planetary system. Journal of Applied Probability, 8(3):470–479, September 1971. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3212171.

Galambos:1972:CCP

[Gal72] János Galambos. Characterization of certain populations by independence of order statistics. Journal of Applied Probability, 9(1):224–230, March 1972. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3212654.

Galambos:1973:DMR

[Gal73] János Galambos. The distribution of the maximum of a random number of random variables with applications. Journal of Applied Probability, 10(1):122–129, March 1973. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3212500.

Galambos:1974:LTA

[Gal74] János Galambos. A limit theorem with applications in order statistics. Journal of Applied Probability, 11(1):219–222, March 1974. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3212603.

Gallot:1993:AFP

[Gal93] S. F. L. Gallot. Absorption and first-passage times for a compound Poisson process in a general upper boundary. Journal of Applied Probability, 30(4):835–850, December 1993. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3214516.

Gani:1965:SMB

[Gan65] J. Gani. Stochastic models for bacteriophage. Journal of Applied Probability, 2(2):225–268, December 1965. CODEN JPRBAM.

ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/3212193.

Gani:1971:OSE

[Gan71] J. Gani. Obituary: Sir Edward Collingwood. Journal of Applied Probability, 8(1):211+213–214, March 1971. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/3211852.

Gani:1982:PGF

[Gan82] J. Gani. On the probability generating function of the sum of Markov Bernoulli random variables. Journal of Applied Probability, 19A:321–326, 1982. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3213571.

Gani:1984:ODN

[Gan84a] J. Gani. Obituary: Dr. Norma Ruth McArthur. Journal of Applied Probability, 21(3):669–672, September 1984. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3213629.

Gani:1984:ORB

[Gan84b] J. Gani. Obituary: Richard Bellman. Journal of Applied Probability, 21(4):935–936, December 1984. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3213711.

Gani:1986:P

[Gan86] J. Gani. Preface. Journal of Applied Probability, 23A:v–vi, 1986. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/3214337.

Gani:1988:AAP

[Gan88a] J. Gani. Adventures in applied probability. Journal of Applied Probability, 25A:3–23, 1988. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.

org/stable/3214143.

Gani:1988:CRW

[Gan88b] J. Gani. A correlated random walk for the transport and sedimentation of particles. Journal of Applied Probability, 25A:

335–346, 1988. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3214167.

Gani:1994:SDE

[Gan94] J. Gani. A simple deterministic epidemic in several related growing communities. Journal of Applied Probability, 31A:17–25, 1994. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3214945.

Gani:1998:OJE

[Gan98] J. Gani. Obituary: José Enrique Moyal. Journal of Applied Probability, 35(4):1012–1017, December 1998. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3215538.

Gani:2002:OMS

[Gan02] Joe Gani. Obituary: Maurice Stevenson Bartlett. Journal of Applied Probability, 39(3):664–670, September 2002. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3216089.

Gani:2004:RAU

[Gan04a] J. Gani. Random-allocation and urn models. Journal of Applied Probability, 41A:313–320, 2004. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.

org/stable/3215985.

Gani:2004:OMH

[Gan04b] Joe Gani. Obituary: Mavis Hitchcock. Journal of Applied Probability, 41(2):597–600, June 2004. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3216041.

Gani:2010:FSY

[Gan10] J. Gani. Forty-six years of the Applied Probability Trust. Journal of Applied Probability, 47(1):??, March 2010. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/25662497.

Gao:2004:MLR

[Gao04] Jiti Gao. Modelling long-range-dependent Gaussian processes with application in continuous-time financial models. Journal of Applied Probability, 41(2):467–482, June 2004. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3216029.

Gapeev:2007:DOS

[Gap07] Pavel V. Gapeev. Discounted optimal stopping for maxima of some jump-diffusion processes. Journal of Applied Probability, 44(3):713–731, September 2007. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.

org/stable/27595877.

Garwood:1974:VRF

[Gar74] F. Garwood. The vectorial representation of the frequency of encounters of freely flowing vehicles. Journal of Applied Probability, 11(4):797–808, December 1974. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3212562.

Gastwirth:1964:TTS

[Gas64] Joseph L. Gastwirth. On a telephone traffic system with several kinds of service distributions. Journal of Applied Probability, 1(1):77–84, June 1964. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3212062.

Gaasemyr:1998:IAM

[Gås98] Jørund Gåsemyr. On identifiability in the autopsy model of reliability systems of lifemonitored and conditionally lifemonitored components. Journal of Applied Probability, 35(2):427–437, June 1998. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3215696.

Gonzalez-Arevalo:2003:BCL

[GAS03] Bárbara González-Arévalo and Gennady Samorodnitsky. Buffer content of a leaky-bucket system with long-range dependent input traffic. Journal of Applied Probability, 40(3):581–601, September 2003. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3215937.

Gates:1982:RTQ

[Gat82a] J. Gates. Recognition of triangles and quadrilaterals by chord length distribution. Journal of Applied Probability, 19(4):873– 879, December 1982. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3213843.

Gates:1982:NIR

[Gat82b] John Gates. The number of intersections of random chords to a circle: Third and fourth moments. Journal of Applied Probability, 19(2):355–372, June 1982. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3213487.

Gates:1984:BPC

[Gat84a] John Gates. Bounds for the probability of complete intersection of random chords in a circle. Journal of Applied Probability, 21 (2):419–424, June 1984. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3213651.

Gates:1984:NVS

[Gat84b] John Gates. A note on the volume of small random simplexes. Journal of Applied Probability, 21(3):651–653, September 1984. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/3213626.

Gates:1987:SPC

[Gat87] J. Gates. Some properties of chord length distributions. Journal of Applied Probability, 24(4):863–874, December 1987. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3214211.

Gaunt:2022:BCS

[Gau22] Robert E. Gaunt. Bounds for the chi-square approximation of the power divergence family of statistics. Journal of Applied Probability, 59(4):1059–1080, December 2022. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL https://www.cambridge.org/core/journals/journal-
of-applied-probability/article/bounds-for-the-chisquare-
approximation-of-the-power-divergence-family-of-statistics/
8E138125CCEB92DB51FECD72A91FFBF3.

Gaver:1968:DAM

[Gav68] D. P. Gaver, Jr. Diffusion approximations and models for certain congestion problems. Journal of Applied Probability, 5(3): 607–623, December 1968. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3211925.

Gaver:1969:HDR

[Gav69] Donald P. Gaver, Jr. Highway delays resulting from flowstopping incidents. Journal of Applied Probability, 6(1):137– 153, April 1969. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3212281.

Gaver:1976:RRM

[Gav76] Donald P. Gaver. Random record models. Journal of Applied Probability, 13(3):538–547, September 1976. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/3212473.

Gay:2004:PRW

[Gay04] Roger Gay. Pricing risk when distributions are fat tailed. Journal of Applied Probability, 41A:157–175, 2004. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/3215975.

Garcia:2002:TAS

[GBG02] Jean-Marie Garcia, Olivier Brun, and David Gauchard. Transient analytical solution of M/D/1/N queues. Journal of Applied Probability, 39(4):853–864, December 2002. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/3216008.

Gopalsamy:1975:CPD

[GBR75] K. Gopalsamy and A. T. Bharucha-Reid. On a class of parabolic differential equations driven by stochastic point processes. Journal of Applied Probability, 12(1):98–106, March 1975. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3212411.

Glazebrook:1992:CRH

[GBW92] K. D. Glazebrook, Michael P. Bailey, and Lyn R. Whitaker. Cost rate heuristics for semi-Markov decision processes. Journal of Applied Probability, 29(3):633–644, September 1992. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3214900.

Gibson:1971:TSS

[GC71a] A. E. Gibson and B. W. Conolly. On a three-state sojourn time problem. Journal of Applied Probability, 8(4):716–723, December 1971. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3212235.

Gibson:1971:CUL

[GC71b] A. E. Gibson and B. W. Conolly. On certain unrestricted, linear, unit step, continuous time random walks. Journal of Applied Probability, 8(2):374–380, June 1971. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3211907.

Gazdag:1974:NSF

[GC74] Jenö Gazdag and José Canosa. Numerical solution of Fisher’s equation. Journal of Applied Probability, 11(3):445–457, September 1974. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3212689.

Grigorescu:2007:OSV

[GCS07] Ilie Grigorescu, Robert Chen, and Larry Shepp. Optimal strategy for the Vardi Casino with interest payments. Journal of Applied Probability, 44(1):199–211, March 2007. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/27595831.

Ganuza:1974:MSA

[GD74] Edgar Z. Ganuza and S. D. Durham. Mean-square and almostsure convergence of supercritical age-dependent branching processes. Journal of Applied Probability, 11(4):678–686, December 1974. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3212551.

Grunbaum:2018:SFD

[GdlI18] F. Alberto Grünbaum and Manuel D. de la Iglesia. Stochastic LU factorizations, Darboux transformations and urn models. Journal of Applied Probability, 55(3):862–886, September 2018. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL https://www.cambridge.org/core/journals/
journal-of-applied-probability/article/stochastic-lu-
factorizations-darboux-transformations-and-urn-models/
4F32741651B8486FA0A69955B18AE310.

Gomes:2004:JEA

[GDP04] M. Ivette Gomes, Laurens De Haan, and Dinis Pestana. Joint exceedances of the ARCH process. Journal of Applied Probability, 41(3):919–926, September 2004. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/4141362. See correction [Iks04].

Ghosh:1982:SML

[GE82] Malay Ghosh and Nader Ebrahimi. Shock models leading to increasing failure rate and decreasing mean residual life survival. Journal of Applied Probability, 19(1):158–166, March 1982. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/3213925.

Geiger:1999:ENP

[Gei99] Jochen Geiger. Elementary new proofs of classical limit theorems for Galton–Watson processes. Journal of Applied Probability, 36 (2):301–309, June 1999. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3215457.

Geller:1977:LDO

[Gel77] Nancy L. Geller. On limit distributions for one- and two-sample Kolmogorov–Smirnov type statistics. Journal of Applied Probability, 14(3):538–547, September 1977. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3213456.

Gelenbe:1991:PFQ

[Gel91] Erol Gelenbe. Product-form queueing networks with negative and positive customers. Journal of Applied Probability, 28(3): 656–663, September 1991. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3214499.

Gelenbe:1993:NTC

[Gel93] Erol Gelenbe. G-networks with triggered customer movement. Journal of Applied Probability, 30(3):742–748, September 1993. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/3214781.

Genadot:2023:ASF

[Gén23] Alexandre Génadot. Averaging for slow-fast piecewise deterministic Markov processes with an attractive boundary. Journal of Applied Probability, 60(4):1439–1468, December 2023. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL https://www.cambridge.org/core/journals/journal-
of-applied-probability/article/averaging-for-slowfast-
piecewise-deterministic-markov-processes-with-an-attractive-
boundary/F7779E98A35CECC2FB4598CD2919A911.

George:1987:SRP

[Geo87] Edward I. George. Sampling random polygons. Journal of Applied Probability, 24(3):557–573, September 1987. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3214089.

Gertsbach:1976:SOC

[Ger76] I. Gertsbach. Sufficient optimality conditions for control-limit policy in a semi-Markov process. Journal of Applied Probability, 13(2):400–406, June 1976. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3212848.

Gertsbakh:1977:EPR

[Ger77] I. B. Gertsbakh. Epidemic process on a random graph: Some preliminary results. Journal of Applied Probability, 14(3):427– 438, September 1977. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3213446.

Gerrard:1983:RCS

[Ger83] Russell Gerrard. Regularity conditions for semi-Markov and Markov chains in continuous time. Journal of Applied Probability, 20(3):505–512, September 1983. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3213887.

Gerontidis:1990:CAN

[Ger90] Ioannis I. Gerontidis. On certain aspects of non-homogeneous Markov systems in continuous time. Journal of Applied Probability, 27(3):530–544, September 1990. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3214539.

Gerontidis:1991:PSE

[Ger91] Ioannis I. Gerontidis. Periodic strong ergodicity in nonhomogeneous Markov systems. Journal of Applied Probability, 28(1):58–73, March 1991. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/ stable/3214740. See correction [Ano91g].

Geman:1995:CNC

[GER95] Hélyette Geman, Nicole El Karoui, and Jean-Charles Rochet. Changes of numéraire. Changes of probability measure and option pricing. Journal of Applied Probability, 32(2):443–458, June 1995. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3215299.

Gerhold:2011:HWD

[Ger11] Stefan Gerhold. The Hartman–Watson distribution revisited: Asymptotics for pricing Asian options. Journal of Applied Probability, 48(3):892–899, September 2011. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/23065946.

Gerencser:2023:ANR

[Ger23] Balázs Gerencsér. Analysis of a non-reversible Markov chain speedup by a single edge. Journal of Applied Probability, 60(2):642–660, June 2023. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL https:/

/www.cambridge.org/core/journals/journal-of-applied-
probability/article/analysis-of-a-nonreversible-markov-
chain-speedup-by-a-single-edge/B70F02AFD69E88F65E046E3F25FE27C9.

Geske:1979:VCO

[Ges79] R. Geske. The valuation of compound options. Journal of Financial Economics, 7(??):63–81, ???? 1979. CODEN JFECDT. ISSN 0304-405X (print), 1879-2774 (electronic). See comment [FS98].

Ghitany:1995:CGC

[GESK95] M. E. Ghitany, M. A. El-Saidi, and Z. Khalil. Characterization of a general class of life-testing models. Journal of Applied Probability, 32(2):548–553, June 1995. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/3215308.

Galbraith:1974:ISP

[GG74] R. F. Galbraith and J. I. Galbraith. On the inverses of some patterned matrices arising in the theory of stationary time series. Journal of Applied Probability, 11(1):63–71, March 1974. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/3212583.

Gittins:1977:BMS

[GG77a] J. C. Gittins and K. D. Glazebrook. On Bayesian models in stochastic scheduling. Journal of Applied Probability, 14(3): 556–565, September 1977. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3213458.

Guess:1977:DAL

[GG77b] Harry A. Guess and John H. Gillespie. Diffusion approximations to linear stochastic difference equations with stationary coefficients. Journal of Applied Probability, 14(1):58–74, March 1977. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3213260.

Gzyl:1978:EED

[GG78] Cristina Gzyl and Henryk Gzyl. On entrance-exit distributions of Markov processes. Journal of Applied Probability, 15(1):78– 86, March 1978. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3213238.

Gupta:1986:CPP

[GG86] Pushpa Lata Gupta and Ramesh C. Gupta. A characterization of the Poisson process. Journal of Applied Probability, 23(1):233– 235, March 1986. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3214134.

Garey:1987:NCS

[GG87] L. E. Garey and R. D. Gupta. A note on continuous search algorithms. Journal of Applied Probability, 24(1):277–280, March 1987. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3214081.

Gadag:1991:PPA

[GG91a] V. G. Gadag and R. P. Gupta. On the properties of processes associated with a Markov branching process. Journal of Applied Probability, 28(3):520–528, September 1991. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/3214488.

Glasserman:1991:TCT

[GG91b] Paul Glasserman and Wei-Bo Gong. Time-changing and truncating K-capacity queues from one K to another. Journal of Applied Probability, 28(3):647–655, September 1991. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3214498.

Galambos:1994:P

[GG94a] J. Galambos and J. Gani. Preface. Journal of Applied Probability, 31A:v, 1994. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3214941.

Gani:1994:LTA

[GG94b] J. Gani and J. Galambos. Lajos Takács: An appreciation. Journal of Applied Probability, 31A:xi–xiii, 1994. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/3214942.

Glazebrook:1995:TIG

[GG95] K. D. Glazebrook and S. Greatrix. On transforming an index for generalised bandit problems. Journal of Applied Probability, 32 (1):168–182, March 1995. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3214927.

Glynn:2000:BLC

[GG00] John E. Glynn and Peter W. Glynn. On the behaviour of a long cascade of linear reservoirs. Journal of Applied Probability, 37 (2):417–428, June 2000. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3215718.

Gupta:2010:REB

[GG10] Ramesh C. Gupta and Rameshwar D. Gupta. Random effect bivariate survival models and stochastic comparisons. Journal of Applied Probability, 47(2):426–440, June 2010. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/20778489.

Gnoatto:2014:ELT

[GG14] Alessandro Gnoatto and Martino Grasselli. The explicit Laplace transform for the Wishart process. Journal of Applied Probability, 51(3):640–656, September 2014. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http:/

/projecteuclid.org/euclid.jap/1409932664.

Gelenbe:1991:QNA

[GGS91] Erol Gelenbe, Peter Glynn, and Karl Sigman. Queues with negative arrivals. Journal of Applied Probability, 28(1):245– 250, March 1991. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3214756.

Geske:1995:CPA

[GGS^{+}95] Mark X. Geske, Anant P. Godbole, Andrew A. Schaffner, Allison M. Skolnick, and Garrick L. Wallstrom. Compound Poisson approximations for word patterns under Markovian hypotheses. Journal of Applied Probability, 32(4):877–892, December 1995. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/3215201.

Gotze:2019:RAS

[GGZ19] Friedrich Götze, Anna Gusakova, and Dmitry Zaporozhets. Random affine simplexes. Journal of Applied Probability, 56(1):39–51, March 2019. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL https:/

/www.cambridge.org/core/journals/journal-of-applied-
probability/article/random-affine-simplexes/369BF3F5B32CF1D6BB5A58A73CFF51E3.

Gani:1982:P

[GH82] J. Gani and E. J. Hannan. Preface. Journal of Applied Probability, 19A:??, 1982. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3213544.

Grassmann:1990:EDB

[GH90] Winfried K. Grassmann and Daniel P. Heyman. Equilibrium distribution of block-structured Markov chains with repeating rows. Journal of Applied Probability, 27(3):557–576, September 1990. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3214541.

Gani:1992:RMH

[GH92] J. Gani and C. C. Heyde. On the retirement of mavis Hitchcock. Journal of Applied Probability, 29(2):484–486, June 1992. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/3214588.

Galambos:1994:USC

[GH94] Janos Galambos and Charles Hagwood. An unreliable server characterization of the exponential distribution. Journal of Applied Probability, 31(1):274–279, March 1994. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3215255.

Grubel:2009:GDR

[GH09] Rudolf Grübel and Paweł Hitczenko. Gaps in discrete random samples. Journal of Applied Probability, 46(4):1038–1051, December 2009. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
25662479.

Ganti:2010:LTC

[GH10] Radha Krishna Ganti and Martin Haenggi. Limit of the transport capacity of a dense wireless network. Journal of Applied Probability, 47(3):886–892, September 2010. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/20778523.

Grubel:2013:PDR

[GH13] Rudolf Grübel and Paweł Hitczenko. Pruned discrete random samples. Journal of Applied Probability, 50(2):542–556, June 2013. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://projecteuclid.org/euclid.jap/
1371648960.

Gudmundsson:2014:MCM

[GH14] Thorbjörn Gudmundsson and Henrik Hult. Markov chain Monte Carlo for computing rare-event probabilities for a heavy-tailed random walk. Journal of Applied Probability, 51(2):359–376, June 2014. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-

6072 (electronic). URL http://projecteuclid.org/euclid.
jap/1402578630.

Grosjean:2016:CPB

[GH16] Nicolas Grosjean and Thierry Huillet. On a coalescence process and its branching genealogy. Journal of Applied Probability, 53(4):1156–1165, December 2016. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http:/

/projecteuclid.org/euclid.jap/1481132843.

Guo:2021:RSA

[GH21] Xin Guo and Yonghui Huang. Risk-sensitive average continuoustime Markov decision processes with unbounded transition and cost rates. Journal of Applied Probability, 58(2):523–550, June 2021. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-

6072 (electronic). URL https://www.cambridge.org/core/
journals/journal-of-applied-probability/article/risksensitive-
average-continuoustime-markov-decision-processes-with-
unbounded-transition-and-cost-rates/FAA72A0D80967207D5C808E5060B242F.

Ghahramani:1986:FMP

[Gha86] Saeed Ghahramani. Finiteness of moments of partial busy periods for M/G/C queues. Journal of Applied Probability, 23 (1):261–264, March 1986. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3214139.

Ghahramani:1989:FMV

[Gha89] Saeed Ghahramani. Finiteness of moments of virtual work for GI/G/c queues. Journal of Applied Probability, 26(2):423– 425, June 1989. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3214050.

Ghahramani:1990:RFB

[Gha90] Saeed Ghahramani. On remaining full busy periods of GI/G/c queues and their relation to stationary point processes. Journal of Applied Probability, 27(1):232–236, March 1990. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3214612.

Gijbels:2004:IBE

[GHK04] Irène Gijbels, Peter Hall, and Aloïs Kneip. Interval and band estimation for curves with jumps. Journal of Applied Probability, 41A:65–79, 2004. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3215969.

Gross:1971:SIM

[GHL71] Donald Gross, Carl M. Harris, and James A. Lechner. Stochastic inventory models with bulk demand and state-dependent leadtimes. Journal of Applied Probability, 8(3):521–534, September 1971. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3212175.

Guo:2003:ZSG

[GHL03] Xianping Guo and Onésimo Hernández-Lerma. Zero-sum games for continuous-time Markov chains with unbounded transition and average payoff rates. Journal of Applied Probability, 40 (2):327–345, June 2003. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3215794.

Guo:2005:NSG

[GHL05] Xianping Guo and Onésimo Hernández-Lerma. Nonzero-sum games for continuous-time Markov chains with unbounded discounted payoffs. Journal of Applied Probability, 42(2):303– 320, June 2005. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
30040794.

Gorostiza:1995:PCS

[GHW95] Luis G. Gorostiza, Kenneth J. Hochberg, and Anton Wakolbinger. Persistence of a critical super-2 process. Journal of Applied Probability, 32(2):534–540, June 1995. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/3215306.

Gyorfi:2019:LDM

[GHW19] László Györfi, Norbert Henze, and Harro Walk. The limit distribution of the maximum probability nearest-neighbour ball. Journal of Applied Probability, 56(2):574–589, June 2019. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL https://www.cambridge.org/core/journals/journal-
of-applied-probability/article/limit-distribution-of-
the-maximum-probability-nearestneighbour-ball/D88CEE916CC915ADDB6B468381CD6185.

Gouet:2022:AOU

[GHW22] Raul Gouet, Paweł Hitczenko, and Jacek Wesołowski. Asymptotics of the overflow in urn models. Journal of Applied Probability, 59(3):797–824, September 2022. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL https:/

/www.cambridge.org/core/journals/journal-of-applied-
probability/article/asymptotics-of-the-overflow-in-urn-
models/035FA2EE35C38F44E88E8EAA30A7936F.

Gnedin:2011:MRS

[GI11] Alexander Gnedin and Alexander Iksanov. Moments of random sums and Robbins’ problem of optimal stopping. Journal of Applied Probability, 48(4):1197–1199, December 2011. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/23066452.

Gilson:1967:SRP

[Gil67] J. G. Gilson. Statistical relativistic particle. Journal of Applied Probability, 4(2):389–396, August 1967. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http:/

/www.jstor.org/stable/3212032.

Gilson:1968:FIP

[Gil68] J. G. Gilson. Feynman integral and phase space probability. Journal of Applied Probability, 5(2):375–386, August 1968. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/3212259.

Gilchrist:1976:IAT

[Gil76] Robert Gilchrist. The inter-arrival times of accepted customers in an M/G/1 queue with finite capacity. Journal of Applied Probability, 13(3):633–638, September 1976. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/3212488.

Gilbert:1977:ED

[Gil77] E. N. Gilbert. Estimators for distances. Journal of Applied Probability, 14(2):260–271, June 1977. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3212997.

Gilat:1987:BOO

[Gil87] David Gilat. On the best order of observation in optimal stopping problems. Journal of Applied Probability, 24(3):773–778, September 1987. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3214107.

Gnedin:2008:AEC

[GIM08] Alexander Gnedin, Alex Iksanov, and Martin Möhle. On asymptotics of exchangeable coalescents with multiple collisions. Journal of Applied Probability, 45(4):1186–1195, December 2008. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/27596017.

Gnedin:2011:CDC

[GIM11] Alexander Gnedin, Alexander Iksanov, and Alexander Marynych. On Λ-coalescents with dust component. Journal of Applied Probability, 48(4):1133–1151, December 2011. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/23066447.

Girling:1982:AVA

[Gir82] A. J. Girling. Approximate variances associated with random configurations of hard spheres. Journal of Applied Probability, 19(3):588–596, September 1982. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.

org/stable/3213516.

Giroux:1986:ARN

[Gir86] Gaston Giroux. Asymptotic results for non-linear processes of the McKean tagged-molecule type. Journal of Applied Probability, 23(1):42–51, March 1986. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3214115.

Gittins:1972:SPS

[Git72] J. C. Gittins. Some problems of stochastic resource allocation. Journal of Applied Probability, 9(2):360–369, June 1972. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/3212804.

Gittins:1973:ORS

[Git73] J. C. Gittins. Observations on a result in sequential analysis by P. Whittle. Journal of Applied Probability, 10(2):456–460, June 1973. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3212364.

Gittins:1974:RAS

[Git74] J. C. Gittins. Resource allocation in speculative chemical research. Journal of Applied Probability, 11(2):255–265, June 1974. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/3212747.

Gittins:1981:MSJ

[Git81] J. C. Gittins. Multiserver scheduling of jobs with increasing completion rates. Journal of Applied Probability, 18(1):321– 324, March 1981. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3213196.

Gani:1972:CGS

[GJ72] J. Gani and D. Jerwood. The cost of a general stochastic epidemic. Journal of Applied Probability, 9(2):257–269, June 1972. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/3212797.

Gaver:1978:NHP

[GJ78] Donald P. Gaver and Patricia A. Jacobs. Non-homogeneously paced random records and associated extremal processes. Journal of Applied Probability, 15(3):552–559, September 1978. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/3213118.

Gilbert:1980:AJC

[GJ80] Linda P. Gilbert and A. M. Johnson. An application of the Jordan canonical form to the epidemic problem. Journal of Applied Probability, 17(2):313–323, June 1980. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/3213021.

Glaz:1988:BCM

[GJ88] J. Glaz and B. Johnson. Boundary crossing for moving sums. Journal of Applied Probability, 25(1):81–88, March 1988. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/3214235.

Gerhold:2021:PLD

[GJP^{+}21] Stefan Gerhold, Antoine Jacquier, Mikko Pakkanen, Henry Stone, and Thomas Wagenhofer. Pathwise large deviations for the rough Bergomi model: Corrigendum. Journal of Applied Probability, 58(3):849–850, September 2021. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL https:/

/www.cambridge.org/core/journals/journal-of-applied-
probability/article/pathwise-large-deviations-for-the-
rough-bergomi-model-corrigendum/E852AD3A74D46F2D94BEE077E8A60183.

See [JPS18].

Ge:2006:REP

[GJQ06] Hao Ge, Da-Quan Jiang, and Min Qian. Reversibility and entropy production of inhomogeneous Markov chains. Journal of Applied Probability, 43(4):1028–1043, December 2006. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/27595797.

Gard:1976:SDE

[GK76] T. C. Gard and D. Kannan. On a stochastic differential equation modeling of prey-predator evolution. Journal of Applied Probability, 13(3):429–443, September 1976. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/3212463.

Gawel:1996:IGW

[GK96] Boleslaw Gawel and Marek Kimmel. The iterated Galton– Watson process. Journal of Applied Probability, 33(4):949–959, December 1996. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3214976.

Georgii:1997:SCP

[GK97] Hans-Otto Georgii and Torsten Küneth. Stochastic comparison of point random fields. Journal of Applied Probability, 34(4): 868–881, December 1997. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3215003.

Granovsky:2012:CPG

[GK12] Boris L. Granovsky and Alexander V. Kryvoshaev. Coagulation processes with Gibbsian time evolution. Journal of Applied Probability, 49(3):612–626, September 2012. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/41713793.

Gan:2018:ADT

[GK18] H. L. Gan and Eric D. Kolaczyk. Approximation of the difference of two Poisson-like counts by Skellam. Journal of Applied Probability, 55(2):416–430, June 2018. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL https:/

/www.cambridge.org/core/journals/journal-of-applied-
probability/article/approximation-of-the-difference-
of-two-poissonlike-counts-by-skellam/B5FCC47AFF5ADBD5D9B393A6A74A8EA9.

Gupta:2022:LCO

[GK22] Ramesh C. Gupta and S. N. U. A. Kirmani. Log-concavity and other concepts of bivariate increasing failure rate distributions. Journal of Applied Probability, 59(2):325–337, June 2022. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL https://www.cambridge.org/core/journals/
journal-of-applied-probability/article/logconcavity-
and-other-concepts-of-bivariate-increasing-failure-rate-
distributions/8D54B098BEA286C44E2B9F7BFBD0076A.

Grosswald:1980:CED

[GKJ80] E. Grosswald, Samuel Kotz, and N. L. Johnson. Characterizations of the exponential distribution by relevation-type equations. Journal of Applied Probability, 17(3):874–877, September 1980. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3212984.

Guttorp:1985:CPW

[GKL85] Peter Guttorp, Reg Kulperger, and Richard Lockhart. A coupling proof of weak convergence. Journal of Applied Probability, 22(2):447–453, June 1985. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3213788.

Giraitis:2001:TLM

[GKL01] Liudas Giraitis, Piotr Kokoszka, and Remigijus Leipus. Testing for long memory in the presence of a general trend. Journal of Applied Probability, 38(4):1033–1054, December 2001. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3215782.

Grabchak:2020:OPR

[GKP20] Michael Grabchak, Mark Kelbert, and Quentin Paris. On the occupancy problem for a regime-switching model. Journal of Applied Probability, 57(1):53–77, March 2020. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL https:/

/www.cambridge.org/core/journals/journal-of-applied-
probability/article/on-the-occupancy-problem-for-a-regimeswitching-
model/45C64B4D08F7939C9F3B92A7D3D79DD7.

Glaz:2006:GTW

[GKPS06] Joseph Glaz, Martin Kulldorff, Vladimir Pozdnyakov, and J. Michael Steele. Gambling teams and waiting times for patterns in two-state Markov chains. Journal of Applied Probability, 43(1):127–140, March 2006. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/27595715.

Gyllenberg:1994:NUP

[GKRV94] Mats Gyllenberg, Timo Koski, Edwin Reilink, and Martin Verlaan. Non-uniqueness in probabilistic numerical identification of bacteria. Journal of Applied Probability, 31(2):542–548, June 1994. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3215044.

Grace:2014:ASD

[GKS14] Adam W. Grace, Dirk P. Kroese, and Werner Sandmann. Automated state-dependent importance sampling for Markov jump processes via sampling from the zero-variance distribution. Journal of Applied Probability, 51(3):741–755, September 2014. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://projecteuclid.org/euclid.jap/1409932671.

Gilles:1967:SIG

[GL67] D. C. Gilles and P. A. W. Lewis. The spectrum of intervals of a geometric branching Poisson process. Journal of Applied Probability, 4(1):201–205, April 1967. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3212313.

Gani:1971:ART

[GL71] J. Gani and J. Lehoczky. An asymptotic result in traffic theory. Journal of Applied Probability, 8(4):815–820, December 1971. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/3212245.

Gaver:1976:GAS

[GL76] Donald P. Gaver and John P. Lehoczky. Gaussian approximations to service problems: A communication system example. Journal of Applied Probability, 13(4):768–780, December 1976. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/3212532.

Grasman:1983:ADA

[GL83] J. Grasman and D. Ludwig. The accuracy of the diffusion approximation to the expected time to extinction for some discrete stochastic processes. Journal of Applied Probability, 20(2):305– 321, June 1983. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3213804.

Gottlieb:1984:ORS

[GL84a] Gary Gottlieb and Benny Levikson. Optimal replacement for self-repairing shock models with general failure rate. Journal of Applied Probability, 21(1):108–119, March 1984. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3213669.

Gupta:1984:DDM

[GL84b] Ramesh C. Gupta and Eric S. Langford. On the determination of a distribution by its median residual life function: A functional equation. Journal of Applied Probability, 21(1):120–128, March 1984. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3213670.

Galambos:1994:FSB

[GL94] Janos Galambos and Min-Young Lee. Further studies of bivariate Bonferroni-type inequalities. Journal of Applied Probability, 31A:63–69, 1994. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3214949.

Gittenberger:1999:BEM

[GL99] Bernhard Gittenberger and Guy Louchard. The Brownian excursion multi-dimensional local time density. Journal of Applied Probability, 36(2):350–373, June 1999. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/3215461.

Girardin:2003:ESM

[GL03] Valerie Girardin and Nikolaos Limnios. On the entropy for semi-Markov processes. Journal of Applied Probability, 40(4):1060– 1068, December 2003. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3216060.

Guo:2005:SMG

[GL05] X. Guo and J. Liu. Stopping at the maximum of geometric Brownian motion when signals are received. Journal of Applied Probability, 42(3):826–838, September 2005. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/30040859.

Gladstien:1977:SPC

[Gla77] Keith Gladstien. Subdivided populations: The characteristic values and rate of loss of alleles. Journal of Applied Probability, 14(2):241–248, June 1977. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3212995.

Glazebrook:1978:CNM

[Gla78] K. D. Glazebrook. On a class of non-Markov decision processes. Journal of Applied Probability, 15(4):689–698, December 1978. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/3213426.

Glazebrook:1979:STE

[Gla79a] K. D. Glazebrook. Scheduling tasks with exponential service times on parallel processors. Journal of Applied Probability, 16 (3):685–689, September 1979. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.

org/stable/3213099.

Glazebrook:1979:SFA

[Gla79b] K. D. Glazebrook. Stoppable families of alternative bandit processes. Journal of Applied Probability, 16(4):843–854, December 1979. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3213150.

Glazebrook:1980:SSP

[Gla80] K. D. Glazebrook. On stochastic scheduling with precedence relations and switching costs. Journal of Applied Probability, 17 (4):1016–1024, December 1980. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.

org/stable/3213211.

Glaz:1981:CES

[Gla81] Joseph Glaz. Clustering of events in a stochastic process. Journal of Applied Probability, 18(1):268–275, March 1981. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3213188.

Glazebrook:1982:SCS

[Gla82a] K. D. Glazebrook. On a sufficient condition for superprocesses due to Whittle. Journal of Applied Probability, 19(1):99–110, March 1982. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3213920.

Glazebrook:1982:ESS

[Gla82b] K. D. Glazebrook. On the evaluation of suboptimal strategies for families of alternative bandit processes. Journal of Applied Probability, 19(3):716–722, September 1982. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/3213534.

Glazebrook:1990:PES

[Gla90] K. D. Glazebrook. Procedures for the evaluation of strategies for resource allocation in a stochastic environment. Journal of Applied Probability, 27(1):215–220, March 1990. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3214609.

Glazebrook:1991:BDS

[Gla91] K. D. Glazebrook. Bounds for discounted stochastic scheduling problems. Journal of Applied Probability, 28(4):791–801, December 1991. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3214682.

Glasserman:1992:PAI

[Gla92] Paul Glasserman. Processes with associated increments. Journal of Applied Probability, 29(2):313–333, June 1992. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3214569.

Glazebrook:1993:PPS

[Gla93] K. D. Glazebrook. On permutation policies for the scheduling of deteriorating stochastic jobs on a single machine. Journal of Applied Probability, 30(1):184–193, March 1993. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3214631.

Glickman:2000:BCP

[Gli00] Hagit Glickman. A best-choice problem with multiple selectors. Journal of Applied Probability, 37(3):718–735, September 2000. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/3215608.

Glover:2023:RSU

[Glo23] Kristoffer Glover. With or without replacement? Sampling uncertainty in Shepp’s urn scheme. Journal of Applied Probability, 60(2):661–675, June 2023. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL https:/

/www.cambridge.org/core/journals/journal-of-applied-
probability/article/with-or-without-replacement-sampling-
uncertainty-in-shepps-urn-scheme/662590943BC0644C5937E67FC32A3E43.

Gouet:2000:ECV

[GLS00] Raúl Gouet, F. Javier López, and Gerardo Sanz. Estimation of critical values in interacting particle systems. Journal of Applied Probability, 37(1):118–125, March 2000. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http:/

/www.jstor.org/stable/3215663.

Geiss:2019:FET

[GLS19] Christel Geiss, Antti Luoto, and Paavo Salminen. On first exit times and their means for Brownian bridges. Journal of Applied Probability, 56(3):701–722, September 2019. CODEN JPRBAM.

ISSN 0021-9002 (print), 1475-6072 (electronic). URL https:/

/www.cambridge.org/core/journals/journal-of-applied-
probability/article/on-first-exit-times-and-their-means-
for-brownian-bridges/7A0011FA7BA83B94428C85EBDBBD45FC.

Glynn:1989:DTS

[Gly89] John E. Glynn. A discrete-time storage process with a general release rule. Journal of Applied Probability, 26(3):566–583, September 1989. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3214414.

Gittins:1973:GIB

[GM73] J. C. Gittins and M. J. Maher. A generalised incomplete beta function and its application to multi-line stock control. Journal of Applied Probability, 10(4):748–760, December 1973. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3212378.

Ghurye:1984:SPA

[GM84a] S. G. Ghurye and Albert W. Marshall. Shock processes with aftereffects and multivariate lack of memory. Journal of Applied Probability, 21(4):786–801, December 1984. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/3213696.

Goodman:1984:NEJ

[GM84b] Jonathan B. Goodman and William A. Massey. The non-ergodic Jackson network. Journal of Applied Probability, 21(4):860–869, December 1984. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3213702.

Gani:1991:CBE

[GM91] J. Gani and Gy. Michaletzky. A carrier-borne epidemic with multiple stages of infection. Journal of Applied Probability, 28(1): 1–8, March 1991. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3214735.

Guillemin:1994:BED

[GM94] Fabrice Guillemin and Ravi Mazumdar. The Beneš equations for the distribution of excursions for general storage processes.

Journal of Applied Probability, 31(2):418–429, June 1994. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/3215035.

Gonzalez:1996:LBS

[GM96] M. Gonzalez and M. Molina. On the limit behaviour of a superadditive bisexual Galton–Watson branching process. Journal of Applied Probability, 33(4):960–967, December 1996. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3214977.

Gonzalez:1997:CSB

[GM97] M. González and M. Molina. On the L^{2}-convergence of a superadditive bisexual Galton–Watson branching process. Journal of Applied Probability, 34(3):575–582, September 1997. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3215085.

Gamarnik:2005:TRA

[GM05] David Gamarnik and Petar Momcilović. A transposition rule analysis based on a particle process. Journal of Applied Probability, 42(1):235–246, March 2005. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/30040782.

Gnedin:2007:WRF

[GM07] Alexander V. Gnedin and Denis I. Miretskiy. Winning rate in the full-information best-choice problem. Journal of Applied Probability, 44(2):560–565, June 2007. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/27595862.

Glynn:2011:SBC

[GM11] Peter W. Glynn and Michel Mandjes. Simulation-based computation of the workload correlation function in a Lévy–driven queue. Journal of Applied Probability, 48(1):114–130, March 2011. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/29777449.

Guillemin:2015:CST

[GM15] Fabrice M. Guillemin and Ravi R. Mazumdar. Conditional sojourn times and the volatility of payment schemes for bandwidth sharing in packet networks. Journal of Applied Probability, 52(4):

962–980, December 2015. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://projecteuclid.

org/euclid.jap/1450802746.

Ghosh:2018:ACB

[GM18] Shyamal Ghosh and Murari Mitra. On an ageing class based on the moment generating function order. Journal of Applied Probability, 55(2):402–415, June 2018. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL https://www.cambridge.org/core/journals/
journal-of-applied-probability/article/on-an-ageing-
class-based-on-the-moment-generating-function-order/
FF6A8E33EB110E68F043C78483F0B25D.

Grusea:2020:IDM

[GM20] S. Grusea and S. Mercier. Improvements on the distribution of maximal segmental scores in a Markovian sequence. Journal of Applied Probability, 57(1):29–52, March 2020. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL https://www.cambridge.org/core/journals/
journal-of-applied-probability/article/improvements-
on-the-distribution-of-maximal-segmental-scores-in-a-
markovian-sequence/AAF95443BE10F0FD0EF48ECBA13A81A4.

Gonzalez:2002:CCB

[GMD02] M. González, M. Molina, and I. Del Puerto. On the class of controlled branching processes with random control functions. Journal of Applied Probability, 39(4):804–815, December 2002. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/3216004.

Gonzalez:2003:GGC

[GMD03] M. González, M. Molina, and I. Del Puerto. On the geometric growth in controlled branching processes with random control function. Journal of Applied Probability, 40(4):995–1006, December 2003. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3216056.

Gonzalez:2005:ABC

[GMD05] M. González, M. Molina, and I. Del Puerto. Asymptotic behaviour of critical controlled branching processes with random control functions. Journal of Applied Probability, 42(2):463– 477, June 2005. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
30040804.

Guillemin:2003:AFW

[GMDB03] Fabrice Guillemin, Ravi Mazumdar, Alain Dupuis, and Jacqueline Boyer. Analysis of the fluid weighted fair queueing system. Journal of Applied Probability, 40(1):180–199, March 2003. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/3215848.

Gonzalez:2021:CBP

[GMdP^{+}21] Miguel González, Manuel Molina, Ines del Puerto, Nikolay M. Yanev, and George P. Yanev. Controlled branching processes with continuous time. Journal of Applied Probability, 58(3):830–848, September 2021. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL https:/

/www.cambridge.org/core/journals/journal-of-applied-
probability/article/controlled-branching-processes-with-
continuous-time/0AD11110A3EC498AEB0A507AE4DE4339.

Gonzalez:2005:GGC

[GMM05] M. González, R. Martínez, and M. Mota. On the geometric growth in a class of homogeneous multitype Markov chain. Journal of Applied Probability, 42(4):1015–1030, December 2005. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/30040825.

Gonzalez:2006:RGC

[GMM06] M. González, R. Martínez, and M. Mota. Rates of growth in a class of homogeneous multidimensional Markov chains. Journal of Applied Probability, 43(1):159–174, March 2006. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/27595717.

Guillemin:2007:SOP

[GMRY07] Fabrice M. Guillemin, Ravi R. Mazumdar, Catherine P. Rosenberg, and Yu Ying. A stochastic ordering property for leaky bucket regulated flows in packet networks. Journal of Applied Probability, 44(2):332–348, June 2007. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/27595845.

Guillemin:1996:HTA

[GMS96] Fabrice M. Guillemin, Ravi R. Mazumdar, and Alain D. Simonian. On heavy traffic approximations for transient characteristics of M/M/∞ queues. Journal of Applied Probability, 33 (2):490–506, June 1996. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3215073.

Gonzalez:2010:SMC

[GMSB10] M. González, R. Martínez, and M. Slavtchova-Bojkova. Stochastic monotonicity and continuity properties of the extinction time of Bellman–Harris branching processes: an application to epidemic modelling. Journal of Applied Probability, 47(1):58– 71, March 2010. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
25662500.

Gillett:2006:MAB

[GMV06] Alexis Gillett, Ronald Meester, and Peter Van Der Wal. Maximal avalanches in the Bak–Sneppen model. Journal of Applied Probability, 43(3):840–851, September 2006. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/27595776.

Gupta:1990:TDE

[GMW90] Vijay K. Gupta, Oscar J. Mesa, and E. Waymire. Treedependent extreme values: The exponential case. Journal of Applied Probability, 27(1):124–133, March 1990. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3214600.

Gripenberg:1996:PFB

[GN96] Gustaf Gripenberg and Ilkka Norros. On the prediction of fractional Brownian motion. Journal of Applied Probability, 33 (2):400–410, June 1996. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3215063.

Geluk:2006:TBN

[GN06] Jaap Geluk and Kai W. Ng. Tail behaviour of negatively associated heavy-tailed sums. Journal of Applied Probability, 43 (2):587–593, June 2006. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/27595752.

Gaasemyr:2017:IAB

[GN17] Jørund Gåsemyr and Bent Natvig. Improved availability bounds for binary and multistate monotone systems with independent component processes. Journal of Applied Probability, 54(3):750–762, September 2017. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL https:/

/www.cambridge.org/core/journals/journal-of-applied-
probability/article/improved-availability-bounds-for-
binary-and-multistate-monotone-systems-with-independent-
component-processes/28A3143C684B207BCC6C8F5FABD349B5.

Gnedin:1994:BCP

[Gne94] Alexander V. Gnedin. On a best-choice problem with dependent criteria. Journal of Applied Probability, 31(1):221–234, March 1994. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3215248.

Gnedin:1996:FIB

[Gne96] Alexander V. Gnedin. On the full information best-choice problem. Journal of Applied Probability, 33(3):678–687, September 1996. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3215349.

Gnedin:1998:CES

[Gne98] Alexander V. Gnedin. On convergence and extensions of sizebiased permutations. Journal of Applied Probability, 35(3):642– 650, September 1998. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3215639.

Gnedin:1999:SSI

[Gne99] Alexander V. Gnedin. Sequential selection of an increasing subsequence from a sample of random size. Journal of Applied Probability, 36(4):1074–1085, December 1999. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/3215579.

Gneiting:2000:PLC

[Gne00] Tilmann Gneiting. Power-law correlations, related models for long-range dependence and their simulation. Journal of Applied Probability, 37(4):1104–1109, December 2000. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3215499.

Gnedin:2007:OSR

[Gne07] Alexander V. Gnedin. Optimal stopping with rank-dependent loss. Journal of Applied Probability, 44(4):996–1011, December 2007. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/27595903.

Giorno:1985:SMF

[GNN85] V. Giorno, C. Negri, and A. G. Nobile. A solvable model for a finite-capacity queueing system. Journal of Applied Probability, 22(4):903–911, December 1985. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.

org/stable/3213957.

Giorno:1986:RIN

[GNN86] V. Giorno, C. Negri, and A. G. Nobile. A reply to Iversen and Nielsen’s letter. Journal of Applied Probability, 23(4):1049–1050, December 1986. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/

3214481. See [GNN86, IN86].

Giorno:1989:SBC

[GNR89] V. Giorno, A. G. Nobile, and L. M. Ricciardi. A symmetrybased constructive approach to probability densities for onedimensional diffusion processes. Journal of Applied Probability, 26(4):707–721, December 1989. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3214376.

Giorno:1986:SRR

[GNRS86] V. Giorno, A. G. Nobile, L. M. Ricciardi, and L. Sacerdote. Some remarks on the Rayleigh process. Journal of Applied Probability, 23(2):398–408, June 1986. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/ stable/3214182. See comments [IN86] and reply [GNN86].

Glaz:1994:PAD

[GNRW94] Joseph Glaz, Joseph Naus, Małgorzata Roos, and Sylvan Wallenstein. Poisson approximations for the distribution and moments of ordered m-spacings. Journal of Applied Probability, 31A:271–281, 1994. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3214961.

Godolphin:1977:ASS

[God77] E. J. Godolphin. On the autocorrelation structure for seasonal moving average models and its implications for the Cramér– Wold factorization. Journal of Applied Probability, 14(4):785– 794, December 1977. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3213351.

Goldstein:1974:SIV

[Gol74] Matthew Goldstein. Some inequalities on variances. Journal of Applied Probability, 11(2):409–412, June 1974. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3212765.

Goldstein:1975:SRP

[Gol75] M. Goldstein. Some remarks on probability inequalities for sums of bounded convex random variables. Journal of Applied Probability, 12(1):155–158, March 1975. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3212418.

Goldstein:1976:BBP

[Gol76] Matthew Goldstein. A bound for bivariate probability of large deviations. Journal of Applied Probability, 13(2):380–384, June 1976. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3212845.

Goldie:1977:LSL

[Gol77] Charles M. Goldie. Lanchester square-law battles: Transient and terminal distributions. Journal of Applied Probability, 14 (3):604–610, September 1977. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.

org/stable/3213464.

Goldie:1978:SDD

[Gol78a] Charles M. Goldie. Subexponential distributions and dominatedvariation tails. Journal of Applied Probability, 15(2):440–442, June 1978. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3213416.

Goldstein:1978:ULT

[Gol78b] Martin I. Goldstein. A uniform limit theorem and exponential limit law for critical multitype age-dependent branching processes. Journal of Applied Probability, 15(2):235–242, June 1978. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/3213397.

Goldie:1995:ALD

[Gol95] Charles M. Goldie. Asymptotics of the Luria–Delbrück distribution. Journal of Applied Probability, 32(3):840–841, September 1995. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3215135.

Goldstein:2005:BEB

[Gol05] Larry Goldstein. Berry–Esseen bounds for combinatorial central limit theorems and pattern occurrences, using zero and size biasing. Journal of Applied Probability, 42(3):661–683, September 2005. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/30040848.

Goldman:2023:NRC

[Gol23] André Goldman. A note related to the CS decomposition and the BK inequality for discrete determinantal processes. Journal of Applied Probability, 60(1):189–203, March 2023. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL https://www.cambridge.org/core/journals/
journal-of-applied-probability/article/note-related-
to-the-cs-decomposition-and-the-bk-inequality-for-discrete-
determinantal-processes/0D91D94D8164E5FF73D3833ED6426CAD.

Garra:2015:SDF

[GOP15] R. Garra, E. Orsingher, and F. Polito. State-dependent fractional point processes. Journal of Applied Probability, 52(1): 18–36, March 2015. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://projecteuclid.

org/euclid.jap/1429282604.

Gordon:1970:MCA

[Gor70] Richard Gordon. On Monte Carlo algebra. Journal of Applied Probability, 7(2):373–387, August 1970. CODEN JPRBAM.

ISSN 0021-9002 (print), 1475-6072 (electronic). URL http:/

/www.jstor.org/stable/3211971.

Gorostiza:1975:CTP

[Gor75] Luis G. Gorostiza. Convergence of transport processes with radially symmetric direction changes, and chain molecules. Journal of Applied Probability, 12(4):812–816, December 1975. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3212732.

Gordy:2014:FDD

[Gor14] Michael B. Gordy. Finite-dimensional distributions of a squareroot diffusion. Journal of Applied Probability, 51(4):930–942, December 2014. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-

6072 (electronic). URL http://projecteuclid.org/euclid.
jap/1421763319.

Gottlieb:1980:FDS

[Got80] Gary Gottlieb. Failure distributions of shock models. Journal of Applied Probability, 17(3):745–752, September 1980. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3212968.

Gottlieb:1985:OBS

[Got85] Gary Gottlieb. An optimal betting strategy for repeated games. Journal of Applied Probability, 22(4):787–795, December 1985. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/3213946.

Gottlieb:1987:NRS

[Got87] Gary Gottlieb. Non-random shuffling for multiple decks. Journal of Applied Probability, 24(2):402–410, June 1987. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3214264.

Gouet:1997:SCP

[Gou97] Raúl Gouet. Strong convergence of proportions in a multicolor Pólya urn. Journal of Applied Probability, 34(2):426–435, June 1997. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3215382.

Gouweleeuw:1998:MOS

[Gou98] José M. Gouweleeuw. A minimax optimal stop rule in reliability checking. Journal of Applied Probability, 35(3):748–761, September 1998. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3215649.

Guo:1991:NRT

[GP91] Meihui Guo and Joseph D. Petruccelli. On the null recurrence and transience of a first-order SETAR model. Journal of Applied Probability, 28(3):584–592, September 1991. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/3214493.

Graversen:1997:WTO

[GP97] S. E. Graversen and G. Peskir. On Wald-type optimal stopping for Brownian motion. Journal of Applied Probability, 34(1):66– 73, March 1997. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3215175.

Graversen:1998:OSM

[GP98a] S. E. Graversen and G. Peskir. Optimal stopping and maximal inequalities for geometric Brownian motion. Journal of Applied Probability, 35(4):856–872, December 1998. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/3215522.

Guillemin:1998:CFA

[GP98b] Fabrice Guillemin and Didier Pinchon. Continued fraction analysis of the duration of an excursion in an M/M/∞ system. Journal of Applied Probability, 35(1):165–183, March 1998. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3215555.

Guillemin:1999:EBD

[GP99] Fabrice Guillemin and Didier Pinchon. Excursions of birth and death processes, orthogonal polynomials, and continued fractions. Journal of Applied Probability, 36(3):752–770, September 1999. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3215439.

Genton:2004:TDR

[GP04a] Marc G. Genton and Olivier Perrin. On a time deformation reducing nonstationary stochastic processes to local stationarity. Journal of Applied Probability, 41(1):236–249, March 2004. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/3215828.

Guillemin:2004:AGP

[GP04b] Fabrice Guillemin and Didier Pinchon. Analysis of generalized processor-sharing systems with two classes of customers and exponential services. Journal of Applied Probability, 41(3):832– 858, September 2004. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/4141357.

Guillotin-Plantard:2005:GRW

[GP05] Nadine Guillotin-Plantard. Gillis’s random walks on graphs. Journal of Applied Probability, 42(1):295–301, March 2005. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/30040789.

Garner:2013:CAV

[GP13] William Garner and Dimitris N. Politis. The correct asymptotic variance for the sample mean of a homogeneous Poisson marked point process. Journal of Applied Probability, 50(3): 889–892, September 2013. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://projecteuclid.

org/euclid.jap/1378401244.

Gannon:2016:RWQ

[GPSY16] M. Gannon, E. Pechersky, Y. Suhov, and A. Yambartsev. Random walks in a queueing network environment. Journal of Applied Probability, 53(2):448–462, June 2016. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

projecteuclid.org/euclid.jap/1466172866.

Germain:2022:RCP

[GPW22] Maximilien Germain, Huyên Pham, and Xavier Warin. Rate of convergence for particle approximation of PDEs in Wasserstein space. Journal of Applied Probability, 59 (4):992–1008, December 2022. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL https:/

/www.cambridge.org/core/journals/journal-of-applied-
probability/article/rate-of-convergence-for-particle-
approximation-of-pdes-in-wasserstein-space/3EB8DE6D82B8A63530860C63EBD865B5.

Gray:2005:EUI

[GPZ05] Brenton Gray, Phil Pollett, and Hanjun Zhang. On the existence of uni-instantaneous Q-processes with a given finite μ-invariant measure. Journal of Applied Probability, 42(3):713–725, September 2005. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/30040851.

Guo:2017:NDC

[GPZ17] Xin Guo, Alexey Piunovskiy, and Yi Zhang. Note on discounted continuous-time Markov decision processes with a lower bounding function. Journal of Applied Probability, 54 (4):1071–1088, December 2017. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL https:/

/www.cambridge.org/core/journals/journal-of-applied-
probability/article/note-on-discounted-continuoustime-
markov-decision-processes-with-a-lower-bounding-function/
33B7475387AD97BBCC960FA1EE199CD8.

Grinstein:1974:QRS

[GR74] Jacob Grinstein and Michael Rubinovitch. Queues with random service output: The case of Poisson arrivals. Journal of Applied Probability, 11(4):771–784, December 1974. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/3212560.

Gadag:1987:MPB

[GR87] V. G. Gadag and M. B. Rajarshi. On multitype processes based on progeny length of particles of a supercritical Galton–Watson process. Journal of Applied Probability, 24(1):14–24, March 1987. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3214055.

Gadag:1989:MBP

[GR89] V. G. Gadag and M. B. Rajarshi. Multitype branching processes based on exact progeny lengths of particles in a Galton– Watson branching process. Journal of Applied Probability, 26(1): 1–8, March 1989. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3214311.

Granovsky:1994:TBN

[GR94] Boris L. Granovsky and Liat Rozov. On transient behaviour of a nearest-neighbour birth-death process on a lattice. Journal of Applied Probability, 31(2):549–553, June 1994. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3215045.

Goldstein:1996:MNA

[GR96] Larry Goldstein and Yosef Rinott. Multivariate normal approximations by Stein’s method and size bias couplings. Journal of Applied Probability, 33(1):1–17, March 1996. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3215259.

Garren:1998:GCC

[GR98] Steven T. Garren and Donald St. P. Richards. General conditions for comparing the reliability functions of systems of components sharing a common environment. Journal of Applied Probability, 35(1):124–135, March 1998. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3215552.

Grubel:2001:TTS

[GR01] Rudolf Grübel and Anke Reimers. On the total time spent in records by a discrete uniform sequence. Journal of Applied Probability, 38(3):768–775, September 2001. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/3216129.

Grubel:2005:REE

[GR05] Rudolf Grübel and Marcus Reich. Rarity and exponentiality: An extension of Keilson’s theorem, with applications. Journal of Applied Probability, 42(2):393–406, June 2005. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/30040800.

Ghamami:2012:IAK

[GR12] Samim Ghamami and Sheldon M. Ross. Improving the Asmussen–Kroese-type simulation estimators. Journal of Applied Probability, 49(4):1188–1193, December 2012. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/41713837.

Goldstein:2013:SMB

[GR13] Larry Goldstein and Gesine Reinert. Stein’s method for the Beta distribution and the Pólya–Eggenberger urn. Journal of Applied Probability, 50(4):1187–1205, December 2013. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://projecteuclid.org/euclid.jap/1389370107.

Gapeev:2014:OSP

[GR14] Pavel V. Gapeev and Neofytos Rodosthenous. Optimal stopping problems in diffusion-type models with running maxima and drawdowns. Journal of Applied Probability, 51(3):799– 817, September 2014. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://projecteuclid.

org/euclid.jap/1409932675.

Gallo:2018:FMT

[GR18] Sandro Gallo and Pablo M. Rodriguez. Frog models on trees through renewal theory. Journal of Applied Probability, 55(3):887–899, September 2018. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL https:/

/www.cambridge.org/core/journals/journal-of-applied-
probability/article/frog-models-on-trees-through-renewal-
theory/E3AD0C1E5F619383F42CECAD4F36BEA8.

Galeotti:2021:CSV

[GR21] Marcello Galeotti and Giovanni Rabitti. On the comparison of Shapley values for variance and standard deviation games. Journal of Applied Probability, 58(3):609–620, September 2021. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL https://www.cambridge.org/core/journals/journal-
of-applied-probability/article/on-the-comparison-of-
shapley-values-for-variance-and-standard-deviation-games/
E5B9F62E98195075619902F27CFA914E.

Galeotti:2024:TVA

[GR24] Marcello Galeotti and Giovanni Rabitti. Tail variance allocation, Shapley value, and the majorization problem. Journal of Applied Probability, 61(1):153–171, March 2024. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL https:/

/www.cambridge.org/core/journals/journal-of-applied-

probability/article/tail-variance-allocation-shapley-
value-and-the-majorization-problem/36D868E22E2D0B46C74ABA98FF7CF90D.

Grandell:1971:NLE

[Gra71] Jan Grandell. A note on linear estimation of the intensity in a doubly stochastic Poisson field. Journal of Applied Probability, 8(3):612–614, September 1971. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.

org/stable/3212184.

Grandell:1972:EIS

[Gra72] Jan Grandell. On the estimation of intensities in a stochastic process generated by a stochastic intensity sequence. Journal of Applied Probability, 9(3):542–556, September 1972. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3212324.

Grassmann:1983:CMQ

[Gra83] W. Grassmann. The convexity of the mean queue size of the M/M/c queue with respect to the traffic intensity. Journal of Applied Probability, 20(4):916–919, December 1983. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3213605.

Gravey:1985:SCU

[Gra85] A. Gravey. A simple construction of an upper bound for the mean of the maximum of n identically distributed random variables. Journal of Applied Probability, 22(4):844–851, December 1985. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3213952.

Gravner:1999:RRD

[Gra99] Janko Gravner. Recurrent ring dynamics in two-dimensional excitable cellular automata. Journal of Applied Probability, 36 (2):492–511, June 1999. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3215470.

Graham:2000:CPS

[Gra00] Carl Graham. Chaoticity on path space for a queueing network with selection of the shortest queue among several. Journal of Applied Probability, 37(1):198–211, March 2000. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3215668.

Graham:2008:CMS

[Gra08] Carl Graham. Chaoticity for multiclass systems and exchangeability within classes. Journal of Applied Probability, 45(4): 1196–1203, December 2008. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.

org/stable/27596018.

Grabchak:2012:NCT

[Gra12] Michael Grabchak. On a new class of tempered stable distributions: Moments and regular variation. Journal of Applied Probability, 49(4):1015–1035, December 2012. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/41713826.

Grabchak:2018:DAP

[Gra18] Michael Grabchak. Domains of attraction for positive and discrete tempered stable distributions. Journal of Applied Probability, 55(1):30–42, March 2018. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL https:/

/www.cambridge.org/core/journals/journal-of-applied-
probability/article/domains-of-attraction-for-positive-
and-discrete-tempered-stable-distributions/1B499829EE062B079871B6CA3974C364.

Gyorfi:2000:DFC

[GRD^{+}00] László Györfi, András Rácz, Ken Duffy, John T. Lewis, and Fergal Toomey. Distribution-free confidence intervals for measurement of effective bandwidth. Journal of Applied Probability, 37 (1):224–235, March 2000. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3215670.

Greenberg:1969:SDR

[Gre69] Irwin Greenberg. Some duality results in the theory of queues. Journal of Applied Probability, 6(1):99–121, April 1969. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3212279.

Grey:1974:ABC

[Gre74] D. R. Grey. Asymptotic behaviour of continuous time, continuous state-space branching processes. Journal of Applied Probability, 11(4):669–677, December 1974. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3212550.

Green:1976:MTA

[Gre76a] P. J. Green. The maximum and time to absorption of a leftcontinuous random walk. Journal of Applied Probability, 13(3): 444–454, September 1976. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3212464.

Green:1976:PAM

[Gre76b] Richard F. Green. Partial attraction of maxima. Journal of Applied Probability, 13(1):159–163, March 1976. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/3212678. See correction

[Ano77e].

Grey:1976:PRG

[Gre76c] D. R. Grey. On possible rates of growth of age-dependent branching processes with immigration. Journal of Applied Probability, 13(1):138–143, March 1976. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3212674.

Green:1977:CLT

[Gre77a] P. J. Green. Conditional limit theorems for general branching processes. Journal of Applied Probability, 14(3):451–463, September 1977. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3213448.

Grey:1977:ASC

[Gre77b] D. R. Grey. Almost sure convergence in Markov branching processes with infinite mean. Journal of Applied Probability, 14(4): 702–716, December 1977. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3213344.

Grey:1978:SGW

[Gre78] D. R. Grey. Supercritical Galton–Watson branching processes: Corrections to a paper of Foster and Goettge. Journal of Applied Probability, 15(3):660, September 1978. CODEN JPRBAM.

ISSN 0021-9002 (print), 1475-6072 (electronic). URL http:/

/www.jstor.org/stable/3213132. See [FG76].

Greenberg:1980:MCL

[Gre80a] Irwin Greenberg. The moments of coverage of a linear set. Journal of Applied Probability, 17(3):865–868, September 1980. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/3212982.

Grey:1980:FCC

[Gre80b] D. R. Grey. The fair charge on a car ferry. Journal of Applied Probability, 17(3):654–661, September 1980. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/3212959.

Green:1981:MYC

[Gre81] P. J. Green. Modelling yeast cell growth using stochastic branching processes. Journal of Applied Probability, 18(4):799–808, December 1981. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3213055.

Grey:1984:NNM

[Gre84] D. R. Grey. Non-negative matrices, dynamic programming and a harvesting problem. Journal of Applied Probability, 21(4): 685–694, December 1984. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/ stable/3213687. See comment [Gre86].

Grey:1986:CSE

[Gre86] D. R. Grey. Comment on survey at earlier work [that] omitted the paper of Mendelssohn. Journal of Applied Probability, 23 (4):1050, December 1986. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/ stable/3214482. See [Gre84, Men76].

Greenberg:1989:MGQ

[Gre89] Betsy S. Greenberg. M/G/1 queueing systems with returning customers. Journal of Applied Probability, 26(1):152–163, March 1989. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3214325.

Grey:2001:NCP

[Gre01] D. R. Grey. A note on convergence of probability measures. Journal of Applied Probability, 38(4):1055–1058, December 2001. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/3215783.

Griffiths:1972:BBD

[Gri72] D. A. Griffiths. A bivariate birth-death process which approximates to the spread of a disease involving a vector. Journal of Applied Probability, 9(1):65–75, March 1972. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3212637.

Griffiths:1973:MBD

[Gri73] D. A. Griffiths. Multivariate birth-and-death processes as approximations to epidemic processes. Journal of Applied Probability, 10(1):15–26, March 1973. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3212492. See correction [Ano74e].

Griffeath:1975:UCN

[Gri75a] David Griffeath. Uniform coupling of non-homogeneous Markov chains. Journal of Applied Probability, 12(4):753–762, December 1975. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3212726.

Grimmett:1975:NMB

[Gri75b] G. R. Grimmett. A non-Markovian birth process with logarithmic growth. Journal of Applied Probability, 12(4):673–683, December 1975. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3212718.

Griffith:1980:MRM

[Gri80] William S. Griffith. Multistate reliability models. Journal of Applied Probability, 17(3):735–744, September 1980. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3212967.

Griffiths:1981:TDN

[Gri81] R. C. Griffiths. Transient distribution of the number of segregating sites in a neutral infinite-sites model with no recombination.

Journal of Applied Probability, 18(1):42–51, March 1981. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/3213165.

Griffiths:1988:DPP

[Gri88] R. C. Griffiths. On the distribution of points in a Poisson Dirichlet process. Journal of Applied Probability, 25(2):336–345, June 1988. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3214441.

Griffiths:1992:DNA

[Gri92] R. C. Griffiths. Distribution of the number of alleles in multigene families. Journal of Applied Probability, 29(4):759–769, December 1992. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3214710.

Grishechkin:1998:IMA

[Gri98] Sergei Grishechkin. Inequalities for multinomial allocations with application to DNA fingerprinting. Journal of Applied Probability, 35(3):707–717, September 1998. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3215645.

Groemer:1980:ADB

[Gro80] H. Groemer. The average distance between two convex sets. Journal of Applied Probability, 17(2):415–422, June 1980. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/3213030.

Gutierrez:1997:FPT

[GRRT97] R. Gutiérrez, L. M. Ricciardi, P. Román, and F. Torres. Firstpassage-time densities for time-non-homogeneous diffusion processes. Journal of Applied Probability, 34(3):623–631, September 1997. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3215089.

Grubel:1985:TBL

[Grü85] Rudolf Grübel. Tail behaviour of ladder-height distributions in random walks. Journal of Applied Probability, 22(3):705–709, September 1985. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3213873.

Grubel:1989:SMF

[Grü89] Rudolf Grübel. Stochastic models as functionals: Some remarks on the renewal case. Journal of Applied Probability, 26(2):296– 303, June 1989. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3214036.

Grubel:1998:HSA

[Grü98] Rudolf Grübel. Hoare’s selection algorithm: A Markov chain approach. Journal of Applied Probability, 35(1):36–45, March 1998. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3215544.

Ghosh:1972:NCU

[GS72] P. Ghosh and D. N. Shanbhag. A note on the characterization of a U-shaped probability density function. Journal of Applied Probability, 9(3):684–685, September 1972. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/3212339.

Gani:1976:PIB

[GS76] J. Gani and I. W. Saunders. On the parity of individuals in a branching process. Journal of Applied Probability, 13(2):219– 230, June 1976. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3212825.

Gross:1979:LAD

[GS79] S. Gross and W. L. Steiger. Least absolute deviation estimates in autoregression with infinite variance. Journal of Applied Probability, 16(1):104–116, March 1979. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3213379.

Griffith:1985:NSM

[GS85] William S. Griffith and C. Srinivasan. A note on survival models under a Markov process. Journal of Applied Probability, 22(4): 982–988, December 1985. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3213969.

Gibson:1987:ATI

[GS87] Diana Gibson and E. Seneta. Augmented truncations of infinite stochastic matrices. Journal of Applied Probability, 24(3): 600–608, September 1987. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3214092.

Getoor:1994:ASL

[GS94] R. K. Getoor and M. J. Sharpe. On the arc-sine laws for Lévy processes. Journal of Applied Probability, 31(1):76–89, March 1994. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3215236.

Gruet:1995:SAR

[GS95] J.-C. Gruet and Z. Shi. Some asymptotic results for exponential functionals of Brownian motion. Journal of Applied Probability, 32(4):930–940, December 1995. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.

org/stable/3215206.

Gouet:1999:POM

[GS99] Raúl Gouet and Jaime San Martín. Partial orders and minimization of records in a sequence of independent random variables. Journal of Applied Probability, 36(4):965–973, December 1999. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/3215570.

Guo:2001:SOS

[GS01] Xin Guo and Larry Shepp. Some optimal stopping problems with nontrivial boundaries for pricing exotic options. Journal of Applied Probability, 38(3):647–658, September 2001. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3216118.

Gruber:2006:DLG

[GS06] Urs Gruber and Martin Schweizer. A diffusion limit for generalized correlated random walks. Journal of Applied Probability, 43(1):60–73, March 2006. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/27595710.

Gani:2008:OCC

[GS08] Joe Gani and Eugene Seneta. Obituary: Christopher Charles Heyde, AM, DSc, FAA, FASSA. Journal of Applied Probability, 45(3):587–592, September 2008. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.

org/stable/27595971.

Gava:2014:SPP

[GS14] Renato Jacob Gava and Danilo Salotti. Stopping probabilities for patterns in Markov chains. Journal of Applied Probability, 51 (1):287–292, March 2014. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://projecteuclid.

org/euclid.jap/1395771430.

Guillemin:2015:VDL

[GS15] Fabrice Guillemin and Bruno Sericola. Volume and duration of losses in finite buffer fluid queues. Journal of Applied Probability, 52(3):826–840, September 2015. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http:/

/projecteuclid.org/euclid.jap/1445543849.

Goffard:2019:ECR

[GS19] Pierre-Olivier Goffard and Andrey Sarantsev. Exponential convergence rate of ruin probabilities for level-dependent Lévydriven risk processes. Journal of Applied Probability, 56 (4):1244–1268, December 2019. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL https:/

/www.cambridge.org/core/journals/journal-of-applied-
probability/article/exponential-convergence-rate-of-
ruin-probabilities-for-leveldependent-levydriven-risk-
processes/840E17B2480AC08C0444EFAC9F648DB8.

Gut:2021:VER

[GS21] Allan Gut and Ulrich Stadtmüller. Variations of the elephant random walk. Journal of Applied Probability, 58 (3):805–829, September 2021. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL https:/

/www.cambridge.org/core/journals/journal-of-applied-
probability/article/variations-of-the-elephant-random-
walk/18D67A1FB908CB6EA2DD07B309BB0239.

Gertsbakh:2011:SCS

[GSS11] Ilya Gertsbakh, Yoseph Shpungin, and Fabio Spizzichino. Signatures of coherent systems built with separate modules. Journal of Applied Probability, 48(3):843–855, September 2011. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/23065941.

Gertsbakh:2012:TDS

[GSS12] Ilya Gertsbakh, Yoseph Shpungin, and Fabio Spizzichino. Twodimensional signatures. Journal of Applied Probability, 49(2): 416–429, June 2012. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/41713776.

Gerardi:2000:FEC

[GST00] A. Gerardi, F. Spizzichino, and B. Torti. Filtering equations for the conditional law of residual lifetimes from a heterogeneous population. Journal of Applied Probability, 37(3):823–834, September 2000. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3215616.

Gasbarra:2011:CFS

[GSV11] Dario Gasbarra, Tommi Sottinen, and Harry Van Zanten. Conditional full support of Gaussian processes with stationary increments. Journal of Applied Probability, 48(2):561–568, June 2011. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/23065902.

Gergely:1974:BPD

[GT74] T. Gergely and T. L. Török. On the busy period of discrete-time queues. Journal of Applied Probability, 11(4):853–857, December 1974. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3212571.

Gani:1985:CCT

[GT85] J. Gani and Pyke Tin. On a class of continuous-time Markov processes. Journal of Applied Probability, 22(4):804–815, December 1985. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/

3213948. See comment [Boy87].

Gani:1986:NTS

[GT86] J. Gani and Pyke Tin. A note on the two-sex population process. Journal of Applied Probability, 23A:335–344, 1986. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/3214362.

Gerardi:2005:HPD

[GT05] A. Gerardi and P. Tardelli. Heterogeneous population dynamical model: A filtering problem. Journal of Applied Probability, 42 (2):346–361, June 2005. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/30040797.

Ganesh:2006:CRP

[GT06] Ayalvadi Ganesh and Giovanni Luca Torrisi. A class of risk processes with delayed claims: Ruin probability estimates under heavy tail conditions. Journal of Applied Probability, 43(4): 916–926, December 2006. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/27595789.

Gerardi:2008:CHC

[GT08] A. Gerardi and P. Tardelli. Controlled heterogeneous collection: The role of occupation numbers. Journal of Applied Probability, 45(3):595–609, September 2008. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.

org/stable/27595973.

Geiger:2014:LMC

[GT14] Bernhard C. Geiger and Christoph Temmel. Lumpings of Markov chains, entropy rate preservation, and higher-order lumpability. Journal of Applied Probability, 51(4):1114–1132, December 2014. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://projecteuclid.org/
euclid.jap/1421763331.

Goodman:2018:ELM

[GT18] B. Goodman and P. F. Tupper. Effects of limiting memory capacity on the behaviour of exemplar dynamics. Journal of Applied Probability, 55(2):390–401, June 2018. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL https:/

/www.cambridge.org/core/journals/journal-of-applied-

probability/article/effects-of-limiting-memory-capacity-
on-the-behaviour-of-exemplar-dynamics/3231BC7D89C3B8D666A08D0A2C1D3A43.

Gong:1992:MSG

[GU92] Wei-Bo Gong and H. U. The MacLaurin series for the GI/G/1 queue. Journal of Applied Probability, 29(1):176–184, March 1992. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3214801.

Guess:1981:IPS

[Gue81] Harry A. Guess. An invariance principle for solutions to stochastic difference equations. Journal of Applied Probability, 18 (2):548–553, June 1981. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3213304.

Guerry:1991:MPS

[Gue91] M. A. Guerry. Monotonicity property of t-step maintainable structures in three-grade manpower systems: A counterexample. Journal of Applied Probability, 28(1):221–224, March 1991. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/3214752. See

correction [Ano91f].

Guerry:1997:PCP

[Gue97] M. A. Guerry. Properties of calculated predictions of grade sizes and the associated integer valued vectors. Journal of Applied Probability, 34(1):94–100, March 1997. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/3215178.

Guerry:1999:UFS

[Gue99] M. A. Guerry. Using fuzzy sets in manpower planning. Journal of Applied Probability, 36(1):155–162, March 1999. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3215411.

Guerry:2013:EPD

[Gue13] Marie-Anne Guerry. On the embedding problem for discretetime Markov chains. Journal of Applied Probability, 50(4):918– 930, December 2013. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://projecteuclid.

org/euclid.jap/1389370090.

Guibert:1994:OPU

[Gui94] Jacky Guibert. Overflow probability upper bound in fluid queues with general on-off sources. Journal of Applied Probability, 31 (4):1134–1139, December 1994. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor. org/stable/3215338. See correction [Ano95e].

Guillas:2002:DSH

[Gui02] Serge Guillas. Doubly stochastic Hilbertian processes. Journal of Applied Probability, 39(3):566–580, September 2002. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3216079.

Guo:2001:ESO

[Guo01] Xin Guo. An explicit solution to an optimal stopping problem with regime switching. Journal of Applied Probability, 38 (2):464–481, June 2001. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3215900.

Guo:2002:SRM

[Guo02] Xin Guo. Some risk management problems for firms with internal competition and debt. Journal of Applied Probability, 39(1): 55–69, March 2002. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3215917.

Gupta:1984:RBO

[Gup84] Ramesh C. Gupta. Relationships between order statistics and record values and some characterization results. Journal of Applied Probability, 21(2):425–430, June 1984. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/3213652.

Gupta:2013:SCR

[Gup13] Nitin Gupta. Stochastic comparisons of residual lifetimes and inactivity times of coherent systems. Journal of Applied Probability, 50(3):848–860, September 2013. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http:/

/projecteuclid.org/euclid.jap/1378401240.

Gut:1975:WCF

[Gut75] Allan Gut. Weak convergence and first passage times. Journal of Applied Probability, 12(2):324–332, June 1975. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3212446.

Gut:2003:MPR

[Gut03] Allan Gut. On the moment problem for random sums. Journal of Applied Probability, 40(3):797–802, September 2003. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3215953.

Gut:2010:LTG

[Gut10] Allan Gut. Limit theorems for a generalized St. Petersburg Game. Journal of Applied Probability, 47(3):752–760, September 2010. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/20778514.

Garavaglia:2018:TGC

[GvdH18] A. Garavaglia and R. van der Hofstad. From trees to graphs: collapsing continuous-time branching processes. Journal of Applied Probability, 55(3):900–919, September 2018. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL https:/

/www.cambridge.org/core/journals/journal-of-applied-
probability/article/from-trees-to-graphs-collapsing-
continuoustime-branching-processes/E60C37266CBDA4AA7FCC0EA7EBCDEADC.

Galbraith:1976:TDB

[GW76] R. F. Galbraith and D. Walley. On a two-dimensional binary process. Journal of Applied Probability, 13(3):548–557, September 1976. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3212474.

Galbraith:1980:EPT

[GW80] R. F. Galbraith and D. Walley. Ergodic properties of a twodimensional binary process. Journal of Applied Probability, 17 (1):124–133, March 1980. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3212930.

Galbraith:1982:FPU

[GW82] R. F. Galbraith and D. Walley. Further properties for unilateral binary processes. Journal of Applied Probability, 19(2):332– 343, June 1982. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3213485.

Gates:1985:AAR

[GW85] D. J. Gates and M. Westcott. Accurate and asymptotic results for distributions of scan statistics. Journal of Applied Probability, 22(3):531–542, September 1985. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.

org/stable/3213858.

Greenberg:1987:UBP

[GW87] Betsy S. Greenberg and Ronald W. Wolff. An upper bound on the performance of queues with returning customers. Journal of Applied Probability, 24(2):466–475, June 1987. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3214270.

Gates:1994:RTW

[GW94a] David Gates and Mark Westcott. Replacement of train wheels: An application of dynamic reversal of a Markov process. Journal of Applied Probability, 31(1):1–8, March 1994. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3215230.

Glynn:1994:LAS

[GW94b] Peter W. Glynn and Ward Whitt. Logarithmic asymptotics for steady-state tail probabilities in a single-server queue. Journal of Applied Probability, 31A:131–156, 1994. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/3214953.

Ghahramani:2005:FMF

[GW05] Saeed Ghahramani and Ronald W. Wolff. On finite moments of full busy periods of GI/G/c queues. Journal of Applied Probability, 42(1):275–278, March 2005. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/30040786.

Grubel:2011:MTE

[GW11] Rudolf Grübel and Hendrik Wegener. Matchmaking and testing for exponentiality in the M/G/∞ queue. Journal of Applied Probability, 48(1):131–144, March 2011. CODEN JPRBAM.

ISSN 0021-9002 (print), 1475-6072 (electronic). URL http:/

/www.jstor.org/stable/29777450.

Gerber:2017:SRI

[GW17] Mathieu Gerber and Nick Whiteley. Stability with respect to initial conditions in V -norm for nonlinear filters with ergodic observations. Journal of Applied Probability, 54(1):118–133, March 2017. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL https:/

/www.cambridge.org/core/journals/journal-of-applied-
probability/article/stability-with-respect-to-initial-
conditions-in-vnorm-for-nonlinear-filters-with-ergodic-
observations/C12FF4FC7DF6D91AC7FF9A9C68F2D8AE.

Gail:1979:SGB

[GWMO79] M. H. Gail, G. H. Weiss, N. Mantel, and S. J. O’Brien. A solution to the generalized birthday problem with application to allozyme screening for cell culture contamination. Journal of Applied Probability, 16(2):242–251, June 1979. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/3212893.

Gray:1980:LBC

[GWS80] Lawrence Gray, John C. Wierman, and R. T. Smythe. Lower bounds for the critical probability in percolation models with oriented bonds. Journal of Applied Probability, 17(4):979–986, December 1980. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3213207.

George:2012:EOA

[GXS12] David K. George, Cathy H. Xia, and Mark S. Squillante. Exactorder asymptotic analysis for closed queueing networks. Journal of Applied Probability, 49(2):503–520, June 2012. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/41713782.

Gong:2018:RAS

[GXY18] Ming Gong, Min Xie, and Yaning Yang. Reliability assessment of system under a generalized run shock model. Journal of Applied Probability, 55(4):1249–1260, December 2018. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL https://www.cambridge.org/core/journals/journal-
of-applied-probability/article/reliability-assessment-
of-system-under-a-generalized-run-shock-model/C630F3546FBFCE12F7814929A1BEDD26.

Gani:1965:SBD

[GY65] J. Gani and G. F. Yeo. Some birth-death and mutation models for phage reproduction. Journal of Applied Probability, 2(1):150– 161, June 1965. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3211880.

Gani:1995:EBD

[GY95a] J. Gani and Sid Yakowitz. Error bounds for deterministic approximations to Markov processes, with applications to epidemic models. Journal of Applied Probability, 32(4):1063–1076, December 1995. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3215220.

Glasserman:1995:SVD

[GY95b] Paul Glasserman and David D. Yao. Stochastic vector difference equations with stationary coefficients. Journal of Applied Probability, 32(4):851–866, December 1995. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/3215199.

Glasserman:2004:OCT

[GY04] Paul Glasserman and David D. Yao. Optimal couplings are totally positive and more. Journal of Applied Probability, 41A: 321–332, 2004. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3215986.

Gao:2008:FLD

[GY08] Fuqing Gao and Jun Yan. Functional large deviations and moderate deviations for Markov-modulated risk models with reinsurance. Journal of Applied Probability, 45(3):800–817, September 2008. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/27595987.

Grey:1994:FLP

[GZ94] D. R. Grey and Lu Zhunwei. The fractional linear probability generating function in the random environment branching process. Journal of Applied Probability, 31(1):38–47, March 1994.

CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/3215233.

Guo:2002:DSC

[GZ02] Xianping Guo and Weiping Zhu. Denumerable-state continuoustime Markov decision processes with unbounded transition and reward rates under the discounted criterion. Journal of Applied Probability, 39(2):233–250, June 2002. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3216092.

Guo:2006:AOM

[GZ06] Xianping Guo and Quanxin Zhu. Average optimality for Markov decision processes in Borel spaces: A new condition and approach. Journal of Applied Probability, 43(2):318–334, June 2006. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/27595734.

Gao:2016:LTS

[GZ16] Zhenlong Gao and Yanhua Zhang. Limit theorems for a supercritical Poisson random indexed branching process. Journal of Applied Probability, 53(1):307–314, March 2016. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://projecteuclid.org/euclid.jap/1457470577.

Gzyl:1977:RSS

[Gzy77] Cristina Gzyl. Regenerative sets and subordinators. Journal of Applied Probability, 14(1):217–220, March 1977. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3213277.

Haukaas:1996:GFD

[HA96] Harald Haukås and Terje Aven. A general formula for the downtime distribution of a parallel system. Journal of Applied Probability, 33(3):772–785, September 1996. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3215358.

Haas:1992:MMR

[Haa92] Peter J. Haas. The maximum and mean of a random length sequence. Journal of Applied Probability, 29(2):460–466, June 1992. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3214583.

Hadidi:1969:STD

[Had69] Nasser Hadidi. On the service time distribution and the waiting time process of a potentially infinite capacity queueing system. Journal of Applied Probability, 6(3):594–603, December 1969. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/3212105.

Hadidi:1974:BPQ

[Had74] Nasser Hadidi. Busy period of queues with state dependent arrival and service rates. Journal of Applied Probability, 11(4): 842–848, December 1974. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3212569.

Hadjiliadis:2005:OCD

[Had05] Olympia Hadjiliadis. Optimality of the 2-CUSUM drift equalizer rules for detecting two-sided alternatives in the Brownian motion model. Journal of Applied Probability, 42(4):1183–1193, December 2005. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
30040837.

Haigh:1970:RRT

[Hai70] John Haigh. The recovery of the root of a tree. Journal of Applied Probability, 7(1):79–88, April 1970. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/3212150.

Haigh:1978:ABD

[Hai78] John Haigh. The asymptotic behaviour of a divergent linear birth and death process. Journal of Applied Probability, 15 (1):187–191, March 1978. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3213250.

Haigh:1981:REI

[Hai81] John Haigh. Random exchanges of information. Journal of Applied Probability, 18(3):743–746, September 1981. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3213330.

Haigh:1983:MMS

[Hai83] John Haigh. Maintainability of manpower structures: Counterexamples, results and conjectures. Journal of Applied Probability, 20(3):700–705, September 1983. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3213905. See correction [Ano86h].

Haigh:1988:DES

[Hai88] John Haigh. The distribution of evolutionarily stable strategies. Journal of Applied Probability, 25(2):233–246, June 1988. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/3214432.

Haigh:1989:HLS

[Hai89] John Haigh. How large is the support of an ESS? Journal of Applied Probability, 26(1):164–170, March 1989. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3214326.

Haigh:1990:RPR

[Hai90] John Haigh. Random polymorphisms and random evolutionarily stable strategies: A comparison. Journal of Applied Probability, 27(4):737–755, December 1990. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.

org/stable/3214819.

Haigh:1992:FCM

[Hai92] John Haigh. Further counterexamples to the monotonicity property of t-step maintainable structures. Journal of Applied Probability, 29(2):441–447, June 1992. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3214580.

Haigh:1996:MPW

[Hai96] John Haigh. More on n-point, win-by-k games. Journal of Applied Probability, 33(2):382–387, June 1996. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/3215061.

Haigh:2003:IIR

[Hai03] John Haigh. An intuitive insight into a result of Cowan. Journal of Applied Probability, 40(1):271–272, March 2003. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3215855.

Hajek:1982:BDP

[Haj82] Bruce Hajek. Birth-and-death processes on the integers with phases and general boundaries. Journal of Applied Probability, 19(3):488–499, September 1982. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.

org/stable/3213508.

Hall:1969:ESW

[Hal69] W. J. Hall. Embedding submartingales in Wiener processes with drift, with applications to sequential analysis. Journal of Applied Probability, 6(3):612–632, December 1969. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/3212107.

Hall:1970:WEC

[Hal70] W. J. Hall. On Wald’s equations in continuous time. Journal of Applied Probability, 7(1):59–68, April 1970. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/3212148.

Hall:1978:RLT

[Hal78] Peter Hall. Representations and limit theorems for extreme value distributions. Journal of Applied Probability, 15(3):639–644, September 1978. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3213128.

Halfin:1979:SPM

[Hal79a] S. Halfin. A singular property of multiserver systems with constant service time. Journal of Applied Probability, 16(2):362– 372, June 1979. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3212903.

Hall:1979:RCN

[Hal79b] Peter Hall. On the rate of convergence of normal extremes. Journal of Applied Probability, 16(2):433–439, June 1979. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3212912.

Hall:1981:DRL

[Hal81] Peter Hall. On the distribution of random lines. Journal of Applied Probability, 18(3):606–616, September 1981. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3213315.

Hall:1982:ATB

[Hal82a] Glen Richard Hall. Acute triangles in the n-ball. Journal of Applied Probability, 19(3):712–715, September 1982. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3213533.

Hall:1982:SVE

[Hal82b] Peter Hall. On Starr and Vardi’s estimates of the number of transmission sources. Journal of Applied Probability, 19(1):52– 63, March 1982. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3213915.

Hall:1983:IRF

[Hal83a] Peter Hall. On the interpretation of random fluctuations in competing chemical systems. Journal of Applied Probability, 20(4): 877–883, December 1983. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3213599.

Hall:1983:RBP

[Hal83b] Peter Hall. On the roles of the Bessel and Poisson distributions in chemical kinetics. Journal of Applied Probability, 20(3):585–599, September 1983. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3213895.

Hall:1984:MVV

[Hal84] Peter Hall. Mean and variance of vacancy for distribution of k-dimensional spheres within k-dimensional space. Journal of Applied Probability, 21(4):738–752, December 1984. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3213692.

Halfin:1985:SQP

[Hal85] Shlomo Halfin. The shortest queue problem. Journal of Applied Probability, 22(4):865–878, December 1985. CODEN JPRBAM.

ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/3213954.

Halfin:1989:NFP

[Hal89] Shlomo Halfin. Next-fit bin packing with random piece sizes. Journal of Applied Probability, 26(3):503–511, September 1989. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/3214408.

Hammersley:1982:CPS

[Ham82] J. M. Hammersley. Critical phenomena in semi-infinite systems. Journal of Applied Probability, 19A:327–331, 1982. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3213572.

Hammersley:1988:MPS

[Ham88] J. M. Hammersley. Mesoadditive processes and the specific conductivity of lattices. Journal of Applied Probability, 25A: 347–358, 1988. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3214168.

Hambly:1992:LDS

[Ham92] Ben Hambly. On the limiting distribution of a supercritical branching process in a random environment. Journal of Applied Probability, 29(3):499–518, September 1992. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/3214889.

Hambly:1995:CTB

[Ham95] B. M. Hambly. On constant tail behaviour for the limiting random variable in a supercritical branching process. Journal of Applied Probability, 32(1):267–273, March 1995. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3214935.

Hammarlid:2005:TEF

[Ham05] Ola Hammarlid. Tools to estimate the first passage time to a convex barrier. Journal of Applied Probability, 42(1):61–81, March 2005. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/30040770.

Hannan:1965:GRA

[Han65] E. J. Hannan. Group representations and applied probability. Journal of Applied Probability, 2(1):1–68, June 1965. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3211873.

Hannan:1967:MWS

[Han67] E. J. Hannan. Measurement of a wandering signal amid noise. Journal of Applied Probability, 4(1):90–102, April 1967. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3212301.

Hannan:1971:NLT

[Han71] E. J. Hannan. Non-linear time series regression. Journal of Applied Probability, 8(4):767–780, December 1971. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/3212240. See correction

[Ano64c].

Hannan:1973:ATL

[Han73a] E. J. Hannan. The asymptotic theory of linear time-series models. Journal of Applied Probability, 10(1):130–145, March 1973. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/3212501. See

correction [Ano73g].

Hannan:1973:EF

[Han73b] E. J. Hannan. The estimation of frequency. Journal of Applied Probability, 10(3):510–519, September 1973. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/3212772.

Hannan:1982:TAA

[Han82] E. J. Hannan. Testing for autocorrelation and Akaike’s criterion. Journal of Applied Probability, 19A:403–412, 1982. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/3213579.

Hanschke:1987:EFC

[Han87] Thomas Hanschke. Explicit formulas for the characteristics of the M/M/2/2 queue with repeated attempts. Journal of Applied Probability, 24(2):486–494, June 1987. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3214272.

Hannan:1989:OSM

[Han89] E. J. Hannan. Obituary: Some memories of Pat Moran. Journal of Applied Probability, 26(1):215–218, March 1989. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/3214334. See correction

[Ano89f].

Hansen:1990:FCL

[Han90] Jennie C. Hansen. A functional central limit theorem for the Ewens sampling formula. Journal of Applied Probability, 27(1): 28–43, March 1990. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3214593.

Hanschke:1992:MCG

[Han92] Thomas Hanschke. Markov chains and generalized continued fractions. Journal of Applied Probability, 29(4):838–849, December 1992. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3214716.

Hyytia:2012:VFM

[HAPV12] Esa Hyytiä, Samuli Aalto, Aleksi Penttinen, and Jorma Virtamo. On the value function of the M/G/1 FCFS and LCFS queues. Journal of Applied Probability, 49(4):1052–1071, December 2012. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/41713828.

Harris:1965:DCB

[Har65] T. E. Harris. Diffusion with “collisions” between particles. Journal of Applied Probability, 2(2):323–338, December 1965. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/3212197.

Harper:1967:NCM

[Har67] L. H. Harper. A necessary condition on minimal cube numberings. Journal of Applied Probability, 4(2):397–401, August 1967. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/3212033.

Harrison:1973:ALQ

[Har73a] J. Michael Harrison. Assembly-like queues. Journal of Applied Probability, 10(2):354–367, June 1973. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3212352.

Harrison:1973:HTA

[Har73b] J. Michael Harrison. The heavy traffic approximation for single server queues in series. Journal of Applied Probability, 10(3): 613–629, September 1973. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3212781.

Harrison:1973:LTP

[Har73c] J. Michael Harrison. A limit theorem for priority queues in heavy traffic. Journal of Applied Probability, 10(4):907–912, December 1973. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3212396.

Hartley:1980:SPW

[Har80] Roger Hartley. A simple proof of Whittle’s bridging condition in dynamic programming. Journal of Applied Probability, 17(4): 1114–1116, December 1980. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3213223.

Harrison:1981:TBQ

[Har81] P. G. Harrison. Transient behaviour of queueing networks. Journal of Applied Probability, 18(2):482–490, June 1981. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3213294.

Harel:1990:CRS

[Har90a] Arie Harel. Convexity results for single-server queues and for multiserver queues with constant service times. Journal of Applied Probability, 27(2):465–468, June 1990. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/3214668.

Harel:1990:HBE

[Har90b] Arie Harel. On honest Bernoulli excursions. Journal of Applied Probability, 27(2):475–476, June 1990. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3214670.

Harrison:1990:LTI

[Har90c] Peter G. Harrison. Laplace transform inversion and passage-time distributions in Markov processes. Journal of Applied Probability, 27(1):74–87, March 1990. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.

org/stable/3214596.

Hartfiel:1991:SLM

[Har91] D. J. Hartfiel. Sequential limits in Markov set-chains. Journal of Applied Probability, 28(4):910–913, December 1991. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3214695.

Hartfiel:1994:HMC

[Har94] D. J. Hartfiel. Homogeneous Markov chains with bounded transition matrix. Journal of Applied Probability, 31(2):362– 372, June 1994. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3215029.

Hartvigsen:2009:AGE

[Har09] David Hartvigsen. The action gambler and equal-sized wagering. Journal of Applied Probability, 46(1):35–54, March 2009. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/25662406.

Hasofer:1964:SSQ

[Has64] A. M. Hasofer. On the single-server queue with nonhomogeneous Poisson input and general service time. Journal of Applied Probability, 1(2):369–384, December 1964. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3211866.

Hasofer:1965:SPR

[Has65] A. M. Hasofer. Some perturbation results for the single-server queue with Poisson input. Journal of Applied Probability, 2(2): 462–466, December 1965. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3212207.

Hasofer:1976:MNM

[Has76] A. M. Hasofer. The mean number of maxima above high levels in Gaussian random fields. Journal of Applied Probability, 13 (2):377–379, June 1976. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3212844.

Hasofer:1979:UPE

[Has79] A. M. Hasofer. A uniqueness problem for the envelope of an oscillatory process. Journal of Applied Probability, 16(4):822–829, December 1979. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3213148.

Haslett:1982:NBT

[Has82a] J. Haslett. New bounds for the thermal energy storage process with stationary input. Journal of Applied Probability, 19(4): 894–899, December 1982. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3213847.

Hasofer:1982:STM

[Has82b] A. M. Hasofer. Simple trigonometric models for narrow-band stationary processes. Journal of Applied Probability, 19A: 333–343, 1982. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3213573.

Hautphenne:2012:EPS

[Hau12] Sophie Hautphenne. Extinction probabilities of supercritical decomposable branching processes. Journal of Applied Probability, 49(3):639–651, September 2012. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.

org/stable/41713795.

Haviv:1991:TSP

[Hav91] Moshe Haviv. Two sufficient properties for the insensitivity of a class of queueing models. Journal of Applied Probability, 28(3): 664–672, September 1991. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3214500.

Hawkes:1968:GAR

[Haw68] A. G. Hawkes. Gap-acceptance in road traffic. Journal of Applied Probability, 5(1):84–92, April 1968. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3212079.

Hawkes:1970:BSM

[Haw70] A. G. Hawkes. Bunching in a semi-Markov process. Journal of Applied Probability, 7(1):175–182, April 1970. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3212158.

Hayakawa:2000:NCP

[Hay00] Yu Hayakawa. A new characterisation property of mixed Poisson processes via Berman’s theorem. Journal of Applied Probability, 37(1):261–268, March 2000. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3215674.

Hardin:1975:FRK

[HB75] Jay C. Hardin and Thomas J. Brown. Further results on Kendall’s autoregressive series. Journal of Applied Probability, 12(1):180–182, March 1975. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/ stable/3212426. See corrections [Ano75f].

Hines:1983:ESS

[HB83a] W. G. S. Hines and D. T. Bishop. Evolutionarily stable strategies in diploid populations with general inheritance patterns. Journal of Applied Probability, 20(2):395–399, June 1983. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/3213812.

Hines:1983:LES

[HB83b] W. G. S. Hines and D. T. Bishop. On learning and the evolutionarily stable strategy. Journal of Applied Probability, 20(3): 689–695, September 1983. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3213903.

Hines:1984:LSE

[HB84] W. G. S. Hines and D. T. Bishop. On the local stability of an evolutionarily stable strategy in a diploid population. Journal of Applied Probability, 21(2):215–224, June 1984. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3213634.

Huggins:1999:EBA

[HB99] R. M. Huggins and I. V. Basawa. Extensions of the bifurcating autoregressive model for cell lineage studies. Journal of Applied Probability, 36(4):1225–1233, December 1999. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3215591.

Huisman:2002:STD

[HB02] Tijs Huisman and Richard J. Boucherie. The sojourn time distribution in an infinite server resequencing queue with dependent interarrival and service times. Journal of Applied Probability, 39 (3):590–603, September 2002. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.

org/stable/3216081.

Hernandez-Bustos:2023:PMU

[HBHH23] Diego Hernández-Bustos and Daniel Hernández-Hernández. Portfolio management under drawdown constraint in discretetime financial markets. Journal of Applied Probability, 60(1):127–147, March 2023. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL https:/

/www.cambridge.org/core/journals/journal-of-applied-
probability/article/portfolio-management-under-drawdown-
constraint-in-discretetime-financial-markets/1C6B3E7D6BB8C56F1CF82F90BC849642.

Haug:2007:ECT

[HC07] Stephan Haug and Claudia Czado. An exponential continuoustime GARCH process. Journal of Applied Probability, 44(4): 960–976, December 2007. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/27595900.

Hwang:2002:WTA

[HCK02] Gang Uk Hwang, Bong Dae Choi, and Jae-Kyoon Kim. The waiting time analysis of a discrete-time queue with arrivals as a discrete autoregressive process of order 1. Journal of Applied Probability, 39(3):619–629, September 2002. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/3216083.

Hsiau:2021:SOB

[HCY21] Shoou-Ren Hsiau, May-Ru Chen, and Yi-Ching Yao. Stochastic ordering for birth–death processes with killing. Journal of Applied Probability, 58(3):708–720, September 2021. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL https://www.cambridge.org/core/journals/journal-
of-applied-probability/article/stochastic-ordering-for-
birthdeath-processes-with-killing/628E99F1574211B4269C70E18A30F90D.

Hall:1971:FQP

[HD71] William K. Hall and Ralph L. Disney. Finite queues in parallel under a generalized channel selection rule. Journal of Applied Probability, 8(2):413–416, June 1971. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3211912.

Haezendonck:1980:CCE

[HD80] J. Haezendonck and F. De Vylder. A comparison criterion for explosions in point processes. Journal of Applied Probability, 17 (4):1102–1107, December 1980. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.

org/stable/3213221.

He:2001:PMN

[He01] Xinyu He. A probabilistic method for Navier–Stokes vortices. Journal of Applied Probability, 38(4):1059–1066, December 2001. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/3215784.

He:2003:CMG

[He03] Qi-Ming He. The classification of matrix GI/M/1-Type Markov chains with a tree structure and its applications to queueing. Journal of Applied Probability, 40(4):1087–1102, December 2003. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/3216062.

Hitz:2016:OCR

[HE16] Adrien Hitz and Robin Evans. One-component regular variation and graphical modeling of extremes. Journal of Applied Probability, 53(3):733–746, September 2016. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

projecteuclid.org/euclid.jap/1476370773.

He:2021:BMS

[He21] Qi-Ming He. Bounds on the mean and squared coefficient of variation of phase-type distributions. Journal of Applied Probability, 58(4):880–889, December 2021. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL https:/

/www.cambridge.org/core/journals/journal-of-applied-
probability/article/bounds-on-the-mean-and-squared-coefficient-
of-variation-of-phasetype-distributions/5F080D4AE32CA0A138ECDD976DEA96F2.

Heathcote:1965:MQG

[Hea65] C. R. Heathcote. On the maximum of the queue GI/M/1. Journal of Applied Probability, 2(1):206–214, June 1965. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3211885.

Heathcote:1967:CEC

[Hea67] C. R. Heathcote. Complete exponential convergence and some related topics. Journal of Applied Probability, 4(2):217–256, August 1967. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3212019.

Heathcote:1982:LRF

[Hea82] C. R. Heathcote. Linear regression by functional least squares. Journal of Applied Probability, 19A:225–239, 1982. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3213563.

Heidergott:2001:WDA

[Hei01] Bernd Heidergott. A weak derivative approach to optimization of threshold parameters in a multicomponent maintenance system. Journal of Applied Probability, 38(2):386–406, June 2001. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/3215895.

Heinzmann:2009:ETM

[Hei09] Dominik Heinzmann. Extinction times in multitype Markov branching processes. Journal of Applied Probability, 46(1):296– 307, March 2009. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
25662422.

Helms:1967:BFB

[Hel67] L. L. Helms. Biharmonic functions and Brownian motion. Journal of Applied Probability, 4(1):130–136, April 1967. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3212305.

Helgert:1970:SRV

[Hel70] H. J. Helgert. On sums of random variables defined on a twostate Markov chain. Journal of Applied Probability, 7(3):761– 765, December 1970. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3211953.

Hellman:1971:OSR

[Hel71] Olavi Hellman. Optimal search for a randomly moving object in a special case. Journal of Applied Probability, 8(3):606–611, September 1971. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3212183.

Helms:1973:SFH

[Hel73] L. L. Helms. Stochastic formation of hierarchies. Journal of Applied Probability, 10(1):27–38, March 1973. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/3212493.

Hemker:1990:NST

[Hem90] Josef Hemker. A note on sojourn times in queuing networks with multiserver nodes. Journal of Applied Probability, 27(2):469– 474, June 1990. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3214669.

Hendricks:1972:SDI

[Hen72] W. J. Hendricks. The stationary distribution of an interesting Markov chain. Journal of Applied Probability, 9(1):231– 233, March 1972. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3212655.

Hendricks:1973:ETC

[Hen73a] W. J. Hendricks. An extension of a theorem concerning an interesting Markov chain. Journal of Applied Probability, 10(4):

886–890, December 1973. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3212392.

Henke:1973:EVS

[Hen73b] M. Henke. Expectations and variances of stopping variables in sequential selection processes. Journal of Applied Probability, 10 (4):786–806, December 1973. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.

org/stable/3212382.

Henderson:1979:SCB

[Hen79] W. Henderson. A solution of the carrier-borne epidemic. Journal of Applied Probability, 16(3):641–645, September 1979. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3213091.

Henery:1981:PPN

[Hen81] R. J. Henery. Place probabilities in normal order statistics models for horse races. Journal of Applied Probability, 18(4):839–852, December 1981. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3213059.

Henze:1982:LDM

[Hen82] Norbert Henze. The limit distribution for maxima of ’weighted’ R th-nearest-neighbour distances. Journal of Applied Probability, 19(2):344–354, June 1982. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.

org/stable/3213486.

Henderson:1983:NSI

[Hen83a] W. Henderson. Non-standard insensitivity. Journal of Applied Probability, 20(2):288–296, June 1983. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3213802.

Henery:1983:PPG

[Hen83b] Robert J. Henery. Permutation probabilities for gamma random variables. Journal of Applied Probability, 20(4):822–834, December 1983. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3213593.

Henze:1983:RTC

[Hen83c] Norbert Henze. Random triangles in convex regions. Journal of Applied Probability, 20(1):111–125, March 1983. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3213725.

Henstridge:1986:DSP

[Hen86a] John Henstridge. Data structures: The problem of time series computing. Journal of Applied Probability, 23A:407–411, 1986. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/3214368.

Henze:1986:MVR

[Hen86b] Norbert Henze. On the moments of vacancy of random arcs on the circle. Journal of Applied Probability, 23(3):837–840, September 1986. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3214022.

Henze:1986:PRP

[Hen86c] Norbert Henze. On the probability that a random point is the jth nearest neighbour to its own kth nearest neighbour. Journal of Applied Probability, 23(1):221–226, March 1986. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3214132.

Henderson:1993:QNN

[Hen93] W. Henderson. Queueing networks with negative customers and negative queue lengths. Journal of Applied Probability, 30(4): 931–942, December 1993. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3214523.

Henze:2018:CST

[Hen18] Norbert Henze. On the consistency of the spacings test for multivariate uniformity, including on manifolds. Journal of Applied Probability, 55(2):659–665, June 2018. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL https:/

/www.cambridge.org/core/journals/journal-of-applied-
probability/article/on-the-consistency-of-the-spacings-
test-for-multivariate-uniformity-including-on-manifolds/
4173745E8D6B756B8A96CAA1B87A8AD3.

Herbert:1972:IDD

[Her72] H. G. Herbert. An infinite discrete dam with dependent inputs. Journal of Applied Probability, 9(2):404–413, June 1972. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/3212808.

Herbert:1974:SAR

[Her74] H. G. Herbert. Some approximate results for storage systems with continuous inputs. Journal of Applied Probability, 11(2): 332–344, June 1974. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3212754.

Herrndorf:1985:FPP

[Her85] Norbert Herrndorf. First-passage percolation processes with finite height. Journal of Applied Probability, 22(4):766–775, December 1985. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3213944.

Herkenrath:1994:GAE

[Her94] Ulrich Herkenrath. Generalized adaptive exponential smoothing procedures. Journal of Applied Probability, 31(3):673–690, September 1994. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3215147.

Herkenrath:1999:GAE

[Her99] Ulrich Herkenrath. Generalized adaptive exponential smoothing of observations from an ergodic hidden Markov model. Journal of Applied Probability, 36(4):987–998, December 1999. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3215572.

Herkenrath:2003:UEM

[Her03] Ulrich Herkenrath. On the uniform ergodicity of Markov processes of order 2. Journal of Applied Probability, 40(2):455– 472, June 2003. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3215803.

Heunis:1987:NLF

[Heu87] A. J. Heunis. Non-linear filtering of rare events with large signalto-noise ratio. Journal of Applied Probability, 24(4):929–948, December 1987. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3214217.

Heyde:1964:SWT

[Hey64] C. C. Heyde. On the stationary waiting time distribution in the queue GI/G/1. Journal of Applied Probability, 1(1):173– 176, June 1964. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3212069.

Heyde:1967:LTR

[Hey67] C. C. Heyde. A limit theorem for random walks with drift. Journal of Applied Probability, 4(1):144–150, April 1967. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/3212307.

Heyde:1968:EHR

[Hey68a] C. C. Heyde. An extension of the Hájek–Rényi inequality for the case without conditions. Journal of Applied Probability, 5 (2):481–483, August 1968. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/ stable/3212269. See correction [Ano71e].

Heyde:1968:CIL

[Hey68b] C. C. Heyde. On the converse to the iterated logarithm law. Journal of Applied Probability, 5(1):210–215, April 1968. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/3212088.

Heyde:1969:MSR

[Hey69] C. C. Heyde. On the maximum of sums of random variables and the supremum functional for stable processes. Journal of Applied Probability, 6(2):419–429, August 1969. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http:/

/www.jstor.org/stable/3212011.

Heyde:1970:RCR

[Hey70] C. C. Heyde. A rate of convergence result for the super-critical Galton–Watson process. Journal of Applied Probability, 7(2):

451–454, August 1970. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3211980.

Heyde:1971:SCL

[Hey71] C. C. Heyde. Some central limit analogues for super-critical Galton–Watson processes. Journal of Applied Probability, 8(1): 52–59, March 1971. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3211837.

Heyde:1973:ILR

[Hey73] C. C. Heyde. An iterated logarithm result for martingales and its application in estimation theory for autoregressive processes. Journal of Applied Probability, 10(1):146–157, March 1973. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/3212502.

Heyman:1974:ABP

[Hey74] D. P. Heyman. An approximation for the busy period of the M/G/1 queue using a diffusion model. Journal of Applied Probability, 11(1):159–169, March 1974. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3212592.

Heyde:1975:SSL

[Hey75] C. C. Heyde. A supplement to the Strong Law of Large Numbers. Journal of Applied Probability, 12(1):173–175, March 1975. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/3212424.

Heyde:1977:CLI

[Hey77] C. C. Heyde. On central limit and iterated logarithm supplements to the martingale convergence theorem. Journal of Applied Probability, 14(4):758–775, December 1977. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3213349.

Heyde:1980:PAF

[Hey80] C. C. Heyde. On a probabilistic analogue of the Fibonacci sequence. Journal of Applied Probability, 17(4):1079–1082, December 1980. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3213217.

Heyde:1981:FLB

[Hey81] C. C. Heyde. On Fibonacci (or lagged Bienaymé–Galton– Watson) branching processes. Journal of Applied Probability, 18 (3):583–591, September 1981. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.

org/stable/3213313.

Heyde:1982:EDR

[Hey82a] C. C. Heyde. The effect of differential reproductive rates on the survival of a gene represented in a founder population. Journal of Applied Probability, 19A:19–25, 1982. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/3213547.

Heyman:1982:RCA

[Hey82b] D. P. Heyman. On Ross’s conjectures about queues with nonstationary Poisson arrivals. Journal of Applied Probability, 19 (1):245–249, March 1982. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3213936.

Heyde:1986:UTS

[Hey86] C. C. Heyde. On the use of time series representations of population models. Journal of Applied Probability, 23A:345–353, 1986. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3214363.

Heyde:1988:STS

[Hey88] C. C. Heyde. Some thoughts on stationary processes and linear time series analysis. Journal of Applied Probability, 25A: 309–318, 1988. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3214165.

Heyman:1991:ASD

[Hey91] Daniel P. Heyman. Approximating the stationary distribution of an infinite stochastic matrix. Journal of Applied Probability, 28(1):96–103, March 1991. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3214743.

Heyde:1994:QLA

[Hey94] C. C. Heyde. A quasi-likelihood approach to estimating parameters in diffusion-type processes. Journal of Applied Probability, 31A:283–290, 1994. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3214962.

Heyman:1995:DTI

[Hey95] Daniel P. Heyman. A decomposition theorem for infinite stochastic matrices. Journal of Applied Probability, 32(4):893–901, December 1995. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3215202.

Heyde:1999:RAM

[Hey99] C. C. Heyde. A risky asset model with strong dependence through fractal activity time. Journal of Applied Probability, 36(4):1234–1239, December 1999. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3215592.

Heyde:2002:MLR

[Hey02] C. C. Heyde. On modes of long-range dependence. Journal of Applied Probability, 39(4):882–888, December 2002. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3216010.

Henderson:1970:NQS

[HF70] J. Henderson and P. D. Finch. A note on the queueing system E_{k} /G/1. Journal of Applied Probability, 7(2):473–475, August 1970. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3211985.

Hochman:1971:SRP

[HF71] Howard G. Hochman and Stephen E. Fienberg. Some renewal process models for single neuron discharge. Journal of Applied Probability, 8(4):802–808, December 1971. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/3212243.

Hansen:1991:SMP

[HF91] B. G. Hansen and J. B. G. Frenk. Some monotonicity properties of the delayed renewal function. Journal of Applied Probability, 28(4):811–821, December 1991. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.

org/stable/3214684.

Hazra:2019:CLC

[HF19] Nil Kamal Hazra and Maxim Finkelstein. Comparing lifetimes of coherent systems with dependent components operating in random environments. Journal of Applied Probability, 56(3):937–957, September 2019. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL https:/

/www.cambridge.org/core/journals/journal-of-applied-
probability/article/comparing-lifetimes-of-coherent-
systems-with-dependent-components-operating-in-random-
environments/A54608A2241BEA0E53D71F6BA82359EC.

Hertzberg:1980:FOS

[HG80] Richard C. Hertzberg and Vincent F. Gallucci. First-order stochastic chemical reactions and oscillations in the variance. Journal of Applied Probability, 17(4):1087–1093, December 1980. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/3213219.

Hill:1981:RWA

[HG81] James M. Hill and Chandra M. Gulati. The random walk associated with the game of roulette. Journal of Applied Probability, 18(4):931–936, December 1981. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.

org/stable/3213067.

Hanqin:1990:SAM

[HGR90] Zhang Hanqin, Hsu Guanghui, and Wang Rongxin. Strong approximations for multiple channel queues in heavy traffic. Journal of Applied Probability, 27(3):658–670, September 1990. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/3214549.

Holst:1984:RCC

[HH84] L. Holst and J. Hüsler. On the random coverage of the circle. Journal of Applied Probability, 21(3):558–566, September 1984. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/3213617.

Hill:1985:SOO

[HH85a] Theodore P. Hill and Arie Hordijk. Selection of order of observation in optimal stopping problems. Journal of Applied Probability, 22(1):177–184, March 1985. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3213757.

Hines:1985:NLE

[HH85b] W. G. S. Hines and John Haigh. Non-linear ESS models and polymorphism. Journal of Applied Probability, 22(4):747–756, December 1985. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3213942.

Ho:1996:AJD

[HH96] Hwai-Chung Ho and Tailen Hsing. On the asymptotic joint distribution of the sum and maximum of stationary normal random variables. Journal of Applied Probability, 33(1):138–145, March 1996. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3215271.

Han:2003:SLW

[HH03] Qing Han and Katuomi Hirano. Sooner and later waiting time problems for patterns in Markov dependent trials. Journal of Applied Probability, 40(1):73–86, March 2003. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3215842.

He:2007:RPI

[HH07] Xiaoxia He and Yijun Hu. Ruin probability for the integrated Gaussian process with force of interest. Journal of Applied Probability, 44(3):685–694, September 2007. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http:/

/www.jstor.org/stable/27595875.

Hao:2023:GDS

[HH23] Nannan Hao and Markus Heydenreich. Graph distances in scale-free percolation: the logarithmic case. Journal of Applied Probability, 60(1):295–313, March 2023. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL https:/

/www.cambridge.org/core/journals/journal-of-applied-
probability/article/graph-distances-in-scalefree-percolation-
the-logarithmic-case/4FC4925135F1BBF85D4EF01288A41170.

Herbert:2001:SRT

[HI01] Julian Herbert and Valerie Isham. A study of the role of the transmission mechanism in macroparasite aggregation. Journal of Applied Probability, 38A:249–262, 2001. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/3215884.

Hiesmayr:2020:ARH

[HI20] Ella Hiesmayr and Ümit Islak. Asymptotic results on Hoppe trees and their variations. Journal of Applied Probability, 57(2):441–457, June 2020. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL https:/

/www.cambridge.org/core/journals/journal-of-applied-
probability/article/asymptotic-results-on-hoppe-trees-
and-their-variations/2DE4132127D3A6333ABF23F2CD0C3F27.

Hickey:1982:NMR

[Hic82] Raymond J. Hickey. A note on the measurement of randomness. Journal of Applied Probability, 19(1):229–232, March 1982. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/3213933.

Hickey:1983:MRS

[Hic83] Raymond J. Hickey. Majorisation, randomness and some discrete distributions. Journal of Applied Probability, 20(4):897– 902, December 1983. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3213602.

Hickey:1984:CMR

[Hic84] Raymond J. Hickey. Continuous majorisation and randomness. Journal of Applied Probability, 21(4):924–929, December 1984. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/3213709.

Hickey:1986:CDD

[Hic86] Raymond J. Hickey. Concepts of dispersion in distributions: A comparative note. Journal of Applied Probability, 23(4):914–921, December 1986. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3214465.

Hildebrand:1967:SFQ

[Hil67] David K. Hildebrand. Stability of finite queue, tandem server systems. Journal of Applied Probability, 4(3):571–583, December 1967. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3212222.

Hines:1980:SSC

[Hin80a] W. G. S. Hines. Strategy stability in complex populations. Journal of Applied Probability, 17(3):600–610, September 1980. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/3212954.

Hines:1980:TCP

[Hin80b] W. G. S. Hines. Three characterizations of population strategy stability. Journal of Applied Probability, 17(2):333–340, June 1980. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3213023.

Hines:1981:MSP

[Hin81] W. G. S. Hines. Multi-species population models and evolutionarily stable strategies. Journal of Applied Probability, 18 (2):507–513, June 1981. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3213297.

Hines:1982:MPE

[Hin82a] W. G. S. Hines. Mutations, perturbations and evolutionarily stable strategies. Journal of Applied Probability, 19(1):204– 209, March 1982. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/

3213929. See comment [Hin95].

Hines:1982:SSC

[Hin82b] W. G. S. Hines. Strategy stability in complex randomly mating diploid populations. Journal of Applied Probability, 19(3): 653–659, September 1982. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3213522.

Hines:1995:CBP

[Hin95] W. G. S. Hines. Comment on “Mutations, Perturbations and Evolutionarily Stable Strategies”. Journal of Applied Probability, 32(3):836–839, September 1995. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.

org/stable/3215134. See [Hin82a].

Hinz:2004:RET

[Hin04] Juri Hinz. A revenue-equivalence theorem for electricity auctions. Journal of Applied Probability, 41(2):299–312, June 2004. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/3216017.

Herkenrath:2003:NIP

[HIR03] Ulrich Herkenrath, Marius Iosifescu, and Andreas Rudolph. A note on invariance principles for iterated random functions. Journal of Applied Probability, 40(3):834–837, September 2003. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/3215959.

Hirsch:2016:BHP

[Hir16] Christian Hirsch. Bounded-hop percolation and wireless communication. Journal of Applied Probability, 53(3):833–845, September 2016. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-

6072 (electronic). URL http://projecteuclid.org/euclid.
jap/1476370779.

Hitchcock:1986:EPP

[Hit86] S. E. Hitchcock. Extinction probabilities in predator-prey models. Journal of Applied Probability, 23(1):1–13, March 1986. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/3214112.

Hitchcock:1988:OGA

[Hit88] Mavis Hitchcock. The origin and growth of the applied probability trust. Journal of Applied Probability, 25A:47–64, 1988. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/3214146.

Hitczenko:2010:TP

[Hit10] Paweł Hitczenko. On tails of perpetuities. Journal of Applied Probability, 47(4):1191–1194, December 2010. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/29777436.

Haggstrom:1999:PTR

[HJ99] Olle Häggström and Johan Jonasson. Phase transition in the random triangle model. Journal of Applied Probability, 36(4): 1101–1115, December 1999. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.

org/stable/3215581.

Hansen:2002:CRM

[HJ02] Jennie C. Hansen and Jerzy Jaworski. Compound random mappings. Journal of Applied Probability, 39(4):712–729, December 2002. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3215998.

Hobolth:2011:SSE

[HJ11] Asger Hobolth and Jens Ledet Jensen. Summary statistics for endpoint-conditioned continuous-time Markov chains. Journal of Applied Probability, 48(4):911–924, December 2011. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/23066433.

Hansen:2013:PSR

[HJ13a] Jennie C. Hansen and Jerzy Jaworski. Predecessors and successors in random mappings with exchangeable in-degrees. Journal of Applied Probability, 50(3):721–740, September 2013. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://projecteuclid.org/euclid.jap/1378401232.

Hu:2013:JRP

[HJ13b] Zechun Hu and Bin Jiang. On joint ruin probabilities of a two-dimensional risk model with constant interest rate. Journal of Applied Probability, 50(2):309–322, June 2013. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://projecteuclid.org/euclid.jap/1371648943.

Hashorva:2014:APT

[HJ14] Enkelejd Hashorva and Lanpeng Ji. Approximation of passage times of γ-reflected processes with FBM input. Journal of Applied Probability, 51(3):713–726, September 2014. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://projecteuclid.org/euclid.jap/1409932669.

Hua:2014:RBH

[HJL14] Lei Hua, Harry Joe, and Haijun Li. Relations between hidden regular variation and the tail order of copulas. Journal of Applied Probability, 51(1):37–57, March 2014. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

projecteuclid.org/euclid.jap/1395771412.

Horvath:2019:ABR

[HJL19] Blanka Horvath, Antoine Jacquier, and Chloé Lacombe. Asymptotic behaviour of randomised fractional volatility models. Journal of Applied Probability, 56(2):496–523, June 2019. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL https://www.cambridge.org/core/journals/journal-
of-applied-probability/article/asymptotic-behaviour-
of-randomised-fractional-volatility-models/101FC224B38AB53A94EF4DF91FD3F387.

Ho:2021:GPR

[HJL21] Man-Wai Ho, Lancelot F. James, and John W. Lau. Gibbs partitions, Riemann–Liouville fractional operators, Mittag-Leffler functions, and fragmentations derived from stable subordinators. Journal of Applied Probability, 58(2):314–334, June 2021. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL https://www.cambridge.org/core/journals/
journal-of-applied-probability/article/gibbs-partitions-
riemannliouville-fractional-operators-mittagleffler-
functions-and-fragmentations-derived-from-stable-subordinators/
CBA4ADB1382096F6455FD9C14874BD08.

Hannan:1977:API

[HK77] E. J. Hannan and Marek Kanter. Autoregressive processes with infinite variance. Journal of Applied Probability, 14(2):411– 415, June 1977. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3213015.

Hooghiemstra:1985:CED

[HK85] G. Hooghiemstra and M. Keane. Calculation of the equilibrium distribution for a solar energy storage model. Journal of Applied Probability, 22(4):852–864, December 1985. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/3213953.

Hill:1990:OSP

[HK90] T. P. Hill and D. P. Kennedy. Optimal stopping problems with generalized objective functions. Journal of Applied Probability, 27(4):828–838, December 1990. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.

org/stable/3214826.

Huillet:1994:PBD

[HK94] Thierry Huillet and Andrzej Klopotowski. Sur une procédure de branchement déterministe et ses dérivees aléatoires. (French) [On a deterministic branching procedure and its random derivatives]. Journal of Applied Probability, 31(2):333–347, June 1994. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/3215027.

Hamza:1995:CIM

[HK95] K. Hamza and F. C. Klebaner. Conditions for integrability of Markov chains. Journal of Applied Probability, 32(2):541– 547, June 1995. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3215307.

Hurd:2009:FPT

[HK09] T. R. Hurd and A. Kuznetsov. On the first passage time for Brownian motion subordinated by a Lévy process. Journal of Applied Probability, 46(1):181–198, March 2009. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/25662415.

Holst:2015:RCT

[HK15] Lars Holst and Takis Konstantopoulos. Runs in coin tossing: a general approach for deriving distributions for functionals. Journal of Applied Probability, 52(3):752–770, September 2015. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://projecteuclid.org/euclid.jap/1445543844.

Houdre:2019:RCL

[HK19] C. Houdré and G. Kerchev. On the rate of convergence for the length of the longest common subsequences in hidden Markov models. Journal of Applied Probability, 56(2):558–573, June 2019. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL https:/

/www.cambridge.org/core/journals/journal-of-applied-
probability/article/on-the-rate-of-convergence-for-the-
length-of-the-longest-common-subsequences-in-hidden-
markov-models/CD081E05841F08F6AC988F03446E070C.

Holmes:2022:CLF

[HK22] Mark Holmes and Daniel Kious. Coexistence of lazy frogs on Z. Journal of Applied Probability, 59(3):702–713, September 2022. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL https://www.cambridge.org/core/journals/
journal-of-applied-probability/article/coexistence-of-
lazy-frogs-on-mathbbz/D408AF28B51E92831A7DB524604675CE.

Hirayama:1989:FRD

[HKN89] Tetsuji Hirayama, Masaaki Kijima, and Shoichi Nishimura. Further results for dynamic scheduling of multiclass G/G/1 queues. Journal of Applied Probability, 26(3):595–603, September 1989. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/3214416.

Handa:1994:PID

[HKO94] Kenji Handa, Prasansa Kalukottege, and Yukio Ogura. A probabilistic interpretation of the degree of fuzziness. Journal of Applied Probability, 31(4):1034–1048, December 1994. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3215327.

Holst:1988:RPC

[HKQ88] L. Holst, J. E. Kennedy, and M. P. Quine. Rates of Poisson convergence for some coverage and urn problems using coupling. Journal of Applied Probability, 25(4):717–724, December 1988. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/3214292.

Heerten:2024:PTC

[HKT24] Nils Heerten, Julia Krecklenberg, and Christoph Thäle. The proportion of triangles in a class of anisotropic Poisson line tessellations. Journal of Applied Probability, 61 (1):214–229, March 2024. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL https:/

/www.cambridge.org/core/journals/journal-of-applied-

probability/article/proportion-of-triangles-in-a-class-
of-anisotropic-poisson-line-tessellations/6E04CAD676D5EE668F24809D9BF62CF1.

Hanin:1996:RMM

[HKY96] L. G. Hanin, L. B. Klebanov, and A. Yu. Yakovlev. Randomized multihit models and their identification. Journal of Applied Probability, 33(2):458–471, June 1996. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/3215071.

Hwang:1971:GRS

[HL71] F. K. Hwang and S. Lin. On generating a random sequence. Journal of Applied Probability, 8(2):366–373, June 1971. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/3211906.

Harrison:1976:VAW

[HL76] J. Michael Harrison and Austin J. Lemoine. On the virtual and actual waiting time distributions of a GI/G/1 queue. Journal of Applied Probability, 13(4):833–836, December 1976. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3212543.

Harrison:1977:LTP

[HL77] J. Michael Harrison and Austin J. Lemoine. Limit theorems for periodic queues. Journal of Applied Probability, 14(3):566–576, September 1977. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3213459.

Harrison:1981:NNI

[HL81a] J. Michael Harrison and Austin J. Lemoine. A note on networks of infinite-server queues. Journal of Applied Probability, 18(2):561–567, June 1981. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3213306.

Harrison:1981:SBM

[HL81b] J. Michael Harrison and Austin J. Lemoine. Sticky Brownian motion as the limit of storage processes. Journal of Applied Probability, 18(1):216–226, March 1981. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3213181.

Hong:1999:CLT

[HL99] Wen-Ming Hong and Zeng-Hu Li. A central limit theorem for super-Brownian motion with super-Brownian immigration. Journal of Applied Probability, 36(4):1218–1224, December 1999. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/3215590.

Hou:2004:ECS

[HL04] Zhenting Hou and Yuanyuan Liu. Explicit criteria for several types of ergodicity of the embedded M/G/1 and GI/M/n queues. Journal of Applied Probability, 41(3):778–790, September 2004. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/4141353.

Hu:2005:SRB

[HL05] Chin-Yuan Hu and Gwo Dong Lin. Sharp reliability bounds for the L-like distributions, a class in which ageing is based on Laplace transform inequalities. Journal of Applied Probability, 42(4):1204–1208, December 2005. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/30040839.

Herve:2013:GSM

[HL13a] Loïc Hervé and James Ledoux. Geometric rho-mixing property of the interarrival times of a stationary Markovian arrival process. Journal of Applied Probability, 50(2):598–601, June 2013. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://projecteuclid.org/euclid.jap/
1371648964.

Hu:2013:DPE

[HL13b] Yaozhong Hu and Chihoon Lee. Drift parameter estimation for a reflected fractional Brownian motion based on its local time. Journal of Applied Probability, 50(2):592–597, June 2013. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://projecteuclid.org/euclid.jap/1371648963.

Hitczenko:2014:DQB

[HL14] Paweł Hitczenko and Gérard Letac. Dirichlet and quasi-Bernoulli laws for perpetuities. Journal of Applied Probability, 51(2):400–416, June 2014. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://projecteuclid.

org/euclid.jap/1402578633.

He:2016:DJC

[HL16a] Xin He and Zenghu Li. Distributions of jumps in a continuousstate branching process with immigration. Journal of Applied Probability, 53(4):1166–1177, December 2016. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://projecteuclid.org/euclid.jap/1481132844.

Herve:2016:CBS

[HL16b] Loïc Hervé and James Ledoux. Computable bounds of an l^{2}- spectral gap for discrete Markov chains with band transition matrices. Journal of Applied Probability, 53(3):946–952, September 2016. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://projecteuclid.org/euclid.jap/
1476370789.

Herve:2023:RIF

[HL23] Loïc Hervé and James Ledoux. Robustness of iterated function systems of Lipschitz maps. Journal of Applied Probability, 60(3):921–944, September 2023. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL https:/

/www.cambridge.org/core/journals/journal-of-applied-
probability/article/robustness-of-iterated-function-
systems-of-lipschitz-maps/118D3335A4006015A146F63C4AB95C57.

Hu:2019:APR

[HLF19] Zhishui Hu, Zheng Li, and Qunqiang Feng. Accessibility percolation on random rooted labeled trees. Journal of Applied Probability, 56(2):533–545, June 2019. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL https:/

/www.cambridge.org/core/journals/journal-of-applied-
probability/article/accessibility-percolation-on-random-
rooted-labeled-trees/2772FF8110EF1F7DC825590979C8F4EF.

Harang:2021:SEM

[HLMOL21] Fabian A. Harang, Marc Lagunas-Merino, and Salvador Ortiz-Latorre. Self-exciting multifractional processes. Journal of Applied Probability, 58(1):22–41, March 2021. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL https://www.cambridge.org/core/journals/
journal-of-applied-probability/article/selfexciting-
multifractional-processes/785A4762D563BBB7B4B5B5E4C7D434A0.

Huang:1993:SCP

[HLS93] Wen-Jang Huang, Shun-Hwa Li, and Jyh-Cherng Su. Some characterizations of the Poisson process and geometric renewal process. Journal of Applied Probability, 30(1):121–130, March 1993. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/3214626.

Hordijk:2000:SES

[HLT00] Arie Hordijk, Zhen Liu, and Don Towsley. Smoothing effect of the superposition of homogeneous sources in tandem networks. Journal of Applied Probability, 37(3):900–913, September 2000. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/3215622.

Hurtado-Lange:2022:LHT

[HLVM22] Daniela Hurtado-Lange, Sushil Mahavir Varma, and Siva Theja Maguluri. Logarithmic heavy traffic error bounds in generalized switch and load balancing systems. Journal of Applied Probability, 59(3):652–669, September 2022. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL https:/

/www.cambridge.org/core/journals/journal-of-applied-
probability/article/logarithmic-heavy-traffic-error-
bounds-in-generalized-switch-and-load-balancing-systems/
01DA4B718BDBDFB5DDC51A9BBA0A1CA0.

Hou:2005:SRC

[HLZ05] Zhenting Hou, Yuanyuan Liu, and Hanjun Zhang. Subgeometric rates of convergence for a class of continuous-time Markov process. Journal of Applied Probability, 42(3):698–712, September 2005. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/30040850.

Homble:1995:WLR

[HM95] Patrick Homble and William P. McCormick. Weak limit results for the extremes of a class of shot noise processes. Journal of Applied Probability, 32(3):707–726, September 1995. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3215124.

Hooghiemstra:1997:EIS

[HM97] Gerard Hooghiemstra and Ludolf E. Meester. The extremal index in 10 seconds. Journal of Applied Probability, 34(3):818–822, September 1997. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3215109.

Ho:1999:ADS

[HM99] Hwai-Chung Ho and William P. McCormick. Asymptotic distribution of sum and maximum for Gaussian processes. Journal of Applied Probability, 36(4):1031–1044, December 1999. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3215576.

Huillet:2009:DAC

[HM09] Thierry Huillet and Martin Möhle. Duality and asymptotics for a class of nonneutral discrete Moran models. Journal of Applied Probability, 46(3):866–893, September 2009. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/25662465.

He:2014:LTC

[HM14] Hui He and Rugang Ma. Limit theorems for continuous-time branching flows. Journal of Applied Probability, 51(2):317–332, June 2014. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-

6072 (electronic). URL http://projecteuclid.org/euclid.
jap/1402578627.

Huber:2015:MDF

[HM15] Mark Huber and Nevena Marić. Multivariate distributions with fixed marginals and correlations. Journal of Applied Probability, 52(2):602–608, June 2015. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://projecteuclid.

org/euclid.jap/1437658619.

Huang:2017:SMT

[HM17] Lu-Jing Huang and Yong-Hua Mao. On some mixing times for nonreversible finite Markov chains. Journal of Applied Probability, 54(2):627–637, June 2017. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL https://www.cambridge.org/core/journals/journal-
of-applied-probability/article/on-some-mixing-times-
for-nonreversible-finite-markov-chains/CDAFF2D0FD19E5AF3BC4EF585F128819.

Hoyle:2020:GLP

[HM20] Edward Hoyle and Levent Ali Menguturk. Generalised Liouville processes and their properties. Journal of Applied Probability, 57(4):1088–1110, December 2020. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL https:/

/www.cambridge.org/core/journals/journal-of-applied-
probability/article/generalised-liouville-processes-
and-their-properties/6959FA1773CA87D7B2FE260E1C4C3D2A.

Hollander:2023:BEE

[HM23] Frank Den Hollander and Maarten Markering. Breaking of ensemble equivalence for dense random graphs under a single constraint. Journal of Applied Probability, 60(4): 1181–1200, December 2023. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL https:/

/www.cambridge.org/core/journals/journal-of-applied-
probability/article/breaking-of-ensemble-equivalence-
for-dense-random-graphs-under-a-single-constraint/D53757DA350CB9E6511D413A17BD8827.

Hong:2014:WBA

[HMBL14] Dawei Hong, Shushuang Man, Jean-Camille Birget, and Desmond S. Lun. A wavelet-based almost-sure uniform approximation of fractional Brownian motion with a parallel algorithm. Journal of Applied Probability, 51(1):1–18, March 2014. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://projecteuclid.org/euclid.jap/1395771410.

Hendi:1993:CNC

[HMM93] M. I. Hendi, A. F. Mashhour, and M. A. Montasser. Closure of the NBUC class under formation of parallel systems. Journal of Applied Probability, 30(4):975–978, December 1993. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3214527.

Higueras:2003:UMD

[HMPS03] I. Higueras, J. Moler, F. Plo, and M. San Miguel. Urn models and differential algebraic equations. Journal of Applied Probability, 40(2):401–412, June 2003. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3215799.

Higueras:2006:CLT

[HMPS06] I. Higueras, J. Moler, F. Plo, and M. San Miguel. Central limit theorems for generalized Pólya urn models. Journal of Applied Probability, 43(4):938–951, December 2006. CODEN JPRBAM.

ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/27595791.

Hyrien:2017:SSB

[HMY17] Ollivier Hyrien, Kosto V. Mitov, and Nikolay M. Yanev. Subcritical Sevastyanov branching processes with nonhomogeneous Poisson immigration. Journal of Applied Probability, 54(2):569–587, June 2017. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL https:/

/www.cambridge.org/core/journals/journal-of-applied-
probability/article/subcritical-sevastyanov-branching-
processes-with-nonhomogeneous-poisson-immigration/F9F73C2B5B332200306FD11214AAE3A9.

Haji:1971:RBS

[HN71] Rasoul Haji and Gordon F. Newell. A relation between stationary queue and waiting time distributions. Journal of Applied Probability, 8(3):617–620, September 1971. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/3212186.

Helland:1976:GRE

[HN76] Inge S. Helland and Trygve S. Nilsen. On a general random exchange model. Journal of Applied Probability, 13(4):781–790, December 1976. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3212533.

Hofmann:2000:RPP

[HN00] Glenn Hofmann and H. N. Nagaraja. Random and Poisson paced record models in the F^{α} setup. Journal of Applied Probability, 37(2):374–388, June 2000. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3215714.

He:2001:CLC

[HN01] Qi-Ming He and Marcel F. Neuts. On the convergence and limits of certain matrix sequences arising in quasi-birth-and-death Markov chains. Journal of Applied Probability, 38(2):519– 541, June 2001. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3215904.

Hjort:1985:ATM

[HNF85] Nils Lid Hjort, Bent Natvig, and Espen Funnemark. The association in time of a Markov process with application to multistate reliability theory. Journal of Applied Probability, 22(2):473– 479, June 1985. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3213792.

Hu:1993:NAI

[HNG93] Jian-Qiang Hu, Soracha Nananukul, and Wei-Bo Gong. A new approach to (s,s) inventory systems. Journal of Applied Probability, 30(4):898–912, December 1993. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3214521.

Hawkes:1974:CPR

[HO74] Alan G. Hawkes and David Oakes. A cluster process representation of a self-exciting process. Journal of Applied Probability, 11 (3):493–503, September 1974. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.

org/stable/3212693.

Hu:2007:OSP

[HØ07] Yaozhong Hu and Bernt Øksendal. Optimal smooth portfolio selection for an insider. Journal of Applied Probability, 44(3): 742–752, September 2007. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/27595879.

Hogan:1986:CBD

[Hog86] Michael L. Hogan. Comment on “Corrected Diffusion Approximations in Certain Random Walk Problems”. Journal of Applied Probability, 23(1):89–96, March 1986. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3214118. See [Sie79].

Hokstad:1975:GMM

[Hok75] Per Hokstad. The G/M/m queue with finite waiting room. Journal of Applied Probability, 12(4):779–792, December 1975. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/3212728.

Hokstad:1977:ABG

[Hok77] Per Hokstad. Asymptotic behaviour of the E_{k} /G/1 queue with finite waiting room. Journal of Applied Probability, 14(2):358– 366, June 1977. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3213006.

Holgate:1966:MSF

[Hol66] P. Holgate. A mathematical study of the founder principle of evolutionary genetics. Journal of Applied Probability, 3(1):115– 128, June 1966. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3212041.

Holgate:1967:DPP

[Hol67] P. Holgate. Divergent population processes and mammal outbreaks. Journal of Applied Probability, 4(1):1–8, April 1967. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/3212295.

Holgate:1968:IBM

[Hol68] P. Holgate. Interaction between migration and breeding studied by means of genetic algebras. Journal of Applied Probability, 5 (1):1–8, April 1968. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3212072.

Holgate:1976:TNH

[Hol76] P. Holgate. The total number of heterozygotes before fixation. Journal of Applied Probability, 13(2):201–207, June 1976. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/3212823.

Holland:1978:SPL

[Hol78] Charles J. Holland. Stochastically perturbed limit cycles. Journal of Applied Probability, 15(2):311–320, June 1978. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3213404.

Holgate:1979:NIP

[Hol79a] P. Holgate. Notes on information in population genetics. Journal of Applied Probability, 16(3):473–481, September 1979. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3213077.

Holst:1979:UAL

[Hol79b] Lars Holst. A unified approach to limit theorems for urn models. Journal of Applied Probability, 16(1):154–162, March 1979. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/3213383.

Holst:1980:MCC

[Hol80a] Lars Holst. On multiple covering of a circle with random arcs. Journal of Applied Probability, 17(1):284–290, March 1980. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/3212948.

Holst:1980:LPS

[Hol80b] Lars Holst. On the lengths of the pieces of a stick broken at random. Journal of Applied Probability, 17(3):623–634, September 1980. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3212956.

Holgate:1981:LLF

[Hol81a] P. Holgate. Linked loci in finite populations. Journal of Applied Probability, 18(4):789–798, December 1981. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/3213054.

Holst:1981:SOP

[Hol81b] Lars Holst. On sequential occupancy problems. Journal of Applied Probability, 18(2):435–442, June 1981. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/3213289.

Holgate:1992:BCA

[Hol92] Philip Holgate. Biometric and chromosome algebras. Journal of Applied Probability, 29(2):247–254, June 1992. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3214563.

Holden:2000:CMC

[Hol00] Lars Holden. Convergence of Markov chains in the relative supremum norm. Journal of Applied Probability, 37(4):1074–1083, December 2000. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3215495.

Holst:2007:CFS

[Hol07] Lars Holst. Counts of failure strings in certain Bernoulli sequences. Journal of Applied Probability, 44(3):824–830, September 2007. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/27595884.

Holst:2008:NEC

[Hol08a] Lars Holst. A note on embedding certain Bernoulli sequences in marked Poisson processes. Journal of Applied Probability, 45(4): 1181–1185, December 2008. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/27596016.

Holst:2008:NTC

[Hol08b] Lars Holst. The number of two consecutive successes in a Hoppe–Pólya urn. Journal of Applied Probability, 45(3):901–906, September 2008. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
27595993.

Holst:2009:CRC

[Hol09] Lars Holst. On consecutive records in certain Bernoulli sequences. Journal of Applied Probability, 46(4):1201–1208, December 2009. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
25662490.

Holst:2013:PPE

[Hol13] Lars Holst. Probabilistic proofs of Euler identities. Journal of Applied Probability, 50(4):1206–1212, December 2013. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://projecteuclid.org/euclid.jap/1389370108.

Hong:2002:MDS

[Hon02] Wen-Ming Hong. Moderate deviation for super-Brownian motion with super-Brownian immigration. Journal of Applied Probability, 39(4):829–838, December 2002. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http:/

/www.jstor.org/stable/3216006.

Hong:2004:NTL

[Hon04] Wen-Ming Hong. A note on two-level superprocesses. Journal of Applied Probability, 41(1):202–210, March 2004. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3215825.

Hong:2013:CSB

[Hon13] Jyy-I Hong. Coalescence in subcritical Bellman-Harris agedependent branching processes. Journal of Applied Probability, 50(2):576–591, June 2013. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://projecteuclid.

org/euclid.jap/1371648962.

Hong:2016:CCS

[Hon16] Jyy-I Hong. Coalescence on critical and subcritical multitype branching processes. Journal of Applied Probability, 53(3):802– 817, September 2016. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://projecteuclid.

org/euclid.jap/1476370777.

Honnappa:2017:RET

[Hon17] Harsha Honnappa. Rare events of transitory queues. Journal of Applied Probability, 54(3):943–962, September 2017. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL https://www.cambridge.org/core/journals/journal-
of-applied-probability/article/rare-events-of-transitory-
queues/3A46A766B9DC5FAD77C9EC9183045660.

Hooke:1970:SLT

[Hoo70] John A. Hooke. On some limit theorems for the GI/G/1 queue. Journal of Applied Probability, 7(3):634–640, December 1970. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/3211942.

Hooghiemstra:1983:CLT

[Hoo83] G. Hooghiemstra. Conditioned limit theorems for waiting-time processes of the M/G/1 queue. Journal of Applied Probability, 20(3):675–688, September 1983. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.

org/stable/3213902.

Hooghiemstra:1987:FAR

[Hoo87] G. Hooghiemstra. On functionals of the adjusted range process. Journal of Applied Probability, 24(1):252–257, March 1987. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/3214077.

Hook:2013:CPS

[Hoo13] James Hook. Critical path statistics of max-plus linear systems with Gaussian noise. Journal of Applied Probability, 50(3): 654–670, September 2013. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://projecteuclid.

org/euclid.jap/1378401228.

Hoppe:1975:SMM

[Hop75] Fred Hoppe. Stationary measures for multitype branching processes. Journal of Applied Probability, 12(2):219–227, June 1975. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/3212435.

Hoppe:1976:RRV

[Hop76] Fred M. Hoppe. On a result of Rubin and vere-Jones concerning subcritical branching processes. Journal of Applied Probability, 13(4):804–808, December 1976. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.

org/stable/3212537.

Hopfner:1982:LLT

[Höp82] R. Höpfner. Local limit theorems for non-critical Galton–Watson processes with or without immigration. Journal of Applied Probability, 19(2):262–271, June 1982. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3213479.

Hopfner:1985:NPE

[Höp85a] R. Höpfner. A note on the probability of extinction in a class of population-size-dependent Galton–Watson processes. Journal of Applied Probability, 22(4):920–925, December 1985. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3213959.

Hopfner:1985:SCP

[Höp85b] Reinhard Höpfner. On some classes of population-size-dependent Galton–Watson processes. Journal of Applied Probability, 22(1):25–36, March 1985. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3213745.

Hopfner:1986:SRP

[Höp86a] Reinhard Höpfner. Some results on population-size-dependent Galton–Watson processes. Journal of Applied Probability, 23 (2):297–306, June 1986. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3214174.

Hoppe:1986:SBF

[Hop86b] Fred M. Hoppe. Size-biased filtering of Poisson–Dirichlet samples with an application to partition structures in genetics. Journal of Applied Probability, 23(4):1008–1012, December 1986. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/3214473.

Horowitz:1965:PRP

[Hor65] Maurice Horowitz. Probability of random paths across elementary geometrical shapes. Journal of Applied Probability, 2 (1):169–177, June 1965. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/ stable/3211882. See correction [Ano65e].

Hori:1979:WPP

[Hor79] Motoo Hori. On Weiner’s proof of the Palásti conjecture. Journal of Applied Probability, 16(3):702–706, September 1979. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3213105.

Hori:1980:CSL

[Hor80] Motoo Hori. Comments on the second letter of Weiner [11]. Journal of Applied Probability, 17(3):888–890, September 1980. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/3212988. See

reply [Wei80b].

Horvath:1998:DAR

[Hor98] Lajos Horváth. Diffusion approximation for random walks on anisotropic lattices. Journal of Applied Probability, 35(1):206– 212, March 1998. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3215558.

Horst:2001:SEN

[Hor01] Ulrich Horst. The stochastic equation Y_{t+1} = A_{t}Y_{t} + B_{t} with non-stationary coefficients. Journal of Applied Probability, 38 (1):80–94, March 2001. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3215743.

Horfelt:2003:GBC

[Hör03] Per Hörfelt. Geometric bounds on certain sublinear functionals of geometric Brownian motion. Journal of Applied Probability, 40(4):893–905, December 2003. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.

org/stable/3216049.

Horbacz:2004:RDS

[Hor04] Katarzyna Horbacz. Random dynamical systems with jumps. Journal of Applied Probability, 41(3):890–910, September 2004. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/4141360.

Horfelt:2005:MPS

[Hör05] Per Hörfelt. The moment problem for some Wiener functionals: Corrections to previous proofs (with an appendix by H. L. Pedersen). Journal of Applied Probability, 42(3):851–860, September 2005. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/30040861.

Hosoya:1986:STP

[Hos86] Yuzo Hosoya. A simultaneous test in the presence of nested alternative hypotheses. Journal of Applied Probability, 23A:187–200, 1986. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3214352.

Hoshino:2006:DMD

[Hos06] Nobuaki Hoshino. A discrete multivariate distribution resulting from the law of small numbers. Journal of Applied Probability, 43(3):852–866, September 2006. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.

org/stable/27595777.

Hosaka:2009:ACS

[Hos09] Kenshi Hosaka. An alternative condition for stochastic domination. Journal of Applied Probability, 46(4):1198–1200, December 2009. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/25662489.

Hossjer:2019:STP

[Hös19] Ola Hössjer. A spatio-temporal point process model for particle growth. Journal of Applied Probability, 56(1):23–38, March 2019. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL https://www.cambridge.org/core/journals/
journal-of-applied-probability/article/spatiotemporal-
point-process-model-for-particle-growth/687ED3E469FCA1F7BF108AE51BDDC6EC.

Hougaard:1991:MHS

[Hou91] Philip Hougaard. Modelling heterogeneity in survival data. Journal of Applied Probability, 28(3):695–701, September 1991. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/3214503.

House:2012:LAS

[Hou12] Thomas House. Lie algebra solution of population models based on time-inhomogeneous Markov chains. Journal of Applied Probability, 49(2):472–481, June 2012. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/41713780.

Howard:2000:LBP

[How00a] C. Douglas Howard. Lower bounds for point-to-point wandering exponents in Euclidean first-passage percolation. Journal of Applied Probability, 37(4):1061–1073, December 2000. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3215494.

Howard:2000:ZTI

[How00b] C. Douglas Howard. Zero-temperature Ising spin dynamics on the homogeneous tree of degree three. Journal of Applied Probability, 37(3):736–747, September 2000. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3215609.

Howard:2001:DME

[How01] C. Douglas Howard. Differentiability and monotonicity of expected passage time in Euclidean first-passage percolation. Journal of Applied Probability, 38(4):815–827, December 2001. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/3215766.

Harrison:1993:STS

[HP93] P. G. Harrison and E. Pitel. Sojourn times in single-server queues with negative customers. Journal of Applied Probability, 30(4): 943–963, December 1993. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3214524.

Harrison:1995:RTD

[HP95] Peter G. Harrison and Edwige Pitel. Response time distributions in tandem G-networks. Journal of Applied Probability, 32 (1):224–246, March 1995. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3214932.

Haggstrom:1998:FPP

[HP98a] Olle Häggström and Robin Pemantle. First passage percolation and a model for competing spatial growth. Journal of Applied Probability, 35(3):683–692, September 1998. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/3215643.

Haviv:1998:BOC

[HP98b] Moshe Haviv and Martin L. Puterman. Bias optimality in controlled queueing systems. Journal of Applied Probability, 35 (1):136–150, March 1998. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3215553.

Haslegrave:2017:RCC

[HP17] John Haslegrave and Mate Puljiz. Reaching consensus on a connected graph. Journal of Applied Probability, 54(1):181–198, March 2017. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL https:/

/www.cambridge.org/core/journals/journal-of-applied-
probability/article/reaching-consensus-on-a-connected-
graph/64FC223E06ACA92DD22E117D2E7323E2.

Howlett:2014:NLU

[HPP14] Phil Howlett, Charles Pearce, and Julia Piantadosi. A new look at urban water storage in a series of connected dams. Journal of Applied Probability, 51(1):118–135, March 2014. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://projecteuclid.org/euclid.jap/1395771418.

Hubbard:1986:DPM

[HPR86] John R. Hubbard, Claude Dennis Pegden, and Matthew Rosenshine. The departure process for the M/M/1 queue. Journal of Applied Probability, 23(1):249–255, March 1986. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3214137.

Huzak:2004:RPC

[HPSV04] Miljenko Huzak, Mihael Perman, Hrvoje Sikić, and Zoran Vondracek. Ruin probabilities for competing claim processes. Journal of Applied Probability, 41(3):679–690, September 2004. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/4141346.

Hordijk:1987:SIO

[HR87] Arie Hordijk and Ad Ridder. Stochastic inequalities for an overflow model. Journal of Applied Probability, 24(3):696–708, September 1987. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3214100.

Hannan:1988:WSW

[HR88a] E. J. Hannan and J. Rissanen. The width of a spectral window. Journal of Applied Probability, 25A:301–307, 1988. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/3214164.

Hordijk:1988:IBS

[HR88b] Arie Hordijk and Ad Ridder. Insensitive bounds for the stationary distribution of non-reversible Markov chains. Journal of Applied Probability, 25(1):9–20, March 1988. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/3214229.

Haigh:2000:OSD

[HR00] John Haigh and Markus Roters. Optimal strategy in a dice game. Journal of Applied Probability, 37(4):1110–1116, December 2000. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3215500.

Haji:2015:QLM

[HR15] Babak Haji and Sheldon M. Ross. A queueing loss model with heterogeneous skill based servers under idle time ordering policies. Journal of Applied Probability, 52(1):269–277, March 2015. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://projecteuclid.org/euclid.jap/
1429282621.

Henderson:1983:NRC

[HRF83] Robin Henderson, Eric Renshaw, and David Ford. A note on the recurrence of a correlated random walk. Journal of Applied Probability, 20(3):696–699, September 1983. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/3213904.

Henderson:1984:CRW

[HRF84] Robin Henderson, Eric Renshaw, and David Ford. A correlated random walk model for two-dimensional diffusion. Journal of Applied Probability, 21(2):233–246, June 1984. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3213636.

Henderson:1985:A

[HRF85] Robin Henderson, Eric Renshaw, and David Ford. Acknowledgement. Journal of Applied Probability, 22(1):251, March 1985. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/3213767.

Hansen:2007:SPS

[HRZ07] Jennie Hansen, Cian Reynolds, and Stan Zachary. Stability of processor sharing networks with simultaneous resource requirements. Journal of Applied Probability, 44(3):636–651, September 2007. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/27595871.

Heathcote:1966:IBP

[HS66] C. R. Heathcote and E. Seneta. Inequalities for branching processes. Journal of Applied Probability, 3(1):261–267, June 1966. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/3212050. See

correction [Ano67e].

Hardin:1969:NAP

[HS69a] J. C. Hardin and A. L. Sweet. A note on absorption probabilities for a random walk between a reflecting and an absorbing barrier. Journal of Applied Probability, 6(1):224–226, April 1969. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/3212291.

Holmes:1969:NWT

[HS69b] Paul T. Holmes and William E. Strawderman. A note on the waiting times between record observations. Journal of Applied Probability, 6(3):711–714, December 1969. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/3212118.

Heyde:1972:ETG

[HS72] C. C. Heyde and E. Seneta. Estimation theory for growth and immigration rates in a multiplicative process. Journal of Applied Probability, 9(2):235–256, June 1972. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3212796. See notes [HS74b].

Hafner:1974:SAR

[HS74a] R. Hafner and W. Sendler. Some asymptotic results on multiple matching. Journal of Applied Probability, 11(3):479–492, September 1974. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/

3212692. See corrections [Ano74g].

Heyde:1974:NBT

[HS74b] C. C. Heyde and E. Seneta. Notes on “Estimation Theory for Growth and Immigration Rates in a Multiplicative Process”. Journal of Applied Probability, 11(3):572–577, September 1974. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/3212702. See

[HS72].

Hudson:1977:NSB

[HS77] Irene L. Hudson and E. Seneta. A note on simple branching processes with infinite mean. Journal of Applied Probability, 14 (4):836–842, December 1977. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.

org/stable/3213356.

Heyde:1978:UBP

[HS78] C. C. Heyde and H.-J. Schuh. Uniform bounding of probability generating functions and the evolution of reproduction rates in birds. Journal of Applied Probability, 15(2):243–250, June 1978. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3213398.

Hooghiemstra:1987:LE

[HS87a] Gerard Hooghiemstra and Wolfgang Stadje. Letter to the Editor. Journal of Applied Probability, 24(4):1012–1013, December 1987. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3214224.

Huffer:1987:PCC

[HS87b] F. W. Huffer and L. A. Shepp. On the probability of covering the circle by random arcs. Journal of Applied Probability, 24 (2):422–429, June 1987. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3214266.

Huang:1993:SCE

[HS93] Wen-Jang Huang and Jyh-Ming Shoung. Some characterizations of the exponential and geometric distributions arising in a queueing system with an unreliable server. Journal of Applied Probability, 30(4):985–990, December 1993. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/3214529.

Hebert:1994:VTM

[HS94] Jaimie L. Hebert and John W. Seaman, Jr. The variance of a truncated mixed exponential process. Journal of Applied Probability, 31(1):167–179, March 1994. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3215244.

Huang:1996:RVM

[HS96a] Dawei Huang and N. M. Spencer. On a random vibration model. Journal of Applied Probability, 33(4):1141–1158, December 1996. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/3214992.

Husler:1996:NPP

[HS96b] J. Hüsler and M. Schmidt. A note on the point processes of rare events. Journal of Applied Probability, 33(3):654–663, September 1996. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3215347.

Heath:2000:MVP

[HS00] David Heath and Martin Schweizer. Martingales versus PDEs in finance: An equivalence result with examples. Journal of Applied Probability, 37(4):947–957, December 2000. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/3215486.

Henderson:2004:ASU

[HS04] Shane G. Henderson and Burt Simon. Adaptive simulation using perfect control variates. Journal of Applied Probability, 41(3): 859–876, September 2004. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/4141358.

Hult:2010:LDP

[HS10] Henrik Hult and Gennady Samorodnitsky. Large deviations for point processes based on stationary sequences with heavy tails. Journal of Applied Probability, 47(1):1–40, March 2010. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/25662498.

Huffer:2012:JDC

[HS12] Fred W. Huffer and Jayaram Sethuraman. Joint distributions of counts of strings in finite Bernoulli sequences. Journal of Applied Probability, 49(3):758–772, September 2012. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/41713804.

Huber:2014:RCI

[HS14] Mark Huber and Sarah Schott. Random construction of interpolating sets for high-dimensional integration. Journal of Applied Probability, 51(1):92–105, March 2014. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

projecteuclid.org/euclid.jap/1395771416.

Hollering:2022:ESC

[HS22] Benjamin Hollering and Seth Sullivant. Exchangeable and sampling-consistent distributions on rooted binary trees. Journal of Applied Probability, 59(1):60–80, March 2022. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL https://www.cambridge.org/core/journals/
journal-of-applied-probability/article/exchangeable-
and-samplingconsistent-distributions-on-rooted-binary-
trees/8FB14D33F51C14A5AAAC268F5ED02DBA.

Hsia:1976:JPD

[Hsi76] Wei Shen Hsia. The joint probability density function of the occupation time of a three-state problem. Journal of Applied Probability, 13(1):57–64, March 1976. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3212665.

Hsiau:1997:CPL

[Hsi97] Shoou-Ren Hsiau. Compound Poisson limit theorems for Markov chains. Journal of Applied Probability, 34(1):24–34, March 1997. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/3215171.

Hordijk:1975:ABM

[HST75] Arie Hordijk, Paul J. Schweitzer, and Henk Tijms. The asymptotic behaviour of the minimal total expected cost for the denumerable state Markov decision model. Journal of Applied Probability, 12(2):298–305, June 1975. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3212443.

Harlow:1983:LTA

[HST83] D. G. Harlow, R. L. Smith, and H. M. Taylor. Lower tail analysis of the distribution of the strength of load-sharing systems. Journal of Applied Probability, 20(2):358–367, June 1983. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/3213808.

Hsu:1979:NEE

[Hsu79] Yu-Sheng Hsu. A note on exchangeable events. Journal of Applied Probability, 16(3):662–664, September 1979. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3213095.

Hannan:1971:SIN

[HT71] E. J. Hannan and P. J. Thomson. Spectral inference over narrow bands. Journal of Applied Probability, 8(1):157–169, March 1971. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3211845.

Hill:1978:AFQ

[HT78] James M. Hill and Keith P. Tognetti. The autocorrelation function of the queue length process for the two-server Poisson queue. Journal of Applied Probability, 15(2):447–451, June 1978. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/3213418.

Henderson:1989:IPI

[HT89] W. Henderson and P. Taylor. Insensitivity of processes with interruptions. Journal of Applied Probability, 26(2):242–258, June 1989. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3214032.

Henderson:1991:SNR

[HT91] W. Henderson and P. G. Taylor. Some new results on queueing networks with batch movement. Journal of Applied Probability, 28(2):409–421, June 1991. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3214876.

Hao:2012:ARP

[HT12] Xuemiao Hao and Qihe Tang. Asymptotic ruin probabilities for a bivariate Lévy–driven risk model with heavy-tailed claims and risky investments. Journal of Applied Probability, 49(4): 939–953, December 2012. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/41713820.

Hall:1992:ACL

[HTV92] Peter Hall, Jozef L. Teugels, and Ann Vanmarcke. The abscissa of convergence of the Laplace transform. Journal of Applied Probability, 29(2):353–362, June 1992. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3214572.

Haghighi-Talab:1973:DRF

[HTW73] Dariush Haghighi-Talab and Christopher Wright. On the distribution of records in a finite sequence of observations, with an application to a road traffic problem. Journal of Applied Probability, 10(3):556–571, September 1973. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3212776.

Hao:2009:MES

[HTW09] Xuemiao Hao, Qihe Tang, and Li Wei. On the maximum exceedance of a sequence of random variables over a renewal threshold. Journal of Applied Probability, 46(2):559–570, June 2009. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/25662443.

Hu:1996:SCO

[Hu96] Taizhong Hu. Stochastic comparisons of order statistics under multivariate imperfect repair. Journal of Applied Probability, 33 (1):156–163, March 1996. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3215273.

Hu:2016:GEB

[Hu16] Yanqing Hu. Generalized Efron’s biased coin design and its theoretical properties. Journal of Applied Probability, 53(2):327–340, June 2016. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-

6072 (electronic). URL http://projecteuclid.org/euclid.
jap/1466172857.

Huang:1974:CED

[Hua74a] J. S. Huang. Characterizations of the exponential distribution by order statistics. Journal of Applied Probability, 11(3):605–608, September 1974. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3212708.

Huang:1974:TAR

[Hua74b] J. S. Huang. On a theorem of ahsanullah and Rahman. Journal of Applied Probability, 11(1):216–218, March 1974. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3212602.

Huang:1989:SCP

[Hua89] Wen-Jang Huang. On some characterizations of the Poisson process. Journal of Applied Probability, 26(1):176–181, March 1989. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3214328.

Hua:1995:CML

[Hua95] Fan Ai Hua. Sur la L^{p} convergence des martingales liées au recouvrement. (French) [On the L^{p} convergence martingales related to recovery]. Journal of Applied Probability, 32(3):668–678, September 1995. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3215121.

Huang:2016:MMM

[Hua16] Chunmao Huang. Moments for multidimensional Mandelbrot cascades. Journal of Applied Probability, 53(1):1–21, March 2016. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://projecteuclid.org/euclid.jap/
1457470554.

Hudson:1973:QMC

[Hud73] R. L. Hudson. A quantum-mechanical central limit theorem for anti-commuting observables. Journal of Applied Probability, 10 (3):502–509, September 1973. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.

org/stable/3212771.

Huffer:1986:VOR

[Huf86] Fred Huffer. Variability orderings related to coverage problems on the circle. Journal of Applied Probability, 23(1):97– 106, March 1986. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3214119.

Huffer:1987:IMG

[Huf87] Fred W. Huffer. Inequalities for the M/G/∞ queue and related shot noise processes. Journal of Applied Probability, 24(4): 978–989, December 1987. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3214220.

Huffer:1988:DDJ

[Huf88a] Fred Huffer. Divided differences and the joint distribution of linear combinations of spacings. Journal of Applied Probability, 25(2):346–354, June 1988. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3214442.

Huffer:1988:SRC

[Huf88b] Fred W. Huffer. Some results concerning random arcs on the circle. Journal of Applied Probability, 25(4):833–838, December 1988. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3214306.

Hull:1982:NCE

[Hul82] David M. Hull. A necessary condition for extinction in those bisexual Galton–Watson branching processes governed by superadditive mating functions. Journal of Applied Probability, 19(4): 847–850, December 1982. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3213838.

Hull:1984:CEC

[Hul84] David M. Hull. Conditions for extinction in certain bisexual Galton–Watson branching processes. Journal of Applied Probability, 21(2):414–418, June 1984. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3213650.

Hull:2001:RLE

[Hul01] David M. Hull. A reconsideration of Lotka’s extinction probability using a bisexual branching process. Journal of Applied Probability, 38(3):776–780, September 2001. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/3216130.

Hunter:1976:UBP

[Hun76a] David Hunter. An upper bound for the probability of a union. Journal of Applied Probability, 13(3):597–603, September 1976. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/3212481.

Huntington:1976:MRT

[Hun76b] Raymond J. Huntington. Mean recurrence times for k successes within m trials. Journal of Applied Probability, 13(3):604–607, September 1976. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3212482.

Hunter:1986:NRN

[Hun86] Jeffrey J. Hunter. The non-renewal nature of the quasi-input process in the M/G/1/∞ queue. Journal of Applied Probability, 23(3):803–811, September 1986. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.

org/stable/3214017.

Hunt:1995:PBL

[Hun95] P. J. Hunt. Pathological behaviour in loss networks. Journal of Applied Probability, 32(2):519–533, June 1995. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3215305.

Husler:1982:RCC

[Hüs82] J. Hüsler. Random coverage of the circle and asymptotic distributions. Journal of Applied Probability, 19(3):578–587, September 1982. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3213515.

Husler:1986:EVN

[Hüs86a] Jürg Hüsler. Extreme values of non-stationary random sequences. Journal of Applied Probability, 23(4):937–950, December 1986. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3214467.

Husler:1986:PPC

[Hüs86b] Jürg Hüsler. On point processes on the circle. Journal of Applied Probability, 23(2):322–331, June 1986. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3214176.

Husler:1988:MNU

[Hüs88] Jürg Hüsler. Maximal, non-uniform spacings and the coverage problem. Journal of Applied Probability, 25(3):519–528, September 1988. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3213981.

Husler:1993:NER

[Hüs93] J. Hüsler. A note on exceedances and rare events of nonstationary sequences. Journal of Applied Probability, 30(4): 877–888, December 1993. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3214519.

Husler:1995:NMR

[Hüs95] J. Hüsler. A note on a model for restoring a destroyed region. Journal of Applied Probability, 32(3):756–767, September 1995. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/3215127.

Hutton:1990:NNT

[Hut90] Jane L. Hutton. Non-negative time series models for dry river flow. Journal of Applied Probability, 27(1):171–182, March 1990. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/3214604.

Hwang:1979:BFL

[HW79] Tea-Yuan Hwang and Nae-Sheng Wang. On best fractional linear generating function bounds. Journal of Applied Probability, 16(2):449–453, June 1979. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3212915.

Heathcote:1983:REA

[HW83] C. R. Heathcote and A. H. Welsh. The robust estimation of autoregressive processes by functional least squares. Journal of Applied Probability, 20(4):737–753, December 1983. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3213586.

Helm:1984:OCA

[HW84a] Werner E. Helm and Karl-Heinz Waldmann. Optimal control of arrivals to multiserver queues in a random environment. Journal of Applied Probability, 21(3):602–615, September 1984. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3213621.

Heyman:1984:ABQ

[HW84b] Daniel P. Heyman and Ward Whitt. The asymptotic behavior of queues with time-varying arrival rates. Journal of Applied Probability, 21(1):143–156, March 1984. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3213672.

Hollander:1988:RCR

[HW88] W. Th. F. Den Hollander and G. H. Weiss. On the range of a constrained random walk. Journal of Applied Probability, 25(3): 451–463, September 1988. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3213975.

Hsu:1996:DBA

[HW96] Ivy Hsu and Jean Walrand. Dynamic bandwidth allocation for ATM switches. Journal of Applied Probability, 33(3):758–771, September 1996. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3215357.

Henderson:2002:EFF

[HW02] Vicky Henderson and Rafal Wojakowski. On the equivalence of floating and fixed-strike Asian options. Journal of Applied Probability, 39(2):391–394, June 2002. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3216104.

Hajek:2018:RHC

[HWX18] Bruce Hajek, Yihong Wu, and Jiaming Xu. Recovering a hidden community beyond the Kesten–Stigum threshold in O(|E| log^{∗} |V |) time. Journal of Applied Probability, 55(2):325–352, June 2018. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL https:/

/www.cambridge.org/core/journals/journal-of-applied-
probability/article/recovering-a-hidden-community-beyond-
the-kestenstigum-threshold-in-oelogv-time/2CBFA9F1203FC00529C3C26E3372EB8E.

Hu:2002:PCP

[HX02] Taizhong Hu and Huiliang Xie. Proofs of the closure property of NBUC and NBU(2) under convolution. Journal of Applied Probability, 39(1):224–227, March 2002. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3215929.

Hsu:1994:TSD

[HY94] Guang-Hui Hsu and Xue-Ming Yuan. Transient solutions for denumerable-state Markov processes. Journal of Applied Probability, 31(3):635–645, September 1994. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3215144.

Heyde:1997:DLR

[HY97] C. C. Heyde and Y. Yang. On defining long-range dependence. Journal of Applied Probability, 34(4):939–944, December 1997. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/3215008.

Hsiau:2002:SLS

[HY02] Shoou-Ren Hsiau and Jiing-Ru Yang. Selecting the last success in Markov-dependent trials. Journal of Applied Probability, 39 (2):271–281, June 2002. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3216095.

Hyndman:1994:ABC

[Hyn94] Rob J. Hyndman. Approximations and boundary conditions for continuous-time threshold autoregressive processes. Journal of Applied Probability, 31(4):1103–1109, December 1994. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3215333.

Hwang:1982:LTG

[HZ82] Jiunn Tzon Hwang and James V. Zidek. Limit theorems for outguesses with mean-guided second guessing. Journal of Applied Probability, 19(2):321–331, June 1982. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3213484.

Harel:1987:CGP

[HZ87] Arie Harel and Paul Zipkin. The convexity of a general performance measure for multiserver queues. Journal of Applied Probability, 24(3):725–736, September 1987. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/3214102.

Helfer:1992:GIW

[HZ92] Adam D. Helfer and Zhongxin Zhao. Gaussian integrals on Wiener spaces. Journal of Applied Probability, 29(1):46–55, March 1992. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3214790.

Hobson:2020:RRS

[HZ20] David Hobson and Matthew Zeng. Randomised rules for stopping problems. Journal of Applied Probability, 57(3): 981–1004, September 2020. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL https:/

/www.cambridge.org/core/journals/journal-of-applied-
probability/article/randomised-rules-for-stopping-problems/
00E06A93020A6A295595363D7FA1F62C.

He:2022:ECQ

[HZ22a] Guoman He and Hanjun Zhang. Exponential convergence to a quasi-stationary distribution for birth-death processes with an entrance boundary at infinity. Journal of Applied Probability, 59(4):983–991, December 2022. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL https:/

/www.cambridge.org/core/journals/journal-of-applied-
probability/article/exponential-convergence-to-a-quasistationary-
distribution-for-birthdeath-processes-with-an-entrance-
boundary-at-infinity/70410FB5F0337360182F7C7EFDCEBBE6.

Houdebert:2022:EDU

[HZ22b] Pierre Houdebert and Alexander Zass. An explicit Dobrushin uniqueness region for Gibbs point processes with repulsive interactions. Journal of Applied Probability, 59(2):541–555, June 2022. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-

6072 (electronic). URL https://www.cambridge.org/core/
journals/journal-of-applied-probability/article/an-explicit-
dobrushin-uniqueness-region-for-gibbs-point-processes-
with-repulsive-interactions/C4FB9090574FD2C97C83C40B24B5D37D.

Hong-Zhi:1986:IAP

[HZZG86] An Hong-Zhi and Chen Zhao-Guo. The identification of ARMA processes. Journal of Applied Probability, 23A:75–87, 1986. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/3214344.

Isaacson:1978:SEC

[IA78] Dean Isaacson and Barry Arnold. Strong ergodicity for continuous-time Markov chains. Journal of Applied Probability, 15(4):699–706, December 1978. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3213427.

Inoue:2007:GFW

[IA07] Kiyoshi Inoue and Sigeo Aki. On generating functions of waiting times and numbers of occurrences of compound patterns in a sequence of multistate trials. Journal of Applied Probability, 44 (1):71–81, March 2007. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/27595823.

Ibragimov:2007:TSD

[Ibr07] Rustam Ibragimov. Thou shalt not diversify: Why ’two of every sort’? Journal of Applied Probability, 44(1):58–70, March 2007. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/27595822.

Isaacson:1982:LNP

[IC82] Dean Isaacson and Peter Colwell. Levels of null persistency for Markov chains. Journal of Applied Probability, 19(2):425– 429, June 1982. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3213495.

Iacobelli:2019:TSV

[IFN19] Giulio Iacobelli, Daniel R. Figueiredo, and Giovanni Neglia. Transient and slim versus recurrent and fat: Random walks and the trees they grow. Journal of Applied Probability, 56 (3):769–786, September 2019. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL https:/

/www.cambridge.org/core/journals/journal-of-applied-
probability/article/transient-and-slim-versus-recurrent-
and-fat-random-walks-and-the-trees-they-grow/F9425D01BEE19BB1D6ABF8ECC916AEAF.

Irle:2001:DWO

[IG01] Albrecht Irle and Joseph Gani. The detection of words and an ordering for Markov chains. Journal of Applied Probability, 38A:66–77, 2001. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3215869.

Iglehart:1965:LDA

[Igl65] Donald L. Iglehart. Limiting diffusion approximations for the many server queue and the repairman problem. Journal of Applied Probability, 2(2):429–441, December 1965. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3212203.

Iglehart:1969:DAC

[Igl69] Donald L. Iglehart. Diffusion approximations in collective risk theory. Journal of Applied Probability, 6(2):285–292, August 1969. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3211999.

Iglehart:1974:RWN

[Igl74] Donald L. Iglehart. Random walks with negative drift conditioned to stay positive. Journal of Applied Probability, 11(4): 742–751, December 1974. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3212557.

Iksanov:2018:FLT

[IJB18] Alexander Iksanov, Wissem Jedidi, and Fethi Bouzeffour. Functional limit theorems for the number of busy servers in a G/G/∞ queue. Journal of Applied Probability, 55(1):15–29, March 2018. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL https://www.cambridge.org/core/journals/
journal-of-applied-probability/article/functional-limit-
theorems-for-the-number-of-busy-servers-in-a-gg-queue/
661261DA4B4C004428D68C9CC1DB8B7D.

Iglehart:1970:WCA

[IK70] Donald L. Iglehart and Douglas P. Kennedy. Weak convergence of the average of flag processes. Journal of Applied Probability, 7(3):747–753, December 1970. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.

org/stable/3211951.

Ignatov:2004:FTR

[IK04] Zvetan G. Ignatov and Vladimir K. Kaishev. A finite-time ruin probability formula for continuous claim severities. Journal of Applied Probability, 41(2):570–578, June 2004. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3216037.

Ignatov:2006:IHP

[IK06] Zvetan G. Ignatov and Vladimir K. Kaishev. On the infinitehorizon probability of (non)ruin for integer-valued claims. Journal of Applied Probability, 43(2):535–551, June 2006. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/27595749.

Ichiba:2011:EEO

[IK11] Tomoyuki Ichiba and Constantinos Kardaras. Efficient estimation of one-dimensional diffusion first passage time densities via Monte Carlo simulation. Journal of Applied Probability, 48(3): 699–712, September 2011. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/23065930.

Iksanov:2016:CLT

[IK16] Alexander Iksanov and Zakhar Kabluchko. A central limit theorem and a law of the iterated logarithm for the Biggins martingale of the supercritical branching random walk. Journal of Applied Probability, 53(4):1178–1192, December 2016. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://projecteuclid.org/euclid.jap/1481132845.

Iksanov:2018:WCN

[IK18] Alexander Iksanov and Zakhar Kabluchko. Weak convergence of the number of vertices at intermediate levels of random recursive trees. Journal of Applied Probability, 55 (4):1131–1142, December 2018. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL https:/

/www.cambridge.org/core/journals/journal-of-applied-
probability/article/weak-convergence-of-the-number-of-
vertices-at-intermediate-levels-of-random-recursive-
trees/6BFF627E04B546212968FB8C6B3AFF36.

Irle:2001:AEO

[IKP01] A. Irle, O. Kubillus, and V. Paulsen. An asymptotic expansion for the optimal stopping boundary in problems with non-linear costs of observation. Journal of Applied Probability, 38(1):67– 79, March 2001. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3215742.

Ivkovic:1970:RFC

[IKS70] Z. Ivkovic, M. Kocher, and C. V. Stanojevic. On random nfold coincidences. Journal of Applied Probability, 7(3):785–787, December 1970. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3211958.

Iksanov:2004:IDI

[Iks04] Aleksander M. Iksanov. Infinite divisibility of infinite sums of lower records: A simple proof. Journal of Applied Probability, 41(4):1187–1190, December 2004. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/4141390. See correction [Ano06e].

Isaacson:1974:PCS

[IM74] Dean Isaacson and Richard Madsen. Positive columns for stochastic matrices. Journal of Applied Probability, 11(4):829– 835, December 1974. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3212567.

Itoh:2003:OSV

[IM03] Yoshiaki Itoh and Hosam M. Mahmoud. One-sided variations on interval trees. Journal of Applied Probability, 40(3):654–670, September 2003. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3215942.

Iksanov:2010:ERA

[IM10] Alexander Iksanov and Matthias Meiners. Exponential rate of almost-sure convergence of intrinsic martingales in supercritical branching random walks. Journal of Applied Probability, 47 (2):513–525, June 2010. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/20778494.

Imhof:1974:RDT

[Imh74] J. P. Imhof. Runs of a discrete time regenerative phenomenon. Journal of Applied Probability, 11(3):588–593, September 1974. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/3212705.

Imhof:1978:OSE

[Imh78] J. P. Imhof. One-sided excursions of Brownian motion, and an application. Journal of Applied Probability, 15(3):635–638, September 1978. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3213127.

Imhof:1984:DFB

[Imh84] J.-P. Imhof. Density factorizations for Brownian motion, meander and the three-dimensional Bessel process, and applications. Journal of Applied Probability, 21(3):500–510, September 1984. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/3213612.

Imhof:1999:SEL

[Imh99] J. P. Imhof. On some equalities of laws for Brownian motion with drift. Journal of Applied Probability, 36(3):682–687, September 1999. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3215433.

Imoto:2001:PPM

[Imo01] Masajiro Imoto. Point process modelling of reservoir-induced seismicity. Journal of Applied Probability, 38A:232–242, 2001. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/3215882.

Itoh:1998:ESI

[IMS98] Yoshiaki Itoh, Colin Mallows, and Larry Shepp. Explicit sufficient invariants for an interacting particle system. Journal of Applied Probability, 35(3):633–641, September 1998. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3215638.

Itoh:2000:OGG

[IMT00] Yoshiaki Itoh, Hiroshi Maehara, and Norihide Tokushige. Oriented graphs generated by random points on a circle. Journal of Applied Probability, 37(2):534–539, June 2000. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3215727.

Iversen:1986:CPV

[IN86] Villy Baek Iversen and Bo Friis Nielsen. Comments on a paper by V. Giorno, C. Negri and A. G. Nobile [1]. Journal of Applied Probability, 23(4):1047–1048, December 1986. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3214480. See [GNN86] and re-

ply [GNN86].

Inoue:2019:CRM

[Ino19] Yoshiaki Inoue. Comparison results for M/G/1 queues with waiting and sojourn time deadlines. Journal of Applied Probability, 56(2):524–532, June 2019. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL https:/

/www.cambridge.org/core/journals/journal-of-applied-
probability/article/comparison-results-for-mg1-queues-
with-waiting-and-sojourn-time-deadlines/686A7E4AD3CC8163D3D5D0C89688A157.

Inoue:2023:ESU

[Ino23] Kiyoshi Inoue. Exactly solvable urn models under random replacement schemes and their applications. Journal of Applied Probability, 60(3):835–854, September 2023. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL https://www.cambridge.org/core/journals/journal-
of-applied-probability/article/exactly-solvable-urn-
models-under-random-replacement-schemes-and-their-applications/
3DD3DBA0E5AF7D48697049B458D67400.

Iossif:1986:RPC

[Ios86] Gillian Iossif. Return probabilities for correlated random walks. Journal of Applied Probability, 23(1):201–207, March 1986. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/3214129.

Ibarrola:2011:CII

[IPR11] Ricardo Vélez Ibarrola and Tomás Prieto-Rumeau. Conditionally independent increment point processes. Journal of Applied Probability, 48(2):490–513, June 2011. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/23065898.

Ispany:2003:AIN

[IPV03] M. Ispány, G. Pap, and M. C. A. Van Zuijlen. Asymptotic inference for nearly unstable INAR(1) models. Journal of Applied Probability, 40(3):750–765, September 2003. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/3215948.

Iksanov:2020:FLT

[IR20] Alexander Iksanov and Bohdan Rashytov. A functional limit theorem for general shot noise processes. Journal of Applied Probability, 57(1):280–294, March 2020. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL https:/

/www.cambridge.org/core/journals/journal-of-applied-
probability/article/functional-limit-theorem-for-general-
shot-noise-processes/68C3F4476B32A6A9BA6F1524F9C9A7A3.

Irle:2003:SOC

[Irl03] A. Irle. Stochastic ordering for continuous-time processes. Journal of Applied Probability, 40(2):361–375, June 2003. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3215796.

Irle:2006:FAS

[Irl06] Albrecht Irle. A forward algorithm for solving optimal stopping problems. Journal of Applied Probability, 43(1):102–113, March 2006. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/27595713.

Iksanov:2023:RTI

[IRS23] Alexander Iksanov, Bohdan Rashytov, and Igor Samoilenko. Renewal theory for iterated perturbed random walks on a general branching process tree: Early generations. Journal of Applied Probability, 60(1):45–67, March 2023. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL https:/

/www.cambridge.org/core/journals/journal-of-applied-
probability/article/renewal-theory-for-iterated-perturbed-
random-walks-on-a-general-branching-process-tree-early-
generations/AC69E51D020A2CD9BC4EB28C3E8C0B5F.

Isobe:1983:WHE

[IS83] Etsuo Isobe and Shunsuke Sato. Wiener–Hermite expansion of a process generated by an Itô stochastic differential equation. Journal of Applied Probability, 20(4):754–765, December 1983. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/3213587.

Ivanoff:1985:CBP

[IS85] B. Gail Ivanoff and E. Seneta. The critical branching process with immigration stopped at zero. Journal of Applied Probability, 22(1):223–227, March 1985. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3213762.

Itoh:1986:RSC

[IS86] Yoshiaki Itoh and Herbert Solomon. Random sequential coding by Hamming distance. Journal of Applied Probability, 23(3): 688–695, September 1986. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3214007.

Isham:1974:EIE

[Ish74] Valerie Isham. The effect of intermittent exposure to risk on failure time distributions. Journal of Applied Probability, 11 (2):310–319, June 1974. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3212752.

Isham:1975:PPI

[Ish75] Valerie Isham. On a point process with independent locations. Journal of Applied Probability, 12(3):435–446, September 1975. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/3212858.

Isham:1977:MCM

[Ish77] Valerie Isham. A Markov construction for a multidimensional point process. Journal of Applied Probability, 14(3):507–515, September 1977. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3213453.

Isham:1980:DTP

[Ish80] Valerie Isham. Dependent thinning of point processes. Journal of Applied Probability, 17(4):987–995, December 1980. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3213208.

Ishizaki:2004:DPD

[Ish04] Fumio Ishizaki. Decomposition property in a discrete-time queue with multiple input streams and service interruptions. Journal of Applied Probability, 41(2):524–534, June 2004. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3216033.

Igaki:1995:AMR

[ISK95] Nobuko Igaki, Ushio Sumita, and Masashi Kowada. Analysis of Markov renewal shock models. Journal of Applied Probability, 32(3):821–831, September 1995. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.

org/stable/3215132.

Izadi:2018:NNC

[ISK18] Muhyiddin Izadi, Maryam Sharafi, and Baha-Eldin Khaledi. New nonparametric classes of distributions in terms of mean time to failure in age replacement. Journal of Applied Probability, 55(4):1238–1248, December 2018. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL https:/

/www.cambridge.org/core/journals/journal-of-applied-
probability/article/new-nonparametric-classes-of-distributions-
in-terms-of-mean-time-to-failure-in-age-replacement/
E2049BAE20411EBA6333ACC75E6C0A33.

Isokawa:1982:IPA

[Iso82] Y. Isokawa. An identification problem in almost and asymptotically almost periodically correlated processes. Journal of Applied Probability, 19(2):456–462, June 1982. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/3213501.

Isaacson:1978:CSE

[IT78] Dean Isaacson and Richard L. Tweedie. Criteria for strong ergodicity of Markov chains. Journal of Applied Probability, 15(1): 87–95, March 1978. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3213239.

Ito:1977:SDP

[Ito77] Yoshifusa Ito. Superposition of distinguishable point processes. Journal of Applied Probability, 14(1):200–204, March 1977. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/3213274.

Ito:1978:SDS

[Ito78] Yoshifusa Ito. Superposition and decomposition of stationary point processes. Journal of Applied Probability, 15(3):481–493, September 1978. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3213111.

Itoh:1979:RCM

[Ito79] Yoshiaki Itoh. Random collision models in oriented graphs. Journal of Applied Probability, 16(1):36–44, March 1979. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/3213372.

Itoh:1980:MGG

[Ito80] Yoshiaki Itoh. On the minimum of gaps generated by onedimensional random packing. Journal of Applied Probability, 17 (1):134–144, March 1980. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3212931.

Itoh:1989:CCR

[Ito89] Yoshiaki Itoh. A certain configuration of random points on a circle associated with a generalized Lotka–Volterra equation. Journal of Applied Probability, 26(4):898–900, December 1989. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/3214396.

Ivanoff:1980:BDI

[Iva80] B. G. Ivanoff. The branching diffusion with immigration. Journal of Applied Probability, 17(1):1–15, March 1980. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3212919.

Ivan:1981:WCS

[Iva81] Constantin Ivan. Weak convergence of the simple birth-and-death process. Journal of Applied Probability, 18(1):245–252, March 1981. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3213184.

Ivanov:2007:PAO

[Iva07] Roman V. Ivanov. On the pricing of American options in exponential Lévy markets. Journal of Applied Probability, 44(2):409– 419, June 2007. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
27595850.

Ivanovs:2010:MMB

[Iva10] Jevgenijs Ivanovs. Markov-modulated Brownian motion with two reflecting barriers. Journal of Applied Probability, 47(4): 1034–1047, December 2010. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.

org/stable/29777424.

Ivanovs:2014:PMO

[Iva14] Jevgenijs Ivanovs. Potential measures of one-sided Markov additive processes with reflecting and terminating barriers. Journal of Applied Probability, 51(4):1154–1170, December 2014. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://projecteuclid.org/euclid.jap/1421763333.

Ivanovs:2016:SAI

[Iva16] Jevgenijs Ivanovs. Sparre Andersen identity and the last passage time. Journal of Applied Probability, 53(2):600–605, June 2016. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://projecteuclid.org/euclid.jap/
1466172878.

Jacod:1975:TDP

[Jac75] Jean Jacod. Two dependent Poisson processes whose sum is still a Poisson process. Journal of Applied Probability, 12(1):170– 172, March 1975. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3212423.

Jacobs:1978:CQN

[Jac78] Patricia A. Jacobs. A cyclic queueing network with dependent exponential service times. Journal of Applied Probability, 15(3): 573–589, September 1978. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3213120.

Jacka:1999:KSA

[Jac99] S. D. Jacka. Keeping a satellite aloft: Two finite fuel stochastic control models. Journal of Applied Probability, 36(1):1– 20, March 1999. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3215398.

Jacka:2009:MCC

[Jac09] Saul Jacka. Markov chains conditioned never to wait too long at the origin. Journal of Applied Probability, 46(3):812–826, September 2009. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
25662462.

Jacobsen:2012:TRS

[Jac12] Martin Jacobsen. The time to ruin in some additive risk models with random premium rates. Journal of Applied Probability, 49 (4):915–938, December 2012. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.

org/stable/41713819.

Jagers:1969:PID

[Jag69] P. Jagers. The proportions of individuals of different kinds in two-type populations. A branching process problem arising in biology. Journal of Applied Probability, 6(2):249–260, August 1969. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3211996.

Jagers:1974:CGB

[Jag74a] Peter Jagers. Convergence of general branching processes and functionals thereof. Journal of Applied Probability, 11(3):471– 478, September 1974. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3212691.

Jagers:1974:GWP

[Jag74b] Peter Jagers. Galton–Watson processes in varying environments. Journal of Applied Probability, 11(1):174–178, March 1974. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/3212594.

Jagers:1992:SIP

[Jag92] Peter Jagers. Stabilities and instabilities in population dynamics. Journal of Applied Probability, 29(4):770–780, December 1992. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3214711.

James:1966:CEV

[Jam66] M. F. James. On the conditional expected value of contributions from a renewal process. Journal of Applied Probability, 3(2):

464–480, December 1966. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3212132.

Janson:1978:RDC

[Jan78] Svante Janson. On random divisions of a convex set. Journal of Applied Probability, 15(3):645–649, September 1978. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3213129.

Janson:1981:UBV

[Jan81] Svante Janson. An upper bound for the velocity of firstpassage percolation. Journal of Applied Probability, 18(1):256– 262, March 1981. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3213186.

Janson:1983:LTS

[Jan83] Svante Janson. Limit theorems for some sequential occupancy problems. Journal of Applied Probability, 20(3):545–553, September 1983. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3213891.

Janson:1988:RCC

[Jan88] Svante Janson. Random coverings of the circle by arcs with restricted endpoints. Journal of Applied Probability, 25(1):215– 219, March 1988. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3214248.

Janson:2001:MCC

[Jan01] Svante Janson. Moment convergence in conditional limit theorems. Journal of Applied Probability, 38(2):421–437, June 2001. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/3215897.

Janson:2015:DSD

[Jan15] Svante Janson. On degenerate sums of m-dependent variables. Journal of Applied Probability, 52(4):1146–1155, December 2015. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-

6072 (electronic). URL http://projecteuclid.org/euclid.
jap/1450802758.

Janson:2020:RCT

[Jan20] Svante Janson. Rate of convergence for traditional Pólya urns. Journal of Applied Probability, 57(4):1029–1044, December 2020. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL https://www.cambridge.org/core/journals/
journal-of-applied-probability/article/rate-of-convergence-
for-traditional-polya-urns/B818B666350B8659D7DB4066B73F15A4.

Jansen:2021:WCS

[Jan21] H. M. Jansen. Weak convergence of stochastic integrals with respect to the state occupation measure of a Markov chain. Journal of Applied Probability, 58(2):372–393, June 2021. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL https://www.cambridge.org/core/journals/journal-
of-applied-probability/article/weak-convergence-of-stochastic-
integrals-with-respect-to-the-state-occupation-measure-
of-a-markov-chain/76DCBF531AEED799ED58E229BB792593.

Jasra:2015:BBI

[Jas15] Ajay Jasra. On the behaviour of the backward interpretation of Feynman–Kac formulae under verifiable conditions. Journal of Applied Probability, 52(2):339–359, June 2015. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://projecteuclid.org/euclid.jap/1437658602.

Jaworski:1984:RMP

[Jaw84] Jerzy Jaworski. On a random mapping (T, P_{j}). Journal of Applied Probability, 21(1):186–191, March 1984. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3213677.

Jaworski:1999:EPD

[Jaw99] Jerzy Jaworski. Epidemic processes on digraphs of random mappings. Journal of Applied Probability, 36(3):780–798, September 1999. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3215441.

James:2006:ATM

[JC06a] Huw W. James and E. J. Collins. An analysis of transient Markov decision processes. Journal of Applied Probability, 43 (3):603–621, September 2006. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.

org/stable/27595759.

Jimenez:2006:LLA

[JC06b] J. C. Jimenez and F. Carbonell. Local linear approximations of jump diffusion processes. Journal of Applied Probability, 43 (1):185–194, March 2006. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/27595719.

Jiang:1998:ORC

[JCM98] Xiaoyue Jiang, Kan Cheng, and Viliam Makis. On the optimality of repair-cost-limit policies. Journal of Applied Probability, 35(4):936–949, December 1998. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.

org/stable/3215530.

Janvresse:2004:UDB

[JdlR04] Élise Janvresse and Thierry de la Rue. From uniform distributions to Benford’s Law. Journal of Applied Probability, 41(4): 1203–1210, December 2004. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://projecteuclid.

org/getRecord?id=euclid.jap/1101840566;   http://www.
jstor.org/stable/4141393;   https://www.univ-rouen.fr/
LMRS/Persopage/Delarue/Publis/PDF/uniform_distribution_
to_Benford_law.pdf.

Jenkins:1966:SJD

[Jen66] J. H. Jenkins. Stationary joint distributions arising in the analysis of the M/G/1 queue by the method of the imbedded Markov chain. Journal of Applied Probability, 3(2):512–520, December 1966. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3212135.

Jenkins:1972:RED

[Jen72] John H. Jenkins. The relative efficiency of direct and maximum likelihood estimates of mean waiting time in the simple queue, M/M/1. Journal of Applied Probability, 9(2):396–403, June 1972. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3212807.

Jensen:1974:SSD

[Jen74] Louis Jensen. Solving a singular diffusion equation occurring in population genetics. Journal of Applied Probability, 11(1): 1–15, March 1974. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3212578.

Jerwood:1970:NCS

[Jer70] D. Jerwood. A note on the cost of the simple epidemic. Journal of Applied Probability, 7(2):440–443, August 1970. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3211977.

Jerwood:1974:CCB

[Jer74] D. Jerwood. The cost of a carrier-borne epidemic. Journal of Applied Probability, 11(4):642–651, December 1974. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3212548.

Jiang:2016:QSE

[JF16] Guangxin Jiang and Michael C. Fu. Quantile sensitivity estimation for dependent sequences. Journal of Applied Probability, 53(3):715–732, September 2016. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http:/

/projecteuclid.org/euclid.jap/1476370772.

Javier:2023:SRM

[JF23] Kayla Javier and Brian Fralix. On the study of the running maximum and minimum level of level-dependent quasi-birth-death processes and related models. Journal of Applied Probability, 60(1):14–29, March 2023. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL https://www.cambridge.org/core/journals/
journal-of-applied-probability/article/on-the-study-
of-the-running-maximum-and-minimum-level-of-leveldependent-
quasibirthdeath-processes-and-related-models/064CDF6EFE379C1D5C1D8A81263DB2CA.

Jasso-Fuentes:2009:BOC

[JFHL09] Héctor Jasso-Fuentes and Onésimo Hernández-Lerma. Blackwell optimality for controlled diffusion processes. Journal of Applied Probability, 46(2):372–391, June 2009. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/25662431.

Jedynak:2012:TQN

[JFS12] Bruno Jedynak, Peter I. Frazier, and Raphael Sznitman. Twenty questions with noise: Bayes optimal policies for entropy loss.

Journal of Applied Probability, 49(1):114–136, March 2012. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/41713755.

Jensen:1985:SEM

[JG85] E. B. Jensen and H. J. G. Gundersen. The stereological estimation of moments of particle volume. Journal of Applied Probability, 22(1):82–98, March 1985. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3213750.

Jiamez:1991:CFP

[JGR91] R. Gutiérrez Jíamez, A. Juan Gonzalez, and P. Román Román. Construction of first-passage-time densities for a diffusion process which is not necessarily time-homogeneous. Journal of Applied Probability, 28(4):903–909, December 1991. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3214694.

Johnson:1988:CSE

[JI88] Jean Johnson and Dean Isaacson. Conditions for strong ergodicity using intensity matrices. Journal of Applied Probability, 25(1):34–42, March 1988. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3214231.

Jiang:2012:BEB

[Jia12] Hui Jiang. Berry–Esseen bounds and the law of the iterated logarithm for estimators of parameters in an Ornstein–Uhlenbeck process with linear drift. Journal of Applied Probability, 49(4): 978–989, December 2012. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/41713823.

Jiang:2015:ODP

[Jia15] Zhengjun Jiang. Optimal dividend policy when cash reserves follow a jump-diffusion process under Markov-regime switching. Journal of Applied Probability, 52(1):209–223, March 2015. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://projecteuclid.org/euclid.jap/1429282616.

Jin:1988:BER

[Jin88] Xing Jin. On Berry–Esseen rate for queue length of the GI/G/K system in heavy traffic. Journal of Applied Probability, 25(3): 596–611, September 1988. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3213987.

Jirina:1976:ENH

[Jir76] Miloslav Jirina. Extinction of non-homogeneous Galton–Watson processes. Journal of Applied Probability, 13(1):132–137, March 1976. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3212673.

Jensen:1971:DMP

[JK71] Arne Jensen and David Kendall. Denumerable Markov processes with bounded generators: A routine for calculating p_{ij}(∞). Journal of Applied Probability, 8(2):423–427, June 1971. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/3211914.

Johnson:1976:MGO

[JK76] N. L. Johnson and Samuel Kotz. On a multivariate generalized occupancy model. Journal of Applied Probability, 13(2):392– 399, June 1976. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3212847.

Janson:1981:MID

[JK81] Svante Janson and Olav Kallenberg. Maximizing the intersection density of fibre processes. Journal of Applied Probability, 18(4): 820–828, December 1981. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3213057.

Johnson:1987:MRS

[JK87] Norman L. Johnson and Samuel Kotz. A modified repair strategy for two-component systems with revealed and unrevealed faults. Journal of Applied Probability, 24(1):178–185, March 1987. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3214069.

See correction [Ano87e].

Jagers:2004:BPN

[JK04] Peter Jagers and Fima Klebaner. Branching processes in nearcritical random environments. Journal of Applied Probability, 41A:17–23, 2004. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3215965.

Jaffe:2020:RNN

[JKL^{+}20] Ariel Jaffe, Yuval Kluger, George C. Linderman, Gal Mishne, and Stefan Steinerberger. Randomized near-neighbor graphs, giant components and applications in data science. Journal of Applied Probability, 57(2):458–476, June 2020. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL https:/

/www.cambridge.org/core/journals/journal-of-applied-
probability/article/randomized-nearneighbor-graphs-giant-
components-and-applications-in-data-science/7A3CD4B815489F96B69F63A079037B06.

Jacod:2012:FRB

[JKM12] Jean Jacod, Claudia Klüppelberg, and Gernot Müller. Functional relationships between price and volatility jumps and their consequences for discretely observed data. Journal of Applied Probability, 49(4):901–914, December 2012. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/41713818.

Jakubowski:2011:PBA

[JKP11] Adam Jakubowski and Joanna Karlowska-Pik. Processes with block-associated increments. Journal of Applied Probability, 48 (2):514–526, June 2011. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/23065899.

Jelenkovic:1998:SAM

[JL98] Predrag R. Jelenković and Aurel A. Lazar. Subexponential asymptotics of a Markov-modulated random walk with queueing applications. Journal of Applied Probability, 35(2):325–347, June 1998. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3215689.

Joffe:2006:MLF

[JL06] A. Joffe and G. Letac. Multitype linear fractional branching processes. Journal of Applied Probability, 43(4):1091–1106, December 2006. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/27595802.

Janson:2012:PAP

[JL12] Svante Janson and Svante Linusson. The probability of the Alabama paradox. Journal of Applied Probability, 49(3):773–794, September 2012. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
41713805.

Ji:2022:CAS

[JL22] Lina Ji and Zenghu Li. Construction of age-structured branching processes by stochastic equations. Journal of Applied Probability, 59(3):670–684, September 2022. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL https://www.cambridge.org/core/journals/
journal-of-applied-probability/article/construction-
of-agestructured-branching-processes-by-stochastic-equations/
F086D601E262FE0CC68AE457826CD00A.

Janson:2010:LCP

[JLN10] Svante Janson, Tomasz Luczak, and Ilkka Norros. Large cliques in a power-law random graph. Journal of Applied Probability, 47(4):1124–1135, December 2010. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/29777431.

Ji:2022:MIS

[JLX22] Lina Ji, Huili Liu, and Jie Xiong. Mutually interacting superprocesses with migration. Journal of Applied Probability, 59(3):904–929, September 2022. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL https:/

/www.cambridge.org/core/journals/journal-of-applied-
probability/article/mutually-interacting-superprocesses-
with-migration/45B42E29494BBACA118ABAD95CB9A295.

Jensen:1986:SVI

[JM86] E. B. Jensen and J. Møller. Stereological versions of integral geometric formulae for n-dimensional ellipsoids. Journal of Applied Probability, 23(4):1031–1037, December 1986. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3214477.

Junior:2011:RPS

[JMZ11] Valdivino V. Junior, Fábio P. Machado, and Mauricio Zuluaga. Rumor processes on N. Journal of Applied Probability, 48(3): 624–636, September 2011. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/23065925.

Jasinski:2009:BVL

[JNR09] Krzysztof Jasiński, Jorge Navarro, and Tomasz Rychlik. Bounds on variances of lifetimes of coherent and mixed systems. Journal of Applied Probability, 46(3):894–908, September 2009. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/25662466.

Jelenkovic:2013:CRI

[JOC13] Predrag R. Jelenković and Mariana Olvera-Cravioto. Convergence rates in the implicit renewal theorem on trees. Journal of Applied Probability, 50(4):1077–1088, December 2013. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://projecteuclid.org/euclid.jap/1389370100.

Joe:1996:TSM

[Joe96] Harry Joe. Time series models with univariate margins in the convolution-closed infinitely divisible class. Journal of Applied Probability, 33(3):664–677, September 1996. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/3215348.

Joffe:1977:RPL

[Jof77] Anatole Joffe. Remark on a paper by S. Litvin. Advances in Applied Probability, 9(2):219–220, June 1977. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

www.jstor.org/stable/1426363. See [Lit69].

Joffe:1982:MPU

[Jof82] Anatole Joffe. Un modèle probabiliste utile en biologie moléculaire. (French) [Probabilistic model useful in molecular biology]. Journal of Applied Probability, 19(3):660–663, September 1982. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3213523.

Joffe:1989:OMM

[Jof89] A. Joffe. Obituary: Michel Métivier 1931–1988. Journal of Applied Probability, 26(2):426–429, June 1989. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/3214051.

Johnson:2006:CLT

[Joh06] Oliver Johnson. A central limit theorem for non-overlapping return times. Journal of Applied Probability, 43(1):32–47, March 2006. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/27595708.

Johnson:2024:ITC

[Joh24] Oliver Johnson. Information-theoretic convergence of extreme values to the Gumbel distribution. Journal of Applied Probability, 61(1):244–254, March 2024. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL https://www.cambridge.org/core/journals/journal-
of-applied-probability/article/informationtheoretic-
convergence-of-extreme-values-to-the-gumbel-distribution/
043C49254E64621B051D08464B986DF0.

Jones:1986:TSR

[Jon86] Richard H. Jones. Time series regression with unequally spaced data. Journal of Applied Probability, 23A:89–98, 1986. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3214345.

Jonasson:1999:RTM

[Jon99a] Johan Jonasson. The random triangle model. Journal of Applied Probability, 36(3):852–867, September 1999. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/3215446.

Jones:1999:CMT

[Jon99b] Owen Dafydd Jones. Continuity for multi-type branching processes with varying environments. Journal of Applied Probability, 36(1):139–145, March 1999. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3215409.

Jones:2004:LDS

[Jon04] Owen Dafydd Jones. Large deviations for supercritical multitype branching processes. Journal of Applied Probability, 41(3): 703–720, September 2004. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/4141348.

Jones:2019:RRT

[Jon19] Owen Dafydd Jones. Runoff on rooted trees. Journal of Applied Probability, 56(4):1065–1085, December 2019. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL https://www.cambridge.org/core/journals/journal-
of-applied-probability/article/runoff-on-rooted-trees/
9A6581FCFEC5079E99DBE3299DA3287E.

Jordan:2018:PAG

[Jor18]  Jonathan Jordan.  Preferential attachment graphs with

co-existing types of different fitnesses. Journal of Applied Probability, 55(4):1211–1227, December 2018. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL https://www.cambridge.org/core/journals/
journal-of-applied-probability/article/preferential-
attachment-graphs-with-coexisting-types-of-different-
fitnesses/700AD59174E74DFE69DD90C04AD202CD.

Joyce:1994:LRI

[Joy94] Paul Joyce. Likelihood ratios for the infinite alleles model. Journal of Applied Probability, 31(3):595–605, September 1994. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/3215140.

Joyce:1995:RES

[Joy95] Paul Joyce. Robustness of the Ewens sampling formula. Journal of Applied Probability, 32(3):609–622, September 1995. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3215116.

Joyce:1998:PSS

[Joy98] Paul Joyce. Partition structures and sufficient statistics. Journal of Applied Probability, 35(3):622–632, September 1998. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3215637.

Jensen:1969:RSA

[JP69] Louis Jensen and Edward Pollak. Random selective advantages of a gene in a finite population. Journal of Applied Probability, 6(1):19–37, April 1969. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3212274.

Jayakumar:1993:FOA

[JP93] K. Jayakumar and R. N. Pillai. The first-order autoregressive Mittag-Leffler process. Journal of Applied Probability, 30 (2):462–466, June 1993. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3214855.

Jalali:1995:MMT

[JP95] Assad Jalali and John Pemberton. Mixture models for time series. Journal of Applied Probability, 32(1):123–138, March 1995. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/3214925.

Jensen:1999:OUT

[JP99] Jens Ledet Jensen and Jan Pedersen. Ornstein–Uhlenbeck type processes with non-normal distribution. Journal of Applied Probability, 36(2):389–402, June 1999. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3215463.

Janson:2009:ISP

[JP09] Svante Janson and Niclas Petersson. The integral of the supremum process of Brownian motion. Journal of Applied Probability, 46(2):593–600, June 2009. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/25662446.

Jakubowski:2016:MCA

[JP16] J. Jakubowski and A. Pytel. The Markov consistency of Archimedean survival processes. Journal of Applied Probability, 53(2):392–409, June 2016. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://projecteuclid.

org/euclid.jap/1466172862.

Jacod:2018:MND

[JP18] Jean Jacod and Mark Podolskij. On the minimal number of driving Lévy motions in a multivariate price model. Journal of Applied Probability, 55(3):823–833, September 2018. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL https://www.cambridge.org/core/journals/journal-
of-applied-probability/article/on-the-minimal-number-
of-driving-levy-motions-in-a-multivariate-price-model/
E4289F4E836AE88B5F22611194D1B8AD.

Ji:2022:EMD

[JP22] Lanpeng Ji and Xiaofan Peng. Extrema of multi-dimensional Gaussian processes over random intervals. Journal of Applied Probability, 59(1):81–104, March 2022. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL https:/

/www.cambridge.org/core/journals/journal-of-applied-
probability/article/extrema-of-multidimensional-gaussian-
processes-over-random-intervals/D210413A43747BC3369DB1FE3BD7ED8F.

Jakeman:1995:EMP

[JPR95] Eric Jakeman, Sean Phayre, and Eric Renshaw. The evolution and measurement of a population of pairs. Journal of Applied Probability, 32(4):1048–1062, December 1995. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3215219.

Jacquier:2018:PLD

[JPS18] Antoine Jacquier, Mikko S. Pakkanen, and Henry Stone. Pathwise large deviations for the rough Bergomi model. Journal of Applied Probability, 55(4):1078–1092, December 2018. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL https://www.cambridge.org/core/journals/journal-
of-applied-probability/article/pathwise-large-deviations-
for-the-rough-bergomi-model/312D56AAED931FB7553EE6857D1CA945.

See corrigendum [GJP^{+}21].

Janssen:1982:SDR

[JR82a] J. Janssen and J. M. Reinhard. Some duality results for a class of multivariate semi-Markov processes. Journal of Applied Probability, 19(1):90–98, March 1982. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3213919.

Jewell:1982:CPR

[JR82b] Nicholas P. Jewell and Joseph P. Romano. Coverage problems and random convex hulls. Journal of Applied Probability, 19(3): 546–561, September 1982. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3213513.

Jacka:1995:WCC

[JR95] S. D. Jacka and G. O. Roberts. Weak convergence of conditioned processes on a countable state space. Journal of Applied Probability, 32(4):902–916, December 1995. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/3215203.

Jasinski:2016:IVO

[JR16] Krzysztof Jasiński and Tomasz Rychlik. Inequalities for variances of order statistics originating from urn models. Journal of Applied Probability, 53(1):162–173, March 2016. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://projecteuclid.org/euclid.jap/1457470566.

Jaimez:1995:NVI

[JRR95] R. Gutiérrez Jáimez, P. Román Román, and F. Torres Ruiz. A note on the Volterra integral equation for the first-passage-time probability density. Journal of Applied Probability, 32(3): 635–648, September 1995. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3215118.

Jelenkovic:2006:CSL

[JRS06] Predrag R. Jelenković, Ana Radovanović, and Mark S. Squillante. Critical sizing of LRU caches with dependent requests. Journal of Applied Probability, 43(4):1013–1027, December 2006. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/27595796.

Jain:1996:PKF

[JS96] Gautam Jain and Karl Sigman. A Pollaczek–Khintchine formula for M/G/1 queues with disasters. Journal of Applied Probability, 33(4):1191–1200, December 1996. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3214996.

Jorgensen:1998:STS

[JS98] Bent Jørgensen and Peter Xue-Kun Song. Stationary time series models with exponential dispersion model margins. Journal of Applied Probability, 35(1):78–92, March 1998. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/3215548. See correction

[Ano98f].

Jagers:2000:GGP

[JS00] Peter Jagers and Serik Sagitov. The growth of general population-size-dependent branching processes year by year. Journal of Applied Probability, 37(1):1–14, March 2000. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/3215655.

Jagers:2004:CCP

[JS04a] Peter Jagers and Serik Sagitov. Convergence to the coalescent in populations of substantially varying size. Journal of Applied Probability, 41(2):368–378, June 2004. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3216022.

Johnstone:2004:BCW

[JS04b] Iain M. Johnstone and Bernard W. Silverman. Boundary coiflets for wavelet shrinkage in function estimation. Journal of Applied Probability, 41A:81–98, 2004. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.

org/stable/3215970.

Jacka:2008:NVV

[JS08] Saul Jacka and Marcus Sheehan. The noisy veto-voter model: A recursive distributional equation on [0,1]. Journal of Applied Probability, 45(3):670–688, September 2008. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/27595979.

Janssen:2014:MTC

[JS14] A. Janssen and J. Segers. Markov tail chains. Journal of Applied Probability, 51(4):1133–1153, December 2014. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://projecteuclid.org/euclid.jap/1421763332.

Jacquier:2020:STM

[JS20] Antoine Jacquier and Fangwei Shi. Small-time moderate deviations for the randomised Heston model. Journal of Applied Probability, 57(1):19–28, March 2020. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL https:/

/www.cambridge.org/core/journals/journal-of-applied-
probability/article/smalltime-moderate-deviations-for-
the-randomised-heston-model/4D2B0FE8AF791AB608B8E21641D106F7.

Jiang:2019:SER

[JSW19] Yiming Jiang, Shiyu Song, and Yongjin Wang. Some explicit results on one kind of sticky diffusion. Journal of Applied Probability, 56(2):398–415, June 2019. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL https:/

/www.cambridge.org/core/journals/journal-of-applied-
probability/article/some-explicit-results-on-one-kind-
of-sticky-diffusion/B67FB33AA398544204479957C8993C92.

Joyce:1992:CTS

[JT92] Paul Joyce and Simon Tavaré. A convergence theorem for symmetric functionals of random partitions. Journal of Applied Probability, 29(2):280–290, June 1992. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3214566.

Jarner:2001:LCI

[JT01] S. F. Jarner and R. L. Tweedie. Locally contracting iterated functions and stability of Markov chains. Journal of Applied Probability, 38(2):494–507, June 2001. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3215902.

Jiang:2013:HTQ

[JTW^{+}13] Bo Jiang, Jian Tan, Wei Wei, Ness Shroff, and Don Towsley. Heavy tails in queueing systems: impact of parallelism on tail performance. Journal of Applied Probability, 50(1):127–150, March 2013. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-

6072 (electronic). URL http://projecteuclid.org/euclid.
jap/1363784429.

Jun:1994:CLD

[Jun94] Cai Jun. Characterizations of life distributions by moments of extremes and sample mean. Journal of Applied Probability, 31 (1):148–155, March 1994. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3215242.

Joffe:1982:EDK

[JW82] A. Joffe and W. A. O’N. Waugh. Exact distributions of kin numbers in a Galton–Watson process. Journal of Applied Probability, 19(4):767–775, December 1982. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3213829.

Joffe:1985:KNP

[JW85] A. Joffe and W. A. O’N. Waugh. The kin number problem in a multitype Galton–Watson population. Journal of Applied Probability, 22(1):37–47, March 1985. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3213746.

Jin:1990:SCQ

[JW90] Xing Jin and Rongxin Wang. On the speed of convergence for the queue length process of the GI/G/K system in heavy traffic. Journal of Applied Probability, 27(2):417–424, June 1990. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/3214661.

Johnson:2011:TPI

[JW11] Torrey Johnson and Edward C. Waymire. Tree polymers in the infinite volume limit at critical strong disorder. Journal of Applied Probability, 48(3):885–891, September 2011. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/23065945.

Jiang:2013:NTR

[JX13] Yunjiang Jiang and Weijun Xu. On the number of turns in reduced random lattice paths. Journal of Applied Probability, 50(2):499–515, June 2013. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://projecteuclid.

org/euclid.jap/1371648957.

Jiang:2021:ABW

[JY21] Hui Jiang and Qingshan Yang. Asymptotic behavior of the weak approximation to a class of Gaussian processes. Journal of Applied Probability, 58(3):693–707, September 2021. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL https://www.cambridge.org/core/journals/journal-
of-applied-probability/article/asymptotic-behavior-of-
the-weak-approximation-to-a-class-of-gaussian-processes/
1ACFE66B30B2C71F25BB7AB91C0B0D37.

Jagers:2002:BPD

[JZ02] Peter Jagers and Lu Zhunwei. Branching processes with deteriorating random environments. Journal of Applied Probability, 39(2):395–401, June 2002. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3216105.

Jiang:2011:RFU

[JZG11] Jianping Jiang, Sanguo Zhang, and Tiande Guo. Russo’s formula, uniqueness of the infinite cluster, and continuous differentiability of free energy for continuum percolation. Journal of Applied Probability, 48(3):597–610, September 2011. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/23065923.

Kyriakidis:1989:OPC

[KA89] E. G. Kyriakidis and Andris Abakuks. Optimal pest control through catastrophes. Journal of Applied Probability, 26(4): 873–879, December 1989. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3214392.

Koutras:1997:SWT

[KA97] M. V. Koutras and V. A. Alexandrou. Sooner waiting time problems in a sequence of trinary trials. Journal of Applied Probability, 34(3):593–609, September 1997. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/3215087.

Kurushima:2003:NFI

[KA03] Aiko Kurushima and Katsunori Ano. A note on the fullinformation Poisson arrival selection problem. Journal of Applied Probability, 40(4):1147–1154, December 2003. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3216066.

Kurushima:2009:MED

[KA09] Aiko Kurushima and Katsunori Ano. Maximizing the expected duration of owning a relatively best object in a Poisson process with rankable observations. Journal of Applied Probability, 46 (2):402–414, June 2009. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/25662433.

Khoolenjani:2016:EBI

[KA16] Nayereh Bagheri Khoolenjani and Mohammad Hossein Alamatsaz. Extension of de Bruijn’s identity to dependent non-Gaussian noise channels. Journal of Applied Probability, 53(2):360–368, June 2016. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-

6072 (electronic). URL http://projecteuclid.org/euclid.
jap/1466172859.

Kelkinnama:2022:ORA

[KA22] Maryam Kelkinnama and Majid Asadi. Optimal redundancy allocation in coherent systems with heterogeneous dependent components. Journal of Applied Probability, 59 (4):1144–1177, December 2022. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL https:/

/www.cambridge.org/core/journals/journal-of-applied-
probability/article/optimal-redundancy-allocation-in-
coherent-systems-with-heterogeneous-dependent-components/
4C38DF7D22D6707F529A7C8333C018EE.

Kabaila:1981:UPI

[Kab81] Paul V. Kabaila. The use of prior information in the estimation of the optimal interpolator parameter. Journal of Applied Probability, 18(1):302–308, March 1981. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3213193.

Kabaila:1983:PVA

[Kab83] Paul Kabaila. Parameter values of ARMA models minimising the one-step-ahead prediction error when the true system is not in the model set. Journal of Applied Probability, 20(2):405– 408, June 1983. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3213814.

Kabluchko:2012:PEB

[Kab12] Zakhar Kabluchko. Persistence and equilibria of branching populations with exponential intensity. Journal of Applied Probability, 49(1):226–244, March 2012. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/41713762.

Kadane:1974:CTA

[Kad74] Joseph B. Kadane. A characterization of triangular arrays which are expectations of order statistics. Journal of Applied Probability, 11(2):413–416, June 1974. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3212766.

KhudaBukhsh:2019:ALM

[KADK19] Wasiur R. KhudaBukhsh, Arnab Auddy, Yann Disser, and Heinz Koeppl. Approximate lumpability for Markovian agentbased models using local symmetries. Journal of Applied Probability, 56(3):647–671, September 2019. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL https:/

/www.cambridge.org/core/journals/journal-of-applied-
probability/article/approximate-lumpability-for-markovian-
agentbased-models-using-local-symmetries/A4B603163980ECC39DAC52A2477D6412.

Kallenberg:1975:LCT

[Kal75] Olav Kallenberg. Limits of compound and thinned point processes. Journal of Applied Probability, 12(2):269–278, June 1975. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/3212440.

Kalamatianou:1988:MRP

[Kal88] A. G. Kalamatianou. A model for responding to promotion blockages. Journal of Applied Probability, 25(2):268–278, June 1988. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3214435.

Kalpazidou:1989:RDM

[Kal89] S. Kalpazidou. Representation of denumerable Markov chains with multiple states by weighted circuits. Journal of Applied Probability, 26(1):23–35, March 1989. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3214313.

Kalpazidou:1990:ABS

[Kal90] S. Kalpazidou. Asymptotic behaviour of sample weighted circuits representing recurrent Markov chains. Journal of Applied Probability, 27(3):545–556, September 1990. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/3214540.

Kalpazidou:1991:ISP

[Kal91a] S. Kalpazidou. Invariant stochastic properties of a class of directed circuits. Journal of Applied Probability, 28(4):727–736, December 1991. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3214676.

Kalpazidou:1991:BIT

[Kal91b] S. Kalpazidou. On Beurling’s inequality in terms of thermal power. Journal of Applied Probability, 28(1):104–115, March 1991. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3214744.

Kalpazidou:1992:WCS

[Kal92] Sophia Kalpazidou. On the weak convergence of sequences of circuit processes: A probabilistic approach. Journal of Applied Probability, 29(2):374–383, June 1992. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3214574.

Kalpazidou:1993:LTC

[Kal93] S. Kalpazidou. On Levy’s theorem concerning positiveness of transition probabilities of Markov processes: The circuit processes case. Journal of Applied Probability, 30(1):28–39, March 1993. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3214619.

Kalamkar:1995:MPD

[Kal95] V. A. Kalamkar. Minification processes with discrete marginals. Journal of Applied Probability, 32(3):692–706, September 1995. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/3215123.

Kalashnikov:1996:OBV

[Kal96] Vladimir Kalashnikov. Obituary: Boris Vladimirovich Gnedenko. Journal of Applied Probability, 33(2):592–599, June 1996.

CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/3215081.

Kamerud:1979:RMG

[Kam79] Dana B. Kamerud. Repeated matrix games. Journal of Applied Probability, 16(4):830–842, December 1979. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/3213149.

Kampke:1987:OSP

[Käm87] Thomas Kämpke. On the optimality of static priority policies in stochastic scheduling on parallel machines. Journal of Applied Probability, 24(2):430–448, June 1987. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3214267.

Kammerle:1989:LFB

[Käm89] K. Kämmerle. Looking forwards and backwards in a bisexual Moran model. Journal of Applied Probability, 26(4):880–885, December 1989. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3214393.

Kammerle:1991:EPD

[Käm91] K. Kämmerle. The extinction probability of descendants in bisexual models of fixed population size. Journal of Applied Probability, 28(3):489–502, September 1991. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3214485.

Kamatani:2017:EMC

[Kam17] K. Kamatani. Ergodicity of Markov chain Monte Carlo with reversible proposal. Journal of Applied Probability, 54(2):638–654, June 2017. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL https:/

/www.cambridge.org/core/journals/journal-of-applied-
probability/article/ergodicity-of-markov-chain-monte-
carlo-with-reversible-proposal/298978348EEB96A6C521559F1DD84B39.

Kanter:1975:ADP

[Kan75] Marek Kanter. Autoregression for discrete processes mod 2. Journal of Applied Probability, 12(2):371–375, June 1975. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/3212453.

Kanter:1985:LBP

[Kan85] Marek Kanter. Lower bounds for the probability of overload in certain queueing networks. Journal of Applied Probability, 22 (2):429–436, June 1985. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3213785.

Kandji:2023:ECT

[Kan23] Baye Matar Kandji. Exponential control of the trajectories of iterated function systems with application to semistrong GARCH(P, Q) models. Journal of Applied Probability, 60(4):1501–1515, December 2023. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL https:/

/www.cambridge.org/core/journals/journal-of-applied-
probability/article/exponential-control-of-the-trajectories-
of-iterated-function-systems-with-application-to-semistrong-
garch-boldsymbolp-q-models/30A25E8B6C596FCCCAA6886A4136D6F3.

Kao:1974:NFT

[Kao74] Edward P. C. Kao. A note on the first two moments of times in transient states in a semi-Markov process. Journal of Applied Probability, 11(1):193–198, March 1974. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http:/

/www.jstor.org/stable/3212598.

Kao:1978:LDR

[Kao78] Peichuen Kao. Limiting diffusion for random walks with drift conditioned to stay positive. Journal of Applied Probability, 15 (2):280–291, June 1978. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3213401.

Kapadia:1971:GSF

[Kap71] Asha Seth Kapadia. The general N server finite queue. Journal of Applied Probability, 8(4):828–834, December 1971. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3212247.

Kaplan:1972:TCR

[Kap72] Norman Kaplan. A theorem on compositions of random probability generating functions and applications to branching processes with random environments. Journal of Applied Probability, 9(1):1–12, March 1972. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3212631.

Kaplan:1974:MAD

[Kap74] Norman Kaplan. Multidimensional age-dependent branching processes allowing immigration: The limiting distribution. Journal of Applied Probability, 11(2):225–236, June 1974. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3212744.

Kaplan:1975:NAD

[Kap75a] Norman Kaplan. A note on the age-dependent branching process with immigration. Journal of Applied Probability, 12(1):130– 134, March 1975. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3212414.

Kaplan:1975:NSG

[Kap75b] Norman Kaplan. A note on the supercritical generalized agedependent branching process. Journal of Applied Probability, 12 (2):341–345, June 1975. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3212448.

Kaplan:1977:GRE

[Kap77] Norman Kaplan. A generalization of a result of Erdős and Rényi. Journal of Applied Probability, 14(1):212–216, March 1977. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/3213276.

Kaplan:1978:LTR

[Kap78] Norman Kaplan. A limit theorem for random coverings of a circle which do not quite cover. Journal of Applied Probability, 15(2):443–446, June 1978. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3213417.

Kaplan:1981:ALT

[Kap81] Norman L. Kaplan. Another look at the two-lift problem. Journal of Applied Probability, 18(3):697–706, September 1981. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/3213323.

Kaplan:1982:NBR

[Kap82] Norman Kaplan. A note on the branching random walk. Journal of Applied Probability, 19(2):421–424, June 1982. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3213494.

Kaplun:2010:CTE

[Kap10] A. Kaplun. The continuous-time Ehrenfest process in term structure modelling. Journal of Applied Probability, 47(3):693–712, September 2010. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
20778511.

Karlin:1968:EBPa

[Kar68a] Samuel Karlin. Equilibrium behavior of population genetic models with non-random mating. Journal of Applied Probability, 5 (2):231–313, August 1968. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3212254.

Karlin:1968:EBPb

[Kar68b] Samuel Karlin. Equilibrium behavior of population genetic models with non-random mating: Part II: Pedigrees, homozygosity and stochastic models. Journal of Applied Probability, 5(3): 487–566, December 1968. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3211920.

Karlsson:1974:SMT

[Kar74] Jan-Erik Karlsson. A stochastic model for time lag in reporting of claims. Journal of Applied Probability, 11(2):382–387, June 1974. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3212761.

Karmeshu:1976:MPV

[Kar76a] Karmeshu. Motion of a particle in a velocity-dependent random force. Journal of Applied Probability, 13(4):684–695, December 1976. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3212523.

Karr:1976:SOD

[Kar76b] Alan F. Karr. Stability of one-dimensional systems of colliding particles. Journal of Applied Probability, 13(1):155–158, March 1976. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3212677.

Karr:1976:TEV

[Kar76c] Alan F. Karr. Two extreme value processes arising in hydrology. Journal of Applied Probability, 13(1):190–194, March 1976. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/3212683.

Karni:1977:PDD

[Kar77] Edi Karni. The probability distribution of the duration of the game in the classical ruin problem. Journal of Applied Probability, 14(2):416–420, June 1977. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3213016. See correction [Ano78h].

Karnin:1983:FRP

[Kar83] Ehud D. Karnin. The first repetition of a pattern in a symmetric Bernoulli sequence. Journal of Applied Probability, 20(2):413– 418, June 1983. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3213816.

Karlin:1988:CPA

[Kar88] Samuel Karlin. Coincident probabilities and applications in combinatorics. Journal of Applied Probability, 25A:185–200, 1988. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/3214156.

Karpowicz:2009:DOS

[Kar09] Anna Karpowicz. Double optimal stopping in the fishing problem. Journal of Applied Probability, 46(2):415–428, June 2009. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/25662434.

Kim:2011:MQH

[KAR11] Jung Hyun Kim, Hyun-Soo Ahn, and Rhonda Righter. Managing queues with heterogeneous servers. Journal of Applied Probability, 48(2):435–452, June 2011. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/23065895.

Kastenbaum:1966:DSO

[Kas66] Marvin A. Kastenbaum. A dialysis system with one absorbing and one semi-reflecting state. Journal of Applied Probability, 3 (2):365–374, December 1966. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.

org/stable/3212125.

Katz:1977:RDE

[Kat77a] Leo Katz. Representation of disease etiologies by certain stochastic models. Journal of Applied Probability, 14(4):663–674, December 1977. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3213341.

Katz:1977:ACD

[Kat77b] Richard W. Katz. An application of chain-dependent processes to meteorology. Journal of Applied Probability, 14(3):598–603, September 1977. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3213463.

Katayama:2002:NLC

[Kat02] Tsuyoshi Katayama. A note on level-crossing analysis for the excess, age, and spread distributions. Journal of Applied Probability, 39(4):896–900, December 2002. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3216012.

Katona:2005:WSF

[Kat05] Zsolt Katona. Width of a scale-free tree. Journal of Applied Probability, 42(3):839–850, September 2005. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/30040860.

Katayama:2011:SRV

[Kat11] Tsuyoshi Katayama. Some results for vacation systems with sojourn time limits. Journal of Applied Probability, 48(3):679–687, September 2011. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
23065928.

Katayama:2012:NMG

[Kat12] Tsuyoshi Katayama. A note on M/G/1 vacation systems with sojourn time limits. Journal of Applied Probability, 49(4):1194– 1199, December 2012. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/41713838.

Kim:1987:RIC

[KB87] Chul Kim and Laurence A. Baxter. Reliability importance for continuum structure functions. Journal of Applied Probability, 24(3):779–785, September 1987. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.

org/stable/3214108.

Konstantopoulos:1991:CPS

[KB91] P. Konstantopoulos and F. Baccelli. On the cut-off phenomenon in some queueing systems. Journal of Applied Probability, 28(3): 683–694, September 1991. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3214502.

Kella:2013:UMS

[KB13] Offer Kella and Onno Boxma. Useful martingales for stochastic storage processes with Lévy-type input. Journal of Applied Probability, 50(2):439–449, June 2013. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

projecteuclid.org/euclid.jap/1371648952.

Kella:2020:SLQ

[KB20] Offer Kella and Onno Boxma. Synchronized Lévy queues. Journal of Applied Probability, 57(4):1222–1233, December 2020. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL https://www.cambridge.org/core/journals/
journal-of-applied-probability/article/synchronized-
levy-queues/52E705840584E0DE83FA68F7BBB2B091.

Kantorovitz:2007:ABW

[KBBW07] M. R. Kantorovitz, H. S. Booth, C. J. Burden, and S. R. Wilson. Asymptotic behavior of k-word matches between two uniformly distributed sequences. Journal of Applied Probability, 44 (3):788–805, September 2007. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.

org/stable/27595882.

Kella:2006:LPR

[KBM06] Offer Kella, Onno Boxma, and Michel Mandjes. A Lévy process reflected at a Poisson age process. Journal of Applied Probability, 43(1):221–230, March 2006. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/27595722.

Khan:2021:CLD

[KBM21] Ruhul Ali Khan, Dhrubasish Bhattacharyya, and Murari Mitra. On classes of life distributions based on the mean time to failure function. Journal of Applied Probability, 58(2):289–313, June 2021. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL https:/

/www.cambridge.org/core/journals/journal-of-applied-
probability/article/on-classes-of-life-distributions-
based-on-the-mean-time-to-failure-function/D88E027E1922C7E422DC04792272B8C5.

Kim:1978:BDS

[KD78] Geung Ho Kim and H. T. David. Bivariate distributions as saddle points of mutual information. Journal of Applied Probability, 15(3):523–530, September 1978. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3213115.

Keats:1967:NBB

[Kea67] R. G. Keats. A note on the black box problem. Journal of Applied Probability, 4(1):113–122, April 1967. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/3212303.

Kebir:1997:LTR

[Keb97] Y. Kebir. Laplace transforms and the renewal equation. Journal of Applied Probability, 34(2):395–403, June 1997. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3215379.

Keilson:1964:AWH

[Kei64a] J. Keilson. An alternative to Wiener–Hopf methods for the study of bounded processes. Journal of Applied Probability, 1(1):85– 120, June 1964. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/

3212063. See correction [Ano64c].

Keilson:1964:RTB

[Kei64b] J. Keilson. A review of transient behavior in regular diffusion and birth-death processes. Journal of Applied Probability, 1(2): 247–266, December 1964. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/ stable/3211858. See correction [Ano65d].

Keilson:1965:RTB

[Kei65] J. Keilson. A review of transient behavior in regular diffusion and birth-death processes. Part II. Journal of Applied Probability, 2(2):405–428, December 1965. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3212202.

Keilson:1971:LCL

[Kei71] Julian Keilson. Log-concavity and log-convexity in passage time densities of diffusion and birth-death processes. Journal of Applied Probability, 8(2):391–398, June 1971. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/3211909.

Keilson:1978:EST

[Kei78] J. Keilson. Exponential spectra as a tool for the study of serversystems with several classes of customers. Journal of Applied Probability, 15(1):162–170, March 1978. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http:/

/www.jstor.org/stable/3213246.

Keilson:1982:DCS

[Kei82] J. Keilson. On the distribution and covariance structure of the present value of a random income stream. Journal of Applied Probability, 19(1):240–244, March 1982. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http:/

/www.jstor.org/stable/3213935.

Kelly:1975:NQC

[Kel75] F. P. Kelly. Networks of queues with customers of different types. Journal of Applied Probability, 12(3):542–554, September 1975. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3212869.

Kelly:1976:SPM

[Kel76] F. P. Kelly. On stochastic population models in genetics. Journal of Applied Probability, 13(1):127–131, March 1976. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3212672.

Kelly:1977:ABI

[Kel77] F. P. Kelly. The asymptotic behaviour of an invasion process. Journal of Applied Probability, 14(3):584–590, September 1977. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/3213461.

Kellerer:1983:NCR

[Kel83] A. M. Kellerer. On the number of clumps resulting from the overlap of randomly placed figures in a plane. Journal of Applied Probability, 20(1):126–135, March 1983. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http:/

/www.jstor.org/stable/3213726.

Kellerer:1985:CFP

[Kel85] A. M. Kellerer. Counting figures in planar random configurations. Journal of Applied Probability, 22(1):68–81, March 1985. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/3213749.

Kellerer:1986:VPP

[Kel86] A. M. Kellerer. The variance of a Poisson process of domains. Journal of Applied Probability, 23(2):307–321, June 1986. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/3214175.

Kella:1992:CRL

[Kel92] Offer Kella. Concavity and reflected Lévy processes. Journal of Applied Probability, 29(1):209–215, March 1992. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3214806.

Kella:1997:SSN

[Kel97] Offer Kella. Stochastic storage networks: Stationarity and the feedforward case. Journal of Applied Probability, 34(2):498– 507, June 1997. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3215388.

Kella:2000:NPF

[Kel00] Offer Kella. Non-product form of two-dimensional fluid networks with dependent Lévy inputs. Journal of Applied Probability, 37 (4):1117–1122, December 2000. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.

org/stable/3215501.

Kella:2009:GCP

[Kel09] Offer Kella. On growth-collapse processes with stationary structure and their shot-noise counterparts. Journal of Applied Probability, 46(2):363–371, June 2009. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/25662430.

Kella:2012:CDA

[Kel12] Offer Kella. The class of distributions associated with the generalized Pollaczek–Khinchine formula. Journal of Applied Probability, 49(3):883–887, September 2012. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/41713813.

Kemperman:1979:VTP

[Kem79] J. H. B. Kemperman. On the variance of a trimmed Poisson distribution. Journal of Applied Probability, 16(1):220–225, March 1979. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3213391.

Kemp:1992:SSM

[Kem92] A. W. Kemp. Steady-state Markov chain models for the Heine and Euler distributions. Journal of Applied Probability, 29(4): 869–876, December 1992. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3214719.

Kemp:1994:CLD

[Kem94] Adrienne W. Kemp. Comments on the Luria–Delbrück distribution. Journal of Applied Probability, 31(3):822–828, September 1994. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3215159.

Kemp:1998:ASA

[Kem98] Adrienne W. Kemp. Absorption sampling and the absorption distribution. Journal of Applied Probability, 35(2):489–494, June 1998. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3215702.

Kendall:1970:OAR

[Ken70] David Kendall. Obituary: Alfréd Rényi. Journal of Applied Probability, 7(2):509–522, August 1970. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http:/

/www.jstor.org/stable/3211992.

Kennedy:1972:CSS

[Ken72a] Douglas P. Kennedy. The continuity of the single server queue. Journal of Applied Probability, 9(2):370–381, June 1972. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/3212805.

Kennedy:1972:NNB

[Ken72b] Douglas P. Kennedy. A note on the number of busy servers in a GI/G/s queue in light traffic. Journal of Applied Probability, 9(4):868–869, December 1972. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.

org/stable/3212625.

Kennedy:1974:LDC

[Ken74] Douglas P. Kennedy. Limiting diffusions for the conditioned M/G/1 queue. Journal of Applied Probability, 11(2):355–362, June 1974. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3212756.

Kennedy:1975:GWP

[Ken75a] Douglas P. Kennedy. The Galton–Watson process conditioned on the total progeny. Journal of Applied Probability, 12(4): 800–806, December 1975. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3212730.

Kent:1975:WCT

[Ken75b] John T. Kent. A weak convergence theorem for the empirical characteristic function. Journal of Applied Probability, 12(3): 515–523, September 1975. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3212866.

Kennedy:1976:DMB

[Ken76] Douglas P. Kennedy. The distribution of the maximum Brownian excursion. Journal of Applied Probability, 13(2):371– 376, June 1976. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3212843.

Kennedy:1978:TFO

[Ken78] Douglas P. Kennedy. On the time to first overflow in dams with inputs forming a Markov chain. Journal of Applied Probability, 15(1):171–178, March 1978. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3213247.

Kendall:1982:SP

[Ken82a] David G. Kendall. The scale of perfection. Journal of Applied Probability, 19A:125–138, 1982. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.

org/stable/3213555.

Kennedy:1982:COS

[Ken82b] D. P. Kennedy. On a constrained optimal stopping problem. Journal of Applied Probability, 19(3):631–641, September 1982. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/3213520.

Kendall:1983:CMS

[Ken83] Wilfrid S. Kendall. Coupling methods and the storage equation. Journal of Applied Probability, 20(2):436–441, June 1983. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/3213820.

Kendall:1988:APP

[Ken88] David G. Kendall. Applied probability: Past, present and future. Journal of Applied Probability, 25A:25–30, 1988. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3214144.

Kendall:1990:SMP

[Ken90] W. S. Kendall. A spatial Markov property for nearest-neighbour Markov point processes. Journal of Applied Probability, 27(4): 767–778, December 1990. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3214821.

Kendall:1991:MDS

[Ken91] David G. Kendall. The Mardia–Dryden shape distribution for triangles: A stochastic calculus approach. Journal of Applied Probability, 28(1):225–230, March 1991. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http:/

/www.jstor.org/stable/3214753.

Kendall:1992:RPI

[Ken92] Wilfrid S. Kendall. A remark on the proof of Itô’s formula for C^{2} functions of continuous semimartingales. Journal of Applied Probability, 29(1):216–221, March 1992. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http:/

/www.jstor.org/stable/3214807.

Kendall:1993:ECP

[Ken93] Wilfrid S. Kendall. On the empty cells of Poisson histograms. Journal of Applied Probability, 30(3):561–574, September 1993. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/3214766.

Kennedy:1994:UWH

[Ken94] Joanne Kennedy. Understanding the Wiener–Hopf factorization for the simple random walk. Journal of Applied Probability, 31 (2):561–563, June 1994. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3215047.

Kersting:1986:RTG

[Ker86] G. Kersting. On recurrence and transience of growth models. Journal of Applied Probability, 23(3):614–625, September 1986.

CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/3214001.

Kersting:2020:UAB

[Ker20] Götz Kersting. A unifying approach to branching processes in a varying environment. Journal of Applied Probability, 57(1):196–220, March 2020. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL https:/

/www.cambridge.org/core/journals/journal-of-applied-
probability/article/unifying-approach-to-branching-processes-
in-a-varying-environment/79CAEC894042773600E6DD6BD57C652C.

Kessler:2008:ESS

[Kes08] David A. Kessler. Epidemic size in the SIS model of endemic infections. Journal of Applied Probability, 45(3):757–778, September 2008. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/27595985.

Kevei:2017:IRT

[Kev17] Péter Kevei. Implicit renewal theory in the arithmetic case. Journal of Applied Probability, 54(3):732–749, September 2017. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL https://www.cambridge.org/core/journals/
journal-of-applied-probability/article/implicit-renewal-
theory-in-the-arithmetic-case/B5D27B461C0263B2B75808FB60CD6200.

Karlin:1970:CET

[KF70] Samuel Karlin and Marcus W. Feldman. Convergence to equilibrium of the two locus additive viability model. Journal of Applied Probability, 7(2):262–271, August 1970. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http:/

/www.jstor.org/stable/3211963.

Kharoufeh:2006:API

[KFM06] Jeffrey P. Kharoufeh, Daniel E. Finkelstein, and Dusting G. Mixon. Availability of periodically inspected systems with Markovian wear and shocks. Journal of Applied Probability, 43 (2):303–317, June 2006. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/27595733.

Khan:1968:IDI

[KG68] M. S. Ali Khan and J. Gani. Infinite dams with inputs forming a Markov chain. Journal of Applied Probability, 5(1):72–83, April 1968. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3212078.

Kulkarni:2002:OAS

[KG02] V. G. Kulkarni and K. D. Glazebrook. Output analysis of a single-buffer multiclass queue: FCFS service. Journal of Applied Probability, 39(2):341–358, June 2002. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3216100.

Kotlarski:1975:CTU

[KH75] Ignacy I. Kotlarski and William E. Hinds. Characterization theorems using conditional expectations. Journal of Applied Probability, 12(2):400–406, June 1975. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3212458.

Kalpakam:1983:QSD

[KH83] S. Kalpakam and M. A. Shahul Hameed. Quasi-stationary distribution of a two-unit warm-standby redundant system. Journal of Applied Probability, 20(2):429–435, June 1983. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3213819.

Karatzas:1996:OSC

[KH96] I. Karatzas and C. C. Heyde. Obituary: Stamatis Cambanis (1943–1995). Journal of Applied Probability, 33(1):284, March 1996. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3215285.

Khazanie:1968:IAN

[Kha68] R. G. Khazanie. An indication of the asymptotic nature of the Mendelian Markov process. Journal of Applied Probability, 5 (2):350–356, August 1968. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3212257.

Khan:1970:FDI

[Kha70] M. S. Ali Khan. Finite dams with inputs forming a Markov chain. Journal of Applied Probability, 7(2):291–303, August 1970. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3211965.

Khan:1977:IDD

[Kha77] M. S. Ali Khan. Infinite dams with discrete additive inputs. Journal of Applied Probability, 14(1):170–180, March 1977. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/3213270.

Khmaladze:2011:CPC

[Khm11] Estáte V. Khmaladze. Convergence properties in certain occupancy problems including the Karlin–Rouault Law. Journal of Applied Probability, 48(4):1095–1113, December 2011. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/23066445.

Kohatsu-Higa:2003:MTB

[KHY03] Arturo Kohatsu-Higa and Makoto Yamazato. On moments and tail behaviors of storage processes. Journal of Applied Probability, 40(4):1069–1086, December 2003. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3216061.

Kijima:1987:SSF

[Kij87] Masaaki Kijima. Spectral structure of the first-passage-time densities for classes of Markov chains. Journal of Applied Probability, 24(3):631–643, September 1987. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3214095. See correction [Ano91i].

Kijima:1988:PCP

[Kij88] Masaaki Kijima. On passage and conditional passage times for Markov chains in continuous time. Journal of Applied Probability, 25(2):279–290, June 1988. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3214436.

Kijima:1989:SRR

[Kij89] Masaaki Kijima. Some results for repairable systems with general repair. Journal of Applied Probability, 26(1):89–102, March 1989. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3214319.

Kijima:1992:EDP

[Kij92a] Masaaki Kijima. Evaluation of the decay parameter for some specialized birth-death processes. Journal of Applied Probability, 29(4):781–791, December 1992. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.

org/stable/3214712.

Kijima:1992:EQS

[Kij92b] Masaaki Kijima. On the existence of quasi-stationary distributions in denumerable R-transient Markov chains. Journal of Applied Probability, 29(1):21–36, March 1992. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/3214788. See correction [Ano93f].

Kijima:1993:QLD

[Kij93] Masaaki Kijima. Quasi-limiting distributions of Markov chains that are skip-free to the left in continuous time. Journal of Applied Probability, 30(3):509–517, September 1993. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3214761.

Kijima:1998:HRR

[Kij98] Masaaki Kijima. Hazard rate and reversed hazard rate monotonicities in continuous-time Markov chains. Journal of Applied Probability, 35(3):545–556, September 1998. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/3215630.

Kimura:1964:DMP

[Kim64] Motoo Kimura. Diffusion models in population genetics. Journal of Applied Probability, 1(2):177–232, December 1964. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3211856.

Kim:1970:TCT

[Kim70a] C. W. Kim. Two classification theorems of states of Markov chains. Journal of Applied Probability, 7(2):490–496, August 1970. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3211988.

Kimbleton:1970:SPR

[Kim70b] Stephen R. Kimbleton. A simple proof of a random stable limit theorem. Journal of Applied Probability, 7(2):502–504, August 1970. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3211990.

Kim:1994:MBN

[Kim94a] Tae Yoon Kim. Moment bounds for non-stationary dependent sequences. Journal of Applied Probability, 31(3):731–742, September 1994. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3215151.

Kimber:1994:ARP

[Kim94b] Alan C. Kimber. An application of random packing to the multiple fracture of composite materials. Journal of Applied Probability, 31(2):564–569, June 1994. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3215048.

Kim:2000:USC

[Kim00] Sun-Tsung Kim. A use of the Stein–Chen method in time series analysis. Journal of Applied Probability, 37(4):1129–1136, December 2000. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3215503.

Kim:2016:OIS

[Kim16] Je Guk Kim. Optimal importance sampling for the Laplace transform of exponential Brownian functionals. Journal of Applied Probability, 53(2):531–542, June 2016. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

projecteuclid.org/euclid.jap/1466172872.

Kim:2017:CQR

[Kim17] Je Guk Kim. On the convergence of the quasi-regression method: polynomial chaos and regularity. Journal of Applied Probability, 54(2):424–443, June 2017. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL https:/

/www.cambridge.org/core/journals/journal-of-applied-
probability/article/on-the-convergence-of-the-quasiregression-
method-polynomial-chaos-and-regularity/8F4D10C330F1E7D350CD5329B5CADFBE.

Kingman:1965:MFP

[Kin65] J. F. C. Kingman. Mean free paths in a convex reflecting region. Journal of Applied Probability, 2(1):162–168, June 1965.

CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/3211881.

Kingman:1966:AQ

[Kin66] J. F. C. Kingman. On the algebra of queues. Journal of Applied Probability, 3(2):285–326, December 1966. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/3212122.

Kingman:1969:MPP

[Kin69a] J. F. C. Kingman. Markov population processes. Journal of Applied Probability, 6(1):1–18, April 1969. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/3212273.

Kingman:1969:RSC

[Kin69b] J. F. C. Kingman. Random secants of a convex body. Journal of Applied Probability, 6(3):660–672, December 1969. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3212110.

Kingman:1977:RSD

[Kin77] J. F. C. Kingman. Remarks on the spatial distribution of a reproducing population. Journal of Applied Probability, 14(3): 577–583, September 1977. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3213460.

Kingman:1978:SMB

[Kin78] J. F. C. Kingman. A simple model for the balance between selection and mutation. Journal of Applied Probability, 15(1): 1–12, March 1978. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3213231.

Kingman:1982:GLP

[Kin82] J. F. C. Kingman. On the genealogy of large populations. Journal of Applied Probability, 19A:27–43, 1982. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/3213548.

Kingman:1988:TPM

[Kin88] J. F. C. Kingman. Typical polymorphisms maintained by selection at a single locus. Journal of Applied Probability, 25A:

113–125, 1988. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3214150.

Kingman:2002:SAL

[Kin02] J. F. C. Kingman. Stochastic aspects of Lanchester’s theory of warfare. Journal of Applied Probability, 39(3):455–465, September 2002. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3216070.

Kingman:2004:EPR

[Kin04] J. F. C. Kingman. Extremal problems for regenerative phenomena. Journal of Applied Probability, 41A:333–346, 2004. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3215987.

Karmeshu:1981:MIM

[KJ81] Karmeshu and N. K. Jaiswal. A machine interference model with threshold effect. Journal of Applied Probability, 18(2):491– 498, June 1981. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3213295.

Koskela:2015:CIB

[KJS15] Jere Koskela, Paul Jenkins, and Dario Spanò. Computational inference beyond Kingman’s coalescent. Journal of Applied Probability, 52(2):519–537, June 2015. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http:/

/projecteuclid.org/euclid.jap/1437658613.

Kemp:1969:BCM

[KK69] Adrienne W. Kemp and C. D. Kemp. Branching and clustering models associated with the ‘lost-games distribution’. Journal of Applied Probability, 6(3):700–703, December 1969. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3212115.

Kofman:1995:SEP

[KK95] D. Kofman and H. Korezlioglu. Some EATA properties for marked point processes. Journal of Applied Probability, 32(4): 922–929, December 1995. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3215205.

Karpelevitch:1996:AAQ

[KK96a] F. I. Karpelevitch and A. Ya. Kreinin. Asymptotic analysis of queueing systems with identical service. Journal of Applied Probability, 33(1):267–281, March 1996. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http:/

/www.jstor.org/stable/3215283.

Kaspi:1996:SFF

[KK96b] Haya Kaspi and Offer Kella. Stability of feed-forward fluid networks with Lévy input. Journal of Applied Probability, 33 (2):513–522, June 1996. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3215075.

Kennedy:1997:PII

[KK97] D. P. Kennedy and R. P. Kertz. A prophet inequality for independent random variables with finite variances. Journal of Applied Probability, 34(4):945–958, December 1997. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3215009.

Khaledi:2000:SNR

[KK00] Baha-Eldin Khaledi and Subhash Kochar. Some new results on stochastic comparisons of parallel systems. Journal of Applied Probability, 37(4):1123–1128, December 2000. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3215502.

Komorowski:2001:RWR

[KK01] T. Komorowski and G. Krupa. Random walk in a random environment with correlated sites. Journal of Applied Probability, 38(4):1018–1032, December 2001. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3215781.

Kaj:2003:CPP

[KK03] Ingemar Kaj and Stephen M. Krone. The coalescent process in a population with stochastically varying size. Journal of Applied Probability, 40(1):33–48, March 2003. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3215840.

Kluppelberg:2006:ERB

[KK06] C. Klüppelberg and A. E. Kyprianou. On extreme ruinous behaviour of Levy insurance risk processes. Journal of Applied Probability, 43(2):594–598, June 2006. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/27595753.

Katzur:2016:OSM

[KK16] Alexander Katzur and Udo Kamps. Order statistics with memory: a model with reliability applications. Journal of Applied Probability, 53(4):974–988, December 2016. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

projecteuclid.org/euclid.jap/1481132830.

Kim:2018:HRO

[KK18] Bara Kim and Jeongsim Kim. Hazard rate ordering of the largest order statistics from geometric random variables. Journal of Applied Probability, 55(2):652–658, June 2018. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL https://www.cambridge.org/core/journals/journal-
of-applied-probability/article/hazard-rate-ordering-
of-the-largest-order-statistics-from-geometric-random-
variables/F81C0601573B3D829F8760D77F299B2F.

Klump:2024:EAI

[KK24] Alexander Klump and Martin Kolb. An elementary approach to the inverse first-passage-time problem for softkilled Brownian motion. Journal of Applied Probability, 61(1):279–300, March 2024. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL https:/

/www.cambridge.org/core/journals/journal-of-applied-
probability/article/an-elementary-approach-to-the-inverse-
firstpassagetime-problem-for-softkilled-brownian-motion/
669EDC0E3C318959335E34F02309F8F7.

Koroljuk:1982:LE

[KKC82] V. S. Koroljuk, Vidyadhar G. Kulkarni, and B. W. Conolly. Letters to the Editor. Journal of Applied Probability, 19(4): 901–905, December 1982. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3213849.

Kim:2007:TAQ

[KKK07] Jerim Kim, Bara Kim, and Sung-Seok Ko. Tail asymptotics for the queue size distribution in an M/G/1 retrial queue. Journal of Applied Probability, 44(4):1111–1118, December 2007. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/27595912.

Kaj:2001:CTS

[KKL01] Ingemar Kaj, Stephen M. Krone, and Martin Lascoux. Coalescent theory for seed bank models. Journal of Applied Probability, 38(2):285–300, June 2001. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3215889.

Karmarkar:1986:PAO

[KKLO86] Narendra Karmarkar, Richard M. Karp, George S. Lueker, and Andrew M. Odlyzko. Probabilistic analysis of optimum partitioning. Journal of Applied Probability, 23(3):626–645, September 1986. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3214002.

Kondratiev:2010:GRF

[KKP10] Yuri Kondratiev, Yuri Kozitsky, and Tanja Pasurek. Gibbs random fields with unbounded spins on unbounded degree graphs. Journal of Applied Probability, 47(3):856–875, September 2010. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/20778521.

Klebanov:2012:CDS

[KKRT12] L. B. Klebanov, A. V. Kakosyan, S. T. Rachev, and G. Temnov. On a class of distributions stable under random summation. Journal of Applied Probability, 49(2):303–318, June 2012. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/41713769.

Kim:2015:FET

[KKS15] Yoora Kim, Irem Koprulu, and Ness B. Shroff. First exit time of a Lévy flight from a bounded region in R^{N}. Journal of Applied Probability, 52(3):649–664, September 2015. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

projecteuclid.org/euclid.jap/1445543838.

Kiessler:2002:API

[KKY02] Peter C. Kiessler, Georgia-Ann Klutke, and Yoonjung Yang. Availability of periodically inspected systems subject to Markovian degradation. Journal of Applied Probability, 39(4):700–711, December 2002. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3215997.

Kimeldorf:1977:APN

[KL77] George Kimeldorf and John Patrick Lang. Asymptotic properties of non-discrete duels. Journal of Applied Probability, 14 (1):153–161, March 1977. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3213268.

Kimeldorf:1978:APD

[KL78] George Kimeldorf and John Patrick Lang. Asymptotic properties of discrete duels. Journal of Applied Probability, 15(2):374– 383, June 1978. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3213408.

Kryscio:1989:ESL

[KL89a] Richard J. Kryscio and Claude Lefévre. On the extinction of the S-I-S stochastic logistic epidemic. Journal of Applied Probability, 26(4):685–694, December 1989. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.

org/stable/3214374.

Kupka:1989:HVM

[KL89b] Joseph Kupka and Sonny Loo. The hazard and vitality measures of ageing. Journal of Applied Probability, 26(3):532–542, September 1989. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3214411.

Konstantopoulos:2000:DPS

[KL00] Takis Konstantopoulos and Günter Last. On the dynamics and performance of stochastic fluid systems. Journal of Applied Probability, 37(3):652–667, September 2000. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/3215603.

Kazakevicius:2003:NTE

[KL03a] Vytautas Kazakevicius and Remigijus Leipus. A new theorem on the existence of invariant distributions with applications to arch processes. Journal of Applied Probability, 40(1):147–162, March 2003. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3215846.

Khoshnevisan:2003:ORS

[KL03b] Davar Khoshnevisan and Thomas M. Lewis. Optimal reward on a sparse tree with random edge weights. Journal of Applied Probability, 40(4):926–945, December 2003. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/3216051.

Kordzakhia:2006:EMP

[KL06] George Kordzakhia and Steven P. Lalley. Ergodicity and mixing properties of the northeast model. Journal of Applied Probability, 43(3):782–792, September 2006. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/27595772.

Kassberger:2016:AAC

[KL16] Stefan Kassberger and Thomas Liebmann. An alternative axiomatic characterisation of pricing operators. Journal of Applied Probability, 53(4):1257–1264, December 2016. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://projecteuclid.org/euclid.jap/1481132850.

Koutras:2017:NMI

[KL17] Markos V. Koutras and Demetrios P. Lyberopoulos. New maximal inequalities for N-demimartingales with scan statistic applications. Journal of Applied Probability, 54(2):363–378, June 2017. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-

6072 (electronic). URL https://www.cambridge.org/core/
journals/journal-of-applied-probability/article/new-
maximal-inequalities-for-ndemimartingales-with-scan-
statistic-applications/140CBCA41B498FE751FCE855F8642070.

Klatt:2022:SRH

[KL22] Michael Andreas Klatt and Günter Last. On strongly rigid hyperfluctuating random measures. Journal of Applied Probability, 59(4):948–961, December 2022. CODEN JPRBAM.

ISSN 0021-9002 (print), 1475-6072 (electronic). URL https:/

/www.cambridge.org/core/journals/journal-of-applied-
probability/article/on-strongly-rigid-hyperfluctuating-
random-measures/818B510BF11BF85FBA9E03B381E338CB.

Klaassen:1994:DCS

[Kla94] Chris A. J. Klaassen. Dixie cups: Sampling with replacement from a finite population. Journal of Applied Probability, 31(4): 940–948, December 1994. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3215319.

Klar:2002:NCL

[Kla02] B. Klar. A note on the L-class of life distributions. Journal of Applied Probability, 39(1):11–19, March 2002. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3215914.

Klefsjo:1981:SUP

[Kle81] Bengt Klefsjö. Survival under the pure birth shock model. Journal of Applied Probability, 18(2):554–560, June 1981. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3213305.

Klebaner:1983:PSD

[Kle83a] F. C. Klebaner. Population-size-dependent branching process with linear rate of growth. Journal of Applied Probability, 20 (2):242–250, June 1983. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3213798.

Klefsjo:1983:UAP

[Kle83b] Bengt Klefsjö. A useful ageing property based on the Laplace transform. Journal of Applied Probability, 20(3):615–626, September 1983. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3213897.

Klebaner:1984:GRG

[Kle84] F. C. Klebaner. Geometric rate of growth in population-size-dependent branching processes. Journal of Applied Probability, 21(1):40–49, March 1984. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3213662.

Klebaner:1985:LTP

[Kle85] F. C. Klebaner. A limit theorem for population-size-dependent branching processes. Journal of Applied Probability, 22(1):48– 57, March 1985. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3213747.

Klebaner:1988:CFA

[Kle88] Fima C. Klebaner. Conditions for fixation of an allele in the density-dependent Wright–Fisher models. Journal of Applied Probability, 25(2):247–256, June 1988. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3214433.

Klebaner:1989:LGN

[Kle89] Fima C. Klebaner. Linear growth in near-critical populationsize-dependent multitype Galton–Watson processes. Journal of Applied Probability, 26(3):431–445, September 1989. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/3214402. See correction

[Ano92g].

Klebaner:1991:ABN

[Kle91] Fima C. Klebaner. Asymptotic behavior of near-critical multitype branching processes. Journal of Applied Probability, 28(3): 512–519, September 1991. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/ stable/3214487. See corrections [Ano91h, Ano92f].

Klebaner:1994:ABM

[Kle94] F. C. Klebaner. Asymptotic behaviour of Markov population processes with asymptotically linear rate of change. Journal of Applied Probability, 31(3):614–625, September 1994. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3215142.

Klimmek:2014:PDO

[Kli14] Martin Klimmek. Parameter dependent optimal thresholds, indifference levels and inverse optimal stopping problems. Journal of Applied Probability, 51(2):492–511, June 2014. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://projecteuclid.org/euclid.jap/1402578639.

Kovacs:2015:BEA

[KLL15] Mihály Kovács, Stig Larsson, and Fredrik Lindgren. On the backward Euler approximation of the stochastic Allen–Cahn equation. Journal of Applied Probability, 52(2):323–338, June 2015. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://projecteuclid.org/euclid.jap/
1437658601.

Kluppelberg:2004:CTG

[KLM04] Claudia Klüppelberg, Alexander Lindner, and Ross Maller. A continuous-time GARCH process driven by a Lévy process: Stationarity and second-order behaviour. Journal of Applied Probability, 41(3):601–622, September 2004. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/4141341.

Korilis:1999:ABP

[KLO99] Yannis A. Korilis, Aurel A. Lazar, and Ariel Orda. Avoiding the Braess paradox in non-cooperative networks. Journal of Applied Probability, 36(1):211–222, March 1999. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http:/

/www.jstor.org/stable/3215415.

Kyprianou:2012:OCA

[KLP12] Andreas E. Kyprianou, Ronnie Loeffen, and José-Luis Pérez. Optimal control with absolutely continuous strategies for spectrally negative Lévy processes. Journal of Applied Probability, 49 (1):150–166, March 2012. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/41713757.

Klein:2019:CBE

[KLP19] Thierry Klein, Agnés Lagnoux, and Pierre Petit. A conditional Berry–Esseen inequality. Journal of Applied Probability, 56(1):76–90, March 2019. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL https:/

/www.cambridge.org/core/journals/journal-of-applied-
probability/article/conditional-berryesseen-inequality/
6B5F9E7D69C5E27B968C3CB7784038F1.

Klein:2022:LDR

[KLP22] Thierry Klein, Agnès Lagnoux, and Pierre Petit. Large-deviation results for triangular arrays of semiexponential random variables. Journal of Applied Probability, 59(2):399–420, June 2022. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL https://www.cambridge.org/core/journals/
journal-of-applied-probability/article/largedeviation-
results-for-triangular-arrays-of-semiexponential-random-
variables/680ED1EF1189B2495343F5C4DC50F78F.

Kluppelberg:1988:SDI

[Klü88] Claudia Klüppelberg. Subexponential distributions and integrated tails. Journal of Applied Probability, 25(1):132–141, March 1988. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3214240.

Konstantopoulos:2016:LTA

[KLY16] Takis Konstantopoulos, Zhenxia Liu, and Xiangfeng Yang. Laplace transform asymptotics and large deviation principles for longest success runs in Bernoulli trials. Journal of Applied Probability, 53(3):747–764, September 2016. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

projecteuclid.org/euclid.jap/1476370774.

Karlin:1965:EUM

[KM65] Samuel Karlin and James McGregor. Ehrenfest urn models. Journal of Applied Probability, 2(2):352–376, December 1965. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/3212199.

Khazanie:1966:MMP

[KM66] R. G. Khazanie and H. E. McKean. A Mendelian Markov process with multinomial transition probabilities. Journal of Applied Probability, 3(2):353–364, December 1966. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/3212124.

Kendall:1987:RMV

[KM87] Wilfrid S. Kendall and Joseph Mecke. The range of the meanvalue quantities of planar tessellations. Journal of Applied Probability, 24(2):411–421, June 1987. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3214265.

Konig:1988:RDD

[KM88] D. König and M. Miyazawa. Relationships and decomposition in the delayed Bernoulli feedback queueing system. Journal of Applied Probability, 25(1):169–183, March 1988. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3214243.

Kijima:1990:IES

[KM90] Masaaki Kijima and Naoki Makimoto. On interchangeability for exponential single-server queues in tandem. Journal of Applied Probability, 27(2):459–464, June 1990. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3214667.

Kovats:1992:APC

[KM92] Antal Kováts and Tamás F. Móri. Ageing properties of certain dependent geometric sums. Journal of Applied Probability, 29(3): 655–666, September 1992. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3214902.

Kino:1993:SWS

[KM93] Issei Kino and Masakiyo Miyazawa. The stationary work in system of a G/G/1 gradual input queue. Journal of Applied Probability, 30(1):207–222, March 1993. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3214633.

Kluppelberg:1997:LDH

[KM97] C. Klüppelberg and T. Mikosch. Large deviations of heavy-tailed random sums with applications in insurance and finance. Journal of Applied Probability, 34(2):293–308, June 1997. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3215371.

Kella:2001:PFQ

[KM01] Offer Kella and Masakiyo Miyazawa. Parallel fluid queues with constant inflows and simultaneous random reductions. Journal of Applied Probability, 38(3):609–620, September 2001. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3216115.

Klar:2003:CCL

[KM03] Bernhard Klar and Alfred Müller. Characterizations of classes of lifetime distributions generalizing the NBUE class. Journal of Applied Probability, 40(1):20–32, March 2003. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3215839.

Kesten:2004:SET

[KM04] Harry Kesten and Ross Maller. Some effects of trimming on the law of the iterated logarithm. Journal of Applied Probability, 41A:253–271, 2004. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3215981.

Kalinkin:2015:LTW

[KM15a] A. V. Kalinkin and A. V. Mastikhin. A limit theorem for a Weiss epidemic process. Journal of Applied Probability, 52(1):247–257, March 2015. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-

6072 (electronic). URL http://projecteuclid.org/euclid.
jap/1429282619.

Kosinski:2015:LAM

[KM15b] K. M. Kosiński and M. Mandjes. Logarithmic asymptotics for multidimensional extremes under nonlinear scalings. Journal of Applied Probability, 52(1):68–81, March 2015. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://projecteuclid.org/euclid.jap/1429282607.

Kella:2023:RLP

[KM23] Offer Kella and Michel Mandjes. From reflected Lévy processes to stochastically monotone Markov processes via generalized inverses and supermodularity. Journal of Applied Probability, 60(1):68–84, March 2023. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL https:/

/www.cambridge.org/core/journals/journal-of-applied-
probability/article/from-reflected-levy-processes-to-
stochastically-monotone-markov-processes-via-generalized-
inverses-and-supermodularity/82EE96C5E66C70D71DAF6B5039A0AAEE.

Kolkovska:2019:DAB

[KMG19] Ekaterina T. Kolkovska and Ehyter M. Martín-González. The distribution and asympotic behaviour of the negative Wiener– Hopf factor for Lévy processes with rational positive jumps.

Journal of Applied Probability, 56(4):1086–1105, December 2019. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL https://www.cambridge.org/core/journals/
journal-of-applied-probability/article/distribution-
and-asympotic-behaviour-of-the-negative-wienerhopf-factor-
for-levy-processes-with-rational-positive-jumps/CC872656B4125B3BF2F1623ECC704693.

Konheim:1981:PST

[KMM81] Alan G. Konheim, Isaac Meilijson, and Abraham Melkman. Processor-sharing of two parallel lines. Journal of Applied Probability, 18(4):952–956, December 1981. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3213071.

Keenay:1975:CPA

[KMR75] G. A. Keenay, R. W. Morgan, and K. H. Ray. Convergence properties of age distributions. Journal of Applied Probability, 12(4):684–691, December 1975. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.

org/stable/3212719.

Klebanov:2000:NRC

[KMRV00] Lev B. Klebanov, Stefan Mittnik, Svetlozar T. Rachev, and Vladimir E. Volkovich. A new representation for the characteristic function of strictly geo-stable vectors. Journal of Applied Probability, 37(4):1137–1142, December 2000. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3215504.

Konig:1983:IPA

[KMS83] Dieter König, Masakiyo Miyazawa, and Volker Schmidt. On the identification of Poisson arrivals in queues with coinciding timestationary and customer-stationary state distributions. Journal of Applied Probability, 20(4):860–871, December 1983. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3213597.

Kyprianou:2012:ABD

[KMSP12] A. E. Kyprianou, A. Murillo-Salas, and J. L. Pérez. An application of the backbone decomposition to supercritical super-Brownian motion with a barrier. Journal of Applied Probability, 49(3):671–684, September 2012. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.

org/stable/41713797.

Kalpathy:2011:APL

[KMW11] Ravi Kalpathy, Hosam M. Mahmoud, and Mark Daniel Ward. Asymptotic properties of a leader election algorithm. Journal of Applied Probability, 48(2):569–575, June 2011. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/23065903.

Keiding:1975:BPV

[KN75] Niels Keiding and John E. Nielsen. Branching processes with varying and random geometric offspring distributions. Journal of Applied Probability, 12(1):135–141, March 1975. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3212415.

Kannappan:1980:FEM

[KN80] PL. Kannappan and C. T. Ng. On functional equations and measures of information. II. Journal of Applied Probability, 17 (1):271–277, March 1980. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3212946.

Kemp:1990:CSD

[KN90] A. W. Kemp and J. Newton. Certain state-dependent processes for dichotomised parasite populations. Journal of Applied Probability, 27(2):251–258, June 1990. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3214644.

Krone:2000:SMR

[KN00] Stephen M. Krone and Claudia Neuhauser. A spatial model of range-dependent succession. Journal of Applied Probability, 37 (4):1044–1060, December 2000. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.

org/stable/3215493.

Kim:2001:OED

[KN01] Hyung Geun Kim and Dougu Nam. Optimal estimation of diffusion processes hidden by general obstacles. Journal of Applied Probability, 38(4):1067–1073, December 2001. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3215785.

Kiessler:2005:LRS

[KN05] Peter C. Kiessler and Kanoktip Nimitkiatklai. The lifetime of a random set. Journal of Applied Probability, 42(2):566–580, June 2005. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/30040810.

Knight:1971:SCP

[Kni71] Frank B. Knight. Some condensation processes of McKean type. Journal of Applied Probability, 8(2):399–406, June 1971. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/3211910.

Koole:2005:EST

[KNR05] Ger Koole, Misja Nuyens, and Rhonda Righter. The effect of service time variability on maximum queue lengths in M^{X}/G/1 queues. Journal of Applied Probability, 42(3):883–891, September 2005. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/30040866.

Katz:1973:MAR

[KNS73] S. Katz, P. Naor, and R. Shinnar. Moments of age and remaining life in renewal processes. Journal of Applied Probability, 10 (2):307–316, June 1973. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3212348.

Konstantinides:2010:PAR

[KNT10] Dimitrios G. Konstantinides, Kai W. Ng, and Qihe Tang. The probabilities of absolute ruin in the renewal risk model with constant force of interest. Journal of Applied Probability, 47 (2):323–334, June 2010. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/20778482.

Karasu:1989:NNP

[KÖ89] Izi Karasu and Süleyman Özekici. NBUE and NWUE properties of increasing Markov processes. Journal of Applied Probability, 26(4):827–834, December 1989. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.

org/stable/3214387.

Kolesnik:2005:PRM

[KO05] Alexander D. Kolesnik and Enzo Orsingher. A planar random motion with an infinite number of directions controlled by the damped wave equation. Journal of Applied Probability, 42(4): 1168–1182, December 2005. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.

org/stable/30040836.

Kyprianou:2012:SNL

[KO12] Andreas E. Kyprianou and Curdin Ott. Spectrally negative Lévy processes perturbed by functionals of their running supremum. Journal of Applied Probability, 49(4):1005–1014, December 2012. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/41713825.

Koc:1995:RCS

[Koç95] Cemal Koç. Recurring-with-carry sequences. Journal of Applied Probability, 32(4):966–971, December 1995. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/3215210.

Kim:2023:OSU

[KOH23] Mun-Chol Kim, Hun O, and Ho-Jin Hwang. Optimal stopping under g-expectation with L exp(μ\sqrt{2 log(1 + L)})- integrable reward process. Journal of Applied Probability, 60(1):241–252, March 2023. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL https:/

/www.cambridge.org/core/journals/journal-of-applied-
probability/article/optimal-stopping-under-gexpectation-
with-lexpbiglmusqrt2log1textbflbigr-integrable-reward-
process/35AF56961773EE1DE08FC02984DBC8DA.

Kollerstrom:1974:HTT

[Köl74] Julian Köllerström. Heavy traffic theory for queues with several servers. I. Journal of Applied Probability, 11(3):544–552, September 1974. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3212698.

Kollerstrom:1979:HTT

[Köl79] Julian Köllerström. Heavy traffic theory for queues with several servers. II. Journal of Applied Probability, 16(2):393–401, June 1979. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3212906.

Kolonko:1982:SCE

[Kol82] Michael Kolonko. Strongly consistent estimation in a controlled Markov renewal model. Journal of Applied Probability, 19(3): 532–545, September 1982. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3213512.

Kolonko:1995:PMM

[Kol95] M. Kolonko. A piecewise Markovian model for simulated annealing with stochastic cooling schedules. Journal of Applied Probability, 32(3):649–658, September 1995. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/3215119.

Kolesnik:1998:EMR

[Kol98] Alexander Kolesnik. The equations of Markovian random evolution on the line. Journal of Applied Probability, 35(1):27– 35, March 1998. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3215543.

Kolassa:2000:EBG

[Kol00] John E. Kolassa. Explicit bounds for geometric convergence of Markov chains. Journal of Applied Probability, 37(3):642–651, September 2000. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3215602.

Kolesnik:2006:FDR

[Kol06] Alexander D. Kolesnik. A four-dimensional random motion at finite speed. Journal of Applied Probability, 43(4):1107–1118, December 2006. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
27595803.

Kolesnik:2007:NPR

[Kol07] Alexander D. Kolesnik. A note on planar random motion at finite speed. Journal of Applied Probability, 44(3):838–842, September 2007. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
27595886.

Kolokoltsov:2015:SMD

[Kol15] Vassili N. Kolokoltsov. Stochastic monotonicity and duality of k-th order with application to put–call symmetry of powered options. Journal of Applied Probability, 52(1):82–101, March 2015. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://projecteuclid.org/euclid.jap/
1429282608.

Konstantopoulos:1996:LPS

[Kon96] Takis Konstantopoulos. A local proof of the Swiss Army formula of Palm calculus. Journal of Applied Probability, 33(3):909–914, September 1996. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3215368.

Koole:2004:LEF

[Koo04] Ger Koole. Letter to the Editor: A formula for tail probabilities of Cox distributions. Journal of Applied Probability, 41(3): 935–938, September 2004. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/4141364.

Kotz:1966:RRI

[Kot66] Samuel Kotz. Recent results in information theory. Journal of Applied Probability, 3(1):1–93, June 1966. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/3212039.

Kotler:1994:DVS

[Kot94] Mitchell Kotler. Drift vectors are not sufficient to determine recurrence of a Markov chain on Z_{+}^{3}. Journal of Applied Probability, 31(4):1098–1102, December 1994. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/3215332.

Koutras:1996:MCA

[Kou96] M. V. Koutras. On a Markov chain approach for the study of reliability structures. Journal of Applied Probability, 33(2):357– 367, June 1996. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3215059.

Kulkarni:1983:TPS

[KP83] M. V. Kulkarni and A. G. Pakes. The total progeny of a simple branching process with state-dependent immigration. Journal of Applied Probability, 20(3):472–481, September 1983. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3213885.

Kuo:1987:RLT

[KP87] Chia-Chyuan Kuo and S. Leigh Phoenix. Recursions and limit theorems for the strength and lifetime distributions of a fibrous composite. Journal of Applied Probability, 24(1):137–159, March 1987. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3214066.

Kaspi:1989:DBN

[KP89] Raya Kaspi and David Perry. On a duality between a non-Markovian storage/production process and a Markovian dam process with state-dependent input and output. Journal of Applied Probability, 26(4):835–844, December 1989. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3214388.

Katzenbeisser:1992:NTW

[KP92] W. Katzenbeisser and W. Panny. On the number of times where a simple random walk reaches its maximum. Journal of Applied Probability, 29(2):305–312, June 1992. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3214568.

Konstant:1993:EVC

[KP93] D. G. Konstant and V. I. Piterbarg. Extreme values of the cyclostationary Gaussian random process. Journal of Applied Probability, 30(1):82–97, March 1993. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3214623.

Katzenbeisser:1996:SRW

[KP96] W. Katzenbeisser and W. Panny. Simple random walk statistics. Part I: Discrete time results. Journal of Applied Probability, 33 (2):311–330, June 1996. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3215056.

Kyprianou:2007:DSF

[KP07] A. E. Kyprianou and Z. Palmowski. Distributional study of de Finetti’s dividend problem for a general Lévy insurance risk process. Journal of Applied Probability, 44(2):428–443, June 2007. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/27595852.

Kyprianou:2008:CSB

[KP08] A. E. Kyprianou and J. C. Pardo. Continuous-state branching processes and self-similarity. Journal of Applied Probability, 45 (4):1140–1160, December 2008. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.

org/stable/27596014.

Kabanov:2022:RPI

[KP22] Yuri Kabanov and Nikita Pukhlyakov. Ruin probabilities with investments: smoothness, inegro-differential and ordinary differential equations, asymptotic behavior. Journal of Applied Probability, 59(2):556–570, June 2022. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL https:/

/www.cambridge.org/core/journals/journal-of-applied-
probability/article/ruin-probabilities-with-investments-
smoothness-inegrodifferential-and-ordinary-differential-
equations-asymptotic-behavior/39392ED200F1CDFB0FDAF05479349691.

Kerner:2023:BEC

[KP23] Joachim Kerner and Maximilian Pechmann. Bose–Einstein condensation for particles with repulsive short-range pair interactions in a Poisson random external potential in R^{d}. Journal of Applied Probability, 60(2):382–393, June 2023. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL https://www.cambridge.org/core/journals/
journal-of-applied-probability/article/boseeinstein-
condensation-for-particles-with-repulsive-shortrange-
pair-interactions-in-a-poisson-random-external-potential-
in-mathbbrd/23F5CC18E910D51740206EFFE3181030.

Koutras:1995:RC

[KPP95] M. V. Koutras, G. K. Papadopoulos, and S. G. Papastavridis. Runs on a circle. Journal of Applied Probability, 32(2):396– 404, June 1995. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3215295.

Kimeldorf:1981:SCP

[KPS81] George Kimeldorf, Detlef Plachky, and Allan R. Sampson. A simultaneous characterization of the Poisson and Bernoulli distributions. Journal of Applied Probability, 18(1):316–320, March 1981. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3213195.

Kelly:2004:EDO

[KPS04] Leah Kelly, Eckhard Platen, and Michael Sørensen. Estimation for discretely observed diffusions using transform functions. Journal of Applied Probability, 41A:99–118, 2004. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3215971.

Krieger:2008:BMS

[KPSC08] Abba M. Krieger, Moshe Pollak, and Ester Samuel-Cahn. Beat the mean: Sequential selection by better than average rules. Journal of Applied Probability, 45(1):244–259, March 2008. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/27595937.

Kloeden:1996:EDE

[KPSS96] P. E. Kloeden, E. Platen, H. Schurz, and M. Sørensen. On effects of discretization on estimators of drift parameters for diffusion processes. Journal of Applied Probability, 33(4):1061–1076, December 1996. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3214986.

Koehl:1999:HDT

[KPT99] Pierre-F. Koehl, Huyên Pham, and Nizar Touzi. Hedging in discrete time under transaction costs and continuous-time limit. Journal of Applied Probability, 36(1):163–178, March 1999. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/3215412.

Kyprianou:2021:DHL

[KPV21] Andreas E. Kyprianou, Juan Carlos Pardo, and Matija Vidmar. Double hypergeometric Lévy processes and self-similarity. Journal of Applied Probability, 58(1):254–273, March 2021. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL https://www.cambridge.org/core/journals/journal-
of-applied-probability/article/double-hypergeometric-
levy-processes-and-selfsimilarity/6736BBD7373CA59D5F405FC1BCC4B038.

Konstantopoulos:1994:FAT

[KPW94] Takis Konstantopoulos, Spyros N. Papadakis, and Jean Walrand. Functional approximation theorems for controlled renewal processes. Journal of Applied Probability, 31(3):765–776, September 1994. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3215154.

Keilson:1970:PCD

[KR70] Julian Keilson and S. Subba Rao. A process with chain dependent growth rate. Journal of Applied Probability, 7(3):699–711, December 1970. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3211947.

Kan:1980:OLO

[KR80] Y. C. Kan and S. M. Ross. Optimal list order under partial memory constraints. Journal of Applied Probability, 17(4):1004–1015, December 1980. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3213210.

Kalashnikov:1986:CIP

[KR86] V. V. Kalashnikov and S. T. Rachev. Characterizations of inverse problems in queueing and their stability. Journal of Applied Probability, 23(2):459–473, June 1986. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/3214187.

Keilson:1988:BDT

[KR88] Julian Keilson and Ravi Ramaswamy. The backlog and depletion-time process for M/G/1 vacation models with exhaustive service discipline. Journal of Applied Probability, 25(2):404– 412, June 1988. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3214448.

Kaufmann:1992:PAI

[KR92] E. Kaufmann and R.-D. Reiss. Poisson approximation of intermediate empirical processes. Journal of Applied Probability, 29 (4):825–837, December 1992. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.

org/stable/3214715.

Kratz:1997:RCE

[KR97] Marie F. Kratz and Holger Rootzén. On the rate of convergence for extremes of mean square differentiable stationary normal processes. Journal of Applied Probability, 34(4):908–923, December 1997. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3215006.

Kuhne:2000:OSD

[KR00] Robert Kühne and Ludger Rüschendorf. Optimal stopping with discount and observation costs. Journal of Applied Probability, 37(1):64–72, March 2000. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3215659.

Kluppelberg:2008:PCA

[KR08] Claudia Klüppelberg and Sidney I. Resnick. The Pareto copula, aggregation of risks, and the Emperor’s socks. Journal of Applied Probability, 45(1):67–84, March 2008. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/27595924.

Kim:2018:CLT

[KR18] Steven S. Kim and Kavita Ramanan. A conditional limit theorem for high-dimensional ℓ^{p}-spheres. Journal of Applied Probability, 55(4):1060–1077, December 2018. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL https:/

/www.cambridge.org/core/journals/journal-of-applied-
probability/article/conditional-limit-theorem-for-highdimensional-
spheres/8BDD234B7DE7B7222F159AB6EC653C2F.

Kopfer:2021:MCM

[KR21] Benedikt Köpfer and Ludger Rüschendorf. The martingale comparison method for Markov processes. Journal of Applied Probability, 58(1):164–176, March 2021. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL https:/

/www.cambridge.org/core/journals/journal-of-applied-
probability/article/martingale-comparison-method-for-
markov-processes/79B71BFA3658EFDCB5A1964F20C961BD.

Kraft:1964:CDF

[Kra64] Charles H. Kraft. A class of distribution function processes which have derivatives. Journal of Applied Probability, 1(2): 385–388, December 1964. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/ stable/3211867. See correction [Ano65c].

Kroese:1997:HTA

[Kro97] Dirk P. Kroese. Heavy traffic analysis for continuous polling models. Journal of Applied Probability, 34(3):720–732, September 1997. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3215097.

Kruk:2004:LDM

[Kru04] Lukasz Kruk. Limiting distributions for minimum relative entropy calibration. Journal of Applied Probability, 41(1):35–50, March 2004. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3215813.

Kryscio:1972:TPE

[Kry72] Richard J. Kryscio. The transition probabilities of the extended simple stochastic epidemic model and the haskey model. Journal of Applied Probability, 9(3):471–485, September 1972. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3212320.

Kryscio:1975:TPG

[Kry75] Richard J. Kryscio. The transition probabilities of the general stochastic epidemic model. Journal of Applied Probability, 12(3): 415–424, September 1975. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3212856.

Kamien:1971:EPR

[KS71] Morton I. Kamien and Nancy L. Schwartz. Expenditure patterns for risky R and D projects. Journal of Applied Probability, 8(1): 60–73, March 1971. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3211838.

Kambo:1972:BAV

[KS72] N. S. Kambo and Samar Singh. Best asymptotic value of the error exponent with random coding on discrete memoryless channels. Journal of Applied Probability, 9(2):327–336, June 1972. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/3212802.

Kryscio:1976:NCC

[KS76] Richard J. Kryscio and Roy Saunders. A note on the cost of carrier-borne, right-shift, epidemic models. Journal of Applied Probability, 13(4):652–661, December 1976. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/3212520.

Karni:1977:TTO

[KS77] Edi Karni and Aba Schwartz. Two theorems on optimal stopping with backward solicitation. Journal of Applied Probability, 14(4): 869–875, December 1977. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3213361.

Katzoff:1978:DLR

[KS78] M. J. Katzoff and R. H. Shumway. Distributed lag regression with an almost periodic design matrix. Journal of Applied Probability, 15(4):759–773, December 1978. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3213432.

Konig:1980:INI

[KS80a] D. König and V. Schmidt. Imbedded and non-imbedded stationary characteristics of queueing systems with varying service rate and point processes. Journal of Applied Probability, 17(3): 753–767, September 1980. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3212969.

Konig:1980:SIB

[KS80b] D. König and V. Schmidt. Stochastic inequalities between customer-stationary and time-stationary characteristics of queueing systems with point processes. Journal of Applied Probability, 17(3):768–777, September 1980. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/3212970.

Klebaner:1982:CBL

[KS82] F. C. Klebaner and H.-J. Schuh. A connection between the limit and the maximum random variable of a branching process in varying environments. Journal of Applied Probability, 19(3): 681–684, September 1982. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3213527.

Konig:1983:PTG

[KS83a] Dieter König and Volker Schmidt. The Palm-type grain size distribution in stationary grain models. Journal of Applied Probability, 20(3):574–584, September 1983. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3213894.

Kryscio:1983:IDP

[KS83b] Richard J. Kryscio and Roy Saunders. On interpoint distances for planar Poisson cluster processes. Journal of Applied Probability, 20(3):513–528, September 1983. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3213888.

Konig:1985:SQL

[KS85] Dieter König and Volker Schmidt. Stationary queue-length characteristics in queues with delayed feedback. Journal of Applied Probability, 22(2):394–407, June 1985. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3213782.

Kammerle:1986:MCG

[KS86a] K. Kämmerle and H.-J. Schuh. The maximum in critical Galton– Watson and birth and death processes. Journal of Applied Probability, 23(3):601–613, September 1986. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http:/

/www.jstor.org/stable/3214000.

Keilson:1986:ORW

[KS86b] J. Keilson and L. D. Servi. Oscillating random walk models for GI/G/1 vacation systems with Bernoulli schedules. Journal of Applied Probability, 23(3):790–802, September 1986. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3214016.

Kijima:1986:UGR

[KS86c] Masaaki Kijima and Ushio Sumita. A useful generalization of renewal theory: Counting processes governed by non-negative Markovian increments. Journal of Applied Probability, 23(1):71– 88, March 1986. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3214117.

Kounias:1989:BBR

[KS89] Stratis Kounias and Kiki Sotirakoglou. Bonferroni bounds revisited. Journal of Applied Probability, 26(2):233–241, June 1989. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/3214031.

Kijima:1991:SRQ

[KS91a] Masaaki Kijima and E. Seneta. Some results for quasi-stationary distributions of birth-death processes. Journal of Applied Probability, 28(3):503–511, September 1991. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3214486.

Kwiecinski:1991:CCS

[KS91b] A. Kwieciński and R. Szekli. Compensator conditions for stochastic ordering of point processes. Journal of Applied Probability, 28(4):751–761, December 1991. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3214678.

Kao:1993:RPR

[KS93a] Edward P. C. Kao and Marion Spokony Smith. On renewal processes relating to counter models: The case of phase-type interarrival times. Journal of Applied Probability, 30(1):175– 183, March 1993. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3214630.

Krafft:1993:MPT

[KS93b] Olaf Krafft and Martin Schaefer. Mean passage times for tridiagonal transition matrices and a two-parameter Ehrenfest urn model. Journal of Applied Probability, 30(4):964–970, December 1993. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3214525.

Keilson:1994:NNH

[KS94a] J. Keilson and L. D. Servi. Networks of non-homogeneous M/G/∞. Journal of Applied Probability, 31A:157–168, 1994. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/3214954.

Kella:1994:CMF

[KS94b] Offer Kella and Michail Sverchkov. On concavity of the mean function and stochastic ordering for reflected processes with stationary increments. Journal of Applied Probability, 31(4):1140– 1142, December 1994. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3215339.

Kang:1999:EVP

[KS99] Sungyeol Kang and Richard F. Serfozo. Extreme values of phasetype and mixed random variables with parallel-processing examples. Journal of Applied Probability, 36(1):194–210, March 1999. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3215414.

Kagan:2001:EUC

[KS01a] Y. Y. Kagan and F. Schoenberg. Estimation of the upper cutoff parameter for the tapered Pareto distribution. Journal of Applied Probability, 38A:158–175, 2001. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3215876.

Kella:2001:HTC

[KS01b] Offer Kella and Wolfgang Stadje. On hitting times for compound Poisson dams with exponential jumps and linear release rate. Journal of Applied Probability, 38(3):781–786, September 2001. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3216131.

Kulik:2001:SCL

[KS01c] Rafal Kulik and Ryszard Szekli. Sufficient conditions for longrange count dependence of stationary point processes on the real line. Journal of Applied Probability, 38(2):570–581, June 2001. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3215907.

Kella:2002:MML

[KS02a] Offer Kella and Wolfgang Stadje. Markov-modulated linear fluid networks with Markov additive input. Journal of Applied Probability, 39(2):413–420, June 2002. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3216108.

Klebaner:2002:AGW

[KS02b] F. C. Klebaner and S. Sagitov. The age of a Galton–Watson population with a geometric offspring distribution. Journal of Applied Probability, 39(4):816–828, December 2002. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3216005.

Kella:2004:BMT

[KS04] Offer Kella and Wolfgang Stadje. A Brownian motion with two reflecting barriers and Markov-modulated speed. Journal of Applied Probability, 41(4):1237–1242, December 2004. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/4141398.

Katsikopoulos:2005:ODS

[KS05] Konstantinos V. Katsikopoulos and Özgür Simsek. Optimal doubling strategy against a suboptimal opponent. Journal of Applied Probability, 42(3):867–872, September 2005. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/30040863.

Kawazu:2006:LTD

[KS06] Kiyoshi Kawazu and Yuki Suzuki. Limit theorems for a diffusion process with a one-sided Brownian potential. Journal of Applied Probability, 43(4):997–1012, December 2006. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/27595795.

Kella:2008:CPR

[KS08] Offer Kella and Wolfgang Stadje. A Collector’s problem with renewal arrival processes. Journal of Applied Probability, 45(3): 610–620, September 2008. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/27595974.

Khandwawala:2010:OMF

[KS10] Mustafa Khandwawala and Rajesh Sundaresan. Optimal multicommodity flow through the complete graph with random edge capacities. Journal of Applied Probability, 47(1):201–215, March 2010. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/25662510.

Kella:2013:AEN

[KS13] Offer Kella and Wolfgang Stadje. Asymptotic expected number of passages of a random walk through an interval. Journal of Applied Probability, 50(1):288–294, March 2013. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://projecteuclid.org/euclid.jap/1363784439.

Kranakis:2016:SAP

[KS16] Evangelos Kranakis and Gennady Shaikhet. Sensor allocation problems on the real line. Journal of Applied Probability, 53(3): 667–687, September 2016. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://projecteuclid.

org/euclid.jap/1476370769.

Kuba:2017:MTS

[KS17] Markus Kuba and Henning Sulzbach. On martingale tail sums in affine two-color urn models with multiple drawings. Journal of Applied Probability, 54(1):96–117, March 2017. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL https://www.cambridge.org/core/journals/journal-
of-applied-probability/article/on-martingale-tail-sums-
in-affine-twocolor-urn-models-with-multiple-drawings/
6CFB304B4E4F9B442F3BA8740B02BB18.

Kuck:2018:ARP

[KS18] Fabian Kück and Dominic Schuhmacher. On the age of a randomly picked individual in a linear birth-and-death process. Journal of Applied Probability, 55(1):82–93, March 2018. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL https://www.cambridge.org/core/journals/journal-
of-applied-probability/article/on-the-age-of-a-randomly-
picked-individual-in-a-linear-birthanddeath-process/
FF48B30CF5ED5FBF41610CA3307D313F.

Kamihigashi:2019:UST

[KS19] Takashi Kamihigashi and John Stachurski. A unified stability theory for classical and monotone Markov chains. Journal of Applied Probability, 56(1):1–22, March 2019. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL https:/

/www.cambridge.org/core/journals/journal-of-applied-
probability/article/unified-stability-theory-for-classical-
and-monotone-markov-chains/8364B8375BA698F3BB42A09244113B95.

Kluppelberg:2020:ERC

[KS20] Claudia Klüppelberg and Miriam Isabel Seifert. Explicit results on conditional distributions of generalized exponential mixtures. Journal of Applied Probability, 57(3): 760–774, September 2020. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL https:/

/www.cambridge.org/core/journals/journal-of-applied-
probability/article/explicit-results-on-conditional-
distributions-of-generalized-exponential-mixtures/446410589851DF8BEB2F9488BFC9913A.

Kulik:2021:LRD

[KS21] Rafal Kulik and Evgeny Spodarev. Long range dependence of heavy-tailed random functions. Journal of Applied Probability, 58(3):569–593, September 2021. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL https:/

/www.cambridge.org/core/journals/journal-of-applied-
probability/article/long-range-dependence-of-heavytailed-
random-functions/33659B9A3F05CBF8123A02B08C359C7F.

Kamatani:2023:NRG

[KS23a] Kengo Kamatani and Xiaolin Song. Non-reversible guided Metropolis kernel. Journal of Applied Probability, 60(3): 955–981, September 2023. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL https:/

/www.cambridge.org/core/journals/journal-of-applied-
probability/article/nonreversible-guided-metropolis-
kernel/C93570C9C1D8D1BFD898C13C7DF26541.

Kaur:2023:IUF

[KS23b] Gursharn Kaur and Neeraja Sahasrabudhe. Interacting urns on a finite directed graph. Journal of Applied Probability, 60(1):166–188, March 2023. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL https:/

/www.cambridge.org/core/journals/journal-of-applied-

probability/article/interacting-urns-on-a-finite-directed-
graph/570F048902BA79492E26A7A6FCBD9A31.

Kapodistria:2023:WAT

[KSBK23] Stella Kapodistria, Mayank Saxena, Onno J. Boxma, and Offer Kella. Workload analysis of a two-queue fluid polling model. Journal of Applied Probability, 60(3):1003–1030, September 2023. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-

6072 (electronic). URL https://www.cambridge.org/core/
journals/journal-of-applied-probability/article/workload-
analysis-of-a-twoqueue-fluid-polling-model/4B10FCBB38F16194FB36324D47C0A877.

Koops:2018:ISQ

[KSBM18] D. T. Koops, M. Saxena, O. J. Boxma, and M. Mandjes. Infiniteserver queues with Hawkes input. Journal of Applied Probability, 55(3):920–943, September 2018. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL https:/

/www.cambridge.org/core/journals/journal-of-applied-
probability/article/infiniteserver-queues-with-hawkes-
input/29A985198A710AB1F7636DCEAEA9C0E6.

Krieger:2012:NSP

[KSC12] Abba M. Krieger and Ester Samuel-Cahn. The noisy secretary problem and some results on extreme concomitant variables. Journal of Applied Probability, 49(3):821–837, September 2012. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/41713808.

Krieger:2013:GBF

[KSC13] Abba M. Krieger and Ester Samuel-Cahn. Generalized bomber and fighter problems: offline optimal allocation of a discrete asset. Journal of Applied Probability, 50(2):403–418, June 2013. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://projecteuclid.org/euclid.jap/
1371648949.

Kryscio:1980:NAB

[KSF80] Richard J. Kryscio, Roy Saunders, and Gerald M. Funk. Normal approximations for binary lattice systems. Journal of Applied Probability, 17(3):674–685, September 1980. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/3212961.

Kurano:1998:CMS

[KSHH98] Masami Kurano, Jinjie Song, Masanori Hosaka, and Youqiang Huang. Controlled Markov set-chains with discounting. Journal of Applied Probability, 35(2):293–302, June 1998. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3215686.

Ksir:1989:SSM

[Ksi89] Brahim Ksir. Shepp statistic for Markov chains application to a long-run average cost criterion. Journal of Applied Probability, 26(4):767–775, December 1989. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.

org/stable/3214381.

Kaplan:1975:BPD

[KSN75] Norman Kaplan, Aidan Sudbury, and Trygve S. Nilsen. A branching process with disasters. Journal of Applied Probability, 12(1):47–59, March 1975. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3212406.

Kozakai:2016:CSC

[KSN16] Ryouta Kozakai, Akinobu Shimizu, and Morihiro Notohara. Convergence to the structured coalescent process. Journal of Applied Probability, 53(2):502–517, June 2016. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

projecteuclid.org/euclid.jap/1466172870.

Katehakis:2016:CAS

[KSS16] Michael N. Katehakis, Laurens C. Smit, and Floske M. Spieksma. A comparative analysis of the successive lumping and the lattice path counting algorithms. Journal of Applied Probability, 53(1):106–120, March 2016. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http:/

/projecteuclid.org/euclid.jap/1457470562.

Koole:1999:MRT

[KST99] Ger Koole, Panayotis D. Sparaggis, and Don Towsley. Minimizing response times and queue lengths in systems of parallel queues. Journal of Applied Probability, 36(4):1185–1193, December 1999. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3215587.

Karlin:1982:LBD

[KT82] Samuel Karlin and Simon Tavaré. Linear birth and death processes with killing. Journal of Applied Probability, 19(3):477– 487, September 1982. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3213507.

Kimeldorf:1983:JCM

[KT83] George Kimeldorf and Peter F. Thall. A joint characterization of the multinomial distribution and the Poisson process. Journal of Applied Probability, 20(1):202–208, March 1983. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3213737.

Kipouridis:2001:SOSb

[KT01a] I. Kipouridis and G. Tsaklidis. The size order of the state vector of a continuous-time homogeneous Markov system with fixed size. Journal of Applied Probability, 38(3):635–646, September 2001. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3216117.

Kipouridis:2001:SOSa

[KT01b] I. Kipouridis and G. Tsaklidis. The size order of the state vector of discrete-time homogeneous Markov systems. Journal of Applied Probability, 38(2):357–368, June 2001. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/3215893.

Ko:2008:SDN

[KT08] Bangwon Ko and Qihe Tang. Sums of dependent nonnegative random variables with subexponential tails. Journal of Applied Probability, 45(1):85–94, March 2008. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/27595925.

Kawanishi:2015:NVW

[KT15] Ken’ichi Kawanishi and Tetsuya Takine. A note on the virtual waiting time in the stationary PH/M/c+D queue. Journal of Applied Probability, 52(3):899–903, September 2015. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://projecteuclid.org/euclid.jap/1445543855.

Kotiah:1969:UED

[KTW69] T. C. T. Kotiah, J. W. Thompson, and W. A. O’N. Waugh. Use of Erlangian distributions for single-server queueing systems. Journal of Applied Probability, 6(3):584–593, December 1969. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3212104.

Kubota:2024:DBL

[Kub24] Naoki Kubota. Differences between Lyapunov exponents for the simple random walk in Bernoulli potentials. Journal of Applied Probability, 61(1):82–103, March 2024. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL https:/

/www.cambridge.org/core/journals/journal-of-applied-
probability/article/differences-between-lyapunov-exponents-
for-the-simple-random-walk-in-bernoulli-potentials/123D61898709348DF747F4430046ADD9.

Kuczek:1982:ASL

[Kuc82] Thomas Kuczek. Almost sure limit results for the supercritical Bellman–Harris process. Journal of Applied Probability, 19(3): 668–674, September 1982. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3213525.

Kuhn:1985:RSH

[Kuh85] Jonathan Kuhn. The risk-sensitive homing problem. Journal of Applied Probability, 22(4):796–803, December 1985. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3213947.

Kulkarni:1983:QSR

[Kul83] Vidyadhar G. Kulkarni. On queueing systems with retrials. Journal of Applied Probability, 20(2):380–389, June 1983. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/3213810.

Kulkarni:1986:EWT

[Kul86] V. G. Kulkarni. Expected waiting times in a multiclass batch arrival retrial queue. Journal of Applied Probability, 23(1):144– 154, March 1986. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3214123.

Kulperger:1987:SRR

[Kul87] R. J. Kulperger. Some remarks on regression with autoregressive errors and their residual processes. Journal of Applied Probability, 24(3):668–678, September 1987. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3214098.

Kulik:2006:LTM

[Kul06] Rafal Kulik. Limit theorems for moving averages with random coefficients and heavy-tailed noise. Journal of Applied Probability, 43(1):245–256, March 2006. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/27595724.

Kunsch:1986:DBM

[Kün86] H. Künsch. Discrimination between monotonic trends and longrange dependence. Journal of Applied Probability, 23(4):1025– 1030, December 1986. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3214476.

Kunwai:2022:FLC

[Kun22] Khwanchai Kunwai. On Foster–Lyapunov criteria for exponential ergodicity of regime-switching jump diffusion processes with countable regimes. Journal of Applied Probability, 59(1):167–186, March 2022. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL https:/

/www.cambridge.org/core/journals/journal-of-applied-
probability/article/on-fosterlyapunov-criteria-for-exponential-
ergodicity-of-regimeswitching-jump-diffusion-processes-
with-countable-regimes/F82012D07BC8A4D3410355778AA30ED1.

Kupka:1990:DMN

[Kup90] Joseph Kupka. The distribution and moments of the number of components of a random function. Journal of Applied Probability, 27(1):202–207, March 1990. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3214607.

Kurtz:1970:SOD

[Kur70] Thomas G. Kurtz. Solutions of ordinary differential equations as limits of pure jump Markov processes. Journal of Applied Probability, 7(1):49–58, April 1970. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3212147.

Kurtz:1971:LTS

[Kur71] T. G. Kurtz. Limit theorems for sequences of jump Markov processes approximating ordinary differential processes. Journal of Applied Probability, 8(2):344–356, June 1971. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3211904.

Kurano:1987:LAM

[Kur87] Masami Kurano. Learning algorithms for Markov decision processes. Journal of Applied Probability, 24(1):270–276, March 1987. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3214080.

Kurauskas:2022:LWL

[Kur22] Valentas Kurauskas. On local weak limit and subgraph counts for sparse random graphs. Journal of Applied Probability, 59(3):755–776, September 2022. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL https:/

/www.cambridge.org/core/journals/journal-of-applied-
probability/article/on-local-weak-limit-and-subgraph-
counts-for-sparse-random-graphs/65812C212EBC471C2AE8CF4178DC78D1.

Kuulasmaa:1982:SGE

[Kuu82] Kari Kuulasmaa. The spatial general epidemic and locally dependent random graphs. Journal of Applied Probability, 19(4): 745–758, December 1982. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3213827.

Kuznetsov:2010:WHF

[Kuz10] A. Kuznetsov. Wiener–Hopf factorization for a family of Lévy processes related to theta functions. Journal of Applied Probability, 47(4):1023–1033, December 2010. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/29777423.

Keilson:1978:OBM

[KW78] Julian Keilson and Jon A. Wellner. Oscillating Brownian motion. Journal of Applied Probability, 15(2):300–310, June 1978.

CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/3213403.

Kumar:1985:IOR

[KW85] P. R. Kumar and J. Walrand. Individually optimal routing in parallel systems. Journal of Applied Probability, 22(4):989–995, December 1985. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3213970.

Konstantopoulos:1989:SSF

[KW89] Panagiotis Konstantopoulos and Jean Walrand. Stationary and stability of fork-join networks. Journal of Applied Probability, 26 (3):604–614, September 1989. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.

org/stable/3214417.

Kella:1992:UMS

[KW92] Offer Kella and Ward Whitt. Useful martingales for stochastic storage processes with Lévy input. Journal of Applied Probability, 29(2):396–403, June 1992. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3214576.

Kella:1996:SSP

[KW96] Offer Kella and Ward Whitt. Stability and structural properties of stochastic storage networks. Journal of Applied Probability, 33(4):1169–1180, December 1996. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3214994.

Kella:1999:LSF

[KW99] Offer Kella and Ward Whitt. Linear stochastic fluid networks. Journal of Applied Probability, 36(1):244–260, March 1999. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/3215418.

Kirmani:2003:TSB

[KW03] S. N. U. A. Kirmani and Jacek Wesolowski. Time spent below a random threshold by a Poisson driven sequence of observations. Journal of Applied Probability, 40(3):807–814, September 2003. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3215955.

Kugler:2013:UBM

[KW13] Johannes Kugler and Vitali Wachtel. Upper bounds for the maximum of a random walk with negative drift. Journal of Applied Probability, 50(4):1131–1146, December 2013. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://projecteuclid.org/euclid.jap/1389370104.

Kontorovich:2014:UCD

[KW14] Aryeh Kontorovich and Roi Weiss. Uniform Chernoff and Dvoretzky–Kiefer-Wolfowitz-type inequalities for Markov chains and related processes. Journal of Applied Probability, 51(4): 1100–1113, December 2014. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://projecteuclid.

org/euclid.jap/1421763330.

Kwerel:1975:MSB

[Kwe75] Seymour M. Kwerel. Most stringent bounds on the probability of the union and intersection of m events for systems partially specified by S_{1}, S_{2}, ⋯ S_{k}, 2 ≤ k < m. Journal of Applied Probability, 12(3):612–619, September 1975. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/3212879.

Kochar:2009:CPS

[KX09] Subhash Kochar and Maochao Xu. Comparisons of parallel systems according to the convex transform order. Journal of Applied Probability, 46(2):342–352, June 2009. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/25662428.

Kochar:2011:SOS

[KX11] Subhash Kochar and Maochao Xu. On the skewness of order statistics in multiple-outlier models. Journal of Applied Probability, 48(1):271–284, March 2011. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/29777459.

Kella:2017:UDL

[KY17a] Offer Kella and Marc Yor. Unifying the Dynkin and Lebesgue–Stieltjes formulae. Journal of Applied Probability, 54(1):252–266, March 2017. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL https:/

/www.cambridge.org/core/journals/journal-of-applied-
probability/article/unifying-the-dynkin-and-lebesguestieltjes-
formulae/137D06E4E013E2AF7422A8544CB916B5.

Kenyon:2017:ACE

[KY17b] Richard Kenyon and Mei Yin. On the asymptotics of constrained exponential random graphs. Journal of Applied Probability, 54(1):165–180, March 2017. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL https:/

/www.cambridge.org/core/journals/journal-of-applied-
probability/article/on-the-asymptotics-of-constrained-
exponential-random-graphs/FAC12C6E4E71C4C0C7A35B3851BE4B7D.

Konstantopoulos:2017:PIG

[KY17c] Takis Konstantopoulos and Linglong Yuan. A probabilistic interpretation of the Gaussian binomial coefficients. Journal of Applied Probability, 54(4):1295–1298, December 2017. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL https://www.cambridge.org/core/journals/journal-
of-applied-probability/article/probabilistic-interpretation-
of-the-gaussian-binomial-coefficients/17E0DA0F32D4A1C25EC20BFED616DA99.

Kyprianou:1971:QSD

[Kyp71] E. K. Kyprianou. On the quasi-stationary distribution of the virtual waiting time in queues with Poisson arrivals. Journal of Applied Probability, 8(3):494–507, September 1971. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3212173.

Kyprianou:1972:QSDa

[Kyp72a] E. K. Kyprianou. On the quasi-stationary distributions of the GI/M/1 queue. Journal of Applied Probability, 9(1):117–128, March 1972. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3212641.

Kyprianou:1972:QSDb

[Kyp72b] E. K. Kyprianou. The quasi-stationary distributions of queues in heavy traffic. Journal of Applied Probability, 9(4):821–831, December 1972. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3212618.

Kyprianou:1998:SVU

[Kyp98] A. E. Kyprianou. Slow variation and uniqueness of solutions to the functional equation in the branching random walk. Journal of Applied Probability, 35(4):795–801, December 1998. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3215516.

Kyriakidis:1999:OCT

[Kyr99] E. G. Kyriakidis. Optimal control of a truncated general immigration process through total catastrophes. Journal of Applied Probability, 36(2):461–472, June 1999. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/3215468.

Kuulasmaa:1984:SGS

[KZ84] Kari Kuulasmaa and Stan Zachary. On spatial general spidemics and bond percolation processes. Journal of Applied Probability, 21(4):911–914, December 1984. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.

org/stable/3213706.

Keilson:1988:HRC

[KZ88] J. Keilson and M. Zachmann. Homogeneous row-continuous bivariate Markov chains with boundaries. Journal of Applied Probability, 25A:237–256, 1988. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3214160.

Konstantopoulos:1995:DTP

[KZ95] Takis Konstantopoulos and Michael Zazanis. A discrete-time proof of Neveu’s exchange formula. Journal of Applied Probability, 32(4):917–921, December 1995. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3215204.

Kyprianou:2009:GTS

[KZ09] Andreas E. Kyprianou and Xiaowen Zhou. General tax structures and the Lévy insurance risk model. Journal of Applied Probability, 46(4):1146–1156, December 2009. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/25662486.

Kella:2005:STD

[KZB05] Offer Kella, Bert Zwart, and Onno Boxma. Some timedependent properties of symmetric M/G/1 queues. Journal of Applied Probability, 42(1):223–234, March 2005. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/30040781.

Kaminska-Zabierowska:2015:EMR

[KZN15] Maria Kamińska-Zabierowska and Jorge Navarro. Erratum to: Mixture representations of residual lifetimes of used systems. Journal of Applied Probability, 52(4):1183–1186, December 2015. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-

6072 (electronic). URL http://projecteuclid.org/euclid.
jap/1450802761. See [NBS08].

Lee:1998:PPD

[LA98] Eui Yong Lee and Sung K. Ahn. P_{λ}^{M}-policy for a dam with input formed by a compound Poisson process. Journal of Applied Probability, 35(2):482–488, June 1998. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3215701.

Litvak:2001:TTC

[LA01] Nelly Litvak and Ivo Adan. The travel time in carousel systems under the nearest item heuristic. Journal of Applied Probability, 38(1):45–54, March 2001. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3215740.

Lachal:1993:LDM

[Lac93] Aimé Lachal. L’intégrale du mouvement Brownien. (French) [The integral of Brownian motion]. Journal of Applied Probability, 30(1):17–27, March 1993. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3214618.

Lachaud:2005:CHT

[Lac05] B. Lachaud. Cut-off and hitting times of a sample of Ornstein– Uhlenbeck processes and its average. Journal of Applied Probability, 42(4):1069–1080, December 2005. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/30040829.

Ladret:2005:AHT

[Lad05] Véronique Ladret. Asymptotic hitting time for a simple evolutionary model of protein folding. Journal of Applied Probability, 42(1):39–51, March 2005. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/30040768.

Lewis:2002:BOA

[LAF02] Mark E. Lewis, Hayriye Ayhan, and Robert D. Foley. Bias optimal admission control policies for a multiclass nonstationary queueing system. Journal of Applied Probability, 39(1):20– 37, March 2002. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3215915.

Lageraas:2005:RPT

[Lag05] Andreas Nordvall Lagerås. A renewal-process-type expression for the moments of inverse subordinators. Journal of Applied Probability, 42(4):1134–1144, December 2005. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/30040834.

Lai:1977:TDI

[Lai77] C. D. Lai. A two-dimensional ‘immigration-branching’ model with application to earthquake occurrence times and energies. Journal of Applied Probability, 14(3):464–474, September 1977. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/3213449.

Lai:1978:EWP

[Lai78] C. D. Lai. An example of Wold’s point processes with Markovdependent intervals. Journal of Applied Probability, 15(4):748– 758, December 1978. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3213431.

Lai:1987:LE

[Lai87] C. D. Lai. Letter to the Editor. Journal of Applied Probability, 24(1):288–289, March 1987. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3214083.

Lalley:1986:FPP

[Lal86] S. Lalley. A first-passage problem for a two-dimensional controlled random walk. Journal of Applied Probability, 23(3): 670–678, September 1986. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3214005.

Lampard:1968:SPW

[Lam68] D. G. Lampard. A stochastic process whose successive intervals between events form a first order Markov chain: I. Journal of Applied Probability, 5(3):648–668, December 1968. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3211928.

Lamperti:1970:MBP

[Lam70] John Lamperti. Maximal branching processes and ‘long-range percolation’. Journal of Applied Probability, 7(1):89–98, April 1970. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3212151.

Lam:1984:CSO

[Lam84] K. Lam. Comparison of self-organizing linear search rules. Journal of Applied Probability, 21(4):763–776, December 1984. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/3213694.

Lam:1991:JSB

[Lam91] C. Y. Teresa Lam. Joint simulation of backward and forward recurrence times in a superposition of independent renewal processes. Journal of Applied Probability, 28(4):930–933, December 1991. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3214699.

Lam:1992:NBT

[Lam92] C. Y. Teresa Lam. New better than used in expectation processes. Journal of Applied Probability, 29(1):116–128, March 1992. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3214796.

Lam:2014:QCL

[Lam14] Hoang-Chuong Lam. A quenched central limit theorem for reversible random walks in a random environment on Z. Journal of Applied Probability, 51(4):1051–1064, December 2014. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://projecteuclid.org/euclid.jap/1421763327.

Lange:1979:CSG

[Lan79a] Kenneth Lange. On Cohen’s stochastic generalization of the strong ergodic theorem of demography. Journal of Applied Probability, 16(3):496–504, September 1979. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http:/

/www.jstor.org/stable/3213079.

Lanska:1979:MCE

[Lán79b] Vera Lánska. Minimum contrast estimation in diffusion processes. Journal of Applied Probability, 16(1):65–75, March 1979. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/3213375.

Langberg:1981:NCC

[Lan81] Naftali A. Langberg. On the normal convergence of a class of simple batch epidemics. Journal of Applied Probability, 18(1): 31–41, March 1981. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3213164.

Lane:1984:CLT

[Lan84] John A. Lane. The central limit theorem for the Poisson shot-noise process. Journal of Applied Probability, 21(2):287– 301, June 1984. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3213640.

Langaris:1986:CQI

[Lan86a] C. Langaris. A correlated queue with infinitely many servers. Journal of Applied Probability, 23(1):155–165, March 1986. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/3214124. See [Ano87h].

Langaris:1986:WTP

[Lan86b] C. Langaris. The waiting-time process of a queueing system with gamma-type input and blocking. Journal of Applied Probability, 23(1):166–174, March 1986. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3214125.

Langaris:1987:BPA

[Lan87] Christos Langaris. Busy-period analysis of a correlated queue with exponential demand and service. Journal of Applied Probability, 24(2):476–485, June 1987. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3214271.

Langberg:1988:CRP

[Lan88] Naftali A. Langberg. Comparisons of replacement policies. Journal of Applied Probability, 25(4):780–788, December 1988. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/3214299.

Lanchier:2005:MCP

[Lan05] Nicolas Lanchier. A multitype contact process with frozen sites: A spatial model of allelopathy. Journal of Applied Probability, 42(4):1109–1119, December 2005. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/30040832.

Lanchier:2011:CPD

[Lan11] N. Lanchier. Contact process with destruction of cubes and hyperplanes: Forest fires versus tornadoes. Journal of Applied Probability, 48(2):352–365, June 2011. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/23065890.

Larget:1998:CRA

[Lar98] Bret Larget. A canonical representation for aggregated Markov processes. Journal of Applied Probability, 35(2):313–324, June 1998. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3215688.

Laslett:1978:MCP

[Las78] G. M. Laslett. Mixing of cluster point processes. Journal of Applied Probability, 15(4):715–725, December 1978. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3213429.

Lasserre:1994:NPI

[Las94a] J. B. Lasserre. A new policy iteration scheme for Markov decision processes using Schweitzer’s formula. Journal of Applied Probability, 31(1):268–273, March 1994. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http:/

/www.jstor.org/stable/3215254.

Lasserre:1994:DON

[Las94b] Jean B. Lasserre. Detecting optimal and non-optimal actions in average-cost Markov decision processes. Journal of Applied Probability, 31(4):979–990, December 1994. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/3215322.

Lasserre:1994:FSP

[Las94c] Jean B. Lasserre. A formula for singular perturbations of Markov chains. Journal of Applied Probability, 31(3):829–833, September 1994. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3215160.

Latouche:1981:MQW

[Lat81a] Guy Latouche. On a Markovian queue with weakly correlated interarrival times. Journal of Applied Probability, 18(1):190– 203, March 1981. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3213179.

Latouche:1981:QPC

[Lat81b] Guy Latouche. Queues with paired customers. Journal of Applied Probability, 18(3):684–696, September 1981. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3213322.

Laue:1980:RRB

[Lau80] G. Laue. Remarks on the relation between fractional moments and fractional derivatives of characteristic functions. Journal of Applied Probability, 17(2):456–466, June 1980. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3213035.

Lawrance:1970:SIP

[Law70a] A. J. Lawrance. Selective interaction of a Poisson and renewal process: First-order stationary point results. Journal of Applied Probability, 7(2):359–372, August 1970. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http:/

/www.jstor.org/stable/3211970.

Lawrance:1970:SIS

[Law70b] A. J. Lawrance. Selective interaction of a stationary point process and a renewal process. Journal of Applied Probability, 7 (2):483–489, August 1970. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3211987.

Lawrance:1971:SIPa

[Law71a] A. J. Lawrance. Selective interaction of a Poisson and renewal process: The dependency structure of the intervals between responses. Journal of Applied Probability, 8(1):170–183, March 1971. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3211846.

Lawrance:1971:SIPb

[Law71b] A. J. Lawrance. Selective interaction of a Poisson and renewal process: The spectrum of the intervals between responses. Journal of Applied Probability, 8(4):731–744, December 1971. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/3212237.

Lawrance:1979:RVF

[Law79] A. J. Lawrance. The rate and variance functions arising from the interaction of two stationary point processes. Journal of Applied Probability, 16(2):287–296, June 1979. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3212897.

Lawrance:1980:MES

[Law80] A. J. Lawrance. The mixed exponential solution to the firstorder autoregressive model. Journal of Applied Probability, 17 (2):546–552, June 1980. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3213044.

Lawler:1983:CCL

[Law83] Gregory F. Lawler. A connective constant for loop-erased selfavoiding random walk. Journal of Applied Probability, 20(2): 264–276, June 1983. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3213800.

Lawrance:1992:UDF

[Law92] A. J. Lawrance. Uniformly distributed first-order autoregressive time series models and multiplicative congruential random number generators. Journal of Applied Probability, 29(4):896–903, December 1992. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3214722.

Liu:1988:GBT

[LB88] Jian Liu and Peter J. Brockwell. On the general bilinear time series model. Journal of Applied Probability, 25(3):553–564, September 1988. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3213984.

Lal:1989:SER

[LB89] Ram Lal and U. Narayan Bhat. Some explicit results for correlated random walks. Journal of Applied Probability, 26(4): 757–766, December 1989. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3214380.

Liang:2023:MCA

[LB23] Vincent Liang and Konstantin Borovkov. On Markov chain approximations for computing boundary crossing probabilities of diffusion processes. Journal of Applied Probability, 60 (4):1386–1415, December 2023. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL https:/

/www.cambridge.org/core/journals/journal-of-applied-
probability/article/on-markov-chain-approximations-for-
computing-boundary-crossing-probabilities-of-diffusion-
processes/27900CF3863D0951CC69AEDDCF4142BC.

Lange:1981:MCS

[LBC81] Kenneth Lange, Michael Boehnke, and Richard Carson. Moment computations for subcritical branching processes. Journal of Applied Probability, 18(1):52–64, March 1981. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/3213166.

Lim:2005:OUP

[LBL05] Kyung Eun Lim, Jee Seon Baek, and Eui Yong Lee. Optimization under the P_{λ,τ}^{M} policy of a finite dam with both continuous and jumpwise inputs. Journal of Applied Probability, 42(2):587– 594, June 2005. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
30040812.

Lund:1999:PSN

[LBP99] Robert B. Lund, Ronald W. Butler, and Robert L. Paige. Prediction of shot noise. Journal of Applied Probability, 36(2):374– 388, June 1999. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3215462.

Lee:1983:NCP

[LC83] Hau Leung Lee and Morris A. Cohen. A note on the convexity of performance measures of M/M/c queueing systems. Journal of Applied Probability, 20(4):920–923, December 1983. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3213606.

Langaris:1984:WTT

[LC84] Christos Langaris and Brian Conolly. On the waiting time of a two-stage queueing system with blocking. Journal of Applied Probability, 21(3):628–638, September 1984. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/3213623.

Lee:2016:PPA

[LC16] Hyunju Lee and Ji Hwan Cha. Point process approach to modeling and analysis of general cascading failure models. Journal of Applied Probability, 53(1):174–186, March 2016. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://projecteuclid.org/euclid.jap/1457470567.

Li:2010:OAA

[LD10] Xiaohu Li and Weiyong Ding. Optimal allocation of active redundancies to k-out-of-n systems with heterogeneous components. Journal of Applied Probability, 47(1):254–263, March 2010. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/25662513.

Lee:1995:CRG

[LDL95] Shiowjen Lee, S. Durham, and J. Lynch. On the calculation of the reliability of general load sharing systems. Journal of Applied Probability, 32(3):777–792, September 1995. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/3215129.

Le:1994:BMS

[Le94] Huiling Le. Brownian motions on shape and size-and-shape spaces. Journal of Applied Probability, 31(1):101–113, March 1994. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3215238.

Le:2014:CTB

[Le14] V. Le. Coalescence times for the Bienaymé-Galton–Watson process. Journal of Applied Probability, 51(1):209–218, March 2014. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://projecteuclid.org/euclid.jap/
1395771424.

Ledoux:1997:GIW

[Led97] James Ledoux. A geometric invariant in weak lumpability of finite Markov chains. Journal of Applied Probability, 34(4): 847–858, December 1997. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3215001.

Ledoux:2005:RFP

[Led05] James Ledoux. Recursive filters for partially observable finite Markov chains. Journal of Applied Probability, 42(3):684–697, September 2005. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
30040849.

Lee:1980:CST

[Lee80] P. A. Lee. Correlation structure of teletraffic measurements from randomly-scanned calls arrival data. Journal of Applied Probability, 17(1):203–217, March 1980. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3212937.

Lee:1997:ATQ

[Lee97] Duan-Shin Lee. Analysis of a two-queue model with Bernoulli schedules. Journal of Applied Probability, 34(1):176–191, March 1997. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3215185.

Lee:2011:GDI

[Lee11a] Chihoon Lee. A geometric drift inequality for a reflected fractional Brownian motion process on the positive orthant. Journal of Applied Probability, 48(3):820–831, September 2011. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/23065939.

Lee:2011:RTR

[Lee11b] Chihoon Lee. On the return time for a reflected fractional Brownian motion process on the positive orthant. Journal of Applied Probability, 48(1):145–153, March 2011. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http:/

/www.jstor.org/stable/29777451.

Lee:2020:NSM

[Lee20] Hyunju Lee. On a new stochastic model for cascading failures. Journal of Applied Probability, 57(4):1150–1161, December 2020. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL https://www.cambridge.org/core/journals/
journal-of-applied-probability/article/on-a-new-stochastic-
model-for-cascading-failures/43DB9F67075164FDB41E78A92D62A8C0.

Lefevre:1978:EUS

[Lef78] Claude Lefèvre. The expected ultimate size of a carrier-borne epidemic. Journal of Applied Probability, 15(2):414–419, June 1978. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3213412.

Lefevre:1979:AUT

[Lef79] Claude Lefèvre. The area under the trajectory of carriers in a carrier-borne epidemic. Journal of Applied Probability, 16 (2):231–241, June 1979. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3212892.

Lefebvre:1989:COD

[Lef89] Mario Lefebvre. Commande optimale du processus de Wiener. (French) [Optimal control of Wiener processes]. Journal of Applied Probability, 26(1):50–57, March 1989. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/3214315.

Lefebvre:1995:PPP

[Lef95] Mario Lefebvre. Problèmes de premier passage résolubles par la méthode de séparation des variables. (French) [First-passage problems solvable by the method of separation of variables]. Journal of Applied Probability, 32(2):337–348, June 1995. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/3215292.

Lefebvre:1997:IFH

[Lef97] Mario Lefebvre. On the inverse of the first hitting time problem for bidimensional processes. Journal of Applied Probability, 34 (3):610–622, September 1997. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.

org/stable/3215088.

Lefebvre:2006:FPP

[Lef06] Mario Lefebvre. First passage problems for asymmetric Wiener processes. Journal of Applied Probability, 43(1):175–184, March 2006. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/27595718.

Lefebvre:2024:FPP

[Lef24] Mario Lefebvre. A first-passage-place problem for integrated diffusion processes. Journal of Applied Probability, 61(1):55–67, March 2024. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL https:/

/www.cambridge.org/core/journals/journal-of-applied-
probability/article/firstpassageplace-problem-for-integrated-
diffusion-processes/441BE227317E9B3468070DA996FC62EC.

Lehoczky:1971:NFE

[Leh71] J. P. Lehoczky. A note on the first emptiness time of an infinite reservoir with inputs forming a Markov chain. Journal of Applied Probability, 8(2):276–284, June 1971. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3211898.

Lehoczky:1972:TIC

[Leh72] J. P. Lehoczky. Traffic intersection control and zero-switch queues under conditions of Markov chain dependence input. Journal of Applied Probability, 9(2):382–395, June 1972. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/3212806.

Lehoczky:1973:FEP

[Leh73] J. P. Lehoczky. First-emptiness problems in applied probability under first-order dependent input and general output conditions. Journal of Applied Probability, 10(2):330–342, June 1973. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/3212350.

Lehtonen:1986:OTQ

[Leh86] Tapani Lehtonen. On the ordering of tandem queues with exponential servers. Journal of Applied Probability, 23(1):115– 129, March 1986. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3214121.

Lehtonen:1988:SJE

[Leh88] Tapani Lehtonen. Scheduling jobs with exponential processing times on parallel machines. Journal of Applied Probability, 25(4): 752–762, December 1988. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3214296.

Lehec:2014:CTG

[Leh14] Joseph Lehec. Cover times and generic chaining. Journal of Applied Probability, 51(1):247–261, March 2014. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://projecteuclid.org/euclid.jap/1395771427.

Lehtomaa:2017:LDM

[Leh17] Jaakko Lehtomaa. Large deviations of means of heavy-tailed random variables with finite moments of all orders. Journal of Applied Probability, 54(1):66–81, March 2017. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL https://www.cambridge.org/core/journals/journal-
of-applied-probability/article/large-deviations-of-means-
of-heavytailed-random-variables-with-finite-moments-
of-all-orders/C6F4D413F3CA51A03AC5619436322B59.

Lehtomaa:2021:ETI

[Leh21] Jaakko Lehtomaa. Estimating tails of independently stopped random walks using concave approximations of hazard functions. Journal of Applied Probability, 58(3):773–793, September 2021.

CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL https://www.cambridge.org/core/journals/
journal-of-applied-probability/article/estimating-tails-
of-independently-stopped-random-walks-using-concave-
approximations-of-hazard-functions/DEFB848507E8D894657A934B261F3E21.

Lei:2023:EMG

[Lei23] Xiaoyu Lei. An efficient method for generating a discrete uniform distribution using a biased random source. Journal of Applied Probability, 60(3):1069–1078, September 2023. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL https://www.cambridge.org/core/journals/
journal-of-applied-probability/article/an-efficient-
method-for-generating-a-discrete-uniform-distribution-
using-a-biased-random-source/B88F3E117BF2D8AA9707D9F96750E380.

Lelarge:2005:FLG

[Lel05] Marc Lelarge. Fluid limit of generalized Jackson queueing networks with stationary and ergodic arrivals and service times. Journal of Applied Probability, 42(2):491–512, June 2005. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/30040806.

Lelarge:2007:TAM

[Lel07] Marc Lelarge. Tail asymptotics for monotone-separable networks. Journal of Applied Probability, 44(2):306–320, June 2007. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/27595843.

Lemoine:1974:TSD

[Lem74] Austin J. Lemoine. On two stationary distributions for the stable GI/G/1 queue. Journal of Applied Probability, 11(4):849–852, December 1974. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3212570.

Lemoine:1981:QPP

[Lem81] Austin J. Lemoine. On queues with periodic Poisson input. Journal of Applied Probability, 18(4):889–900, December 1981. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/3213063.

Lemoine:1987:STJ

[Lem87] Austin J. Lemoine. On sojourn time in Jackson networks of queues. Journal of Applied Probability, 24(2):495–510, June 1987. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3214273.

Lemoine:1989:WTW

[Lem89] Austin J. Lemoine. Waiting time and workload in queues with periodic Poisson input. Journal of Applied Probability, 26(2): 390–397, June 1989. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3214044.

Lenz:1981:EPR

[Len81] Norbert Lenz. Extinction probability, regularity and asymptotic growth of Markovian populations. Journal of Applied Probability, 18(1):1–18, March 1981. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.

org/stable/3213162.

Leslie:1967:RCE

[Les67] R. T. Leslie. Recurrent composite events. Journal of Applied Probability, 4(1):34–61, April 1967. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3212298.

Leslie:1969:RTC

[Les69] R. T. Leslie. Recurrence times of clusters of Poisson points. Journal of Applied Probability, 6(2):372–388, August 1969. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/3212007.

Leslie:1973:SLT

[Les73] J. R. Leslie. Some limit theorems for Markov branching processes. Journal of Applied Probability, 10(2):299–306, June 1973. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/3212347.

Lesanovsky:1986:CTR

[Les86] Antonín Lesanovský. Comparison of two replacement policies. Journal of Applied Probability, 23(3):759–769, September 1986.

CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/3214013.

Leslie:1989:NCU

[Les89] J. R. Leslie. On the non-closure under convolution of the subexponential family. Journal of Applied Probability, 26(1):58–66, March 1989. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3214316.

Leskela:2006:SOQ

[Les06] Lasse Leskelä. Stabilization of an overloaded queueing network using measurement-based admission control. Journal of Applied Probability, 43(1):231–244, March 2006. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http:/

/www.jstor.org/stable/27595723.

Lessard:2010:REP

[Les10] Sabin Lessard. Recurrence equations for the probability distribution of sample configurations in exact population genetics models. Journal of Applied Probability, 47(3):732–751, September 2010. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/20778513.

Letac:1974:TRI

[Let74] Gérard Letac. Transience and recurrence of an interesting Markov chain. Journal of Applied Probability, 11(4):818–824, December 1974. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3212565.

Letac:1978:CCT

[Let78] Gérard Letac. Chaînes colorées: Trois extensions d’une formule de P. Nelson. (French) [Colored chains: three extensions of a formula of P. Nelson]. Journal of Applied Probability, 15(2):321– 339, June 1978. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3213405.

Letac:1994:FES

[Let94] Gérard Letac. Les familles exponentielles statistiques invariantes par les groupes du Cône et du paraboloïde de révolution. (French) [exponential statistical invariant families groups of Cône and the paraboloid of revolution]. Journal of Applied Probability, 31A:71–95, 1994. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3214950.

Letac:2019:SDP

[Let19] Gérard Letac. Is the Sibuya distribution a progeny? Journal of Applied Probability, 56(1):52–56, March 2019. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL https://www.cambridge.org/core/journals/journal-
of-applied-probability/article/is-the-sibuya-distribution-
a-progeny/38DD4534D45699DEFE5ABBD0EAF88B18.

Levikson:1977:ADM

[Lev77] Benny Levikson. The age distribution of Markov processes. Journal of Applied Probability, 14(3):492–506, September 1977. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/3213452.

Lewis:1964:IFM

[Lew64] P. A. W. Lewis. Implications of a failure model for the use and maintenance of computers. Journal of Applied Probability, 1(2): 347–368, December 1964. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3211865.

Lewis:1967:FRD

[Lew67] T. Lewis. The factorisation of the rectangular distribution. Journal of Applied Probability, 4(3):529–542, December 1967. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/3212219.

Lewis:1969:APE

[Lew69] P. A. W. Lewis. Asymptotic properties and equilibrium conditions for branching Poisson processes. Journal of Applied Probability, 6(2):355–371, August 1969. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3212006.

Lewis:2001:AOP

[Lew01] Mark E. Lewis. Average optimal policies in a controlled queueing system with dual admission control. Journal of Applied Probability, 38(2):369–385, June 2001. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3215894.

Leite:2008:DAS

[LF08] Saul C. Leite and Marcelo D. Fragoso. Diffusion approximation of state-dependent G-networks under heavy traffic. Journal of Applied Probability, 45(2):347–362, June 2008. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/27595949.

Lehoczky:1981:DAC

[LG81] J. P. Lehoczky and D. P. Gaver. Diffusion approximations for the cooperative service of Voice and data Messages. Journal of Applied Probability, 18(3):660–671, September 1981. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3213320.

Lathrop:2016:NGE

[LGC16] Eve D. Lathrop, Isaac H. Goldstein, and Yung-Pin Chen. A note on a generalized Ehrenfest urn model: another look at the mean transition times. Journal of Applied Probability, 53(2):630–632, June 2016. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-

6072 (electronic). URL http://projecteuclid.org/euclid.
jap/1466172882.

Launay:2020:ESD

[LGD20] Claire Launay, Bruno Galerne, and Agnès Desolneux. Exact sampling of determinantal point processes without eigendecomposition. Journal of Applied Probability, 57(4):1198–1221, December 2020. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-

6072 (electronic). URL https://www.cambridge.org/core/
journals/journal-of-applied-probability/article/exact-
sampling-of-determinantal-point-processes-without-eigendecomposition/
5033B1D3138145CBBDFC6FA7B8239832.

Lange:1981:SSP

[LH81] Kenneth Lange and William Holmes. Stochastic stable population growth. Journal of Applied Probability, 18(2):325–334, June 1981. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3213280.

Lopez-Herrero:2002:DNC

[LH02] M. J. Lopez-Herrero. Distribution of the number of customers served in an M/G/1 retrial queue. Journal of Applied Probability, 39(2):407–412, June 2002. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3216107.

Liu:2006:STE

[LH06] Yuanyuan Liu and Zhenting Hou. Several types of ergodicity for M/G/1-type Markov chains and Markov processes. Journal of Applied Probability, 43(1):141–158, March 2006. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/27595716.

Lindell:2007:DLE

[LH07] Andreas Lindell and Lars Holst. Distributions of the longest excursions in a tied down simple random walk and in a Brownian bridge. Journal of Applied Probability, 44(4):1056–1067, December 2007. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/27595907.

Lefevre:1986:FLB

[LHN86] Claude Lefevre, Marc Hallin, and Prakash Narayan. On fractional linear bounds for probability generating functions. Journal of Applied Probability, 23(4):904–913, December 1986. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/3214464.

Luo:2017:ASD

[LHT17] Chunling Luo, Zhou He, and Chin Hon Tan. Almost stochastic dominance under inconsistent utility and loss functions. Journal of Applied Probability, 54(3):763–772, September 2017. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL https://www.cambridge.org/core/journals/journal-
of-applied-probability/article/almost-stochastic-dominance-
under-inconsistent-utility-and-loss-functions/533436F00220275B58F77D01D5EFE1B3.

Li:1996:FLF

[Li96] Jingwen Li. From fifo to lifo: A functional ordering of service delay via arrival discipline. Journal of Applied Probability, 33 (2):507–512, June 1996. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3215074.

Li:2000:SMD

[Li00a] Haijun Li. Stochastic models for dependent life lengths induced by common pure jump shock environments. Journal of Applied Probability, 37(2):453–469, June 2000. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3215720.

Li:2000:OUT

[Li00b] Zeng-Hu Li. Ornstein–Uhlenbeck type processes and branching processes with immigration. Journal of Applied Probability, 37 (3):627–634, September 2000. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.

org/stable/3215600.

Li:2004:SPA

[Li04] Xiaohu Li. Some properties of ageing notions based on the moment-generating-function order. Journal of Applied Probability, 41(3):927–934, September 2004. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/4141363.

Li:2005:ABH

[Li05] Yulin Li. Asymptotic baseline of the hazard rate function of mixtures. Journal of Applied Probability, 42(3):892–901, September 2005. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/30040867.

Li:2006:CSP

[Li06a] Yuqiang Li. On a continuous-state population-size-dependent branching process and its extinction. Journal of Applied Probability, 43(1):195–207, March 2006. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/27595720.

Li:2006:LTD

[Li06b] Zenghu Li. A limit theorem for discrete Galton–Watson branching processes with immigration. Journal of Applied Probability, 43(1):289–295, March 2006. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/27595728.

Li:2009:WLT

[Li09] Yuqiang Li. A weak limit theorem for generalized Jiřina processes. Journal of Applied Probability, 46(2):453–462, June 2009. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/25662436.

Li:2012:MDS

[Li12] Yuqiang Li. Moderate deviations for stable random walks in random scenery. Journal of Applied Probability, 49(1):280–294, March 2012. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
41713765.

Liberman:1985:OSC

[Lib85] Uri Liberman. An order statistic characterization of the Poisson renewal process. Journal of Applied Probability, 22(3):717–722, September 1985. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3213875.

Liebetrau:1977:WCC

[Lie77] A. M. Liebetrau. On the weak convergence of a class of estimators of the variance-time curve of a weakly stationary point process. Journal of Applied Probability, 14(1):114–126, March 1977. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3213264.

Liebetrau:1978:WCC

[Lie78] A. M. Liebetrau. The weak convergence of a class of estimators of the variance function of a two-dimensional Poisson process. Journal of Applied Probability, 15(2):433–439, June 1978. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/3213415.

Lillestol:1968:AAS

[Lil68] Jostein Lillestøl. Another approach to some Markov chain models in population genetics. Journal of Applied Probability, 5(1): 9–20, April 1968. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3212073.

Lindvall:1972:CCG

[Lin72] Torgny Lindvall. Convergence of critical Galton–Watson branching processes. Journal of Applied Probability, 9(2):445–450, June 1972. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3212815.

Lindvall:1973:WCP

[Lin73] Torgny Lindvall. Weak convergence of probability measures and random functions in the function space D[0, ∞). Journal of Applied Probability, 10(1):109–121, March 1973. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3212499.

Lindvall:1979:NCB

[Lin79] Torgny Lindvall. A note on coupling of birth and death processes. Journal of Applied Probability, 16(3):505–512, September 1979. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3213080.

Lindvall:1982:CCT

[Lin82] Torgny Lindvall. On coupling of continuous-time renewal processes. Journal of Applied Probability, 19(1):82–89, March 1982. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/3213918.

Lindvall:1983:CDP

[Lin83] Torgny Lindvall. On coupling of diffusion processes. Journal of Applied Probability, 20(1):82–93, March 1983. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3213722.

Lindqvist:1987:MAM

[Lin87] Bo Henry Lindqvist. Monotone and associated Markov chains, with applications to reliability theory. Journal of Applied Probability, 24(3):679–695, September 1987. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3214099.

Lindgren:1989:SMP

[Lin89] Georg Lindgren. Slepian models for χ^{2}-processes with dependent components with application to envelope upcrossings. Journal of Applied Probability, 26(1):36–49, March 1989. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3214314.

Lin:1994:NDH

[Lin94] Gwo Dong Lin. A note on distributions having the almostlack-of-memory property. Journal of Applied Probability, 31(3): 854–856, September 1994. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3215164.

Lin:1995:WCW

[Lin95] Gwo Dong Lin. On weak convergence within a Lorenz ordering family of distributions. Journal of Applied Probability, 32(4): 1028–1031, December 1995. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.

org/stable/3215216.

Lin:1996:TCD

[Lin96] Xiaodong Lin. Tail of compound distributions and excess time. Journal of Applied Probability, 33(1):184–195, March 1996. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/3215276.

Lin:1997:WCW

[Lin97] Gwo Dong Lin. On weak convergence within a family of life distributions. Journal of Applied Probability, 34(3):823–826, September 1997. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3215110.

Ling:1999:PPD

[Lin99] Shiqing Ling. On the probabilistic properties of a double threshold ARMA conditional heteroskedastic model. Journal of Applied Probability, 36(3):688–705, September 1999. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3215434.

Linetsky:2004:SRB

[Lin04] Vadim Linetsky. The spectral representation of Bessel processes with constant drift: Applications in queueing and finance. Journal of Applied Probability, 41(2):327–344, June 2004. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3216019.

Lin:2008:SBT

[Lin08] Jianxi Lin. Some Blackwell-type renewal theorems for weighted renewal functions. Journal of Applied Probability, 45(4):972–993, December 2008. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
27596003.

Lindberg:2014:PTO

[Lin14] Carl Lindberg. Pairs trading with opportunity cost. Journal of Applied Probability, 51(1):282–286, March 2014. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://projecteuclid.org/euclid.jap/1395771429.

Lin:2016:GMS

[Lin16] Yi-Shen Lin. A generalization of the Mabinogion sheep problem of D. Williams. Journal of Applied Probability, 53(4):1240– 1256, December 2016. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://projecteuclid.

org/euclid.jap/1481132849.

Lipow:1977:LDP

[Lip77] Carla Lipow. Limiting diffusions for population-size dependent branching processes. Journal of Applied Probability, 14(1):14– 24, March 1977. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3213257.

Litwin:1969:DRR

[Lit69] Samuel Litwin. The distribution of radioactive recovery in randomly cut and sedimented DNA. Journal of Applied Probability, 6(2):275–284, August 1969. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/ stable/3211998. See correction [Jof77] and remark [Jof77].

Litwin:1974:DSD

[Lit74] Samuel Litwin. The distribution on the shortest distance between random cuts on opposite strands of DNA. Journal of Applied Probability, 11(2):363–368, June 1974. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/3212757.

Litwin:1979:LFC

[Lit79] Samuel Litwin. The lengths of fragments covering a fixed marker in randomly cut linear or circular DNA. Journal of Applied Probability, 16(4):721–731, December 1979. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/3213139.

Littlejohn:1992:DMP

[Lit92a] R. P. Littlejohn. Discrete minification processes and reversibility. Journal of Applied Probability, 29(1):82–91, March 1992. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/3214793.

Littlejohn:1992:OWI

[Lit92b] R. P. Littlejohn. An operation which inverts Bernoulli multiplication and associated stationary reversible Markov processes. Journal of Applied Probability, 29(1):234–238, March 1992. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/3214810.

Littlejohn:1994:RRT

[Lit94] R. P. Littlejohn. A reversibility relationship for two Markovian time series models with stationary exponential tailed distribution. Journal of Applied Probability, 31(2):575–581, June 1994. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/3215050.

Litvak:2001:SPE

[Lit01] Nelly Litvak. Some peculiarities of exponential random variables. Journal of Applied Probability, 38(3):787–792, September 2001. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3216132.

Liu:1994:CMM

[Liu94] Huaigao Liu. On Cowan and Mecke’s Markov chain. Journal of Applied Probability, 31(2):554–560, June 1994. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3215046.

Liu:2004:DCH

[Liu04] Shuangzhe Liu. On diagnostics in conditionally heteroskedastic time series models under elliptical distributions. Journal of Applied Probability, 41A:393–405, 2004. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3215992.

Liu:2011:AFD

[Liu11] Yuanyuan Liu. Additive functionals for discrete-time Markov chains with applications to birth-death processes. Journal of Applied Probability, 48(4):925–937, December 2011. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/23066434.

Lange:1978:RPR

[LJ78] Kenneth Lange and Norman J. Johnson. Renewal processes with random numbers of delays: Application to a conception and birth model. Journal of Applied Probability, 15(2):209–224, June 1978. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3213395.

Lee:1977:ASA

[LK77] T. S. Lee and F. Kozin. Almost sure asymptotic likelihood theory for diffusion processes. Journal of Applied Probability, 14(3): 527–537, September 1977. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3213455.

Li:2001:SNR

[LK01] Xiaohu Li and Subhash C. Kochar. Some new results involving the NBU(2) class of life distributions. Journal of Applied Probability, 38(1):242–247, March 2001. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3215756.

Liu:1990:GGS

[LKT90] L. Liu, B. R. K. Kashyap, and J. G. C. Templeton. On the GI^{X} /G/∞ system. Journal of Applied Probability, 27(3):671–683, September 1990. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3214550.

Lawrance:1975:PBD

[LL75] A. J. Lawrance and P. A. W. Lewis. Properties of the bivariate delayed Poisson process. Journal of Applied Probability, 12 (2):257–268, June 1975. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3212439.

Lawrance:1977:EMA

[LL77] A. J. Lawrance and P. A. W. Lewis. An exponential movingaverage sequence and point process (EMA1). Journal of Applied Probability, 14(1):98–113, March 1977. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/3213263.

Latouche:1978:RTN

[LL78] G. Latouche and G. Louchard. Return times in nearlycompletely decomposable stochastic processes. Journal of Applied Probability, 15(2):251–267, June 1978. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/3213399.

Lacayo:1982:ESF

[LL82] H. Lacayo and Naftali A. Langberg. The exact and symptotic formulas for the state probabilities in simple epidemics with m kinds of susceptibles. Journal of Applied Probability, 19(1):1– 9, March 1982. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3213910.

Louchard:1990:GBI

[LL90] Guy Louchard and Guy Latouche. Geometric bounds on iterative approximations for nearly completely decomposable Markov chains. Journal of Applied Probability, 27(3):521–529, September 1990. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3214538.

Lee:1993:MSS

[LL93] Eui Yong Lee and Jiyeon Lee. A model for a system subject to random shocks. Journal of Applied Probability, 30(4):979–984, December 1993. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3214528.

Lefevre:2013:MMD

[LL13] Claude Lefèvre and Stéphane Loisel. On multiply monotone distributions, continuous or discrete, with applications. Journal of Applied Probability, 50(3):827–847, September 2013. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://projecteuclid.org/euclid.jap/1378401239.

Li:2024:RRC

[LL24] Chen Li and Xiaohu Li. On relevation redundancy to coherent systems at component and system levels. Journal of Applied Probability, 61(1):104–120, March 2024. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL https:/

/www.cambridge.org/core/journals/journal-of-applied-
probability/article/on-relevation-redundancy-to-coherent-
systems-at-component-and-system-levels/D4EE97D58679B4734B65BA4834F76DE3.

Liu:2020:GAT

[LLL20] Yuanyuan Liu, Wendi Li, and Xiuqin Li. On geometric and algebraic transience for block-structured Markov chains. Journal of Applied Probability, 57(4):1313–1338, December 2020. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL https://www.cambridge.org/core/journals/
journal-of-applied-probability/article/on-geometric-
and-algebraic-transience-for-blockstructured-markov-
chains/5197FFD2D587438930565E555406F3C7.

Lloyd:1971:NTD

[Llo71] Emlyn H. Lloyd. A note on the time-dependent and the stationary behaviour of a semi-infinite reservoir subject to a combination of Markovian inflows. Journal of Applied Probability, 8(4): 708–715, December 1971. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3212234.

Lloyd:1989:SFR

[Llo89] Christopher J. Lloyd. Some further results on random splitting of an interval. Journal of Applied Probability, 26(3):466–476, September 1989. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3214405.

Lee:1994:AMG

[LLPC94] Ho Woo Lee, Soon Seok Lee, Jeong Ok Park, and K. C. Chae. Analysis of the M^{X} /G/1 queue with N-policy and multiple vacations. Journal of Applied Probability, 31(2):476–496, June 1994. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3215040.

Lam:1984:SOF

[LLS84] Kin Lam, Ming-Ying Leung, and Man-Keung Siu. Selforganizing files with dependent accesses. Journal of Applied Probability, 21(2):343–359, June 1984. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3213645.

Lee:2022:ESE

[LLT^{+}22] Young Lee, Patrick J. Laub, Thomas Taimre, Hongbiao Zhao, and Jiancang Zhuang. Exact simulation of extrinsic stress-release processes. Journal of Applied Probability, 59(1):105–117, March 2022. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL https:/

/www.cambridge.org/core/journals/journal-of-applied-
probability/article/exact-simulation-of-extrinsic-stressrelease-
processes/B474D0F5423DFB09C13528F1877D82E0.

Landriault:2015:FDB

[LLZ15] David Landriault, Bin Li, and Hongzhong Zhang. On the frequency of drawdowns for Brownian motion processes. Journal of Applied Probability, 52(1):191–208, March 2015. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://projecteuclid.org/euclid.jap/1429282615.

Landriault:2017:UAD

[LLZ17] David Landriault, Bin Li, and Hongzhong Zhang. A unified approach for drawdown (drawup) of time-homogeneous Markov processes. Journal of Applied Probability, 54(2):603–626, June 2017. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-

6072 (electronic). URL https://www.cambridge.org/core/
journals/journal-of-applied-probability/article/unified-
approach-for-drawdown-drawup-of-timehomogeneous-markov-
processes/38780024016D4D432728645496538AAB.

Lefevre:1988:CCB

[LM88] Claude Lefevre and Marie-Pierre Malice. Comparisons for carrier-borne epidemics in heterogeneous and homogeneous populations. Journal of Applied Probability, 25(4):663–674, December 1988. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3214287.

Lefevre:1989:SCJ

[LM89] Claude Lefèvre and Marie-Pierre Malice. On a system of components with joint lifetimes distributed as a mixture of independent exponential laws. Journal of Applied Probability, 26 (1):202–208, March 1989. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/ stable/3214332. See correction [Ano90h].

Lefevre:1990:IDL

[LM90] Claude Lefèvre and György Michaletzky. Interparticle dependence in a linear death process subjected to a random environment. Journal of Applied Probability, 27(3):491–498, September 1990. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3214535.

Lewis:1991:MPT

[LM91] Peter A. W. Lewis and Ed McKenzie. Minification processes and their transformations. Journal of Applied Probability, 28(1):45– 57, March 1991. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3214739.

Likhanov:1999:CLA

[LM99] Nikolay Likhanov and Ravi R. Mazumdar. Cell loss asymptotics for buffers fed with a large number of independent stationary sources. Journal of Applied Probability, 36(1):86–96, March 1999. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3215404.

Lin:2004:LPE

[LM04a] Daming Lin and Viliam Makis. On-line parameter estimation for a failure-prone system subject to condition monitoring. Journal of Applied Probability, 41(1):211–220, March 2004. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3215826.

Ling:2004:OTB

[LM04b] Yibei Ling and Jie Mi. An optimal trade-off between content freshness and refresh cost. Journal of Applied Probability, 41(3): 721–734, September 2004. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/4141349.

Lewis:2008:WHF

[LM08] Alan L. Lewis and Ernesto Mordecki. Wiener-hope factorization for Lévy processes having positive jumps with rational transforms. Journal of Applied Probability, 45(1):118–134, March 2008. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/27595928.

Lamberton:2012:SFP

[LM12] Damien Lamberton and Mohammed Mikou. The smooth-fit property in an exponential Lévy model. Journal of Applied Probability, 49(1):137–149, March 2012. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/41713756.

Lambert:2015:CPM

[LM15] Amaury Lambert and Chunhua Ma. The coalescent in peripatric metapopulations. Journal of Applied Probability, 52(2):538–557, June 2015. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-

6072 (electronic). URL http://projecteuclid.org/euclid.
jap/1437658614.

Lyon:2020:TGU

[LM20] Merritt R. Lyon and Hosam M. Mahmoud. Trees grown under young-age preferential attachment. Journal of Applied Probability, 57(3):911–927, September 2020. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL https:/

/www.cambridge.org/core/journals/journal-of-applied-
probability/article/trees-grown-under-youngage-preferential-
attachment/558FA5AD65C8ED6F96AF9211840375EB.

Lifshits:2024:EMA

[LM24] Mikhail Lifshits and Gilles Mordant. Extrema of a multinomial assignment process. Journal of Applied Probability, 61(1):198–213, March 2024. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL https:/

/www.cambridge.org/core/journals/journal-of-applied-
probability/article/extrema-of-a-multinomial-assignment-
process/A77777174012E33BBD304DFFD1F31BEB.

Lageraas:2006:GSB

[LML06] Andreas Nordvall Lagerås and Anders Martin-Löf. Genealogy for supercritical branching processes. Journal of Applied Probability, 43(4):1066–1076, December 2006. CODEN JPRBAM.

ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/27595800.

Lebensztayn:2010:NRW

[LMM10] Elcio Lebensztayn, Fábio Prates Machado, and Mauricio Zuluaga Martinez. Nonhomogeneous random walks systems on Z. Journal of Applied Probability, 47(2):562–571, June 2010. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/20778497.

Laska:1972:CTO

[LMS72] E. Laska, M. Meisner, and C. Siegel. Contributions to the theory of optimal resource allocation. Journal of Applied Probability, 9(2):337–359, June 1972. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3212803.

Letac:1976:NLC

[LMS76] Gérard Letac, Pierre Mazet, and Gérard Schiffmann. A note on left-continuous random walks. Journal of Applied Probability, 13 (4):814–817, December 1976. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.

org/stable/3212539.

Lasmar:2018:MDM

[LMS18] Nabil Lasmar, Cécile Mailler, and Olfa Selmi. Multiple drawing multi-colour urns by stochastic approximation. Journal of Applied Probability, 55(1):254–281, March 2018. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL https://www.cambridge.org/core/journals/journal-
of-applied-probability/article/multiple-drawing-multicolour-
urns-by-stochastic-approximation/0BDE8B742FF5F03FC03C6AA48AC3D981.

Lopez-Mimbela:1998:LGW

[LMW98] J. Alfredo López-Mimbela and Anton Wakolbinger. Length of Galton–Watson trees and blow-up of semilinear systems. Journal of Applied Probability, 35(4):802–811, December 1998. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3215517.

Lopez-Mimbela:2000:PPN

[LMW00] J. Alfredo López-Mimbela and Anton Wakolbinger. A probabilistic proof of non-explosion of a non-linear PDE system. Journal of Applied Probability, 37(3):635–641, September 2000. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/3215601.

Lund:2004:LPP

[LMX04] Robert Lund, William P. McCormick, and Yuanhui Xiao. Limiting properties of Poisson shot noise processes. Journal of Applied Probability, 41(3):911–918, September 2004. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/4141361.

Lucantoni:1994:SPS

[LN94] David M. Lucantoni and Marcel F. Neuts. Simpler proofs of some properties of the fundamental period of the MAP/G/1 queue. Journal of Applied Probability, 31(1):235–243, March 1994. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3215249.

Lanchier:2007:VMB

[LN07] N. Lanchier and C. Neuhauser. Voter model and biased voter model in heterogeneous environments. Journal of Applied Probability, 44(3):770–787, September 2007. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/27595881.

Leckey:2013:AAH

[LN13] Kevin Leckey and Ralph Neininger. Asymptotic analysis of Hoppe trees. Journal of Applied Probability, 50(1):228–238, March 2013. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-

6072 (electronic). URL http://projecteuclid.org/euclid.
jap/1363784435.

Leskela:2019:FPP

[LN19] Lasse Leskelä and Hoa Ngo. First passage percolation on sparse random graphs with boundary weights. Journal of Applied Probability, 56(2):458–471, June 2019. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL https:/

/www.cambridge.org/core/journals/journal-of-applied-
probability/article/first-passage-percolation-on-sparse-
random-graphs-with-boundary-weights/AE47C7B3A499C3E42FC260F9EA5627AD.

Lifshits:2023:LDP

[LN23] Mikhail Lifshits and Sergei Nikitin. Large deviations of Poisson Telecom processes. Journal of Applied Probability, 60(1):267–283, March 2023. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL https:/

/www.cambridge.org/core/journals/journal-of-applied-
probability/article/large-deviations-of-poisson-telecom-
processes/E83892F798F975E1B79D096FB813327F.

Liu:1999:ABM

[LNTZ99] Zhen Liu, Philippe Nain, Don Towsley, and Zhi-Li Zhang. Asymptotic behavior of a multiplexer fed by a long-range dependent process. Journal of Applied Probability, 36(1):105– 118, March 1999. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3215406.

Lo:1986:ABH

[Lo86] Gane Samb Lo. Asymptotic behavior of Hill’s estimate and applications. Journal of Applied Probability, 23(4):922–936, December 1986. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3214466.

Li:1995:CIP

[LO95] Jingwen Li and Jihong Ou. Characterizing the idle-period distribution of GI/G/1 queues. Journal of Applied Probability, 32 (1):247–255, March 1995. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3214933.

Lockhart:1982:NEC

[Loc82] Richard Lockhart. On the non-existence of consistent estimates in Galton–Watson processes. Journal of Applied Probability, 19 (4):842–846, December 1982. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.

org/stable/3213837.

Locatelli:1996:CPS

[Loc96] M. Locatelli. Convergence properties of simulated annealing for continuous global optimization. Journal of Applied Probability, 33(4):1127–1140, December 1996. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3214991.

Loeffen:2009:ODP

[Loe09] R. L. Loeffen. An optimal dividends problem with a terminal value for spectrally negative Lévy processes with a completely monotone jump density. Journal of Applied Probability, 46(1): 85–98, March 2009. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
25662409.

Loisel:2005:DSF

[Loi05] Stéphane Loisel. Differentiation of some functionals of risk processes, and optimal reserve allocation. Journal of Applied Probability, 42(2):379–392, June 2005. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/30040799.

Londono:2009:SDU

[Lon09] Jaime A. Londoño. State-dependent utility. Journal of Applied Probability, 46(1):55–70, March 2009. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/25662407.

Lopes:2014:IBR

[Lop14] Fabio Lopes. Invariant bipartite random graphs on R^{d}. Journal of Applied Probability, 51(3):769–779, September 2014. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://projecteuclid.org/euclid.jap/1409932673.

Lopez:2016:CTB

[Lóp16] Sergio I. López. Convergence of tandem Brownian queues. Journal of Applied Probability, 53(2):585–592, June 2016. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://projecteuclid.org/euclid.jap/1466172876.

Leao:2019:DTA

[LOR19] Dorival Leão, Alberto Ohashi, and Francesco Russo. Discretetype approximations for non-Markovian optimal stopping problems: Part I. Journal of Applied Probability, 56(4): 981–1005, December 2019. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL https:/

/www.cambridge.org/core/journals/journal-of-applied-
probability/article/discretetype-approximations-for-
nonmarkovian-optimal-stopping-problems-part-i/F285FFCB345FF17B49A62C2A751D31B5.

Lotwick:1984:SMM

[Lot84] H. W. Lotwick. Some models for multitype spatial point processes with remarks on analysing multitype patterns. Journal of Applied Probability, 21(3):575–582, September 1984. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3213619.

Loulou:1973:MCQ

[Lou73] Richard Loulou. Multi-channel queues in heavy traffic. Journal of Applied Probability, 10(4):769–777, December 1973. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3212380.

Loulou:1978:EUB

[Lou78] Richard Loulou. An explicit upper bound for the mean busy period in a GI/G/1 queue. Journal of Applied Probability, 15 (2):452–455, June 1978. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3213419.

Louchard:1984:KFL

[Lou84] G. Louchard. Kac’s formula, Levy’s local time and Brownian excursion. Journal of Applied Probability, 21(3):479–499, September 1984. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3213611.

Loynes:1988:SAW

[Loy88] R. M. Loynes. Some applications of weak convergence to statistics. Journal of Applied Probability, 25A:201–211, 1988. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3214157.

Leitnaker:1985:CMT

[LP85] Mary G. Leitnaker and Peter Purdue. Compartmental models with transfer delays: A semi-Markov approach. Journal of Applied Probability, 22(3):570–582, September 1985. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3213861.

Lefevre:1994:DFE

[LP94a] Claude Lefevre and Philippe Picard. Distribution of the final extent of a rumour process. Journal of Applied Probability, 31 (1):244–249, March 1994. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3215250.

Luczak:1994:CNR

[LP94b] Tomasz Luczak and Zbigniew Palka. On the clique number of a random overlap graph. Journal of Applied Probability, 31 (2):582–588, June 1994. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3215051.

Lewis:2000:NBO

[LP00] Mark E. Lewis and Martin L. Puterman. A note on bias optimality in controlled queueing systems. Journal of Applied Probability, 37(1):300–305, March 2000. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3215679.

Lefevre:2005:NRR

[LP05] Claude Lefèvre and Philippe Picard. Nonstationarity and randomization in the Reed–Frost epidemic model. Journal of Applied Probability, 42(4):950–963, December 2005. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/30040820.

Letac:2014:DRW

[LP14] Gérard Letac and Mauro Piccioni. Dirichlet random walks. Journal of Applied Probability, 51(4):1081–1099, December 2014. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://projecteuclid.org/euclid.jap/
1421763329.

Lefevre:2017:OED

[LP17] Claude Lefèvre and Philippe Picard. On the outcome of epidemics with detections. Journal of Applied Probability, 54 (3):890–904, September 2017. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL https:/

/www.cambridge.org/core/journals/journal-of-applied-
probability/article/on-the-outcome-of-epidemics-with-
detections/B2D5AD375E018297E7832A17EF1AFF0C.

Lefevre:2018:GAI

[LP18] Claude Lefèvre and Philippe Picard. A general approach to the integral functionals of epidemic processes. Journal of Applied Probability, 55(2):593–609, June 2018. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL https:/

/www.cambridge.org/core/journals/journal-of-applied-
probability/article/general-approach-to-the-integral-
functionals-of-epidemic-processes/0CE2EA6F6C455BF92D58AFF8D3F4AF8B.

Leman:2021:ECI

[LP21] H. Leman and J. C. Pardo. Extinction and coming down from infinity of continuous-state branching processes with competition in a Lévy environment. Journal of Applied Probability, 58(1):128–139, March 2021. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL https:/

/www.cambridge.org/core/journals/journal-of-applied-
probability/article/extinction-and-coming-down-from-
infinity-of-continuousstate-branching-processes-with-
competition-in-a-levy-environment/0517EA5B1A4CB806E86D5A7C7C65335A.

Li:2023:SAH

[LP23] Bo Li and Guodong Pang. On the splitting and aggregating of Hawkes processes. Journal of Applied Probability, 60(2):676–692, June 2023. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL https:/

/www.cambridge.org/core/journals/journal-of-applied-
probability/article/on-the-splitting-and-aggregating-
of-hawkes-processes/1ACB3AB50238FDE5112B7839140AB50D.

Li:2008:LBD

[LPLG08] Yangrong Li, Anthony G. Pakes, Jia Li, and Anhui Gu. The limit behavior of dual Markov branching processes. Journal of Applied Probability, 45(1):176–189, March 2008. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/27595932.

Liao:2013:TPA

[LPN13] Xin Liao, Zuoxiang Peng, and Saralees Nadarajah. Tail properties and asymptotic expansions for the maximum of the logarithmic skew-normal distribution. Journal of Applied Probability, 50(3):900–907, September 2013. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http:/

/projecteuclid.org/euclid.jap/1378401246.

Longla:2012:FCL

[LPP12] Martial Longla, Costel Peligrad, and Magda Peligrad. On the functional central limit theorem for reversible Markov chains with nonlinear growth of the variance. Journal of Applied Probability, 49(4):1091–1105, December 2012. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/41713830.

Leipus:2006:RCA

[LPS06] Remigijus Leipus, Vygantas Paulauskas, and Donatas Surgailis. On a random-coefficient AR(1) process with heavy-tailed renewal switching coefficient and heavy-tailed noise. Journal of Applied Probability, 43(2):421–440, June 2006. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/27595741.

Lefevre:2017:ERI

[LPS17] Claude Lefèvre, Philippe Picard, and Matthieu Simon. Epidemic risk and insurance coverage. Journal of Applied Probability, 54(1):286–303, March 2017. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL https:/

/www.cambridge.org/core/journals/journal-of-applied-
probability/article/epidemic-risk-and-insurance-coverage/
15FCE01DDBEC6A0DDA04FF9AD5A91C3E.

Lenin:2000:FBD

[LPSV00] R. B. Lenin, P. R. Parthasarathy, W. R. W. Scheinhardt, and E. A. Van Doorn. Families of birth-death processes with similar time-dependent behaviour. Journal of Applied Probability, 37(3): 835–849, September 2000. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3215617.

Lefevre:2020:BMA

[LPU20] Claude Lefèvre, Philippe Picard, and Sergey Utev. On branching models with alarm triggerings. Journal of Applied Probability, 57(3):734–759, September 2020. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL https://www.cambridge.org/core/journals/
journal-of-applied-probability/article/on-branching-
models-with-alarm-triggerings/9AAD10FE4CB8EB1FD29D8768C3D82925.

Lavenberg:1980:SSP

[LR80] S. S. Lavenberg and M. Reiser. Stationary state probabilities at arrival instants for closed queueing networks with multiple types of customers. Journal of Applied Probability, 17(4):1048–1061, December 1980. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3213214.

Lau:1990:CED

[LR90] Ka Sing Lau and B. L. S. Prakasa Rao. Characterization of the exponential distribution by the relevation transform. Journal of Applied Probability, 27(3):726–729, September 1990. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3214557.

Lau:1992:CED

[LR92] K. A. Sing Lau and B. L. S. Prakasa Rao. Characterization of the exponential distribution by the relevation transform. Journal of Applied Probability, 29(4):1003–1004, December 1992. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3214731.

Latouche:1993:LRA

[LR93] Guy Latouche and V. Ramaswami. A logarithmic reduction algorithm for quasi-birth-death processes. Journal of Applied Probability, 30(3):650–674, September 1993. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/3214773.

Loskot:1995:LTS

[LR95] Krzysztof Loskot and Ryszard Rudnicki. Limit theorems for stochastically perturbed dynamical systems. Journal of Applied Probability, 32(2):459–469, June 1995. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3215300.

Lindvall:1996:CRW

[LR96] Torgny Lindvall and L. C. G. Rogers. On coupling of random walks and renewal processes. Journal of Applied Probability, 33 (1):122–126, March 1996. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3215269.

Lamberton:2000:OSE

[LR00] Damien Lamberton and L. C. G. Rogers. Optimal stopping and embedding. Journal of Applied Probability, 37(4):1143–1148, December 2000. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3215505.

Lin:2004:OAC

[LR04] Kyle Y. Lin and Sheldon M. Ross. Optimal admission control for a single-server loss queue. Journal of Applied Probability, 41 (2):535–546, June 2004. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3216034.

Lu:2007:ABP

[LR07] Yingdong Lu and Ana Radovanović. Asymptotic blocking probabilities in loss networks with subexponential demands. Journal of Applied Probability, 44(4):1088–1102, December 2007. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/27595910.

Lagnoux-Renaudie:2009:TSB

[LR09] Agnès Lagnoux-Renaudie. A two-step branching splitting model under cost constraint for rare event analysis. Journal of Applied Probability, 46(2):429–452, June 2009. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/25662435.

Lopez:2012:OPD

[LR12] Oscar López and Nikita Ratanov. Option pricing driven by a telegraph process with random jumps. Journal of Applied Probability, 49(3):838–849, September 2012. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/41713809.

Lee:2013:MEM

[LR13] Young Lee and Thorsten Rheinländer. The minimal entropy martingale measure for exponential Markov chains. Journal of Applied Probability, 50(2):344–358, June 2013. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://projecteuclid.org/euclid.jap/1371648945.

Latuszynski:2014:CCA

[LR14a] Krzysztof Latuszyński and Jeffrey S. Rosenthal. The containment condition and AdapFail algorithms. Journal of Applied Probability, 51(4):1189–1195, December 2014. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://projecteuclid.org/euclid.jap/1421763335.

Lopez:2014:ATP

[LR14b] Oscar López and Nikita Ratanov. On the asymmetric telegraph processes. Journal of Applied Probability, 51(2):569–589, June 2014. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://projecteuclid.org/euclid.jap/
1402578644.

Lee:2023:CTH

[LR23] Young Lee and Thorsten Rheinländer. On the cumulant transforms for Hawkes processes. Journal of Applied Probability, 60(2):528–541, June 2023. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL https:/

/www.cambridge.org/core/journals/journal-of-applied-
probability/article/on-the-cumulant-transforms-for-hawkes-
processes/0C803CC30B31D3629D76A79D1426F8D5.

Li:2024:MDI

[LR24] Jialin Li and Ilya O. Ryzhov. Moderate deviations inequalities for Gaussian process regression. Journal of Applied Probability, 61(1):172–197, March 2024. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL https:/

/www.cambridge.org/core/journals/journal-of-applied-
probability/article/moderate-deviations-inequalities-
for-gaussian-process-regression/0A2685C86CC946B2C5D9C2D4FFD943BC.

Leonenko:2023:SFS

[LRM23] N. N. Leonenko and M. D. Ruiz-Medina. Sojourn functionals for spatiotemporal Gaussian random fields with long memory. Journal of Applied Probability, 60(1):148–165, March 2023. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL https://www.cambridge.org/core/journals/journal-
of-applied-probability/article/sojourn-functionals-for-
spatiotemporal-gaussian-random-fields-with-long-memory/
3D01EEAD85FC8EA1A5CF97F8B9784E86.

Ledoux:1994:EAA

[LRS94] James Ledoux, Gerardo Rubino, and Bruno Sericola. Exact aggregation of absorbing Markov processes using the quasistationary distribution. Journal of Applied Probability, 31(3): 626–634, September 1994. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3215143.

Lippman:2007:WLM

[LRS07] Steven A. Lippman, Sheldon M. Ross, and Sridhar Seshadri. A weakest link marked stopping problem. Journal of Applied Probability, 44(4):843–851, December 2007. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/27595891.

Liu:2009:CCS

[LRS09] Rong-Li Liu, Yan-Xia Ren, and Renming Song. L log L criterion for a class of superdiffusions. Journal of Applied Probability, 46 (2):479–496, June 2009. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/25662438.

Logan:1974:BDM

[LS74] B. F. Logan and L. A. Shepp. A birth and death model of neuron firing. Journal of Applied Probability, 11(2):369–373, June 1974. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3212758.

Lloyd:1979:NSM

[LS79] E. H. Lloyd and S. D. Saleem. A note on seasonal Markov chains with gamma or gamma-like distributions. Journal of Applied Probability, 16(1):117–128, March 1979. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http:/

/www.jstor.org/stable/3213380.

Lindley:1986:MDL

[LS86] Dennis V. Lindley and Nozer D. Singpurwalla. Multivariate distributions for the life lengths of components of a system sharing a common environment. Journal of Applied Probability, 23 (2):418–431, June 1986. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3214184.

Langberg:1987:MAM

[LS87] Naftali A. Langberg and David S. Stoffer. Moving-average models with bivariate exponential and geometric distributions. Journal of Applied Probability, 24(1):48–61, March 1987. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3214058.

Liese:1991:API

[LS91] Friedrich Liese and Volker Schmidt. Asymptotic properties of intensity estimators for Poisson shot-noise processes. Journal of Applied Probability, 28(3):568–583, September 1991. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3214492.

Lal:1992:ZST

[LS92a] Arbind K. Lal and Sagnik Sinha. Zero-sum two-person semi-Markov games. Journal of Applied Probability, 29(1):56–72, March 1992. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3214791.

Liu:1992:SSE

[LS92b] Jian Liu and Ed Susko. On strict stationarity and ergodicity of a non-linear ARMA model. Journal of Applied Probability, 29 (2):363–373, June 1992. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3214573.

Liu:1996:BPG

[LS96] Liming Liu and Ding-Hua Shi. Busy period in GI^{X} /G/∞. Journal of Applied Probability, 33(3):815–829, September 1996. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/3215361.

Li:1997:AFP

[LS97a] Haijum Li and Moshe Shaked. Ageing first-passage times of Markov processes: A matrix approach. Journal of Applied Probability, 34(1):1–13, March 1997. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.

org/stable/3215169.

Lund:1997:CCR

[LS97b] Robert Lund and Walter Smith. A comparison of convergence rates for three models in the theory of dams. Journal of Applied Probability, 34(1):74–83, March 1997. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3215176.

Last:1998:SCR

[LS98] Günter Last and Ryszard Szekli. Stochastic comparison of repairable systems by coupling. Journal of Applied Probability, 35 (2):348–370, June 1998. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3215690.

Lladser:2000:DAQ

[LS00] Manuel Lladser and Jaime San Martín. Domain of attraction of the quasi-stationary distributions for the Ornstein–Uhlenbeck process. Journal of Applied Probability, 37(2):511–520, June 2000. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3215725.

Lin:2002:MDD

[LS02a] Gwo Dong Lin and Jordan Stoyanov. On the moment determinacy of the distributions of compound geometric sums. Journal of Applied Probability, 39(3):545–554, September 2002. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3216077.

Lopez:2002:MCS

[LS02b] F. Javier López and Gerardo Sanz. Markovian couplings staying in arbitrary subsets of the state space. Journal of Applied Probability, 39(1):197–212, March 2002. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3215927.

Li:2003:IRM

[LS03] Haijun Li and Moshe Shaked. Imperfect repair models with preventive maintenance. Journal of Applied Probability, 40(4): 1043–1059, December 2003. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3216059.

Lepeltier:2004:BST

[LS04] Jean-Pierre Lepeltier and Jaime San Martín. Backward SDEs with two barriers and continuous coefficient: An existence result. Journal of Applied Probability, 41(1):162–175, March 2004. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/3215822.

Leipus:2007:LRD

[LS07] Remigijus Leipus and Donatas Surgailis. On long-range dependence in regenerative processes based on a general on/off scheme. Journal of Applied Probability, 44(2):379–392, June 2007. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/27595848.

Lageraas:2008:RBP

[LS08a] Andreas N. Lagerås and Serik Sagitov. Reduced branching processes with very heavy tails. Journal of Applied Probability, 45 (1):190–200, March 2008. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/27595933.

Lehrer:2008:TRB

[LS08b] Ehud Lehrer and Eran Shmaya. Two remarks on Blackwell’s theorem. Journal of Applied Probability, 45(2):580–586, June 2008. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/27595967.

See [Bla53].

Li:2009:TDH

[LS09a] Haijun Li and Yannan Sun. Tail dependence for heavy-tailed scale mixtures of multivariate distributions. Journal of Applied Probability, 46(4):925–937, December 2009. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/25662472.

Liggett:2009:SMP

[LS09b] Thomas M. Liggett and Rinaldo B. Schinazi. A stochastic model for phylogenetic trees. Journal of Applied Probability, 46(2):601– 607, June 2009. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
25662447.

Lin:2009:LSN

[LS09c] Gwo Dong Lin and Jordan Stoyanov. The logarithmic skewnormal distributions are moment-indeterminate. Journal of Applied Probability, 46(3):909–916, September 2009. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/25662467.

Lopker:2011:HTR

[LS11] Andreas Löpker and Wolfgang Stadje. Hitting times and the running maximum of Markovian growth-collapse processes. Journal of Applied Probability, 48(2):295–312, June 2011. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/23065886.

Laraki:2017:CQD

[LS17] R. Laraki and W. Sudderth. A continuity question of Dubins and Savage. Journal of Applied Probability, 54(2):462–473, June 2017. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-

6072 (electronic). URL https://www.cambridge.org/core/
journals/journal-of-applied-probability/article/continuity-
question-of-dubins-and-savage/0F568A0A6724EC39DF1291AB61CEE9A0.

Last:2019:NAS

[LS19] Günter Last and Ryszard Szekli. On negative association of some finite point processes on general state spaces. Journal of Applied Probability, 56(1):139–152, March 2019. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL https:/

/www.cambridge.org/core/journals/journal-of-applied-
probability/article/on-negative-association-of-some-
finite-point-processes-on-general-state-spaces/C33E161ADD7D8BB68DFACBECAD42426D.

Liao:2020:LTB

[LS20] Zhongwei Liao and Jinghai Shao. Long-time behavior of Lévydriven Ornstein–Uhlenbeck processes with regime switching. Journal of Applied Probability, 57(1):266–279, March 2020. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL https://www.cambridge.org/core/journals/journal-
of-applied-probability/article/longtime-behavior-of-
levydriven-ornsteinuhlenbeck-processes-with-regime-switching/
BE65B7A2A7586D36A8CCA14DF1C40F7F.

Li:1997:RAM

[LSC97] Wei Li, Dinghua Shi, and Xiuli Chao. Reliability analysis of M/G/1 queueing systems with server breakdowns and vacations. Journal of Applied Probability, 34(2):546–555, June 1997. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/3215393.

Lin:2004:RQS

[LSC04] Y.-X. Lin, D. Steel, and R. L. Chambers. Restricted quasi-score estimating functions for sample survey data. Journal of Applied Probability, 41A:119–130, 2004. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.

org/stable/3215972.

Leonenko:2019:LTF

[LST19] Nikolai Leonenko, Enrico Scalas, and Mailan Trinh. Limit theorems for the fractional nonhomogeneous Poisson process. Journal of Applied Probability, 56(1):246–264, March 2019. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL https://www.cambridge.org/core/journals/journal-
of-applied-probability/article/limit-theorems-for-the-
fractional-nonhomogeneous-poisson-process/C87D20A2DC9B36AD0B39C9E8C1769994.

Lindqvist:2019:PMR

[LSW19] Bo H. Lindqvist, Francisco J. Samaniego, and Nana Wang. Preservation of the mean residual life order for coherent and mixed systems. Journal of Applied Probability, 56(1):153–173, March 2019. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL https:/

/www.cambridge.org/core/journals/journal-of-applied-
probability/article/preservation-of-the-mean-residual-
life-order-for-coherent-and-mixed-systems/AC8E13ED72A1D991F87DDA5006FBD20A.

Liu:2020:IPL

[LSWX20] Yuelin Liu, Vladas Sidoravicius, Longmin Wang, and Kainan Xiang. An invariance principle and a large deviation principle for the biased random walk on Z^{d}. Journal of Applied Probability, 57(1):295–313, March 2020. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL https:/

/www.cambridge.org/core/journals/journal-of-applied-
probability/article/an-invariance-principle-and-a-large-
deviation-principle-for-the-biased-random-walk-on-mathbbzlowercasetextbftextitd/
CB4A2FEA0343A8A2F9CEE944ECBBD893.

Loris-Teghem:1971:WTD

[LT71] Jacqueline Loris-Teghem. On the waiting time distribution in a generalized GI/G/1 queueing system. Journal of Applied Probability, 8(2):241–251, June 1971. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.

org/stable/3211894.

Loris-Teghem:1972:WTD

[LT72] Jacqueline Loris-Teghem. On the waiting time distribution in a generalized queueing system with uniformly bounded sojourn times. Journal of Applied Probability, 9(3):642–649, September 1972. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3212332.

Loris-Teghem:1973:AAW

[LT73] Jacqueline Loris-Teghem. An algebraic approach to the waiting time process in GI/M/S. Journal of Applied Probability, 10 (1):181–191, March 1973. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3212505.

Lee:1976:FRU

[LT76] Larry Lee and W. A. Thompson, Jr. Failure rate: A unified approach. Journal of Applied Probability, 13(1):176–182, March 1976. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3212681.

Letac:1980:RWD

[LT80] G. Letac and L. Takács. Random walks on a dodecahedron. Journal of Applied Probability, 17(2):373–384, June 1980. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/3213027.

Lewis:1981:DDC

[LT81] Toby Lewis and J. W. Thompson. Dispersive distributions, and the connection between dispersivity and strong unimodality. Journal of Applied Probability, 18(1):76–90, March 1981. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/3213168.

Loris-Teghem:1990:DRC

[LT90] Jacqueline Loris-Teghem. On a decomposition result for a class of vacation queueing systems. Journal of Applied Probability, 27 (1):227–231, March 1990. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3214611.

Liu:1994:ORR

[LT94] Zhen Liu and Don Towsley. Optimality of the round-robin routing policy. Journal of Applied Probability, 31(2):466–475, June 1994. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3215039.

Li:1995:GNC

[LT95] Xunjing Li and Shanjian Tang. General necessary conditions for partially observed optimal stochastic controls. Journal of Applied Probability, 32(4):1118–1137, December 1995. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3215225.

Latouche:2000:LPI

[LT00] Guy Latouche and P. G. Taylor. Level-phase independence for GI/M/1-Type Markov chains. Journal of Applied Probability, 37 (4):984–998, December 2000. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.

org/stable/3215489.

Latouche:2004:MRF

[LT04a] Guy Latouche and Tetsuya Takine. Markov-renewal fluid queues. Journal of Applied Probability, 41(3):746–757, September 2004. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/4141351.

Loris-Teghem:2004:NWT

[LT04b] Jacqueline Loris-Teghem. A note on the waiting time in M^{[X]} /G/1 queueing systems with a removable server. Journal of Applied Probability, 41(1):287–291, March 2004. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3215834.

Lambert:2013:STS

[LT13] Amaury Lambert and Pieter Trapman. Splitting trees stopped when the first clock rings and Vervaat’s transformation. Journal of Applied Probability, 50(1):208–227, March 2013. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://projecteuclid.org/euclid.jap/1363784434.

Levin:2023:FMR

[LT23] David A. Levin and Chandan Tankala. Fast mixing of a randomized shift-register Markov chain. Journal of Applied Probability, 60(1):253–266, March 2023. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL https:/

/www.cambridge.org/core/journals/journal-of-applied-
probability/article/fast-mixing-of-a-randomized-shiftregister-
markov-chain/BFDCDCF101AE03E95895D8CDDD79F401.

Li:2013:THD

[LTZ13] Bin Li, Qihe Tang, and Xiaowen Zhou. A time-homogeneous diffusion model with tax. Journal of Applied Probability, 50 (1):195–207, March 2013. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://projecteuclid.

org/euclid.jap/1363784433.

Lukins:2018:MPC

[LTZ18] James Lukins, Roger Tribe, and Oleg Zaboronski. Multipoint correlations for two-dimensional coalescing or annihilating random walks. Journal of Applied Probability, 55 (4):1158–1185, December 2018. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL https:/

/www.cambridge.org/core/journals/journal-of-applied-
probability/article/multipoint-correlations-for-twodimensional-
coalescing-or-annihilating-random-walks/779E7E9C05E4A86E3F84DC2CBC336BD2.

Lefevre:1995:CNE

[LU95] Claude Lefévre and Sergey Utev. On a conjecture for the nonexistence of the expectation of randomly stopped sums. Journal of Applied Probability, 32(4):1138–1141, December 1995. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/3215226.

Lefevre:1996:CSE

[LU96] Claude Lefèvre and Sergey Utev. Comparing sums of exchangeable Bernoulli random variables. Journal of Applied Probability, 33(2):285–310, June 1996. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3215055.

Lu:2010:RPC

[Lu10] Bing Lu. Recovering a piecewise constant volatility from perpetual put option prices. Journal of Applied Probability, 47(3):

680–692, September 2010. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/20778510.

Ludwig:1973:SAG

[Lud73] Donald Ludwig. Stochastic approximation for the general epidemic. Journal of Applied Probability, 10(2):263–276, June 1973. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/3212344.

Lukacs:1969:CSP

[Luk69] Eugene Lukacs. A characterization of stable processes. Journal of Applied Probability, 6(2):409–418, August 1969. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3212010.

Lund:1994:DSI

[Lun94] Robert B. Lund. A dam with seasonal input. Journal of Applied Probability, 31(2):526–541, June 1994. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3215043.

Lund:1996:SSM

[Lun96] Robert B. Lund. The stability of storage models with shot noise input. Journal of Applied Probability, 33(3):830–839, September 1996. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3215362.

Lund:1997:GCR

[Lun97] Robert B. Lund. The geometric convergence rate of a Lindley random walk. Journal of Applied Probability, 34(3):806–811, September 1997. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3215107.

Luo:2000:PAS

[Luo00] Shunlong Luo. Probabilistic aspects of the SU(2)-Harmonic oscillator. Journal of Applied Probability, 37(1):306–314, March 2000. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3215680.

Luo:2014:SBG

[Luo14] Shangzhen Luo. Stochastic Brownian game of absolute dominance. Journal of Applied Probability, 51(2):436–452, June 2014. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://projecteuclid.org/euclid.jap/
1402578635.

Littlewood:1972:EDP

[LV72] B. Littlewood and J. L. Verrall. Equilibrium distributions of polymers on sites which are linearly arrayed. Journal of Applied Probability, 9(1):54–64, March 1972. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3212636.

Lopker:2008:TMT

[LV08] Andreas H. Löpker and Johans S. H. Van Leeuwaarden. Trasient moments of the TCP window size process. Journal of Applied Probability, 45(1):163–175, March 2008. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http:/

/www.jstor.org/stable/27595931.

Leskela:2012:JEP

[LV12a] Lasse Leskelä and Harri Varpanen. Juggler’s exclusion process. Journal of Applied Probability, 49(1):266–279, March 2012. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/41713764.

Li:2012:PCB

[LV12b] Wenbo V. Li and Vladislav V. Vysotsky. Probabilities of competing binomial random variables. Journal of Applied Probability, 49(3):731–744, September 2012. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/41713802.

Li:2019:EPG

[LVZ19] Bo Li, Nhat Linh Vu, and Xiaowen Zhou. Exit problems for general draw-down times of spectrally negative Lévy processes. Journal of Applied Probability, 56(2):441–457, June 2019. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL https://www.cambridge.org/core/journals/journal-
of-applied-probability/article/exit-problems-for-general-
drawdown-times-of-spectrally-negative-levy-processes/
448DA9A2E0C4A312E982CFB484B51466.

Lee:1993:SPD

[LW93] Mei-Ling Ting Lee and G. Alex Whitmore. Stochastic processes directed by randomized time. Journal of Applied Probability, 30 (2):302–314, June 1993. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3214840.

Lenthe:2019:CAS

[LW19] Pierre Lenthe and Philip Weissmann. Completely asymmetric stable processes conditioned to avoid an interval. Journal of Applied Probability, 56(4):1187–1197, December 2019. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL https://www.cambridge.org/core/journals/journal-
of-applied-probability/article/completely-asymmetric-
stable-processes-conditioned-to-avoid-an-interval/C82356D6AAC2D2140C635BAD0E601D62.

Ledoux:2008:RBL

[LWB08] James Ledoux, Langford B. White, and Gary D. Brushe. On the relationships between lumpability and filtering of finite stochastic systems. Journal of Applied Probability, 45(3):650–669, September 2008. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
27595978.

Li:2020:SCL

[LWM20] Lu Li, Qinyu Wu, and Tiantian Mao. Stochastic comparisons of largest-order statistics for proportional reversed hazard rate model and applications. Journal of Applied Probability, 57(3):832–852, September 2020. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL https:/

/www.cambridge.org/core/journals/journal-of-applied-
probability/article/stochastic-comparisons-of-largestorder-
statistics-for-proportional-reversed-hazard-rate-model-
and-applications/33600895AFBBECCF4091F7C1DFF7EDB5.

Li:2018:TAP

[LWW18] Bin Li, Gordon E. Willmot, and Jeff T. Y. Wong. A temporal approach to the Parisian risk model. Journal of Applied Probability, 55(1):302–317, March 2018. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL https://www.cambridge.org/core/journals/journal-
of-applied-probability/article/temporal-approach-to-
the-parisian-risk-model/353A4346C50A5B9A54591C4622F45CFC.

Li:2000:DSB

[LX00] Haijun Li and Susan H. Xu. On the dependence structure and bounds of correlated parallel queues and their applications to synchronized stochastic systems. Journal of Applied Probability, 37(4):1020–1043, December 2000. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3215492.

Li:2018:UES

[LX18] Xiaoou Li and Gongjun Xu. Uniformly efficient simulation for extremes of Gaussian random fields. Journal of Applied Probability, 55(1):157–178, March 2018. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL https:/

/www.cambridge.org/core/journals/journal-of-applied-
probability/article/uniformly-efficient-simulation-for-
extremes-of-gaussian-random-fields/465B56715C13477ECBAB4D40555F6BF5.

Liu:2020:MDP

[LX20] Qingwei Liu and Aihua Xia. On moderate deviations in Poisson approximation. Journal of Applied Probability, 57 (3):1005–1027, September 2020. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL https:/

/www.cambridge.org/core/journals/journal-of-applied-
probability/article/on-moderate-deviations-in-poisson-
approximation/34E48F71D4584B1EA68D0DCF9A617D38.

Lee:1995:MDA

[LY95] Chinsan Lee and Grace L. Yang. A multitype decomposable age-dependent branching process and its applications. Journal of Applied Probability, 32(3):591–608, September 1995. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3215115.

Liu:2020:SSA

[LY20] Xin Liu and Lei Ying. Steady-state analysis of loadbalancing algorithms in the sub-Halfin–Whitt regime. Journal of Applied Probability, 57(2):578–596, June 2020. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL https://www.cambridge.org/core/journals/journal-
of-applied-probability/article/steadystate-analysis-
of-loadbalancing-algorithms-in-the-subhalfinwhitt-regime/
2031145F0EFD11E476AF693E7866FEDC.

Liang:2022:DPE

[LY22] Xiaoqing Liang and Virginia R. Young. Discounted probability of exponential Parisian ruin: Diffusion approximation. Journal of Applied Probability, 59(1):17–37, March 2022. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL https://www.cambridge.org/core/journals/journal-
of-applied-probability/article/discounted-probability-
of-exponential-parisian-ruin-diffusion-approximation/
5B76272A02797EAD64B846C3B3FAFAB7.

Liang:2023:OPR

[LY23] Xiaoqing Liang and Virginia R. Young. Optimal proportional reinsurance to maximize an insurer’s exponential utility under unobservable drift. Journal of Applied Probability, 60(3):874–894, September 2023. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL https:/

/www.cambridge.org/core/journals/journal-of-applied-
probability/article/optimal-proportional-reinsurance-
to-maximize-an-insurers-exponential-utility-under-unobservable-
drift/E2DC7E603E8738B31B3CEFBEC4945627.

Lyle:1988:HMH

[Lyl88] Kathleen M. Lyle. Hot metal to hard disc. Journal of Applied Probability, 25A:65–75, 1988. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.

org/stable/3214147.

Lynch:1987:ADM

[Lyn87] James Lynch. Antistarshapedness, dispersiveness and mixtures. Journal of Applied Probability, 24(1):265–269, March 1987. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/3214079.

Lynch:2000:GWP

[Lyn00] James D. Lynch. The Galton–Watson process revisited: Some martingale relationships and applications. Journal of Applied Probability, 37(2):322–328, June 2000. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3215710.

Li:2017:LET

[LYZ17] Yingqiu Li, Chuancun Yin, and Xiaowen Zhou. On the last exit times for spectrally negative Lévy processes. Journal of Applied Probability, 54(2):474–489, June 2017. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL https:/

/www.cambridge.org/core/journals/journal-of-applied-
probability/article/on-the-last-exit-times-for-spectrally-
negative-levy-processes/EFDC0D6179F0972EDD48B341985F951E.

Li:1994:NAG

[LZ94] Huan Li and Yixin Zhu. A new approach to the G/G/1 queue with generalized setup time and exhaustive service. Journal of Applied Probability, 31(4):1083–1097, December 1994. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3215331.

Li:2002:BSN

[LZ02] Xiaohu Li and Ming J. Zuo. On the behaviour of some new ageing properties based upon the residual life of k-out-of-n systems. Journal of Applied Probability, 39(2):426–433, June 2002. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/3216110.

Liu:2014:EAR

[LZ14] Jingchen Liu and Xiang Zhou. Extreme analysis of a random ordinary differential equation. Journal of Applied Probability, 51(4):1021–1036, December 2014. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

projecteuclid.org/euclid.jap/1421763325.

Lund:2006:MRM

[LZK06a] Robert Lund, Ying Zhao, and Peter C. Kiessler. A monotonicity in reversible Markov chains. Journal of Applied Probability, 43 (2):486–499, June 2006. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/27595745.

Lund:2006:SSA

[LZK06b] Robert Lund, Ying Zhao, and Peter C. Kiessler. Shapes of stationary autocovariances. Journal of Applied Probability, 43(4): 1186–1193, December 2006. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/27595810.

Mine:1969:EAR

[MA69] Hisashi Mine and Tatsuyuki Asakura. The effect of an age replacement to a standby redundant system. Journal of Applied Probability, 6(3):516–523, December 1969. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/3212099.

Ma:1997:NSO

[Ma97] Chunsheng Ma. A note on stochastic ordering of order statistics. Journal of Applied Probability, 34(3):785–789, September 1997. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3215103.

Ma:2002:CML

[Ma02] Chunsheng Ma. Correlation models with long-range dependence. Journal of Applied Probability, 39(2):370–382, June 2002. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/3216102.

Ma:2003:LMC

[Ma03a] Chunsheng Ma. Long-memory continuous-time correlation models. Journal of Applied Probability, 40(4):1133–1146, December 2003. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3216065.

Ma:2003:PLC

[Ma03b] Chunsheng Ma. Power-law correlations and other models with long-range dependence on a lattice. Journal of Applied Probability, 40(3):690–703, September 2003. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3215944.

Ma:2004:UVC

[Ma04] Chunsheng Ma. The use of the variogram in construction of stationary time series models. Journal of Applied Probability, 41 (4):1093–1103, December 2004. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.

org/stable/4141381.

Matsui:2014:NLB

[MA14] Tomomi Matsui and Katsunori Ano. A note on a lower bound for the multiplicative odds theorem of optimal stopping. Journal of Applied Probability, 51(3):885–889, September 2014. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://projecteuclid.org/euclid.jap/1409932681.

Matsui:2017:CRS

[MA17] Tomomi Matsui and Katsunori Ano. Compare the ratio of symmetric polynomials of odds to one and stop. Journal of Applied Probability, 54(1):12–22, March 2017. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL https://www.cambridge.org/core/journals/journal-
of-applied-probability/article/compare-the-ratio-of-
symmetric-polynomials-of-odds-to-one-and-stop/6A98E3E54304386183A9F60EC1E66085.

MacGillivray:1980:MRP

[Mac80] H. L. MacGillivray. The moment ratios of particle size distributions in some simple growth models. Journal of Applied Probability, 17(4):956–967, December 1980. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3213205.

MacGillivray:1981:NNM

[Mac81] H. L. MacGillivray. A note on the normalised moments of distributions with non-monotonic hazard rate. Journal of Applied Probability, 18(2):530–535, June 1981. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3213301.

Macci:2001:CTM

[Mac01] Claudio Macci. Continuous-time Markov additive processes: Composition of large deviations principles and comparison between exponential rates of convergence. Journal of Applied Probability, 38(4):917–931, December 2001. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/3215774.

Madan:1975:CTN

[Mad75a] K. C. Madan. On a certain type of network of queues. Journal of Applied Probability, 12(1):195–200, March 1975. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3212429.

Madsen:1975:DSK

[Mad75b] Richard W. Madsen. Decidability of α(P^{k}) > 0 for some k. Journal of Applied Probability, 12(2):333–340, June 1975. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/3212447.

Madan:1976:LE

[Mad76] K. C. Madan. Letter to the Editor. Journal of Applied Probability, 13(2):427, June 1976. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3212853.

Maehara:1980:RSP

[Mae80] Hiroshi Maehara. On random simplices in product distributions. Journal of Applied Probability, 17(2):553–558, June 1980. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/3213045.

Mahmoud:2003:MDS

[Mah03] Hosam M. Mahmoud. Mixed distributions in Sattolo’s algorithm for cyclic permutations via randomization and derandomization. Journal of Applied Probability, 40(3):790–796, September 2003. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/3215952.

Mahmoud:2020:MSF

[Mah20] Hosam Mahmoud. A model for the spreading of fake news. Journal of Applied Probability, 57(1):332–342, March 2020. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL https://www.cambridge.org/core/journals/journal-
of-applied-probability/article/model-for-the-spreading-
of-fake-news/70EEC37DCA24C06C47B7D5F9C0CCA781.

Makowski:1994:ECI

[Mak94] Armand M. Makowski. On an elementary characterization of the increasing convex ordering, with an application. Journal of Applied Probability, 31(3):834–840, September 1994. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3215161.

Makri:2010:OFS

[Mak10] Frosso S. Makri. On occurrences of F-S strings in linearly and circularly ordered binary sequences. Journal of Applied Probability, 47(1):157–178, March 2010. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/25662507.

Mallows:1967:LPN

[Mal67] C. L. Mallows. Linear processes are nearly Gaussian. Journal of Applied Probability, 4(2):313–329, August 1967. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3212026.

Mallows:1974:ADG

[Mal74] C. L. Mallows. Another definition of genealogical distance. Journal of Applied Probability, 11(1):179–183, March 1974. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3212595.

Mamer:1987:MSG

[Mam87] John W. Mamer. Monotone stopping games. Journal of Applied Probability, 24(2):386–401, June 1987. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3214263.

Mandl:1969:IMR

[Man69] Petr Mandl. An identity for Markovian replacement processes. Journal of Applied Probability, 6(2):348–354, August 1969. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/3212005.

Mandl:1971:SRM

[Man71] Petr Mandl. Some results on Markovian replacement processes. Journal of Applied Probability, 8(2):357–365, June 1971. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/3211905.

Mandelbrot:1974:PBM

[Man74] Benoît Mandelbrot. A population birth-and-mutation process, I: Explicit distributions for the number of mutants in an old culture of bacteria. Journal of Applied Probability, 11(3):437–444, September 1974. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3212688.

Mann:1977:LDS

[Man77] J. E. Mann. On the length distribution of stochastic subarcs of a rectifiable arc. Journal of Applied Probability, 14(3):654–660, September 1977. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3213473.

Mannion:1983:SRD

[Man83] David Mannion. Sequential random displacements of points in an interval. Journal of Applied Probability, 20(2):251–263, June 1983. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3213799.

Mano:2013:AGB

[Man13a] Shuhei Mano. Ancestral graph with bias in gene conversion. Journal of Applied Probability, 50(1):239–255, March 2013. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://projecteuclid.org/euclid.jap/1363784436.

Mano:2013:DBT

[Man13b] Shuhei Mano. Duality between the two-locus Wright–Fisher diffusion model and the ancestral process with recombination. Journal of Applied Probability, 50(1):256–271, March 2013. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://projecteuclid.org/euclid.jap/1363784437.

Mao:2002:SEM

[Mao02] Yong-Hua Mao. Strong ergodicity for Markov processes by coupling methods. Journal of Applied Probability, 39(4):839–852, December 2002. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3216007.

Mao:2004:EIC

[Mao04] Yong-Hua Mao. The eigentime identity for continuous-time ergodic Markov chains. Journal of Applied Probability, 41(4):1071– 1080, December 2004. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/4141379.

Maruyama:1969:GCS

[Mar69] Takeo Maruyama. Genetic correlation in the stepping stone model with non-symmetrical migration rates. Journal of Applied Probability, 6(3):463–477, December 1969. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/3212095.

Marcus:1973:TNF

[Mar73a] Allan H. Marcus. Traffic noise as a filtered Markov renewal process. Journal of Applied Probability, 10(2):377–386, June 1973. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/3212354.

Marsh:1973:EDF

[Mar73b] L. M. Marsh. An expansion for the distribution function of a random sum. Journal of Applied Probability, 10(4):869–874, December 1973. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3212389.

Martin:1986:NAE

[Mar86] R. J. Martin. A note on the asymptotic eigenvalues and eigenvectors of the dispersion matrix of a second-order stationary process on a d-dimensional lattice. Journal of Applied Probability, 23(2):529–535, June 1986. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3214196.

Marcheselli:2000:GDM

[Mar00] Marzia Marcheselli. A generalized delta method with applications to intrinsic diversity profiles. Journal of Applied Probability, 37(2):504–510, June 2000. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3215724.

Margolius:2005:TST

[Mar05] B. H. Margolius. Transient solution to the time-dependent multiserver Poisson queue. Journal of Applied Probability, 42(3): 766–777, September 2005. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/30040855.

Markstrom:2007:NAD

[Mar07] Klas Markström. Negative association does not imply logconcavity of the rank sequence. Journal of Applied Probability, 44(4):1119–1121, December 2007. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/27595913. See correction [Ano10e].

Marouby:2011:MDT

[Mar11] Matthieu Marouby. Micropulses and different types of Brownian motion. Journal of Applied Probability, 48(3):792–810, September 2011. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/23065937.

Marty:2013:HPM

[Mar13] Renaud Marty. From Hermite polynomials to multifractional processes. Journal of Applied Probability, 50(2):323–343, June 2013. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://projecteuclid.org/euclid.jap/
1371648944.

Marichal:2015:AFC

[Mar15] Jean-Luc Marichal. Algorithms and formulae for conversion between system signatures and reliability functions. Journal of Applied Probability, 52(2):490–507, June 2015. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

projecteuclid.org/euclid.jap/1437658611.

Martinez:2016:SDA

[Mar16a] Servet Martínez. A stationary distribution associated to a set of laws whose initial states are grouped into classes. An application in genomics. Journal of Applied Probability, 53(2):315–326, June 2016. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-

6072 (electronic). URL http://projecteuclid.org/euclid.
jap/1466172856.

Martyr:2016:SFT

[Mar16b] Randall Martyr. Solving finite time horizon Dynkin games by optimal switching. Journal of Applied Probability, 53(4):957– 973, December 2016. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://projecteuclid.

org/euclid.jap/1481132829.

Martinez:2021:EKR

[Mar21]  Servet Martínez.  Entropy of killed-resurrected station-

ary Markov chains. Journal of Applied Probability, 58

(1):177–196, March 2021.  CODEN JPRBAM.  ISSN

0021-9002 (print), 1475-6072 (electronic). URL https:/

/www.cambridge.org/core/journals/journal-of-applied-
probability/article/entropy-of-killedresurrected-stationary-
markov-chains/48EE7490904391A8AA8BD7D84D49D80A.

Masry:1977:NGP

[Mas77] Elias Masry. A note on Gaussian processes and zero-memory non-linear transformations. Journal of Applied Probability, 14 (4):857–861, December 1977. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.

org/stable/3213359.

Mase:1982:APS

[Mas82] Shigeru Mase. Asymptotic properties of stereological estimators of volume fraction for stationary random sets. Journal of Applied Probability, 19(1):111–126, March 1982. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http:/

/www.jstor.org/stable/3213921.

Massey:1984:OAA

[Mas84] William A. Massey. An operator-analytic approach to the Jackson network. Journal of Applied Probability, 21(2):379–393, June 1984. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3213647.

Mase:1986:PFS

[Mas86a] Shigeru Mase. On the possible form of size distributions for Gibbsian processes of mutually non-intersecting balls. Journal of Applied Probability, 23(3):646–659, September 1986. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3214003.

Massey:1986:FBT

[Mas86b] William A. Massey. A family of bounds for the transient behavior of a Jackson network. Journal of Applied Probability, 23 (2):543–549, June 1986. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3214198.

Massey:1987:CET

[Mas87] William A. Massey. Calculating exit times for series Jackson networks. Journal of Applied Probability, 24(1):226–234, March 1987. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3214073.

Masuda:1993:POS

[Mas93] Yasushi Masuda. Partially observable semi-Markov reward processes. Journal of Applied Probability, 30(3):548–560, September 1993. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3214765.

Masse:1995:IPD

[Mas95] Bruno Massé. Invariance principle for the deviation between the probability content and the interior point proportion of a random convex hull. Journal of Applied Probability, 32(4):1041– 1047, December 1995. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3215218.

Matheron:1974:TUP

[Mat74] G. Matheron. Un théorème d’unicité pour les hyperplans poissoniens. (French) [A uniqueness theorem for Poissonian hyperplanes]. Journal of Applied Probability, 11(1):184–189, March 1974. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3212596.

Matheron:1976:FCP

[Mat76] Georges Matheron. La formule de Crofton pour les sections Épaisses. (French) [The Crofton formula for thick sections]. Journal of Applied Probability, 13(4):707–713, December 1976. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/3212525.

Matheron:1978:FSP

[Mat78] G. Matheron. La formule de Steiner pour les Érosions. (French) [Steiner’s erosion formula]. Journal of Applied Probability, 15 (1):126–135, March 1978. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3213242.

Mathar:1984:LBM

[Mat84] Rudolf Mathar. The limit behaviour of the maximum of random variables with applications to outlier-resistance. Journal of Applied Probability, 21(3):646–650, September 1984. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3213625.

Matthews:1990:MRB

[Mat90] Peter Matthews. Mixing rates for Brownian motion in a convex polyhedron. Journal of Applied Probability, 27(2):259–268, June 1990. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3214645.

Matthews:1992:SST

[Mat92a] Peter Matthews. Strong stationary times and eigenvalues. Journal of Applied Probability, 29(1):228–233, March 1992. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3214809.

Matus:1992:EMP

[Mat92b] F. Matús. On equivalence of Markov properties over undirected graphs. Journal of Applied Probability, 29(3):745–749, September 1992. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3214910.

Mathe:2004:NIU

[Mat04] Peter Mathé. Numerical integration using V -uniformly ergodic Markov chains. Journal of Applied Probability, 41(4):1104–1112, December 2004. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
4141382.

Mayster:2005:ABP

[May05] Penka Mayster. Alternating branching processes. Journal of Applied Probability, 42(4):1095–1108, December 2005. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/30040831.

Mazziotto:1986:AOS

[Maz86] G. Mazziotto. Approximations of the optimal stopping problem in partial observation. Journal of Applied Probability, 23(2):341– 354, June 1986. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3214178.

McKinlay:2016:EFS

[MB16] S. McKinlay and K. Borovkov. On explicit form of the stationary distributions for a class of bounded Markov chains. Journal of Applied Probability, 53(1):231–243, March 2016. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://projecteuclid.org/euclid.jap/1457470571.

Moyal:2021:PFG

[MBM21] Pascal Moyal, Ana Busić, and Jean Mairesse. A product form for the general stochastic matching model. Journal of Applied Probability, 58(2):449–468, June 2021. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL https://www.cambridge.org/core/journals/journal-
of-applied-probability/article/product-form-for-the-
general-stochastic-matching-model/BBE189CE5EFC761DBD444B0633CFC46D.

Mukherjee:2016:ULB

[MBvLW16] Debankur Mukherjee, Sem C. Borst, Johan S. H. van Leeuwaarden, and Philip A. Whiting. Universality of load balancing schemes on the diffusion scale. Journal of Applied Probability, 53(4):1111–1124, December 2016. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http:/

/projecteuclid.org/euclid.jap/1481132840.

Mallows:1970:LID

[MC70] C. L. Mallows and J. M. C. Clark. Linear-intercept distributions do not characterize plane sets. Journal of Applied Probability, 7(1):240–244, April 1970. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/ stable/3212164. See correction [MC71].

Mallows:1971:CBI

[MC71] C. L. Mallows and J. M. C. Clark. Corrections to “Linear-Intercept Distributions Do Not Characterize Plane Sets”. Journal of Applied Probability, 8(1):208–209, March 1971. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3211851. See [MC70].

Marcus:1978:MET

[MC78] Allan H. Marcus and Stanley Czajkowski, Jr. Moments of the exit-time distribution for carboxyhemoglobin due to cigarette smoking. Journal of Applied Probability, 15(1):13–25, March 1978. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3213232.

McNamara:1990:JSP

[MC90] J. M. McNamara and E. J. Collins. The job search problem as an employer-candidate game. Journal of Applied Probability, 27(4): 815–827, December 1990. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3214825.

Martin:2011:DCS

[MC11] Donald E. K. Martin and Deidra A. Coleman. Distribution of clump statistics for a collection of words. Journal of Applied Probability, 48(4):1049–1059, December 2011. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/23066442.

Mercier:2013:MIR

[MC13] Sophie Mercier and I. T. Castro. On the modelling of imperfect repairs for a continuously monitored gamma wear process through age reduction. Journal of Applied Probability, 50(4): 1057–1076, December 2013. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://projecteuclid.

org/euclid.jap/1389370099.

Moller:2018:PCF

[MC18] Jesper Møller and Andreas D. Christoffersen. Pair correlation functions and limiting distributions of iterated cluster point processes. Journal of Applied Probability, 55 (3):789–809, September 2018. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL https:/

/www.cambridge.org/core/journals/journal-of-applied-
probability/article/pair-correlation-functions-and-limiting-
distributions-of-iterated-cluster-point-processes/B8B9F35BBD1B8AB7A258802B8363D00C.

McCabe:1974:SOD

[McC74] Bernard J. McCabe. Searching for a one-dimensional random Walker. Journal of Applied Probability, 11(1):86–93, March 1974. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3212585.

McClean:1976:CTP

[McC76] Sally I. McClean. A continuous-time population model with Poisson recruitment. Journal of Applied Probability, 13(2):348– 354, June 1976. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3212838.

McClean:1978:CTS

[McC78] Sally I. McClean. Continuous-time stochastic models of a multigrade population. Journal of Applied Probability, 15(1):26– 37, March 1978. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3213233.

McClean:1980:SMM

[McC80] Sally I. McClean. A semi-Markov model for a multigrade population with Poisson recruitment. Journal of Applied Probability, 17(3):846–852, September 1980. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.

org/stable/3212978.

McCullagh:1984:RP

[McC84] Peter McCullagh. Recurrence processes. Journal of Applied Probability, 21(1):167–172, March 1984. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http:/

/www.jstor.org/stable/3213674.

McCormick:1985:WSL

[McC85] William P. McCormick. Weak and strong law results for a function of the spacings. Journal of Applied Probability, 22(3): 543–555, September 1985. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3213859.

McCormick:1997:ESN

[McC97] William P. McCormick. Extremes for shot noise processes with heavy tailed amplitudes. Journal of Applied Probability, 34(3): 643–656, September 1997. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3215091.

Mercier:2003:IAA

[MCC03] S. Mercier, D. Cellier, and D. Charlot. An improved approximation for assessing the statistical significance of molecular sequence features. Journal of Applied Probability, 40(2):427– 441, June 2003. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3215801.

McDunnough:1978:SAS

[McD78] Philip McDunnough. Some aspects of the Smoluchowski process. Journal of Applied Probability, 15(4):663–674, December 1978. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3213424.

McDunnough:1979:ELR

[McD79] Philip McDunnough. Estimating the law of randomly moving particles by counting. Journal of Applied Probability, 16(1):25– 35, March 1979. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3213371.

McDiarmid:1980:PSS

[McD80] Colin McDiarmid. Percolation on subsets of the square lattice. Journal of Applied Probability, 17(1):278–283, March 1980. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/3212947.

McKenzie:1981:ECS

[McK81] Ed McKenzie. Extending the correlation structure of exponential autoregressive-moving-average processes. Journal of Applied Probability, 18(1):181–189, March 1981. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http:/

/www.jstor.org/stable/3213178.

McKenzie:1982:PAT

[McK82] Ed McKenzie. Product autoregression: A time-series characterization of the gamma distribution. Journal of Applied Probability, 19(2):463–468, June 1982. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3213502.

McKeague:1984:EDP

[McK84] Ian W. McKeague. Estimation for diffusion processes under misspecified models. Journal of Applied Probability, 21(3):511–520, September 1984. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3213613.

McKenzie:1988:DSF

[McK88] Ed McKenzie. The distributional structure of finite movingaverage processes. Journal of Applied Probability, 25(2):313– 321, June 1988. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3214439.

McKenna:1989:GLL

[McK89] James McKenna. A generalization of Little’s law to moments of queue lengths and waiting times in closed, product-form queueing networks. Journal of Applied Probability, 26(1):121– 133, March 1989. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3214322.

McKinlay:2014:CTR

[McK14] S. McKinlay. A characterisation of transient random walks on stochastic matrices with Dirichlet distributed limits. Journal of Applied Probability, 51(2):542–555, June 2014. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://projecteuclid.org/euclid.jap/1402578642.

McLeish:1976:FRC

[McL76] D. L. McLeish. Functional and random central limit theorems for the Robbins-munro process. Journal of Applied Probability, 13(1):148–154, March 1976. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3212676.

McNeil:1968:SFC

[McN68] Donald R. McNeil. A solution to the fixed-cycle traffic light problem for compound Poisson arrivals. Journal of Applied Probability, 5(3):624–635, December 1968. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3211926. See note [McQ67].

McNeil:1970:TAC

[McN70] Donald R. McNeil. A theoretical analysis of congestion in a two-Lane freeway. Journal of Applied Probability, 7(2):304–315, August 1970. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3211966.

McNeil:1973:DLC

[McN73] Donald R. McNeil. Diffusion limits for congestion models. Journal of Applied Probability, 10(2):368–376, June 1973. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3212353.

McNickle:1974:NDS

[McN74] D. C. McNickle. The number of departures from a semi-Markov queue. Journal of Applied Probability, 11(4):825–828, December 1974. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3212566.

McNeil:1982:TSM

[McN82] Don McNeil. Time series modelling: An integrated approach. Journal of Applied Probability, 19A:413–425, 1982. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3213580.

McNamara:1985:OSP

[McN85] J. M. McNamara. An optimal sequential policy for controlling a Markov renewal process. Journal of Applied Probability, 22 (2):324–335, June 1985. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3213776.

McQuarrie:1967:SAC

[McQ67] Donald A. McQuarrie. Stochastic approach to chemical kinetics. Journal of Applied Probability, 4(3):413–478, December 1967. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/3212214. See

correction [Ano68d].

McVinish:2008:SER

[McV08] R. McVinish. On the structure and estimation of reflection positive processes. Journal of Applied Probability, 45(1):150– 162, March 2008. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
27595930.

Milligan:1996:MSA

[MD96] P. J. M. Milligan and D. Y. Downham. Models of superinfection and acquired immunity to multiple parasite strains. Journal of Applied Probability, 33(4):915–932, December 1996. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3214973.

Miao:2018:ABR

[MD18]  Yu Miao and Manru Dong.  Asymptotic behavior for

the Robbins–Monro process. Journal of Applied Probability, 55(2):559–570, June 2018. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL https:/

/www.cambridge.org/core/journals/journal-of-applied-
probability/article/asymptotic-behavior-for-the-robbinsmonro-
process/AA50550341560F4875470ED73A042305.

Misra:2009:OAA

[MDG09] Neeraj Misra, Ishwari D. Dhariyal, and Nitin Gupta. Optimal allocation of active spares in series systems and comparison of component and system redundancies. Journal of Applied Probability, 46(1):19–34, March 2009. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/25662405.

Meester:1995:ECD

[Mee95] Ronald Meester. Equality of critical densities in continuum percolation. Journal of Applied Probability, 32(1):90–104, March 1995. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3214923.

Mehr:1967:CCS

[Meh67] C. B. Mehr. Characterization of a class of second-order density functions. Journal of Applied Probability, 4(1):123–129, April 1967. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3212304.

Mehlmann:1977:MMM

[Meh77a] Alexander Mehlmann. Markovian manpower models in continuous time. Journal of Applied Probability, 14(2):249–259, June 1977. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3212996.

Mehlmann:1977:NLB

[Meh77b] Alexander Mehlmann. A note on the limiting behaviour of discrete-time Markovian manpower models with inhomogeneous independent Poisson input. Journal of Applied Probability, 14 (3):611–613, September 1977. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.

org/stable/3213465.

Mehlmann:1979:SMM

[Meh79] Alexander Mehlmann. Semi-Markovian manpower models in continuous time. Journal of Applied Probability, 16(2):416– 422, June 1979. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3212909.

Meilijson:1981:ELD

[Mei81] Isaac Meilijson. Estimation of the lifetime distribution of the parts from the autopsy statistics of the machine. Journal of Applied Probability, 18(4):829–838, December 1981. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3213058.

Meise:1999:SGE

[Mei99] Christian Meise. On spectral gap estimates of a Markov chain via hitting times and coupling. Journal of Applied Probability, 36(2):310–319, June 1999. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3215458.

Meiners:2010:ASR

[Mei10] Matthias Meiners. An almost-sure renewal theorem for branching random walks on the line. Journal of Applied Probability, 47 (3):811–825, September 2010. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.

org/stable/20778518.

Mejzler:1965:CPS

[Mej65] D. Mejzler. On a characteristic property of stationary streams without aftereffect. Journal of Applied Probability, 2(2):455–461, December 1965. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3212206.

Mejzler:1968:GEF

[Mej68] D. Mejzler. A generalisation of Erlang’s formulas in queueing theory. Journal of Applied Probability, 5(1):143–157, April 1968. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/3212082.

Mendelssohn:1976:OPA

[Men76] Roy Mendelssohn. Optimization problems associated with a Leslie matrix. American Naturalist, 110(973):339–349, May/ June 1976. CODEN AMNTA4. ISSN 0003-0147 (print),

1537-5323 (electronic). URL http://www.jstor.org/stable/

2459758. See comment [Gre86].

Mendelson:1982:BBP

[Men82] Haim Mendelson. A batch-ballot problem and applications. Journal of Applied Probability, 19(1):144–157, March 1982. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/3213924.

Meng:1996:MOA

[Men96] Fan C. Meng. More on optimal allocation of components in coherent systems. Journal of Applied Probability, 33(2):548– 556, June 1996. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3215078.

Menguturk:2023:TCR

[Men23] Levent Ali Mengütürk. Time-convergent random matrices from mean-field pinned interacting eigenvalues. Journal of Applied Probability, 60(2):394–417, June 2023. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL https:/

/www.cambridge.org/core/journals/journal-of-applied-
probability/article/timeconvergent-random-matrices-from-
meanfield-pinned-interacting-eigenvalues/9DB6E1656004BB8138E942EF1368460B.

Mergenthaler:1980:SDL

[Mer80] Wolfgang Mergenthaler. A state-dependent lifetime process of individuals subject to external perturbations. Journal of Applied Probability, 17(4):922–938, December 1980. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/3213202.

Metzler:2013:LTH

[Met13] Adam Metzler. The Laplace transform of hitting times of integrated geometric Brownian motion. Journal of Applied Probability, 50(1):295–299, March 2013. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http:/

/projecteuclid.org/euclid.jap/1363784440.

Matis:2001:TAS

[MF01] Timothy I. Matis and Richard M. Feldman. Transient analysis of state-dependent queueing networks via cumulant functions. Journal of Applied Probability, 38(4):841–859, December 2001. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/3215768. See

correction [Ano05e].

Marchal:1976:MEA

[MH76a] William G. Marchal and Carl M. Harris. A modified Erlang approach to approximating GI/G/1 queues. Journal of Applied Probability, 13(1):118–126, March 1976. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http:/

/www.jstor.org/stable/3212671.

Morgan:1976:AMW

[MH76b] Byron J. T. Morgan and John P. Hinde. On an approximation made when analysing stochastic processes. Journal of Applied Probability, 13(4):672–683, December 1976. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/3212522.

Menon:1983:ANN

[MI83] V. V. Menon and N. K. Indira. On the asymptotic normality of the number of replications of a paired comparison. Journal of Applied Probability, 20(3):554–562, September 1983. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3213892.

Mi:1994:MSP

[Mi94] Jie Mi. Maximization of a survival probability and its application. Journal of Applied Probability, 31(4):1026–1033, December 1994. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3215326.

Mi:1995:LBS

[Mi95] Jie Mi. Limiting behavior of some measures of system availability. Journal of Applied Probability, 32(2):482–493, June 1995. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/3215302.

Mi:1999:OAR

[Mi99] Jie Mi. Optimal active redundancy allocation in k-out-of-n system. Journal of Applied Probability, 36(3):927–933, September 1999. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3215452.

Mi:2002:ARP

[Mi02a] Jie Mi. Age-replacement policy and optimal work size. Journal of Applied Probability, 39(2):296–311, June 2002. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3216097.

Mi:2002:BSO

[Mi02b] Jie Mi. On bounds for some optimal policies in reliability. Journal of Applied Probability, 39(3):491–502, September 2002. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/3216073.

Miao:2013:ADO

[Mia13] Daniel Wei-Chung Miao. Analysis of the discrete Ornstein– Uhlenbeck process caused by the tick size effect. Journal of Applied Probability, 50(4):1102–1116, December 2013. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://projecteuclid.org/euclid.jap/1389370102.

Magdon-Ismail:2004:MDB

[MIAPAM04] Malik Magdon-Ismail, Amir F. Atiya, Amrit Pratap, and Yaser S. Abu-Mostafa. On the maximum drawdown of a Brownian motion. Journal of Applied Probability, 41(1):147–161, March 2004. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3215821.

Miller:1967:NPT

[Mil67] H. D. Miller. A note on passage times and infinitely divisible distributions. Journal of Applied Probability, 4(2):402–405, August 1967. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3212034.

Miller:1976:OSP

[Mil76] Douglas R. Miller. Order statistics, Poisson processes and repairable systems. Journal of Applied Probability, 13(3):519–529, September 1976. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3212471.

Miles:1979:SNI

[Mil79] R. E. Miles. Some new integral geometric formulae, with stochastic applications. Journal of Applied Probability, 16(3):592–606, September 1979. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3213087.

Miller:1981:CSA

[Mil81] Bruce L. Miller. Countable-state average-cost regenerative stopping problems. Journal of Applied Probability, 18(2):361– 377, June 1981. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3213283.

Milian:1992:TPD

[Mil92] A. Milian. Transition probability density of a certain diffusion process concentrated on a finite spatial interval. Journal of Applied Probability, 29(2):334–342, June 1992. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/3214570.

Miller:1994:MEA

[Mil94] James W. Miller. A matrix equation approach to solving recurrence relations in two-dimensional random walks. Journal of Applied Probability, 31(3):646–659, September 1994. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3215145.

Miller:2008:BSP

[Mil08] Joel C. Miller. Bounding the size and probability of epidemics on networks. Journal of Applied Probability, 45(2):498–512, June 2008. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/27595960.

Minh:1978:DTS

[Min78] Do Le Minh. The discrete-time single-server queue with timeinhomogeneous compound Poisson input and general service time distribution. Journal of Applied Probability, 15(3):590–601, September 1978. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3213121.

Minh:1979:AWT

[Min79] Do Le Minh. The actual waiting time of each customer in a GI/G/1 queue. Journal of Applied Probability, 16(4):910–916, December 1979. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3213157.

Mitrophanov:2003:SEC

[Mit03] A. Yu. Mitrophanov. Stability and exponential convergence of continuous-time Markov chains. Journal of Applied Probability, 40(4):970–979, December 2003. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.

org/stable/3216054.

Mitrophanov:2004:NZB

[Mit04a] A. Yu. Mitrophanov. Note on Zeifman’s bounds on the rate of convergence for birth-death processes. Journal of Applied Probability, 41(2):593–596, June 2004. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3216040. See [Zei95].

Mitrophanov:2004:SGP

[Mit04b] A. Yu. Mitrophanov. The spectral gap and perturbation bounds for reversible continuous-time Markov chains. Journal of Applied Probability, 41(4):1219–1222, December 2004. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/4141395.

Mitrophanov:2005:SCU

[Mit05] A. Yu. Mitrophanov. Sensitivity and convergence of uniformly ergodic Markov chains. Journal of Applied Probability, 42(4): 1003–1014, December 2005. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.

org/stable/30040824.

Miura:1977:SAR

[Miu77] M. Miura. Stochastic analysis of the raw silk reeling process. Journal of Applied Probability, 14(3):548–555, September 1977. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/3213457.

Miyazawa:1977:TCP

[Miy77] Masakiyo Miyazawa. Time and customer processes in queues with stationary inputs. Journal of Applied Probability, 14(2): 349–357, June 1977. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3213005.

Miyazawa:1979:FAQ

[Miy79] Masakiyo Miyazawa. A formal approach to queueing processes in the steady state and their applications. Journal of Applied Probability, 16(2):332–346, June 1979. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3212901.

Miyazawa:1982:SDI

[Miy82] Masakiyo Miyazawa. Simple derivations of the invariance relations and their applications. Journal of Applied Probability, 19 (1):183–194, March 1982. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3213927.

Miyazawa:1985:ICL

[Miy85] Masakiyo Miyazawa. The intensity conservation law for queues with randomly changed service rate. Journal of Applied Probability, 22(2):408–418, June 1985. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3213783. See comment [Bré91].

Miyazawa:1986:AQL

[Miy86] Masakiyo Miyazawa. Approximation of the queue-length distribution of an M/GI/s queue by the basic equations. Journal of Applied Probability, 23(2):443–458, June 1986. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3214186.

Miyazawa:1990:CGF

[Miy90] Masakiyo Miyazawa. Complementary generating functions for the M^{X} /GI/1/k and GI/M^{Y} /1/k queues and their application to the comparison of loss probabilities. Journal of Applied Probability, 27(3):684–692, September 1990. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/3214551.

Miyazawa:1992:SQL

[Miy92] Masakiyo Miyazawa. On the system queue length distributions of LCFS-P queues with arbitrary acceptance and restarting policies. Journal of Applied Probability, 29(2):430–440, June 1992. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/3214579.

Miyazawa:1994:TDR

[Miy94] Masakiyo Miyazawa. Time-dependent rate conservation laws for a process defined with a stationary marked point process and their applications. Journal of Applied Probability, 31(1):114– 129, March 1994. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3215239.

Miyazawa:1995:NGM

[Miy95] Masakiyo Miyazawa. Note on generalizations of Mecke’s formula and extensions of H = λG. Journal of Applied Probability, 32 (1):105–122, March 1995. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3214924.

Miyoshi:2001:SWD

[Miy01] Naoto Miyoshi. On the stationary workload distribution of workconserving single-server queues: A general formula via stochastic intensity. Journal of Applied Probability, 38(3):793–798, September 2001. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3216133.

Mizuno:2006:RBP

[Miz06] Toshihide Mizuno. A relation between positive dependence of signal and the variability of conditional expectation given signal. Journal of Applied Probability, 43(4):1181–1185, December 2006. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/27595809.

MacNeill:1985:RPN

[MJ85] Ian B. MacNeill and V. K. Jandhyala. The residual process for non-linear regression. Journal of Applied Probability, 22(4): 957–963, December 1985. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3213965.

Mode:1987:SPP

[MJP87] Charles J. Mode, Marc E. Jacobson, and Gary T. Pickens. A stochastic population projection system based on general agedependent branching processes. Journal of Applied Probability, 24(1):1–13, March 1987. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3214054.

Melnick:1973:AMD

[MK73] E. L. Melnick and S. Kullback. An application of the minimum discrimination information estimate to compute log-likelihood ratios. Journal of Applied Probability, 10(2):469–474, June 1973. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/3212367.

McDonald:2015:MIA

[MK15] Thomas O. McDonald and Marek Kimmel. A multitype infiniteallele branching process with applications to cancer evolution. Journal of Applied Probability, 52(3):864–876, September 2015. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://projecteuclid.org/euclid.jap/
1445543852.

McDonald:2017:MIA

[MK17] Thomas O. McDonald and Marek Kimmel. Multitype infiniteallele branching processes in continuous time. Journal of Applied Probability, 54(2):550–568, June 2017. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL https:/

/www.cambridge.org/core/journals/journal-of-applied-
probability/article/multitype-infiniteallele-branching-
processes-in-continuous-time/D434A4508FEB034F1148972D44485B4F.

Muhle-Karbe:2011:STA

[MKN11] Johannes Muhle-Karbe and Marcel Nutz. Small-time asymptotics of option prices and first absolute moments. Journal of Applied Probability, 48(4):1003–1020, December 2011. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/23066439.

Mazumdar:1991:RCN

[MKR91] Ravi Mazumdar, Raghavan Kannurpatti, and Catherine Rosenberg. On rate conservation for non-stationary processes. Journal of Applied Probability, 28(4):762–770, December 1991. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3214679.

McFadden:1967:MNI

[ML67] J. A. McFadden and James L. Lewis. Multivariate normal integrals for highly correlated samples from a Wiener process. Journal of Applied Probability, 4(2):303–312, August 1967. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3212025.

Morgan:1977:MBG

[ML77] Byron J. T. Morgan and B. Leventhal. A model for blue-green algae and gorillas. Journal of Applied Probability, 14(4):675–688, December 1977. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3213342.

Martin-Lof:1985:LTW

[ML85] Anders Martin-Löf. A limit theorem which clarifies the ’petersburg paradox’. Journal of Applied Probability, 22(3):634–643, September 1985. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3213866.

Martin-Lof:1986:SSP

[ML86] Anders Martin-Löf. Symmetric sampling procedures, general epidemic processes and their threshold limit theorems. Journal of Applied Probability, 23(2):265–282, June 1986. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3214172.

Martin-Lof:1998:FSN

[ML98] Anders Martin-Löf. The final size of a nearly critical epidemic, and the first passage time of a Wiener process to a parabolic barrier. Journal of Applied Probability, 35(3):671–682, September 1998. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3215642.

Mitrophanov:2005:SHM

[MLB05] Alexander Yu. Mitrophanov, Alexandre Lomsadze, and Mark Borodovsky. Sensitivity of hidden Markov models. Journal of Applied Probability, 42(3):632–642, September 2005. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/30040846.

Malmros:2016:RDS

[MLB16] J. Malmros, F. Liljeros, and T. Britton. Respondent-driven sampling and an unusual epidemic. Journal of Applied Probability, 53(2):518–530, June 2016. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://projecteuclid.

org/euclid.jap/1466172871.

Marliss:1971:CLD

[MM71] G. S. Marliss and J. R. McGregor. The construction of limiting distributions of response probabilities. Journal of Applied Probability, 8(4):757–766, December 1971. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/3212239.

Miles:1982:BSV

[MM82] R. E. Miles and R. J. Maillardet. The basic structures of Voronoi and generalized Voronoi polygons. Journal of Applied Probability, 19A:97–111, 1982. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3213553.

Merad:1994:OFR

[MM94] S. Merad and J. M. McNamara. Optimal foraging of a reproducing animal as a discounted reward problem. Journal of Applied Probability, 31(2):287–300, June 1994. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/3215024.

Marckert:2003:LVI

[MM03a] Jean-Franiçois Marckert and Abdelkader Mokkadem. Ladder variables, internal structure of Galton–Watson trees and finite branching random walks. Journal of Applied Probability, 40(3): 671–689, September 2003. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3215943.

Muller:2003:FTC

[MM03b] Johannes Müller and Martin Möhle. Family trees of continuoustime birth-and-death processes. Journal of Applied Probability, 40(4):980–994, December 2003. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.

org/stable/3216055.

Mallows:2006:SS

[MM06] Colin L. Mallows and Jean Meloche. Searching for searchers. Journal of Applied Probability, 43(3):899–904, September 2006. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/27595783.

Ma:2009:TSB

[MM09] S. Ma and M. Molina. Two-sex branching processes with offspring and mating in a random environment. Journal of Applied Probability, 46(4):993–1004, December 2009. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/25662476.

Matsui:2010:PPC

[MM10] Muneya Matsui and Thomas Mikosch. Prediction in a Poisson cluster model. Journal of Applied Probability, 47(2):350– 366, June 2010. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
20778484.

Mairesse:2016:SSM

[MM16] Jean Mairesse and Pascal Moyal. Stability of the stochastic matching model. Journal of Applied Probability, 53(4):1064– 1077, December 2016. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://projecteuclid.

org/euclid.jap/1481132837.

Misra:2011:SRA

[MMD11] Neeraj Misra, Amit Kumar Misra, and Ishwari Dutt Dhariyal. Standby redundancy allocations in series and parallel systems. Journal of Applied Probability, 48(1):43–55, March 2011. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/29777445.

Molina:2002:BGW

[MMR02] M. Molina, M. Mota, and A. Ramos. Bisexual Galton–Watson branching process with population-size-dependent mating. Journal of Applied Probability, 39(3):479–490, September 2002. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/3216072.

Molina:2007:SCT

[MMR07] M. Molina, M. Mota, and A. Ramos. Some contributions to the theory of near-critical bisexual branching processes. Journal of Applied Probability, 44(2):492–505, June 2007. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/27595856.

Meilijson:1979:CMA

[MN79] Isaac Meilijson and Arthur Nádas. Convex majorization with an application to the length of critical paths. Journal of Applied Probability, 16(3):671–677, September 1979. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/3213097.

Murthy:1982:NEB

[MN82] D. N. P. Murthy and D. G. Nguyen. A note on extended block replacement policy with used items. Journal of Applied Probability, 19(4):885–889, December 1982. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3213845.

Matsumoto:2013:LTG

[MN13] Keisuke Matsumoto and Toshio Nakata. Limit theorems for a generalized Feller game. Journal of Applied Probability, 50 (1):54–63, March 2013. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://projecteuclid.

org/euclid.jap/1363784424.

Mohle:2016:ECT

[MN16] Martin Möhle and Morihiro Notohara. An extension of a convergence theorem for Markov chains arising in population genetics. Journal of Applied Probability, 53(3):953–956, September 2016. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-

6072 (electronic). URL http://projecteuclid.org/euclid.
jap/1476370790.

Mercankosk:2001:CTM

[MNS01] G. Mercankosk, G. M. Nair, and W. J. Soet. Combinatorial techniques for M/G/1-type queues. Journal of Applied Probability, 38(3):722–736, September 2001. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3216124.

Marshall:1967:GBE

[MO67] Albert W. Marshall and Ingram Olkin. A generalized bivariate exponential distribution. Journal of Applied Probability, 4(2): 291–302, August 1967. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3212024.

Mine:1970:LDI

[MO70] Hisashi Mine and Katsuhisa Ohno. Low density inhomogeneous traffic flow. Journal of Applied Probability, 7(3):573–598, December 1970. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3211939.

Mine:1971:TLQ

[MO71] Hisashi Mine and Katsuhisa Ohno. Traffic light queues as a generalization to queueing theory. Journal of Applied Probability, 8(3):480–493, September 1971. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3212172.

Maier:1992:CCP

[MO92] Robert S. Maier and Colm Art O’Cinneide. A closure characterisation of phase-type distributions. Journal of Applied Probability, 29(1):92–103, March 1992. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3214794.

Marshall:1993:BLD

[MO93] Albert W. Marshall and Ingram Olkin. Bivariate life distributions from Pólya’s urn model for contagion. Journal of Applied Probability, 30(3):497–508, September 1993. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/3214760.

Mallor:2001:SRR

[MO01] F. Mallor and E. Omey. Shocks, runs and random sums. Journal of Applied Probability, 38(2):438–448, June 2001. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3215898.

Moller:2021:CDP

[MO21] Jesper Møller and Eliza O’Reilly. Couplings for determinantal point processes and their reduced Palm distributions with a view to quantifying repulsiveness. Journal of Applied Probability, 58(2):469–483, June 2021. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL https:/

/www.cambridge.org/core/journals/journal-of-applied-
probability/article/couplings-for-determinantal-point-
processes-and-their-reduced-palm-distributions-with-
a-view-to-quantifying-repulsiveness/35162EE3993745A01096A285410E3ED0.

Martirosyan:2023:IPF

[MO23] Davit Martirosyan and Victor Ohanyan. On intersection probabilities of four lines inside a planar convex domain. Journal of Applied Probability, 60(2):504–527, June 2023. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL https://www.cambridge.org/core/journals/journal-
of-applied-probability/article/on-intersection-probabilities-
of-four-lines-inside-a-planar-convex-domain/61EC12A8E8E34A0EB1548EED2D9F24BE.

Mode:1966:SCA

[Mod66] Charles J. Mode. A stochastic calculus and its application to some fundamental theorems of natural selection. Journal of Applied Probability, 3(2):327–352, December 1966. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/3212123. See corrections

[Mod67b].

Mode:1967:RDT

[Mod67a] Charles J. Mode. A renewal density theorem in the multidimensional case. Journal of Applied Probability, 4(1):62–76, April 1967. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3212299.

Mode:1967:SCM

[Mod67b] Charles J. Mode. Some corrections to my paper “A Stochastic Calculus and Its Application to Some Fundamental Theorems of Natural Selection”. Journal of Applied Probability, 4(3):609–611, December 1967. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/

3212225. See [Mod66].

Mohanty:1972:QIB

[Moh72] S. G. Mohanty. On queues involving batches. Journal of Applied Probability, 9(2):430–435, June 1972. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3212812.

Mohan:1976:LDN

[Moh76] N. R. Mohan. Limit distributions of the number of renewals and waiting times. Journal of Applied Probability, 13(2):301– 312, June 1976. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3212833.

Mohle:1994:FBP

[Möh94] M. Möhle. Forward and backward processes in bisexual models with fixed population sizes. Journal of Applied Probability, 31 (2):309–332, June 1994. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3215026.

Mohle:1997:FBM

[Möh97] M. Möhle. Fixation in bisexual models with variable population sizes. Journal of Applied Probability, 34(2):436–448, June 1997. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3215383.

Mohle:1998:RRC

[Möh98] M. Möhle. Robustness results for the coalescent. Journal of Applied Probability, 35(2):438–447, June 1998. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/3215697.

Mohle:1999:WCC

[Möh99] M. Möhle. Weak convergence to the coalescent in neutral population models. Journal of Applied Probability, 36(2):446–460, June 1999. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3215467.

Mohle:2005:CRC

[Möh05] M. Möhle. Convergence results for compound Poisson distributions and applications to the standard Luria–Delbrück distribution. Journal of Applied Probability, 42(3):620–631, September 2005. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/30040845.

Mohle:2010:LFB

[Möh10] M. Möhle. Looking forwards and backwards in the multi-allelic neutral Cannings population model. Journal of Applied Probability, 47(3):713–731, September 2010. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/20778512.

Mohle:2021:RPC

[Möh21] Martin Möhle. A restaurant process with cocktail bar and relations to the three-parameter Mittag-Leffler distribution. Journal of Applied Probability, 58(4):978–1006, December 2021. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL https://www.cambridge.org/core/journals/journal-
of-applied-probability/article/restaurant-process-with-
cocktail-bar-and-relations-to-the-threeparameter-mittagleffler-
distribution/319060AF1866F56C8C6B651146A3E0B5.

Molter:1986:TMS

[Mol86] Ursula Maria Molter. Tangential measure on the set of convex infinite cylinders. Journal of Applied Probability, 23(4):961–972, December 1986. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3214469.

Moller:1988:SAP

[Møl88] J. Møller. Stereological analysis of particles of varying ellipsoidal shape. Journal of Applied Probability, 25(2):322–335, June 1988. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/3214440.

Moller:1995:SDE

[Møl95] Christian Max Møller. Stochastic differential equations for ruin probabilities. Journal of Applied Probability, 32(1):74–89, March 1995. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3214922.

Moller:1996:IEC

[Møl96] Christian Max Møller. Integral equations for compound distribution functions. Journal of Applied Probability, 33(2):388– 399, June 1996. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3215062.

Montroll:1967:NPI

[Mon67] Elliott W. Montroll. On nonlinear processes involving population growth and diffusion. Journal of Applied Probability, 4(2): 281–290, August 1967. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3212023.

Monahan:1982:OSP

[Mon82] George E. Monahan. Optimal stopping in a partially observable binary-valued Markov chain with costly perfect information. Journal of Applied Probability, 19(1):72–81, March 1982. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/3213917.

Moore:1973:NSR

[Moo73] Marc Moore. Non-stationary random set processes with application to pattern reconstruction. Journal of Applied Probability, 10(4):857–863, December 1973. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.

org/stable/3212387.

Moore:1974:ARM

[Moo74] Marc Moore. Anisotropically random mosaics. Journal of Applied Probability, 11(2):374–376, June 1974. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/3212759.

Moran:1966:NRR

[Mor66a] P. A. P. Moran. A note on recent research in geometrical probability. Journal of Applied Probability, 3(2):453–463, December 1966. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3212131.

Morton:1966:ENA

[Mor66b] R. R. A. Morton. The expected number and angle of intersections between random curves in a plane. Journal of Applied Probability, 3(2):559–562, December 1966. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/3212140.

Moran:1967:DSC

[Mor67] P. A. P. Moran. Dams in series with continuous release. Journal of Applied Probability, 4(2):380–388, August 1967. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3212031.

Morton:1968:DPA

[Mor68] Richard Morton. On the dynamic programming approach to Pontriagin’s maximum principle. Journal of Applied Probability, 5(3):679–692, December 1968. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.

org/stable/3211930.

Moran:1969:TDC

[Mor69] P. A. P. Moran. A theory of dams with continuous input and a general release rule. Journal of Applied Probability, 6(1):88– 98, April 1969. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3212278.

Morgan:1971:SDE

[Mor71a] Byron J. T. Morgan. On the solution of differential equations arising in some attachment models of virology. Journal of Applied Probability, 8(2):215–221, June 1971. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/3211891.

Morton:1971:OCSa

[Mor71b] R. Morton. On the optimal control of stationary diffusion processes with inaccessible boundaries and no discounting. Journal of Applied Probability, 8(3):551–560, September 1971. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3212178.

Morton:1971:OCSb

[Mor71c] R. Morton. On the optimal control of stationary diffusion processes with natural boundaries and discounted cost. Journal of Applied Probability, 8(3):561–572, September 1971. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3212179.

Moran:1973:CLT

[Mor73a] P. A. P. Moran. A central limit theorem for exchangeable variates with geometric applications. Journal of Applied Probability, 10(4):837–846, December 1973. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.

org/stable/3212385.

Moran:1973:GMP

[Mor73b] P. A. P. Moran. A Gaussian Markovian process on a square lattice. Journal of Applied Probability, 10(1):54–62, March 1973. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/3212495.

Moran:1973:NCM

[Mor73c] P. A. P. Moran. Necessary conditions for Markovian processes on a lattice. Journal of Applied Probability, 10(3):605–612, September 1973. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3212780.

Moran:1973:OYV

[Mor73d] P. A. P. Moran. Obituary: Yu. V. Linnik. Journal of Applied Probability, 10(1):245–246, March 1973. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http:/

/www.jstor.org/stable/3212515.

Moran:1973:RVI

[Mor73e] P. A. P. Moran. The random volume of interpenetrating spheres in space. Journal of Applied Probability, 10(2):483–490, June 1973. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3212370.

Morgan:1974:DIM

[Mor74] Byron J. T. Morgan. On the distribution of inanimate marks over a linear birth-and-death process. Journal of Applied Probability, 11(3):423–436, September 1974. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3212687. See correction [Ano74f].

Mori:1975:EIR

[Mor75] Toshio Mori. Ergodicity and identifiability for random translations of stationary point processes. Journal of Applied Probability, 12(4):734–743, December 1975. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3212724.

Morrison:1978:LIM

[Mor78] Donald G. Morrison. On linearly increasing mean residual lifetimes. Journal of Applied Probability, 15(3):617–620, September 1978. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3213124.

Moran:1979:AAL

[Mor79] P. A. P. Moran. Asymptotic analysis and local lattice central limit theorems. Journal of Applied Probability, 16(3):541–553, September 1979. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3213083.

Mori:1980:RBQ

[Mor80] Masao Mori. Relations between queue-size and waiting-time distributions. Journal of Applied Probability, 17(3):822–830, September 1980. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3212975.

Moran:1986:OPG

[Mor86] P. A. P. Moran. Orthant probabilities and Gaussian Markov processes. Journal of Applied Probability, 23A:413–417, 1986. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/3214369.

Morgan:1993:ESD

[Mor93] Byron J. T. Morgan. Expected size distributions in models of group dynamics. Journal of Applied Probability, 30(1):1– 16, March 1993. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3214617.

Mori:1996:BID

[Mór96] Tamás F. Móri. Bonferroni inequalities and deviations of discrete distributions. Journal of Applied Probability, 33(1):115– 121, March 1996. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3215268.

Morrison:2006:UMA

[Mor06] John A. Morrison. An urn model arising from all-optical networks. Journal of Applied Probability, 43(4):952–966, December 2006. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/27595792.

Mostovyi:2018:OCM

[Mos18] Oleksii Mostovyi. Optimal consumption of multiple goods in incomplete markets. Journal of Applied Probability, 55 (3):810–822, September 2018. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL https:/

/www.cambridge.org/core/journals/journal-of-applied-
probability/article/optimal-consumption-of-multiple-
goods-in-incomplete-markets/5B2165FF9FC9E3213B08F425AC836554.

Mountford:1993:CFP

[Mou93] T. S. Mountford. A coupling of finite particle systems. Journal of Applied Probability, 30(1):258–262, March 1993. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3214638.

Moustakides:1999:EWF

[Mou99] George V. Moustakides. Extension of Wald’s first lemma to Markov processes. Journal of Applied Probability, 36(1):48– 59, March 1999. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3215401.

Moyal:1964:MPP

[Moy64] J. E. Moyal. Multiplicative population processes. Journal of Applied Probability, 1(2):267–283, December 1964. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3211859.

Moyal:1965:IDM

[Moy65] J. E. Moyal. Incomplete discontinuous Markov processes. Journal of Applied Probability, 2(1):69–78, June 1965. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3211874.

Moyal:2008:SPS

[Moy08] Pascal Moyal. Stability of a processor-sharing queue with varying throughput. Journal of Applied Probability, 45(4):953–962, December 2008. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
27596001.

Moyal:2010:QIC

[Moy10] Pascal Moyal. The queue with impatience: Construction of the stationary workload under fifo. Journal of Applied Probability, 47(2):498–512, June 2010. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/20778493.

Marshall:1970:MLS

[MP70] Albert W. Marshall and Frank Proschan. Mean life of series and parallel systems. Journal of Applied Probability, 7(1):165– 174, April 1970. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3212157.

Makowski:1992:SCS

[MP92a] Armand M. Makowski and Thomas K. Philips. Stochastic convexity of sums of I.I.D. non-negative random variables with applications. Journal of Applied Probability, 29(1):156–167, March 1992. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3214799.

McCormick:1992:AAE

[MP92b] William P. McCormick and You Sung Park. Asymptotic analysis of extremes from autoregressive negative binomial processes. Journal of Applied Probability, 29(4):904–920, December 1992. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/3214723.

Mountford:1993:RSI

[MP93] T. S. Mountford and S. C. Port. Random splittings of an interval. Journal of Applied Probability, 30(1):131–152, March 1993. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/3214627.

Machado:2000:ODB

[MP00] F. P. Machado and S. Yu. Popov. One-dimensional branching random walks in a Markovian random environment. Journal of Applied Probability, 37(4):1157–1163, December 2000. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3215507.

Morais:2004:NAC

[MP04] Manuel Cabral Morais and António Pacheco. A note on the ageing character of the run length of Markov-type quality-control schemes. Journal of Applied Probability, 41(4):1243–1247, December 2004. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
4141399.

McVinish:2012:CLT

[MP12] R. McVinish and P. K. Pollett. A central limit theorem for a discrete-time SIS model with individual variation. Journal of Applied Probability, 49(2):521–530, June 2012. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/41713783.

McVinish:2014:LBH

[MP14] R. McVinish and P. K. Pollett. The limiting behaviour of Hanski’s incidence function metapopulation model. Journal of Applied Probability, 51(2):297–316, June 2014. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

projecteuclid.org/euclid.jap/1402578626.

Mendoza-Perez:2010:AND

[MPHL10] Armando F. Mendoza-Pérez and Onésimo Hernández-Lerma. Asymptotic normality of discrete-time Markov control processes. Journal of Applied Probability, 47(3):778–795, September 2010. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/20778516.

Mijatovic:2015:RRM

[MPS15] Aleksandar Mijatović, Martijn R. Pistorius, and Johannes Stolte. Randomisation and recursion methods for mixedexponential Lévy models, with financial applications. Journal of Applied Probability, 52(4):1076–1096, December 2015. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://projecteuclid.org/euclid.jap/1450802754.

Mangoubi:2021:SCM

[MPS21] Oren Mangoubi, Natesh Pillai, and Aaron Smith. Simple conditions for metastability of continuous Markov chains. Journal of Applied Probability, 58(1):83–105, March 2021. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL https://www.cambridge.org/core/journals/journal-
of-applied-probability/article/simple-conditions-for-
metastability-of-continuous-markov-chains/6EF80575C90D4DF6288943EAD4883D84.

Munoz-Perez:1990:DOD

[MPSG90] J. Muñoz-Perez and A. Sanchez-Gomez. Dispersive ordering by dilation. Journal of Applied Probability, 27(2):440–444, June 1990. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3214664.

McCormick:2000:ADS

[MQ00] W. P. McCormick and Y. Qi. Asymptotic distribution for the sum and maximum of Gaussian processes. Journal of Applied Probability, 37(4):958–971, December 2000. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/3215487.

Muller:2001:OSV

[MR01] Alfred Müller and Ludger Rüschendorf. On the optimal stopping values induced by general dependence structures. Journal of Applied Probability, 38(3):672–684, September 2001. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3216120.

Mohan:2002:SPT

[MR02] N. R. Mohan and S. Ravi. Some properties of two families of classes of life distributions. Journal of Applied Probability, 39(3): 581–589, September 2002. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3216080.

Mercier:2003:AFR

[MR03] Sophie Mercier and Michel Roussignol. Asymptotic failure rate of a Markov deteriorating system with preventive maintenance. Journal of Applied Probability, 40(1):1–19, March 2003. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3215838.

Munsonius:2011:LTD

[MR11] Götz Olaf Munsonius and Ludger Rüschendorf. Limit theorems for depths and distances in weighted random b-ary recursive trees. Journal of Applied Probability, 48(4):1060–1080, December 2011. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/23066443.

Miziula:2015:EDS

[MR15] Patryk Miziula and Tomasz Rychlik. Extreme dispersions of semicoherent and mixed system lifetimes. Journal of Applied Probability, 52(1):117–128, March 2015. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http:/

/projecteuclid.org/euclid.jap/1429282610.

Malinovsky:2023:TND

[MR23] Yaakov Malinovsky and Yosef Rinott. On tournaments and negative dependence. Journal of Applied Probability, 60 (3):945–954, September 2023. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL https:/

/www.cambridge.org/core/journals/journal-of-applied-
probability/article/on-tournaments-and-negative-dependence/
A58CEE28D991F16DD6766E7C9664225E.

MacPhee:2004:SAT

[MRP04] I. M. MacPhee, Jonathan Rougier, and G. H. Pollard. Server advantage in tennis matches. Journal of Applied Probability, 41 (4):1182–1186, December 2004. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.

org/stable/4141389.

Maulik:2002:AIN

[MRR02] Krishanu Maulik, Sidney Resnick, and Holger Rootzén. Asymptotic independence and a network traffic model. Journal of Applied Probability, 39(4):671–699, December 2002. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3215996.

Mocquard:2021:SAA

[MRSA21] Yves Mocquard, Frédérique Robin, Bruno Séricola, and Emmanuelle Anceaume. Stochastic analysis of averagebased distributed algorithms. Journal of Applied Probability, 58(2):394–410, June 2021. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL https:/

/www.cambridge.org/core/journals/journal-of-applied-
probability/article/stochastic-analysis-of-averagebased-
distributed-algorithms/5471E18EB73AE2D9328DDC86FDFAACFF.

Moore:1979:BPS

[MS79] T. Moore and J. L. Snell. A branching process showing a phase transition. Journal of Applied Probability, 16(2):252–260, June 1979. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3212894.

Mori:1985:NBS

[MS85] Tamás F. Móri and Gábor J. Székely. A note on the background of several Bonferroni–Galambos-type inequalities. Journal of Applied Probability, 22(4):836–843, December 1985. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3213951.

Murphree:1986:TRP

[MS86] Emily Murphree and Walter L. Smith. On transient regenerative processes. Journal of Applied Probability, 23(1):52–70, March 1986. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3214116.

Masuda:1991:MRP

[MS91] Yasushi Masuda and Ushio Sumita. A multivariate reward process defined on a semi-Markov process and its first-passage-time distributions. Journal of Applied Probability, 28(2):360– 373, June 1991. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3214872.

Muche:1992:CCL

[MS92] Lutz Muche and Dietrich Stoyan. Contact and chord length distributions of the Poisson Voronoi tessellation. Journal of Applied Probability, 29(2):467–471, June 1992. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/3214584. See correction [CW91].

McCormick:1993:SMD

[MS93] William P. McCormick and Jiayang Sun. Sums and maxima of discrete stationary processes. Journal of Applied Probability, 30 (4):863–876, December 1993. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor. org/stable/3214518. See correction [Ano94f].

Martinez:1994:QSD

[MS94] Servet Martinez and Jaime San Martín. Quasi-stationary distributions for a Brownian motion with drift and associated limit laws. Journal of Applied Probability, 31(4):911–920, December 1994. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3215316.

Massoulie:1999:LBA

[MS99] Laurent Massoulie and Alain Simonian. Large buffer asymptotics for the queue with fractional Brownian input. Journal of Applied Probability, 36(3):894–906, September 1999. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3215449.

Mazalov:2002:HSP

[MS02] Vladimir V. Mazalov and Vesa Saario. The house-selling problem with reward rate criterion. Journal of Applied Probability, 39(3):644–649, September 2002. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.

org/stable/3216086.

Meerschaert:2004:LTC

[MS04a] Mark M. Meerschaert and Hans-Peter Scheffler. Limit theorems for continuous-time random walks with infinite mean waiting times. Journal of Applied Probability, 41(3):623–638, September 2004. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/4141342.

Morters:2004:MSB

[MS04b] Peter Mörters and Narn-Rueih Shieh. On the multifractal spectrum of the branching measure on a Galton–Watson tree. Journal of Applied Probability, 41(4):1223–1229, December 2004. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/4141396.

Mallows:2005:S

[MS05a] Colin Mallows and Larry Shepp. B-stability. Journal of Applied Probability, 42(2):581–586, June 2005. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/30040811.

Mountford:2005:NBR

[MS05b] Thomas Mountford and Rinaldo B. Schinazi. A note on branching random walks on finite sets. Journal of Applied Probability, 42(1):287–294, March 2005. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/30040788.

Macci:2006:LDR

[MS06] Claudio Macci and Gabriele Stabile. Large deviations for risk processes with reinsurance. Journal of Applied Probability, 43(3): 713–728, September 2006. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/27595767.

Mallows:2008:NP

[MS08] Colin Mallows and Larry Shepp. The necklace process. Journal of Applied Probability, 45(1):271–278, March 2008. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/27595939.

Mossel:2010:BPA

[MS10] Elchanan Mossel and Arnab Sen. Branching process approach for 2-SAT thresholds. Journal of Applied Probability, 47(3): 796–810, September 2010. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/20778517.

Mountford:2020:FDG

[MS20] Thomas Mountford and Jacques Saliba. Flooding and diameter in general weighted random graphs. Journal of Applied Probability, 57(3):956–980, September 2020. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL https://www.cambridge.org/core/journals/
journal-of-applied-probability/article/flooding-and-
diameter-in-general-weighted-random-graphs/034666BE14FFFE21525797D622C718E1.

Meitz:2021:SEM

[MS21] Mika Meitz and Pentti Saikkonen. Subgeometric ergodicity and β-mixing. Journal of Applied Probability, 58 (3):594–608, September 2021. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL https:/

/www.cambridge.org/core/journals/journal-of-applied-

probability/article/subgeometric-ergodicity-and-mixing/
D75C1D4A0175B88ACDEF292C98BDC16B.

McKeague:2023:SMA

[MS23] Ian W. McKeague and Yvik Swan. Stein’s method and approximating the multidimensional quantum harmonic oscillator. Journal of Applied Probability, 60(3):855–873, September 2023. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL https://www.cambridge.org/core/journals/
journal-of-applied-probability/article/steins-method-
and-approximating-the-multidimensional-quantum-harmonic-
oscillator/32C49CDD84B2738B53F7438BD60BDD57.

Mertens:1978:NSC

[MSCZ78] Jean-François Mertens, Ester Samuel-Cahn, and Shmuel Zamir. Necessary and sufficient conditions for recurrence and transience of Markov chains, in terms of inequalities. Journal of Applied Probability, 15(4):848–851, December 1978. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/3213440.

Ma:1992:ALD

[MSS92] W. T. Ma, G. Vh. Sandri, and S. Sarkar. Analysis of the Luria– Delbrück distribution using discrete convolution powers. Journal of Applied Probability, 29(2):255–267, June 1992. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3214564.

Mendivil:2009:GCG

[MSS09] F. Mendivil, R. Shonkwiler, and M. C. Spruill. Geometric convergence of genetic algorithms under tempered random restart. Journal of Applied Probability, 46(4):960–974, December 2009. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/25662474.

Mikosch:2013:FMS

[MST13] Thomas Mikosch, Gennady Samorodnitsky, and Laleh Tafakori. Fractional moments of solutions to stochastic recurrence equations. Journal of Applied Probability, 50(4):969–982, December 2013. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://projecteuclid.org/euclid.jap/
1389370094.

Martinez:2014:EUQ

[MSV14] Servet Martínez, Jaime San Martín, and Denis Villemonais. Existence and uniqueness of a quasistationary distribution for Markov processes with fast return from infinity. Journal of Applied Probability, 51(3):756–768, September 2014. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://projecteuclid.org/euclid.jap/1409932672.

Morrow:1994:CCP

[MSZ94] Gregory J. Morrow, Rinaldo B. Schinazi, and Yu Zhang. The critical contact process on a homogeneous tree. Journal of Applied Probability, 31(1):250–255, March 1994. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3215251.

Molchanov:2004:CGS

[MSZ04] Ilya Molchanov, Vadim Shcherbakov, and Sergei Zuyev. Critical growth of a semi-linear process. Journal of Applied Probability, 41(2):355–367, June 2004. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3216021.

Mine:1970:LPC

[MT70] Hisashi Mine and Yoshio Tabata. Linear programming and continuous Markovian decision problems. Journal of Applied Probability, 7(3):657–666, December 1970. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3211945.

Marlow:1995:SGC

[MT95] Norman A. Marlow and Michael Tortorella. Some general characteristics of two-state reliability models for maintained systems. Journal of Applied Probability, 32(3):805–820, September 1995. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/3215131.

McDonald:2000:CLP

[MT00] David McDonald and François Théberge. Cell loss probability for M/G/1 and time-slotted queues. Journal of Applied Probability, 37(4):1149–1156, December 2000. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/3215506.

Miyazawa:2002:MEF

[MT02] Masakiyo Miyazawa and Hiroyuki Takada. A matrix exponential form for hitting probabilities and its application to a Markovmodulated fluid queue with downward jumps. Journal of Applied Probability, 39(3):604–618, September 2002. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/3216082.

Mazalov:2006:EFO

[MT06] Vladimir V. Mazalov and Mitsushi Tamaki. An explicit formula for the optimal gain in the full-information problem of owning a relatively best object. Journal of Applied Probability, 43(1):87– 101, March 2006. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
27595712.

Martin:2012:OPL

[MT12] Ryan Martin and Omkar Tilak. On ϵ-optimality of the pursuit learning algorithm. Journal of Applied Probability, 49(3): 795–805, September 2012. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/41713806.

Manuceau:1999:ECP

[MTV99] Jérôme Manuceau, Marylène Troupé, and Jean Vaillant. On an entropy conservation principle. Journal of Applied Probability, 36(2):607–610, June 1999. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3215479.

Maehara:2000:WTJ

[MU00] Hiroshi Maehara and Sumie Ueda. On the waiting time in a janken game. Journal of Applied Probability, 37(2):601– 605, June 2000. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3215733.

Mijatovic:2011:NPW

[MU11] Aleksandar Mijatović and Mikhail Urusov. A note on a paper by Wong and Heyde. Journal of Applied Probability, 48(3):811–819, September 2011. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/

23065938. See [WH04, Won11].

Muciek:2002:OSR

[Muc02] Bogdan Krzysztof Muciek. Optimal stopping of a risk process: Model with interest rates. Journal of Applied Probability, 39 (2):261–270, June 2002. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3216094.

Mullikin:1968:SPD

[Mul68a] T. W. Mullikin. Some probability distributions for neutron transport in a half-space. Journal of Applied Probability, 5(2): 357–374, August 1968. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3212258.

Mullooly:1968:ODR

[Mul68b] John P. Mullooly. A one dimensional random space-filling problem. Journal of Applied Probability, 5(2):427–435, August 1968. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/3212263.

Mullooly:1972:MLE

[Mul72] John P. Mullooly. Maximum likelihood estimation for stochastic R th-order reactions. Journal of Applied Probability, 9(1):32– 42, March 1972. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3212634.

Mullooly:1973:MLE

[Mul73] John P. Mullooly. Maximum likelihood estimation for stochastic r th-order reactions. II. Journal of Applied Probability, 10 (2):441–446, June 1973. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3212361.

Munsonius:2012:TBR

[Mun12] Götz Olaf Munsonius. On tail bounds for random recursive trees. Journal of Applied Probability, 49(2):566–581, June 2012. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/41713786.

Murthy:1980:CAT

[Mur80] K. Rama Murthy. Convergence of age and type distributions in multitype critical Bellman–Harris processes. Journal of Applied Probability, 17(4):948–955, December 1980. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/3213204.

Murphree:1987:TRP

[Mur87] Emily S. Murphree. Transient renewal processes in the subexponential case. Journal of Applied Probability, 24(1):88–96, March 1987. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3214061.

Murphree:1989:SNR

[Mur89] Emily S. Murphree. Some new results on the subexponential class. Journal of Applied Probability, 26(4):892–897, December 1989. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3214395.

Muselli:2000:NIB

[Mus00] Marco Muselli. New improved bounds for reliability of consecutive-k-out-of-n:f systems. Journal of Applied Probability, 37(4):1164–1170, December 2000. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3215508.

Mutafchiev:1987:RMS

[Mut87] Ljuben R. Mutafchiev. On random mappings with a single attracting centre. Journal of Applied Probability, 24(1):258– 264, March 1987. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3214078.

Mellein:1986:KMP

[MV86] B. Mellein and J. L. Vicente. On the kinetics of multitype polymer reactions. Journal of Applied Probability, 23(1):14– 27, March 1986. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3214113.

Martinez:1995:MCA

[MV95] Servet Martínez and Maria Eulália Vares. A Markov chain associated with the minimal quasi-stationary distribution of birthdeath chains. Journal of Applied Probability, 32(1):25–38, March 1995. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3214918.

Maheshwari:2016:LRD

[MV16] A. Maheshwari and P. Vellaisamy. On the long-range dependence of fractional Poisson and negative binomial processes. Journal of Applied Probability, 53(4):989–1000, December 2016. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://projecteuclid.org/euclid.jap/1481132831.

Miclo:2021:FAS

[MV21]  Laurent Miclo and Stéphane Villeneuve.  On the for-

ward algorithm for stopping problems on continuous-time Markov chains. Journal of Applied Probability, 58(4): 1043–1063, December 2021. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL https:/

/www.cambridge.org/core/journals/journal-of-applied-
probability/article/on-the-forward-algorithm-for-stopping-
problems-on-continuoustime-markov-chains/C5138583716F17561C16B960F541C2E0.

Marshall:1971:CAT

[MW71] Kneale T. Marshall and Ronald W. Wolff. Customer average and time average queue lengths and waiting times. Journal of Applied Probability, 8(3):535–542, September 1971. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3212176.

Melamed:1990:AST

[MW90a] Benjamin Melamed and Ward Whitt. On arrivals that see time averages: A martingale approach. Journal of Applied Probability, 27(2):376–384, June 1990. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.

org/stable/3214656.

Miyazawa:1990:FRA

[MW90b] Masakiyo Miyazawa and Ronald W. Wolff. Further results on ASTA for general stationary processes and related problems. Journal of Applied Probability, 27(4):792–804, December 1990. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/3214823.

Mortimer:1991:CMR

[MW91] T. M. Mortimer and David Williams. Change of measure up to a random time: Theory. Journal of Applied Probability, 28(4): 914–918, December 1991. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3214696.

McBeth:1996:FTO

[MW96] Douglas W. McBeth and Ananda P. N. Weerasinghe. Finitetime optimal control of a process leaving an interval. Journal of Applied Probability, 33(3):714–728, September 1996. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3215353.

Martin:1997:PLM

[MW97] R. J. Martin and A. M. Walker. A power-law model and other models for long-range dependence. Journal of Applied Probability, 34(3):657–670, September 1997. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3215092.

Mahmoud:2008:ACA

[MW08] Hosam M. Mahmoud and Mark Daniel Ward. Average-case analysis of cousins in m-ary tries. Journal of Applied Probability, 45 (3):888–900, September 2008. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.

org/stable/27595992.

Ma:2010:EVC

[MW10] Chunhua Ma and Longmin Wang. On estimation of the variances for critical branching processes with immigration. Journal of Applied Probability, 47(2):526–542, June 2010. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/20778495.

Mahmoud:2015:APP

[MW15] Hosam M. Mahmoud and Mark D. Ward. Asymptotic properties of protected nodes in random recursive trees. Journal of Applied Probability, 52(1):290–297, March 2015. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://projecteuclid.org/euclid.jap/1429282623.

Mao:2021:LTC

[MW21] Yong-Hua Mao and Tao Wang. Lyapunov-type conditions for non-strong ergodicity of Markov processes. Journal of Applied Probability, 58(1):238–253, March 2021. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL https://www.cambridge.org/core/journals/
journal-of-applied-probability/article/lyapunovtype-
conditions-for-nonstrong-ergodicity-of-markov-processes/
E2FE7FDE8F0EF69273FDE5DE9EC48B35.

Mo:2001:OSA

[MWCY01] Charles T. C. Mo, Samuel S. Wu, Robert Chen, and Mark C. K. Yang. Optimal sequential allocation with imperfect feedback information. Journal of Applied Probability, 38(1):248–254, March 2001. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3215757.

Mao:2022:FPF

[MWH22] Tiantian Mao, Qinyu Wu, and Taizhong Hu. Further properties of fractional stochastic dominance. Journal of Applied Probability, 59(1):202–223, March 2022. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL https:/

/www.cambridge.org/core/journals/journal-of-applied-
probability/article/further-properties-of-fractional-
stochastic-dominance/3DFA3B4F05D385E0EBDAB8D33B2F2DB5.

Mao:2019:SSU

[MWW19] Tiantian Mao, Bin Wang, and Ruodu Wang. Sums of standard uniform random variables. Journal of Applied Probability, 56(3):918–936, September 2019. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL https:/

/www.cambridge.org/core/journals/journal-of-applied-
probability/article/sums-of-standard-uniform-random-
variables/313A647F5C6874B5ACCAA8A1C840EA88.

Mitov:1984:CGW

[MY84] K. V. Mitov and N. M. Yanev. Critical Galton–Watson processes with decreasing state-dependent immigration. Journal of Applied Probability, 21(1):22–39, March 1984. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/3213661.

Mitov:1985:BHB

[MY85] K. V. Mitov and N. M. Yanev. Bellman–Harris branching processes with state-dependent immigration. Journal of Applied Probability, 22(4):757–765, December 1985. CODEN JPRBAM.

ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/3213943.

Miyazawa:1988:BES

[MY88] Masakiyo Miyazawa and Genji Yamazaki. The basic equations for a supplemented GSMP and its applications to queues. Journal of Applied Probability, 25(3):565–578, September 1988. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/3213985.

Miyazawa:1991:COA

[MY91] Masakiyo Miyazawa and Genji Yamazaki. Convex ordering of the attained waiting times in single-server queues and related problems. Journal of Applied Probability, 28(2):433–445, June 1991. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3214878.

Mitov:2001:RPI

[MY01] K. V. Mitov and N. M. Yanev. Regenerative processes in the infinite mean cycle case. Journal of Applied Probability, 38(1):165– 179, March 2001. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3215749.

Myers:1981:SFI

[Mye81] Lawrence E. Myers. Survival functions induced by stochastic covariate processes. Journal of Applied Probability, 18(2):523– 529, June 1981. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3213300.

Maller:1993:PLO

[MZ93] R. A. Maller and S. Zhou. The probability that the largest observation is censored. Journal of Applied Probability, 30(3): 602–615, September 1993. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3214769.

Ma:2001:RQW

[MZ01] Li Ma and Jiancang Zhuang. Relative quiescence within the Jiashi swarm in Xinjiang, China: An application of the ETAS point process model. Journal of Applied Probability, 38A: 213–221, 2001. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3215880.

Mao:2004:EES

[MZ04] Yong-Hua Mao and Yu-Hui Zhang. Exponential ergodicity for single-birth processes. Journal of Applied Probability, 41(4): 1022–1032, December 2004. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.

org/stable/4141375.

Muratov:2016:BFT

[MZ16] Anton Muratov and Sergei Zuyev. Bit flipping and time to recover. Journal of Applied Probability, 53(3):650–666, September 2016. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://projecteuclid.org/euclid.jap/
1476370768.

Mu:2020:STO

[MZ20] Y. X. Mu and Y. Zhang. On some threshold-one attractive interacting particle systems on homogeneous trees. Journal of Applied Probability, 57(3):866–898, September 2020. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL https://www.cambridge.org/core/journals/journal-
of-applied-probability/article/on-some-thresholdone-
attractive-interacting-particle-systems-on-homogeneous-
trees/1FC82B6551B25C12DCA8257F0B48503A.

Ma:2024:ECS

[MZ24] Rugang Ma and Xiaowen Zhou. Explosion of continuous-state branching processes with competition in a Lévy environment. Journal of Applied Probability, 61(1):68–81, March 2024. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL https://www.cambridge.org/core/journals/
journal-of-applied-probability/article/explosion-of-
continuousstate-branching-processes-with-competition-
in-a-levy-environment/A412A5EDB675073D0706E14A664583AF.

Murciano:2002:PDN

[MZLSRS02] Antonio Murciano, Javier Zamora, Jesus Lopez-Sanchez, and Emilia Rodriguez-Santamaria. Probability distribution of the number of deceits in collective robotics. Journal of Applied Probability, 39(3):657–663, September 2002. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/3216088.

Mao:2016:SCU

[MZZ16] Y. H. Mao, C. Zhang, and Y. H. Zhang. Separation cutoff for upward skip-free chains. Journal of Applied Probability, 53 (1):299–306, March 2016. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://projecteuclid.

org/euclid.jap/1457470576.

Neuts:2004:PFM

[NA04] Marcel F. Neuts and Attahiru Sule Alfa. Pair formation in a Markovian arrival process with two event labels. Journal of Applied Probability, 41(4):1124–1137, December 2004. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/4141384.

Nadas:1971:TIP

[Nád71] Arthur Nádas. Times to intermittent and to permanent failures as Brownian crossing times. Journal of Applied Probability, 8(4): 838–840, December 1971. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3212249.

Nagaraja:1982:RVE

[Nag82a] H. N. Nagaraja. Record values and extreme value distributions. Journal of Applied Probability, 19(1):233–239, March 1982. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/3213934.

Nagaraja:1982:SAR

[Nag82b] H. N. Nagaraja. Some asymptotic results for the induced selection differential. Journal of Applied Probability, 19(2):253– 261, June 1982. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3213478.

Nagel:1993:ODC

[Nag93] W. Nagel. Orientation-dependent chord length distributions characterize convex polygons. Journal of Applied Probability, 30 (3):730–736, September 1993. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.

org/stable/3214779.

Nagaev:2001:AFB

[Nag01] A. V. Nagaev. An asymptotic formula for the Bayes risk in discriminating between two Markov chains. Journal of Applied Probability, 38A:131–141, 2001. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.

org/stable/3215874.

Nair:1971:SST

[Nai71] Sreekantan S. Nair. A single server tandem queue. Journal of Applied Probability, 8(1):95–109, March 1971. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3211840.

Nakamura:1974:SLT

[Nak74] Masao Nakamura. Some limit theorems for a class of network problems as related to finite Markov chains. Journal of Applied Probability, 11(1):94–101, March 1974. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/3212586.

Nakagawa:1979:RPP

[Nak79a] T. Nakagawa. Replacement problem of a parallel system in random environment. Journal of Applied Probability, 16(1):203– 205, March 1979. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3213388.

Nakagawa:1979:FRR

[Nak79b] Toshio Nakagawa. Further results of replacement problem of a parallel system in random environment. Journal of Applied Probability, 16(4):923–926, December 1979. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/3213159.

Nakagawa:1984:ONU

[Nak84] Toshio Nakagawa. Optimal number of units for a parallel system. Journal of Applied Probability, 21(2):431–436, June 1984. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/3213653.

Nakagawa:1986:PSP

[Nak86a] Toshio Nakagawa. Periodic and sequential preventive maintenance policies. Journal of Applied Probability, 23(2):536– 542, June 1986. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3214197.

Nakai:1986:SGO

[Nak86b] Teruhisa Nakai. A search game with one object and two searchers. Journal of Applied Probability, 23(3):696–707, September 1986. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3214008.

Nakatsuka:1986:SEC

[Nak86c] Toshinao Nakatsuka. The substability and ergodicity of complicated queueing systems. Journal of Applied Probability, 23 (1):193–200, March 1986. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3214128.

Nakai:1991:OSP

[Nak91] Teruhisa Nakai. An optimal stopping problem in the excavation of archaeological remains. Journal of Applied Probability, 28(4): 924–929, December 1991. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3214698.

Nakatsuka:1998:APR

[Nak98] Toshinao Nakatsuka. Absorbing process in recursive stochastic equations. Journal of Applied Probability, 35(2):418–426, June 1998. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3215695.

Nakatsuka:2006:UEM

[Nak06] Toshinao Nakatsuka. The untraceable events method for absorbing processes. Journal of Applied Probability, 43(3):652–664, September 2006. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
27595762.

Nakata:2008:PAO

[Nak08] Toshio Nakata. A Poisson approximation for an occupancy problem with collisions. Journal of Applied Probability, 45(2):430– 439, June 2008. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
27595955.

Nakata:2009:NPP

[Nak09] Toshio Nakata. Necklace processes via Pólya urns. Journal of Applied Probability, 46(1):284–295, March 2009. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/25662421.

Nanthi:1979:ECP

[Nan79] K. Nanthi. Estimation of the criticality parameter of a superitical branching process with random environments. Journal of Applied Probability, 16(4):890–896, December 1979. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3213154.

Nanthi:1982:EVM

[Nan82] K. Nanthi. Estimation of the variance for the multitype Galton– Watson process. Journal of Applied Probability, 19(2):408– 414, June 1982. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3213492.

Narasimham:1968:NAD

[Nar68] T. Narasimham. A note on the asymptotic distribution of the traffic-time-average in a GI/G/∞ with bulk arrivals. Journal of Applied Probability, 5(2):476–480, August 1968. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3212268.

Nash:2008:CHA

[Nas08] L. Nash. Chris Heyde: An appreciation. Journal of Applied Probability, 45(3):593–594, September 2008. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/27595972.

Natvig:1974:TSP

[Nat74] Bent Natvig. On the transient state probabilities for a queueing model where potential customers are discouraged by queue length. Journal of Applied Probability, 11(2):345–354, June 1974. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/3212755.

Natvig:1975:WTB

[Nat75] Bent Natvig. On the waiting-time and busy period distributions for a general birth-and-death queueing model. Journal of Applied Probability, 12(3):524–532, September 1975. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3212867.

Natvig:1977:RIO

[Nat77] Bent Natvig. On the reversibility of the input and output processes for a general birth-and-death queueing model. Journal of Applied Probability, 14(4):876–883, December 1977. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3213362.

Natvig:1982:RRS

[Nat82] Bent Natvig. On the reduction in remaining system lifetime due to the failure of a specific component. Journal of Applied Probability, 19(3):642–652, September 1982. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/3213521. See correction [Ano83f] and

note [Nor86].

Natvig:1990:IBM

[Nat90] Bent Natvig. On information-based minimal repair and the reduction in remaining system lifetime due to the failure of a specific module. Journal of Applied Probability, 27(2):365–375, June 1990. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3214655.

Natvig:2014:DNM

[Nat14] Bent Natvig. On the deterioration of nonrepairable multistate strongly coherent systems. Journal of Applied Probability, 51 (1):69–81, March 2014. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://projecteuclid.

org/euclid.jap/1395771414.

Nawijn:1983:NMS

[Naw83] W. M. Nawijn. A note on many-server queueing systems with ordered entry, with an application to conveyor theory. Journal of Applied Probability, 20(1):144–152, March 1983. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3213728.

Nayak:1987:MLD

[Nay87] Tapan Kumar Nayak. Multivariate Lomax distribution: Properties and usefulness in reliability theory. Journal of Applied Probability, 24(1):170–177, March 1987. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3214068.

Nagaraja:2003:CUR

[NB03] H. N. Nagaraja and Gadi Barlevy. Characterizations using record moments in a random record model and applications. Journal of Applied Probability, 40(3):826–833, September 2003. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/3215958.

Navarro:2008:MRR

[NBS08] Jorge Navarro, N. Balakrishnan, and Francisco J. Samaniego. Mixture representations of residual lifetimes of used systems. Journal of Applied Probability, 45(4):1097–1112, December 2008. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/27596011. See

erratum [KZN15].

Neuts:1972:ISQ

[NC72] Marcel F. Neuts and Shun-Zer Chen. The infinite server queue with semi-Markovian arrivals and negative exponential services. Journal of Applied Probability, 9(1):178–184, March 1972. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/3212646.

Nath:1981:OSR

[NE81] G. B. Nath and E. G. Enns. Optimal service rates in the multiserver loss system with heterogeneous servers. Journal of Applied Probability, 18(3):776–781, September 1981. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/3213336.

Nath:1982:ROS

[NE82] G. B. Nath and E. G. Enns. Retraction: Optimal service rates in the multiserver loss system with heterogeneous servers. Journal of Applied Probability, 19(1):250–251, March 1982. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3213938.

Navarro:2007:MRL

[NE07] Jorge Navarro and Serkan Eryilmaz. Mean residual lifetimes of consecutive-k-out-of-n systems. Journal of Applied Probability, 44(1):82–98, March 2007. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/27595824.

Neal:2003:SEB

[Nea03] Peter Neal. SIR epidemics on a Bernoulli random graph. Journal of Applied Probability, 40(3):779–782, September 2003. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3215950.

Neal:2005:CPL

[Nea05] Peter Neal. Compound Poisson limits for household epidemics. Journal of Applied Probability, 42(2):334–345, June 2005. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/30040796.

Neal:2007:CTS

[Nea07] Peter Neal. Coupling of two SIR epidemic models with variable susceptibilities and infectivities. Journal of Applied Probability, 44(1):41–57, March 2007. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/27595821.

Neal:2008:GCC

[Nea08a] Peter Neal. The generalised coupon collector problem. Journal of Applied Probability, 45(3):621–629, September 2008. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/27595975.

Neal:2008:SGC

[Nea08b] Peter Neal. The SIS great circle epidemic model. Journal of Applied Probability, 45(2):513–530, June 2008. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/27595961.

Neal:2016:HSE

[Nea16] Peter Neal. A household SIR epidemic model incorporating time of day effects. Journal of Applied Probability, 53(2):489–501, June 2016. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-

6072 (electronic). URL http://projecteuclid.org/euclid.
jap/1466172869.

Neininger:2001:RCP

[Nei01] Ralph Neininger. Rates of convergence for products of random stochastic 2 × 2 matrices. Journal of Applied Probability, 38(3): 799–806, September 2001. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3216134.

Nelson:1977:SSL

[Nel77] Peter R. Nelson. Single-shelf library-type Markov chains with infinitely many books. Journal of Applied Probability, 14(2):298– 308, June 1977. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3213000.

Nelson:1982:TRP

[Nel82] Peter R. Nelson. The transposition replacement policy with a partial memory. Journal of Applied Probability, 19(3):733–736, September 1982. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3213537.

Nerman:1984:SPC

[Ner84] Olle Nerman. The stable pedigrees of critical branching populations. Journal of Applied Probability, 21(3):447–463, September 1984. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3213609.

Neuts:1966:SSQ

[Neu66] Marcel F. Neuts. The single server queue with Poisson input and semi-Markov service times. Journal of Applied Probability, 3(1):202–230, June 1966. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3212047.

Neuts:1967:WBR

[Neu67] Marcel F. Neuts. Waitingtimes between record observations. Journal of Applied Probability, 4(1):206–208, April 1967. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/3212314.

Neuts:1968:JDV

[Neu68a] Marcel F. Neuts. The joint distribution of the virtual waitingtime and the residual busy period for the M/G/1 queue. Journal of Applied Probability, 5(1):224–229, April 1968. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3212091.

Neuts:1968:TQS

[Neu68b] Marcel F. Neuts. Two queues in series with a finite, intermediate waitingroom. Journal of Applied Probability, 5(1):123–142, April 1968. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3212081.

Neuts:1978:SMA

[Neu78] Marcel F. Neuts. The second moments of the absorption times in the Markov renewal branching process. Journal of Applied Probability, 15(4):707–714, December 1978. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/3213428.

Neuts:1979:VMP

[Neu79] Marcel F. Neuts. A versatile Markovian point process. Journal of Applied Probability, 16(4):764–779, December 1979. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3213143.

Neuts:1980:PSR

[Neu80] Marcel F. Neuts. The probabilistic significance of the rate matrix in matrix-geometric invariant vectors. Journal of Applied Probability, 17(1):291–296, March 1980. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3212949.

Neuts:1981:SWT

[Neu81] Marcel F. Neuts. Stationary waiting-time distributions in the GI/PH/1 queue. Journal of Applied Probability, 18(4):901–912, December 1981. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3213064.

Neumann:1982:PPR

[Neu82] B. H. Neumann. Plane polygons revisited. Journal of Applied Probability, 19A:113–122, 1982. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.

org/stable/3213554.

Neuts:1988:CEA

[Neu88] Marcel F. Neuts. Computer experimentation in applied probability. Journal of Applied Probability, 25A:31–43, 1988. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3214145.

Neuts:1994:ESP

[Neu94a] Marcel F. Neuts. An explicit solution to a particular Markov chain of M/G/1 type. Journal of Applied Probability, 31A: 337–342, 1994. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3214966.

Neuts:1994:IRW

[Neu94b] Marcel F. Neuts. An interesting random walk on the nonnegative integers. Journal of Applied Probability, 31(1):48–58, March 1994. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3215234.

Neumann:2023:NMC

[Neu23] Berenice Anne Neumann. Nonlinear Markov chains with finite state space: invariant distributions and long-term behaviour. Journal of Applied Probability, 60(1):30–44, March 2023. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL https://www.cambridge.org/core/journals/journal-
of-applied-probability/article/nonlinear-markov-chains-
with-finite-state-space-invariant-distributions-and-
longterm-behaviour/C7AFF2EA3E4280E59636FECD32CB2982.

Newell:1966:EPD

[New66] G. F. Newell. Equilibrium probability distributions for low density highway traffic. Journal of Applied Probability, 3(1):247– 260, June 1966. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3212049.

Newell:1968:QTDb

[New68a] G. F. Newell. Queues with time-dependent arrival rates: II. The maximum queue and the return to equilibrium. Journal of Applied Probability, 5(3):579–590, December 1968. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3211923.

Newell:1968:QTDc

[New68b] G. F. Newell. Queues with time-dependent arrival rates: III. A mild rush hour. Journal of Applied Probability, 5(3):591–606, December 1968. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3211924.

Newell:1968:QTDa

[New68c] G. F. Newell. Queues with time-dependent arrival rates i:the transition through saturation. Journal of Applied Probability, 5 (2):436–451, August 1968. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3212264.

Newman:1970:NFP

[New70] David S. Newman. A new family of point processes which are characterized by their second moment properties. Journal of Applied Probability, 7(2):338–358, August 1970. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3211969.

Newton:1986:SMF

[New86] J. Newton. A solvable model for a finite-capacity queueing system. Journal of Applied Probability, 23(4):1045–1047, December 1986. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3214479.

Novikov:1999:ABC

[NFK99] Alex Novikov, Volf Frishling, and Nino Kordzakhia. Approximations of boundary crossing probabilities for a Brownian motion. Journal of Applied Probability, 36(4):1019–1030, December 1999. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3215575.

Navarro:2020:ESB

[NFS20] Jorge Navarro and Juan Fernández-Sánchez. On the extension of signature-based representations for coherent systems with dependent non-exchangeable components. Journal of Applied Probability, 57(2):429–440, June 2020. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL https:/

/www.cambridge.org/core/journals/journal-of-applied-
probability/article/on-the-extension-of-signaturebased-

representations-for-coherent-systems-with-dependent-
nonexchangeable-components/BFA014A32538F5A1E90D15656F9DBF29.

Nananukul:1995:MWT

[NG95] Soracha Nananukul and Wei-Bo Gong. The mean waiting time of a GI/G/1 queue in light traffic via random thinning. Journal of Applied Probability, 32(1):256–266, March 1995. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3214934.

Nguyen:1990:PCR

[Ngu90] Bao Gia Nguyen. Percolation of coalescing random walks. Journal of Applied Probability, 27(2):269–277, June 1990. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3214646.

Najock:1982:NTV

[NH82] D. Najock and C. C. Heyde. On the number of terminal vertices in certain random trees with an application to stemma construction in philology. Journal of Applied Probability, 19(3):675–680, September 1982. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3213526.

Niculescu:1984:ADM

[Nic84] Stefan P. Niculescu. On the asymptotic distribution of multivariate renewal processes. Journal of Applied Probability, 21(3): 639–645, September 1984. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3213624.

Nicholls:1986:BJA

[Nic86] D. F. Nicholls. The Box–Jenkins approach to random coefficient autoregressive modelling. Journal of Applied Probability, 23A: 231–240, 1986. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3214355.

Niemi:1985:NSM

[Nie85] Seppo Niemi. On non-singular Markov renewal processes with an application to a growth-catastrophe model. Journal of Applied Probability, 22(2):253–266, June 1985. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3213770.

Niemiro:1995:TES

[Nie95] Wojciech Niemiro. Tail events of simulated annealing Markov chains. Journal of Applied Probability, 32(4):867–876, December 1995. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3215200.

Niese:2006:MCP

[Nie06] Birgit Niese. A martingale characterization of Pólya–Lundberg processes. Journal of Applied Probability, 43(3):741–754, September 2006. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
27595769.

Nishimura:1974:RCPa

[Nis74a] Shoichi Nishimura. Random collision processes and their limiting distribution using the discrimination information. Journal of Applied Probability, 11(2):266–280, June 1974. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3212748.

Nishimura:1974:RCPb

[Nis74b] Shoichi Nishimura. Random collision processes with transition probabilities belonging to the same type of distribution. Journal of Applied Probability, 11(4):703–714, December 1974. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3212554.

Nishimura:1975:IMR

[Nis75] Shoichi Nishimura. An inequality for a metric in a random collision process. Journal of Applied Probability, 12(2):239–247, June 1975. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3212437.

Niu:1980:BED

[Niu80] Shun-Chen Niu. Bounds for the expected delays in some tandem queues. Journal of Applied Probability, 17(3):831–838, September 1980. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3212976.

Niu:1981:CWT

[Niu81] Shun-Chen Niu. On the comparison of waiting times in tandem queues. Journal of Applied Probability, 18(3):707–714, September 1981. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3213324.

Nanda:1996:CSP

[NJS96a] Asok K. Nanda, Kanchan Jain, and Harshinder Singh. On closure of some partial orderings under mixtures. Journal of Applied Probability, 33(3):698–706, September 1996. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3215351.

Nanda:1996:PMO

[NJS96b] Asok K. Nanda, Kanchan Jain, and Harshinder Singh. Properties of moments for s-order equilibrium distributions. Journal of Applied Probability, 33(4):1108–1111, December 1996. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3214989.

Nino-Mora:2000:AIC

[NMG00] José Niño-Mora and Kevin D. Glazebrook. Assessing an intuitive condition for stability under a range of traffic conditions via a generalised Lu-kumar network. Journal of Applied Probability, 37(3):890–899, September 2000. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.

org/stable/3215621.

Nguyen:2020:ASI

[NNY20] Dang H. Nguyen, Nhu N. Nguyen, and George Yin. Analysis of a spatially inhomogeneous stochastic partial differential equation epidemic model. Journal of Applied Probability, 57(2):613–636, June 2020. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL https:/

/www.cambridge.org/core/journals/journal-of-applied-
probability/article/analysis-of-a-spatially-inhomogeneous-
stochastic-partial-differential-equation-epidemic-model/
E552DE646F77B0A01FCF5F4386097539.

Nakagawa:1974:ORP

[NO74] Toshio Nakagawa and Shunji Osaki. Optimum replacement policies with delay. Journal of Applied Probability, 11(1):102– 110, March 1974. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3212587.

Niculescu:1985:EON

[NO85] Stefan P. Niculescu and Edward Omey. On the exact order of normal approximation in multivariate renewal theory. Journal of Applied Probability, 22(2):280–287, June 1985. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3213772.

Norman:1975:ETE

[Nor75] M. Frank Norman. An ergodic theorem for evolution in a random environment. Journal of Applied Probability, 12(4):661–672, December 1975. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3212717.

Norman:1977:EDT

[Nor77] M. Frank Norman. Ergodicity of diffusion and temporal uniformity of diffusion approximation. Journal of Applied Probability, 14(2):399–404, June 1977. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3213013.

Norros:1986:NNM

[Nor86] Ilkka Norros. Notes on Natvig’s measure of importance of system components. Journal of Applied Probability, 23(3):736–747, September 1986. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/

3214011. See [Nat82].

Norberg:1997:TMC

[Nor97] Tommy Norberg. On the time a Markov chain spends in a lumped state. Journal of Applied Probability, 34(2):340– 345, June 1997. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3215374.

Notohara:2000:PMS

[Not00] Morihiro Notohara. A perturbation method for the structured coalescent with strong migration. Journal of Applied Probability, 37(1):148–167, March 2000. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3215665.

Notohara:2001:SCP

[Not01] Morihiro Notohara. The structured coalescent process with weak migration. Journal of Applied Probability, 38(1):1–17, March 2001. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3215737.

Nowik:1990:IPC

[Now90] Shmuel Nowik. Identifiability problems in coherent systems. Journal of Applied Probability, 27(4):862–872, December 1990. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/3214829.

Neuts:1991:SID

[NP91] Marcel F. Neuts and Charles E. M. Pearce. The superposition of independent discrete Markovian packet streams. Journal of Applied Probability, 28(1):84–95, March 1991. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/3214742.

Nuzman:2000:LES

[NP00] Carl J. Nuzman and H. Vincent Poor. Linear estimation of selfsimilar processes via Lamperti’s transformation. Journal of Applied Probability, 37(2):429–452, June 2000. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/3215719.

Niemiro:2009:FPM

[NP09] Wojciech Niemiro and Piotr Pokarowski. Fixed precision MCMC estimation by median of products of averages. Journal of Applied Probability, 46(2):309–329, June 2009. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/25662426.

Nguyen:2022:RSC

[NP22] Giang T. Nguyen and Oscar Peralta. Rate of strong convergence to Markov-modulated Brownian motion. Journal of Applied Probability, 59(1):1–16, March 2022. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL https:/

/www.cambridge.org/core/journals/journal-of-applied-
probability/article/rate-of-strong-convergence-to-markovmodulated-
brownian-motion/4B9F3AEF32C243E2D6248D55120FC634.

Noba:2018:OPD

[NPYY18] Kei Noba, José-Luis Pérez, Kazutoshi Yamazaki, and Kouji Yano. On optimal periodic dividend and capital injection strategies for spectrally negative Lévy models. Journal of Applied Probability, 55(4):1272–1286, December 2018. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL https://www.cambridge.org/core/journals/journal-
of-applied-probability/article/on-optimal-periodic-dividend-
and-capital-injection-strategies-for-spectrally-negative-
levy-models/26CE7B4C328CA1E7B9FD160FD07AE590.

Nozaki:1978:AFC

[NR78] Shirley A. Nozaki and Sheldon M. Ross. Approximations in finite-capacity multi-server queues with Poisson arrivals. Journal of Applied Probability, 15(4):826–834, December 1978. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3213437.

Neuts:1984:SMW

[NR84] Marcel F. Neuts and M. F. Ramalhoto. A service model in which the server is required to search for customers. Journal of Applied Probability, 21(1):157–166, March 1984. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3213673.

Nishide:2011:OTE

[NR11] Katsumasa Nishide and L. C. G. Rogers. Optimal time to exchange two baskets. Journal of Applied Probability, 48(1):21– 30, March 2011. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
29777443.

Nouiehed:2018:CFB

[NR18] Maher Nouiehed and Sheldon M. Ross. A conjecture on the Feldman bandit problem. Journal of Applied Probability, 55(1):318–324, March 2018. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL https:/

/www.cambridge.org/core/journals/journal-of-applied-
probability/article/conjecture-on-the-feldman-bandit-
problem/B957CCD9069EC13D9839519E1EC9B743.

Norvaisa:2020:UAN

[NR20] Rimas Norvaisa and Alfredas Rackauskas. Uniform asymptotic normality of weighted sums of short-memory linear processes. Journal of Applied Probability, 57(1):174–195, March 2020. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL https://www.cambridge.org/core/journals/journal-
of-applied-probability/article/uniform-asymptotic-normality-
of-weighted-sums-of-shortmemory-linear-processes/C90925A84677D1DD96AC98E3E7AA3816.

Nguyen:2022:ECM

[NR22] Quang Huy Nguyen and Christian Y. Robert. Efficient conditional Monte Carlo simulations for the exponential integrals of Gaussian random fields. Journal of Applied Probability, 59(2):366–383, June 2022. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL https:/

/www.cambridge.org/core/journals/journal-of-applied-
probability/article/efficient-conditional-monte-carlo-
simulations-for-the-exponential-integrals-of-gaussian-
random-fields/AD58F49A4F56CCAF97DEA2A9C0A71187.

Nobile:1985:ETF

[NRS85a] A. G. Nobile, L. M. Ricciardi, and L. Sacerdote. Exponential trends of first-passage-time densities for a class of diffusion processes with steady-state distribution. Journal of Applied Probability, 22(3):611–618, September 1985. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3213864.

Nobile:1985:ETO

[NRS85b] A. G. Nobile, L. M. Ricciardi, and L. Sacerdote. Exponential trends of Ornstein–Uhlenbeck first-passage-time densities. Journal of Applied Probability, 22(2):360–369, June 1985. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3213779.

Nobile:1985:NFP

[NRS85c] A. G. Nobile, L. M. Ricciardi, and L. Sacerdote. A note on firstpassage time and some related problems. Journal of Applied Probability, 22(2):346–359, June 1985. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3213778.

Nain:1980:TPM

[NS80] R. B. Nain and Kanwar Sen. Transition probability matrices for correlated random walks. Journal of Applied Probability, 17 (1):253–258, March 1980. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3212943.

Natvig:1984:SSB

[NS84] Bent Natvig and Arnfried Streller. The steady-state behaviour of multistate monotone systems. Journal of Applied Probability, 21(4):826–835, December 1984. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.

org/stable/3213699.

Norberg:2005:WTV

[NS05] Ragnar Norberg and Mogens Steffensen. What is the time value of a stream of investments? Journal of Applied Probability, 42 (3):861–866, September 2005. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.

org/stable/30040862.

Navarro:2006:HRO

[NS06] Jorge Navarro and Moshe Shaked. Hazard rate ordering of order statistics and systems. Journal of Applied Probability, 43(2):391– 408, June 2006. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
27595739.

Nikitin:2022:GLT

[NS22] Yakov Nikitin and Ekaterina Simarova. Generalized limit theorems for U-max statistics. Journal of Applied Probability, 59(3):825–848, September 2022. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL https:/

/www.cambridge.org/core/journals/journal-of-applied-
probability/article/generalized-limit-theorems-for-umax-
statistics/449A9553E06781C2BBF724748F1252CF.

Navarro:2010:JSC

[NSB10] Jorge Navarro, Francisco J. Samaniego, and N. Balakrishnan. The joint signature of coherent systems with shared components. Journal of Applied Probability, 47(1):235–253, March 2010. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/25662512.

Navarro:2011:SBR

[NSB11] Jorge Navarro, Francisco J. Samaniego, and N. Balakrishnan. Signature-based representations for the reliability of systems with heterogeneous components. Journal of Applied Probability, 48(3):856–867, September 2011. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/23065942.

Nair:2008:TTT

[NSK08] N. Unnikrishnan Nair, P. G. Sankaran, and B. Vinesh Kumar. Total time on test transforms of order n and their implications in reliability analysis. Journal of Applied Probability, 45(4):1126– 1139, December 2008. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/27596013.

Nakayama:2004:FEM

[NSS04] Marvin K. Nakayama, Perwez Shahabuddin, and Karl Sigman. On finite exponential moments for branching processes and busy periods for queues. Journal of Applied Probability, 41A:273–280, 2004. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3215982.

Ng:2004:ABT

[NT04] Kai W. Ng and Qihe Tang. Asymptotic behavior of tail and local probabilities for sums of subexponential random variables. Journal of Applied Probability, 41(1):108–116, March 2004. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/3215818.

Ng:2004:PLD

[NTaYY04] Kai W. Ng, Qihe Tang, Jia an Yan, and Hailiang Yang. Precise large deviations for sums of random variables with consistently varying tails. Journal of Applied Probability, 41(1):93– 107, March 2004. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3215817.

Nain:1989:IAS

[NTW89] Philippe Nain, Pantelis Tsoucas, and Jean Walrand. Interchange arguments in stochastic scheduling. Journal of Applied Probability, 26(4):815–826, December 1989. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3214386.

Ng:2003:PLD

[NTYY03] Kai W. Ng, Qihe Tang, Jiaan Yan, and Hailiang Yang. Precise large deviations for the prospective-loss process. Journal of Applied Probability, 40(2):391–400, June 2003. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/3215798.

Nuel:2008:PMC

[Nue08] Grégory Nuel. Pattern Markov chains: Optimal Markov chain embedding through deterministic finite automata. Journal of Applied Probability, 45(1):226–243, March 2008. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/27595936.

Nuel:2010:FMR

[Nue10] G. Nuel. On the first k moments of the random count of a pattern in a multistate sequence generated by a Markov source. Journal of Applied Probability, 47(4):1105–1123, December 2010. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/29777430.

Nguyen:2020:CRW

[NV20] Tuan-Minh Nguyen and Stanislav Volkov. On a class of random walks in simplexes. Journal of Applied Probability, 57(2):409–428, June 2020. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL https:/

/www.cambridge.org/core/journals/journal-of-applied-
probability/article/on-a-class-of-random-walks-in-simplexes/
D7AB5D916FC7EB316B930C5205FBE6C7.

Natarajan:2024:WCP

[NvdBOdI24] Abhinav Natarajan, Willem van den Boom, Kristoforus Bryant Odang, and Maria de Iorio. On a wider class of prior distributions for graphical models. Journal of Applied Probability, 61(1):230–243, March 2024. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL https:/

/www.cambridge.org/core/journals/journal-of-applied-
probability/article/on-a-wider-class-of-prior-distributions-
for-graphical-models/0B9D4F33C7747508A8688FBAC442FAB6.

Nelsen:1970:RDR

[NW70] Roger B. Nelsen and Trevor Williams. Random displacements of regularly spaced events. Journal of Applied Probability, 7 (1):183–195, April 1970. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3212159.

Nagaev:2007:CGW

[NW07] S. V. Nagaev and V. Wachtel. The critical Galton–Watson process without further power moments. Journal of Applied Probability, 44(3):753–769, September 2007. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/27595880.

Nyquist:2017:MDP

[Nyq17] Pierre Nyquist. Moderate deviation principles for importance sampling estimators of risk measures. Journal of Applied Probability, 54(2):490–506, June 2017. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL https:/

/www.cambridge.org/core/journals/journal-of-applied-
probability/article/moderate-deviation-principles-for-
importance-sampling-estimators-of-risk-measures/63E604C94B06D09A0CECCF5962107318.

Nyrhinen:1994:RLR

[Nyr94] Harri Nyrhinen. Rough limit results for level-crossing probabilities. Journal of Applied Probability, 31(2):373–382, June 1994. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/3215030.

Nyrhinen:1999:LDT

[Nyr99] Harri Nyrhinen. Large deviations for the time of ruin. Journal of Applied Probability, 36(3):733–746, September 1999. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3215437.

Osaki:1970:TUS

[OA70] Shunji Osaki and Tatsuyuki Asakura. A two-unit standby redundant system with repair and preventive maintenance. Journal of Applied Probability, 7(3):641–648, December 1970. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3211943.

Oguz:2012:HIF

[OA12] Barlas Oguz and Venkat Anantharam. Hurst index of functions of long-range-dependent Markov chains. Journal of Applied Probability, 49(2):451–471, June 2012. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/41713779.

Oakes:1975:MSE

[Oak75] David Oakes. The Markovian self-exciting process. Journal of Applied Probability, 12(1):69–77, March 1975. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3212408.

Oakes:1979:MPI

[Oak79] David Oakes. Matched pairs in independent Poisson processes. Journal of Applied Probability, 16(4):780–793, December 1979. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/3213144.

OBrien:1974:LTS

[O’B74] G. L. O’Brien. Limit theorems for sums of chain-dependent processes. Journal of Applied Probability, 11(3):582–587, September 1974. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3212704.

OBrien:1975:IQD

[O’B75] G. L. O’Brien. Inequalities for queues with dependent interarrival and service times. Journal of Applied Probability, 12(3): 653–656, September 1975. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3212887.

OBrien:1980:LTS

[O’B80] G. L. O’Brien. A limit theorem for sample maxima and heavy branches in Galton–Watson trees. Journal of Applied Probability, 17(2):539–545, June 1980. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.

org/stable/3213043.

OConnell:1993:YPA

[O’C93] Neil O’Connell. Yule process approximation for the skeleton of a branching process. Journal of Applied Probability, 30(3):

725–729, September 1993. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3214778.

OConnell:1997:LDD

[O’C97] Neil O’Connell. Large deviations for departures from a shared buffer. Journal of Applied Probability, 34(3):753–766, September 1997. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3215100.

OConnell:1998:LDQ

[O’C98] Neil O’Connell. Large deviations for queue lengths at a multibuffered resource. Journal of Applied Probability, 35(1):240– 245, March 1998. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3215563.

Ovaskainen:2003:BMB

[OC03] Otso Ovaskainen and Stephen J. Cornell. Biased movement at a boundary and conditional occupancy times for diffusion processes. Journal of Applied Probability, 40(3):557–580, September 2003. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3215936.

Olvera-Cravioto:2022:SCS

[OC22] Mariana Olvera-Cravioto. Strong couplings for static locally tree-like random graphs. Journal of Applied Probability, 59(4):1261–1285, December 2022. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL https:/

/www.cambridge.org/core/journals/journal-of-applied-
probability/article/strong-couplings-for-static-locally-
treelike-random-graphs/36991B796EB21674736CC8226070FD77.

Oesting:2018:ERM

[Oes18] Marco Oesting. Equivalent representations of max-stable processes via ℓ_{p}-norms. Journal of Applied Probability, 55(1):54–68, March 2018. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL https:/

/www.cambridge.org/core/journals/journal-of-applied-
probability/article/equivalent-representations-of-maxstable-
processes-via-pnorms/62359A2442FF5C9230D1BAD5EB19FB69.

OFlaherty:1989:SRT

[O’F89] Brendan O’Flaherty. Some results on two-armed bandits when both projects vary. Journal of Applied Probability, 26(3):655– 658, September 1989. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3214424.

Ogata:1980:MLE

[Oga80] Yosihiko Ogata. Maximum likelihood estimates of incorrect Markov models for time series and the derivation of AIC. Journal of Applied Probability, 17(1):59–72, March 1980. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3212924.

Ogata:2001:EAE

[Oga01] Yosihiko Ogata. Exploratory analysis of earthquake clusters by likelihood-based trigger models. Journal of Applied Probability, 38A:202–212, 2001. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3215879.

Ohtsubo:1991:DTN

[Oht91] Yoshio Ohtsubo. On a discrete-time non-zero-sum Dynkin problem with monotonicity. Journal of Applied Probability, 28(2): 466–472, June 1991. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3214881.

Ozaki:2001:IAN

[OI01] Tohru Ozaki and Mitsunori Iino. An innovation approach to non-Gaussian time series analysis. Journal of Applied Probability, 38A:78–92, 2001. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3215870.

Ogata:1986:PPM

[OK86] Yosihiko Ogata and Koichi Katsura. Point-process models with linearly parametrized intensity for application to earthquake data. Journal of Applied Probability, 23A:291–310, 1986. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/3214359.

Ohnishi:1986:OIR

[OKM86] Masamitsu Ohnishi, Hajime Kawai, and Hisashi Mine. An optimal inspection and replacement policy for a deteriorating system. Journal of Applied Probability, 23(4):973–988, December 1986. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3214470.

Odoom:1965:NED

[OL65] S. Odoom and E. H. Lloyd. A note on the equilibrium distribution of levels in a semi-infinite reservoir subject to Markovian inputs and unit withdrawals. Journal of Applied Probability, 2 (1):215–222, June 1965. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3211886.

Olkin:1977:OLK

[Olk77] Ingram Olkin. Obituary: Leo Katz. Journal of Applied Probability, 14(4):891–896, December 1977. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3213364.

Olofsson:1996:GBP

[Olo96] Peter Olofsson. General branching processes with immigration. Journal of Applied Probability, 33(4):940–948, December 1996. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/3214975.

Olofsson:1998:CGB

[Olo98] Peter Olofsson. The x log x condition for general branching processes. Journal of Applied Probability, 35(3):537–544, September 1998. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3215629.

Olshen:1967:APP

[Ols67] Richard A. Olshen. Asymptotic properties of the periodogram of a discrete stationary process. Journal of Applied Probability, 4(3):508–528, December 1967. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.

org/stable/3212218.

Olson:1977:NUB

[Ols77] William H. Olson. Non-uniform breakage-mechanism branching processes and degradation of long-chain polymers. Journal of Applied Probability, 14(1):1–13, March 1977. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3213256.

Olson:1978:MCM

[Ols78] William H. Olson. A Markov chain model for the kinetics of reactant isolation. Journal of Applied Probability, 15(4):835–841, December 1978. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3213438.

Ohno:1972:TLQ

[OM72] Katsuhisa Ohno and Hisashi Mine. Traffic light queues with dependent arrivals as a generalization to queueing theory. Journal of Applied Probability, 9(3):630–641, September 1972. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3212331.

ONeill:1996:SAS

[O’N96] Philip O’Neill. Strong approximations for some open population epidemic models. Journal of Applied Probability, 33(2):448– 457, June 1996. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3215070.

Oertel:2007:UBS

[OO07] Frank Oertel and Mark Owen. On utility-based superreplication prices of contingent claims with unbounded payoffs. Journal of Applied Probability, 44(4):880–888, December 2007. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/27595894.

OCinneide:1986:MMQ

[OP86] C. A. O’Cinneide and P. Purdue. The M/M/∞ queue in a random environment. Journal of Applied Probability, 23(1):175– 184, March 1986. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3214126.

Orive:1979:DCV

[Ori79] Luis-M. Cruz Orive. Distortion of certain Voronoi tessellations when one particle moves. Journal of Applied Probability, 16 (1):95–103, March 1979. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3213378.

Orkin:1972:RMG

[Ork72] Michael Orkin. Recursive matrix games. Journal of Applied Probability, 9(4):813–820, December 1972. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/3212617.

Orriss:1969:EDS

[Orr69] J. Orriss. Equilibrium distributions for systems of chemical reactions with applications to the theory of molecular adsorption. Journal of Applied Probability, 6(3):505–515, December 1969. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/3212098.

Orsingher:1986:PRM

[Ors86] Enzo Orsingher. A planar random motion governed by the twodimensional telegraph equation. Journal of Applied Probability, 23(2):385–397, June 1986. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3214181.

Orsingher:1987:MRF

[Ors87a] Enzo Orsingher. On the maximum of random fields represented by stochastic integrals over circles. Journal of Applied Probability, 24(3):574–585, September 1987. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3214090.

Orsingher:1987:SMS

[Ors87b] Enzo Orsingher. Stochastic motions on the 3-sphere governed by wave and heat equations. Journal of Applied Probability, 24 (2):315–327, June 1987. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3214256.

Orsingher:1989:MGF

[Ors89] Enzo Orsingher. On the maximum of Gaussian Fourier series emerging in the analysis of random vibrations. Journal of Applied Probability, 26(1):182–188, March 1989. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3214329.

Ortisi:2008:PRC

[Ort08] Matteo Ortisi. Polynomial-rate convergence to the stationary state for the continuum-time limit of the minority game. Journal of Applied Probability, 45(2):376–387, June 2008. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/27595951.

Olshen:1970:GU

[OS70] Richard A. Olshen and Leonard J. Savage. A generalized unimodality. Journal of Applied Probability, 7(1):21–34, April 1970. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/3212145.

Osborne:1986:AEF

[OS86] M. R. Osborne and G. K. Smyth. An algorithm for exponential fitting revisited. Journal of Applied Probability, 23A:419–430, 1986. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3214370.

Ott:1991:BPP

[OS91] Teunis J. Ott and J. George Shanthikumar. On a buffer problem for packetized Voice with an N-periodic strongly interchangeable input process. Journal of Applied Probability, 28(3):630–646, September 1991. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3214497.

OHely:2001:AP

[OS01] Martin O’Hely and Aidan Sudbury. The annihilating process. Journal of Applied Probability, 38(1):223–231, March 2001. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/3215753.

Ott:2007:ABG

[OS07] Teunis J. Ott and Jason Swanson. Asymptotic behavior of a generalized TCP congestion avoidance algorithm. Journal of Applied Probability, 44(3):618–635, September 2007. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/27595870.

Olofsson:2014:CMJ

[OS14] Peter Olofsson and Suzanne S. Sindi. A Crump–Mode–Jagers branching process model of prion loss in yeast. Journal of Applied Probability, 51(2):453–465, June 2014. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

projecteuclid.org/euclid.jap/1402578636.

Osawa:1985:RMC

[Ōsa85] Hideo Ōsawa. Reversibility of Markov chains with applications to storage models. Journal of Applied Probability, 22(1):123– 137, March 1985. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3213752.

Osborne:1982:FAR

[Osb82] M. R. Osborne. A finite algorithm for the rank regression problem. Journal of Applied Probability, 19A:241–252, 1982. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3213564.

Osei:1977:SOR

[OT77] G. K. Osei and J. W. Thompson. The supersession of one rumour by another. Journal of Applied Probability, 14(1):127– 134, March 1977. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3213265.

Orsingher:2015:CPB

[OT15] Enzo Orsingher and Bruno Toaldo. Counting processes with Bernstein intertimes and random jumps. Journal of Applied Probability, 52(4):1028–1044, December 2015. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://projecteuclid.org/euclid.jap/1450802751.

Olsson:2016:EAD

[OT16] Fredrik Olsson and Tatyana S. Turova. Ergodicity of agedependent inventory control systems. Journal of Applied Probability, 53(3):688–699, September 2016. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http:/

/projecteuclid.org/euclid.jap/1476370770.

Odor:2021:SMD

[ÓT21] Gergely Ódor and Patrick Thiran. Sequential metric dimension for random graphs. Journal of Applied Probability, 58(4):909–951, December 2021. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL https:/

/www.cambridge.org/core/journals/journal-of-applied-
probability/article/sequential-metric-dimension-for-
random-graphs/A6EA67BB64E7B645F7040384048E76EF.

Ott:1979:SMR

[Ott79] Teunis J. Ott. Some more results for the stable M/G/1 queue in heavy traffic. Journal of Applied Probability, 16(1):187– 197, March 1979. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3213386.

Ott:1984:MGQ

[Ott84a] Teunis J. Ott. On the M/G/1 queue with additional inputs. Journal of Applied Probability, 21(1):129–142, March 1984. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/3213671.

Ott:1984:STD

[Ott84b] Teunis J. Ott. The sojourn-time distribution in the M/G/1 queue with processor sharing. Journal of Applied Probability, 21(2):360–378, June 1984. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3213646.

Ottolini:2020:OOS

[Ott20] Andrea Ottolini. Oscillations for order statistics of some discrete processes. Journal of Applied Probability, 57(3): 703–719, September 2020. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL https:/

/www.cambridge.org/core/journals/journal-of-applied-
probability/article/oscillations-for-order-statistics-
of-some-discrete-processes/3B534CBBD4C6A14959173CC597602471.

Olsen:2003:PSP

[OV03] Tava Lennon Olsen and R. D. Van Der Mei. Polling systems with periodic server routeing in heavy traffic: Distribution of the delay. Journal of Applied Probability, 40(2):305–326, June 2003. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3215793.

Ovaskainen:2001:QDS

[Ova01] Otso Ovaskainen. The quasistationary distribution of the stochastic logistic model. Journal of Applied Probability, 38(4):

898–907, December 2001. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3215772.

Oxley:1979:SPR

[OW79] J. G. Oxley and D. J. A. Welsh. On some percolation results of J. M. Hammersley. Journal of Applied Probability, 16(3):526–540, September 1979. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3213082.

Olivier:1994:EFD

[OW94] C. Olivier and J. Walrand. On the existence of finite-dimensional filters for Markov-modulated traffic. Journal of Applied Probability, 31(2):515–525, June 1994. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3215042.

Owada:2023:LTU

[Owa23] Takashi Owada. Limit theory for U-statistics under geometric and topological constraints with rare events. Journal of Applied Probability, 60(1):314–340, March 2023. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL https://www.cambridge.org/core/journals/
journal-of-applied-probability/article/limit-theory-
for-ustatistics-under-geometric-and-topological-constraints-
with-rare-events/DB2FC7940F955178769216306D98983F.

Ozaki:1980:NLT

[Oza80] T. Ozaki. Non-linear time series models for non-linear random vibrations. Journal of Applied Probability, 17(1):84–93, March 1980. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3212926.

Ozaki:1981:NLT

[Oza81] Tohru Ozaki. Non-linear threshold autoregressive models for non-linear random vibrations. Journal of Applied Probability, 18(2):443–451, June 1981. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3213290.

Ozaki:1986:LGM

[Oza86] Tohru Ozaki. Local Gaussian modelling of stochastic dynamical systems in the analysis of non-linear random vibrations. Journal of Applied Probability, 23A:241–255, 1986. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/3214356.

Pagano:1971:SAP

[Pag71] M. Pagano. Some asymptotic properties of a two-dimensional periodogram. Journal of Applied Probability, 8(4):841–847, December 1971. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3212250.

Pakes:1970:NQS

[Pak70] A. G. Pakes. A note on the queueing systems GI/D/1 and D/G/1. Journal of Applied Probability, 7(2):465–468, August 1970. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3211983.

Pakes:1971:BPI

[Pak71] A. G. Pakes. Branching processes with immigration. Journal of Applied Probability, 8(1):32–42, March 1971. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3211835.

Pakes:1973:CLT

[Pak73a] A. G. Pakes. Conditional limit theorems for a left-continuous random walk. Journal of Applied Probability, 10(1):39–53, March 1973. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3212494.

Pakes:1973:DMI

[Pak73b] A. G. Pakes. On dams with Markovian inputs. Journal of Applied Probability, 10(2):317–329, June 1973. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/3212349.

Pakes:1973:BPM

[Pak73c] A. G. Pakes. On the busy period of the modified GI/G/1 queue. Journal of Applied Probability, 10(1):192–197, March 1973. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/3212506.

Pakes:1974:SGW

[Pak74] A. G. Pakes. On supercritical Galton–Watson processes allowing immigration. Journal of Applied Probability, 11(4):814–817, December 1974. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3212564.

Pakes:1975:TWT

[Pak75] A. G. Pakes. On the tails of waiting-time distributions. Journal of Applied Probability, 12(3):555–564, September 1975. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3212870.

Pakes:1976:SLT

[Pak76] A. G. Pakes. Some limit theorems for a supercritical branching process allowing immigration. Journal of Applied Probability, 13(1):17–26, March 1976. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3212661.

Pakes:1978:ADM

[Pak78a] Anthony G. Pakes. On the age distribution of a Markov chain. Journal of Applied Probability, 15(1):65–77, March 1978. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/3213237.

Pakes:1978:MAT

[Pak78b] Anthony G. Pakes. On the maximum and absorption time of left-continuous random walk. Journal of Applied Probability, 15 (2):292–299, June 1978. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3213402.

Pakes:1983:RMC

[Pak83] Anthony G. Pakes. Remarks on a model of competitive bidding for employment. Journal of Applied Probability, 20(2):349– 357, June 1983. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3213807.

Pakes:1987:RMM

[Pak87] Anthony G. Pakes. Remarks on the maxima of a martingale sequence with application to the simple critical branching process. Journal of Applied Probability, 24(3):768–772, September 1987. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3214106.

Pakes:1993:RLD

[Pak93] Anthony G. Pakes. Remarks on the Luria–Delbrück distribution. Journal of Applied Probability, 30(4):991–994, December 1993. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3214530.

Pakes:2004:CEI

[Pak04] Anthony G. Pakes. Convolution equivalence and infinite divisibility. Journal of Applied Probability, 41(2):407–424, June 2004. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/3216025. See

corrections [Pak07].

Pakes:2007:CEI

[Pak07] Anthony G. Pakes. Convolution equivalence and infinite divisibility: Corrections and corollaries. Journal of Applied Probability, 44(2):295–305, June 2007. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/27595842. See [Pak04].

Pakes:2008:TSR

[Pak08] Anthony G. Pakes. Tails of stopped random products: The factoid and some relatives. Journal of Applied Probability, 45 (4):1161–1180, December 2008. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.

org/stable/27596015.

Pakkanen:2010:SIC

[Pak10] Mikko S. Pakkanen. Stochastic integrals and conditional full support. Journal of Applied Probability, 47(3):650–667, September 2010. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/20778508.

Papoulis:1971:HDS

[Pap71] A. Papoulis. High density shot noise and Gaussianity. Journal of Applied Probability, 8(1):118–127, March 1971. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3211842.

Papangelou:1994:DBA

[Pap94] F. Papangelou. On a distributional bound arising in autoregressive model fitting. Journal of Applied Probability, 31(2):401– 408, June 1994. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3215033.

Papangelou:1996:LDB

[Pap96] F. Papangelou. Large deviations and the Bayesian estimation of higher-order Markov transition functions. Journal of Applied Probability, 33(1):18–27, March 1996. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3215260.

Papavasiliou:2005:UCA

[Pap05] Anastasia Papavasiliou. A uniformly convergent adaptive particle filter. Journal of Applied Probability, 42(4):1053–1068, December 2005. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
30040828.

Papangelou:2010:SMR

[Pap10] F. Papangelou. A simple model for random oscillations. Journal of Applied Probability, 47(4):1164–1173, December 2010. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/29777434.

Papadatos:2017:SEM

[Pap17] Nickos Papadatos. On sequences of expected maxima and expected ranges. Journal of Applied Probability, 54(4): 1144–1166, December 2017. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL https:/

/www.cambridge.org/core/journals/journal-of-applied-
probability/article/on-sequences-of-expected-maxima-
and-expected-ranges/C877BDE36F5AA88DB3302AD254C157F2.

Park:1975:LIL

[Par75] W. J. Park. The law of the iterated logarithm for Brownian sheets. Journal of Applied Probability, 12(4):840–844, December 1975. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3212737.

Park:1985:BNJ

[Par85] C. Park. Boundaries with negative jumps for the Brownian motion. Journal of Applied Probability, 22(4):964–970, December 1985. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3213966.

Parzen:1986:QSA

[Par86] Emanuel Parzen. Quantile spectral analysis and long-memory time series. Journal of Applied Probability, 23A:41–54, 1986. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/3214341.

Parthasarathy:1988:OAN

[Par88a] K. R. Parthasarathy. Obituary: Andrei Nikolaevich Kolmogorov. Journal of Applied Probability, 25(2):445–450, June 1988. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3214455.

Parthasarathy:1988:PRW

[Par88b] K. R. Parthasarathy. The passage from random walk to diffusion in quantum probability. Journal of Applied Probability, 25A: 151–166, 1988. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3214153.

Parmigiani:1994:ITS

[Par94] Giovanni Parmigiani. Inspection times for stand-by units. Journal of Applied Probability, 31(4):1015–1025, December 1994. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/3215325.

Parthasarathy:2001:ICQ

[Par01] K. R. Parthasarathy. An inducing construction of quantum codes from classical error correcting codes. Journal of Applied Probability, 38A:27–32, 2001. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.

org/stable/3215864.

Patie:2013:AOU

[Pat13] P. Patie. Asian options under one-sided Lévy models. Journal of Applied Probability, 50(2):359–373, June 2013. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://projecteuclid.org/euclid.jap/1371648946.

Pauwels:1987:SFP

[Pau87] E. J. Pauwels. Smooth first-passage densities for onedimensional diffusions. Journal of Applied Probability, 24(2): 370–377, June 1987. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3214261.

Paulsen:1997:MF

[Pau97] Volkert Paulsen. The moments of FIND. Journal of Applied Probability, 34(4):1079–1082, December 1997. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3215021.

Paulsen:2001:BAP

[Pau01] V. Paulsen. Bounds for the American perpetual put on a stock index. Journal of Applied Probability, 38(1):55–66, March 2001. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/3215741.

Pawlowski:1989:PTN

[Paw89] Janusz Pawlowski. Poisson theorem for non-homogeneous Markov chains. Journal of Applied Probability, 26(3):637–642, September 1989. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3214421.

Park:1983:SBW

[PB83] C. Park and J. A. Beekman. Stochastic barriers for the Wiener process. Journal of Applied Probability, 20(2):338–348, June 1983. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3213806.

Parvardeh:2014:CRL

[PB14] A. Parvardeh and N. Balakrishnan. On the conditional residual life and inactivity time of coherent systems. Journal of Applied Probability, 51(4):990–998, December 2014. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

projecteuclid.org/euclid.jap/1421763323.

Pitzen:2020:RCI

[PB20] S. Pitzen and M. Burkschat. Representations of component importance for coherent systems with exchangeable components.

Journal of Applied Probability, 57(2):385–406, June 2020. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL https://www.cambridge.org/core/journals/journal-
of-applied-probability/article/representations-of-component-
importance-for-coherent-systems-with-exchangeable-components/
B21A380AC6B9AB91A7608228E5004DA6.

Parker:1976:SPL

[PC76] Philip Parker and Richard Cowan. Some properties of line segment processes. Journal of Applied Probability, 13(1):96– 107, March 1976. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3212669.

Pinedo:1985:IBS

[PC85] Michael Pinedo and Zvi Chechner. Inequalities and bounds for the scheduling of stochastic jobs on parallel machines. Journal of Applied Probability, 22(3):739–744, September 1985. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3213879.

Plazzotta:2016:AFS

[PC16] Giacomo Plazzotta and Caroline Colijn. Asymptotic frequency of shapes in supercritical branching trees. Journal of Applied Probability, 53(4):1143–1155, December 2016. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://projecteuclid.org/euclid.jap/1481132842.

Pavithra:2020:MFS

[PD20]  Celeste R. Pavithra and T. G. Deepak.  Multivariate

finite-support phase-type distributions. Journal of Applied Probability, 57(4):1260–1275, December 2020. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL https://www.cambridge.org/core/journals/
journal-of-applied-probability/article/multivariate-
finitesupport-phasetype-distributions/D21901290A596B70F983E2A2FBE9F3EC.

Pearce:1965:QSN

[Pea65] C. Pearce. A queueing system with non-recurrent input and batch servicing. Journal of Applied Probability, 2(2):442–448, December 1965. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3212204.

Pearce:1967:QMA

[Pea67] C. Pearce. Queues with moving average service times. Journal of Applied Probability, 4(3):553–570, December 1967. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3212221.

Pedler:1971:OTT

[Ped71] P. J. Pedler. Occupation times for two state Markov chains. Journal of Applied Probability, 8(2):381–390, June 1971. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/3211908.

Pedersen:2000:DOS

[Ped00] Jesper Lund Pedersen. Discounted optimal stopping problems for the maximum process. Journal of Applied Probability, 37(4): 972–983, December 2000. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3215488.

Pedersen:2002:POU

[Ped02] Jan Pedersen. Periodic Ornstein–Uhlenbeck processes driven by Lévy processes. Journal of Applied Probability, 39(4):748–763, December 2002. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3216000.

Pegg:1973:SRI

[Peg73] P. A. Pegg. Some ratio identities for a class of skip free random walks. Journal of Applied Probability, 10(1):213–217, March 1973. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3212509.

Pegram:1975:MMT

[Peg75a] G. G. S. Pegram. A multinomial model for transition probability matrices. Journal of Applied Probability, 12(3):498–506, September 1975. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3212864.

Pegram:1975:RTD

[Peg75b] G. G. S. Pegram. Recurrence times of draft-patterns from reservoirs. Journal of Applied Probability, 12(3):647–652, September 1975. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3212886.

Pegram:1980:AMM

[Peg80] G. G. S. Pegram. An autoregressive model for multilag Markov chains. Journal of Applied Probability, 17(2):350–362, June 1980. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/3213025.

Pekoz:1996:SMG

[Pek96] Erol A. Peköz. Stein’s method for geometric approximation. Journal of Applied Probability, 33(3):707–713, September 1996. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/3215352.

Pekoz:1999:NRB

[Pek99] Erol A. Peköz. A note on reliability bounds via conditional inequalities. Journal of Applied Probability, 36(4):1251–1254, December 1999. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3215595.

Pekoz:2003:SMB

[Pek03] Erol A. Peköz. Some memoryless bandit policies. Journal of Applied Probability, 40(1):250–256, March 2003. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3215852.

Pekoz:2006:CPA

[Pek06] Erol A. Peköz. A compound Poisson approximation inequality. Journal of Applied Probability, 43(1):282–288, March 2006. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/27595727.

Pellerey:1994:SMU

[Pel94] Franco Pellerey. Shock models with underlying counting process. Journal of Applied Probability, 31(1):156–166, March 1994. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/3215243.

Pellerey:2007:CRB

[Pel07] Franco Pellerey. Comparison results for branching processes in random environments. Journal of Applied Probability, 44(1):

142–150, March 2007. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/27595827.

Peng:1996:SRT

[Pen96a] Nan Fu Peng. Spectral representations of the transition probability matrices for continuous time finite Markov chains. Journal of Applied Probability, 33(1):28–33, March 1996. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3215261.

Penrose:1996:SEL

[Pen96b] Mathew D. Penrose. Spatial epidemics with large finite range. Journal of Applied Probability, 33(4):933–939, December 1996. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/3214974.

Penrose:2016:SGR

[Pen16] Mathew D. Penrose. The strong giant in a random digraph. Journal of Applied Probability, 53(1):57–70, March 2016. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://projecteuclid.org/euclid.jap/1457470558.

Penrose:2018:IRG

[Pen18] Mathew D. Penrose. Inhomogeneous random graphs, isolated vertices, and Poisson approximation. Journal of Applied Probability, 55(1):112–136, March 2018. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL https://www.cambridge.org/core/journals/journal-
of-applied-probability/article/inhomogeneous-random-
graphs-isolated-vertices-and-poisson-approximation/B08892B47709552306A27AD20C87AB1F.

Peralta:2023:MJP

[Per23] Oscar Peralta. A Markov jump process associated with the matrix-exponential distribution. Journal of Applied Probability, 60(1):1–13, March 2023. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL https:/

/www.cambridge.org/core/journals/journal-of-applied-
probability/article/markov-jump-process-associated-with-
the-matrixexponential-distribution/E3386C4A89EEC7A54B5BC304DF375C6B.

Pestalozzi:1968:QMD

[Pes68] Gerold Pestalozzi. A queue with Markov-dependent service times. Journal of Applied Probability, 5(2):461–466, August 1968. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3212266.

Petruccelli:1981:BCP

[Pet81] Joseph D. Petruccelli. Best-choice problems involving uncertainty of selection and recall of observations. Journal of Applied Probability, 18(2):415–425, June 1981. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/3213287.

Petruccelli:1983:BCP

[Pet83] Joseph D. Petruccelli. On the best-choice problem when the number of observations is random. Journal of Applied Probability, 20(1):165–171, March 1983. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3213731.

Pereira:2006:LCP

[PF06] L. Pereira and H. Ferreira. Limiting crossing probabilities of random fields. Journal of Applied Probability, 43(3):884–891, September 2006. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
27595781.

Pfeifer:1982:CED

[Pfe82a] Dietmar Pfeifer. Characterizations of exponential distributions by independent non-stationary record increments. Journal of Applied Probability, 19(1):127–135, March 1982. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/3213922. See correction

[Ano82g].

Pfeifer:1982:SEP

[Pfe82b] Dietmar Pfeifer. The structure of elementary pure birth processes. Journal of Applied Probability, 19(3):664–667, September 1982. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3213524.

Pfeifer:1989:EPS

[Pfe89] Dietmar Pfeifer. Extremal processes, secretary problems and the 1/e law. Journal of Applied Probability, 26(4):722–733, December 1989. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3214377.

Puri:1979:PMQ

[PG79] Prem S. Puri and Charles M. Goldie. Poisson mixtures and quasi-infinite divisibility of distributions. Journal of Applied Probability, 16(1):138–153, March 1979. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http:/

/www.jstor.org/stable/3213382. See corrections [Ano79f].

Pihlsgaard:2013:DST

[PG13] Mats Pihlsgård and Peter W. Glynn. On the dynamics of semimartingales with two reflecting barriers. Journal of Applied Probability, 50(3):671–685, September 2013. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

projecteuclid.org/euclid.jap/1378401229.

Pakes:1982:OEC

[PH82] Anthony G. Pakes and C. C. Heyde. Optimal estimation of the criticality parameter of a supercritical branching process having random environments. Journal of Applied Probability, 19 (2):415–420, June 1982. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3213493.

Priestley:1986:INL

[PH86] M. B. Priestley and S. M. Heravi. Identification of non-linear systems using general state-dependent models. Journal of Applied Probability, 23A:257–272, 1986. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3214357.

Pfeifer:1987:MCM

[PH87] Dietmar Pfeifer and Ursula Heller. A martingale characterization of mixed Poisson processes. Journal of Applied Probability, 24(1):246–251, March 1987. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3214076.

Papastavridis:1988:FRC

[PH88] S. Papastavridis and I. Hadzichristos. Formulas for the reliability of a consecutive-k-out-of-n: F system. Journal of Applied Probability, 25(4):772–779, December 1988. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/3214298.

Phatarfod:1969:NFE

[Pha69] R. M. Phatarfod. A note on the first emptiness problem of a finite dam with Poisson type inputs. Journal of Applied Probability, 6(1):227–230, April 1969. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3212292.

Phatarfod:1974:NRC

[Pha74] R. M. Phatarfod. Note on the reversible counters system of Lampard. Journal of Applied Probability, 11(3):624–628, September 1974. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3212712.

Phatarfod:1975:RCS

[Pha75] R. M. Phatarfod. On the reversible counters system of Lampard. Journal of Applied Probability, 12(3):639–646, September 1975. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3212885.

Phatarfod:1979:NID

[Pha79] R. M. Phatarfod. A note on the infinitely deep dam with a Markovian input. Journal of Applied Probability, 16(4):917–922, December 1979. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3213158.

Phatarfod:1988:SWS

[Pha88] Ravindra M. Phatarfod. Sums and weighted sums of a gamma Markov sequence. Journal of Applied Probability, 25(1):204– 209, March 1988. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3214246.

Phatarfod:1991:MOL

[Pha91] R. M. Phatarfod. On the matrix occurring in a linear search problem. Journal of Applied Probability, 28(2):336–346, June 1991. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3214870.

Phatarfod:1993:GLD

[Pha93] R. M. Phatarfod. The geometricity of the limiting distributions in queues and dams. Journal of Applied Probability, 30(2):438– 445, June 1993. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3214852.

Phatarfod:1994:TPM

[Pha94] R. M. Phatarfod. On the transition probabilities of the moveto-front scheme. Journal of Applied Probability, 31(2):570– 574, June 1994. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3215049.

Phelan:1996:BDB

[Phe96a] Michael J. Phelan. Birth and death on a Brownian flow: A Feller semigroup and its generator and a martingale problem. Journal of Applied Probability, 33(1):71–87, March 1996. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/3215265.

Phelan:1996:GTB

[Phe96b] Michael J. Phelan. A Girsanov transformation for birth and death on a Brownian flow. Journal of Applied Probability, 33 (1):88–100, March 1996. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3215266.

Phillips:1981:CNE

[Phi81] M. J. Phillips. A characterization of the negative exponential distribution with application to reliability theory. Journal of Applied Probability, 18(3):652–659, September 1981. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3213319.

Phillips:2001:DVJ

[Phi01] Brian Phillips. David Vere-Jones’s influence on statistical education. Journal of Applied Probability, 38A:6–19, 2001. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3215862.

Piau:2000:ABR

[Pia00a] Didier Piau. Addendum to ‘Quasi-Renewal Estimates’ for modelling DNA replication. Journal of Applied Probability, 37(4): 1171–1172, December 2000. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor. org/stable/3215509. See comment [Pia00b].

Piau:2000:QRE

[Pia00b] Didier Piau. Quasi-renewal estimates. Journal of Applied Probability, 37(1):269–275, March 2000. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3215675. See addendum [Pia00a].

Piau:2008:APB

[Pia08] Didier Piau. Asymptotics of posteriors for binary branching processes. Journal of Applied Probability, 45(3):727–742, September 2008. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/27595983.

Piau:2009:AIB

[Pia09] Didier Piau. Asymptotics of iterated branching processes. Journal of Applied Probability, 46(3):917–924, September 2009. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/25662468.

Pickands:1971:TDP

[Pic71] James Pickands III. The two-dimensional Poisson process and extremal processes. Journal of Applied Probability, 8(4):745–756, December 1971. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3212238.

Pickard:1976:AII

[Pic76] D. K. Pickard. Asymptotic inference for an Ising lattice. Journal of Applied Probability, 13(3):486–497, September 1976. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3212468.

Pickard:1977:CBL

[Pic77] D. K. Pickard. A curious binary lattice process. Journal of Applied Probability, 14(4):717–731, December 1977. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3213345.

Pickard:1979:AII

[Pic79] David K. Pickard. Asymptotic inference for an Ising lattice. III. Non-zero field and ferromagnetic states. Journal of Applied Probability, 16(1):12–24, March 1979. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3213370.

Picard:1980:AMT

[Pic80] Philippe Picard. Applications of martingale theory to some epidemic models. Journal of Applied Probability, 17(3):583–599, September 1980. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3212953.

Pickard:1982:IGI

[Pic82a] David K. Pickard. Inference for general Ising models. Journal of Applied Probability, 19A:345–357, 1982. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/3213574.

Pickard:1982:INN

[Pic82b] David K. Pickard. Isolated nearest neighbors. Journal of Applied Probability, 19(2):444–449, June 1982. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3213499.

Picard:1984:AMT

[Pic84] Philippe Picard. Applications of martingale theory to some epidemic models, II. Journal of Applied Probability, 21(4):677–684, December 1984. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3213686.

Pievatolo:2007:DDA

[Pie07] Antonio Pievatolo. The downtime distribution after a failure of a system with multistate independent components. Journal of Applied Probability, 44(4):1068–1077, December 2007. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/27595908.

Pihlsgaard:2015:LMT

[Pih15] Mats Pihlsgård. Local martingales with two reflecting barriers. Journal of Applied Probability, 52(4):1062–1075, December 2015. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://projecteuclid.org/euclid.jap/
1450802753.

Pillai:1991:SPP

[Pil91] R. N. Pillai. Semi-Pareto processes. Journal of Applied Probability, 28(2):461–465, June 1991. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3214880. See correction [Ebr90].

Pimentel:2014:LMC

[Pim14] Leandro P. R. Pimentel. On the location of the maximum of a continuous stochastic process. Journal of Applied Probability, 51 (1):152–161, March 2014. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://projecteuclid.

org/euclid.jap/1395771420.

Pinsky:1971:RMP

[Pin71] Mark Pinsky. Recurrence for Markov processes on N lines. Journal of Applied Probability, 8(4):724–730, December 1971. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/3212236.

Pinedo:1980:SSE

[Pin80] Michael Pinedo. Scheduling spares with exponential lifetimes in a two-component parallel system. Journal of Applied Probability, 17(4):1025–1032, December 1980. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3213212.

Pinelis:2011:ELB

[Pin11] Iosif Pinelis. Exact lower bounds on the exponential moments of truncated random variables. Journal of Applied Probability, 48(2):547–560, June 2011. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/23065901.

Pirlot:1980:SVP

[Pir80] Marc Pirlot. A strong variational principle for continuous spin systems. Journal of Applied Probability, 17(1):47–58, March 1980. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3212923.

Pistorius:2006:MLP

[Pis06] Martijn Pistorius. On maxima and ladder processes for a dense class of Lévy process. Journal of Applied Probability, 43(1):208– 220, March 2006. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
27595721.

Pittenger:1977:NIP

[Pit77] A. O. Pittenger. A note on the inverse problem for reducible Markov chains. Journal of Applied Probability, 14(3):621–625, September 1977. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3213467.

Pittenger:1980:SPS

[Pit80] A. O. Pittenger. Success probabilities for second guessers. Journal of Applied Probability, 17(4):1133–1137, December 1980. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/3213227.

Pittel:1987:UMC

[Pit87] B. Pittel. An urn model for cannibal behavior. Journal of Applied Probability, 24(2):522–526, June 1987. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/3214275.

Pittel:1990:DKM

[Pit90] B. Pittel. On a daley-Kendall model of random rumours. Journal of Applied Probability, 27(1):14–27, March 1990. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3214592.

Pittel:2018:RQF

[Pit18] Boris Pittel. On random quadratic forms: supports of potential local maxima. Journal of Applied Probability, 55 (4):1113–1130, December 2018. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL https:/

/www.cambridge.org/core/journals/journal-of-applied-
probability/article/on-random-quadratic-forms-supports-
of-potential-local-maxima/1488B8263FF1089B83B797AD00BAC1A0.

Pekoz:2002:PTF

[PJ02] Erol A. Peköz and Nitindra Joglekar. Poisson traffic flow in a general feedback queue. Journal of Applied Probability, 39(3): 630–636, September 2002. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3216084.

Pakes:1974:SBH

[PK74] A. G. Pakes and Norman Kaplan. On the subcritical Bellman– Harris process with immigration. Journal of Applied Probability, 11(4):652–668, December 1974. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.

org/stable/3212549.

Peterson:1980:RDD

[PK80] Arthur V. Peterson, Jr. and Richard A. Kronmal. A representation for discrete distributions by equiprobable mixtures. Journal of Applied Probability, 17(1):102–111, March 1980. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3212928.

Packham:2016:ABM

[PKO16] N. Packham, M. Kalkbrener, and L. Overbeck. Asymptotic behaviour of multivariate default probabilities and default correlations under stress. Journal of Applied Probability, 53(1):71–81, March 2016. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-

6072 (electronic). URL http://projecteuclid.org/euclid.
jap/1457470559.

Proudfoot:1972:RWP

[PL72] A. D. Proudfoot and D. G. Lampard. A random walk problem with correlation. Journal of Applied Probability, 9(2):436– 440, June 1972. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3212813.

Proudfoot:1973:DPO

[PL73] A. D. Proudfoot and D. G. Lampard. The divided pooled output of several strictly-periodic point processes. Journal of Applied Probability, 10(2):461–463, June 1973. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3212365.

Purdue:1981:ISQ

[PL81] P. Purdue and D. Linton. An infinite-server queue subject to an extraneous phase process and related models. Journal of Applied Probability, 18(1):236–244, March 1981. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http:/

/www.jstor.org/stable/3213183.

Picard:2003:PRE

[PL03] Philippe Picard and Claude Lefèvre. Probabilité de ruine Éventuelle dans un modèle de risque à temps discret. (French) [Probability of eventual ruin in a discrete-time risk model]. Journal of Applied Probability, 40(3):543–556, September 2003. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/3215935.

Pan:2017:ICO

[PL17] Xiaoqing Pan and Xiaohu Li. Increasing convex order on generalized aggregation of SAI random variables with applications. Journal of Applied Probability, 54(3):685–700, September 2017. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL https://www.cambridge.org/core/journals/
journal-of-applied-probability/article/increasing-convex-
order-on-generalized-aggregation-of-sai-random-variables-
with-applications/08071F4DB22A52171A799BBAF58A0C57.

Platz:1983:CLT

[Pla83] Ole Platz. A central limit theorem for a security of supply model. Journal of Applied Probability, 20(1):220–225, March 1983. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/3213740.

Platz:1985:SUP

[Pla85] Ole Platz. A sharp upper probability bound for the occurrence of at least m out of n events. Journal of Applied Probability, 22(4): 978–981, December 1985. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3213968.

Platen:2004:CCB

[Pla04] Eckhard Platen. A class of complete benchmark models with intensity-based jumps. Journal of Applied Probability, 41(1):19– 34, March 2004. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3215812.

Picard:2003:PRT

[PLC03] Philippe Picard, Claude Lefèvre, and Ibrahim Coulibaly. Problèmes de ruine en Théorie du risque à temps discret avec horizon fini. (French) [Problems in ruin theory of risk with finite discrete time horizon]. Journal of Applied Probability, 40(3): 527–542, September 2003. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3215934.

Pliska:1975:DPM

[Pli75] Stanley R. Pliska. A diffusion process model for the optimal operation of a reservoir system. Journal of Applied Probability, 12(4):859–863, December 1975. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.

org/stable/3212740.

Phi:2023:KDC

[PLRB23] Tien Cuong Phi, Eva Löcherbach, and Patricia Reynaud-Bouret. Kalikow decomposition for counting processes with stochastic intensity and application to simulation algorithms. Journal of Applied Probability, 60(4):1469–1500, December 2023. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL https://www.cambridge.org/core/journals/journal-
of-applied-probability/article/kalikow-decomposition-
for-counting-processes-with-stochastic-intensity-and-
application-to-simulation-algorithms/570156637E02C52762C8B0E344ADEFAA.

Phatarfod:1973:SRD

[PM73] R. M. Phatarfod and K. V. Mardia. Some results for dams with Markovian inputs. Journal of Applied Probability, 10(1):166– 180, March 1973. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3212504.

Perrin:1999:INS

[PM99] Olivier Perrin and Wendy Meiring. Identifiability for nonstationary spatial structure. Journal of Applied Probability, 36 (4):1244–1250, December 1999. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.

org/stable/3215594.

Perrin:2003:NSO

[PM03] Olivier Perrin and Wendy Meiring. Nonstationarity in R^{n} is second-order stationarity in R^{2n}. Journal of Applied Probability, 40(3):815–820, September 2003. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.

org/stable/3215956.

Podkaminer:1972:DNC

[Pod72] Leon Podkaminer. On the distribution of the number of cyclically occurring random events. Journal of Applied Probability, 9(3):681–683, September 1972. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.

org/stable/3212338.

Pollak:1966:SGS

[Pol66] Edward Pollak. On the survival of a gene in a subdivided population. Journal of Applied Probability, 3(1):142–155, June 1966. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/3212043.

Pollard:1967:HPM

[Pol67] J. H. Pollard. Hierarchical population models with Poisson recruitment. Journal of Applied Probability, 4(1):209–213, April 1967. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3212315.

Pollak:1968:RGD

[Pol68] Edward Pollak. On random genetic drift in a subdivided population. Journal of Applied Probability, 5(2):314–333, August 1968. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3212255.

Pollak:1969:BCB

[Pol69] Edward Pollak. Bounds for certain branching processes. Journal of Applied Probability, 6(1):201–204, April 1969. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3212287.

Pollak:1971:SPE

[Pol71] Edward Pollak. On survival probabilities and extinction times for some branching processes. Journal of Applied Probability, 8 (4):633–654, December 1971. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.

org/stable/3212229.

Pollaczek:1975:OSP

[Pol75] Felix Pollaczek. Order statistics of partial sums of mutually independent random variables. Journal of Applied Probability, 12(2):390–395, June 1975. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/ stable/3212456. See correction [Ano75e].

Pollak:1978:ENV

[Pol78] Moshe Pollak. The expected number of visits of a sequence of normal partial sums to each side of a boundary. Journal of Applied Probability, 15(1):192–198, March 1978. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3213251.

Pollard:1982:BHA

[Pol82] David Pollard. Beyond the heuristic approach to Kolmogorov– Smirnov theorems. Journal of Applied Probability, 19A:359–365, 1982. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3213575.

Pollett:1990:NCP

[Pol90] P. K. Pollett. A note on the classification of Q-processes when Q is not regular. Journal of Applied Probability, 27(2):278– 290, June 1990. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3214647.

Pollett:1994:ICT

[Pol94] P. K. Pollett. On the identification of continuous-time Markov chains with a given invariant measure. Journal of Applied Probability, 31(4):897–910, December 1994. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3215315.

Pollett:1999:QDC

[Pol99] P. K. Pollett. Quasistationary distributions for continuous time Markov chains when absorption is not certain. Journal of Applied Probability, 36(1):268–272, March 1999. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3215420.

Pollett:2001:SMC

[Pol01] P. K. Pollett. Similar Markov chains. Journal of Applied Probability, 38A:53–65, 2001. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3215868.

Pollak:2010:CTM

[Pol10] Edward Pollak. Coalescent theory for a monoecious random mating population with a varying size. Journal of Applied Probability, 47(1):41–57, March 2010. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/25662499.

Pontiggia:2007:NSR

[Pon07] Laura Pontiggia. Nonconstant sum red-and-black games with bet-dependent win probability function. Journal of Applied Probability, 44(2):547–553, June 2007. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/27595860.

Posfai:2009:PAP

[Pós09] Anna Pósfai. Poisson approximation in a Poisson limit theorem inspired by coupon collecting. Journal of Applied Probability, 46(2):585–592, June 2009. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/25662445.

Pourbabai:1990:NDG

[Pou90] Behnam Pourbabai. A note on a D/G/K loss system with retrials. Journal of Applied Probability, 27(2):385–392, June 1990. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/3214657.

Pourahmadi:1992:API

[Pou92] Mohsen Pourahmadi. Alternating projections and interpolation of stationary processes. Journal of Applied Probability, 29(4): 921–931, December 1992. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3214724.

Percus:1969:DNZ

[PP69] Ora Engelberg Percus and Jerome K. Percus. On the distribution of the number of zero partial sums. Journal of Applied Probability, 6(1):162–176, April 1969. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3212283.

Percus:1970:ECC

[PP70] Ora Engleberg Percus and Jerome K. Percus. Extended criterion for comparison of empirical distributions. Journal of Applied Probability, 7(1):1–20, April 1970. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3212144.

Percus:1971:GFC

[PP71] Ora Engelberg Percus and Jerome K. Percus. Generating functions for a class of one-dimensional random walks. Journal of Applied Probability, 8(1):110–117, March 1971. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/3211841.

Paranjape:1973:DST

[PP73a] S. R. Paranjape and C. Park. Distribution of the supremum of the two-parameter Yeh–Wiener process on the boundary. Journal of Applied Probability, 10(4):875–880, December 1973. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/3212390.

Pledger:1973:SCR

[PP73b] Gordon Pledger and Frank Proschan. Stochastic comparisons of random processes, with applications in reliability. Journal of Applied Probability, 10(3):572–585, September 1973. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3212777.

Pegg:1977:DAI

[PP77] P. A. Pegg and R. M. Phatarfod. Dams with additive inputs revisited. Journal of Applied Probability, 14(2):367–374, June 1977. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3213007.

Percus:1988:CTR

[PP88] O. E. Percus and J. K. Percus. Coin tossing, revisited. Journal of Applied Probability, 25(1):70–80, March 1988. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3214234.

Phatarfod:1998:SMR

[PP98] R. M. Phatarfod and A. J. Pryde. On some multi-request move-to-front heuristics. Journal of Applied Probability, 35(4): 911–918, December 1998. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3215527.

Pitts:2007:JDS

[PP07] Susan M. Pitts and Konstadinos Politis. The joint density of the surplus before and after ruin in the Sparre Andersen model. Journal of Applied Probability, 44(3):695–712, September 2007. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/27595876.

Percus:2014:MSN

[PP14] Ora E. Percus and Jerome K. Percus. The maximum of a symmetric next neighbor walk on the nonnegative integers. Journal of Applied Probability, 51(1):162–173, March 2014. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://projecteuclid.org/euclid.jap/1395771421.

Percus:1999:APD

[PPBT99] J. K. Percus, O. E. Percus, W. J. Bruno, and D. C. Torney. Asymptotics of pooling design performance. Journal of Applied Probability, 36(4):951–964, December 1999. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/3215569.

Phatarfod:1997:MFS

[PPD97] R. M. Phatarfod, A. J. Pryde, and David Dyte. On the moveto-front scheme with Markov dependent requests. Journal of Applied Probability, 34(3):790–794, September 1997. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3215104.

Park:2003:CSP

[PPF03] Byung Chun Park, Jae Young Park, and Robert D. Foley. Carousel system performance. Journal of Applied Probability, 40 (3):602–612, September 2003. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.

org/stable/3215938.

Posner:1967:UFA

[PR67] Edward C. Posner and Howard Rumsey, Jr. An uncertainty function arising in sequential search theory. Journal of Applied Probability, 4(1):137–143, April 1967. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3212306.

Pekoz:1995:SDE

[PR95] Erol A. Peköz and Sheldon M. Ross. A simple derivation of exact reliability formulas for linear and circular consecutive-k-of-n: F systems. Journal of Applied Probability, 32(2):554–557, June 1995. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3215309.

Platen:1996:PMF

[PR96] Eckhard Platen and Rolando Rebolledo. Principles for modelling financial markets. Journal of Applied Probability, 33(3):601–613, September 1996. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3215342.

Pekoz:2008:RTC

[PR08] Erol A. Peköz and Sheldon M. Ross. Relating time and customer averages for queues using ’forward’ coupling from the past. Journal of Applied Probability, 45(2):568–574, June 2008. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/27595965.

Puccetti:2013:SBS

[PR13] Giovanni Puccetti and Ludger Rüschendorf. Sharp bounds for sums of dependent risks. Journal of Applied Probability, 50 (1):42–53, March 2013. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://projecteuclid.

org/euclid.jap/1363784423.

Prabhu:1964:TDR

[Pra64] N. U. Prabhu. Time-dependent results in storage theory. Journal of Applied Probability, 1(1):1–46, June 1964. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/3212059.

Prabhu:1968:SNR

[Pra68] N. U. Prabhu. Some new results in storage theory. Journal of Applied Probability, 5(2):452–460, August 1968. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3212265.

Prabhu:1970:LTS

[Pra70] N. U. Prabhu. Limit theorems for the single server queue with traffic intensity one. Journal of Applied Probability, 7(1):227– 233, April 1970. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3212162.

Prajneshu:1980:TDS

[Pra80] Prajneshu. Time-dependent solution of the logistic model for population growth in random environment. Journal of Applied Probability, 17(4):1083–1086, December 1980. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3213218.

Prasad:1982:CBT

[Pra82] V. Rajendra Prasad. Comment on “Scheduling Tasks with Exponential Service Times on Non-Identical Processors to Minimize Various Cost Functions”. Journal of Applied Probability, 19(3):741, September 1982. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3213539. See [WP80].

Prabhu:1988:TSP

[Pra88] N. U. Prabhu. Theory of semiregenerative phenomena. Journal of Applied Probability, 25A:257–274, 1988. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/3214161.

Pekoz:2009:TPB

[PRCS09] Erol A. Peköz, Adrian Röllin, Vydas Cekanavicius, and Michael Shwartz. A three-parameter binomial approximation. Journal of Applied Probability, 46(4):1073–1085, December 2009. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/25662481.

Prekopa:1994:ECE

[Pré94] András Prékopa. Estimation of cause-effect relationship under noise. Journal of Applied Probability, 31A:343–350, 1994. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/3214967.

Preater:1997:MMT

[Pre97a] J. Preater. M/M/∞ transience revisited. Journal of Applied Probability, 34(4):1061–1067, December 1997. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3215018.

Preater:1997:PMM

[Pre97b] J. Preater. A perpetuity and the M/M/∞ ranked server system. Journal of Applied Probability, 34(2):508–513, June 1997. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/3215389.

Preater:2002:SMM

[Pre02] J. Preater. On the severity of M/M/∞ congested episodes. Journal of Applied Probability, 39(1):228–230, March 2002. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3215930.

Preater:2006:LSA

[Pre06] J. Preater. On-line selection of an acceptable pair. Journal of Applied Probability, 43(3):729–740, September 2006. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/27595768.

Prieto-Rumeau:2012:DCT

[PRHL12] Tomás Prieto-Rumeau and Onésimo Hernández-Lerma. Discounted continuous-time controlled Markov chains: Convergence of control models. Journal of Applied Probability, 49(4):1072– 1090, December 2012. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/41713829.

Privault:2000:HTS

[Pri00] Nicolas Privault. Hypothesis testing and Skorokhod stochastic integration. Journal of Applied Probability, 37(2):560–574, June 2000. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3215730.

Privault:2019:MHC

[Pri19] Nicolas Privault. Moments of k-hop counts in the randomconnection model. Journal of Applied Probability, 56(4): 1106–1121, December 2019. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL https:/

/www.cambridge.org/core/journals/journal-of-applied-
probability/article/moments-of-khop-counts-in-the-randomconnection-
model/4230AD542A7A96B4DCE514D95BB5FB50.

Prorok:1982:BRT

[Pro82] Philip C. Prorok. Bounded recurrence times and lead time in the design of a repetitive screening program. Journal of Applied Probability, 19(1):10–19, March 1982. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3213911. See correction [Ano82f].

Prodinger:1996:ALD

[Pro96] Helmut Prodinger. Asymptotics of the Luria–Delbrück distribution via singularity analysis. Journal of Applied Probability, 33 (1):282–283, March 1996. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3215284.

Profeta:2018:RPS

[Pro18] Christophe Profeta. A remark on positive sojourn times of symmetric processes. Journal of Applied Probability, 55(1):69–81, March 2018. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL https:/

/www.cambridge.org/core/journals/journal-of-applied-
probability/article/remark-on-positive-sojourn-times-
of-symmetric-processes/575BAE001657E7AEE0D930A9284B63D6.

Pekoz:2008:HND

[PRS08] Erol A. Peköz, Sheldon M. Ross, and Sridhar Seshadri. How nearly do arriving customers see time-average behavior? Journal of Applied Probability, 45(4):963–971, December 2008. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/27596002.

Pruscha:1983:LMC

[Pru83] Helmut Pruscha. Learning models with continuous time parameter and multivariate point processes. Journal of Applied Probability, 20(4):884–890, December 1983. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/3213600.

Pruscha:1985:PIM

[Pru85] Helmut Pruscha. Parametric inference in Markov branching processes with time-dependent random immigration rate. Journal of Applied Probability, 22(3):503–517, September 1985. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3213856.

Pekoz:2003:CLD

[PRX03] Erol A. Peköz, Rhonda Righter, and Cathy H. Xia. Characterizing losses during busy periods in finite buffer systems. Journal of Applied Probability, 40(1):242–249, March 2003. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3215851.

Park:1976:EBC

[PS76a] C. Park and F. J. Schuurmann. Evaluations of barrier-crossing probabilities of Wiener paths. Journal of Applied Probability, 13 (2):267–275, June 1976. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3212830.

Port:1976:RMT

[PS76b] Sidney C. Port and Charles J. Stone. Random measures and their application to motion in an incompressible fluid. Journal of Applied Probability, 13(3):498–506, September 1976. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3212469.

Park:1980:EAP

[PS80] C. Park and F. J. Schuurmann. Evaluations of absorption probabilities for the Wiener process on large intervals. Journal of Applied Probability, 17(2):363–372, June 1980. CODEN JPRBAM.

ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/3213026.

Park:1983:PBA

[PS83] C. Park and F. J. Schuurmann. Partial barrier-absorption probabilities for the Wiener process. Journal of Applied Probability, 20(1):103–110, March 1983. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3213724.

Parthasarathy:1984:SNT

[PS84] P. R. Parthasarathy and M. Sharafali. Some notes on the total service time in certain loss systems. Journal of Applied Probability, 21(3):616–627, September 1984. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3213622.

Pollak:1986:CQS

[PS86a] Moshe Pollak and David Siegmund. Convergence of quasistationary to stationary distributions for stochastically monotone Markov processes. Journal of Applied Probability, 23(1): 215–220, March 1986. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3214131.

Pourbabai:1986:SUF

[PS86b] Behnam Pourbabai and David Sonderman. Server utilization factors in queueing loss systems with ordered entry and heterogeneous servers. Journal of Applied Probability, 23(1):236– 242, March 1986. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3214135.

Parthasarathy:1989:TSM

[PS89] P. R. Parthasarathy and M. Sharafali. Transient solution to the many-server Poisson queue: A simple approach. Journal of Applied Probability, 26(3):584–594, September 1989. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3214415.

Pflug:1992:CES

[PS92] G. Ch. Pflug and W. Schachermayer. Coefficients of ergodicity for stochastically monotone Markov chains. Journal of Applied Probability, 29(4):850–860, December 1992. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/3214717.

Perry:1999:HTA

[PS99] David Perry and Wolfgang Stadje. Heavy traffic analysis of a queueing system with bounded capacity for two types of customers. Journal of Applied Probability, 36(4):1155–1166, December 1999. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3215585.

Perry:2001:BCR

[PS01a] David Perry and Wolfgang Stadje. The busy cycle of the reflected superposition of Brownian motion and a compound Poisson process. Journal of Applied Probability, 38(1):255–261, March 2001. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/3215758.

Pritchard:2001:ECR

[PS01b] Geoffrey Pritchard and David J. Scott. Empirical convergence rates for continuous-time Markov chains. Journal of Applied Probability, 38(1):262–269, March 2001. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http:/

/www.jstor.org/stable/3215759.

Pellerey:2002:ASC

[PS02a] Franco Pellerey and Patrizia Semeraro. Ageing and stochastic comparisons for a covariate failure model. Journal of Applied Probability, 39(2):421–425, June 2002. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3216109.

Pollett:2002:PIC

[PS02b] P. K. Pollett and V. T. Stefanov. Path integrals for continuoustime Markov chains. Journal of Applied Probability, 39(4):901– 904, December 2002. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3216013.

Pritchard:2004:EET

[PS04] Geoffrey Pritchard and David J. Scott. The eigenvalues of the empirical transition matrix of a Markov chain. Journal of Applied Probability, 41A:347–360, 2004. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3215988.

Pardoux:2009:HLA

[PS09] Étienne Pardoux and Majid Salamat. On the height and length of the ancestral recombination graph. Journal of Applied Probability, 46(3):669–689, September 2009. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/25662454.

Pilipauskaite:2016:ASR

[PS16] Vytaute Pilipauskaite and Donatas Surgailis. Anisotropic scaling of the random grain model with application to network traffic. Journal of Applied Probability, 53(3):857–879, September 2016. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://projecteuclid.org/euclid.jap/
1476370781.

Perzina:2018:HMM

[PS18] Vít Perzina and Jan M. Swart. How much market making does a market need? Journal of Applied Probability, 55(3):667–681, September 2018. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL https:/

/www.cambridge.org/core/journals/journal-of-applied-
probability/article/how-much-market-making-does-a-market-
need/B8CB9EADE1B4F2C6C4B89D975AEF1440.

Portier:2019:MCI

[PS19] FranÇois Portier and Johan Segers. Monte Carlo integration with a growing number of control variates. Journal of Applied Probability, 56(4):1168–1186, December 2019. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL https://www.cambridge.org/core/journals/journal-
of-applied-probability/article/monte-carlo-integration-
with-a-growing-number-of-control-variates/65FA0C32E9D87297C6C0C05CA65BA117.

Psarrakos:2022:HPA

[Psa22] Georgios Psarrakos. How a probabilistic analogue of the mean value theorem yields stein-type covariance identities. Journal of Applied Probability, 59(2):350–365, June 2022. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-

6072 (electronic). URL https://www.cambridge.org/core/
journals/journal-of-applied-probability/article/how-

a-probabilistic-analogue-of-the-mean-value-theorem-yields-
steintype-covariance-identities/0249A691C5A80EA84EF7484C1769A313.

Pardoux:2017:LDP

[PSK17] Etienne Pardoux and Brice Samegni-Kepgnou. Large deviation principle for epidemic models. Journal of Applied Probability, 54(3):905–920, September 2017. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL https:/

/www.cambridge.org/core/journals/journal-of-applied-
probability/article/large-deviation-principle-for-epidemic-
models/4A3137B42A47143209DF65B22B37D331.

Phatarfod:1971:NRW

[PSW71] R. M. Phatarfod, T. P. Speed, and A. M. Walker. A note on random walks. Journal of Applied Probability, 8(1):198–201, March 1971. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3211849.

Perry:2004:FRT

[PSZ04] D. Perry, W. Stadje, and S. Zacks. The first rendezvous time of Brownian motion and compound Poisson-type processes. Journal of Applied Probability, 41(4):1059–1070, December 2004. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/4141378.

Perry:2013:DAQ

[PSZ13] D. Perry, W. Stadje, and S. Zacks. A duality approach to queues with service restrictions and storage systems with statedependent rates. Journal of Applied Probability, 50(3):612– 631, September 2013. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://projecteuclid.

org/euclid.jap/1378401226.

Padgett:1971:SIE

[PT71] W. J. Padgett and C. P. Tsokos. On a stochastic integral equation of the Volterra type in telephone traffic theory. Journal of Applied Probability, 8(2):269–275, June 1971. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3211897.

Pickard:1980:CWW

[PT80] David K. Pickard and E. M. Tory. A critique of Weiner’s work on Palásti’s conjecture. Journal of Applied Probability, 17(3):

880–884, September 1980. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3212986.

Pham:1981:FOB

[PT81] Tuan Dinh Pham and Lanh Tat Tran. On the first-order bilinear time series model. Journal of Applied Probability, 18(3):617–627, September 1981. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/

3213316. See correction [Ano82h].

Penm:1986:DMA

[PT86] J. H. W. Penm and R. D. Terrell. The ‘derived’ moving-average model and its role in causality. Journal of Applied Probability, 23A:99–111, 1986. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3214346.

Prabhu:1994:MMS

[PT94] N. U. Prabhu and L. C. Tang. Markov-modulated single-server queueing systems. Journal of Applied Probability, 31A:169–184, 1994. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3214955.

Petakos:1997:RAS

[PT97] Kyriakos Petakos and Theothosios Tsapelas. Reliability analysis for systems in a random environment. Journal of Applied Probability, 34(4):1021–1031, December 1997. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/3215014.

Pollak:2011:FET

[PT11] Moshe Pollak and Alexander G. Tartakovsky. On the first exit time of a nonnegative Markov process started at a quasistationary distribution. Journal of Applied Probability, 48(2):589– 595, June 2011. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
23065906.

Peigne:2022:ABD

[PT22] Marc Peigné and Tat Dat Tran. On the asymptotic behavior of the Diaconis-Freedman chain in a multi-dimensional simplex.

Journal of Applied Probability, 59(2):505–526, June 2022. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL https://www.cambridge.org/core/journals/journal-
of-applied-probability/article/on-the-asymptotic-behavior-
of-the-diaconisfreedman-chain-in-a-multidimensional-
simplex/424E758CFBC9721318E8308C691A4888.

Pianoforte:2023:MPP

[PT23a] Federico Pianoforte and Riccardo Turin. Multivariate Poisson and Poisson process approximations with applications to Bernoulli sums an U-statistics. Journal of Applied Probability, 60(1):223–240, March 2023. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL https:/

/www.cambridge.org/core/journals/journal-of-applied-
probability/article/multivariate-poisson-and-poisson-
process-approximations-with-applications-to-bernoulli-
sums-and-u-statistics/CDDCD72F5BFE59699A6360CE20427B7A.

Pojer:2023:RPM

[PT23b] Simon Pojer and Stefan Thonhauser. Ruin probabilities in a Markovian shot-noise environment. Journal of Applied Probability, 60(2):542–556, June 2023. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL https:/

/www.cambridge.org/core/journals/journal-of-applied-
probability/article/ruin-probabilities-in-a-markovian-
shotnoise-environment/E91A9CA515810F57ADA77951A19AFE84.

Peligrad:2016:IPR

[PU16] Magda Peligrad and Sergey Utev. On the invariance principle for reversible Markov chains. Journal of Applied Probability, 53 (2):593–599, June 2016. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://projecteuclid.

org/euclid.jap/1466172877.

Puri:1968:IBD

[Pur68] Prem S. Puri. Interconnected birth and death processes. Journal of Applied Probability, 5(2):334–349, August 1968. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3212256.

Puri:1969:SNR

[Pur69] Prem S. Puri. Some new results in the mathematical theory of phage-reproduction. Journal of Applied Probability, 6(3):

493–504, December 1969. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3212097.

Puri:1971:MSI

[Pur71] Prem S. Puri. A method for studying the integral functionals of stochastic processes with applications: I. Markov chain case. Journal of Applied Probability, 8(2):331–343, June 1971. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/3211903.

Purdue:1973:NLM

[Pur73] Peter Purdue. Non-linear matrix integral equations of Volterra type in queueing theory. Journal of Applied Probability, 10(3): 644–651, September 1973. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3212783.

Purdue:1975:QPI

[Pur75a] Peter Purdue. A queue with Poisson input and semi-Markov service times: Busy period analysis. Journal of Applied Probability, 12(2):353–357, June 1975. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3212450.

Puri:1975:LBD

[Pur75b] Prem S. Puri. A linear birth and death process under the influence of another process. Journal of Applied Probability, 12(1): 1–17, March 1975. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3212402.

Puri:1978:LTP

[Pur78] Prem S. Puri. A limit theorem for point processes with applications. Journal of Applied Probability, 15(4):726–747, December 1978. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3213430.

Puri:1982:CPP

[Pur82] Prem S. Puri. On the characterization of point processes with the order statistic property without the moment condition. Journal of Applied Probability, 19(1):39–51, March 1982. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3213914.

Puri:1987:ASC

[Pur87] Prem S. Puri. On almost sure convergence of an erosion process due to Todorovic and Gani. Journal of Applied Probability, 24 (4):1001–1005, December 1987. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.

org/stable/3214222.

Pandelis:2012:SPO

[PV12] Dimitrios G. Pandelis and Mark P. Van Oyen. Sample path optimal policies for serial lines with flexible workers. Journal of Applied Probability, 49(2):582–589, June 2012. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/41713787.

Papaderou-Vogiatzaki:1988:CPP

[PVS88] Irini Papaderou-Vogiatzaki and Rolf Schneider. A collision probability problem. Journal of Applied Probability, 25(3):617–623, September 1988. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3213989.

Pettigrew:1967:ECL

[PW67] Hugh M. Pettigrew and George H. Weiss. Epidemics with carriers: The large population approximation. Journal of Applied Probability, 4(2):257–263, August 1967. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http:/

/www.jstor.org/stable/3212020.

Puri:1981:GSM

[PW81] Prem S. Puri and Samuel W. Woolford. On a generalized storage model with moment assumptions. Journal of Applied Probability, 18(2):473–481, June 1981. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3213293.

Petruccelli:1984:TAM

[PW84] Joseph D. Petruccelli and Samuel W. Woolford. A threshold AR(1) model. Journal of Applied Probability, 21(2):270– 286, June 1984. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3213639.

Peterson:1996:HTA

[PW96] William P. Peterson and Lawrence M. Wein. Heavy traffic analysis of a transportation network model. Journal of Applied Probability, 33(3):870–885, September 1996. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3215365.

Potzelberger:2001:BCP

[PW01] Klaus Pötzelberger and Liqun Wang. Boundary crossing probability for Brownian motion. Journal of Applied Probability, 38 (1):152–164, March 2001. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3215748.

Prince:2007:FCB

[PW07a] Thomas Prince and Neville Weber. Fixation in conditional branching process models in population genetics. Journal of Applied Probability, 44(4):1103–1110, December 2007. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/27595911.

Privault:2007:IPP

[PW07b] Nicolas Privault and Xiao Wei. Integration by parts for point processes and Monte Carlo estimation. Journal of Applied Probability, 44(3):806–823, September 2007. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http:/

/www.jstor.org/stable/27595883.

Pedarsani:2016:SMQ

[PW16] Ramtin Pedarsani and Jean Walrand. Stability of multiclass queueing networks under longest-queue and longest-dominating-queue scheduling. Journal of Applied Probability, 53(2):421–433, June 2016. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-

6072 (electronic). URL http://projecteuclid.org/euclid.
jap/1466172864.

Puccetti:2012:ACM

[PWW12] Giovanni Puccetti, Bin Wang, and Ruodu Wang. Advances in complete mixability. Journal of Applied Probability, 49(2):430– 440, June 2012. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
41713777.

Pycke:2021:GRS

[Pyc21] Jean-Renaud Pycke. On a generalization of the Rényi– Srivastava characterization of the Poisson law. Journal of Applied Probability, 58(1):68–82, March 2021. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL https:/

/www.cambridge.org/core/journals/journal-of-applied-
probability/article/on-a-generalization-of-the-renyisrivastava-
characterization-of-the-poisson-law/1D87DBE0804C8410347D0108CFC384F0.

Pyke:1994:MOS

[Pyk94] Ronald Pyke. Minimax one-sided Kolmogorov-type distributionfree tests. Journal of Applied Probability, 31A:291–308, 1994. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/3214963.

Posner:1990:ORP

[PZ90] M. J. M. Posner and D. Zuckerman. Optimal R&D programs in a random environment. Journal of Applied Probability, 27 (2):343–350, June 1990. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3214653.

Palmowski:2007:TAS

[PZ07] Zbigniew Palmowski and Bert Zwart. Tail asymptotics of the supremum of a regenerative process. Journal of Applied Probability, 44(2):349–365, June 2007. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/27595846.

Palmowski:2010:PRW

[PZ10] Zbigniew Palmowski and Bert Zwart. On perturbed random walks. Journal of Applied Probability, 47(4):1203–1204, December 2010. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/29777438.

Pollett:2007:NET

[PZC07] Phil Pollett, Hanjun Zhang, and Benjamin J. Cairns. A note on extinction times for the general birth, death and catastrophe process. Journal of Applied Probability, 44(2):566–569, June 2007. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/27595863.

Quine:1977:EMB

[QD77] M. P. Quine and P. Durham. Estimation for multitype branching processes. Journal of Applied Probability, 14(4):829–835, December 1977. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3213355.

Qu:2021:ESO

[QDZ21] Yan Qu, Angelos Dassios, and Hongbiao Zhao. Exact simulation of Ornstein–Uhlenbeck tempered stable processes. Journal of Applied Probability, 58(2):347–371, June 2021. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL https://www.cambridge.org/core/journals/journal-
of-applied-probability/article/exact-simulation-of-ornsteinuhlenbeck-
tempered-stable-processes/B96BF4090B32779D155A2AC3D53B7198.

Qiu:1991:NOI

[Qiu91] Yuping Qiu. A note on optimal inspection policy for stochastically deteriorating series systems. Journal of Applied Probability, 28(4):934–939, December 1991. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3214700.

Quine:1996:ERS

[QL96] M. P. Quine and J. S. Law. Exact results for a secretary problem. Journal of Applied Probability, 33(3):630–639, September 1996. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3215345.

Quine:1999:MRL

[QL99] M. P. Quine and J. S. Law. Modelling random linear nucleation and growth by a Markov chain. Journal of Applied Probability, 36(1):273–278, March 1999. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3215421.

Quinn:1986:NAD

[QM86] B. G. Quinn and H. L. MacGillivray. Normal approximations to discrete unimodal distributions. Journal of Applied Probability, 23(4):1013–1018, December 1986. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3214474.

Quine:1990:LRG

[QR90] M. P. Quine and J. Robinson. A linear random growth model. Journal of Applied Probability, 27(3):499–509, September 1990. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/3214536.

Quine:1999:EPL

[QS99] M. P. Quine and W. Szczotka. Existence and positivity of the limit in processes with a branching structure. Journal of Applied Probability, 36(1):132–138, March 1999. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http:/

/www.jstor.org/stable/3215408.

Quine:1970:MTG

[Qui70] M. P. Quine. The multi-type Galton–Watson process with immigration. Journal of Applied Probability, 7(2):411–422, August 1970. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3211974.

Quine:1976:BEP

[Qui76] M. P. Quine. Bounds for the extinction probability of a simple branching process. Journal of Applied Probability, 13(1): 9–16, March 1976. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3212660.

Quigg:1978:LLT

[Qui78] David L. Quigg. A local limit theorem for the critical agedependent branching process. Journal of Applied Probability, 15(1):46–53, March 1978. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3213235.

Quinn:1986:TPS

[Qui86] B. G. Quinn. Testing for the presence of sinusoidal components. Journal of Applied Probability, 23A:201–210, 1986. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3214353.

Quine:1991:MRI

[Qui91] M. P. Quine. A model for random instantaneous growth on an interval. Journal of Applied Probability, 28(3):529–538, September 1991. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3214489.

Quine:1984:RGD

[QW84] M. P. Quine and D. F. Watson. Radial generation of ndimensional Poisson procesess. Journal of Applied Probability, 21(3):548–557, September 1984. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.

org/stable/3213616.

Qi:2019:RSP

[QZ19] Weinan Qi and Mahmoud Zarepour. On resampling schemes for polytopes. Journal of Applied Probability, 56(4):959–980, December 2019. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-

6072 (electronic). URL https://www.cambridge.org/core/
journals/journal-of-applied-probability/article/on-resampling-
schemes-for-polytopes/46A7496135D28DB100864284E5A363CF.

Raaijmakers:1980:LSS

[Raa80] F. J. M. Raaijmakers. The lifetime of a standby system of units having the Birnbaum and Saunders distribution. Journal of Applied Probability, 17(2):490–497, June 1980. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/3213038.

Rabehasaina:2006:MMM

[Rab06] Landy Rabehasaina. Moments of a Markov-modulated, irreducible network of fluid queues. Journal of Applied Probability, 43(2):510–522, June 2006. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/27595747.

Radcliffe:1972:ISC

[Rad72a] J. Radcliffe. An immigration super-critical branching diffusion process. Journal of Applied Probability, 9(1):13–23, March 1972. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/3212632.

Raade:1972:LTT

[Råd72b] L. Råde. Limit theorems for thinning of renewal point processes. Journal of Applied Probability, 9(4):847–851, December 1972. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/3212621.

Raade:1972:MIP

[Råd72c] Lennart Råde. A model for interaction of a Poisson and a renewal process and its relation with queuing theory. Journal of Applied Probability, 9(2):451–456, June 1972. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3212816.

Radcliffe:1973:AFT

[Rad73a] J. Radcliffe. The asymptotic frequencies of the types in a multitype age-dependent branching process allowing immigration. Journal of Applied Probability, 10(3):652–658, September 1973. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/3212784.

Radcliffe:1973:IGS

[Rad73b] J. Radcliffe. The initial geographical spread of host-vector and carrier-borne epidemics. Journal of Applied Probability, 10(4): 703–717, December 1973. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3212374.

Radcliffe:1974:NRY

[Rad74a] J. Radcliffe. A note on the recurrence of yellow fever epidemics in urban populations. Journal of Applied Probability, 11(1):170– 173, March 1974. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3212593.

Radcliffe:1974:PEM

[Rad74b] J. Radcliffe. The periodicity of endemic malaria. Journal of Applied Probability, 11(3):562–567, September 1974. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3212700.

Radcliffe:1976:CPD

[Rad76a] J. Radcliffe. The convergence of a position-dependent branching process used as an approximation to a model describing the spread of an epidemic. Journal of Applied Probability, 13(2):338– 344, June 1976. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3212836.

Radcliffe:1976:SVH

[Rad76b] J. Radcliffe. The severity of a viral host-vector epidemic. Journal of Applied Probability, 13(4):791–794, December 1976. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3212534.

Raade:1976:RSR

[Råd76c] L. Råde. Reliability systems in random environment. Journal of Applied Probability, 13(2):407–410, June 1976. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3212849.

Raffety:1973:REB

[Raf73] Donald C. Raffety. Random environment branching processes with equal environmental extinction probabilities. Journal of Applied Probability, 10(3):659–665, September 1973. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3212785.

Rahimov:1998:MPR

[Rah98] Ibrahim Rahimov. Multitype processes with reproductiondependent immigration. Journal of Applied Probability, 35(2): 281–292, June 1998. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3215685.

Rahimov:2017:SIP

[Rah17] Ibrahim Rahimov. Statistical inference for partially observed branching processes with immigration. Journal of Applied Probability, 54(1):82–95, March 2017. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL https:/

/www.cambridge.org/core/journals/journal-of-applied-
probability/article/statistical-inference-for-partially-
observed-branching-processes-with-immigration/61216B91ABF434A12560A8DFAA7CA180.

Rajarshi:1974:SRT

[Raj74] M. B. Rajarshi. Success runs in a two-state Markov chain. Journal of Applied Probability, 11(1):190–192, March 1974. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/3212597. See correction

[Ano77f].

Rajarshi:1981:SPT

[Raj81] M. B. Rajarshi. Simpler proofs of two threshold theorems for a general stochastic epidemic. Journal of Applied Probability, 18 (3):721–724, September 1981. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.

org/stable/3213326.

Ramaswami:1981:ACR

[Ram81] V. Ramaswami. Algorithms for a continuous-review (s, s) inventory system. Journal of Applied Probability, 18(2):461–472, June 1981. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3213292.

Ramaswami:1984:STG

[Ram84] V. Ramaswami. The sojourn time in the GI/M/1 queue with processor sharing. Journal of Applied Probability, 21(2):437– 442, June 1984. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3213654.

Rao:1966:ASP

[Rao66] J. Sudarsana Rao. An application of stationary point processes to queueing theory and textile research. Journal of Applied Probability, 3(1):231–246, June 1966. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.

org/stable/3212048.

Rao:1979:CDT

[Rao79] B. L. S. Prakasa Rao. Characterizations of distributions through some identities. Journal of Applied Probability, 16(4):903–909, December 1979. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3213156.

Rapoo:2008:VSS

[Rap08] E. Rapoo. A variable step size Riemannian sum for an Itô integral. Journal of Applied Probability, 45(2):551–567, June 2008. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/27595964.

Rathie:1970:GEC

[Rat70] Pushpa N. Rathie. On a generalized entropy and a coding theorem. Journal of Applied Probability, 7(1):124–133, April 1970.

CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/3212154.

Rathie:1992:VDT

[Rat92] P. N. Rathie. On the volume distribution of the typical Poisson– Delaunay cell. Journal of Applied Probability, 29(3):740–744, September 1992. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3214909.

Raynaud:1970:ECN

[Ray70] H. Raynaud. Sur l’enveloppe convexe des nuages de points aléatoires dans R^{n}. I. (French) [On the convex envelope of clouds of random points in R^{n}. I]. Journal of Applied Probability, 7(1): 35–48, April 1970. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3212146.

Rodgers:1991:NMC

[RB91] R. P. C. Rodgers and A. J. Baddeley. Nested Monte Carlo study of random packing on the sphere. Journal of Applied Probability, 28(3):539–552, September 1991. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.

org/stable/3214490.

Richardson:1998:RQW

[RB98] Daniel Richardson and Wayne Burton. A relatively quick way to simulate local random processes on a lattice. Journal of Applied Probability, 35(3):770–775, September 1998. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/3215651.

Rieder:2005:POU

[RB05] Ulrich Rieder and Nicole Bäuerle. Portfolio optimization with unobservable Markov-modulated drift process. Journal of Applied Probability, 42(2):362–378, June 2005. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/30040798.

Raman:1973:SMP

[RC73] S. Raman and C. L. Chiang. On a solution of the migration process and the application to a problem in epidemiology. Journal of Applied Probability, 10(4):718–727, December 1973. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3212375.

Rubinovitch:1980:LCS

[RC80] Michael Rubinovitch and J. W. Cohen. Level crossings and stationary distributions for general dams. Journal of Applied Probability, 17(1):218–226, March 1980. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3212938.

Ramirez-Cobo:2012:NDS

[RCC12] Pepa Ramírez-Cobo and Emilio Carrizosa. A note on the dependence structure of the two-state Markovian arrival process. Journal of Applied Probability, 49(1):295–302, March 2012. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/41713766.

Ramirez-Cobo:2010:NTS

[RCLW10] Pepa Ramírez-Cobo, Rosa E. Lillo, and Michael P. Wiper. Nonidentifiability of the two-state Markovian arrival process. Journal of Applied Probability, 47(3):630–649, September 2010. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/20778507.

Rosenkrantz:1980:LTM

[RD80] Walter A. Rosenkrantz and C. C. Y. Dorea. Limit theorems for Markov processes via a variant of the Trotter–Kato theorem. Journal of Applied Probability, 17(3):704–715, September 1980. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3212964.

Renshaw:1997:RRR

[RD97] Eric Renshaw and Yonglong Dai. Regularity and reversibility results for birth-death-migration processes. Journal of Applied Probability, 34(3):685–697, September 1997. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/3215094.

Robin:1999:EDW

[RD99] S. Robin and J. J. Daudin. Exact distribution of word occurrences in a random sequence of letters. Journal of Applied Probability, 36(1):179–193, March 1999. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3215413.

Ratanov:2019:PDP

[RDM19] Nikita Ratanov, Antonio Di Crescenzo, and Barbara Martinucci. Piecewise deterministic processes following two alternating patterns. Journal of Applied Probability, 56(4): 1006–1019, December 2019. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL https:/

/www.cambridge.org/core/journals/journal-of-applied-
probability/article/piecewise-deterministic-processes-
following-two-alternating-patterns/1D7534A24F519658190C7B7820D5E71E.

Rothmann:2001:SPI

[RE01] Mark D. Rothmann and Hammou El Barmi. Stochastic processes involving random deletion. Journal of Applied Probability, 38(1):95–107, March 2001. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3215744.

Reder:2003:TBG

[Red03] Christine Reder. Transient behaviour of a Galton–Watson process with a large number of types. Journal of Applied Probability, 40(4):1007–1030, December 2003. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3216057.

Reh:1977:GER

[Reh77] W. Reh. Generalizations of the elementary renewal theorem to distributions defined by concave recurrence relations. Journal of Applied Probability, 14(3):516–526, September 1977. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3213454.

Reh:1979:FPP

[Reh79] Wolfgang Reh. First-passage percolation under weak moment conditions. Journal of Applied Probability, 16(4):750–763, December 1979. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3213142.

Reid:1967:MMD

[Rei67] D. H. Reid. A mathematical model for delays caused by rightturning vehicles at an uncontrolled intersection. Journal of Applied Probability, 4(1):180–191, April 1967. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/3212310.

Reker:2023:STB

[Rek23] Jana Reker. Short-time behavior of solutions to Lévy-driven stochastic differential equations. Journal of Applied Probability, 60(3):765–780, September 2023. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL https:/

/www.cambridge.org/core/journals/journal-of-applied-
probability/article/shorttime-behavior-of-solutions-
to-levydriven-stochastic-differential-equations/8594EB0A2DBB5B72D1600923B1367B3E.

Renyi:1964:TMM

[Rén64] A. Rényi. On two mathematical models of the traffic on a divided highway. Journal of Applied Probability, 1(2):311–320, December 1964. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3211862.

Renshaw:1973:IPP

[Ren73] E. Renshaw. Interconnected population processes. Journal of Applied Probability, 10(1):1–14, March 1973. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3212491.

Renshaw:1974:SSM

[Ren74] E. Renshaw. Stepping stone models for population growth. Journal of Applied Probability, 11(1):16–31, March 1974. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3212579.

Renshaw:1977:VPS

[Ren77] Eric Renshaw. Velocities of propagation for stepping-stone models of population growth. Journal of Applied Probability, 14(3): 591–597, September 1977. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3213462.

Renshaw:1979:WVN

[Ren79] Eric Renshaw. Waveforms and velocities for non-nearest-neighbour contact distributions. Journal of Applied Probability, 16(1):1–11, March 1979. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3213369.

Renshaw:1980:SDT

[Ren80] Eric Renshaw. The spatial distribution of Tribolium Confusum. Journal of Applied Probability, 17(4):895–911, December 1980. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/3213200. See

correction [Ano81f].

Renshaw:1981:WVM

[Ren81] Eric Renshaw. Waveforms and velocities for models of spatial infection. Journal of Applied Probability, 18(3):715–720, September 1981. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3213325.

Rennie:1982:GF

[Ren82] B. C. Rennie. On generalized functions. Journal of Applied Probability, 19A:139–156, 1982. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.

org/stable/3213556.

Renshaw:1987:DUO

[Ren87] Eric Renshaw. The discrete Uhlenbeck–Ornstein process. Journal of Applied Probability, 24(4):908–917, December 1987. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/3214215.

Renshaw:1988:HOA

[Ren88] Eric Renshaw. The high-order autocovariance structure of the telegraph wave. Journal of Applied Probability, 25(4):744–751, December 1988. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3214295.

Renaud:2014:TSR

[Ren14] Jean-François Renaud. On the time spent in the red by a refracted Lévy risk process. Journal of Applied Probability, 51(4): 1171–1188, December 2014. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://projecteuclid.

org/euclid.jap/1421763334.

Restrepo:1966:CMC

[Res66] Rodrigo A. Restrepo. A combinatorial method for a class of matrix games. Journal of Applied Probability, 3(2):495–511, December 1966. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3212134.

Resnick:1971:PDF

[Res71a] Sidney I. Resnick. Products of distribution functions attracted to extreme value laws. Journal of Applied Probability, 8(4):781–793, December 1971. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3212241.

Resnick:1971:TEA

[Res71b] Sidney I. Resnick. Tail equivalence and its applications. Journal of Applied Probability, 8(1):136–156, March 1971. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3211844.

Resnick:1973:EPR

[Res73] Sidney I. Resnick. Extremal processes and record value times. Journal of Applied Probability, 10(4):864–868, December 1973. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/3212388.

Resnick:2004:FMH

[Res04] Sidney Resnick. On the foundations of multivariate heavy-tail analysis. Journal of Applied Probability, 41A:191–212, 2004. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/3215977.

Reuter:1982:APD

[Reu82] Harry Reuter. Appreciation: Professor D. G. Kendall, FRS. Journal of Applied Probability, 19(4):900, December 1982. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/3213848.

Revyakov:1993:CAD

[Rev93] Mikhail Revyakov. Component allocation for a distributed system: Reliability maximization. Journal of Applied Probability, 30(2):471–477, June 1993. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3214857.

Revesz:2004:PPB

[Rév04] P. Révész. A prediction problem of the branching random walk. Journal of Applied Probability, 41A:25–31, 2004. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3215966.

Reynolds:1968:ASF

[Rey68] John F. Reynolds. On the autocorrelation and spectral functions of queues. Journal of Applied Probability, 5(2):467–475, August 1968. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3212267.

Reynolds:1970:TDT

[Rey70] John F. Reynolds. A theorem on discrete time branching processes allowing immigration. Journal of Applied Probability, 7 (2):446–450, August 1970. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3211979.

Reynolds:1972:LRP

[Rey72a] John F. Reynolds. On linearly regressive processes. Journal of Applied Probability, 9(1):208–213, March 1972. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3212651.

Reynolds:1972:STT

[Rey72b] John F. Reynolds. Some theorems on the transient covariance of Markov chains. Journal of Applied Probability, 9(1):214– 218, March 1972. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3212652.

Reynolds:1972:TMB

[Rey72c] John F. Reynolds. A theorem on Markov branching processes. Journal of Applied Probability, 9(3):667–670, September 1972. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/3212335.

Reeves:1988:GCS

[RF88] J. H. Reeves and V. F. Flack. A generalization of the classical secretary problem: Dependent arrival sequences. Journal of Applied Probability, 25(1):97–105, March 1988. CODEN JPRBAM.

ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/3214237.

Rangan:1988:NMM

[RG88] A. Rangan and R. Esther Grace. A non-Markov model for the optimum replacement of self-repairing systems subject to shocks. Journal of Applied Probability, 25(2):375–382, June 1988. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/3214445.

Renshaw:1981:CRW

[RH81a] Eric Renshaw and Robin Henderson. The correlated random walk. Journal of Applied Probability, 18(2):403–414, June 1981. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/3213286.

Richardson:1981:VSC

[RH81b] Sylvia Richardson and Denis Hemon. On the variance of the sample correlation between two independent lattice processes. Journal of Applied Probability, 18(4):943–948, December 1981. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/3213069.

Rhee:1998:TBR

[Rhe98] Wansoo T. Rhee. Tight bound of random interval packing. Journal of Applied Probability, 35(4):990–997, December 1998. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/3215534.

Resing:1990:MGP

[RHK90] J. A. C. Resing, G. Hooghiemstra, and M. S. Keane. The M/G/1 processor sharing queue as the almost sure limit of feedback queues. Journal of Applied Probability, 27(4):913–918, December 1990. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3214834.

Ridder:1996:FSM

[Rid96] Ad Ridder. Fast simulation of Markov fluid models. Journal of Applied Probability, 33(3):786–803, September 1996. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3215359.

Rieder:1976:OPM

[Rie76] Ulrich Rieder. On optimal policies and martingales in dynamic programming. Journal of Applied Probability, 13(3):507–518, September 1976. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3212470.

Riedel:1980:RCF

[Rie80] M. Riedel. Representation of the characteristic function of a stochastic integral. Journal of Applied Probability, 17(2):448– 455, June 1980. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3213034.

Righter:1990:SMN

[Rig90] Rhonda Righter. Stochastically maximizing the number of successes in a sequential assignment problem. Journal of Applied Probability, 27(2):351–364, June 1990. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3214654.

Righter:1996:OPS

[Rig96] Rhonda Righter. Optimal policies for scheduling repairs and allocating heterogeneous servers. Journal of Applied Probability, 33(2):536–547, June 1996. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3215077.

Righter:1997:SST

[Rig97] Rhonda Righter. Stochastic scheduling for a two-machine open shop. Journal of Applied Probability, 34(3):733–744, September 1997. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3215098.

Righter:1999:BMH

[Rig99a] Rhonda Righter. A brokered market with heterogeneous suppliers and consumers. Journal of Applied Probability, 36(2):512– 522, June 1999. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3215471.

Righter:1999:NLM

[Rig99b] Rhonda Righter. A note on losses in M/GI/1/n queues. Journal of Applied Probability, 36(4):1240–1243, December 1999. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/3215593.

Righter:2011:SCD

[Rig11] Rhonda Righter. Stochastic comparison of discounted rewards. Journal of Applied Probability, 48(1):293–294, March 2011. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/29777461.

Ripley:1976:SOA

[Rip76] B. D. Ripley. The second-order analysis of stationary point processes. Journal of Applied Probability, 13(2):255–266, June 1976. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/3212829.

Rissanen:1986:OEA

[Ris86] Jorma Rissanen. Order estimation by accumulated prediction errors. Journal of Applied Probability, 23A:55–61, 1986. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3214342.

Rising:1995:GMC

[Ris95] William Rising. Geometric Markov chains. Journal of Applied Probability, 32(2):349–374, June 1995. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3215293.

Ritter:1980:CTA

[Rit80] Grant Ritter. A continuous-time analogue of random walk in a random environment. Journal of Applied Probability, 17(1):259– 264, March 1980. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3212944.

Russell:1965:MAC

[RJ65] Allan M. Russell and Nora S. Josephson. Measurement of area by counting. Journal of Applied Probability, 2(2):339–351, December 1965. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3212198.

Roberts:1994:WCC

[RJ94] G. O. Roberts and S. D. Jacka. Weak convergence of conditioned birth and death processes. Journal of Applied Probability, 31 (1):90–100, March 1994. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3215237.

Roberts:1997:NEL

[RJP97] G. O. Roberts, S. D. Jacka, and P. K. Pollett. Non-explosivity of limits of conditioned birth and death processes. Journal of Applied Probability, 34(1):35–45, March 1997. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/3215172.

Ramaswami:1988:MSW

[RL88] V. Ramaswami and D. M. Lucantoni. Moments of the stationary waiting time in the GI/PH/1 queue. Journal of Applied Probability, 25(3):636–641, September 1988. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3213992.

Ray:1976:IPR

[RM76] W. D. Ray and F. Margo. The inverse problem in reducible Markov chains. Journal of Applied Probability, 13(1):49–56, March 1976. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3212664.

Roy:1988:GME

[RM88] Dilip Roy and S. P. Mukherjee. Generalised mixtures of exponential distributions. Journal of Applied Probability, 25(3): 510–518, September 1988. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3213980.

Ruiz-Medina:1997:ORR

[RMV97] M. D. Ruiz-Medina and M. J. Valderrama. Orthogonal representations of random fields and an application to geophysics data. Journal of Applied Probability, 34(2):458–476, June 1997. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/3215385.

Ramaswami:1980:SEF

[RN80] V. Ramaswami and Marcel F. Neuts. Some explicit formulas and computational methods for infinite-server queues with phase-type arrivals. Journal of Applied Probability, 17(2):498– 514, June 1980. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3213039.

Ronn-Nielsen:2016:TAS

[RNJ16] Anders Rønn-Nielsen and Eva B. Vedel Jensen. Tail asymptotics for the supremum of an infinitely divisible field with convolution equivalent Lévy measure. Journal of Applied Probability, 53 (1):244–261, March 2016. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://projecteuclid.

org/euclid.jap/1457470572.

Ronn-Nielsen:2017:ESI

[RNJ17] Anders Rønn-Nielsen and Eva B. Vedel Jensen. Excursion sets of infinitely divisible random fields with convolution equivalent Lévy measure. Journal of Applied Probability, 54(3):833–851, September 2017. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL https:/

/www.cambridge.org/core/journals/journal-of-applied-
probability/article/excursion-sets-of-infinitely-divisible-
random-fields-with-convolution-equivalent-levy-measure/
7575A5B16C823A29AD339AA06F234833.

Ronn-Nielsen:2019:CLT

[RNJ19] Anders Rønn-Nielsen and Eva B. Vedel Jensen. Central limit theorem for mean and variogram estimators in Lévy-based models. Journal of Applied Probability, 56 (1):209–222, March 2019. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL https:/

/www.cambridge.org/core/journals/journal-of-applied-
probability/article/central-limit-theorem-for-mean-and-
variogram-estimators-in-levybased-models/BEF4CA02164880757D318EE88AEA0147.

Rychlik:2018:EPV

[RNR18] Tomasz Rychlik, Jorge Navarro, and Rafael Rubio. Effective procedure of verifying stochastic ordering of system lifetimes. Journal of Applied Probability, 55(4):1261–1271, December 2018. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL https://www.cambridge.org/core/journals/

journal-of-applied-probability/article/effective-procedure-
of-verifying-stochastic-ordering-of-system-lifetimes/
7F9DC1B5D6C0790C9CB247AC71F1777D.

Rachev:1999:MTP

[RO99] S. T. Rachev and I. Olkin. Mass transportation problems with capacity constraints. Journal of Applied Probability, 36(2):433– 445, June 1999. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3215466.

Robinson:1972:NLR

[Rob72] P. M. Robinson. Non-linear regression for multiple time-series. Journal of Applied Probability, 9(4):758–768, December 1972. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/3212613.

Robinson:1980:OCM

[Rob80] D. R. Robinson. Optimality conditions for a Markov decision chain with unbounded costs. Journal of Applied Probability, 17 (4):996–1003, December 1980. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.

org/stable/3213209.

Robinson:1986:MTS

[Rob86] P. M. Robinson. On a model for a time series of cross-sections. Journal of Applied Probability, 23A:113–125, 1986. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3214347.

Robert:1990:ECT

[Rob90] Claudine Robert. An entropy concentration theorem: Applications in artificial intelligence and descriptive statistics. Journal of Applied Probability, 27(2):303–313, June 1990. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3214649.

Roberts:1991:CTC

[Rob91] G. O. Roberts. A comparison theorem for conditioned Markov processes. Journal of Applied Probability, 28(1):74–83, March 1991. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3214741.

Roberts:1992:RRV

[Rob92] Leigh A. Roberts. On ratios of random variables and generalised mortality rates. Journal of Applied Probability, 29(2):268– 279, June 1992. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3214565.

Roberts:1999:NAR

[Rob99] G. O. Roberts. A note on acceptance rate criteria for CLTs for Metropolis–Hastings algorithms. Journal of Applied Probability, 36(4):1210–1217, December 1999. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3215589.

Robert:2005:APE

[Rob05] Christian Y. Robert. Asymptotic probabilities of an exceedance over renewal thresholds with an application to risk theory. Journal of Applied Probability, 42(1):153–162, March 2005. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/30040776.

Robert:2014:VGC

[Rob14] C. Y. Robert. On the De Vylder and Goovaerts conjecture about ruin for equalized claims. Journal of Applied Probability, 51(3): 874–879, September 2014. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://projecteuclid.

org/euclid.jap/1409932679.

Roberts:2022:CTB

[Rob22] Matthew I. Roberts. Cover time for branching random walks on regular trees. Journal of Applied Probability, 59(1):256–277, March 2022. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL https:/

/www.cambridge.org/core/journals/journal-of-applied-
probability/article/cover-time-for-branching-random-
walks-on-regular-trees/A4123669E89D0EDFAF5D01CA8FBF5F40.

Rodrigues:1995:CSS

[Rod95a] Eliane R. Rodrigues. Convergence to stationary state for a Markov move-to-front scheme. Journal of Applied Probability, 32(3):768–776, September 1995. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.

org/stable/3215128.

Rodrigues:1995:PMF

[Rod95b] Eliane R. Rodrigues. The performance of the move-to-front scheme under some particular forms of Markov requests. Journal of Applied Probability, 32(4):1089–1102, December 1995. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/3215222.

Roes:1970:FD

[Roe70a] P. B. M. Roes. The finite dam. Journal of Applied Probability, 7 (2):316–326, August 1970. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3211967.

Roes:1970:FDI

[Roe70b] P. B. M. Roes. The finite dam II. Journal of Applied Probability, 7(3):599–616, December 1970. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.

org/stable/3211940.

Roes:1970:ENC

[Roe70c] P. B. M. Roes. On the expected number of crossings of a level in certain stochastic processes. Journal of Applied Probability, 7(3): 766–770, December 1970. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3211954.

Roerdink:1985:NAP

[Roe85] J. B. T. M. Roerdink. A note on the asymptotic properties of correlated random walks. Journal of Applied Probability, 22(4): 951–956, December 1985. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3213964.

Rogers:1978:PTS

[Rog78] L. C. G. Rogers. The probability that two samples in the plane will have disjoint convex hulls. Journal of Applied Probability, 15(4):790–802, December 1978. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.

org/stable/3213434.

Rogerson:1987:PCA

[Rog87] Peter A. Rogerson. Probabilities of choosing applicants of arbitrary rank in the secretary problem. Journal of Applied Probability, 24(2):527–533, June 1987. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3214276.

Rogers:2000:EFP

[Rog00] L. C. G. Rogers. Evaluating first-passage probabilities for spectrally one-sided Lévy processes. Journal of Applied Probability, 37(4):1173–1180, December 2000. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3215510.

Rogers:2023:BRS

[Rog23] L. C. G. Rogers. The Bruss–Robertson–Steele inequality. Journal of Applied Probability, 60(3):1112–1114, September 2023. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL https://www.cambridge.org/core/journals/
journal-of-applied-probability/article/brussrobertsonsteele-
inequality/C07EF32C6BB654E51337F84D2BEE8F52.

Rohatgi:1988:OEL

[Roh88] Vijay K. Rohatgi. Obituary: Eugene Lukacs. Journal of Applied Probability, 25(3):642–646, September 1988. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/3213993.

Rootzen:1976:GTP

[Roo76a] Holger Rootzén. Gordin’s theorem and the periodogram. Journal of Applied Probability, 13(2):365–370, June 1976. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3212842.

Rootzen:1976:NCL

[Roo76b] Holger Rootzén. A note on the central limit theorem for doubly stochastic Poisson processes. Journal of Applied Probability, 13 (4):809–813, December 1976. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.

org/stable/3212538.

Roos:2003:PAM

[Roo03] Bero Roos. Poisson approximation of multivariate Poisson mixtures. Journal of Applied Probability, 40(2):376–390, June 2003. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/3215797.

Ross:1969:ODP

[Ros69] Sheldon M. Ross. Optimal dispatching of a Poisson process. Journal of Applied Probability, 6(3):692–699, December 1969. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/3212114.

Ross:1970:ACS

[Ros70a] Sheldon M. Ross. Average cost semi-Markov decision processes. Journal of Applied Probability, 7(3):649–656, December 1970. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/3211944.

Ross:1970:IGG

[Ros70b] Sheldon M. Ross. Identifiability in GI/G/k queues. Journal of Applied Probability, 7(3):776–780, December 1970. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3211956.

Ross:1971:GGP

[Ros71] Sheldon M. Ross. Goofspiel: The game of pure strategy. Journal of Applied Probability, 8(3):621–625, September 1971. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3212187.

Ross:1972:MSI

[Ros72] Sheldon M. Ross. On the maximum of a stationary independent increment process. Journal of Applied Probability, 9(3):677–680, September 1972. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3212337.

Rosenlund:1973:MGM

[Ros73] Stig I. Rosenlund. An M/G/1 model with finite waiting room in which a customer remains during part of service. Journal of Applied Probability, 10(4):778–785, December 1973. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3212381.

Ross:1974:BDD

[Ros74] Sheldon M. Ross. Bounds on the delay distribution in GI/G/1 queues. Journal of Applied Probability, 11(2):417–421, June 1974. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3212767.

Rosenlund:1975:BPF

[Ros75] Stig I. Rosenlund. Busy period of a finite queue with phase type service. Journal of Applied Probability, 12(1):201–204, March 1975. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3212430.

Rosenblatt:1976:FIS

[Ros76a] M. Rosenblatt. Fractional integrals of stationary processes and the central limit theorem. Journal of Applied Probability, 13(4): 723–732, December 1976. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3212527.

Rosenlund:1976:REF

[Ros76b] Stig I. Rosenlund. Recurrent emptiness in a finite queue with increasing arrival rate. Journal of Applied Probability, 13(2):423– 426, June 1976. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3212852.

Ross:1978:ADQ

[Ros78] Sheldon M. Ross. Average delay in queues with non-stationary Poisson arrivals. Journal of Applied Probability, 15(3):602–609, September 1978. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3213122.

Rosberg:1980:PRC

[Ros80a] Zvi Rosberg. A positive recurrence criterion associated with multidimensional queueing processes. Journal of Applied Probability, 17(3):790–801, September 1980. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3212972.

Rosenblatt:1980:LPB

[Ros80b] M. Rosenblatt. Linear processes and bispectra. Journal of Applied Probability, 17(1):265–270, March 1980. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3212945.

Rosberg:1981:NEM

[Ros81a] Zvi Rosberg. A note on the ergodicity of Markov chains. Journal of Applied Probability, 18(1):112–121, March 1981. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3213171.

Rosenlund:1981:MCS

[Ros81b] Stig I. Rosenlund. Markov chains in small time intervals. Journal of Applied Probability, 18(3):747–751, September 1981. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3213331.

Ross:1981:RG

[Ros81c] Sheldon M. Ross. A random graph. Journal of Applied Probability, 18(1):309–315, March 1981. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3213194.

Rosberg:1982:SMD

[Ros82] Zvi Rosberg. Semi-Markov decision processes with polynomial reward. Journal of Applied Probability, 19(2):301–309, June 1982. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3213482.

Rosenblatt:1986:PEF

[Ros86] M. Rosenblatt. Parameter estimation for finite-parameter stationary random fields. Journal of Applied Probability, 23A: 311–318, 1986. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3214360.

Rosberg:1987:BEW

[Ros87a] Zvi Rosberg. Bounds on the expected waiting time in a GI/G/1 queue: Upgrading for low traffic intensity. Journal of Applied Probability, 24(3):749–757, September 1987. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/3214104.

Rosenblatt:1987:SRR

[Ros87b] M. Rosenblatt. Scale renormalization and random solutions of the Burgers equation. Journal of Applied Probability, 24(2):328– 338, June 1987. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3214257.

Rosenthal:1993:RWD

[Ros93] Jeffrey S. Rosenthal. Random walks on discrete and continuous circles. Journal of Applied Probability, 30(4):780–789, December 1993. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3214512.

Rossignol:2007:ATW

[Ros07] Raphaël Rossignol. Arbitrary threshold widths for monotone, symmetric properties. Journal of Applied Probability, 44(1):164– 180, March 2007. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
27595829.

Rossberg:2008:LTP

[Ros08] A. G. Rossberg. Laplace transforms of probability distributions and their inversions are easy on logarithmic scales. Journal of Applied Probability, 45(2):531–541, June 2008. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/27595962.

Ross:2012:DMP

[Ros12] Sheldon M. Ross. Distribution of minimal path lengths when edge lengths are independent heterogeneous exponential random variables. Journal of Applied Probability, 49(3):895–900, September 2012. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
41713815.

Ross:2014:OSS

[Ros14a] Sheldon M. Ross. Optimal server selection in a queueing loss model with heterogeneous exponential servers, discriminating arrivals, and arbitrary arrival times. Journal of Applied Probability, 51(3):880–884, September 2014. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http:/

/projecteuclid.org/euclid.jap/1409932680.

Ross:2014:SAS

[Ros14b] Sheldon M. Ross. Simulation analysis of system life when component lives are determined by a marked point process. Journal of Applied Probability, 51(2):377–386, June 2014. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://projecteuclid.org/euclid.jap/1402578631.

Ross:2016:ICS

[Ros16] Sheldon M. Ross. Improved Chen–Stein bounds on the probability of a union. Journal of Applied Probability, 53(4):1265– 1270, December 2016. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://projecteuclid.

org/euclid.jap/1481132851.

Rosenberg:2018:NFM

[Ros18] Josh Rosenberg. The nonhomogeneous frog model on Z. Journal of Applied Probability, 55(4):1093–1112, December 2018. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL https://www.cambridge.org/core/journals/journal-
of-applied-probability/article/nonhomogeneous-frog-model-
on/A0CAE649B8D8462C281FB74397339F33.

Roters:1994:VWE

[Rot94] Markus Roters. On the validity of Wald’s equation. Journal of Applied Probability, 31(4):949–957, December 1994. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3215320.

Roters:1995:OSP

[Rot95] Markus Roters. An optimal stopping problem for random walks with non-zero drift. Journal of Applied Probability, 32(4):956– 959, December 1995. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3215208.

Roters:1998:OSD

[Rot98] Markus Roters. Optimal stopping in a dice game. Journal of Applied Probability, 35(1):229–235, March 1998. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3215561.

Routleff:1982:VWC

[Rou82] C. Routleff. A variation of Weiss’s carrier-borne epidemic model. Journal of Applied Probability, 19(2):403–407, June 1982. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/3213491.

Routledge:1990:SPA

[Rou90] R. D. Routledge. Spatial patterns arising from plant dispersal as modelled by a correlated random walk. Journal of Applied Probability, 27(1):1–13, March 1990. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3214591.

Roughan:1996:AMM

[Rou96] Matthew Roughan. An analysis of a modified M/G/1 queue using a martingale technique. Journal of Applied Probability, 33 (1):224–238, March 1996. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3215280.

Roy:1972:SAR

[Roy72] Roch Roy. Spectral analysis for a random process on the circle. Journal of Applied Probability, 9(4):745–757, December 1972. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/3212612.

Roy:1989:CGB

[Roy89] Dilip Roy. A characterization of Gumbel’s bivariate exponential and Lindley and Singpurwalla’s bivariate Lomax distributions. Journal of Applied Probability, 26(4):886–891, December 1989. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/3214394. See

correction [Ano90f].

RoyChoudhury:2014:ICB

[Roy14] Arindam RoyChoudhury. Identifiability of a coalescent-based population tree model. Journal of Applied Probability, 51(4): 921–929, December 2014. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://projecteuclid.

org/euclid.jap/1421763318.

Roy:2024:BRW

[Roy24] Rishideep Roy. A branching random walk in the presence of a hard wall. Journal of Applied Probability, 61(1):1–17, March 2024. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-

6072 (electronic). URL https://www.cambridge.org/core/
journals/journal-of-applied-probability/article/branching-
random-walk-in-the-presence-of-a-hard-wall/E1A33306384ED3FD32CB1719C30FFEDC.

Roos:2003:DBD

[RP03] Bero Roos and Dietmar Pfeifer. On the distance between the distributions of random sums. Journal of Applied Probability, 40(1):87–106, March 2003. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3215843.

Ridler-Rowe:1967:SME

[RR67] C. J. Ridler-Rowe. On a stochastic model of an epidemic. Journal of Applied Probability, 4(1):19–33, April 1967. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3212297.

Ruben:1973:MGF

[RR73] H. Ruben and W. J. Reed. A more general form of a theorem of Crofton. Journal of Applied Probability, 10(2):479–482, June 1973. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3212369.

Rasmussen:1975:CPU

[RR75a] Will T. Rasmussen and Herbert Robbins. The candidate problem with unknown population size. Journal of Applied Probability, 12(4):692–701, December 1975. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3212720.

Resnick:1975:BNO

[RR75b] Sidney I. Resnick and Michael Rubinovitch. The behavior near the origin of the supremum functional in a process with stationary independent increments. Journal of Applied Probability, 12 (1):159–160, March 1975. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3212419.

Ripley:1977:FEP

[RR77] B. D. Ripley and J.-P. Rasson. Finding the edge of a Poisson forest. Journal of Applied Probability, 14(3):483–491, September 1977. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3213451.

Ridler-Rowe:1978:CBT

[RR78] C. J. Ridler-Rowe. On competition between two species. Journal of Applied Probability, 15(3):457–465, September 1978. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3213109.

Ridler-Rowe:1988:ETC

[RR88] C. J. Ridler-Rowe. Extinction times for certain predatorprey processes. Journal of Applied Probability, 25(3):612–616, September 1988. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3213988.

Resnick:1994:SEP

[RR94] Sidney I. Resnick and Rishin Roy. Super-extremal processes and the argmax process. Journal of Applied Probability, 31(4): 958–978, December 1994. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3215321.

Rothmann:1998:WCR

[RR98] Mark D. Rothmann and Ralph P. Russo. Weak convergence on randomly deleted sets. Journal of Applied Probability, 35(4): 893–902, December 1998. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3215525.

Roberts:2007:CEA

[RR07] Gareth O. Roberts and Jeffrey S. Rosenthal. Coupling and ergodicity of adaptive Markov chain Monte Carlo algorithms. Journal of Applied Probability, 44(2):458–475, June 2007. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/27595854.

Reinert:2010:RSC

[RR10] Gesine Reinert and Adrian Röllin. Random subgraph counts and U-statistics: Multivariate normal approximation via exchangeable pairs and embedding. Journal of Applied Probability, 47 (2):378–393, June 2010. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/20778486.

Romera:2012:RPF

[RR12] R. Romera and W. Runggaldier. Ruin probabilities in a finitehorizon risk model with investment and reinsurance. Journal of Applied Probability, 49(4):954–966, December 2012. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/41713821.

Roberts:2016:CBM

[RR16] Gareth O. Roberts and Jeffrey S. Rosenthal. Complexity bounds for Markov chain Monte Carlo algorithms via diffusion limits. Journal of Applied Probability, 53(2):410–420, June 2016. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://projecteuclid.org/euclid.jap/1466172863.

Roberts:1998:CPP

[RRS98] Gareth O. Roberts, Jeffrey S. Rosenthal, and Peter O. Schwartz. Convergence properties of perturbed Markov chains. Journal of Applied Probability, 35(1):1–11, March 1998. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3215541.

Revesz:2005:LTA

[RRS05] Pál Révész, Jay Rosen, and Zhan Shi. Large-time asymptotics for the density of a branching Wiener process. Journal of Applied Probability, 42(4):1081–1094, December 2005. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/30040830.

Roberts:2022:SBM

[RRT22] Gareth O. Roberts, Jeffrey S. Rosenthal, and Nicholas G. Tawn. Skew Brownian motion and complexity of the ALPS algorithm. Journal of Applied Probability, 59(3): 777–796, September 2022. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL https:/

/www.cambridge.org/core/journals/journal-of-applied-
probability/article/skew-brownian-motion-and-complexity-
of-the-alps-algorithm/34FB3397A93FCFE042C76B332C39FB12.

Robbins:1966:ELW

[RS66] H. Robbins and E. Samuel. An extension of a lemma of Wald. Journal of Applied Probability, 3(1):272–273, June 1966. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/3212052.

Radcliffe:1969:FOC

[RS69] J. Radcliffe and P. J. Staff. First-order conservative processes with multiple latent roots. Journal of Applied Probability, 6 (1):186–194, April 1969. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3212285.

Ricciardi:1983:NEF

[RS83] L. M. Ricciardi and S. Sato. A note on the evaluation of firstpassage-time probability densities. Journal of Applied Probability, 20(1):197–201, March 1983. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3213736.

Ricciardi:1987:PDO

[RS87] L. M. Ricciardi and L. Sacerdote. On the probability densities of an Ornstein–Uhlenbeck process with a reflecting boundary. Journal of Applied Probability, 24(2):355–369, June 1987. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/3214260.

Ricciardi:1988:FPT

[RS88] Luigi M. Ricciardi and Shunsuke Sato. First-passage-time density and moments of the Ornstein–Uhlenbeck process. Journal of Applied Probability, 25(1):43–57, March 1988. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3214232.

Rataj:1989:APA

[RS89a] Jan Rataj and Ivan Saxl. Analysis of planar anisotropy by means of the Steiner compact. Journal of Applied Probability, 26(3): 490–502, September 1989. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3214407.

Rubino:1989:WLM

[RS89b] Gerardo Rubino and Bruno Sericola. On weak lumpability in Markov chains. Journal of Applied Probability, 26(3):446–457, September 1989. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3214403.

Rubino:1989:STF

[RS89c] Gerardo Rubino and Bruno Sericola. Sojourn times in finite Markov processes. Journal of Applied Probability, 26(4):744–756, December 1989. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3214379.

Rhee:1990:DBP

[RS90] Hahn-Kyou Rhee and B. D. Sivazlian. Distribution of the busy period in a controllable M/M/2 queue operating under the triadic (0,k,n,m) policy. Journal of Applied Probability, 27(2):425– 432, June 1990. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3214662.

Righter:1992:EPF

[RS92] Rhonda Righter and J. George Shanthikumar. Extremal properties of the FIFO discipline in queueing networks. Journal of Applied Probability, 29(4):967–978, December 1992. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3214728.

Rogers:1994:CIL

[RS94] L. C. G. Rogers and Z. Shi. Computing the invariant law of a fluid model. Journal of Applied Probability, 31(4):885–896, December 1994. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3215314.

Resnick:1995:CHE

[RS95a] Sidney Resnick and Catalin Starica. Consistency of Hill’s estimator for dependent data. Journal of Applied Probability, 32 (1):139–167, March 1995. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3214926.

Rogers:1995:VAO

[RS95b] L. C. G. Rogers and Z. Shi. The value of an Asian option. Journal of Applied Probability, 32(4):1077–1088, December 1995. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/3215221.

Ross:1999:HTM

[RS99] Sheldon M. Ross and Sridhar Seshadri. Hitting time in an M/G/1 queue. Journal of Applied Probability, 36(3):934–940, September 1999. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3215453.

Ruschendorf:2000:SAP

[RS00] Ludger Rüschendorf and Gernot Sachs. Stochastic analysis of partitioning algorithms for matching problems. Journal of Applied Probability, 37(2):494–503, June 2000. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/3215723.

Rabehasaina:2004:SOM

[RS04] Landy Rabehasaina and Bruno Sericola. A second-order Markov-modulated fluid queue with linear service rate. Journal of Applied Probability, 41(3):758–777, September 2004. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/4141352.

Rogers:2006:CMC

[RS06] L. C. G. Rogers and Larry Shepp. The correlation of the maxima of correlated Brownian motions. Journal of Applied Probability, 43(3):880–883, September 2006. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.

org/stable/27595780.

Rybarczyk:2010:PAN

[RS10] Katarzyna Rybarczyk and Dudley Stark. Poisson approximation of the number of cliques in random intersection graphs. Journal of Applied Probability, 47(3):826–840, September 2010. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/20778519.

Roueff:2015:CSL

[RS15] François Roueff and Philippe Soulier. Convergence to stable laws in the space D. Journal of Applied Probability, 52(1):1–17, March 2015. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-

6072 (electronic). URL http://projecteuclid.org/euclid.
jap/1429282603.

Richey:2022:IRS

[RS22] Jacob Richey and Amites Sarkar. Intersections of random sets. Journal of Applied Probability, 59(1):131–151, March 2022. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL https://www.cambridge.org/core/journals/journal-
of-applied-probability/article/intersections-of-random-
sets/F94BF856EBE488909E2132C3BEBD0DDC.

Ricciardi:1984:IEF

[RSS84] L. M. Ricciardi, L. Sacerdote, and S. Sato. On an integral equation for first-passage-time probability densities. Journal of Applied Probability, 21(2):302–314, June 1984. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/3213641.

Righter:1990:ESD

[RSY90] Rhonda Righter, J. George Shanthikumar, and Genji Yamazaki. On extremal service disciplines in single-stage queueing systems. Journal of Applied Probability, 27(2):409–416, June 1990. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/3214660.

Ross:2005:IFR

[RSZ05] Sheldon M. Ross, J. George Shanthikumar, and Zegang Zhu. On increasing-failure-rate random variables. Journal of Applied Probability, 42(3):797–809, September 2005. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/30040857.

Resnick:1973:ASS

[RT73] Sidney I. Resnick and R. J. Tomkins. Almost sure stability of maxima. Journal of Applied Probability, 10(2):387–401, June 1973. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3212355.

Rachev:1990:RCS

[RT90] S. Rachev and P. Todorovic. On the rate of convergence of some functionals of a stochastic process. Journal of Applied Probability, 27(4):805–814, December 1990. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/3214824.

Rhee:1991:NSR

[RT91] Wansoo Rhee and Michel Talagrand. A note on the selection of random variables under a sum constraint. Journal of Applied Probability, 28(4):919–923, December 1991. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/3214697.

Roberts:2000:RCS

[RT00] G. O. Roberts and R. L. Tweedie. Rates of convergence of stochastically monotone and continuous time Markov models. Journal of Applied Probability, 37(2):359–373, June 2000. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/3215713.

Roberts:2001:GCE

[RT01] Gareth O. Roberts and Richard L. Tweedie. Geometric L^{2} and L^{1} convergence are equivalent for reversible Markov chains. Journal of Applied Probability, 38A:37–41, 2001. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3215866.

Rudolf:2018:CHR

[RU18] Daniel Rudolf and Mario Ullrich. Comparison of hit-and-run, slice sampler and random walk Metropolis. Journal of Applied Probability, 55(4):1186–1202, December 2018. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL https://www.cambridge.org/core/journals/journal-
of-applied-probability/article/comparison-of-hitandrun-
slice-sampler-and-random-walk-metropolis/9E32DA02A6E30F0D405C4D1115759BD3.

Ruben:1964:GCF

[Rub64] Harold Ruben. Generalised concentration fluctuations under diffusion equilibrium. Journal of Applied Probability, 1(1):47– 68, June 1964. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3212060.

Ruben:1977:VRS

[Rub77] Harold Ruben. The volume of a random simplex in an n-ball is asymptotically normal. Journal of Applied Probability, 14(3): 647–653, September 1977. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3213472.

Ruben:1979:VIR

[Rub79] Harold Ruben. The volume of an isotropic random parallelotope. Journal of Applied Probability, 16(1):84–94, March 1979. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/3213377.

Rubinovitch:1985:SSPa

[Rub85a] Michael Rubinovitch. The slow server problem. Journal of Applied Probability, 22(1):205–213, March 1985. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3213760.

Rubinovitch:1985:SSPb

[Rub85b] Michael Rubinovitch. The slow server problem: A queue with stalling. Journal of Applied Probability, 22(4):879–892, December 1985. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3213955.

Rubenthaler:2023:CLT

[Rub23] Sylvain Rubenthaler. A central limit theorem for conservative fragmentation chains. Journal of Applied Probability, 60(3):1031–1068, September 2023. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL https:/

/www.cambridge.org/core/journals/journal-of-applied-
probability/article/central-limit-theorem-for-conservative-
fragmentation-chains/49E3C43C41D2A96ED311CD47FD9EAFB2.

Rudemo:1975:SNS

[Rud75] Mats Rudemo. Some non-stationary point processes with stationary forward recurrence time distribution. Journal of Applied Probability, 12(1):167–169, March 1975. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http:/

/www.jstor.org/stable/3212422.

Rudolfer:1976:NRM

[Rud76a] S. M. Rudolfer. A note on roots of Markov shifts. Journal of Applied Probability, 13(3):591–596, September 1976. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3212480.

Rudolfer:1976:DH

[Rud76b] S. M. Rudolfer. On the distribution of hits. Journal of Applied Probability, 13(1):164–168, March 1976. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http:/

/www.jstor.org/stable/3212679.

Ruggiero:2014:SDT

[Rug14] Matteo Ruggiero. Species dynamics in the two-parameter Poisson–Dirichlet diffusion model. Journal of Applied Probability, 51(1):174–190, March 2014. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

projecteuclid.org/euclid.jap/1395771422.

Ruiz:2008:CFH

[Rui08] Víctor Ruiz. A compact framework for hidden Markov chains with applications to fractal geometry. Journal of Applied Probability, 45(3):630–639, September 2008. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/27595976.

Rukhin:2000:AET

[Ruk00] Andrew L. Rukhin. Approximate entropy for testing randomness. Journal of Applied Probability, 37(1):88–100, March 2000. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/3215661.

Ruschendorf:1977:TMP

[Rüs77] Ludger Rüschendorf. On two marked point processes. Journal of Applied Probability, 14(2):320–327, June 1977. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3213002.

Rust:1978:VDS

[Rus78] Philip F. Rust. The variance of duration of stay in an absorbing Markov process. Journal of Applied Probability, 15(2):420– 425, June 1978. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3213413.

Ruschendorf:1982:CPP

[Rüs82] Ludger Rüschendorf. Comparison of percolation probabilities. Journal of Applied Probability, 19(4):864–868, December 1982. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/3213841.

Russell:1983:NUR

[Rus83] K. G. Russell. On the number of uniform random variables which must be added to exceed a given level. Journal of Applied Probability, 20(1):172–177, March 1983. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3213732.

Russell:2001:TSC

[Rus01] K. G. Russell. The teaching of statistical consulting. Journal of Applied Probability, 38A:20–26, 2001. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3215863.

Ruschendorf:2004:CMR

[Rüs04] Ludger Rüschendorf. Comparison of multivariate risks and positive dependence. Journal of Applied Probability, 41(2):391– 406, June 2004. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3216024.

Ruschendorf:2006:SRE

[Rüs06] Ludger Rüschendorf. On stochastic recursive equations of sum and max type. Journal of Applied Probability, 43(3):687–703, September 2006. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
27595765.

Ruschendorf:2008:CRM

[Rüs08] Ludger Rüschendorf. On a comparison result for Markov processes. Journal of Applied Probability, 45(1):279–286, March 2008. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/27595940.

Ruth:1999:FRB

[Rut99] Kevin Ruth. Favorable red and black on the integers with a minimum bet. Journal of Applied Probability, 36(3):837–851, September 1999. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3215445.

Ruzankin:2004:RPP

[Ruz04] P. S. Ruzankin. On the rate of Poisson process approximation to a Bernoulli process. Journal of Applied Probability, 41(1):271– 276, March 2004. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3215831.

Ruzankin:2010:AEU

[Ruz10] Pavel S. Ruzankin. Approximation for expectations of unbounded functions of dependent integer-valued random variables. Journal of Applied Probability, 47(2):594–600, June 2010.

CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/20778500.

Resnick:2000:WCH

[RV00] Sidney Resnick and Eric Van Den Berg. Weak convergence of high-speed network traffic models. Journal of Applied Probability, 37(2):575–597, June 2000. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.

org/stable/3215731.

Rogers:2008:SVA

[RV08] L. C. G. Rogers and L. A. M. Veraart. A stochastic volatility alternative to Sabr. Journal of Applied Probability, 45(4): 1071–1085, December 2008. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.

org/stable/27596009.

Rubin:1968:DAS

[RVJ68] H. Rubin and D. Vere-Jones. Domains of attraction for the subcritical Galton–Watson branching process. Journal of Applied Probability, 5(1):216–219, April 1968. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3212089.

Reichel:1964:TRI

[RW64] A. Reichel and C. A. Wilkins. On the theory of resonance integrals in the statistical region. Journal of Applied Probability, 1(2):335–346, December 1964. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.

org/stable/3211864.

Robbins:1979:SRP

[RW79] Herbert Robbins and John Whitehead. Sequential rank and the Pólya urn. Journal of Applied Probability, 16(1):213–219, March 1979. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3213390.

Rabehasaina:2021:MBP

[RW21] Landy Rabehasaina and Jae-Kyung Woo. Multitype branching process with non-homogeneous Poisson and contagious Poisson immigration. Journal of Applied Probability, 58 (4):1007–1042, December 2021. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL https:/

/www.cambridge.org/core/journals/journal-of-applied-
probability/article/multitype-branching-process-with-
nonhomogeneous-poisson-and-contagious-poisson-immigration/
5A4C27C246ACBCF51596366296BD3A87.

Richman:2002:CLS

[RWN02] Fred Richman and Katarzyna Winkowska-Nowak. Computing limiting stationary distributions of small noisy networks. Journal of Applied Probability, 39(1):161–178, March 2002. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3215925.

Reed:2013:GDS

[RWZ13] Josh Reed, Amy Ward, and Dongyuan Zhan. On the generalized drift Skorokhod problem in one dimension. Journal of Applied Probability, 50(1):16–28, March 2013. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

projecteuclid.org/euclid.jap/1363784421.

Rahimov:1999:MFS

[RY99] Ibrahim Rahimov and George P. Yanev. On maximum family size in branching processes. Journal of Applied Probability, 36 (3):632–643, September 1999. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.

org/stable/3215429.

Reinert:2020:BRC

[RY20] G. Reinert and C. Yang. A bound on the rate of convergence in the central limit theorem for renewal processes under second moment conditions. Journal of Applied Probability, 57(1):343–360, March 2020. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL https://www.cambridge.org/core/journals/journal-
of-applied-probability/article/bound-on-the-rate-of-
convergence-in-the-central-limit-theorem-for-renewal-
processes-under-second-moment-conditions/5CB3479AAD2A8FA28AD05C0C7F53DE83.

Rychlik:1987:JDS

[Ryc87] Igor Rychlik. Joint distribution of successive zero crossing distances for stationary Gaussian processes. Journal of Applied Probability, 24(2):378–385, June 1987. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3214262.

Ryden:1996:IOC

[Ryd96] Tobias Rydén. On identifiability and order of continuoustime aggregated Markov chains, Markov-modulated Poisson processes, and phase-type distributions. Journal of Applied Probability, 33(3):640–653, September 1996. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3215346.

Renaud:2007:DPV

[RZ07] Jean-François Renaud and Xiaowen Zhou. Distribution of the present value of dividend payments in a Lévy risk model. Journal of Applied Probability, 44(2):420–427, June 2007. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/27595851.

Ross:2008:SSC

[RZ08] Sheldon M. Ross and Zegang Zhu. On the structure of a swing contracts optimal value and optimal strategy. Journal of Applied Probability, 45(1):1–15, March 2008. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/27595919.

Sahin:1972:LBB

[SA72] Izzet Sahin and Oussama Achou. On the limiting behaviour of a basic stochastic process. Journal of Applied Probability, 9 (1):202–207, March 1972. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3212650.

Sharma:1973:CGI

[SA73] Bhu Dev Sharma and Ram Autar. On characterization of a generalized inaccuracy measure in information theory. Journal of Applied Probability, 10(2):464–468, June 1973. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/3212366. See correction

[Ano73f].

Sadeghi:2001:CDM

[Sad01] Ali A. Sadeghi. Convergence in distribution of the multidimensional Kohonen algorithm. Journal of Applied Probability, 38(1):136–151, March 2001. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3215747.

Shiva:1973:OPM

[SAG73] S. G. S. Shiva, N. U. Ahmed, and N. D. Georganas. Order preserving measures of information. Journal of Applied Probability, 10(3):666–670, September 1973. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.

org/stable/3212786.

Sagitov:1999:GCA

[Sag99] Serik Sagitov. The general coalescent with asynchronous mergers of ancestral lines. Journal of Applied Probability, 36(4):1116– 1125, December 1999. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3215582.

Sagitov:2003:CCS

[Sag03] Serik Sagitov. Convergence to the coalescent with simultaneous multiple mergers. Journal of Applied Probability, 40(4):839–854, December 2003. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3216045.

Sagitov:2020:PFT

[Sag20] Serik Sagitov. Perron–Frobenius theory for kernels and Crump– Mode–Jagers processes with macro-individuals. Journal of Applied Probability, 57(3):720–733, September 2020. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL https://www.cambridge.org/core/journals/journal-
of-applied-probability/article/perronfrobenius-theory-
for-kernels-and-crumpmodejagers-processes-with-macroindividuals/
1E25D712CB04208E76ADB6185A93D017.

Sagitov:2023:CBP

[Sag23] Serik Sagitov. Critical branching as a pure death process coming down from infinity. Journal of Applied Probability, 60(2):607–628, June 2023. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL https:/

/www.cambridge.org/core/journals/journal-of-applied-
probability/article/critical-branching-as-a-pure-death-
process-coming-down-from-infinity/0F4652DDA7F1354508A1E488FC0560E1.

Sahin:1971:EBS

[Sah71a] Izzet Sahin. Equilibrium behavior of a stochastic system with secondary input. Journal of Applied Probability, 8(2):252– 260, June 1971. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3211895.

Sahin:1971:SSQ

[Sah71b] Izzet Sahin. On the single server queue with preemptive service interruptions. Journal of Applied Probability, 8(4):835–837, December 1971. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3212248.

Sahin:1978:CCS

[Sah78] Izzet Sahin. Cumulative constrained sojourn times in semi-Markov processes with an application to pensionable service. Journal of Applied Probability, 15(3):531–542, September 1978. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/3213116.

Sahin:1983:CRI

[Sah83] Izzet Sahin. On the continuous-review (s, s) inventory model under compound renewal demand and random lead times. Journal of Applied Probability, 20(1):213–219, March 1983. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3213739.

Sahasrabudhe:2016:SFT

[Sah16] Neeraja Sahasrabudhe. Synchronization and fluctuation theorems for interacting Friedman urns. Journal of Applied Probability, 53(4):1221–1239, December 2016. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

projecteuclid.org/euclid.jap/1481132848.

Scott:1982:SEC

[SAI82] Mark Scott, Barry C. Arnold, and Dean L. Isaacson. Strong ergodicity for continuous-time, non-homogeneous Markov chains. Journal of Applied Probability, 19(3):692–694, September 1982. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/3213529.

Salez:2013:JDD

[Sal13] Justin Salez. Joint distribution of distances in large random regular networks. Journal of Applied Probability, 50(3):861– 870, September 2013. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://projecteuclid.

org/euclid.jap/1378401241.

Samuel:1967:EER

[Sam67] Ester Samuel. On the existence of the expectation of randomly stopped sums. Journal of Applied Probability, 4(1):197– 200, April 1967. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3212312.

Samuel:1970:SPO

[Sam70] Ester Samuel. On some problems in operations research. Journal of Applied Probability, 7(1):157–164, April 1970. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3212156.

Samuels:1971:DBP

[Sam71] M. L. Samuels. Distribution of the branching-process population among generations. Journal of Applied Probability, 8(4):655–667, December 1971. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3212230.

Samuels:1974:CPP

[Sam74] S. M. Samuels. A characterization of the Poisson process. Journal of Applied Probability, 11(1):72–85, March 1974. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3212584.

Sampson:2006:SCN

[Sam06] Koffi Y. Sampson. Structured coalescent with nonconservative migration. Journal of Applied Probability, 43(2):351–362, June 2006. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/27595736.

Sankoff:1971:BPT

[San71] David Sankoff. Branching processes with terminal types: Application to context-free grammars. Journal of Applied Probability, 8(2):233–240, June 1971. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3211893.

Sanghvi:1980:CSA

[San80] Arun P. Sanghvi. Convergence of successive approximations to a stochastic fixed point. Journal of Applied Probability, 17(1):297– 299, March 1980. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3212950.

Santacroce:2006:DPO

[San06] Marina Santacroce. Derivatives pricing via p-optimal martingale measures: Some extreme cases. Journal of Applied Probability, 43(3):634–651, September 2006. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.

org/stable/27595761.

Saperstein:1975:NCP

[Sap75] B. Saperstein. Note on a clustering problem. Journal of Applied Probability, 12(3):629–632, September 1975. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/3212882.

Saporu:1987:SCP

[Sap87] F. W. O. Saporu. On the solution of compartmental processes and their use in modelling stages of infection of a disease with reference to onchocerciasis. Journal of Applied Probability, 24 (2):293–303, June 1987. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3214254.

Saretsalo:1973:OST

[Sar73] Lauri Saretsalo. On the optimal search for a target whose motion is a Markov process. Journal of Applied Probability, 10(4): 847–856, December 1973. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3212386.

Sarkar:1997:CCP

[Sar97] Anish Sarkar. Continuity and convergence of the percolation function in continuum percolation. Journal of Applied Probability, 34(2):363–371, June 1997. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3215376.

Stoyanov:2004:PCL

[SAS04] Jordan Stoyanov and M. H. M. Al-Sadi. Properties of classes of life distribution based on the conditional variance. Journal of Applied Probability, 41(4):953–960, December 2004. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/4141369.

Sato:1977:EFP

[Sat77] Shunsuke Sato. Evaluation of the first-passage time probability to a square root boundary for the Wiener process. Journal of Applied Probability, 14(4):850–856, December 1977. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3213358.

Sato:1983:LDS

[Sat83] Ken-Iti Sato. Limit diffusions of some stepping-stone models. Journal of Applied Probability, 20(3):460–471, September 1983. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/3213884.

Saunders:1975:CPS

[Sau75a] Roy Saunders. Conservative processes with stochastic rates. Journal of Applied Probability, 12(3):447–456, September 1975. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/3212859.

Saunders:1975:LRC

[Sau75b] Roy Saunders. Limiting results for carrier-borne epidemics. Journal of Applied Probability, 12(4):793–799, December 1975. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/3212729.

Saunders:1976:MBC

[Sau76a] D. I. Saunders. Malthusian behaviour of the critical and subcritical age-dependent branching processes with arbitrary state space. Journal of Applied Probability, 13(3):455–465, September 1976. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3212465.

Saunders:1976:CTP

[Sau76b] I. W. Saunders. A convergence theorem for parities in a birthand-death process. Journal of Applied Probability, 13(2):231– 238, June 1976. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3212826.

Saunders:1976:JEC

[Sau76c] Roy Saunders. On joint exchangeability and conservative processes with stochastic rates. Journal of Applied Probability, 13 (3):584–590, September 1976. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.

org/stable/3212479.

Saunders:2023:NTC

[Sau23] John C. Saunders. The number of K-tons in the coupon collector problem. Journal of Applied Probability, 60(3): 723–736, September 2023. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL https:/

/www.cambridge.org/core/journals/journal-of-applied-
probability/article/number-of-ktons-in-the-coupon-collector-
problem/41842B5D6FDE42F3DE8F7E7CAAB22121.

Savits:1985:MIC

[Sav85] Thomas H. Savits. A multivariate IFR class. Journal of Applied Probability, 22(1):197–204, March 1985. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http:/

/www.jstor.org/stable/3213759.

Savits:1988:CRB

[Sav88a] Thomas H. Savits. A cost relationship between age and block replacement policies. Journal of Applied Probability, 25(4):789– 796, December 1988. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3214300.

Savits:1988:SMD

[Sav88b] Thomas H. Savits. Some multivariate distributions derived from a non-fatal shock model. Journal of Applied Probability, 25 (2):383–390, June 1988. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3214446.

Savits:2003:PGB

[Sav03] Thomas H. Savits. Preservation of generalized bathtub-shaped functions. Journal of Applied Probability, 40(2):473–484, June 2003. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3215804.

Sawyer:1977:PHA

[Saw77] Stanley Sawyer. On the past history of an allele now known to have frequency p. Journal of Applied Probability, 14(3):439–450, September 1977. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3213447.

Sawyer:1979:LTP

[Saw79] Stanley Sawyer. A limit theorem for patch sizes in a selectivelyneutral migration model. Journal of Applied Probability, 16(3): 482–495, September 1979. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3213078.

Sawyer:1983:SPE

[Saw83] Stanley Sawyer. A stability property of the Ewens sampling formula. Journal of Applied Probability, 20(3):449–459, September 1983. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3213883.

Silverman:1978:SDF

[SB78] Bernard Silverman and Tim Brown. Short distances, flat triangles and Poisson limits. Journal of Applied Probability, 15(4): 815–825, December 1978. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3213436.

Swan:2006:MAA

[SB06] Yvik C. Swan and F. Thomas Bruss. A matrix-analytic approach to the N-player ruin problem. Journal of Applied Probability, 43 (3):755–766, September 2006. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.

org/stable/27595770.

Socoli:2009:LLA

[SB09] Sanda N. Socoli and A. D. Barbour. Local limit approximations for Markov population processes. Journal of Applied Probability, 46(3):690–708, September 2009. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.

org/stable/25662455.

Suvinthra:2017:LDS

[SB17] M. Suvinthra and K. Balachandran. Large deviations for the stochastic predator-prey model with nonlinear functional response. Journal of Applied Probability, 54(2):507–521, June 2017. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-

6072 (electronic). URL https://www.cambridge.org/core/
journals/journal-of-applied-probability/article/large-
deviations-for-the-stochastic-predatorprey-model-with-
nonlinear-functional-response/757C26EDA6C0983CA36017DCD34A08C0.

Suh:1970:DMS

[SBG70] M. W. Suh, B. B. Bhattacharyya, and A. Grandage. On the distribution and moments of the strength of a bundle of filaments. Journal of Applied Probability, 7(3):712–720, December 1970. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3211948.

Sparks:2022:CPH

[SBM22] Joshua Sparks, Srinivasan Balaji, and Hosam Mahmoud. The containment profile of hyper-recursive trees. Journal of Applied Probability, 59(1):278–296, March 2022. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL https:/

/www.cambridge.org/core/journals/journal-of-applied-
probability/article/containment-profile-of-hyperrecursive-
trees/04D100CB1E1F0BE2CFA75C9747D951CB.

Sudbury:1973:SRG

[SC73] Aidan Sudbury and Peter Clifford. Some results for general cellsize-dependent branching processes. Journal of Applied Probability, 10(2):289–298, June 1973. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3212346.

Samuel-Cahn:1975:ADC

[SC75] Ester Samuel-Cahn. Asymptotic distribution for the couponcollector’s and sampling-tagging problems, when tagging affects catchability. Journal of Applied Probability, 12(3):625–628, September 1975. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3212881.

Sudbury:1981:DAC

[SC81] Aidan Sudbury and Peter Clifford. The determination of the age of a clone from characteristics of its geographical distribution. Journal of Applied Probability, 18(3):725–731, September 1981. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3213327.

Smith:1982:SMP

[SC82] Charles E. Smith and Loren Cobb. The stationary moments of Poisson-driven non-linear dynamical systems. Journal of Applied Probability, 19(3):702–706, September 1982. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/3213531.

Seneta:2002:EFO

[SC02] Eugene Seneta and John Tuhao Chen. On explicit and Fréchet– Optimal lower bounds. Journal of Applied Probability, 39(1):81– 90, March 2002. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3215919.

Scarsini:1998:MCO

[Sca98] Marco Scarsini. Multivariate convex orderings, dependence, and stochastic equality. Journal of Applied Probability, 35(1):93– 103, March 1998. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3215549.

Schweitzer:1968:PTF

[Sch68] Paul J. Schweitzer. Perturbation theory and finite Markov chains. Journal of Applied Probability, 5(2):401–413, August 1968. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3212261.

Schweder:1970:CMP

[Sch70] Tore Schweder. Composable Markov processes. Journal of Applied Probability, 7(2):400–410, August 1970. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3211973.

Schassberger:1972:WLP

[Sch72] R. Schassberger. On the work load process in a general preemptive resume priority queue. Journal of Applied Probability, 9(3):

588–603, September 1972. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3212328.

Schassberger:1975:BAH

[Sch75a] R. Schassberger. Basic analysis of a head-of-the-line priority queue with renewal type input. Journal of Applied Probability, 12(2):346–352, June 1975. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3212449.

Schreuer:1975:SPE

[Sch75b] Mynda Schreuer. Stochastic pursuit-evasion games with information lag: I. Perfect observation. Journal of Applied Probability, 12(2):248–256, June 1975. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3212438.

Schreuer:1976:SPE

[Sch76] Mynda Schreuer. Stochastic pursuit-evasion games with information lag: II. Observation with error. Journal of Applied Probability, 13(2):313–328, June 1976. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3212834.

Schroeter:1980:AHT

[Sch80a] Gerhard Schroeter. The application of the Hankel transformation to spherically symmetric coverage problems. Journal of Applied Probability, 17(4):1121–1126, December 1980. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3213225.

Schurger:1980:AGB

[Sch80b] Klaus Schürger. On the asymptotic geometrical behaviour of percolation processes. Journal of Applied Probability, 17(2):385– 402, June 1980. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3213028.

Schuh:1982:SDR

[Sch82] H.-J. Schuh. Sums of I.I.D. random variables and an application to the explosion criterion for Markov branching processes. Journal of Applied Probability, 19(1):29–38, March 1982. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3213913. See note [Don84a].

Schneider:1985:IRF

[Sch85] Rolf Schneider. Inequalities for random flats meeting a convex body. Journal of Applied Probability, 22(3):710–716, September 1985. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3213874.

Schmidt:1986:SIS

[Sch86] Gerd Schmidt. The Spitzer identity for some bivariate processes in queueing theory: An algebraic approach. Journal of Applied Probability, 23(4):1000–1007, December 1986. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3214472.

Schassberger:1988:SSD

[Sch88a] R. Schassberger. The steady-state distribution of spent service times present in the M/G/1 foreground-background processorsharing queue. Journal of Applied Probability, 25(1):194–203, March 1988. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3214245.

Scheinberg:1988:WTM

[Sch88b] Stephen Scheinberg. Waiting times for multiple occurrences of several events. Journal of Applied Probability, 25(3):624–629, September 1988. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3213990.

Schmitt:1991:EDS

[Sch91a] Michel Schmitt. Estimation of the density in a stationary Boolean model. Journal of Applied Probability, 28(3):702–708, September 1991. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3214504.

Schmuland:1991:RIM

[Sch91b] B. Schmuland. A result on the infinitely many neutral alleles diffusion model. Journal of Applied Probability, 28(2):253– 267, June 1991. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3214864.

Scheike:1992:BCR

[Sch92a] Thomas H. Scheike. A boundary-crossing result for Brownian motion. Journal of Applied Probability, 29(2):448–453, June 1992. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3214581.

Scheike:1992:GRP

[Sch92b] Thomas H. Scheike. A general risk process and its properties. Journal of Applied Probability, 29(1):73–81, March 1992. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/3214792.

Schwarz:1992:WPB

[Sch92c] Wolfgang Schwarz. The Wiener process between a reflecting and an absorbing barrier. Journal of Applied Probability, 29(3): 597–604, September 1992. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3214896.

Scheutzow:1993:MAP

[Sch93] Michael Scheutzow. A martingale approach to the Pasta property. Journal of Applied Probability, 30(1):252–257, March 1993. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/3214637.

Schmidli:1996:LIC

[Sch96] Hanspeter Schmidli. Lundberg inequalities for a Cox model with a piecewise constant intensity. Journal of Applied Probability, 33 (1):196–210, March 1996. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3215277.

Schottl:1998:OSR

[Sch98] A. Schöttl. Optimal stopping of a risk reserve process with interest and cost rates. Journal of Applied Probability, 35(1):115– 123, March 1998. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3215551.

Schroder:2003:BEP

[Sch03] Michael Schröder. Brownian excursions and Parisian barrier options: A note. Journal of Applied Probability, 40(4):855–864, December 2003. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3216046.

Schweinsberg:2005:IBP

[Sch05] Jason Schweinsberg. Improving on bold play when the gambler is restricted. Journal of Applied Probability, 42(2):321–333, June 2005. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/30040795.

Schlemm:2011:MTK

[Sch11] Eckhard Schlemm. On the Markov transition kernels for first passage percolation on the ladder. Journal of Applied Probability, 48(2):366–388, June 2011. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/23065891.

Schneider:2019:IPH

[Sch19] Rolf Schneider. Interaction of Poisson hyperplane processes and convex bodies. Journal of Applied Probability, 56(4): 1020–1032, December 2019. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL https:/

/www.cambridge.org/core/journals/journal-of-applied-
probability/article/interaction-of-poisson-hyperplane-
processes-and-convex-bodies/DD1D78E565A0F52F6F58E6EA5C81F3C2.

Scott:1982:CTT

[Sco82] Elizabeth L. Scott. Comparison of two treatments when there may be an initial effect. Journal of Applied Probability, 19A: 253–264, 1982. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3213565.

Samuel-Cahn:1977:ACI

[SCZ77] Ester Samuel-Cahn and Shmuel Zamir. Algebraic characterization of infinite Markov chains where movement to the right is limited to one step. Journal of Applied Probability, 14(4): 740–747, December 1977. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3213347.

Solari:1976:NSL

[SD76] M. E. Solari and J. E. A. Dunnage. A note on Spitzer’s lemma and its application to the maximum of partial sums of dependent random variables. Journal of Applied Probability, 13(2):361– 364, June 1976. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3212841.

Schassberger:1987:STQ

[SD87] R. Schassberger and H. Daduna. Sojourn times in queuing networks with multiserver modes. Journal of Applied Probability, 24(2):511–521, June 1987. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3214274.

Smith:1988:HBE

[SD88] Laurel Smith and Persi Diaconis. Honest Bernoulli excursions. Journal of Applied Probability, 25(3):464–477, September 1988. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/3213976.

Sengupta:1994:SRR

[SD94] Debasis Sengupta and Jayant V. Deshpande. Some results on the relative ageing of two life distributions. Journal of Applied Probability, 31(4):991–1003, December 1994. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http:/

/www.jstor.org/stable/3215323.

Stenzel:2012:WCL

[SD12] Ole Stenzel and Hans Daduna. Weak convergence limits for closed cyclic networks of queues with multiple bottleneck nodes. Journal of Applied Probability, 49(1):60–83, March 2012. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/41713752.

Szekli:1994:MGQ

[SDH94] R. Szekli, R. L. Disney, and S. Hur. MR/GI/1 queues with positively correlated arrival stream. Journal of Applied Probability, 31(2):497–514, June 1994. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3215041.

Sommer:2016:NJN

[SDH16] Jennifer Sommer, Hans Daduna, and Bernd Heidergott. Nonergodic Jackson networks with infinite supply-local stabilization and local equilibrium analysis. Journal of Applied Probability, 53(4):1125–1142, December 2016. CODEN JPRBAM.

ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

projecteuclid.org/euclid.jap/1481132841.

Shen:2012:ODD

[SE12] Leo Shen and Robert Elliott. Optimal design of dynamic default risk measures. Journal of Applied Probability, 49(4):967–977, December 2012. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
41713822.

Secchi:1997:TPR

[Sec97] Piercesare Secchi. Two-person red-and-black stochastic games. Journal of Applied Probability, 34(1):107–126, March 1997. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/3215180.

Segers:2006:RET

[Seg06] Johan Segers. Rare events, temporal dependence, and the extremal index. Journal of Applied Probability, 43(2):463– 485, June 2006. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
27595744.

Seifert:2016:WIA

[Sei16] Miriam Isabel Seifert. Weakening the independence assumption on polar components: limit theorems for generalized elliptical distributions. Journal of Applied Probability, 53(1):130–145, March 2016. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-

6072 (electronic). URL http://projecteuclid.org/euclid.
jap/1457470564.

Sellke:1983:ADS

[Sel83] Thomas Sellke. On the asymptotic distribution of the size of a stochastic epidemic. Journal of Applied Probability, 20(2):390– 394, June 1983. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3213811.

Seleznjev:1991:LTM

[Sel91] O. V. Seleznjev. Limit theorems for maxima and crossings of a sequence of Gaussian processes and approximation of random processes. Journal of Applied Probability, 28(1):17–32, March 1991. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3214737.

Selim:1997:TDS

[Sel97] Shokri Z. Selim. Time-dependent solution and optimal control of a bulk service queue. Journal of Applied Probability, 34(1):258– 266, March 1997. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3215192.

Seneta:1967:GWP

[Sen67] E. Seneta. The Galton–Watson process with mean one. Journal of Applied Probability, 4(3):489–495, December 1967. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3212216.

Seneta:1971:IMS

[Sen71] E. Seneta. On invariant measures for simple branching processes. Journal of Applied Probability, 8(1):43–51, March 1971. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3211836.

Sen:1973:AET

[Sen73a] Pranab Kumar Sen. An asymptotically efficient test for the bundle strength of filaments. Journal of Applied Probability, 10 (3):586–596, September 1973. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.

org/stable/3212778.

Seneta:1973:SBP

[Sen73b] E. Seneta. The simple branching process with infinite mean. I. Journal of Applied Probability, 10(1):206–212, March 1973. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/3212508.

Sen:1981:WCI

[Sen81] Pranab Kumar Sen. Weak convergence of an iterated renewal process. Journal of Applied Probability, 18(1):291–296, March 1981. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3213191.

Seneta:1982:EMM

[Sen82a] E. Seneta. Entropy and martingales in Markov chain models. Journal of Applied Probability, 19A:367–381, 1982. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3213576.

Sengupta:1982:ER

[Sen82b] Bhaskar Sengupta. An exponential riddle. Journal of Applied Probability, 19(3):737–740, September 1982. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/3213538.

Sennott:1987:CNE

[Sen87] Linn I. Sennott. Conditions for the non-ergodicity of Markov chains with application to a communication system. Journal of Applied Probability, 24(2):339–346, June 1987. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3214258.

Sengupta:1989:PMS

[Sen89] Bhaskar Sengupta. A perturbation method for solving some queues with processor sharing discipline. Journal of Applied Probability, 26(1):209–214, March 1989. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http:/

/www.jstor.org/stable/3214333. See comment [Bré91].

Sengupta:1994:ALM

[Sen94] Debasis Sengupta. Another look at the moment bounds on reliability. Journal of Applied Probability, 31(3):777–787, September 1994. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3215155.

Sengupta:1995:RBC

[Sen95] Debasis Sengupta. Reliability bounds for the L-class and Laplace order. Journal of Applied Probability, 32(3):832–835, September 1995. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3215133.

Seneta:2001:COP

[Sen01] E. Seneta. Characterization by orthogonal polynomial systems of finite Markov chains. Journal of Applied Probability, 38A:42– 52, 2001. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3215867.

Seneta:2004:FVG

[Sen04] Eugene Seneta. Fitting the variance-gamma model to financial data. Journal of Applied Probability, 41A:177–187, 2004. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/3215976.

Seneta:2017:OJM

[Sen17] Eugene Seneta. Obituary: Joseph Mark Gani AM, DSc, FAA, FASSA. Journal of Applied Probability, 54(1):1–11, March 2017. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL https://www.cambridge.org/core/journals/
journal-of-applied-probability/article/joseph-mark-gani-
am-dsc-faa-fassa/D1A14F7BF0FB70FCB4282EBA4270EA5B.

Serfozo:1972:CPP

[Ser72a] Richard F. Serfozo. Conditional Poisson processes. Journal of Applied Probability, 9(2):288–302, June 1972. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3212799.

Serfozo:1972:PCS

[Ser72b] Richard F. Serfozo. Processes with conditional stationary independent increments. Journal of Applied Probability, 9(2):303– 315, June 1972. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3212800.

Serfozo:1980:HLE

[Ser80] Richard Serfozo. High-level exceedances of regenerative and semi-stationary processes. Journal of Applied Probability, 17 (2):423–431, June 1980. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3213031.

Serfozo:1984:RCP

[Ser84] Richard F. Serfozo. Rarefactions of compound point processes. Journal of Applied Probability, 21(4):710–719, December 1984. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/3213689.

Sericola:1990:CFS

[Ser90] Bruno Sericola. Closed-form solution for the distribution of the total time spent in a subset of states of a homogeneous Markov process during a finite observation period. Journal of Applied Probability, 27(3):713–719, September 1990. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/3214555.

Serfozo:1992:AKR

[Ser92] Richard F. Serfozo. Applications of the key renewal theorem: Crudely regenerative processes. Journal of Applied Probability, 29(2):384–395, June 1992. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3214575.

Serra:2006:WTE

[Ser06] Maria Conceição Serra. On the waiting time to escape. Journal of Applied Probability, 43(1):296–302, March 2006. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/27595729.

Seshadri:1996:SPA

[Ses96] Sridhar Seshadri. A sample path analysis of the delay in the M/G/C system. Journal of Applied Probability, 33(1):256–266, March 1996. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3215282.

Sethuraman:2003:CLS

[Set03] Sunder Sethuraman. A clustering law for some discrete order statistics. Journal of Applied Probability, 40(1):226–241, March 2003. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3215850.

Severo:1967:TTS

[Sev67] Norman C. Severo. Two theorems on solutions of differentialdifference equations and applications to epidemic theory. Journal of Applied Probability, 4(2):271–280, August 1967. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3212022.

Severo:1969:RTS

[Sev69] Norman C. Severo. A recursion theorem on solving differentialdifference equations and applications to some stochastic processes. Journal of Applied Probability, 6(3):673–681, December 1969. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3212111.

Severo:1970:NEM

[Sev70] Norman C. Severo. A note on the Ehrenfest multiurn model. Journal of Applied Probability, 7(2):444–445, August 1970. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/3211978.

Seva:1995:LLT

[Sév95] Marc Séva. On the local limit theorem for non-uniformly ergodic Markov chains. Journal of Applied Probability, 32(1):52– 62, March 1995. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3214920.

Sezer:2018:AEB

[Sez18] Ali Devin Sezer. Approximation of excessive backlog probabilities of two tandem queues. Journal of Applied Probability, 55(3):968–997, September 2018. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL https:/

/www.cambridge.org/core/journals/journal-of-applied-
probability/article/approximation-of-excessive-backlog-
probabilities-of-two-tandem-queues/E551507ACB60AB4729EB3701CD8368E8.

Saunders:1977:PLC

[SF77] Roy Saunders and Gerald M. Funk. Poisson limits for a clustering model of Strauss. Journal of Applied Probability, 14(4): 776–784, December 1977. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3213350.

Sawyer:1983:IDH

[SF83] Stanley Sawyer and Joseph Felsenstein. Isolation by distance in a hierarchically clustered population. Journal of Applied Probability, 20(1):1–10, March 1983. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.

org/stable/3213715.

Sagitov:2017:LTP

[SF17] Serik Sagitov and Thibaut France. Limit theorems for pure death processes coming down from infinity. Journal of Applied Probability, 54(3):720–731, September 2017. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL https:/

/www.cambridge.org/core/journals/journal-of-applied-
probability/article/limit-theorems-for-pure-death-processes-
coming-down-from-infinity/99C4EA2CDDE73E1F965DF0CFEB0E1A9F.

Svoronos:1988:CRS

[SG88] Antony Svoronos and Linda Green. A convexity result for singleserver exponential loss systems with non-stationary arrivals. Journal of Applied Probability, 25(1):224–227, March 1988. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/3214250.

Santana:1989:PPO

[SG89] Pedro Saavedra Santana and Miguel Sanchez Garcia. A population process with offspring distribution depending on both population size and generation. Journal of Applied Probability, 26(4):866–872, December 1989. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3214391.

Sheu:1992:MIR

[SG92] Shey-Huei Sheu and William S. Griffith. Multivariate imperfect repair. Journal of Applied Probability, 29(4):947–956, December 1992. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3214726.

Sparaggis:1993:OBA

[SG93] Panayotis D. Sparaggis and Wei-Bo Gong. Optimal buffer allocation in a two-stage queueing system. Journal of Applied Probability, 30(2):478–482, June 1993. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3214858.

Singh:2016:DCA

[SGC16] Gagandeep Singh, U. C. Gupta, and M. L. Chaudhry. Detailed computational analysis of queueing-time distributions of the BMAP/G/1 queue using roots. Journal of Applied Probability, 53(4):1078–1097, December 2016. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

projecteuclid.org/euclid.jap/1481132838.

Sgibnev:1996:IVS

[Sgi96] M. S. Sgibnev. An infinite variance solidarity theorem for Markov renewal functions. Journal of Applied Probability, 33 (2):434–438, June 1996. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3215067.

Sgibnev:1999:ESM

[Sgi99] M. S. Sgibnev. On the existence of submultiplicative moments for the stationary distributions of some Markovian random walks. Journal of Applied Probability, 36(1):78–85, March 1999. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3215403.

Sgibnev:2001:AEE

[Sgi01] M. S. Sgibnev. An asymptotic expansion for the expectation of an age-dependent branching process with a submultiplicative estimate of the remainder. Journal of Applied Probability, 38(3): 807–814, September 2001. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3216135.

Schulte-Geers:2015:MPP

[SGS15] Ernst Schulte-Geers and Wolfgang Stadje. Maximal percentages in Pólya’s urn. Journal of Applied Probability, 52(1):180–190, March 2015. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-

6072 (electronic). URL http://projecteuclid.org/euclid.
jap/1429282614.

Schulte-Geers:2017:SDL

[SGS17] Ernst Schulte-Geers and Wolfgang Stadje. Small drift limit theorems for random walks. Journal of Applied Probability, 54(1):199–212, March 2017. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL https:/

/www.cambridge.org/core/journals/journal-of-applied-
probability/article/small-drift-limit-theorems-for-random-
walks/6339CC37548CD42F8BB0CE23989BD270.

Strawderman:1970:LIL

[SH70a] William E. Strawderman and Paul T. Holmes. On the law of the iterated logarithm for inter-record times. Journal of Applied Probability, 7(2):432–439, August 1970. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http:/

/www.jstor.org/stable/3211976.

Sweet:1970:SSD

[SH70b] A. L. Sweet and J. C. Hardin. Solutions for some diffusion processes with two barriers. Journal of Applied Probability, 7(2): 423–431, August 1970. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3211975.

Siegel:1982:CCR

[SH82] Andrew F. Siegel and Lars Holst. Covering the circle with random arcs of random sizes. Journal of Applied Probability, 19 (2):373–381, June 1982. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3213488.

Shmerling:2003:SJP

[SH03] Efraim Shmerling and Kenneth J. Hochberg. Solution of jump parameter systems of differential and difference equations with semi-Markov coefficients. Journal of Applied Probability, 40 (2):442–454, June 2003. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3215802.

Schilling:2021:TPL

[SH21] Judith Schilling and Norbert Henze. Two Poisson limit theorems for the coupon collector’s problem with group drawings. Journal of Applied Probability, 58(4):966–977, December 2021. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-

6072 (electronic). URL https://www.cambridge.org/core/
journals/journal-of-applied-probability/article/two-
poisson-limit-theorems-for-the-coupon-collectors-problem-
with-group-drawings/14AC31D4C06D477DD4A613B846AE3CAE.

Shanbhag:1966:ISQ

[Sha66] D. N. Shanbhag. On infinite server queues with batch arrivals. Journal of Applied Probability, 3(1):274–279, June 1966. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/3212053.

Shaw:1971:BPV

[Sha71] Leonard Shaw. Busy periods in vehicle queues. Journal of Applied Probability, 8(2):261–268, June 1971. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/3211896.

Shanbhag:1972:LE

[Sha72a] D. N. Shanbhag. Letter to the Editor. Journal of Applied Probability, 9(2):470, June 1972. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3212820.

Shanbhag:1972:SCB

[Sha72b] D. N. Shanbhag. Some characterizations based on the Bhattacharya matrix. Journal of Applied Probability, 9(3):580–587, September 1972. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3212327.

Sharma:1973:MQS

[Sha73] O. P. Sharma. A model for queues in series. Journal of Applied Probability, 10(3):691–696, September 1973. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/3212791.

Shanbhag:1974:EPR

[Sha74a] D. N. Shanbhag. An elementary proof for the Rao–Rubin characterization of the Poisson distribution. Journal of Applied Probability, 11(1):211–215, March 1974. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3212601.

Shanbhag:1974:SET

[Sha74b] D. N. Shanbhag. Some extensions of Takác’s limit theorems. Journal of Applied Probability, 11(4):752–761, December 1974. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/3212558.

Sharma:1975:CDT

[Sha75] S. D. Sharma. On a continuous/discrete time queueing system with arrivals in batches of variable size and correlated departures. Journal of Applied Probability, 12(1):115–129, March 1975. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3212413.

Shanbhag:1977:ERR

[Sha77] D. N. Shanbhag. An extension of the Rao–Rubin characterization of the Poisson distribution. Journal of Applied Probability, 14(3):640–646, September 1977. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.

org/stable/3213471.

Sharpe:1979:LCC

[Sha79] Ken Sharpe. Level crossings of a constructed process. Journal of Applied Probability, 16(1):206–212, March 1979. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3213389.

Shanthikumar:1980:SAC

[Sha80] J. G. Shanthikumar. Some analyses on the control of queues using level crossings of regenerative processes. Journal of Applied Probability, 17(3):814–821, September 1980. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/3212974.

Shaked:1981:BDM

[Sha81] Moshe Shaked. Bounds on the distance of a mixture from its parent distribution. Journal of Applied Probability, 18(4):853–863, December 1981. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3213060.

Shaked:1982:DOD

[Sha82] Moshe Shaked. Dispersive ordering of distributions. Journal of Applied Probability, 19(2):310–320, June 1982. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/3213483. See correction

[Ano85e].

Shanthikumar:1987:SCR

[Sha87] J. George Shanthikumar. On stochastic comparison of random vectors. Journal of Applied Probability, 24(1):123–136, March 1987. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3214065.

Sharma:1988:SCS

[Sha88] Vinod Sharma. Stability and continuity for slotted Aloha with stationary non-independent input traffic. Journal of Applied Probability, 25(4):797–807, December 1988. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/3214301.

Shamsiev:1989:CSN

[Sha89a] Rahim Shamsiev. Circuit-switching networks with priorities on a one-dimensional lattice. Journal of Applied Probability, 26(4):

901–905, December 1989. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3214397.

Sharma:1989:NMT

[Sha89b] S. D. Sharma. On a non-Markovian time-dependent queueing system. Journal of Applied Probability, 26(1):142–151, March 1989. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3214324.

Shapiro:2013:CSE

[Sha13] Alexander Shapiro. Consistency of sample estimates of risk averse stochastic programs. Journal of Applied Probability, 50 (2):533–541, June 2013. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://projecteuclid.

org/euclid.jap/1371648959.

Shang:2023:CRR

[Sha23] Yilun Shang. On connectivity and robustness of random graphs with inhomogeneity. Journal of Applied Probability, 60(1):284–294, March 2023. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL https:/

/www.cambridge.org/core/journals/journal-of-applied-
probability/article/on-connectivity-and-robustness-of-
random-graphs-with-inhomogeneity/63E67543B2CC627844B0E62DBD5599BA.

Shepp:1979:JDM

[She79] L. A. Shepp. The joint density of the maximum and its location for a Wiener process with drift. Journal of Applied Probability, 16(2):423–427, June 1979. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3212910.

Sheu:1992:OBR

[She92] Shey-Huei Sheu. Optimal block replacement policies with multiple choice at failure. Journal of Applied Probability, 29(1):129– 141, March 1992. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3214797.

Sheu:1996:MBR

[She96] Shey-Huei Sheu. A modified block replacement policy with two variables and general random minimal repair cost. Journal of Applied Probability, 33(2):557–572, June 1996. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3215079.

Sherlock:2013:OSR

[She13] Chris Sherlock. Optimal scaling of the random walk Metropolis: general criteria for the 0.234 acceptance rule. Journal of Applied Probability, 50(1):1–15, March 2013. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

projecteuclid.org/euclid.jap/1363784420.

Shinozuka:1965:TSB

[Shi65] Masanobu Shinozuka. On the two sided barrier problem. Journal of Applied Probability, 2(1):79–87, June 1965. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3211875.

Shibata:1986:CMS

[Shi86] Ritei Shibata. Consistency of model selection and parameter estimation. Journal of Applied Probability, 23A:127–141, 1986. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/3214348.

Shinoda:2002:EPT

[Shi02] Masato Shinoda. Existence of phase transition of percolation on Sierpiński carpet lattices. Journal of Applied Probability, 39(1): 1–10, March 2002. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3215913.

Shiozawa:2006:EBS

[Shi06] Yuichi Shiozawa. Extinction of branching symmetric α-stable processes. Journal of Applied Probability, 43(4):1077–1090, December 2006. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
27595801.

Shorrock:1972:LTI

[Sho72a] R. W. Shorrock. A limit theorem for inter-record times. Journal of Applied Probability, 9(1):219–223, March 1972. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/3212653. See correction

[Tam71].

Shorrock:1972:RVR

[Sho72b] R. W. Shorrock. On record values and record times. Journal of Applied Probability, 9(2):316–326, June 1972. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3212801.

Shorrock:1973:RVI

[Sho73] R. W. Shorrock. Record values and inter-record times. Journal of Applied Probability, 10(3):543–555, September 1973. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3212775.

Shorrock:1975:EPR

[Sho75] R. W. Shorrock. Extremal processes and random measures. Journal of Applied Probability, 12(2):316–323, June 1975. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/3212445.

Scott:1983:PIS

[SI83] Mark Scott and Dean L. Isaacson. Proportional intensities and strong ergodicity for Markov processes. Journal of Applied Probability, 20(1):185–190, March 1983. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3213734.

Siegel:1978:RAC

[Sie78a] Andrew F. Siegel. Random arcs on the circle. Journal of Applied Probability, 15(4):774–789, December 1978. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/3213433.

Siegel:1978:RSF

[Sie78b] Andrew F. Siegel. Random space filling and moments of coverage in geometrical probability. Journal of Applied Probability, 15 (2):340–355, June 1978. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3213406.

Siegmund:1979:CDA

[Sie79] D. Siegmund. Corrected diffusion approximations in certain random walk problems. Advances in Applied Probability, 11(4): 701–719, December 1979. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/ stable/1426855. See correction [Hog86].

Siedersleben:1981:DOR

[Sie81] Johannes Siedersleben. Dynamically optimized replacement with a Markovian renewal process. Journal of Applied Probability, 18(3):641–651, September 1981. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3213318.

Siegrist:1985:EOS

[Sie85] Kyle Siegrist. Estimation and optimal stopping in a debugging model with masking. Journal of Applied Probability, 22(2):336– 345, June 1985. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3213777.

Siegrist:1986:MCB

[Sie86] Kyle Siegrist. Markov chains with binomial time change. Journal of Applied Probability, 23(2):519–523, June 1986. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3214194.

Siegrist:1989:PWG

[Sie89] Kyle Siegrist. n-Point, win-by-k games. Journal of Applied Probability, 26(4):807–814, December 1989. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/3214385.

Sigman:1988:RTQ

[Sig88] Karl Sigman. Regeneration in tandem queues with multiserver stations. Journal of Applied Probability, 25(2):391–403, June 1988. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3214447.

Sigman:1989:NSC

[Sig89] Karl Sigman. Notes on the stability of closed queueing networks. Journal of Applied Probability, 26(3):678–682, September 1989. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/3214428. See

correction [Ano90g].

Silverman:1978:DCT

[Sil78] B. W. Silverman. Distances on circles, toruses and spheres. Journal of Applied Probability, 15(1):136–143, March 1978. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3213243.

Sim:1990:FOA

[Sim90] C. H. Sim. First-order autoregressive models for gamma and exponential processes. Journal of Applied Probability, 27(2):325– 332, June 1990. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3214651.

Sim:1993:FOA

[Sim93] C. H. Sim. First-order autoregressive logistic processes. Journal of Applied Probability, 30(2):467–470, June 1993. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3214856.

Simonot:1997:CRD

[Sim97] F. Simonot. Convergence rate for the distributions of GI/M/1/n and M/GI/1/n as n tends to infinity. Journal of Applied Probability, 34(4):1049–1060, December 1997. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/3215017.

Simonot:1998:CSD

[Sim98] F. Simonot. A comparison of the stationary distributions of GI/M/c/n and GI/M/c. Journal of Applied Probability, 35 (2):510–515, June 1998. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3215705.

Simonot:2000:MBA

[Sim00] F. Simonot. Moments based approximation for the stationary distribution of a random walk in Z_{+} with an application to the M/GI/1/n queueing system. Journal of Applied Probability, 37 (1):290–299, March 2000. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3215678.

Simatos:2014:CLO

[Sim14] Florian Simatos. Coupling limit order books and branching random walks. Journal of Applied Probability, 51(3):625– 639, September 2014. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://projecteuclid.

org/euclid.jap/1409932663.

Simard:2019:MDS

[Sim19] Clarence Simard. Martingale decomposition of an L^{2} space with nonlinear stochastic integrals. Journal of Applied Probability, 56(4):1231–1243, December 2019. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL https:/

/www.cambridge.org/core/journals/journal-of-applied-
probability/article/martingale-decomposition-of-an-l2-
space-with-nonlinear-stochastic-integrals/E88B22B8C4E5343DA54EC1D6FFCC7753.

Simper:2021:RAU

[Sim21] Mackenzie Simper. Random additions in urns of integers. Journal of Applied Probability, 58(2):335–346, June 2021. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL https://www.cambridge.org/core/journals/journal-
of-applied-probability/article/random-additions-in-urns-
of-integers/EC0358BF7291641D9FD998512E9C048A.

Simon:2023:SED

[Sim23] Matthieu Simon. SIR epidemics driven by Feller processes. Journal of Applied Probability, 60(4):1293–1316, December 2023. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-

6072 (electronic). URL https://www.cambridge.org/core/
journals/journal-of-applied-probability/article/sir-
epidemics-driven-by-feller-processes/FB2C7350D4499E97F786643D093C4452.

Siskind:1970:FCT

[Sis70] Victor Siskind. The fixed cycle traffic light problem: A note on a paper by McNeil. Journal of Applied Probability, 7(1):245– 248, April 1970. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/

3212165. See [McN68].

Soltani:1986:RPM

[SK86] A. Reza Soltani and A. Khodadadi. Run probabilities and the motion of a particle on a given path. Journal of Applied Probability, 23(1):28–41, March 1986. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3214114.

Shanthikumar:1990:UCS

[SK90] J. George Shanthikumar and Hui-Wen Koo. On uniform conditional stochastic order conditioned on planar regions. Journal of Applied Probability, 27(1):115–123, March 1990. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3214599.

Solovyev:1991:REC

[SK91] A. D. Solovyev and D. G. Konstant. Reliability estimation of a complex renewable system with an unbounded number of repair units. Journal of Applied Probability, 28(4):833–842, December 1991. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3214686.

Soltani:1998:RPS

[SK98] A. Reza Soltani and K. Khorshidian. Reward processes for semi-Markov processes: Asymptotic behaviour. Journal of Applied Probability, 35(4):833–842, December 1998. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/3215520.

Soja-Kukiela:2019:MSF

[SK19] N. Soja-Kukieła. On maxima of stationary fields. Journal of Applied Probability, 56(4):1217–1230, December 2019. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL https://www.cambridge.org/core/journals/
journal-of-applied-probability/article/on-maxima-of-
stationary-fields/8DB6F2A8D398422CF74D6FA5600CBDB4.

Shioda:2023:DCB

[SK23] Shigeo Shioda and Dai Kato. Distribution of consensus in a broadcast-based consensus algorithm with random initial opinions. Journal of Applied Probability, 60 (4):1416–1438, December 2023. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL https:/

/www.cambridge.org/core/journals/journal-of-applied-
probability/article/distribution-of-consensus-in-a-broadcastbased-
consensus-algorithm-with-random-initial-opinions/918E24C8A0D1EC261A7F7F37AA19ADAD.

Skalsky:1971:NEM

[Ska71] Michael Skalsky. A note on the estimation of the mean of a homogeneous random field. Journal of Applied Probability, 8(3): 626–629, September 1971. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3212188.

Saunders:1979:LRA

[SKF79] Roy Saunders, Richard J. Kryscio, and Gerald M. Funk. Limiting results for arrays of binary random variables on rectangular lattices under sparseness conditions. Journal of Applied Probability, 16(3):554–566, September 1979. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3213084.

Skibinsky:1967:RTM

[Ski67] Morris Skibinsky. The range of the (n + 1)th moment for distributions on [0, 1]. Journal of Applied Probability, 4(3):543–552, December 1967. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3212220.

Skibinsky:1968:ENM

[Ski68] Morris Skibinsky. Extreme nth moments for distributions on [0, 1] and the inverse of a moment space map. Journal of Applied Probability, 5(3):693–701, December 1968. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/3211931.

Skipper:2012:PAP

[Ski12] Max Skipper. A Pólya approximation to the Poisson-binomial law. Journal of Applied Probability, 49(3):745–757, September 2012. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/41713803.

Sofronov:2006:OSP

[SKK06] G. Sofronov, Jonathan M. Keith, and Dirk P. Kroese. An optimal sequential procedure for a buying-selling problem with independent observations. Journal of Applied Probability, 43(2):454– 462, June 2006. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
27595743.

Skovgaard:1987:SEC

[Sko87] Ib M. Skovgaard. Saddlepoint expansions for conditional distributions. Journal of Applied Probability, 24(4):875–887, December 1987. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3214212.

Skoulakis:2000:GSM

[Sko00] Georgios Skoulakis. A general shock model for a reliability system. Journal of Applied Probability, 37(4):925–935, December 2000. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3215484.

Saarinen:2021:ECD

[SL21] Harto Saarinen and Jukka Lempa. Ergodic control of diffusions with random intervention times. Journal of Applied Probability, 58(1):1–21, March 2021. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL https:/

/www.cambridge.org/core/journals/journal-of-applied-
probability/article/ergodic-control-of-diffusions-with-
random-intervention-times/3D9F3F2E263C50FEDB8DB94FBA5565C7.

Saunders:1979:NEN

[SLK79] Roy Saunders, Claude Lefévre, and Richard J. Kryscio. A note on the expected number of survivors in supercritical carrierborne epidemics. Journal of Applied Probability, 16(3):646–650, September 1979. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3213092.

Slud:1977:RET

[Slu77] Eric Slud. Relative extinction times for independent branching processes. Journal of Applied Probability, 14(2):382–386, June 1977. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3213009.

Sly:2007:IFW

[Sly07] Allan Sly. Integrated fractional white noise as an alternative to multifractional Brownian motion. Journal of Applied Probability, 44(2):393–408, June 2007. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/27595849.

Shi:2001:ECA

[SLZ01] Yaolin Shi, Jie Liu, and Guomin Zhang. An evaluation of Chinese annual earthquake predictions, 1990–1998. Journal of Applied Probability, 38A:222–231, 2001. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3215881.

Srinivasan:1972:SMP

[SM72] S. K. Srinivasan and K. M. Mehata. A stochastic model for polymer degradation. Journal of Applied Probability, 9(1):43– 53, March 1972. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3212635.

Saunders:1978:QGD

[SM78] I. W. Saunders and P. A. P. Moran. On the quantiles of the gamma and F distributions. Journal of Applied Probability, 15 (2):426–432, June 1978. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3213414.

Stulman:1982:NDM

[SM82] A. Stulman and A. Mehrez. A note describing the marginal benefits of spares on availability. Journal of Applied Probability, 19(4):890–893, December 1982. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.

org/stable/3213846.

Singh:1994:RAS

[SM94a] Harshinder Singh and Neeraj Misra. On redundancy allocations in systems. Journal of Applied Probability, 31(4):1004–1014, December 1994. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3215324.

Soltani:1994:DSR

[SM94b] A. Reza Soltani and R. Moeanaddin. On dispersion of stable random vectors and its application in the prediction of multivariate stable processes. Journal of Applied Probability, 31(3): 691–699, September 1994. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3215148.

Sumita:1997:TQB

[SM97] Ushio Sumita and Yasushi Masuda. Tandem queues with bulk arrivals, infinitely many servers and correlated service times. Journal of Applied Probability, 34(1):248–257, March 1997. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/3215191.

Sakuma:2005:EFB

[SM05] Yutaka Sakuma and Masakiyo Miyazawa. On the effect of finite buffer truncation in a two-node Jackson network. Journal of Applied Probability, 42(1):199–222, March 2005. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/30040780.

Smith:1972:NBC

[Smi72] C. S. Smith. A note on boundary-crossing probabilities for the Brownian motion. Journal of Applied Probability, 9(4):857–861, December 1972. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3212623.

Smith:1973:SNG

[Smi73] Woolcott Smith. Shot noise generated by a semi-Markov process. Journal of Applied Probability, 10(3):685–690, September 1973. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3212790.

Smith:1975:SPU

[Smi75] M. H. Smith. A secretary problem with uncertain employment. Journal of Applied Probability, 12(3):620–624, September 1975. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/3212880.

Smith:1982:CRF

[Smi82] Cedric A. B. Smith. Correlations and regressions in ‘family’ data. Journal of Applied Probability, 19A:45–57, 1982. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3213549.

Smith:1992:EIM

[Smi92] Richard L. Smith. The extremal index for a Markov chain. Journal of Applied Probability, 29(1):37–45, March 1992. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3214789.

Sibuya:1994:ALP

[SMS94] Masaaki Sibuya, Norihiko Miyawaki, and Ushio Sumita. Aspects of Lagrangian probability distributions. Journal of Applied Probability, 31A:185–197, 1994. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3214956.

Smythe:1976:RRT

[Smy76] R. T. Smythe. Remarks on renewal theory for percolation processes. Journal of Applied Probability, 13(2):290–300, June 1976. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/3212832.

Smythe:2011:GCC

[Smy11] R. T. Smythe. Generalized coupon collection: the superlinear case. Journal of Applied Probability, 48(1):189–199, March 2011. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/29777454.

Sabnis:1997:CSU

[SN97] S. V. Sabnis and Mini R. Nair. Coherent structures and unimodality. Journal of Applied Probability, 34(3):812–817, September 1997. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3215108.

Snell:2005:OJL

[Sne05] J. Laurie Snell. Obituary: Joseph Leonard Doob. Journal of Applied Probability, 42(1):247–256, March 2005. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/30040783.

Spencer:2011:PCM

[SO11] Simon E. F. Spencer and Philip D. O’Neill. The probability of containment for multitype branching process models for emerging epidemics. Journal of Applied Probability, 48(1):173– 188, March 2011. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
29777453.

Sobel:1967:PND

[Sob67] Eugene Sobel. Prediction of a noise-distorted, multivariate, nonstationary signal. Journal of Applied Probability, 4(2):330–342, August 1967. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3212027.

Sobel:1973:CSG

[Sob73] Matthew J. Sobel. Continuous stochastic games. Journal of Applied Probability, 10(3):597–604, September 1973. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3212779.

Sobel:1982:VDM

[Sob82] Matthew J. Sobel. The variance of discounted Markov decision processes. Journal of Applied Probability, 19(4):794–802, December 1982. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3213832.

Sobel:1990:TML

[Sob90] Matthew J. Sobel. Throughput maximization in a loss queueing system with heterogeneous servers. Journal of Applied Probability, 27(3):693–700, September 1990. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3214552.

Solo:1984:CLS

[Sol84] Victor Solo. Consistency for the least squares estimator in a transfer function model. Journal of Applied Probability, 21(1): 88–97, March 1984. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3213667.

Solo:1986:ITS

[Sol86] Victor Solo. Identifiability of time series models with errors in variables. Journal of Applied Probability, 23A:63–71, 1986. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/3214343.

Soltani:1996:RPN

[Sol96] A. Reza Soltani. Reward processes with nonlinear reward functions. Journal of Applied Probability, 33(4):1011–1017, December 1996. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3214982.

Solo:2004:AAP

[Sol04] Victor Solo. Averaging analysis of a point process adaptive algorithm. Journal of Applied Probability, 41A:361–372, 2004. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/3215989.

Sonin:1996:IRM

[Son96] Isaac M. Sonin. Increasing the reliability of a machine reduces the period of its work. Journal of Applied Probability, 33(1):217– 223, March 1996. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3215279.

Sonmez:2023:MFS

[Sön23] Ercan Sönmez. On mixed fractional stochastic differential equations with discontinuous drift coefficient. Journal of Applied Probability, 60(2):589–606, June 2023. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL https:/

/www.cambridge.org/core/journals/journal-of-applied-
probability/article/on-mixed-fractional-stochastic-differential-
equations-with-discontinuous-drift-coefficient/BEDB435F64CFA1ECAF15683F9A62331F.

Stanford:1992:SCN

[SP92] D. A. Stanford and B. Pagurek. On the serial correlation for number in system in the stationary GI/M/1 bulk arrival and GI/E_{m} /1 queues. Journal of Applied Probability, 29(2):404– 417, June 1992. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3214577.

Stadje:1999:CSM

[SP99] Wolfgang Stadje and P. R. Parthasarathy. On the convergence to stationarity of the many-server Poisson queue. Journal of Applied Probability, 36(2):546–557, June 1999. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/3215473.

Sordo:2017:SCI

[SP17] Miguel A. Sordo and Georgios Psarrakos. Stochastic comparisons of interfailure times under a relevation replacement policy.

Journal of Applied Probability, 54(1):134–145, March 2017. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL https://www.cambridge.org/core/journals/journal-
of-applied-probability/article/stochastic-comparisons-
of-interfailure-times-under-a-relevation-replacement-
policy/7042EEA6662154E0AA1F47FD805D01CE.

Simper:2022:ASM

[SP22] Mackenzie Simper and Julia A. Palacios. An adjacent-swap Markov chain on coalescent trees. Journal of Applied Probability, 59(4):1243–1260, December 2022. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL https:/

/www.cambridge.org/core/journals/journal-of-applied-
probability/article/an-adjacentswap-markov-chain-on-
coalescent-trees/8EE9C1313BEEB7FCF3D8897B64360DC9.

Spataru:1991:MSC

[Spa91] Aurel Spataru. A maximum sequence in a critical multitype branching process. Journal of Applied Probability, 28(4):893– 897, December 1991. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3214692.

Speed:1973:NRW

[Spe73a] T. P. Speed. A note on random walks. II. Journal of Applied Probability, 10(1):218–222, March 1973. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http:/

/www.jstor.org/stable/3212510.

Speed:1973:SRR

[Spe73b] T. P. Speed. Some remarks on a result of Blomqvist. Journal of Applied Probability, 10(1):229–232, March 1973. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3212512. See [Blo73].

Speed:1986:AMR

[Spe86] T. P. Speed. ANOVA models with random effects: An approach via symmetry. Journal of Applied Probability, 23A:355–368, 1986. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3214364.

Spencer:1994:GBI

[Spe94] Thomas H. Spencer. Generalized Bonferroni inequalities. Journal of Applied Probability, 31(2):409–417, June 1994. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3215034.

Spiliopoulos:2015:NPA

[Spi15] Konstantinos Spiliopoulos. Nonasymptotic performance analysis of importance sampling schemes for small noise diffusions. Journal of Applied Probability, 52(3):797–810, September 2015. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://projecteuclid.org/euclid.jap/1445543847.

Spiro:2022:APH

[Spi22] Sam Spiro. An averaging process on hypergraphs. Journal of Applied Probability, 59(2):495–504, June 2022. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL https://www.cambridge.org/core/journals/journal-
of-applied-probability/article/an-averaging-process-
on-hypergraphs/0DBB08D1BE1504BF97A16320F28B64CD.

Spouge:2001:FSC

[Spo01] John L. Spouge. Finite-size corrections to Poisson approximations of rare events in renewal processes. Journal of Applied Probability, 38(2):554–569, June 2001. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3215906.

Spouge:2004:PRI

[Spo04] John L. Spouge. Path reversal, islands, and the gapped alignment of random sequences. Journal of Applied Probability, 41(4): 975–983, December 2004. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/4141371.

Spouge:2007:MAP

[Spo07] John L. Spouge. Markov additive processes and repeats in sequences. Journal of Applied Probability, 44(2):514–527, June 2007. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/27595858.

See correction [Pak04].

Sathe:1980:IPO

[SPS80] Y. S. Sathe, Meena Pradhan, and S. P. Shah. Inequalities for the probability of the occurrence of at least m out of n events. Journal of Applied Probability, 17(4):1127–1132, December 1980. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/3213226.

Srinivasan:1970:SIB

[SR70a] S. K. Srinivasan and G. Rajamannar. Selective interaction between two independent stationary recurrent point processes. Journal of Applied Probability, 7(2):476–482, August 1970. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/3211986.

Srinivasan:1970:ADS

[SR70b] S. K. Srinivasan and A. Rangan. Age dependent stochastic models for phage reproduction. Journal of Applied Probability, 7(2): 251–261, August 1970. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3211962.

Sumita:1988:FPT

[SR88] Ushio Sumita and Maria Rieders. First passage times and lumpability of semi-Markov processes. Journal of Applied Probability, 25(4):675–687, December 1988. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.

org/stable/3214288.

Sreehari:1983:CGD

[Sre83] M. Sreehari. A characterization of the geometric distribution. Journal of Applied Probability, 20(1):209–212, March 1983. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/3213738.

Srivastava:1967:SAS

[Sri67] R. C. Srivastava. Some aspects of the stochastic model for the attachment of phages to bacteria. Journal of Applied Probability, 4(1):9–18, April 1967. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3212296.

Srivastava:1971:CPP

[Sri71] R. C. Srivastava. On a characterization of the Poisson process. Journal of Applied Probability, 8(3):615–616, September 1971. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/3212185.

Srinivasan:1974:ASF

[Sri74] S. K. Srinivasan. Analytic solution of a finite dam governed by a general input. Journal of Applied Probability, 11(1):134– 144, March 1974. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3212590.

Srivastava:1986:CGD

[Sri86] R. C. Srivastava. On characterizations of the geometric distribution by independence of functions of order statistics. Journal of Applied Probability, 23(1):227–232, March 1986. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3214133.

Srinivasan:1991:SCD

[Sri91] Rengarajan Srinivasan. Stochastic comparisons of density profiles for the road-hog process. Journal of Applied Probability, 28 (4):852–863, December 1991. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.

org/stable/3214688.

Sriram:1998:AEA

[Sri98] T. N. Sriram. Asymptotic expansions for array branching processes with applications to bootstrapping. Journal of Applied Probability, 35(1):12–26, March 1998. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3215542.

Stehr:2021:TAI

[SRN21] Mads Stehr and Anders Rønn-Nielsen. Tail asymptotics of an infinitely divisible space-time model with convolution equivalent Lévy measure. Journal of Applied Probability, 58(1):42–67, March 2021. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL https:/

/www.cambridge.org/core/journals/journal-of-applied-
probability/article/tail-asymptotics-of-an-infinitely-

divisible-spacetime-model-with-convolution-equivalent-
levy-measure/FBC766F9A6CFB801E9E4EECFD48C414A.

Samorodnitsky:2016:NRV

[SRT^{+}16] Gennady Samorodnitsky, Sidney Resnick, Don Towsley, Richard Davis, Amy Willis, and Phyllis Wan. Nonstandard regular variation of in-degree and out-degree in the preferential attachment model. Journal of Applied Probability, 53(1):146–161, March 2016. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://projecteuclid.org/euclid.jap/
1457470565.

Srivastava:1970:CPD

[SS70] R. C. Srivastava and A. B. L. Srivastava. On a characterization of Poisson distribution. Journal of Applied Probability, 7(2):497– 501, August 1970. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3211989.

Shepp:1976:FPT

[SS76] L. A. Shepp and D. Slepian. First-passage time for a particular stationary periodic Gaussian process. Journal of Applied Probability, 13(1):27–38, March 1976. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3212662.

Solomon:1980:ADG

[SS80] H. Solomon and M. A. Stephens. Approximations to densities in geometric probability. Journal of Applied Probability, 17(1):145– 153, March 1980. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3212932.

Shanthikumar:1983:GSM

[SS83a] J. G. Shanthikumar and U. Sumita. General shock models associated with correlated renewal sequences. Journal of Applied Probability, 20(3):600–614, September 1983. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/3213896.

Siegel:1983:MPS

[SS83b] Andrew F. Siegel and George Sugihara. Moments of particle size distributions under sequential breakage with applications to species abundance. Journal of Applied Probability, 20(1):158– 164, March 1983. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3213730.

Shanthikumar:1985:BPD

[SS85a] J. George Shanthikumar and Ushio Sumita. On the busy-period distributions of M/G/1/K queues with state-dependent arrivals and FCFS/LCFS-P service disciplines. Journal of Applied Probability, 22(4):912–919, December 1985. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http:/

/www.jstor.org/stable/3213958.

Shukla:1985:SAD

[SS85b] Rakesh Shukla and R. C. Srivastava. The statistical analysis of direct repeats in nucleic acid sequences. Journal of Applied Probability, 22(1):15–24, March 1985. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3213744. See note [ZS92].

Solomon:1986:SSR

[SS86] H. Solomon and C. Sutton. A simulation study of random caps on a sphere. Journal of Applied Probability, 23(4):951–960, December 1986. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3214468.

Shanthikumar:1987:COS

[SS87] J. George Shanthikumar and Ushio Sumita. Convex ordering of sojourn times in single-server queues: Extremal properties of FIFO and LIFO service disciplines. Journal of Applied Probability, 24(3):737–748, September 1987. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3214103.

Shaked:1988:MCH

[SS88a] Moshe Shaked and J. George Shanthikumar. Multivariate conditional hazard rates and the MIFRA and MIFR properties. Journal of Applied Probability, 25(1):150–168, March 1988. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3214242.

Shaked:1988:FPT

[SS88b] Moshe Shaked and J. George Shanthikumar. On the first-passage times of pure jump processes. Journal of Applied Probability, 25 (3):501–509, September 1988. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.

org/stable/3213979.

Shanthikumar:1989:MLP

[SS89] J. George Shanthikumar and Ushio Sumita. Modified Lindley process with replacement: Dynamic behavior, asymptotic decomposition and applications. Journal of Applied Probability, 26 (3):552–565, September 1989. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.

org/stable/3214413.

Shaked:1991:DMM

[SS91] Moshe Shaked and J. George Shanthikumar. Dynamic multivariate mean residual life functions. Journal of Applied Probability, 28(3):613–629, September 1991. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.

org/stable/3214496.

Shaked:1993:DCM

[SS93] Moshe Shaked and J. George Shanthikumar. Dynamic conditional marginal distributions in reliability theory. Journal of Applied Probability, 30(2):421–428, June 1993. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/3214850.

Simonot:1996:CCR

[SS96] F. Simonot and Y. Q. Song. Characterization of convergence rates for the approximation of the stationary distribution of infinite monotone stochastic matrices. Journal of Applied Probability, 33(4):974–985, December 1996. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3214979.

Singh:1997:ASC

[SS97] Harshinder Singh and R. S. Singh. On allocation of spares at component level versus system level. Journal of Applied Probability, 34(1):283–287, March 1997. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3215194.

Scarsini:1999:STP

[SS99] Marco Scarsini and Fabio Spizzichino. Simpson-type paradoxes, dependence, and ageing. Journal of Applied Probability, 36(1): 119–131, March 1999. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3215407.

Siegrist:2001:SG

[SS01] Kyle Siegrist and John Steele. Sequential games. Journal of Applied Probability, 38(4):1006–1017, December 2001. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3215780.

Szabados:2003:EFR

[SS03] Tamás Szabados and Balázs Székely. An exponential functional of random walks. Journal of Applied Probability, 40(2):413– 426, June 2003. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3215800.

Schied:2007:RDH

[SS07] Alexander Schied and Mitja Stadje. Robustness of delta hedging for path-dependent options in local volatility models. Journal of Applied Probability, 44(4):865–879, December 2007. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/27595893.

Sabourin:2017:MSU

[SS17] Anne Sabourin and Johan Segers. Marginal standardization of upper semicontinuous processes. with application to max-stable processes. Journal of Applied Probability, 54 (3):773–796, September 2017. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL https:/

/www.cambridge.org/core/journals/journal-of-applied-
probability/article/marginal-standardization-of-upper-
semicontinuous-processes-with-application-to-maxstable-
processes/7DD68D7235CE93DAC300CD19DBF894C3.

Steele:1987:NLE

[SSE87] J. Michael Steele, Lawrence A. Shepp, and William F. Eddy. On the number of leaves of a Euclidean minimal spanning tree. Journal of Applied Probability, 24(4):809–826, December 1987.

CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/3214207.

Shaked:2010:CLI

[SSSL10] Moshe Shaked, Miguel A. Sordo, and Alfonso Suárez-Llorens. A class of location-independent variability orders, with applications. Journal of Applied Probability, 47(2):407–425, June 2010. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/20778488.

Santacroce:2018:EMO

[SST18] Marina Santacroce, Paola Siri, and Barbara Trivellato. Exponential models by Orlicz spaces and applications. Journal of Applied Probability, 55(3):682–700, September 2018. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL https://www.cambridge.org/core/journals/journal-
of-applied-probability/article/exponential-models-by-
orlicz-spaces-and-applications/3F22EB33454B4BF74404825B6D767A71.

Srinivasan:1972:CFQ

[SSV72] S. K. Srinivasan, R. Subramanian, and R. Vasudevan. Correlation functions in queueing theory. Journal of Applied Probability, 9(3):604–616, September 1972. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3212329.

Shanbhag:1973:EFS

[ST73] D. N. Shanbhag and D. G. Tambouratzis. Erlang’s formula and some results on the departure process for a loss system. Journal of Applied Probability, 10(1):233–240, March 1973. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3212513.

Seneta:1983:NMU

[ST83] E. Seneta and S. Tavaré. A note on models using the branching process with immigration stopped at zero. Journal of Applied Probability, 20(1):11–18, March 1983. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3213716.

Seneta:1985:MSQ

[ST85] E. Seneta and R. L. Tweedie. Moments for stationary and quasistationary distributions of Markov chains. Journal of Applied Probability, 22(1):148–155, March 1985. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http:/

/www.jstor.org/stable/3213754.

Scalia-Tomba:1986:AFS

[ST86a] Gianpaolo Scalia-Tomba. Asymptotic final size distribution of the multitype Reed-frost process. Journal of Applied Probability, 23(3):563–584, September 1986. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.

org/stable/3213998.

Steele:1986:BDD

[ST86b] J. Michael Steele and Luke Tierney. Boundary domination and the distribution of the largest nearest-neighbor link in higher dimensions. Journal of Applied Probability, 23(2):524–528, June 1986. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3214195.

Smith:1988:EIP

[ST88] Cedric A. B. Smith and Robert Thomson. Estimation of inbreeding from population samples. Journal of Applied Probability, 25A:127–135, 1988. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3214151.

Sansom:2001:FHS

[ST01] John Sansom and Peter Thomson. Fitting hidden semi-Markov models to breakpoint rainfall data. Journal of Applied Probability, 38A:142–157, 2001. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3215875.

Shieh:2002:MSB

[ST02] Narn-Rueih Shieh and S. James Taylor. Multifractal spectra of branching measure on a Galton–Watson tree. Journal of Applied Probability, 39(1):100–111, March 2002. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http:/

/www.jstor.org/stable/3215921.

Shah:2008:PQ

[ST08] Devavrat Shah and John N. Tsitsiklis. Bin packing with queues. Journal of Applied Probability, 45(4):922–939, December 2008.

CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/27595999.

Sheu:2012:OSP

[ST12] Yuan-Chung Sheu and Ming-Yao Tsai. On optimal stopping problems for matrix-exponential jump-diffusion processes. Journal of Applied Probability, 49(2):531–548, June 2012. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/41713784.

Schluter:2016:WCS

[ST16] Christian Schluter and Mark Trede. Weak convergence to the Student and Laplace distributions. Journal of Applied Probability, 53(1):121–129, March 2016. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http:/

/projecteuclid.org/euclid.jap/1457470563.

Starr:1974:OAS

[Sta74] Norman Starr. Optimal and adaptive stopping based on capture times. Journal of Applied Probability, 11(2):294–301, June 1974. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3212750.

Stam:1977:RLR

[Sta77] A. J. Stam. The reversed ladder of a random walk. Journal of Applied Probability, 14(1):190–194, March 1977. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3213272.

Stam:1981:NRL

[Sta81] A. J. Stam. A non-random limit for the volume covered k times. Journal of Applied Probability, 18(1):148–156, March 1981. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/3213175.

Stam:1982:LTU

[Sta82] A. J. Stam. Limit theorems for uniform distributions on spheres in high-dimensional Euclidean spaces. Journal of Applied Probability, 19(1):221–228, March 1982. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3213932.

Stadje:1985:BPQ

[Sta85] Wolfgang Stadje. The busy period of the queueing system M/G/∞. Journal of Applied Probability, 22(3):697–704, September 1985. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3213872.

Stadje:1987:A

[Sta87a] W. Stadje. Acknowledgement. Journal of Applied Probability, 24(1):290, March 1987. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3214084.

Stadje:1987:ABS

[Sta87b] Wolfgang Stadje. Asymptotic behaviour of a stopping time related to cumulative sum procedures and single-server queues. Journal of Applied Probability, 24(1):200–214, March 1987. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/3214071.

Stadje:1991:NCT

[Sta91] Wolfgang Stadje. A new continuous-time search model. Journal of Applied Probability, 28(4):771–778, December 1991. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3214680.

Stadje:1993:NLM

[Sta93] W. Stadje. A new look at the Moran dam. Journal of Applied Probability, 30(2):489–495, June 1993. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3214860.

Stadje:1999:SSP

[Sta99] Wolfgang Stadje. Stationarity of a stochastic population flow model. Journal of Applied Probability, 36(1):291–294, March 1999. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3215424.

Startsev:2001:AAG

[Sta01] A. N. Startsev. Asymptotic analysis of the general stochastic epidemic with variable infectious periods. Journal of Applied Probability, 38(1):18–35, March 2001. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3215738.

Stanke:2004:SSR

[Sta04] Mario Stanke. Sequential selection of random vectors under a sum constraint. Journal of Applied Probability, 41(1):131– 146, March 2004. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3215820.

Stadje:2020:TRD

[Sta20] Mitja Stadje. Two results on dynamic extensions of deviation measures. Journal of Applied Probability, 57(2):531–540, June 2020. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-

6072 (electronic). URL https://www.cambridge.org/core/
journals/journal-of-applied-probability/article/two-
results-on-dynamic-extensions-of-deviation-measures/
6709A5B3A8817A75CB30A6270F91BCEF.

Sparaggis:1993:EPS

[STC93] P. D. Sparaggis, D. Towsley, and C. G. Cassandras. Extremal properties of the shortest/longest non-full queue policies in finite-capacity systems with state-dependent service rates. Journal of Applied Probability, 30(1):223–236, March 1993. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3214634.

Steinebach:1978:SLE

[Ste78] Josef Steinebach. A strong law of Erdős–Rényi type for cumulative processes in renewal theory. Journal of Applied Probability, 15(1):96–111, March 1978. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3213240.

Steinebach:1981:SGP

[Ste81] Josef Steinebach. The stochastic geyser problem for first-passage times. Journal of Applied Probability, 18(1):91–103, March 1981. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/3213169.

Stefanov:1985:MSF

[Ste85] Valeri T. Stefanov. On the moments of some first-passage times. Journal of Applied Probability, 22(2):461–466, June 1985. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/3213790.

Steutel:1986:AEG

[Ste86] F. W. Steutel. Avalanches of electrons in a gas. Journal of Applied Probability, 23(4):867–879, December 1986. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3214461.

Stefanov:1995:MPT

[Ste95] Valeri T. Stefanov. Mean passage times for tridiagonal transition matrices. Journal of Applied Probability, 32(3):846–849, September 1995. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3215137.

Stefanov:1999:OCE

[Ste99] Valeri T. Stefanov. On the occurrence of composite events and clusters of points. Journal of Applied Probability, 36(4):1012– 1018, December 1999. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3215574.

Stefanov:2000:SWT

[Ste00] Valeri T. Stefanov. On some waiting time problems. Journal of Applied Probability, 37(3):756–764, September 2000. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3215611.

Stefanov:2003:IDB

[Ste03] Valeri T. Stefanov. The intersite distances between pattern occurrences in strings generated by general discrete- and continuous-time models: An algorithmic approach. Journal of Applied Probability, 40(4):881–892, December 2003. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3216048.

Stefanov:2004:DAG

[Ste04] Valeri T. Stefanov. Distribution of the amount of genetic material from a chromosomal segment surviving to the following generation. Journal of Applied Probability, 41(2):345–354, June 2004. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3216020.

See correction [Ano04g].

Stefanov:2006:EDR

[Ste06] Valeri T. Stefanov. Exact distributions for reward functions on semi-Markov and Markov additive processes. Journal of Applied Probability, 43(4):1053–1065, December 2006. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/27595799.

Stegehuis:2019:DCS

[Ste19]  Clara Stegehuis.  Degree correlations in scale-free ran-

dom graph models. Journal of Applied Probability, 56 (3):672–700, September 2019. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL https:/

/www.cambridge.org/core/journals/journal-of-applied-
probability/article/degree-correlations-in-scalefree-
random-graph-models/95BF1562C63CDA63020315F923A21D56.

Stidham:1994:SPA

[Sti94] Shaler Stidham, Jr. A sample-path approach to Palm probabilities. Journal of Applied Probability, 31(2):430–437, June 1994. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/3215036.

Stormer:1974:DAO

[Stö74] Horand Störmer. On the distribution of attributes in organizations. Journal of Applied Probability, 11(3):529–543, September 1974. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3212697.

Stockbridge:1991:MAS

[Sto91] Richard H. Stockbridge. A martingale approach to the slow server problem. Journal of Applied Probability, 28(2):480– 486, June 1991. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3214883.

Stoyan:1999:OKM

[Sto99] D. Stoyan. Obituary: Klaus Matthes. Journal of Applied Probability, 36(4):1255–1257, December 1999. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/3215596.

Stoyanov:2004:SCM

[Sto04] Jordan Stoyanov. Stieltjes classes for moment-indeterminate probability distributions. Journal of Applied Probability, 41A: 281–294, 2004. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3215983.

Stolyar:2008:DDS

[Sto08] Alexander L. Stolyar. Dynamic distributed scheduling in random access networks. Journal of Applied Probability, 45(2):297– 313, June 2008. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
27595946.

Strauss:1975:ABL

[Str75a] David J. Strauss. Analysing binary lattice data with the nearestneighbor property. Journal of Applied Probability, 12(4):702– 712, December 1975. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3212721.

Streit:1975:RIR

[Str75b] Franz Streit. Results on the intersection of randomly located sets. Journal of Applied Probability, 12(4):817–823, December 1975. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3212733.

Strauss:1977:CCL

[Str77] David J. Strauss. Clustering on coloured lattices. Journal of Applied Probability, 14(1):135–143, March 1977. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3213266.

Sigman:1994:NER

[STW94] Karl Sigman, Hermann Thorisson, and Ronald W. Wolff. A note on the existence of regeneration times. Journal of Applied Probability, 31(4):1116–1122, December 1994. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/3215335.

Subelman:1979:DBS

[Sub79] Eduardo J. Subelman. The dependence of betting strategies on the probability of winning. Journal of Applied Probability, 16(4):

855–866, December 1979. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3213151.

Subelman:1981:HSG

[Sub81] Eduardo J. Subelman. A hide-search game. Journal of Applied Probability, 18(3):628–640, September 1981. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/3213317.

Sudbury:1976:SRO

[Sud76] Aidan Sudbury. The size of the region occupied by one type in an invasion process. Journal of Applied Probability, 13(2):355– 356, June 1976. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3212839.

Sudbury:1977:CEM

[Sud77] Aidan Sudbury. Clumping effects in models of isolation by distance. Journal of Applied Probability, 14(2):391–395, June 1977. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/3213011.

Sudbury:1981:EPS

[Sud81] Aidan Sudbury. The expected population size in a cell-size-dependent branching process. Journal of Applied Probability, 18 (1):65–75, March 1981. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3213167.

Sudbury:1983:TTP

[Sud83] Aidan Sudbury. The time taken for a population to grow from size m to size km. Journal of Applied Probability, 20(4):728–736, December 1983. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3213585.

Sudbury:1985:PPN

[Sud85] Aidan Sudbury. The proportion of the population never hearing a rumour. Journal of Applied Probability, 22(2):443–446, June 1985. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3213787.

Sudbury:2001:RLB

[Sud01] Aidan Sudbury. Rigorous lower bounds for the critical infection rate in the diffusive contact process. Journal of Applied Probability, 38(4):1074–1078, December 2001. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/3215786.

Sudbury:2002:IEM

[Sud02] Aidan Sudbury. Inclusion-exclusion methods for treating annihilating and deposition processes. Journal of Applied Probability, 39(3):466–478, September 2002. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3216071.

Sudbury:2004:APR

[Sud04] Aidan Sudbury. The annihilating process on random trees and the square lattice. Journal of Applied Probability, 41(3):816–831, September 2004. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
4141356.

Sudbury:2005:RLB

[Sud05] Aidan Sudbury. Rigorous lower bounds for extinction probabilities of the contact process. Journal of Applied Probability, 42(3): 902–904, September 2005. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/30040868.

Sudbury:2012:MVS

[Sud12] Aidan Sudbury. Means and variances in stochastic multistage cancer models. Journal of Applied Probability, 49(2):590– 594, June 2012. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
41713788.

Sumita:1984:CPT

[Sum84] Ushio Sumita. On conditional passage time structure of birthdeath processes. Journal of Applied Probability, 21(1):10–21, March 1984. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3213660.

Sumita:1987:LBO

[Sum87] U. Sumita. On limiting behavior of ordinary and conditional first-passage times for a class of birth-death processes. Journal of Applied Probability, 24(1):235–240, March 1987. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3214074.

Sun:1992:NVI

[Sun92] M. Sun. Nested variational inequalities and related optimal starting-stopping problems. Journal of Applied Probability, 29 (1):104–115, March 1992. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3214795.

Suomela:1979:IMT

[Suo79] P. Suomela. Invariant measures of time-reversible Markov chains. Journal of Applied Probability, 16(1):226–229, March 1979. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3213392.

Surya:2008:ESF

[Sur08] B. A. Surya. Evaluating scale functions of spectrally negative Lévy processes. Journal of Applied Probability, 45(1):135– 149, March 2008. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
27595929.

Sutton:1989:PDR

[Sut89] Clifton Sutton. Packing densities of randomly constructed codes. Journal of Applied Probability, 26(3):512–523, September 1989. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/3214409.

Singull:2021:LDA

[SUY21] Martin Singull, Denise Uwamariya, and Xiangfeng Yang. Largedeviation asymptotics of condition numbers of random matrices. Journal of Applied Probability, 58(4):1114–1130, December 2021. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL https://www.cambridge.org/core/journals/
journal-of-applied-probability/article/largedeviation-
asymptotics-of-condition-numbers-of-random-matrices/
CEA792BD504A46E75B3BEC9D29CF2C55.

Suzuki:2010:MCA

[Suz10] Joe Suzuki. A Markov chain analysis of genetic algorithms: Large deviation principle approach. Journal of Applied Probability, 47(4):967–975, December 2010. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/29777419.

Staff:1971:SDO

[SV71] P. J. Staff and M. K. Vagholkar. Stationary distributions of open Markov processes in discrete time with application to hospital planning. Journal of Applied Probability, 8(4):668–680, December 1971. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3212231.

Schrijner:1997:WCC

[SV97] Pauline Schrijner and Erik A. Van Doorn. Weak convergence of conditioned birth-death processes in discrete time. Journal of Applied Probability, 34(1):46–53, March 1997. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3215173.

Solan:2003:PMC

[SV03a] Eilon Solan and Nicolas Vieille. Perturbed Markov chains. Journal of Applied Probability, 40(1):107–122, March 2003. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3215844.

Suyono:2003:MCT

[SV03b] Suyono and J. A. M. Van Der Weide. A method for computing total downtime distributions in repairable systems. Journal of Applied Probability, 40(3):643–653, September 2003. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3215941.

Shcherbakov:2019:BEC

[SV19] Vadim Shcherbakov and Stanislav Volkov. Boundary effect in competition processes. Journal of Applied Probability, 56(3):750–768, September 2019. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL https:/

/www.cambridge.org/core/journals/journal-of-applied-
probability/article/boundary-effect-in-competition-processes/
3B174BF41D4B6C7F68EBC23050638FC7.

Svensson:1995:ASD

[Sve95] Åke Svensson. On the asymptotic size and duration of a class of epidemic models. Journal of Applied Probability, 32(1):11– 24, March 1995. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3214917.

Seneta:1966:QSD

[SVJ66] E. Seneta and D. Vere-Jones. On quasi-stationary distributions in discrete-time Markov chains with a denumerable infinity of states. Journal of Applied Probability, 3(2):403–434, December 1966. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3212128.

Smith:1967:DT

[SW67] W. Smith and G. S. Watson. Diffusion out of a triangle. Journal of Applied Probability, 4(3):479–488, December 1967. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3212215.

Smith:1987:LDT

[SW87] Richard L. Smith and Ishay Weissman. Large deviations of tail estimators based on the Pareto approximation. Journal of Applied Probability, 24(3):619–630, September 1987. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3214094.

Szynal:1988:SSI

[SW88] D. Szynal and S. Wedrychowicz. On solutions of a stochastic integral equation of the Volterra type with applications for chemotherapy. Journal of Applied Probability, 25(2):257– 267, June 1988. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3214434.

Surgailis:1994:BEN

[SW94] Donatas Surgailis and Wojbor A. Woyczynski. Burgers’ equation with non-local shot noise data. Journal of Applied Probability, 31A:351–362, 1994. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3214968.

Shaked:1997:SCR

[SW97a] Moshe Shaked and Tityik Wong. Stochastic comparisons of random minima and maxima. Journal of Applied Probability, 34 (2):420–425, June 1997. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3215381.

Shaked:1997:SOB

[SW97b] Moshe Shaked and Tityik Wong. Stochastic orders based on ratios of Laplace transforms. Journal of Applied Probability, 34 (2):404–419, June 1997. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3215380.

Schiefermayr:2005:CSO

[SW05] Klaus Schiefermayr and Josef Weichbold. A complete solution for the optimal stochastic scheduling of a two-stage tandem queue with two flexible servers. Journal of Applied Probability, 42(3):778–796, September 2005. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/30040856.

Sigman:2011:HTL

[SW11] Karl Sigman and Ward Whitt. Heavy-traffic limits for nearly deterministic queues. Journal of Applied Probability, 48(3):657– 678, September 2011. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/23065927.

Song:2020:FPT

[SW20] Shiyu Song and Yongjin Wang. On first passage times of sticky reflecting diffusion processes with double exponential jumps. Journal of Applied Probability, 57(1):221–236, March 2020. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL https://www.cambridge.org/core/journals/journal-
of-applied-probability/article/on-first-passage-times-
of-sticky-reflecting-diffusion-processes-with-double-
exponential-jumps/52EAF720EE25DFB04A69D9919B8D5CB0.

Swensen:1988:NRO

[Swe88] Anders Rygh Swensen. A note on a randomized occupancy problem. Journal of Applied Probability, 25(4):725–732, December 1988. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3214293.

Sweet:1994:OSU

[Swe94] Ted Sweet. Optimizing a single uncertain selection with recall of observations. Journal of Applied Probability, 31(3):660–672, September 1994. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3215146.

Switzer:1976:GMS

[Swi76] Paul Switzer. Geometrical measures of the smoothness of random functions. Journal of Applied Probability, 13(1):86–95, March 1976. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3212668.

Swift:2008:SMC

[Swi08] Andrew W. Swift. Stochastic models of cascading failures. Journal of Applied Probability, 45(4):907–921, December 2008. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/27595998.

Sass:2021:DAR

[SWW21] Jörn Sass, Dorothee Westphal, and Ralf Wunderlich. Diffusion approximations for randomly arriving expert opinions in a financial market with Gaussian drift. Journal of Applied Probability, 58(1):197–216, March 2021. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL https:/

/www.cambridge.org/core/journals/journal-of-applied-
probability/article/diffusion-approximations-for-randomly-
arriving-expert-opinions-in-a-financial-market-with-
gaussian-drift/F500CD4068E28CDA72B611960AC15130.

Sloothaak:2018:FPT

[SWZ18] Fiona Sloothaak, Vitali Wachtel, and Bert Zwart. First-passage time asymptotics over moving boundaries for random walk bridges. Journal of Applied Probability, 55(2):627–651, June 2018. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-

6072 (electronic). URL https://www.cambridge.org/core/
journals/journal-of-applied-probability/article/firstpassage-
time-asymptotics-over-moving-boundaries-for-random-walk-
bridges/0DCBC569D072145418AD2E49E6490247.

Sang:2018:EML

[SX18] Hailin Sang and Yimin Xiao. Exact moderate and large deviations for linear random fields. Journal of Applied Probability, 55(2):431–449, June 2018. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL https:/

/www.cambridge.org/core/journals/journal-of-applied-
probability/article/exact-moderate-and-large-deviations-
for-linear-random-fields/234E7E50D39A443D93FF690994168EFA.

Santalo:1972:APF

[SY72] L. A. Santaló and I. Yañez. Averages for polygons formed by random lines in Euclidean and hyperbolic planes. Journal of Applied Probability, 9(1):140–157, March 1972. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/3212643.

Shanthikumar:1986:EIS

[SY86] J. George Shanthikumar and David D. Yao. The effect of increasing service rates in a closed queueing network. Journal of Applied Probability, 23(2):474–483, June 1986. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/3214188.

Shanthikumar:1989:SMG

[SY89] J. George Shanthikumar and David D. Yao. Stochastic monotonicity in general queueing networks. Journal of Applied Probability, 26(2):413–417, June 1989. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3214048.

Shanthikumar:1991:SSC

[SY91] J. George Shanthikumar and David D. Yao. Strong stochastic convexity: Closure properties and applications. Journal of Applied Probability, 28(1):131–145, March 1991. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3214746.

Syski:1994:SIP

[Sys94] Ryszard Syski. A simple increasing process. Journal of Applied Probability, 31A:363–371, 1994. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.

org/stable/3214969.

Stadje:1991:OMS

[SZ91] Wolfgang Stadje and Dror Zuckerman. Optimal maintenance strategies for repairable systems with general degree of repair. Journal of Applied Probability, 28(2):384–396, June 1991. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/3214874.

Shaked:1992:SRB

[SZ92] Moshe Shaked and Haolong Zhu. Some results on block replacement policies and renewal theory. Journal of Applied Probability, 29(4):932–946, December 1992. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3214725.

Steinebach:1993:NSL

[SZ93] Josef Steinebach and Hanqin Zhang. Note on the strong limiting behaviour of busy periods in GI/G/1 queues under heavy traffic. Journal of Applied Probability, 30(2):483–488, June 1993. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/3214859.

Stadje:2004:TPR

[SZ04] W. Stadje and S. Zacks. Telegraph processes with random velocities. Journal of Applied Probability, 41(3):665–678, September 2004. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/4141345.

Spicer:2006:UOS

[SZ06] Scott Spicer and Ilze Ziedins. User-optimal state-dependent routeing in parallel tandem queues with loss. Journal of Applied Probability, 43(1):274–281, March 2006. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http:/

/www.jstor.org/stable/27595726.

Szasz:1974:LTS

[Szá74] Domokos Szász. A limit theorem for semi-Markov processes. Journal of Applied Probability, 11(3):521–528, September 1974. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/3212696.

Szczotka:1997:ASQ

[Szc97] Władysław Szczotka. Asymptotic stationarity of queueing processes. Journal of Applied Probability, 34(4):1041–1048, December 1997. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3215016.

Szekli:1986:CWT

[Sze86] R. Szekli. On the concavity of the waiting-time distribution in some GI/G/I queues. Journal of Applied Probability, 23(2):555– 561, June 1986. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3214200.

Szpankowski:1985:SSC

[Szp85] Wojciech Szpankowski. Some sufficient conditions for nonergodicity of Markov chains. Journal of Applied Probability, 22 (1):138–147, March 1985. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3213753.

Sirl:2007:CBD

[SZP07] David Sirl, Hanjun Zhang, and Phil Pollett. Computable bounds for the decay parameter of a birth-death process. Journal of Applied Probability, 44(2):476–491, June 2007. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/27595855.

Siu:2019:MAA

[SZY19a] Tak Kuen Siu, Jinxia Zhu, and Hailiang Yang. A martingale approach for asset allocation with derivative security and hidden economic risk. Journal of Applied Probability, 56(3):723–749, September 2019. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL https:/

/www.cambridge.org/core/journals/journal-of-applied-
probability/article/martingale-approach-for-asset-allocation-
with-derivative-security-and-hidden-economic-risk/52F547A677FD61EB3410F2E2A5262D63.

Szymkowiak:2019:MAT

[Szy19b] Magdalena Szymkowiak. Measures of ageing tendency. Journal of Applied Probability, 56(2):358–383, June 2019. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL https://www.cambridge.org/core/journals/journal-
of-applied-probability/article/measures-of-ageing-tendency/
4B1E4B59C8D584FD4584BFCFA7955E85.

Taib:1995:BPF

[Tai95] Ziad Taib. Branching processes and functional differential equations determining steady-size distributions in cell populations. Journal of Applied Probability, 32(1):1–10, March 1995. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/3214916.

Taibi:1996:GDM

[Taï96] Djaouad Taïbi. Une généralisation du modèle de diffusion de Bernoulli–Laplace. (French) [A generalization of the Bernoulli– Laplace diffusion model]. Journal of Applied Probability, 33(3): 688–697, September 1996. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3215350.

Takacs:1964:CMT

[Tak64a] Lajos Takács. Combinatorial methods in the theory of dams. Journal of Applied Probability, 1(1):69–76, June 1964. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3212061.

Takacs:1964:FRS

[Tak64b] Lajos Takács. Fluctuations in the ratio of scores in counting a ballot. Journal of Applied Probability, 1(2):393–396, December 1964. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3211869.

Takacs:1964:UBT

[Tak64c] Lajos Takács. The use of a ballot theorem in order statistics. Journal of Applied Probability, 1(2):389–392, December 1964. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/3211868.

Takacs:1967:DCF

[Tak67] Lajos Takács. The distribution of the content of finite dams. Journal of Applied Probability, 4(1):151–161, April 1967. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/3212308.

Takacs:1970:FER

[Tak70] Lajos Takács. On the fluctuations of election returns. Journal of Applied Probability, 7(1):114–123, April 1970. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3212153.

Takacs:1971:DQO

[Tak71a] Lajos Takács. Discrete queues with one server. Journal of Applied Probability, 8(4):691–707, December 1971. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3212233.

Takacs:1971:CTE

[Tak71b] Lajos Takács. On the comparison of a theoretical and an empirical distribution function. Journal of Applied Probability, 8 (2):321–330, June 1971. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3211902.

Takacs:1974:SSQ

[Tak74] Lajos Takács. A single-server queue with limited virtual waiting time. Journal of Applied Probability, 11(3):612–617, September 1974. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3212710.

Takacs:1976:SRC

[Tak76] Lajos Takács. Some remarks on a counter process. Journal of Applied Probability, 13(3):623–627, September 1976. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3212486.

Takacs:1977:OPS

[Tak77] Lajos Takács. On the ordered partial sums of real random variables. Journal of Applied Probability, 14(1):75–88, March 1977. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/3213261.

Takacs:1988:LDN

[Tak88] Lajos Takács. On the limit distribution of the number of cycles in a random graph. Journal of Applied Probability, 25A:359–376, 1988. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3214169.

Takacs:1989:BQR

[Tak89] Lajos Takács. Ballots, queues and random graphs. Journal of Applied Probability, 26(1):103–112, March 1989. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3214320.

Takacs:1992:THR

[Tak92a] Lajos Takacs. On the total heights of random rooted trees. Journal of Applied Probability, 29(3):543–556, September 1992. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/3214892.

Takagi:1992:TDP

[Tak92b] Hideaki Takagi. Time-dependent process of M/G/1 vacation models with exhaustive service. Journal of Applied Probability, 29(2):418–429, June 1992. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3214578.

Takacs:1995:LDB

[Tak95] Lajos Takács. Limit distributions for the Bernoulli meander. Journal of Applied Probability, 32(2):375–395, June 1995. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/3215294.

Tallis:1969:IMD

[Tal69] G. M. Tallis. The identifiability of mixtures of distributions. Journal of Applied Probability, 6(2):389–398, August 1969. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/3212008.

Talwalker:1980:NGR

[Tal80] Sheela Talwalker. A note on the generalized Rao–Rubin condition and characterization of certain discrete distributions. Journal of Applied Probability, 17(2):563–569, June 1980. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3213047.

Tambouratzis:1968:PVW

[Tam68] D. G. Tambouratzis. On a property of the variance of the waiting time of a queue. Journal of Applied Probability, 5(3):702–703, December 1968. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3211932.

Tambouratzis:1970:NBL

[Tam70] D. G. Tambouratzis. A note on Belyaev’s limiting distribution of the intervals between losses in an n-server system. Journal of Applied Probability, 7(2):457–464, August 1970. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3211982.

Tambouratzis:1971:SSM

[Tam71] D. G. Tambouratzis. Study of a special M/M/1 queue. Journal of Applied Probability, 8(3):630–631, September 1971. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/3212189. See correction

[Ano72g].

Tamaki:1979:RBM

[Tam79] M. Tamaki. Recognizing both the maximum and the second maximum of a sequence. Journal of Applied Probability, 16(4): 803–812, December 1979. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3213146.

Tamaki:1982:OPP

[Tam82] Mitsushi Tamaki. An optimal parking problem. Journal of Applied Probability, 19(4):803–814, December 1982. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3213833.

Tamaki:1985:AAS

[Tam85] Mitsushi Tamaki. Adaptive approach to some stopping problems. Journal of Applied Probability, 22(3):644–652, September 1985. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3213867.

Tamaki:1986:FIB

[Tam86] Mitsushi Tamaki. A full-information best-choice problem with finite memory. Journal of Applied Probability, 23(3):718–735, September 1986. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3214010.

Tamaki:1988:OSP

[Tam88] Mitsushi Tamaki. Optimal stopping in the parking problem with U-turn. Journal of Applied Probability, 25(2):363–374, June 1988. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3214444.

Tamaki:2001:OST

[Tam01] Mitsushi Tamaki. Optimal stopping on trajectories and the ballot problem. Journal of Applied Probability, 38(4):946–959, December 2001. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3215776.

Tamaki:2009:OCB

[Tam09] Mitsushi Tamaki. Optimal choice of the best available applicant in full-information models. Journal of Applied Probability, 46(4): 1086–1099, December 2009. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/25662482.

Tamaki:2010:SMO

[Tam10] Mitsushi Tamaki. Sum the multiplicative odds to one and stop. Journal of Applied Probability, 47(3):761–777, September 2010. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/20778515.

Tamaki:2015:OSP

[Tam15] Mitsushi Tamaki. On the optimal stopping problems with monotone thresholds. Journal of Applied Probability, 52(4): 926–940, December 2015. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://projecteuclid.

org/euclid.jap/1450802744.

Tango:1978:EBR

[Tan78] Toshiro Tango. Extended block replacement policy with used items. Journal of Applied Probability, 15(3):560–572, September 1978. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3213119.

Tan:1979:AMB

[Tan79a] Harry H. Tan. Another martingale bound on the waiting-time distribution in GI/G/1 queues. Journal of Applied Probability, 16(2):454–457, June 1979. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3212916.

Tanemura:1979:PCB

[Tan79b] M. Tanemura. Has the Palásti conjecture been proved?: A criticism of a paper by H. J. Weiner. Journal of Applied Probability, 16(3):697–698, September 1979. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor. org/stable/3213103. See [Wei78] and reply [Wei79].

Taniguchi:1979:EPG

[Tan79c] Masanobu Taniguchi. On estimation of parameters of Gaussian stationary processes. Journal of Applied Probability, 16(3): 575–591, September 1979. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3213086.

Tanemura:1980:SCL

[Tan80a] M. Tanemura. Some comments on the letters by H. J. Weiner. Journal of Applied Probability, 17(3):884–887, September 1980. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/3212987. See

[Wei79].

Taniguchi:1980:EIC

[Tan80b] Masanobu Taniguchi. On estimation of the integrals of certain functions of spectral density. Journal of Applied Probability, 17 (1):73–83, March 1980. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3212925.

Tan:1982:FFP

[Tan82] Choon-Peng Tan. A functional form for a particular coefficient of ergodicity. Journal of Applied Probability, 19(4):858–863, December 1982. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3213840.

Tan:1983:CER

[Tan83a] Choon-Peng Tan. Coefficients of ergodicity with respect to vector norms. Journal of Applied Probability, 20(2):277–287, June 1983. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3213801.

Tanner:1983:PFR

[Tan83b] J. C. Tanner. Polygons formed by random lines in a plane: Some further results. Journal of Applied Probability, 20(4):778–787, December 1983. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3213589.

Tanner:1983:PQF

[Tan83c] J. C. Tanner. The proportion of quadrilaterals formed by random lines in a plane. Journal of Applied Probability, 20(2):400– 404, June 1983. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3213813.

Tanaka:1986:AET

[Tan86] Katsuto Tanaka. Asymptotic expansions for time series statistics. Journal of Applied Probability, 23A:211–227, 1986. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3214354.

Tanemura:1993:BSP

[Tan93] Hideki Tanemura. Behavior of the supercritical phase of a continuum percolation model on R^{d}. Journal of Applied Probability, 30(2):382–396, June 1993. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3214847.

Tang:1994:DPM

[Tan94] Yinghui Tang. The departure process of the M/G/1 queueing model with server vacation and exhaustive service discipline. Journal of Applied Probability, 31(4):1070–1082, December 1994. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/3215330.

Tanikawa:1997:RCB

[Tan97] Akio Tanikawa. On the rate of convergence of Borovkov’s multidimensional ergodic Markov chain. Journal of Applied Probability, 34(2):328–339, June 1997. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3215373.

Tang:2005:FTR

[Tan05] Qihe Tang. The finite-time ruin probability of the compound Poisson model with constant interest force. Journal of Applied Probability, 42(3):608–619, September 2005. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/30040844.

Tang:2007:HTD

[Tan07] Qihe Tang. Heavy tails of discounted aggregate claims in the continuous-time renewal model. Journal of Applied Probability, 44(2):285–294, June 2007. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/27595841.

Tankov:2011:IFB

[Tan11] Peter Tankov. Improved Fréchet bounds and model-free pricing of multi-asset options. Journal of Applied Probability, 48(2):389– 403, June 2011. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/

23065892. See note [BJV12].

Tan:2017:LLE

[Tan17] Zhongquan Tan. Limit laws on extremes of nonhomogeneous Gaussian random fields. Journal of Applied Probability, 54(3):811–832, September 2017. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL https:/

/www.cambridge.org/core/journals/journal-of-applied-
probability/article/limit-laws-on-extremes-of-nonhomogeneous-
gaussian-random-fields/D5AB6956695FED31C6FBCCEC8A0A777F.

Tardelli:2015:PII

[Tar15] P. Tardelli. Partially informed investors: hedging in an incomplete market with default. Journal of Applied Probability, 52(3): 718–735, September 2015. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://projecteuclid.

org/euclid.jap/1445543842.

Tasche:1997:SBC

[Tas97] Dirk Tasche. On the second Borel–Cantelli lemma for strongly mixing sequences of events. Journal of Applied Probability, 34 (2):381–394, June 1997. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3215378.

Tavare:1980:TRA

[Tav80a] S. Tavaré. Time reversal and age distributions, I. Discretetime Markov chains. Journal of Applied Probability, 17(1):33– 46, March 1980. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3212922.

Tavares:1980:EMS

[Tav80b] L. Valadares Tavares. An exponential Markovian stationary process. Journal of Applied Probability, 17(4):1117–1120, December 1980. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3213224.

Tavares:1981:DMC

[Tav81] L. Valadares Tavares. The distribution of the maximum of i consecutive values of a process. Journal of Applied Probability, 18(4):957, December 1981. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3213072.

Tavangar:2020:BFR

[Tav20] Mahdi Tavangar. On the behavior of the failure rate and reversed failure rate in engineering systems. Journal of Applied Probability, 57(3):899–910, September 2020. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL https:/

/www.cambridge.org/core/journals/journal-of-applied-
probability/article/on-the-behavior-of-the-failure-rate-
and-reversed-failure-rate-in-engineering-systems/D5960DFC7A7038709FB0C8C1D7D6D133.

Tavare:2022:NST

[Tav22] Simon Tavaré. A note on the Screaming Toes game. Journal of Applied Probability, 59(1):118–130, March 2022. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL https://www.cambridge.org/core/journals/journal-
of-applied-probability/article/note-on-the-screaming-
toes-game/A19778FD0F8C4163A4401526992F7A97.

Taylor:1971:MMS

[Tay71] G. C. Taylor. Moments in Markovian systems with lumped states. Journal of Applied Probability, 8(3):599–605, September 1971. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3212182.

Taylor:1979:ESS

[Tay79] Peter D. Taylor. Evolutionarily stable strategies with two types of player. Journal of Applied Probability, 16(1):76–83, March 1979. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3213376.

Taylor:1999:FMC

[Tay99] Howard M. Taylor. The fundamental matrix for a certain random walk. Journal of Applied Probability, 36(2):320–333, June 1999. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3215459.

Tehranchi:2009:AIV

[Teh09] Michael R. Tehranchi. Asymptotics of implied volatility far from maturity. Journal of Applied Probability, 46(3):629–650, September 2009. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/25662452.

Tehranchi:2010:CAC

[Teh10] Michael R. Tehranchi. Characterizing attainable claims: a new proof. Journal of Applied Probability, 47(4):1013–1022, December 2010. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/29777422.

Tehranchi:2013:UMC

[Teh13] Michael R. Tehranchi. On the uniqueness of martingales with certain prescribed marginals. Journal of Applied Probability, 50 (2):557–575, June 2013. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://projecteuclid.

org/euclid.jap/1371648961.

Tenenbein:1975:DMP

[Ten75] Aaron Tenenbein. Distribution of minimum path. Journal of Applied Probability, 12(1):164–166, March 1975. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3212421.

Tessera:1984:PPA

[Tes84] Abebe Tessera. Population processes allowing emigration of families. Journal of Applied Probability, 21(2):225–232, June 1984. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/3213635.

Tetzlaff:1998:NCC

[Tet98] Ulrich A. W. Tetzlaff. On normalization constants for closed queueing networks with finite local buffers. Journal of Applied Probability, 35(3):600–607, September 1998. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/3215635.

Tetzlaff:2002:BRR

[Tet02] Guillermo Tomás Tetzlaff. Breakage and restoration in recursive trees. Journal of Applied Probability, 39(2):383–390, June 2002. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3216103.

Teugels:1968:EED

[Teu68a] Jozef L. Teugels. Exponential ergodicity in derived Markov chains. Journal of Applied Probability, 5(3):669–678, December 1968. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3211929.

Teugels:1968:EEM

[Teu68b] Jozef L. Teugels. Exponential ergodicity in Markov renewal processes. Journal of Applied Probability, 5(2):387–400, August 1968. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3212260.

Teugels:1972:EGE

[Teu72] Jozef L. Teugels. An example on geometric ergodicity of a finite Markov chain. Journal of Applied Probability, 9(2):466– 469, June 1972. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3212819.

Todorovic:1987:MEE

[TG87] P. Todorovic and J. Gani. Modeling the effect of erosion on crop production. Journal of Applied Probability, 24(4):787–797, December 1987. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3214205.

Tufekci:1999:IAS

[TG99] Tolga Tüfekçi and Refik Güllü. An iterative approximation scheme for repetitive Markov processes. Journal of Applied Probability, 36(3):654–667, September 1999. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/3215431.

Takacs:1971:LE

[TH71] L. Takács and John A. Hooke. Letters to the Editor. Journal of Applied Probability, 8(4):848–849, December 1971. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3212251.

Tsokos:1972:SAE

[TH72] Chris P. Tsokos and M. A. Hamdan. Stochastic asymptotic exponential stability of stochastic integral equations. Journal of Applied Probability, 9(1):169–177, March 1972. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/3212645.

Tan:2011:SAM

[TH11] Chin Hon Tan and Joseph C. Hartman. Sensitivity analysis in Markov decision processes with uncertain reward parameters. Journal of Applied Probability, 48(4):954–967, December 2011. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/23066436.

Thall:1981:CSM

[Tha81] Peter F. Thall. Cluster shock models. Journal of Applied Probability, 18(1):104–111, March 1981. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3213170.

Toomaj:2023:ECD

[THB23] Abdolsaeed Toomaj, Majid Hashempour, and Narayanaswamy Balakrishnan. Extropy: Characterizations and dynamic versions. Journal of Applied Probability, 60(4):1333–1351, December 2023. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-

6072 (electronic). URL https://www.cambridge.org/core/
journals/journal-of-applied-probability/article/extropy-
characterizations-and-dynamic-versions/B719445063A491F78A02E63D7494C03B.

Thedeen:1969:RTF

[The69] Torbjörn Thedéen. On road traffic with free overtaking. Journal of Applied Probability, 6(3):524–549, December 1969. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3212100.

Thedeen:1979:RBO

[The79] Torbjörn Thedeen. On the relation between offered, critical and accepted gap distributions. Journal of Applied Probability, 16 (1):54–64, March 1979. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3213374.

Thomas:1975:SMF

[Tho75a] Marlin U. Thomas. Some mean first-passage time approximations for the Ornstein–Uhlenbeck process. Journal of Applied Probability, 12(3):600–604, September 1975. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/3212877.

Thompson:1975:NLD

[Tho75b] J. W. Thompson. A note on the Lévy distance. Journal of Applied Probability, 12(2):412–414, June 1975. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/3212460.

Thorisson:1985:REP

[Tho85] Hermann Thorisson. On regenerative and ergodic properties of the k-server queue with non-stationary Poisson arrivals. Journal of Applied Probability, 22(4):893–902, December 1985. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3213956.

Thomson:1986:BLS

[Tho86] P. J. Thomson. Band-limited spectral estimation of autoregressive: Moving-average processes. Journal of Applied Probability, 23A:143–155, 1986. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3214349.

Thomas:2011:LBF

[Tho11] Peter J. Thomas. A lower bound for the first passage time density of the suprathreshold Ornstein–Uhlenbeck process. Journal of Applied Probability, 48(2):420–434, June 2011. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/23065894.

Thornblad:2016:DCI

[Thö16] Erik Thörnblad. The dominating colour of an infinite Pólya urn model. Journal of Applied Probability, 53(3):914–924, September 2016. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-

6072 (electronic). URL http://projecteuclid.org/euclid.
jap/1476370785.

Tan:2012:AMS

[THP12] Zhongquan Tan, Enkelejd Hashorva, and Zuoxiang Peng. Asymptotics of maxima of strongly dependent Gaussian processes. Journal of Applied Probability, 49(4):1106–1118, December 2012. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/41713831.

Tin:1985:QSM

[Tin85] Pyke Tin. A queueing system with Markov-dependent arrivals. Journal of Applied Probability, 22(3):668–677, September 1985. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/3213869.

Titterington:1974:EPT

[Tit74] D. M. Titterington. An example of a perturbation-type extremum controller. Journal of Applied Probability, 11(3):513– 520, September 1974. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3212695.

Troutman:1984:EWF

[TK84] Brent M. Troutman and Michael R. Karlinger. On the expected width function for topologically random channel networks. Journal of Applied Probability, 21(4):836–849, December 1984. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/3213700.

Torrado:2015:SOR

[TK15] Nuria Torrado and Subhash C. Kochar. Stochastic order relations among parallel systems from Weibull distributions. Journal of Applied Probability, 52(1):102–116, March 2015. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://projecteuclid.org/euclid.jap/1429282609.

Tallis:1969:GMM

[TL69] G. M. Tallis and R. T. Leslie. General models for r-molecular reactions. Journal of Applied Probability, 6(1):74–87, April 1969. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/3212277.

Turkman:1992:LLM

[TO92] K. F. Turkman and M. F. Oliveira. Limit laws for the maxima of chain-dependent sequences with positive extremal index. Journal of Applied Probability, 29(1):222–227, March 1992. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3214808.

Todorovic:1985:LE

[Tod85] P. Todorovic. Letter to the Editor. Journal of Applied Probability, 22(4):996, December 1985. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3213971.

Tokarev:2008:GIT

[Tok08] Daniel Tokarev. Growth of integral transforms and extinction in critical Galton–Watson processes. Journal of Applied Probability, 45(2):472–480, June 2008. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/27595958.

Tollar:1988:CGD

[Tol88] Eric S. Tollar. A characterization of the gamma distribution from a random difference equation. Journal of Applied Probability, 25(1):142–149, March 1988. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3214241.

Tomkins:1971:IH

[Tom71] R. J. Tomkins. On an inequality of Heyde. Journal of Applied Probability, 8(2):428–429, June 1971. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3211915.

Tomko:1972:RCL

[Tom72] Joseph Tomko. The rate of convergence in limit theorems for service systems with finite queue capacity. Journal of Applied Probability, 9(1):87–102, March 1972. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3212639.

Tong:1973:SCS

[Ton73] H. Tong. Some comments on spectral representations of nonstationary stochastic processes. Journal of Applied Probability, 10(4):881–885, December 1973. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.

org/stable/3212391.

Tong:1974:TDL

[Ton74] H. Tong. On time-dependent linear transformations of nonstationary stochastic processes. Journal of Applied Probability, 11(1):53–62, March 1974. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3212582.

Tong:1975:DOM

[Ton75] H. Tong. Determination of the order of a Markov chain by Akaike’s information criterion. Journal of Applied Probability, 12(3):488–497, September 1975. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.

org/stable/3212863.

Tong:1985:RIL

[Ton85] Y. L. Tong. A rearrangement inequality for the longest run, with an application to network reliability. Journal of Applied Probability, 22(2):386–393, June 1985. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3213781.

Tong:1997:SMO

[Ton97] Y. L. Tong. Some majorization orderings of heterogeneity in a class of epidemics. Journal of Applied Probability, 34(1):84– 93, March 1997. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3215177.

Topkis:1986:RHR

[Top86] Donald M. Topkis. Reordering heuristics for routing in communications networks. Journal of Applied Probability, 23(1):130– 143, March 1986. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3214122.

Tortrat:1969:TLD

[Tor69] A. Tortrat. Sur un théorème de Lewis et la décomposition en facteurs premiers de la loi rectangulaire. (French) [On a theorem of Lewis and decomposition in prime factors of the rectangular law]. Journal of Applied Probability, 6(1):177–185, April 1969. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/3212284.

Torrez:1979:CEP

[Tor79] William C. Torrez. Calculating extinction probabilities for the birth and death chain in a random environment. Journal of Applied Probability, 16(4):709–720, December 1979. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3213138.

Torrisi:2002:CIM

[Tor02] G. L. Torrisi. A class of interacting marked point processes: Rate of convergence to equilibrium. Journal of Applied Probability, 39 (1):137–160, March 2002. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3215924.

Tortorella:2010:RPS

[Tor10] Michael Tortorella. On reliability prediction and semi-renewal processes. Journal of Applied Probability, 47(2):394–406, June 2010. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/20778487.

Toyoizumi:1997:SIR

[Toy97] Hiroshi Toyoizumi. Sengupta’s invariant relationship and its application to waiting time inference. Journal of Applied Probability, 34(3):795–799, September 1997. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3215105.

Tin:1974:IDI

[TP74] Pyke Tin and R. M. Phatarfod. On infinite dams with inputs forming a stationary process. Journal of Applied Probability, 11 (3):553–561, September 1974. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.

org/stable/3212699.

Tin:1976:SER

[TP76] Pyke Tin and R. M. Phatarfod. Some exact results for dams with Markovian inputs. Journal of Applied Probability, 13(2):329– 337, June 1976. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/

3212835. See correction [Bal78].

Tory:1979:SCB

[TP79] Elmer M. Tory and David K. Pickard. Some comments on ’Sequential Random Packing in the Plane’ by H. J. Weiner. Journal of Applied Probability, 16(3):699–702, September 1979. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/3213104. See comment

[Wei78] and reply [Wei79].

TenHoopen:1965:SIT

[TR65] M. Ten Hoopen and H. A. Reuver. Selective interaction of two independent recurrent processes. Journal of Applied Probability, 2(2):286–292, December 1965. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.

org/stable/3212195.

TenHoopen:1967:FPP

[TR67] M. Ten Hoopen and H. A. Reuver. On a first passage problem in stochastic storage systems with total release. Journal of Applied Probability, 4(2):409–412, August 1967. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3212036.

Terdik:1989:WIR

[TR89] Gy. Terdik and T. Subba Rao. On Wiener–Itô representation and the best linear predictors for bilinear time series. Journal of Applied Probability, 26(2):274–286, June 1989. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3214034.

Trajstman:1981:BGP

[Tra81] A. C. Trajstman. A bounded growth population subjected to emigrations due to population pressure. Journal of Applied Probability, 18(3):571–582, September 1981. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/3213312.

Trieb:1992:PUM

[Tri92] Gudrun Trieb. A Pólya urn model and the coalescent. Journal of Applied Probability, 29(1):1–10, March 1992. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3214786.

Troutman:1983:WCA

[Tro83] Brent M. Troutman. Weak convergence of the adjusted range of cumulative sums of exchangeable random variables. Journal of Applied Probability, 20(2):297–304, June 1983. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3213803.

Truffet:2000:RTD

[Tru00] Laurent Truffet. Reduction techniques for discrete-time Markov chains on totally ordered state space using stochastic comparisons. Journal of Applied Probability, 37(3):795–806, September 2000. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3215614.

Tsaklidis:1994:EAS

[Tsa94] George M. Tsaklidis. The evolution of the attainable structures of a homogeneous Markov system with fixed size. Journal of Applied Probability, 31(2):348–361, June 1994. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/3215028.

Tsaklidis:1996:EAS

[Tsa96] George M. Tsaklidis. The evolution of the attainable structures of a continuous time homogeneous Markov system with fixed size. Journal of Applied Probability, 33(1):34–47, March 1996. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/3215262.

Tsaklidis:1999:STE

[Tsa99] George M. Tsaklidis. The stress tensor and the energy of a continuous time homogeneous Markov system with fixed size. Journal of Applied Probability, 36(1):21–29, March 1999. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/3215399.

Touyar:2008:PAN

[TSCD08] Narjiss Touyar, Sophie Schbath, Dominique Cellier, and Hélène Dauchel. Poisson approximation for the number of repeats in a stationary Markov chain. Journal of Applied Probability, 45 (2):440–455, June 2008. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/27595956.

Tsitsiashvili:2007:MRA

[Tsi07] G. S. Tsitsiashvili. Moments of random allocation processes reaching a boundary. Journal of Applied Probability, 44(4): 990–995, December 2007. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/27595902.

Toomaj:2017:SPC

[TSN17] A. Toomaj, S. M. Sunoj, and J. Navarro. Some properties of the cumulative residual entropy of coherent and mixed systems. Journal of Applied Probability, 54(2):379–393, June 2017. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL https://www.cambridge.org/core/journals/journal-
of-applied-probability/article/some-properties-of-the-
cumulative-residual-entropy-of-coherent-and-mixed-systems/
492442B841FC4F96AD0446BF2EE1E3C1.

Tsokos:1971:MVM

[Tso71] Chris P. Tsokos. The method of V. M. Popov for differential systems with random parameters. Journal of Applied Probability, 8(2):298–310, June 1971. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3211900.

Tsoucas:1992:RES

[Tso92] Pantelis Tsoucas. Rare events in series of queues. Journal of Applied Probability, 29(1):168–175, March 1992. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3214800.

Tsukuda:2017:ELM

[Tsu17] Koji Tsukuda. Estimating the large mutation parameter of the Ewens sampling formula. Journal of Applied Probability, 54(1):42–54, March 2017. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL https:/

/www.cambridge.org/core/journals/journal-of-applied-
probability/article/estimating-the-large-mutation-parameter-
of-the-ewens-sampling-formula/81515B68DC3F04B9275EC81B6B1427DB.

See correction [Tsu18].

Tsukuda:2018:ELM

[Tsu18] K. Tsukuda. Estimating the large mutation parameter of the Ewens sampling formula — correction. Journal of Applied Probability, 55(3):998–999, September 2018. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL https://www.cambridge.org/core/journals/journal-
of-applied-probability/article/estimating-the-large-
mutation-parameter-of-the-ewens-sampling-formula-correction/
3CE84F5757567FC0131934CC578B7B91. See [Tsu17].

Tuominen:1979:EEM

[TT79] Pekka Tuominen and Richard L. Tweedie. Exponential ergodicity in Markovian queueing and dam models. Journal of Applied Probability, 16(4):867–880, December 1979. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/3213152.

Thavaneswaran:1986:OES

[TT86] A. Thavaneswaran and M. E. Thompson. Optimal estimation for semimartingales. Journal of Applied Probability, 23(2):409– 417, June 1986. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3214183.

Takine:1991:STV

[TTH91] T. Takine, H. Takagi, and T. Hasegawa. Sojourn times in vacation and polling systems with Bernoulli feedback. Journal of Applied Probability, 28(2):422–432, June 1991. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/3214877.

Takine:1993:AMG

[TTH93] T. Takine, H. Takagi, and T. Hasegawa. Analysis of an M/G/1/K/N queue. Journal of Applied Probability, 30(2):446– 454, June 1993. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3214853.

Tomita:2022:RPG

[TTI22] Masashi Tomita, Koichiro Takaoka, and Motokazu Ishizaka. On the ruin probability of a generalized Cramér–Lundberg model driven by mixed Poisson processes. Journal of Applied Probability, 59(3):849–859, September 2022. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL https:/

/www.cambridge.org/core/journals/journal-of-applied-
probability/article/on-the-ruin-probability-of-a-generalized-
cramerlundberg-model-driven-by-mixed-poisson-processes/
D1721DD0ECA7C2CDDC36F42CCF616376.

Tu:2019:AOF

[Tu19] Eddie Tu. On association and other forms of positive dependence for Feller processes. Journal of Applied Probability, 56(2):624–646, June 2019. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL https:/

/www.cambridge.org/core/journals/journal-of-applied-
probability/article/on-association-and-other-forms-of-
positive-dependence-for-feller-processes/F739F2C1767DE54FB82A97197BB58A90.

Tuckwell:1976:FET

[Tuc76] Henry C. Tuckwell. On the first-exit time problem for temporally homogeneous Markov processes. Journal of Applied Probability, 13(1):39–48, March 1976. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3212663.

Tuffin:1999:BNA

[Tuf99] Bruno Tuffin. Bounded normal approximation in simulations of highly reliable Markovian systems. Journal of Applied Probability, 36(4):974–986, December 1999. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3215571.

Turnbull:1972:CLI

[Tur72] Bruce W. Turnbull. Chebyshev-like inequalities for dam models. Journal of Applied Probability, 9(3):617–629, September 1972. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/3212330.

Turkman:1992:OJT

[Tur92] Maria Antónia Amaral Turkman. Obituary: José Tiago de Oliveira. Journal of Applied Probability, 29(4):1005, December 1992. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3214732.

Turner:2000:JSQ

[Tur00] Stephen R. E. Turner. A join the shorter queue model in heavy traffic. Journal of Applied Probability, 37(1):212–223, March 2000. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3215669.

Turkman:2006:NEI

[Tur06] K. F. Turkman. A note on the extremal index for space-time processes. Journal of Applied Probability, 43(1):114–126, March 2006. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/27595714.

Tsaklidis:1988:APV

[TV88] G. Tsaklidis and P.-C. G. Vassiliou. Asymptotic periodicity of the variances and covariances of the state sizes in nonhomogeneous Markov systems. Journal of Applied Probability, 25(1):21–33, March 1988. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3214230.

Tsantas:1993:NHM

[TV93] N. Tsantas and P.-C. G. Vassiliou. The non-homogeneous Markov system in a stochastic environment. Journal of Applied Probability, 30(2):285–301, June 1993. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/3214839.

Teugels:2004:BCT

[TV04] Jef L. Teugels and Giovanni Vanroelen. Box–Cox transformations and heavy-tailed distributions. Journal of Applied Probability, 41A:213–227, 2004. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3215978.

Thomson:1975:NHC

[TW75] P. J. Thomson and K. M. Wilkinson. A note on Harris’s condition for recurrence in random walk processes. Journal of Applied Probability, 12(3):633–634, September 1975. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/3212883.

Turkman:1983:LLM

[TW83] K. F. Turkman and A. M. Walker. Limit laws for the maxima of a class of quasi-stationary sequences. Journal of Applied Probability, 20(4):814–821, December 1983. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/3213592.

Tuckwell:1984:FPT

[TW84] Henry C. Tuckwell and Frederic Y. M. Wan. First-passage time of Markov process to moving barriers. Journal of Applied Probability, 21(4):695–709, December 1984. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3213688.

Tsoucas:1989:MTN

[TW89] Pantelis Tsoucas and Jean Walrand. Monotonicity of throughput in non-Markovian networks. Journal of Applied Probability, 26(1):134–141, March 1989. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3214323.

Telcs:1999:BTI

[TW99] András Telcs and Nicholas C. Wormald. Branching and tree indexed random walks on fractals. Journal of Applied Probability, 36(4):999–1011, December 1999. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3215573.

Tang:2001:LMR

[TW01] Mengxiang Tang and Michael S. Waterman. Local matching of random restriction maps. Journal of Applied Probability, 38 (2):335–356, June 2001. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3215892.

Tweedie:1971:TPN

[Twe71] Richard L. Tweedie. Truncation procedures for non-negative matrices. Journal of Applied Probability, 8(2):311–320, June 1971. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/3211901.

Tweedie:1973:CLP

[Twe73] Richard L. Tweedie. The calculation of limit probabilities for denumerable Markov processes from infinitesimal properties. Journal of Applied Probability, 10(1):84–99, March 1973. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3212497.

Tweedie:1974:QSD

[Twe74] Richard L. Tweedie. Quasi-stationary distributions for Markov chains on a general state space. Journal of Applied Probability, 11(4):726–741, December 1974. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.

org/stable/3212556.

Tweedie:1975:RPR

[Twe75a] Richard L. Tweedie. The robustness of positive recurrence and recurrence of Markov chains under perturbations of the transition probabilities. Journal of Applied Probability, 12(4):744–752, December 1975. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3212725.

Tweedie:1975:TAL

[Twe75b] Richard L. Tweedie. Truncation approximation of the limit probabilities for denumerable semi-Markov processes. Journal of Applied Probability, 12(1):161–163, March 1975. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3212420.

Tweedie:1981:CEE

[Twe81] R. L. Tweedie. Criteria for ergodicity, exponential ergodicity and strong ergodicity of Markov processes. Journal of Applied Probability, 18(1):122–130, March 1981. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3213172.

Tweedie:1983:EMS

[Twe83] R. L. Tweedie. The existence of moments for stationary Markov chains. Journal of Applied Probability, 20(1):191–196, March 1983. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3213735.

Tweedie:1988:IMM

[Twe88] R. L. Tweedie. Invariant measures for Markov chains with no irreducibility assumptions. Journal of Applied Probability, 25A: 275–285, 1988. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3214162.

Tweedie:1998:TAI

[Twe98] R. L. Tweedie. Truncation approximations of invariant measures for Markov chains. Journal of Applied Probability, 35(3): 517–536, September 1998. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3215628.

Tzioufas:2011:GOD

[Tzi11] A. Tzioufas. On the growth of the one-dimensional reverse immunization contact processes. Journal of Applied Probability, 48 (3):611–623, September 2011. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.

org/stable/23065924.

Tzioufas:2016:ECR

[Tzi16] Achillefs Tzioufas. The entirely coupled region of supercritical contact processes. Journal of Applied Probability, 53(3): 925–929, September 2016. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://projecteuclid.

org/euclid.jap/1476370786.

Uppuluri:1974:APN

[UB74] V. R. R. Uppuluri and W. J. Blot. Asymptotic properties of the number of replications of a paired comparison. Journal of Applied Probability, 11(1):43–52, March 1974. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/3212581.

Unkelbach:1979:CET

[Unk79] Hans Dieter Unkelbach. Convergence to equilibrium in a traffic model with restricted passing. Journal of Applied Probability, 16(4):881–889, December 1979. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.

org/stable/3213153.

Unkelbach:1974:NMF

[UW74] Hans D. Unkelbach and Helmut Wegmann. The number of meetings in free Poisson traffic. Journal of Applied Probability, 11 (2):320–331, June 1974. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3212753.

Uwamariya:2023:LDE

[UY23] Denise Uwamariya and Xiangfeng Yang. Large deviations of extremal eigenvalues of sample covariance matrices. Journal of Applied Probability, 60(4):1275–1280, December 2023. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL https://www.cambridge.org/core/journals/journal-
of-applied-probability/article/large-deviations-of-extremal-
eigenvalues-of-sample-covariance-matrices/4D79A81EEBE91FE19765843476BE7404.

Vazquez-Abad:1992:EDS

[VAK92] Felisa J. Vázquez-Abad and Harold J. Kushner. Estimation of the derivative of a stationary measure with respect to a control parameter. Journal of Applied Probability, 29(2):343–352, June 1992. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3214571.

VanDoorn:1981:TSP

[Van81] Erik A. Van Doorn. The transient state probabilities for a queueing model where potential customers are discouraged by queue length. Journal of Applied Probability, 18(2):499–506, June 1981. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/3213296.

VanAssche:1986:PSM

[Van86] Walter Van Assche. Products of 2 × 2 stochastic matrices with random entries. Journal of Applied Probability, 23(4):1019–1024, December 1986. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3214475.

VanDijk:1989:PSI

[Van89a] Nico M. Van Dijk. A proof of simple insensitive bounds for a pure overflow system. Journal of Applied Probability, 26(1):113– 120, March 1989. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3214321.

VanWel:1989:CHU

[Van89b] Barthold F. Van Wel. The convex hull of a uniform sample from the interior of a simple d-polytope. Journal of Applied Probability, 26(2):259–273, June 1989. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3214033.

VanDijk:1990:QSR

[Van90] Nico M. Van Dijk. Queueing systems with restricted workload: An explicit solution. Journal of Applied Probability, 27(2):393– 400, June 1990. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3214658.

VanDijk:1991:NEU

[Van91a] Nico M. Van Dijk. A note on extended uniformization for nonexponential stochastic networks. Journal of Applied Probability, 28(4):955–961, December 1991. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.

org/stable/3214704.

VanDijk:1991:SRS

[Van91b] Nico M. Van Dijk. On ‘stop = repeat’ servicing for nonexponential queueing networks with blocking. Journal of Applied Probability, 28(1):159–173, March 1991. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http:/

/www.jstor.org/stable/3214748.

VanDenBerg:1996:NDO

[Van96] J. Van Den Berg. A note on disjoint-occurrence inequalities for marked Poisson point processes. Journal of Applied Probability, 33(2):420–426, June 1996. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3215065.

VanDerMei:1999:DPS

[Van99] R. D. Van Der Mei. Delay in polling systems with large switchover times. Journal of Applied Probability, 36(1):232–243, March 1999. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3215417.

Vanniasegaram:2006:HSD

[Van06] Sithparran Vanniasegaram. Le Her with s suits and d denominations. Journal of Applied Probability, 43(1):1–15, March 2006. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/27595706.

VanHoudt:2022:SAS

[Van22] Benny Van Houdt. Simple analytical solutions for the M^{b}/E_{k}/1/m, E_{k}/M^{b}/1/m, and related queues. Journal of Applied Probability, 59(4):1129–1143, December 2022. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL https://www.cambridge.org/core/journals/journal-
of-applied-probability/article/simple-analytical-solutions-
for-the-mathbfmbmathbfek1textbfm-mathbfekmathbfmb1textbfm-
and-related-queues/9FB3A3C700C3320C9B72FD06DEA188A8.

Vardi:1980:STS

[Var80] Yehuda Vardi. On a stopping time of Starr and its use in estimating the number of transmission sources. Journal of Applied Probability, 17(1):235–242, March 1980. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3212940.

Vassiliou:1981:SNH

[Vas81] P.-C. G. Vassiliou. Stability in a non-homogeneous Markov chain model in manpower systems. Journal of Applied Probability, 18 (4):924–930, December 1981. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.

org/stable/3213066.

Vassiliou:1982:ABM

[Vas82a] P.-C. G. Vassiliou. Asymptotic behavior of Markov systems. Journal of Applied Probability, 19(4):851–857, December 1982. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/3213839.

Vassiliou:1982:LBN

[Vas82b] P.-C. G. Vassiliou. On the limiting behaviour of a nonhomogeneous Markovian manpower model with independent Poisson input. Journal of Applied Probability, 19(2):433–438, June 1982. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3213497.

Vassiliou:1984:CBA

[Vas84] P.-C. G. Vassiliou. Cyclic behaviour and asymptotic stability of non-homogeneous Markov systems. Journal of Applied Probability, 21(2):315–325, June 1984. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3213642.

Vlasiou:2004:LTE

[VAW04] M. Vlasiou, I. J. B. F. Adan, and J. Wessels. A Lindleytype equation arising from a carousel problem. Journal of Applied Probability, 41(4):1171–1181, December 2004. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/4141388.

VanHoudt:2002:DDT

[VB02] B. Van Houdt and C. Blondia. The delay distribution of a type k customer in a first-come-first-served MMAP[K]/PH[K]/1 queue. Journal of Applied Probability, 39(1):213–223, March 2002. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/3215928.

Vellaisamy:1996:PCP

[VC96] P. Vellaisamy and B. Chaudhuri. Poisson and compound Poisson approximations for random sums of random variables. Journal of Applied Probability, 33(1):127–137, March 1996. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3215270.

Vlach:1969:DPG

[VD69] Thomas L. Vlach and Ralph L. Disney. The departure process from the GI/G/1 queue. Journal of Applied Probability, 6(3): 704–707, December 1969. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3212116.

vanDijk:1989:LFB

[vD89] Nico M. van Dijk. An LCFS finite buffer model with finite source batch input. Journal of Applied Probability, 26(2):372–380, June 1989. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3214042.

vanDoorn:2015:RDP

[vD15] Erik A. van Doorn. Representations for the decay parameter of a birth-death process based on the Courant–Fischer theorem. Journal of Applied Probability, 52(1):278–289, March 2015. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://projecteuclid.org/euclid.jap/1429282622.

Vatutin:2017:PSC

[VD17] Vladimir Vatutin and Elena Dyakonova. Path to survival for the critical branching processes in a random environment. Journal of Applied Probability, 54(2):588–602, June 2017. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL https://www.cambridge.org/core/journals/journal-
of-applied-probability/article/path-to-survival-for-
the-critical-branching-processes-in-a-random-environment/
9D13134FA94DF1574F134FAA2087ECF0.

vanDoorn:1987:CPJ

[vDB87] Erik A. van Doorn and Alan J. Branford. Comments on a paper by A. J. Branford. Journal of Applied Probability, 24(2):552– 554, June 1987. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/

3214280. See [Bra86].

vanderHofstad:2023:UAS

[vdHKPS23] Remco van der Hofstad, Stella Kapodistria, Zbigniew Palmowski, and Seva Shneer. Unified approach for solving exit problems for additive-increase and multiplicative-decrease processes. Journal of Applied Probability, 60(1):85–105, March 2023. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL https://www.cambridge.org/core/journals/journal-
of-applied-probability/article/unified-approach-for-
solving-exit-problems-for-additiveincrease-and-multiplicativedecrease-
processes/7C2AC329C8521F2E5A5160B46C8CBDB6.

vanDoorn:2005:EPB

[vDZ05] Erik A. van Doorn and Alexander I. Zeifman. Extinction probability in a birth-death process with killing. Journal of Applied Probability, 42(1):185–198, March 2005. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http:/

/www.jstor.org/stable/30040779.

VanDenBerg:1998:DCP

[VE98] J. Van Den Berg and A. Ermakov. On a distributed clustering process of Coffman, Courtois, Gilbert and Piret. Journal of Applied Probability, 35(4):919–924, December 1998. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3215528.

Veestraeten:2015:RFM

[Vee15] Dirk Veestraeten. A recursion formula for the moments of the first passage time of the Ornstein–Uhlenbeck process. Journal of Applied Probability, 52(2):595–601, June 2015. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://projecteuclid.org/euclid.jap/1437658618.

Vellaisamy:2004:PAE

[Vel04] P. Vellaisamy. Poisson approximation for (k_{1}, k_{2})-events via the Stein–Chen method. Journal of Applied Probability, 41(4):1081– 1092, December 2004. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/4141380.

Venezia:1979:OIP

[Ven79] Itzhak Venezia. Optimal insurance premium rates when the distribution of claims is unknown. Journal of Applied Probability, 16(3):678–684, September 1979. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.

org/stable/3213098.

Verhagen:1969:RST

[Ver69] A. M. V. Verhagen. Random size trees on the one and two dimensional integer lattice. Journal of Applied Probability, 6 (2):301–308, August 1969. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3212001.

Vermet:1998:PTL

[Ver98] Franck Vermet. Phase transition and law of large numbers for a non-symmetric one-dimensional random walk with selfinteractions. Journal of Applied Probability, 35(1):55–63, March 1998. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3215546.

Vergara:2014:ABT

[Ver14] Vicente Vergara. Asymptotic behaviour of the time-fractional telegraph equation. Journal of Applied Probability, 51(3):890– 893, September 2014. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://projecteuclid.

org/euclid.jap/1409932682.

Vassiliou:1985:VCG

[VG85] P.-C. G. Vassiliou and I. Gerontidis. Variances and covariances of the grade sizes in manpower systems. Journal of Applied Probability, 22(3):583–597, September 1985. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/3213862.

VanDenBerg:1992:LBT

[VG92] J. Van Den Berg and A. Gandolfi. LRU is better than FIFO under the independent reference model. Journal of Applied Probability, 29(1):239–243, March 1992. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3214811.

Vassiliou:1990:CAV

[VGT90] P.-C. G. Vassiliou, A. C. Georgiou, and N. Tsantas. Control of asymptotic variability in non-homogeneous Markov systems. Journal of Applied Probability, 27(4):756–766, December 1990. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/3214820.

vanHoorn:1986:AGG

[vHS86] M. H. van Hoorn and L. P. Seelen. Approximations for the GI/G/c queue. Journal of Applied Probability, 23(2):484– 494, June 1986. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3214189.

Vaze:2014:PIT

[VI14] Rahul Vaze and Srikanth Iyer. Percolation on the informationtheoretically secure signal to interference ratio graph. Journal of Applied Probability, 51(4):910–920, December 2014. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://projecteuclid.org/euclid.jap/1421763317.

Viaud:1994:SLC

[Via94] Daniel Pierre Loti Viaud. A strong law and a central limit theorem for controlled Galton–Watson processes. Journal of Applied Probability, 31(1):22–37, March 1994. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3215232.

Viaud:1995:RPR

[Via95] Daniel Pierre Loti Viaud. Random perturbations of recursive sequences with an application to an epidemic model. Journal of Applied Probability, 32(3):559–578, September 1995. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3215113.

Vidmar:2019:FPU

[Vid19] Matija Vidmar. First passage upwards for state-dependent-killed spectrally negative Lévy processes. Journal of Applied Probability, 56(2):472–495, June 2019. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL https:/

/www.cambridge.org/core/journals/journal-of-applied-
probability/article/first-passage-upwards-for-statedependentkilled-
spectrally-negative-levy-processes/832AA9D88EADE564851F5B76F08A7177.

Vidmar:2020:TFM

[Vid20] Matija Vidmar. A temporal factorization at the maximum for certain positive self-similar Markov processes. Journal of Applied Probability, 57(4):1045–1069, December 2020. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL https://www.cambridge.org/core/journals/journal-
of-applied-probability/article/temporal-factorization-
at-the-maximum-for-certain-positive-selfsimilar-markov-
processes/8A5011D6D2979E192A9B359EEBAA0833.

Videla:2023:TPG

[Vid23] Leonardo Videla. A transport process on graphs and its limiting distributions. Journal of Applied Probability, 60(1):341–357, March 2023. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL https:/

/www.cambridge.org/core/journals/journal-of-applied-
probability/article/transport-process-on-graphs-and-
its-limiting-distributions/7C547B573ADF214662685E2264D180C0.

Villeneuve:2007:TSS

[Vil07] Stéphane Villeneuve. On threshold strategies and the smoothfit principle for optimal stopping problems. Journal of Applied Probability, 44(1):181–198, March 2007. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http:/

/www.jstor.org/stable/27595830.

Viscolani:1986:AMQ

[Vis86] Bruno Viscolani. An adaptive multistage queueing system. Journal of Applied Probability, 23(2):495–503, June 1986. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3214190.

Vit:1974:ECT

[Vit74] Pedro Vit. On the equivalence of certain truncated point processes. Journal of Applied Probability, 11(3):601–604, September 1974. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3212707.

Vit:1976:ECM

[Vit76] Pedro Vit. On the equivalence of certain Markov chains. Journal of Applied Probability, 13(2):357–360, June 1976. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3212840.

Vere-Jones:1982:EFP

[VJ82] D. Vere-Jones. On the estimation of frequency in point-process data. Journal of Applied Probability, 19A:383–394, 1982. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/3213577.

Vere-Jones:2001:MSS

[VJ01] David Vere-Jones. The marriage of statistics and seismology. Journal of Applied Probability, 38A:1–5, 2001. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3215861.

Vere-Jones:1984:MSC

[VJO84] D. Vere-Jones and Y. Ogata. On the moments of a self-correcting process. Journal of Applied Probability, 21(2):335–342, June 1984. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3213644.

See corrections [Ano87i].

VanDenBerg:1985:IAP

[VK85] J. Van Den Berg and H. Kesten. Inequalities with applications to percolation and reliability. Journal of Applied Probability, 22(3): 556–569, September 1985. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3213860.

VanDijk:2000:SAM

[VK00] N. M. Van Dijk and H. Korezlioglu. Sensitivity analysis for Markov reward structures until entrance times. Journal of Applied Probability, 37(1):45–63, March 2000. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/3215658.

Vonta:2010:GMD

[VK10] Filia Vonta and Alex Karagrigoriou. Generalized measures of divergence in survival analysis and reliability. Journal of Applied Probability, 47(1):216–234, March 2010. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http:/

/www.jstor.org/stable/25662511.

Vasudeva:2014:ABE

[VKR14] R. Vasudeva, J. Vasantha Kumari, and S. Ravi. On the asymptotic behaviour of extremes and near maxima of random observations from the general error distributions. Journal of Applied Probability, 51(2):528–541, June 2014. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

projecteuclid.org/euclid.jap/1402578641.

Vatutin:2015:LTD

[VL15] Vladimir Vatutin and Quansheng Liu. Limit theorems for decomposable branching processes in a random environment. Journal of Applied Probability, 52(3):877–893, September 2015. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://projecteuclid.org/euclid.jap/1445543853.

vanLeeuwaarden:2013:RWR

[vLR13] Johan S. H. van Leeuwaarden and Kilian Raschel. Random walks reaching against all odds the other side of the quarter plane.

Journal of Applied Probability, 50(1):85–102, March 2013. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://projecteuclid.org/euclid.jap/1363784426.

Vasantam:2021:SMF

[VM21] Thirupathaiah Vasantam and Ravi R. Mazumdar. Sensitivity of mean-field fluctuations in Erlang loss models with randomized routing. Journal of Applied Probability, 58(2):428–448, June 2021. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL https:/

/www.cambridge.org/core/journals/journal-of-applied-
probability/article/sensitivity-of-meanfield-fluctuations-
in-erlang-loss-models-with-randomized-routing/ED84156F1ADD3A8A9F67C1E2A9FF61A1.

Voit:1996:ADE

[Voi96] Michael Voit. Asymptotic distributions for the Ehrenfest urn and related random walks. Journal of Applied Probability, 33 (2):340–356, June 1996. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3215058.

Voit:2000:NRC

[Voi00] Michael Voit. A note on the rate of convergence to equilibrium for Erlang’s model in the subcritical case. Journal of Applied Probability, 37(3):918–923, September 2000. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/3215624.

Voorn:1987:CLL

[Voo87] W. J. Voorn. Characterization of the logistic and loglogistic distributions by extreme value related stability with random sample size. Journal of Applied Probability, 24(4):838–851, December 1987. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3214209.

Voorn:1989:SER

[Voo89] W. J. Voorn. Stability of extremes with random sample size. Journal of Applied Probability, 26(4):734–743, December 1989. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/3214378.

Valiveti:1995:ALL

[VOZ95] R. S. Valiveti, B. J. Oommen, and J. R. Zgierski. Adaptive linear list reorganization under a generalized query system. Journal of Applied Probability, 32(3):793–804, September 1995. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3215130.

Vassiliou:1992:NHS

[VP92] P.-C. G. Vassiliou and A. A. Papadopoulou. Non-homogeneous semi-Markov systems and maintainability of the state sizes. Journal of Applied Probability, 29(3):519–534, September 1992. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/3214890.

Vellaisamy:2001:NBD

[VP01] P. Vellaisamy and Abraham P. Punnen. On the nature of the binomial distribution. Journal of Applied Probability, 38(1):36– 44, March 2001. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3215739.

Vlasiou:2010:TAR

[VP10] Maria Vlasiou and Zbigniew Palmowski. Tail asymptotics for a random sign Lindley recursion. Journal of Applied Probability, 47(1):72–83, March 2010. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/25662501.

VanOyen:1992:OIP

[VPT92] Mark P. Van Oyen, Dimitrios G. Pandelis, and Demosthenis Teneketzis. Optimality of index policies for stochastic scheduling with switching penalties. Journal of Applied Probability, 29(4): 957–966, December 1992. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3214727.

Vasdekis:2022:NPE

[VR22] Giorgos Vasdekis and Gareth O. Roberts. A note on the polynomial ergodicity of the one-dimensional zig-zag process. Journal of Applied Probability, 59(3):895–903, September 2022. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL https://www.cambridge.org/core/journals/journal-
of-applied-probability/article/note-on-the-polynomial-
ergodicity-of-the-onedimensional-zigzag-process/EE7214902A7325A94DDFF14CAC28E724.

Vrbik:1985:ADC

[Vrb85] Jan Vrbik. Asymptotic distribution of the coefficient of crossassociation based on a simple Markov-chain model. Journal of Applied Probability, 22(4):946–950, December 1985. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3213963.

VanDijk:1990:NEQ

[VS90] Nico M. Van Dijk and Eric Smeitink. A non-exponential queueing system with independent arrivals and batch servicing. Journal of Applied Probability, 27(2):401–408, June 1990. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3214659.

VanHarn:1991:CED

[VS91] K. Van Harn and F. W. Steutel. On a characterization of the exponential distribution. Journal of Applied Probability, 28(4): 947–949, December 1991. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3214702.

VanDijk:2000:NUD

[VS00] Nico M. Van Dijk and Karel Sladký. A note on uniformization for dynamic non-negative systems. Journal of Applied Probability, 37(2):329–341, June 2000. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3215711.

VanDijk:2006:TRV

[VS06] Nico M. Van Dijk and Karel Sladký. On the total reward variance for continuous-time Markov reward chains. Journal of Applied Probability, 43(4):1044–1052, December 2006. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/27595798.

VanLeeuwaarden:2009:QBD

[VSW09] Johan S. H. Van Leeuwaarden, Mark S. Squillante, and Erik M. M. Winands. Quasi-birth-and-death processes, lattice path counting, and hypergeometric functions. Journal of Applied Probability, 46(2):507–520, June 2009. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/25662440.

Veraverbeke:1975:ERC

[VT75] N. Veraverbeke and J. L. Teugels. The exponential rate of convergence of the distribution of the maximum of a random walk. Journal of Applied Probability, 12(2):279–288, June 1975. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/3212441.

Veraverbeke:1976:ERC

[VT76] N. Veraverbeke and J. L. Teugels. The exponential rate of convergence of the distribution of the maximum of a random walk. Part II. Journal of Applied Probability, 13(4):733–740, December 1976. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3212528.

Vassiliou:1989:RCV

[VT89] P.-C. G. Vassiliou and G. Tsaklidis. The rate of convergence of the vector of variances and covariances in non-homogeneous Markov systems. Journal of Applied Probability, 26(4):776–783, December 1989. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3214382.

Veillette:2011:TCP

[VT11] Mark S. Veillette and Murad S. Taqqu. A technique for computing the pdfs and cdfs of nonnegative infinitely divisible random variables. Journal of Applied Probability, 48(1):217–237, March 2011. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/29777456.

Vellaisamy:2009:SCN

[VU09] P. Vellaisamy and N. S. Upadhye. On the sums of compound negative binomial and gamma random variables. Journal of Applied Probability, 46(1):272–283, March 2009. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http:/

/www.jstor.org/stable/25662420.

Veitch:1986:RCT

[VW86] J. G. Veitch and G. S. Watson. Random cyclic transformations of points. Journal of Applied Probability, 23A:369–382, 1986.

CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/3214365.

Vecer:2004:MCT

[VX04] Jan Vecer and Mingxin Xu. The mean comparison theorem cannot be extended to the Poisson case. Journal of Applied Probability, 41(4):1199–1202, December 2004. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/4141392.

Waksman:1988:ISC

[Wak88] Peter Waksman. Inverting the statistics of the circle transform. Journal of Applied Probability, 25(4):708–716, December 1988. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/3214291.

Wakuta:1991:OCM

[Wak91] Kazuyoshi Wakuta. Optimal control of an M/G/1 queue with imperfectly observed queue length when the input source is finite. Journal of Applied Probability, 28(1):210–220, March 1991. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/3214751.

Walkup:1976:PSB

[Wal76] David W. Walkup. Pólya sequences, binomial convolution and the union of random sets. Journal of Applied Probability, 13(1): 76–85, March 1976. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3212667.

Walker:1981:CBI

[Wal81] A. M. Walker. On the classical Bonferroni inequalities and the corresponding galambos inequalities. Journal of Applied Probability, 18(3):757–763, September 1981. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3213333.

Walrand:1982:EFN

[Wal82] J. Walrand. On the equivalence of flows in networks of queues. Journal of Applied Probability, 19(1):195–203, March 1982. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/3213928.

Walrand:1983:NNT

[Wal83a] J. Walrand. A note on Norton’s theorem for queuing networks. Journal of Applied Probability, 20(2):442–444, June 1983. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/3213821.

Walrand:1983:DTQ

[Wal83b] Jean Walrand. A discrete-time queueing network. Journal of Applied Probability, 20(4):903–909, December 1983. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3213603.

Walker:1998:ETU

[Wal98] D. M. Walker. The expected time until absorption when absorption is not certain. Journal of Applied Probability, 35(4): 812–823, December 1998. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3215518.

Walton:2011:UOC

[Wal11] N. S. Walton. Utility optimization in congested queueing networks. Journal of Applied Probability, 48(1):68–89, March 2011. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/29777447.

Wang:1976:COQ

[Wan76a] Robert C. Wang. Computing optimal quality control policies: Two actions. Journal of Applied Probability, 13(4):826–832, December 1976. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3212542.

Wang:1976:ETM

[Wan76b] Robert C. Wang. The extinction time of Markov branching processes. Journal of Applied Probability, 13(2):345–347, June 1976. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/3212837.

Wang:1976:FEA

[Wan76c] Y. H. Wang. A functional equation and its application to the characterization of the Weibull and stable distributions. Journal of Applied Probability, 13(2):385–391, June 1976. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3212846.

Wang:1977:GAS

[Wan77a] Frank J. S. Wang. Gaussian approximation of some closed stochastic epidemic models. Journal of Applied Probability, 14 (2):221–231, June 1977. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3212993.

Wang:1977:ORP

[Wan77b] Robert C. Wang. Optimal replacement policy with unobservable states. Journal of Applied Probability, 14(2):340–348, June 1977. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3213004.

Wang:1980:CBB

[Wan80] Frank J. S. Wang. The convergence of a branching Brownian motion used as a model describing the spread of an epidemic. Journal of Applied Probability, 17(2):301–312, June 1980. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/3213020.

Wang:1981:LMB

[Wan81] Y. H. Wang. On the limit of the Markov binomial distribution. Journal of Applied Probability, 18(4):937–942, December 1981. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3213068.

Wang:1990:SAB

[Wan90] Suojin Wang. Saddlepoint approximations for bivariate distributions. Journal of Applied Probability, 27(3):586–597, September 1990. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3214543.

Wang:1992:AOT

[Wan92] Y. H. Wang. Approximating k-th-order two-state Markov chains. Journal of Applied Probability, 29(4):861–868, December 1992. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3214718.

Wang:1993:SAC

[Wan93] Suojin Wang. Saddlepoint approximations in conditional inference. Journal of Applied Probability, 30(2):397–404, June 1993. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/3214848.

Wang:1997:PPC

[Wan97] Yong-Jin Wang. A proof of the persistence criterion for a class of superprocesses. Journal of Applied Probability, 34(2):559– 563, June 1997. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3215395.

Wang:1999:TDM

[Wan99a] Chia-Li Wang. On the transient delays of M/G/1 queues. Journal of Applied Probability, 36(3):882–893, September 1999. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/3215448.

Wang:1999:EPS

[Wan99b] Han-Xing Wang. Extinction of population-size-dependent branching processes in random environments. Journal of Applied Probability, 36(1):146–154, March 1999. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http:/

/www.jstor.org/stable/3215410.

Wang:2012:EEL

[Wan12] Jian Wang. On the exponential ergodicity of Lévy–driven Ornstein–Uhlenbeck processes. Journal of Applied Probability, 49(4):990–1004, December 2012. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.

org/stable/41713824.

Wang:2014:ABD

[Wan14] Ruodu Wang. Asymptotic bounds for the distribution of the sum of dependent random variables. Journal of Applied Probability, 51(3):780–798, September 2014. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http:/

/projecteuclid.org/euclid.jap/1409932674.

Wang:2015:ESS

[Wan15] Hanchao Wang. The Euler scheme for a stochastic differential equation driven by pure jump semimartingales. Journal of Applied Probability, 52(1):149–166, March 2015. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://projecteuclid.org/euclid.jap/1429282612.

Wardrop:1976:SAR

[War76] Robert L. Wardrop. Some asymptotic results for optimal stopping based on capture times. Journal of Applied Probability, 13 (4):741–750, December 1976. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.

org/stable/3212529.

Warren:1999:FHT

[War99] Di Warren. The Frobenius–Harper technique in a general recurrence model. Journal of Applied Probability, 36(1):30–47, March 1999. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3215400.

Wasan:1969:SCF

[Was69] M. T. Wasan. Sufficient conditions for a first passage time process to be that of Brownian motion. Journal of Applied Probability, 6(1):218–223, April 1969. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3212290.

Watterson:1964:ADT

[Wat64] G. A. Watterson. The application of diffusion theory to two population genetic models of Moran. Journal of Applied Probability, 1(2):233–246, December 1964. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3211857.

Watson:1972:ESP

[Wat72] R. K. Watson. On an epidemic in a stratified population. Journal of Applied Probability, 9(3):659–666, September 1972. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3212334.

Watterson:1976:SDI

[Wat76] G. A. Watterson. The stationary distribution of the infinitelymany neutral alleles diffusion model. Journal of Applied Probability, 13(4):639–651, December 1976. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3212519. See correction [Ano77g].

Watson:1980:SDS

[Wat80a] Ray Watson. On the size distribution for some epidemic models. Journal of Applied Probability, 17(4):912–921, December 1980. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3213201.

Watson:1980:URT

[Wat80b] Ray Watson. A useful random time-scale transformation for the standard epidemic model. Journal of Applied Probability, 17 (2):324–332, June 1980. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3213022.

Watson:1982:DCS

[Wat82a] Geoffrey S. Watson. Distributions on the circle and sphere. Journal of Applied Probability, 19A:265–280, 1982. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3213566.

Watterson:1982:STM

[Wat82b] G. A. Watterson. Substitution times for mutant nucleotides. Journal of Applied Probability, 19A:59–70, 1982. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3213550.

Watson:1988:SDD

[Wat88] Geoffrey S. Watson. Some discrete distributions associated with orthogonal polynomials. Journal of Applied Probability, 25A: 167–171, 1988. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3214154.

Watkins:1997:MMC

[Wat97] Joseph C. Watkins. Mechanical models for cell movement: Locomotion, translocation, migration. Journal of Applied Probability, 34(4):827–846, December 1997. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3215000.

Watson:2023:GFM

[Wat23] Alexander R. Watson. A growth-fragmentation model connected to the ricocheted stable process. Journal of Applied Probability, 60(2):493–503, June 2023. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL https:/

/www.cambridge.org/core/journals/journal-of-applied-
probability/article/growthfragmentation-model-connected-
to-the-ricocheted-stable-process/CA211A6BC11AD5745F4F6062AAA120FE.

Waugh:1972:TES

[Wau72] W. A. O’N. Waugh. Taboo extinction, sojourn times, and asymptotic growth for the Markovian birth and death process. Journal of Applied Probability, 9(3):486–506, September 1972. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/3212321.

Waugh:1974:AGC

[Wau74a] W. A. O’N. Waugh. Asymptotic growth of a class of size-and-age-dependent birth processes. Journal of Applied Probability, 11(2):248–254, June 1974. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3212746.

Waugh:1974:MGC

[Wau74b] W. A. O’N. Waugh. Modes of growth of counting processes with increasing arrival rates. Journal of Applied Probability, 11 (2):237–247, June 1974. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3212745.

Ward:2003:SQN

[WB03] Amy R. Ward and Nicholas Bambos. On stability of queueing networks with job deadlines. Journal of Applied Probability, 40 (2):293–304, June 2003. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3215792.

Wilkinson:2016:DKM

[WBS16] Robert R. Wilkinson, Frank G. Ball, and Kieran J. Sharkey. The deterministic Kermack–McKendrick model bounds the general stochastic epidemic. Journal of Applied Probability, 53(4): 1031–1040, December 2016. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://projecteuclid.

org/euclid.jap/1481132834.

Weissman:1995:EIC

[WC95] Ishay Weissman and Uri Cohen. The extremal index and clustering of high values for derived stationary sequences. Journal of Applied Probability, 32(4):972–981, December 1995. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3215211.

Willmot:2004:ARL

[WC04] Gordon E. Willmot and Jun Cai. On applications of residual lifetimes of compound geometric convolutions. Journal of Applied Probability, 41(3):802–815, September 2004. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/4141355.

Wang:2003:CTL

[WCC03] Y. H. Wang, Hsiao-Fen Chang, and Shun-Yi Chen. Convergence theorems for the lengths of consecutive successes of Markov Bernoulli sequences. Journal of Applied Probability, 40(3):741– 749, September 2003. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3215947.

Wang:2005:CLS

[WCW05] Yuebao Wang, Dongya Cheng, and Kaiyong Wang. The closure of a local subexponential distribution class under convolution roots, with applications to the compound Poisson process. Journal of Applied Probability, 42(4):1194–1203, December 2005. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/30040838.

Walker:1999:CSE

[WD99] Stephen Walker and Paul Damien. Characterizations, stochastic equations, and the Gibbs sampler. Journal of Applied Probability, 36(3):747–751, September 1999. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3215438.

Wu:2009:DSE

[WDLC09] Xian-Yuan Wu, Zhao Dong, Ke Liu, and Kai-Yuan Cai. On the degree sequence of an evolving random graph process and its critical phenomenon. Journal of Applied Probability, 46(4): 1213–1220, December 2009. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/25662492.

Weber:1978:OAC

[Web78] Richard R. Weber. On the optimal assignment of customers to parallel servers. Journal of Applied Probability, 15(2):406– 413, June 1978. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3213411.

Weber:1979:IMQ

[Web79] Richard R. Weber. The interchangeability of · /M/1 queues in series. Journal of Applied Probability, 16(3):690–695, September 1979. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3213100.

Weber:1980:MAC

[Web80] N. C. Weber. A martingale approach to central limit theorems for exchangeable random variables. Journal of Applied Probability, 17(3):662–673, September 1980. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3212960.

Weber:1982:SJS

[Web82] Richard R. Weber. Scheduling jobs with stochastic processing requirements on parallel machines to minimize makespan or flowtime. Journal of Applied Probability, 19(1):167–182, March 1982. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3213926.

Weber:1986:WFM

[Web86a] N. C. Weber. The Wright–Fisher model with varying selection. Journal of Applied Probability, 23(2):504–508, June 1986. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/3214191.

Weber:1986:OSR

[Web86b] R. R. Weber. Optimal search for a randomly moving object. Journal of Applied Probability, 23(3):708–717, September 1986. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/3214009.

Weba:1999:BTV

[Web99] Michael Weba. Bounds for the total variation distance between the binomial and the Poisson distribution in case of mediumsized success probabilities. Journal of Applied Probability, 36 (1):97–104, March 1999. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3215405.

Weber:2007:DIR

[Web07] Stefan Weber. Distribution-invariant risk measures, entropy, and large deviations. Journal of Applied Probability, 44(1):16– 40, March 2007. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
27595820.

Wegmann:1976:CIN

[Weg76] Helmut Wegmann. The capacity of an intersection with nonstationary traffic. Journal of Applied Probability, 13(2):418– 422, June 1976. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3212851.

Weiner:1966:ADB

[Wei66a] Howard Weiner. On age dependent branching processes. Journal of Applied Probability, 3(2):383–402, December 1966. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3212127.

Weiner:1966:AAD

[Wei66b] Howard J. Weiner. Applications of the age distribution in age dependent branching processes. Journal of Applied Probability, 3(1):179–201, June 1966. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3212046.

Weiner:1969:SLA

[Wei69a] Howard J. Weiner. Sums of lifetimes in age dependent branching processes. Journal of Applied Probability, 6(1):195–200, April 1969. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3212286.

Weiss:1969:JAD

[Wei69b] Lionel Weiss. The joint asymptotic distribution of the k-smallest sample spacings. Journal of Applied Probability, 6(2):442–448, August 1969. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3212013.

Weiner:1970:MTC

[Wei70] H. J. Weiner. On a multi-type critical age-dependent branching process. Journal of Applied Probability, 7(3):523–543, December 1970. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3211936.

Weissner:1971:MBP

[Wei71] Edward W. Weissner. Multitype branching processes in random environments. Journal of Applied Probability, 8(1):17–31, March 1971. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3211834.

Weiner:1972:MTC

[Wei72] Howard J. Weiner. A multi-type critical age-dependent branching process with immigration. Journal of Applied Probability, 9 (4):697–706, December 1972. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.

org/stable/3212609.

Weiner:1974:TPM

[Wei74] Howard J. Weiner. Total progeny in a multitype critical age dependent branching process with immigration. Journal of Applied Probability, 11(3):458–470, September 1974. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/3212690.

Weiner:1975:CMC

[Wei75a] Howard J. Weiner. Conditional moments in a critical agedependent branching process. Journal of Applied Probability, 12 (3):581–587, September 1975. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.

org/stable/3212874.

Weiss:1975:TRL

[Wei75b] Gideon Weiss. Time-reversibility of linear stochastic processes. Journal of Applied Probability, 12(4):831–836, December 1975.

CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/3212735.

Weissman:1975:MEP

[Wei75c] Ishay Weissman. Multivariate extremal processes generated by independent non-identically distributed random variables. Journal of Applied Probability, 12(3):477–487, September 1975. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/3212862.

Weiner:1976:APC

[Wei76] Howard J. Weiner. Asymptotic probabilities in a critical agedependent branching process. Journal of Applied Probability, 13 (3):476–485, September 1976. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor. org/stable/3212467. See correction [Ano78f].

Weiner:1978:SRP

[Wei78] Howard J. Weiner. Sequential random packing in the plane. Journal of Applied Probability, 15(4):803–814, December 1978. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/3213435. See

criticism [Tan79b] and comments [TP79].

Weiner:1979:RLM

[Wei79] Howard Weiner. Reply to letters of M. Tanemura and E. M. Tory and D. K. Pickard. Journal of Applied Probability, 16(3): 706–707, September 1979. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/ stable/3213106. See [Tan79b, TP79, Tan80a].

Weiner:1980:AAL

[Wei80a] Howard Weiner. An alternative argument for lemma 3 and theorem 1 of [9]. Journal of Applied Probability, 17(3):878–880, September 1980. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3212985.

Weiner:1980:RRP

[Wei80b] Howard Weiner. Reply to remarks of Professor M. Hori. Journal of Applied Probability, 17(3):890–892, September 1980. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3212989. See [Hor80].

Weissman:1982:ASE

[Wei82] Ishay Weissman. Asymptotic sample extremes are not functions of an extremal process: A counterexample. Journal of Applied Probability, 19(2):475–476, June 1982. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3213504.

Weiner:1984:MMC

[Wei84] Howard Weiner. Moments of the maximum in a critical branching process. Journal of Applied Probability, 21(4):920–923, December 1984. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3213708.

Weil:1989:CPC

[Wei89] Wolfgang Weil. Collision probabilities for convex sets. Journal of Applied Probability, 26(3):649–654, September 1989. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3214423.

Weissman:1995:RPM

[Wei95] Ishay Weissman. Records from a power model of independent observations. Journal of Applied Probability, 32(4):982–990, December 1995. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3215212.

Weinberg:2000:SFB

[Wei00] Graham V. Weinberg. Stein factor bounds for random variables. Journal of Applied Probability, 37(4):1181–1187, December 2000. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/3215511.

Weiss:2005:JNU

[Wei05] Gideon Weiss. Jackson networks with unlimited supply of work. Journal of Applied Probability, 42(3):879–882, September 2005. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/30040865.

Wenocur:1986:BMQ

[Wen86] Michael L. Wenocur. Brownian motion with quadratic killing and some implications. Journal of Applied Probability, 23(4):

893–903, December 1986. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3214463.

Wenocur:1990:OUP

[Wen90] Michael L. Wenocur. Ornstein–Uhlenbeck process with quadratic killing. Journal of Applied Probability, 27(3):707–712, September 1990. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3214554.

Westcott:1973:RAE

[Wes73] M. Westcott. Results in the asymptotic and equilibrium theory of Poisson cluster processes. Journal of Applied Probability, 10 (4):807–823, December 1973. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.

org/stable/3212383.

Westcott:1977:NRT

[Wes77a] Mark Westcott. A note on record times. Journal of Applied Probability, 14(3):637–639, September 1977. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/3213470.

Westcott:1977:NNH

[Wes77b] Mark Westcott. A note on the non-homogeneous Poisson cluster process. Journal of Applied Probability, 14(2):396–398, June 1977. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3213012.

Westcott:1981:CED

[Wes81] Mark Westcott. Characterizing the exponential distribution. Journal of Applied Probability, 18(2):568, June 1981. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3213307.

Westcott:1982:AHC

[Wes82] Mark Westcott. Approximations to hard-core models and their application to statistical analysis. Journal of Applied Probability, 19A:281–292, 1982. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3213567.

Wesollowski:1999:PPM

[Wes99] Jacek Wesollowski. Poisson process via martingale and related characteristics. Journal of Applied Probability, 36(3):919–926, September 1999. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3215451.

Whittaker:1980:MGB

[WF80] Craig Whittaker and Richard M. Feldman. Moments for a general branching process in a semi-Markovian environment. Journal of Applied Probability, 17(2):341–349, June 1980. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3213024.

Wang:1999:ABP

[WF99] Han-Xing Wang and Dafan Fang. Asymptotic behaviour of population-size-dependent branching processes in Markovian random environments. Journal of Applied Probability, 36(2): 611–619, June 1999. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3215480.

Wheelwright:1993:DBC

[WG93] A. V. Wheelwright and C. A. Glasbey. Distances between censored intersections between a square lattice and a random smooth path. Journal of Applied Probability, 30(1):269–274, March 1993. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3214640.

Wu:2015:FPO

[WG15] Xiao Wu and Xianping Guo. First passage optimality and variance minimisation of Markov decision processes with varying discount factors. Journal of Applied Probability, 52(2):441–456, June 2015. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-

6072 (electronic). URL http://projecteuclid.org/euclid.
jap/1437658608.

Wu:2013:DCC

[WGF13] Tung-Lung Wu, Joseph Glaz, and James C. Fu. Discrete, continuous and conditional multiple window scan statistics. Journal of Applied Probability, 50(4):1089–1101, December 2013. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://projecteuclid.org/euclid.jap/1389370101.

Woodroofe:1975:ZL

[WH75] Michael Woodroofe and Bruce Hill. On Zipf’s law. Journal of Applied Probability, 12(3):425–434, September 1975. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3212857.

Wong:2004:MPS

[WH04] Bernard Wong and C. C. Heyde. On the martingale property of stochastic exponentials. Journal of Applied Probability, 41(3): 654–664, September 2004. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/ stable/4141344. See note [MU11, Won11].

Whittle:1968:EDO

[Whi68] P. Whittle. Equilibrium distributions for an open migration process. Journal of Applied Probability, 5(3):567–571, December 1968. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3211921.

Whitt:1971:WCF

[Whi71a] Ward Whitt. Weak convergence of first passage time processes. Journal of Applied Probability, 8(2):417–422, June 1971. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/3211913.

Whitt:1971:WCT

[Whi71b] Ward Whitt. Weak convergence theorems for priority queues: Preemptive-resume discipline. Journal of Applied Probability, 8 (1):74–94, March 1971. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3211839.

Whitt:1972:CHT

[Whi72a] Ward Whitt. Complements to heavy traffic limit theorems for the GI/G/1 queue. Journal of Applied Probability, 9(1):185– 191, March 1972. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3212647.

Whitt:1972:ERP

[Whi72b] Ward Whitt. Embedded renewal processes in the GI/G/s queue. Journal of Applied Probability, 9(3):650–658, September 1972. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/3212333.

Whitt:1972:LSD

[Whi72c] Ward Whitt. Limits for the superposition of m-dimensional point processes. Journal of Applied Probability, 9(2):462–465, June 1972. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3212818.

Whittle:1973:ACO

[Whi73] P. Whittle. An approximate characterisation of optimal stopping boundaries. Journal of Applied Probability, 10(1):158– 165, March 1973. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3212503.

Whittle:1979:SCR

[Whi79] P. Whittle. A simple condition for regularity in negative programming. Journal of Applied Probability, 16(2):305–318, June 1979. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3212899.

Whitt:1980:EVG

[Whi80a] Ward Whitt. The effect of variability in the GI/G/s queue. Journal of Applied Probability, 17(4):1062–1071, December 1980. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/3213215. See

correction [Ano84e].

Whitt:1980:UCS

[Whi80b] Ward Whitt. Uniform conditional stochastic order. Journal of Applied Probability, 17(1):112–123, March 1980. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3212929.

Whittle:1980:SCC

[Whi80c] P. Whittle. Stability and characterisation conditions in negative programming. Journal of Applied Probability, 17(3):635–645, September 1980. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3212957.

Whitt:1982:MML

[Whi82a] Ward Whitt. Multivariate monotone likelihood ratio and uniform conditional stochastic order. Journal of Applied Probability, 19(3):695–701, September 1982. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3213530.

Whittle:1982:IRF

[Whi82b] P. Whittle. Identities relating forward and backward treatments of optimisation. Journal of Applied Probability, 19A:395–399, 1982. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3213578.

Whitt:1985:RPS

[Whi85a] Ward Whitt. The renewal-process stationary-excess operator. Journal of Applied Probability, 22(1):156–167, March 1985. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/3213755.

Whitt:1985:UCV

[Whi85b] Ward Whitt. Uniform conditional variability ordering of probability distributions. Journal of Applied Probability, 22(3):619– 633, September 1985. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3213865.

Whittle:1985:PBI

[Whi85c] P. Whittle. Partial balance and insensitivity. Journal of Applied Probability, 22(1):168–176, March 1985. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http:/

/www.jstor.org/stable/3213756.

Whittle:1986:RSC

[Whi86] Peter Whittle. The risk-sensitive certainty equivalence principle. Journal of Applied Probability, 23A:383–388, 1986. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3214366.

Whittle:1988:RBA

[Whi88] P. Whittle. Restless bandits: Activity allocation in a changing world. Journal of Applied Probability, 25A:287–298, 1988. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/3214163.

Whittle:1990:CMM

[Whi90] P. Whittle. A construction for multi-modal processes, and a potential memory device. Journal of Applied Probability, 27 (1):146–155, March 1990. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3214602.

Whittle:2001:SPB

[Whi01] Peter Whittle. On the structure of proper Black–Scholes formulae. Journal of Applied Probability, 38A:243–248, 2001. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3215883.

Whittle:2005:TPU

[Whi05] P. Whittle. Tax problems in the undiscounted case. Journal of Applied Probability, 42(3):754–765, September 2005. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/30040854.

Wickwire:1975:NOC

[Wic75] K. Wickwire. A note on the optimal control of carrier-borne epidemics. Journal of Applied Probability, 12(3):565–568, September 1975. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3212871.

Wickwire:1979:EFP

[Wic79] K. Wickwire. The expected first-passage time for a sufficient statistic arising in the Poisson disorder problem. Journal of Applied Probability, 16(2):274–286, June 1979. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/3212896.

Wierman:1979:FVS

[Wie79] John C. Wierman. The front velocity of the simple epidemic. Journal of Applied Probability, 16(2):409–415, June 1979. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/3212908.

Wierman:1980:WMC

[Wie80] John C. Wierman. Weak moment conditions for time coordinates in first-passage percolation models. Journal of Applied Probability, 17(4):968–978, December 1980. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/3213206.

Wierman:1984:MPS

[Wie84] John C. Wierman. Mixed percolation on the square lattice. Journal of Applied Probability, 21(2):247–259, June 1984. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3213637.

Wiens:1989:BPD

[Wie89] Douglas P. Wiens. On the busy period distribution of the M/G/2 queueing system. Journal of Applied Probability, 26(4):858–865, December 1989. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3214390.

Wilkinson:1969:CEP

[Wil69] William E. Wilkinson. On calculating extinction probabilities for branching processes in random environments. Journal of Applied Probability, 6(3):478–492, December 1969. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/3212096.

Wilson:1977:LEC

[Wil77] S. Wilson. Letter to the Editor: On clumping effects in models of isolation by distance. Journal of Applied Probability, 14(3):662, September 1977. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3213475.

Williams:1982:SCC

[Wil82a] E. J. Williams. Some classes of conditional inference procedures. Journal of Applied Probability, 19A:293–303, 1982. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/3213568. See correction

[Ano82i].

Willie:1982:CTS

[Wil82b] J. S. Willie. Covariation of a time series and a point process. Journal of Applied Probability, 19(3):609–618, September 1982. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/3213518.

Willie:1982:MAT

[Wil82c] J. S. Willie. Measuring the association of a time series and a point process. Journal of Applied Probability, 19(3):597–608, September 1982. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3213517.

Wilson:1982:H

[Wil82d] Susan R. Wilson. Heritability. Journal of Applied Probability, 19A:71–85, 1982. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3213551.

Willekens:1988:HMP

[Wil88a] Eric Willekens. On higher moments of the population size in a subcritical branching process. Journal of Applied Probability, 25(2):413–417, June 1988. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3214449.

Willmot:1988:LE

[Wil88b] Gordon E. Willmot. Letter to the Editor. Journal of Applied Probability, 25(4):839, December 1988. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/3214307.

Willmot:1988:NEM

[Wil88c] Gordon E. Willmot. A note on the equilibrium M/G/1 queue length. Journal of Applied Probability, 25(1):228–231, March 1988. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3214251.

Williams:1992:AVP

[Wil92] R. J. Williams. Asymptotic variance parameters for the boundary local times of reflected Brownian motion on a compact interval. Journal of Applied Probability, 29(4):996–1002, December 1992. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3214730.

Willmot:1996:NEG

[Wil96] Gordon E. Willmot. A non-exponential generalization of an inequality arising in queueing and insurance risk. Journal of Applied Probability, 33(1):176–183, March 1996. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3215275.

Willie:1997:NSS

[Wil97] Helmut Willie. A note on single server loss systems with a superposition of inputs. Journal of Applied Probability, 34(1):213– 222, March 1997. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3215188.

Willmot:1999:LTR

[Wil99] Gordon E. Willmot. A Laplace transform representation in a class of renewal queueing and risk processes. Journal of Applied Probability, 36(2):570–584, June 1999. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3215475.

Willmot:2002:HOP

[Wil02] Gordon E. Willmot. On higher-order properties of compound geometric distributions. Journal of Applied Probability, 39(2):324– 340, June 2002. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3216099.

Winston:1977:OSL

[Win77] Wayne Winston. Optimality of the shortest line discipline. Journal of Applied Probability, 14(1):181–189, March 1977. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3213271.

Winter:1979:CRP

[Win79] B. B. Winter. Convergence rate of perturbed empirical distribution functions. Journal of Applied Probability, 16(1):163– 173, March 1979. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3213384.

Winter:1989:JSB

[Win89] B. B. Winter. Joint simulation of backward and forward recurrence times in a renewal process. Journal of Applied Probability, 26(2):404–407, June 1989. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3214046.

Winkler:1990:ETS

[Win90] Gerhard Winkler. An ergodic L^{2}-theorem for simulated annealing in Bayesian image reconstruction. Journal of Applied Probability, 27(4):779–791, December 1990. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3214822.

Winkel:2005:EFE

[Win05] Matthias Winkel. Electronic foreign-exchange markets and passage events of independent subordinators. Journal of Applied Probability, 42(1):138–152, March 2005. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http:/

/www.jstor.org/stable/30040775.

Windridge:2015:ETS

[Win15] Peter Windridge. The extinction time of a subcritical branching process related to the SIR epidemic on a random graph. Journal of Applied Probability, 52(4):1195–1201, December 2015. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://projecteuclid.org/euclid.jap/1450802763.

Wiroonsri:2018:SMN

[Wir18] Nathakhun Wiroonsri. Stein’s method for negatively associated random variables with applications to second-order stationary random fields. Journal of Applied Probability, 55(1):196–215, March 2018. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL https:/

/www.cambridge.org/core/journals/journal-of-applied-
probability/article/steins-method-for-negatively-associated-
random-variables-with-applications-to-secondorder-stationary-
random-fields/6704BC1A87E73C60E86C58F964F7BD74.

Wisniewolski:2020:IFU

[Wiś20] Maciej Wiśniewolski. Integral functionals under the excursion measure. Journal of Applied Probability, 57(1):137–155, March 2020. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-

6072 (electronic). URL https://www.cambridge.org/core/
journals/journal-of-applied-probability/article/integral-
functionals-under-the-excursion-measure/F065B407B971E6780FBF1C2E45126EF6.

Witte:1990:USA

[Wit90] Hans-Jürgen Witte. A unification of some approaches to Poisson approximation. Journal of Applied Probability, 27(3):611–621, September 1990. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3214545.

Witte:1993:SCE

[Wit93] Hans-Jürgen Witte. Some characterizations of exponential or geometric distributions in a non-stationary record value model. Journal of Applied Probability, 30(2):373–381, June 1993. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/3214846.

Wittmann:1996:SII

[Wit96] Rainer Wittmann. Superprophet inequalities for independent random variables. Journal of Applied Probability, 33(3):904–908, September 1996. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3215367.

Wang:1995:LTN

[WJ95] Y. H. Wang and Shuixin Ji. Limit theorems for the number of occurrences of consecutive k successes in n Markovian trials. Journal of Applied Probability, 32(3):727–735, September 1995. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3215125.

Weiss:1965:MAS

[WK65] George H. Weiss and Motoo Kimura. A mathematical analysis of the stepping stone model of genetic correlation. Journal of Applied Probability, 2(1):129–149, June 1965. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/3211879.

Wortman:1994:MSO

[WK94] Martin A. Wortman and Georgia-Ann Klutke. On maintained systems operating in a random environment. Journal of Applied Probability, 31(2):589–594, June 1994. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3215052.

Willmot:1994:LBT

[WL94] Gordon E. Willmot and Xiaodong Lin. Lundberg bounds on the tails of compound distributions. Journal of Applied Probability, 31(3):743–756, September 1994. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.

org/stable/3215152.

Willmot:1997:SBT

[WL97] G. E. Willmot and Xiaodong Lin. Simplified bounds on the tails of compound distributions. Journal of Applied Probability, 34 (1):127–133, March 1997. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3215181.

Welberry:1977:ATD

[WM77] T. R. Welberry and G. H. Miller. An approximation to a twodimensional binary process. Journal of Applied Probability, 14 (4):862–868, December 1977. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.

org/stable/3213360.

Wolff:1970:WCP

[Wol70] Ronald W. Wolff. Work-conserving priorities. Journal of Applied Probability, 7(2):327–337, August 1970. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http:/

/www.jstor.org/stable/3211968.

Wolfowitz:1975:SGC

[Wol75] J. Wolfowitz. Signalling over a Gaussian channel with feedback and autoregressive noise. Journal of Applied Probability, 12(4): 713–723, December 1975. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3212722.

Wolff:1977:UBM

[Wol77a] Ronald W. Wolff. An upper bound for multi-channel queues. Journal of Applied Probability, 14(4):884–888, December 1977. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/3213363.

Wolfson:1977:LTS

[Wol77b] David B. Wolfson. Limit theorems for sums of a sequence of random variables defined on a Markov chain. Journal of Applied Probability, 14(3):614–620, September 1977. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/3213466.

Wolfson:1984:SMR

[Wol84] Barry Wolfson. Some moment results for certain tandem and multiple-server queues. Journal of Applied Probability, 21(4): 901–910, December 1984. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3213705.

Wolff:1987:UBW

[Wol87] Ronald W. Wolff. Upper bounds on work in system for multichannel queues. Journal of Applied Probability, 24(2):547– 551, June 1987. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3214279.

Wolff:1988:SPD

[Wol88] Ronald W. Wolff. Sample-path derivations of the excess, age, and spread distributions. Journal of Applied Probability, 25 (2):432–436, June 1988. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3214453.

Wolff:2002:LPC

[Wol02] Ronald W. Wolff. Losses per cycle in a single-server queue. Journal of Applied Probability, 39(4):905–909, December 2002. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/3216014.

Wong:1997:PMS

[Won97] Tityik Wong. Preservation of multivariate stochastic orders under multivariate Poisson shock models. Journal of Applied Probability, 34(4):1009–1020, December 1997. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/3215013.

Wong:2011:WH

[Won11] Bernard Wong. On Wong and Heyde (2004). Journal of Applied Probability, 48(4):1200, December 2011. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http:

//www.jstor.org/stable/23066453. See comment [WH04,

MU11].

Woods:1982:FPT

[Woo82] A. J. Woods. First-passage times for the partial sums of a sequence of geometric distributions. Journal of Applied Probability, 19(2):430–432, June 1982. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.

org/stable/3213496.

Woodward:1983:LPP

[Woo83a] Mark Woodward. The limiting properties of population distributions with particular application to manpower planning. Journal of Applied Probability, 20(1):19–30, March 1983. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/3213717.

Woolford:1983:GFC

[Woo83b] Samuel W. Woolford. On a generalized finite-capacity storage model. Journal of Applied Probability, 20(3):663–674, September 1983. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3213901.

Weiss:1980:STE

[WP80] Gideon Weiss and Michael Pinedo. Scheduling tasks with exponential service times on non-identical processors to minimize various cost functions. Journal of Applied Probability, 17(1):187– 202, March 1980. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/

3212936. See comment [Pra82].

Wang:1997:BCP

[WP97] Liqun Wang and Klaus Pötzelberger. Boundary crossing probability for Brownian motion and general boundaries. Journal of Applied Probability, 34(1):54–65, March 1997. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3215174.

Woodside:1980:DAC

[WPN80] C. M. Woodside, B. Pagurek, and G. F. Newell. A diffusion approximation for correlation in queues. Journal of Applied Probability, 17(4):1033–1047, December 1980. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/3213213.

Whittle:1974:OLS

[WR74] P. Whittle and J. Rudge. The optimal linear solution of a symmetric team control problem. Journal of Applied Probability, 11 (2):377–381, June 1974. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3212760.

Woodham:1997:CSR

[WR97] Sara-Anne Woodham and Donald St. P. Richards. Comparison of system reliability functions under laboratory and common operating environments. Journal of Applied Probability, 34 (2):536–545, June 1997. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3215392.

Warren:1996:PEN

[WS96] D. I. Warren and E. Seneta. Peaks and Eulerian numbers in a random sequence. Journal of Applied Probability, 33(1):101– 114, March 1996. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3215267.

Wu:2004:LTI

[WS04] Wei Biao Wu and Xiaofeng Shao. Limit theorems for iterated random functions. Journal of Applied Probability, 41(2):425– 436, June 2004. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3216026.

Wu:2005:PSR

[WT05] De-An Wu and Hideaki Takagi. Processor-sharing and randomservice queues with semi-Markovian arrivals. Journal of Applied Probability, 42(2):478–490, June 2005. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/30040805.

Wu:1968:SDO

[Wu68] P. M. Wu. Storage with deterministic outputs and inputs subject to breakdown. Journal of Applied Probability, 5(3):636–647, December 1968. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3211927.

Wu:1993:CBP

[Wu93] C. Chris Wu. Critical behavior of percolation and Markov fields on branching planes. Journal of Applied Probability, 30(3):538– 547, September 1993. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3214764.

Wubker:2012:STW

[Wüb12] Achim Wübker. Spectral theory for weakly reversible Markov chains. Journal of Applied Probability, 49(1):245–265, March 2012. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/41713763.

Weber:1986:SJS

[WVW86] R. R. Weber, P. Varaiya, and J. Walrand. Scheduling jobs with stochastically ordered processing times on parallel machines to minimize expected flowtime. Journal of Applied Probability, 23 (3):841–847, September 1986. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.

org/stable/3214023.

Wolff:1976:EEL

[WW76] Ronald W. Wolff and Charles W. Wrightson. An extension of Erlang’s loss formula. Journal of Applied Probability, 13(3): 628–632, September 1976. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3212487.

Weber:1990:IPR

[WW90] Richard R. Weber and Gideon Weiss. On an index policy for restless bandits. Journal of Applied Probability, 27(3):637–648, September 1990. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3214547.

Wu:2000:CLT

[WW00] Wei Biao Wu and Michael Woodroofe. A central limit theorem for iterated random functions. Journal of Applied Probability, 37 (3):748–755, September 2000. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.

org/stable/3215610.

Wolff:2002:CLP

[WW02] Ronald W. Wolff and Chia-Li Wang. On the convexity of loss probabilities. Journal of Applied Probability, 39(2):402– 406, June 2002. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3216106.

Wang:2004:SLT

[WW04] Hai-Bin Wang and Bo-Cheng Wei. Separable lower triangular bilinear model. Journal of Applied Probability, 41(1):221– 235, March 2004. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3215827.

Wan:2006:TSB

[WW06a] Yat-Wah Wan and Ronald W. Wolff. On the time spent below a random threshold by a system driven by a general counting process. Journal of Applied Probability, 43(2):599–601, June 2006. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/27595754.

Wang:2006:ADS

[WW06b] Yuebao Wang and Kaiyong Wang. Asymptotics of the density of the supremum of a random walk with heavy-tailed increments. Journal of Applied Probability, 43(3):874–879, September 2006. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/27595779.

Wang:2007:PLD

[WW07] Shijie Wang and Wensheng Wang. Precise large deviations for sums of random variables with consistently varying tails in multirisk models. Journal of Applied Probability, 44(4):889–900, December 2007. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
27595895.

Wendler:2020:CLT

[WW20] Martin Wendler and Wei Biao Wu. Central limit theorems for nearly long range dependent subordinated linear processes. Journal of Applied Probability, 57(2):637–656, June 2020. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL https://www.cambridge.org/core/journals/journal-
of-applied-probability/article/central-limit-theorems-
for-nearly-long-range-dependent-subordinated-linear-
processes/961895E181C7382BB843022C26ECDB9C.

Wang:2008:NBA

[WX08] Xiaoxin Wang and Aihua Xia. On negative binomial approximation to k-runs. Journal of Applied Probability, 45(2):456– 471, June 2008. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
27595957.

Watson:1993:BCP

[WY93] Ray Watson and Paul Yip. A bivariate counting process. Journal of Applied Probability, 30(2):353–364, June 1993. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3214844.

Wang:2022:TTL

[WY22] Hua-Ming Wang and Huizi Yao. Two-type linear-fractional branching processes in varying environments with asymptotically constant mean matrices. Journal of Applied Probability, 59(1):224–255, March 2022. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL https:/

/www.cambridge.org/core/journals/journal-of-applied-
probability/article/twotype-linearfractional-branching-
processes-in-varying-environments-with-asymptotically-
constant-mean-matrices/A2BA7D8F3DAE786ADC792545E7CD9644.

Wynn:1973:LSM

[Wyn73] H. P. Wynn. Limiting second moments for transient states of Markov chains. Journal of Applied Probability, 10(4):891–894, December 1973. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3212393.

Weiss:1965:SMM

[WZ65] George H. Weiss and Marvin Zelen. A semi-Markov model for clinical trials. Journal of Applied Probability, 2(2):269–285, December 1965. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3212194.

Wu:2007:CMP

[WZ07] Bo Wu and Yu-Hui Zhang. A class of multidimensional Qprocesses. Journal of Applied Probability, 44(1):226–237, March 2007. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/27595833.

Wu:2013:OBP

[WZ13] Xianyi Wu and Xian Zhou. Open bandit processes with uncountable states and time-backward effects. Journal of Applied Probability, 50(2):388–402, June 2013. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

projecteuclid.org/euclid.jap/1371648948.

Wang:2018:GDB

[WZ18] Wenyuan Wang and Xiaowen Zhou. General drawdown-based de Finetti optimization for spectrally negative Lévy risk processes. Journal of Applied Probability, 55(2):513–542, June 2018. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL https://www.cambridge.org/core/journals/journal-
of-applied-probability/article/general-drawdownbased-
de-finetti-optimization-for-spectrally-negative-levy-
risk-processes/45C6D61C3BED702980373AA77A430DD0.

Xu:2011:CHB

[XB11] Maochao Xu and N. Balakrishnan. On the convolution of heterogeneous Bernoulli random variables. Journal of Applied Probability, 48(3):877–884, September 2011. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/23065944.

Xu:2011:SIL

[XH11] Maochao Xu and Taizhong Hu. Some inequalities of linear combinations of independent random variables. I. Journal of Applied Probability, 48(4):1179–1188, December 2011. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/23066450.

Xia:1997:RPP

[Xia97] Aihua Xia. On the rate of Poisson process approximation to a Bernoulli process. Journal of Applied Probability, 34(4):898–907, December 1997. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3215005.

Xia:1999:PPS

[Xia99] Aihua Xia. A probabilistic proof of Stein’s factors. Journal of Applied Probability, 36(1):287–290, March 1999. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3215423.

Xie:1998:RBC

[XL98] M. Xie and C. D. Lai. On reliability bounds via conditional inequalities. Journal of Applied Probability, 35(1):104–114, March 1998. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3215550.

Xu:1990:SDM

[XMKW90] Susan H. Xu, Pitu B. Mirchandani, Srikanta P. R. Kumar, and Richard R. Weber. Stochastic dispatching of multi-priority jobs to heterogeneous processors. Journal of Applied Probability, 27 (4):852–861, December 1990. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.

org/stable/3214828.

Xue:2003:ABA

[XS03] J. Xue and Attahiru Sulealfa. Accurate bounds for the asymptotic constant in a statistical multiplexer with homogeneous generalized binary Markov sources. Journal of Applied Probability, 40(2):273–292, June 2003. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3215791.

Xu:2011:GCC

[XT11] Weiyu Xu and A. Kevin Tang. A generalized coupon collector problem. Journal of Applied Probability, 48(4):1081–1094, December 2011. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
23066444.

Xu:1998:CED

[Xu98] Jian-Lun Xu. Characterizations of the exponential distribution via the blocking time in a queueing system with an unreliable server. Journal of Applied Probability, 35(1):236–239, March 1998. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3215562.

Xu:2009:CLT

[Xu09] Fang Xu. A central limit theorem associated with the transformed two-parameter Poisson–Dirichlet distribution. Journal of Applied Probability, 46(2):392–401, June 2009. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/25662432.

Xu:2017:SOL

[Xu17] Fangjun Xu. Second-order limit laws for occupation times of fractional Brownian motion. Journal of Applied Probability, 54(2):444–461, June 2017. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL https:/

/www.cambridge.org/core/journals/journal-of-applied-
probability/article/secondorder-limit-laws-for-occupation-
times-of-fractional-brownian-motion/69701A3E49463215E084B531ED5981D5.

Xing:2005:ECP

[XW05] Yongsheng Xing and Yongjin Wang. On the extinction of a class of population-size-dependent bisexual branching processes. Journal of Applied Probability, 42(1):175–184, March 2005. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/30040778.

Xian-Yin:1992:RSR

[XY92] Zhou Xian-Yin. On the recurrence of simple random walks on some fractals. Journal of Applied Probability, 29(2):454– 459, June 1992. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3214582.

Xi:2009:APM

[XY09] Fubao Xi and G. Yin. Asymptotic properties of a meanfield model with a continuous-state-dependent switching process. Journal of Applied Probability, 46(1):221–243, March 2009. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/25662417.

Xing:2009:SDT

[XZW09] Xiaoyu Xing, Wei Zhang, and Yongjin Wang. The stationary distributions of two classes of reflected Ornstein–Uhlenbeck processes. Journal of Applied Probability, 46(3):709–720, September 2009. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/25662456.

Yashin:1988:NRI

[YA88] Anatoli Yashin and Elja Arjas. A note on random intensities and conditional survival functions. Journal of Applied Probability, 25 (3):630–635, September 1988. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.

org/stable/3213991.

Yadin:1970:QAP

[Yad70] Micha Yadin. Queueing with alternating priorities, treated as random walk on the lattice in the plane. Journal of Applied Probability, 7(1):196–218, April 1970. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3212160.

Yahav:1975:FPT

[Yah75] Joseph A. Yahav. On a fixed point theorem and its stochastic equivalent. Journal of Applied Probability, 12(3):605–611, September 1975. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3212878.

Yamada:1989:LTJ

[Yam89] Keigo Yamada. Limit theorems for jump shock models. Journal of Applied Probability, 26(4):793–806, December 1989. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3214384.

Yamamoto:2017:CLT

[Yam17] Ken Yamamoto. Central limit theorem for the Horton–Strahler bifurcation ratio of general branch order. Journal of Applied Probability, 54(4):1111–1124, December 2017. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL https://www.cambridge.org/core/journals/journal-
of-applied-probability/article/central-limit-theorem-
for-the-hortonstrahler-bifurcation-ratio-of-general-
branch-order/34285F834370E252E9FB5E39BE54D092.

Yamato:2022:UAN

[Yam22] Kosuke Yamato. A unifying approach to non-minimal quasistationary distributions for one-dimensional diffusions. Journal of Applied Probability, 59(4):1106–1128, December 2022. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL https://www.cambridge.org/core/journals/journal-
of-applied-probability/article/unifying-approach-to-
nonminimal-quasistationary-distributions-for-onedimensional-
diffusions/9B5D3E16FBAC1C765E13C07E598496A9.

Yang:1972:BPA

[Yan72] Y. S. Yang. On branching processes allowing immigration. Journal of Applied Probability, 9(1):24–31, March 1972. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3212633.

Yang:1973:APS

[Yan73] Y. S. Yang. Asymptotic properties of the stationary measure of a Markov branching process. Journal of Applied Probability, 10(2):447–450, June 1973. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3212362.

Yang:1974:RMR

[Yan74] Mark C. K. Yang. Recognizing the maximum of a random sequence based on relative rank with backward solicitation. Journal of Applied Probability, 11(3):504–512, September 1974. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/3212694.

Yang:1975:DIR

[Yan75] Mark C. K. Yang. On the distribution of the inter-record times in an increasing population. Journal of Applied Probability, 12 (1):148–154, March 1975. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3212417.

Yannaros:1988:ITR

[Yan88a] Nikos Yannaros. The inverses of thinned renewal processes. Journal of Applied Probability, 25(4):822–828, December 1988. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/3214304.

Yannaros:1988:CPG

[Yan88b] Nikos Yannaros. On Cox processes and gamma renewal processes. Journal of Applied Probability, 25(2):423–427, June 1988. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/3214451.

Yang:1990:RLS

[Yan90] Grace L. Yang. A renewal look at switching rules in the MIL-STD-105D sampling system. Journal of Applied Probability, 27 (1):183–192, March 1990. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3214605.

Yao:1985:FPT

[Yao85] David D. Yao. First-passage-time moments of Markov processes. Journal of Applied Probability, 22(4):939–945, December 1985. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/3213962.

Yao:1993:BCP

[Yao93] Qiwei Yao. Boundary-crossing probabilities of some random fields related to likelihood ratio tests for epidemic alternatives. Journal of Applied Probability, 30(1):52–65, March 1993. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/3214621.

Yao:1996:KCF

[Yao96] Yi-Ching Yao. On Kingman’s characteristic functional approach to Rényi’s characterization of Poisson processes. Journal of Applied Probability, 33(2):439–442, June 1996. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/3215068.

Yao:2007:EOS

[Yao07a] Yi-Ching Yao. Explicit optimal strategy for the Vardi Casino with limited playing time. Journal of Applied Probability, 44(3): 831–837, September 2007. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/27595885.

Yao:2007:OBP

[Yao07b] Yi-Ching Yao. On optimality of bold play for discounted dubins-Savage gambling problems with limited playing times. Journal of Applied Probability, 44(1):212–225, March 2007. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/27595832.

Yao:2013:DRB

[Yao13] Yi-Ching Yao. A duality relation between the workload and attained waiting time in FCFS G/G/s queues. Journal of Applied Probability, 50(1):300–307, March 2013. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http:/

/projecteuclid.org/euclid.jap/1363784441.

Yap:2009:SSC

[Yap09] V. B. Yap. Similar states in continuous-time Markov chains. Journal of Applied Probability, 46(2):497–506, June 2009. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/25662439.

Yeh:1988:PP

[YAR88] Hsiaw-Chan Yeh, Barry C. Arnold, and Christopher A. Robertson. Pareto processes. Journal of Applied Probability, 25(2):291– 301, June 1988. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3214437.

Yasuda:1983:SPI

[Yas83] Masami Yasuda. On a stopping problem involving refusal and forced stopping. Journal of Applied Probability, 20(1):71–81, March 1983. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3213721.

Yasuda:1984:ARB

[Yas84] Masami Yasuda. Asymptotic results for the best-choice problem with a random number of objects. Journal of Applied Probability, 21(3):521–536, September 1984. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.

org/stable/3213614.

Yasuda:1988:OVM

[Yas88] Masami Yasuda. The optimal value of Markov stopping problems with one-step look ahead policy. Journal of Applied Probability, 25(3):544–552, September 1988. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3213983.

Yashin:1993:ECM

[Yas93] A. I. Yashin. An extension of the Cameron–Martin result. Journal of Applied Probability, 30(1):247–251, March 1993. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3214636.

Yi:2023:OSS

[YBL23] He Yi, Narayanaswamy Balakrishnan, and Xiang Li. On ordered system signature and its dynamic version for coherent systems with applications. Journal of Applied Probability, 60(3):982–1002, September 2023. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL https:/

/www.cambridge.org/core/journals/journal-of-applied-
probability/article/on-ordered-system-signature-and-
its-dynamic-version-for-coherent-systems-with-applications/
819571E5F1F4D50E917804DC3BB6C853.

Yang:1982:ITS

[YC82] Grace Yang and C. L. Chiang. On interarrival times in simple stochastic epidemic models. Journal of Applied Probability, 19 (4):835–841, December 1982. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.

org/stable/3213836.

Yang:2001:NGM

[YC01] Won S. Yang and Kyung C. Chae. A note on the GI/M/1 queue with Poisson negative arrivals. Journal of Applied Probability, 38(4):1081–1085, December 2001. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3215788.

Yao:2008:SOB

[YC08] Yi-Ching Yao and May-Ru Chen. Strong optimality of bold play for discounted dubins-Savage gambling problems with timedependent parameters. Journal of Applied Probability, 45(2): 403–416, June 2008. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/27595953.

Yao:2011:LDG

[YCG11] Chang-Long Yao, Ge Chen, and Tian-De Guo. Large deviations for the graph distance in supercritical continuum percolation.

Journal of Applied Probability, 48(1):154–172, March 2011. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/29777452.

Yao:2015:OCP

[YCW15] Dacheng Yao, Xiuli Chao, and Jingchen Wu. Optimal control policy for a Brownian inventory system with concave ordering cost. Journal of Applied Probability, 52(4):909–925, December 2015. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://projecteuclid.org/euclid.jap/
1450802743.

Ye:2007:PAC

[Ye07] Heng-Qing Ye. A paradox for admission control of multiclass queueing network with differentiated service. Journal of Applied Probability, 44(2):321–331, June 2007. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/27595844.

Yeh:1985:OCF

[Yeh85] Lam Yeh. Optimal control of a finite dam: Average-cost case. Journal of Applied Probability, 22(2):480–484, June 1985. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/3213793.

Yeh:1991:ORR

[Yeh91] Lam Yeh. An optimal repairable replacement model for deteriorating systems. Journal of Applied Probability, 28(4):843–851, December 1991. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3214687.

Yeh:1995:OIR

[Yeh95] Lam Yeh. An optimal inspection-repair-replacement policy for standby systems. Journal of Applied Probability, 32(1):212– 223, March 1995. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3214931.

Yeh:1997:ROF

[Yeh97] Lam Yeh. The rate of occurrence of failures. Journal of Applied Probability, 34(1):234–247, March 1997. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http:/

/www.jstor.org/stable/3215190.

Yeh:2003:IRR

[Yeh03] Lam Yeh. An inspection-repair-replacement model for a deteriorating system with unobservable state. Journal of Applied Probability, 40(4):1031–1042, December 2003. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/3216058.

Yeh:2005:MPM

[Yeh05] Lam Yeh. A monotone process maintenance model for a multistate system. Journal of Applied Probability, 42(1):1–14, March 2005. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/30040765.

Yeo:1974:FDE

[Yeo74] G. F. Yeo. A finite dam with exponential release. Journal of Applied Probability, 11(1):122–133, March 1974. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3212589.

Yeo:1975:FDV

[Yeo75] G. F. Yeo. A finite dam with variable release rate. Journal of Applied Probability, 12(1):205–211, March 1975. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3212431.

Yeo:1997:DSP

[Yeo97] Geoffrey F. Yeo. Duration of a secretary problem. Journal of Applied Probability, 34(2):556–558, June 1997. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/3215394.

Yuen:2005:URD

[YGN05] Kam C. Yuen, Junyi Guo, and Kai W. Ng. On ultimate ruin in a delayed-claims risk model. Journal of Applied Probability, 42 (1):163–174, March 2005. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/30040777.

Yeh:1987:OCF

[YH87] Lam Yeh and Lou Jiann Hua. Optimal control of a finite dam: Wiener process input. Journal of Applied Probability, 24(1):186– 199, March 1987. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3214070.

Yang:2004:MDP

[YH04] Jiing-Ru Yang and Shoou-Ren Hsiau. On Markov-dependent parking problems. Journal of Applied Probability, 41(2):579– 586, June 2004. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3216038.

Yao:1999:IND

[YI99] Yi-Ching Yao and Hari Iyer. On an inequality for the normal distribution arising in bioequivalence studies. Journal of Applied Probability, 36(1):279–286, March 1999. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http:/

/www.jstor.org/stable/3215422.

Yin:2021:SOM

[Yin21] Chuancun Yin. Stochastic orderings of multivariate elliptical distributions. Journal of Applied Probability, 58(2):551–568, June 2021. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-

6072 (electronic). URL https://www.cambridge.org/core/
journals/journal-of-applied-probability/article/stochastic-
orderings-of-multivariate-elliptical-distributions/626C0740E497D8FA0BE0C660C2182CA8.

You:2016:OSP

[YL16] Yinping You and Xiaohu Li. Ordering scalar products with applications in financial engineering and actuarial science. Journal of Applied Probability, 53(1):47–56, March 2016. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://projecteuclid.org/euclid.jap/1457470557.

Yamazaki:1991:EWT

[YM91] Genji Yamazaki and Masakiyo Miyazawa. The equality of the workload and total attained waiting time in average. Journal of Applied Probability, 28(1):238–244, March 1991. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3214755.

Yao:2017:CDA

[YML17] Yi-Ching Yao, Daniel Wei-Chung Miao, and Xenos Chang-Shuo Lin. Corrected discrete approximations for the conditional and unconditional distributions of the continuous scan statistic.

Journal of Applied Probability, 54(1):304–319, March 2017. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL https://www.cambridge.org/core/journals/journal-
of-applied-probability/article/corrected-discrete-approximations-
for-the-conditional-and-unconditional-distributions-
of-the-continuous-scan-statistic/2DB8D57F568C8DA501300A021B6AA70F.

Yor:1992:CFE

[Yor92] Marc Yor. Sur certaines fonctionnelles exponentielles du mouvement brownien réel. (French) [Some functional features of real exponential Brownian motion]. Journal of Applied Probability, 29(1):202–208, March 1992. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3214805.

Yor:1995:DBQ

[Yor95] Marc Yor. The distribution of Brownian quantiles. Journal of Applied Probability, 32(2):405–416, June 1995. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3215296.

Yoshida:1984:PMA

[Yos84] Minoru Yoshida. Probability maximizing approach to a secretary problem with random change-point of the distribution law of the observed process. Journal of Applied Probability, 21(1):98– 107, March 1984. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3213668.

Yang:2017:PRD

[YSL17] Chen Yang, Kristian P. Sendova, and Zhong Li. On the Parisian ruin of the dual Lévy risk model. Journal of Applied Probability, 54(4):1193–1212, December 2017. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL https://www.cambridge.org/core/journals/journal-
of-applied-probability/article/on-the-parisian-ruin-
of-the-dual-levy-risk-model/51232043AD2C82F91F60ADE3B66C661B.

Yamazaki:1991:CLS

[YSS91] Genji Yamazaki, Hirotaka Sakasegawa, and J. George Shanthikumar. A conservation law for single-server queues and its applications. Journal of Applied Probability, 28(1):198–209, March 1991. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3214750.

Yeh:1983:ACM

[YT83] Lam Yeh and L. C. Thomas. Adaptive control of M/M/1 queues: Continuous-time Markov decision process approach. Journal of Applied Probability, 20(2):368–379, June 1983. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/3213809.

Yu:1974:SBH

[Yu74] Oliver S. Yu. Stochastic bounds for heterogeneous-server queues with Erlang service times. Journal of Applied Probability, 11(4): 785–796, December 1974. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3212561.

Yu:2009:SOE

[Yu09] Yaming Yu. Stochastic ordering of exponential family distributions and their mixtures. Journal of Applied Probability, 46 (1):244–254, March 2009. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/25662418.

Yu:2011:CRF

[Yu11] Yaming Yu. Concave renewal functions do not imply DFR interrenewal times. Journal of Applied Probability, 48(2):583– 588, June 2011. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
23065905.

Yu:2017:MVC

[Yu17] Yaming Yu. Minimum variance in the coupon collector’s problem. Journal of Applied Probability, 54(2):655–656, June 2017. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL https://www.cambridge.org/core/journals/
journal-of-applied-probability/article/minimum-variance-
in-the-coupon-collectors-problem/595441C616360F7A5EE442D4D6405700.

Yukich:2006:USS

[Yuk06] J. E. Yukich. Ultra-small scale-free geometric networks. Journal of Applied Probability, 43(3):665–677, September 2006. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/27595763.

Yun:2000:DCF

[Yun00] Seokhoon Yun. The distributions of cluster functionals of extreme events in a d th-order Markov chain. Journal of Applied Probability, 37(1):29–44, March 2000. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3215657.

Yeo:1964:DRT

[YW64] G. F. Yeo and B. Weesakul. Delays to road traffic at an intersection. Journal of Applied Probability, 1(2):297–310, December 1964. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3211861.

Yu:2021:UBB

[YW21] Gaoran Yu and John C. Wierman. An upper bound for the bond percolation threshold of the cubic lattice by a growth process approach. Journal of Applied Probability, 58 (3):677–692, September 2021. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL https:/

/www.cambridge.org/core/journals/journal-of-applied-
probability/article/an-upper-bound-for-the-bond-percolation-
threshold-of-the-cubic-lattice-by-a-growth-process-approach/
C1FB01EDD1DB4B08DA84153F79FF2B20.

Yang:2014:UAD

[YWK14] Yang Yang, Kaiyong Wang, and Dimitrios G. Konstantinides. Uniform asymptotics for discounted aggregate claims in dependent risk models. Journal of Applied Probability, 51(3):669– 684, September 2014. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://projecteuclid.

org/euclid.jap/1409932666.

Yang:1982:OSR

[YWR82] Mark C. K. Yang, Dennis D. Wackerly, and Andrew Rosalsky. Optimal stopping rules in proofreading. Journal of Applied Probability, 19(3):723–729, September 1982. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/3213535.

Yano:2011:REM

[YY11] Kouji Yano and Kenji Yasutomi. Realization of an ergodic Markov chain as a random walk subject to a synchronizing road coloring. Journal of Applied Probability, 48(3):766–777, September 2011. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/23065935.

Yam:2009:TRB

[YYZ09] S. C. P. Yam, S. P. Yung, and W. Zhou. Two rationales behind the ‘Buy-and-Hold Or Sell-at-Once’ strategy. Journal of Applied Probability, 46(3):651–668, September 2009. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/25662453.

Yadin:1982:RCC

[YZ82] M. Yadin and S. Zacks. Random coverage of a circle with applications to a shadowing problem. Journal of Applied Probability, 19(3):562–577, September 1982. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.

org/stable/3213514.

Yadin:1985:VSM

[YZ85] M. Yadin and S. Zacks. The visibility of stationary and moving targets in the plane subject to a Poisson field of shadowing elements. Journal of Applied Probability, 22(4):776–786, December 1985. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3213945.

Yin:2006:NAS

[YZ06] Chuancun Yin and Junsheng Zhao. Nonexponential asymptotics for the solutions of renewal equations, with applications. Journal of Applied Probability, 43(3):815–824, September 2006. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/27595774.

Yan:2021:OAR

[YZZ21] Rongfang Yan, Jiandong Zhang, and Yiying Zhang. Optimal allocation of relevations in coherent systems. Journal of Applied Probability, 58(4):1152–1169, December 2021. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL https://www.cambridge.org/core/journals/journal-
of-applied-probability/article/optimal-allocation-of-
relevations-in-coherent-systems/D2909E8A40E6128C9939747BF5E2DF2C.

Zachary:2000:DLU

[Zac00] Stan Zachary. Dynamics of large uncontrolled loss networks. Journal of Applied Probability, 37(3):685–695, September 2000. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/3215605.

Zacks:2004:GIT

[Zac04] S. Zacks. Generalized integrated telegraph processes and the distribution of related stopping times. Journal of Applied Probability, 41(2):497–507, June 2004. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3216031.

Zachary:2007:NIS

[Zac07] Stan Zachary. A note on insensitivity in stochastic networks. Journal of Applied Probability, 44(1):238–248, March 2007. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/27595834.

Ziya:2006:MRG

[ZAFP06] Serhan Ziya, Hayriye Ayhan, Robert D. Foley, and Erol Peköz. A monotonicity result for a G/GI/c queue with balking or reneging. Journal of Applied Probability, 43(4):1201–1205, December 2006. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/27595812.

Zahle:1990:WCP

[Zäh90] M. Zähle. Wicksell’s corpuscle problem in spherical spaces. Journal of Applied Probability, 27(3):701–706, September 1990. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/3214553.

Zajic:1995:LEU

[Zaj95] Tim Zajic. Large exceedences for uniformly recurrent Markovadditive processes and strong-mixing stationary processes. Journal of Applied Probability, 32(3):679–691, September 1995. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/3215122.

Zeephongsekul:1995:OSR

[ZC95] P. Zeephongsekul and C. Chiera. Optimal software release policy based on a two-person game of timing. Journal of Applied Probability, 32(2):470–481, June 1995. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3215301.

Zhou:2014:ODR

[ZC14] Ming Zhou and Jun Cai. Optimal dynamic risk control for insurers with state-dependent income. Journal of Applied Probability, 51(2):417–435, June 2014. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://projecteuclid.

org/euclid.jap/1402578634.

Zeifman:1989:QEN

[Zei89] A. I. Zeifman. Quasi-ergodicity for non-homogeneous continuoustime Markov chains. Journal of Applied Probability, 26(3):643– 648, September 1989. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3214422.

Zeifman:1991:SER

[Zei91] A. I. Zeifman. Some estimates of the rate of convergence for birth and death processes. Journal of Applied Probability, 28 (2):268–277, June 1991. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3214865.

Zeifman:1995:EPB

[Zei95] A. I. Zeifman. On the estimation of probabilities for birth and death processes. Journal of Applied Probability, 32(3):623–634, September 1995. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/

3215117. See note [Mit04a].

Zeitouni:1998:SDG

[Zei98] O. Zeitouni. Superexponential decay for the gem process. Journal of Applied Probability, 35(3):776–781, September 1998. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/3215652.

Zeng:2010:SDR

[Zen10] Xudong Zeng. A stochastic differential reinsurance game. Journal of Applied Probability, 47(2):335–349, June 2010. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/20778483.

Ziezold:1988:CIR

[ZG88] Herbert Ziezold and Christian Grillenberger. On the critical infection rate of the one-dimensional basic contact process: Numerical results. Journal of Applied Probability, 25(1):1–8, March 1988. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3214228.

Zeevi:2004:RPA

[ZG04] Assaf Zeevi and Peter W. Glynn. Recurrence properties of autoregressive processes with super-heavy-tailed innovations. Journal of Applied Probability, 41(3):639–653, September 2004. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/4141343.

Zhang:1992:SPO

[Zha92] Yuan Lin Zhang. Some problems on a one-dimensional correlated random walk with various types of barrier. Journal of Applied Probability, 29(1):196–201, March 1992. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http:/

/www.jstor.org/stable/3214804.

Zhang:1994:BOR

[Zha94] Yuan Lin Zhang. A bivariate optimal replacement policy for a repairable system. Journal of Applied Probability, 31(4):1123– 1127, December 1994. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3215336.

Zhang:1996:CPP

[Zha96] Yu Zhang. Continuity of percolation probability in ∞ + 1 dimensions. Journal of Applied Probability, 33(2):427–433, June 1996. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3215066.

Zhang:2001:SES

[Zha01] Yu-Hui Zhang. Strong ergodicity for single-birth processes. Journal of Applied Probability, 38(1):270–277, March 2001. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3215760.

Zhang:2004:LDS

[Zha04] Mei Zhang. Large deviations for super-Brownian motion with immigration. Journal of Applied Probability, 41(1):187–201, March 2004. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3215824.

Zhang:2005:LDP

[Zha05] Mei Zhang. A large deviation principle for a Brownian immigration particle system. Journal of Applied Probability, 42(4): 1120–1133, December 2005. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/30040833.

Zhang:2010:MRI

[Zha10] Zhengcheng Zhang. Mixture representations of inactivity times of conditional coherent systems and their applications. Journal of Applied Probability, 47(3):876–885, September 2010. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/20778522.

Zhang:2014:AOC

[Zha14] Yi Zhang. Average optimality for continuous-time Markov decision processes under weak continuity conditions. Journal of Applied Probability, 51(4):954–970, December 2014. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://projecteuclid.org/euclid.jap/1421763321.

Zhang:2021:TOL

[Zha21] Yiying Zhang. On transform orders for largest claim amounts. Journal of Applied Probability, 58(4):1064–1085, December 2021. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL https://www.cambridge.org/core/journals/
journal-of-applied-probability/article/on-transform-
orders-for-largest-claim-amounts/F6119DBA3849DA7E40EFD00D6375D023.

Zhang:2022:LDP

[Zha22] Shuxiong Zhang. On large-deviation probabilities for the empirical distribution of branching random walks with heavy tails. Journal of Applied Probability, 59(2):471–494, June 2022. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL https://www.cambridge.org/core/journals/journal-
of-applied-probability/article/on-largedeviation-probabilities-
for-the-empirical-distribution-of-branching-random-walks-
with-heavy-tails/17FD6DC949C9264C774BC8CCB8B7BBB5.

Zheng:1991:SPO

[Zhe91] Yu-Sheng Zheng. A simple proof for optimality of (s, s) policies in infinite-horizon inventory systems. Journal of Applied Probability, 28(4):802–810, December 1991. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/3214683.

Zheng:2009:RAL

[Zhe09] Qi Zheng. Remarks on the asymptotics of the Luria–Delbrück and related distributions. Journal of Applied Probability, 46(4): 1221–1224, December 2009. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/25662493.

Zhou:2004:SFI

[Zho04] Xiaowen Zhou. Some fluctuation identities for Lévy processes with jumps of the same sign. Journal of Applied Probability, 41 (4):1191–1198, December 2004. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.

org/stable/4141391.

Zhou:2007:EPS

[Zho07] Xiaowen Zhou. Exit problems for spectrally negative Lévy processes reflected at either the supremum or the infimum. Journal of Applied Probability, 44(4):1012–1030, December 2007. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/27595904.

Zhou:2008:CBD

[Zho08] Yang Zhou. Correction to Bingham and Doney (1988). Journal of Applied Probability, 45(4):1204–1206, December 2008. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/27596019. See [BD88b].

Zhou:2015:EIT

[Zho15] Youzhou Zhou. Ergodic inequality of a two-parameter infinitelymany-alleles diffusion model. Journal of Applied Probability, 52 (1):238–246, March 2015. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://projecteuclid.

org/euclid.jap/1429282618.

Zhou:2023:TDI

[Zho23] Youzhou Zhou. Transition density of an infinite-dimensional diffusion with the jack parameter. Journal of Applied Probability, 60(3):797–811, September 2023. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL https:/

/www.cambridge.org/core/journals/journal-of-applied-
probability/article/transition-density-of-an-infinitedimensional-
diffusion-with-the-jack-parameter/01554EBF0C18CCC75987FD76745253E1.

Zhu:2013:CLT

[Zhu13] Lingjiong Zhu. Central limit theorem for nonlinear Hawkes processes. Journal of Applied Probability, 50(3):760–771, September 2013. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-

6072 (electronic). URL http://projecteuclid.org/euclid.
jap/1378401234.

Zhu:2014:LTC

[Zhu14] Lingjiong Zhu. Limit theorems for a Cox–Ingersoll–Ross process with Hawkes jumps. Journal of Applied Probability, 51(3): 699–712, September 2014. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://projecteuclid.

org/euclid.jap/1409932668.

Zhu:2023:LDT

[Zhu23] Yaping Zhu. A large deviation theorem for a supercritical super-Brownian motion with absorption. Journal of Applied Probability, 60(4):1249–1274, December 2023. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL https:/

/www.cambridge.org/core/journals/journal-of-applied-
probability/article/large-deviation-theorem-for-a-supercritical-
superbrownian-motion-with-absorption/CAE3C715A979319549944FCE981F9007.

Ziedins:1987:QSD

[Zie87] Ilze Ziedins. Quasi-stationary distributions and one-dimensional circuit-switched networks. Journal of Applied Probability, 24(4): 965–977, December 1987. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3214219.

Zijlstra:1983:CGD

[Zij83] Mynt Zijlstra. Characterizations of the geometric distribution by distributional properties. Journal of Applied Probability, 20(4):

843–850, December 1983. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3213595.

Zipkin:2011:DSD

[Zip11] P. Zipkin. On Durbin’s series for the density of first passage times. Journal of Applied Probability, 48(3):713–722, September 2011. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/23065931.

Zhunwei:2004:NAB

[ZJ04] Lu Zhunwei and Peter Jagers. A note on the asymptotic behaviour of the extinction probability in supercritical populationsize-dependent branching processes with independent and identically distributed random environments. Journal of Applied Probability, 41(1):176–186, March 2004. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3215823.

Zhao:1996:CMC

[ZL96] Y. Quennel Zhao and Danielle Liu. The censored Markov chain and the best augmentation. Journal of Applied Probability, 33 (3):623–629, September 1996. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.

org/stable/3215344.

Zhang:2010:SNR

[ZL10] Shuhong Zhang and Xiaohu Li. Some new results on the moment generating function order and related life distributions. Journal of Applied Probability, 47(4):923–933, December 2010. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/29777416.

Zhang:2015:MMC

[ZL15] Lina Zhang and Junping Li. The M/M/c queue with mass exodus and mass arrivals when empty. Journal of Applied Probability, 52(4):990–1002, December 2015. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http:/

/projecteuclid.org/euclid.jap/1450802748.

Zhao:1999:DRB

[ZLA99] Yiqiang Q. Zhao, Wei Li, and Attahiru Sule Alfa. Duality results for block-structured transition matrices. Journal of Applied Probability, 36(4):1045–1057, December 1999. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3215577.

Zhao:2008:COS

[ZLB08] Peng Zhao, Xiaohu Li, and N. Balakrishnan. Conditional ordering of k-out-of-n systems with independent but nonidentical components. Journal of Applied Probability, 45(4):1113–1125, December 2008. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
27596012.

Zhang:2006:SLP

[ZLW06] Yi Zhang, Zhengyan Lin, and Chengguo Weng. Some limiting properties of the bounds of the present value function of a life insurance portfolio. Journal of Applied Probability, 43(4): 1155–1164, December 2006. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/27595806.

Zuoxiang:2010:CRE

[ZNF10] Peng Zuoxiang, Saralees Nadarajah, and Lin Fuming. Convergence rate of extremes for the general error distribution. Journal of Applied Probability, 47(3):668–679, September 2010. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/20778509.

Zhu:2008:BOO

[ZPR08] Quanxin Zhu and Tomás Prieto-Rumeau. Bias and overtaking optimality for continuous-time jump Markov decision processes in Polish spaces. Journal of Applied Probability, 45(2):417– 429, June 2008. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
27595954.

Zheng:2001:EIV

[ZR01] Xiaogu Zheng and James Renwick. Estimating interannual variability arising from weather events. Journal of Applied Probability, 38A:274–288, 2001. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/3215886.

Zaludova:1968:SSR

[ZRU68] A. H. Zaludová, Z. Rezný, and M. Ullrich. A survey of some recent Czechoslovak work in automatic statistical process control. Journal of Applied Probability, 5(1):43–54, April 1968. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3212076.

Zhang:1992:NBS

[ZS92] Jianliang Zhang and Rakesh Shukla. A note on “The Statistical Analysis of Direct Repeats in Nucleic Acid Sequences”. Journal of Applied Probability, 29(3):750–752, September 1992. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3214911. See [SS85b].

Zhang:2004:BMM

[ZS04] Zhengjun Zhang and Richard L. Smith. The behavior of multivariate maxima of moving maxima processes. Journal of Applied Probability, 41(4):1113–1123, December 2004. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/4141383.

Zhang:2001:SDP

[ZT01] Zhiqiang Zhang and Howell Tong. On some distributional properties of a first-order nonnegative bilinear time series model. Journal of Applied Probability, 38(3):659–671, September 2001. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/3216119.

Zhang:2021:PDD

[ZTML21] Zhenzhong Zhang, Jinying Tong, Qingting Meng, and You Liang. Population dynamics driven by truncated stable processes with Markovian switching. Journal of Applied Probability, 58(2):505–522, June 2021. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL https:/

/www.cambridge.org/core/journals/journal-of-applied-
probability/article/population-dynamics-driven-by-truncated-
stable-processes-with-markovian-switching/6E341C139EE236C788DA8B68495979CD.

Zuckerman:1977:TSO

[Zuc77] Dror Zuckerman. Two-stage output procedure of a finite dam. Journal of Applied Probability, 14(2):421–425, June 1977. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://www.jstor.org/stable/3213017.

Zuckerman:1978:OSS

[Zuc78] Dror Zuckerman. Optimal stopping in a semi-Markov shock model. Journal of Applied Probability, 15(3):629–634, September 1978. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3213126.

Zuckerman:1980:DPM

[Zuc80a] Dror Zuckerman. A diffusion process model for the optimal investment strategies of an R&D project. Journal of Applied Probability, 17(3):646–653, September 1980. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/3212958.

Zuckerman:1980:IRP

[Zuc80b] Dror Zuckerman. Inspection and replacement policies. Journal of Applied Probability, 17(1):168–177, March 1980. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3212934.

Zuckerman:1981:COC

[Zuc81] Dror Zuckerman. Characterization of the optimal class of output policies in a control model of a finite dam. Journal of Applied Probability, 18(4):913–923, December 1981. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/3213065.

Zuckerman:1983:JSC

[Zuc83] Dror Zuckerman. Job search: The continuous case. Journal of Applied Probability, 20(3):637–648, September 1983. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3213899.

Zuckerman:1986:OSC

[Zuc86] Dror Zuckerman. Optimal stopping in a continuous search model. Journal of Applied Probability, 23(2):514–518, June 1986. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL http://www.jstor.org/stable/3214193.

Zuckerman:1989:OIP

[Zuc89] Dror Zuckerman. Optimal inspection policy for a multi-unit machine. Journal of Applied Probability, 26(3):543–551, September 1989. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3214412.

Zhang:2002:TDN

[ZW02] Chunsheng Zhang and Rong Wu. Total duration of negative surplus for the compound Poisson process that is perturbed by diffusion. Journal of Applied Probability, 39(3):517–532, September 2002. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3216075.

Zhou:2017:DRL

[ZW17] Jiang Zhou and Lan Wu. The distribution of refracted Lévy processes with jumps having rational Laplace transforms. Journal of Applied Probability, 54(4):1167–1192, December 2017. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (elec-

tronic). URL https://www.cambridge.org/core/journals/
journal-of-applied-probability/article/distribution-
of-refracted-levy-processes-with-jumps-having-rational-
laplace-transforms/6B959B09A2C6718E2D60F44D5CFCD214.

Zhao:2010:CLT

[ZWV10] Ou Zhao, Michael Woodroofe, and Dalibor Volný. A central limit theorem for reversible processes with nonlinear growth of variance. Journal of Applied Probability, 47(4):1195–1202, December 2010. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/29777437.

Zeng:2017:OBT

[ZX17] Yun Zeng and Cathy Honghui Xia. Optimal bulking threshold of batch service queues. Journal of Applied Probability, 54(2):409–423, June 2017. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL https:/

/www.cambridge.org/core/journals/journal-of-applied-
probability/article/optimal-bulking-threshold-of-batch-
service-queues/C1DD49670E8DFF9B2E55F82BF049BC29.

Zacks:1970:ACO

[ZY70] S. Zacks and M. Yadin. Analytic characterization of the optimal control of a queueing system. Journal of Applied Probability, 7 (3):617–633, December 1970. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.

org/stable/3211941.

Zhu:2008:EAR

[ZY08] Jinxia Zhu and Hailiang Yang. Estimates for the absolute ruin probability in the compound Poisson risk model with credit and debit interest. Journal of Applied Probability, 45(3):818–830, September 2008. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
27595988.

Zachary:1999:LNM

[ZZ99] Stan Zachary and Ilze Ziedins. Loss networks and Markov random fields. Journal of Applied Probability, 36(2):403–414, June 1999. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3215464.

Zhu:2004:MGQ

[ZZ04] Yijun Zhu and Zhe George Zhang. M/GI/1 queues with services of both positive and negative customers. Journal of Applied Probability, 41(4):1157–1170, December 2004. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/4141387.

Zhang:2009:TLT

[ZZ09] Shibin Zhang and Xinsheng Zhang. On the transition law of tempered stable Ornstein–Uhlenbeck processes. Journal of Applied Probability, 46(3):721–731, September 2009. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/25662457.

Zhang:2013:DAQ

[ZZ13] Hanjun Zhang and Yixia Zhu. Domain of attraction of the quasistationary distribution for birth-and-death processes. Journal of Applied Probability, 50(1):114–126, March 2013. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://projecteuclid.org/euclid.jap/1363784428.

Zang:2016:GLB

[ZZ16] Qing-Pei Zang and Li-Xin Zhang. A general lower bound of parameter estimation for reflected Ornstein–Uhlenbeck processes. Journal of Applied Probability, 53(1):22–32, March 2016. CO-DEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL http://projecteuclid.org/euclid.jap/1457470555.